A Complete Bibliography of Publications in Ecological Modelling (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

05 October 2023 Version 1.07

Title word cross-reference

1 [AHNK13, MBUM12]. $129.95 [Jör11a]. $19.99 [Jør14a]. 2 [AHNK13, Con11, Con12, VLF^{+}17]. $29.99 [Jør14a]. 3 [BVGR^{+}17, BCB^{+}11, CPS^{+}14, DBP16, IFC14, KPS17, LMT13, MSGG14, MRB14, MBUM12, MJGC13, NGR^{+}11, PMM11, SRHL11, SMH16]. $3.99 [Jør14a]. 3/4 [SIS^{+}14]. $61.42 [Jør14b]. ^{+} [HBA16a]. ^{1} [Mob18]. ^{13} [EFM^{+}10, Vav12, VRL18]. ^{14} [HKlL14, KHW^{+}16]. ^{15} [TGM^{+}21]. ^{N+} [HAA^{+}21]. _{2} [CGY11, CSH11, CSH13, CN13, CR11, CR12b, FCG^{+}19, FCZ13, GKW^{+}13, HMS^{+}11, Hol13, KK12b, KMM^{+}13, LKLK15, LYJ^{+}15, LHP16, LBdA^{+}19, LSM19, MKF^{+}11, mNkGyT^{+}17, PJ16, PD15, QlSeZ^{+}10, RML^{+}15, SLSS15, TIC^{+}14, VRL18, WRM^{+}11, ZJP^{+}11]. _{3} [MKTO19]. _{4} [DAW^{+}10, GKW^{+}13, HMS^{+}11, Vav12, VRL18]. _{wetland} [LHZS10]. a [KHJ^{+}19, LGY10, MMM^{+}16, WQD^{+}13, ZZT13]. α [Sci11a]. β [Sci11a]. C [LKKK^{+}14, SVL12, WR18]. : [JS12, LWGH13]. Δ [HKlL14, EFM^{+}10, Vav12]. G [MSG^{+}13]. K [ZJB19]. λ_{1} [Log13b]. M_{b} [NP15]. N [GHS^{+}16a, LKKK^{+}14, OTA16, SVL12, WR18, ZML^{+}12]. R∗ [JD10]. R^{2} [GO15a]. ∼ [BCT15]. θ [MCA18]. × [MWD^{+}15].

-aggregated [ZJB19]. -based [GO15a]. -concentration [CN13]. -D [Con12, IFC14, MBUM12, VLF^{+}17, Con11, MBUM12]. -exponent [SIS^{+}14]. -logistic [MCA18].

/H [KMM^{+}13].

0D [MG18, MHG18].

1 [Ano14u, Ano15q, Ano15p, Ano19a, Ano19b, Ano19c, Ano19-36, Ano19-33, Ano19-39, Ano19-38, Ano19-46, Ano19-41, Ano19-44, Ano19d, BLD^{+}10a, DFM12a]. 1-144 [Ano15p]. 1-246 [Ano15q]. 1-60 [Ano14u]. 10 [Ano12-37, Ano12-28, Ano12-30, Ano12y, Ano12z, Ano12x, Ano12-31, Ano12-32, Ano12-35, Ano13-39, Ano13-32, Ano13-33, Ano13-29, Ano13y, Ano13-35, Ano13-37, Ano13x, Ano13-27, Ano13-30, Ano14u, Ano14e, Ano14r, Ano14j, Ano14m, Ano14q, Ano14b, Ano14d, Ano14i, Ano14g, Ano15q, Ano15a, Ano15b, Ano15l, Ano15m, Ano15f, Ano15c, Ano15h, Ano15o, Ano16t, Ano16a, Ano16q, Ano16e, Ano16d, Ano16f, Ano16c, Ano16l, Ano16n, Ano16g, Ano17-38, Ano17-35, Ano17x, Ano17-29, Ano17z, Ano17u, Ano17-39, Ano17y, Ano17-32, Ano17-31, Ano17-34, Ano18-41, Ano18-46, Ano18-34, Ano18y, Ano18-42, Ano18-39, Ano18-44, Ano18-27, Ano18-36, Ano18-31, Ano18-30, Ano18-29, Ano19-42, Ano21b]. 10-year [GZK^{+}11]. 144 [Ano15p]. 15 [Ano19e, Ano19f, Ano19g, Ano19h]. 156 [Ano12-37]. 180 [Och10]. 1950s [KHL^{+}16]. 1980s [KHL^{+}16].

2 [BVGR^{+}17, BlMG^{+}11, DFM12b, MAO19]. 2-methylisoborneol [HD14, HD16]. 2000s [KHL^{+}16]. 2009 [LYP^{+}11, LPA^{+}11]. 2011 [Ano11p, Ano11q]. 2012 [Ano12-37, Ano12-28, Ano12-29, Ano12-30, Ano12-36, Ano12y, Ano12-33, Ano12-34, Ano12v, Ano12z, Ano12x, Ano12-31, Ano12u, Ano12-32, Ano12-35, Ano12-27]. 2013 [Ano13-39, Ano13-38, Ano13-32, Ano13-36, Ano13-33, Ano13-31, Ano13-29, Ano13y, Ano13-34, Ano13-35, Ano13-40, Ano13w, Ano13-37, Ano13v, Ano13x, Ano13-27, Ano13z, Ano13-30, Ano13-28]. 2014 [Ano14u, Ano14e, Ano14c, Ano14r, Ano14o, Ano14f, Ano14j, Ano14s, Ano14k, Ano14m, Ano14p, Ano14q, Ano14n, Ano14b, Ano14t, Ano14d, Ano14i, Ano14l, Ano14g, Ano14h]. 2015 [Ano15q, Ano15p, Ano15a, Ano15j, Ano15b, Ano15d, Ano15l, Ano15m, Ano15k, Ano15f, Ano15e, Ano15c, Ano15n, Ano15g, Ano15h, Ano15o, Ano15i, Ano21b]. 2016 [Ano16t, Ano16a, Ano16u, Ano16q, Ano16m, Ano16e, Ano16h, Ano16o, Ano16d, Ano16r, Ano16b, Ano16f, Ano16p, Ano16c, Ano16i, Ano16l, Ano16k, Ano16n, Ano16j, Ano16g, Ano16s]. 2017 [Ano17-38, Ano17r, Ano17-35, Ano17-36, Ano17x, Ano17-37, Ano17-29, Ano17s, Ano17z, Ano17-40, Ano17-30, Ano17u, Ano17-28, Ano17-39, Ano17w, Ano17y, Ano17-27, Ano17-32, Ano17v, Ano17-31, Ano17t, Ano17-34, Ano17-33]. 2018 [Ano18-41, Ano18-38, Ano18-46, Ano18-47, Ano18-34, Ano18-48, Ano18y, Ano18-40, Ano18-42, Ano18-33, Ano18-39, Ano18-37, Ano18-44, Ano18-35, Ano18-27, Ano18-43, Ano18-36, Ano18-45, Ano18-31, Ano18-28, Ano18-30, Ano18z, Ano18-29, Ano18-32]. 2019 [Ano19a, Ano19e, Ano19b, Ano19f, Ano19c, Ano19g, Ano19-36, Ano19-47, Ano19-33, Ano19-37, Ano19-39, Ano19-45, Ano19-42, Ano19-40, Ano19-38, Ano19-34, Ano19-46, Ano19-35, Ano19-41, Ano19-48, Ano19-44, Ano19-43, Ano19d, Ano19h]. 20th [Fel12, NMS^{+}14, WTC^{+}18]. 21st [Fel12]. 220 [BRV11, YTT^{+}11]. 221 [Bar11, Lin16, Pat10, WLS^{+}11]. 222 [BMF^{+}11a, Con12, CR12b, ERGW^{+}11a, GORJ12, WGA^{+}12]. 225 [ABR12b]. 227 [Ano12a]. 24 [Ano12-29, Ano12-33, Ano12-34, Ano12v, Ano12u, Ano12-27, Ano13-38, Ano13-31, Ano13-34, Ano13-40, Ano13w, Ano13v, Ano13z, Ano13-28, Ano14c, Ano14o, Ano14f, Ano14s, Ano14k, Ano14p, Ano14n, Ano14t, Ano14l, Ano14h, Ano15p, Ano15j, Ano15d, Ano15k, Ano15e, Ano15n, Ano15g, Ano15i, Ano16u, Ano16m, Ano16h, Ano16o, Ano16r, Ano16b, Ano16p, Ano16i, Ano16k, Ano16j, Ano16s, Ano17r, Ano17-36, Ano17-37, Ano17s, Ano17-40, Ano17-30, Ano17-28, Ano17w, Ano17-27, Ano17v, Ano17t, Ano17-33, Ano18-38, Ano18-47, Ano18-48, Ano18-40, Ano18-33, Ano18-37, Ano18-35, Ano18-43, Ano18-45, Ano18-28, Ano18z, Ano18-32, Ano19-47, Ano19-37, Ano19-45, Ano19-40, Ano19-34, Ano19-35, Ano19-48, Ano19-43]. 246 [Ano15q]. 247 [KMS^{+}13]. 250 [HLLO13a]. 258 [MJL^{+}14]. 273 [MSCH14a]. 295 [Jør15a]. 297 [CCW^{+}15a, EH19, Str15a]. 299 [mLcLJ16].

3 [FT16, GS19c, JNC^{+}11]. 3-PG [FT16, GS19c]. 315 [ABAG16]. 345 [ZLJM17b]. 361 [KS17a]. 377 [LSJ^{+}19]. 387 [HM19]. 393 [SMT19b]. 3PG [ABK^{+}16].

4 [BVGR^{+}17]. 4-D [BVGR^{+}17]. 408 [DYMYGO19a]. 4th [FKDG14].

50-year [NB11].

60 [Ano14u].

70th [GB16b].

9 [Kat11]. 978 [Kat11]. 978-1-59947-154-9 [Kat11].

= [BKP10]. =Bufo [Bar11].

A. [HH12]. A1 [Ano12-37, Ano14u, Ano15q, Ano15p]. A1-A2 [Ano15q, Ano15p]. A1-A4 [Ano14u]. A2 [Ano12-37, Ano15q, Ano15p]. A4 [Ano14u]. ABA [HSD^{+}15]. abalone [LJ15b, LJWN14]. abatement [BGPJ10]. Abies [KKP^{+}17, JL18]. abiotic [GSSM13, MHL^{+}19, PVB16, RDD^{+}18, SLB14]. able [DAF^{+}11]. ABM [KvTL^{+}18]. Abolition [LJ10]. Aborigines [GWR11]. above [SBC^{+}11, VHM^{+}17, VALC12, XGH^{+}17]. above- [SBC^{+}11]. above-ground [VHM^{+}17, VALC12, XGH^{+}17]. aboveground [CLWH15, SLLW15]. abrasion [LSL10]. Abrupt [PMTKC17, JDZC14]. absence [CMV^{+}16, IBH^{+}15, LPT^{+}18, MMPG16, PÓF^{+}18, SHS11, UG12]. absoluta [RSCS19]. absolute [NU11, PC15]. abundance [AGQP^{+}14, ATCJ^{+}19, AWSS14, BCTL16, BDG^{+}14, BV15, BAJ^{+}11, CTB^{+}12, DOB^{+}15, FTK^{+}14, dSGGBW12, HCP10, HLLP^{+}10, JSB^{+}14, KKH^{+}17, KCP^{+}15, MF12, PWGS^{+}17, PLL^{+}19, RLHD11, RH17, SW19, SLB^{+}18, TVI^{+}13, WJS^{+}18, ZZ17, ZYXT18, vPTOW18]. abundances [YAGMM^{+}14]. Acacia [KR14]. Academic [Jør14b]. Acanthocybium [MROGVR12]. Acari [WGT12, dOMZ13]. Acartia [CEC10]. Accelerated [ZW12, TB11, TB12a]. acceleration [DLL^{+}19]. accessibility [FN15]. accessible [BBJV^{+}11]. accession [LCBNP19]. according [CCSF11, FML^{+}17]. account [BPD^{+}19, vVBHZ11]. Accounting [ADF^{+}17, BLW^{+}17, CS17, IT17, LCMG18, ST13, TB12b, ZBSF17, ABB^{+}14, BFPR16, BPT^{+}19, BCT15, BHPF14, DLD^{+}11, EMH10, FBU14, FBDR17, FLA19, GRRU15, KØ13, LDJZC15, LSJ^{+}18, LSJ^{+}19, LLTQ14, LLTQ15, LG17, PBO19, PMH17, PME^{+}16, SWC13, Til14, Til21, VPB^{+}17, WO14, Sci11a]. accounts [LFH^{+}15, VLZ^{+}19]. accumulation [GPK^{+}13, HD14, HD16, LBdA^{+}19, NBD^{+}13, Vav10]. Accuracy [BTPG10, DHL^{+}12, Jör11a, KCP^{+}15, DPLC13, For18a, LLKY18, MdMS^{+}19, MLG^{+}18b, PC17, PRH17, RFZB12, Slo11, VHM^{+}17, vVBHZ11, vVHZHvD13]. accurate [Wol11]. accurately [MBHW12]. acetate [VRC17]. achieve [Bus17, LJR14]. achieved [PTML14]. achievement [Bin17]. Achieving [BA17]. Acid [ZF11, JLN10, NP15, SAAD16]. acidic [CL10]. acidification [FLH17, HKM^{+}18, HOŠ^{+}12, PPDF14, RCST11, THC18]. Acipenser [HMP^{+}19b]. acknowledging [PDBP11]. acoustic [CLD^{+}17, FBH^{+}14, FYR^{+}18, Md18]. acquisition [KP10, Lee14]. acreages [YDD^{+}13]. Acropora [NGRDTHRL19, RMvW13]. across [AHNK13, BKP10, Bar11, CBB19, CK12, DDM^{+}17, DPPI14, DS12, DFD11, EZ14, ENC^{+}15, Fai10, FMA^{+}12, For18b, GE17, HV15, HBA16b, KTV^{+}15, KDK^{+}10, LLQ^{+}15, LXS^{+}16, MML^{+}11, MSB^{+}17, MJLC14, OTA16, PMTKC17, RWFD13, SVD^{+}13, SBC^{+}11, SFB^{+}15, TLL^{+}15, WBC^{+}12, WLY^{+}15, ZZB^{+}18]. Acting [Gia14]. action [KK15a]. actions [SM11]. active [CLL14, DCMK11, EZ14, dMHGA10]. Activities [PNHL14, AYASR^{+}17, BHPF14, DYU^{+}12, GHS^{+}16a]. activity [CKS^{+}17, CSS^{+}17, FBA^{+}11, HFMM^{+}12, KCB^{+}17, LBE^{+}18, TTL^{+}18]. actor [GBL^{+}13]. actual [VLZ^{+}19, VLF^{+}17]. acute [FPO18]. ad [Pér18]. ad-hoc [Pér18]. adaptability [ACDC17]. adaptable [CF17]. Adaptation [CZR^{+}14, MNK^{+}21, Ana17, CDL^{+}11, DZS12, FS14b, HHK19, RVL^{+}18, SBRD10, UDP17, Wal16]. adaptations [EAJG16, ZW11]. adapted [GDW12]. Adapting [BFM11, CCL^{+}13, WhFxGZ16]. Adaptive [DPR10, OJM^{+}12, POB18, URJ10, WEB11, ANMJ12, BFM11, CRUF13, CH15, GVGC16, GLA11, HS13a, LPT^{+}18, Pra12, RFV^{+}12, RPM13, SAT^{+}15, WYHC15, Wil12, aGSCMF15]. adaptively [KK13]. add [AFC12]. added [RRBI14, XHC^{+}15]. addiction [RMGU11]. Adding [Ban11, Lin10, Lin16, RHS^{+}16, FHAG12]. addition [TIC^{+}14]. Additional [RDD^{+}18, TF13]. additive [MROGVR12]. additivity [WC19]. address [SGB10, VASH^{+}16]. Addressing [SG11b, KNU^{+}15]. Adélie [PLS^{+}16]. adequacy [AAFH15]. adhesion [CWLC13]. adjacencies [FT10]. adjacent [FS14a, MRG13b]. Adjunctive [MG16]. adjuncts [TV15]. Adjusting [CMSR19, JSW^{+}13, MH11, LJR14]. adjustment [vDHH^{+}13]. adjustments [MCR^{+}15]. administrative [TCGC16]. admixed [CHG^{+}15]. admixing [RGH^{+}13]. Adriatic [CQS15, RPP^{+}15, TPL^{+}15, VAK^{+}11]. adult [BS10c, LCD^{+}16, NWAR11, SF10, TGB^{+}17]. advanced [Nak17, NP18, ZDG^{+}13]. Advances [BH10, THBF^{+}17, SBC13]. Advantages [LLB^{+}19]. adverse [CJM^{+}12]. advice [CMCASS12]. advisories [WGD^{+}17]. Aedes [AOD^{+}13, EPA^{+}10a, EPA^{+}10b, MD16, SSC15, VSC18, ZYXT18]. aegypti [AOD^{+}13, MD16, SSC15, VSC18]. aeration [YWA^{+}17]. aerial [FDP^{+}10]. aerobic [BdCS11]. aeroecology [WGE^{+}19]. aeruginosa [MUP^{+}13, MvdWU^{+}16]. AEZ [TZS^{+}14]. affect [CHG^{+}15, DPvBvA11, DKJH13, GVV17, JDZC14, LHW^{+}15, LS13, MMRS^{+}19, RHW17, SPWR17, Str15a, Str15b, ZSSR18]. affected [CVPH^{+}14, CWMC^{+}12, NP15, NPH^{+}17, vSGR16]. Affecting [WYPF15, AFC11, HKlL14, MSCH14a, MSCH14b, WZL^{+}16]. affective [AA12]. affects [CTC^{+}18, DCCS17, GdSHA10, LLL19, MF12, MCA18, PBFS14, SK10]. affinis [DJMBD^{+}13]. Africa [KCF18, BCB^{+}11, DVL11, HBSF16, JMdW^{+}10, KCF17, OTA16, SS16, WCS^{+}14, WSW^{+}16]. African [BSFR17, BTS19, BMMC21, BBC^{+}19, KMRW16, RAF17, SJGB16, SWS16, WCS^{+}14, WSW^{+}16]. africana [MMS12]. after [BHC^{+}15, GDT^{+}14, Gat18, KKR19, LLL^{+}10, PDD^{+}18, RZF^{+}13, SS11, TVF^{+}10, VASH^{+}16]. against [For18b, LBR^{+}17, MLK17, RWP11, SM10, WYHC15]. age [AG15, BNC^{+}18, dÁGP14, CAL13, CD15, Dil10, FNR11, GHST19, GS19a, HKlL14, IKL14, JB11, JDZC14, KHW^{+}16, Lem16, MM19, MPK18, MH11, Oga17, Oga18, TKP14, TM18, VHDM17, WAGM17, ZBMBDU^{+}17]. age-related [Oga17, Oga18]. age-specific [Dil10, GHST19, TM18]. age-structured [BNC^{+}18, dÁGP14, GS19a, Lem16, MM19, TKP14, WAGM17]. Aged [Jen12]. Agent [ABS17, BFA11, CLA^{+}12, KvTL^{+}18, An12, AD16, AD18, ACSR^{+}13, BHD17, BMMC21, BTB17, BPR^{+}11, BRM^{+}14, BFNH14, BSCG10, CGLA10, CVPH^{+}14, CLBG15, CF17, CHvdW^{+}10, CLT^{+}11, CCG15, DVL11, FS18, GH13, GLC^{+}18, GPZB18, GLS10b, GJB^{+}11, HM18a, HM19, HZSH17, IN16, JX15, JHT^{+}14, KH15, KR14, LDM15b, LBGM19, MD16, MLG18a, MHS^{+}17, MSMM11, MSM^{+}17, MRES17, MJÁ^{+}12, PRH17, PNHL14, RMP^{+}18, Rin14, SGBD10, TB11, TB12a, TOQW11, THO10b, VLF^{+}17, WFW^{+}17, WNF^{+}15, YRGH16, ZZJ^{+}16, ZVS18, SMBGH^{+}19]. Agent-based [ABS17, BFA11, CLA^{+}12, An12, ACSR^{+}13, BHD17, BTB17, BFNH14, BSCG10, CGLA10, CVPH^{+}14, CLBG15, CF17, CLT^{+}11, CCG15, FS18, GH13, GPZB18, GLS10b, GJB^{+}11, HM18a, HM19, IN16, JHT^{+}14, KH15, LDM15b, LBGM19, MD16, MLG18a, MHS^{+}17, MSMM11, MSM^{+}17, MRES17, MJÁ^{+}12, PRH17, PNHL14, RMP^{+}18, Rin14, SGBD10, TB11, TB12a, THO10b, WFW^{+}17, WNF^{+}15, YRGH16, ZVS18, SMBGH^{+}19]. agents [HOKH18, KR14, PD12, RE10, WGTH16]. agglomeration [GWZ^{+}15, HZZ^{+}18, ZLZ17b]. AggModel [SDD^{+}13]. aggregate [KCG18, MJY17]. aggregated [ZJB19]. Aggregation [Log17, PJ17, CNI11, FSB13, Gar17, NYH15, PMRGF10, SW11]. aging [GAV^{+}11, KH15]. agri [BvHR^{+}15, EST^{+}19]. agri-environment [EST^{+}19]. agri-environmental [BvHR^{+}15]. Agricultural [Ano12a, LTL^{+}19, AYASR^{+}17, ADPS10, ABGB15, BPR^{+}11, BRM^{+}14, Che17, CB19, CTC^{+}18, DK17, DPBK10, DSM^{+}12, DNFG18, DSA^{+}18, EGMW12, FCH^{+}17, FB15, FYWA12, FDHL17, GZVU14, Hoa11, JAKJ10, KAC^{+}12a, KAC^{+}12b, LSGT13, LCPB^{+}10, MC11, MDOJ14, NNP19, OQH^{+}13, PHvB^{+}18, PTML14, PvVC15, RW10, SvBH^{+}14, SWZ^{+}12, TTCH^{+}14, UFK^{+}10, WAS^{+}16, WS15, WHNR19, YWH15, YDD^{+}13, YDZ^{+}21, vDSC^{+}10]. agriculture [CL15b, GE17, Hoa11, LCBNP19, LDL12, LWL^{+}17a, SJS^{+}13]. Agrilus [KKR19]. Agro [TZS^{+}14, Ana17, ABC^{+}17, CLBD^{+}11, GPM^{+}16a, HFMM^{+}12, LYJ^{+}15, MFESBG^{+}13, OTA16, QCLW18, SGB10, WLY^{+}15, HGN^{+}16, DAW^{+}10]. AGRO-BGC [HGN^{+}16, DAW^{+}10]. Agro-Ecological [TZS^{+}14, OTA16]. agro-ecosystem [LYJ^{+}15]. agro-ecosystems [Ana17, CLBD^{+}11, GPM^{+}16a, SGB10, WLY^{+}15]. agro-environmental [HFMM^{+}12]. agro-forests [ABC^{+}17]. agro-landscapes [MFESBG^{+}13]. agro-pastoral [QCLW18]. Agrodiversity [SGB10]. agroecosystem [BHD17, DAA^{+}12, DSK^{+}11, HGN^{+}16, MRTW18, RCC^{+}11, RCG^{+}13, RCGC17, SDT10, XSN^{+}18, ZSKI17]. AgroIBIS [ZSKI17]. AHC [XSN^{+}18]. AHP [GH13]. AHP-TOPSIS [GH13]. aided [AS11]. Ailuropoda [GFW^{+}17]. aimed [ABC^{+}17]. air [KBS^{+}15, KPTC14]. airborne [CLLB12, KHS^{+}19, MdMS^{+}19, STSA10, SM10]. al [HBT^{+}18, Och10, WSS^{+}16, GVV^{+}16, NP15]. Alabama [BPC^{+}17]. Alaska [FHL^{+}13, FAS^{+}19]. alba [dÁGP14]. albedo [LSR13]. albedos [KLS^{+}14]. Albemarle [JMB^{+}11]. Alberta [HMPC11]. albifrons [JJKJ11]. albopictus [EPA^{+}10a, EPA^{+}10b, ZYXT18]. Alburnus [MMVAHMC16]. albus [DHC16]. Alcon [CWJ^{+}13]. Aleppo [aGSCMF15]. Alexander [SGCB15]. Alfacs [GDDFT16]. alga [GPJL15]. Algae [SM13, BPN12, FCS13, GSJ^{+}10, IKK^{+}12, LHO^{+}16, yLWhMhG11, PBP^{+}15, Sin17, TVT12, YZG16, ZZZW15]. Algal [KLS17, BS18, CRO13, CBB19, CMLY19, DG11, DA19, GRB12, GDDFT16, JZG10, KSM^{+}18, LAB^{+}10, MWY^{+}18, PSB^{+}17, PLS^{+}10, SELB10, SQWS19, TGM^{+}21, WX11, XJY^{+}16, YXZ^{+}21, ZWRK13]. algebra [BMF^{+}11a, BMF^{+}11b, CT12, LLCR10, Pat12, Pat14b]. Algebraic [PBP^{+}15]. algorithm [BCL^{+}14, CRO13, CN11, DAF^{+}11, DVT^{+}10, HSLC^{+}19, HKVD16, QWJL19, RFV^{+}12, SZSV13, SAT^{+}15, SPR^{+}18, UG12, VZF^{+}11, YZPL16, ZAW^{+}18]. algorithmic [RKL^{+}15]. algorithms [CRB16, CGP12, GMC^{+}19, LWY^{+}11, MEC14, PSWW11, SG16, SMI^{+}19]. alien [CMM^{+}10, DPV^{+}16, FTYI13, KR14, PGZ^{+}19, SZV14]. Aligning [FN18]. Allee [BRF^{+}11, CHZ18, dSCdA18, Lab11, MF16, MP13, SBCB13, SSS^{+}11, VCE17, Ver10]. Allee-like [MP13]. Allegheny [FWRZ10]. allelopathy [GRB12]. Allen [HKW^{+}18]. alleviate [DFW^{+}18]. alligator [MSKS13]. allocate [Yan11]. allocation [ADF^{+}17, BMP^{+}19, CVV^{+}13, DSA^{+}18, GTC16, JBE^{+}15, LHP16, LOLA13, PDBP11, WHB12, dWcWbZ15, ZFZ^{+}14, ZZJ^{+}16]. allocations [YM21]. Allogamus [PGS16]. Allometric [KP10, PVL^{+}14, WK11, WUB15]. allometry [SW15]. along [CRA^{+}11, DBG14b, DBG14a, FGPR10, GLRT16, HKL^{+}11, LJCG15, MSG^{+}13, MBW14, MKF^{+}11, OBLJ13, TMC^{+}16, WWC^{+}18, WMK^{+}13, XWC^{+}11, ZLW^{+}15]. Alouatta [WFDDR10]. alpine [BVC^{+}18, CSS^{+}19, DVCM^{+}10, FMG^{+}16, FBDZ12, KJVL^{+}17, LLM^{+}13, MSP^{+}13, MSV12, SLZ^{+}17, ZPNM13]. Alps [FCFT16, MBW14]. Alta [LDM15a]. alter [DWWS15, FAMD^{+}12, MHL^{+}19]. alteration [KHJ^{+}19]. altering [PTML14, SM10]. Alternanthera [MSKS13]. alternate [HCS^{+}10, PAT17, dAdSC15, Log19a]. alternated [AdMC^{+}13]. alternative [AAFH15, BM11, BFJZ^{+}11, BGD^{+}15, DUASLB^{+}12, DLL18, DAP18, GRYM11, GVV17, HMP19a, KAJØ19, KKI^{+}14, Lem16, MDJ^{+}11, MJL^{+}13, MJL^{+}14, McC16, PLHG14, RL15, RvKM10, RCP17, YWA^{+}17, vLTdBP16]. alterniflora [CZCL18, TSP19, WJL^{+}13]. alters [NA11]. altimetry [MCMT^{+}16]. altitude [AKLB19, MSP^{+}13]. altitudinal [DVCM^{+}10]. altivelis [MSO10, FSN^{+}15, MSO10]. always [XM13]. AMAPstudio [GdC14]. Amazon [CGY11, dSD14]. Amazonas [IGR^{+}18, dSdCF10]. Amazonian [dSD18]. ambient [CGG^{+}18, KKK^{+}19]. ambiguity [SG16]. Amboseli [MNHW^{+}13]. amended [OTA16, SFvB10]. America [BEPC15, CW11, GPA^{+}13, HBA16b, LN11, LN16, SGM^{+}14, SBC^{+}11, TTL^{+}18, VASL^{+}15]. American [CR12b, BB18, CR11, DNFG18, HPW^{+}12, LAR^{+}16, LAD11, MRD^{+}12, MTJF11b, PGZ^{+}19, TCW11, Tho10a]. americanus [MSH^{+}19]. amino [JLN10]. Ammersee [WV12]. Ammonia [LNB17, MTL^{+}10]. Ammonia-uptake [LNB17]. ammonium [DWP13, MCH15, VRL14]. among [BCP16, DVF^{+}19, GGZ14, KKI^{+}14, LSH15, NYH15, PLB16, SDKE18, ZZA^{+}18]. amount [Tie16]. amphibian [BKP10, Cha12, DA17, ZZA^{+}18]. amphidromous [OOH^{+}12]. amphipod [GCC^{+}10]. Amur [LKG^{+}17]. anaerobic [VRL18]. ANAFORE [DCMK11]. analog [ZZT13]. analyse [FLB^{+}15, PHvB^{+}18]. analyses [FM11, HAOS17, HB14, JAKJ10, KDHC^{+}18, LM16, MCB14, OOFVBC16, PICR13, PPBD13, ROG^{+}16, SM14, VPM^{+}18]. Analysing [FT16, NMC^{+}14, DPV^{+}16, DGKSM13, JZ12, RPPW16, VSB13]. Analysis [AAW^{+}15, APWT15, BS18, BB18, BFS11, CHvdW^{+}10, DSK^{+}11, DZW^{+}11, HAA^{+}21, LLW^{+}15, LPV11, LLC11, MK13, Pat12, PSS16, RAR15, SNS^{+}12, TLL^{+}17, TL10, VPS10, XGH^{+}17, ZLF^{+}11, ZLF^{+}15, ZLZ17c, ZKSF17, AADA14, AYASR^{+}17, Ano12s, Ano12t, ACC11, BB10, BKM^{+}18, BS10a, BSMA18, BTPG10, BSZ^{+}11, BMU10, CCS13, Cam15, CC12, COBF13, CZR^{+}14, CMA^{+}13, Cho13, CAMCK^{+}12, CCC12, CMCASS12, CDT10, CCG15, CPS15, CBB^{+}10, DAA^{+}12, DLH^{+}14, DBK^{+}13, DVCM^{+}10, DSM^{+}12, DCZ^{+}15, DFM12b, DBJ13, FH11, FT10, FE14b, Fre16, GPM^{+}16a, GB16a, GZS^{+}18, GDT^{+}14, GLR^{+}19, GGSP14, GZVU14, GU11, GGVB11, GAF^{+}14, GB16c, GWZ^{+}15, GYL^{+}13, HGPF15, HCM^{+}15, HYG14, HZZ^{+}18, HCS^{+}10, Hei11, HLK^{+}12, HHvO^{+}17, HL19, HhLY^{+}19, HBA16a, HBL18, HBB15, IN16, JA15a, JLZ^{+}18, JSdSdL^{+}18, KB10, KM12a, KRE^{+}18, KYvdPS14, KUK11, LKMCW14, LST15]. analysis [LTL15, LCM11, LB17, LCH12, LCBNP19, LkSXhX17, LWGH13, LCL^{+}13, LVLS14, LHP16, LHL18, LWTZ18, LHB^{+}18, LHH12, LKAL^{+}13, LGY10, LYCZ13, LJLO11, MW12, MBBA17, MY11, MCW13, MWY^{+}18, MGR^{+}18, MHTC15, MXC13, MDJ^{+}11, MJL^{+}13, MJL^{+}14, MTJF11b, MDL^{+}10, MJC^{+}11, MSCH14a, MSCH14b, MATS14, mNkGyT^{+}17, NCMF10, NLM^{+}19, Oga11, OCCF17, OKS^{+}11, OCB^{+}13, PC16, PMH17, PSBH16, PMCKH13, PLB16, PLS11, QFG^{+}13, RHvdTS14, RBM^{+}17, RAF17, RHZ^{+}18, RPPW16, RMLTCQ^{+}19, RPA12, SELB10, SSV^{+}12, SZ14, SØB14, SLLW15, SBPZ12, SDSH15, SP10, SH18, TPL^{+}15, TRD^{+}15, Ten15b, Tho10a, TVF^{+}10, TB19, VMP^{+}19, WMF^{+}13, WGA^{+}13, WLY^{+}15, WSSM15, WAM13, WPB14, WFSH15, WAGM17, WSC^{+}19, WPZ^{+}14, XYTS19, XZCL16, YDD^{+}13, YZPL16, YWA^{+}17, YM21, YZG16, YLW^{+}11, YZLX14, ZRCNLI^{+}15, ZYFL10, ZH11, ZYLY11, ZZY^{+}13, ZLZS14, ZLF14, ZZF15, ZLFZ16, ZLZ17b, ZWF18]. analysis [ZBS14, ZTL^{+}17, ZSF^{+}12, ZSG^{+}18, ZMH19, dSD18, SSF14]. analyst [MRC19]. Analytic [CBGK15, YLC13, ZLZS14]. Analytical [BP10, GS19a, MH13, PL10b, TT12, BC11, Md18, MV17, RHR^{+}15b, RHS^{+}15, RHD^{+}16, RDV^{+}17]. analyze [BFKS16, EBL^{+}11, WLY^{+}15]. Analyzing [ABD^{+}18, GCW^{+}16, XZSL17, ZZ17, LDL12, MVC^{+}17, SAR11, vDHH^{+}14]. Anammox [MFD^{+}15]. Anas [KK12a]. Anaxyrus [BKP10, Bar11, MBG17]. anchovy [GPH17, JBCÑ11, KKH^{+}17, OAPP12, PRL^{+}13, PHP15, WJS^{+}18, XCR^{+}13]. ancient [CK12, LJAD17]. ancient/recent [LJAD17]. ancillary [HS11]. Andes [MRN12, SRV^{+}12]. anemotaxis [GKKW10]. angle [Md18]. Anglesey [JWLS17]. angling [LCD^{+}16]. angular [RAM10]. angustifolium [BHD17]. Anhui [YLW^{+}11]. animal [ALKB10, ADF13, BEZ19, BM11, BBHP13, BCJ^{+}12, EZ14, GMK^{+}19, HA18, HB13, MGD^{+}12, SJGB16, SHCR10, VR12, WR13a, XSL18]. animals [LB15, PC15, SFR^{+}17, WBG^{+}14, Wil11b, vdVJS16]. animated [DBP16]. anisotropic [HHW11, SBM^{+}11]. anisotropy [dLB19]. Ankole [MGW^{+}21].

Annelid [DLL^{+}19]. Annotated [BPT^{+}19]. Announcement [Ano11a, Ano11r, Ano11-37, Ano13a, Ano13b, Ano13u, Ano17b, Ano17c, Ano17d]. announcements [Fat15a]. annual [BMD18, CDG15, GLT10, HLS^{+}14, qHLlZ^{+}16, HGN^{+}16, LPS^{+}10, MK12, SW19, SMKM19, XLL15, ZTJ^{+}10, ZYXT18]. anoxia [SHM^{+}17]. anoxic [GS10]. Ansai [ZZLZ14]. answer [ACSR^{+}13]. Answering [DFH^{+}14]. ant [FHG12, YZL12]. antagonistic [KMKL19]. Antarctic [CR12a, GJK^{+}15, MRCW10, PLS^{+}16, RMTM^{+}16, VJR^{+}16]. Antarctica [HLLP^{+}10]. anthill [KOBK13]. ANTHRO [MCWG21]. ANTHRO-BGC [MCWG21]. Anthropogenic [YLW^{+}11, ARV^{+}16, GHP12, HSM^{+}12, MJC^{+}11, PVM12, PVM13, REB^{+}16, RDD^{+}18, RCS^{+}19, SRV^{+}12, SFL^{+}19]. anthropological [SGBD10]. anti [BEBA19]. anti-predator [BEBA19]. antibiotics [MWM^{+}16]. Antikythera [BC11]. antillarum [KK13]. antimony [SSW^{+}18]. Antipathes [JCKG15]. Antofagasta [OBC^{+}21]. anuran [VHDM17]. Anzali [SZSV12a, SZSV12b, SZSV13, SZV14]. APECOSM [DFM12a, DFM12b]. APECOSM-E [DFM12a, DFM12b]. aphid [BWH13, BGGL16, DFF17, GHS^{+}10, GOSH^{+}12, WPZ^{+}14]. aphids [FDP^{+}10, PZM13]. APHIDSim [PZM13]. Apicomplexa [LJB12]. Apis [AOW13, RBH13]. Apodemus [LSGT13]. app [MRC19]. Appalachia [DDBK15]. Appalachian [KMP14]. apparent [LMDG14]. appearance [GSJ^{+}10, RH17]. apple [BCP15, DHC14, TWL11]. Applicability [FRM^{+}19, LLKY18]. applicable [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q]. Application [AB13, BFP^{+}18, CCS13, CLCC11, CRQW18, DC10, DLCR13, DVT^{+}10, DJMBD^{+}13, Fai10, Flo18, GWHL15, GOSH^{+}12, HLS^{+}14, HFB^{+}11, Hoa11, JBB^{+}11, JMB^{+}11, KSSA13, KDMPC10, LBSB^{+}14, LLL^{+}10, NS15, PDAL14, PGK^{+}13, PFSF14, RHB14, RML^{+}13, SZSV12a, SZSV13, TAT^{+}16b, TLS^{+}16, VZF^{+}11, WZGH18, ZGJ10, ZCSL18, ZXD^{+}13, ZDG^{+}13, BAB^{+}16, BLB^{+}17, BKM^{+}18, BRS^{+}18, BGD^{+}15, CGLA10, CGV^{+}15, EEB^{+}17, EEK^{+}18, FDHL17, FSS^{+}17, GBB13, GPM16b, GV11, GS19b, HYG14, HPO^{+}15, HSVL12, Hu16, JGF14, JZG10, JHT^{+}14, KHW^{+}16, KDK^{+}10, KP13, LWKS10, LkSXhX17, LSZC10, LZL^{+}18, LJWN14, MGD^{+}12, MH10, MCPB15, OWMK10, PHC10, RSC12, RSC13, RSCS19, SELB10, STD15, SW15, SXZ^{+}15, TZS^{+}14, WB14, WHH^{+}18, WK17, WPZ^{+}14, XSN^{+}18, ZBMC14, BDG^{+}14, FE14a, GSJ^{+}10]. Applications [BH10, FKDG14, dSGGBW12, GCIB10, JBH^{+}14, Mon13, ZW10, AD12, BB10, BLW^{+}17, BCT15, CVG^{+}15, CGP12, DP15, GB16c, LJBS12, LWL^{+}17a, MRWN19, SBC13, SZVV10]. applied [BDL^{+}14, BKN14, BPM^{+}17, BMU10, CGAB17, DWNN10, FHL^{+}13, FCA18, FSM19b, GGV13, MSHC13, MCB14, NPJ^{+}17, PES19, Rau11, RMHE15]. Applying [CLL^{+}15, FBH^{+}14, HMP^{+}19b, WZD^{+}14, YJW15, ZLLZ13, BFR17, FBdP^{+}16, LLM^{+}13, ZLJM17b, ZLJM17a]. approach [ALG16, AWF10, AWL^{+}13, ABR12a, ACDC17, Ano12s, AAFH15, ABC^{+}17, ACSR^{+}13, AVG14, BRR10, BBRJ15, BMN11, BC13, BG15, BLD10b, BBC^{+}19, BCL^{+}14, IHSM12, BRS^{+}18, BvHR^{+}15, BK13, BAT^{+}15, BAJ^{+}11, BHPF14, BPR^{+}15, CF14, CLBG15, CHR^{+}13, CVS^{+}13, CS18, CLHH19, CSF11, CVG^{+}15, CRTS15, Con14, CTP^{+}17, CSFB17, DZS12, DBSA16, DAF^{+}11, DSM^{+}12, DMH17, DWP16, DSBS15, EFM^{+}10, EKB18, F.11, FKH^{+}15, FCG^{+}19, FT10, FCFT16, Fre16, FAT^{+}15b, FHC15, GSQ17, GPH17, GBG^{+}15, GGZ^{+}16, GKF^{+}14, GCTS15, GHS^{+}16b, GDDFT16, GEB10, HCK18, HHC^{+}13, HBS^{+}16a, HJ17, HMS^{+}11, HS13a, HSHF^{+}13, HMA10a, HPG^{+}10, HEJP17, IBD^{+}19, JR16, JM15, JHHS17, KHW^{+}16, KGM12, KR10b, KPLS15, KB18, KJH^{+}13, Kos17, LMT13, LHL^{+}17, LZS^{+}17, LLP17, LAK^{+}18, LBC^{+}15, LPD14, LC13, LCML11, LRB12, MDW^{+}12, MY11]. approach [MCW13, MGR^{+}18, MHS^{+}17, MSKS13, MAFG^{+}19, MKSL^{+}17, MSM^{+}17, MRZU14, Mer12, MOMC17, MRM^{+}17, MRR^{+}14, MHC16, MN12, MFI13, MRTW18, Ned15, NJCF11, OFK^{+}12, ODK16, PC17, PLFO^{+}12, PAB13a, PVL^{+}14, PRL^{+}13, Pra12, PDG^{+}14, RKL^{+}15, RW10, RPA^{+}18, RIMPBT19, RSCS19, RBRK16, SBA^{+}19, SGBD10, SvdMKS11, SFvdM^{+}17, SJGB16, SBR^{+}19, SHX^{+}12, SKH^{+}13, SA13, SQWS19, SVL12, SDEP18, SNZ^{+}19, SB15, TAC^{+}14, THBF^{+}17, THOA10, VR12, VSB13, VBS14, VPC11, VASH^{+}16, VQS^{+}18, WSW^{+}13, WFW^{+}17, WJS^{+}18, WCS^{+}14, WSS^{+}16, WKF^{+}17, WHH^{+}18, Yan11, YYQ11, YJW15, YZPL16, YZLX14, ZH11, ZFZ^{+}14, ZLG16, ZF11, ZVS18, vVHZHvD13, Wei19]. Approaches [BH10, CXL^{+}19, CPS^{+}19, DAP18, HKL^{+}11, HKVR13, HKG^{+}17, JGF14, JDP^{+}12, KR10b, KCF17, KCF18, KBL10, LCD^{+}16, LP17, LB18b, MBD^{+}12, PM16, Wil11a]. appropriate [GSQ17, KVU^{+}10, LCCL15, MU10, San11]. appropriation [CLWH15]. Approximate [LJLC15, vdVBJS15, CDK^{+}17, RFV^{+}12]. Approximating [LM13, SAT^{+}15]. approximation [NSIiT17, TT12]. approximations [MRES17]. April [Ano10g, Ano11v, Ano11w, Ano12-28, Ano12-29, Ano13-39, Ano13-38, Ano14e, Ano14c, Ano15a, Ano15j, Ano16t, Ano17-38, Ano17r, Ano18-41, Ano18-38, Ano19-36, Ano19-47]. APSIM [ZBS14]. APSIM-wheat [ZBS14]. Aquaculture [HD14, AJDS16, BLCPB11, GL10, HCZ^{+}17, IFC14, KTM^{+}16, MHRB19, Mur13, RSDP^{+}12, RPC^{+}17, TPT^{+}11]. Aquatic [AdMC^{+}13, ALC19, AS14, ČTT^{+}12, CCC12, EAB11, FSM19b, HB12, HB10a, JDGJ13, LJWN14, MY11, MNK^{+}21, MU10, PHP^{+}14, Pea15, RPSA12, RE10, SZV14, SGAS14, SBR^{+}19, STD15, SGHW^{+}17, TM12, TŠAD12, TM14, VMSM10, Wil11b]. AQUATOX [BPC^{+}17, CBP^{+}17, NGWZ16, YLY^{+}18, ZLLZ13, ZCSL18]. Aquifer [MGWW13]. Ara [HBT^{+}18, MFB17]. Arabian [SPM^{+}18]. arbuscular [VH12]. archaea [LNB17]. archaeological [Mur12]. archipelagic [CXW19]. archipelago [BSO13, GPC14]. ArchiSimple [PBB^{+}14]. architectural [GGR^{+}13]. architecture [BBL11b, COBF13, EAJG16, GdC14, PBB^{+}14, TBL^{+}15]. architectures [dLB19]. Architeuthis [CMEP15]. Arctic [CR12a, DMB^{+}19, JJMW14, LSR16, WR18]. Area [SRXP14, BLH10, BBJV^{+}11, BVVS15, CZC^{+}15, CGEdM12, DKJH13, FS14a, FGH^{+}10, FJH15, GWHL15, Gav15, GZK^{+}11, GBB13, GDBB19, GdSHA10, HA18, IBD^{+}19, KKP^{+}17, MWK19, MSF18, MS16, NMA^{+}19, OOFVBC16, RML^{+}15, ROSC15, SFL^{+}19, VVA^{+}15, WGD^{+}17, ZLZS14]. area-based [VVA^{+}15]. area-restricted [BLH10]. area-wide [GBB13]. areas [ASP^{+}15, BS13, BFPR16, BC13, BRL12, BJJ12, BAT^{+}15, CL15b, CLHH19, CLCC11, DLH^{+}14, FHG12, FYWA12, FBDR17, FPG11, FPWG14, GWR11, HCE^{+}19, HSM^{+}12, IGBSASSJ16, LSZC10, LDYL18, MD16, NKSA13, OAPP12, OCD^{+}13, PMBM11, PPB^{+}18, Pra12, SWZ^{+}12, VR12, VPB^{+}17]. Arenicola [DLL^{+}19]. Argentina [GLT10, CCF13a]. Argentine [BPR^{+}11, FHG12]. Argentinean [BRM^{+}14, DSS11]. Argentinian [FB15]. arid [AHF^{+}13, BC13, Ber13, Che17, CV19, FS18, GLRT16, GWF^{+}18, HYL^{+}17, KF16, LZZ^{+}14, LLTQ14, LLTQ15, LHP16, LWL^{+}17b, OES16, PLL^{+}19, PLHG14, SJGB16, UR10, WYS12, ZLLC13, aGSCMF15, vLTdBP16]. aridity [PLL^{+}19]. arising [Kal19, SWR^{+}16]. Arizona [BHD16]. armadillos [OLC^{+}17]. Armenia [ABS17]. armigera [BNC^{+}18, FGH^{+}10, PVB16]. aromatic [XWW^{+}11, XQZ^{+}13, ZCSL18]. arrangement [SXNL18]. arrested [SU15]. art [CCF13b]. artedi [MYJ^{+}14]. artefact [Pau14]. Artemisia [SLB14]. Artificial [GDDFT16, BAT^{+}15, CRBC11, GO15a, HDL^{+}16, HWH^{+}18, JBCÑ11, KHJ^{+}19, LPX^{+}12, MFS18, NNP19, PSWW11, RUSL19, SSW^{+}18, TLZ17, WLR^{+}11, WZD^{+}14, ZZ17]. artificially [Yan11]. artisanal [ROSC15]. ash [KKR19, VCSSC12]. Asia [ACK^{+}14, HLZ^{+}18, RNE^{+}12, SSY14]. Asian [GJM^{+}13, LSJG17, LLC^{+}21, ZLLC13, ZSG^{+}18]. Asiatic [DFK^{+}11]. aside [LJ10]. aspects [DH18, Kal19, VMP^{+}19]. aspen [ARM^{+}12, Gar17, MRR^{+}14]. asper [CMGF14]. assemblage [BGC10, HCP10, MYHN17]. assemblages [BRMW18, CRB16, HJWG12, Pen11, PGKG12, TS11, WRMB18]. assembling [LCTH12]. assembly [AWL^{+}13, BND19, GDTP13, XZZ^{+}13]. Assesment [GJÁ^{+}10]. assess [CTB^{+}12, CVPR^{+}16, CPD^{+}11, CGEdM12, GDT^{+}14, GJB^{+}11, HGBA^{+}13, JTB^{+}19, KBS^{+}15, KEM^{+}16, LDS14, MSB^{+}17, Mur11, SGM^{+}14, TAGC^{+}16, vVHZHvD13]. assessed [RWS^{+}16, VLF^{+}17]. Assessing [AAFH15, BFJ^{+}11, BRW^{+}12, BPM^{+}15, BPU11, BU16a, BHPF14, DPBS10, FLX^{+}14, FLH17, Fie12, FMLM11, Gre12, HRSKW15, HHC^{+}13, HBA16b, KPT11, LCP14, LVV^{+}15, LGDW12, MJY17, MBM^{+}18, NL12, NMS^{+}14, NPP^{+}18, PBO19, PDD^{+}18, PSKO16, QFG^{+}13, RHW17, RMR^{+}16, RV15, SMT19a, SMT19b, SDEP18, TLS^{+}16, VPB^{+}17, XCL^{+}11, YSB11, BC13, CLH17, CSR15, CLNN14, GSA^{+}19, HKG^{+}17, JCKG15, JBW^{+}17, LCBNP19, MSCM18, PKB^{+}18, Pec12, RSD^{+}15, SGDP19, TMT^{+}18, VG19, XYY^{+}16, dJGM^{+}18]. Assessment [AAO10, CBGK15, GSJ^{+}10, GMLRF10, HT14, JKS^{+}10, LLKY18, MSHC13, PSWM^{+}10, RDD^{+}17, SFB^{+}15, TSN17, TLL^{+}17, UAB^{+}11, WLZC11, WB16, YLC13, ZH11, AALP18, BCC^{+}11, BPC^{+}17, BCD12, BU16b, CHCC15, CZW10, CCC11, CCF13b, CLWH15, CLCC11, CAY^{+}10, CBP^{+}17, CRTS15, CKH15, DFL^{+}16, DLD^{+}11, EBL^{+}11, FB15, FtHvdB^{+}14, Fre16, FAC^{+}19b, GHWG^{+}14, GRRU15, GZK^{+}11, GL10, GFP^{+}16, GT14, GFC15, GB15, HA18, Hal11, HLW15, HFU11, HF14, HTL^{+}14, HHL17, IPSH14, IS13, JZCL14, JXL^{+}15, JDP^{+}12, JL18, KSSA13, LDJZC15, LDG14, LSP^{+}10, LB15, MLY^{+}19, MY11, MSS^{+}19, MR17, MPF^{+}14, MDKK15, dMHGA10, MAMNL^{+}15, NGGFM19, OCB^{+}13, OBC^{+}21, PMS^{+}17, PPB^{+}18, PMH17, PAB^{+}13b, PBD^{+}17, QCB^{+}11, RHS^{+}15, RHD^{+}16, SBM^{+}12, SSV^{+}12, SMI^{+}19, SWZ^{+}13, SZVV10, TNA^{+}17, TA11, UFK^{+}10, VHM^{+}17, VMSM10, XYCZ11, YSH^{+}10, YZPL16, ZGJ10, ZLLZ13]. assessment [ZLY^{+}13, ZLJM17b, ZLZ^{+}17a, ZLJM17a, ZCSL18, ZWG^{+}17, dSD18, vDSC^{+}10, vVBHZ11, vGKB10]. assessments [HKL^{+}11, HCM^{+}18, JMB^{+}11, KB18, PRH17]. assets [BLOJ^{+}19]. Assigning [PSWM^{+}10]. Assimilating [LZZ^{+}14, ZCS13]. assimilation [CC14, HKlL14, MF12, VRL14]. assimilative [XYY^{+}16]. assist [LBR^{+}17]. assisted [CMM^{+}10, DSM^{+}12, SBM^{+}11]. assisting [MLAT^{+}18]. associated [DBVS19, GPR^{+}13, HS11, KTK^{+}15, KDT^{+}17, LJ10, LYL^{+}17, Mat13, PLHG14, PANL11, RGH^{+}13]. association [KKP^{+}11, San11]. associations [BTPG10, HM18a, HM19, LBE^{+}13, MDH15, WGSR19]. assumption [BDG^{+}14, WSE12, WLR18]. assumptions [BG11, PC16, TF13]. Astroni [IBD^{+}19]. asymmetric [RW17, TYEH10]. asymmetry [BA17]. asymptotical [WAM13]. asynchrony [CH14]. at-risk [BSW18]. Athabasca [MMRS^{+}19]. Athene [EGMW12]. Atlantic [CSFB17, GQNMR18, PMF^{+}13, WHRR15, AGQP^{+}14, BRS^{+}18, BSH^{+}10, CRA^{+}11, DMR13, DWP11, FAPM12, FCA18, GRL18, HSHF^{+}13, HCM^{+}18, JNHF18, LCD^{+}16, LJ13, MSHC13, MSM^{+}16, PAGCV^{+}10, PP12, PGA^{+}11]. Atlantis [IT17, OCCF17, PWP^{+}19, WLF^{+}16]. atmosphere [LPD14, MKF^{+}11, WSZT10]. atmospheric [ACN14, GBG^{+}15, Hol13, LY16, TRE^{+}12]. attached [FCS13]. attacks [DHHB15]. attenuation [WBS13]. attitudes [BRW^{+}12, BWP16]. attraction [GJB^{+}11, RUSL19, SW11]. attribute [CL10]. attribute-regulated [CL10]. attributes [BPGSS^{+}18, HMA10b, NBD^{+}13, PBO19]. August [Ano10o, Ano10p, Ano11-30, Ano12-30, Ano13-32, Ano14r, Ano14o, Ano16a, Ano16u, Ano17-35, Ano17-36, Ano18-46, Ano18-47, Ano19-33, Ano19-37, Ano21b, BMF^{+}11a]. aurata [PAGCV^{+}10]. Aurelia [MIC14, Ned15]. auricollis [PGS16]. aurita [MIC14]. Australia [CRO13, KSZW16, DLW^{+}17, DLvGGNC16, For11, GPR^{+}13, HPO^{+}15, LS10, LWNG13, LMBL12, NJM11, NB11, NBD^{+}13, PKB^{+}18, SMC11, TL15, WRSL11, WEJ11]. Australian [GPR^{+}13, ADPS10, BTBM17, GWR11, LDG14, TBA19]. australis [WJL^{+}13]. Austrian [HRSKW15]. auto [GM10]. auto-facilitation [GM10]. autocatalytic [GHK17]. autochthonous [KH11]. autocorrelation [DF13, HMD^{+}16, HFMM^{+}12, MSV^{+}11]. automata [Cra10, DAF^{+}11, HHL17, JBE^{+}15, MJÁ^{+}12, PD12, YZL12]. Automated [WKF^{+}17, ČTT^{+}12]. automatic [GWC^{+}10, Tat16a]. automaton [GLI^{+}21, ZJL^{+}16, ZHLZ17]. autonomous [SJTS13]. autotrophic [GBR13b]. autumn [DDBK15, qHLlZ^{+}16]. availability [BQF^{+}18, FS18, mGZK^{+}10, GPCG15, HMP^{+}19b, MOMC17, NWAR12, PVL^{+}14, ROG^{+}16, SFHK12]. available [HRWT12, Pie10a]. Aveiro [BPGSS^{+}18, DVD^{+}14]. avellana [BOF^{+}16]. Avena [BCL^{+}14]. avenae [DFF17]. averaging [BSW18, Log19b]. averted [Bus17]. averted-loss [Bus17]. avian [CMSPJ11, DWWS15, Rin14, WHNR19, ZWA19]. aviculare [MBBA17]. avoidance [GJB^{+}11]. avoiding [SGSR10]. away [BL15]. Ayu [KUI^{+}18, FSN^{+}15]. Azolla [SZSV12a, SZSV12b, SZSV13, SZV14]. Azorean [PES19]. Azores [MAC^{+}12]. azoricus [HSvO^{+}18].

B [Wei19]. babassu [SFL^{+}19]. back [CRQW18, Jør14a, JGX13]. back-calculation [CRQW18]. background [JTM^{+}17]. backscatter [BW12]. bacteria [KPSC10, LNB17, MFD^{+}15, MJÁ^{+}12, WB17]. bacterial [Mur13]. Bactrocera [ABR12a]. badger [AGQP^{+}14]. bag [RCS^{+}19]. Bahia [MZLCSZdP11]. Bai [XWW^{+}11]. Bai-Yang-Dian [XWW^{+}11]. baiting [MS14]. Baiyangdian [XQZ^{+}13, MY11, XYCZ11, XYCZ13, YYQ11, YM21, ZLLZ13]. Baja [CMCASS12]. balance [BLCPB11, CW17, CZ11, DCZ^{+}15, FE14a, GCB^{+}11, HYL^{+}17, JOHW17, JAKJ10, KZR^{+}13, LJR14, LY11, Lin10, Lin16, MDW^{+}12, MDL^{+}10, MSO10, MEC14, NOOM^{+}16, Pie10b, Pie10a, SGAS14, URUNAM12, WZG15, XCG^{+}16, ZKR^{+}13]. balanced [CRUF13, JHH^{+}12, OBC^{+}21, PLR^{+}13, SS11]. balances [BC13]. Balancing [YY11]. baleen [RMTM^{+}16]. Balgzand [SFvdM^{+}17]. Balkan [KTV^{+}15]. Balogh [Wei19]. Baltic [BH13b, MMM^{+}16, BH13a, BSO13, Hei11, HSM^{+}12, IKL13, MML^{+}11, PWK^{+}17, TNHB12, WBS13, WB14]. Baltimore [SHS^{+}12]. bamboo [GRL18, RBRK16]. bamboo-dominated [GRL18]. ban [ZKSF17]. banana [LBR^{+}17, Och10]. Band [KHKC^{+}18, PSKO16]. banded [OLC^{+}17]. Bandwidth [KUK11]. Banff [MAM^{+}10]. Bangladesh [BAA^{+}14, BKM^{+}18, HSCS14]. Bank [BWRA16, GDC12, HLS^{+}14, vLTdBP16]. banking [RHB14]. banksiana [SGG14]. Barataria [LdMS^{+}16]. bare [MLG^{+}18b]. bare-earth [MLG^{+}18b]. Barents [BSH17]. bark [FH11, HOKH18, KSI14, LTGD16, OJ10, VCSSC12]. Barkedji [BVVS15]. barley [JG10, LSS^{+}15]. barn [dÁGP14]. barnacle [SSD^{+}17]. barramundi [HD16]. Barrier [CDK^{+}17, HBR19, TLW^{+}14, DWP11, GKF^{+}14]. barriers [MMCBP19]. base [WXB10]. Based [BMCG12, KvTL^{+}18, LXS^{+}16, AAJB21, ABS17, AD12, AMM^{+}14, An12, AAW^{+}15, AD16, AD18, Ano12s, Ano12t, ARTBB14, ABC^{+}17, ACSR^{+}13, Ash10, ADF13, BFA11, BVP^{+}13, BAAB18, BHD17, BM11, BTS19, BP15, BPB^{+}18, BG15, BTB17, BJJ12, BPR^{+}11, BRM^{+}14, BLB^{+}17, BFNH14, BBGJ13, BBL11a, IHSM12, BSCG10, BDG^{+}14, BQF^{+}18, BPM^{+}15, BRS^{+}18, CGLA10, CRBC11, CC12, CVPH^{+}14, CLBG15, CVS^{+}13, CLLB12, CdMMF16, CF17, CC10, CWL^{+}16, CL10, CCC11, CLA^{+}12, CDK^{+}17, CZCL18, CS18, CLT^{+}11, CDL^{+}11, CPS^{+}14, CTC^{+}18, CLBSG11, CRTS15, CCG15, CBA10, CT12, CPJ^{+}13, CKH15, CDG15, CSL^{+}12, DHC14, DJT^{+}19, DLD^{+}11, DVT^{+}10, DFD11, DJMBD^{+}13, DDR^{+}14, EKB18, EST^{+}19, FCH^{+}17, FN15, FB15, FBA^{+}11, FB13, FCV^{+}12, FS18, GH13, GLC^{+}18, GEEC14, GGM^{+}12, GDTP13, GO15a, GFW^{+}17, GCTS15, GPZB18, GGVB11, GOSH^{+}12, GLS10b, GJB^{+}11]. based [GHS^{+}16b, GLS^{+}17, GS19c, GCvDM12, GCM^{+}13, HMP19a, HCW^{+}17, HM18a, HM19, HLW15, HZSH17, HWS11, HS13a, HTL^{+}14, HSHF^{+}13, HHC15, HGAT17, HCL^{+}16, HHK19, HR11, HVCP16, HRWT12, HDL^{+}16, HhLY^{+}19, HYL^{+}17, HSD^{+}15, HBB15, IFC14, IN16, JD18, JKS^{+}10, JTB^{+}19, JC11a, JOHW17, JHT^{+}14, KFWM19, KMRW16, KF16, KCG18, KCF17, KCF18, KTM^{+}16, KB18, KH15, KGD14, KWJH16, KKI^{+}14, KBMM15, LJLC15, mLcLJ15, mLcLJ16, LDG14, LLB^{+}19, LPF11, LJ15a, LSA^{+}19, LTL^{+}13, LLT^{+}14, LJ14, LYSW15, LJ15c, LHL18, LPS^{+}10, LDM15b, LKP13, LSGT13, LYCZ13, LYCU14, LBS^{+}14, LYC^{+}15a, LKAY15, LAK^{+}18, LAEAG12, LBGM19, LGDW12, LSL10, LRB12, MK13, MHL^{+}19, MLY^{+}19, MG16, MD16, MY11, MCW13, MWY^{+}18, MLG18a, MWK19, MHS^{+}17, MSH^{+}19, MBRR13, MSMM11, MSM^{+}17, MGW^{+}13, MvdWU^{+}16, MAP^{+}13, MPF^{+}14, MAB19, MRES17, MBM^{+}18, MRR^{+}14, MGSB11, MMGK15]. based [MJÁ^{+}12, MATS14, MSM^{+}16, MMMGLN^{+}18, NL12, NMT^{+}19, NCEGB16, NPJ^{+}17, NNP19, NMA^{+}19, NMdR16, OGG^{+}16, OHS^{+}17, PC17, PCTW17, PHP^{+}14, Pér18, PLB16, PRL^{+}13, PWP^{+}19, PGZ^{+}19, PLL^{+}19, PRH17, PP12, PNHL14, PZM13, PLS11, PGG14, PSSH11, QWJL19, RVL^{+}18, RHvdTS14, RMP^{+}18, RLHD11, RHZ^{+}18, Rin14, RDJ19, RDD^{+}17, RR10, SBCB13, SGBD10, SJGB16, SM11, SMT19a, SMT19b, SZ14, SRSS12, SWC13, SLY^{+}17, SVL12, SSK14, SZVV10, SLH^{+}11, Sil11, Slo11, SBPZ12, SBA^{+}13, SGSR10, SGWV15, SvdR10, SS11, SH18, SYY^{+}19, SB15, SLD^{+}19, TB13, TB11, TB12a, TLL^{+}17, TMPF18, TAT^{+}16b, TSM^{+}15, TVF^{+}10, THO10b, TTCC^{+}19, Uch16, UG12, VLZ^{+}19, VSF10, VR12, VBS14, Vav12, VRL14, VVA^{+}15, VGR^{+}11, Wal17, WLZC11, WMW^{+}12, WLDL12, WC13a, WC13b, WGA^{+}13, dWcWbZ15, WSL17, WFW^{+}17, WLC^{+}19, WJS^{+}18, WR13a, WNF^{+}15]. based [WR17, WAM13, WV12, WSS^{+}16, WP14, WHC^{+}13, XLR11, XZC^{+}17, XCL^{+}11, XYCZ11, XPY^{+}14, XGH^{+}17, YqWB^{+}15, YL17, YZL12, YLC13, YXZ^{+}21, YSA^{+}14, YRGH16, YXD^{+}18, ZSS17, ZMS19, ZHC^{+}16, ZLZS14, ZQC15, ZLFZ16, ZLJM17b, ZLZ^{+}17a, ZDL^{+}17, ZLJM17a, ZWW^{+}18, ZZT13, ZVS18, ZWJ^{+}17, ZPNM13, dÁSMRT^{+}19, dJHMS19, dOMZ13, dOAC18, vPTOW18, vdVBJS15, vdVJS16, SMBGH^{+}19, GCCB18, mNkGyT^{+}17, PD12, PMS11, SWZ^{+}12, WYS12, WFN^{+}16]. based-model [BRM^{+}14]. baseline [BPC^{+}17, BU10, BU16a, BCUF16, BCDU16, Cam16, DBSA16, RVB^{+}10]. BASGRA [HHvO^{+}17, HVCP16]. Basic [WGSR19, SM12]. Basin [FCH^{+}17, FBB^{+}18, GLS^{+}17, HG17, LY16, LC11, RZF^{+}13, SXZ^{+}15, WZG15, YYL13, YLC13, YM21, YDL^{+}15, ZLZ^{+}17a, dFMdSA18, AD12, Con11, Con12, GLC^{+}18, GYL^{+}13, HSM^{+}12, MMRS^{+}19, MSO10, OWMK10, SWZ^{+}12, SGRAdJ17, YXD^{+}18, ZWRK13, CL15b, HZSH17, MY11, PWK^{+}17, WFSH15]. basin-level [OWMK10]. basins [JMB^{+}11, JBW^{+}17, MLH^{+}17]. basis [KAF^{+}14, LV17, TNHB12]. bass [LBW18, MLS^{+}11, WFN^{+}16]. bat [CESC19, HCP10, HBSF16]. bathymetric [NMA^{+}19]. Bathymodiolus [HSvO^{+}18]. Batrachochytrium [ACCM16]. bats [RTE^{+}15]. battle [Rin18]. Bay [BN10, BLCPB11, GSQ17, GL10, KR10a, KSZW16, LdMS^{+}16, MLS^{+}11, NJCF11, WEJ11, XZC^{+}17, BSO13, DCZ^{+}15, GSJ^{+}10, CN10a, CMLY19, DC10, FMLM11, GPH17, GDDFT16, HPC13a, HPC13b, JX17, KH11, KZR^{+}13, KPLS15, LBSB^{+}14, LZW^{+}17, MTZ^{+}10, PHP15, PPDF14, SSN^{+}18, TSN15, TSN17, URUNAM12, ZKR^{+}13, ZW10, ZSF^{+}12]. Bayes [SFL^{+}19]. Bayesian [KCF18, WLS^{+}11, ALG16, AHS^{+}17, ABL^{+}13, BSC^{+}15a, BVC^{+}18, BSW18, BLFS18, BFR17, BDG^{+}14, BAJ^{+}11, BLAV18, CSOS17, CDK^{+}17, CPS^{+}19, CRQW18, FDP^{+}10, FHL^{+}13, FLB^{+}15, FCV^{+}12, GKF^{+}14, HPJ15, HSK17, HMS^{+}11, HlLB12, HHvO^{+}17, HEJP17, JMdW^{+}10, KSSA13, KCF17, LJLC15, LLB^{+}19, LBE^{+}13, LYJ^{+}15, LQW^{+}19, LLKY18, MMD^{+}17, Mar12, Mar17, MR17, MDH15, MJLC14, MHC16, MFI13, NF11, PLR^{+}13, PES19, PBD19, RLB^{+}12, RFV^{+}12, RWS^{+}16, RAR15, RRvdG18, RKNO17, SML12, SRM11, SP12, TOHT14, VRM^{+}10, VPB^{+}12, WC13a, WC13b, WC13c, WLS^{+}10, YZPL16, ZLG16, vdVBJS15]. Bayesian-network [RRvdG18]. bays [OBC^{+}21]. BC [MSH^{+}19]. be [BU11a, Che16, CHG^{+}15, dSCdA19, FMBD15, FE14a, GVGC16, GAB12, HSLC^{+}19, Har11, Kos17, MBLC14, PTML14, PLS^{+}16, RA13, RPGSCF16, TCGC16, vdVJS16]. beach [LDO^{+}18]. beaches [LBD10]. BEAMS [SNS^{+}12]. bear [DSBM^{+}18, DNFG18, DFK^{+}11, JSW^{+}13, LBWW14, LPSC16, MSH^{+}19, OJM^{+}12]. bearing [DJMBD^{+}13]. bears [AMM^{+}14, LARZN13, MHS^{+}17]. beavers [PGZ^{+}19]. become [BGGL16]. bed [MBUM12]. Bee [DDR^{+}14, GAG^{+}12, AOW12, AOW13, BBB^{+}17, BGK^{+}16, KEHG17, KDHC^{+}18, MHTC15, OBSL15, RBH13, SGM^{+}14, TWL11]. bee-pollinated [TWL11]. beech [BLB^{+}17, CVV^{+}13, EKB18, TGB^{+}17]. BEEHAVE [BGK^{+}16]. Beer [dLB19]. bees [GLSA^{+}11, GPA^{+}13, KEHG17, SGM^{+}14]. BEESCOUT [BGK^{+}16]. beet [BDL^{+}13, BDL^{+}14, CDT10]. Beetle [SRT16, BA14b, CHK^{+}12, FH11, GEB10, HOKH18, KKR19, KSI14, LTGD16, PB14]. beetles [OJ10]. before [KKR19, SA16, SS11]. beginners [MTA19]. begins [Nie16]. Begomovirus [JSKK15]. behave [FL11]. behavior [AAA13, APWT15, BCP15, BAAB18, CWM12, DFL^{+}16, DWWS15, DDR^{+}14, GGTP^{+}17, HBB15, MHTC15, MAM^{+}10, NG11, PMM11, PF12, Rin14, SAT^{+}15, SHCR10, Sim11, SK15, TSN15, URJ10, WSZT10, vDHH^{+}14]. Behavioral [SGSR10, BSP12, BCJ^{+}12, DDR^{+}14, GLS10b, LLC11, PNHL14, SMH^{+}12, ZHV14]. Behaviour [TV15, BGK^{+}16, BLD^{+}10a, DMBS11, EIG^{+}10, HMP19a, HVL^{+}19, LDS14, LCD^{+}16, LS10, MvDL^{+}10, MRC19, PC17, POB18, VRM^{+}10, WMCM12]. Behaviour-manipulating [TV15]. behavioural [CHR^{+}13, MSM^{+}17]. behavioural-strategy [MSM^{+}17]. behaviourally [CRUF13]. behaviours [MBM^{+}18, RHO11]. Behind [RJG^{+}13]. Beijing [CCC11, DLD^{+}11, FCZ13, HCM^{+}15, HLW15, HZZ^{+}18, HFB^{+}11, HM18b, JMW11, LLW^{+}15, XHC^{+}15, ZYLY11, ZLC13, ZZY^{+}13, ZLF14, ZLF^{+}15, ZLFZ16, ZLZ17b, ZMH19]. being [SSS^{+}11]. Bejan [KOBK13]. Belgium [BVP^{+}13, EBL^{+}11, GVV^{+}16]. Belief [KCF17, KCF18, LLKY18, FLB^{+}15, VPB^{+}12]. bell [HS11]. below [PLB16, SBC^{+}11, YKL^{+}12]. below-ground [PLB16, SBC^{+}11]. belowground [KPMJ14]. Belt [MR17, ADPS10, Hal11]. beluga [MDVG^{+}15]. benchmark [JFS^{+}12]. benefit [CSF11, CMD12, CLNN14, Jag14]. Benefits [GCCJ12, BFJ^{+}11, DLZ^{+}17, DDR^{+}14, KK10, LDJZC15, MJY17, VQS^{+}18, WSE10]. Bengal [URUNAM12]. Benguela [BPB14, HT16, OCCF17, WJS^{+}18]. Bennett [HH12]. Benning [DÖWB12]. Benthic [HPRNP12, ADJ^{+}17, BNMP11, BH13b, BCO17, BPK^{+}16, CWLC13, HFH^{+}19, HKM^{+}18, KL12, KR16, LBE^{+}13, MG18, MHG18, OdlFDP15, OBC^{+}21, PLFO^{+}12, RPC^{+}17, RCP17, SAR11, SLY^{+}17, Sin17, ZSJ15]. benthos [MSCH14a, MSCH14b]. Bernard [GCTS15]. Beroe [AALP18]. berry [GPM16b, RCG^{+}13, RCGC17]. Bertalanffy [CSOS17, KDMPC10, SIS^{+}14]. Best [HCL^{+}16, ABC^{+}17, RSD^{+}15]. beta [KCG18, SCHGM16, CDT10]. better [ACE^{+}17, GAF^{+}14, MBLC14, MPF^{+}14, OGOWW16, RRJVA13, SG11b]. Between [HTVE18, AR11, ALC19, AAW^{+}15, BFPR16, BLD10b, BMCG12, BSS^{+}14, BDTV11, BPN12, CT16, CRA^{+}11, CBB19, CAMNN16, CBA10, DdSCG11, DZW^{+}11, DBG14b, DBG14a, DSS11, DLW^{+}17, DGKSM13, DMBS11, GDC11, GVV17, GJB^{+}11, GPP^{+}19, GAV^{+}11, HB12, HCS^{+}10, Hei11, HMA10a, HWL17, JZCL14, KMMS17, LV17, LSJ^{+}18, LSJ^{+}19, LMT13, LB11, LY11, LWGH13, LLQ^{+}15, LZS^{+}17, LCPB^{+}10, LSR13, MHS^{+}17, MTA19, MP13, MDH15, NWAR11, NGRDTHRL19, Pau14, PGZ^{+}19, PSWW11, PBP^{+}15, RHR^{+}15a, RDD^{+}18, RH15, SDT10, San11, SRTS18, SCG^{+}13, Str15a, Str15b, SGKR10, TMPF18, TVT12, TQZ12, TNTH11, WN17, WFA10, WNF^{+}15, WK11, XKI^{+}11, YZ18, YLC^{+}18, YLL^{+}12, ZPZ^{+}15, ZW11, ZFZ^{+}14, ZBS14, ZWG^{+}17, dAdSC15]. Between-group [HTVE18]. Beyond [BRL12, GMTC^{+}17, HLK^{+}12, KPS17, SS10, THJ^{+}12, TB12b, TPT^{+}11, Kat11, WK11]. BGC [DAW^{+}10, HBH^{+}12, HGN^{+}16, MCWG21, OWMK10, RHvdTS14, SLZ^{+}17]. BGC-ES [OWMK10]. bi [MLMH18, vDHH^{+}14]. bi-level [vDHH^{+}14]. bi-stable [MLMH18]. Bias [DAP18, FM11, AG11, ACC11, BOGA14, CGS^{+}17, GU11, KHW^{+}16, LW17, PPBD13, YAGMM^{+}14, dSD18, vPTOW18]. biased [TB12b]. biasing [BOW^{+}12]. BIBDM [HFH^{+}19]. Bibliometric [BSMA18]. Bidens [DZS12]. biennial [WTC^{+}18]. Bifurcation [CCG15]. big [GMTC^{+}17, MN12, SLB14]. Bight [FSB13, GPR^{+}13]. Bijar [HA18]. bilateral [Ju19]. Bilberry [LM19, Log19a]. billed [JHHS17]. billion [SvBH^{+}14]. bin [YWA^{+}17]. binary [Sto10]. binding [SWS16]. Binhai [YLHZ14]. bio [CMSPJ11, IFC14, IS13, ISG13, LDCB21]. bio-economic [IS13, ISG13, LDCB21]. bio-inspired [CMSPJ11]. bio-physics [IFC14]. Bioaccumulation [TPL^{+}15, BCC^{+}14, XPY^{+}14]. biochemical [MH13, TMT^{+}18]. bioclimate [WRB^{+}12]. bioclimatic [CSS^{+}19, OBLJ13, PKW^{+}10]. biocontrol [KR14]. biodegradation [BFJ^{+}11]. biodiesel [JC11b]. Biodiversity [Bus17, DFH^{+}14, GHK17, KKBZ17, CT16, Che15b, Che16, Con11, Con12, DDR^{+}18, DLW^{+}17, FL11, KK16, LL17, MRP14, NJM11, PK12, RE10, SRTS18, SBC^{+}15, SB15, WLS^{+}10, WLS^{+}11, dSJHG10]. bioeconomic [KMMS17]. bioenergetic [BWC17, CW17, FAT^{+}15b]. bioenergetically [HHUD14]. bioenergetics [BTB17, DC10, DJT^{+}19, DMB^{+}19, DHC16, FE14a, JSH15, PHP15]. bioenergy [CSH11, CSH13, HGN^{+}16, KØ13, LCPB^{+}10, PMS^{+}17, PANL11, WMF^{+}13]. bioethanol [SØB14]. biofilm [DLCR13]. biofilms [MFD^{+}15]. biofuel [LRM^{+}14, TFH12, YDD^{+}13]. Biogenic [CSH13, OC15]. Biogeochemical [OWMK10, AL16, BAB^{+}16, BLLJ17, CLL14, DWP13, GS10, GCM^{+}13, HPG^{+}10, KTGH19, MSGG14, MU10, MBUM12, MJC^{+}11, MF12, Nak17, NP18, PCTW17, RPSA12, RML^{+}13, SPM^{+}18, SFC^{+}12, SG11a, WB17, ZPZ^{+}15]. biogeochemistry [GBG^{+}15, PHP^{+}14, SEHD18]. Biogeographical [AR11]. biogeography [BLD10b, CWMC^{+}12, HKVD16, TPO^{+}18, VPC11]. biogeophysical [LHZS10]. bioinspired [CMGF14]. Biological [AG15, MCT10, PFP^{+}16, WJL^{+}13, AHM^{+}16a, AD16, BMN11, CHvdW^{+}10, CTC^{+}18, CKH15, CRVS13, Dol10, FBA^{+}11, GPH17, Gre12, GJK^{+}15, HYG14, HDL^{+}16, IGR^{+}18, JB11, JX15, LCAJGFNS14, LCR12, LZB17, MWY^{+}18, MJB^{+}11b, MvdWU^{+}16, MLK17, Mon13, SR14, VRL14, VG19, VJR^{+}16, VPM^{+}18, YLW14, ZLG16]. biologically [CLL14]. biology [HLS^{+}14, LBM16]. biomanipulation [LLKY18, OHS^{+}17]. BIOMAS [IS13]. Biomass [RKC^{+}15, RHD^{+}13, BDL^{+}14, BWC17, BSO13, CSR15, CRD^{+}16, CTP^{+}17, FMLM11, GPK^{+}13, HFH^{+}19, HLLP^{+}10, JDGJ13, LBE^{+}13, LBD10, LdMS^{+}16, LSL10, MBD^{+}18, MSCH14a, MSCH14b, MWD^{+}15, NBD^{+}13, Oga18, PCTW17, PP16, PANL11, RMHE15, SGAS14, SAI10, SSY14, SLLW15, VHM^{+}17, Vav10, VALC12, WMF^{+}13, WCG^{+}11, XGH^{+}17, YZ18, ZWRK13]. Biomass-dispersal [RHD^{+}13]. biome [FLY13, RHvdTS14, HBH^{+}12, SLZ^{+}17]. BIOME-BGC [RHvdTS14, HBH^{+}12, SLZ^{+}17]. biomes [SBC^{+}11, WBC^{+}12]. biometric [Rem11]. Biophysical [BAP18, JXL^{+}15, BAB^{+}16, BU11b, DK17, FC11, LB18a, PBO19, THC18, VPB^{+}17, ZH13]. biophysics [RDD^{+}18]. biopore [GPK^{+}13]. bioresource [Ret12]. biosecurity [GKKW10]. Biosphere [HV15, San17, COY^{+}10, PSSH11, PSS16, SNS^{+}12, SHC^{+}15]. biospheric [HPW^{+}12]. biota [BNGP19, CKS^{+}17, CSS^{+}17, DKJH13, KCB^{+}17]. Biotic [AL16, Els10, FM10, FMW^{+}19, GSSM13, LMR16, MHL^{+}19, MSV12, NGL13, PGS16, RDD^{+}18, SGM^{+}14, VPC11]. Biotope [CB19]. biotransformations [GBB^{+}11]. bioturbation [RGU^{+}15]. biphasic [AB13]. biphenyl [ZLLZ13]. biphenyls [SHS^{+}12]. birch [HKN17]. bird [BBRJ15, CPR10, IBD^{+}19, KHKC^{+}18, PK14, RJ11, SDKE18, TRP12, TDH12, Tho10a]. birds [CHR^{+}13, EGMW12, EEFE^{+}11, FMW^{+}19, GSF^{+}16, Kos17, MLK^{+}11, SJGB16, TZD^{+}15]. birth [JDZC14, NYF16]. birth-and-death [JDZC14]. birthday [GB16b]. Biscay [GPH17, LBSB^{+}14, PHP15]. bistability [RCP17]. bistable [Ber13]. bisulcus [LAEAG12]. bivalve [CN10a, IFC14, MG18, MHG18]. bivalves [FP17, MSR^{+}14, SvdMKS11, SFvdM^{+}17]. black [BFJZ^{+}11, BRKP16, CAC^{+}16, DSBM^{+}18, DBG14b, DBG14a, DNFG18, DFK^{+}11, JCKG15, LBR^{+}17, LBWW14, MSH^{+}19, MCG19, TVF^{+}10, THO10b, ZDL^{+}17, AALP18, GS10]. black-crowned [BFJZ^{+}11]. blacklip [LJWN14]. blackwater [WB16]. blacodes [CRQW18]. Blanc [CRA^{+}11]. bleaching [BMR^{+}18, YSB11]. bleak [MMVAHMC16]. blight [PHvB^{+}18]. blind [RV13]. bloom [CIHL14, HHC15, KLS17, NvGZ^{+}16, SELB10, SLW^{+}14, WBS13, WFW^{+}17, XJY^{+}16]. bloom-dominated [SLW^{+}14]. bloom-forming [WFW^{+}17]. blooming [WX11]. Bloomington [Jør14a]. blooms [BS18, CMLY19, DG11, DA19, DWP16, GRB12, GDDFT16, MWY^{+}18, MHC16, SQWS19, YXZ^{+}21, ZYZ17]. Blue [HBT^{+}18, MFB17, CWJ^{+}13, KK12a, MTZ^{+}10, WCG^{+}11]. Blue-throated [HBT^{+}18, MFB17]. blue-winged [KK12a]. blueberry [BHD17]. bluefin [PAGCV^{+}10]. bluegill [GV11]. BMP [GWHL15]. BMPs [MJY17]. boar [RCS^{+}19]. Board [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano14a, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r]. Board [Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano21a]. boat [CLD^{+}17]. body [CW17, CB15, SMT19a, SMT19b, SCHGM16]. bog [WRM^{+}11]. Bogotá [RMLTCQ^{+}19]. Bohai [DSW^{+}17, LLP17]. Bolboschoenus [AML13]. bole [Gav15]. Bolinao [GL10]. Bolivia [SAKB^{+}10]. Bolivian [TSSL11]. bollworm [FGH^{+}10]. bomb [HKlL14, KHW^{+}16]. bomb-produced [HKlL14, KHW^{+}16]. bonariensis [BCC^{+}14]. Book [Jør14a, Kat11, Pat14a]. Boom [PMB^{+}17, ABAA21]. Boophilus [WTG^{+}16]. boosted [MROGVR12]. boosts [FN18]. bootstrapped [SMT19a, SMT19b]. Boran [TT17b]. Boreal [Hol13, MPL^{+}16, ARM^{+}12, AMR16, BHC^{+}15, BS16b, BS17, CAC^{+}16, CSH11, DBG14b, DBG14a, FAMD^{+}12, GZK^{+}11, GWK^{+}12, GCB^{+}11, HMS^{+}10, HCS^{+}10, KRL^{+}16, KLS^{+}14, LKKK^{+}14, LKLK15, LM19, Log19a, LSR13, MK12, MCG19, SSK14, SPG11, TLRL11, Vav12, WGA^{+}13, WMK^{+}13, WJK11, XKI^{+}11, ZJM^{+}11, ZZB^{+}18]. boreas [BKP10, Bar11]. borer [KKR19, MBD^{+}12, RCG^{+}13, RCGC17, VCSSC12]. borne [LBR16, WGTH16, WGE^{+}19]. Bornean [KK12b]. Bornholm [PWK^{+}17]. Botswana [LMCA^{+}12]. bottlenose [BWC17, MMCBP19, PNHL14]. bottom [BWML15, FCS13, FRP18, TSN17, ZZL^{+}19]. bottom-attached [FCS13]. bottom-up [BWML15, FRP18]. bounce [DKM^{+}14]. boundaries [ALKB10, BM11, LFH^{+}15, Mur13, PFA^{+}16, TCGC16]. boundary [HRSKW15, YCF15]. bounty [JSW^{+}13]. Bourget [BGPJ10, BMD18, KJVL^{+}17]. Bow [NFU19, WR13b]. box [LMT13, THO10b]. Brachycephalidae [PMF^{+}13]. Brachycephalus [PMF^{+}13]. brackish [CHPS17]. branch [ZGF14]. brasiliensis [KMM^{+}13]. Brazil [CSFB17, dFMdSA18, AAO10, BSS^{+}14, DBK^{+}13, GAG^{+}12, GRL18, IGR^{+}18, RCCN11, dHR18]. Brazilian [CGY11, LBE^{+}18, PMF^{+}13, PGA^{+}11, dSD14]. breadth [LJ15a, SWS16]. break [RMGU11]. Breeding [MGB13, BVVS15, DHS^{+}16, HPS10, HSTS19, JHHS17, LDS14, MSCM18, MBM^{+}18, SGA^{+}16, SSQ^{+}10, VHDM17]. brevis [DA19, HHC15]. bridge [FWRZ10, KMMS17, LMT13, WBG^{+}14]. bridge-replacement [FWRZ10]. bridges [GPP^{+}19]. Bridging [DLW^{+}17, HB12, LSJ^{+}18, LSJ^{+}19, SCG^{+}13]. brigalow [NJM11]. brings [TYEH10]. Bristol [PPDF14]. British [Gar17, MCS^{+}18, RLS11, WFA10]. broad [BGD^{+}16, RHZ^{+}18, ZLB14]. broad-leaved [RHZ^{+}18]. broad-scale [BGD^{+}16, ZLB14]. broadcast [CH14]. broadened [VSB13]. broadleaved [CZC^{+}15]. Brood [BHHM10]. brooding [MATS14]. brook [FB13]. brown [AMM^{+}14, CSR15, CESC19, FB13, KSZW16, LMGW16, LARZN13, MMMGLN^{+}18, OJM^{+}12, RTE^{+}15, LTL15, Ten15b]. Brownian [WBG^{+}14]. browsing [CHG^{+}15, DDM^{+}17]. bryophyte [LKLK15]. Bt [FPO18, GV15, HLLO13a, HLLO13b, HQF^{+}17, PPZ^{+}17]. Bt-crops [HQF^{+}17]. Bt-maize [HLLO13a, HLLO13b]. bud [NS15]. budburst [FCV^{+}12]. Buddleja [PKD11]. Budget [GMvV17, SKTG11, EIG^{+}10, HMP^{+}19b, JR16, JNHF18, JHT^{+}14, PGK^{+}13, ZLG16, DLL^{+}19, GPH17, JZ12, LLC^{+}17, PLS^{+}10]. budgets [DJS10, ENM11, GS10, JBH^{+}14]. buffalo [KMRW16, MMS12]. buffer [SRXP14]. buffers [ZLJM17b, ZLJM17a]. Bufo [BKP10]. build [BMR^{+}18, CMEP15, vdVJS16]. build-up [BMR^{+}18]. Building [Ano12a, GPP^{+}19, KAC^{+}12a, KAC^{+}12b, LB17, SBC^{+}16a]. built [GA15, MGS15, SLH^{+}11, SBC^{+}15]. built-in [GA15]. Bulk [KMRW16]. bullfinch [MAC^{+}12]. bullhead [KDMPC10]. bumble [BBB^{+}17]. buoyancy [MUP^{+}13, SAI10]. buoyant [BM16, GPJL15]. burdens [CW17, MGCL17]. Burm. [MAO19]. burn [dMHGA10]. burning [CPD^{+}11, SSY14]. burrowing [EGMW12]. Burullus [ESA12]. bush [BG11]. bush-crickets [BG11]. business [FTTV^{+}18]. bust [ABAA21, PMB^{+}17]. bustard [MAM^{+}12, SSQ^{+}10]. butterflies [DPBK10, FJH15]. Butterfly [WPZ^{+}14, CJM^{+}12, CWJ^{+}13, EST^{+}19, GQRC^{+}17, GPZB18, RJG^{+}13, RKBG14]. Butterworth [WSS^{+}16]. by-catch [NNST^{+}14]. bycatch [LJ13].

C [Jør14b, GPM^{+}16a, DCMK11, DAW^{+}10, DVD^{+}14, EFM^{+}10, HKlL14, JS12, KHW^{+}16, MKTO19, TTCH^{+}14]. C-TOOL [TTCH^{+}14]. C. [AR11, PL10a]. C3 [HBP18, YqWB^{+}15]. C4 [HBP18, YqWB^{+}15]. CA [TLLX19, dWcWbZ15, YZC14, GVV^{+}16]. cabbage [HWH^{+}18]. cackling [WK17]. cadastral [VVA^{+}15]. Cadiz [TCH^{+}13]. cadmium [BFP^{+}18]. CAFFEE [SM13]. caffer [MMS12]. CAGD [AS11]. cage [CHCC15, RPC^{+}17, TPT^{+}11]. Calamagrostis [Log13a]. Calanus [JSH15]. calcarifer [HD16]. calcification [MB14]. calcite [HC11]. calculate [Jør15a, Jør15b]. Calculating [BLCPB11, CCS13, LY11, SBDA16, LLCR10, Pat14b, MBRR13]. Calculation [JMW11, PPF11, Cam16, CRQW18, Sci11a]. calculations [FSB13, Fie12]. calculus [LTL15, Ten15b]. Calcutta [MGMR15]. calibrate [vdVJS16]. calibrated [HMS^{+}11, KPS19, PBP^{+}15]. Calibrating [PWP^{+}19, JAKJ10]. Calibration [GGM^{+}12, MPL^{+}16, PBB^{+}14, vdVBJS15, BVC^{+}18, BDV^{+}19, CDK^{+}17, CBA16, GdSHA10, GBP^{+}10, HHvO^{+}17, JSL^{+}10, KHW^{+}17, LH15, LYJ^{+}15, Log13a, MA10, MBUM12, NGWZ16, OGOWW16, OCD^{+}13, RKL^{+}15, RCCN11, SFC^{+}12, XBvdV^{+}14, ZCYL15]. California [CMCASS12, RBGB16, DUASLB^{+}12, GEZ^{+}13, HKM^{+}18, KKH^{+}17, KEM^{+}16, LVZ^{+}11, MKL14, MJK^{+}17, TPG14]. californicus [dOMZ13]. call [LBGM19]. callarias [IKL13]. callitris [NBD^{+}13]. Callorhinus [ZKSF17]. calls [WSS^{+}16]. Calotriton [CMGF14]. calvescens [PMS11]. CAM [HBP18]. Cambodia [KMM^{+}13]. camera [CGS^{+}17]. Campaign [LLZ^{+}16]. Campania [MRZU14]. Can [BU11a, CVV^{+}13, CJM^{+}12, FE14a, GAB12, Har11, MCRB17, MMMCAHM21, PTML14, RMGU11, RPGSCF16, SBM^{+}11, SVH^{+}19, TKW17, TRP12, TBL^{+}15, WKZ^{+}19, WR13b, YLW14, YHW^{+}12, YHJ^{+}17, ZKR^{+}13, BOGA14, Che16, CHZ18, CHG^{+}15, dSCdA19, FS14b, FTYI13, GVGC16, Gav15, HSLC^{+}19, Lee14, MBLC14, NZL13, NYH15, POS^{+}19, Rin14, SS16, SA16, SJB^{+}16, VEMS12, dAdSC15]. Canada [CLD^{+}17, LMY^{+}11, LYP^{+}11, LPA^{+}11, MSHC13, ABK^{+}16, BHC^{+}15, CAC^{+}16, ENC^{+}15, HMS^{+}10, KZR^{+}13, LYD^{+}17, MAM^{+}10, SHM^{+}14, VCSSC12, WFA10, ZJM^{+}11, ZKR^{+}13]. canadensis [PGZ^{+}19]. Canadian [HSK17, HCS^{+}10, SKTG11, SPG11, YDD^{+}13]. canals [ASL13, FBL^{+}10]. Canaria [CMCC15]. Canary [MGGLBIBD13]. candida [HFSL14]. canine [BG15]. Canis [CWL^{+}16]. cannibalism [GVGC16]. cannibals [SKPHW16]. canola [SFHK12]. Canonical [JG10]. canopies [For14, MJGC13]. canopy [AS11, AS12, Bai18, DKM^{+}14, FAC^{+}19b, GZK^{+}11, GGR^{+}13, KGS11, KKL^{+}12, KMM^{+}13, LKLK15, LB18a, MK14, Pie10b, Pie10a, SMSAN^{+}14, SRV^{+}12, XKI^{+}11, dLB19]. Canterbury [JBB^{+}11]. Cantin [CRA^{+}11]. capabilities [SBC^{+}16a]. Capacity [VEK12, BDL^{+}14, BLCPB11, hCjHMW18, CN19, EMH10, FHAG12, GLA11, GL10, IFC14, JL18, KTM^{+}16, LDG14, NCK10, PW10b, SWS16, TT17b, WB16, XCL^{+}11, XYY^{+}16, YZ16, ZLZS14, ZZL^{+}19]. Cape [CRA^{+}11, HB14]. capita [CTC^{+}18]. capital [BPD^{+}19, DYU^{+}12, FBDR17, Ken17, PPB^{+}18, PBD^{+}17, VPB^{+}17]. Capra [AR11]. capricious [CN19]. Capture [SCHGM16, CGAB17, HVL^{+}19, RGHS11, vPTOW18]. capture-mark-recapture [HVL^{+}19]. Capturing [JSH15, BCJ^{+}12, GHWG^{+}14]. Carapa [KCKdOW12]. carbohydrate [SBC16b]. Carbon [ABK^{+}16, CSS^{+}19, FDR^{+}17, Fel12, GS10, HPW^{+}12, MPM^{+}15, SKK^{+}13, SKH^{+}13, SKTG11, SS11, VJR^{+}16, Ant14, ACN14, BMMC21, CVV^{+}13, CZC^{+}15, COBF13, CAC^{+}16, CS18, CHVH13, CQS15, DJS10, DFRT14, DLL18, FBdP^{+}16, For11, GTC16, GHP12, GE17, GCB^{+}11, GGVB11, GRD^{+}14, HMS^{+}10, HWS11, HHF^{+}10, HBA16a, HBA16b, HAA^{+}21, JWX18, JDT^{+}17, JN15b, JAKJ10, JNC^{+}11, KH11, KJMA11, KRL^{+}16, KHW^{+}17, KVS17, KPS19, LZLC13, LLZ^{+}16, LHP16, LLT^{+}17, LHL18, LVZ^{+}11, LHR^{+}15, LFH^{+}15, LBC^{+}15, LBdA^{+}19, LDYL18, LB18a, MLB^{+}16, MKF^{+}11, MJP^{+}12, MVC^{+}17, MTL^{+}10, MDF^{+}18, MCG19, MJF19, MLH^{+}17, MB14, MRG13b, Nak17, NP18, NJM11, Oga11, OKS^{+}11, OLG^{+}13, OQH^{+}13, PAS^{+}10, PP16, PMB^{+}17, PDAL14, PD15, PLFO^{+}12, PBFS14, PGK^{+}13, RHZ^{+}18, SHB^{+}11a, SKMB11, SHM^{+}14, SYL^{+}13, SSY14, SEHD18, SCG^{+}16, TTCH^{+}14, TLL^{+}15, TRE^{+}12, UKI^{+}11, Vav12]. carbon [VRC17, VPM^{+}18, Wan19, WO14, WMK^{+}13, WJF^{+}11, WLH^{+}10, WZD^{+}14, WR18, WJK11, WJvdL^{+}13, XZSL17, XGH^{+}17, YYH19, ZZY^{+}13, ZLLC13, ZJM^{+}13, ZJL^{+}14, ZLF^{+}15, ZLW^{+}15, ZWF18, ZZLZ14, ZHCX15]. Carbon-based [SS11]. carbon-nitrogen [KPS19, PMB^{+}17]. carbonate [SG11a]. carcass [Pér18]. cardinal [Pat16a]. Caretta [WHRR15]. Caribbean [ABAG14, KTK^{+}15, KDT^{+}17]. caribou [BMSLC16, SMH^{+}12]. Carlo [BWA17, CTMS10, HT14, Ken19, KGM12, MA10, Qia12]. carnivore [CGV^{+}15, MV11]. Carp [CPS15, AHS^{+}16, CEWW18, CSN^{+}19, EEB^{+}17, EEK^{+}18, GJM^{+}13, SLB^{+}18, TW18, ZSG^{+}18]. carpio [CEWW18]. Carposina [KL10]. Carposinidae [KL10]. carps [WKZ^{+}19]. Carry [RH17]. Carry-over [RH17]. Carrying [CN19, KTM^{+}16, NCK10, BLCPB11, GL10, IFC14, LDG14, SWS16, TT17b, XCL^{+}11, YZ16, ZLZS14, ZZL^{+}19]. carrying-capacity [SWS16]. Cartesian [HHW11]. cascade [BFP^{+}18, CD17, CSFB17, DFW^{+}18, XC14]. cascaded [CCL^{+}13, WhFxGZ16]. cascades [FAMD^{+}12]. cascading [SH18]. Case [Ana17, Ano12a, GWHL15, GDDFT16, HBR19, HFSL14, KSSA13, KHJ^{+}19, LLL^{+}10, LYC^{+}15b, MRD^{+}12, SJS^{+}13, WX11, AGQP^{+}14, ABS17, ACM15, AHM^{+}16b, BVC^{+}18, BRV11, BLD10b, BC11, BG11, CMT^{+}11, CVV^{+}13, CZC^{+}15, CRA^{+}11, CBGK15, CWLL15, CL15b, CXL^{+}19, CAMNN16, CGEdM12, CRTS15, CMSPJ11, CBB^{+}10, CWA^{+}15, DLL18, DSW^{+}17, DLD^{+}11, DLvGGNC16, EGMW12, FKH^{+}15, FCZ13, FCH^{+}17, FP17, FCFT16, FBDR17, Fre16, GHWG^{+}14, GZS^{+}18, GDT^{+}14, Gia14, GPA^{+}13, GMTC^{+}17, GS19b, GGTP^{+}17, GB15, HRSKW15, HJ17, HM18b, HYL^{+}17, JMW11, JJYM19, KZR^{+}13, KAC^{+}12a, KAC^{+}12b, KCKdOW12, Kos17, KL12, LDG14, LSH15, LHL^{+}17, LYSW15, LLZ^{+}16, LHL18, LDM15b, LBS^{+}14, LYC^{+}15a, LML17, LMBL12, LGDW12, LB18b, MTZ^{+}10, MM19, MAM^{+}12, MAC14, MRZU14, MDVG^{+}15, MMVAHMC16, MOM^{+}11, Mur13, MWD^{+}15, mNkGyT^{+}17, NXZ18, NGGFM19, NPP^{+}18, OHS^{+}17, Oga11, OTA16, PLR^{+}13, PHvB^{+}18]. case [PMS^{+}17, PPB^{+}18, PWP^{+}19, PBD^{+}17, PWK^{+}17, PLS11, PGGMMA19, RSCS19, SMC11, SSV^{+}12, SK11, SGRAdJ17, SDKE18, SAKB^{+}10, SGP18, SSM^{+}13, SGKR10, TPL^{+}15, THJ^{+}12, TPG14, TAC^{+}14, TOQW11, TVV13, UDNSLCG16, VVH12, VKdW^{+}10, WALCY12, WSW^{+}13, WSSM15, WFN^{+}16, WØ15, YYQ11, YJW15, YLHZ14, YBA^{+}16, ZPZ^{+}15, ZBAY14, ZLY16, ZYFL10, ZFT^{+}10, ZLC13, ZKR^{+}13, ZLF14, ZLF^{+}15, ZZF15, ZLFZ16, ZZJ^{+}16, ZDL^{+}17, ZZLZ14, ZMH19, dSdCF10]. Case-study [MRD^{+}12]. cases [Flo14, LBGM19]. Caspian [SZSV12b]. Castor [PGZ^{+}19]. Catalonia [GPM^{+}16a]. catastrophe [PZM13, WLZC11, WPZ^{+}14, YL17]. catastrophes [VEK12]. catastrophic [CLCC11, SK15, ZW12]. catch [DMH17, LCD^{+}16, MROGVR12, MTGD17, NNST^{+}14, ZBMBDU^{+}17]. catch-and-release [LCD^{+}16]. catch-at-age [ZBMBDU^{+}17]. catchability [BDG^{+}14, KSZW16]. catches [CJR^{+}16]. Catchment [SMC11, GMP13, HPO^{+}15, KW18, KSSA13, KSG^{+}14, OGOWW16, PTE^{+}12, WFM16]. catchments [JNC^{+}11]. categorization [CPJ^{+}13, DPV^{+}16]. categorizing [VBS14]. caterpillar [BFA11]. catfish [dABLBAdOR17]. cations [NP15]. CaTMAS [BMMC21]. cats [IN16]. cattail [TXM^{+}10]. cattle [LCH12, TT17b, WTG^{+}16]. Causal [FRP18, Fai10, KdVRdW17, RPA12, SBM^{+}12]. causality [KS17a, KS17b]. cause [BAAB18, KWGG19]. cause-effect [KWGG19]. caused [CLCC11, FPO18, HLLO13a, HLLO13b, RCBPMN10, SGWV15, SH18, WSZT10]. causes [BCV13, ODK16]. causing [CBGK15]. caution [Rau11]. CBM [ABK^{+}16, HSK17, MK12, PGK^{+}13, SKTG11]. CBM-CFS3 [ABK^{+}16, HSK17, MK12, PGK^{+}13, SKTG11]. Cd [HFSL14]. CDOM [UDRA13]. CDPD [Ben12]. CE [ZSJ15]. CE-QUAL-W2 [ZSJ15]. cell [HB15, KH15]. cellular [Cra10, DAF^{+}11, GDT^{+}14, GLI^{+}21, HHL17, MJÁ^{+}12, PD12, YZL12, ZJL^{+}16, ZHLZ17]. cellulase [ABAG15]. cellulose [ABAG16]. censored [WSC^{+}19]. census [PC17, PC15]. Central [BWRA16, DUASLB^{+}12, FCFT16, FBDR17, YLW^{+}11, BG15, BBL11a, BPWB14, DDBK15, DBG14b, DBG14a, JMW11, KMM^{+}13, LDS14, MLC^{+}16, MJP^{+}12, MK19, PAS^{+}10, RCBPMN10, RR10, CLCC11, CPS^{+}14, GB15, HLZ^{+}18, LLC^{+}21, SKMB11, SK11, SBC^{+}11, TNA^{+}17, TKH17, WGSR19, ZLLC13]. Centralized [KBMM15]. centric [LD16]. Centris [GPA^{+}13]. centroids [WFA^{+}13]. centuries [Fel12]. century [ENC^{+}15, NMS^{+}14, WMF^{+}13]. Cerastoderma [RGU^{+}15]. cereal [KKBZ17]. cerevisiae [PGG14]. Cerrado [LBE^{+}18]. Cerro [RR10]. certain [NXZ18]. cervicornis [NGRDTHRL19]. Cervus [RR10]. cessation [WJ15]. CFS [LAD11]. CFS3 [ABK^{+}16, HSK17, MK12, PGK^{+}13, SKTG11]. CH [Vav12, VRL18, GKW^{+}13, HMS^{+}11]. CH4MOD [LHZS10]. Chain [BWA17, CTMS10, KGM12, BFP^{+}18, CSF11, CLCC11, JC11b, PSWM^{+}10, Pea15, Qia12, TM12, TVV13, YZL12, FCH^{+}17]. chains [GLS10a, JLN10, LSP^{+}10, XYTS19, vGKB10]. challenge [LJAD17, SS10]. Challenges [HT16, TSB^{+}18, BMSLC16, Bec10, LMY^{+}11, RHR^{+}15c, SGB10, WHH^{+}18]. chalumnae [CPE13]. Chamaecyparis [Oga11]. chamber [MDW^{+}12]. chamise [PFSF14]. chance [RRJVA13]. Changbai [CZC^{+}15]. Change [LWKS10, AW16, ABAG14, ACDC17, AGB^{+}13, ACK^{+}14, AH17, ARV^{+}16, BUB^{+}12, BTS19, BCTL16, BTB17, BKM^{+}18, BGPJ10, CLH17, CY13, CWLL15, CJR^{+}16, CSR15, CMGF14, CJM^{+}12, CWJ^{+}13, CDM13, DK17, DWD19, DWWS15, DPR10, DDR^{+}18, DAF^{+}11, DMR13, DP15, DBP16, FLX^{+}14, FTK^{+}14, FCG^{+}19, FMS13, FCFT16, FAS^{+}19, FP13, FBKS17, GLA11, GBG^{+}15, GGZ^{+}16, GAG^{+}12, GCCB18, GPR^{+}13, GLI^{+}21, HHC^{+}13, HHL17, HV12, HGE^{+}10, HSR^{+}15, HPC13a, HPC13b, JDT^{+}16, KTV^{+}15, KFvSH15, KBS^{+}15, KPTC14, KPK^{+}17, LLZ^{+}15, LWL^{+}17b, LQW^{+}19, LS10, LDP14, LVZ^{+}11, LC11, LLZ12, LKAY15, LYL^{+}17, LZZW17, LBC^{+}15, Loe18, LCDJ^{+}12, MLC^{+}16, MLK^{+}11, MTPC13, MCS^{+}18, MDOJ14, MDKK15, MRM^{+}17, MCG19, MHC16, MSR^{+}14, MMRS^{+}19, MMMGLN^{+}18, NA11, Oga11, OJ10, PH18, PD15, PAB^{+}13b, Phi11, PW10b, Pra12, RJG^{+}13, RVL^{+}18, Rem11]. change [RCSS13, Rin18, RA13, RWvDM15, RBGB16, SvBH^{+}14, SS16, SKMB11, SKK^{+}13, SBRD10, SJB^{+}16, SSN^{+}18, SDY^{+}15, SS11, SB15, TCW11, TAT^{+}16b, TNA^{+}17, TZD^{+}15, TT17b, WO14, WNT^{+}13, WMK^{+}13, XZSL17, YYL13, YHS^{+}11, YZL12, YZC14, YFC^{+}11, YDL^{+}15, YKvWK16, YHJ^{+}17, ZLLC13, ZLD^{+}19, ZJP^{+}11, dFMdSA18, vDSC^{+}10, vVBHZ11]. change/variability [LC11]. Changes [KHL^{+}16, MLB^{+}16, MMRS^{+}19, PVL^{+}14, ACC^{+}10, ACM15, ADF^{+}17, AH17, BYR18, BTS19, BPR^{+}11, BCI^{+}17, BAJ^{+}11, CZ11, CWM12, CSL^{+}12, FAG10, FLY13, Fel12, FG11, FTTV^{+}18, GAP^{+}19, GJÁ^{+}10, GE17, GWK^{+}12, GKW^{+}13, HTL^{+}14, Hoa11, HPC13a, JDT^{+}17, KKK^{+}19, LDS14, LY16, LDO^{+}18, LCBNP19, LSJG17, LLT^{+}17, LLC11, MBD^{+}18, MBLC14, MNEG11, NP18, NMC^{+}14, NMS^{+}14, OGG^{+}16, Oga17, OLG^{+}13, SGAS14, SJGB16, SBF15, SED15, TAGC^{+}16, VPC11, WGTH16, WKZ^{+}19, WYY^{+}19, WJS^{+}18, WNT^{+}13, WV12, WJvdL^{+}13, WFM16, XYY^{+}16, XC14, YDD^{+}13, YHK^{+}19]. changing [Ana17, BFKS16, CGLA10, CW11, GHD^{+}19, mGZK^{+}10, GWK^{+}12, GKW^{+}13, GJK^{+}15, Hei11, KWJH16, LKH^{+}18, LdMS^{+}16, RJH^{+}10, RBGB16, TB12b, TWS^{+}11, WP14, dSJHG10]. Changjiang [FS14a]. Changsha [ZZJ^{+}16]. Changzhou [LYSW15, TLL^{+}15]. channel [CZW10, CCL^{+}13, Jag14, MH13]. Chaoborus [SKPHW16]. Chaohu [KJH^{+}13, NXZ18]. Chaos [MCN16, RE10]. chaotic [AAA13]. Characterisation [Mon13]. Characteristics [BS15, BV15, CNI11, DCCS17, GO15b, GLY^{+}15, LBH^{+}15, Lin15a, LYCZ11, MBW14, RKBG14, RBGB16, TMT^{+}18, YXZ^{+}21, ZLC13, ZLF^{+}15, ZLZ17c].

Characterization [BMN11, Ing10, KCLL14]. characterized [GRB12, Sol19]. Characterizing [BSW18, KHS^{+}19, MTJF11a, TCSG15, ALG16, BGK^{+}16]. Charadrius [GKF^{+}14]. chart [DFW^{+}18]. Cheat [ZF11]. cheetah [JMdW^{+}10, OMDP11]. Chemical [TA11, AJDS16, CGG^{+}18, EGMW12, FtHvdB^{+}14, HD16, LBW18, Vav12]. chemical-induced [LBW18]. chemicals [GHWG^{+}14, GT14, HFSL14, MPF^{+}14]. chemistry [HKL^{+}11, LGY10, MB14]. chemostat [AAA13, CJLV11, F.11, FHC15]. chemostats [CBB19]. Chesapeake [KH11, CN10a, DC10, JX17, KPLS15, MTZ^{+}10]. chestnut [GPTR13, PFP^{+}16, PGGMMA19]. Chile [BCO17, CRQW18, LAEAG12, MNEG11, ROSC15]. Chilean [MTPC13, Bae18]. chilling [GLR^{+}19]. chimeric [POS^{+}19]. China [LCC^{+}14, NvGZ^{+}16, PCTW17, ZYZ17, AA12, BCC^{+}11, CZC^{+}15, CZ11, CL15b, CLWH15, CMLY19, DZW^{+}11, DYU^{+}12, DLZ^{+}17, DLL18, DSW^{+}17, FLY13, FLX^{+}14, FDR^{+}17, FCG^{+}19, FGH^{+}10, FCZ13, GYL^{+}15, GSQ17, GZS^{+}18, GLS^{+}17, GCL^{+}19, HJ17, HFB^{+}11, HZT^{+}13, HG17, HYL^{+}17, JWX18, JLZ^{+}18, JJYM19, JC11b, Ju19, KJH^{+}13, LkSXhX17, LLL^{+}10, LSZC10, LLY^{+}13, LCY^{+}13, LZLC13, LLM^{+}13, LYSW15, LLQ^{+}15, LZZP16, LWL^{+}17a, LKG^{+}17, LWL^{+}17b, LWTZ18, LHB^{+}18, LZJ^{+}19, LCZ11, LC11, LCS11, LLZ12, LYC^{+}15b, LML17, LYL^{+}17, LLP17, LP17, LCFD19, yLWhMhG11, LGDW12, MLW^{+}15, MLB^{+}16, MWY^{+}18, MLH^{+}17, NXZ18, PHC10, QCLW18, QCB^{+}11, RZF^{+}13, SRXP14, SLW^{+}14, SWC13, SLZ^{+}18, SAW^{+}11, SDY^{+}15, SM14, SXZ^{+}15, SCG^{+}16, TLL^{+}15, TJYZ11, TQZ12, TZS^{+}14, WALCY12, WSW^{+}13, WYHC15, WLY^{+}15, WWC^{+}18, WKZ^{+}19, WYY^{+}19, WJF^{+}11, WZD^{+}14, WZG15, XYTS19, XTLM15, XLS11, XZC^{+}17, XCL^{+}11, XWW^{+}11, XYCZ13, XQZ^{+}13, YYL13, YWH15, YLY^{+}18, Yan11, YYQ11, YC11, YLC13, YC13, YZPL16, YDZ^{+}21]. China [YM21, YSA^{+}14, YLW^{+}11, YDL^{+}15, ZFT^{+}10, ZLLZ13, ZLZS14, ZLF14, ZQC15, ZLW^{+}15, ZLFZ16, ZZJ^{+}16, ZLZ^{+}17a, ZCSL18, ZWA19, ZXD^{+}13, ZZLZ14, ZJP^{+}11]. Chinese [AAW^{+}15, BRHD15, KHL^{+}16, LWL^{+}17a, LPX^{+}12, WKZ^{+}19, WZD^{+}14, XWC^{+}11, YWY10, YHZ19, ZYFL10]. Chl [MMM^{+}16]. Chlorella [YLC^{+}18]. chlorophyll [CS18, KHJ^{+}19, LGY10, MSGG14, TLZ17, WQD^{+}13, ZZT13]. chlorophyll-to-carbon [CS18]. Choice [DKJH13, NA11, HMD^{+}16, HS13b, Loe12, ROG^{+}16, RVB^{+}10, VF16]. choices [ZAW^{+}18]. Chongqing [yLWhMhG11]. choose [Ken19]. Choosing [POBH10]. Chozas [MJB^{+}11a]. Chronic [PMP^{+}13, HMPC11]. chronosequences [Phi15]. chub [LCC^{+}14]. chytridiomycosis [ACCM16]. ciguatera [KTK^{+}15, KDT^{+}17]. ciguatera-associated [KDT^{+}17]. ciliata [DSS11]. cinerea [ACC^{+}10]. cinnamomi [DLvGGNC16]. circadian [FS14b, Hel10]. circular [FDR^{+}17]. circulation [GL10, MB14, WLS^{+}10, WLS^{+}11, ZW10]. cisco [JFS^{+}12, MYJ^{+}14]. CISTA [Ana17]. CISTA-A [Ana17]. cited [BS15]. CiteSpace [HhLY^{+}19]. Cities [BBA12, YZLX14, BK13, GLC^{+}18, HLW15, YHZ19, ZYFL10, ZF19]. citizen [RCS^{+}19]. City [LMY^{+}11, LYP^{+}11, LPA^{+}11, LLZ12, WL13, LYC^{+}15a, NPP^{+}18, WYY^{+}19, WGD^{+}17]. Cl [GVV^{+}16]. clam [BBGJ13, DWP16]. Clark [KRE^{+}18, MSM^{+}17]. class [BSH^{+}10, HKW^{+}18, RV15, RWFD13, HBA16a, HAA^{+}21]. CLASS-CTEM [HBA16a, HAA^{+}21]. classes [POBH10]. classic [LMT13]. classical [BDL^{+}14, RHR^{+}15a]. classification [ABL^{+}13, EBL^{+}11, FIG^{+}10, GCIB10, HJWG12, NMS10, SZSV12a, SZSV13, Sil11, ZPNM13]. classification-then-modelling [HJWG12]. classified [BG11, CKMD14]. classifier [RAR15]. classifiers [RRvdG18]. Classifying [CWA^{+}15]. Claude [RHQ^{+}16]. clay [GPK^{+}13]. clear [MJB^{+}11a]. clearcutting [Gat18]. clever [SGSR10]. Climate [AW16, DBP16, GHST19, GZK^{+}11, KPTC14, MHC16, TNA^{+}17, TZD^{+}15, ABAG14, ACM15, ACDC17, Ana17, AGB^{+}13, ARM^{+}12, ACK^{+}14, BTS19, BCTL16, BGD^{+}16, BKM^{+}18, BFH14, BGG^{+}16b, BSHG16, BBH^{+}17, BGPJ10, BFKS16, CTB^{+}12, CLLB12, CCW^{+}15a, CCW^{+}15b, CY13, CAC^{+}16, CBB19, CSH11, CSH13, CJR^{+}16, CSR15, CW11, CJM^{+}12, CWJ^{+}13, CDM13, DWWS15, DDR^{+}18, DAF^{+}11, DMR13, DP15, DLW^{+}17, DSA^{+}18, EFM^{+}10, FLX^{+}14, FTK^{+}14, FCG^{+}19, FMS13, Fel12, FBA^{+}11, FAS^{+}19, FP13, FBKS17, GAP^{+}19, GHD^{+}19, GLA11, GVGZ^{+}21, GGM^{+}12, GBG^{+}15, GLR^{+}19, mGZK^{+}10, GAG^{+}12, GCCB18, GWK^{+}12, GKW^{+}13, HHC^{+}13, HV12, HlLB12, HGE^{+}10, HSR^{+}15, HPC13a, HPC13b, HRWT12, HBL18, HEJP17, JVY19, JSB^{+}14, JFS^{+}12, JL18, KTV^{+}15, KCF^{+}13, KMF^{+}21, KM12b, KBS^{+}15, KPK^{+}17, KKP^{+}17, LDO^{+}18, LZLC13, LHP16, LWL^{+}17b, LDP14, LC11, LKAY15, LCFD19, Loe14, Loe18]. climate [MLC^{+}16, MLK^{+}11, MBD^{+}12, MU17, MTPC13, MCS^{+}18, MGS16, MDKK15, MRM^{+}17, MCG19, MLH^{+}17, MSR^{+}14, MMRS^{+}19, MSM^{+}16, MMMGLN^{+}18, NZL13, NA11, NW18, OJ10, OES16, PH18, PVM12, PVM13, PVBC14, PD15, PW10b, Pra12, RJG^{+}13, RVL^{+}18, RXSK14, Rem11, RCSS13, RKC^{+}15, Rin18, RA13, RWvDM15, RVB^{+}10, RBGB16, SvBH^{+}14, SLSS15, SS16, SKMB11, SKK^{+}13, SBF15, SBRD10, SSN^{+}18, SAKB^{+}10, SB15, TCW11, TT17b, UKI^{+}11, VASL^{+}15, WGT12, WTG^{+}16, WRB^{+}12, WNT^{+}13, WV12, WRSL11, WYS12, WSBP13, WFM16, XC14, YYL13, YFC^{+}11, YKvWK16, YHJ^{+}17, ZLLC13, ZLW^{+}15, ZLD^{+}19, ZJP^{+}11, ZSSR18]. climate-based [CLLB12]. climate-change [WNT^{+}13]. climate-driven [CTB^{+}12, KM12b]. climate-induced [ACM15, BGD^{+}16, XC14]. climate-smart [DSA^{+}18]. climate/CO [FCG^{+}19]. climates [AOD^{+}13, KWJH16, LLTQ14, LLTQ15, PAS^{+}10]. climatic [AVL^{+}11, CHR^{+}13, CMGF14, FHAG12, FBA^{+}11, GCR18, MDB19, PRP17, PLHG14, SJGB16, SM13, SAKB^{+}10, YDL^{+}15]. climb [TOHT14]. CLM [SLZ^{+}18]. CLM-Crop [SLZ^{+}18]. clock [FS14b]. clocks [GJK^{+}15, Hel10]. clonal [EGA12, HW12, IWW12, MGSB11, OMH12, VMG^{+}12, VAO^{+}17, YGL18]. close [RMLTCQ^{+}19]. Closed [TM14, BO14, BS13, BDG^{+}14, DLH^{+}14, GLS10a, HCZ^{+}17, HWH^{+}18, LDJZC15]. closed-loop [GLS10a, LDJZC15]. closure [KCG18, SM13, ZCR16]. closures [LMBL12]. cloud [Bre12]. clover [KAJØ19]. CLUE [JDT^{+}17]. CLUE-S [JDT^{+}17]. clues [RHR^{+}15c]. Clustering [PMRGF10, Con11, Con12, DSK^{+}11, DPLC13, LKDD15, MLMH18, RWS^{+}15]. clusters [JSKK15, Pie10a, ZGF14]. CMCC [CPS^{+}14]. CMIP5 [SBF15]. CO [CR12b, VRL18, DMB^{+}18, FM10, FM11, KØ13, LTL15, PBO19, SZ14, Ten15a, Ten15b, CGY11, CSH11, CSH13, CN13, CR11, FCG^{+}19, FCZ13, GKW^{+}13, HMS^{+}11, Hol13, KK12b, KMM^{+}13, LKLK15, LHP16, LBdA^{+}19, MKF^{+}11, mNkGyT^{+}17, PD15, QlSeZ^{+}10, RML^{+}15, SLSS15, TIC^{+}14, WRM^{+}11, ZJP^{+}11]. co-emergy [LTL15, Ten15a, Ten15b]. co-evolution [SZ14]. co-location [DMB^{+}18]. co-occurrence [FM10, FM11]. co-product [KØ13]. co-registration [PBO19]. coal [BPU11, LCT^{+}17]. coalition [Ret12]. coarse [GGV10, SHC^{+}15]. coarse-scale [GGV10, SHC^{+}15]. coast [ADL^{+}10, CRA^{+}11, CGEdM12, DWP11, JVY19, NCK10, OCB^{+}13, ROSC15]. Coastal [BH10, RLS11, BPC^{+}17, BPGSS^{+}18, CEWW18, DJS10, DVD^{+}14, FHL^{+}13, GSF^{+}16, HBS^{+}16a, HB10b, HR11, HSCS14, IGBSASSJ16, JDT^{+}16, LSZC10, LLQ^{+}15, LRB12, MDL^{+}10, MMK15, MZLCSZdP11, NvGZ^{+}16, PRHK19, PK14, PICR13, RRHH10, RCST11, RPA12, RBGB16, SFMA15, SSN^{+}18, TBM^{+}14, URUNAM12, WLZC11, WGA^{+}11, WGA^{+}12, WEJ11, ZDG^{+}13, dMLM^{+}17]. Coco [FAC^{+}19a]. cocoa [ABAA21]. coconut [MdMS^{+}19]. Cocos [MdMS^{+}19]. cod [CAL13, Hei11, IKL13, JNHF18, MM19, PWK^{+}17, SLRC19]. code [MEC14, SBC^{+}16a]. Codium [GPJL15]. coefficient [Qia12]. Coevolution [NdA16]. Coexistence [MdlP15, ASS14, DJH10, FC12, qHLlZ^{+}16, KK11, LM19, MN12, NSIiT17, RDH^{+}13, RHR^{+}15b, SFM^{+}18, SFHK12, YLL^{+}17]. coffee [RVL^{+}18, RJ11, RCC^{+}11, RCG^{+}13, RCGC17]. cognitive [BRW^{+}12, CLT^{+}11, MRM16, RFV^{+}12, SGRAdJ17, VSB13, VBS14]. coherence [RWH^{+}14]. Cohort [Gar17, CH15, GCTS15, Och10]. cohort-based [GCTS15]. Cold [VCSSC12, ALC19, NW18, RJH^{+}10, SCG^{+}16]. cold-climate [NW18]. cold-region [ALC19]. coli [JA15b, KPSC10, KH15, SFvB10]. coliforms [RCCN11]. COLIWAVE [SFvB10]. collaboration [BME14, GBL^{+}13]. collaborative [GHD^{+}19]. collapse [ALG16, MM19]. collar [ACC11]. collection [BAJ^{+}11]. collections [PTA^{+}12]. collective [LTGD16, NG11]. Collision [KHKC^{+}18, ACSR^{+}13]. collisions [GJB^{+}11, KHKC^{+}18]. colobus [BSCG10]. Colombia [AYASR^{+}17, NCMF10, ZPZ^{+}15]. Colonial [TSP19, GDW12, MG18, MHG18, SSS^{+}11]. colonies [CESC19, KDHC^{+}18]. colonisation [FDP^{+}10]. colonization [DRS^{+}16, GVGZ^{+}21, MJÁ^{+}12, Str15a, Str15b]. colonized [JXL^{+}15]. colony [AOW12, AOW13, BHHM10, DDR^{+}14, MG18, MHG18, RBH13, YZL12]. colony-forming [MG18, MHG18]. Color [LS13, Jør14b, PH18]. coloration [LCFD19]. colouring [OJ15]. Columba [OFK^{+}12]. Columbia [WFA10, FBB^{+}18, Gar17, HSTS19, MCS^{+}18, RLS11]. column [CW17, GS10, LWGH13, WRMB18]. combination [AMM^{+}14, LCP15, MJÁ^{+}12, RDJ19, RRH^{+}19, SHB^{+}11a, YZPL16]. combinations [FKK^{+}18]. Combined [CN13, DGKSM13, GBG^{+}15, TSN15, UR10, WLC^{+}19, CSF11, FTTV^{+}18, HL19, KCKdOW12, RM13, SGBD10, VGR^{+}11, WFW^{+}17]. Combining [CPE13, FYR^{+}18, HSD^{+}15, KF16, LS10, LOLA13, MPF^{+}14, MPP^{+}15, MMMGLN^{+}18, NF11, SML12, SKM^{+}16, dHR18, IHSM12, MPT^{+}11, RCS^{+}19, ZPNM13]. come [CEWW18]. Comment [GCIB10, Par10, Pat10, Ula10]. Comments [GORJ12, NMS10, Och10, WM14, RHD^{+}16]. commercial [CMCASS12, HD16, JDP^{+}12, MTGD17, WPZ^{+}11]. Common [CPS15, Mar17, MTA19, AW15, AHS^{+}16, BCC^{+}14, ODK16, PMH17, PKG16, SZ14, SDEP18, LMY^{+}11]. commonly [Löp11]. commonness [CWA^{+}15]. Communicating [CBC^{+}16]. communities [ABC^{+}14, ADJ^{+}17, CVS^{+}13, DJH10, FAG10, FM10, For18a, FTTV^{+}18, GGZ14, GCL^{+}19, KCG18, KR16, LM13, LMSK12, LG12, LS18, LCR12, MSM^{+}17, MRR^{+}14, MBA^{+}19, NJM11, OMH12, OCB^{+}13, PES19, PL10b, PCZ^{+}16, RPM13, RSP^{+}17, SDKE18, TDH12, VG19, WJ15, ZLFI11, ZYKS15]. Community [HW12, LKDD15, SYL^{+}13, SJB^{+}16, TGKSJ18, Agu11, BND19, CRBC11, CHR^{+}13, CZCL18, Con14, CSFR15, CGM^{+}14, DA19, EAB11, FGPR10, Fuj16, GGZ^{+}16, GDTP13, GRB12, GCW^{+}16, GPM16c, HLL^{+}19, HKG^{+}17, IBD^{+}19, KvOP10, KMKL19, KNM^{+}10, LY16, LS18, LCP14, LCP15, MGGLBIBD13, MLMFN^{+}15, MDJ^{+}11, MJL^{+}13, MJL^{+}14, MYHN17, NGL13, PSB^{+}17, PES19, ROC^{+}13, SAR11, SRV^{+}12, SDKE18, SMG^{+}12, SGH16, SSM^{+}13, SMLH^{+}11, TKW17, wTA12, TM12, Vav10, VJR^{+}16, WR12, WP14, XZZ^{+}13, XTPT11]. Community-level [HW12, CHR^{+}13]. Comparative [BRPK14, MROGVR12, OBC^{+}21, RBM^{+}17, WPB14, Ano12t, FE14b, FKK^{+}18, KCF^{+}13, LYCZ13, MBD^{+}18, OCB^{+}13, YHZ19]. compare [PWBG16]. Comparing [IHSM12, CRD^{+}16, CKMD14, DCZ^{+}15, DSBS15, GPH17, GSSM13, HMS^{+}10, LBD10, LLT^{+}14, SVD^{+}13, THC18, TM14, TBM^{+}14, WB14, VBS14, YBA^{+}16]. Comparison [AKvO^{+}10, ALC19, BND19, BBH^{+}17, CBB^{+}10, HJWG12, HKVR13, KHKC^{+}18, KPS19, LTL^{+}13, MBD^{+}12, PES19, RvKM10, SFR^{+}17, VRA^{+}17, WBS13, ZLF^{+}11, ZJM^{+}11, ZZFH15, BDL^{+}14, FCV^{+}12, GGV11, HHD12, HB14, HKP10, IPSH14, MDL^{+}10, MRES17, OLG^{+}13, PSWW11, SBVA^{+}13, UFK^{+}10, ZPZ^{+}15]. comparisons [ZLG16]. Compartment [WMW^{+}12, CHvdW^{+}10, WPB14]. Compartment-based [WMW^{+}12]. compartmentalised [Mur13]. compensate [CJM^{+}12]. compensation [LMDG14, MTL^{+}10, Rai13]. compensatory [TBJ^{+}15, VEK12]. competing [ABC^{+}14, BYR18, BPM^{+}15, CN10b, SRV^{+}12, SFHK12, YHW19]. Competition [BRF^{+}11, MYHN17, NYH15, BM16, BRMW18, BMCG12, BEBA19, BKB10, BvBB10, BLAV18, BPN12, Che10, EGA12, FMW^{+}19, FC12, For18b, GvdWBS11, GGSP14, GRYM11, KSM^{+}18, LV17, LAK^{+}21, LLH^{+}13, MSG^{+}13, MPR13, MdlP15, MG18, MHG18, OFK^{+}12, PLB16, PDG^{+}14, RW17, Rin18, RE10, SFM^{+}18, wTA12, TYEH10, UDP17, VLF^{+}17, YLC^{+}18, dVLS11]. competitive [CLMA19, CRA^{+}11, GO15b, HBL18, KSS11, PGZ^{+}19, PL10b, RHS^{+}15, RHD^{+}16, ŠK14, ZW11]. competitively [VBPY11]. competitiveness [HGE^{+}10, KKBZ17]. competitor [MdlP15]. competitors [XCG^{+}16]. complacency [Pér18]. Complementarity [CPS^{+}19]. complementary [WSS^{+}16, GAB12]. complementation [MGB13]. complete [HSHF^{+}13, ZSKI17]. completely [BCJ^{+}12]. Complex [DSW^{+}17, FNR11, KGYG15, AD18, AOD^{+}13, BAB^{+}16, Ban11, BS17, CBC^{+}16, CPS^{+}14, CGdCCPCPM14, DWP11, FBdP^{+}16, Gar17, HCP10, IT17, KPM^{+}14, KPS17, LAD11, LS10, LS13, LDB14, LCDJ^{+}12, LM16, MTJF11a, MH10, ROG^{+}16, SWGC16, VASH^{+}16, WJL14, WLH^{+}10, YYL13, YZ16, YZ18, YXZ^{+}21, ZLF^{+}11]. Complexity [AHM^{+}16b, LVPF15, AA12, Ano17a, BS16a, DG16, GCA16, LMSK12, Mat17, MU10, MAMNL^{+}15, Mur12, NYIH17, Nie16, PdM11, PMBB^{+}16, SA13, VFB19, WHH^{+}18, ZYKS15]. Compliance [JZCL14]. component [BSH^{+}10, GKKW10]. components [BDTV11, GGVB11, Löp11, MGW^{+}13, SYL^{+}13, WAS^{+}16]. composite [DHHB15]. composition [Ant14, CTP^{+}17, GLRT16, Pen11, SGH16]. compositional [GGZ14, MWML13a, VPC11]. compositions [MAC14]. composting [BT16]. compound [HLS^{+}14, WSL17, ZHC^{+}16]. compounds [CLL14, LKKK^{+}14]. Comprehensive [SAI10, ZLC13, BRM^{+}14, KFvSH15, VVA^{+}15, WHH^{+}18, ZWG^{+}17]. Computation [vdVBJS15, CDK^{+}17, CT12, KPM^{+}14, LJLC15, LT15, ROC^{+}13, ZRC^{+}15]. Computational [FCA18, DHC14, DWNN10, DBP16, GB16c, SWZ^{+}12]. computationally [AADA14, CS18, WHC^{+}13]. compute [DB15]. Computer [LMSK12, AS11, BRW^{+}12, RGHS11, YHS^{+}11, YHK^{+}19]. Computing [DBSA16, CMSPJ11, MGSB11, ZB15]. concentration [ASL13, BLOJ^{+}19, CGG^{+}18, CN13, PD15]. Concentrations [GVV^{+}16, BGPJ10, CW17, FE14b, JNC^{+}11, KPSC10, MM16, NW18, ZJM^{+}13]. concept [GDBB19, JZCL14, LJAD17, MDKK15]. concepts [HM17, SFM^{+}18, VMG^{+}12, WAM13]. Conceptual [Ana17, EEFE^{+}11, GLC^{+}18, GGZ14, GJK^{+}15, KPT11, WS15]. concerning [OGOWW16]. concolor [CGV^{+}15]. Concurrent [CH12, GQRC^{+}17]. condition [PSG19, TGB^{+}17]. conditional [Loe12]. Conditions [San19, AHS14, AVL^{+}11, BMN11, BDG^{+}14, CVV^{+}13, CHR^{+}13, CTNH13, CCSF11, CN13, CBA10, GSPSP15, GWK^{+}12, GKW^{+}13, GJK^{+}15, HWE^{+}18, JL18, KD13, LMSK12, LKH^{+}18, LCY^{+}13, LWY^{+}11, yLWhMhG11, MBSW12, MTPC13, MTL^{+}10, MKTO19, RDR^{+}17, RP11, SHT^{+}21, SLB14, Tie16, UKI^{+}11, ZWW^{+}18, dBGS^{+}14]. conductance [GZK^{+}11, HSD^{+}15, MEC14]. conductivities [LPD14]. cone [CMT^{+}11]. cones [YY13]. Conference [Ano13u, LMY^{+}11, LYP^{+}11, LPA^{+}11, WTC^{+}18]. configuration [HSLC^{+}19].

Configurational [BCI^{+}17]. configurations [vSGR16]. Configuring [LM16]. conflict [CWM12, FSM^{+}19a]. conflicting [KNU^{+}15]. conflicts [DSBM^{+}18, MHS^{+}17, TAT^{+}16b]. confluences [JD18]. congeneric [DZS12]. conger [TSN15]. conifer [HAA^{+}21]. coniferous [FPG^{+}10, GZK^{+}11, MPL^{+}16, OKS^{+}11]. conifers [KHS^{+}19]. Conjoint [GZS^{+}18]. conjugation [CJC11]. Connected [KK11, BCV13]. connection [WSZT10]. Connectivity [LYC^{+}15b, ZZA^{+}18, AMM^{+}14, BH13b, BJJ12, BS10a, CXW19, CLDRJ18, CD15, CL12b, DMDJ13, DRS^{+}16, JM15, KBB^{+}11, KB18, KNPH^{+}17, LML17, MHG18, MRM^{+}17, PTML14, RBRK16, SSBJ12, SDEP18, TMPF18, TLW^{+}14, TDF^{+}19, URR^{+}17, VPC^{+}14, YJW15, dHR18, vSGR16]. conquer [LM16]. Consequence [RCP17]. Consequences [Ash10, HKM^{+}18, NBSGW19, YMCV21, BGGL16, BWML15, CLDRJ18, GSF^{+}16, GV15, LJ15a, LCD^{+}16, MSCM18, MAP^{+}13, MCPB15, PTL14, RHR^{+}15b, RHS^{+}15, RHD^{+}16, RDV^{+}17, SBCB13, SSS^{+}11, TGKSJ18]. Conservation [NCLP10, AR11, ASP^{+}15, ABC^{+}17, BFPR16, BEPC15, BDWP16, BM10, BvHR^{+}15, Cho13, CB19, CTC^{+}18, DUASLB^{+}12, DA17, DS12, DLW^{+}17, DG16, EMH10, ESP^{+}17, FN18, Flo18, FLA19, GWR11, GFW^{+}17, GBL^{+}13, HCK18, HCS^{+}10, HAOS17, HBT^{+}18, KBB^{+}11, LV17, LDM15a, LM12, LA11, LWNG13, LG17, LAEAG12, MFB17, MRP14, OOFVBC16, PK12, PSM^{+}12, RMR^{+}16, SDT10, TPG14, WSS^{+}16, WhFxGZ16, ZFT^{+}10]. consider [WAM13]. consideration [HGE^{+}10, Oga18, PGS16]. considerations [EEFE^{+}11, VGR^{+}11]. considered [DLL18, Kos17]. Considering [XYY^{+}16, LCFD19, TLLX19, ZML^{+}12]. considers [ECO^{+}14]. Consistency [GEPCB13]. conspecific [FMW^{+}19, TMPF18]. Constancy [SW19]. constant [CEC10, WB14]. constituent [FL11]. Constraining [HSK17, PSB^{+}17, LHR^{+}15, RKL^{+}15]. constraint [Log13a]. Constraints [OCD^{+}13, PSJ11, hCjHMW18, DDR^{+}18, ELPK12, FP17, GDBB19, LCR12, SHB11b, TCSG15, TSSL11]. construct [CN19]. constructal [KOBK13]. Constructed [NXZ18, HZT^{+}13, LL18b, MB17, OSS13, PLFO^{+}12, SWZ^{+}13]. Constructing [SSKJ18]. Construction [WPZ^{+}14, ZWW^{+}18, KFDL^{+}19, LYC^{+}15a, PFSF14, RZF^{+}13]. Consumer [dVLS11, CD16, CN18, KMS^{+}12, KMS^{+}13]. consumer-resource [CD16, CN18, KMS^{+}12, KMS^{+}13]. Consumption [Ben12, CD16, BWC17, FCZ13, FE14a, LLW^{+}15, LYCU14, MRM16, XHC^{+}15]. contact [HMPC11, SM10, TMPF18, WGD^{+}17]. contagious [SM10, SM12]. contaminant [CZW10, GPD^{+}13, LSR16, MGCL17, ZC11, ZWJ^{+}12]. contaminants [BFP^{+}18]. contaminated [TPL^{+}15]. contamination [TWL11, ZLLZ13]. contemporary [CF14]. content [LKP13]. conterminous [LXS^{+}16, SBF15]. context [Ash10, AFC12, CY13, CPJ^{+}13, GMC^{+}19, GEPCB13, KK16, LPT^{+}18, LKAY15, MMRS^{+}19, SSQ^{+}10, SLD^{+}19, TFH12, ZLLC13]. contiguous [SHC^{+}15]. Continent [LLZ^{+}16]. continental [BlMG^{+}11, CQS15, GC14, HBSF16, LBSB^{+}14, OC15, PFA^{+}16, RAF17, ZRCNLI^{+}15]. continental-scale [RAF17]. continual [BP15]. Continuous [BKV17, Sto10, BP16, BP17, BP19, CGAB17, HB13, KMMS17, MPR13, SDSH15, ZFG13]. continuous-time [CGAB17, HB13]. continuously [GS19a]. continuum [PES19]. contracts [LA11]. contrasted [BMCG12]. Contrasting [BBHP13, CVV^{+}13, GBP^{+}10, KYvdPS14, LBD10, PICR13]. contribute [DSJ19, EKB18, XM13]. contributes [YHW19]. Contribution [BPT^{+}19, TWP^{+}14, HHL17, IPSH14, JAH^{+}10, RHQ^{+}16, ZJL^{+}14]. Contributions [BCO17, LW17, Fat14, Fis21, GDTP13, GSSM13, LNB17]. Control [PF12, AD16, AdMC^{+}13, BCP15, BJK^{+}16, BJR16, BMP^{+}19, BNP10, CTB^{+}12, CHvdW^{+}10, CTC^{+}18, CESC19, ETG^{+}18, EEB^{+}17, FSM19b, GLS10a, GWHL15, GHS^{+}10, GDW12, JX15, JWLS17, LDM15a, LBR^{+}17, LDM15b, LCAJGFNS14, LZB17, MBLC14, MWY^{+}18, MS14, MLK17, Mur13, MS16, MFI13, NvGZ^{+}16, PHvB^{+}18, PKB^{+}18, PTML14, PFP^{+}16, Pec12, PvVC13, PMBB^{+}16, PME^{+}16, RJ11, RFV^{+}12, RPA12, STSA10, SRXP14, TV15, WPG15, WSW^{+}13, WPB14, YLW14, YWA^{+}17, YRGH16, dÁSMRT^{+}19]. controlled [LZ10, MKF^{+}11, RPGSCF16, WR13b]. controlling [CRAP13, HJS^{+}16, MS14, SM10]. Controls [PPO12, CHVH13, DBG14a, GZK^{+}11, HBA16b, HAA^{+}21, IN16, JSB^{+}14, KGYG15, LZLC13, LHP16]. controversial [ACE^{+}17]. convection [SSW^{+}18]. convection-dispersion [SSW^{+}18]. conventional [SBC13]. Convergence [Abe15, KA17, WR18, PMH17]. Conversion [ZZLZ14]. Converting [SBC13]. Convolution [CC14]. cool [FBA^{+}11]. Cooperation [BSC^{+}15b, BEBA19, MR10]. cooperative [TLP17]. coordinates [HHW11]. copepod [JSH15, RPM13]. copepods [BVJE18, DJMBD^{+}13]. Coping [LPF11]. copper [BRHD15, KSM^{+}18, RCST11]. coppice [PGGMMA19]. Coquito [ZPZ^{+}15]. Coral [LBE^{+}17, ABAG14, BMR^{+}18, BPN12, CCCP17, CTMS10, GBP^{+}16, GRYM11, GBR13b, HJ10b, JCKG15, KvOP10, KR16, MTJF11a, MTJF11b, MB14, MATS14, SM13, wTA12, VSF10, YMCV21, ZLFI11, MLAT^{+}18]. coralline [BPN12]. Corallium [GBP^{+}16]. corals [ENM11, TGM^{+}21, YSB11]. Corbicula [GCDA16, HFH^{+}19]. cordgrass [CZCL18]. core [BVGR^{+}17, VR12, DLH^{+}14]. Coregonus [MYJ^{+}14, VVH12]. cork [ABC^{+}17]. Corn [MR17, BB10, LYD^{+}17, MBD^{+}12, Tat16a, YC13]. Corophium [GCC^{+}10]. Correct [GMS11]. correlated [HM12]. Correlating [LML17]. correlation [LCM11]. Correlations [WFA10]. correlative [IHSM12]. corridor [LLL^{+}10]. corridors [DPBK10, DDBK15, DFK^{+}11, RH15, SGDP19, VPC^{+}14]. Corrigendum [ABAG16, BMF^{+}11a, CCW^{+}15a, Con12, CR12b, DYMYGO19a, EH19, HM19, HLLO13a, KS17a, mLcLJ16, Lin16, MJL^{+}14, SMT19b, Str15a, WGA^{+}12, WLS^{+}11, ZLJM17b]. Corylus [BOF^{+}16]. Cosmic [CCC11, JC11a]. CoSMo [Con14]. cost [CEWW18, CSF11, CMD12, CPAG15, LLL^{+}10]. cost-effective [CPAG15]. cost/benefit [CMD12]. Costa [FAC^{+}19a]. costal [GSJ^{+}10]. costs [BMU10, BHPF14, EZ14, LZB17, vDHH^{+}13]. Cotter [SMC11]. Cotton [LBC14, AS11, FGH^{+}10, PVB16, ZML^{+}12]. Cottus [KDMPC10]. Cotylorhiza [RPA12]. cougar [Gli14]. cougars [LN11, LN16]. Could [PLS^{+}16, FMBD15]. count [DAP18, FDP^{+}10, ZJB19]. counterpart [PCB11]. countries [KBMM15, SAKB^{+}10, ZWG^{+}17]. County [GPM^{+}16a, YLW^{+}11, ZZLZ14]. Coupled [CTMS10, AHM^{+}16a, An12, ALC12, CAL13, CSS^{+}17, DJT^{+}19, FRM^{+}19, GGZ^{+}16, GHS^{+}10, GCM^{+}13, HBA16a, HBA16b, HG17, Kei12, KCB^{+}17, KVS17, KMS^{+}12, KMS^{+}13, LZZ^{+}14, LB18a, MSGG14, PD15, PHP15, RBGB16, WALCY12, WR18, YY11, vSGR16]. Coupling [GHWG^{+}14, IFC14, KSAS12, KWGW19, LKLK15, MLG18a, MUP^{+}13, PRHH13, PD12, YDZ^{+}21, AL16, CCL14, CKS^{+}17, GS10, HPRNP12, LDS14, Nak17, NP18, TVT12, TZS^{+}14, YLL^{+}12, ZWG^{+}17, MHG18]. CoupModel [HJS^{+}16]. course [GBB^{+}11, PF18, WJK11]. covariance [WBC^{+}12]. Cover [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18x, ADF^{+}17, BFKS16, CKMD14, DK17, FLY13, FCFT16, FC11, Jør14a, LLZ^{+}16, LWL^{+}17b, LDP14, LVZ^{+}11, LCDJ^{+}12, TSB^{+}18, XLR11, dFMdSA18, dJGM^{+}18]. Coverage [LCAJGFNS14, FBDZ12]. cow [GBR^{+}13a, GBW^{+}15]. cowberry [LM19, Log19a]. cows [Mat17]. CQESTR [GE17]. crab [MTZ^{+}10, PPDF14]. crafting [RIMPBT19]. Crassostrea [GBB^{+}11, LLC^{+}17]. crater [IBD^{+}19]. crayfish [MJB^{+}11a, MJK^{+}17, ZLY16]. CRC [Jör11a]. creaons [AG14]. creates [WYS12]. Creating [LF11, HPC13a]. creator [KKBZ17]. credit [OE12]. Creek [NFU19]. Crepidula [MG18, MHG18]. crickets [BG11]. crisis [BU11b, SS10, YL17]. criteria [MDG10, RLB^{+}12, SHX^{+}12, VFFP16, ZWW^{+}18]. criterion [MU10, ZZT13]. Critical [SEHD18, Bae18, Flo14, JSL^{+}10, NKSA13, PMBM11, RBRK16, Sci11b, SWZ^{+}12, SA16, SW15, SB15, XYTS19, ZSKI17]. critically [HBT^{+}18, MFB17]. critique [HBT^{+}18, Pat10, PGW19]. croaker [DMR13]. Crocosphaera [GRD^{+}14]. crop [AVL^{+}11, AWL^{+}13, BPM^{+}17, CLLB12, CC14, CB19, CDT10, Con14, CBA16, DBJ13, FYWA12, GDT^{+}14, GCCB18, GCK^{+}15, GBP^{+}10, GLS^{+}17, HGE^{+}10, KCLL14, LRM^{+}14, LWKS10, LZZ^{+}14, LBC14, MBLC14, MGB13, MBA^{+}19, MJGC13, MWD^{+}15, Nei11, NBK^{+}11, Och10, OQH^{+}13, PHvB^{+}18, PRP17, Pie10b, Pie10a, PANL11, PVB16, SX11, SMS^{+}13, SBRD10, Tat16a, TAT^{+}16b, THBF^{+}17, YWH15, YDD^{+}13, ZCS13, ZTL^{+}17, SLZ^{+}18]. crop-disease [PHvB^{+}18]. crop-wild [GDT^{+}14]. cropland [LLT^{+}14, LLZ^{+}16, LLT^{+}17, QlSeZ^{+}10]. croplands [KHW^{+}17]. cropped [GNB^{+}14]. cropping [HGSJ15, KKBZ17, SVL12, WTL17]. crops [HQF^{+}17, MCWG21, STSA10, TWL11]. Cross [LBR16, Pie10a, ZLG16, LJ10, MAMNL^{+}15, SLB^{+}18, TZS^{+}14]. cross-farm [LJ10]. Cross-lake [ZLG16]. Cross-row [Pie10a]. Cross-scale [LBR16, TZS^{+}14]. cross-temporal [MAMNL^{+}15]. cross-wavelet [SLB^{+}18]. Crossing [dS18, DTS^{+}12]. Crossing-effect [dS18]. crossings [PGATS13]. Crow [KRE^{+}18]. Crown [UDP17, BLB^{+}17, VALC12, CDK^{+}17]. Crown-of-Thorns [CDK^{+}17]. crowned [BFJZ^{+}11]. crowns [CLBSG11, EKB18]. crucial [BBJV^{+}11]. crude [BPU11]. crust [SBDA16]. Crustacean [COK^{+}11]. crustal [DB15]. crustose [BPN12]. cryptic [NBSGW19]. crystallinus [SKPHW16]. CSR [BBL11b]. CTEM [HBA16a, HAA^{+}21]. ctenophores [AALP18]. cuckoo [JHHS17]. Cucurbita [GLSA^{+}11]. Cuenca [HGBA^{+}13]. Culex [JSB^{+}14]. cultivar [ZBS14]. cultivars [KYvdPS14]. cultivating [HWH^{+}18]. cultivation [BAP18, FPO18, GCR18, MJF19, OQH^{+}13]. cultural [MBLC14]. culturally [GWR11]. culture [Abe10, Abe13, CHCC15, RRHH10, ZZL^{+}19]. culvert [FN18]. Cumulative [TBJ^{+}15, CSS^{+}16, Hoa11, LYSW15, SHX^{+}12, TdS12, XPY^{+}14, YKCN10]. cunicularia [EGMW12]. Cunninghamia [WZD^{+}14, ZXD^{+}13]. Current [SBC13, EEFE^{+}11, MNK^{+}21, MK19, SvBH^{+}14, SLSS15, SA16, ZLFI11, ZZZW15, GEZ^{+}13, HKM^{+}18, KKH^{+}17, KEM^{+}16, MKL14, RBGB16]. current-driven [ZLFI11]. currents [HRWT12, MLW^{+}15]. curve [CBCB17, CL10, CLW^{+}19, MLK^{+}11]. curves [HM10]. cusk [CRQW18]. cusk-eel [CRQW18]. customary [GWR11]. cutaneous [RMP^{+}18]. cutting [CAC^{+}16]. cyanobacteria [CRB16, Hel10, HB10b, HB10a, MHC16, ZRC^{+}15]. cyanobacterial [SAI10, SLW^{+}14]. cyanobacterium [GRD^{+}14, GMvV17, LZYH19]. cybernetics [Nie16]. Cycle [JMdW^{+}10, UAB^{+}11, Abe13, BH13b, BFM11, BRKP16, DLL^{+}19, DLZ^{+}17, DLD^{+}11, DRS^{+}16, GRRU15, qHLlZ^{+}16, HN14, HSHF^{+}13, HPGM18, HB10a, HB15, HBB15, JWX18, KK12b, KBMM15, LJB12, MJP^{+}12, MBRR13, MRG13b, Nak17, NP18, PMH17, RSCS19, RPGSCF16, TPO^{+}18, WGE^{+}19, Wol11, ZHCX15]. cycles [AL16, BB10, CLL14, DCMK11, FS14b, HPG^{+}10, KVS17, PP16, TNTH11, VWTA16, WR18, ZPZ^{+}15, ZLLC13]. Cycling [ZB15, CT11, CGM^{+}14, HB14, JN15b, JDGJ13, KS18, KWS^{+}13, KJMA11, LLZ^{+}15, LZL^{+}18, MJC^{+}11, SSF14, Til21, TRE^{+}12, WB17, WJF^{+}11]. cyclops [KR14]. cygnus [LMBL12]. cypress [UKI^{+}11, Oga11]. Cyprinus [CEWW18]. Cystoseira [FAC^{+}19b]. Czech [HV15].

D [CPS^{+}14, Con12, IFC14, Jør14b, MBUM12, NGR^{+}11, VLF^{+}17, AHNK13, BVGR^{+}17, BCB^{+}11, Con11, DBP16, KPS17, LMT13, MSGG14, MRB14, MBUM12, MJGC13, PMM11, SRHL11, SMH16]. D-CMCC [CPS^{+}14]. D. [Pat10]. daily [BSC^{+}15a, Sim11, SPG11, ZZB^{+}18]. dairy [GBR^{+}13a, GBW^{+}15, SFvB10]. Dajiuhu [HZT^{+}13]. Dall [FYR^{+}18]. dam [BCTL16, FN18, ZVS18]. damage [BPS^{+}11, BPC^{+}17, CBP^{+}17, DYMYGO19a, DYMYGO19b, KH15, LJBS12, SRT16]. Dams [BAAB18, dSGGBW12]. Dana [MRCW10]. Danajon [BWRA16]. Danaus [GPZB18]. dangers [BS16a]. Danio [HTL^{+}14, KNU^{+}15, LLC11]. Danish [JN15b, KRCVF16, KvTL^{+}18, NNST^{+}14, WØ15]. Daoxiang [WX11]. Daphnia [ECO^{+}14, PLS^{+}10]. Daphnia-algal [PLS^{+}10]. daphnid [LHO^{+}16]. daphnid-algae [LHO^{+}16]. Darjeeling [RBRK16, SGP18]. dark [FS14b]. darter [MGWW13]. Darwin [Kat11]. Data [CC14, KKI^{+}14, LST15, AHS^{+}17, AB10, AB13, ASV19, ACC11, BLFS18, BG19, BMSLC16, BBL^{+}17, BNC^{+}18, BGMJ^{+}12, BOW^{+}12, BAJ^{+}11, CGG^{+}18, CGAB17, CKMD14, CMEP15, CMV^{+}16, CLNN14, CGP12, DSK^{+}11, DOB^{+}15, DDB^{+}12, DAP18, FDP^{+}10, FKH^{+}15, FN15, FBH^{+}14, FYWA12, FBA^{+}11, GCA16, GCIB10, HMD^{+}16, HPJ15, HVL^{+}19, HSK17, HMS^{+}11, HS11, HCL^{+}14, HlLB12, HCP10, HJ10b, HRWT12, HhLY^{+}19, JZ12, JG10, JSCS19, JJYM19, JBCÑ11, JGX13, JAKJ10, KMRW16, KSZW16, KKL^{+}12, KHS^{+}19, KNPH^{+}17, KPS19, LJLC15, LLB^{+}19, LBSB^{+}14, LPF11, LBH^{+}15, LBE^{+}13, LH15, LF11, LCL^{+}13, LJ15c, LVZ^{+}11, LW17, MAO19, MRCW10, MHTC15, MU17, MCBM12, MVC^{+}17, MAC14, dMHGA10, MK12, MRWN19, MSR^{+}14, MAMNL^{+}15, MDV^{+}12, MMMGLN^{+}18, MYJ^{+}14, PMBM11, PJ17, Pat14b, Pér18, PGZ^{+}19, PÓF^{+}18, Pog11, PSS16, ROG^{+}16, RHZ^{+}18]. data [RWVK11, RUSL19, RMHE15, RCBPMN10, RRvdG18, RKNO17, RPGSCF16, RCS^{+}19, RWP11, SZV14, SHT^{+}21, SMI^{+}19, SHM^{+}14, SVH^{+}19, SQWS19, SSKJ18, SFL^{+}19, SHB11b, SCG^{+}13, SP12, TRP12, Tul14, TAGC^{+}16, VRM^{+}10, VALC12, VB10, WB10, WTWS19, WLC^{+}19, WBG^{+}14, WC13c, XGH^{+}17, YAGMM^{+}14, YSB11, ZSS17, ZLZ^{+}17a, ZKSF17, ZJB19]. data-derived [LF11]. data-driven [DDB^{+}12, MDV^{+}12, MMMGLN^{+}18, SZV14, SQWS19]. data-poor [DOB^{+}15, RKNO17]. database [ARTBB14, BSC^{+}15b, HhLY^{+}19]. dataset [CRD^{+}16, TNHB12]. datasets [GCDA16, SBC13, Tul14]. dates [LCFD19]. davidii [PKD11]. Day [MSB^{+}17]. DAYCENT [COPP13, RKL^{+}15]. DE-CA [dWcWbZ15]. de-watered [CCL^{+}13]. deal [WHH^{+}18]. Dealing [GRRU15, MMCBP19, WRM^{+}11]. DeAngelis [GB16b]. death [JDZC14]. Deaths [VSF10]. DEB [GMvV17]. debris [HHUD14, RWFD13]. debt [Lab11]. decadal [GVGZ^{+}21, KD13, SBC^{+}11, WJvdL^{+}13]. decades [GDC^{+}19, SW19, ZMS19]. decay [HSK17, RWFD13]. December [Ano10v, Ano11-35, Ano14f, Ano15b, Ano16q, Ano16m, Ano17x, Ano17-37, Ano18-34, Ano18-48, Ano19a, Ano19e, Ano12-36, Ano13-36]. deciduous [CPS^{+}14, SMSAN^{+}14, WJvdL^{+}13]. decimate [POS^{+}19]. Decision [Gia14, LKAL^{+}13, BGG16a, BFNH14, BPT^{+}19, CVV^{+}13, CGL15, DWD19, DTS^{+}12, EAJG16, GCCJ12, GAF^{+}14, KRE^{+}18, KK15a, LWKS10, LAEAG12, NCLP10, ST13, SHX^{+}12, SWGC16, TAT^{+}16b, WC13c, vGKB10]. decision-making [CGL15, EAJG16, GCCJ12, NCLP10, SHX^{+}12, SWGC16, WC13c]. decisions [An12, ALC12, CAMNN16, JTM^{+}17, SM11, VSPHS15, WALCY12, ZKR^{+}13]. declaring [ALG16]. decline [ENC^{+}15, KPTC14, MKTO19, MHRB19, MDVG^{+}15, NJCF11, Oga18, ODK16]. declines [HKM^{+}18]. decomposer [Kei12]. decomposing [Hoa11, LKKK^{+}14]. decomposition [BdCS11, DFRT14, GZVU14, HKVR13, Lec17, LCY^{+}13, LW14, LKP13, MBG^{+}10, SKTG11, TLRL11, ZTJ^{+}10, ZZY^{+}13, ZTL^{+}17, GAC^{+}14]. Deconstructing [GB16c]. Decrease [MJB^{+}11a, GAG^{+}12]. decreased [HLZ^{+}18, RGH^{+}13]. decreases [Mat13, RWVK11]. Decreasing [Flo18, LDP14, PLS^{+}10]. dedicated [GB16b]. deep [FRM^{+}19, FRP18, GB16a, KJVL^{+}17, LZF^{+}14, MSGG14, RIMPBT19, SAI10]. Deepwater [BPC^{+}17, CBP^{+}17]. deer [HMPC11, LVLS14, LAEAG12, MWML13b, PMP^{+}13, RFV^{+}12, WTG^{+}16, WR13b]. defence [JL18]. defense [LB15]. deficit [KS17a, KS17b]. define [MCB14]. defined [CDL^{+}11, MMD^{+}17]. Defining [JDT^{+}16]. definition [MWB^{+}10, YqWB^{+}15]. Definitions [RRH^{+}19]. defoliation [BPS^{+}11, CAC^{+}16]. deforestation [dSD14, dSD18]. defragmenting [MAB19]. degradation [BVGR^{+}17, CBGK15, DFW^{+}18, HBS16b, MMMGLN^{+}18, OES16, TAGC^{+}16, ZC11]. degraded [HRSKW15]. degree [CRBC11, LS13]. Degrees [RRH^{+}19, CWA^{+}15]. Deiva [VB10]. del [FAC^{+}19a]. Delaware [FMLM11]. delay [BBB^{+}17, ZW12]. delayed [LL17, NYF16]. delimitation [SS10]. Delineating [DNFG18, GFW^{+}17, VR12]. delineatus [BGC10]. Delphinapterus [MDVG^{+}15]. Delta [LMCA^{+}12, XLS11, XCL^{+}11, LJ15c, AH17, MLS^{+}11, TJYZ11, Yan11]. deltaic [DJS10]. demand [AS14, MH13, SGAS14]. demand-of-energy [AS14]. demersal [NMA^{+}19, RMR^{+}16]. DemGenTrout [FB13]. demo [PP12]. demo-genetic [PP12]. demogenetic [BAAB18, NMT^{+}19]. demogenetics [FB13]. Demographic [CMA^{+}13, ECO^{+}14, Fuj16, LJ15b, NBSGW19, VSF10]. demography [TNTH11]. demonstrated [CRO13, HD14]. demonstrates [WJvdL^{+}13]. Demonstration [FSB13, SSN^{+}18]. DEMs [MLG^{+}18b]. dendritic [PW10a]. dendrobatidis [ACCM16]. Dendroctonus [PB14]. dengue [EPA^{+}10a, JVY19, MD16, ZYXT18]. denitrification [IMW^{+}17, KGYG15, SGKR10, GAC^{+}14]. Denmark [CVPR^{+}16]. dens [LPSC16]. densities [BM11]. Density [KUI^{+}18, LB17, Rai13, WUB15, AGJ^{+}12, Agu11, ACDC17, AOW12, AOW13, BSW18, BM11, BLB^{+}17, CVS^{+}13, CGS^{+}17, FT16, HMPC11, HFSL14, HKP10, IWW12, JSB^{+}14, JDZC14, KSS11, KSM^{+}18, Kos17, KUK11, LSR13, MLW^{+}15, MHTC15, MIC14, MCA18, NZL13, NYF16, PC15, PKMH16, PB14, RHO11, RW10, SC18, SK10, SSQ^{+}10, SED15, SU15, SKPHW16, SW15, TMPF18, TQZ12, WGTH16, WFA^{+}13, YZ16, ZSSR18]. density-based [BLB^{+}17]. density-dependence [ZSSR18]. Density-dependent [KUI^{+}18, WUB15, HFSL14, HKP10, JSB^{+}14, MIC14, PKMH16, PB14, RHO11, RW10, SU15, SKPHW16]. density-independent [SK10]. depend [MPG18, RCSS13]. Dependence [FSB13, AGJ^{+}12, ABB^{+}14, BSW18, BPN12, CGdCCPCPM14, MCA18, Nak17, QCLW18, SC18, ZSSR18]. dependencies [Ju19, PGS16]. dependent [BEPC15, DFF17, FC11, FML^{+}17, GHH^{+}15, HFSL14, HD16, HKP10, Jen12, JSB^{+}14, KSS11, KUI^{+}18, KDMPC10, KMS^{+}12, KMS^{+}13, LDB14, MBBA17, MSCM18, MIC14, NYF16, PKMH16, PB14, PDG^{+}14, Rai13, RHO11, RW10, SED15, SBA^{+}13, SU15, SKPHW16, WBC^{+}12, WUB15, WCG^{+}11, YL18]. depending [BDG^{+}14]. depends [HBL18, Log19b]. depensation [CAL13]. depleted [BWML15, ZCR16]. depletion [BS13, BKN21]. deployment [JMW11]. deposition [CL10, GBG^{+}15, LY16, TRE^{+}12]. depression [LBW18, YY13]. depth [MFS18, NMA^{+}19, PBFS14, WFN^{+}16]. depth-based [NMA^{+}19]. depth-resolved [MFS18]. Derivation [HLL^{+}19, WSSM15]. derived [AK17, GLC^{+}15, HJS^{+}16, LF11, MAO19, MHMRAHHS19, NMdR16, PLR^{+}13, SBVA^{+}13, ZJL^{+}14]. Deriving [LSS^{+}15, SHBG10, SWGC16]. describe [AMF12, DFF17, GMvV17, MHMRAHHS19, MCPB13, RSCS19, TF13, WXB10]. described [BO14]. Describing [BC12, YL18, FPG11, TMC^{+}16, VRL18]. Description [KSAS12, SM13, AHS14, BP17, BKB10, GBR^{+}13a, IMW^{+}17, KS17a, KS17b, KCB^{+}17, RSCS19, WWYO18, XSN^{+}18, ZLF^{+}15, ZLZ17c]. Descriptive [TS11, VAK^{+}11]. descriptors [BGA17]. desert [LLC^{+}21, SYY^{+}19]. desertification [CLMA19, CpW16]. desiccation [ALeSD17]. Design [LYC^{+}15b, AS11, hCjHMW18, CR11, CR12b, DÖWB12, KRE^{+}18, Ken19, KCP^{+}15, KBMM15, LBR^{+}17, LLL17, MRP14, SR14, SHS11, TB12b, YHS^{+}11, YZLX14]. Designing [GOSH^{+}12, HCE^{+}19, LYC^{+}15a]. designs [FMF12, WN17]. desman [BRV11]. despite [BAJ^{+}11]. destruction [PL10b, dSJHG10]. detail [AVL^{+}11]. detailed [GMTC^{+}17]. detectability [LW17]. detected [VRA^{+}17]. Detecting [LW17, MLK^{+}11, OGG^{+}16, PVB16, RWS^{+}15, San11]. Detection [HGPF15, BMP^{+}19, BOW^{+}12, GMK^{+}19, HTB13, HTB17, KKP^{+}11, MdMS^{+}19, TB12b, VB10, WSC^{+}19, ZR19]. detection-censored [WSC^{+}19]. detections [FYR^{+}18]. determinant [BND19]. determinants [VEK12]. Determination [KHJ^{+}19, LCCL15, San11, FML^{+}17, KUK11, LYCZ11]. determine [BSHG16, CD17, CMGF14, FSN^{+}15, Fis21, GO15b, GJK^{+}15, LPD14, PTL14, VEMS12]. determined [GK14]. determines [SXNL18, SSR16, WYS12]. Determining [IKL14, KvOP10, PSWW11, ROC^{+}13, ZLLZ13]. Deterministic [F.11, HBT^{+}18, DR17, FM10, MYHN17]. deterrent [ZHV14]. detoxification [GBB^{+}11]. detrital [DVD^{+}14]. Detritus [LW14]. devaluation [GKL18]. Devaluing [LR16]. develop [FCA18, HLS^{+}14]. developed [BBH^{+}17, JN15b]. Developing [DLL18, KEM^{+}16, LHB^{+}18, PWK^{+}17, SDSH15, SLRC19, SB15, BFR17, KBMM15, LHZN11, Sol19, ZLB14]. Development [AAJB21, AD12, BWC17, BKM^{+}18, BOF^{+}16, CAY^{+}10, DWD19, DAW^{+}10, DBJ13, GJM^{+}13, GBW^{+}15, HBH^{+}12, JMdW^{+}10, KK13, KWS^{+}13, LAD11, LWKS10, LkSXhX17, MKSL^{+}17, MJLC14, MBA^{+}19, PGGMMA19, RR10, SHT^{+}21, SAR11, THO10b, VSB13, WPZ^{+}11, WXB10, YY11, ZLZS14, AK17, ALeSD17, AOD^{+}13, BBA12, BHC^{+}15, BFM11, CHPS17, CLLB12, CC10, DLZ^{+}17, Els10, GZK^{+}11, GHH^{+}15, GAF^{+}14, GB15, HS13a, HSVL12, HOŠ^{+}12, JZG10, LBA14, LBA15, LWL^{+}17a, LYCU14, LLP17, LBGM19, MDW^{+}12, MVC11, MGW^{+}21, Nak17, NvGZ^{+}16, Oga11, OKS^{+}11, RML^{+}15, RHB14, RCC^{+}11, SFC^{+}12, TCSG15, TVF^{+}10, TSSL11, WSW^{+}13, YYQ11, YLHZ14, YXD^{+}18, ZGJ10, ZYFL10, ZLC13, ZWG^{+}13, ZLF14, ZFZ^{+}14]. developmental [NU11]. developments [EEFE^{+}11, GPM16c, Jør10b, LB17, VASH^{+}16]. Deviations [ZNS^{+}18]. Diadromous [RDD^{+}14]. diagenesis [ZSJ15]. diagnostic [SNS^{+}12]. diagnostics [DMDJ13, DAP18, Lin10, Lin16]. diagram [LT15, SvdR10]. Diameter [RLS11, WK11]. diamondback [SFHK12]. Dian [XWW^{+}11]. Dianchi [SLW^{+}14]. diapause [MPGCS14]. diatom [CWLC13, HB15, KNM^{+}10, VRL14]. diatoms [ZYZ17]. diazotrophic [GRD^{+}14]. Dicentrarchus [PAGCV^{+}10]. dichotomous [GTM^{+}19]. dictate [FE14b]. did [Sci11b]. dieback [HS11, VAO^{+}17]. Diel [TPO^{+}18, BVJE18, HBB15]. diet [DGKSM13, MAC14, PLR^{+}13, SWS16, ZAW^{+}18]. differ [FML^{+}17]. differences [AR11, DSJ19, DDR^{+}14, KYvdPS14, SKL^{+}13, SM14, YKZZ14]. Different [VMSM10, AHS14, AKvO^{+}10, AKLB19, BCP15, BMN11, BC13, BKMK10, CNH^{+}11, DAA^{+}12, FCV^{+}12, GHWG^{+}14, GCDA16, GSPSP15, GWF^{+}18, HHvO^{+}17, HWL17, IKL13, JDT^{+}17, JDP^{+}12, KPLS15, KLS^{+}14, LS10, LHR^{+}15, LL18b, MPR13, MRM^{+}17, NYIH17, OCB^{+}13, PDV17, RBM^{+}17, RLB^{+}12, SZ14, SKMB11, SXNL18, SWR^{+}16, TDH12, Tie16, dHR18]. differential [BBB^{+}17, MMC11]. differentially [BRMW18, WRMB18]. differentiation [GGZ14, qHLlZ^{+}16, ZWG^{+}17, SJP14]. diffuse [PMBM11, RLHD11]. diffusion [BPK^{+}16, BJR16, BRF^{+}11, EGA12, HYG14, PvVC13, SSW^{+}18, TB11, TB12a]. diffusive [AMF12]. diflubenzuron [MHRB19]. digesta [GBW^{+}15]. digital [YZLX14]. digraphs [Pat15]. dilemma [DML15, RUSL19, RHS^{+}15, RHD^{+}16]. dilution [CHZ18]. dimensional [AHM^{+}16a, BLH10, BKB10, DBK^{+}13, FRM^{+}19, HFB^{+}11, KB11, LkSXhX17, LZYH19, MBSF14, MH10, PDAL14, PBK14, SELB10, WPG15, WQD^{+}13, XCR^{+}13, YWY10, YZG16, ZSJ15]. dimensions [CXL^{+}19, Flo18, LB18b, SG11b]. dimming [GLS^{+}17]. dinoflagellate [HHC15]. dinoflagellates [KTK^{+}15, KDT^{+}17, ZYZ17]. dioecious [MRB14]. dioxide [ACN14, GHP12, KRL^{+}16, KVS17, MKF^{+}11, MPM^{+}15]. diploproa [GC11]. Diptera [KTV^{+}15, LIM^{+}16]. Direct [LMDG14, BRB14, DPBS10, FG11, LCR12, MDH15, RSP^{+}17, ZCSL18]. directed [DDB^{+}12, FMA^{+}12]. directing [YY13]. direction [CD17, KK12a]. Directional [SG16, SMKM19]. directions [FCLFB18, GPM16c, KR10b, LTL^{+}19]. Directive [BMU10]. dis [RH17]. dis- [RH17]. disaster [WLZC11]. disasters [HYG14]. discards [MMD^{+}17]. Discerning [ODK16]. discharge [ER10, JCKG15, ZZZW15]. disciplinary [LP17]. discors [KK12a]. Discovering [HSCS14, TPL^{+}15]. Discovery [Jør14a]. discrepancy [PBP^{+}15]. Discrete [MPR13, BP19, CC10, dSCdA19, EEK^{+}18, KMMS17, Log16, PES19, ZJB19]. discrete-structured [Log16]. discrete-time [KMMS17]. Discretizing [RRvdG18]. Discriminating [CAMNN16]. discrimination [SVH^{+}19]. Discussion [CXL^{+}19, SW11]. discussions [LMY^{+}11, PL10b]. disease [BJK^{+}16, BG15, BPK^{+}16, BSCG10, DVL11, HMPC11, HM18a, HM19, HGSJ15, LBR^{+}17, LBR16, MLK17, Mur13, NPJ^{+}17, NTK14, OFGH^{+}19, PHvB^{+}18, PSBH16, PKW^{+}10, PMP^{+}13, SM12, TF13, WGSR19, YLW14, YSB11]. diseases [STSA10, SM10, SDS^{+}21]. disentangle [MYHN17, MDH15, VHDM17]. Disentangling [CHNdÁP18, EAB11, PNT^{+}19]. Disequilibrium [Loe18]. Dispersal [BSHG16, BRF^{+}11, CH19, FC12, KBB^{+}11, LCC^{+}14, SSR16, AAME11, ABGB15, BFJ^{+}11, BKV17, BEPC15, BBGJ13, BG11, CMM^{+}10, CHNdÁP18, CH12, Che15b, CV19, CDG15, DWNN10, DCCS17, DDB^{+}12, FDP^{+}10, FMA^{+}12, FPG11, GPJL15, GVGZ^{+}21, GPTR13, GTM^{+}19, HRWT12, Joh10, KSI14, LM13, LLL^{+}10, LLH^{+}13, MPKC12, MBG17, MdlP15, MLMH18, MMPG16, NLMdL10, NPJ^{+}17, PGATS13, PKMH16, PSWW11, PvVC15, PB14, RHW17, RHO11, RRJVA13, RHD^{+}13, SSBJ12, STSA10, SBM^{+}11, Slo11, SPWR17, SMKM19, SRM11, SGDP19, Str15a, Str15b, TKN14, TDF^{+}19, VF16, VCE17, VAO^{+}17, YDT^{+}13, ZW11]. Dispersal-mediated [FC12]. dispersal/colonization [Str15a, Str15b]. dispersed [BFNH14]. dispersion [GJM^{+}13, OAPP12, RAM10, RPC^{+}17, SSW^{+}18, WCZ11, ZC11, ZWJ^{+}12]. displacement [FC12]. disrupt [ZHV14]. Disruption [KL12, MBM^{+}18]. Dissecting [TNTH11]. dissimilarity [BS19]. dissipation [MPT^{+}11, YFL^{+}18]. dissipative [HWH^{+}18]. Dissolved [JNC^{+}11, NFU19, ASL13, BA14a, CGdCCPCPM14, HN14, KH11, LKKK^{+}14, MH13, MGW^{+}13, OSS13, PJ16, RRR^{+}10, YFL^{+}18, ZJM^{+}13]. dissolved-oxygen [MGW^{+}13]. disstria [BFA11]. distance [CGS^{+}17, GPJL15, GPTR13, KB18, KK12a, MBG17, PWGS^{+}17, SSR16, Str15a, Str15b, THJ^{+}12]. distance-based [KB18]. distance-invariant [THJ^{+}12]. distances [KRL11, RDJ19]. distemper [BG15]. distilled [Jen12]. distinct [SM12]. Distributed [GCK^{+}15, AADA14, AD12, GCB^{+}11, KW18, LKMCW14, MHG18, SEHD18, WPB14]. Distribution [BRL12, OT15, RHR^{+}15a, RDD^{+}14, AMRH19, AFC12, BCC^{+}11, BEZ19, BGA17, BS15, BTS19, BP16, BP17, BBJV^{+}11, BGD^{+}16, BS16a, BGG16a, BGMJ^{+}12, BPM^{+}15, BG11, COBF13, CKMD14, CMEP15, CPAG15, CD15, DCMS18, DKJH13, DLvGGNC16, EZ14, FLX^{+}14, FMW^{+}19, FJH15, FCLFB18, FAPM12, FCA18, FAT^{+}15b, FKK^{+}18, GCDA16, GGV10, GGV13, GA15, GdSHA10, GMLRF10, GMC^{+}19, GC14, GJK^{+}15, GLY^{+}15, HSLC^{+}19, HMD^{+}16, HFH^{+}19, HHD12, HM10, HCM^{+}18, HJN12, HM17, IBH^{+}15, JVY19, JTM^{+}17, JCKG15, JGX13, KTK^{+}15, KCG18, KMP14, KSG^{+}14, LMR16, LBE^{+}13, LSJG17, LHW^{+}15, LPT^{+}18, LDM15b, LDC^{+}16, LCS11, LWNG13, LRB12, MAO19, MU17, MMCBP19, MMS12, MCS^{+}18, MHMRAHHS19, MUP^{+}13, MDH15, MK19, MPP^{+}15, MMC^{+}19, MAMNL^{+}15, MHY^{+}15, MSO10, MDG10, NP15, OOFVBC16, PPZ^{+}17, PSG19, PMB^{+}19, PK14, PMF^{+}13, Pie10a, PGS16]. distribution [PLHG14, PMS11, PBK14, RWVK11, RvKM10, RRJVA13, RHR^{+}15b, RDR^{+}17, RR10, RCS^{+}19, SZSV12a, SZSV12b, SBM^{+}12, SGA^{+}16, SAI10, SGM^{+}14, SSKJ18, SKL^{+}13, SP10, SP11, SSM^{+}13, SHS11, SLD^{+}19, TAC^{+}14, TJYZ11, TQZ12, Tul14, VVH12, VZF^{+}11, WJS^{+}18, WBB^{+}21, WXB10, WKF^{+}17, WZL^{+}16, XCR^{+}13, ZSG12, dSD14]. distribution-based [KCG18]. distributional [PC16]. distributions [AG11, AFC11, BBL11a, CDR^{+}13, CBR^{+}16, CS17, CPS^{+}19, CDM13, FMF12, FPG11, FBKS17, GCA16, GGV11, GVV^{+}16, GSSM13, HEJP17, JDP^{+}12, KFWM19, KPK^{+}17, MWK19, MSV12, QDH^{+}10, RvS16, RKS^{+}10, RVNB11, SW19, SJB^{+}16, TCSG15, VLF^{+}17, ZJB19]. District [YLHZ14, KUK11, TNA^{+}17]. Disturbance [GO15b, PBG^{+}17, BFJZ^{+}11, BSS^{+}14, CNH^{+}11, FBDZ12, GHST19, Gat18, HMS^{+}10, HOKH18, JGB15, KK13, KMF^{+}21, LB11, LKG^{+}17, LLH^{+}17, LVZ^{+}11, MSCM18, PD12, PDD^{+}18, PNHL14, SHB^{+}11a, SM13, XCG^{+}16, YHW19, dSJHG10]. Disturbance-grazer-vegetation [PBG^{+}17]. disturbances [DVF^{+}19, FTYI13, HM12, LAK^{+}21, SFF^{+}11, YLZ15, YBA^{+}16]. disturbed [MCL^{+}17]. diurnal [GBR^{+}13a, KK12b, MvdWU^{+}16]. divergence} [Abe15]. Divergent [SWR^{+}16, SK15]. Diverging [MCS^{+}18]. diverse [FS14b, For18b, KYvdPS14]. diversification [MKH^{+}17]. diversified [SJS^{+}13]. Diversion [GCL^{+}19, CCL^{+}13, YY11]. diversion-type [CCL^{+}13]. diversions [WCL^{+}17, dMLM^{+}17]. diversity [ABR12b, ABR12c, Ban11, BRMW18, BME14, CHR^{+}13, CHZ18, CBB19, DDR^{+}14, Fuj16, GEEC14, GEZ^{+}13, GCCJ12, Gre12, GWF^{+}18, GEB10, HBSF16, KPMJ14, LLQN13, LZF^{+}14, MBG^{+}10, PGR^{+}16, PBG^{+}17, PPO12, RKBG14, RHD^{+}13, RHR^{+}15a, TBL^{+}15, TPO^{+}18, VCD^{+}17, VH12, Wal15, WRMB18, ZJG10]. Divide [LM16]. divided [MAB19]. division [BHHM10, VVA^{+}15]. DMSP [WLC^{+}19]. DMSP-OLS [WLC^{+}19]. DNA [Abe13]. DNDC [GAC^{+}14, LWL^{+}17a, LYD^{+}17]. Do [FL11, GVV17, MSV12, NLMdL10, WRB^{+}12, XM13, CHR^{+}13, SHC^{+}15, SPWR17, TL15, VPM^{+}18]. DOC [JNC^{+}11]. DOC-3 [JNC^{+}11]. documentation [SAKB^{+}10, THO10b]. Documenting [GAF^{+}14]. Does [FHAG12, FE14b, Log19b, MPT^{+}11, MHS^{+}17, RCSS13, RVB^{+}10, SP10, VVA^{+}15, BL15, CLHH19, Sil11, ZKSF17]. dog [PKB^{+}18, YRGH16]. Dogs [BG15]. dolphin [BWC17, MMCBP19, PNHL14]. dolphins [SGSR10, dJHMS19]. domain [LNB17, LJB12, PWK^{+}17]. domain-level [LNB17]. domain-specific [PWK^{+}17]. Dome [YZ18, KOBK13]. Dome-shaped [YZ18]. domesticated [GLSA^{+}11]. dominance [CD16, FRP18, MSG^{+}13, SB11]. dominant [HGPF15, IKK^{+}12]. Dominants [HM18b]. dominated [Ano12s, CPS15, GRL18, JHH^{+}12, LBWW14, MCW13, NFU19, OCB^{+}13, SLW^{+}14, XYCZ11, XYCZ13, XYY^{+}16, ZWJ^{+}12, ZLY^{+}13]. domingensis [ESA12]. dominica [YWA^{+}17]. Donald [GB16b]. donkeys [SU15]. Donor [GFC15]. dorée [LPPA15]. dormancy [BCL^{+}14, MBBA17, SMLH^{+}11]. dormant [ASS14]. dose [BPK^{+}16]. Dosidicus [ZBMBDU^{+}17]. Doubtful [PNHL14]. dougallii [GQNMR18]. Douro [ABD14]. dowels [SBC^{+}11]. down [FMF12, MBD^{+}18, PMBB^{+}16]. down-sampling [FMF12]. downed [RWFD13]. downwind [GKKW10]. DPSIR [SDSH15]. drainage [MKSL^{+}17, ZF11]. drained [HJS^{+}16]. Dreissena [KVP16, Lin15b]. dreissenid [JXL^{+}15]. dreissenid-colonized [JXL^{+}15]. drier [HWE^{+}18]. Drift [GJM^{+}13, AHNK13]. drifting [CVPH^{+}14]. drive [CHPS17, Rai13]. driven [BMR^{+}18, Ben12, BNC^{+}18, CTB^{+}12, DDB^{+}12, FKH^{+}15, FTK^{+}14, GPJL15, HDL^{+}16, KM12b, KdVRdW17, KWGW19, LDO^{+}18, LLZ^{+}15, LMCA^{+}12, MJLC14, MDV^{+}12, MMMGLN^{+}18, PHP^{+}14, PSBH16, SZV14, SQWS19, TSM^{+}15, ZLFI11]. Driver [CpW16, GHST19, STJ15]. Driver-system [CpW16]. drivers [ABD^{+}18, ACK^{+}14, CSS^{+}16, HBS16b, HSM^{+}12, JXL^{+}15, KSI14, LSH15, MPG18, MCP16, PDAL14, RPP^{+}15, SHM^{+}17, TNTH11, VPM^{+}18, WCS^{+}14]. drives [Gra14, HKM^{+}18, PGA^{+}11]. Driving [CWLL15, SCG^{+}13, BU10, BU16a, DMZ^{+}19, MNEG11, VEMS12]. drones [AOW13]. droplets [DKM^{+}14]. Drosophila [CMH14, LAR^{+}16]. Drossel [ZJG10]. drought [FMS13, JWX18, LCH12, LZF^{+}14, MBW14, MEC14, SLZ^{+}13, TGB^{+}17, WYHC15]. droughts [XC14]. drowned [BdCS11]. dry [HBG^{+}16, HSVL12, JG10, ZCYL15]. dryland [CNH^{+}11, GE17, LZLC13]. drylands [CpW16, STJ15]. DSSAT [DJA^{+}10, DBJ13]. dual [CXW19, JHHS17, Par10, Ula10]. due [BMR^{+}18, CCL^{+}13, CBA16, dSCdA19, MJB^{+}11a, OJ10, PTA^{+}12, PVL^{+}14, PL10b, RGH^{+}13, SM14, vPTOW18]. dune [HCW^{+}17]. Dunker [dÁSMRT^{+}19]. durability [HGSJ15]. duration [CGB^{+}13, TCSG15]. during [BdCS11, CAMNN16, CMLY19, CLDRJ18, Fel12, HKN17, KHL^{+}16, LSJG17, LBC^{+}15, PMB^{+}17, RJ17, SSY14, SBC16b, Vav12, YSA^{+}14, ZLW^{+}15]. dusky [SGSR10]. dust [SH18]. Dutch [HV12]. dux [CMEP15]. dyad [KKP^{+}11]. Dyer [WSW^{+}16]. Dynamic [BGMFES15, FCH^{+}17, JBH^{+}14, JA15a, LSJG17, MJB^{+}11a, MDF^{+}18, PBEL17, RBH13, Til21, VRC17, VRL18, WN17, WO14, ZR19, AYASR^{+}17, AD18, BSH17, CW17, Che10, CSFR15, DAA^{+}12, DA17, EPA^{+}10a, FSV^{+}12, FS18, GvdWBS11, GBW^{+}15, HFH^{+}19, HPS10, Hei11, HCM^{+}18, HAOS17, HBL18, JWX18, KM12a, KJH^{+}13, KP10, LJBS12, LBZA19, LJR14, LH15, LWGH13, LLZ^{+}15, LZS^{+}17, LLH^{+}17, LYCU14, LPD14, LJCG15, MBLC14, MBHW12, MLMFN^{+}15, MRG13b, NMC^{+}14, NF11, PBFS14, PLL^{+}19, PGGMMA19, RFZB12, RR12, SBC13, SMLH^{+}11, TLLX19, TŠAD12, TJYZ11, Til14, TLS^{+}16, WR17, YJW15, YZG16, YKvWK16, ZGJ10, ZSF^{+}12, dMSL^{+}16, vDSC^{+}10, vDHH^{+}14, DLL^{+}19, GPH17, GMvV17, JZ12, LLC^{+}17, PLS^{+}10]. Dynamical [BPK^{+}12, LDM15a, BRHD15, DAF^{+}11, DH18, GLT10, JSKK15, OBC^{+}21, PRP17, TAR13, WLR18]. Dynamics [Ano12a, BO14, DPR10, NS15, NYF16, NSN17, OLC^{+}17, RDD^{+}14, VGR^{+}11, ZFG13, AD13, ACCM16, AAME11, ACKS13, AD12, ACM15, AGB^{+}13, AWSS14, Ant14, ARV^{+}16, BFA11, BCB^{+}11, BBB^{+}17, BKV17, BMMC21, Ben12, BLB^{+}17, BT16, BSC^{+}15b, BVVS15, BPK^{+}16, Bin17, BNC^{+}18, BKMK10, BCI^{+}17, BSCG10, BD18, BQF^{+}18, BPM^{+}15, BPM^{+}17, BCJ^{+}12, BvBB10, BRF^{+}11, CC12, CLBG15, CHPS17, CMH14, COBF13, CSN^{+}19, CHCC15, CAC^{+}16, Che17, CBO16, CDT10, CD16, Con14, CLDRJ18, CHK^{+}12, CLL^{+}15, CTP^{+}17, DWNN10, DJT^{+}19, DTC^{+}10, DdSCG11, DHC16, DLCR13, DMR13, DBVS19, DSD13, DS12, DHS^{+}16, DBP16, DFM12a, DFM12b, DFF17, DRS^{+}16, ESA12, ER10, ESLJ18, EAB11, FGH^{+}10, FCZ13, FS14b, FBdP^{+}16, FT16, FC11, FNR11, Fuj16, FPG^{+}10, GCC^{+}10, GHS^{+}16a, GQRC^{+}17, GQNMR18, GDC12, GC11]. dynamics [GPM16b, GPCG15, GV15, GGVB11, GHS^{+}10, GRD^{+}14, GMvV17, GRB12, GSS^{+}14, GVC10, GDDFT16, GCM^{+}13, HLS^{+}14, HJ17, HS13c, HS13b, HSR^{+}15, HBG^{+}16, HMA10a, HOKH18, HHF^{+}10, HBA16a, IGBSASSJ16, JM15, JX15, JXL^{+}15, JDT^{+}16, JX17, JASW18, KSS11, KK11, KW18, KCF^{+}13, KP13, KF16, KAF^{+}14, KAC^{+}12a, KAC^{+}12b, KPLS15, KWGG19, KHHN17, KWJH16, KJS13, KMS^{+}12, KMS^{+}13, KvTL^{+}18, LKKK^{+}14, LJAD17, LHO^{+}16, LAR^{+}16, LBR16, LLB^{+}19, LAD11, LCH12, LBWW14, LAB^{+}10, LHZ10, LDL12, LZLC13, LVLS14, LLZ^{+}16, LZ10, LWC^{+}18, LW14, LKP13, LOLA13, LSGT13, LZZW17, LZYH19, LPD14, LCML11, MRG13a, MGMR15, MD16, MLG18a, MXC13, MM19, MJL^{+}13, MJL^{+}14, MVC^{+}17, MYHN17, MKSL^{+}17, MFD^{+}15, MRD^{+}12, MGW^{+}13, MG18, MPK18, MJF19, Mon13, MCA18, MP17, MNHW^{+}13, MPG18, NMT^{+}19, Ned15, NWAR12, NBK^{+}11, NJM11, NB11, NBD^{+}13]. dynamics [NP15, NPH^{+}17, NTK14, PRHH13, PMB^{+}17, Par10, PLFO^{+}12, PMRGF10, PK14, PNHL14, PZM13, PVB16, PMTKC17, PME^{+}16, QL15, QFG^{+}13, RJG^{+}13, RPSA12, ROC^{+}13, RKC^{+}15, RDR^{+}13, RRH^{+}19, RSC12, RSC13, RV15, RAR15, RPGSCF16, SHBG10, SVD^{+}13, SMSAN^{+}14, SSBJ12, SHB^{+}11a, SK10, SBR^{+}19, SAR11, SRSS12, SKMB11, SKH^{+}13, SXNL18, SVL12, SSK14, SLH^{+}11, SBRD10, STD15, Sim11, SSC15, SJB^{+}16, SFL^{+}19, SBC^{+}11, SDKE18, SMG^{+}12, SU15, SBC16b, SKPHW16, SGBG19, SMBGH^{+}19, SW15, STJ15, TSP19, TM14, TDC^{+}11, TT12, TF13, TXM^{+}10, TOQW11, TLZ17, Tie16, Til15, TNHB12, Uch16, URJ10, Vav12, VRL14, VJR^{+}16, VASH^{+}16, VHDM17, VPS10, VPM^{+}18, WAS^{+}16, WGL^{+}12, WJL^{+}13, WMK^{+}13, WAM13, WLH^{+}10, WCS^{+}14, WSW^{+}16, WRSL11, WTL17, WFDDR10, WPZ^{+}14, XLS11, XZC^{+}17, XTPT11, XCR^{+}13, XSL18, XGH^{+}17, YL18, ZSS17]. dynamics [ZJM^{+}11, ZLZS14, ZLB14, ZLW^{+}15, ZJL^{+}16, ZWW^{+}18, ZBSF17, ZZLZ14, ZSSR18, dBGS^{+}14, dJHMS19, dAdSC15, vLTdBP16, RCG^{+}13]. DYRESM [WV12].

E-book [Jør14a]. e-flows [YYQ11]. e-waste [KS18]. E. [JA15b, KPSC10, KH15, SFvB10]. Early [Bae18, OFGH^{+}19, GLS10b, GCvDM12, GCM^{+}13, LSZC10, MCMT^{+}16, PGA^{+}11]. Earth [PVM13, MLG^{+}18b, SFC^{+}12, Cam16, FvBM^{+}11, PVM12, SBDA16]. earthquake [CLCC11, LLL^{+}10]. earthworm [CSS^{+}17, JHT^{+}14]. earthworms [HHF^{+}10]. East [LY16, LCC^{+}14, LSJG17, MGMR15, PSKO16, SHB11b, ZYZ17]. Eastern [FAC^{+}19a, BHC^{+}15, CPR10, CAC^{+}16, DLW^{+}17, EH15, EH19, GPR^{+}13, HMPC11, LLC^{+}17, NB11, PWK^{+}17, RWFD13, RTE^{+}15, SM14, TL15, WCL^{+}17, XKI^{+}11, ZLB14, MR17, NBD^{+}13]. easy [DR14]. eating [DF13]. Eco [WL13, AMG19, ARV^{+}16, BC13, Bev15, BPD^{+}19, CHTC11, CZCL18, CJC11, DR17, DLZ^{+}17, JN14, LDL12, LTL^{+}13, LB18a, MJB^{+}11a, Nak17, NP18, Pat10, RHQ^{+}16, TKW17, VFB19, XYCZ11, YXD^{+}18, YHZ19, ZGJ10, ZZF15, ZWG^{+}17]. eco-agriculture [LDL12]. Eco-City [WL13]. eco-efficiency [YHZ19]. eco-environment [ZWG^{+}17]. eco-evolutionary [AMG19, ARV^{+}16, RHQ^{+}16]. eco-exergy [BPD^{+}19, CZCL18, CJC11, JN14, MJB^{+}11a, VFB19, XYCZ11, ZGJ10]. eco-genetic [DR17]. eco-hydrologic [Nak17, NP18]. eco-hydrological [BC13, LB18a, YXD^{+}18]. eco-industrial [DLZ^{+}17, ZZF15]. eco-physiological [LTL^{+}13, TKW17]. eco-resilience [CHTC11]. eco-systems [Pat10]. EcoBase [CVG^{+}15]. ecodynamic [JZG10]. Ecodynamics [Tie11]. ecohydrologic [KF16, KJS13]. Ecohydrological [QWJL19, AAFH15, GCK^{+}15, GLRT16, HGPF15, LZS^{+}17, YYL13]. Ecol [ABR12b, ABAG16, Ano12a, BRV11, Bar11, BMF^{+}11a, CCW^{+}15a, Con12, CR12b, DYMYGO19a, ERGW^{+}11a, GORJ12, HM19, HLLO13a, Jør15a, KS17a, KMS^{+}13, mLcLJ16, LSJ^{+}19, Lin16, MJL^{+}14, MSCH14a, Och10, Pat10, SMT19b, Str15a, WGA^{+}12, WLS^{+}11, YTT^{+}11, ZLJM17b]. ecologic [Fai10]. Ecological [Ano13u, ADL^{+}10, BFS11, CWLC13, CC11, CCF13b, CPE13, DBG14a, EH19, GWR11, GLSA^{+}11, GPR^{+}13, GWZ^{+}15, KA17, KAF^{+}14, KEHG17, LYP^{+}11, LPA^{+}11, LBK^{+}13, LHL18, LYCZ11, LYC^{+}15b, LLP17, MDG10, NGR^{+}11, PBL^{+}15, PAEE19, QCB^{+}11, ROC^{+}13, RHO^{+}12, SELB10, Sol19, TLL^{+}17, TM14, TZS^{+}14, VMG^{+}12, YLZ15, ZC11, ZYFL10, ZZF15, ZYKS15, ZLZ17b, ZLZ^{+}17a, ZMH19, ABS17, AG15, AD18, Ano12t, ASP^{+}15, ACK^{+}14, ABL^{+}13, AVG14, BCC^{+}11, BGMFES15, BBJV^{+}11, BS16a, BPB^{+}18, BGG16a, BA14b, BOGA14, BB18, BSMA18, BRHD15, BFJL15, BLCPB11, CNH^{+}11, CBC^{+}16, CGV^{+}15, CGGL16, CZW10, CCC11, CCL^{+}13, Che15a, CXW19, ChHC12, CLD^{+}17, CD16, CPJ^{+}13, CLL^{+}15, CK12, CGP12, DWNN10, DFW^{+}18, DNFG18, DFK^{+}11, EBL^{+}11, FMBD15, FLLB11, FP17, FRM^{+}19, FLB^{+}15, FDHL17, FBU14, FPG11, FPWG14, GPM^{+}16a, GSQ17, GSF^{+}16, GGV13, GdTCFG10]. ecological [GK14, GCV14, GS19b, GFW^{+}17, GCCJ12, GFP^{+}16, GT14, GB16b, GB16c, GB16d, GLC^{+}15, GCIB10, GEPCB13, HRSKW15, Hal11, HBF18, HKL^{+}11, HLW15, HF14, HS13a, HM10, HGBA^{+}13, HDL^{+}16, HWH^{+}18, HhLY^{+}19, JGF14, JTB^{+}19, JMW11, JBW^{+}17, JN14, JSdSdL^{+}18, KK16, KM12a, KK10, Ken19, KPM^{+}14, KR10b, KMKL19, KWGW19, LL17, LKMCW14, mLcLJ15, mLcLJ16, LM13, LMY^{+}11, LA11, LF11, LAB^{+}10, LSZC10, LHZ10, LHZN11, LCCL15, LYSW15, LLQ^{+}15, LWTZ18, LS13, Lin10, Lin16, LMGM11, LYCZ13, LAC10, MLY^{+}19, MG16, MPR13, MRM16, MY11, MJY17, MWY^{+}18, MRN12, MDJ^{+}11, MWB^{+}10, MCR^{+}15, MPF^{+}14, MHC16, MLG^{+}18b, MSM^{+}16, MMMCAHM21, NCMF10, Nie16, NMdR16, OTA16, OBC^{+}21, OCD^{+}13, PW10a, PBO19, Pat12, Pat14b, PMH17, PHP^{+}14, PJ15, PMCKH13, PWK^{+}17, PLS11, REB^{+}16, RR12, RJ17, RNS^{+}19, RKW12, RGRMRR18, RDH^{+}13, RHS^{+}16]. ecological [RDV^{+}17, RVB^{+}10, RH17, RH15, SBVA^{+}13, SBC13, SSV^{+}12, SKHH19, SBD^{+}19, SMH^{+}12, SWC13, SWZ^{+}13, SA13, SLH^{+}11, Sil11, Ske13, SHB11b, Sma13, SBA^{+}13, SDSH15, SYY^{+}19, SXZ^{+}15, TVT12, TBA19, TLW^{+}14, TVV13, TTL^{+}18, TWP^{+}14, URR^{+}17, VMP^{+}19, VSB13, VG19, VQS^{+}18, WJL14, WSE12, WLR18, WAM13, WSS^{+}16, WR12, WK17, WKH19, WWYO17, WWYO18, XCG^{+}16, XYY^{+}16, YZ16, YZ18, YSH^{+}10, YJW15, YJL11, ZFT^{+}10, ZLLZ13, ZLZS14, ZFZ^{+}14, ZLFZ16, ZCSL18, ZF19, dHR18, vdKGV11, SSF14, WTC^{+}18, BH10]. Ecological-economic [KEHG17, HF14, VQS^{+}18]. ecological-economic-social [SXZ^{+}15]. ecological-hydrodynamic [FRM^{+}19]. ecological-network-analysis [MY11, MWY^{+}18]. ecologically [CHTC11, KVU^{+}10, MAFG^{+}19]. ecologies [AA12, KRE^{+}18]. ecologist [PC17, SDEP18]. Ecology [BFW14, BME14, HKW^{+}18, PF18, RvS16, Wei19, Bin17, FBU14, FS18, HBF18, HB12, Ji15, JNF16, KS18, LCCL15, LHH12, MSMM11, MGS15, Pat14a, Pat16b, Pau14, PES19, PBT^{+}11, PÓF^{+}18, RHR^{+}15a, RHR^{+}15c, RHR^{+}15b, SHCR10, VGR^{+}11, WHB^{+}16, Wal17, XZCL16, ZMS19, ZDL^{+}17]. econometric [Sci11a]. Economic [HFU11, Ano12t, BB10, BD18, BU11b, CRTS15, DK17, DZW^{+}11, GK14, HF14, HKM^{+}18, IS13, ISG13, KEHG17, LA11, LHZN11, LZL^{+}18, LYCZ11, LYCZ13, LDCB21, MSM^{+}16, NXZ18, Pat12, Pat14b, Rau11, RGRMRR18, RGH^{+}13, SXZ^{+}15, TLLX19, VQS^{+}18, WFSH15, ZFZ^{+}14]. economic-ecological [LHZN11]. economic-ecological-social [RGRMRR18]. economics [ABS17, Wal16]. Economy [GLC^{+}18, BCP16, BCO17, CC11, FDR^{+}17, HCM^{+}15, HZSH17, Ji15, MH11, NLM^{+}19, XHC^{+}15]. Economy-pollution [GLC^{+}18]. Ecopath [HCL^{+}16, HKL^{+}11, AW15, CVG^{+}15, DLH^{+}14, ERGW^{+}11a, ERGW^{+}11b, Fre16, FMLM11, GLC^{+}15, HMA10b, LRZ12, LSR16, NJCF11, SBC^{+}16a, URUNAM12, VASH^{+}16]. Ecopath-derived [GLC^{+}15]. ecophysiological [GTC16, KYvdPS14]. ecophysiology [IFC14]. Ecoregion [LLT^{+}17]. Ecosim [AW15, CVG^{+}15, ERGW^{+}11a, ERGW^{+}11b, FMLM11, HCL^{+}16, HMA10b, LRZ12, SBC^{+}16a, VASH^{+}16, WLDL12]. Ecospace [AMH16, RMHE15]. ECOSSE [LYJ^{+}15]. ecosys [Gra14]. Ecosystem [AHM^{+}16b, CMCASS12, CPS^{+}14, GAP^{+}19, GB15, HF14, Jør14b, KB10, LF11, LXS^{+}16, LYL^{+}17, MML^{+}11, MSS^{+}19, PBL^{+}15, PCB11, SS16, TNHB12, TPT^{+}11, VLZ^{+}19, WNT^{+}13, WLF^{+}16, XYCZ11, YHK^{+}19, AAJB21, ALC19, AMH16, Ano17a, ADL^{+}10, AJDS16, BHD16, BDV^{+}19, BSFR17, BSH17, BJPB11, BS19, BS10b, Bre12, BDTV11, BPGSS^{+}18, BFM11, BRPK14, BRKP16, CLMA19, CLBD^{+}11, CR12a, CRUF13, CBR^{+}16, CSS^{+}19, CHvdW^{+}10, CSN^{+}19, CLH17, CGGL16, CLA^{+}12, CWLL15, COPP13, CSF11, CCC12, CSS^{+}16, CMCC15, CGM^{+}14, DAW^{+}10, DUASLB^{+}12, DBVS19, DBG14a, DYU^{+}12, DML15, DMBS11, EBA^{+}19, ESP^{+}17, ER13, FLLG13, FLH17, Fie12, FHL^{+}13, FBdP^{+}16, FD19, FBKS17, GYL^{+}15, GMTC^{+}17, GPR^{+}13, GANB11, GFP^{+}16, GFC15, GPP^{+}19, GYL^{+}13, GPC14, Hal11, HBR19, HBS^{+}16a, HKL^{+}11, HCZ^{+}17, HV15, HWS11, HJ10a, HCL^{+}14, HJS^{+}16, HS13c]. ecosystem [HS13b, HB10a, HWE^{+}18, HCL^{+}16, HT16, HPRNP12, HPC13a, HPC13b, HKG^{+}17, HMA10b, HDL^{+}16, HWH^{+}18, HBA16b, IT17, JBV^{+}15, JSL^{+}10, JC11a, JDT^{+}17, JOHW17, JMB^{+}11, JBW^{+}17, JL18, KM12a, KK10, KSAS12, KHW^{+}17, KTM^{+}16, KEM^{+}16, KHL^{+}16, KVU^{+}10, KL12, KGYG15, KVP16, LSJ^{+}18, LSJ^{+}19, LBA14, LBA15, LBSB^{+}14, LB11, Lee14, LBD10, LDO^{+}18, LdMS^{+}16, LLM^{+}13, LYJ^{+}15, LZZP16, LWL^{+}17b, LLC^{+}21, LS10, Lin15a, LVZ^{+}11, LLZ12, LC13, LG17, Löp11, LCML11, LMBL12, MTZ^{+}10, MK13, MPT^{+}11, MRG13a, MGR^{+}18, MJB^{+}11b, MKL14, MGCL17, MGS16, MK12, MK14, MPL^{+}16, MZLCSZdP11, MSCH14a, MSCH14b, NCEGB16, NMC^{+}14, NXZ18, NPP^{+}18, NGWZ16, NJCF11, OGG^{+}16, OWMK10, OCCF17, OCB^{+}13, PLR^{+}13, Par10, PWP^{+}19, PGV10, PPO12, RGU^{+}15, RJH^{+}10, RRHH10, RSDP^{+}12, RKBG14, RDH^{+}13, RDD^{+}17, RRH^{+}19, RA13, RSD^{+}15, RWP11, RBGB16, SLW17]. ecosystem [SVD^{+}13, SB11, San17, SNS^{+}12, SSV^{+}12, SS10, SM13, SRSS12, SLW^{+}14, SHM^{+}14, SAW^{+}11, Sin17, SSN^{+}18, SDY^{+}15, SMG^{+}12, SCG^{+}13, SS11, SFB^{+}15, SH18, SGKR10, TLLX19, TM18, TLL^{+}15, TŠAD12, TXM^{+}10, TRE^{+}12, TCH^{+}13, TAGC^{+}16, Ula10, URUNAM12, UDRA13, VCD^{+}17, VC17a, VVA^{+}15, VFB19, VASL^{+}15, WGA^{+}11, WGA^{+}12, WHB12, WDL^{+}15, WO14, WYS12, WRM^{+}11, WJK11, WJvdL^{+}13, XZC^{+}17, XTPT11, XYCZ13, XLL15, XC14, YL17, YLY^{+}18, YLC13, YLF^{+}15, YLHZ14, YLL^{+}12, YDL^{+}15, ZHC^{+}16, ZRCNLI^{+}15, ZGJ10, ZWG^{+}13, ZLY^{+}13, ZLLC13, ZZB^{+}18, ZZL^{+}19, ZSF^{+}12, ZZLZ14, dMSL^{+}16, dMLM^{+}17, FSB13, HPC13b]. ecosystem-based [HCL^{+}16, KTM^{+}16, NCEGB16]. ecosystem-level [WRM^{+}11]. ecosystem-scale [LF11]. Ecosystems [BH10, PSJ11, SJP14, Ana17, BO14, BSC^{+}15a, BCO17, BBC^{+}19, BBA12, BPD^{+}19, BRB14, CLBD^{+}11, ČTT^{+}12, CY13, CSR15, CGEdM12, CMSPJ11, CPJ^{+}13, DWD19, DYMYGO19a, DYMYGO19b, EPM^{+}14, GPM^{+}16a, GLA11, HB12, HBH^{+}12, HJ10b, yHAB12, HYL^{+}17, JS12, KRL^{+}16, KBC^{+}12, LM13, LKG10, LBE^{+}18, LBK^{+}13, LLY^{+}13, LZLC13, LLTQ14, LLTQ15, LXS^{+}16, LVPF15, LPX^{+}12, LHR^{+}15, LKAY15, LJLO11, LD16, MSP^{+}13, MY11, MBSF14, MCT10, MDF^{+}18, Mer12, MJC^{+}11, NJM11, NBD^{+}13, PRHK19, PNT^{+}19, Pat10, PCB11, QL15, RDD^{+}18, RBGB16, SMSAN^{+}14, SFF^{+}11, SKMB11, STD15, Sol19, SGB10, SGHW^{+}17, SPP18, Tie11, TYEH10, UDNSLCG16, UDNSLC^{+}18, VLZ^{+}19, Vav12, VASH^{+}16, WJL14, WLY^{+}15, WR13b, WPB14, YCF15, YLL^{+}17, YFC^{+}11, ZW12, ZJL^{+}14, ZYKS15, ZJP^{+}11, Pat10]. ecotone [DBG14b, DBG14a, JDT^{+}16, PRHK19, PR11]. ecotones [ACM15, DVCM^{+}10]. Ecotourism [PKG16]. ecotoxicity [JZ12]. ecotoxicological [PAEE19]. ecotoxicology [JBH^{+}14]. Ecotracer [LSR16]. ecotypic [MDB19]. Ecuador [Bre12, HGBA^{+}13, JVY19, SRV^{+}12]. eddy [MYHN17]. eddy-resolving [MYHN17]. edge [DNFG18, FJH15, FCA18, LdMS^{+}16]. editing [GdC14]. editor [Fat10]. Editorial [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Fat12, FGG^{+}14, Fat16a, FKDG14, Jor11b, Jør16, Las10, Ano10a, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano14a, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l]. Editorial [Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z, Ano19-27, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano21a]. Eds [Wei19]. Education [DSBM^{+}18, FMBD15, MHS^{+}17, Mob18, WHH^{+}18]. educational [CXL^{+}19, LB18b, SGB10]. edule [RGU^{+}15]. Edward [GPC14]. Edwards [MGWW13]. EEHS [DFL^{+}16]. eel [CRQW18, TSN15]. Eelgrass [CVPH^{+}14, CVPR^{+}16, FRV^{+}16, KR10a, KRCVF16, KvTL^{+}18, LZW^{+}17]. EFDC [KLS17, LL18b, WX11]. Effect [BRKP16, CGLA10, CL12a, CZCL18, FDM^{+}11, KPSC10, LLTQ14, LLTQ15, NMA^{+}19, PGS16, SHS11, TLP17, WC19, ZWJ^{+}12, ACM15, AML13, BGA17, BB18, CZC^{+}15, CEC10, CHNdÁP18, CHZ18, COY^{+}10, CN13, CTC^{+}18, dSCdA18, DWWS15, DMH17, DWP16, FGPR10, FtHvdB^{+}14, FG11, FAPM12, FAC^{+}19b, GHWG^{+}14, GBG^{+}15, GORJ12, GLRT16, Hol13, JVY19, JGB15, Kal19, KWGG19, KPMJ14, KR14, KMM^{+}13, LJR14, Lec17, LHZ10, LJ13, LPS^{+}10, LPT^{+}18, MLW11, MW12, MW15, MSP^{+}13, MSS^{+}19, MF16, MP13, MMC^{+}19, PJ15, RGU^{+}15, RSE10, RW17, SBCB13, SCHQA16, SBDS13, SSC15, SLZ^{+}13, SGWV15, UR10, VRL18, Ver10, WM14, WDL^{+}15, WWYO17, YFL^{+}18, ZBSF17, dS18]. Effective [NSIiT17, ASP^{+}15, CCS13, CPAG15, DLH^{+}14, GDBB19, KUI^{+}18, LBA14, LBA15, MS14, NLMdL10, RPA^{+}18, SAT^{+}15, VF16]. Effectively [NYIH17]. effectiveness [EST^{+}19, FMF12, HGBA^{+}13, LCP14, LRB12, PHvB^{+}18, RGHS11, RAM10, TLS^{+}16]. Effects [AVL^{+}11, AAME11, AKLB19, BFP^{+}18, BH13a, BGG^{+}16b, CRBC11, hCjHMW18, CSH11, CMLY19, CMCC15, DA19, DG16, EZ14, FBA^{+}11, Fuj16, GVGZ^{+}21, mGZK^{+}10, GPCG15, GGR^{+}13, GLY^{+}15, HPC13a, HPC13b, IWW12, JJYM19, JL18, KSS11, KTK^{+}15, KM12b, LRM^{+}14, LBZA19, LCY^{+}13, LKG^{+}17, LLQN13, LLH^{+}13, MPK18, MHRB19, MMPG16, NNST^{+}14, PL10a, PMM11, Pea15, PGV10, PLHG14, PMBB^{+}16, RHO11, RW10, Rem11, RDH^{+}13, SGAS14, SAI10, SPB^{+}17, SFHK12, TM18, Tat16a, TDH12, WR12, YAGMM^{+}14, dMLM^{+}17, ABAG14, BFA11, BNMP11, BNL^{+}13, BR16, BMCG12, BFS11, BvHR^{+}15, BGPJ10, BWML15, BRB14, CCCP17, CDR^{+}13, CLA^{+}12, CGB^{+}13, CSS^{+}16, CJM^{+}12, CWJ^{+}13, CHK^{+}12, DSJ19, DPPI14, DMR13, DBVS19, EGMW12, EAB11, ERGW^{+}11a, ERGW^{+}11b, FLH17, FT16, FKSK^{+}10, FCA18, FAT^{+}15b, FTTV^{+}18, FPWG14, GHWG^{+}14, GPK^{+}13, GCC^{+}10, GB16a]. effects [GCA16, GSJ^{+}10, GV11, GLS^{+}17, HMPC11, HHC^{+}13, HFSL14, HMP^{+}19b, HZT^{+}13, HW12, HHW11, HR11, JTM^{+}17, JCKG15, JX17, JHT^{+}14, KKH^{+}17, KFWM19, KSZW16, KBS^{+}15, KPLS15, KHS^{+}19, KJH^{+}13, KPTC14, KPK^{+}17, Lab11, LLC^{+}17, LMGW16, LJ10, LLT^{+}17, LBW18, LDP14, LHR^{+}15, LCFD19, LBC^{+}15, LLFL12, LCR12, LMBL12, LSL10, LHLW15, LSM19, MK13, MKH^{+}17, MCH15, MRCW10, MWK19, MJL^{+}17, MKL14, MYHN17, MBG^{+}10, MGCL17, MJK^{+}17, MJC^{+}11, MB14, MSR^{+}14, MGWW13, OLC^{+}17, OAPP12, PC17, PAT17, PSG19, PVBC14, PCZ^{+}16, RPSA12, ROSC15, RLS11, REB^{+}16, RSP^{+}17, RMR^{+}16, RCST11, RSC12, RSC13, RWvDM15, RSD^{+}15, RH17, RBGB16, SMC11, SB11, SSBJ12, SHB^{+}11a, SSS^{+}11, SHX^{+}12, SKTG11, SGH16, TdS12, TWH^{+}14, VG19, VASL^{+}15, WGA^{+}11, WGA^{+}12, WJL^{+}13, WGTH16, WKZ^{+}19, WPB14, WCG^{+}11, XLS11, XYY^{+}16, YKCN10, YSB11, YGL18, YYH19]. effects [ZBAY14, ZYKS15, ZZ17, ZCSL18, ZLD^{+}19, ZML^{+}12, ZZLZ14, ZJP^{+}11, ZSKI17, dMSL^{+}16, dSJHG10]. efficacy [HL19, LZB17, ZCR16]. efficiencies [Pea15]. Efficiency [BVC^{+}18, PP19, PDV17, BSC^{+}15a, CPD^{+}11, CGEdM12, DHC14, DMBS11, GCV14, GDW12, HLZ^{+}18, HFMM^{+}12, Hoa11, HNWT10, LGMLD19, LY11, LCL^{+}13, MF12, MDV^{+}12, Oga11, SRTS18, TM12, TBL^{+}15, UB12, YHZ19, ZZFH15, ZJP^{+}11]. Efficient [CN11, CS18, LM16, LZB17, Pat12, WHC^{+}13]. effluent [NFU19]. effluent-dominated [NFU19]. Effort [BWRA16, BMP^{+}19, BPWB14, CGLA10, FN15, RMHE15, VRM^{+}10]. efforts [BA17, SKM^{+}16]. Egg [GJM^{+}13, CEC10, CC10, DJMBD^{+}13, GPZB18, MLS^{+}11, ZSG^{+}18]. egg-bearing [DJMBD^{+}13]. egg-laying [GPZB18]. eggs [GJM^{+}13, PPZ^{+}17]. Egypt [ESA12]. eigenfunction [LJ15c]. eigenfunction-based [LJ15c]. elaphus [RR10]. elasmobranch [NCMF10]. elasticities [EC10]. Elbe [ZZT13]. Eldana [PvVC13]. electricity [HFU11]. electrofishing [LLB^{+}19]. ELEFAN [SMT19b, SMT19a]. elegans [BGA17]. elementary [BP19, GB16a]. elements [CLL14]. Elephant [RWS^{+}16, BQF^{+}18, FKSK^{+}10, MMS12, RWS^{+}15]. elephants [SWS16]. elevated [ZJP^{+}11]. elevation [HJN12, KR10a, OT15, RSC12, RSC13]. elevational [ENC^{+}15, MLK^{+}11]. Eliciting [CLT^{+}11, PSM^{+}12]. eliminate [Log17]. elk [RR10]. Ellenberg [KGD14]. Elsevier [Jør14b]. Elsheik [Jør14a]. elucidated [ROC^{+}13]. Elucidating [BSP12, KPLS15]. Elucidation [HMA10b]. elusive [CGV^{+}15]. embayment [WEJ11]. Embedding [dBGS^{+}14]. Embodied [HCM^{+}15, JC11b, SWZ^{+}13, ZQC15, AAW^{+}15, JC11a, JLN10]. emerald [KKR19, VCSSC12]. emerge [GDTP13]. Emergence [CGM^{+}14, TTCC^{+}19, BCL^{+}14, GDC12, GM10, GB16d, LPF11]. Emergent [Ban11, BR16, RPM13, RNS^{+}19, SK15, BA14b, EAJG16, KGM12]. emerges [LDB14]. Emerging [Gia14, Cam15, DBH^{+}17, JD18, KR10b, PSJ11, SJP14, Pat10]. Emergy [Abe13, BMF^{+}11a, BMF^{+}11b, BCO17, BCT15, BU16b, CLL14, Cam15, Cam16, FBU14, Har11, LDJZC15, LT15, LLCR10, LYCU14, MBRR13, MCPB15, Rau11, Ten15a, ZYLY11, ALL11, Ano12t, ARTBB14, BPT^{+}19, BLOJ^{+}19, BU10, BU11a, BU16a, BCUF16, BCDU16, BPD^{+}19, CT16, CLHH19, Cho13, CT12, DB15, DBSA16, DLD^{+}11, FMBD15, FB15, FAC^{+}19b, GRRU15, GZVU14, GAB12, GFC15, HT14, Ing10, JZCL14, JC11b, KØ13, KRE^{+}18, KBMM15, LTL15, LB17, LCBNP19, LLCR11, LLTQ14, LLTQ15, LYCZ13, MCPB13, MCB14, Pat12, PMH17, PPF11, PCB11, RRBI14, Ten15b, Til14, Til15, Til21, UB12, WLC^{+}19, WO14, YKZZ14, ZB15, AAO10, MRZU14]. Emergy-based [LYCU14, WLC^{+}19]. emigration [HKP10]. emini [MSG^{+}13]. emission [CGY11, CZW10, GNB^{+}14, HJS^{+}16, WC13a, WC13b]. emissions [CC11, CSH11, DZW^{+}11, FCZ13, KGS11, LHZS10, LHL18, LSM19, mNkGyT^{+}17, SSY14, TB19, WSL17, YC13, ZZY^{+}13]. Empirical [RHR^{+}15c, BCV13, DWD19, GDC^{+}19, HB12, KHKC^{+}18, KCC16, MPL^{+}16, MYJ^{+}14, RDD^{+}18, VPB^{+}12, WYHC15, WTWS19, dSD18]. empirically [ADF13, FG11, TVF^{+}10]. Employing [LdMS^{+}16, KGD14]. enable [LZB17, RP11]. enables [DJH10]. encounter [KNPH^{+}17]. encounters [CF14]. end [BDV^{+}19, CBC^{+}16, DBVS19, FLH17, HBS^{+}16a, MSCH14a, MSCH14b, RBGB16]. end-to-end [BDV^{+}19, DBVS19, FLH17, HBS^{+}16a, MSCH14a, MSCH14b, RBGB16]. endangered [CGV^{+}15, DHC16, DHS^{+}16, HBT^{+}18, KWGG19, LJ15b, LAEAG12, MFB17, PCRP12, TAC^{+}14, TL15, WAGM17, WSW^{+}16]. endemic [DUASLB^{+}12, LTGD16, MAO19]. endemism [HBSF16]. Endl [Oga11]. endosymbionts [TGM^{+}21]. endotherms [BM10]. endozoochory [DDB^{+}12]. enemies [BWH13]. Energetic [MSM^{+}17, ABAG15, ABAG16, LB11]. Energetics [LARZN13, SM13, ATL18, BKP10, Bar11, CB15, HKW^{+}18, LDS14, MAC^{+}12, Ske13]. Energy [Abe10, BB10, BM10, BMU10, DLL^{+}19, FAC^{+}19a, GPM^{+}16a, GPH17, GMvV17, JZ12, Ju19, LLC^{+}17, LCBNP19, PLS^{+}10, SJS^{+}13, SØB14, XYTS19, Abe15, AMH16, AS14, BLY11, BSASSJ^{+}13, BWC17, BK13, BPD^{+}19, CZC^{+}15, EIG^{+}10, ENM11, FKWZ19, FCZ13, FG11, HMP^{+}19b, HZZ^{+}18, HM18b, HBA16a, HBA16b, JBH^{+}14, JR16, JNHF18, JHH^{+}12, JHT^{+}14, JLN10, Jør15a, Jør15b, JC11b, KAJØ19, KL12, KBMM15, LKLK15, LT15, LTL^{+}13, LLW^{+}15, LZJ^{+}19, LDYL18, MPT^{+}11, MEC14, MH11, NLM^{+}19, Pie10a, RMGU11, RHR^{+}15b, RDR^{+}17, SGAS14, SWZ^{+}13, SG11b, UAB^{+}11, VFB19, WWC^{+}18, WYY^{+}19, WHNR19, WWYO18, XHC^{+}15, ZYFL10, ZLF^{+}11, ZZY^{+}13, ZQC15, ZLZ17b, ZMH19, JN15b]. energy-budget [HMP^{+}19b, JR16, JNHF18]. energy-related [LDYL18, ZZY^{+}13]. energy-saving [FG11]. energy-water [NLM^{+}19]. engineer [OCB^{+}13]. Englishmen [NGGFM19]. Engraulis [XCR^{+}13]. enhance [ACDC17, Hel10, KRL11, LWC^{+}18, LQW^{+}19, MCRB17].

Enhanced [CRO13]. enhancement [RAB^{+}10, WKZ^{+}19]. enhances [KCG18]. Enhancing [AADA14, KPS17, LAB^{+}10, RFZB12, VHM^{+}17]. Enigma [Pat16b]. enough [SP11]. enriched [TXM^{+}10, XTPT11]. enrichment [F.11, LLFL12]. enrollment [CLA^{+}12]. Ensemble [HG17, RXSK14, ZCS13, CGG^{+}18, CDM13, Fie12, GLY^{+}15, HGLZ13, PAS^{+}10, SGP18, YHJ^{+}17]. ensembles [MMMCAHM21, STD15, ZZ17]. ensure [MDW^{+}12]. entire [vGKB10]. entities [LM13]. entropic [Ske13]. entropy [BND19, DS21, GGV11, HLW15, Lin15a, MMS12, MK14, QL15, RHR^{+}15b, Sma13]. envelope [BBH^{+}17, WRB^{+}12]. envirogram [KPT11]. Environ [TLL^{+}17, Ano12s, MCW13, MJC^{+}11, WPB14]. environment [BCP16, BF11, CMD12, EST^{+}19, FSV^{+}12, FS18, HZSH17, Hei11, IKL13, JDZC14, LBWW14, LWGH13, LSJG17, LHH12, MSG^{+}13, MAO19, MRB14, MCL^{+}17, MGS15, NGRDTHRL19, Pat16a, PLHG14, RPGA15, WGE^{+}19, YKvWK16, ZHC^{+}16, ZZZW15, ZWG^{+}17]. environment-economy-society [HZSH17]. environment-sensitive [YKvWK16]. Environmental [AA12, FBU14, FLA19, FAT^{+}15b, JSB^{+}14, LG17, MJL^{+}17, SC18, WCZ11, XCR^{+}13, ABAG15, ABAG16, AFC11, ARV^{+}16, BRV11, Bec10, BMN11, BTB17, BDG^{+}14, BRW^{+}12, BvHR^{+}15, BCT15, BU16a, BHPF14, CPMP13, CdOR10, CBCB17, CBGK15, CZ11, CSS^{+}16, CRTS15, CDG15, CSL^{+}12, DLD^{+}11, DFD11, EMH10, EC10, FGPR10, Fre16, GCDA16, GdTCFG10, GJK^{+}15, HFU11, HKlL14, HFMM^{+}12, Hu16, IFC14, JMB^{+}11, KPS17, KK15a, KPT11, KWGW19, LAK^{+}21, LDS14, LJ15a, LST15, LHL^{+}17, LMGW16, LS13, LYCU14, LCT^{+}17, MTL^{+}10, MHMRAHHS19, MDH15, MLK17, MFI13, OCD^{+}13, PPB^{+}18, PMH17, PRL^{+}13, PANL11, RIMPBT19, RNE^{+}12, RRvdG18, SF10, San19, SHT^{+}21, SSV^{+}12, SBD^{+}19, SGA^{+}16, SJB^{+}16, SKM^{+}16, SWGC16, TCSG15, Tie16, TSSL11, UAB^{+}11, UDP17, VPB^{+}17, VG19, VKdW^{+}10, Wal16, WMK^{+}13, WRSL11, Yan11, YM21, YY11]. environmental [YTT^{+}11, ZLY16, ZBS14, ZAW^{+}18, dBGS^{+}14]. environmentalism [MRM16, Pat14a, Pat16b]. environmentally [BGG^{+}16b]. environments [ACKS13, ASS14, BRMW18, Che17, CV19, FMA^{+}12, FAC^{+}19a, HHvO^{+}17, HQF^{+}17, KYvdPS14, KASS12, KP10, Log13b, Mon13, RKRB16, RPGSCF16, SAT^{+}15, ŠK14, VP12, WR13a, WR17, ZBMC14, ZCL^{+}16]. Environs [Pat10, AG14, BF11]. enzyme [ABAG15, ABAG16]. EPIC [IMW^{+}17, QCLW18, XBvdV^{+}14]. epidemic [JSKK15, PTML14]. epidemics [BA14b, PKW^{+}10, SCBH18]. epidemiological [FBB^{+}18, RKW12]. epidemiology [DVL11, LPPA15, TCGC16]. epilimnion [CT11, HC11]. Epinephelus [CNI11]. Epistasis [CdOR10]. Epistemic [CWMC^{+}12]. equally [Pat12]. Equation [Vav10, CL10, MLG18a, MV17, MMC11, PC15, RHO^{+}12, RDH^{+}13, RHS^{+}16, RDD^{+}17, SIS^{+}14, SCHGM16, SGKR10]. equation-based [CL10, MLG18a]. equations [ACKS13, BBB^{+}17, HSD^{+}15, JOHW17, Sak13, SFR^{+}17, VRC17, VRL18, WSSM15]. equifinality [WC13b]. equilibria [DH18]. Equilibrium [MM19, GLS10a, LJAD17, PBP^{+}15, Wan19]. equity [RvS16]. Equivalence [BS19, BFS11, SBDA16, YTT^{+}11]. equivalent [GBP^{+}10]. eradication [ATCJ^{+}19, VBPY11, WTG^{+}16, YHK^{+}19]. Erie [JXL^{+}15]. Erik [Fat16b]. ERIMO [SAR11]. ERIMO-I [SAR11]. EROEI [BLY11]. EROI [GPM^{+}16a]. Erratum [ABR12b, BRV11, Bar11, ERGW^{+}11a, Jør15a, KMS^{+}13, LSJ^{+}19, MSCH14a, YTT^{+}11]. error [HTB13, HTB17, SP12]. errors [AGJ^{+}12, CBA10, PBO19, WBC^{+}12]. escapes [IGBSASSJ16]. Escaping [SDSH15]. esculentus [KSZW16]. ESEM [Ano11p]. Esla [SGRAdJ17]. Esox [BVP^{+}13, MMVAHMC16]. essential [LDG14]. Establishing [BPC^{+}17, CXW19, CBP^{+}17]. Establishment [PHC10, PR11, ABR12a, CVPH^{+}14, HKN17, RJ17, TLP17]. Estimate [NSN17, BMSLC16, BQF^{+}18, FG11, GYL^{+}15, LZZ^{+}14, NLMdL10, Pér18, PC15, PGK^{+}13, SBM^{+}11, SML12, SSN^{+}18, WV12]. Estimated [AHS^{+}17, SFL^{+}19, CGS^{+}17, COY^{+}10, DCZ^{+}15, MK12]. Estimates [QDH^{+}10, RWFD13, AOW13, BPC^{+}17, BV15, FE14a, FYWA12, FML^{+}17, GTC16, LLT^{+}14, LBS^{+}14, OMDP11, OKS^{+}11, PLHG14, RMTM^{+}16, SMT19a, SMT19b, SHM^{+}14, Tho10a, VHM^{+}17, WGA^{+}13]. Estimating [AGJ^{+}12, BM11, BAJ^{+}11, CWL^{+}16, CGB^{+}13, CW11, CMV^{+}16, FPO18, JSL^{+}10, KKP^{+}11, LVZ^{+}11, Log16, MSF18, MSO10, PPZ^{+}17, PWGS^{+}17, PKW^{+}10, RVNB11, RTE^{+}15, SRXP14, SA13, SSQ^{+}10, SG11a, Vav12, Wan19, Wol11, BWC17, CCCP17, CCW^{+}15a, CCW^{+}15b, CC10, Cho13, HVL^{+}19, Lec17, LLCR11, LWY^{+}11, Loe14, MGW^{+}13, MDKK15, PR13, RRHH10, SM12, SED15, SLZ^{+}17, VRM^{+}10, YCL^{+}14, ZZFH15, ZZL^{+}19, dLB19, vPTOW18]. Estimation [CCSF11, FN15, KW18, KBS^{+}15, LCL^{+}13, LLY^{+}13, LZZP16, LPS^{+}10, PBT^{+}11, PSSH11, VALC12, ZZY^{+}13, BSC^{+}15a, BDG^{+}14, BWA17, BMU10, CGAB17, DVT^{+}10, DFM12b, DG16, FHL^{+}13, KSZW16, KCP^{+}15, LBM16, MMC11, Ned15, PBO19, Qia12, RGHS11, RRG^{+}15, SBVA^{+}13, TŠAD12, Wol11, YLL^{+}12, ZCS13]. estimations [FvBM^{+}11, Til15]. estimator [BM11]. estimators [DHL^{+}12]. estuaries [CVPH^{+}14, FRV^{+}16, KM12b, KRCVF16, RML^{+}13]. Estuarine [BH10, DCMS18, BNMP11, BSFR17, DJS10, DG11, DWD19, DVD^{+}14, FPG11, FPWG14, HB12, IT17, KVU^{+}10, LMT13, LdMS^{+}16, MSFR15, RGU^{+}15, RBM^{+}17, RSC12, RSC13, San17, VASL^{+}15, WJL^{+}13, WSW^{+}13, YLY^{+}18, YKCN10, ZJL^{+}14, ZZ17]. estuarine-shelf [DJS10]. Estuary [BCB^{+}11, CLD^{+}17, PLR^{+}13, WSW^{+}13, WPB14, ABD14, BN10, DJS10, DWP13, DJMBD^{+}13, FS14a, GJÁ^{+}10, JJKJ11, KvTL^{+}18, MRG13b, TRD^{+}15, WQD^{+}13, DWP16, HB14, MDVG^{+}15, WLDL12, WDL^{+}15]. ethanol [ABAG15, ABAG16, BB10, YC13]. Etheostoma [MGWW13]. Ethiopia [FSML11, TW18]. ETM [VB10]. EU27 [MCPB15]. eucalypt [HS11]. Eucalyptus [CN11, NB11]. Eulerian [DFL^{+}16, IFC14]. Eulerian-Eulerian-Habitat-Selection [DFL^{+}16]. Euphausia [GJK^{+}15, MRCW10]. Euphausiid [TWH^{+}14]. euphotic [UDRA13]. Eurasian [AGQP^{+}14, PGZ^{+}19, SGA^{+}16, TAC^{+}14]. europaeus [RPGA15]. Europe [GPTR13, GCK^{+}15, GBP^{+}10, GB15, IPSH14, MCWG21, OJ15, PKD11, SLSS15, SFB^{+}15, TAC^{+}14, TNA^{+}17, TKH17, WGSR19, ZZB^{+}18].

European [HLLO13a, BLB^{+}17, BKMK10, BMU10, HGE^{+}10, HLLO13b, LCBNP19, MBD^{+}12, OAPP12, ODK16, PRL^{+}13, PAGCV^{+}10, PSKO16, VPC^{+}14, WFN^{+}16]. Eurytemora [DJMBD^{+}13]. euthanasia [MCT10]. eutrophic [ALC19, CZR^{+}14, CPS15, CTP^{+}17, LCY^{+}13, MGMR15, SØB14, SHM^{+}17, WB17]. eutrophication [BNMP11, CKH15, DMZ^{+}19, HFB^{+}11, HhLY^{+}19, LkSXhX17, LL18b, NGGFM19, PAB13a, WSW^{+}13, XWC^{+}11, XPY^{+}14, ZCSL18, ZDG^{+}13]. evaluate [BGD^{+}15, DHC16, FRV^{+}16, FP13, GJM^{+}13, GLS10b, HHF^{+}10, LBSB^{+}14, LCD^{+}16, LL18b, PRH17, RR10, SHX^{+}12, TAR13, WSE10, WYPF15, dMLM^{+}17, vdVJS16]. Evaluated [CVPH^{+}14, MSGG14]. Evaluating [BvHR^{+}15, BCD12, Cha12, CZ11, CAC^{+}16, CDM13, DLD^{+}11, FSM^{+}19a, FMF12, GGV13, GPD^{+}13, HCZ^{+}17, HSM^{+}12, KKR19, KK10, LLZ^{+}16, LHLW15, MFB17, MCP16, MWD^{+}15, PWBG16, PGW19, PMCKH13, PPDF14, RMHE15, WGA^{+}11, WGA^{+}12, WR13a, YKZZ14, ZJP^{+}11, dLB19, CSF11, DZS12, EEB^{+}17, GZVU14, KHW^{+}16, Lin10, Lin16, LC13, Mar12, MGCL17, NMT^{+}19, OWMK10, SXZ^{+}15, SLRC19, VPB^{+}12, WTWS19, WWYO17]. Evaluation [AAO10, CPS^{+}14, GS19b, KJS13, LJ15b, LYSW15, LYD^{+}17, LDYL18, MH10, NL12, Pat14b, PICR13, ROSC15, REB^{+}16, SFMA15, SHM^{+}14, SLZ^{+}18, SLY^{+}17, WC13a, WC13b, XGH^{+}17, YHK^{+}19, ZZB^{+}18, Abe13, ALL11, AFC11, AVG14, BTM^{+}10, BDV^{+}19, BFJZ^{+}11, BCO17, BOF^{+}16, BU10, FMBD15, FKK^{+}18, FDM^{+}11, GH13, GBB13, GCV14, GHH^{+}15, GHS^{+}16b, HPJ15, HHL17, IS13, ISG13, IMW^{+}17, JC11a, JGB15, JC11b, KWS^{+}13, KVP16, LRZ12, LHL^{+}17, LHB^{+}18, MCWG21, MRZU14, MCPB15, MGW^{+}21, MJGC13, OGOWW16, PBB^{+}14, QWJL19, SC21, SSY14, SA16, SMG^{+}12, SLRC19, TPT^{+}11, VHM^{+}17, VFFP16, WZGH18, XPY^{+}14, ZLC13, ZWG^{+}13, ZLZS14, ZHV14, vdVBJS15]. Evaluations [NiSK21, CLL14, HT14, JGF14, MCPB13]. evaludation [AVG14, CLL^{+}15]. evapotranspiration [AMRH19, GZK^{+}11, KBS^{+}15, LLM^{+}13, LLC^{+}21, LPS^{+}10, LPX^{+}12, SHT^{+}21, TYEH10, WJK11]. Event [Bev15, BNMP11, WGD^{+}17]. Event-persistence [Bev15]. events [dSCdA19, CTP^{+}17, GCC^{+}10, Jam16, KUK11, NL12, PCTW17, RMvW13, WB16]. Everglades [WMW^{+}12, BGD^{+}15, JDGJ13, MM16, TXM^{+}10, XTPT11]. evergreen [Bae18, SMSAN^{+}14]. Evidence [BAAB18, BS10b, GYL^{+}13, McC16, SB11, YKL^{+}12, Pér18, PAB^{+}13b, RDD^{+}18, SBCB13, GDC^{+}19]. evident [BGGL16]. evolution [BBL11b, CH12, CH19, DR17, Els10, GAC^{+}14, HS13a, HM18b, HQF^{+}17, MV17, PKMH16, RNS^{+}19, SGBD10, SZ14, Sin17, SPWR17, Vav12]. Evolutionary [GQRC^{+}17, GMC^{+}19, RKRB16, AMG19, ARV^{+}16, CRO13, CRB16, DVT^{+}10, GKL18, KPM^{+}14, KMKL19, PSWW11, ROC^{+}13, RHR^{+}15b, RHQ^{+}16, RDV^{+}17, Tie11, ZRC^{+}15]. Evolving [DHHB15]. EwE [HMA10b]. ex [ESA12]. Exact [CT12, SJTS13, Log16]. Examination [RPSA12, GCvDM12]. examined [HW12]. Examining [BEZ19, SGKR10, PFA^{+}16]. example [AWL^{+}13, BMD18, CLBD^{+}11, DÖWB12, GWC^{+}10, qHLlZ^{+}16, HB12, JTM^{+}17, JBCÑ11, LJ15c, Ned15, PWP^{+}19, PTL14, PPDF14, SSK14, SHB11b, SGHW^{+}17, WAGM17]. examples [FDHL17]. excess [ENM11, ZJB19]. exchange [BDTV11, KRL^{+}16, KHW^{+}17, KVU^{+}10, KK12b, MKTO19, QlSeZ^{+}10, RML^{+}15, Wan19]. exchanges [DJS10, HBA16b, HAA^{+}21, WRM^{+}11]. excitations [DG11]. exclosure [LLKY18]. Exclusion [LM19, CEWW18, GO15b, MdlP15, RHS^{+}15, RHD^{+}16]. executable [VC17a]. exemplified [KØ13, WØ15]. exercise [HKW^{+}18, KS18, LBZA19]. exergies [XWC^{+}11]. Exergoecology [TVV13]. Exergy [Sci11a, SWC13, AAW^{+}15, BU10, BPD^{+}19, CCC11, CZCL18, CJC11, CRTS15, Hoa11, JZCL14, JC11a, JMSM12, JN14, LBK^{+}13, LRB12, MPT^{+}11, MJB^{+}11a, SZ14, TA11, VFB19, XYCZ11, ZGJ10]. exergy-based [CRTS15, LRB12, SZ14]. existence [SG11b]. existing [EH15, EH19, HLS^{+}14, HKG^{+}17, KR10b, LTL^{+}19, SMBGH^{+}19]. exotic [CZCL18, EBL^{+}11, MJB^{+}11a, SZSV12a, SGM^{+}14, TW18, XCG^{+}16]. expanding [JJMW14, WJL^{+}13]. expansion [GSJ^{+}10, LYL^{+}17, LCPB^{+}10, RHO11, RRJVA13]. expansions [LRM^{+}14]. expected [FL11]. experience [BRM^{+}14, DLL^{+}19]. experiment [BGGL16, Gra14, HH12, JG10, MSG^{+}13, RFV^{+}12, TNA^{+}17]. Experimental [HD16, Ken19, LWTZ18, MG16]. experimentally [MCH15]. experiments [BVGR^{+}17, BS10a, DDR^{+}14, For18b, HHK19, LBZA19, LLKY18, MHRB19, PLS^{+}10, PGR^{+}16, RDJ19]. Expert [SSD^{+}17, BBH^{+}17, Cha12, CPE13, JBV^{+}15, JKS^{+}10, KCF17, KCF18, KdVRdW17, MSHC13, PSM^{+}12]. expert-based [KCF17, KCF18]. expert-driven [KdVRdW17]. expert-selected [BBH^{+}17]. experts [AWF10, MMD^{+}17]. explain [FHAG12, FS14b, TBL^{+}15, VAO^{+}17]. explained [BMF^{+}11a, BMF^{+}11b, BCC^{+}14, KMKL19]. Explaining [MMS12, War10]. explains [ACE^{+}17, Loe12, POB18, VH12, WHNR19]. explanation [vLTdBP16]. explanatory [KPS17]. explicit [ARTBB14, BSP12, BEPC15, BJR16, BBL11b, BBH^{+}11, BS16b, BS17, BPN12, CRB16, Che17, DC10, DVF^{+}19, DVCM^{+}10, DVL11, FBKS17, GPZB18, HSHF^{+}13, HHC15, JOHW17, KR10a, LF11, LYC^{+}15a, MLG18a, McC16, MLAT^{+}18, NMT^{+}19, NPJ^{+}17, PKB^{+}18, RMP^{+}18, SBC13, SYL^{+}13, Sto10, Str15a, Str15b, TVF^{+}10, WGT12, WFW^{+}17, WHRR15, WNF^{+}15, WHC^{+}13, ZRC^{+}15, dMSL^{+}16, vdKGV11]. exploitation [BAA^{+}14, BLY11, RAdA12, RCP17]. Exploitative [BRMW18]. exploited [BRL12, CD15, DR17, DML15, MZLCSZdP11, PK16]. exploiting [DNFG18, LKDD15]. Exploration [Til14, CBB19, Cho13, WLDL12]. explorations [Ben12]. explore [HM18a, HM19, HMP^{+}19b, JDGJ13, LKH^{+}18, MTZ^{+}10, MV11, MD16, MP17, PDAL14, PTE^{+}12, PHP15, RAB^{+}10, RDD^{+}14, SJGB16, SGWV15]. explored [CBA10]. explores [WNT^{+}13]. Exploring [Ano12s, ChHC12, CAMCK^{+}12, HGE^{+}10, KCF^{+}13, LLZ^{+}15, LGY10, LMBL12, LB18b, MHL^{+}19, MCW13, MDJ^{+}11, MRR^{+}14, Mur12, NOOM^{+}16, PÓF^{+}18, PME^{+}16, RVL^{+}18, Rem11, RHQ^{+}16, RDV^{+}17, WAS^{+}16, WCG^{+}11, dMSL^{+}16, LZS^{+}17, YJW15]. exponent [SIS^{+}14, SSR16]. export [BH17, LRM^{+}14, SKH^{+}13, XHC^{+}15]. exports [DVD^{+}14]. exposed [Rai13]. Exposure [CKH15, dABLBAdOR17, EGMW12, ECO^{+}14, GV11, KSM^{+}18, LSGT13, LBS^{+}14, MAP^{+}13]. expressed [MSO10]. extend [HAOS17]. Extended [MLMH18, WSSM15, CN18, FR11, Mer12, Sci11a]. Extending [KM12a, LKKK^{+}14, Fat14, GAB12]. extensible [VC17a]. Extension [RDR^{+}17, AMG19, CBCB17, DDM^{+}17, KWS^{+}13]. extensions [SBC^{+}16a]. Extensive [FDHL17]. extent [FYWA12, VPM^{+}18]. extents [YHS^{+}11]. external [MMM^{+}16, SCG^{+}13]. extinction [BCV13, BSC^{+}15b, CHvdW^{+}10, CMGF14, CRVS13, Flo14, GVGZ^{+}21, HBS16b, Lab11, LKG10, LLH^{+}13, LYC^{+}15a, MM19, Mat13, Pec12, PTA^{+}12, PLHG14, PL10b, San19, SW15]. extinctions [SRTS18]. extrapolating [HFSL14]. Extrapolation [MHRB19]. extraterrestrial [KBS^{+}15]. extreme [BNMP11, CVV^{+}13, CMGF14, CTP^{+}17, GCC^{+}10, Jam16, YYH19, ZHLZ17]. extremes [BTS19, DFW^{+}18]. Exurban [PTA^{+}12]. eyes [MPF^{+}14].

F. [YWA^{+}17, MAO19]. F1 [KYvdPS14]. face [BMP^{+}19, DHS^{+}16, LLL17, MRCW10, Rin18, RTE^{+}15]. facet [Mat17]. Facets [Pat14a, Pat16a]. facilitate [KK11]. facilitated [LMY^{+}11]. Facilitation [STJ15, BPN12, GM10, XM13, XZZ^{+}13]. facing [DBS18, ISG13]. factor [TPO^{+}18, YqWB^{+}15, ZZY^{+}13]. Factors [FR11, HJS^{+}16, MSCH14a, MSCH14b, AFC11, CHPS17, CBGK15, CHG^{+}15, CKH15, DK17, DH18, DMZ^{+}19, FBA^{+}11, GSSM13, GCR18, HKlL14, KPT11, LMR16, MHL^{+}19, MHMRAHHS19, MR17, MFI13, PSWW11, PCZ^{+}16, PVB16, QlSeZ^{+}10, RDD^{+}18, SBM^{+}12, Sci11a, UB12, UDP17, VEMS12, XCG^{+}16, CRAP13]. facultative [SRHL11]. failed [GHWG^{+}14]. failure [RGH^{+}13, RBH13]. false [BLW^{+}17]. family [TT17a]. Fanjingshan [AA12, WALCY12]. farm [AHS^{+}17, BSASSJ^{+}13, DMB^{+}18, GPM^{+}16a, GWC^{+}10, LJ10, TB19]. farmed [HD14, HD16, KGM12, PAGCV^{+}10]. farmed-fish [HD14]. Farmer [Fai10, KHHN17, VBS14, WYHC15]. farming [BBC^{+}19, DMBS11, FLLB11, FDHL17, IGBSASSJ16, SGBD10]. farmland [HLLO13a, HLLO13b, KEHG17, ZLW^{+}15]. farmlands [MKSL^{+}17]. farms [AHS^{+}17, Ken17, RJ11]. farmyard [MRP^{+}11]. Fast [LH15, WGL^{+}12, HNWT10, MdlP15]. fast-growing [HNWT10]. fastigata [CN11]. fate [BWA17, CAMCK^{+}12, FtHvdB^{+}14, HGWL11, MLS^{+}11, XQZ^{+}13]. fates [MDF^{+}18]. fatty [SAAD16]. fatua [BCL^{+}14]. fauna [PLFO^{+}12]. faunal [CZCL18]. favorable [HHUD14]. favourability [BRV11]. Fayle [GU11]. Fear [HB14, TGKSJ18]. feature [MBHW12, PLB16, TTL^{+}18, WFW^{+}17, WKF^{+}17]. features [BTPG10, CLBSG11, FS14b, GKF^{+}14]. February [Ano10w, Ano10x, Ano11-34, Ano11-36, Ano12y, Ano12-33, Ano13-33, Ano13-31, Ano14j, Ano14s, Ano15q, Ano16e, Ano16h, Ano17-29, Ano17s, Ano18y, Ano18-40, Ano19-39, Ano19-45, Jör11a]. fecundities [Dil10]. fecundity [DP19, EEFE^{+}11, NLMdL10, PSG19]. fed [AJDS16, MTPC13]. federally [GKF^{+}14, WZL^{+}16]. federally-threatened [GKF^{+}14, WZL^{+}16]. feed [KAJØ19]. feedback [AHM^{+}16a, FTYI13, MDJ^{+}11, MCA18, PP16]. Feedbacks [BSS^{+}14, Agu11, BGP11, CGM^{+}14, DBVS19, KR16, LL17, NWAR11]. feeder [KMRW16]. feeders [BPK^{+}16, CN10a]. Feeding [SvBH^{+}14, AHS14, FS18, JNHF18, KC18, PKT15, SvdMKS11, SAI10]. feedstock [SØB14]. feedstocks [HGN^{+}16]. felled [CRD^{+}16]. felled-tree [CRD^{+}16]. felling [OJ10]. fen [DBG14b, DBG14a, WMK^{+}13]. fencing [ACSR^{+}13]. feral [AOW12, IN16]. fermentation [PJ16]. fermentations [PGG14]. fern [SZSV12b, SZV14]. fertilisation [RMTM^{+}16]. Fertility [MS14]. fertilization [CH14, GHS^{+}10, LBdA^{+}19, PRP17, RLS11]. fertilizer [BKN14]. fever [JVY19]. FFSM [LDCB21]. FIA [QWJL19]. fiber [ZML^{+}12, PGZ^{+}19]. fidelity [MBG17]. Field [Ano12a, PF18, BCL^{+}14, BvHR^{+}15, DTS^{+}12, DPBK10, FDP^{+}10, GNB^{+}14, GCvDM12, HVL^{+}19, HhLY^{+}19, JG10, JAKJ10, LBM16, NPH^{+}17, RW10]. fields [BKN14, BC11, GCK^{+}15, MKTO19, SNS^{+}12]. fifteen [CW11]. Figured [AA12]. figures [Jør14b]. filiculoides [SZSV12a, SZSV12b, SZSV13, SZV14]. Filling [MWML13a, SBC16b]. Filter [HG17, ZCS13, CN10a, HGLZ13, SG11a]. Filtering [CC14, BOGA14, FBA^{+}11, LJ15a]. filters [LJ15c]. filtration [EH15, EH19, KTGH19]. final [MPR13]. financial [VMP^{+}19]. Finding [GMP13]. findings [SW11]. Fine [MGR^{+}18, ABD^{+}18, CCLX17, GGV10, KMRW16, MCR^{+}15, WMF^{+}13, YHS^{+}11]. Fine-scale [MGR^{+}18, ABD^{+}18, GGV10, KMRW16]. Finite [MEC14, DBP16]. Finland [LSJ^{+}19, mGZK^{+}10, GWK^{+}12, GKW^{+}13, HFU11, LSJ^{+}18]. finless [MCL^{+}17]. Finnish [KLS^{+}14]. fiordensis [JCKG15]. fir [KKP^{+}17, LPX^{+}12, WZD^{+}14]. fire [AD13, AWWD16, ASP^{+}15, BRR10, BGP11, CADB19, CMSR19, CPD^{+}11, CCL14, CAY^{+}10, CSFR15, DWP11, DSBS15, DPBS10, FMS13, FAMD^{+}12, GLT10, GTM^{+}19, HWE^{+}18, KK10, KDK^{+}10, KCF^{+}13, KHS^{+}19, LM19, LCML11, LHLW15, dMHGA10, MRC19, MMGK15, PMM11, PW10b, RKC^{+}15, SKHH19, SVH^{+}19, SBF15, SGG14, TL10, TT12, TLS^{+}16, UR10, VB10, ZHLZ17, RR10]. fire-killed [HWE^{+}18]. fire-maintained [LCML11]. FireBGCv2 [RKC^{+}15]. FireBGCv2/GrazeBGC [RKC^{+}15]. fireflies [NiSK21]. firefly [BNL^{+}13]. fireline [MRC19]. fires [DBR12, GdTCFG10, TXM^{+}10, WFA10, WSZT10, XTPT11]. First [Ano13u, BLH10, CGV^{+}15, GAC^{+}14, SFvdM^{+}17, ARTBB14, ABGB15, GBD^{+}10, VMP^{+}19, WHC^{+}13]. first-order-based [WHC^{+}13]. Fish [DTC^{+}10, FIG^{+}10, MR10, PC17, Pen11, QWJL19, SM13, BAAB18, BAA^{+}14, BSASSJ^{+}13, BR16, BGC10, BPB14, BRS^{+}18, BSO13, CRBC11, CW17, CRUF13, CBCB17, CWA^{+}15, CD15, CSL^{+}12, DFL^{+}16, DC10, DATI10, DPPI14, DR17, DA19, DGKSM13, FE14a, FTTV^{+}18, FDM^{+}11, FPG11, FPWG14, GDTP13, GV11, GMTC^{+}17, GHH^{+}15, GCL^{+}19, HD14, HTL^{+}14, IPSH14, IGBSASSJ16, JDP^{+}12, KFWM19, KHW^{+}16, KUI^{+}18, KTK^{+}15, KBL10, KEM^{+}16, KK15b, LHZ10, LCCL15, LWY^{+}11, LMCA^{+}12, LCTH12, MSP^{+}13, MWK19, MAC14, MGCL17, Md18, MMRS^{+}19, MSCH14a, MSCH14b, MDV^{+}12, MCP16, MN12, MMMCAHM21, NL12, ODK16, PH18, PGKG12, PGATS13, RHW17, RKRB16, RMHE15, RSD^{+}15, RDD^{+}14, RPP^{+}15, SAAD16, SFC^{+}12, SGH16, SSM^{+}13, TSN15, TL15, TTCC^{+}19, VSF10, WL13, WhFxGZ16, WAGM17, WFN^{+}16, WZL^{+}16, YDT^{+}13, ZSG12, ZZ17, ZHV14, ZVS18]. fish [dMSL^{+}16, dMLM^{+}17, vDHH^{+}13, vDHH^{+}14]. Fish-Fisheries [SM13]. Fisheries [SM13, AdMC^{+}13, AHM^{+}16b, BWRA16, CGLA10, CF17, CJR^{+}16, CMCASS12, CMCC15, DPR10, FSM^{+}19a, HCE^{+}19, IGR^{+}18, IS13, IGBSASSJ16, KSAS12, KMMS17, MTZ^{+}10, MMD^{+}17, MKL14, MSM^{+}16, NOOM^{+}16, NCEGB16, PBEL17, PGV10, ROSC15, RvS16, RUSL19, RKNO17, SLW17, TLP17, VQS^{+}18, WHRR15, ZCR16, dMSL^{+}16, dMLM^{+}17, vDHH^{+}14]. Fishermen [vDHH^{+}14, WMCM12]. Fishery [WLDL12, ACE^{+}17, CRA^{+}11, DBS18, GPR^{+}13, GYL^{+}13, IKL13, ISG13, JBCÑ11, LDO^{+}18, LJ13, LMBL12, MROGVR12, MTGD17, PLS^{+}16, PPDF14, RAB^{+}10, TSN15, TSN17, WKZ^{+}19, WMCM12]. fishes [FBB^{+}18, FBH^{+}14, JD18, KC18, NCMF10, NMA^{+}19, NiSK21, PWBG16, SMT19a, SMT19b, VEK12, WCF^{+}16, WC19, XWW^{+}11]. Fishing [dSdCF10, BWRA16, BS13, CGLA10, CSF11, CSS^{+}16, FLLG13, FTTV^{+}18, GH13, HPC13a, HPC13b, JBCÑ11, PKG16, RMR^{+}16, RMHE15, SGH16, TSN15, TCH^{+}13, VRM^{+}10, WDL^{+}15]. fit [BG19, SP10, WZL^{+}16]. fitness [DSJ19, Hel10, KH15, LJB12, Log16, ŠK14, YTT^{+}11]. fitted [CRD^{+}16]. Fitting [DAP18, HTB13, HTB17, ČTT^{+}12, YSB11]. Five [GSA^{+}19, BDL^{+}14, GdSHA10, SW19, TBM^{+}14, YSA^{+}14]. Fixed [LJ13]. fjord [JCKG15, MTPC13]. FL [RWvDM15]. Flanders [EBL^{+}11]. flat [COK^{+}11, MB14]. flatwoods [WSBP13]. flavescence [LPPA15]. fleet [ISG13]. Flexible [CSOS17, HB13, MMD^{+}17]. flies [ABR12a, DATI10]. flight [RDJ19, ZSS17]. floating [MXC13, McC16]. flock [LJ14, ZSS17]. flocking [KKI^{+}14, ZSS17]. FloMan [DA17]. FloMan-MF [DA17]. flood [BNMP11, CSN^{+}19, CGGL16, CMLY19, DVF^{+}19, LMCA^{+}12, WP14, XLS11]. flood-pulse [CSN^{+}19, CGGL16]. flooded [BBDB13, KRL^{+}16, SKH^{+}13]. flooding [CLH17, HR11, LCP14, MGR^{+}18, WYS12]. Floodplain [DA17, BRKP16, DVF^{+}19, GAP^{+}19, LKH^{+}18, WYS12]. floodplains [WRSL11]. floods [CHTC11]. floor [JAH^{+}10, ZJM^{+}11]. floral [GPA^{+}13, MHTC15]. Florianópolis [dHR18]. Florida [WMW^{+}12, CAMCK^{+}12, DA19, DWP11, MM16, YSB11]. Flow [CNI11, AAW^{+}15, AHNK13, BPK^{+}12, CT16, CD17, CZW10, CCL^{+}13, CDT10, DLCR13, DWP13, FAC^{+}19a, GHH^{+}15, JD18, LBD10, LBW18, LWY^{+}11, MLY^{+}19, MNK^{+}21, MB17, MGWW13, NL12, SBVA^{+}13, SLB^{+}18, TNHB12, VLZ^{+}19, WWC^{+}18, WZL^{+}16, YM21, YY11, YLW^{+}11, ZWJ^{+}12]. flow-dependent [GHH^{+}15]. flow-quantified [BPK^{+}12]. Flowering [KR14, JX15, KKL^{+}12, RPGA15]. flowers [SGM^{+}14]. flows [ALL11, BPD^{+}19, CZ11, CGEdM12, GRRU15, Jag14, JHH^{+}12, KAJØ19, LKLK15, LCT^{+}17, MGR^{+}18, MRZU14, PPB^{+}18, Pat15, PCB11, SS11, WFM16, Yan11, YYQ11, ZPZ^{+}15]. Fluctuating [vDHH^{+}13, CEC10, GEB10]. fluctuation [QWJL19]. fluctuations [BLRN15, PRL^{+}13, RPGA15]. fluid [DWNN10, DBP16, HSvO^{+}18]. fluminea [GCDA16, HFH^{+}19]. Fluvial [GJM^{+}13, OOH^{+}12, WL13]. Flux [KHKC^{+}18, LXS^{+}16, Bai18, CSS^{+}19, GKW^{+}13, HCL^{+}14, LYJ^{+}15, LHR^{+}15, MPL^{+}16, PD15, QlSeZ^{+}10, WZD^{+}14, XKI^{+}11, XGH^{+}17, ZZB^{+}18]. Flux-Based [LXS^{+}16]. fluxes [BSASSJ^{+}13, CSH13, HMS^{+}11, JD10, JNC^{+}11, KMM^{+}13, Lin15a, MLH^{+}17, RHZ^{+}18, SMH16, WR18, WJK11, WJvdL^{+}13, ZZ17]. fly [LDM15b, MRES17, Pec12]. Flycatcher [BSP12, HPS10]. flying [RDJ19, TDF^{+}19]. focal [BPK^{+}16]. focus [CGL15, GRB12, HFMM^{+}12, OOFVBC16, VRL18]. focused [LHH12]. focusing [VRC17, WBS13]. focussed [LVB^{+}15]. foliage [CRD^{+}16]. foliar [LBR^{+}17]. following [KWGG19, LCH12, OGG^{+}16, PvVC15, RR10, WFM16, YHK^{+}19, ZKSF17]. Folsomia [HFSL14]. fonticola [MGWW13]. Food [BRPK14, CGGL16, DGKSM13, ESLJ18, FSML11, TCH^{+}13, ABAG14, AdMC^{+}13, BNMP11, BFP^{+}18, BCV13, BSH17, BQF^{+}18, BSZ^{+}11, BPK^{+}12, CEC10, hCjHMW18, CKS^{+}17, CPS^{+}19, CPS15, DCZ^{+}15, FE14b, Fuj16, GLC^{+}15, GYL^{+}13, HBS^{+}16a, HMP^{+}19b, HCL^{+}16, Ju19, KKH^{+}17, KSAS12, KEM^{+}16, KHL^{+}16, KHHN17, KCC16, LJR14, LSR16, LBSB^{+}14, LLKY18, LCTH12, LLFL12, MKL14, MDL^{+}10, MSFR15, NOOM^{+}16, NCMF10, NCEGB16, NMC^{+}14, NMS^{+}14, OGG^{+}16, OHS^{+}17, PJ17, Pea15, PLS^{+}10, PTL14, PICR13, PMBB^{+}16, RGU^{+}15, RBM^{+}17, RSE10, RMR^{+}16, SVD^{+}13, SBVA^{+}13, SGAS14, SBR^{+}19, SFHK12, SKM^{+}16, SCG^{+}13, SS11, TPL^{+}15, TM12, TRD^{+}15, TNHB12, TVV13, VASH^{+}16, WLY^{+}15, WTWS19, WYY^{+}19, WSE10, XZCL16, XYTS19]. Food-Energy-Water [Ju19]. food-for-protection [WSE10]. Food-web [BRPK14, TCH^{+}13, BSH17, CPS^{+}19, GLC^{+}15, HCL^{+}16, LJR14, LBSB^{+}14, MDL^{+}10, OGG^{+}16, OHS^{+}17, PTL14, PICR13, SKM^{+}16, TRD^{+}15]. foods [KNU^{+}15]. foot [DVL11]. foot-and-mouth [DVL11]. footed [CAMNN16]. Footprint [NGR^{+}11, FCH^{+}17, GWZ^{+}15, LZJ^{+}19, SWC13, WFSH15, ZTL^{+}17]. footprinting [PMH17]. footprints [FDR^{+}17, LDYL18]. Foping [LCS11]. forage [AHF^{+}13, KMRW16, KEM^{+}16, KK15b, KLGL19, WCF^{+}16, WC19, YDT^{+}13, dJGM^{+}18]. Forage-SAFE [dJGM^{+}18]. forager [LDS14, RAdA12]. foragers [BPWB14]. Foraging [BPM^{+}17, MCN16, MYJ^{+}14, BPM^{+}15, BPWB14, CAMNN16, DHC16, EIG^{+}10, HFMM^{+}12, KKI^{+}14, LGMLD19, LJ14, LDB14, LTGD16, MHTC15, MSH^{+}19, MvDL^{+}10, MRES17, OE12, OBSL15, RJ11, Rin14, SRP18, SGA^{+}16, TGKSJ18, VSPHS15, War10, WCG^{+}11]. foraging-bioenergetics [DHC16]. force [CWLL15, LJ14]. forces [JGB15, PBD19, SGBD10]. forcing [HPRNP12, JSB^{+}14, MBD^{+}18, MGS16, MCN16]. forecast [ABR12a, CHG^{+}15, RXSK14]. Forecasting [CVM^{+}18, DFH^{+}14, Hui11, CRB16, FCH^{+}17, MPP^{+}15, RSC12, RSC13, ZRC^{+}15]. forecasts [CDM13]. foreland [LJCG15]. Forest [CPS^{+}14, CSFB17, DMDJ13, FMF12, FAPM12, HSK17, PP16, SKTG11, VZF^{+}11, ZHLZ17, AD13, AMF12, ABD^{+}18, ACM15, AD16, Ant14, ARM^{+}12, BFA11, Bae18, BSC^{+}15a, BVC^{+}18, BPS^{+}11, BLD^{+}10a, BlMG^{+}11, Ber13, BSS^{+}14, BS16b, BW12, CZC^{+}15, CPR10, CAC^{+}16, CSH11, CLL^{+}15, DDM^{+}17, DVF^{+}19, DCMK11, DYMYGO19a, DYMYGO19b, DBG14b, DBG14a, DH12, DLL18, ENC^{+}15, FAMD^{+}12, FBdP^{+}16, FSS^{+}17, FAS^{+}19, FCA18, GLA11, Gar17, GBG^{+}15, GZK^{+}11, GVV^{+}16, GCK^{+}15, Gra14, GRL18, GSS^{+}14, GS19c, HMS^{+}10, HS11, HJS^{+}16, HBG^{+}16, HR11, HKN17, Hol13, HWL17, HOŠ^{+}12, HNWT10, HAA^{+}21, JAH^{+}10, JNC^{+}11, KWH14, KAF^{+}14, KUK11, KLS^{+}14, LJLC15, LJAD17, LAD11, LKLK15, LTL^{+}13, LZ10, LHW^{+}15, LLH^{+}17, LDP14, LCZ11, LNB17, LSP^{+}10, LL18a, LHR^{+}15, LDCB21, LM19, Log19a, LSR13, LHLW15, MKH^{+}17, MBHW12, MBSF14]. forest [MAC^{+}12, MWML13b, MdMS^{+}19, MP17, MOM^{+}11, NB11, NBD^{+}13, Oga11, Oga17, Oga18, OWMK10, OKS^{+}11, OLG^{+}13, PSWM^{+}10, PMS^{+}17, PBO19, PMB^{+}19, PHC10, PD12, PGK^{+}13, PLS11, PFR15, PGA^{+}11, RFV^{+}12, RHZ^{+}18, RKC^{+}15, RV15, SMSAN^{+}14, SHT^{+}21, SSV^{+}12, SHB^{+}11a, SPR^{+}18, SFF^{+}11, SRSS12, SHX^{+}12, SKMB11, SHM^{+}14, SVH^{+}19, SSK14, SMH16, SBPZ12, SPG11, SYY^{+}19, SMBGH^{+}19, TCSG15, TL10, TT12, TNA^{+}17, TYEH10, TLS^{+}16, UKI^{+}11, VAB^{+}18, VHM^{+}17, VALC12, WFA10, WPZ^{+}11, WGA^{+}11, WGA^{+}12, WGA^{+}13, WSL17, WJvdL^{+}13, XKI^{+}11, YHS^{+}11, YYH19, YBA^{+}16, YZA^{+}18, ZBAY14, ZJM^{+}11, ZLB14, ZXD^{+}13, aGSCMF15, dBGS^{+}14, vGKB10, Kos17]. forest-fire [TL10]. forest-floor [ZJM^{+}11]. forest-grass [SYY^{+}19]. forest-grassland [RKC^{+}15]. forest-wood-chain [PSWM^{+}10]. forest-wood-chains [LSP^{+}10]. forested [DBG14b, DBG14a, GTC16, LW14, ZJM^{+}13]. ForestMAS [KGD14]. forestry [BHPF14, GGSP14, HCS^{+}10, MDKK15]. forests [AKvO^{+}10, AMR16, ABC^{+}17, BHD16, BFH14, BS17, Bre12, CVV^{+}13, CNH^{+}11, CCL14, CPS^{+}14, DBG14b, DBG14a, DH12, ESP^{+}17, EKB18, FBdP^{+}16, FT16, FPG^{+}10, Gat18, HMP19a, HMS^{+}10, HSVL12, KGYG15, LBdA^{+}19, MLC^{+}16, MVC11, MK12, MWML13a, MPL^{+}16, MCG19, MJF19, NB11, PDD^{+}18, PGA^{+}11, RJH^{+}10, RWFD13, SSK14, SNZ^{+}19, THJ^{+}12, TDH12, TLRL11, ZLB14, SHB^{+}11a]. Fork [MSB^{+}17]. form [LLCR11, MSG^{+}13, XZSL17]. formaldehyde [LLC11]. formalism [RHR^{+}15b]. formation [BH17, BdCS11, BEBA19, CKS^{+}17, CSS^{+}17, KCB^{+}17, Mer12, PLS11, VRC17, VRL18, VAO^{+}17, WZWL10, XJY^{+}16, ZML^{+}12]. forming [JGB15, MG18, MHG18, WFW^{+}17]. Formosan [BCTL16]. forms [BMCG12, OOH^{+}12, Rai13]. formulate [GDBB19]. formulation [DFM12a, FDM^{+}11, KTGH19, KCB^{+}17]. formulations [ZJM^{+}11]. fornicata [MG18, MHG18]. fortunei [dÁSMRT^{+}19]. Foshan [HJ17]. fossil [Cam15, RMGU11]. foster [BMD18]. fountain [MGWW13]. Four [LFH^{+}15, FKK^{+}18, FCV^{+}12, LBZA19, MCR^{+}15, SBDA16, WKZ^{+}19, ZYFL10, ZZFH15, ZQC15]. Fowler [MBG17]. fowleri [MBG17]. foxes [BG15, MS14]. FPAs [HCE^{+}19]. fractal [Flo18, GM10]. fraction [dSCdA19]. fractional [JA15a]. fractionation [Vav12]. fragile [BCV13, GPJL15]. fragility [Kei12]. Fragment [VAO^{+}17]. Fragmentation [GTM^{+}19, PGA^{+}11, CPR10, CLDRJ18, LLH^{+}13, RMvW13, YGL18, YMCV21]. fragmented [CHNdÁP18, KK11, KCAP17, PGATS13, RCSS13, RKBG14, TdS12, VSPHS15, XSL18]. fragments [MP17, PGA^{+}11]. frame [WJS^{+}18]. frame-based [WJS^{+}18]. framed [ZPZ^{+}15]. Framework [BMU10, ACKS13, ACE^{+}17, Ano12t, BEPC15, CF17, CB15, CAY^{+}10, CSR15, CLNN14, CN18, DWNN10, FP17, GLC^{+}18, GPM16b, GDW12, HFU11, HOKH18, IBH^{+}15, JSH15, JMB^{+}11, LZL^{+}18, LL18a, LYCZ13, LYC^{+}15a, LB18a, MRD^{+}12, MK19, MJLC14, NMT^{+}19, OJ15, PWBG16, PRP17, PDD^{+}18, RJ17, SWZ^{+}12, SML12, SLRC19, TMPF18, TMT^{+}18, UB12, WHH^{+}18, dOAC18]. France [BMD18, BGPJ10, DJMBD^{+}13, GSPSP15, PLR^{+}13, SB15]. Francisco [DWP16, MLS^{+}11]. Fraxinus [KKR19]. free [CCSF11, JLN10, MGW^{+}13, MK19, MHY^{+}15, OSS13, RWvDM15, YRGH16, ZWJ^{+}12]. free-living [CCSF11, RWvDM15]. free-roaming [YRGH16]. free-surface [ZWJ^{+}12]. free-water [MGW^{+}13]. freedom [RRH^{+}19]. freeze [OQH^{+}13]. French [DBS18]. frequency [AB10, BBRJ15, CLNN14, PSB^{+}17, ZPZ^{+}15]. fresh [CN10a, CTNH13, ZW10]. freshwater [ABD14, CRB16, CLW^{+}19, CSR15, DWNN10, DWP13, FP17, HCE^{+}19, JTB^{+}19, JCKG15, KFWM19, KM12b, LKAY15, LZYH19, MWK19, McC16, MMMCAHM21, NiSK21, PCRP12, PNT^{+}19, SAAD16, SQWS19, TL15, VEK12, VJR^{+}16, WKZ^{+}19, XWW^{+}11, ZJL^{+}14, ZZ17]. Friction [LJ14]. frog [ACCM16, DA17]. frogs [NiSK21, PMF^{+}13]. frondosa [DZS12]. Front [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano18x]. fronted [LSJG17]. frontier [FAT^{+}15b]. fronts [PMTKC17, Rin18]. frost [RJH^{+}10]. frosted [WSBP13]. frugal [AADA14]. Frugivory [MLMH18, BNGP19]. fruit [ABR12a, BHD17, BGGL16, KL10, MCBM12, PDG^{+}14]. fry [RLHD11]. Ft. [DÖWB12]. fuel [BRR10, Cam15]. fuels [DSBS15, PMM11]. Fukuyoa [KDT^{+}17]. fulfillment [RHR^{+}15c]. full [CXW19, DMB^{+}19]. fully [WAGM17]. fully-stochasticized [WAGM17]. Function [CLBSG11, BPS^{+}11, CBR^{+}16, Fis21, KB18, LCM11, Lem16, MSG^{+}13, SBVA^{+}13, SGG14, SH18, SGKR10, YLC13]. Functional [MY11, TM12, ADJ^{+}17, BBL^{+}17, BSO13, CBB19, CN10b, EFM^{+}10, GANB11, GWF^{+}18, HBR19, HGAT17, HBA16b, KCLL14, KEM^{+}16, KB18, KLGL19, LDO^{+}18, LHZ10, LBK^{+}13, LDB14, MV11, MYHN17, NOOM^{+}16, Rai13, RHD^{+}13, SJGB16, SA16, SGB10, VCD^{+}17, VPC^{+}14, WZWL10, WUB15, WR18, WJvdL^{+}13, YJW15, ZPZ^{+}15, dVLS11]. functional-structural [BBL^{+}17]. functionality [AYASR^{+}17, CRO13]. functionally [RMLTCQ^{+}19]. functioning [ALC19, Bre12, KHL^{+}16, KL12, OGG^{+}16, OCCF17, PTE^{+}12, SMG^{+}12, TGB^{+}17]. functions [AAJB21, BJPB11, CLLB12, LVB^{+}15, LKP13, SBVA^{+}13, TF13, VSF10, VFFP16, Wil12]. Fundamental [MLC^{+}16, JSCS19, WAS^{+}16]. Fundamentals [Jør14a, Jør14b]. fungal [BCD12, CLLB12, COBF13, MBLC14, STSA10]. Further [Ben12]. fusing [CCLX17, CGG^{+}18]. fusion [RHZ^{+}18]. Future [RCS^{+}19, SP11, BKP10, Bar11, BUB^{+}12, BWP16, BGPJ10, BFKS16, CCF13b, CHK^{+}12, DAF^{+}11, EEFE^{+}11, FMBD15, FRM^{+}19, FCLFB18, GAP^{+}19, HGE^{+}10, HPC13b, JVY19, JFS^{+}12, KBS^{+}15, LRM^{+}14, LTL^{+}19, MBD^{+}12, PAS^{+}10, PKD11, PGS16, Pra12, RKW12, SLSS15, SBC13, SWR^{+}16, VLF^{+}17, WZG15, ZJP^{+}11]. futures [MTJF11b, SBF15]. Fuyang [YLC13]. Fuzzy [AWF10, CMH14, CPAG15, GCV14, HA18, dAEPE16, CC10, CCSF11, FDM^{+}11, GEPCB13, HLW15, HL19, JKS^{+}10, LGDW12, MSHC13, MMMGLN^{+}18, PAEE19, Pra12, RFV^{+}12, SGRAdJ17, vVHZHvD13, CBGK15].

G [Jør14b, Pat10]. GA [DÖWB12]. Gabes [HBS^{+}16a]. Gadus [IKL13, JNHF18, PWK^{+}17, SLRC19]. gain [Sim11, WHNR19]. gaining [DPLC13]. gains [FMLM11]. Galemys [BRV11]. Galing [LHL^{+}17]. gall [GPTR13, PFP^{+}16]. gallery [RV15]. Gambierdiscus [KDT^{+}17]. Game [MRM16, Che10, CB15, GKL18, HLL^{+}19, LR16, MN12, RE10, SM11, ZMMHR10]. Game-theoretic [MRM16]. games [SGSR10]. Gammarus [KMS^{+}12, KMS^{+}13]. Gangetic [SVL12]. Gap [HSVL12, AS12, BTPG10, BS16b, DLW^{+}17, FBdP^{+}16, FSS^{+}17, HB12, KWH14, LSJ^{+}18, LSJ^{+}19, LAD11, LMT13, MWML13a, RFZB12, SvBH^{+}14, SCG^{+}13]. gap-induced [AS12]. Garden [YZLX14]. Garonne [GSPSP15]. Gas [MKTO19, BPU11, CC11, IKH^{+}11, TB19, WC13a, WC13b, YC13, YFL^{+}18]. gasoline [BB10]. gastro [RWvDM15]. gastro-intestinal [RWvDM15]. gastropod [MG18, MHG18]. Gauging [AMH16]. Gaussian [MDH15, VVH12]. GCM [PAS^{+}10]. gDefrag [MAB19]. GDV [YL18]. gear [vPTOW18]. Geese [LSJG17, CAMNN16, SSD^{+}17]. GEM [SELB10]. Gemini [SMG^{+}12]. Gen [ARV^{+}16]. gene [CJC11, CDT10, CDT10]. general [BLD^{+}10a, BRS^{+}18, JGF14, KWGW19, LBGM19, Mob18, Pat16a, SR14, SAAD16, SMS^{+}13]. Generalised [LJWN14]. generalist [VWTA16]. generalized [For18a, MROGVR12, MK19, PRHH13, PHP^{+}14, SBA^{+}19]. Generalizing [AGQP^{+}14]. generally [KB10]. generate [dAdSC15]. generated [LJ15a]. Generating [SG11b, BPU11, HKVD16]. Generation [Dil10, VC17a, GB16b, GB16d, IBH^{+}15, JDT^{+}16, SLW17, SJB^{+}16]. generators [LTL^{+}19]. Generic [CLLB12, SELB10, BCD12, Con14, JSH15, LAB^{+}10, MBA^{+}19, SH14]. genetic [BAAB18, BCL^{+}14, CN11, DR17, FDM^{+}11, KvOP10, LWY^{+}11, PP12, SZSV13, SNZ^{+}19, YZPL16]. genetically [CDL^{+}11]. genetics [CS17, MPP^{+}15]. genome [LMSK12]. genomes [GHD^{+}19]. genons [AG14]. genospecies [WGSR19]. genospecies-host [WGSR19]. genotype [PDG^{+}14]. genotype-dependent [PDG^{+}14]. genotypic [KCLL14]. genus [KTV^{+}15]. Genypterus [CRQW18]. geo [ZCL^{+}16]. geo-process [ZCL^{+}16]. geobiosphere [BU10, BPU11, BU16a, BCUF16, BCDU16, DBSA16]. geoduck [HlLB12]. geographic [BRV11, CAY^{+}10, DNFG18, LdMS^{+}16, LGDW12, SCG^{+}13, ZCL^{+}16]. Geographical [AAO10, Kos17, WFA^{+}13]. geography [WLC^{+}19, NGR^{+}11]. geolocate [WFN^{+}16]. geolocation [NMA^{+}19]. Geometric [WHC^{+}13, AS11, YTT^{+}11]. geometrical [PFSF14, RSE10]. geometry [FCA18]. geomorphic [LDC^{+}16]. geomorphological [Bev15]. GeoPetri [SH18]. geophysical [TQZ12]. geosmin [HD14, HD16]. Geospatial [AD16, LZ10]. geostatistical [BL15, HEJP17]. germination [BCL^{+}14, GDC11, MOMC17, WXB10]. Getting [Pat14a, Pat16b]. Ghana [KCF17, KCF18]. Giant [GFW^{+}17, YL17, ATCJ^{+}19, CLW^{+}19, FLX^{+}14, LLL^{+}10, LCS11]. giganteus [MWD^{+}15]. gigas [GBB^{+}11, ZBMBDU^{+}17]. gilthead [PAGCV^{+}10]. GIS [ARTBB14, CVPR^{+}16, COK^{+}11, FTK^{+}14, FRV^{+}16, LdMS^{+}16, LLH^{+}17, LGDW12, MRZU14, PD12]. GIS-based [PD12]. GIS-maps [ARTBB14]. GIS-tool [FRV^{+}16]. glacier [LJCG15, MTPC13]. glacier-fed [MTPC13]. glaucogularis [HBT^{+}18, MFB17]. Glaucomys [TDF^{+}19]. gleaned [LDCB21]. Global [CVG^{+}15, DAA^{+}12, HBL18, LKMCW14, LWKS10, MSCH14a, MSCH14b, RDD^{+}14, WMF^{+}13, YZG16, ZSF^{+}12, Abe10, ARTBB14, BLD^{+}10a, BlMG^{+}11, BH13a, BSC^{+}15b, BPU11, BU11b, BU16a, BU16b, CLL14, CY13, CJR^{+}16, CAMCK^{+}12, CN13, DMR13, DSM^{+}12, DPLC13, DBS18, GCDA16, GB16a, GHP12, GPP^{+}19, GLS^{+}17, HN14, HHK19, HBA16a, HEJP17, JDT^{+}16, JMSM12, KVS17, LDP14, LCPB^{+}10, LSM19, MDOJ14, MLG^{+}18b, NP18, OQH^{+}13, PD15, PMCKH13, PBFS14, RvS16, RKS^{+}10, RA13, SVD^{+}13, SLLW15, VPS10, WTC^{+}18, WLS^{+}10, WLS^{+}11, XHC^{+}15, XBvdV^{+}14, XGH^{+}17, YqWB^{+}15, YJL11, YFC^{+}11, ZJL^{+}14]. globalizing [KR10b]. globally [RKL^{+}15, ZWG^{+}17]. GLOF [MTPC13]. GLOWORM [RWvDM15]. GLOWORM-FL [RWvDM15]. GLUE [OKS^{+}11]. GNESFA [CGV^{+}15]. goal [Bin17]. goals [BA17]. gobio [KDMPC10]. Going [TWS^{+}11, RA13, ZCL^{+}16, dSJHG10]. golden [BCP15, LML17, dÁSMRT^{+}19]. Gompertz [Ned15]. gone [AG14, TWS^{+}11]. Gonghu [GSQ17]. good [SKM^{+}16]. goose [JJMW14, WK17]. goral [SK11]. Gorges [CMLY19, MLW^{+}15, SRXP14]. gorse [RPGA15]. Gossypium [AS11, ZML^{+}12]. Gotelli [FM11]. governance [MSM^{+}16]. governed [CTP^{+}17]. government [PKT15]. Gower [MdlP15]. GPS [ACC11, PR13, WBG^{+}14, ZSS17]. GR3D [RDD^{+}14]. gradient [ENC^{+}15, FGPR10, GA15, GLRT16, LLQ^{+}15, LJCG15, MML^{+}11, MKF^{+}11, OBLJ13, TLL^{+}15, WMK^{+}13, ZLW^{+}15]. gradients [CBB19, DFD11, MSG^{+}13, MBW14, MJLC14, PICR13, SFB^{+}15, XWC^{+}11]. Graham [Hat14]. grain [DFF17, GAHE11, LVB^{+}15, SBC16b, WTL17]. Gran [CMCC15]. Grand [Bec10]. Grande [RR10, WLH^{+}10]. grani [CEC10]. grape [GPM16b, PDG^{+}14]. grapevine [JASW18]. graph [AMM^{+}14, CC12, DMDJ13, KP13, MAB19, RBRK16, TLW^{+}14, VBS14]. graph-based [AMM^{+}14, MAB19]. graph-structured [CC12]. graphics [TB11, TB12a, vdKGV11]. graphs [KPS17]. Grass [BGP11, AWWD16, EEB^{+}17, JBB^{+}11, JBB^{+}12, KAJØ19, PLL^{+}19, SMKM19, SYY^{+}19, YLL^{+}17]. grass-clover [KAJØ19]. grasses [DAW^{+}10, TBA19]. grasshopper [ZJL^{+}16]. grasshoppers [BG11]. grassland [BSS^{+}14, EST^{+}19, FDHL17, FBDZ12, GvdWBS11, GCW^{+}16, HYG14, HW12, HVCP16, KPMJ14, LLM^{+}13, LWL^{+}17b, RSP^{+}17, RKC^{+}15, SDT10, SHBG10, SAW^{+}11, SLH^{+}11, SRV^{+}12, SMG^{+}12, TFH12, WJ15, XLL15, YSA^{+}14]. grasslands [BRPK14, CRAP13, LKH^{+}18, MPG18, MBA^{+}19, TDC^{+}11]. Grassy [DPBK10]. gravity [PMLL11]. gray [SGDP19, YLC13]. GrazeBGC [RKC^{+}15]. grazer [KMRW16, PRHH13, PBG^{+}17]. Grazing [CRAP13, HLZ^{+}18, PLL^{+}19, ABR12b, ABR12c, DWP16, EMH10, FBDZ12, GBR^{+}13a, GLRT16, GWF^{+}18, PAT17, RVNB11, RP11, SDT10, WJ15, ZAW^{+}18, vLTdBP16]. Great [KBS^{+}15, LRM^{+}14, LRZ12, SSF14, MAM^{+}12, CDK^{+}17, DS12, GPR^{+}13, HBR19, KPT11, NJCF11, SSM^{+}13, TLW^{+}14]. Greater [LSJG17, GO15a, KDT^{+}17]. greatest [HKM^{+}18]. Greece [BC11]. greedy [SZSV13]. Green [ACC^{+}10, MSS^{+}19, HMP^{+}19b, HJ10a, KSS11, KKR19, LCP14, LCP15, MGR^{+}18, MRP^{+}11]. greenery [MPM^{+}15]. Greenhouse [CC11, YC13, LHB^{+}18, TB19, WC13a, WC13b]. greenhouses [LCAJGFNS14]. gregarine [LJB12]. Grey [JWLS17, THOA10]. Grid [BHD17, Slo11, VLF^{+}17]. grid-based [Slo11]. Grids [Gia14]. grip [BPM^{+}17]. griseifrons [TDF^{+}19]. gross [BSC^{+}15a, GGZ^{+}16, LLY^{+}13, LZZP16, RHvdTS14, SMC11, YqWB^{+}15, YLY^{+}18, YCL^{+}14, ZZFH15, ZZB^{+}18]. ground [HSCS14, MTA19, MCBM12, MVC^{+}17, MVC11, PKG16, PLB16, SHM^{+}14, SBC^{+}11, VHM^{+}17, VALC12, XGH^{+}17, YKL^{+}12]. grounded [BGG16a]. groundfish [MTGD17]. grounds [ZSG^{+}18]. Groundwater [SPR^{+}18, BKB10, BvBB10, LZS^{+}17, XLS11, YL18, ZSKI17]. groundwater-dependent [YL18]. groundwater-mediated [ZSKI17]. Group [TAT^{+}16b, Bin17, BEBA19, HTVE18, MYHN17]. Group-based [TAT^{+}16b]. grouper [CNI11]. grouping [DFD11]. groups [ABGB15, BSO13, DBH^{+}17, FAC^{+}19a, GCIB10, HBR19, KJVL^{+}17, NMS10, RV13]. growing [HNWT10, NP15, YSA^{+}14]. grown [GLS^{+}17, KYvdPS14]. Growth [BBA12, FKDG14, GCvDM12, AMF12, ANMJ12, AWL^{+}13, ACE^{+}17, AK17, ACM15, ARM^{+}12, AMR16, AB13, AS12, ASV19, BDL^{+}14, BAB^{+}16, BC12, BBL11a, BMCG12, BFH14, BOF^{+}16, CCS13, CCCP17, CADB19, CRO13, CEWW18, CBCB17, CSOS17, CL10, CS18, CLW^{+}19, CN13, CPS^{+}14, CRQW18, DZW^{+}11, DHC16, DLCR13, Dil10, DS21, Dol10, DSS11, ESA12, EC10, FKH^{+}15, FAT^{+}15b, GPK^{+}13, GBP^{+}16, GPR^{+}13, GJK^{+}15, GS19c, GCM^{+}13, HJ17, HTB13, HTB17, HlLB12, HW12, HBT^{+}18, HVCP16, Hos11, JR16, JG10, JASW18, KTK^{+}15, KDT^{+}17, KDMPC10, KPT11, LBZA19, LLB^{+}19, LBE^{+}17, LSA^{+}19, LKH^{+}18, LWKS10, LTL^{+}13, LZZ^{+}14, LHP16, LJWN14, Log19b, yLWhMhG11, MDW^{+}12, MJL^{+}17, MTA19, MKTO19, MRB14, MIC14, MSF18, MPGCS14, NYIH17, NBK^{+}11, NJM11, NB11, NBD^{+}13, OE12, OMH12, PAB13a, PRL^{+}13, PM16, PGGMMA19]. growth [PFR15, RFZB12, RLS11, RBH13, Sak13, San19, SBDS13, SMT19a, SMT19b, SLZ^{+}18, SMS^{+}13, SIS^{+}14, SCHGM16, SFR^{+}17, SML12, SMH16, SBPZ12, SV13, TCSG15, TF13, TKH17, TJYZ11, TT17a, VMG^{+}12, VFFP16, WPZ^{+}11, WCL^{+}17, ZRC^{+}15, ZZZW15, ZWJ^{+}17, dAdSC15, RCC^{+}11]. growth/decline [MKTO19]. Gruza [RRR^{+}10]. GSM [HD14, HD16]. GSS [WHC^{+}13]. Guacu [CSFB17]. Guadalquivir [RML^{+}13]. guanaco [ZSSR18]. Guangdong [LHL18]. Guanting [HFB^{+}11]. guaranteed [AdMC^{+}13]. Guatemala [CPR10, LCDJ^{+}12]. Guiana [DBS18]. guianensis [KCKdOW12]. guide [FCLFB18, TL15, WK17, ZKR^{+}13]. guided [HS13a]. Guidelines [SED15, PGR^{+}16]. guiding [CF17]. guilds [HKL^{+}11, LBK^{+}13]. Guiyang [FDR^{+}17]. Gulf [DUASLB^{+}12, BDV^{+}19, BLLJ17, DBVS19, FHL^{+}13, HBS^{+}16a, HHC15, KWGG19, KK15b, LLC^{+}17, LMGW16, MAC14, NCMF10, PDAL14, SLW17, SLRC19, TCH^{+}13, dMSL^{+}16]. guttatus [CNI11]. GWLF [NKSA13].

H [KMM^{+}13, Wei19, LKLK15, PJ16]. H5N1 [TZD^{+}15]. H7 [SFvB10]. H7N9 [ZWA19]. Habitat [FHG12, HPJ15, HBS16b, MCL^{+}17, PMS^{+}17, AMM^{+}14, AWSS14, BKP10, Bar11, BJJ12, BBGJ13, BS10a, BTPG10, BOW^{+}12, CRBC11, CEWW18, CH12, COK^{+}11, CLDRJ18, Cra10, DFL^{+}16, DWD19, DPvBvA11, DFK^{+}11, DKJH13, DP19, DHS^{+}16, DF13, FLX^{+}14, FSV^{+}12, FYR^{+}18, FDM^{+}11, FPG11, FPWG14, FWRZ10, GCDA16, GM10, Gli14, GAHE11, HHUD14, HPS10, Hat14, HSTS19, HHL17, JJKJ11, JFS^{+}12, JHHS17, JGB15, KK13, KC18, KBL10, KNM^{+}10, KASS12, KPTC14, KMP14, KK12a, LN11, LM13, LMSK12, LBE^{+}13, LJ10, LCCL15, LLH^{+}13, LPSC16, LML17, LLL17, LZW^{+}17, LLL19, Loe12, LGDW12, MTZ^{+}10, MAM^{+}12, MAC^{+}12, MRD^{+}12, MHMRAHHS19, MP13, MSV^{+}11, MGB13, MRR^{+}14, MSHC13, MMRS^{+}19, MDV^{+}12, MMMGLN^{+}18, NWAR12, NiSK21, PCRP12, Pec12, PGR^{+}16, PBG^{+}17, PL10b, PPBD13, RKBG14, RH15, SZSV13, SZV14, SC21, SCHQA16, SGA^{+}16, SMH^{+}12]. habitat [SLY^{+}17, SK11, SFHK12, SGDP19, TRP12, VRA^{+}17, VWTA16, WL13, WTG^{+}16, WBG^{+}14, WhFxGZ16, YKL^{+}12, YWY10, YGL18, ZCY^{+}19, ZPNM13, dSJHG10, vSGR16]. habitat-based [MRR^{+}14]. habitat-forming [JGB15]. habitation [NWAR11]. habitats [BPC^{+}17, BvHR^{+}15, BAT^{+}15, CTC^{+}18, FAC^{+}19b, GAG^{+}12, HAOS17, KK11, KL12, SBA^{+}19, SRP18, TRD^{+}15, WNF^{+}15, WLR^{+}11, YDT^{+}13, ZJL^{+}16]. Hablehrood [KSSA13]. Haematopus [SGA^{+}16]. Haemorrhagic [KGM12]. Haihe [YLC13, MLH^{+}17, WFSH15]. Hainan [YDZ^{+}21]. half [MH11].

Haliotis [LJWN14]. Hall [Wei19]. hampei [RCGC17]. Han [KLS17]. Harbor [SHS^{+}12]. harbour [NNST^{+}14]. hard [HC11, Jør14a]. hard-water [HC11]. hardbound [Jör11a]. hardiness [RJH^{+}10]. hardwood [KGYG15, LCML11, MLC^{+}16, MWML13a, MWML13b]. hare [LV17, TNTH11]. Harizaj [BU11a]. harmful [DA19, ENM11, GDDFT16]. Harvard [Gra14]. harvest [BNL^{+}13, dSCdA19, GDBB19, GS19a, HMS^{+}10, JSW^{+}13, LHLW15, MRP14, MRD^{+}12, Och10, PP16, PME^{+}16, RAB^{+}10, SHB^{+}11a, WBW14, YWA^{+}17, ZKSF17]. harvested [FYWA12, IKL14, ZJL^{+}14]. harvesters [DS12]. Harvesting [FvBM^{+}11, BHC^{+}15, LBZA19, MV17, TKP14, UKI^{+}11, WCF^{+}16, WMF^{+}13]. harvests [KCKdOW12, TVF^{+}10]. hatching [AOD^{+}13]. hatchlings [HGWL11]. having [GCK^{+}15]. Hawai’i [Ano12a, KAC^{+}12a, KAC^{+}12b]. Hawaiian [RKW12]. Hayq [FSML11]. hazard [CPD^{+}11, KDK^{+}10]. hazelnut [BOF^{+}16]. HEA [CRO13]. headwater [LW14]. health [CGGL16, CZCL18, HKL^{+}11, MLY^{+}19, SM11, XYCZ11, XWW^{+}11, XYCZ13, ZGJ10, ZLY^{+}13]. healthy [IKL14]. Heat [MSS^{+}19, DBP16, SBDA16]. heat-producing [SBDA16]. heating [Gra14]. heavy [BCC^{+}11, BCC^{+}14, GVC10, HZT^{+}13, LGMLD19, MAP^{+}13, MB17, RM13, TLL^{+}17]. heavy-tailed [LGMLD19]. Hebei [HLW15, HZZ^{+}18, ZLZ17b]. HEC [WZGH18]. HEC-RAS [WZGH18]. hedges [SMKM19]. height [ARM^{+}12, RFZB12, WK11]. Heihe [FCH^{+}17, JJYM19, WZG15]. held [LMY^{+}11]. Helgoland [SW19]. Helicoverpa [BNC^{+}18, FGH^{+}10, PPZ^{+}17, PVB16]. Heliothis [BNC^{+}18]. help [MAB19, PJ17, RRH^{+}19, VPM^{+}18, VEMS12]. helvetica [ZPNM13]. Hemiptera [FTK^{+}14]. Heptapteridae [dABLBAdOR17]. herbaceous [HBH^{+}12, vLTdBP16]. herbage [GBR^{+}13a]. herbicide [SPWR17, WTL17]. herbicide-tolerant [WTL17]. herbicides [RSP^{+}17]. herbivore [AKLB19, EMH10, FAG10, FBKS17, FS18, KSS11, LMDG14, MNHW^{+}13, TKP14]. Herbivores [RP11, AB10, SRP18, TM12, WHNR19]. herbivory [DDM^{+}17, KPMJ14, RKC^{+}15, ZCY^{+}19]. herons [BFJZ^{+}11]. herring [Hei11, SBCB13]. heterogeneities [YHZ19]. Heterogeneity [SDKE18, BL15, CdOR10, CIHL14, DAP18, EKB18, EAB11, GdTCFG10, GCW^{+}16, GLRT16, GWF^{+}18, KMKL19, LDCB21, LCMG18, MAM^{+}12, MGB13, NMA^{+}19, OMDP11, PMM11, PJ15, SX11, SPWR17, Sto10, WTG^{+}16, WSSM15, WYS12, ZLB14, dLB19]. heterogeneous [BRMW18, BPM^{+}15, DML15, FMA^{+}12, For14, FS18, GCV14, HB13, HQF^{+}17, KKP^{+}11, KASS12, KP10, MBSF14, MAP^{+}13, MJGC13, OBSL15, PvVC15, RKRB16, RV13, RH15, VBPY11, VP12, WGE^{+}19]. heterogenising [GVV17]. heterotroph [FD19]. heterotroph-phototroph [FD19]. heterotrophic [GBR13b]. heuristic [BPB^{+}18, GEB10, TŠAD12]. Hevea [KMM^{+}13]. hibernating [BM10]. hibernation [LARZN13]. hickory [LL18a]. Hidden [VRM^{+}10, LLC11, LCMG18]. Hierarchical [CCCP17, FDP^{+}10, JN15a, LYJ^{+}15, SRM11, TMC^{+}16, VSPHS15, WLS^{+}10, WLS^{+}11, AHS^{+}17, ABL^{+}13, AB13, AHS^{+}16, BDL^{+}13, BMN11, BAJ^{+}11, CPS^{+}19, HlLB12, KASS12, LLB^{+}19, LBE^{+}13, LJ15b, LL18a, MAFG^{+}19, PWGS^{+}17, RE10, SML12, SFL^{+}19, TMT^{+}18, ZWG^{+}13, ZLG16]. hierarchically [MCPB15]. hierarchies [Abe15, ESLJ18]. hierarchy [Abe10, AG14, JZCL14, PK12, YLC13, ZLZS14]. High [IHSM12, CLNN14, Jör11a, ARTBB14, AB10, BVJE18, DMB^{+}19, DBVS19, FMS13, GCK^{+}15, HMS^{+}10, HBSF16, KHS^{+}19, KEM^{+}16, LM12, LB17, LZB17, MBSW12, MSP^{+}13, MOM^{+}11, NL12, PSB^{+}17, RML^{+}13, WFN^{+}16, ZCY^{+}19]. high-altitude [MSP^{+}13]. high-boreal [HMS^{+}10]. high-density [LB17]. High-frequency [CLNN14, AB10]. high-latitude [BVJE18]. high-level [LM12]. high-order [ZCY^{+}19]. High-resolution [IHSM12, ARTBB14, BVJE18, HBSF16, KHS^{+}19, MBSW12, WFN^{+}16]. high-rise [LB17]. high-severity [FMS13]. higher [YZ18]. highland [FSML11]. highlands [TSSL11]. Highlighting [DLZ^{+}17]. highlights [BND19]. highly [BCV13, HSLC^{+}19, POS^{+}19, RCS^{+}19, SMH^{+}12]. hillslope [JM15]. hilly [TKN14]. Hilsa [BAA^{+}14, HSCS14]. Himalayan [MU17, RXSK14]. Himalayas [SK11, RBRK16, SGP18, URR^{+}17]. hind [CNI11]. Hinoki [Oga11]. Hippocamelus [LAEAG12]. hirsutum [ZML^{+}12]. hispanica [AR11]. historic [CDR^{+}13, RWVK11, WSBP13]. Historical [CL15b, BKMK10, BRHD15, Cam16, CAC^{+}16, GMP13, GRB12, LdMS^{+}16, LW17]. histories [WFN^{+}16]. history [AWWD16, BLD10b, BBL11b, DCCS17, HMP^{+}19b, qHLlZ^{+}16, JSH15, JB11, Joh10, MJL^{+}17, SW15]. HMM [WFN^{+}16]. HMM-based [WFN^{+}16]. Hoare [HLLP^{+}10]. hoc [AFC12, Pér18]. Hohokam [Mur12]. Hokkaido [HMA10b, OJM^{+}12]. holes [BM11]. holistic [Pat14a]. Holoecology [Pat16a]. Holons [AG14]. home [BM11, BRL12, DHL^{+}12, GM10, WSC^{+}19]. homeostasis [WSE12, WLR18]. homing [BAAB18, PP19]. hominids [GLS10b, GLS10b]. homogeneous [ASS14, BAJ^{+}11, DHL^{+}12, DP19]. Homogenised [DLCR13]. homogenising [GVV17]. homogenization [BAAB18, BS10b, VPC11]. honey [AOW12, AOW13, KDHC^{+}18, RBH13]. honeybee [HFMM^{+}12]. honeybees [BHHM10, MPK18]. honeycreeper [RKW12]. Hooghly [MRG13b, RBM^{+}17]. horizon [Rem11, BPC^{+}17, CBP^{+}17]. horizontal [DKM^{+}14]. horizontally [For14]. horn [GKL18]. horned [HSTS19]. horns [LR16]. Host [CWJ^{+}13, ASS14, BWH13, BLAV18, CH12, CHZ18, HM18a, HM19, HMA10a, KSS11, KSI14, MLK17, NPJ^{+}17, OLC^{+}17, PHvB^{+}18, RMP^{+}18, WGT12, WGTH16, WGSR19]. host-parasitoid [CH12]. host-plant [KSS11]. host-tick [HM18a, HM19]. hosts [BPK^{+}16, CHZ18, LLFL12]. hot [BSCG10, ZPZ^{+}15]. hotspot [WLS^{+}10, WLS^{+}11]. hotspots [DOB^{+}15, RWS^{+}15]. household [LLW^{+}15, LDYL18]. households [FG11]. Howard [GDC^{+}19]. howler [WFDDR10]. Hoze [SC21]. Huaihe [YXD^{+}18]. Huangqihai [CL15b]. huberi [SGM^{+}14]. Hübner [BFA11, FGH^{+}10, PVB16]. Hudson [HPC13a, HPC13b]. huemul [LAEAG12]. Human [Abe10, An12, ALC12, ADPS10, BGA17, BFJZ^{+}11, CMM^{+}10, CLWH15, CXL^{+}19, CLT^{+}11, DSBM^{+}18, Dol10, DYU^{+}12, GHS^{+}16a, GHST19, JSW^{+}13, KK13, LDM15a, LBWW14, LKG^{+}17, LHH12, LP17, LCDJ^{+}12, LB18b, MNEG11, MHS^{+}17, MSH^{+}19, MF16, OBC^{+}21, Rau11, RNE^{+}12, RCBPMN10, SHCR10, SRM11, SGWV15, WALCY12, WSZT10, YY11, vSGR16]. human-assisted [CMM^{+}10]. human-caused [RCBPMN10, SGWV15, WSZT10]. human-dominated [LBWW14]. human-mediated [SRM11]. humans [MHS^{+}17, MAM^{+}10, YY13]. Humboldt [NMC^{+}14, NMS^{+}14]. Humid [AD13, AWWD16, FBA^{+}11, LLTQ14, LLTQ15, MLH^{+}17]. humidity [KBS^{+}15, WSZT10]. humpback [PWGS^{+}17]. Hunan [ZXD^{+}13]. Hunchun [LKG^{+}17]. Hungarian [HRSKW15]. hunting [HPC13a, HPC13b, RCS^{+}19, WFDDR10]. hunts [WR13b]. Huron [LJR14]. hurricanes [XC14]. hybrid [CRO13, CRB16, IFC14, JBE^{+}15, LOLA13, OFK^{+}12, RAR15, SBR^{+}19, SSKJ18, SGHW^{+}17, WN17, XJY^{+}16]. hybridization [DPPI14, GDT^{+}14, LV17, NMT^{+}19]. hydra [dSCdA18, Kal19]. Hydraulic [FWRZ10, AHNK13, HGBA^{+}13, LPD14, TGB^{+}17, ZC11]. hydro [Bev15, Jam16, WFSH15]. hydro-eco-geomorphological [Bev15]. hydro-economic [WFSH15]. hydro-meteorological [Jam16]. Hydrobal [BC13]. hydrocarbons [XWW^{+}11, XQZ^{+}13, ZCSL18]. Hydrodynamic [YSH^{+}10, AHM^{+}16a, FRM^{+}19, GCM^{+}13, HGLZ13, KLS17, KvOP10, LPV11, yLWhMhG11, MSGG14, Md18, QDH^{+}10, WQD^{+}13, WFW^{+}17, WV12, Wil11b, YLC^{+}18, ZWRK13, ZSJ15]. hydrodynamic-biogeochemical [MSGG14]. hydrodynamic-marsh [AHM^{+}16a]. hydrodynamic-phytoplankton [HGLZ13]. hydrodynamics [MUP^{+}13, WMW^{+}12, YKCN10]. hydroelectric [dSGGBW12, WLH^{+}10]. hydroepidemiological [MC11]. hydrograph [LCCL15]. hydrologic [AK17, Ano12s, JJYM19, MCW13, Nak17, NP18, SFMA15, SEHD18]. Hydrological [CHVH13, HGAT17, BC13, CHR^{+}13, CZ11, DFW^{+}18, GJÁ^{+}10, GSPSP15, GDBB19, KHJ^{+}19, LKH^{+}18, LB18a, MLMFN^{+}15, MSO10, WKZ^{+}19, XTPT11, XYY^{+}16, YLF^{+}15, YXD^{+}18]. hydrology [CTNH13, LZZ^{+}14, LVPF15, WRSL11]. hydrology-crop [LZZ^{+}14]. hydropower [APWT15, ACK^{+}14, CCL^{+}13, DFW^{+}18]. hydrothermal [HSvO^{+}18, WXB10]. Hyla [ACC^{+}10]. Hyperbolic [FS14b, Dol10]. hypereutrophic [ALC19]. Hyperparameter [SMI^{+}19]. hyperspectral [CCLX17, LSS^{+}15]. hypolimnetic [SHM^{+}17]. Hypothenemus [RCGC17]. Hypotheses [BFM11, GAV^{+}11, HM10, LJ15a, MJB^{+}11b, MCMT^{+}16, Pat16a]. hypothesis [Gav15, VHM^{+}17, Pat16b]. hypoxia [BLLJ17, BH13a, DJS10, ER10, dMSL^{+}16]. hypoxic [WB16]. hypugaea [EGMW12]. hysteresis [BM16, LBN^{+}18]. hystrix [KvOP10, MATS14].

Iberia [RML^{+}13]. Iberian [BRV11, AR11, ABC^{+}17, BGA17, GVGZ^{+}21]. IBIS [CZC^{+}15]. ICBM [FBA^{+}11]. ice [AHS^{+}16, CR12a, TVT12, TSB^{+}18]. ice-shelf [CR12a]. ideal [MK19, MHY^{+}15]. ideas [JD18]. identifiability [BLLJ17, MDW^{+}12, WC13b]. Identification [CHTC11, QlSeZ^{+}10, IPSH14, PMBM11, QWJL19]. identifies [DDBK15]. identify [CCC12, GWR11, KGM12, LLL^{+}10, MU10, RKL^{+}15, THC18]. Identifying [BFPR16, BRL12, CDT10, DFK^{+}11, GH13, HHC15, JHHS17, LSH15, LJLO11, NKSA13, NJCF11, VRM^{+}10, XYTS19, ZWA19, JFS^{+}12, JBCÑ11, MRTW18, SWZ^{+}12]. identity [SS10]. if [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q]. Igapó [RCCN11]. ignition [RCBPMN10]. IGP [SVL12]. ii [Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano16v, Ano16w, Ano16x, Ano16y, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16-37, Ano16-38, Ano16-39, Ano16-40, Ano16-41, Ano16-42, Ano16-43, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, BvBB10, CKS^{+}17, DDM^{+}17, DBVS19, FPWG14, HPC13b, KWS^{+}13, Log19a, Pat16b, RCG^{+}13, SLLW15, XTPT11, dBGS^{+}14]. iii [Ano16v, Ano16w, Ano16x, Ano16y, Ano16z, Ano16-27, Ano16-28, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, CSS^{+}17, RCGC17]. IKONOS [ZPNM13]. ilisha [BAA^{+}14, HSCS14]. illegal [BWRA16]. Illinois [CDR^{+}13, SCH^{+}17, ZSG^{+}18]. illumination [JGX13]. illustrated [JKS^{+}10]. illustrates [DDB^{+}12]. Illustration [LJLC15, MTA19]. image [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, SHBG10]. imagery [Hat14, PK14, VALC12, XLR11, ZPNM13]. images [CCLX17, LAB^{+}10, TQZ12]. imaging [LSS^{+}15]. imbalance [Kei12]. imbalanced [SVH^{+}19]. immature [DdSCG11]. Impact [ENM11, FYWA12, HS13b, HV12, HPG^{+}10, LG10, LG12, MRG13a, MF16, MK14, MAMNL^{+}15, MDV^{+}12, NP18, OQH^{+}13, PSWM^{+}10, SAKB^{+}10, VSC18, WP14, XYCZ13, AMRH19, AHF^{+}13, BSFR17, BS13, BJK^{+}16, BPS^{+}11, BCTL16, BH17, BKM^{+}18, dÁGP14, BEBA19, BRHD15, CPR10, CSH11, CRTS15, CSFR15, DH12, DYU^{+}12, EBL^{+}11, FLX^{+}14, FTK^{+}14, Fre16, dSGGBW12, HHL17, HCM^{+}18, IKL14, LDS14, LKH^{+}18, Lin15b, LSP^{+}10, MDB19, MDKK15, NZL13, NGL13, OHS^{+}17, PC16, PH18, PFA^{+}16, PGKG12, PB14, PPDF14, ROG^{+}16, RPC^{+}17, SGBD10, SZVV10, SBRD10, TW18, TNA^{+}17, TPT^{+}11, WHRR15, Wol11, YLC^{+}18, dJGM^{+}18]. Impacts [ACK^{+}14, Che15b, DMBS11, FCG^{+}19, HBG^{+}16, KG13, LVLS14, LC11, Nei11, PVM12, PVM13, PK16, PD15, RJH^{+}10, SDY^{+}15, SPG11, SRT16, AW16, ACE^{+}17, ATCJ^{+}19, AGB^{+}13, ADL^{+}10, BTB17, BSH17, BHPF14, CN10a, CAC^{+}16, CLD^{+}17, CSR15, CKH15, DLD^{+}11, DVD^{+}14, DBP16, ER13, FMS13, FP13, HCZ^{+}17, HSR^{+}15, HKM^{+}18, HC11, HHF^{+}10, IFC14, JDT^{+}17, JBW^{+}17, KDHC^{+}18, KVP16, LHL^{+}17, LJ10, LBK^{+}13, LYCU14, LCT^{+}17, yLWhMhG11, MTZ^{+}10, MBM^{+}18, NL12, OBC^{+}21, PANL11, RLHD11, THC18, TLLX19, TCW11, TBJ^{+}15, TCH^{+}13, TDF^{+}19, VWTA16, WCF^{+}16, WMF^{+}13, WCL^{+}17, WO14, WNT^{+}13, WFDDR10, XC14, ZLLC13, ZLW^{+}15]. imperfect [BMP^{+}19, GMK^{+}19, KKP^{+}11]. imperiled [LBM16]. Implementation [BDV^{+}19, BAT^{+}15, DBVS19, FLLB11, GBP^{+}10, OGOWW16, PGKG12, ZBMBDU^{+}17, AG11, ALeSD17, DWD19, KB11, SR14]. implemented [Che10, RPGSCF16]. Implementing [ESP^{+}17, KLGL19, LJBS12]. implicates [BSCG10]. implication [ACCM16, Ju19]. Implications [AR11, ACC11, Bin17, DJS10, FAS^{+}19, HMPC11, HKlL14, Joh10, KCP^{+}15, LV17, LDM15a, MGS16, Md18, Mur13, Pec12, SHB11b, WRSL11, WLS^{+}10, WLS^{+}11, WEJ11, ZLB14, BAA^{+}14, BLD10b, BBA12, DLZ^{+}17, mGZK^{+}10, HSM^{+}12, KL12, MAC^{+}12, NCEGB16, PPZ^{+}17, PME^{+}16, F.11]. implicit [AS12, Cho13]. Importance [ABB^{+}14, ER10, FAS^{+}19, HPRNP12, MRWN19, ZW11, AWWD16, CW17, CdMMF16, CD15, DWP13, DG16, GCDA16, GO15a, JB11, KGS11, LPD14, LDCB21, MHL^{+}19, MMM^{+}16, NCMF10, TCSG15, VLF^{+}17, XCG^{+}16]. important [Agu11, ACK^{+}14, HMD^{+}16, QlSeZ^{+}10, TF13, WJvdL^{+}13]. imported [KAJØ19]. Imposed [Cra10]. impounded [GCL^{+}19]. impounding [CMLY19]. impoundment [MGMR15, PGKG12]. imprecise [dSCdA19]. improve [BOGA14, CEWW18, FHAG12, GGSP14, GMP13, MRP14, MPP^{+}15, SVH^{+}19, SJB^{+}16, TRP12, TPG14, WRB^{+}12, XBvdV^{+}14, YHJ^{+}17]. Improved [KCC16, SSV^{+}12, YqWB^{+}15, ZCYL15, dSD18, BU11a, DFF17, FE14a, GAB12, GBW^{+}15, GYL^{+}13, Har11, HS13a, HG17, IBH^{+}15, JGX13, LKDD15, LCCL15, LS18, SWZ^{+}12, SV13, SLZ^{+}17, Til15, YLC13]. Improvement [KPM^{+}14, ZLJM17b, ZLJM17a, ZZB^{+}18]. Improvements [HHW11]. improves [FMW^{+}19, TFMP11, ZBSF17, ZJB19]. Improving [BT16, CZC^{+}15, CMEP15, JSdSdL^{+}18, KZR^{+}13, LLCR10, LYC^{+}15b, PPBD13, RAM10, SLMS14, SCG^{+}16, TZS^{+}14, WB16, WZL^{+}16, YLL^{+}12, CHTC11, RJ17, ZCS13]. in-lake [PSB^{+}17]. in-situ [LLKY18]. in-stream [BA14a, HKVR13]. Inachis [HLLO13a, HLLO13b, PAB^{+}13b]. inapplicable [SMS^{+}13]. incidence [BRR10, JVY19]. incidental [CCS13, MROGVR12]. includes [NiSK21]. Including [DCMK11, SKTG11, GLRT16, LSA^{+}19, MCH15, MFD^{+}15, MDL^{+}10, RLS11, SB15]. inclusion [dSdCF10]. income [MSCM18]. income-breeding [MSCM18]. incompatibilities [BFPR16]. incomplete [For18b, PJ17]. inconstant [BNGP19]. incorporated [Che16]. Incorporating [Ano12t, AS12, BTS19, BG11, CH14, DÖWB12, FMW^{+}19, Fuj16, GB16a, GGSP14, GLC^{+}15, LYCZ13, LBS^{+}14, RJH^{+}10, RWVK11, SWS16, SMH^{+}12, SMLH^{+}11, WGA^{+}13, BRS^{+}18, CSOS17, CB19, SFM^{+}18, TM18, WSSM15]. Incorporation [KHW^{+}16, PLR^{+}13, PPBD13]. Increase [dAdSC15, ACE^{+}13, Ano17a, CMGF14, ESP^{+}17, Lee14, MPT^{+}11, NYH15, PRH17, VEK12]. Increased [HLLO13a, HLLO13b, CEWW18, CN13]. Increases [Mat13, PLS^{+}16, AG11, DBS18, PTA^{+}12, RGH^{+}13, VCD^{+}17]. Increasing [BCV13, KB18, Pra12, Slo11, CJC11, GLR^{+}19, KKP^{+}17]. increment [HlLB12, NBD^{+}13]. incubation [MRP^{+}11, SDD^{+}13]. independence [GMS11]. independent [LG10, SK10]. Index [Har11, ZTL^{+}17, ZPNM13, AKvO^{+}10, ALKB10, BU11a, CZC^{+}15, CSS^{+}19, DGKSM13, GAB12, LCCL15, NiSK21, RHB14, VPB^{+}12, YKvWK16]. indexes [TAR13, YZC14]. India [ADL^{+}10, BG15, CBGK15, DCMS18, MRG13a, MGMR15, MSKS13, MRG13b, RBM^{+}17, San17, SK11]. Indian [DFM12a, DFM12b]. indicator [BAJ^{+}11, CBA10, GLA11, KFvSH15, KGD14, LBK^{+}13, MC11, SJS^{+}13, SLH^{+}11, TA11]. Indicators [Wal15, Ana17, BGMFES15, BCP16, BFJL15, FLLG13, GLC^{+}15, KB10, KK15a, KCC16, LSJ^{+}18, LSJ^{+}19, MCPB15, Rau11, SG11b, VBS14, VFB19, WWYO18, vGKB10]. indices [FL11, FAPM12, HlLB12, MV11, MSFR15, SXZ^{+}15]. indicum [YKL^{+}12]. Indirect [DWWS15, WPB14, BFS11, BRB14, FG11, GV15, Jag14, LCR12, MK13, MDH15, MJC^{+}11, PTL14, RSP^{+}17, SB11, TAR13, XM13, ZCSL18, ZSKI17]. indiscriminate [HCE^{+}19]. INDISIM [PGG14]. INDISIM-Saccha [PGG14]. Individual [BMCG12, BQF^{+}18, CRBC11, CDL^{+}11, DJMBD^{+}13, GGVB11, HR11, KSI14, LAEAG12, MG16, MDB19, MSH^{+}19, OMDP11, Uch16, AHS^{+}17, Ash10, ADF13, BVP^{+}13, BAAB18, Bin17, BS10a, BRS^{+}18, CC12, CdMMF16, CWL^{+}16, CDK^{+}17, CJC11, CTC^{+}18, CRQW18, CDG15, DJT^{+}19, DVT^{+}10, EKB18, EST^{+}19, FB13, GHWG^{+}14, GDTP13, GHS^{+}16b, GCvDM12, GCM^{+}13, HMP19a, HCW^{+}17, HFSL14, HVL^{+}19, HS13a, HTL^{+}14, HSHF^{+}13, HHC15, HGAT17, HBB15, IFC14, JD18, KWJH16, KKI^{+}14, LJLC15, LBR16, LPF11, LSA^{+}19, LSGT13, LBS^{+}14, LAK^{+}18, MHL^{+}19, MCRB17, MAP^{+}13, MPF^{+}14, MBM^{+}18, MdMS^{+}19, MGSB11, MATS14, NMT^{+}19, NMdR16, PC17, PLB16, PRL^{+}13, PP12, PGG14, RR10, SJGB16, SRSS12, SSK14, SML12, Sim11, Sma13, SGSR10, SGWV15, TB13, TMPF18, TSM^{+}15, TTCC^{+}19, VSF10, VR12, VGR^{+}11, Wal17, WSL17, WR13a, WR17, WP14]. individual [WFN^{+}16, XZC^{+}17, ZDL^{+}17, dJHMS19, dOMZ13, vdVBJS15, vdVJS16]. individual- [GHWG^{+}14]. Individual-based [BQF^{+}18, CRBC11, CDL^{+}11, DJMBD^{+}13, GGVB11, HR11, LAEAG12, MG16, MSH^{+}19, BVP^{+}13, BAAB18, BRS^{+}18, CC12, CDG15, DJT^{+}19, DVT^{+}10, EKB18, EST^{+}19, FB13, GDTP13, GHS^{+}16b, GCvDM12, GCM^{+}13, HMP19a, HCW^{+}17, HS13a, HTL^{+}14, HSHF^{+}13, HHC15, HGAT17, HBB15, IFC14, JD18, KWJH16, KKI^{+}14, LJLC15, LSA^{+}19, LSGT13, LBS^{+}14, LAK^{+}18, MHL^{+}19, MAP^{+}13, MPF^{+}14, MBM^{+}18, MGSB11, MATS14, NMT^{+}19, NMdR16, PC17, PLB16, PRL^{+}13, PP12, PGG14, RR10, SJGB16, SRSS12, SSK14, SGSR10, SGWV15, TB13, TMPF18, TSM^{+}15, TTCC^{+}19, Uch16, VSF10, VR12, VGR^{+}11, Wal17, WSL17, WR13a, WR17, WP14, XZC^{+}17, ZDL^{+}17, dJHMS19, dOMZ13, vdVBJS15, vdVJS16]. individual-level [HFSL14]. individuals [CGS^{+}17, GCW^{+}16]. Indo [SVL12]. Indo-Gangetic [SVL12]. Indomitable [Fat16b]. Indonesia [SHB11b]. induced [ACM15, AS12, BGD^{+}16, BPB14, LBW18, MBM^{+}18, MMMGLN^{+}18, NP15, SS16, SS11, VAO^{+}17, WNT^{+}13, XC14]. induction [MBBA17]. industrial [DLZ^{+}17, GWHL15, KS18, LHH12, XZCL16, ZZF15]. industrialized [BS10c, GPM^{+}16a, SMH^{+}12]. industries [AAW^{+}15]. industry [KGM12, mNkGyT^{+}17]. IndVal [DGKSM13]. inequality [RW17]. inexact [LHZN11]. inexact-stochastic [LHZN11]. infected [ASS14, BJK^{+}16, WGTH16]. infecting [CC10]. infection [BCD12, LLFL12]. infections [CHZ18, SM10]. infectious [BJK^{+}16, SDS^{+}21, WGTH16]. Infective [BPK^{+}16]. infectivity [CCSF11]. inference [BWA17, WC13a, WC13b, ZBSF17]. Inferences [DLL^{+}19]. inferential [ZDL^{+}17]. inferred [KWJH16, WBC^{+}12]. Inferring [BPB^{+}18, BGMJ^{+}12, FDP^{+}10, GMK^{+}19]. infestation [AD16, BGGL16, KSI14, NPJ^{+}17]. infestations [FH11, RPA^{+}18]. infinite [KB11]. infinite-dimensional [KB11]. inflated [CMT^{+}11, LBE^{+}13, LJ15c]. inflow [ABD14, BGD^{+}15, KM12b, ZW10]. Influence [ABAG15, ABAG16, AYASR^{+}17, ABD14, CQS15, GHP12, GAHE11, HMA10a, MAM^{+}12, PLS^{+}10, SMH16, ABR12b, ABR12c, Agu11, AHNK13, BA14a, BEZ19, BV15, BKB10, CAL13, CHR^{+}13, ČTT^{+}12, CIHL14, DF13, FMW^{+}19, HN14, KSM^{+}18, LMR16, LDC^{+}16, LGY10, OdlFDP15, PLS11, PSWW11, QDH^{+}10, RVNB11, SFMA15, SSQ^{+}10, Tie16, TFMP11, VF16, VPM^{+}18, WRMB18, vLTdBP16]. influenced [ATL18, BCO17, SMKM19]. Influences [KK12b, LCP15, XKI^{+}11, BRMW18, GTC16, LLZ^{+}16, MWK19, WRSL11, XCR^{+}13, YZ16, YXD^{+}18]. influencing [HLLP^{+}10, IT17, PVB16]. influential [CCC12]. influenza [ZWA19]. inform [MCMT^{+}16, PSM^{+}12]. Information [AAO10, HKP10, TLL^{+}17, VFB19, WYHC15, Abe13, IHSM12, BDTV11, CAY^{+}10, DPLC13, GDC^{+}19, GPP^{+}19, HYG14, HDL^{+}16, JJMW14, JJS10, JLN10, Jør15a, Jør15b, LdMS^{+}16, LZZ^{+}14, LGDW12, PLR^{+}13, PKMH16, PVB16, SCG^{+}13, WEB11, YHW^{+}12, ZCS13]. Information-based [TLL^{+}17]. informed [TCGC16]. Infrastructure [MSS^{+}19, MGR^{+}18]. infrastructures [LCP14, LCP15, MAB19]. ingenuous [BMF^{+}11a, BMF^{+}11b]. inhabiting [AR11, GFW^{+}17]. inherent [Cra10, MJC^{+}11]. inhibiting [FR11]. inhibition [MCH15, PJ16]. inhibitors [SBDS13, TB19]. inhomogeneous [CVS^{+}13, DHL^{+}12]. initial [Ant14, ABGB15, IKL14, JDZC14, KSM^{+}18, PSG19]. initialise [WHC^{+}13]. initializing [HWS11]. injuries [CBP^{+}17]. INLA [CF14]. Inner [HYL^{+}17, RNE^{+}12, YSA^{+}14, ZJL^{+}16, LWL^{+}17b, SAW^{+}11]. innovative [MKSL^{+}17, YHS^{+}11]. Input [SCH^{+}17, BB10, BCP16, BFS11, CC11, hCjHMW18, Cho13, Fis21, FYWA12, HCM^{+}15, LHL18, MU17, Pog11, SZSV13, XYTS19, ZQC15]. Input-Output [SCH^{+}17, XYTS19]. input-state-output [BCP16]. input-variables [Fis21]. inputs [BJK^{+}16, GHS^{+}16a, GJÁ^{+}10, HSM^{+}12, KØ13, LJR14, MCB14, YDD^{+}13]. insect [AD16, BLRN15, BNC^{+}18, CCL14, CAC^{+}16, DRS^{+}16, ETG^{+}18, FPO18, HQF^{+}17, LLH^{+}17, LJCG15, PD12, PvVC13, RSCS19, TMC^{+}16, THBF^{+}17, WBW14]. insect-resistant [FPO18]. insecticidal [YWA^{+}17]. insecticide [SDSH15]. insects [CMM^{+}10, EAB11, MNK^{+}21, RDJ19]. insight [GO15a, NMdR16]. Insights [ALeSD17, BGGL16, CHPS17, KH15, LDCB21, MHL^{+}19, MBD^{+}18, MJL^{+}13, MJL^{+}14, MDVG^{+}15, PES19, PLL^{+}19, TRD^{+}15, WHNR19, AMM^{+}14, BPWB14, FT10, GAB12, HMP19a, LLB^{+}19, MRM16, OCCF17, VASH^{+}16, BAT^{+}15]. inspired [CMSPJ11, ZVS18]. instability [FR11, dAdSC15]. installations [AMH16, KL12]. instance [Con14]. INSTAR [SMBGH^{+}19]. InSTREAM [ARV^{+}16]. InSTREAM-Gen [ARV^{+}16]. insular [BGMFES15, TVI^{+}13]. insurance [VMP^{+}19]. intake [GBR^{+}13a, RSE10, Sim11]. integer [BDWP16, EBA^{+}19]. integral [EEB^{+}17, EEK^{+}18, TWP^{+}14]. integrals [FAPM12]. integrate [VMP^{+}19]. Integrated [CLH17, CHCC15, Che15a, EIG^{+}10, HCM^{+}18, JC11a, LBA14, LBA15, MA10, TCW11, VPC^{+}14, WHB12, WGE^{+}19, ZLY^{+}13, vDSC^{+}10, AGJ^{+}12, DUASLB^{+}12, EBA^{+}19, GH13, GOSH^{+}12, HJ17, HOKH18, JMB^{+}11, JBW^{+}17, KPLS15, LZS^{+}17, LLP17, MKSL^{+}17, MRD^{+}12, MDVG^{+}15, QL15, RSDP^{+}12, RGRMRR18, SFvdM^{+}17, SHB11b, SDSH15, TPG14, UFK^{+}10, VKdW^{+}10, XSN^{+}18, ZLJM17b, ZLJM17a]. integrates [SMBGH^{+}19]. Integrating [BG19, BS10a, CS17, DK17, FLLB11, FtHvdB^{+}14, FS18, HGBA^{+}13, KJMA11, KSG^{+}14, LBM16, LLC^{+}17, LLH^{+}17, MRP14, NMdR16, PBD^{+}17, RRBI14, RKC^{+}15, RKBG14, SHX^{+}12, SGM^{+}14, VHDM17, YZC14, vGKB10, DWNN10, KK10, LCL^{+}13]. Integration [ASL13, CMM^{+}10, DC10, KK12a, Sak13, ZSJ15, BGMFES15, GCCJ12, GLI^{+}21, Ken17, KNPH^{+}17, MCBM12, XZZ^{+}13, YGL18]. Integrative [HKL^{+}11]. integrity [SLY^{+}17, SFB^{+}15]. intellect [Fat16b]. intelligent [PD12]. Intensification [Ano12a, KAC^{+}12a, KAC^{+}12b, LCPB^{+}10, ZBAY14]. intensity [DWWS15, DSM^{+}12, RHB14, SHB^{+}11a, SGG14]. Intensive [LLZ^{+}16, HCZ^{+}17, SVL12]. intensively [BRL12]. Inter [BMD18, MK12, GGSP14, LCT^{+}17, NLM^{+}19, RDR^{+}17, dVLS11]. Inter-annual [BMD18, MK12]. inter-regional [LCT^{+}17]. inter-sectoral [NLM^{+}19]. inter-specific [dVLS11]. inter-taxocene [RDR^{+}17]. inter-tree [GGSP14]. Interacting [VP12, dSJHG10, BG19, MB14, PKG16, SSBJ12]. Interaction [FPG11, HOKH18, MP13, ACCM16, AAME11, BWH13, BNGP19, CpW16, DBH^{+}17, DMZ^{+}19, F.11, KB18, LZS^{+}17, NG11, SCB13, YZ16, ZFG13, dOMZ13]. Interactions [GPA^{+}13, Hei11, TNTH11, ZBS14, AALP18, AD13, AAA13, AKLB19, ADPS10, Ban11, CBB19, CLT^{+}11, CCG15, DVF^{+}19, DKM^{+}14, DMBS11, ELPK12, FSML11, FGPR10, FSM^{+}19a, FD19, FT16, FAS^{+}19, FKSK^{+}10, FRP18, GVGZ^{+}21, GDTP13, GVC10, HLL^{+}19, HBL18, HSvO^{+}18, KSS11, KMF^{+}21, LSH15, LMDG14, LCM11, LP17, MKF^{+}11, MSH^{+}19, MvdWU^{+}16, MSV12, NGRDTHRL19, PHvB^{+}18, PLS^{+}10, PBG^{+}17, PGZ^{+}19, PNHL14, PS13, PDBP11, QFG^{+}13, RNE^{+}12, SLW^{+}14, SGM^{+}14, SPP18, TW18, Tie16, TAR13, Ver10, WGT12, WJL^{+}13, WTG^{+}16, WNF^{+}15, XM13, XSL18, YZ18, ZCY^{+}19]. Interactive [LLFL12, BC12, JWX18, KHHN17]. interacts [VF16]. Interannual [LY16, NZL13, NWAR11, WGA^{+}11, WGA^{+}12, XCR^{+}13]. Interception [DKM^{+}14, DHC14, For14, MJGC13, dLB19]. interchangeability [GBR13b]. intercomparison [HPW^{+}12, PTE^{+}12, WGA^{+}11, WGA^{+}12]. interconnected [LT15, MY11]. Interdisciplinary [CK12, Wei19, ZH13]. interest [HLS^{+}14]. interface [OdlFDP15]. interference [dSCdA18, DMH17, F.11, RSE10, RAdA12, dVLS11]. interim [HPW^{+}12]. interindividual [BDL^{+}13]. interior [FAS^{+}19]. interlaced [CL15b]. interlinked [WTWS19]. Intermediate [AHM^{+}16b, SRP18, LLFL12]. intermittent [SAR11]. Internal [HS13c, WWYO18, BFNH14, KZR^{+}13, LWGH13, LLZ^{+}15, Wan19]. International [FKDG14, LMY^{+}11, LYP^{+}11, LPA^{+}11, Wei19, Ano13u]. internesting [DBP16, HVL^{+}19]. Interplay [BSC^{+}15b, TMPF18, SK10, dOAC18]. interpolation [ChHC12]. interpretation [JTM^{+}17, OCD^{+}13, TWP^{+}14]. Interpreting [SX11]. interregional [KR10b]. Interspecific [YZ16, BLAV18, KSM^{+}18, KKBZ17, LS13, MPR13, VLF^{+}17]. Intertidal [ALeSD17, OCB^{+}13, WBHM11, XCL^{+}11]. Intertype [LCM11]. interval [HWE^{+}18]. interventions [DSA^{+}18]. intestinal [RWvDM15]. Intra [BLAV18, GLT10, DDR^{+}14, EFM^{+}10, RPGA15, SW19, XLL15, dVLS11]. intra- [dVLS11]. Intra-annual [GLT10, SW19, XLL15]. intra-colony [DDR^{+}14]. Intra-host [BLAV18]. intra-seasonal [EFM^{+}10]. intra-specific [RPGA15]. intraguild [DdSCG11, URJ10]. intransitive [BRMW18, WRMB18]. intransitivities [GV15]. intraspecific [qHLlZ^{+}16, MCA18, OFK^{+}12]. intrinsic [RDD^{+}18, SCBH18]. introduced [MSP^{+}13, MGCL17, PTL14, TW18]. introducing [LCCL15, LDCB21]. Introduction [BFW14, BH10, HBF18, VMG^{+}12, DPV^{+}16, GT14, KGM12, MJB^{+}11a]. introductions [MF16]. introductory [PF18]. introgressive [DPPI14]. inundation [DVF^{+}19]. invaded [BGC10]. invaders [BRMW18, YHW19]. invading [TCGC16]. invariance [LCAJGFNS14]. invariant [THJ^{+}12]. Invasion [VBPY11, Agu11, CVM^{+}18, CGM^{+}14, FTYI13, GDT^{+}14, GdTCFG10, KKR19, KVP16, LAR^{+}16, MJB^{+}11b, MG18, MHG18, MMVAHMC16, SR14, SGP18, SS11, VHDM17, XCG^{+}16, YLW14]. invasion-induced [SS11]. invasions [AWF10, GQRC^{+}17, JB11, LRZ12, MCT10, PMLL11, SRM11]. Invasive [WPG15, AALP18, BGA17, BJR16, BRKP16, CMM^{+}10, CPAG15, ETG^{+}18, EEK^{+}18, FP17, FHAG12, FKK^{+}18, GPJL15, GPTR13, HHF^{+}10, JBB^{+}11, JWLS17, KR14, MKH^{+}17, MF16, MPP^{+}15, MWD^{+}15, PK16, POS^{+}19, PKD11, PMS11, RvKM10, RBRK16, SZSV12b, SDS^{+}21, TBA19, YHK^{+}19, ZLY16, ZVS18]. invasives [BMP^{+}19]. inventories [MBRR13]. inventory [LLT^{+}14, OLG^{+}13, VALC12, WGA^{+}13, HSK17, SHM^{+}14]. inventory-based [WGA^{+}13]. Inverse [NLMdL10, BSZ^{+}11, DVT^{+}10, SBVA^{+}13, THJ^{+}12]. inversion [HSK17]. invertebrate [AHNK13, HFH^{+}19]. invertebrates [BH13b]. InVEST [JDT^{+}17]. investigate [BLH10, BCP16, CD16, IBD^{+}19, LdMS^{+}16, Str15a, Str15b]. Investigating [BSH17, DBVS19]. investigation [BG15, CT16, GS19a, HB15, PGKG12, TQZ12]. investing [MPGCS14]. Investment [GPM^{+}16a, FN18]. involving [BWH13, BRW^{+}12, SGSR10]. io [HLLO13a, HLLO13b, PAB^{+}13b]. ion [NP15, NPH^{+}17]. Ionian [PGV10]. Ips [OFH^{+}19]. Iran [AHF^{+}13, SZSV12a, SZSV12b, SZSV13, SZV14]. Irish [DOB^{+}15]. iron [RMTM^{+}16]. irregular [BM11, VAO^{+}17]. irrigated [MFESBG^{+}13]. irrigation [DAA^{+}12, FBL^{+}10, KS17a, KS17b, PRP17, XLS11]. ISBN [Kat11]. ISBN-13 [Kat11]. Ise [TSN15, TSN17]. ISEM [Ano13u, LYP^{+}11, LPA^{+}11, WTC^{+}18, Ano11q, Ano11r]. ISI [HhLY^{+}19]. Isla [FAC^{+}19a]. Island [JN15b, WCS^{+}14, BP10, CN11, DWP11, GKF^{+}14, KKP^{+}17, NCK10, PP19, VPC11, ATCJ^{+}19, BLCPB11, GPC14, MSS^{+}19, ZKSF17]. Islands [CNI11, MGGLBIBD13, MRG13a, WSW^{+}16, FBDR17]. isolated [dS18]. isolation [CHNdÁP18]. isometrically [Gav15]. isoprenoid [KGS11]. isotope [DLH^{+}14, DCZ^{+}15, MDL^{+}10, PLR^{+}13, Vav10, Vav12, VRC17, VRL18]. isotopes [SBDA16]. isotopic [VRL14]. Israel [ROC^{+}13]. Issue [Ano11-37, Ano14-40, FKDG14, BFW14, BH10, GT14, GB16b, GB16d, HBF18, KA17, LP17, VMG^{+}12, WTC^{+}18]. Issues [SSK14, Ano11-38, FCLFB18]. Italian [BSC^{+}15a, GZVU14, PPF11, XWC^{+}11]. Italy [FRM^{+}19, FBDR17, MJP^{+}12, CPS^{+}14, CRTS15, DAF^{+}11, IBD^{+}19, JSB^{+}14, MSS^{+}19, MRZU14, MCPB15, PGK^{+}13, TL10, TAC^{+}14, VB10, VZF^{+}11]. Italy/Southern [FRM^{+}19]. Iterated [CLBSG11]. Iterative [JMdW^{+}10]. iUniverse [Jør14a]. Ivlev [WZWL10]. Ivlev-type [WZWL10]. Ixodes [HM18a, HM19, HMA10a, LVLS14]. Ixodidae [WGT12].

jack [MK12]. jackknife [SA13]. jaguar [BEPC15, WNF^{+}15]. James [SQWS19]. Janthinobacterium [ACCM16]. January [Ano10i, Ano11s, Ano12w, Ano12-34, Ano13-29, Ano14k, Ano15d, Ano16o, Ano17z, Ano17-40, Ano18-42, Ano18-33, Ano19-42, Ano19-40, WLS^{+}11]. Janus [Pat16b]. Japan [DFK^{+}11, HMA10b, IKK^{+}12, Ju19, LDYL18, MSO10, OJM^{+}12]. Japanese [LSA^{+}19]. japonica [TZS^{+}14]. japonicus [DFK^{+}11, LCC^{+}14]. javanica [MAO19]. jellyfish [RPA12]. Ji [HZZ^{+}18, LCT^{+}17, WLC^{+}19, ZLZ17b]. Jialing [yLWhMhG11]. Jiangsu [NvGZ^{+}16]. Jiaozhou [XZC^{+}17]. Jilin [ZLZS14]. Jin [HZZ^{+}18, ZLZ17b, LCT^{+}17, WLC^{+}19]. Jing [HZZ^{+}18, LCT^{+}17, WLC^{+}19, ZLZ17b]. Jing-Jin-Ji [LCT^{+}17, WLC^{+}19].

John [MSB^{+}17]. joint [CRTS15, HHL17, LP17]. Jointly [GK14]. Jørgensen [Fat16b]. July [Ano10m, Ano10n, Ano11-28, Ano11-29, Ano12z, Ano13-35, Ano13-40, Ano14q, Ano14n, Ano15m, Ano15k, Ano16b, Ano17u, Ano17-28, Ano18-44, Ano18-35, Ano19-46, Ano19-35]. jumbo [ZBMBDU^{+}17]. jump [Pau14]. June [Ano10k, Ano10l, Ano11z, Ano11-27, Ano12v, Ano13y, Ano13-34, Ano14m, Ano14p, Ano15l, Ano16d, Ano16r, Ano17-30, Ano18-39, Ano18-37, Ano19-38, Ano19-34, Con12]. juniper [yHAB12, PFSF14]. Juquiá [CSFB17]. Juquiá-Guacu [CSFB17]. justification [BS16a]. juvenile [ACM15, DWNN10, Dil10, FPG11, GCvDM12, HHUD14].

Kachchh [KK15b]. Kalimantan [SHB11b]. Kalman [HGLZ13, HG17, SG11a, ZCS13]. Kappa [vVBHZ11]. Karenia [DA19, HHC15]. Karlodinium [GDDFT16]. Kazakhstan [LBR16, PVBC14]. kelp [ROSC15]. Kemp [KWGG19]. Kenya [JAKJ10, LDM15b, MNHW^{+}13]. Keoladeo [CBGK15]. Kerguelen [MCMT^{+}16]. kernel [BM11, KUK11]. kernels [SBM^{+}11, Slo11]. key [BND19, BSCG10, CDT10, GFW^{+}17, GB15, HKL^{+}11, HBS16b, LJLO11, MBHW12, MRTW18, RCSS13, SBM^{+}12, TPL^{+}15, TPO^{+}18]. Keys [YSB11]. keystone [BSFR17, EPM^{+}14, FAC^{+}19a]. keystoneness [OCB^{+}13]. kidney [Mur13]. killed [HWE^{+}18]. killer [SGSR10]. kin [GVGC16]. KINEROS2 [GPD^{+}13]. KINEROS2/STWIR [GPD^{+}13]. kinetic [PHP^{+}14, VRL14]. Kinetics [BCC^{+}14, GBB^{+}11, LNB17, Sak13]. king [PPDF14]. Kinneret [ROC^{+}13, OHS^{+}17]. Kitaura [IKK^{+}12]. Klausmeier [CV19]. knotweed [LSA^{+}19]. Knowing [BRR10]. knowledge [CPE13, FMBD15, FKH^{+}15, For18b, GCCJ12, JBV^{+}15, JKS^{+}10, PSM^{+}12, SA16, SMBGH^{+}19]. knowledge-and-data-driven [FKH^{+}15]. known [FCLFB18, RGHS11]. Koch [dOMZ13]. Kohala [Ano12a, KAC^{+}12a, KAC^{+}12b]. Koka [TW18]. Kolmogorovian [ACKS13]. Komarov [SGCB15]. Kootenay [MAM^{+}10]. Korea [BS18, KLS17, LP17, NCK10]. Korean [KL10, KPK^{+}17, KKP^{+}17, RHZ^{+}18]. koreana [KKP^{+}17]. kriging [LWY^{+}11]. krill [CWM12, GJK^{+}15, JR16, MRCW10, RMTM^{+}16, TWH^{+}14, WCG^{+}11]. Kronoberg [PMS^{+}17]. Kruger [BFKS16]. Kuantan [LHL^{+}17].

L [Jør14b, KMS^{+}13, PFSF14]. L-systems [PFSF14]. L. [BCL^{+}14, BOF^{+}16, BRKP16, CMT^{+}11, CPMP13, CSR15, HSHF^{+}13, KYvdPS14, KMS^{+}12, MBBA17, MPKC12, MdMS^{+}19, MMVAHMC16, MMMGLN^{+}18, ODK16, UDP17, VVH12, YWA^{+}17, ZML^{+}12]. labelling [TGM^{+}21]. lability [PDAL14]. laboratory [LHO^{+}16, MHRB19, MRP^{+}11]. Labrador [HMS^{+}10]. labrax [PAGCV^{+}10]. Lacs [KVP16]. lacustrine [WL13]. lag [NS15]. Lago [dSdCF10]. Lagocephalus [CVM^{+}18]. lagoon [BSZ^{+}11, BPGSS^{+}18, DVD^{+}14, GMLRF10, LAC10, LVV^{+}15, LRB12, MDL^{+}10, RPA12, SFMA15, VZF^{+}11, ZDG^{+}13]. lagoonal [KM12b]. lagoons [ABL^{+}13, PICR13, TBM^{+}14, ZDG^{+}13]. Lagopus [ZPNM13].

Lagrangian [QDH^{+}10, ZZT13]. Lagrangian-based [ZZT13]. lags [ROC^{+}13]. LAI [ZCS13]. Laizhou [LZW^{+}17]. Lake [CZR^{+}14, CL15b, DATI10, DZW^{+}11, GSQ17, HSR^{+}15, KVP16, MWY^{+}18, MSKS13, MBUM12, NXZ18, NGGFM19, RCCN11, SXZ^{+}15, TW18, XYCZ13, YDL^{+}15, ZLLZ13, ALC19, BCC^{+}11, BM16, BMD18, CdMMF16, CGGL16, CWLL15, CPS15, CLNN14, CTP^{+}17, DMZ^{+}19, FRM^{+}19, GYL^{+}13, HJWG12, HC11, HMA10b, Hu16, JHH^{+}12, Jør10b, KD13, KHL^{+}16, LkSXhX17, LWGH13, LCY^{+}13, LLZ^{+}15, LGY10, MSP^{+}13, MDF^{+}18, MGW^{+}13, MH10, NGWZ16, OGG^{+}16, PSB^{+}17, PMLL11, PBK14, SWZ^{+}12, SHM^{+}17, TCW11, TSB^{+}18, VJR^{+}16, WB17, WFW^{+}17, WKZ^{+}19, XC14, YYQ11, YZPL16, YZG16, ZLY^{+}13, ZWRK13, ZLG16, ZCSL18, ZWW^{+}18, BMD18, BGPJ10, DATI10, ESA12, FRM^{+}19, FSML11, HRSKW15, HFH^{+}19, HLLP^{+}10, Hu16, HG17, IKK^{+}12, Jam16, JXL^{+}15, JLZ^{+}18, Jør10a, KB10, KJVL^{+}17, KNM^{+}10, KJH^{+}13, LJR14, MJB^{+}11a, MYJ^{+}14, NOOM^{+}16, OHS^{+}17, PAB13a, QCB^{+}11, ROC^{+}13, SHB^{+}11a, SLW^{+}14, SGWV15, SS11, WV12]. Lake [WJF^{+}11, WX11, XYCZ11, XWW^{+}11, XQZ^{+}13, ZRC^{+}15]. lakes [AdMC^{+}13, BdCS11, CRB16, CT11, DS12, EBL^{+}11, FRP18, GCL^{+}19, JSL^{+}10, JTB^{+}19, JFS^{+}12, LHZ10, McC16, MUP^{+}13, MHC16, PMBB^{+}16, RWH^{+}14, VRC17, VRL18, XJY^{+}16, XWC^{+}11, XYY^{+}16, DS12, LRM^{+}14, LRZ12, SSF14]. lakeshore [WP14]. Lam [SZSV12a, SZSV12b, SZSV13, SZV14]. lamppost [BL15]. lamprey [NWAR11]. lampreys [NWAR12]. lanceolata [WZD^{+}14, ZXD^{+}13]. Land [FLY13, LWL^{+}17b, LBC^{+}15, SYL^{+}13, ZZLZ14, dFMdSA18, ADF^{+}17, BPR^{+}11, BLRN15, BKMK10, CLH17, CZ11, CB19, CKMD14, DK17, DSM^{+}12, DPLC13, DSA^{+}18, ER13, Fel12, FCFT16, FC11, GAP^{+}19, GMP13, GANB11, GBP^{+}10, GLI^{+}21, HHL17, HWL17, JJYM19, LBE^{+}18, LBK^{+}13, LYSW15, LLZ^{+}16, LZL^{+}18, LVZ^{+}11, LLZ12, LOLA13, LZZW17, LCDJ^{+}12, MNEG11, MDOJ14, RLHD11, SJGB16, SLH^{+}11, SWR^{+}16, SDY^{+}15, SCG^{+}16, Tul14, TAGC^{+}16, TA11, WHB12, WBC^{+}12, dWcWbZ15, WO14, WBHM11, WS15, WFM16, XLR11, YZL12, YZC14, YXD^{+}18, ZFT^{+}10, ZFZ^{+}14, ZZJ^{+}16, vDSC^{+}10, vVBHZ11, vVHZHvD13]. Land-cover [FLY13, ADF^{+}17, DK17, FCFT16]. land-use [ADF^{+}17, BKMK10, DK17, DSM^{+}12, DPLC13, ER13, FCFT16, JJYM19, LBK^{+}13, LZZW17, SDY^{+}15, WO14, WS15, vVBHZ11]. land-use/land-cover [ADF^{+}17]. landcover [AH17]. LANDIS [DDM^{+}17, KWS^{+}13, LHLW15, SHX^{+}12, SLLW15, dBGS^{+}14]. LANDIS-II [KWS^{+}13, SLLW15, dBGS^{+}14, DDM^{+}17]. landlocked [BCTL16]. lands [UR10]. Landsat [Hat14, MHMRAHHS19, TQZ12, VB10]. Landsat-derived [MHMRAHHS19]. Landscape [CPR10, CBO16, KFvSH15, PD12, PGR^{+}16, WZL^{+}16, ABD^{+}18, AMM^{+}14, BC11, BGG^{+}16b, BPM^{+}15, BPM^{+}17, CADB19, CAC^{+}16, CXW19, Con11, Con12, DVL11, DNFG18, FAMD^{+}12, FT10, FtHvdB^{+}14, FAPM12, FCA18, GCB^{+}11, GCK^{+}15, GB15, GWF^{+}18, HMPC11, HRSKW15, HV15, JD10, KWS^{+}13, KCF^{+}13, KMF^{+}21, KGS11, KCAP17, KGYG15, LTL^{+}19, LBH^{+}15, LSZC10, LCCL15, LLQ^{+}15, LHW^{+}15, LLH^{+}17, LDB14, LML17, LYL^{+}17, LHLW15, MNEG11, MSV12, MDH15, MRM^{+}17, MJLC14, MWD^{+}15, NGWZ16, PMS^{+}17, PTML14, PPZ^{+}17, PSM^{+}12, RHB14, RPA^{+}18, RCS^{+}19, SCHQA16, SKHH19, SvWP^{+}15, SRP18, SRSS12, SMH^{+}12, SSQ^{+}10, SDEP18, SGWV15, Sto10, TLLX19, TMPF18, TT12, TdS12, TWS^{+}11, TLS^{+}16, TDF^{+}19, VHDM17, WGA^{+}11, WGA^{+}12, WGT12, WYY^{+}19, WHNR19, YHS^{+}11, YZC14, YHW19, ZJG10, dBGS^{+}14]. Landscape-level [CPR10, PD12, LLH^{+}17]. Landscape-scale [PGR^{+}16, RPA^{+}18, SRSS12]. landscapes [BSP12, BKP10, Bar11, BGK^{+}16, BCI^{+}17, BRHD15, CHNdÁP18, DDM^{+}17, DPBK10, DS12, DML15, EZ14, EGMW12, GTM^{+}19, HRSKW15, LSGT13, MFESBG^{+}13, MAB19, NNP19, OBSL15, PHvB^{+}18, PvVC15, PSKO16, RLHD11, RCSS13, VSPHS15, VBPY11, Wal15, XSL18, ZWG^{+}13, ZNS^{+}18, dHR18]. LandSim [PMS^{+}17]. landslide [CLCC11]. landslides [DH12, VEMS12]. landuse [AH17]. landuse/landcover [AH17]. language [DR14, MCPB13, UDNSLCG16]. languages [PWK^{+}17]. Lanier [KB10]. larch [NS15, XKI^{+}11]. Large [MGSB11, AAO10, ACE^{+}17, AWSS14, AB10, BKN21, BB18, BV15, DBR12, DMB^{+}19, FMA^{+}12, FRM^{+}19, Flo18, FYWA12, FRP18, FBKS17, GLA11, GSPSP15, GPA^{+}13, GWC^{+}10, GLY^{+}15, KHL^{+}16, Kos17, LDC^{+}16, LPV11, LZYH19, MSF18, MNHW^{+}13, NCLP10, NGL13, PGKG12, RPA^{+}18, SvBH^{+}14, SRP18, SSKJ18, WFW^{+}17, WWC^{+}18, YHS^{+}11, YZG16, ZW12, ZWRK13, dMLM^{+}17]. large-area [MSF18]. large-scale [BKN21, FMA^{+}12, FYWA12, GLY^{+}15, SSKJ18, dMLM^{+}17]. largely [CVS^{+}13]. larger [LGMLD19, NYH15]. largest [GPR^{+}13, WKZ^{+}19]. Larix [KWJH16]. lark [HSTS19]. larva [CC10, SLB^{+}18]. larvae [HLLO13a, HLLO13b, LCC^{+}14, OAPP12, PAB^{+}13b, VVH12]. LARVAHS [BBGJ13]. larval [BBGJ13, BLAV18, GCvDM12, MLS^{+}11, MYJ^{+}14, NWAR11, NWAR12, SU15]. laser [HS11, KHS^{+}19, MdMS^{+}19]. last [LBC^{+}15]. Late [Flo14, PHvB^{+}18]. latent [CL10]. lateral [FCS13]. Lates [HD16]. Latimeria [CPE13]. latitude [BVJE18]. Latitudinal [TWH^{+}14, ENC^{+}15]. lattice [BM11, IWW12, IKH^{+}11]. lattice-based [BM11]. laurel [MAC^{+}12]. Laurentian [LRM^{+}14, LRZ12, SSF14]. lavaretus [VVH12]. law [GCW^{+}16, JDZC14, KOBK13, Oga17, Oga18, SC18, SSR16, dLB19]. Lawrence [CLD^{+}17, MDVG^{+}15]. Laws [Jør14a, HLS^{+}14]. layer [LKLK15]. layered [AMR16, JAH^{+}10]. layers [LBE^{+}18]. laying [GPZB18, SG16]. LCA [RRBI14]. LDI [RHB14]. leachate [RM13]. leaching [HOŠ^{+}12, YDD^{+}13]. lead [MKL14, NYH15]. leading [ASS14, JD18]. Leaf [BLB^{+}17, Bai18, BPS^{+}11, CPMP13, CZC^{+}15, FS14b, GZK^{+}11, Hos11, KSS11, KGS11, KMS^{+}12, KMS^{+}13, LBR^{+}17, LCFD19, Oga18, OJ15, RML^{+}15]. leaf-resolving [Bai18]. LEAFC3 [MEC14]. LEAFC3-N [MEC14]. leafy [LCFD19]. leaky [Mur13]. learn [Rin14]. learned [CLL^{+}15, FBdP^{+}16, GMP13]. Learning [KNM^{+}10, STD15, WHH^{+}18, CGP12, DDR^{+}14, FKK^{+}18, GMC^{+}19, MHL^{+}19, MRWN19, NGGFM19, PLB16, RRvdG18, SMI^{+}19, WKF^{+}17, ZHLZ17]. Least [IBD^{+}19, KK13, LLL^{+}10]. least-cost [LLL^{+}10]. leatherback [DBP16]. leaved [RHZ^{+}18]. leaves [AS11, DKM^{+}14, JX15]. Leeward [Ano12a, KAC^{+}12a, KAC^{+}12b]. Legal [CGY11, CGL15]. legend [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q]. leidyi [AALP18]. leishmaniasis [RMP^{+}18]. Lemna [SBDS13]. length [MCH15, Pea15, SMT19a, SMT19b, ZML^{+}12, vPTOW18]. length-based [SMT19a, SMT19b, vPTOW18]. lentic [HKL^{+}11]. León [MJB^{+}11a]. Lepidoptera [FPO18, BTS19, FGH^{+}10, KL10]. LepiX [FPO18]. Lepomis [GV11]. leprosy [OLC^{+}17]. Leslie [CCS13, MdlP15, Mon13]. less [BGD^{+}16, MS14, TYEH10]. Lesser [GCvDM12, GCM^{+}13]. Lessons [BRM^{+}14, CLL^{+}15, FBdP^{+}16, GHWG^{+}14, GMP13, SU15, ZCR16]. lethal [MS14]. leucas [MDVG^{+}15]. Leucaspius [BGC10]. Level [WYPF15, CHR^{+}13, CPR10, CN11, CAMCK^{+}12, CRTS15, CWMC^{+}12, DVD^{+}14, FH11, FtHvdB^{+}14, FD19, For14, FAS^{+}19, FPWG14, GHWG^{+}14, HFSL14, HW12, HLK^{+}12, JD10, KR10a, KHS^{+}19, KH15, LM12, LBW18, LLH^{+}17, LNB17, LKAL^{+}13, LVV^{+}15, LCDJ^{+}12, MAP^{+}13, dMHGA10, NPP^{+}18, OWMK10, PD12, PWK^{+}17, RSC12, RSC13, RMLTCQ^{+}19, SJS^{+}13, TTCC^{+}19, TB19, WCL^{+}17, WKZ^{+}19, WRM^{+}11, YHZ19, vDHH^{+}14]. level-size [TTCC^{+}19]. Levels [LPF11, BS13, DLH^{+}14, DBVS19, HB14, KHHN17, PLS^{+}10, Sol19, XWW^{+}11, XLL15]. Liaohe [HZSH17]. Liaoning [GZS^{+}18]. liberalised [HGE^{+}10]. lice [AHS^{+}17, Mur11, MS16]. LID [MJY17]. LID-BMPs [MJY17]. Lidar [BBL^{+}17, TRP12, Hat14]. LIDET [SBC^{+}11]. liebre [RIMPBT19]. Life [DLZ^{+}17, Jør14a, Kat11, KBMM15, UAB^{+}11, AWWD16, BB10, BH13b, BLD10b, BBL11b, CDT10, DLL^{+}19, DCCS17, DLD^{+}11, GRRU15, GCM^{+}13, HA18, HMP^{+}19b, qHLlZ^{+}16, HSHF^{+}13, HPGM18, HB10a, JSH15, JB11, Joh10, LJB12, MJL^{+}17, MBRR13, MCMT^{+}16, NBSGW19, PMH17, RSCS19, SW15, WGE^{+}19, Jør14a]. life-cycle [HSHF^{+}13]. life-history [BBL11b, MJL^{+}17, SW15]. Life-Organizing [Jør14a]. life-stages [GCM^{+}13]. life-traits [CDT10]. lifecycle [DMB^{+}19, RCSS13]. lifetime [DPvBvA11]. Light [BVC^{+}18, DHC14, BSC^{+}15a, BMR^{+}18, BS17, BKB10, BvBB10, DMZ^{+}19, FS14b, For14, GRD^{+}14, HNWT10, LCL^{+}13, LVV^{+}15, MKTO19, MJGC13, Pea15, TPO^{+}18, WBS13, YLC^{+}18, ZZFH15, dLB19]. light-driven [BMR^{+}18]. lights [WLC^{+}19]. lignocellulosic [ABAG15, ABAG16]. Liguria [PPB^{+}18, VB10]. Liguria-Italy [VB10]. like [MP13]. likelihood [KSZW16]. likely [BGD^{+}16, PC16, PTL14]. Likens [Jør14b]. Limfjorden [CVPR^{+}16]. Limitation [HKVD16, BKB10, BvBB10, Che15b, FC12, LWC^{+}18, Mul11, SFM^{+}18, WEJ11]. limitations [GCCJ12, LB18a, PMBM11, THC18, VF16]. limited [CN13, FMBD15, HPJ15, HJ10b, MRCW10, OCD^{+}13, PKMH16, RP11, YLL^{+}17]. limiting [AL16, TGB^{+}17]. limits [Ban11, FP17, PKD11]. Limnoperna [dÁSMRT^{+}19]. limpid [MSO10]. line [FAPM12, PR11]. linear [AK17, BDWP16, BPK^{+}12, CCCP17, DH18, EBA^{+}19, For18a, GdTCFG10, MAB19, SBVA^{+}13, SXNL18, VRL14, VC17b, WK11]. linearity [FS14b]. lines [SSQ^{+}10, YJW15]. lingering [RRBI14]. Linguistic [HL19]. Link [Pat15, RHR^{+}15a, SBR^{+}19]. Linkage [WQD^{+}13]. linkages [LYCZ11, ZH13]. linked [FTK^{+}14, HHC^{+}13, LST15, SHCR10, ZWRK13]. Linking [CB15, CN18, EFM^{+}10, HGSJ15, HOŠ^{+}12, LSGT13, LG17, MSB^{+}17, PMM11, PK14, RRG^{+}15, SBC^{+}15, ZDL^{+}17, DLCR13, FDP^{+}10, FE14a, JDT^{+}17]. Linyphiid [Joh10]. lions [RRG^{+}15, BDV^{+}19, DBVS19]. lipid [RM13]. Lippia [MAO19]. literature [MMMGLN^{+}18, PMB^{+}19, VG19]. literature-based [MMMGLN^{+}18]. Lithurgus [SGM^{+}14]. Litter [ZTJ^{+}10, DFRT14, KMS^{+}12, KMS^{+}13, MBG^{+}10, MRG13b, ZJM^{+}11]. litterbag [Lec17]. litterfall [MRG13a]. Little [JJKJ11, CESC19, RTE^{+}15, SSQ^{+}10]. lived [GHH^{+}15, SF10]. livelihoods [SHB11b]. lives [SKPHW16]. livestock [Bae18, CCF13a, FDHL17, FS18, LAK^{+}21, RVNB11, SDT10]. lividum [ACCM16]. Living [FJH15, CCSF11, IT17, RWvDM15]. lizard [ABC^{+}14, TVI^{+}13]. lizards [FSV^{+}12]. llama [TSSL11]. LLC [Jør14a]. LMC [PdM11]. load [NXZ18, OGOWW16]. loaded [ELPK12]. loading [JSL^{+}10, REB^{+}16, SQWS19]. loads [CTNH13, MMM^{+}16, MHTC15, RML^{+}13, SB15]. lobster [BPB14, LMBL12]. lobster-induced [BPB14]. Lobsters [EPM^{+}14]. Local [Che16, MWK19, RCCN11, SLD^{+}19, WYS12, AAME11, ARM^{+}12, BND19, BBL11a, BGG^{+}16b, CWM12, CRVS13, DJH10, DBH^{+}17, FCLFB18, GPTR13, GJK^{+}15, IWW12, KAJØ19, KHHN17, Log16, MMM^{+}16, Mat13, MLAT^{+}18, Pec12, PGV10, RPA^{+}18]. localized [KBMM15]. locally [VC17a]. located [LLTQ14, LLTQ15, ZPZ^{+}15]. location [ACC11, DMB^{+}18, RWVK11]. locations [CMV^{+}16, Fai10]. lock [SDSH15, ZVS18]. lock-and-dam [ZVS18]. lock-in [SDSH15]. locust [BRKP16, DBH^{+}17]. Loess [FLY13, ZZLZ14]. logbook [KSZW16]. loggerhead [MJL^{+}17, WHRR15]. logging [Gat18, TDH12]. Logic [BPT^{+}19]. Logistic [MIC14, AMF12, GGV11, MCA18, SJTS13, Sak13, SK11, Ske13]. logistics [LBM16]. Londrina [RCCN11]. Long [Ano12a, ABC^{+}17, BUB^{+}12, HOŠ^{+}12, LDO^{+}18, PGKG12, SVL12, WP14, YSA^{+}14, ZJM^{+}11, AB13, BYR18, BLB^{+}17, BSS^{+}14, EC10, FB15, GPJL15, GPTR13, GV11, GHH^{+}15, GCB^{+}11, KAC^{+}12a, KAC^{+}12b, LBZA19, LA11, LBdA^{+}19, MBG17, MP17, Och10, PWBG16, Pen11, PGZ^{+}19, RUSL19, RDD^{+}18, SF10, Str15a, Str15b, TGM^{+}21, VG19, VASL^{+}15]. long-distance [GPJL15, GPTR13, MBG17]. long-lived [GHH^{+}15, SF10]. long-run [EC10]. Long-Term [Ano12a, ABC^{+}17, BUB^{+}12, HOŠ^{+}12, LDO^{+}18, PGKG12, SVL12, WP14, YSA^{+}14, ZJM^{+}11, AB13, BYR18, BLB^{+}17, BSS^{+}14, FB15, GV11, GCB^{+}11, KAC^{+}12a, KAC^{+}12b, LBZA19, LBdA^{+}19, MP17, Och10, PWBG16, Pen11, PGZ^{+}19, RUSL19, TGM^{+}21, VG19, VASL^{+}15]. Longitudinal [WWC^{+}18]. Longleaf [LCML11]. longline [LJ13]. Looking [BUB^{+}12, MAFG^{+}19, YKL^{+}12]. loop [GLS10a, LDJZC15].

Lophotrochozoans [DLL^{+}19]. Lorenzo [GSJ^{+}10]. Losers [RE10, KTV^{+}15]. loss [Bus17, CR12a, CPR10, CHG^{+}15, CLDRJ18, DPLC13, HBS16b, LLH^{+}13, LLL17, LLL19, SAAD16, SBC^{+}11, VWTA16, YGL18, ZCY^{+}19]. losses [Ano12t, LDJZC15, LYCZ13]. Lotka [ELPK12, For18a, GvdWBS11, Sci11b, TSM^{+}15, WSSM15]. Louisiana [LdMS^{+}16, ER10, MMK15, PFA^{+}16]. Low [For11, DWP16, FBA^{+}11, KKH^{+}17, LM12, LW17, MM16, PLS^{+}10, Pog11, RWVK11, WJK11]. low-level [LM12]. low-pass [FBA^{+}11]. low-resolution [RWVK11]. low-salinity [DWP16]. lowbush [BHD17]. lower [KLS17, LVV^{+}15, WRSL11, GFP^{+}16, HKM^{+}18, WAGM17]. lowland [HG17, Lin15b, PGKG12, TVF^{+}10, WB16]. LOWQM [Jam16]. Loxodonta [MMS12]. LPJ [SM14]. LTER [SFB^{+}15]. Lubei [ZZF15]. LUCC [LCDJ^{+}12]. lucius [BVP^{+}13, MMVAHMC16]. Lujiang [YLW^{+}11]. lumpy [BRMW18, WRMB18]. lupus [CWL^{+}16]. Lygodium [HEJP17]. Lyme [WGSR19].

M [Jør14a]. maackianus [HH12]. MACAROMOD [RPC^{+}17]. Macaronesian [RPC^{+}17]. Macaw [HBT^{+}18, MFB17]. macchia [MVC^{+}17]. Machine [PLB16, CGP12, FKK^{+}18, GMC^{+}19, MHL^{+}19, MRWN19, NGGFM19, SMI^{+}19, WKF^{+}17, ZHLZ17]. machine-learning [MHL^{+}19, SMI^{+}19]. machines [MMVAHMC16, SZSV12b]. mackerel [BRS^{+}18, HCM^{+}18, LCC^{+}14]. Macquarie [ATCJ^{+}19]. macro [BH13b]. macro-invertebrates [BH13b]. macroalgae [BN10, CVPH^{+}14, JGB15, SØB14, YMCV21]. macrobenthic [CZCL18, KM12b, MGGLBIBD13]. Macrobrachium [CLW^{+}19]. macrochirus [GV11]. macroecological [DDR^{+}18]. macroinvertebrate [MR17, SLY^{+}17, TFMP11]. macroinvertebrate-based [SLY^{+}17]. macroinvertebrates [KSG^{+}14, LBE^{+}13, SMT19a, SMT19b]. macromolecular [ANMJ12]. macrophyte [GSQ17, JHH^{+}12, LCY^{+}13, NFU19, VAO^{+}17, ZLY^{+}13, ZJL^{+}14]. macrophyte-derived [ZJL^{+}14]. macrophyte-dominated [ZLY^{+}13]. macrophyte-rich [NFU19]. macrophytes [LHZ10]. macrophytic [GYL^{+}13]. macroscopic [DLCR13]. macrotidal [GJÁ^{+}10]. macrozooplankton [RZF^{+}13]. made [AW15, BdCS11]. Madrid [RCBPMN10]. Maestri [HBT^{+}18]. Maggiore [FRM^{+}19]. MAGI [ZSKI17]. main [GHST19]. Maine [MJLC14, SLRC19]. mainland [BP10]. mainstream [RDV^{+}17]. maintain [PBG^{+}17, SRP18]. maintained [LCML11]. maintains [YZ18]. maintenance [MR10]. maize [FPO18, FYWA12, HLLO13a, HLLO13b, LZZ^{+}14, SLZ^{+}18]. major [HSM^{+}12, HBA16b, MSO10, NP18, SBF15, SBDA16, WKZ^{+}19]. make [SBD^{+}19, SKM^{+}16]. Making [Gia14, BFNH14, BPT^{+}19, CGL15, EAJG16, GCCJ12, KRE^{+}18, KK15a, NCLP10, SHX^{+}12, SWGC16, WC13c]. Malacosoma [BFA11]. malaianus [HH12]. malaria [MAO19, RDR^{+}13].

Malawi [DATI10]. Malaysia [ABAA21, LHL^{+}17]. male [AMM^{+}14, SK10]. males [EEB^{+}17]. Maling [RBRK16, RBRK16, SGP18]. mammal [DMB^{+}19, GGTP^{+}17, WGL^{+}12]. mammals [BFP^{+}18, FAPM12, FCA18, MSCM18, RMR^{+}16]. man [BdCS11, dSdCF10]. man-made [BdCS11]. Manacapuru [dSdCF10]. manage [BMP^{+}19, KØ13, WB16]. managed [GCK^{+}15, HBH^{+}12, MWML13a, MPG18, WB17]. Management [DA17, ETG^{+}18, GHS^{+}16b, Mur12, TCGC16, ACCM16, AHM^{+}16b, ASP^{+}15, ABC^{+}17, AHF^{+}13, Bae18, BS13, BLD^{+}10a, BlMG^{+}11, BGG16a, BFJZ^{+}11, BEPC15, BT16, BKMK10, CRAP13, CF17, CN10a, CRA^{+}11, Che15a, hCjHMW18, CSH11, CLD^{+}17, CAMCK^{+}12, CMCASS12, DLL18, DS12, DML15, EMH10, FH11, FSM^{+}19a, FSM19b, FBKS17, FS18, GH13, GWC^{+}10, GBB13, GPM16b, GOSH^{+}12, GB15, GYL^{+}13, GS19c, HV12, HPO^{+}15, HBT^{+}18, HCL^{+}16, HT16, HGBA^{+}13, Hol13, HSM^{+}12, Hu16, IS13, ISG13, JGF14, JBB^{+}11, JWLS17, KDK^{+}10, KSSA13, KPLS15, KTM^{+}16, KLGL19, LV17, LBE^{+}17, LSA^{+}19, LVLS14, LLT^{+}17, LHB^{+}18, LKAY15, MFB17, MGR^{+}18, MRD^{+}12, MSMM11, MWML13b, MS16, NCEGB16, Nei11, NBD^{+}13, NNP19, OJM^{+}12, OWMK10, PAT17, PPZ^{+}17, PMB^{+}19, PKT15, PGR^{+}16, PW10b, Pra12, RFV^{+}12, Ret12, SS16, SBD^{+}19, SPR^{+}18, SHCR10]. management [SHX^{+}12, SKMB11, SBC^{+}15, SGRAdJ17, SRT16, SLRC19, SB15, SLD^{+}19, TSN17, TLP17, VVA^{+}15, VQS^{+}18, WKZ^{+}19, WRSL11, WEB11, Wil11a, Wil12, WK17, WZL^{+}16, WEJ11, XZCL16, YWA^{+}17, ZBAY14, ZSG12, ZHC^{+}16, ZKR^{+}13, ZWRK13, ZLW^{+}15, ZF11, ZBS14, ZHCX15, aGSCMF15, dMLM^{+}17, vDHH^{+}13, MLAT^{+}18, WPZ^{+}11]. manager [DML15]. managers [KLGL19]. Managing [ACDC17, BHD16, DPV^{+}16, ESP^{+}17, GHS^{+}16a, GCTS15, JJMW14, KK13, RP10]. manganese [WWYO17, WWYO18]. Mangrove [DBS18, CZCL18, GSS^{+}14, MRG13a, MRG13b, PLFO^{+}12, PVL^{+}14, XCL^{+}11, dHR18]. mangrove-based [XCL^{+}11]. Manhattan [LB17]. Manica [GU11]. maniculatus [BB18]. manipulating [TV15]. manipulation [LLFL12]. manner [KHHN17, PDG^{+}14]. Manorina [HS11]. mantis [ADL^{+}10]. mantle [SBDA16]. manual [MJK^{+}17]. manure [GPD^{+}13, MRP^{+}11, SFvB10]. manure-amended [SFvB10]. many [LM12, OOFVBC16, vdVJS16]. map [BFR17, DGKSM13, LLC11, PGKG12, RFV^{+}12, SSM^{+}13]. Mapping [ASP^{+}15, Hat14, HS11, MAO19, SGP18, CCLX17, COK^{+}11, Fai10, JBV^{+}15, KDK^{+}10, KS17a, KS17b, MGGLBIBD13, dMHGA10, MOM^{+}11, NPP^{+}18, VSB13, VBS14, WLR^{+}11]. MAppleT [DHC14]. Maps [BME14, ARTBB14, MMGK15, PLB16, RPPW16, SGRAdJ17]. March [Ano10y, Ano10z, Ano11t, Ano11u, Ano12x, Ano13w, Ano14b, Ano14t, Ano15f, Ano15e, Ano16f, Ano16p, Ano17-39, Ano17w, Ano18-27, Ano18-43, Ano19-41, Ano19-48, mLcLJ16]. Marco [UB12]. Marennes [PLR^{+}13]. margin [RW10]. marginal [MRES17]. marginality [GS19b]. margins [DPBK10, WFA^{+}13]. mariana [SGG14]. Marina [VB10, DLL^{+}19, Hos11, KR10a, KvTL^{+}18]. Marine [PCTW17, ANMJ12, AWF10, AMH16, ADJ^{+}17, AGB^{+}13, ADL^{+}10, BS13, BTB17, BJJ12, BPC^{+}17, BV15, BAT^{+}15, BPD^{+}19, CHCC15, CS17, CDL^{+}11, CBP^{+}17, CH14, CSS^{+}16, CPS^{+}19, CGEdM12, CMV^{+}16, CKH15, dSCdA19, CD15, DMB^{+}18, FLH17, Fie12, FHL^{+}13, FBH^{+}14, FBDR17, GMTC^{+}17, GMvV17, GPP^{+}19, HGWL11, HCZ^{+}17, HVL^{+}19, HN14, HPC13a, HPC13b, IT17, JSH15, JOHW17, KBB^{+}11, LSR16, LMBL12, MFS18, MSCH14a, MSCH14b, NCEGB16, NGL13, PLR^{+}13, PPB^{+}18, PWP^{+}19, PBD^{+}17, PPO12, ROSC15, RKS^{+}10, RMR^{+}16, RA13, SVD^{+}13, SW19, SG11a, TLP17, TM18, UDRA13, VPB^{+}17, VRL14, VASH^{+}16, ZSF^{+}12, MGGLBIBD13]. marinus [GMvV17]. maritime [JAH^{+}10]. maritimus [AML13]. mark [ABB^{+}14, BS10c, GMS11, HVL^{+}19, KNPH^{+}17, LCM11, PS13, RGHS11, RDJ19, TB12b, vPTOW18]. mark-recapture [BS10c, TB12b, vPTOW18]. mark-release-recapture [RDJ19]. marked [GMS11]. market [HGE^{+}10, TSSL11]. Markov [AHS^{+}16, BWA17, CLCC11, CTMS10, FCH^{+}17, GLI^{+}21, KGM12, LLC11, LCMG18, PHC10, Qia12, ST13, VRM^{+}10, YZL12, YZC14]. Markov-CA [YZC14]. Markov-Chain [KGM12, CLCC11]. Markovian [ZGF14]. marsh [AHM^{+}16a, CEWW18, DVD^{+}14, FMLM11, LdMS^{+}16]. Marshes [WYPF15, CHPS17, GGZ^{+}16, MMK15, ZJL^{+}14]. Maryland [WTC^{+}18, TFMP11]. mask [SM10]. Mass [OBC^{+}21, BLCPB11, CW17, CRUF13, DCZ^{+}15, FE14a, JHH^{+}12, KZR^{+}13, LJR14, MDW^{+}12, MDL^{+}10, MSO10, NOOM^{+}16, PLR^{+}13, SBC^{+}11, URUNAM12, ZKR^{+}13]. mass-balance [BLCPB11, DCZ^{+}15, MDL^{+}10]. mass-balanced [CRUF13, PLR^{+}13]. massoniana [ZXD^{+}13]. mast [TB13, VAB^{+}18]. match [SKM^{+}16, BHD17]. Matchstick [BW12]. Mate [VF16, GCR18, MRB14]. Material [UAB^{+}11, BJK^{+}16, LDJZC15, LWTZ18]. Mathematical [DCMS18, DP15, JWLS17, KK15b, LAK^{+}21, LDP14, MWB^{+}10, Oga18, YTT^{+}11, ZHV14, ALL11, BVC^{+}18, BAT^{+}15, CCF13a, CT12, GSJ^{+}10, GMLRF10, LV17, LPPA15, MSKS13, NGRDTHRL19, PFP^{+}16, PM16, REB^{+}16, SAI10, WHB^{+}16, YXZ^{+}21]. mathematics [VMG^{+}12]. mating [AOW12, VCE17]. Matla [RBM^{+}17]. matrices [Log13b, UG12]. Matrix [JBV^{+}15, Log13a, AD13, BG19, CMA^{+}13, FLLB11, HFSL14, IPSH14, KFDL^{+}19, LLCR10, MGD^{+}12, MPF^{+}14, MHG18, Mon13, Pat12, Pat14b, POBH10, San19]. Matsumura [KL10]. matter [BP15, BP16, BP17, BP19, BL15, BKMK10, BBH^{+}11, BPD^{+}19, CKS^{+}17, CSS^{+}17, CGG^{+}18, HSK17, HN14, JG10, KAJØ19, KCB^{+}17, LKP13, MSV12, Nei11, OSS13, RVB^{+}10, SDD^{+}13, SBRD10, TDC^{+}11, WHC^{+}13, YCF15]. matters [DDR^{+}18, MU17, XCG^{+}16, ZZA^{+}18]. mature [MCS^{+}18]. maturity [GAV^{+}11, NJCF11]. Maui [dJHMS19]. MaxEnt [HMD^{+}16, TLLX19, ZLY16, AG11, CDR^{+}13, MAMNL^{+}15]. maxima [SAI10]. maximization [Log13a]. maximizes [LB15]. Maximizing [BK13]. Maximum [KSZW16, LKG10, BH17, BND19, Dil10, DS21, FvBM^{+}11, GGV11, KG13, LG10, LG12, MSGG14, MMS12, RFZB12, RHR^{+}15b, Sci11b, SCHGM16, SKM^{+}16, Til15]. maximus [MG18, MHG18]. May [Ano10h, Ano10j, Ano11x, Ano11y, Ano12-31, Ano12u, Ano13-37, Ano13v, Ano14d, Ano15c, Ano15n, Ano16c, Ano16i, Ano17y, Ano17-27, Ano18-36, Ano18-45, Ano19-44, Ano19-43, WTC^{+}18, BMD18, CEWW18, Log17, PJ17]. MC2 [PBFS14]. McGregor [dOMZ13]. MCMC [WC13a, WC13b]. meadow [CSS^{+}19, FBDZ12, SLZ^{+}17]. meal [KAJØ19]. mealybug [FTK^{+}14]. mean [SBVA^{+}13, YTT^{+}11]. meaning [MCPB15, RHD^{+}13, RHR^{+}15c]. meaningful [MAFG^{+}19, RRH^{+}19]. means [BP17, BMF^{+}11a, BMF^{+}11b, GMLRF10, NMC^{+}14]. measurable [SDD^{+}13]. measure [AOW12, KFvSH15, mLcLJ15, mLcLJ16, PdM11, VPB^{+}12, dHR18]. Measured [KLS^{+}14, HMS^{+}10]. measurement [HTB13, HTB17, SP12]. measurements [Gra14, PSKO16]. Measures [Sma13, AMRH19, BNMP11, BvHR^{+}15, CLD^{+}17, GGZ14, GJB^{+}11, HL19, KM12a, NL12]. Measuring [ACC^{+}10, AWSS14, AS14, DSM^{+}12, GCV14, Hoa11, MPGCS14, MSFR15, TdS12, ZF19, ALKB10, Hos11, MdMS^{+}19]. Mechanics [KR16, RHR^{+}15a]. mechanism [BCC^{+}14, hCjHMW18, GGR^{+}13, HDL^{+}16, KKI^{+}14, LWTZ18, XJY^{+}16, XZCL16]. Mechanisms [CH15, CMM^{+}10, DSW^{+}17, GGSP14, LYCU14, PMF^{+}13, PMP^{+}13, RPSA12, SCHQA16, TBA19, WSE10, YDT^{+}13]. Mechanistic [LHO^{+}16, SBDS13, TKN14, ACC11, BWH13, BMR^{+}18, BCB^{+}11, BPM^{+}17, CHPS17, CKH15, GHWG^{+}14, GBW^{+}15, HB12, HOKH18, LAR^{+}16, PAB13a, PVL^{+}14, POB18, SvdMKS11, SAR11, SJB^{+}16, TS11, TWL11, WHNR19]. mechanistically [SBM^{+}11, THO10b]. mediate [BPN12]. mediated [BMR^{+}18, CWJ^{+}13, FC12, HTL^{+}14, SRM11, ZSKI17]. mediating [ERGW^{+}11a, ERGW^{+}11b]. mediators [GSSM13]. medicinal [GWR11]. mediopunctata [CMH14]. medirostris [HMP^{+}19b]. Mediterranean [BDV^{+}19, DBVS19, PPB^{+}18, WLS^{+}11, AKvO^{+}10, ABGB15, BSASSJ^{+}13, BC11, CCCP17, CRAP13, CHR^{+}13, CB19, CSS^{+}16, CVM^{+}18, CQS15, FAC^{+}19b, FPG^{+}10, GMP13, GDDFT16, HWE^{+}18, IGBSASSJ16, KAF^{+}14, LCAJGFNS14, MSGG14, MFESBG^{+}13, MVC^{+}17, PBD^{+}17, QDH^{+}10, RPA12, SAR11, UDNSLC^{+}18, WLS^{+}10, aGSCMF15]. Mediterranean-type [HWE^{+}18]. medium [NSIiT17]. meets [PLB16]. mega [WYY^{+}19]. megafaunal [Flo14]. megalopolises [LZJ^{+}19]. megalopolitan [TJYZ11]. Meganyctiphanes [JR16]. megapopulation [MLG18a]. Mejillones [ROSC15]. Mejillonesa [OBC^{+}21]. melanoleuca [GFW^{+}17]. melanophys [HS11]. mellifera [RBH13]. mellowed [Jen12]. melodus [GKF^{+}14]. memory [ADF13, EIG^{+}10, GM10, GPZB18, KKI^{+}14]. memory-based [KKI^{+}14]. menhaden [DC10]. mental [KHHN17]. mercury [AML13, CW17, FE14a, FE14b]. Merging [AVG14, MK19, SJGB16]. Merodon [KTV^{+}15]. meromictic [PBK14]. mesh [Lec17]. Mesic [SHB^{+}11a]. Meso [MTJF11b]. Meso-American [MTJF11b]. mesocosm [KPS19, SPP18]. mesocosms [DJT^{+}19, MM16]. mesoFON [GSS^{+}14]. mesophilic [Vav10]. mesophyll [MEC14]. mesopredator [ACE^{+}17]. mesotrophic [CT11]. mesozooplankton [ANMJ12, AHS14]. meta [BBC^{+}19, CR12a, JBB^{+}11, JS12, SLSS15, TŠAD12, WGA^{+}13]. meta-analysis [WGA^{+}13]. meta-ecosystem [CR12a]. meta-ecosystems [BBC^{+}19, JS12]. meta-heuristic [TŠAD12]. meta-model [SLSS15]. meta-population [JBB^{+}11]. metabolic [AYASR^{+}17, BND19, CSL^{+}12, DLL^{+}19, FD19, HZZ^{+}18, HB12, RMLTCQ^{+}19, Wal16, XTLM15, ZPZ^{+}15, ZYFL10, ZLF^{+}11, ZLC13, ZLF14, ZLF^{+}15, ZLFZ16, ZMH19]. metabolism [CdMMF16, Che15a, CHVH13, CLNN14, FKWZ19, GZS^{+}18, GRD^{+}14, Hal11, HM18b, LHL18, LWTZ18, MGW^{+}13, XZSL17, YKZZ14, ZYLY11, ZLZ17c, ZLZ17b, ZWF18, ZF19, ZHCX15]. metabolisms [LG17, MTL^{+}10]. metabolites [SWS16]. metacommunities [BND19, CH19, DRS^{+}16]. metacommunity [Con11, Con12, FGPR10, SBC^{+}15]. Metagenome [RMLTCQ^{+}19]. metagenomes [ZPZ^{+}15]. metagenomics [AYASR^{+}17]. metal [AJDS16, GSPSP15, RM13]. metals [BCC^{+}11, BCC^{+}14, GVC10, HZT^{+}13, MAP^{+}13, MB17, TLL^{+}17]. metamer [GCR18]. Metapopulation [GBB13, NSN17, UM18, BD18, BP10, BNP10, CJM^{+}12, CDG15, KPLS15, LJAD17, LBR16, LW17, MAM^{+}12, MRD^{+}12, MRM^{+}17, PP10, Pec12, RJG^{+}13, Uch16, WPG15]. metapopulations [MP13, RHO11, RAdA12, RP10, SSS^{+}11, WFA^{+}13]. Meteorological [SHM^{+}17, CTP^{+}17, Jam16]. methane [KVS17, LHZS10, VRC17, VRL18]. methanization [Vav10]. methanogenesis [COPP13, PJ16, Vav12]. methanol [Vav10]. Method [CVS^{+}13, HKVD16, AFC12, AH17, Bai18, BLBB13, BRHD15, BHPF14, CGV^{+}15, Cho13, DFL^{+}16, DB15, DDR^{+}18, FHL^{+}13, GH13, GB16a, HMD^{+}16, Jør15a, Jør15b, KDK^{+}10, LSZC10, LPT^{+}18, LWC^{+}18, LZZW17, LP17, MLY^{+}19, OKS^{+}11, Sci11a, SML12, TFMP11, WLR^{+}11, WHC^{+}13, XPY^{+}14, YKL^{+}12]. Methodological [GRRU15, CBR^{+}16, DWNN10, DAF^{+}11, FCLFB18, JTM^{+}17]. methodologies [BCT15]. Methodology [MA10, SBC13, BHD16, BYR18, CLBSG11, KCP^{+}15, MK13, YL18]. Methods [BMCG12, FG11, Lec17, LLCR11, SW11, SZVV10, VG19, AADA14, BTM^{+}10, BCJ^{+}12, BS10c, BP10, BPD^{+}19, Cam15, CGL15, ČTT^{+}12, Cho13, DPLC13, FIG^{+}10, FBH^{+}14, FKK^{+}18, GdSHA10, HLW15, LBR^{+}17, LLCR10, LWY^{+}11, MGW^{+}13, PWBG16, PBO19, Pat12, Pat14b, PMH17, PES19, PBT^{+}11, Pér18, RWS^{+}15, SZV14, SvWP^{+}15, Sil11, THBF^{+}17, TAGC^{+}16, WTG^{+}16, ZLY16, Jör11a]. methylisoborneol [HD14, HD16]. Metric [NG11, GO15a, SLY^{+}17]. Metrics [Mar12, BFS11, FSB13, FCA18, HHD12, Kos17, LBH^{+}15, SvWP^{+}15, SDEP18, WBB^{+}21, WZL^{+}16, dHR18]. Metropolis [CTMS10]. Metropolis-Coupled [CTMS10]. metropolitan [HLW15, LDYL18, TOQW11]. Mexican [FMBD15, MROGVR12]. Mexico [RR10, BLLJ17, CNH^{+}11, CMCASS12, HHC15, KWGG19, LLC^{+}17, LMGW16, MAC14, MZLCSZdP11, PDAL14, RIMPBT19, SLW17, ZRCNLI^{+}15, dMSL^{+}16]. MF [DA17]. MFA [mNkGyT^{+}17]. MFA-based [mNkGyT^{+}17]. Mg [GVV^{+}16]. MIB [HD14, HD16]. Mice [TB13, PME^{+}16]. Michigan [SGDP19]. Michx [ARM^{+}12]. Miconia [PMS11]. micro [BA14b, DCMK11, FSV^{+}12]. micro-organisms [DCMK11]. micro-scale [BA14b]. microalgae [RM13, TIC^{+}14]. microalgal [BAP18, ZYZ17]. microbial [CDL^{+}11, GGVB11, IMW^{+}17, KMKL19, LS18, LD16, MFD^{+}15, SMLH^{+}11, Vav10, Vav12, WHB^{+}16]. microbially [PHP^{+}14]. microbiome [RMLTCQ^{+}19]. microclimate [FP13]. microcosm [LHO^{+}16]. Microcystis [CIHL14, LZYH19, MUP^{+}13, MvdWU^{+}16, WFW^{+}17, YLC^{+}18]. microecosystem [LBN^{+}18]. microendemic [PMF^{+}13]. microhabitats [BPN12]. microorganisms [BVGR^{+}17, SBRD10]. Microparasitic [BPK^{+}16]. microphyllum [HEJP17]. microphytobenthos [HPG^{+}10, HPRNP12, RGU^{+}15]. microplus [WTG^{+}16]. microscale [VPM^{+}18]. microscopic [DLCR13]. microtopography [WRM^{+}11]. microwave [BW12]. microzooplankton [SPP18]. Mid [CAMNN16, LLZ^{+}16, DMR13, Fat10, RPP^{+}15]. mid-2010 [Fat10]. mid-Atlantic [DMR13]. Mid-Continent [LLZ^{+}16]. Mid-Norway [CAMNN16]. mid-water [RPP^{+}15]. middle [GWZ^{+}15, IGR^{+}18, JX17, Mur12, MSB^{+}17]. midlittoral [FAC^{+}19b]. midst [RW10]. Midwest [PANL11, LLT^{+}14]. midwestern [LN11, LN16, CPS15]. migrating [ATL18, ZBMC14]. Migration [ACE^{+}13, AMG19, BVP^{+}13, BLFS18, BS10c, CWLC13, CAMNN16, DDBK15, DP15, FSN^{+}15, Flo18, GGZ14, HMA10a, LCD^{+}16, LSJG17, MvdWU^{+}16, MSO10, NZL13, NWAR11, OAPP12, PR13, PHP15, RCSS13, RH15, SH18, SSW^{+}18]. migrations [BVJE18, MNHW^{+}13]. migratory [AGB^{+}13, JJKJ11, KNPH^{+}17, KK12a, OOH^{+}12, TZD^{+}15]. Milking [Mat17]. mill [RDJ19]. Mille [KVP16]. millennia [CK12]. MINDY [GBR^{+}13a]. mine [ZF11]. Mineral [JZCL14, NPH^{+}17]. mineralisation [OTA16]. mineralization [Gra14, MRP^{+}11, ZTJ^{+}10]. minerals [DB15]. miners [HS11]. mini [LTL^{+}13]. mini-model [LTL^{+}13]. minimal [CCW^{+}15a, CCW^{+}15b, Loe14]. minimally [MCL^{+}17]. minimises [CMD12]. Minimizing [JTM^{+}17]. Minimum [SPM^{+}18, JMW11, LH15, LYSW15, RRH^{+}19, VEK12, VCSSC12]. Mining [WC13c, BRHD15, LST15, NW18, ZFT^{+}10]. Minnesota [JFS^{+}12, KVP16]. minor [LBdA^{+}19]. Miq [HH12]. mirages [UM18]. Miscanthus [MWD^{+}15, Pog11]. misclassification [BMP^{+}19]. misidentification [LCMG18]. Mississippi [DWNN10, DVF^{+}19, PSBH16, BPC^{+}17, SCB13, WCL^{+}17]. Missouri [WAGM17, SHX^{+}12]. mistakes [AW15]. mite [dOMZ13]. mitigate [dSGGBW12, SS16]. mitigated [KPTC14]. mitigation [ACSR^{+}13, ACC11, BNMP11, BG15, MSS^{+}19, RHB14, SSN^{+}18]. mixed [ACM15, BBC^{+}19, CCCP17, FT16, GBP^{+}10, KKBZ17, LAD11, RLS11, RHZ^{+}18, STSA10, SKK^{+}13, SMH16]. mixed-effects [CCCP17]. mixed-forest [ACM15]. mixed-species [FT16, RLS11]. mixedwoods [BHC^{+}15]. mixing [ABR12b, ABR12c, BM16, LY16, PLR^{+}13, PFR15, TGM^{+}21]. mixture [DAP18, ZR19]. mixtures [FDHL17, Gar17, YHW^{+}12]. Mn [GVV^{+}16].

Mnemiopsis [AALP18]. mobile [MRC19, MHY^{+}15, PC15, Sto10]. mobility [PJ15]. mode [KP13]. Model [AHM^{+}16b, Ano12a, BRV11, BdCS11, BMCG12, BS17, CBB19, Con12, CR12b, DFD11, ERGW^{+}11a, GAC^{+}14, GHH^{+}15, HPJ15, HLLO13a, IMW^{+}17, Jam16, Jør15a, KHKC^{+}18, KS17a, mLcLJ16, LSJ^{+}19, LXS^{+}16, Lin16, MTL^{+}10, PVM12, PVM13, Pat10, PTE^{+}12, PLS11, SELB10, SMT19b, SFC^{+}12, SYL^{+}13, SZVV10, SKTG11, WSW^{+}13, WLF^{+}16, WYPF15, YYQ11, YTT^{+}11, ZYFL10, ZLC13, ZLF14, dÁSMRT^{+}19, ATL18, AMG19, AD13, AMF12, ANMJ12, AHS14, ACCM16, ABB^{+}10, ABD^{+}18, AAJB21, AD12, AHS^{+}17, AHM^{+}16a, AK17, ABK^{+}16, Ana17, AHNK13, AD16, AD18, ARM^{+}12, ADF^{+}17, ACN14, AB10, ABL^{+}13, ABC^{+}17, AS12, ACSR^{+}13, ASV19, ACC11, ADF13, AJDS16, BTM^{+}10, BVP^{+}13, BDL^{+}13, BAAB18, BSC^{+}15a, BWH13, BAB^{+}16, BMR^{+}18, BHD17, BCB^{+}11, BSW18, BDV^{+}19, Ban11, BP15, BP16, BP17, BP19, BKV17, BCP16, BKN21, BHHM10]. model [BGK^{+}16, BLLJ17, BH17, BBDB13, BWC17, BR16, BMMC21, BS16a, BLD^{+}10a, BlMG^{+}11, BG15, BTB17, BLBB13, BH13b, BSH17, Ber13, BPR^{+}11, BRM^{+}14, BBL^{+}17, BFNH14, BBGJ13, BKM^{+}18, BNC^{+}18, BJR16, BSCG10, BBL11b, BD18, BPM^{+}17, BSHG16, BBH^{+}11, BS16b, BvHR^{+}15, BCD12, BSH^{+}10, BSZ^{+}11, BKB10, BvBB10, BS10c, BP10, BRF^{+}11, BPN12, BFKS16, BRB14, CGLA10, CAL13, CCS13, CTB^{+}12, CL12a, CC12, CDR^{+}13, CZC^{+}15, CHPS17, CPD^{+}11, CMH14, CRUF13, CVS^{+}13, CCW^{+}15a, CCW^{+}15b, CF17, ČTT^{+}12, CHvdW^{+}10, CLH17, CC10, CCSF11, CCF13a, CWL^{+}16, Cha12, CRA^{+}11, CCL14, Che10, CZW10, CZ11, CT11, CZR^{+}14, CC14, Che15b, CpW16, CDK^{+}17, CLW^{+}19, COPP13, CKS^{+}17, CSS^{+}17, CGG^{+}18, CN11, CLD^{+}17, CLCC11, CCC12, CTMS10, CH14, CDT10, CPS^{+}14, CMSPJ11, CMGF14, CCG15, CBB^{+}10, CBA10, CBA16]. model [CSFR15, CV19, CMEP15, CKH15, dSCdA18, dSCdA19, CLL^{+}15, CDG15, CWM12, CTP^{+}17, CSL^{+}12, CGM^{+}14, DC10, DATI10, DJT^{+}19, DWD19, DLL^{+}19, DAA^{+}12, DTS^{+}12, DCMK11, DdSCG11, DR17, DHC16, DBK^{+}13, DLCR13, DAW^{+}10, DUASLB^{+}12, DVCM^{+}10, DBVS19, DA17, DFRT14, DSM^{+}12, DPLC13, DKM^{+}14, DCZ^{+}15, DVT^{+}10, DFF17, DWP13, DG16, DJMBD^{+}13, DBJ13, EMH10, EGA12, EPA^{+}10a, EPA^{+}10b, ECO^{+}14, EEB^{+}17, EEK^{+}18, FPO18, FTK^{+}14, FLLB11, FSB13, FLH17, FM11, FMA^{+}12, FRM^{+}19, FBB^{+}18, Fie12, FHL^{+}13, FGPR10, FBdP^{+}16, For14, FT16, For18a, For18b, FSS^{+}17, FB13, FAPM12, FP13, FDM^{+}11, FPG11, FPWG14, FBKS17, FS18, GH13, GWHL15, GYL^{+}15, GHS^{+}16a, GLC^{+}18, GCA16, GEEC14, GDT^{+}14, GZK^{+}11, GvdWBS11, GGSP14, GDTP13, GSA^{+}19, GMLRF10, GKL18, GS19a, GS19b, GORJ12, GU11, GO15b, GANB11, GPZB18]. model [GBP^{+}10, GFP^{+}16, GC14, GS10, GBR^{+}13a, GBW^{+}15, GLS10b, GAF^{+}14, GJK^{+}15, GJB^{+}11, GNB^{+}14, GMvV17, GSS^{+}14, GLS^{+}17, GCW^{+}16, GLI^{+}21, GPD^{+}13, GLY^{+}15, GLRT16, GS19c, GCM^{+}13, GPC14, HHC^{+}13, HM18a, HM19, HMD^{+}16, HFSL14, HMP^{+}19b, HKL^{+}11, HLS^{+}14, qHLlZ^{+}16, HZSH17, HFH^{+}19, HHD12, HBP18, HD14, HD16, HSTS19, HTL^{+}14, HFB^{+}11, HCL^{+}14, HHL17, Hei11, HHC15, HW12, HGAT17, HBSF16, HBH^{+}12, HLK^{+}12, HHvO^{+}17, HVCP16, HSVL12, HL19, HOKH18, HPC13b, HRWT12, Hos11, HWL17, HNWT10, HHF^{+}10, HGLZ13, HBA16a, HBA16b, HG17, HYL^{+}17, HAA^{+}21, HBL18, HGN^{+}16, HSD^{+}15, HPW^{+}12, HBB15, IFC14, IT17, IWW12, IGR^{+}18, IN16, IKH^{+}11, JSKK15, JZ12, JNHF18, JBB^{+}11, JSL^{+}10, JTB^{+}19, JZG10, JHH^{+}12, JDZC14, JCKG15, JLZ^{+}18, JOHW17, JHT^{+}14, JJS10, JL18, JN15b, JAKJ10, JDGJ13, JNC^{+}11, KR10a, KK13, KCLL14, KWS^{+}13]. model [KUI^{+}18, KB10, KW18, KWH14, KJMA11, KTGH19, KYvdPS14, KDMPC10, KL10, KLS17, KHJ^{+}19, KPS17, KHW^{+}17, KEM^{+}16, KdVRdW17, KB18, KH15, KGD14, KCB^{+}17, KPT11, KPK^{+}17, KVS17, KK12a, KWJH16, KKI^{+}14, KVP16, KKK^{+}19, KB11, KMS^{+}12, KMS^{+}13, LV17, LJLC15, LJBS12, LKMCW14, LAK^{+}21, LKKK^{+}14, LBR^{+}17, LDG14, LAR^{+}16, LJR14, LDS14, LLB^{+}19, LAD11, LMT13, LSR16, LBSB^{+}14, LJ15a, LKLK15, LSA^{+}19, LLC^{+}17, LBE^{+}13, LH15, LWKS10, LPPA15, LF11, LkSXhX17, LLL^{+}10, LHZS10, LAB^{+}10, LHZ10, LHZN11, LDL12, LTL^{+}13, LCL^{+}13, LLM^{+}13, LZZ^{+}14, LLT^{+}14, LJ14, LCCL15, LLZ^{+}15, LJ15b, LYSW15, LYJ^{+}15, LJ15c, LLZ^{+}16, LHP16, LSJG17, LYD^{+}17, LS18, LZ10, LHB^{+}18, LQW^{+}19, LLQN13, LDP14, LDM15b, LMCA^{+}12, LKP13, LGY10, LLC11, LVZ^{+}11, LCTH12, LSGT13, LBS^{+}14, Loe12, Loe14, Log13a, LIM^{+}16, Löp11, LSL10, LHLW15, LL18b, LSM19, MTZ^{+}10, MLW11]. model [MW12, MW15, MCWG21, MBSW12, MDW^{+}12, MMD^{+}17, MSP^{+}13, MCH15, MGMR15, MD16, MRM16, MBG17, MJB^{+}11a, MPM^{+}15, MRN12, MKL14, MLMFN^{+}15, MdlP15, MJL^{+}13, MJL^{+}14, MYHN17, Mat17, MKSL^{+}17, MTGD17, MKTO19, MFS18, McC16, MU10, MBUM12, MGCL17, MSCM18, MvdWU^{+}16, MTJF11a, MAP^{+}13, MSF18, Md18, MG18, MHG18, MK19, MDL^{+}10, MRES17, MWML13a, MLAT^{+}18, MBM^{+}18, MPL^{+}16, MJF19, MH10, MSHC13, Mon13, MGSB11, MMGK15, MP17, MAMNL^{+}15, MSCH14a, MSCH14b, MDVG^{+}15, MSO10, MCP16, MSFR15, MATS14, MGW^{+}21, MEC14, MJGC13, MMMCAHM21, Mur11, MS16, Nak17, NP18, NOOM^{+}16, Ned15, NBK^{+}11, NB11, NPJ^{+}17, NMA^{+}19, NPH^{+}17, NGRDTHRL19, NG11, NiSK21, OE12, Och10, OJ10, OFH^{+}19, OJ15, OWMK10, OCCF17, OKS^{+}11, OGOWW16, PRHH13, PKB^{+}18, PSB^{+}17, PBB^{+}14, PC17, PCTW17, PFP^{+}16, PDAL14, PGW19]. model [PHP^{+}14, PBD19, PLS^{+}10, PHC10, PD12, PAB13a, PBFS14, PLL^{+}19, PRH17, PMRGF10, PWK^{+}17, PGK^{+}13, PKW^{+}10, PP12, PNHL14, PZM13, PHP15, PGG14, PvVC13, PMS11, PGGMMA19, PMTKC17, PBK14, PPO12, PME^{+}16, Qia12, QCLW18, QDH^{+}10, QFG^{+}13, ROG^{+}16, RKL^{+}15, RVL^{+}18, RML^{+}15, RMP^{+}18, RJH^{+}10, RL15, RSE10, RFZB12, RMTM^{+}16, REB^{+}16, RPM13, RAF17, RRHH10, RSDP^{+}12, RHZ^{+}18, RPC^{+}17, Rin14, RPGA15, RNS^{+}19, RGRMRR18, RRJVA13, RCC^{+}11, RCG^{+}13, RDR^{+}13, RDD^{+}17, RCGC17, RAB^{+}10, RMHE15, RWvDM15, RDD^{+}14, RPGSCF16, RPA12, RML^{+}13, RR10, RWP11, RBGB16, RP11, SZSV12a, SJTS13, SLW17, SRHL11, SBCB13, SMSAN^{+}14, SLSS15, SFM^{+}18, SNS^{+}12, SBM^{+}12, SR14, SHT^{+}21, SAAD16, SKHH19, SW11, SBDS13, SJS^{+}13, SAR11, SM13, SDD^{+}13, SRSS12, SH14, SHCR10, SFvB10, SAI10, SHX^{+}12, SX11, SMS^{+}13, SLY^{+}17, SEHD18, SAW^{+}11]. model [Sil11, SMH16, Sim11, SSKJ18, SCH^{+}17, SFL^{+}19, Ske13, SSN^{+}18, SBPZ12, SBA^{+}13, SMG^{+}12, SGG14, SGSR10, SKM^{+}16, SCG^{+}13, SBC^{+}16a, SBC16b, SGWV15, SvdR10, SGDP19, SRT16, Str15a, Str15b, SGBG19, SP12, SM14, SCG^{+}16, SLZ^{+}17, SSW^{+}18, TTCH^{+}14, THC18, TLLX19, TM12, TGM^{+}21, TCW11, TM18, TŠAD12, TB12b, TOHT14, TT12, TAC^{+}14, TSM^{+}15, TZS^{+}14, TLZ17, TRE^{+}12, TL15, THO10b, TTCC^{+}19, TWH^{+}14, TLRL11, TWL11, Uch16, URUNAM12, UG12, UDNSLCG16, UDNSLC^{+}18, VMSM10, VSF10, VPB^{+}17, VRL14, VH12, VC17b, WB10, WN17, WQD^{+}13, WB14, WLZC11, WPZ^{+}11, WGA^{+}11, WGA^{+}12, WLDL12, WHB12, WGT12, WBC^{+}12, WC13a, WC13b, WMF^{+}13, WGA^{+}13, WJL^{+}13, WM14, dWcWbZ15, WSL17, WTWS19, War10, WHRR15, WO14, WNF^{+}15, WNT^{+}13, WXB10, WV12, WJF^{+}11, WBG^{+}14, WPB14, WCG^{+}11, WS15, WAGM17, WC13c, WLS^{+}10, WLS^{+}11, WFN^{+}16]. model [WHNR19, WZL^{+}16, WX11, WPZ^{+}14, WWYO17, XLR11, XZC^{+}17, XPY^{+}14, XYY^{+}16, XZCL16, XSN^{+}18, YL17, YZL12, YLC13, YZC14, YLF^{+}15, YDZ^{+}21, YXZ^{+}21, YZG16, YY11, YY13, YSA^{+}14, YLL^{+}12, YZLX14, YCL^{+}14, YKvWK16, YHJ^{+}17, ZMMHR10, ZHC^{+}16, ZBMBDU^{+}17, ZTJ^{+}10, ZLF^{+}11, ZJM^{+}11, ZLFI11, ZWG^{+}13, ZWRK13, ZJL^{+}14, ZQC15, ZSJ15, ZJL^{+}16, ZLG16, ZLJM17b, ZLZ^{+}17a, ZLJM17a, ZCSL18, ZML^{+}12, ZCS13, ZXD^{+}13, ZZT13, ZBS14, ZZL^{+}19, ZSF^{+}12, ZHLZ17, ZVS18, ZJG10, ZSKI17, dJGM^{+}18, dMSL^{+}16, dOMZ13, dSD18, vDHH^{+}14, vPTOW18, BRM^{+}14, BW12, BKB10, CPS^{+}14, DFM12a, SMBGH^{+}19]. model- [MSF18]. model-assisted [DSM^{+}12]. Model-based [PLS11, SZVV10, dÁSMRT^{+}19, ABC^{+}17, BG15, SBCB13]. model-building [SBC^{+}16a]. model-data [RHZ^{+}18]. Modeled [GEZ^{+}13, XLL15, MGS16, XGH^{+}17]. Modeling [ABR12b, ABR12c, AS11, An12, AML13, AGB^{+}13, AHNK13, AB10, AHS^{+}16, ABGB15, AHF^{+}13, AOD^{+}13, BCP15, BAA^{+}14, BBB^{+}17, BKP10, Bar11, BNL^{+}13, BVVS15, BMCG12, BCL^{+}14, BLRN15, BNGP19, BA14b, dÁGP14, dABLBAdOR17, BH10, BN10, BPD^{+}19, CPMP13, CLBG15, CTNH13, CESC19, DJS10, DDBK15, DMZ^{+}19, DJA^{+}10, ESA12, ELPK12, EEFE^{+}11, FSN^{+}15, FMS13, FGH^{+}10, FMW^{+}19, FBH^{+}14, FHG12, FCS13, FKDG14, FBDZ12, GPK^{+}13, GHD^{+}19, GQNMR18, GDC12, GdSHA10, GV11, Gli14, GWK^{+}12, GKW^{+}13, GCCJ12, GRB12, GLI^{+}21, HH12, HN14, HPS10, HZT^{+}13, HKlL14, HSR^{+}15, HKN17, HC11, HNWT10, yHAB12, HQF^{+}17, IGBSASSJ16, JHH^{+}12, JWX18, JAH^{+}10, Jör11a, KKH^{+}17, KD13, KH11, Ken17, KSM^{+}18, LRZ12, LBE^{+}13, LHL^{+}17, LBE^{+}18, LBWW14, LHZ10, LZLC13, LJ13, LJ15c, LZL^{+}18, Lin15b, LL18a, LPSC16, MLW^{+}15, MKF^{+}11, MTPC13, MM16, MRB14, MSV12, MDKK15]. Modeling [MJF19, MZLCSZdP11, NP15, NPH^{+}17, Oga17, OBSL15, OdlFDP15, PJ16, PH18, PSBH16, PAB13a, PANL11, RZF^{+}13, RHZ^{+}18, RDJ19, RNE^{+}12, RA13, ŠK14, SCHQA16, San17, SW19, SLB14, SHS^{+}12, SSN^{+}18, STJ15, TW18, TXM^{+}10, TDF^{+}19, TT17b, VSPHS15, WQD^{+}13, WJL14, WDL^{+}15, WB17, WLH^{+}10, WGD^{+}17, WTL17, WSBP13, WFDDR10, WJK11, XCG^{+}16, XTPT11, XC14, YLHZ14, YZA^{+}18, YHZ19, ZBMC14, ZLLC13, ZLW^{+}15, ZZLZ14, ZHCX15, aGSCMF15, ACC^{+}10, BS18, BVJE18, BBRJ15, BBJV^{+}11, Bec10, BMN11, BAP18, BS16a, BEPC15, BT16, BJPB11, BWA17, BLCPB11, CRBC11, CT16, CJLV11, CVPH^{+}14, CWLC13, CdMMF16, CN10a, ČTT^{+}12, CHCC15, CL10, CC11, CLA^{+}12, CB15, CS18, CSF11, CVG^{+}15, CMSPJ11, CPS15, CRQW18, CKMD14, CLNN14, CSFB17, DH18, DSW^{+}17, DSBS15, EKB18, ER13, FKH^{+}15, FCZ13, FP17, FYWA12, FDHL17, FHC15]. modeling [FWRZ10, GWR11, GBG^{+}15, GGZ^{+}16, GCB^{+}11, GGR^{+}13, GCK^{+}15, GHS^{+}16b, GCvDM12, HGPF15, HJ17, HlLB12, HTVE18, HPG^{+}10, HR11, HEJP17, IKK^{+}12, JVY19, JD18, JJKJ11, JC11a, JXL^{+}15, JDT^{+}16, JSCS19, JMB^{+}11, JBW^{+}17, KCF^{+}13, KSSA13, KPSC10, KM12b, KF16, KPLS15, KJS13, LBR16, LMSK12, LZS^{+}17, LPV11, LHR^{+}15, LLP17, LZW^{+}17, LZYH19, LBC^{+}15, LPD14, LVV^{+}15, LAEAG12, LBGM19, LD16, MSG^{+}13, Mar17, MXC13, MROGVR12, MLS^{+}11, MRD^{+}12, MDF^{+}18, MSM^{+}17, MDOJ14, MJLC14, MRR^{+}14, MRG13b, Mul11, Mur12, NFU19, NLMdL10, OHS^{+}17, PCRP12, PNT^{+}19, PFA^{+}16, PRL^{+}13, PMLL11, PSM^{+}12, PDG^{+}14, RJG^{+}13, RJ11, RPSA12, RRR^{+}10, RFZB12, RvKM10, RRG^{+}15, RSD^{+}15, SGBD10, SBR^{+}19, SFMA15, SMH^{+}12, SKH^{+}13, SQWS19, SLH^{+}11, SK11, SKL^{+}13, SCG^{+}13, SV13, TSN15, TB11, TB12a, TMPF18, THJ^{+}12, TZS^{+}14, TKN14, TBL^{+}15, Tul14]. modeling [UKI^{+}11, VVH12, VFFP16, VGR^{+}11, WALCY12, WMW^{+}12, WKF^{+}17, WSE10, XJY^{+}16, XCR^{+}13, XQZ^{+}13, YSH^{+}10, YHS^{+}11, YJW15, YWY10, YLZ15, YDL^{+}15, ZRCNLI^{+}15, ZCR16, ZCL^{+}16, ZJB19, ZDG^{+}13, dAEPE16, dMLM^{+}17]. modelization [VJR^{+}16]. Modell [ABR12b, ABAG16, Bar11, BMF^{+}11a, DYMYGO19a, WLS^{+}11]. modelled [COBF13, EC10, HMS^{+}10, KLS^{+}14, SMC11, ZGF14]. Modelling [AALP18, ABAG14, ACM15, Ano12t, Ano13u, ABAA21, AOW12, AOW13, ARV^{+}16, BNMP11, BDL^{+}13, Bae18, BLY11, BJK^{+}16, BPS^{+}11, BLD^{+}10a, BlMG^{+}11, BFJZ^{+}11, BKMK10, BWP16, BHC^{+}15, BTBM17, BSO13, CMT^{+}11, CEC10, CLBD^{+}11, CGV^{+}15, CSN^{+}19, CY13, CDK^{+}17, CSS^{+}16, CGEdM12, CLBSG11, CK12, DJT^{+}19, DBR12, DOB^{+}15, DYMYGO19a, DYMYGO19b, DBG14b, DH12, DVD^{+}14, DFM12a, DFM12b, DMBS11, EH19, EAJG16, EGMW12, FDP^{+}10, Flo14, FBL^{+}10, FKK^{+}18, GAP^{+}19, GPJL15, GBP^{+}16, GCC^{+}10, GLA11, GSPSP15, GGV11, Gat18, GPTR13, Gra14, GHS^{+}10, GRD^{+}14, GBB^{+}11, GKKW10, GPM16c, GCL^{+}19, HMPC11, HBS^{+}16a, HGWL11, HV15, HSHF^{+}13, HCP10, HPGM18, HB10b, HHK19, HJ10b, HSvO^{+}18, JR16, JX15, JDT^{+}17, JMdW^{+}10, JDP^{+}12, JGB15, JA15b, JASW18, KAJØ19, KSS11, KJVL^{+}17, KBC^{+}12, KPMJ14, KR14, KRCVF16, KvTL^{+}18, LN11, LYP^{+}11, LPA^{+}11]. Modelling [LBE^{+}17, LMR16, LKH^{+}18, LLC^{+}21, LLH^{+}13, LYCZ13, LCFD19, LAK^{+}18, LAC10, LTGD16, MRCW10, MPKC12, MJP^{+}12, MVC^{+}17, MFD^{+}15, MBG^{+}10, MTJF11b, MG18, MHG18, MMC^{+}19, MWML13b, MB17, MRP^{+}11, MWM^{+}16, MNHW^{+}13, MPG18, MOM^{+}11, NW18, OTA16, OES16, OC15, OKS^{+}11, PBL^{+}15, PVBC14, PGZ^{+}19, Pie10b, Pie10a, Pog11, RGU^{+}15, RvS16, RW17, RSP^{+}17, RCST11, RUSL19, RKW12, RIMPBT19, RSC12, RSC13, RBRK16, RPP^{+}15, SDT10, SZV14, SF10, SvdMKS11, SGA^{+}16, SFF^{+}11, SFC^{+}12, SLW^{+}14, SKMB11, SKK^{+}13, SVH^{+}19, SBRD10, SBC^{+}11, SMKM19, SGH16, SKPHW16, TVF^{+}10, TSSL11, TIC^{+}14, TGB^{+}17, VMG^{+}12, VRC17, VCSSC12, VASH^{+}16, WL13, WMK^{+}13, WFS^{+}19, Wil11b, WBW14, WJvdL^{+}13, YLL^{+}17, YLC^{+}18, YFL^{+}18, ZSS17, ZBAY14, ZLFZ16, ZPNM13, dJHMS19, AWF10, AVL^{+}11, ABR12a, AKvO^{+}10, AW15, ABS17, AMH16, ADJ^{+}17, ACDC17]. modelling [AMR16, Ash10, ABD14, Bai18, BGA17, BC13, BKN14, BLD10b, BLB^{+}17, BBC^{+}19, BRW^{+}12, BQF^{+}18, BOF^{+}16, BAJ^{+}11, BLAV18, BNP10, BPR^{+}15, CADB19, CRO13, CHR^{+}13, Che15a, Che17, CLT^{+}11, CAMNN16, CSR15, CB19, CDL^{+}11, CPS^{+}19, CTC^{+}18, dSCdA19, DZS12, DR14, DG11, DCMS18, DVF^{+}19, DDR^{+}18, DP15, DWP16, DSA^{+}18, EFM^{+}10, EIG^{+}10, FCG^{+}19, FCFT16, FD19, FBU14, FAT^{+}15b, GSQ17, GSJ^{+}10, GJÁ^{+}10, Gar17, GPH17, GCCB18, GBB13, GPM16b, GPCG15, GMC^{+}19, GCTS15, GGVB11, GRL18, GdC14, GAF^{+}14, GB16b, GB16d, GDW12, GLY^{+}15, HA18, HBF18, HJWG12, HMD^{+}16, HPJ15, HVL^{+}19, HB13, HCM^{+}18, HKVR13, HMA10a, HJN12, HM17, HOŠ^{+}12, IBH^{+}15, IS13, JBE^{+}15, JG10, JM15, JJYM19, JHHS17, Jør10b, KK16, KGM12, KPM^{+}14, KAF^{+}14, KR10b, KEHG17, KBL10, KL12, KK15b, LHO^{+}16, LBZA19, LMY^{+}11, LBA14]. modelling [LBA15, LY16, LJ10, LWNG13, MHL^{+}19, MML^{+}11, MBD^{+}18, MBD^{+}12, MBSF14, MJY17, MU17, MHS^{+}17, MSH^{+}19, MMS12, MVC11, MAFG^{+}19, Md18, MSV^{+}11, Mer12, MOMC17, MPP^{+}15, MGS15, MMC11, MB14, MLG^{+}18b, MSM^{+}16, Mur13, MYJ^{+}14, Nei11, NMC^{+}14, NXZ18, NJCF11, OFK^{+}12, OMH12, OSS13, ODK16, PMB^{+}19, PLFO^{+}12, PVL^{+}14, PLB16, PDD^{+}18, PGA^{+}11, RW10, RLHD11, RPA^{+}18, RJ17, RDR^{+}17, RRH^{+}19, RCBPMN10, RSCS19, RVB^{+}10, RPGSCF16, RBH13, SBVA^{+}13, SBA^{+}19, SFvdM^{+}17, SJGB16, SvWP^{+}15, SSD^{+}17, SVL12, SA16, SLB^{+}18, Sil11, STD15, SHB11b, Sol19, SAKB^{+}10, SGP18, SRM11, SP10, SP11, SHS11, SGHW^{+}17, SPP18, SGKR10, TPL^{+}15, TB13, TRD^{+}15, TSB^{+}18, TLW^{+}14, TKH17, TNHB12, THBF^{+}17, THOA10, TVV13, TMT^{+}18, TPT^{+}11, TPO^{+}18, UDNSLCG16, VPC^{+}14, VKdW^{+}10, VPB^{+}12, VLF^{+}17, WHB^{+}16, WN17, WAS^{+}16]. modelling [WFW^{+}17, WTC^{+}18, WGE^{+}19, WJS^{+}18, WGSR19, WCS^{+}14, WSW^{+}16, WMCM12, YBA^{+}16, YJL11, YFC^{+}11, ZMS19, ZSG12, ZRC^{+}15, ZWW^{+}18, ZF19, dFMdSA18, dOAC18, vDSC^{+}10, BH10]. Models [Ano12a, CGL15, CPE13, RCS^{+}19, VRM^{+}10, AGJ^{+}12, AGQP^{+}14, AWL^{+}13, AMRH19, AADA14, AMM^{+}14, An12, AG11, Ano17a, AB13, AAFH15, AFC12, AVG14, ALeSD17, BDL^{+}14, BVC^{+}18, BLFS18, BEZ19, BB10, BRV11, BS15, BG19, BTS19, BGD^{+}16, BGG16a, BND19, IHSM12, dÁGP14, BOGA14, BB18, BPWB14, BRS^{+}18, BTPG10, BOW^{+}12, BBH^{+}17, BFJL15, BG11, BGD^{+}15, CCCP17, CMM^{+}10, CBC^{+}16, CLLB12, CBCB17, CSOS17, CCF13b, CMA^{+}13, COY^{+}10, CMCASS12, CRD^{+}16, CVG^{+}15, CBA12, Cra10, CN10b, CN18, DWNN10, DLH^{+}14, DÖWB12, DKJH13, DPV^{+}16, DAP18, DLvGGNC16, EH15, EH19, EEFE^{+}11, EST^{+}19, Fai10, FMW^{+}19, FJH15, FE14a, FYR^{+}18, FtHvdB^{+}14, FCLFB18, FLB^{+}15, FLA19, FMF12, FCV^{+}12, Fuj16, GHWG^{+}14, GCDA16, GCA16, GGV10, GGV11, GGV13, GLR^{+}19, GWC^{+}10, GMTC^{+}17, GAHE11, GGZ14, GT14, GB16c, GPP^{+}19]. models [dSGGBW12, GAV^{+}11, HMP19a, HBR19, HSLC^{+}19, HCW^{+}17, HLS^{+}14, HHD12, HB12, HWS11, HMS^{+}11, HS13a, HTB13, HTB17, HS13c, HS13b, HB10a, HHW11, HBT^{+}18, HCL^{+}16, HB14, HGBA^{+}13, HKP10, Hu16, HhLY^{+}19, IPSH14, JGF14, JKS^{+}10, JDT^{+}17, JWLS17, JSdSdL^{+}18, KFWM19, KM12a, KK10, KMF^{+}21, KSAS12, KFDL^{+}19, Ken19, KKL^{+}12, KZR^{+}13, KP13, KPM^{+}14, KAC^{+}12a, KAC^{+}12b, KCG18, KRL11, KNM^{+}10, KS17a, KS17b, KHHN17, KWGW19, KVU^{+}10, KPS19, KSG^{+}14, KMMS17, LJLC15, LTL^{+}19, LRZ12, LBA14, LBA15, LMT13, LBSB^{+}14, LPF11, LBH^{+}15, Lem16, LBD10, LKDD15, LdMS^{+}16, LLCR11, LZZP16, LHW^{+}15, LLH^{+}17, LPT^{+}18, LS10, Lin10, Lin16, LJWN14, LDCB21, LCDJ^{+}12, LM16, LCMG18, MK13, MG16, MBHW12, MHTC15, Mar12, MGD^{+}12, MLG18a, MWK19, MM19, MA10, MTA19, MROGVR12, MC11, MCR^{+}15, MU10, MSB^{+}17]. models [MSMM11, MGS16, MPF^{+}14, MK19, MSV^{+}11, MF12, MAMNL^{+}15, MCN16, MDG10, MDV^{+}12, MBA^{+}19, MJÁ^{+}12, MMMGLN^{+}18, MRTW18, NYIH17, NS15, Ned15, NMS^{+}14, NBSGW19, NF11, NMdR16, OOFVBC16, OGG^{+}16, OT15, OBLJ13, OLG^{+}13, OBC^{+}21, OCD^{+}13, PLR^{+}13, PMM11, PSG19, PES19, Pea15, PBT^{+}11, PRP17, PD15, PWP^{+}19, Phi11, POBH10, PGS16, PLHG14, PMLL11, PFR15, RLB^{+}12, RFV^{+}12, RWS^{+}16, RW17, RLS11, Rem11, RWVK11, RAM10, RR12, RKBG14, RRJVA13, RV15, RKNO17, SZSV13, SHBG10, SC21, SVD^{+}13, San17, San19, SW11, SWS16, SBD^{+}19, SPR^{+}18, SHC^{+}15, SA13, SSK14, Sil11, SGM^{+}14, STD15, SJB^{+}16, SK11, Slo11, Sma13, SLMS14, SWR^{+}16, SPWR17, SPB^{+}17, SAKB^{+}10, SWGC16, Sto10, SMLH^{+}11, SLD^{+}19, Tat16a, TRP12, TFH12, TPG14, TMC^{+}16, TT17a, TTL^{+}18, TFMP11, TS11, TBM^{+}14]. models [TAGC^{+}16, UDRA13, VAB^{+}18, VC17a, VALC12, VHDM17, VRA^{+}17, VPM^{+}18, VAK^{+}11, VEMS12, WLY^{+}15, WR13a, WRB^{+}12, WBB^{+}21, WUB15, WBHM11, WR12, WKH19, Wil11b, WHC^{+}13, WYPF15, YSB11, YRGH16, ZBMC14, ZGJ10, ZKR^{+}13, ZLB14, ZZFH15, ZDL^{+}17, ZBSF17, ZR19, dJHMS19, dSdCF10, dSD14, vVBHZ11, vdKGV11, vdVBJS15, vdVJS16, Jør10a]. moderate [MMK15]. modes [Log19a, Phi11]. MODFLOW [ZSKI17]. MODFLOW-AgroIBIS [ZSKI17]. modification [SV13]. modifications [PW10a]. modified [KPSC10, KB18, LJ15c, WMF^{+}13, WBHM11]. Modis [dMHGA10, KKL^{+}12, LCL^{+}13, PSS16, PSKO16]. modular [Kal19, LL18a]. modularity [ZRCNLI^{+}15]. modulate [VCE17]. modulates [GTM^{+}19, PDG^{+}14]. Modulation [GCR18, Ver10]. module [ALeSD17, LYJ^{+}15, SEHD18, WZGH18, ZSJ15]. moisture [DS21, FBA^{+}11, LYD^{+}17, LPS^{+}10, NBK^{+}11, WJK11]. moisture-based [FBA^{+}11]. Molinos [REB^{+}16]. Molly [GBW^{+}15]. moment [KCG18]. monarch [GPZB18]. Mongolia [HYL^{+}17, LWL^{+}17b, SAW^{+}11, YSA^{+}14, ZJL^{+}16]. Monica [MJK^{+}17, NBK^{+}11]. monitor [Wal15]. monitored [WC13c]. Monitoring [CN10a, GZVU14, KFvSH15, KPSC10, LWY^{+}11, BJPB11, CPAG15, CR11, CR12b, HVL^{+}19, Hat14, OFH^{+}19, PWBG16, PGZ^{+}19, PRH17, RHS^{+}16, TPG14, YBA^{+}16, ZCS13]. monkey [LML17]. monkeys [WFDDR10]. monocarpic [HCW^{+}17]. monoculture [LYD^{+}17]. monocultures [RW10, YHW^{+}12]. Monod [Ano17a]. monotonicity [ZYKS15]. monoxide [LYJ^{+}15]. montado} [ABC^{+}17]. Montana [RRG^{+}15]. montane [ACM15, DH12, HBG^{+}16, MBSF14, PMF^{+}13, PTE^{+}12]. Monte [BWA17, CTMS10, HT14, Ken19, KGM12, MA10, Qia12]. Monte-Carlo [KGM12]. monthly [LCS11, ZTJ^{+}10, ZJM^{+}13]. Moor [DA17]. moorings [KL12]. moose [GJB^{+}11, JSW^{+}13, Rem11]. moose-vehicle [GJB^{+}11]. mor [LKKK^{+}14]. Moran [NS15]. moratorium [WDL^{+}15]. Morava [SSM^{+}13]. Moreton [KSZW16, WEJ11]. morhua [IKL13, JNHF18, PWK^{+}17, SLRC19]. morph [PH18]. morphodynamics [LBD10]. morphological [CWLL15, CLBSG11, YMCV21]. morphometry [MH10]. Morris [BLBB13]. mortality [BEZ19, BLW^{+}17, dÁGP14, CCS13, CHNdÁP18, CD15, DSS11, FPO18, GHST19, HLLO13a, HLLO13b, KSS11, KSZW16, MBHW12, PSG19, Pér18, RRG^{+}15, SK10, SVH^{+}19, SGWV15, WC19, War10]. mosaic [DSBS15, FPG11, FPWG14, RKC^{+}15, TRD^{+}15, ZWG^{+}13]. mosaics [BSS^{+}14, HGSJ15]. MOSES [AAJB21, TKH17]. mosquito [BVVS15, CTB^{+}12, FKK^{+}18, WKF^{+}17]. most [ACK^{+}14, RRH^{+}19]. Moth [GKKW10, GBB13, GPM16b, GCTS15, KL10, NS15, SFHK12]. motile [BEBA19]. motile-phytoplankton [BEBA19]. motivation [BA17]. motivations [PNHL14]. Motus [BLFS18]. Mount [Hat14]. Mountain [SRT16, AKvO^{+}10, BA14b, BRPK14, GANB11, MVC11, OT15, PGR^{+}16, PBG^{+}17, PB14, RRG^{+}15, TOHT14, TNTH11, CZC^{+}15, MJK^{+}17]. mountainous [Bre12]. Mountains [FLX^{+}14, GFW^{+}17, KPT11, KMP14, LGDW12, KAF^{+}14, CL12b, FSS^{+}17, GB15, HOŠ^{+}12]. mouse [LSGT13, LBS^{+}14]. mouth [DVL11]. move [TL15]. Movement [BPWB14, GGTP^{+}17, KRL11, AHS^{+}16, ADF13, BEZ19, BMSLC16, BV15, CH14, DFL^{+}16, DDBK15, EZ14, GPZB18, GKKW10, HB13, HM17, KNU^{+}15, KK12a, LGMLD19, LB15, MBG17, MLS^{+}11, MHY^{+}15, OC15, PHP15, PPBD13, RR10, Str15a, Str15b, VF16, WR13a, WR17, WSC^{+}19, ZDL^{+}17, ZHV14]. movements [ALKB10, APWT15, BLH10, BKP10, Bar11, BFNH14, CLD^{+}17, FBH^{+}14, FS18, HCM^{+}18, LAK^{+}18, RV13, TDF^{+}19, VSPHS15]. Moving [BL15, MRWN19, WK11]. mowing [LSA^{+}19]. mown [MBA^{+}19]. MPA [CGEdM12]. MRIO [WFSH15]. Mu [HYL^{+}17]. much [DFH^{+}14, ESP^{+}17, MS14]. mud [CSN^{+}19]. Multi [BRHD15, CN11, CTP^{+}17, CL12b, DSA^{+}18, HZZ^{+}18, HLK^{+}12, KHW^{+}17, SGDP19, VFFP16, ABB^{+}14, AMR16, AH17, BKN21, BMMC21, BGGL16, BGMJ^{+}12, BHPF14, CRB16, CF17, CY13, CSFR15, DTS^{+}12, DVL11, EBA^{+}19, FHAG12, FGPR10, FD19, GSQ17, GLC^{+}18, GBL^{+}13, GOSH^{+}12, HZSH17, HMS^{+}11, HCP10, HPO^{+}15, ISG13, JZG10, KGM12, KK15a, KH15, LVB^{+}15, LM13, LSJG17, LYC^{+}15a, LP17, LCDJ^{+}12, MBA^{+}19, NLM^{+}19, NNP19, PK14, PWK^{+}17, RKC^{+}15, SHX^{+}12, SLY^{+}17, Tat16a, TAC^{+}14, TOQW11, TNHB12, WFS^{+}19, WK17, WKH19, Yan11, YLF^{+}15, YHJ^{+}17, ZWRK13, ZQC15, ZZJ^{+}16, ZLZ^{+}17a, ZZT13]. multi-actor [GBL^{+}13]. Multi-agent [ZZJ^{+}16, BMMC21, DVL11, GLC^{+}18, HZSH17, TOQW11]. multi-basin [ZWRK13]. Multi-criteria [VFFP16, SHX^{+}12]. multi-criterion [ZZT13]. multi-disciplinary [LP17]. multi-field [DTS^{+}12]. multi-fleet [ISG13]. multi-grain [LVB^{+}15]. multi-layered [AMR16]. Multi-level [HLK^{+}12, FD19, KH15, LCDJ^{+}12, PWK^{+}17]. Multi-method [BRHD15, AH17, BHPF14, LP17]. multi-metric [SLY^{+}17]. Multi-model [CTP^{+}17, YHJ^{+}17]. Multi-objective [CN11, DSA^{+}18, CRB16, EBA^{+}19, GOSH^{+}12, NNP19, Tat16a, WK17, WKH19, Yan11, ZZJ^{+}16]. multi-purpose [KK15a]. multi-regional [ZQC15]. Multi-scale [HZZ^{+}18, SGDP19, CY13, FHAG12, HCP10, NLM^{+}19, YLF^{+}15]. multi-scaled [WFS^{+}19]. multi-sector [KGM12]. Multi-site [KHW^{+}17, HMS^{+}11]. multi-sources [ZLZ^{+}17a]. multi-species [ABB^{+}14, BKN21, CF17, CSFR15, FGPR10, GSQ17, ISG13, JZG10, LM13, LYC^{+}15a, MBA^{+}19, RKC^{+}15]. Multi-taxa [CL12b]. multi-temporal [LSJG17, PK14, TAC^{+}14]. multi-trophic [TNHB12]. multi-type [BGMJ^{+}12]. multi-year [BGGL16]. multiannual [vDHH^{+}13]. multicriteria [HFU11, ZH11]. multilayer [Fis21, MMMCAHM21]. Multimedia [XQZ^{+}13]. multimodal [PP19]. multinational [HSM^{+}12]. Multiple [dSCdA18, PW10a, AHS^{+}17, BWH13, BG19, BR16, BA17, BPC^{+}17, CZR^{+}14, DH18, Fai10, HL19, IT17, JBW^{+}17, KDK^{+}10, KR16, LZZP16, MAC14, MTJF11b, MSHC13, MMPG16, POS^{+}19, PGATS13, PÓF^{+}18, SFM^{+}18, SG11b, SYY^{+}19, WBC^{+}12, WCS^{+}14, ZBAY14, ZHC^{+}16]. multiple-expert [MSHC13]. Multiplex [PCRP12]. multiplicative [BVC^{+}18]. multiscale [ChHC12, DHC14, MTJF11a]. Multisite [AADA14]. Multisite-multivariable [AADA14]. multispecies [BFPR16, CRD^{+}16, For18a, GDC12, GHS^{+}16b, HW12, MFD^{+}15, NB11, PF12, TTCC^{+}19]. multispectral [KHS^{+}19, PSKO16]. multistage [BP19]. multistate [BS10c, TB12b]. multitrophic [RSDP^{+}12]. multivariable [AADA14]. multivariate [LGY10, WLC^{+}19]. mungbean [BKM^{+}18]. MungGro [BKM^{+}18]. municipal [BWRA16, ZHCX15]. municipalities [ZQC15]. municipality [MSS^{+}19, dHR18]. muntjak [SK11]. Murgia [LDM15a]. Murrumbidgee [WRSL11]. Muscidae [LIM^{+}16]. mussel [DWNN10, DMB^{+}18, FWRZ10, HSvO^{+}18, LSR16, dÁSMRT^{+}19]. Mussels [CPS15, KVP16, Lin15b, PCRP12]. muta [ZPNM13]. mutations [HDL^{+}16]. Mutual [JJS10, dSCdA18, F.11]. mutualism [CN18, IKH^{+}11, JBE^{+}15, MCA18]. mutualisms [WSE10]. mutualistic [BBHP13, LG12, MCA18]. mycorrhizal [JBE^{+}15]. mycorrhizas [VH12]. myiasis [LIM^{+}16]. mykiss [KGM12, MSB^{+}17]. Myrna [Wei19]. Myzus [GOSH^{+}12].

N [DVD^{+}14, DAP18, HMS^{+}11, HBA16a, LWGH13, LYJ^{+}15, LSM19, MRP^{+}11, MEC14, SBC^{+}11, TGM^{+}21, WB14, ZR19]. N-labelling [TGM^{+}21]. N-mixture [DAP18, ZR19]. N/P [WB14]. N14C [TRE^{+}12]. NACP [HPW^{+}12]. Nakdong [BS18, JJKJ11, KHJ^{+}19]. NANI [HSM^{+}12]. NANI/NAPI [HSM^{+}12]. NAPI [HSM^{+}12]. Naples [ZW10]. Narragansett [BN10, BLCPB11]. narrative [AH17]. Nassella [JBB^{+}11]. natal [DCCS17]. National [BFKS16, CBGK15, HSK17, KPT11, LDM15a, MNHW^{+}13, MAM^{+}10, NGR^{+}11, SHM^{+}14, WALCY12, ZPZ^{+}15, OLG^{+}13, YWH15]. nationwide [JVY19]. native [BGC10, CPS15, GQRC^{+}17, MAC^{+}12, PK16, PGZ^{+}19, PTL14, TBA19, XCG^{+}16]. Natural [FBDR17, Kat11, PPB^{+}18, An12, ALC12, ASP^{+}15, BWH13, BPC^{+}17, BPU11, BPD^{+}19, CPMP13, CC11, CBP^{+}17, DYU^{+}12, DWP11, FHG12, GHP12, HMP19a, HMS^{+}10, IBD^{+}19, Ken17, KSZW16, KCG18, LHZS10, LCDJ^{+}12, MNEG11, PVM12, PVM13, Pen11, PBD^{+}17, Pra12, SFF^{+}11, TLLX19, TDC^{+}11, TS11, VPB^{+}17, WALCY12, Wil11a]. natural-disturbance [HMS^{+}10]. naturalized [GV15]. Nature [JJS10, LLL^{+}10, LKG^{+}17, LCS11, LYC^{+}15b, WALCY12, BvHR^{+}15, EMH10, FLA19, LYC^{+}15a, LLL17, LP17, LLL19, Par10, Ula10, VLZ^{+}19, WGTH16, WFS^{+}19, Pat14a]. nature-based [VLZ^{+}19]. navigating [PP19]. NC [DLH^{+}14, JMB^{+}11]. NDVI [DBR12, WYS12]. NDVI-based [WYS12]. near [WLZC11]. near-shore [WLZC11]. nearshore [BPC^{+}17, CBP^{+}17, YSH^{+}10, YKCN10]. necessity [NU11]. nectar [BGK^{+}16]. nectar/pollen [BGK^{+}16]. need [GHD^{+}19, GYL^{+}13, NLMdL10, SBM^{+}12, VMP^{+}19]. needed [Bus17, ESP^{+}17]. negative [TTCC^{+}19, WKZ^{+}19, YZ18]. neglected [RDV^{+}17, STJ15]. neglecting [NBSGW19]. Negro [IGR^{+}18]. neighborhood [GDBB19, HB14]. neighborhoods [Hal11]. neighboring [FHG12]. neighbourhood [RAM10]. neighbouring [LLH^{+}13]. nekton [LdMS^{+}16]. nelsoni [RR10]. nematode [RWvDM15]. nematodes [SU15]. Neolithic [ABGB15]. Neoseiulus [dOMZ13]. Neotropical [BFNH14, GAV^{+}11, HCP10, RV15]. Neotropics [GdSHA10]. NEP [MGS16]. nest [DWWS15, GKF^{+}14, JJKJ11, Rin14]. nested [MCPB13, MCPB15, RDD^{+}18]. nesting [DBP16, FMW^{+}19, HVL^{+}19, KK13]. Net [OLC^{+}17, Bai18, BLY11, Bus17, CPMP13, CSS^{+}19, CLWH15, FCG^{+}19, HSM^{+}12, HNWT10, yHAB12, HSD^{+}15, JMSM12, KHW^{+}17, LCL^{+}13, LLT^{+}14, Pie10b, Pie10a, RHvdTS14, SNS^{+}12, SH18, SM14, SLZ^{+}17, WGA^{+}11, WGA^{+}12, XM13]. Network [AD18, BFS11, BFW14, BME14, CBGK15, CR11, CR12b, JMdW^{+}10, KCF18, LYC^{+}15b, OCB^{+}13, SSF14, SDS^{+}21, TLL^{+}17, ABB^{+}10, AAW^{+}15, Ano12s, BA14a, BGC10, BT16, BS10a, BSMA18, BK13, CD17, CXW19, DPPI14, DSD13, DSW^{+}17, FSB13, FBH^{+}14, FLB^{+}15, GKF^{+}14, GDDFT16, GLC^{+}15, HKL^{+}11, HB14, JBCÑ11, JGX13, KB10, KM12a, KHJ^{+}19, LLL^{+}10, LY11, LHL18, LWTZ18, Lin10, Lin16, LYCZ11, LYC^{+}15a, MY11, MCW13, MWY^{+}18, Mar12, Mar17, MFS18, Mat13, MR17, MJC^{+}11, MHC16, MSFR15, NWAR11, NWAR12, OBC^{+}21, PMS^{+}17, Pat15, Pat14b, Pau14, PICR13, RBM^{+}17, RRR^{+}10, RAR15, RRvdG18, RMLTCQ^{+}19, SSV^{+}12, SM10, SFB^{+}15, SYY^{+}19, SSW^{+}18, TRD^{+}15, TLZ17, TWP^{+}14, URR^{+}17, VC17a, VPB^{+}12, WPG15, WZD^{+}14, WPB14, WC13c, XZCL16, YJW15, YM21, ZRCNLI^{+}15, ZYFL10, ZZF15, ZLFZ16, ZZ17, ZLZ17b, ZMH19, vSGR16, KCF17, LLKY18]. Network-agent [AD18]. network-environ-analysis [Ano12s, MCW13]. Networks [CPE13, MDH15, Pec12, ASL13, AS14, BYR18, BFJ^{+}11, BFR17, BS10b, BNP10, FRP18, GHD^{+}19, GO15a, GMK^{+}19, HPJ15, JN14, JN15a, KSSA13, KvOP10, mLcLJ15, mLcLJ16, LS13, LPX^{+}12, LJCG15, MJLC14, NNP19, NF11, OOFVBC16, PW10a, Pat12, PAEE19, PGATS13, PSWW11, RHW17, SB11, SZNL19, SCBH18, VC17b, WWC^{+}18, WLR^{+}11, YZ16, YZ18, YXZ^{+}21, ZB15, dAEPE16, vSGR16]. Neural [CPE13, RRR^{+}10, ASL13, BT16, GO15a, GDDFT16, JBCÑ11, JGX13, KHJ^{+}19, LPX^{+}12, MFS18, NNP19, PSWW11, SSW^{+}18, TLZ17, WLR^{+}11, WZD^{+}14, ZZ17]. Neuse [WPB14]. Neusiedl [HRSKW15]. Neutral [Con11, WYPF15, BRMW18, Che15b, Che16, KWH14, LS18, VC17b, WR12, WRMB18, XZZ^{+}13, Con12]. Nevada [FMF12]. Nevados [ZPZ^{+}15]. newly [PTL14]. newt [MJK^{+}17]. Newton [Kat11]. Next [GB16b, GB16d, JDT^{+}16, SLW17, SJB^{+}16]. Next-generation [GB16b, GB16d, SLW17]. nexus [CL15a, FKWZ19, GLC^{+}18, Ju19, LZL^{+}18, NLM^{+}19, XYTS19]. Nguruman [LDM15b]. Niche [CPE13, ACE^{+}13, BBJV^{+}11, IHSM12, BOGA14, BB18, CGV^{+}15, CH12, DBR12, DGKSM13, FP17, FHAG12, GWR11, GLSA^{+}11, qHLlZ^{+}16, HGAT17, JSCS19, JSdSdL^{+}18, KFWM19, LJ15a, LPT^{+}18, MWK19, MAFG^{+}19, MMMCAHM21, OCD^{+}13, PNT^{+}19, PKW^{+}10, RHS^{+}15, RHD^{+}16, RVB^{+}10, SBM^{+}12, SA13, Sil11, SGM^{+}14, TTL^{+}18, HKVD16]. niche-based [KFWM19, MWK19]. NicheLim [HKVD16]. niches [MDB19]. Niger [AH17]. night [BFJZ^{+}11]. night-herons [BFJZ^{+}11]. nighttime [WLC^{+}19]. nine [OLC^{+}17]. nine-banded [OLC^{+}17]. nitrate [BA14a, HGPF15, MCH15, VRL14, YDD^{+}13]. nitrification [LNB17]. nitrite [BH17]. Nitrogen [Gra14, SCH^{+}17, WEJ11, BKN14, BRKP16, CS18, DZW^{+}11, Fel12, GZS^{+}18, GBG^{+}15, mGZK^{+}10, GGVB11, GHS^{+}10, GS10, GRD^{+}14, GBR13b, HB14, Hoa11, HSM^{+}12, HBA16a, HBA16b, HAA^{+}21, JWX18, KH11, KJMA11, KPS19, LYJ^{+}15, LHB^{+}18, LDC^{+}16, LBC^{+}15, MRG13a, MTL^{+}10, MKSL^{+}17, MJC^{+}11, NBK^{+}11, NW18, PP16, PMB^{+}17, PBP^{+}15, SKMB11, SYL^{+}13, SEHD18, SSF14, TGM^{+}21, TBA19, TRE^{+}12, VRL14, WPB14, YDD^{+}13, ZTJ^{+}10, ZLFZ16, ZLZ17c]. nitrogen-to-carbon [CS18]. Nitrous [HJS^{+}16, GNB^{+}14, LYJ^{+}15]. nitzschia [GDDFT16]. NME [BPC^{+}17]. No [PAB^{+}13b, Bus17, LYJ^{+}15]. Noctuidae [FGH^{+}10]. nocturnal [BLFS18]. node [HB14, Pat15]. NOE [GNB^{+}14]. Noise [ZW12, LS13, NNST^{+}14]. Non [GdTCFG10, LJAD17, Ned15, VRL14, ACKS13, AK17, AL16, BLY11, BM16, BGC10, BPK^{+}16, CCCP17, CBC^{+}16, CCC12, CPS15, DH18, GLSA^{+}11, JTB^{+}19, KWH14, KCKdOW12, LPT^{+}18, MH13, MPP^{+}15, MHRB19, PTL14, RWS^{+}16, RSP^{+}17, RHS^{+}16, RDR^{+}17, RE10, SJTS13, TLP17, TGB^{+}17, WC19, XCG^{+}16, YDZ^{+}21, ZH11, ZYKS15, dS18]. non-adaptive [LPT^{+}18]. non-additivity [WC19]. non-autonomous [SJTS13]. non-buoyant [BM16]. non-cooperative [TLP17]. non-domesticated [GLSA^{+}11]. Non-equilibrium [LJAD17]. non-focal [BPK^{+}16]. non-hierarchical [RE10]. non-influential [CCC12]. non-invasive [MPP^{+}15]. non-isolated [dS18]. non-Kolmogorovian [ACKS13]. non-limiting [AL16, TGB^{+}17]. Non-linear [GdTCFG10, VRL14, AK17, CCCP17, DH18]. non-monotonicity [ZYKS15]. non-native [BGC10, CPS15, PTL14, XCG^{+}16]. non-neutral [KWH14]. non-point [YDZ^{+}21, ZH11]. non-scientist [CBC^{+}16]. non-spatial [RWS^{+}16]. non-stationary [RHS^{+}16, RDR^{+}17]. non-stratified [JTB^{+}19]. non-symmetric [dS18]. non-target [MHRB19, RSP^{+}17]. non-timber [KCKdOW12]. Non-traditional [Ned15]. non-uniform [MH13]. nonexistent [Ula14]. Nonlinear [YYL13, EGA12, Fis21, Mon13, PBT^{+}11, PBD19, TŠAD12]. nonparametric [SML12]. nonpoint [GPD^{+}13, NKSA13, SRXP14]. nonrandom [LLL19]. norms [CLA^{+}12]. North [BH13a, CW11, CR12b, GSJ^{+}10, GCL^{+}19, LAB^{+}10, MSCH14a, PVM13, SW19, TTL^{+}18, ZLLZ13, MMGK15, QCLW18, RR10, SHM^{+}17, ZDG^{+}13, BH13b, BB18, CAL13, CWA^{+}15, CR11, DUASLB^{+}12, DYU^{+}12, DSW^{+}17, GCvDM12, GCM^{+}13, HKlL14, HBA16b, HPW^{+}12, JBB^{+}11, LN11, LN16, LAR^{+}16, LAD11, LCY^{+}13, MML^{+}11, MRD^{+}12, MSCH14b, PVM12, PGZ^{+}19, QDH^{+}10, RMHE15, SELB10, SBC^{+}11, SDY^{+}15, SGH16, SKM^{+}16, TCW11, Tho10a, WYHC15, YLY^{+}18]. Northeast [FLH17, GC11, MSM^{+}16, NGL13, SLZ^{+}18, TZS^{+}14]. northeastern [DLL18, PGV10, Gli14, JSB^{+}14, SCG^{+}16]. Northern [FRM^{+}19, KCF18, MWML13b, NGGFM19, XWW^{+}11, XQZ^{+}13, ATCJ^{+}19, BHD16, BLLJ17, BCO17, CL15b, CQS15, CR11, CR12b, FGH^{+}10, HKM^{+}18, JR16, JA15b, KHW^{+}16, KGYG15, LLC^{+}17, LMGW16, MRD^{+}12, MWML13a, MP17, MMVAHMC16, PDAL14, ROSC15, SLW17, TDF^{+}19, VAK^{+}11, WGL^{+}12, WLH^{+}10, ZKSF17, dMSL^{+}16, CL12b, GYL^{+}15, HT16, KBS^{+}15, KCF17, LDP14, MPKC12, MTPC13, MM19, RBGB16, WMW^{+}12]. Northwest [JJYM19, WZG15, CW11, GQNMR18, PGGMMA19, SBF15, WHRR15, ZLW^{+}15]. northwestern [GdTCFG10]. norvegica [JR16]. Norway [CAMNN16, mGZK^{+}10, HOKH18, OJ10, TNA^{+}17]. Norwegian [BSH17, SDKE18]. nose [CESC19, RTE^{+}15]. Note [Ano19-49, PdM11, Ano10-27, Ano12-38, Pér18]. Notes [Fat10, PF18]. NOTG [SMH16]. Novel [MTGD17, ZLY16, BHD16, DBSA16, DFW^{+}18, KHW^{+}16, LMT13, RSCS19, WR13a, WLR^{+}11, dBGS^{+}14]. November [Ano10t, Ano10u, Ano12-37, Ano13x, Ano14i, Ano14l, Ano15g, Ano16l, Ano16k, Ano17-32, Ano17v, Ano18-31, Ano18-28, Ano19b, Ano19f, Ano21c]. NPP [SLSS15]. NPZ [TBM^{+}14]. NPZD [HS13c, HS13b, PMTKC17]. NSF [Bec10]. nucifera [MdMS^{+}19]. Nuclear [GSF^{+}16]. nuisance [IN16]. Null [MSV^{+}11, FM11, GU11, UG12]. number [CWL^{+}16, GCW^{+}16, Ken19, Log16, NMS10, OLC^{+}17, PDG^{+}14, SM12, WGSR19]. Numerical [BKN14, DFL^{+}16, IKK^{+}12, LZW^{+}17, yLWhMhG11, PLS^{+}10, TLW^{+}14, FS14a, FRM^{+}19, GJÁ^{+}10, GL10, GS19a, GC14, GBR13b, LZYH19, MH13, PES19, SFMA15, Sin17, XSN^{+}18, ZLFI11, ZYZ17]. nursery [BJK^{+}16, FB13, FPG11, FPWG14, OAPP12, SGBG19]. nursery-pollinator [SGBG19]. nutcrackers [MSM^{+}17]. Nutrient [JD10, WFM16, AD12, AL16, BGPJ10, CGM^{+}14, HKVR13, HSM^{+}12, HOŠ^{+}12, KWS^{+}13, KW18, Kei12, KZR^{+}13, LCY^{+}13, LLZ^{+}15, LWC^{+}18, LW14, LLFL12, LVV^{+}15, MMM^{+}16, Mul11, NPH^{+}17, PFA^{+}16, RPSA12, REB^{+}16, RP11, SQWS19, TBM^{+}14, WMK^{+}13, XYY^{+}16, ZLFI11, ZWW^{+}18, ZML^{+}12]. nutrient-cycling [CGM^{+}14]. nutrient-runoff [KW18]. nutrients [DBVS19, DMZ^{+}19, ER10, FS14a, LLZ^{+}15, LQW^{+}19, Lin15b, MR17, Pea15, YZG16]. nutrients-algae [YZG16]. nutrients-phytoplankton [LQW^{+}19]. nutrition [GBP^{+}16, YZA^{+}18]. nutritional [AHS14, WSE10]. NW [MJB^{+}11a, BDV^{+}19, CSS^{+}16, DBVS19, FAC^{+}19b, GDDFT16, PPB^{+}18]. Nyctaginaceae [RXSK14].

O [HMS^{+}11, KMM^{+}13, LKLK15, LYJ^{+}15, LSM19]. O157 [SFvB10]. oak [ABC^{+}17, KSS11, LL18a]. oak-hickory [LL18a]. oasis [SYY^{+}19]. OBEC [RDD^{+}18]. OBIS [CMV^{+}16]. Object [wTA12, BFR17, KK13, SH18, ZPNM13]. object-based [ZPNM13]. Object-oriented [wTA12, BFR17, KK13]. objective [CRB16, CN11, DSA^{+}18, EBA^{+}19, GOSH^{+}12, HPO^{+}15, NNP19, Tat16a, TA11, Wil12, WK17, WKH19, Yan11, ZZJ^{+}16]. obligate [CN18]. observation [AGJ^{+}12, LH15, PP10, ZKSF17]. observational [CGG^{+}18, HJ10b]. Observations [UKI^{+}11, CBA16, LS18, MPT^{+}11, PSB^{+}17, SHS11, WBC^{+}12, WP14]. Observatory [SEHD18]. observed [GLT10, TL10]. observer [WB10]. obstacles [RRBI14]. obstruct [VVA^{+}15]. obtained [VHM^{+}17]. obtusa [Oga11]. occasion [GB16b]. occupancy [Hui11, LCMG18]. occupation [YMCV21]. occur [HKM^{+}18]. Occurrence [dSGGBW12, ADF^{+}17, CPR10, CWJ^{+}13, DYMYGO19a, DYMYGO19b, DSW^{+}17, FM10, FM11, JDZC14, JSCS19, MMGK15, YAGMM^{+}14]. occurrences [BGA17, SAKB^{+}10]. occurring [OGOWW16]. ocean [ABR12b, ABR12c, BSH17, DVD^{+}14, FLH17, HKM^{+}18, KTK^{+}15, PGV10, PPDF14, RWP11, THC18, BSH^{+}10, DFM12a, DFM12b, FE14b, GQNMR18, GC11, HKlL14, PMTKC17, RMTM^{+}16, WFA10, WHRR15, WCG^{+}11]. oceanic [Flo18, WGD^{+}17, ZJL^{+}14]. October [Ano10r, Ano10s, Ano11-33, Ano12-32, Ano13-27, Ano13z, Ano14g, Ano15p, Ano15h, Ano16n, Ano16j, Ano17-31, Ano17t, Ano18-30, Ano18z, Ano19c, Ano19g, Ano21c, LMY^{+}11, LYP^{+}11, LPA^{+}11]. ODD [GBD^{+}10]. Odocoileus [WTG^{+}16]. Odontesthes [BCC^{+}14]. Odum [GDC^{+}19, JZCL14, LTL15, Ten15b, Til14]. OECD [Hoa11]. off [BLFS18, BBL11a, CRQW18, DZW^{+}11, DP19, GJB^{+}11, NCK10, RHD^{+}13, ZFZ^{+}14]. off-shore [BLFS18]. offs [BLD10b, GK14, HV15, LARZN13, LCPB^{+}10, MWML13b, SDT10, dMLM^{+}17]. offset [WKZ^{+}19]. offsets [Bus17, TBJ^{+}15]. offshore [DMB^{+}18, RPC^{+}17, SS11]. Ohio [MR17]. oil [BPU11, GPA^{+}13, KWGG19, MH11]. Oilseed [GDT^{+}14]. Ojo [RIMPBT19]. Okavango [LMCA^{+}12]. Okeechobee [Jam16]. old [AMR16, BR16]. old-growth [AMR16]. oleae [GCTS15]. Oléron [PLR^{+}13]. oligohaline [CN10a]. oligotrophic [CT11, HPRNP12, HMA10b, JA15b, SPM^{+}18, VJR^{+}16]. olive [GCTS15, MCBM12]. OLS [WLC^{+}19]. ombrotrophic [WRM^{+}11]. Oncorhynchus [KGM12, MSB^{+}17]. one [BKB10, FRM^{+}19, HQF^{+}17, KCKdOW12, MPF^{+}14, OOFVBC16, PDAL14, PBK14, SP10, WPG15, WB14]. one- [HQF^{+}17]. one-dimensional [BKB10, FRM^{+}19, PDAL14, PBK14, WPG15]. Oneida [HSR^{+}15]. ones [NYH15, ZJB19]. ongoing [CVM^{+}18]. Only [EPM^{+}14, FN15, GdSHA10, YHW^{+}12]. onset [KKL^{+}12]. Ontario [KZR^{+}13, ZKR^{+}13, SP10, SS11]. Ontic [NU11]. ontogenetic [SMS^{+}13, SIS^{+}14, SFR^{+}17]. ontogeny [TCSG15]. Open [GLS10a, BAP18, BK13, CBB19, DLH^{+}14, LST15, MSGG14, MH13, MLMFN^{+}15, SRP18, SFL^{+}19]. Open- [GLS10a]. open-source [MLMFN^{+}15]. Opening [THO10b]. openness [NU11]. operating [YY13]. operation [CCL^{+}13, DFW^{+}18, WhFxGZ16, YY11]. operational [IGR^{+}18]. operationalisation [CGL15]. Operationalizing [KK15a]. operations [BWRA16, TSN15]. operators [BO14]. opimus [RMP^{+}18]. opinion [Ano17a, Cha12, TB11, TB12a]. opportunist [DNFG18, MLK17]. opportunistic [CGAB17]. Optical [GANB11]. Optimal [BBRJ15, BJJ12, BMP^{+}19, CMD12, dSCdA19, CRVS13, DML15, KOBK13, NSN17, PR13, TKP14, BFPR16, BM10, CHTC11, ETG^{+}18, GS19a, HFMM^{+}12, LDB14, MAFG^{+}19, MOMC17, PP10, PRP17, RKL^{+}15, SRXP14, SA13, War10, ZFZ^{+}14, vDHH^{+}14]. Optimality [CS18]. Optimality-based [CS18]. optimally [RP10]. optimisation [CN11, EBA^{+}19]. optimising [RSDP^{+}12]. Optimization [HCL^{+}14, JBCÑ11, MCWG21, AK17, BHD16, BSC^{+}15a, CZR^{+}14, DAW^{+}10, FDM^{+}11, GOSH^{+}12, HCK18, LOLA13, LHR^{+}15, NNP19, OFH^{+}19, PHP^{+}14, TŠAD12, Tat16a, TLZ17, TA11, VQS^{+}18, WSW^{+}13, WK17, WKH19, Yan11, YYQ11, YZL12, YTT^{+}11, ZZJ^{+}16, ZF11]. optimize [GCIB10, LBM16]. optimized [CGY11, GDW12, MEC14, dWcWbZ15]. Optimizing [HBR19, MdMS^{+}19, XLR11, YRGH16, AD16, HG17, NMS10]. Options [LSM19, AHM^{+}16b, ABC^{+}17, FSM^{+}19a, FtHvdB^{+}14, For11, GH13, GMP13, KBMM15, NNP19, PMS^{+}17, RvS16]. Opuntia [MRN12]. oral [LPA^{+}11]. Orbetello [GMLRF10]. orchard [KL10, TWL11]. orchards [GCTS15]. Order [dSCdA19, JA15a, Nie16, WHC^{+}13, ZCY^{+}19]. Ore [GB15]. Oregon [MSB^{+}17, RWP11]. Organic [PLFO^{+}12, BP15, BP16, BP17, BP19, BKMK10, BBH^{+}11, CKS^{+}17, CSS^{+}17, GE17, HSK17, HN14, HJS^{+}16, JNC^{+}11, KH11, KCB^{+}17, LLT^{+}17, LKP13, MLB^{+}16, MDF^{+}18, MJF19, Nei11, OSS13, OLG^{+}13, RDD^{+}18, SJS^{+}13, SDD^{+}13, SBRD10, SCG^{+}16, TDC^{+}11, WSL17, WHC^{+}13, ZJM^{+}13, ZJL^{+}14]. organisation [CTC^{+}18, SLZ^{+}13]. organising [DGKSM13, PLB16]. organism [Pat16a, WBHM11]. organisms [DCMK11, GPP^{+}19, JLN10, Jør15a, Jør15b, MC11, Mul11, RKS^{+}10]. organization [GDC^{+}19, LCZ11, MJB^{+}11b, PSS16]. organized [DBH^{+}17, LMY^{+}11, LYP^{+}11]. Organizing [Jør14a, LLC11, PGKG12, RPPW16, SSM^{+}13]. organogenesis [BDL^{+}13]. orientale [GCC^{+}10]. oriented [AAFH15, BFR17, BFS11, CRA^{+}11, CAMNN16, GPCG15, JGF14, JCKG15, KK13, MJL^{+}13, MJL^{+}14, RJ11, SHT^{+}21, SMH^{+}12, wTA12, THOA10]. origin [HMS^{+}10, XCG^{+}16]. Origins [ER10, HHC15]. oriii [Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16-37, Ano16-38, Ano16-39, Ano16-40, Ano16-41, Ano16-42, Ano16-43]. oryzae [YWA^{+}17]. oscillating [RDD^{+}17]. oscillations [FS14b, RH17]. oscillatory [PF12]. OSMOSE [GHS^{+}16b]. osmotic [BCB^{+}11]. ostertagi [CC10, CCSF11, CCF13a]. Ostertagia [CC10, CCSF11, CCF13a]. ostralegus [SGA^{+}16]. other [DLL^{+}19, DVF^{+}19, MR17, PMH17]. otoliths [KHW^{+}16]. otter [TSN17, TAC^{+}14, VPC^{+}14]. our [Ano17a, FHAG12, RDH^{+}13, WB16]. out-crossing [DTS^{+}12]. outbreak [CCL14, PSBH16]. outbreaks [FH11, OFGH^{+}19, RPA12, WFS^{+}19]. outbursts [JSKK15]. outcome [MF12]. outcomes [MPR13, VPB^{+}12, ZAW^{+}18]. outflow [GBW^{+}15]. outflows [GSF^{+}16]. outline [LTL^{+}19]. outlook [JBH^{+}14]. Output [SCH^{+}17, BB10, BCP16, BFS11, CC11, Cho13, HCM^{+}15, KPS17, LHL18, Ske13, XYTS19, ZQC15]. outputs [YDD^{+}13]. Outside-host [MLK17]. outwelling [DVD^{+}14]. ovata [AALP18]. over-harvested [IKL14]. Over-reporting [GU11, FM10]. overabundant [WR13b]. overall [KL12]. Overcoming [BMSLC16, HVL^{+}19]. overestimate [Log19b]. overestimated [dSCdA19]. overexploitation [PBEL17]. overfishing [ACE^{+}17, BPB14]. overfitting [MSV^{+}11, VHM^{+}17]. overhead [SSQ^{+}10]. overlap [mLcLJ15, mLcLJ16]. overlap-based [mLcLJ15, mLcLJ16]. Overlooking [KGS11, PC16]. overstory [XKI^{+}11]. overview [CVG^{+}15, DPV^{+}16, EEFE^{+}11, GRRU15]. overwhelm [PW10b]. overwintering [BKN21, HVCP16]. overyielding [FDHL17]. owing [GAG^{+}12]. owl [dÁGP14, EGMW12]. owls [TPG14]. oxbow [AdMC^{+}13]. oxic [GS10]. oxidation [VRC17]. Oxide [HJS^{+}16, GNB^{+}14, LYJ^{+}15]. Oxygen [SPM^{+}18, TM14, ASL13, BA14a, BMR^{+}18, CGdCCPCPM14, GS10, MH13, MGW^{+}13, NFU19, OdlFDP15, RRR^{+}10, VMSM10]. Oxythermal [JFS^{+}12]. oyster [BGD^{+}15, EH15, EH19, KTGH19, KPLS15, LLC^{+}17, WCL^{+}17]. oystercatchers [SGA^{+}16]. Ozarks [SHX^{+}12].

P [Och10, Wei19, AR11, JS12, LWGH13, WB14]. Pacific [GC11, HKlL14, PVM13, CF17, FE14b, FAC^{+}19a, HlLB12, KSAS12, MGCL17, OCB^{+}13, OBC^{+}21, PVM12, PME^{+}16, ROSC15, SBF15, WFA10, ZLD^{+}19]. packs [CWL^{+}16]. paddies [COPP13]. paddy [BKN14, SNS^{+}12]. Padilla [KR10a]. Page [Ano13-41, Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano16v, Ano16w, Ano16x, Ano16y, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16-37, Ano16-38, Ano16-39, Ano16-40, Ano16-41, Ano16-42, Ano16-43, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano17k, Ano17l, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano18x, Ano21d]. Page/Cover [Ano17m, Ano17n, Ano17o, Ano17p, Ano17q]. Pages [Ano12-37, Ano12-28, Ano12-29, Ano12-30, Ano12-36, Ano12y, Ano12-33, Ano12-34, Ano12v, Ano12z, Ano12x, Ano12-31, Ano12u, Ano12-32, Ano12-35, Ano12-27, Ano13-39, Ano13-38, Ano13-32, Ano13-36, Ano13-33, Ano13-31, Ano13-29, Ano13y, Ano13-34, Ano13-35, Ano13-40, Ano13w, Ano13-37, Ano13v, Ano13x, Ano13-27, Ano13z, Ano13-30, Ano13-28, Ano14u, Ano14e, Ano14c, Ano14r, Ano14o, Ano14f, Ano14j, Ano14s, Ano14k, Ano14m, Ano14p, Ano14q, Ano14n, Ano14b, Ano14t, Ano14d, Ano14i, Ano14l, Ano14g, Ano14h, Ano15q, Ano15p, Ano15a, Ano15j, Ano15b, Ano15d, Ano15l, Ano15m, Ano15k, Ano15f, Ano15e, Ano15c, Ano15n, Ano15g, Ano15h, Ano15o, Ano15i, Ano16t, Ano16a, Ano16u, Ano16q, Ano16m, Ano16e, Ano16h, Ano16o, Ano16d, Ano16r, Ano16b, Ano16f, Ano16p, Ano16c, Ano16i, Ano16l]. Pages [Ano16k, Ano16n, Ano16j, Ano16g, Ano16s, Ano17-38, Ano17r, Ano17-35, Ano17-36, Ano17x, Ano17-37, Ano17-29, Ano17s, Ano17z, Ano17-40, Ano17-30, Ano17u, Ano17-28, Ano17-39, Ano17w, Ano17y, Ano17-27, Ano17-32, Ano17v, Ano17-31, Ano17t, Ano17-34, Ano17-33, Ano18-41, Ano18-38, Ano18-46, Ano18-47, Ano18-34, Ano18-48, Ano18y, Ano18-40, Ano18-42, Ano18-33, Ano18-39, Ano18-37, Ano18-44, Ano18-35, Ano18-27, Ano18-43, Ano18-36, Ano18-45, Ano18-31, Ano18-28, Ano18-30, Ano18z, Ano18-29, Ano18-32, Ano19-36, Ano19-47, Ano19-33, Ano19-37, Ano19-39, Ano19-45, Ano19-42, Ano19-40, Ano19-38, Ano19-34, Ano19-46, Ano19-35, Ano19-41, Ano19-48, Ano19-44, Ano19-43, Ano21b, Ano10g, Ano10o, Ano10p, Ano10v, Ano10w, Ano10x, Ano10i, Ano10k, Ano10l, Ano10m, Ano10n, Ano10y, Ano10z, Ano10h, Ano10j, Ano10t, Ano10u, Ano10r, Ano10s]. Pages [Ano10q, Ano11v, Ano11w, Ano11-30, Ano11-35, Ano11-34, Ano11-36, Ano11s, Ano11z, Ano11-27, Ano11-28, Ano11-29, Ano11t, Ano11u, Ano11x, Ano11y, Ano11-33, Ano11-31, Ano11-32, Ano12w, Ano21c]. Pahang [LHL^{+}17]. PAHs [XQZ^{+}13, ZCSL18]. pairs [MPF^{+}14]. paleobotanical [MAMNL^{+}15]. pallid [DHC16, WAGM17]. palm [SFL^{+}19]. Palma [MGGLBIBD13]. palumbus [OFK^{+}12]. palustris [LCML11]. Pamlico [JMB^{+}11]. Pampas [BPR^{+}11, BRM^{+}14]. panacea [DSBM^{+}18]. panda [FLX^{+}14, LLL^{+}10, YL17]. pandas [GFW^{+}17, LCS11]. Pantanal [AdMC^{+}13]. Panulirus [LMBL12]. Paper [RE10, Jør14a, NSIiT17, Bec10]. Paperback [Jør14b]. papers [BS15, BNP10, LYP^{+}11, LPA^{+}11]. Pará [DBK^{+}13]. Paraconsistent [BPT^{+}19]. paradigm [CL15a]. paradox [ALL11, F.11, HKW^{+}18, KK16, Loe12, WTWS19]. Parallel [TB11, TB12a]. paralytic [GBB^{+}11]. Paramecium [Ned15]. Parameter [BLLJ17, DFM12b, JLZ^{+}18, MMC11, TŠAD12, WC13b, Ash10, BA14a, BWA17, ČTT^{+}12, CZR^{+}14, CTMS10, DVT^{+}10, FHL^{+}13, GBP^{+}10, GLC^{+}15, HGPF15, HS13b, HGLZ13, KB10, LBM16, LHR^{+}15, MGSB11, QCLW18, QFG^{+}13, RMTM^{+}16, SLLW15, Wol11, YLL^{+}12]. parameter-sparse [GBP^{+}10]. Parameterisation [COBF13, CHPS17, CKS^{+}17, CSS^{+}17]. parameterised [SBM^{+}11]. Parameterising [CN10b]. Parameterization [GGVB11, BdCS11, GTC16, JGF14, KKL^{+}12, NMA^{+}19, PMBM11, PP12, SCG^{+}16, WBS13, BDV^{+}19, SMG^{+}12]. parameterizations [GGR^{+}13, THC18]. parameterize [SHT^{+}21]. parameterized [ADF13]. parameters [BLBB13, BRPK14, CCCP17, CDK^{+}17, CCC12, CDL^{+}11, For18b, GDC11, GCCB18, GPD^{+}13, HSK17, HCL^{+}14, HG17, JBB^{+}11, JFS^{+}12, LDG14, LB17, LXS^{+}16, MSO10, MRTW18, Ned15, OKS^{+}11, PBP^{+}15, PSSH11, RKL^{+}15, RCCN11, TQZ12, VG19, YCL^{+}14, ZBS14]. parametric [ST13, ZBMC14]. Parametrization [BDL^{+}14]. parametrizations [KWH14]. parametrize [GANB11]. parametrized [CC10, CCSF11]. paramo [AYASR^{+}17]. Paraná [RCCN11]. parasite [SU15, WGT12]. parasites [BGG^{+}16b, LJB12, LLFL12, RWvDM15, TV15, TNTH11]. Parasitic [CCF13a, SFHK12]. parasitoid [AD16, ASS14, BLAV18, CH12, CHZ18, DPvBvA11]. parasitoids [RAdA12, SFHK12, SGBG19, RCGC17]. parental [CWM12]. Pareto [LZB17]. Pareto-efficient [LZB17]. Paria [RDR^{+}13]. Park [CBGK15, KPT11, MNHW^{+}13, DLD^{+}11, ZZF15, BFKS16, LDM15a, ZPZ^{+}15]. parkinsoni [ZDL^{+}17]. Parks [MAM^{+}10]. parsimonious [RPGSCF16, War10]. parsing [UM18]. Part [FPWG14, PSKO16, BLD^{+}10a, BlMG^{+}11, DFM12a, DFM12b, FPG11, Pat14a, Pat16b]. Partial [CEWW18, IBD^{+}19, ACSR^{+}13, BHC^{+}15, TVF^{+}10]. participatory [EBA^{+}19, KHHN17]. Particle [CC14, BBGJ13, BPK^{+}16, LOLA13]. particulate [CGG^{+}18, RPC^{+}17, YCF15]. partition [Pie10b]. Partitioning [TGM^{+}21, ACE^{+}13, DGKSM13, LLQN13, WBB^{+}21, XKI^{+}11]. partners [CWM12]. partridge [THOA10]. parts [GS10]. parvum [GRB12]. pass [FBA^{+}11]. Passage [BLH10, ZVS18]. Passeriformes [LIM^{+}16]. Past [ENC^{+}15, BUB^{+}12, CS17, GAP^{+}19, HPC13a]. Past-century [ENC^{+}15]. pastoral [MGW^{+}21, QCLW18]. pastoralists [MHY^{+}15]. Pasture [BLBB13, CCSF11, CL15b, HMA10a, PGR^{+}16, PBG^{+}17, dJGM^{+}18]. pasture-woodland [PGR^{+}16]. pasture-woodlands [PBG^{+}17]. pastures [DYU^{+}12, PKT15]. Patagonia [GLT10, GdTCFG10, MTPC13, MMGK15]. Patagonian [MCMT^{+}16, RMR^{+}16]. Patch [RAdA12, XSL18, ALKB10, BS10a, DMDJ13, DSBS15, LDM15a, ZWG^{+}13]. patches [BS10a, DFK^{+}11, GCK^{+}15, Pec12, RBRK16, RH15, SXNL18, ZW11, dS18]. patchwork [IBD^{+}19]. patchy [BCI^{+}17, GCvDM12, JDZC14, OE12]. path [FSN^{+}15, LSR16, LLL^{+}10]. pathogen [HTVE18, JX15, MBLC14, RAF17, WBW14]. pathogens [CLLB12, MKH^{+}17, MC11, Mur13, TCGC16, WGD^{+}17]. paths [CK12, LT15]. Pathway [DPV^{+}16, MK13]. pathway-based [MK13]. pathways [HBP18, Jag14, MSM^{+}16, POS^{+}19, SBA^{+}19, Vav10]. Pattern [JGF14, Mer12, RJ11, WZWL10, AAFH15, CWLC13, CAMNN16, CGdCCPCPM14, GGSP14, GPCG15, HS13a, HSM^{+}12, Hui11, JX15, JCKG15, LPF11, LCS11, MPKC12, MJL^{+}13, MJL^{+}14, PAGCV^{+}10, PKD11, PLS11, SZSV12a, SHBG10, SMH^{+}12, TJYZ11, THOA10, UM18, VAO^{+}17, WYY^{+}19, XZSL17, YZC14]. Pattern-formation [Mer12]. pattern-guided [HS13a]. Pattern-oriented [JGF14, RJ11, AAFH15, CAMNN16, GPCG15, JCKG15, MJL^{+}13, MJL^{+}14, SMH^{+}12, THOA10]. patterning [ZNS^{+}18]. patterns [ABB^{+}10, BEZ19, Ban11, BA14b, BGG^{+}16b, BFKS16, Che15b, CLT^{+}11, Con11, Con12, CKMD14, DBK^{+}13, DS21, DHL^{+}12, FM10, Flo18, GM10, GGZ^{+}16, GLT10, GMS11, GBR^{+}13a, GLS10b, GKKW10, HCM^{+}18, HWL17, KMRW16, KSI14, LMR16, LZLC13, LLM^{+}13, LJ15c, LLQ^{+}15, LPV11, LL18a, LBdA^{+}19, LAC10, MBG17, MBG^{+}10, NP15, OQH^{+}13, PR11, PBEL17, PvVC15, PB14, RHO11, Rin14, RCBPMN10, RR10, SSM^{+}13, UR10, VH12, WCL^{+}17, War10, WHNR19, XTLM15, YYL13, YSA^{+}14]. payments [CLA^{+}12]. PCBs [SHS^{+}12, TPL^{+}15]. PCLake [JSL^{+}10, JTB^{+}19]. PDE [EGA12]. PDP [FCFT16]. pea [AWL^{+}13]. peach [GHS^{+}10, GOSH^{+}12, KL10]. Pearl [WLDL12, WDL^{+}15, XCL^{+}11]. Peat [ZNS^{+}18, HV12]. peatland [QFG^{+}13, RML^{+}15, Vav12]. peatlands [DBG14b, DBG14a, GWK^{+}12, GKW^{+}13]. Pecten [MG18, MHG18]. pelagic [CRUF13, FE14b, HPRNP12, KSAS12, LJ13, Löp11, PSWW11, RPP^{+}15, WMCM12, WFN^{+}16]. Penaeus [KSZW16]. penalised [GGV11]. Penguin [BPR^{+}15, WSS^{+}16, ACE^{+}17, CWM12, WCS^{+}14, WSW^{+}16]. penguins [PLS^{+}16]. Peninsula [RDR^{+}13, AR11, ABC^{+}17, BGA17, KTV^{+}15, KPK^{+}17, ROSC15]. Pennsylvania [FWRZ10, Fel12]. pennsylvanica [KKR19]. people [Abe10, SvBH^{+}14, VLZ^{+}19]. Peponapis [GLSA^{+}11]. per-capita [CTC^{+}18]. Perception [BPM^{+}15, WALCY12, ADF13, LGMLD19]. Perception-based [BPM^{+}15]. perceptions [AADA14]. perceptron [Fis21, MMMCAHM21]. perceptual [GPZB18]. perch [DGKSM13]. perennial [DAW^{+}10, HCW^{+}17, HGN^{+}16, JBB^{+}12, PLL^{+}19, SV13]. Performance [OBLJ13, SvWP^{+}15, SP12, WBB^{+}21, ABAG15, ABAG16, BSFR17, BOGA14, DKJH13, FAPM12, FCA18, GGV13, GdSHA10, HFU11, KFWM19, LM12, Mar12, MWK19, MROGVR12, MDG10, MDV^{+}12, NL12, NMA^{+}19, PJ15, RWVK11, RGH^{+}13, SC21, SBM^{+}12, SMI^{+}19, SDEP18, SPB^{+}17, TZS^{+}14, TBM^{+}14, UAB^{+}11, WR13a, WRB^{+}12, YHW^{+}12]. performances [SXZ^{+}15]. performing [JMB^{+}11]. peri [CLHH19, KJVL^{+}17]. peri-alpine [KJVL^{+}17]. peri-urbanization [CLHH19]. perilous [WSS^{+}16]. period [GZK^{+}11, HVL^{+}19, NMC^{+}14, SSY14]. periods [JFS^{+}12, RBM^{+}17]. periphyton [FBL^{+}10, KW18, LSL10]. Peromyscus [BB18]. perpendicularity [ALKB10]. Pers [ESA12]. persicae [GOSH^{+}12]. persistence [BNL^{+}13, Bev15, BSS^{+}14, BGG^{+}16b, BRF^{+}11, CHNdÁP18, DFH^{+}14, GFW^{+}17, HWE^{+}18, Joh10, MJK^{+}17, MLMH18, RCSS13, Rin18, SZNL19, SNZ^{+}19, TBA19, YZ18]. Persistent [KFDL^{+}19]. personal [SM11]. Perspective [BFW14, LYC^{+}15b, ANMJ12, BGGL16, BU11b, CW17, CBR^{+}16, CAMCK^{+}12, FKWZ19, Flo14, Gre12, HF14, Ji15, Ju19, LB11, MWY^{+}18, NYIH17, Pat12, PFR15, TAR13, UAB^{+}11, VKdW^{+}10, Wal17, XHC^{+}15, XZCL16, YHZ19]. Perspectives [WHB^{+}16, CCF13b, LST15, vGKB10]. perturbations [BRB14, MSFR15]. Peruvian [JBCÑ11, MRN12, XCR^{+}13]. pest [ATCJ^{+}19, AD16, BNC^{+}18, BPM^{+}15, BCD12, GBB13, GPM16b, HL19, KLGL19, LZB17, PDV17, PvVC15, RJ11, RPPW16, SRM11, SMBGH^{+}19, TV15, WFS^{+}19]. pest-parasitoid [AD16]. pesticide [BVGR^{+}17, BWA17, ECO^{+}14, IPSH14, KDHC^{+}18, LSGT13, LBS^{+}14, MHRB19, QCB^{+}11]. Pesticides [KEHG17, FtHvdB^{+}14, JHT^{+}14]. pests [KLGL19, LCAJGFNS14, MKH^{+}17, WGE^{+}19]. pests’an [DPV^{+}16]. pet [BGA17, PTA^{+}12]. petrels [ATCJ^{+}19, ZDL^{+}17]. PG [FT16, GS19c]. pH [CTNH13, HKM^{+}18, NP15, SG11a]. phase [AB10, MJB^{+}11a, TCSG15]. phase-state [AB10]. Phenacoccus [FTK^{+}14]. Phengaris [CWJ^{+}13]. phenological [CJM^{+}12, DBR12, KKK^{+}19, PSG19]. Phenology [PB14, AGB^{+}13, BRR10, CMM^{+}10, CSS^{+}19, DZS12, DJMBD^{+}13, FAG10, FTK^{+}14, GLR^{+}19, GCR18, LSS^{+}15, LBC14, MBD^{+}12, MMK15, OFH^{+}19, OBLJ13, RWH^{+}14, RPGA15, SW19, VFFP16, YHJ^{+}17]. phenomena [ZBMC14]. phenomes [GHD^{+}19]. PhenoPine [KKK^{+}19]. phenotypic [CMD12, JBE^{+}15, KCLL14, MCRB17]. pheromone [GKKW10]. Philadelphia [BLOJ^{+}19]. philippinarum [VZF^{+}11]. Philippines [BWRA16, GL10, LSH15]. Philornis [LIM^{+}16]. philoxeroides [MSKS13]. Phloem [SH14]. Phoenix [TOQW11]. phosphatase [LBE^{+}18]. phosphorus [BGPJ10, CT11, DZW^{+}11, DJA^{+}10, GZS^{+}18, HZT^{+}13, HSM^{+}12, HG17, JSL^{+}10, JAH^{+}10, JDGJ13, KD13, KB10, KJMA11, KZR^{+}13, LRM^{+}14, MGMR15, MM16, OGOWW16, TXM^{+}10, WB17, WEJ11, XTPT11, YZA^{+}18, YLW^{+}11, ZKR^{+}13, ZLG16, TXM^{+}10, XTPT11]. phosphorus-enriched [TXM^{+}10, XTPT11]. photo [LBE^{+}17]. photo-physiology [LBE^{+}17]. Photo3 [HBP18]. photoinhibition [Mul11]. photoperiod [GLR^{+}19]. Photosynthesis [CGdCCPCPM14, CPMP13, HSD^{+}15, MDW^{+}12, MB14, MEC14, SRV^{+}12]. photosynthesis-based [HSD^{+}15]. photosynthesizing [Mul11]. photosynthetic [HBP18, HCL^{+}14, WB17]. phototroph [FD19]. phototropic [BLB^{+}17]. photovoltaic [CRTS15]. Phragmites [WJL^{+}13]. phreatophytes [LZS^{+}17]. phylogenetic [ABB^{+}14]. Physical [NvGZ^{+}16, SCH^{+}17, BTPG10, CIHL14, Els10, HPRNP12, KC18, KWGW19, MMCBP19, MvdWU^{+}16, PCTW17, PCRP12, PCZ^{+}16, RR12, RRH^{+}19, TMT^{+}18, VPM^{+}18, ZLG16]. physical-biogeochemical [PCTW17]. physical-biological [MvdWU^{+}16]. physically [LSL10, RBGB16]. physicochemical [HGBA^{+}13]. physics [IFC14]. Physiochemical [HLLP^{+}10]. physiography [CRAP13, MDH15]. Physiological [CBCB17, FP17, COBF13, COY^{+}10, CIHL14, CDL^{+}11, KWH14, LTL^{+}13, PRHH13, PBP^{+}15, TKW17, XKI^{+}11]. physiologically [NMdR16, ZVS18]. physiology [BBL11b, FAT^{+}15b, LBE^{+}17, LLC^{+}17, PAB13a]. phytoclimatic [GLA11]. phytoflagellate [PBK14]. phytogeographic [SP10]. Phytophthora [DLvGGNC16]. Phytoplankton [SA16, VCD^{+}17, ABR12b, ABR12c, BRMW18, BEBA19, BSH^{+}10, CS18, CTP^{+}17, DCMS18, DBK^{+}13, DWP13, DWP16, GEZ^{+}13, GO15b, GC14, HJWG12, HHK19, HGLZ13, JA15a, JX17, LY16, LWGH13, LLZ^{+}15, LWC^{+}18, LQW^{+}19, Lin15b, MBD^{+}18, MCH15, MYHN17, MFS18, MHMRAHHS19, MMC11, MFI13, NvGZ^{+}16, PCTW17, PCZ^{+}16, ROC^{+}13, SW19, TKW17, TVT12, TPO^{+}18, TS11, TBM^{+}14, VAK^{+}11, WRMB18, SFM^{+}18]. phytoremediation [WWYO18]. phytosanitary [HL19]. Phytoseiidae [dOMZ13]. Picea [JL18, KPT11, KMP14, SGG14]. picta [PL10a]. pig [WØ15]. pigeons [ZSS17]. pike [BVP^{+}13, MMVAHMC16]. pinaster [VALC12]. Pine [GBB13, SRT16, ACDC17, BHD16, BA14b, CMT^{+}11, CHK^{+}12, FMS13, GVV^{+}16, JAH^{+}10, LBdA^{+}19, LM19, Log19a, LCML11, MPKC12, MSM^{+}17, MK12, NPJ^{+}17, PB14, RHZ^{+}18, UDP17, WJK11, aGSCMF15]. pinea [CMT^{+}11, CPMP13, MPKC12]. pink [CAMNN16, CRQW18]. pink-footed [CAMNN16]. pintail [MRD^{+}12]. Pinus [ASV19, CMT^{+}11, CPMP13, KKK^{+}19, LCML11, MPKC12, SGG14, THJ^{+}12, UDP17, VALC12, ZXD^{+}13]. pinyon [yHAB12]. pioneer [WB10]. PIOT [SCH^{+}17]. pipiens [JSB^{+}14]. piping [AWSS14, GKF^{+}14]. piscine [ACE^{+}17]. piscivorous [BR16]. Pitangus [LIM^{+}16]. pitcher [LBN^{+}18]. pitcher-plant [LBN^{+}18]. place [BPWB14, LDS14, LCDJ^{+}12]. placed [SBC^{+}11]. plague [LBR16]. Plain [SDY^{+}15, WYHC15, PSKO16, YLC13]. Plains [KBS^{+}15, MR17, SVL12, HCS^{+}10]. planet [FvBM^{+}11]. planipennis [KKR19]. planktivorous [CW17]. plankton [BCB^{+}11, Ban11, BM16, CN10b, DBVS19, ER10, GRB12, JXL^{+}15, KJVL^{+}17, KK15b, MCN16, PVM12, PVM13, RPSA12, RWP11, WB14, WTWS19, dAEPE16]. planktonic [KSAS12, RGU^{+}15]. planning [BDWP16, DFH^{+}14, GBL^{+}13, HCK18, LWNG13, LLL19, MWB^{+}10, MJLC14, PBD^{+}17, RPA^{+}18, TAT^{+}16b, WSC^{+}19, YXD^{+}18, ZMMHR10]. plans [DLL18]. Plant [FAMD^{+}12, FMA^{+}12, FKDG14, LZF^{+}14, OE12, RCSS13, Agu11, AKLB19, ASV19, BDL^{+}14, Bai18, BJK^{+}16, BBHP13, BdCS11, BNGP19, BBL11b, CC12, CHvdW^{+}10, CLW^{+}19, Con14, CWJ^{+}13, CDM13, CGM^{+}14, DJH10, DPV^{+}16, DJA^{+}10, EGA12, FAG10, FDP^{+}10, FKH^{+}15, FTYI13, GWR11, GQRC^{+}17, GGV13, GANB11, GVC10, GCW^{+}16, dSGGBW12, GLRT16, GWF^{+}18, qHLlZ^{+}16, HLL^{+}19, HW12, HGAT17, HBA16b, JBE^{+}15, JX15, KSS11, LJAD17, LBN^{+}18, LMDG14, LCY^{+}13, LLQN13, LLH^{+}13, LPD14, LJCG15, LB18a, MDW^{+}12, MSG^{+}13, MBW14, MLMFN^{+}15, MTL^{+}10, MSV12, MK14, MGSB11, MBA^{+}19, NYIH17, NPJ^{+}17, OT15, OFGH^{+}19, PES19, PS13, PM16, RAM10, RJ17, RCC^{+}11, RSCS19, RP11, SC21, San11, SFHK12, SMG^{+}12, SRM11, SV13, TKP14, TRE^{+}12, TBL^{+}15, VMG^{+}12, VAO^{+}17, VLF^{+}17, WP14, XYCZ11, XYCZ13, XYY^{+}16, XSL18, YGL18, ZLLC13]. plant-animal [XSL18]. plant-dominated [XYCZ11, XYCZ13, XYY^{+}16]. plant-induced [VAO^{+}17]. plant-mediated [CWJ^{+}13]. plant-soil [Agu11]. plantation [BPS^{+}11, KMM^{+}13, LPX^{+}12, LYL^{+}17, Oga11, WZD^{+}14, aGSCMF15]. plantations [DSS11, MdMS^{+}19]. planted [KMM^{+}13]. plants [BWH13, BMCG12, CKMD14, DF13, FS14b, FSM19b, FTYI13, GPA^{+}13, GANB11, GdC14, HLS^{+}14, HWE^{+}18, HWL17, IWW12, McC16, MLMH18, NYH15, OMH12, PLB16, POB18, PLHG14, RW17, SCHGM16, SFR^{+}17, XYCZ13, YqWB^{+}15, YSA^{+}14]. Plastic [EKB18]. Plasticity [MSG^{+}13, CBA12, CMD12, JBE^{+}15, MCRB17, UDP17, YMCV21]. Plateau [FLY13, MPKC12, SCG^{+}16, BCC^{+}11, ZCSL18, CSS^{+}19, CBO16, GYL^{+}15, MLB^{+}16, MCMT^{+}16, SLZ^{+}17, ZZLZ14]. platform [LBA14, LBA15, RBGB16]. players [TPL^{+}15]. plays [LGMLD19, LBdA^{+}19]. plea [Lin10, Lin16]. Plecoglossus [FSN^{+}15, MSO10]. Pleistocene [Flo14, GLS10b]. plexippus [GPZB18]. plot [MBLC14, NPH^{+}17, SHM^{+}14]. plots [CTC^{+}18, PGW19]. plover [GKF^{+}14]. Plovers [CAMCK^{+}12, AWSS14]. plug [DLCR13]. plugin [KRE^{+}18]. PMA’12 [FKDG14, Ano14-40]. PnET [dBGS^{+}14]. poaching [KdVRdW17, RWS^{+}15, RWS^{+}16]. pocket [MRC19]. podocarp [MP17]. podocarp-tawa [MP17]. point [DPR10, DHL^{+}12, GGSP14, GMS11, LQW^{+}19, LL18a, MTL^{+}10, PLS11, RCSS13, YDZ^{+}21, ZH11]. Poir. [ESA12]. poisoning [KTK^{+}15]. polar [JS12, LPSC16]. polder [EBL^{+}11]. polders [HV12, HG17]. policies [CMCASS12, CRVS13, UFK^{+}10, vDSC^{+}10, vDHH^{+}14]. Policy [UFK^{+}10, BAA^{+}14, CGY11, Fat12, HCS^{+}10, Ju19, KG13, KUI^{+}18, LDL12, RvS16, SHB11b, WLDL12, WYHC15]. Polish [LCBNP19]. political [SDSH15]. Pollen [GGM^{+}12, BGK^{+}16, HLLO13a, HLLO13b, MHTC15]. Pollen-based [GGM^{+}12]. pollinated [TWL11]. pollinating [OBSL15]. Pollination [GAG^{+}12, ZCY^{+}19]. pollination-herbivory [ZCY^{+}19]. pollinator [FAG10, RW10, SGBG19]. pollinators [DF13]. pollutants [LSR16]. polluted [Mon13]. pollution [BFP^{+}18, GLC^{+}18, KPTC14, NKSA13, PMBM11, RAF17, SRXP14, SWZ^{+}12, SB15, WLZC11, YDZ^{+}21, ZH11]. polychaetes [DLL^{+}19]. polychlorinated [SHS^{+}12, ZLLZ13]. polyculture [XCL^{+}11]. Polycyclic [ZCSL18, XWW^{+}11, XQZ^{+}13]. polyethism [MPK18]. Polygonum [MBBA17]. polymictic [HSR^{+}15]. polymorpha [KVP16, Lin15b]. polymorphism [RPGA15]. polyps [MIC14]. polyunsaturated [SAAD16]. POMDP [Wil11a]. pond [HH12, LL18b, MXC13, MWM^{+}16, SRHL11, VHDM17]. pond-breeding [VHDM17]. ponderosa [BHD16, FMS13]. ponderosae [PB14]. ponds [BAP18, CBB19, NW18, SFC^{+}12, SBC^{+}15]. pools [AL16, GJB^{+}11, HWS11, SDD^{+}13]. poor [DOB^{+}15, DBG14b, DBG14a, RKNO17, SAKB^{+}10]. poorly [FCLFB18]. POP [Och10]. populated [PLFO^{+}12]. Population [Ano12a, CHNdÁP18, GC11, GT14, HTL^{+}14, JDZC14, KWGG19, LN16, MAP^{+}13, MCA18, Rin18, TVI^{+}13, TWS^{+}11, ALG16, AGJ^{+}12, ACC^{+}10, ACCM16, ACKS13, ATCJ^{+}19, AGB^{+}13, AHNK13, BFA11, BAA^{+}14, BBB^{+}17, BG19, BKV17, BNL^{+}13, BPB^{+}18, BH13b, Ben12, BSC^{+}15b, BBL11a, BLRN15, BNC^{+}18, BDG^{+}14, BQF^{+}18, BPM^{+}15, BPM^{+}17, BTBM17, BAJ^{+}11, BWML15, CL12a, CLBG15, CGV^{+}15, CEWW18, CMH14, CWL^{+}16, CS17, CMA^{+}13, CJC11, CLDRJ18, dSCdA19, CL12b, DJT^{+}19, DSJ19, DTC^{+}10, DMR13, DHS^{+}16, DFF17, DMBS11, EPA^{+}10a, EPA^{+}10b, ECO^{+}14, FGH^{+}10, FMG^{+}16, FtHvdB^{+}14, FvBM^{+}11, FSM19b, FNR11, GHWG^{+}14, GQRC^{+}17, GQNMR18, GPM16b, GV11, GHS^{+}10, GJK^{+}15, GMvV17, GDDFT16, GDW12, HCW^{+}17, HFSL14, HPS10, HTB13, HTB17, HBS16b, HAOS17, HBT^{+}18, HMA10a, Hui11, IPSH14, IN16, IKL14, IKH^{+}11, IGBSASSJ16, JSKK15]. population [JBH^{+}14, JBB^{+}11, JOHW17, JJMW14, KSS11, KKR19, KUI^{+}18, KFDL^{+}19, KYvdPS14, KL10, KAC^{+}12a, KAC^{+}12b, KH15, LJLC15, LAR^{+}16, LCH12, LBWW14, LBW18, LLH^{+}13, LMCA^{+}12, LZYH19, LW17, Log16, Log19b, MTZ^{+}10, MV11, MCRB17, MPR13, MD16, MBG17, MGD^{+}12, MFD^{+}15, Mat13, MSB^{+}17, MSCM18, MIC14, MG18, MPK18, MBM^{+}18, MHRB19, MGWW13, MDVG^{+}15, MATS14, NNST^{+}14, NZL13, NL12, NMdR16, Och10, OJM^{+}12, OLC^{+}17, PRHH13, PRL^{+}13, PRH17, PMRGF10, PC15, PZM13, PVB16, RJG^{+}13, ROG^{+}16, RGHS11, RR12, RRG^{+}15, RGRMRR18, RMvW13, RH17, SCB13, SJTS13, SBCB13, Sak13, SCHQA16, SK10, SBR^{+}19, SXNL18, STD15, SSC15, SFL^{+}19, Slo11, SPWR17, SvdR10, SKPHW16, SGBG19, SMBGH^{+}19, SW15, TSP19, TPG14, TMC^{+}16, TSM^{+}15, Tho10a, TF13, TL15, TNTH11, TSSL11, VEK12, WHRR15, WSSM15, WFA^{+}13]. population [WCS^{+}14, WRSL11, WFDDR10, WAGM17, WPZ^{+}14, XCR^{+}13, YZ16, YSB11, ZJL^{+}16, ZBSF17, ZKSF17, ZSSR18, dAdSC15]. population-driven [TSM^{+}15]. Population-level [MAP^{+}13, FtHvdB^{+}14, GHWG^{+}14, LBW18]. populations [AAME11, ACE^{+}17, AHS^{+}17, ARV^{+}16, BDL^{+}13, BAAB18, BBRJ15, dÁGP14, BRS^{+}18, BSHG16, CCCP17, CVS^{+}13, CDK^{+}17, CDL^{+}11, CMGF14, CD15, DCCS17, DPPI14, Dil10, DFK^{+}11, GHD^{+}19, GCTS15, JX15, Joh10, JHT^{+}14, KCAP17, KCG18, KEHG17, LG10, LSGT13, LTGD16, MHRB19, Mon13, MPGCS14, MNHW^{+}13, MFI13, NdA16, NS15, NYF16, NTK14, OOH^{+}12, PL10a, PF12, POS^{+}19, PDV17, PP12, Rai13, Rem11, SBCB13, SBA^{+}19, SBDS13, SMT19a, SMT19b, SSQ^{+}10, SED15, Sto10, SNZ^{+}19, THOA10, TT17b, VSC18, VP12, WGL^{+}12, WNF^{+}15, WR13b, WSW^{+}16, WSBP13, ZZA^{+}18, ZW11, ZYKS15]. Populus [ARM^{+}12]. porpoise [FYR^{+}18, MCL^{+}17, NNST^{+}14]. port [GSJ^{+}10]. Portugal [BRV11, BPGSS^{+}18, DVD^{+}14]. Portuguese [DPBS10, TDC^{+}11]. position [AB10]. positions [JBCÑ11]. positive [FGPR10, MCA18, YZ18]. Possible [PMP^{+}13, CAMNN16, CRTS15, SKK^{+}13]. Post [VB10, AFC12, CADB19, LM19, SVH^{+}19, SGG14, YWA^{+}17]. Post-fire [VB10, CADB19, LM19, SVH^{+}19, SGG14]. post-harvest [YWA^{+}17]. post-hoc [AFC12]. poster [LPA^{+}11]. Potamogeton [HH12]. potato [KYvdPS14, PHvB^{+}18]. Potential [GV15, KPK^{+}17, LCS11, RRBI14, RHR^{+}15c, AVL^{+}11, AMRH19, ADL^{+}10, BSH17, CVPR^{+}16, CNH^{+}11, COK^{+}11, CPAG15, DPBK10, DFK^{+}11, DSBS15, ENC^{+}15, FRV^{+}16, FKK^{+}18, GA15, HRSKW15, HHC^{+}13, IKL13, JDT^{+}17, KBS^{+}15, KWGG19, LN11, LBE^{+}17, LJ10, LPS^{+}10, LW17, MU17, MSB^{+}17, MPGCS14, MCP16, OQH^{+}13, Qia12, RRHH10, RTE^{+}15, SJS^{+}13, SGP18, SGDP19, SGKR10, VZF^{+}11, WC19, WSW^{+}13, WLY^{+}15, WXB10, Wol11, XTLM15, YKvWK16]. potentially [LB15]. Potentials [FT10]. Potomac [CTNH13]. power [GSF^{+}16, GCW^{+}16, dSGGBW12, KPS17, Lee14, RDD^{+}17, Sci11b, SSR16, SSQ^{+}10, SGHW^{+}17, Til15]. Poyang [DZW^{+}11, YDL^{+}15]. pp [Jör11a, Jør14a, Jør14b]. practical [AVG14, Mar17]. practice [HCL^{+}16, KB11, Log13b, VBS14]. practices [CSH11, DMBS11, LVLS14, LLT^{+}17, MBLC14, OWMK10, PAT17, RSD^{+}15, VSB13]. prairie [TSB^{+}18, XLL15, ZWJ^{+}17]. Prairies [LLT^{+}17]. prawn [CLW^{+}19, KSZW16]. Prays [GCTS15]. pre [FMG^{+}16, Lin10, Lin16, WPZ^{+}11]. pre-alpine [FMG^{+}16]. pre-balance [Lin10, Lin16]. pre-commercial [WPZ^{+}11]. PREBAL [Lin16, Lin10]. precipitation [BTS19, DS21, GLRT16, GWF^{+}18, HBG^{+}16, HC11, XLL15, YSA^{+}14]. precision [KCP^{+}15, PC17, XLR11]. Predation [LAK^{+}21, BLD10b, BBL11a, BEBA19, CB15, DWWS15, DdSCG11, MW15, MLK17, MYJ^{+}14, SW15, URJ10, VWTA16, ZFG13]. Predator [Rin14, RCP17, AAA13, BKV17, BTB17, BEBA19, BWML15, CL12a, CTC^{+}18, CCG15, dSCdA18, DHHB15, F.11, FR11, FSM^{+}19a, Flo14, GK14, GORJ12, Jen12, JSW^{+}13, KG13, LDM15a, MLW11, MW12, PME^{+}16, RHO11, SGSR10, Ver10, VWTA16, WN17, WCF^{+}16, WZWL10, WM14, WSSM15, ZFG13, dOMZ13, dSdCF10]. predators [BV15, FE14b, GMTC^{+}17, KNU^{+}15, NdA16, NGL13]. predatory [BPB14]. predict [AGQP^{+}14, ASL13, CTP^{+}17, DAF^{+}11, FLB^{+}15, GBP^{+}10, LAD11, MPR13, MRM^{+}17, MMMGLN^{+}18, PB14, QL15, RRJVA13, RV13, SZSV12b, SW15, TKW17, TWL11, VZF^{+}11, WZD^{+}14, YHW^{+}12, dSD14]. predictability [Ban11, Bev15, NBK^{+}11, ZH13]. predicted [BTS19, BP16, HLLO13a, HLLO13b, JGB15, MMVAHMC16, WSBP13]. Predicting [BFNH14, BBGJ13, BBL11a, BNC^{+}18, BGPJ10, COPP13, CPS^{+}19, CSFR15, CDG15, CWM12, FTK^{+}14, FSV^{+}12, FBKS17, GCDA16, GSF^{+}16, GPZB18, HCW^{+}17, KP13, KHKC^{+}18, KJH^{+}13, LS10, LLKY18, LB18a, MSKS13, MJK^{+}17, MRC19, OHS^{+}17, Phi11, RJG^{+}13, RKS^{+}10, SLLW15, Vet12, WCL^{+}17, WBHM11, YAGMM^{+}14, ZZT13, vdVJS16, AWSS14, Ant14, AJDS16, BTB17, BCL^{+}14, BTPG10, CWMC^{+}12, CPAG15, For18a, GKF^{+}14, GC14, GEB10, HJWG12, HFMM^{+}12, Kos17, Mon13, NW18, PMS11, SGRAdJ17, TLZ17, YDD^{+}13, ZXD^{+}13, ZSG^{+}18]. Prediction [GDC11, JDP^{+}12, KC18, WX11, YKL^{+}12, AKvO^{+}10, CC14, FIG^{+}10, JTM^{+}17, KLS17, LAB^{+}10, Log19a, MRWN19, MRC19, SCB13, SZSV13, SELB10, VALC12, VAK^{+}11, YSA^{+}14, YKvWK16]. Predictions [GCM^{+}13, KASS12, WCF^{+}16, BR16, CS17, CBA16, DDR^{+}18, FCFT16, For18b, GCA16, GLR^{+}19, GB16d, LHW^{+}15, MSF18, MSR^{+}14, RLB^{+}12, RR12, RMHE15, SPWR17, SWGC16, WC19, YHJ^{+}17]. predictive [Ano17a, AFC12, BDL^{+}14, BS15, BLB^{+}17, BG11, CRO13, DSK^{+}11, GGV13, HD14, KRL11, KK12a, LKDD15, RKNO17, SC21, SPB^{+}17, TFMP11, dÁSMRT^{+}19, vVHZHvD13]. predictor [HJN12, MRWN19]. predictors [DKJH13, MGGLBIBD13, PLHG14, SGA^{+}16]. predicts [BHD17]. Preface [BPRT13, BNP10, Jør10a]. prefecture [YHZ19]. prefecture-level [YHZ19]. preference [FSN^{+}15, MCL^{+}17, SZV14]. preferences [MHMRAHHS19, RW10]. preferential [MSV^{+}11]. preferred [GH13]. prehispanic [KBC^{+}12]. Prelim [Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano16v, Ano16w, Ano16x, Ano16y, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16-37, Ano16-38, Ano16-39, Ano16-40, Ano16-41, Ano16-42, Ano16-43, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51]. Preliminary [Ano21d, Hal11, RMTM^{+}16]. prerequisite [CZR^{+}14]. prescribed [CPD^{+}11, TXM^{+}10, XTPT11]. presence [AGJ^{+}12, CGB^{+}13, FN15, FJH15, GKF^{+}14, GdSHA10, HMD^{+}16, HTB13, HTB17, LWY^{+}11, LB15, MMPG16, PKT15, PÓF^{+}18, UG12]. presence-absence [PÓF^{+}18]. presence-only [FN15, GdSHA10]. Present [JVY19, WZG15, BAAB18, BKP10, Bar11, GAP^{+}19, JX15, VPC11]. presented [LYP^{+}11, LPA^{+}11]. preservation [LKAL^{+}13, YYQ11]. presettlement [Tul14]. Prespa [KNM^{+}10]. Press [Jör11a, Jør14b, Kat11]. pressure [CGLA10, TBA19]. pressures [MLMFN^{+}15]. Preto [dSdCF10]. prevalence [FMF12, LLFL12, SPB^{+}17, WGTH16]. prevention [ZMMHR10]. Prey [MF12, AHS14, AAA13, BKV17, BWC17, BM16, CL12a, CCG15, dSCdA18, F.11, FR11, FSM^{+}19a, Flo14, GK14, GORJ12, GCvDM12, Jen12, KG13, LDM15a, LKG^{+}17, MLW11, MW12, MW15, NdA16, RHO11, SGA^{+}16, SGSR10, URJ10, Ver10, VWTA16, WN17, WCF^{+}16, WZWL10, WM14, WSSM15, ZFG13, dSdCF10]. Prey-abundance [MF12]. prey-subsidy [BKV17]. Price [Jør14b, Jör11a]. primary [ABD14, BSC^{+}15a, BH17, CLWH15, FCG^{+}19, FTTV^{+}18, GGZ^{+}16, HB12, HLLP^{+}10, HNWT10, yHAB12, LCL^{+}13, LLY^{+}13, LLT^{+}14, LZZP16, LJCG15, MMM^{+}16, MPKC12, MFS18, OdlFDP15, PFA^{+}16, RHvdTS14, RMTM^{+}16, SMC11, SM14, SLZ^{+}17, TMC^{+}16, WBS13, WNT^{+}13, WYS12, YqWB^{+}15, YLY^{+}18, YCL^{+}14, ZZFH15, ZZB^{+}18]. primate [BLD10b]. primates [BFNH14]. primeval [HOŠ^{+}12]. primrose [FLLB11]. Primula [FLLB11]. Prince [GPC14]. Principle [Jør14a, RHS^{+}15, RHD^{+}16, Sci11b]. principles [Ken19, LBGM19, RHR^{+}15c, RRH^{+}19, WSS^{+}16]. printed [KPS17]. priori [SBM^{+}12]. priorities [Bin17, CAMCK^{+}12]. Prioritising [HPO^{+}15]. prioritization [DMDJ13, MSM^{+}17]. prioritize [HAOS17]. prioritizing [DSA^{+}18]. priors [DG16, MMD^{+}17]. pristine [GKW^{+}13]. PRO [LHLW15]. probabilistic [Bil11, HYG14, LWC^{+}18, MAC14, MDKK15, SMSAN^{+}14, YTT^{+}11]. Probabilities [NSN17, Pec12]. probability [KF16, KKP^{+}11, LWY^{+}11, LYC^{+}15a, OMDP11, PTA^{+}12, ZR19]. probability-based [KF16]. problem [Bil11, JSCS19, KdVRdW17, Log13a, MWB^{+}10, Qia12, Ret12, RKW12, RCCN11, KCB^{+}17]. Problems [UM18, BFR17, BDWP16, HCK18, KK16, KFDL^{+}19, PK12, WK17, XTLM15]. procedure [ADF^{+}17, BAJ^{+}11, GCIB10, XBvdV^{+}14]. Procellaria [ZDL^{+}17]. Process [CdMMF16, HVCP16, LVB^{+}15, Pér18, SLH^{+}11, TVT12, ZWJ^{+}17, AAJB21, AD12, BTS19, IHSM12, CdOR10, CMLY19, CPS^{+}14, CLDRJ18, GGSP14, GLS^{+}17, GS19c, HWS11, HYL^{+}17, HGN^{+}16, JTB^{+}19, KWGW19, KRCVF16, LBH^{+}15, LTL^{+}13, LLT^{+}14, LL18a, LVZ^{+}11, PSB^{+}17, PWP^{+}19, RVL^{+}18, SHT^{+}21, SRSS12, SX11, SBA^{+}13, SRT16, SLD^{+}19, UM18, WGA^{+}13, WWYO18, YLC13, ZLZS14, ZCL^{+}16, ZGF14, CBGK15]. Process-based [CdMMF16, HVCP16, Pér18, SLH^{+}11, ZWJ^{+}17, AAJB21, AD12, BTS19, IHSM12, CPS^{+}14, GLS^{+}17, GS19c, HWS11, HYL^{+}17, JTB^{+}19, LTL^{+}13, LLT^{+}14, PWP^{+}19, RVL^{+}18, SBA^{+}13, SLD^{+}19]. Process-focussed [LVB^{+}15]. process-oriented [SHT^{+}21]. process-representation [PSB^{+}17]. process-rich [SRT16]. Processes [ASS14, AWWD16, AAJB21, ADF^{+}17, AJDS16, BPB^{+}18, BSS^{+}14, BA14b, CHPS17, CLLB12, CdMMF16, COBF13, CLT^{+}11, COY^{+}10, CGAB17, CIHL14, DVF^{+}19, DJH10, FDP^{+}10, FMA^{+}12, GGM^{+}12, GdTCFG10, GCK^{+}15, GVV17, HGPF15, HZZ^{+}18, HB10a, HTVE18, LV17, LHW^{+}15, MPR13, MNEG11, MDVG^{+}15, MRTW18, NU11, OGOWW16, Pat12, PLS11, QDH^{+}10, SPM^{+}18, SvdMKS11, ST13, SM13, SFC^{+}12, SYL^{+}13, SEHD18, SKPHW16, VVH12, WRMB18, XSN^{+}18, YLF^{+}15, ZH13, ZLF^{+}15, ZLFZ16, ZLZ17c]. processing [HKP10, TB11, TB12a]. processionary [GBB13]. processors [vdKGV11]. Prochlorococcus [GMvV17, HBB15]. produce [CPE13, PW10a]. produced [HKlL14, KHW^{+}16, RMLTCQ^{+}19]. producer [Kei12]. producing [GPA^{+}13, SBDA16]. product [KØ13, YqWB^{+}15]. production [Abe10, Abe15, ABAG15, ABAG16, ABAA21, AHF^{+}13, BPS^{+}11, BH17, BGGL16, CMT^{+}11, Cam15, CEC10, CCF13a, CLWH15, FDHL17, FPWG14, GPK^{+}13, GGZ^{+}16, GSA^{+}19, GBP^{+}10, GCR18, HB12, HFU11, HLLP^{+}10, HGE^{+}10, HNWT10, yHAB12, JC11b, Ju19, KR10a, KØ13, KM12b, KR14, LMGW16, LLY^{+}13, LLT^{+}14, LZZP16, LHB^{+}18, LCPB^{+}10, MMM^{+}16, MFS18, MK12, MGW^{+}21, OdlFDP15, PFA^{+}16, PAGCV^{+}10, PANL11, QL15, RVL^{+}18, RHvdTS14, RL15, RRHH10, RSDP^{+}12, RM13, RUSL19, SvBH^{+}14, SNS^{+}12, SJS^{+}13, SM13, SM14, TZS^{+}14, TVV13, WAS^{+}16, WBS13, WCL^{+}17, WØ15, YWH15, YLY^{+}18, YC13, YSA^{+}14, YLL^{+}12, YCL^{+}14, YDT^{+}13, ZZFH15, ZXD^{+}13, ZZL^{+}19]. Productivity [Kei12, AW16, ACK^{+}14, ABD14, BSC^{+}15a, BPC^{+}17, BNGP19, BFH14, BWML15, CSS^{+}19, CBB19, DBG14a, DWP13, FCG^{+}19, FTTV^{+}18, mGZK^{+}10, GEZ^{+}13, Gra14, HN14, Hoa11, JL18, KKBZ17, LKH^{+}18, LB11, LCL^{+}13, MSG^{+}13, MBW14, OES16, PVBC14, PGV10, PPO12, RMTM^{+}16, SMC11, San17, SLZ^{+}17, VCD^{+}17, WGA^{+}11, WGA^{+}12, WGA^{+}13, WNT^{+}13, WYS12, YKvWK16, ZZB^{+}18, ZSKI17]. products [GRRU15, GV15, PSWM^{+}10, Vav10]. profile [TTCH^{+}14]. profiles [BP17, DBR12, RPPW16]. profitability [dJGM^{+}18]. profitable [BWRA16]. profits [PPDF14]. prognosis [NGWZ16]. Program [HPW^{+}12, KFvSH15, ZZLZ14]. programming [BDWP16, Che10, EBA^{+}19, vDHH^{+}14]. programs [VC17a, YRGH16]. Progress [SLW17, YJL11, HhLY^{+}19, JNF16]. project [BGD^{+}16, GCL^{+}19]. Projected [SBF15, HEJP17]. Projecting [FPWG14, MNEG11, BvHR^{+}15, CJR^{+}16]. Projection [KMP14, Log13b, DK17, EEB^{+}17, EEK^{+}18, MTGD17, POBH10, ZBSF17]. projections [AS12, DKJH13, FRM^{+}19, HPC13b, KDT^{+}17, NA11, OJ10, SJB^{+}16, SWR^{+}16, TCW11, dBGS^{+}14]. projects [BRW^{+}12, YKCN10]. prokaryotic [LMSK12]. promote [BSS^{+}14, Fat14]. promotes [KMKL19, NSIiT17]. promoting [FR11]. promotion [YFL^{+}18]. promotions [FDR^{+}17]. prone [ADPS10, TLS^{+}16]. pronounced [GPK^{+}13]. Propagation [RR12, CPD^{+}11, IWW12, JGX13]. propagules [GPJL15]. Proper [TNHB12]. properties [BGC10, BFM11, BPK^{+}12, BRPK14, EIG^{+}10, Fie12, GANB11, HLLP^{+}10, KSG^{+}14, LB18a, MBHW12, MMC^{+}19, NOOM^{+}16, OCB^{+}13, PBP^{+}15, PICR13, TL10, UDRA13]. property [KVU^{+}10]. propionate [PJ16]. proportional [UG12]. proposal [CBA10, GSA^{+}19]. Proposed [RSD^{+}15, ROSC15, Sil11, TT17a]. protected [ASP^{+}15, BJJ12, BAT^{+}15, CGEdM12, dSCdA19, FJH15, FBDR17, GAP^{+}19, HA18, HCE^{+}19, PPB^{+}18, Pra12, ROSC15, VPB^{+}17]. protection [BPM^{+}17, CRVS13, NXZ18, RSCS19, WSE10]. protein [KAJØ19, KH15]. proteins [SWS16]. protocol [GBD^{+}10]. provide [BPWB14, SBM^{+}11]. providers [BS10a]. provides [GA15, SBVA^{+}13]. Providing [AMM^{+}14, HHUD14]. Province [GZS^{+}18, MJP^{+}12, YLW^{+}11, ZLZS14, ZXD^{+}13, LHL18, SC21, ZZLZ14]. provision [WYHC15, WLY^{+}15]. provisioning [CWM12]. proxies [VMSM10]. proximate [EAJG16]. proxy [FML^{+}17, RHR^{+}15b, RDR^{+}17]. pruning [GHS^{+}10]. Prymnesium [GRB12]. psammophytic [HYL^{+}17]. pseudo [IBH^{+}15, LPT^{+}18, SHS11, GDDFT16]. pseudo-absence [IBH^{+}15, LPT^{+}18, SHS11]. Pseudo-nitzschia [GDDFT16]. pseudoacacia [BRKP16]. Pseudococcidae [FTK^{+}14]. ptarmigan [ZPNM13]. public [SM11]. publicly [HRWT12]. published [Bec10, Jör11a]. Publisher [Ano10-27, Ano12-38, Ano19-49]. Publishing [Wei19]. Pueblo [KBC^{+}12]. Puerto [HSVL12]. PUFA [SAAD16]. pulex [KMS^{+}12, KMS^{+}13]. pulse [CSN^{+}19, CGGL16, Jag14, Ji15, LMCA^{+}12]. pulsed [CZW10]. pulses [Lee14]. Puma [CGV^{+}15]. pump [CQS15]. purification [HZT^{+}13]. purpose [KK15a, ROG^{+}16]. purse [JBCÑ11, MROGVR12]. purse-seine [MROGVR12]. Pursuing [ZH13]. put [KB11]. Putting [Ula14]. puzzle [SFM^{+}18]. pyramids [RDR^{+}17]. pyrenaica [AR11]. pyrenaicus [BRV11]. Pyrenean [ACM15]. Pyrenees [LBC^{+}15, LBdA^{+}19]. pyrophite [UDNSLC^{+}18].

Qinghai [CBO16, SLZ^{+}17]. Qinghai-Tibet [SLZ^{+}17]. Qinling [FLX^{+}14, GFW^{+}17, LGDW12]. Qom [SC21]. quadrat [San11]. QUAL [ZSJ15]. Qualitative [ADJ^{+}17, BJPB11, GV15, WMCM12, MDJ^{+}11, Pat15]. Quality [Jam16, BA14a, BJJ12, BT16, hCjHMW18, CMEP15, DC10, DPvBvA11, ER13, FLB^{+}15, GMP13, GCL^{+}19, HHL17, HGBA^{+}13, JLZ^{+}18, KSS11, KMRW16, KPM^{+}14, KLS17, LHL^{+}17, LLQ^{+}15, LPV11, MTZ^{+}10, MA10, MP13, MH10, MSHC13, MCP16, PAT17, PANL11, RLB^{+}12, RAF17, RZF^{+}13, RCCN11, SFMA15, SAKB^{+}10, TSB^{+}18, UB12, WMW^{+}12, YZPL16, ZWRK13, ZSJ15, ZLJM17b, ZLJM17a]. quandaries [Mar17]. Quantification [DMB^{+}19, KSZW16]. quantified [BPK^{+}12]. quantify [BYR18, KNPH^{+}17, VSF10]. Quantifying [BRB14, CGS^{+}17, CTMS10, CBA12, CBA16, DMH17, EST^{+}19, For18a, HHUD14, HM12, LY11, LBW18, LMGM11, MMK15, Pat15, SBA^{+}19, VQS^{+}18, ZSKI17, CBA10, LBH^{+}15]. Quantitative [BP17, CLWH15, For18b, MSM^{+}16, HM17, LC13, MBBA17, Pen11, VR12, YZLX14]. quantities [GCV14]. quantity [CRBC11, KMRW16]. Quantum [ABC^{+}14, RHR^{+}15a, RHR^{+}15c]. Quebec [LMY^{+}11, LYP^{+}11, LPA^{+}11, WLH^{+}10, ARM^{+}12]. queen [AOW12]. Queensland [CRO13, HPO^{+}15]. quelen [dABLBAdOR17]. queried [BPR^{+}15]. question [KWGW19]. question-driven [KWGW19]. Quinte [KZR^{+}13, ZKR^{+}13]. quota [PBP^{+}15, vDHH^{+}13, vDHH^{+}14].

R [Kat11, CF14, KB11]. R-INLA [CF14]. Råcksta [WFM16]. radial [AK17]. radiation [AML13, ENM11, JMSM12, KBS^{+}15, LCS11, MRCW10, Pie10b, Pie10a, SMC11, ZML^{+}12]. radiative [Bai18, BBDB13, GGR^{+}13, Pie10b, Pie10a]. radioecology [Mon13]. radiogenic [SBDA16]. radish [GDT^{+}14]. rain [LCZ11]. Rainbow [KGM12]. Rainfall [ADPS10, AKLB19, BVVS15, BMD18, CMGF14, FKSK^{+}10, KK12b, LDP14, Tie16, VSC18]. rainfed [TDC^{+}11]. rainforest [LYL^{+}17, PTE^{+}12, SLZ^{+}13, PMF^{+}13]. rainforests [KK12b, PMS11]. RaMDry [FS18]. Ramsar [MGMR15]. ranavirus [POS^{+}19]. ranch [PAT17]. Random [DG11, FMF12, TTL^{+}18, VZF^{+}11, CCS13, CH12, GGSP14, LJ13, MYHN17, PvVC15, PPBD13, SVH^{+}19, Kos17]. range [AVL^{+}11, BSC^{+}15a, BM11, BRL12, BSHG16, DHL^{+}12, DHS^{+}16, GM10, GPZB18, KPTC14, MPP^{+}15, MRM^{+}17, NA11, RHO11, RRJVA13, SHT^{+}21, WFA^{+}13, WSC^{+}19, XCG^{+}16]. range-wide [DHS^{+}16]. rangeland [AHF^{+}13, EMH10, FS18, KF16, OES16]. rangelands [PLL^{+}19, SC21, vLTdBP16]. ranges [BGD^{+}16, OCD^{+}13, Str15a, Str15b, VCE17]. ranging [MvDL^{+}10]. rank [LKDD15]. Ranking [BS10a, AKvO^{+}10, RP10]. ranks [RPPW16]. rape [GDT^{+}14]. rapid [GGZ^{+}16, LSZC10, LLQ^{+}15, SLSS15]. rapidly [POS^{+}19, SDY^{+}15, WZG15]. Raptor [BFP^{+}18, DDBK15]. rare [DDBK15, HPJ15, LCCL15, LJ15c, PH18]. rarity [ACE^{+}13]. RAS [HD14, HD16, WZGH18]. rate [ACKS13, BP15, CCS13, CMSR19, Dil10, EH15, EH19, FE14a, HMPC11, Hos11, LKP13, MCR^{+}15, NYF16, PR13, SLB^{+}18, Sim11, VSF10, Vet12, WCL^{+}17, dAdSC15]. rates [CEC10, DTS^{+}12, DCCS17, DSS11, DMH17, GVGZ^{+}21, GDC11, OBSL15, PRL^{+}13, SG11a, TWH^{+}14, WC19, dSJHG10]. rather [GHST19, TCGC16]. Ratio [Jen12, LDB14, CMD12, FML^{+}17, GBP^{+}10, HTL^{+}14, MEC14, RAdA12, SK10, TVI^{+}13, TMC^{+}16, WB14]. Ratio-dependent [LDB14]. ratios [CS18, JS12, LWGH13, SBDA16, WWYO18]. ratsnake [DWWS15]. Raugei [UB12]. ray [Bai18]. ray-tracing [Bai18]. rays [DOB^{+}15]. Re [Ula10, WTWS19, AWL^{+}13, CVPH^{+}14, ELPK12, KNPH^{+}17, MJB^{+}11b, HPC13a]. re-assembly [AWL^{+}13]. Re-creating [HPC13a]. re-encounter [KNPH^{+}17]. re-establishment [CVPH^{+}14]. Re-evaluating [WTWS19]. re-loaded [ELPK12]. re-organization [MJB^{+}11b]. reach [JX17, JJYM19, LWY^{+}11, SCHGM16]. Reaction [LD16, BJR16, BRF^{+}11, PvVC13]. Reaction-centric [LD16]. reaction-diffusion [BRF^{+}11]. reactions [Vav12]. reactive [BMR^{+}18, NP15]. reactivity [VMSM10]. reactor [DLCR13]. reaeration [CGdCCPCPM14]. real [BP17, CMSR19, DLL18, KK15a, OMH12, ZSS17]. real-time [CMSR19]. realism [GB16d, PPBD13]. realistic [CLBSG11, JGF14, LBS^{+}14, MHG18]. realities [MH11]. realized [BFS11, HWH^{+}18]. reallocation [BWRA16]. really [Nie16, Sci11b]. reasons [Dol10]. reassessment [Sci11b]. rebound [FG11]. rebuilding [BWML15, ZCR16]. recalibrate [LJLC15]. Recalibration [Jam16]. recapture [ABB^{+}14, BS10c, CGAB17, HVL^{+}19, OMDP11, RGHS11, RDJ19, TB12b, vPTOW18]. recharge [SPR^{+}18]. Recirculating [HD14]. Reckoning [Ula14]. reclaimed [MRG13a]. recolonization [Mat13]. recolonizing [LN16]. recommendations [LDL12]. Reconnecting [BF11]. Reconsider [MW15]. reconstruction [BRHD15, CLBSG11, GGM^{+}12, RMLTCQ^{+}19]. recontextualizing [SS10]. records [Tul14, WSC^{+}19]. recourse [LHZN11]. recover [ZKSF17]. recovery [AMF12, BMU10, CVPR^{+}16, CRAP13, IKL13, IKL14, KRCVF16, LCH12, PRH17, PDD^{+}18, RTE^{+}15, TXM^{+}10, VPC^{+}14]. recreation [VLZ^{+}19, WGD^{+}17]. recreational [AWF10, CMCASS12, CMCC15, GH13]. recruit [HLK^{+}12]. recruitment [ACM15, BBGJ13, CAL13, FIG^{+}10, FC12, GHH^{+}15, Lem16, OAPP12, SGG14, SP12, THJ^{+}12, vPTOW18]. recurrent [BT16]. Recycling [ALL11, KZR^{+}13, LDJZC15]. red [BSCG10, CNI11, DSW^{+}17, GBP^{+}16, Hat14, KPT11, KPTC14, KMP14, PPDF14, WFDDR10]. redesigned [CRO13]. redistribution [XLL15]. redline [GFW^{+}17]. redox [KD13]. reduce [BOGA14, CC14, MHS^{+}17, PLL^{+}19, WR13b]. reduced [CL10, MLMH18, MGWW13]. reduces [LM12]. Reducing [DPLC13, FM10, MGD^{+}12, MSH^{+}19, TB19, Wil12, CPD^{+}11, DSBM^{+}18, DSBS15, GJB^{+}11, PDV17, RPA^{+}18]. reduction [CGY11, DZW^{+}11, DSBS15, GWC^{+}10, HB15, KPTC14, LCP14, LCP15, MGR^{+}18]. reductions [WCG^{+}11]. redundancy [LPF11, RHS^{+}15, RHD^{+}16]. reed [XLS11]. reef [ABAG14, CRBC11, CTMS10, GRYM11, HJ10b, KR16, MTJF11a, MTJF11b, MB14, PWBG16, SM13, wTA12, VSF10, YMCV21, ZLFI11, CDK^{+}17, HBR19, MTJF11b, TLW^{+}14]. reefs [MJL^{+}13, MJL^{+}14, RUSL19, WLR^{+}11, MLAT^{+}18]. reestablishment [FRV^{+}16]. reference [PHC10, SHT^{+}21, UFK^{+}10, ZWW^{+}18]. refine [KHHN17]. refined [KAJØ19]. refinement [BU10]. refinements [Til14]. Refining [SK11]. reflect [Lin15a, OGOWW16]. reflectance [dMHGA10, PSS16]. reflects [MCS^{+}18]. reforestation [Ber13, LSH15, PMB^{+}17]. refractory [KPM^{+}14]. Refuge [SZSV12a, CRBC11, GORJ12, JFS^{+}12, LHZ10, MLW11, MW12, MW15, WM14]. regarding [Md18, YJW15]. regardless [KHJ^{+}19]. regenerating [AW16]. regeneration [CPMP13, CADB19, ENC^{+}15, FMS13, FPG^{+}10, HKN17, MWML13a, MWML13b, RPSA12, SLB14]. Regime [HT16, LBN^{+}18, BPB14, BMD18, CpW16, DWP11, FAMD^{+}12, LBC14, MCP16, OGG^{+}16, WP14, ZW12]. regimes [CNH^{+}11, CN11, DSBS15, GRD^{+}14, GWF^{+}18, HCS^{+}10, LBE^{+}17, LDO^{+}18, RJH^{+}10, RVNB11, SKHH19, SKMB11]. Region [DMR13, MRZU14, San17, TFMP11, ALC19, ADPS10, BLOJ^{+}19, CCF13a, CR11, CR12b, GB15, HRSKW15, HLW15, JVY19, JMW11, LCH12, LLM^{+}13, LLZ^{+}16, LYL^{+}17, LCT^{+}17, MU17, MPL^{+}16, PPB^{+}18, RPC^{+}17, TOQW11, TJYZ11, VB10, WLC^{+}19]. Regional [KDT^{+}17, KKL^{+}12, MTJF11b, YWA^{+}17, ARM^{+}12, BA14a, BCO17, BGG^{+}16b, Cho13, CRTS15, CHK^{+}12, DLZ^{+}17, DS12, DLW^{+}17, FTK^{+}14, GHS^{+}16a, HBS16b, HGE^{+}10, HSM^{+}12, yHAB12, HPW^{+}12, KGYG15, LJBS12, LHZN11, LZZ^{+}14, LWL^{+}17b, LPS^{+}10, LLP17, LCT^{+}17, MGS16, dMHGA10, RJH^{+}10, SM14, SLD^{+}19, WV12, YWH15, YLHZ14, YYH19, YHZ19, ZHC^{+}16, ZQC15]. Regional-scale [MTJF11b, DLW^{+}17]. regions [AHF^{+}13, BM11, CHTC11, CN10a, DK17, HKM^{+}18, LZZ^{+}14, SCG^{+}16, WL13]. registration [PBO19]. regression [Fis21, MROGVR12, SK11]. regressions [GGV11]. Regrowth [JBB^{+}12, VB10]. regular [KMM^{+}13, SZNL19]. regularization [CDR^{+}13]. regulate [SKPHW16]. regulated [CL10, Pen11, TMT^{+}18]. Regulating [MSS^{+}19, GWF^{+}18, HV15]. regulation [AHS14, AL16, BFP^{+}18, BSC^{+}15b, hCjHMW18, LJ10, LPD14, MBBA17, MUP^{+}13, NYF16, WSW^{+}13, ZW10, dÁSMRT^{+}19]. regulations [RIMPBT19, RAB^{+}10]. regulatory [KPS19]. rehabilitation [HPO^{+}15]. Rehd [LGDW12]. Reindeer [PKT15]. reinterpretation [CT12]. reintroduced [SNZ^{+}19]. reintroductions [ERGW^{+}11a, ERGW^{+}11b]. related [BCP16, BGD^{+}15, CGM^{+}14, DCCS17, LDYL18, MSF18, Oga17, Oga18, PCZ^{+}16, Til15, ZZY^{+}13]. Relating [GGZ14]. relation [ALKB10, GCL^{+}19, JX15, LCS11, MTL^{+}10, MGS15, SGA^{+}16]. relational [DTS^{+}12]. relations [BRPK14, FC11, LYCZ11, PVL^{+}14]. Relationship [HWL17, LSR13, CBA10, GAV^{+}11, HZSH17, LB11, LWGH13, LLQ^{+}15, LQW^{+}19, LJCG15, Pat16a, PPO12, PDG^{+}14, RDD^{+}18, SRTS18, SGKR10, TQZ12, WYY^{+}19, ZWG^{+}17]. Relationships [CT16, DSS11, GDC11, ADJ^{+}17, Ano12s, AS14, BCP16, BBHP13, CJC11, FMW^{+}19, GBR13b, HMPC11, KBL10, KWGG19, MV11, MCW13, ROC^{+}13, RZF^{+}13, TTCC^{+}19, WK11, YLL^{+}12]. Relative [PM16, ROG^{+}16, VCE17, ZGF14, BEZ19, CdMMF16, CDT10, CD15, GSSM13, KR10a, KGM12, MHL^{+}19, NGL13, RAB^{+}10, WSZT10, XCG^{+}16]. relaxation [Loe18, OFK^{+}12]. release [BCL^{+}14, GPD^{+}13, KUI^{+}18, KPSC10, LCD^{+}16, RDJ19, YYQ11]. Relevance [Gia14, HTL^{+}14, MSGG14, MRM16, MvDL^{+}10, MDG10, UFK^{+}10]. relevant [LBH^{+}15]. Reliability [RH15, KCG18, LBSB^{+}14]. reliable [Cho13, SBM^{+}11]. relict [KKP^{+}17]. relictual [BNGP19]. relied [NGGFM19]. Reloaded [JBV^{+}15]. relocation [DÖWB12, JMdW^{+}10]. remains [YL17]. remnant [DFK^{+}11]. remobilization [SBC16b]. remote [BQF^{+}18, CZ11, CLHH19, COK^{+}11, CAY^{+}10, CKMD14, GYL^{+}15, JSdSdL^{+}18, LSS^{+}15, LAB^{+}10, LCL^{+}13, LZZ^{+}14, LLH^{+}17, MPT^{+}11, MCBM12, MVC^{+}17, MMK15, PTE^{+}12, PSSH11, SX11, VHM^{+}17, XLR11, ZCS13]. remotely [BRR10, BOW^{+}12, DBR12, KMRW16, VFFP16, VALC12]. removal [BCTL16, BS10a, BDG^{+}14, LLB^{+}19, LDC^{+}16, MJK^{+}17, OSS13, PANL11, RM13]. removals [FN18]. Removing [vPTOW18, NGL13]. renewability [WØ15]. renewable [AMH16, BLY11, BU16b, CT16, JN15b, LLTQ14, LLW^{+}15, LLTQ15, MRZU14]. repeated [MF16]. repeatedly [BRL12]. replacement [FWRZ10]. replicates [Ken19]. Reply [GCIB10, Pat10, Ten15b, WSS^{+}16]. Report [WLF^{+}16]. reporting [FM10, GU11]. Repositioning [RDD^{+}14]. repository [CVG^{+}15]. represent [CBB19]. Representation [PFR15, BAB^{+}16, BS17, GBW^{+}15, HBP18, HB10a, KZR^{+}13, KASS12, MAC14, PSB^{+}17, SLMS14]. representations [GMTC^{+}17, Sto10]. representative [IPSH14]. Representing [ERGW^{+}11a, ERGW^{+}11b, KMF^{+}21, HKVR13]. Reprint [HTB17, KAC^{+}12b, KCF18, LLTQ15, PVM13, RSC13, TB12a]. reproduces [MBG17]. reproducibility [Tat16a]. Reproducing [VAB^{+}18, Con14, NdA16]. reproduction [LBW18, MJL^{+}17, MRWN19, SM12, VAB^{+}18, Vet12, WGSR19, ZFG13]. reproductive [BGA17, CH14, DPvBvA11, Kal19, Log17, LARZN13, OLC^{+}17]. reptile [BSHG16]. Republic [ABS17, HV15]. repulsion [SW11]. required [TF13, YCL^{+}14]. requirement [CCL^{+}13, JMW11]. Requirements [SR14, BWC17, Cho13, SZSV13, YY11]. requires [Wol11]. requiring [PAB^{+}13b]. rerio [HTL^{+}14, KNU^{+}15, LLC11]. resampling [DG16, UG12]. Research [HhLY^{+}19, XJY^{+}16, CAMCK^{+}12, LST15, Pau14, YLW14, ZMS19]. Reserve [HV15, HZT^{+}13, LKG^{+}17, LCS11, LYC^{+}15b, MGGLBIBD13, WALCY12, Bil11, dSCdA19, DÖWB12, KBB^{+}11, LYC^{+}15a, LLL19, MRP14, TLP17, TLS^{+}16, LLL^{+}10]. reserves [BPU11, LLL17, VEK12]. Reservoir [CRO13, CMLY19, HFB^{+}11, MLW^{+}15, SRXP14, AWSS14, CIHL14, DBK^{+}13, GRB12, LPV11, LZYH19, RMP^{+}18, RRR^{+}10, REB^{+}16, YY11, YY13]. reservoirs [APWT15, CHR^{+}13, CCL^{+}13, CSFB17, GAV^{+}11, KRL^{+}16, NP18, RR12, WhFxGZ16, XJY^{+}16]. residence [WQD^{+}13]. resident [MGCL17]. Residual [XWW^{+}11, YDD^{+}13]. residue [Nei11]. residues [QCB^{+}11]. resilience [AS14, CHTC11, IPSH14, LY11, PMCKH13, Pra12, SK15, ZH13]. resiliency [DBS18]. Resistance [MJB^{+}11b, HGSJ15, HQF^{+}17, LYSW15, PHvB^{+}18, PPZ^{+}17, SPWR17, VHDM17]. resistant [FPO18, Mur11]. resolution [ARTBB14, BVJE18, IHSM12, BS17, CCLX17, FYWA12, GCK^{+}15, HBSF16, JJYM19, KHS^{+}19, KEM^{+}16, MBSW12, MMK15, MSR^{+}14, MOM^{+}11, Pog11, RWVK11, WFN^{+}16]. resolutions [YHS^{+}11]. resolve [RUSL19]. resolved [CGG^{+}18, MFS18]. Resolving [Wil11a, Bai18, KK16, MYHN17, TAT^{+}16b]. Resonating [Hel10]. Resource [Lee14, PDBP11, PDG^{+}14, UB12, BTM^{+}10, BPC^{+}17, BSCG10, BS10b, BEBA19, BPM^{+}15, Che10, CBP^{+}17, CD16, CN18, FMBD15, FS18, HCS^{+}10, HAOS17, HWL17, JZCL14, KP10, KMS^{+}12, KMS^{+}13, LVB^{+}15, LDG14, LYCU14, MRM16, MRZU14, OE12, PRHK19, PVL^{+}14, RRG^{+}15, SFM^{+}18, SRTS18, dWcWbZ15, WEB11, WHNR19, dOAC18]. resource-based [LDG14, dOAC18]. resource-limitation [SFM^{+}18]. resource-limited [FMBD15]. resources [BLY11, BMMC21, BdCS11, BSCG10, BPM^{+}15, CC11, DMB^{+}18, IT17, NTK14, SZ14, Sol19, Wil11a, ZMMHR10]. respecting [JZCL14]. respective [TCSG15]. respiration [CGdCCPCPM14, GYL^{+}15, Gav15, LYD^{+}17, LPS^{+}10, MDW^{+}12, Oga17, TWH^{+}14, VRL18, YLY^{+}18, YLL^{+}12]. Response [CSR15, GLRT16, RHD^{+}16, ZFT^{+}10, ZZZW15, ARM^{+}12, BU11a, CPMP13, CLLB12, CSH13, CWM12, CSL^{+}12, FM11, Fel12, GBP^{+}16, GQRC^{+}17, GJÁ^{+}10, GU11, GTM^{+}19, Gra14, HM10, JDT^{+}16, KKK^{+}19, KLGL19, LF11, LdMS^{+}16, LHZ10, LSJG17, LMGM11, LLC11, LDB14, MLK^{+}11, MJB^{+}11b, OFK^{+}12, OES16, PP16, Rem11, SQWS19, SLB^{+}18, Sim11, TKH17, TIC^{+}14, UB12, WZWL10, WPZ^{+}11, WMK^{+}13, WJ15, WSBP13, WUB15, WJK11, XZSL17, XTPT11, YYL13, ZJL^{+}14]. Responses [JWX18, MBD^{+}18, MPG18, PRL^{+}13, AK17, AS12, BMN11, BFJZ^{+}11, CBCB17, CDG15, FLLG13, FBDZ12, LLKY18, Loe18, MLMFN^{+}15, MRB14, MR17, MCG19, PW10a, PCTW17, RFV^{+}12, RA13, SCHQA16, SJGB16, SLB14, SGRAdJ17, TXM^{+}10, TRE^{+}12, WYHC15, XTPT11, XLL15, YGL18, YXD^{+}18, YDL^{+}15, dVLS11]. responsible [BPB14, PLS^{+}16, WSE10]. restitution [HB15]. Restoration [NJM11, NBD^{+}13, BGD^{+}15, CNH^{+}11, CLCC11, FMLM11, GSQ17, GGV13, HRSKW15, JDGJ13, KJH^{+}13, MAC^{+}12, MWB^{+}10, MSB^{+}17, PHC10, RJ17, RSD^{+}15, TBJ^{+}15, TAGC^{+}16, WWYO17, YSH^{+}10, YKCN10, YYQ11]. restorations [SXZ^{+}15]. restored [SNZ^{+}19, SGKR10, Yan11]. restraints [HVL^{+}19]. restricted [BLH10, GdSHA10]. restriction [WKZ^{+}19]. result [CMGF14, MHY^{+}15]. resulting [BFH14]. Results [CLD^{+}17, PTE^{+}12, SMH16, KVS17, MH13, NPH^{+}17, PSWM^{+}10, PGW19, WGA^{+}13, WV12, WR12, ZLF^{+}11, vSGR16]. resurrection [HHK19]. resuspension [RGU^{+}15]. retention [DKM^{+}14, HKVR13, HSM^{+}12, KD13, LDC^{+}16, MM16]. Retrieval [LLM^{+}13, XLR11]. Retrieving [DS21]. retrodictions [KASS12]. Retrospective [BFW14]. Return [GPM^{+}16a, FN18, JSW^{+}13]. return-on-investment [FN18]. Reunification [HLS^{+}14]. revamped [TWH^{+}14]. reveal [BLFS18, DDR^{+}14, FRP18, GKKW10, MSV^{+}11, Vav10, WBB^{+}21]. revealed [SLB^{+}18]. Revealing [Ju19]. reveals [GLR^{+}19, HFMM^{+}12, MGB13, RWH^{+}14, RMLTCQ^{+}19]. reverse [Bai18, DDR^{+}14, ZSG^{+}18]. reverse-learning [DDR^{+}14]. reverse-time [ZSG^{+}18]. Review [An12, FCLFB18, Jør14a, Kat11, ZWF18, AVG14, BSMA18, BU10, Cam16, CCF13b, CGP12, EH15, EH19, FKWZ19, GMC^{+}19, GBD^{+}10, GS19c, HF14, Hu16, JBV^{+}15, Jør10b, LTL^{+}19, LBH^{+}15, LWL^{+}17a, LHH12, OMH12, PMB^{+}19, Pér18, PFR15, SW11, SFF^{+}11, TFH12, TAGC^{+}16, VG19, Wal17, YLZ15, ZMS19]. Reviewing [KR10b]. revised [BU16a, Sci11a]. revisited [KOBK13, Log13a, PM16]. Revisiting [vVBHZ11]. Rhamdia [dABLBAdOR17]. rhino [GKL18, KdVRdW17, LR16]. Rhipicephalus [WTG^{+}16]. rhizodeposits [GA15]. rhizosphere [GA15, NP15, NPH^{+}17]. Rhizospheric [NP15].

Rhode [BLCPB11]. Rhombomys [RMP^{+}18]. rhythmic [MRB14]. Rhyzopertha [YWA^{+}17]. RI [BN10]. Ria [BPGSS^{+}18, DVD^{+}14]. Rica [FAC^{+}19a]. rice [COPP13, CBB^{+}10, DS12, SNS^{+}12, SVL12, SDSH15, SBC16b, TZS^{+}14, XBvdV^{+}14]. rice-based [SVL12]. rich [GWR11, NFU19, SRT16, THO10b]. richness [ACE^{+}13, HCP10, MPG18, SvWP^{+}15, WRMB18]. ricinus [HMA10a, LVLS14]. Ricker [NS15]. Rico [HSVL12]. Ridley [KWGG19]. Rift [TW18]. right [Pat14a, Pat16b, Ula14]. rigor [Lin10, Lin16]. Rim [LLP17]. ring [EFM^{+}10, MK12, VAO^{+}17]. ring-shaped [VAO^{+}17]. ringed [LAK^{+}18]. ringens [XCR^{+}13]. riparian [WZGH18, YLZ15, ZLJM17b, ZLJM17a]. rise [BME14, CAMCK^{+}12, CWMC^{+}12, DVD^{+}14, FPWG14, KR10a, KKK^{+}19, LB17, LKAL^{+}13, RSC12, RSC13, WCL^{+}17]. Risk [MMVAHMC16, RLHD11, TLL^{+}17, WB16, AWF10, BCC^{+}11, BSW18, BCV13, BLD10b, BBL11a, BCD12, dABLBAdOR17, CHvdW^{+}10, CZW10, CCF13b, CAY^{+}10, CMGF14, DPBS10, FtHvdB^{+}14, GHWG^{+}14, GDT^{+}14, GAG^{+}12, GFP^{+}16, GT14, HA18, HYG14, HL19, IPSH14, KDK^{+}10, KGM12, MPF^{+}14, MDKK15, MHRB19, MYJ^{+}14, NL12, PAB^{+}13b, PKW^{+}10, PLHG14, QCB^{+}11, RWS^{+}16, RGH^{+}13, RCBPMN10, SCB13, Sol19, SGWV15, TMPF18, TZD^{+}15, XWW^{+}11, XPY^{+}14, ZLLZ13, ZWA19]. Risk-based [RLHD11, NL12]. risk-growth [BBL11a]. risks [BG15, LL18b]. Risky [FTTV^{+}18]. RITY [OFH^{+}19]. rival [MTA19]. River [BS18, BCB^{+}11, CLD^{+}17, Con11, Con12, CSFB17, FWRZ10, HZSH17, HSTS19, HB14, JJYM19, KLS17, LLM^{+}13, LC11, yLWhMhG11, MSB^{+}17, MLH^{+}17, MSHC13, NWAR11, RZF^{+}13, RML^{+}13, SMC11, TJYZ11, WLDL12, WDL^{+}15, WPB14, WFSH15, WZG15, XLS11, XCL^{+}11, YYL13, Yan11, ZLZ^{+}17a, ALC19, AD12, BVP^{+}13, BGG16a, BS10c, CCL^{+}13, ER10, FSN^{+}15, FB13, GLC^{+}18, GSPSP15, JMW11, JBW^{+}17, JA15b, LHL^{+}17, LBK^{+}13, LDC^{+}16, Lin15b, MLY^{+}19, MSO10, NWAR12, NiSK21, PGKG12, RHW17, SLY^{+}17, SLB^{+}18, SGRAdJ17, TMT^{+}18, VPC^{+}14, WWC^{+}18, YLC13, ZSG12, dMLM^{+}17, CTNH13, DWNN10, DVF^{+}19, FS14a, FCH^{+}17, FBB^{+}18, GSPSP15, GFP^{+}16, GWZ^{+}15, GYL^{+}13, HGBA^{+}13, IGR^{+}18, KSSA13, KHJ^{+}19, MMRS^{+}19, MCP16, PSBH16, RMLTCQ^{+}19, SQWS19, SSM^{+}13, WCL^{+}17, WZGH18, WRSL11, WhFxGZ16, WAGM17, YXD^{+}18, ZF11, ZZT13, ZSG^{+}18]. river-lake [ALC19]. river/nursery [FB13]. Riverine [KPM^{+}14, GHH^{+}15, JD18, JBW^{+}17, LRM^{+}14, OC15, SXNL18, SQWS19, Sin17]. rivers [ARTBB14, FCS13, FLB^{+}15, GJM^{+}13, RAF17, WB16]. roach [ODK16]. road [BFR17, dÁGP14, FN18, FAPM12, FCA18, GJB^{+}11, PGATS13, vSGR16]. Roads [SW19, CHNdÁP18]. roadside [GJB^{+}11]. roaming [YRGH16]. Robben [WSW^{+}16, WCS^{+}14]. Robinia [BRKP16]. Robust [HDL^{+}16, HWH^{+}18, PP10, BCV13, CMEP15, FIG^{+}10, GB16a, HCK18, KB10, OGOWW16, TB12b, dBGS^{+}14]. robust-to-fragile [BCV13].

Robustness [GB16c, ZCY^{+}19, AG11, CR12a, CBA10, HHvO^{+}17, JJMW14, MSFR15, PBO19, Phi15, SBVA^{+}13]. robustus [KCAP17]. Rock [RE10, LMBL12, NSIiT17, ZPNM13]. Rock-Paper-Scissor [RE10]. rock-paper-scissors [NSIiT17]. rockfish [GC11, KHW^{+}16, ROSC15]. Rocky [CL12b, MJL^{+}13, MJL^{+}14, MZLCSZdP11, WLR^{+}11, FSS^{+}17]. rodent [BB18]. Roem [DSS11]. role [AMG19, ADL^{+}10, ASS14, BBJV^{+}11, BDTV11, BN10, CJC11, DCMK11, DCCS17, DJH10, DRS^{+}16, FAC^{+}19a, GRYM11, GPZB18, GWF^{+}18, HHUD14, JDGJ13, KCF^{+}13, KEM^{+}16, LGMLD19, LZF^{+}14, LBdA^{+}19, MRM16, MSMM11, MS16, MWD^{+}15, NOOM^{+}16, PvVC15, QFG^{+}13, RAdA12, RE10, SFvdM^{+}17, SMG^{+}12, SU15, SGBG19, SPP18, WB17, WJvdL^{+}13]. roller [KSS11]. Romaine [MSHC13]. Romanian [PGR^{+}16]. Rome [MSS^{+}19]. Rømø [FSB13]. ROMS [IFC14]. ROMUL [LKKK^{+}14, LKP13]. Romul Hum [CKS^{+}17, CSS^{+}17, KCB^{+}17]. roof [HJ10a]. Rookery [ZW10]. root [BDL^{+}14, GPK^{+}13, KJS13, MCR^{+}15, MJÁ^{+}12, NP15, PBB^{+}14, TBL^{+}15]. root-induced [NP15]. roots [LZF^{+}14, NP15, NPH^{+}17, SLMS14]. roseate [GQNMR18]. rosenbergii [CLW^{+}19]. Ross [ACE^{+}17, PLS^{+}16]. rotational [LYD^{+}17]. Rothc [MJF19]. rough [XLR11]. row [Pie10b, Pie10a]. row-crop [Pie10b, Pie10a]. roxhburghii [KKK^{+}19]. rubber [KMM^{+}13]. rubens [KPT11, KMP14]. rubra [LJWN14]. rubrum [GBP^{+}16]. Ruditapes [VZF^{+}11]. ruffe [DGKSM13]. rule [CC10, Cho13, JKS^{+}10, JD10, MHL^{+}19, PME^{+}16]. rule-based [CC10, JKS^{+}10]. rule-implicit [Cho13]. rules [CVV^{+}13, CT12, GEPCB13, MHY^{+}15, PAEE19, RP10, Til14]. rumen [GBW^{+}15]. ruminants [KP10, RWvDM15]. run [EC10, LZW^{+}17, RDD^{+}18]. run-up [LZW^{+}17]. Running [SA16]. runoff [BT16, GWHL15, JM15, KW18, LPX^{+}12, LCP15, SHC^{+}15, TBA19, TYEH10, WSW^{+}13]. rural [LLW^{+}15, MJLC14, RLHD11, TQZ12]. Russia [BKMK10, PVBC14, PSKO16, SKMB11, SSK14]. rust [Hal11]. Rutilus [ODK16].

S [JDT^{+}17]. S. [FBDR17]. s.l [VVH12]. sabrinus [TDF^{+}19]. Saccha [PGG14]. saccharina [PvVC13]. Saccharomyces [PGG14]. Sacramento [WZGH18]. SAFE [GBP^{+}10, dJGM^{+}18]. safety [hCjHMW18]. sagebrush [AHF^{+}13, SLB14]. Sahel [LCH12]. Sahelian [SGBD10]. SAILHFlood [BBDB13]. Saimaa [LAK^{+}18]. salamander [WSBP13]. salamanders [GPCG15]. salar [HSHF^{+}13, LCD^{+}16, MSHC13]. salinity [AML13, ADPS10, DWP16, FP17, HR11, LLC^{+}17, MML^{+}11, PICR13]. salinity-prone [ADPS10]. Salix [SP10]. SALMO [CZR^{+}14, CSR15, HSHF^{+}13, LCD^{+}16, MSHC13, MMMGLN^{+}18]. salmon [AHS^{+}17, BCTL16, CF17, HHUD14, HSHF^{+}13, Jag14, LCD^{+}16, MGCL17, MSHC13, Mur13, PP12, ZBMC14, ZLD^{+}19]. salmonid [BS10c, FBB^{+}18, LLB^{+}19, RLHD11]. Salpa [HPGM18]. salt [DVD^{+}14, GGZ^{+}16, GJB^{+}11, XLS11]. SALUS [DBJ13]. salvage [SGG14]. Same [Tie16]. Sample [HHD12, BB18, KCP^{+}15, SA13, SHS11, YAGMM^{+}14]. samples [AYASR^{+}17, GGV10]. sampling [AG11, BBRJ15, BPB^{+}18, BDG^{+}14, BOGA14, CGS^{+}17, DG16, EAB11, FN15, FMF12, LW17, MSF18, MSV^{+}11, MDV^{+}12, PJ17, PWGS^{+}17]. sampling-related [MSF18]. sampling/resampling [DG16]. Samsara2 [CLL^{+}15]. Samsø [JN15b]. San [DWP16, GSJ^{+}10, MLS^{+}11]. sand [LSL10]. Sandeel [GCvDM12, GCM^{+}13]. Sandland [HYL^{+}17]. sandy [LBD10, LDO^{+}18]. Sanggou [ZSF^{+}12]. Sanitary [OJ10]. Santa [MJK^{+}17]. Sap [CSN^{+}19]. Sapajus [KCAP17]. saplings [CHG^{+}15]. SAR [Con12, Con11]. sardine [CRA^{+}11, GPH17, KKH^{+}17, PME^{+}16, WJS^{+}18]. Sarg [KPT11]. Sargent [KMP14]. sasakii [KL10]. Saskatchewan [ABK^{+}16, DBG14b, DBG14a]. satellite [BLRN15, BNC^{+}18, KHS^{+}19, LLT^{+}14, MCMT^{+}16, PK14, PSS16, RPGSCF16]. satellite-based [LLT^{+}14]. savanna [AD13, Bae18, GLRT16, STJ15, Tie16, BFKS16]. savannas [AWWD16, BGP11, GWF^{+}18, LS10, SJGB16]. saving [FG11]. say [Sci11b]. Scale [FC11, KFWM19, MWK19, QCLW18, Abe10, ABD^{+}18, BP15, BP19, BKN21, BGD^{+}16, BlMG^{+}11, BA14b, BPM^{+}15, Cha12, CCF13b, CY13, Cho13, CMCC15, DLW^{+}17, EKB18, FMA^{+}12, FHAG12, FtHvdB^{+}14, FCLFB18, FYWA12, GLC^{+}18, GLA11, GGV10, Gav15, GPA^{+}13, GS19b, GNB^{+}14, GLY^{+}15, HCM^{+}15, HZZ^{+}18, HCP10, HFMM^{+}12, IGR^{+}18, IGBSASSJ16, JVY19, JMSM12, KMRW16, KD13, KHHN17, Kos17, KSG^{+}14, KUK11, LBR16, LF11, LLTQ14, LLTQ15, LPS^{+}10, LHW^{+}15, LWY^{+}11, LCP14, LSM19, MBLC14, MGR^{+}18, MSV12, MDOJ14, MTJF11b, MGB13, MGSB11, NLM^{+}19, NPH^{+}17, PGR^{+}16, RKS^{+}10, RAF17, RPA^{+}18, RDR^{+}17, RNE^{+}12, SHC^{+}15, SGA^{+}16, SRSS12, SWZ^{+}12, SSKJ18, SBC^{+}11, SKL^{+}13, SGDP19, TZS^{+}14, UB12, VSF10, YLF^{+}15, ZLB14, ZZT13, dMLM^{+}17, MBRR13]. Scale-dependent [FC11]. Scaled [CVS^{+}13, Nak17, WFS^{+}19]. Scaled-dependence [Nak17]. Scales [WØ15, AHNK13, CRA^{+}11, CXW19, Cra10, ER10, HWL17, KDK^{+}10, KVU^{+}10, KGYG15, MSB^{+}17, MGS16, PAT17, SG11b]. Scaling [vdKGV11, HWS11, Hui11, KGS11, KP10, MvdWU^{+}16, TL10, WSZT10]. Scaling-up [vdKGV11]. scalings [TKW17]. scallop [KTM^{+}16]. Scandinavia [JA15b]. Scandinavian [JSW^{+}13]. scanning [HS11, KHS^{+}19, MdMS^{+}19]. Scaphirhynchus [DHC16]. scapularis [HM18a, HM19]. scarcely [VRA^{+}17]. scarcity [LPF11, UR10]. scatter [PGW19]. scavenger [DMBS11]. scavengers [CMSPJ11]. sceleratus [CVM^{+}18]. scenario [BWRA16, CZR^{+}14, CMGF14, HHL17, LVLS14, MTJF11b, PAS^{+}10, SBRD10, VPC11]. Scenarios [LCPB^{+}10, BFJZ^{+}11, BGD^{+}15, CRTS15, DAF^{+}11, FL11, GOSH^{+}12, HV15, IKL13, JFS^{+}12, JDT^{+}17, KMKL19, LDM15a, MBD^{+}12, MTPC13, MRM^{+}17, MSM^{+}16, OJ10, PSM^{+}12, REB^{+}16, SLSS15, SKMB11, SWR^{+}16, WB14, XTPT11, YHK^{+}19, YXD^{+}18, YFC^{+}11, dFMdSA18]. scenes [RJG^{+}13, SYY^{+}19]. scheduling [MRP14]. scheme [HWS11, HFMM^{+}12, ZCYL15]. schemes [EST^{+}19, LJ10, TDH12]. Schisandra [LGDW12]. Schizas [Pat10]. schools [DHHB15, RKRB16]. Schwabl [ZJG10]. science [Cam16, DLW^{+}17, Pat14a, Pat16b, RCS^{+}19, Ula14, HhLY^{+}19, Pau14, Jør14b]. scientific [BME14, CMV^{+}16]. scientist [CBC^{+}16]. Scissor [RE10]. scissors [NSIiT17]. Scleractinian [GBR13b, ENM11]. Scomber [LCC^{+}14]. scope [CSL^{+}12]. Scots [ACDC17, GVV^{+}16, LM19, Log19a, UDP17, WJK11]. Scottish [KGM12, Mur13]. scouting [BGK^{+}16]. Screening [BLBB13]. scripta [BGA17]. scrublands [MRN12]. SD [TLLX19, WN17, ZLZ^{+}17a]. SD-MaxEnt-CA [TLLX19]. SE [ROSC15, HPO^{+}15, OCB^{+}13, OBC^{+}21]. SE-Queensland [HPO^{+}15]. Sea [BH13a, BH13b, DSW^{+}17, KR10a, MSCH14a, SZSV12b, SW19, CF14, CAMCK^{+}12, CWMC^{+}12, DVD^{+}14, DBP16, FS14a, FP13, FPWG14, KWGG19, LKAL^{+}13, Mur11, MS16, NWAR11, NWAR12, PP19, PRH17, PBD^{+}17, RPC^{+}17, RSC12, RSC13, RCP17, TVT12, VVH12, WFA10, WCL^{+}17, WHRR15, WFN^{+}16, AALP18, ACE^{+}17, BDV^{+}19, BH13a, BH13b, BSO13, CAL13, CWA^{+}15, CVM^{+}18, CQS15, DOB^{+}15, DBVS19, DSW^{+}17, GS10, GCvDM12, GCM^{+}13, Hei11, HSM^{+}12, LY16, LAB^{+}10, LCC^{+}14, MML^{+}11, MMM^{+}16, MSGG14, MSCH14b, PCTW17, PPB^{+}18, PLS^{+}16, PGV10, RMHE15, RPP^{+}15, SELB10, SPM^{+}18, SGH16, SKM^{+}16, TPL^{+}15, WBS13, WB14, WYPF15, ZYZ17]. sea-cage [RPC^{+}17]. sea-level [CAMCK^{+}12, CWMC^{+}12, LKAL^{+}13, RSC12, RSC13, WCL^{+}17]. Sea/Baltic [BH13b]. seabass [PAGCV^{+}10]. seabird [LJ13]. seabirds [SSS^{+}11]. seabream [PAGCV^{+}10]. seagrass [ALeSD17, BAB^{+}16, BGD^{+}15, Hos11]. Seal [CF14, BTB17]. seals [LAK^{+}18, ZKSF17]. SEAMANCORE [MLAT^{+}18]. Search [Jør14a, BLH10, FCLFB18]. seas [HB10b, BSH17]. season [CMLY19, LCFD19, LAC10, YSA^{+}14]. Seasonal [WHNR19, ZYZ17, ACC^{+}10, BVJE18, DP15, EFM^{+}10, ER10, FS14a, FTK^{+}14, KH11, KK15b, LBR16, LDP14, LCZ11, LRB12, MCN16, Nak17, NiSK21, PICR13, RML^{+}15, SAT^{+}15, SLZ^{+}13, SM14, TSB^{+}18, TWH^{+}14, WDL^{+}15, WJK11, XTPT11, ZFG13]. Seasonality [KMMS17, AD13]. seasonally [DTC^{+}10]. seasons [ZCYL15]. SEBAL [LLM^{+}13]. Sebastes [GC11]. Second [Nie16, MH11]. Secondary [CV19, BH17, BPC^{+}17, HSVL12, KM12b, KGD14, MBBA17, PR11, SRHL11, SWS16]. secret [SKPHW16]. Section [yLWhMhG11, Jør10a]. Sector [SKTG11, GSA^{+}19, KGM12, LZJ^{+}19]. sectoral [LYCZ11, NLM^{+}19, ZLF^{+}15]. security [HLW15, LSZC10]. sedges [RML^{+}15]. sediment [HPG^{+}10, KD13, OdlFDP15, RML^{+}13, ZSJ15]. sedimentary [WJL^{+}13]. sediments [BCC^{+}11, PJ16, VMSM10, VRC17, VRL18]. see [Bev15]. seed [BFNH14, BCL^{+}14, CV19, DF13, GDC12, GTM^{+}19, HLS^{+}14, HKN17, KR14, MOMC17, MLMH18, PDG^{+}14, SMKM19, TKN14, TWL11, WXB10, vLTdBP16]. seed-eating [DF13]. seedbank [KvTL^{+}18]. seeding [TB13, VAB^{+}18]. seedling [qHLlZ^{+}16, HKN17, Oga11]. seedlings [JAH^{+}10, MCS^{+}18]. seeds [HRWT12, MBBA17]. Seeking [SGM^{+}14]. Segment [ZLJM17b, ZLJM17a]. Segment-based [ZLJM17b, ZLJM17a]. segmented [BDL^{+}13]. segregating [KYvdPS14]. segregation [HGAT17]. seine [JBCÑ11, MROGVR12, DJMBD^{+}13, TRD^{+}15]. select [GSQ17, KKI^{+}14, SZSV13, SvWP^{+}15, ZLY16]. selected [Ana17, BJPB11, BBH^{+}17, BNP10, GANB11, SBF15, LYP^{+}11, LPA^{+}11]. Selecting [WKH19]. Selection [DFL^{+}16, MU17, BTM^{+}10, BPB^{+}18, BJJ12, Bil11, BBH^{+}17, ČTT^{+}12, GGV13, GGZ14, GCIB10, HMD^{+}16, HAOS17, JTM^{+}17, JJKJ11, KBB^{+}11, LVB^{+}15, LPT^{+}18, Loe12, MAM^{+}12, MU10, MRWN19, MGB13, Mur11, MFI13, NMS10, PBO19, PBD19, RRG^{+}15, SCHQA16, SSD^{+}17, SGDP19, TOHT14, TTL^{+}18, WBG^{+}14, WKF^{+}17]. selections [CDR^{+}13]. selective [Gat18, KMRW16, ZAW^{+}18, vPTOW18]. selenium [GV11]. Self [DBH^{+}17, GDC^{+}19, LCZ11, PLB16, RPPW16, DGKSM13, LLC11, MHL^{+}19, Oga17, Oga18, PGKG12, POB18, SLZ^{+}13, SSM^{+}13]. self-organisation [SLZ^{+}13]. Self-organising [PLB16, DGKSM13]. Self-organization [GDC^{+}19, LCZ11]. Self-organized [DBH^{+}17]. Self-organizing [RPPW16, LLC11, PGKG12, SSM^{+}13]. self-thinning [MHL^{+}19, Oga17, Oga18, POB18]. Selkeh [SZSV12a]. Semantics [UDNSLCG16]. semelparous [SP12]. Semi [GWC^{+}10, AHF^{+}13, BC13, Ber13, CDK^{+}17, Che17, CV19, FS18, GLRT16, GWF^{+}18, HCZ^{+}17, HYL^{+}17, KF16, LHP16, MPL^{+}16, MLH^{+}17, OES16, PLL^{+}19, SJGB16, SRP18, WYS12, aGSCMF15, vLTdBP16]. semi-arid [AHF^{+}13, BC13, Ber13, Che17, CV19, FS18, GLRT16, GWF^{+}18, HYL^{+}17, KF16, LHP16, OES16, PLL^{+}19, SJGB16, WYS12, aGSCMF15, vLTdBP16]. Semi-automatic [GWC^{+}10]. semi-closed [HCZ^{+}17]. semi-empirical [MPL^{+}16]. semi-humid [MLH^{+}17]. semi-individual [CDK^{+}17]. semi-open [SRP18]. seminars [LP17]. semiotics [Nie16]. SEMoLa [DR14]. Senegal [BCB^{+}11, BVVS15, CGEdM12]. seniculus [WFDDR10]. sensed [BRR10, BOW^{+}12, DBR12, KMRW16, VFFP16, VALC12]. sensing [BQF^{+}18, CZ11, COK^{+}11, CAY^{+}10, CKMD14, EAJG16, GYL^{+}15, JSdSdL^{+}18, LSS^{+}15, LAB^{+}10, LCL^{+}13, LZZ^{+}14, LLH^{+}17, MPT^{+}11, MCBM12, MVC^{+}17, MMK15, PSSH11, SX11, VHM^{+}17, XLR11, ZCS13]. sensitive [EC10, HSLC^{+}19, KHJ^{+}19, YKvWK16]. sensitivities [JWX18, Löp11]. Sensitivity [CDT10, FAG10, GLR^{+}19, GCCB18, HHvO^{+}17, JSL^{+}10, KTGH19, KKP^{+}17, KDHC^{+}18, LCH12, LHB^{+}18, MBW14, MJL^{+}13, MJL^{+}14, MCR^{+}15, MGS16, MLH^{+}17, NGWZ16, PC16, RMTM^{+}16, RPPW16, RWP11, SMC11, SELB10, SPM^{+}18, Tho10a, UDRA13, YDD^{+}13, ZBS14, ZAW^{+}18, AADA14, BLLJ17, BS17, CCW^{+}15a, CCW^{+}15b, COBF13, CL10, CAMCK^{+}12, CCC12, CBB^{+}10, DAA^{+}12, DFM12b, DBJ13, GB16a, GGVB11, HGPF15, HBL18, JLZ^{+}18, LKMCW14, LJLO11, Loe14, LB18a, MXC13, MTJF11a, MDKK15, MSCH14a, MSCH14b, MSFR15, OOFVBC16, OCCF17, PMCKH13, PP12, QCLW18, QFG^{+}13, RHvdTS14, RAF17, SLLW15, SBPZ12, SKL^{+}13, SBA^{+}13, TPL^{+}15, UKI^{+}11, VJR^{+}16, VPM^{+}18, WMF^{+}13, WGA^{+}13, YZG16, ZSF^{+}12]. sensor [JGX13]. Sensory [LGMLD19]. Sentinel [MAO19]. Sentinel-2 [MAO19]. separate [LKLK15, PM16, YqWB^{+}15]. separating [ZCYL15]. September [Ano10q, Ano11-31, Ano11-32, Ano12-35, Ano12-27, Ano13-30, Ano13-28, Ano14u, Ano14h, Ano15o, Ano15i, Ano16g, Ano16s, Ano17-34, Ano17-33, Ano18-29, Ano18-32, Ano19d, Ano19h, GORJ12]. Septicaemia [KGM12]. sequence [TL10]. sequences [GLT10]. sequential [DG16, DPBS10]. sequestration [ABK^{+}16, AML13, ACN14, GHP12, MPM^{+}15, NJM11, OQH^{+}13, UKI^{+}11]. Serbia [RRR^{+}10, SSM^{+}13]. Serengeti [War10]. Seriatopora [KvOP10, MATS14]. Series [WHC^{+}13, ACN14, BAJ^{+}11, DSK^{+}11, KA17, RWP11, SHBG10, VPS10, WPB14]. serious [YL17]. service [DML15, DMBS11, LSJ^{+}18, LSJ^{+}19, LLZ12, SDY^{+}15, SFB^{+}15]. Services [GAP^{+}19, LYL^{+}17, MSS^{+}19, AAJB21, BHD16, BMU10, BFM11, CLA^{+}12, DYU^{+}12, EBA^{+}19, ESP^{+}17, GAG^{+}12, GFC15, HRSKW15, HV15, HF14, JBV^{+}15, JMB^{+}11, JBW^{+}17, LC13, MGR^{+}18, NPP^{+}18, OWMK10, PCB11, RDH^{+}13, TLL^{+}15, TAGC^{+}16, VLZ^{+}19, VVA^{+}15, WO14, ZBAY14]. SES [CF17]. sessile [VCE17]. sessions [LPA^{+}11]. Set [BHD17, MCB14, BMF^{+}11a, BMF^{+}11b, CPAG15, JBCÑ11, KYvdPS14, KFvSH15, LJ10, Lin10, Lin16, vVHZHvD13]. set-aside [LJ10]. sets [AHS^{+}17, HL19, MU17, MCPB13, PLHG14, SFL^{+}19, TBJ^{+}15, XLR11]. setting [DTS^{+}12, LCP15, RLB^{+}12, SB15]. settlement [TQZ12, vSGR16]. settling [ZLG16]. seven [San17, WPB14]. seven-compartment [WPB14]. several [HKVR13]. severe [TGB^{+}17]. severity [FMS13]. sewage [AJDS16]. sewage-fed [AJDS16]. Sex [BEPC15, FML^{+}17, HTL^{+}14, RAdA12, SK10, TKP14, TVI^{+}13, TMC^{+}16]. sex-ratio [TVI^{+}13, TMC^{+}16]. sexual [IWW12]. sexually [NdA16]. Shaanxi [ZZLZ14]. shad [HSCS14]. shadows [BFNH14]. shallow [BAB^{+}16, BM16, BN10, CVPH^{+}14, CdMMF16, CZR^{+}14, CTP^{+}17, DH12, FAC^{+}19a, GJÁ^{+}10, GYL^{+}13, HB10b, HPRNP12, JSL^{+}10, JHH^{+}12, KHL^{+}16, KvTL^{+}18, LHZ10, LCY^{+}13, LLZ^{+}15, MH10, SLW^{+}14, TSB^{+}18, WFW^{+}17, YZG16]. shallow-water [BAB^{+}16]. Shandong [ZZF15]. Shannon [PVB16, RHQ^{+}16]. shape [HM10, KOBK13, LJ15a, MAM^{+}10, NWAR12, POB18, San11]. shaped [VAO^{+}17, YZ18]. shapes [MLK^{+}11, NWAR11]. shaping [DRS^{+}16, TWH^{+}14]. share [SFM^{+}18]. Shared [NTK14, DS12]. shares [MTGD17]. Sharing [JSW^{+}13, SZ14]. sharks [BTBM17]. sharp [BCV13]. shatter [DKM^{+}14]. shattercane [WTL17]. shed [GO15a]. sheep [LJ14]. Shelf [DA19, CR12a, CQS15, DJS10, ER10, GC14, LBSB^{+}14, PFA^{+}16, ZRCNLI^{+}15]. SHELL [IFC14]. SHELL-E [IFC14]. shellfish [BLCPB11, GBB^{+}11, HCZ^{+}17, IFC14, RRHH10, ZZL^{+}19]. shelter [BBL11a]. shelterbelts [ABK^{+}16]. sheltered [RRHH10]. shift [BPB14, BSHG16, HH12, LPT^{+}18, MPP^{+}15, MCP16, OGG^{+}16, SK15, TZD^{+}15, XTPT11]. shifting [LQW^{+}19, MJF19, WFA^{+}13]. Shifts [DWWS15, BGD^{+}16, BMD18, CpW16, CJM^{+}12, DAF^{+}11, DVCM^{+}10, FBKS17, HT16, LBN^{+}18, MLC^{+}16, MLK^{+}11, MOMC17, MRM^{+}17, NA11, RH17, SS16, SKK^{+}13, ZW12]. shipping [CLD^{+}17]. shoot [Hos11]. shore [BLFS18, WLZC11]. shorebird [CWMC^{+}12]. Short [GT14, MCG19, CRVS13, DR17, LA11, LAB^{+}10, OBC^{+}21, RTE^{+}15, Str15a, Str15b, ZZT13]. Short-term [MCG19, CRVS13, DR17, LAB^{+}10, OBC^{+}21, RTE^{+}15]. short-time-scale [ZZT13]. Should [Kos17, RA13, SM10, TCGC16]. shrimp [DLH^{+}14, DBS18, LMGW16, MHRB19]. shrimps [ADL^{+}10]. shrinking [KPTC14]. shrub [GdTCFG10, UDNSLC^{+}18]. shrubland [AW16, VALC12]. shrubs [PFSF14]. SIAT [VKdW^{+}10]. Sib [GVGC16]. SIBBORK [BS16b]. Siberia [KWJH16]. Siberian [THJ^{+}12, TWS^{+}11, XKI^{+}11]. Sichuan [LML17]. sidedness [WJL14]. Sidney [KB10]. Sieb [Oga11]. Siena [MJP^{+}12]. sigmoid [SCHGM16]. sigmoidal [TT17a]. signals [Bae18, OFGH^{+}19]. signature [VRL14]. signatures [BPB^{+}18]. significant [PVB16]. silica [OC15]. silico [Hel10]. Silver [SLB^{+}18, dABLBAdOR17, NFU19]. SIM [KB18]. SIMBA [Och10]. SIMBA-POP [Och10]. similarities [GLSA^{+}11]. similarity [VPC11]. Simple [AJDS16, HMS^{+}11, JNHF18, MHY^{+}15, ROG^{+}16, RW17, RP10, AFC12, BVC^{+}18, BP16, BP17, Ber13, Con14, DR14, DBJ13, EZ14, Flo14, FNR11, GORJ12, GC14, IKH^{+}11, MLW11, MDW^{+}12, MMC11, Mur11, MS16, PW10a, PBB^{+}14, Pea15, RML^{+}15, RRJVA13, RML^{+}13, RWP11, SHT^{+}21, Sim11, SWGC16, TTCH^{+}14, TM12, WCF^{+}16, WM14, WK11, WWYO17, ZLF^{+}11, ZKR^{+}13]. Simplified [JZ12, HMP^{+}19b]. simulate [AWL^{+}13, BPR^{+}11, BOF^{+}16, CVV^{+}13, CLLB12, CPS^{+}14, KF16, LKKK^{+}14, LDM15b, MBD^{+}12, SHC^{+}15, SRSS12, SBC16b, SMBGH^{+}19, TRE^{+}12, VAB^{+}18, WO14, WZD^{+}14, YDZ^{+}21]. Simulated [MGWW13, PAT17, WTG^{+}16, WGTH16, AVL^{+}11, CSS^{+}19, CPE13, DSS11, FL11, GLS^{+}17, LHLW15, MBW14, NZL13, PBFS14, RV13, SPM^{+}18, SHM^{+}14, WHRR15, WJ15]. simulates [RML^{+}15]. Simulating [ABB^{+}10, BFH14, Bre12, BWML15, CNH^{+}11, CAMCK^{+}12, CHK^{+}12, DDM^{+}17, DCMK11, DPPI14, DHHB15, DMR13, DH18, DHS^{+}16, DWP11, DRS^{+}16, FB13, FPG^{+}10, GE17, GBL^{+}13, GNB^{+}14, HJ17, HFH^{+}19, HWE^{+}18, IMW^{+}17, KRL^{+}16, KVU^{+}10, LMGW16, LLT^{+}17, NJM11, PHvB^{+}18, PGATS13, PAGCV^{+}10, RWH^{+}14, RBGB16, SRV^{+}12, TOQW11, TS11, WR17, WSC^{+}19, YLY^{+}18, YDT^{+}13, ZZJ^{+}16, ANMJ12, AAJB21, AD18, BMMC21, CLH17, FKH^{+}15, FMA^{+}12, KDHC^{+}18, LHZS10, LTL^{+}13, LYD^{+}17, LZZW17, MBHW12, MJÁ^{+}12, NW18, PKD11, RPGA15, SBDS13, SQWS19, Sim11, SMG^{+}12, SNZ^{+}19, TTCH^{+}14, TLLX19, THOA10, WPZ^{+}11, WHB12, WMF^{+}13, WRM^{+}11, XZC^{+}17, XSN^{+}18, ZZB^{+}18, ZHLZ17, vSGR16]. Simulation [BLBB13, CC12, FH11, FKDG14, GQRC^{+}17, HJ10a, JJKJ11, KWGG19, KVS17, KMM^{+}13, KVP16, LB15, MGMR15, MNEG11, MCBM12, MYHN17, Mur12, NBD^{+}13, NGWZ16, OSS13, SH18, SYY^{+}19, WGT12, dWcWbZ15, YKCN10, YYH19, ZJM^{+}13, ABR12a, AHS^{+}16, BFA11, BHD16, BHD17, BPB^{+}18, BEBA19, BS17, CZC^{+}15, CMM^{+}10, CZR^{+}14, CH14, CDT10, CBA12, DFL^{+}16, DPBK10, DPvBvA11, DDB^{+}12, DA17, DVL11, DFF17, DPBS10, GH13, GLC^{+}18, GOSH^{+}12, GLS10b, GdC14, HZSH17, HHL17, HBH^{+}12, HVCP16, HOKH18, HHF^{+}10, JBB^{+}12, KWS^{+}13, KK10, KMKL19, KPT11, KWGW19, KKK^{+}19, LJLC15, LBR^{+}17, LTL^{+}19, LDS14, LCD^{+}16, LBW18, LLH^{+}17, LL18a, yLWhMhG11, LM16, LCML11, MV11, MD16, MKSL^{+}17, MHG18, MLAT^{+}18, MCG19, MP17, MBA^{+}19, MGW^{+}21, NMT^{+}19, Och10, OLG^{+}13, PWBG16, PAEE19, PD12, PGR^{+}16, RMP^{+}18, RGHS11, RAB^{+}10, RWvDM15, SFvB10]. simulation [SLZ^{+}18, SGB10, wTA12, TT12, TNA^{+}17, TDH12, TYEH10, VC17b, Wal17, WDL^{+}15, WZGH18, WSE10, WP14, YHK^{+}19, ZHC^{+}16, ZYZ17, ZWJ^{+}17]. simulation-based [BPB^{+}18, GOSH^{+}12, ZHC^{+}16]. simulation-to-optimization [BHD16]. Simulations [DDR^{+}14, KCAP17, LPX^{+}12, MvdWU^{+}16, SLD^{+}19, YLF^{+}15, ABK^{+}16, BVGR^{+}17, Bai18, CMSR19, CGG^{+}18, HAOS17, HT14, KC18, KPS17, KTM^{+}16, LJR14, LBM16, MSGG14, MPT^{+}11, MDB19, MPR13, NMdR16, OBC^{+}21, Qia12, SBC13, SRP18, SBC^{+}15, SPG11, WGD^{+}17, XBvdV^{+}14, YqWB^{+}15, ZW10, vVHZHvD13]. Simulator [FMA^{+}12, GJM^{+}13, PMS^{+}17, TSN17, TKH17, ZZB^{+}18]. simulators [Con14]. simultaneously [ACE^{+}13]. since [Ano17a, LCBNP19]. sinensis [PFP^{+}16]. single [BCC^{+}14, Con14, dSCdA19, CRVS13, HHW11, KGD14, YHW^{+}12, dAdSC15]. sinicus [JSH15]. sink [BBC^{+}19, Loe12, SBA^{+}19]. Siping [ZLZS14]. Site [GS19b, AKvO^{+}10, AdMC^{+}13, CNI11, FWRZ10, HMS^{+}11, JJKJ11, KHW^{+}17, LHW^{+}15, MGMR15, MBG17, MGS16, PAS^{+}10, SPM^{+}18, SSD^{+}17, WS15, WWYO18, YZA^{+}18, YKvWK16]. Site- [GS19b]. site-scale [LHW^{+}15]. site-specific [WS15]. sites [BVVS15, BSCG10, FHG12, LXS^{+}16, YWY10, ZZB^{+}18]. Sitobion [DFF17]. Sitophilus [YWA^{+}17]. situ [LLKY18]. situations [GMP13, GBP^{+}10, LVV^{+}15]. six [CLL14, FCV^{+}12, PD15]. Size [BPN12, PCZ^{+}16, AAME11, Ban11, BSH^{+}10, CL12a, CRUF13, CWL^{+}16, CB15, CGM^{+}14, GdSHA10, GV11, GAHE11, GPM16c, HB15, Lec17, LLH^{+}13, LRB12, MGWW13, NLMdL10, POBH10, PDG^{+}14, RW17, RGHS11, RV15, San11, SXNL18, SCHGM16, SFHK12, SGH16, SP10, SHS11, TKW17, TVI^{+}13, TMC^{+}16, TJYZ11, TQZ12, TTCC^{+}19, YAGMM^{+}14, ZCR16, ZGF14, vPTOW18]. size-asymmetric [RW17]. size-based [TTCC^{+}19]. size-different [SXNL18]. Size-related [PCZ^{+}16]. size-selective [vPTOW18]. size-spectral [Ban11]. size-spectrum [GPM16c, ZCR16]. size-structured [AAME11, CRUF13]. sizes [BB18, HHD12, LCP15, MPR13, SA13]. skipjack [DFM12a, DFM12b]. Skylarks [MGB13]. SLAMM [WYPF15]. SLCD [SMSAN^{+}14]. slope [MBSF14, POB18]. sloped [CV19]. Sloping [ZZLZ14]. Slovenia [GMP13, OJ10]. slow [WGL^{+}12]. SLR [DVD^{+}14]. sludge [OTA16]. sludge-amended [OTA16]. Small [XWW^{+}11, XQZ^{+}13, ABB^{+}10, BFP^{+}18, CHR^{+}13, CSN^{+}19, CMCC15, EKB18, FAPM12, FCA18, GGV10, IGR^{+}18, IGBSASSJ16, KHHN17, KSG^{+}14, LZB17, MN12, RPP^{+}15, SGA^{+}16, SA13, WGL^{+}12, YDT^{+}13, ZJM^{+}13]. small-scale [CMCC15, EKB18, IGR^{+}18, IGBSASSJ16, KHHN17, SGA^{+}16]. small-world [ABB^{+}10]. smaller [NYH15]. smallmouth [LBW18]. smart [DSA^{+}18, Gia14]. SmartScape^{T M} [TAT^{+}16b]. Smoky [KPT11]. smolts [APWT15]. snails [BCP15]. snapping [AB13]. snow [DYMYGO19a, DYMYGO19b]. snowdrift [LPSC16]. Snowy [CAMCK^{+}12]. soaring [DDBK15]. Social [BGC10, BYR18, CWLC13, CLA^{+}12, CLD^{+}17, GMK^{+}19, GCCJ12, Ken17, LL17, LJ14, LZL^{+}18, MRM16, MGD^{+}12, PJ15, RGRMRR18, SGBD10, SKHH19, SvdR10, SXZ^{+}15, UAB^{+}11, VSB13, XZCL16]. social-ecological [CLD^{+}17, LL17, MRM16, SKHH19]. societal [CSF11, LG17]. societies [KBC^{+}12]. Society [Ano13u, LMY^{+}11, LYP^{+}11, LPA^{+}11, BCP16, HZSH17]. Socio [GPM^{+}16a, TBA19, BGMFES15, CPJ^{+}13, CK12, DK17, Fai10, Hal11, LA11, MRN12, SDSH15, TLLX19, WAM13, ZFZ^{+}14]. socio-ecologic [Fai10]. Socio-ecological [GPM^{+}16a, TBA19, BGMFES15, CPJ^{+}13, CK12, Hal11, MRN12, WAM13]. socio-economic [DK17, LA11, TLLX19, ZFZ^{+}14]. socio-political [SDSH15]. socioecological [BUB^{+}12, RAR15]. socioeconomic [RCBPMN10]. Software [MBRR13, BGK^{+}16, GdC14, LBA14, LBA15]. Soil [COBF13, FTYI13, LKLK15, OLG^{+}13, PBFS14, TDC^{+}11, vLTdBP16, Agu11, AAJB21, AYASR^{+}17, BVGR^{+}17, BP15, BP16, BP17, BP19, BFP^{+}18, BKMK10, BBH^{+}11, CKS^{+}17, CSS^{+}17, CR11, CR12b, DS21, DJA^{+}10, FBA^{+}11, GPK^{+}13, GBG^{+}15, GE17, Gra14, GVC10, HWS11, HMS^{+}11, HJS^{+}16, HWL17, HOŠ^{+}12, HHF^{+}10, IBD^{+}19, JL18, JAKJ10, KJMA11, KCB^{+}17, KJS13, LBZA19, LBE^{+}18, LYD^{+}17, LLT^{+}17, LPS^{+}10, LKP13, LPD14, MLB^{+}16, MBSW12, MBW14, MGS16, MAP^{+}13, MMC^{+}19, MJF19, Nei11, NBK^{+}11, OKS^{+}11, PAS^{+}10, PP16, Pie10b, Pie10a, PANL11, SDD^{+}13, SFvB10, SVL12, SBRD10, SCG^{+}16, TRE^{+}12, TGB^{+}17, VPM^{+}18, WC13a, WC13b, Wol11, WHC^{+}13, XLS11, YDD^{+}13, YYH19, ZLW^{+}15]. soil-core [BVGR^{+}17]. soil-plant-atmosphere [LPD14]. soils [BL15, DCMK11, FAS^{+}19, GPK^{+}13, GGVB11, GNB^{+}14, JAH^{+}10, LVPF15, LNB17, MOM^{+}11, NP15, OTA16, OKS^{+}11, OLG^{+}13, SSW^{+}18, SGKR10, TTCH^{+}14, TLL^{+}17, YDD^{+}13]. solandri [MROGVR12]. Solanum [KYvdPS14]. solar [AML13, JMSM12, LCS11, PSS16, SBDA16]. sole [KHW^{+}16]. solenopsis [FTK^{+}14]. solid [BNGP19, ZHCX15]. solids [MTPC13]. solitary [CCCP17]. soliton [ACKS13]. soltan [SC21]. Solution [WHC^{+}13, MEC14, PBP^{+}15, Qia12, RHS^{+}15, RHD^{+}16, SJTS13]. Solutions [Gia14, BOF^{+}16, GCCB18, LTL^{+}19, Mar17, MV17]. solved [BLAV18]. Solving [BDWP16, Bil11, KA17, HCK18, vDHH^{+}14]. Some [MJB^{+}11b, RHS^{+}15, RHD^{+}16, SGB10, VRL18]. SOMPROF [BBH^{+}11].

Songtao [YDZ^{+}21]. sorbed [LKKK^{+}14]. sorghum [WTL17]. SOSS [KHKC^{+}18]. Sound [DLH^{+}14, PNHL14, WSS^{+}16]. sounds [FSN^{+}15]. Source [SBC^{+}16a, BBC^{+}19, LLM^{+}13, Loe12, MLMFN^{+}15, NKSA13, PMBM11, SBA^{+}19, SRXP14, SWZ^{+}12, WFM16, YDZ^{+}21, ZH11]. source-sink [BBC^{+}19, Loe12]. sources [BG19, BU16a, BU16b, GBR13b, KH11, KGM12, MAC14, OLG^{+}13, SMT19a, SMT19b, TGM^{+}21, WBB^{+}21, ZLZ^{+}17a]. South [GCL^{+}19, NJCF11, WCS^{+}14, WSW^{+}16, BSS^{+}14, DLW^{+}17, TL15, WTG^{+}16, BSFR17, BEPC15, DVL11, GPA^{+}13, OTA16, PCTW17, SGM^{+}14, TCH^{+}13, VASL^{+}15]. south-eastern [DLW^{+}17, TL15]. South-to-North [GCL^{+}19]. South-western [TCH^{+}13]. southeast [ADL^{+}10, ACK^{+}14, SSY14, WRSL11]. Southeastern [SCB13]. Southern [FRM^{+}19, dHR18, Bre12, CHK^{+}12, FJH15, FSS^{+}17, HKM^{+}18, IBD^{+}19, JMdW^{+}10, LAEAG12, MNEG11, NMC^{+}14, NMS^{+}14, OCCF17, PMS^{+}17, SZSV12b, SRV^{+}12, SKM^{+}16, SGWV15, VASL^{+}15, WTG^{+}16, WJS^{+}18, ZXD^{+}13, mGZK^{+}10, GRL18, KMP14, LAB^{+}10, PHC10, PMTKC17, RMTM^{+}16, WCG^{+}11]. southwest [BTPG10, GSPSP15, LCZ11, PSKO16, DLvGGNC16, LYL^{+}17]. Southwestern [HPS10, LBdA^{+}19, NCK10, YZA^{+}18]. sown [TDC^{+}11]. soy [KAJØ19]. soybean [BWH13, LYD^{+}17]. Space [LCDJ^{+}12, NSIiT17, BLFS18, BRV11, BCJ^{+}12, BPN12, GDW12, KMRW16, LF11, LLQN13, PBT^{+}11, SP12, TMC^{+}16, VR12, WB10, WBG^{+}14, YJW15, YMCV21]. space-for-time [TMC^{+}16]. spaces [NCLP10]. spacing [KMM^{+}13]. Spain [BRV11, GSJ^{+}10, MPKC12, MJB^{+}11a, MGGLBIBD13, NGGFM19, AGQP^{+}14, DLvGGNC16, FJH15, PGGMMA19, RCBPMN10, SGRAdJ17, TCH^{+}13, ZDG^{+}13]. sparse [BMSLC16, GBP^{+}10, TB12b]. Spartina [CZCL18, TSP19, WJL^{+}13]. Sparus [PAGCV^{+}10]. Spatial [AMH16, BCC^{+}11, BA14a, BGG^{+}16b, BOGA14, BLOJ^{+}19, BvBB10, CTC^{+}18, EBA^{+}19, GM10, GA15, HFMM^{+}12, KMKL19, LCDJ^{+}12, MDOJ14, NLM^{+}19, RWS^{+}15, RCBPMN10, SXNL18, UDNSLC^{+}18, Wal17, ZLB14, AMG19, AWWD16, AGQP^{+}14, AHNK13, AD18, ADF^{+}17, AFC12, ADPS10, BKN21, BL15, BMP^{+}19, BGMJ^{+}12, BRS^{+}18, BS17, CMT^{+}11, CMM^{+}10, CF14, CSN^{+}19, CLH17, Cha12, CCLX17, CXW19, CIHL14, CSR15, CSS^{+}16, DBK^{+}13, DDR^{+}18, DVCM^{+}10, DBH^{+}17, DLL18, DS12, DSBS15, DPBS10, FN15, FT16, GGSP14, GPA^{+}13, GMLRF10, GCK^{+}15, GPZB18, GLS10b, GNB^{+}14, GCW^{+}16, HM18a, HM19, HMD^{+}16, HJ17, HSTS19, HCS^{+}10, HCP10, HM12, IN16, JBB^{+}11, JGB15, JGX13, KSS11, KSI14, KB18, LBE^{+}13, LVLS14, LJ15c, LDC^{+}16, LPV11, LL18a, LSGT13, LZZW17, LDCB21, LMBL12, MAO19, MPR13, MD16, MAM^{+}12, MMS12, MSV^{+}11, MJLC14, MRWN19]. spatial [MGB13, MP17, NYIH17, NWAR12, PMM11, PDV17, PJ15, Phi11, PBD^{+}17, PS13, PSS16, RHO11, RWS^{+}16, RMHE15, RKNO17, SNS^{+}12, SMI^{+}19, SX11, SSR16, SEHD18, SMH16, SPWR17, SCG^{+}13, TB11, TB12a, TAT^{+}16b, TT12, TQZ12, VSPHS15, VALC12, VZF^{+}11, VQS^{+}18, WLY^{+}15, WGTH16, WCL^{+}17, WYY^{+}19, XZSL17, XCR^{+}13, YHS^{+}11, YL18, YHW^{+}12, ZH11, ZWF18, ZF11, ZBSF17, ZWG^{+}17, dJHMS19, dSD14, dSD18]. spatial-temporal [WGTH16]. spatialised [TSM^{+}15]. Spatially [BJR16, CRB16, Che17, DVF^{+}19, GCB^{+}11, NPJ^{+}17, PKB^{+}18, STSA10, ZRC^{+}15, ARTBB14, AS12, BSP12, BEPC15, BRMW18, BS16b, BPN12, dSCdA19, DC10, DVL11, EEK^{+}18, FR11, FBKS17, GdSHA10, GPZB18, GVV17, HSHF^{+}13, HHC15, HM12, HKM^{+}18, HQF^{+}17, JOHW17, KR10a, LKMCW14, LYC^{+}15a, MLG18a, MKL14, McC16, MAP^{+}13, Mer12, MMC^{+}19, MLAT^{+}18, NMT^{+}19, NTK14, QL15, RMP^{+}18, SBA^{+}19, SBC13, SZNL19, SEHD18, Sto10, Str15a, Str15b, TVF^{+}10, WGT12, WFW^{+}17, WGE^{+}19, WHRR15, WNF^{+}15, dMSL^{+}16, vdKGV11]. spatially-distributed [SEHD18]. Spatially-explicit [CRB16, PKB^{+}18, ZRC^{+}15, BS16b, FBKS17, GPZB18, JOHW17, McC16, NMT^{+}19, Sto10, vdKGV11]. spatially-heterogeneous [WGE^{+}19]. Spatio [CdMMF16, DVCM^{+}10, FSM19b, GCM^{+}13, MLS^{+}11, MVC11, MvDL^{+}10, SSY14, CHPS17, EMH10, FS18, MPKC12, MLMFN^{+}15, PKW^{+}10, TJYZ11, WLY^{+}15, ZWA19]. spatio-dynamic [MLMFN^{+}15]. Spatio-temporal [CdMMF16, DVCM^{+}10, FSM19b, GCM^{+}13, MLS^{+}11, MVC11, MvDL^{+}10, SSY14, CHPS17, EMH10, FS18, MPKC12, PKW^{+}10, TJYZ11, WLY^{+}15, ZWA19]. Spatiotemporal [BVGR^{+}17, GGZ^{+}16, GHH^{+}15, YHW19, ABD^{+}18, ChHC12, GPCG15, LZLC13, MHMRAHHS19, RWS^{+}15, TOQW11, YZL12, YKZZ14]. Spatiotemporally [CGG^{+}18]. spawner [CH14]. spawning [CEWW18, CNI11, HSCS14, MMMGLN^{+}18, OAPP12, PHP15, QWJL19, VVH12, YWY10, ZSG^{+}18]. Special [Ano11-37, Ano11-38, Ano14-40, FKDG14, LP17, WTC^{+}18, BFW14, BH10, GT14, GB16b, GB16d, HBF18, VMG^{+}12]. specialization [BBHP13, XHC^{+}15]. Speciation [OOH^{+}12, CdOR10, Che16, RCST11]. specie [GSSM13, NZL13, YAGMM^{+}14]. Species [AG11, BGG16a, BOW^{+}12, HSLC^{+}19, LWNG13, PMB^{+}19, RRJVA13, SBR^{+}19, SZNL19, SRTS18, SP10, SHS11, VVH12, ABB^{+}14, AMRH19, ACE^{+}13, AGB^{+}13, AFC11, AFC12, AHF^{+}13, BSW18, BSFR17, BYR18, BS15, BG19, BBJV^{+}11, BKN21, BFPR16, BGD^{+}16, BS16a, BCV13, BM16, BSC^{+}15b, BFNH14, BMCG12, BJR16, CADB19, CDR^{+}13, CR12a, CF17, CBR^{+}16, CHvdW^{+}10, CSN^{+}19, CS17, CHG^{+}15, CPS^{+}19, CSFR15, CW11, CWA^{+}15, CMV^{+}16, CPAG15, CDM13, CRVS13, DAF^{+}11, DUASLB^{+}12, DJH10, DÖWB12, DKJH13, DG16, DFD11, DLvGGNC16, ETG^{+}18, ERGW^{+}11a, ERGW^{+}11b, EBL^{+}11, FM11, FMW^{+}19, FHAG12, FJH15, FL11, FGPR10, FSM^{+}19a, Flo18, FCLFB18, FT16, For18a, FAS^{+}19, FMF12, FKK^{+}18, FCV^{+}12, FML^{+}17, GWR11, GCDA16, GSQ17, GCA16, GVGZ^{+}21, GDC11, GDC12, GEEC14, GGV10, GGV11, GGV13, GvdWBS11, GK14, GLSA^{+}11]. species [GMC^{+}19, GFW^{+}17, GAHE11, GJK^{+}15, dSGGBW12, GLY^{+}15, HJWG12, HPJ15, HH12, HLL^{+}19, HHD12, HCP10, HKM^{+}18, HJN12, HM17, HKVD16, IPSH14, IBH^{+}15, ISG13, JVY19, JTM^{+}17, JZG10, JGB15, JWLS17, KDT^{+}17, KSM^{+}18, KAF^{+}14, KCKdOW12, KKBZ17, KKP^{+}17, Kos17, KSG^{+}14, KLS^{+}14, LV17, mLcLJ15, mLcLJ16, LM13, LJAD17, LRZ12, LBM16, LMR16, LCM11, LKG10, LCY^{+}13, LCCL15, LJ15c, LSJG17, LHW^{+}15, LPT^{+}18, LLQN13, LKAL^{+}13, LYC^{+}15a, LJLO11, LJWN14, Loe18, MGD^{+}12, MU17, MF16, MDH15, MG18, MHG18, MRM^{+}17, MMPG16, MAMNL^{+}15, MPG18, MDG10, MBA^{+}19, MMMCAHM21, MWD^{+}15, NSIiT17, NA11, OOFVBC16, OT15, OCB^{+}13, PR13, PKG16, PK16, PMRGF10, PTL14, PKD11, PGS16, PLHG14, PL10b, PSWW11, PvVC15, PFR15, PDBP11, RXSK14, RLS11, RWVK11, RKC^{+}15, RvKM10, RKBG14, RHD^{+}13, RDH^{+}13, RHR^{+}15a, RHR^{+}15b, RGH^{+}13]. species [SZSV12a, SC21, ŠK14, SSBJ12, SF10, San11, SFM^{+}18, SW19, SvWP^{+}15, SZ14, SA13, SRV^{+}12, SGM^{+}14, SJB^{+}16, SSKJ18, SDKE18, SKL^{+}13, SPB^{+}17, SAKB^{+}10, SvdR10, SDS^{+}21, Str15a, Str15b, SP12, SNZ^{+}19, SLD^{+}19, TAC^{+}14, TW18, TBA19, TLS^{+}16, Tul14, VCE17, VRA^{+}17, VLF^{+}17, VBPY11, WPG15, WBB^{+}21, YGL18, YHK^{+}19, YHW^{+}12, dSD14]. species-by-species [HJWG12]. species-level [FAS^{+}19]. species-rich [GWR11]. Species-specific [AG11, SBR^{+}19, AMRH19]. specific [AMRH19, AG11, Dil10, EMH10, GHST19, HGE^{+}10, KWGW19, PWK^{+}17, RPGA15, SBR^{+}19, TM18, UFK^{+}10, WS15, YCL^{+}14, dVLS11, dAdSC15]. Spectral [GRL18, Ban11]. Spectral-temporal [GRL18]. spectrum [GPM16c, ZCR16]. Speed [Jör11a, GDC11, RHO11]. sperm [CMA^{+}13]. Sphagnum [HZT^{+}13]. sphenanthera [LGDW12]. Spherical [BLH10, Mat17]. spider [Joh10]. spill [KWGG19]. spillover [BG15, CHvdW^{+}10]. Spin [SYL^{+}13]. Spin-up [SYL^{+}13]. spined [MBM^{+}18]. spirochaetes [WGSR19]. splines [CSOS17]. splitnose [GC11]. splitting [CH15]. sponge [GRYM11]. spontaneous [RHQ^{+}16]. SPOT [HS11]. spots [BSCG10]. spotted [TPG14, LAR^{+}16]. Spotted-Wing [LAR^{+}16]. spp [AS11, AOW13, PPZ^{+}17, PMF^{+}13]. sprat [FAT^{+}15b, Hei11, JOHW17]. Sprattus [FAT^{+}15b]. sprawl [PTA^{+}12]. spray [DKM^{+}14]. spraying [SDSH15]. Spread [FMA^{+}12, ABR12a, ADF^{+}17, CMSR19, DPPI14, DSBS15, FHG12, GPJL15, LSR16, MBLC14, MSKS13, MF16, MRC19, MMGK15, MWD^{+}15, PKD11, SM10, SDS^{+}21, SRT16, SH18, ZHLZ17]. spreading [SM12]. Spreng [MAO19]. spring [CAMNN16, DWP16, qHLlZ^{+}16, JG10, JX17, LAC10, MGWW13, PVBC14, WBS13, ZPZ^{+}15]. Springer [Wei19]. springtails [MAP^{+}13]. Spruce [Gar17, ABK^{+}16, CAC^{+}16, DBG14b, DBG14a, mGZK^{+}10, HJS^{+}16, HOKH18, HOŠ^{+}12, KPT11, KPTC14, KMP14, MVC11, MCG19, NPH^{+}17, OJ10, TNA^{+}17, TVF^{+}10]. Squares [IBD^{+}19]. squeeze [HWE^{+}18]. squid [ZBMBDU^{+}17]. squirrel [Hat14, JWLS17]. squirrels [TDF^{+}19]. SRTM [MAO19]. SRTM-derived [MAO19]. St. [CNI11, CLD^{+}17, MDVG^{+}15]. Stability [MW12, PP16, BP19, BFP^{+}18, BBHP13, BH13b, CD17, CD16, HDL^{+}16, HWH^{+}18, KCC16, MBSF14, SU15, VCD^{+}17, ZFZ^{+}14, ZYKS15, dAdSC15]. stabilization [MWM^{+}16]. stabilize [BGP11]. Stable [Ret12, DLH^{+}14, DCZ^{+}15, GVV17, KK11, MDL^{+}10, MLMH18, PLR^{+}13, Vav12]. stacked [GGV13]. stage [AHS^{+}17, ASS14, CCSF11, CCF13a, EPA^{+}10b, MCMT^{+}16, Oga11, SFL^{+}19]. stage-structured [AHS^{+}17, EPA^{+}10b, SFL^{+}19]. stages [DdSCG11, GCM^{+}13, JNHF18, NBSGW19, RWvDM15, SM12, SU15, VVH12]. stakeholder [GCCJ12, MJLC14, XZCL16, vGKB10]. stakeholder-driven [MJLC14]. stakeholders [BRW^{+}12]. Stamou [Pat10]. STAMP [MTL^{+}10]. stand [ACDC17, BHC^{+}15, CADB19, CN11, FH11, For14, FT16, LSA^{+}19, Oga11, Oga17, RGH^{+}13, TVF^{+}10, TKH17, TGB^{+}17, ZXD^{+}13]. stand-level [CN11, For14]. standard [GHWG^{+}14, SSKJ18]. standardisation [PM16]. stands [ACDC17, Gar17, mGZK^{+}10, GCK^{+}15, KLS^{+}14, MPKC12, Oga17, Oga18, RLS11, SKK^{+}13, VALC12]. starch [FS14b]. starting [RDD^{+}18]. starvation [MCH15]. State [BCJ^{+}12, HGLZ13, RDR^{+}13, WBC^{+}12, AYASR^{+}17, AB10, BLFS18, BCP16, CCF13b, CpW16, CGAB17, JTB^{+}19, KP13, LVV^{+}15, MSCM18, NF11, NCLP10, PBT^{+}11, Phi11, RHO^{+}12, RDH^{+}13, RHS^{+}16, SHBG10, SA16, SP12, WB10, Wan19, WCG^{+}11]. state-and-transition [KP13, Phi11, SHBG10]. State-dependent [WBC^{+}12, MSCM18, WCG^{+}11]. state-of-the-art [CCF13b]. State-space [BCJ^{+}12, BLFS18, PBT^{+}11, SP12, WB10]. States [SHB^{+}11a, BS19, GRYM11, GVV17, LF11, Löp11, MDJ^{+}11, MJL^{+}13, MJL^{+}14, McC16, PGA^{+}11, RCP17, TB12b, CRD^{+}16, Gli14, HSR^{+}15, LLT^{+}14, LLT^{+}17, RWFD13, RTE^{+}15, SHC^{+}15, SBF15, SWR^{+}16, SB15]. static [OGG^{+}16, VRC17]. stationarity [RDD^{+}18]. stationary [RHS^{+}16, RDR^{+}17, RV15]. stations [CCL^{+}13]. Statistical [MGS15, BTM^{+}10, BBH^{+}17, LGY10, LSL10, MROGVR12, NYIH17, SMI^{+}19, VEMS12, WLC^{+}19, ZDL^{+}17]. Statistical-thermodynamics [MGS15]. statistics [BLAV18, OGOWW16, SBVA^{+}13]. status [AHS14, ABL^{+}13, MHC16, SKM^{+}16, WSS^{+}16]. StDM [SBC13]. Steady [Löp11, GRYM11]. steel [mNkGyT^{+}17]. steelhead [MSB^{+}17]. steepest [TOHT14]. steepness [HLK^{+}12]. Stefano [FBDR17]. step [OGOWW16]. Stephen [Wei19]. steppe [CSS^{+}19, PVBC14, SAW^{+}11, ZJL^{+}16]. steppe-meadow [CSS^{+}19]. steps [SFvdM^{+}17]. stepwise [SZSV13]. sterile [PvVC13]. Sterna [GQNMR18, JJKJ11]. Sternula [KK13]. Steud. [ESA12]. stickleback [MBM^{+}18]. stimuli [KNU^{+}15]. stitch [SP11]. Stochastic [CJLV11, DSJ19, HFSL14, MGW^{+}21, PMLL11, BEPC15, BP10, CMM^{+}10, CMD12, EAB11, EC10, GDT^{+}14, JBB^{+}11, JM15, Ken19, LSA^{+}19, LBWW14, LHZN11, LJLO11, MBG17, MMC11, MMGK15, RNS^{+}19, RGRMRR18, TT12, TL15, vDHH^{+}14, SBC13]. stochasticity [BRF^{+}11, CDG15, ECO^{+}14, MMPG16, RHW17, SF10, San19]. stochasticized [WAGM17]. stock [BS13, CAL13, HLK^{+}12, Lem16, MTZ^{+}10, MM19, OLG^{+}13, RAB^{+}10, SP12, WKZ^{+}19, YYH19]. stock-recruitment [Lem16, SP12]. stock-recruitment-age-structured [CAL13]. Stockholm [WFM16]. Stocking [FMG^{+}16, LJ15b, RVNB11]. stocks [DR17, HMS^{+}10, LBC^{+}15, MLB^{+}16, MCG19, Nei11, SHM^{+}14, TLL^{+}15, ZCR16]. stoichiometric [ELPK12, PHP^{+}14, Pea15, Vav12]. stoichiometry [LWGH13, WSE12, WLR18]. stomata [MEC14]. stomatal [HSD^{+}15, MTL^{+}10, MEC14]. stomatopods [ADL^{+}10]. stone [CMT^{+}11, MPKC12]. stonefly [CDR^{+}13]. stop [CGB^{+}13]. stopover [SSD^{+}17]. storage [ACN14, CT16, JDT^{+}17, TTCH^{+}14, Wan19, WPB14, ZFT^{+}10]. Storm [YBA^{+}16, LJBS12, YYH19]. stormwater [LCP15, MXC13]. stranding [YDT^{+}13]. strategic [MJLC14]. strategies [Ana17, BS13, BPB^{+}18, BBL11b, CHTC11, CTB^{+}12, CB15, CESC19, EZ14, FMG^{+}16, FLA19, GBB13, HGBA^{+}13, IWW12, JBE^{+}15, Joh10, Kal19, KPLS15, KKI^{+}14, LGMLD19, LJ15b, LLL19, LL18b, LB15, MKH^{+}17, MFB17, MvDL^{+}10, PKB^{+}18, PP19, PJ17, PRP17, PDV17, PRH17, RVL^{+}18, RAdA12, RMR^{+}16, RvKM10, SMH^{+}12, Str15a, Str15b, TBJ^{+}15, TL15, XLS11]. strategy [CGLA10, ESP^{+}17, Flo18, GTC16, GHS^{+}16b, HBT^{+}18, IS13, ISG13, LLH^{+}13, MSM^{+}17, MCN16, RPP^{+}15, SLRC19, UDP17]. stratification [KMKL19, RWH^{+}14]. stratified [CIHL14, FMF12, JTB^{+}19, RGRMRR18, STSA10]. stratified-population [RGRMRR18]. Strayer [Jør14b]. streak [LBR^{+}17]. streaked [HSTS19]. stream [BA14a, CL10, CGdCCPCPM14, DPPI14, DKJH13, DRS^{+}16, GPCG15, GWZ^{+}15, HKW^{+}18, HKVR13, JBW^{+}17, KC18, KSG^{+}14, LW14, MGCL17, Pen11, PGATS13, SAR11, WR12, ZLJM17b, ZLJM17a]. stream-resident [MGCL17]. streambed [KPSC10]. streamflow [GTC16, LC11, MMRS^{+}19, YYL13]. streams [ARTBB14, CDR^{+}13, JNC^{+}11, MNK^{+}21, MR17, MJK^{+}17]. strength [AGJ^{+}12, BSW18, HLL^{+}19, Md18, YZ16, ZML^{+}12]. strengthen [SRTS18]. strengths [WFM16]. stress [BFA11, BCB^{+}11, BEBA19, CBCB17, DSS11, FMG^{+}16, KNU^{+}15, KS17a, KS17b, LVV^{+}15, MEC14, RMvW13, SKTG11, SLZ^{+}13, WFSH15, YqWB^{+}15]. stressors [BR16, IT17, JBW^{+}17, KRCVF16, MTJF11b, Rai13]. strict [WSE12, WLR18]. string [RDD^{+}17]. strip [SRXP14]. striped [MLS^{+}11]. striped-bass [MLS^{+}11]. strong [BRF^{+}11, DJH10]. strongest [HKM^{+}18]. Structural [CL10, CJR^{+}16, GB16d, AAFH15, BPR^{+}11, BBL^{+}17, KCLL14, KJH^{+}13, LDO^{+}18, LYCZ11, NOOM^{+}16, PW10a, SGKR10, VPC^{+}14, Wal15, Wil11a, XWC^{+}11, ZLC13]. structurally [ZGJ10]. Structure [CR12a, LJCG15, ABC^{+}14, AYASR^{+}17, AD18, Ant14, BVC^{+}18, BSASSJ^{+}13, BLD^{+}10a, BND19, BC11, BPGSS^{+}18, ČTT^{+}12, CBR^{+}16, CGGL16, CBO16, CPS^{+}14, DA19, DH12, DCZ^{+}15, FE14b, FSML11, FGPR10, FAC^{+}19a, FNR11, Fuj16, GYL^{+}13, HBS^{+}16a, HKG^{+}17, HWH^{+}18, IWW12, IKL14, JB11, JHH^{+}12, Kal19, KGS11, KP13, KvOP10, KHL^{+}16, KLS^{+}14, LBH^{+}15, LBD10, LKDD15, LJ10, LLFL12, MMD^{+}17, MRB14, MK14, MJGC13, NCEGB16, NWAR11, NWAR12, NJCF11, OBC^{+}21, Pec12, PP12, Ret12, RNS^{+}19, RAR15, RMvW13, SCHQA16, SLW^{+}14, SMH16, SS11, SGKR10, TKW17, TVI^{+}13, TCH^{+}13, TS11, VC17a, XC14, XKI^{+}11, YHW^{+}12, ZLLC13, ZLZ17c]. Structure-dynamic [LJCG15]. structured [AAME11, AHS^{+}17, BNC^{+}18, dÁGP14, CAL13, CC12, CRUF13, dSCdA19, EPA^{+}10b, GS19a, KMS^{+}12, KMS^{+}13, LKLK15, Lem16, Log16, MM19, NMdR16, NTK14, RPM13, SBA^{+}19, SFL^{+}19, TKP14, VHDM17, WAGM17]. structures [DBH^{+}17, LAD11, PÓF^{+}18, SVD^{+}13]. structuring [RAdA12]. Stubai [FCFT16]. stubbornness [ZYXT18]. students [Mob18]. studied [AYASR^{+}17]. studies [ABB^{+}14, BCD12, CL15a, GDDFT16, HKlL14, JTM^{+}17, LCMG18, SDD^{+}13, TVT12, VG19, YC11, YHJ^{+}17]. Study [AAA13, Ano12a, HZSH17, LYC^{+}15b, WLC^{+}19, AGQP^{+}14, ABS17, Ana17, AHM^{+}16b, ABD14, BVC^{+}18, BCB^{+}11, BVJE18, BH17, BEPC15, BLD10b, BC11, BEBA19, BG11, CVV^{+}13, CZC^{+}15, CBGK15, CWLL15, CL15b, CpW16, CTC^{+}18, CGEdM12, CRTS15, CMSPJ11, CBB^{+}10, CWA^{+}15, DHC14, DPBK10, DPvBvA11, DKJH13, DSW^{+}17, DLD^{+}11, EGMW12, F.11, FS14a, FKH^{+}15, FCZ13, FCH^{+}17, FP17, FCFT16, Fre16, GWHL15, GZS^{+}18, GDT^{+}14, GL10, GS19b, GB15, GBR13b, HRSKW15, HBR19, HFSL14, HJ17, HM18b, HKVD16, HYL^{+}17, JBE^{+}15, JMW11, JJYM19, JHT^{+}14, JGX13, KSSA13, KM12b, KZR^{+}13, KHJ^{+}19, KAC^{+}12a, KAC^{+}12b, KCKdOW12, KMKL19, KL12, LDG14, LSH15, LY16, LHL^{+}17, LLL^{+}10, LYSW15, LLZ^{+}16, LHL18, LBW18, LWTZ18, LDM15b, LCP14, LBS^{+}14, LYC^{+}15a, LML17, LVV^{+}15, LMBL12, LGDW12, MTZ^{+}10, MLY^{+}19, MLG18a, MAM^{+}12, MAC14, MRD^{+}12, MRZU14, MCG19, MRP^{+}11, MB14]. study [MGSB11, MMVAHMC16, MOM^{+}11, Mur13, NCMF10, Nei11, NXZ18, NMA^{+}19, NGGFM19, OHS^{+}17, Oga11, OTA16, PLR^{+}13, PMS^{+}17, PPB^{+}18, PFA^{+}16, Pen11, PGKG12, PBD^{+}17, PLS11, PGGMMA19, PGA^{+}11, RBM^{+}17, RSCS19, SMC11, SBCB13, SSV^{+}12, SJS^{+}13, SEHD18, SK11, Sin17, SGRAdJ17, SDKE18, SAKB^{+}10, SGP18, SSM^{+}13, SSW^{+}18, SGKR10, TPL^{+}15, TLL^{+}15, THJ^{+}12, TPG14, TDH12, TLW^{+}14, TOQW11, TYEH10, TBL^{+}15, URR^{+}17, VVH12, WALCY12, WSW^{+}13, WYHC15, WFN^{+}16, WØ15, WX11, XCR^{+}13, YSH^{+}10, YYQ11, YJW15, YLHZ14, YBA^{+}16, ZBAY14, ZLY16, ZYFL10, ZFT^{+}10, ZLC13, ZKR^{+}13, ZLF14, ZLF^{+}15, ZQC15, ZZF15, ZLFZ16, ZZJ^{+}16, ZDL^{+}17, ZZLZ14, ZYZ17, ZMH19, dSdCF10]. sturgeon [DHC16, HMP^{+}19b, WAGM17, YWY10]. STWIR [GPD^{+}13]. styles [BRW^{+}12]. sub [BN10, GPK^{+}13, KUK11, ZNS^{+}18]. sub-district [KUK11]. sub-estuary [BN10]. sub-soil [GPK^{+}13]. sub-tropical [ZNS^{+}18]. subalpine [ACM15, KKP^{+}17]. subarctic [KSAS12]. subject [SP12]. subjected [PICR13]. subjects [RDV^{+}17]. submerged [GSQ17, GMLRF10, JDGJ13, LHZ10, McC16, VAO^{+}17, XYCZ13]. submodels [LM16]. suboptimal [BM10]. subpopulation [CD15]. subset [TTL^{+}18]. subsidies [BEZ19, EZ14, PKT15]. subsidy [BKV17]. Subspaces [CVS^{+}13]. subspecies [AR11]. substantial [NZL13]. subsystem [TM14]. subtidal [JGB15, WLR^{+}11]. subtropical [DSS11, DCZ^{+}15, HSVL12, MDL^{+}10, WQD^{+}13, WWC^{+}18, WZD^{+}14, WEJ11]. suburban [WR13b]. success [CH14, CGM^{+}14, DPvBvA11, JMdW^{+}10, LSH15, Rin14, SK10]. succession [DVF^{+}19, FAMD^{+}12, GRL18, HSVL12, KCF^{+}13, KGD14, Lin15a, LM19, LJCG15, LRB12, LHLW15, MPT^{+}11, MRR^{+}14, Ske13, TMC^{+}16, ZYZ17]. successional [DVF^{+}19, HBL18, PHC10, PGA^{+}11]. such [LM13]. Sucre [RDR^{+}13]. SUCROS LL [GLS^{+}17]. sugar [BDL^{+}13, BDL^{+}14]. sugarcane [PvVC13]. Sugeno [FDM^{+}11]. suggest [PJ17]. suggested [Til14]. suggests [Bin17, TZD^{+}15]. Suitability [LA11, AMM^{+}14, BKP10, Bar11, BBGJ13, BOW^{+}12, BCD12, CKMD14, DKJH13, FHG12, FDM^{+}11, GCDA16, HPJ15, KMP14, LCCL15, LML17, LGDW12, MSV^{+}11, NiSK21, RXSK14, SK11, TRP12, VFFP16, WL13, YXD^{+}18, ZPNM13]. suitability-based [BBGJ13]. suitable [HWE^{+}18, KC18, LYSW15, SMSAN^{+}14]. suite [GdC14, SVD^{+}13, SPP18]. sulfide [GS10]. sulfur [GZS^{+}18, WB17]. sulphuratus [LIM^{+}16]. Sum [RV13]. summary [GT14]. summer [MBD^{+}18]. Summit [WLF^{+}16]. sun [FvBM^{+}11]. sunbleak [BGC10]. Sundarban [MRG13a, San17]. Sundarbans [DCMS18]. sunflower [KCLL14]. SUNLAB [KCLL14]. sunshine [JGX13]. superba [GJK^{+}15, MRCW10]. superior [MdlP15, VBPY11, MBUM12, MYJ^{+}14, SGWV15]. supersaturated [YFL^{+}18]. supervised [ASL13, CGP12, FIG^{+}10]. supplementary [HKG^{+}17, PKT15]. supply [AS14, CEC10, GAP^{+}19, HRSKW15, HKN17, JAH^{+}10, LL18b, SGAS14, XYTS19, YY13]. support [BGG16a, CGL15, DWD19, GAF^{+}14, HRSKW15, KRE^{+}18, KK15a, KKI^{+}14, LWKS10, MMVAHMC16, PMB^{+}19, SZSV12b, SZVV10, SWGC16, TAT^{+}16b, WYHC15, vGKB10]. supporting [ABC^{+}17]. suppress [FTYI13]. suppression [PvVC15, YHW19, ZYXT18]. Surface [YFC^{+}11, YDL^{+}15, BCC^{+}11, BLRN15, Gav15, GPCG15, GANB11, JZG10, LLM^{+}13, LZ10, MBSW12, MJÁ^{+}12, OSS13, Pie10b, Pie10a, RSC12, RSC13, SMC11, SH14, SAI10, SCG^{+}16, WFA10, WBC^{+}12, WBHM11, ZWJ^{+}12, Jör11a]. surfaces [LMGM11]. surplus [RL15]. Surrogate [NNP19]. Surrogate-based [NNP19]. surrounded [IBD^{+}19]. surrounding [ZJL^{+}14]. surveillance [GKKW10]. survey [Bin17, MYJ^{+}14, Tul14]. surveys [CMV^{+}16, HOŠ^{+}12, PC17, Pér18]. survival [BHHM10, CHG^{+}15, CN10b, Dil10, GCvDM12, GCM^{+}13, KR14, LCC^{+}14, MRCW10, OMDP11, RP11, SBCB13, SF10, SFvB10, THJ^{+}12, YL17]. survive [dSJHG10]. susceptibility [MBLC14]. susceptible [HS11]. suspended [MTPC13, Sin17, YCF15]. suspension [BPK^{+}16]. SUSTAIN [GWHL15]. Sustainability [DH18, Har11, IGR^{+}18, MFESBG^{+}13, PSWM^{+}10, PBL^{+}15, WAM13, vGKB10, Ano12t, ABC^{+}17, BGMFES15, BBA12, BU11a, DPR10, FB15, GZVU14, GAB12, GSA^{+}19, GWZ^{+}15, KSSA13, KR10b, LL17, LHZN11, LY11, LSP^{+}10, LYCZ13, RAB^{+}10, SSV^{+}12, SJS^{+}13, WHH^{+}18, YKZZ14, ZHC^{+}16, dSdCF10]. Sustainable [BAA^{+}14, DMB^{+}18, DSD13, Fat14, GK14, CGL15, CHCC15, Che15a, EMH10, GB15, KG13, KRE^{+}18, LG10, LKG10, LG12, LLP17, LP17, PK16, SKM^{+}16, YYQ11, ZF19]. Sustained [GvdWBS11]. sustaining [LWL^{+}17a, RE10, SGB10]. sustains [MSG^{+}13]. suzukii [LAR^{+}16]. Sven [Fat16b]. SVM [PMS11]. SVM-based [PMS11]. SVMs [SZSV12b]. SW [RML^{+}13, GCK^{+}15, RCBPMN10]. swallowtail [PZM13]. swamp [Ano12s, HR11, MCW13]. swamp-dominated [Ano12s, MCW13]. swarm [LOLA13, PD12]. SWAT [KJMA11, KPSC10, MCP16, NKSA13, PMBM11, SWZ^{+}12, SV13, YDZ^{+}21, ZCYL15]. SWAT-based [SWZ^{+}12]. Sweden [NPP^{+}18, JL18, JAKJ10, LJBS12, OBLJ13, PMS^{+}17, SB15, WFM16]. Swedish [OKS^{+}11, OLG^{+}13, RJH^{+}10, YBA^{+}16, YZA^{+}18, ZBAY14]. sweet [PGGMMA19]. swimming [APWT15, Wil11b]. Swiss [KFvSH15, MLK^{+}11, MBW14, PGR^{+}16]. Switching [WN17]. Switzerland [FRM^{+}19, SB15]. Sylt [FSB13]. sylvatic [LBR16]. sylvaticus [LSGT13]. sylvestris [ASV19, THJ^{+}12, UDP17]. symbiosis [DLZ^{+}17, ZZF15]. symbiotic [CLW^{+}19, GBR13b]. symmetric [dS18]. symposia [LMY^{+}11, LYP^{+}11]. Symposium [FKDG14]. Syncerus [MMS12]. synchrony [LS13, Mat13]. syndrome [CESC19, RTE^{+}15]. synergies [GRRU15, HDL^{+}16, MWML13b]. Synergism [ZLF14]. synoptic [WBG^{+}14]. syntax [YJW15]. Synthesis [BCT15, GB16d, HKW^{+}18, LMY^{+}11, BLOJ^{+}19, BCUF16, BCDU16, FB15, HM17, HPW^{+}12, MYJ^{+}14, RLB^{+}12, YKZZ14]. Synthesizing [Mul11, SFM^{+}18]. synthetic [FD19]. syntrophic [VRC17]. Syrphidae [KTV^{+}15]. Sys [CDT10]. System [Ano12a, CLBSG11, DCMS18, FCZ13, GEZ^{+}13, KAC^{+}12a, KAC^{+}12b, KS17a, KS17b, MXC13, PVM12, PVM13, PK12, VGR^{+}11, WSW^{+}16, BSFR17, BPK^{+}12, CN10a, CC10, CHCC15, CGGL16, CCF13b, CpW16, CLW^{+}19, CAY^{+}10, DFW^{+}18, FB15, FB13, FDM^{+}11, GHS^{+}16a, GvdWBS11, GZVU14, GK14, GOSH^{+}12, GCR18, HGSJ15, HWH^{+}18, IS13, JA15a, JMW11, JBW^{+}17, JSW^{+}13, KSI14, KL10, KTM^{+}16, KHHN17, KJS13, LDM15a, LT15, LWKS10, LDL12, LYCZ11, LOLA13, LFH^{+}15, LPD14, LGDW12, MPM^{+}15, MKSL^{+}17, MTJF11a, MTJF11b, MWM^{+}16, MRG13b, MGW^{+}21, NSIiT17, OCCF17, PBB^{+}14, RBM^{+}17, RKC^{+}15, RP11, SJS^{+}13, SSD^{+}17, SBRD10, SG11a, SSM^{+}13, SGBG19, SXZ^{+}15, TKP14, TAT^{+}16b, THBF^{+}17, WAS^{+}16, WZWL10, WYY^{+}19, WCS^{+}14, XCL^{+}11, YLC13, YM21, YSA^{+}14, ZYFL10, ZLF^{+}11, ZLC13, ZLZS14, ZLF14, ZZF15, ZZJ^{+}16, ZWW^{+}18, ZCY^{+}19, ZHCX15, ZMH19, AAO10]. System-of-Systems [PK12]. System-theoretical [KS17a, KS17b]. system-wide [BPK^{+}12]. Systematic [YC11, AWL^{+}13, MRTW18]. systemic [FLLG13]. Systems [ACKS13, BFW14, HKW^{+}18, HD14, Pat14a, Pat16b, PF18, PK12, TM14, Wei19, ALC19, AG15, An12, ALC12, Ana17, AD18, ABAA21, BGMFES15, BUB^{+}12, BFP^{+}18, BBHP13, BPR^{+}11, BRM^{+}14, BLD10b, Bev15, BT16, BBC^{+}19, BK13, CH12, CH19, CBB19, CK12, DG11, Dol10, DH18, Fai10, FR11, Fat14, FBU14, GPM^{+}16a, GSA^{+}19, GBP^{+}10, GPP^{+}19, HBF18, HF14, Ji15, JZG10, JNF16, KG13, Kei12, KHHN17, KBMM15, LL17, LMT13, LdMS^{+}16, LY11, LTL^{+}13, LLW^{+}15, LHH12, LCDJ^{+}12, MG16, MDJ^{+}11, MKSL^{+}17, Mob18, MCPB13, MCB14, MCPB15, Nie16, OC15, OBC^{+}21, PTML14, Pat10, Pat16a, PRP17, PJ15, PFSF14, PDBP11, Rau11, Rem11, RSDP^{+}12, RE10, SvBH^{+}14, SGBD10, SJS^{+}13, SXNL18, SVL12, SZVV10, SK15, SCG^{+}13, TOQW11, TA11, UAB^{+}11, URJ10, VSB13, VBS14, WAS^{+}16, WALCY12]. systems [WAM13, WTL17, WØ15, ZZL^{+}19, ZHV14, dS18]. systems-level [SJS^{+}13].

T. [PFP^{+}16]. Tabasco [ZRCNLI^{+}15]. Table [SCH^{+}17, DBG14b, GWK^{+}12, LLCR11, XLS11]. table-form [LLCR11]. tables [KPS17, LHL18]. tackle [PGG14]. tactics [DHHB15]. tadpole [POS^{+}19]. tag [BLW^{+}17]. Taihu [GSQ17, HG17, JLZ^{+}18, WJF^{+}11, HFH^{+}19, Hu16, QCB^{+}11, SXZ^{+}15, ZRC^{+}15]. tailed [LGMLD19, WTG^{+}16, WR13b]. tailing [WWYO17, WWYO18]. taint [HD14, HD16]. Taiwan [CLCC11]. Taiyuan [LLZ12]. Takagi [FDM^{+}11]. Taking [Ji15]. tale [Tul14]. Tangled [JJS10]. tank [LCTH12]. tannin [SWS16]. tannin-binding [SWS16]. Tanzania [MWM^{+}16]. Tapirus [GGTP^{+}17]. target [FSM^{+}19a, MKL14, Md18, MHRB19, RSP^{+}17]. targeting [BvHR^{+}15, DVCM^{+}10, RPP^{+}15]. targets [LM12]. task [BHHM10]. Tasmanian [WNT^{+}13]. taste [HD14, HD16]. Taurus [KAF^{+}14]. tawa [MP17]. taxa [CL12b, For18b]. taxocene [RDR^{+}17]. taxonomic [KEM^{+}16, LKDD15]. Taylor [JDZC14, SC18, SSR16]. TD-model [SBDS13]. Teaching [Mob18, HBF18, KS18]. teal [KK12a]. technique [BWA17, HYG14, HSK17, LSA^{+}19, PvVC13]. techniques [AKvO^{+}10, BKMK10, CBB^{+}10, FLLB11, GLY^{+}15, LDM15b, NGGFM19, RGHS11]. Technological [DPR10, FG11, UAB^{+}11, UB12]. technologies [CAY^{+}10]. teleconnect [CLHH19]. telemetry [BMSLC16, BLW^{+}17, FBH^{+}14, WBG^{+}14]. Teleostei [dABLBAdOR17]. tell [PS13]. temperate [Ant14, AOD^{+}13, BFH14, BPGSS^{+}18, CVV^{+}13, FCV^{+}12, GBG^{+}15, JSB^{+}14, KW18, LNB17, LCFD19, LBdA^{+}19, LAC10, LVV^{+}15, LMBL12, MJL^{+}13, MJL^{+}14, PWBG16, RWH^{+}14, RKC^{+}15, SHM^{+}17, TTCH^{+}14, UKI^{+}11, WGA^{+}11, WGA^{+}12, WJvdL^{+}13, LLT^{+}17]. Temperature [ATL18, KMS^{+}12, KMS^{+}13, MBBA17, AHS14, ACCM16, AML13, AKLB19, BMR^{+}18, BTS19, BHHM10, BLRN15, CSR15, CGdCCPCPM14, DPvBvA11, DS21, DFF17, FTK^{+}14, FP17, FMG^{+}16, FBA^{+}11, FCV^{+}12, FML^{+}17, GBP^{+}16, GLR^{+}19, GPCG15, HMP^{+}19b, KDMPC10, KSZW16, KBS^{+}15, KKP^{+}17, KKK^{+}19, LYD^{+}17, LBW18, LBC14, MMRS^{+}19, NiSK21, PL10a, PSBH16, RPA12, SSC15, SM14, VCSSC12, VJR^{+}16, VPS10, WFA10, WV12, WFN^{+}16, WJK11, ZML^{+}12, ZZZW15]. temperature- [FBA^{+}11]. temperature-based [FCV^{+}12]. Temperature-dependent [KMS^{+}12, KMS^{+}13, MBBA17, DFF17, FML^{+}17, KDMPC10]. temperature-driven [FTK^{+}14, PSBH16]. Temperature-influenced [ATL18]. temperature-mediated [BMR^{+}18]. temperature-radiation [ZML^{+}12]. temperatures [BCP15, BM10, MOMC17, WBHM11]. template [HCM^{+}18]. Templeton [Kat11]. Temporal [BSFR17, PKD11, RPGA15, WSZT10, WYY^{+}19, BA14a, BRS^{+}18, CMT^{+}11, CHPS17, CdMMF16, CSN^{+}19, Che15b, CSS^{+}16, DBK^{+}13, DVCM^{+}10, EMH10, FMW^{+}19, FT16, FSM19b, FS18, GRL18, GCM^{+}13, HGPF15, HM12, LSJG17, LS18, LPV11, LZZW17, MPKC12, MAM^{+}12, MLS^{+}11, MVC11, MvDL^{+}10, MGB13, MSR^{+}14, MAMNL^{+}15, PDV17, PK14, PAGCV^{+}10, PKW^{+}10, PSS16, SC18, SAR11, SSY14, SML12, SLLW15, SED15, SCG^{+}13, TAC^{+}14, TJYZ11, WLY^{+}15, WGTH16, YSA^{+}14, YL18, ZBMC14, ZLB14, ZWA19, ZF11, ZBSF17, ZR19]. Temporal-spatial [WYY^{+}19]. temporally [WGE^{+}19]. temporally-variable [WGE^{+}19]. temporary [GEB10, YDT^{+}13]. Ten [AW15]. tendency [ZLZS14]. Tennenbaum [LTL15]. tent [BFA11]. Tenualosa [BAA^{+}14, HSCS14]. Term [Ano12a, AB13, ABC^{+}17, BYR18, BUB^{+}12, BLB^{+}17, BSS^{+}14, CRVS13, DR17, FB15, GV11, GCB^{+}11, HOŠ^{+}12, KAC^{+}12a, KAC^{+}12b, LBZA19, LDO^{+}18, LAB^{+}10, LBdA^{+}19, MCG19, MP17, Och10, OBC^{+}21, PWBG16, Pen11, PGKG12, PGZ^{+}19, RUSL19, RTE^{+}15, SVL12, TGM^{+}21, VG19, VASL^{+}15, WP14, YSA^{+}14, ZJM^{+}11]. terminology [AVG14]. Tern [JJKJ11, KK13]. terns [GQNMR18]. Terraced [BC11]. terrain [BS17, TKN14]. Terrarana [PMF^{+}13]. Terrestrial [HPW^{+}12, BBL^{+}17, HKVR13, HBA16a, JWX18, LLY^{+}13, LXS^{+}16, MV11, MLH^{+}17, OC15, PD15, San17, SHC^{+}15, SLMS14, TLL^{+}15, TRE^{+}12, UDNSLC^{+}18, Wan19, YFC^{+}11, ZZFH15, ZZLZ14, ZJP^{+}11]. terrestris [GGTP^{+}17]. territorial [LB15, MCB14]. territories [VC17b]. territory [CLBG15, RAR15]. Test [DHC16, AHM^{+}16b, FAPM12, Gav15, LWKS10, MBUM12, PGR^{+}16, PLHG14, WR18]. test-bed [MBUM12]. testable [LJ15a]. tested [For18b, SG11a]. Testing [Ano12a, FLLG13, GAV^{+}11, HM10, LRB12, MHTC15, MSR^{+}14, NBK^{+}11, RGHS11, SGBD10, SSK14, AAJB21, AD12, GMS11, GAF^{+}14, HKW^{+}18, HHvO^{+}17, KK13, KJMA11, KAC^{+}12a, KAC^{+}12b, KCB^{+}17, PWBG16, THO10b, VHM^{+}17, MPT^{+}11]. Tetranychidae [dOMZ13]. Tetranychus [dOMZ13]. Tetrax [SSQ^{+}10]. Texas [GRB12, KM12b, WTG^{+}16]. Thailand [LDP14]. their [BDL^{+}14, BSFR17, BLD10b, BF11, BV15, BvHR^{+}15, CAMNN16, CLDRJ18, CGP12, DF13, FH11, FCA18, FKK^{+}18, GLR^{+}19, GHS^{+}10, Hu16, IWW12, KBC^{+}12, MCT10, Mar17, OLG^{+}13, RRH^{+}19, SBR^{+}19, SCHGM16, Sil11, STD15, Tie16, WAS^{+}16, WSZT10, WYY^{+}19, WNF^{+}15, XTLM15]. thematic [Ano11-38, Jør10a]. theme [ZNS^{+}18]. theorem [MRES17]. theoretic [Che10, CB15, GKL18, MRM16, RBRK16, URR^{+}17]. Theoretical [LJ15a, Oga11, VGR^{+}11, VWTA16, DMH17, HB15, HWL17, KS17a, KS17b, LR16, MPT^{+}11, MN12, RvS16, RHR^{+}15c, YLW14, YL18]. Theories [BCT15, MTA19, SBVA^{+}13]. Theory [KJMA11, KB11, NSN17, AG15, AG14, AAW^{+}15, BMF^{+}11a, BMF^{+}11b, BR16, BGG16a, Che16, CPAG15, DMDJ13, ETG^{+}18, Els10, For18b, FPG11, GB16c, HLL^{+}19, HB12, IKH^{+}11, Jen12, KA17, KP13, Log13b, LBGM19, MPT^{+}11, MCB14, OBSL15, Pat16a, PHP^{+}14, POB18, PMRGF10, PZM13, PVB16, RHQ^{+}16, TLW^{+}14, VBS14, WCF^{+}16, WLZC11, XZZ^{+}13, YXZ^{+}21, YZLX14, ZH13]. there [PMH17]. thermal [FSV^{+}12, HSR^{+}15, MPT^{+}11, RMvW13, SBRD10, ZZZW15]. thermal-stress [RMvW13]. Thermodynamic [LHH12, Lin15a, AG15, Ano12t, ASV19, CZCL18, Cho13, HJ10a, LYCZ13, MCT10, PHP^{+}14, PSSH11, PSKO16, WLC^{+}19, WWYO17]. thermodynamic-based [Ano12t, LYCZ13]. Thermodynamics [CPJ^{+}13, HWH^{+}18, MGS15, Tie11]. Thermodynamics-based [CPJ^{+}13]. Thermostatistical [RHR^{+}15b, RDR^{+}17]. thibetanus [DFK^{+}11]. think [PMH17]. Thinking [Gia14, Jag14, Ana17, Fat14, Mob18]. thinning [CN11, MHL^{+}19, Oga17, Oga18, POB18, RLS11, WPZ^{+}11]. Third [Kat11]. thistle [HCW^{+}17]. Thomas [CNI11]. thompsoni [HPGM18]. Thorns [CDK^{+}17]. thought [RDV^{+}17]. threat [MCT10]. threatened [BBRJ15, GKF^{+}14, dSGGBW12, KKP^{+}17, MSB^{+}17, PTA^{+}12, TLS^{+}16, WZL^{+}16]. Three [DBK^{+}13, GDC^{+}19, LZYH19, MBSF14, BTM^{+}10, BLH10, BSH^{+}10, CVV^{+}13, FJH15, Hal11, HCM^{+}15, HFB^{+}11, LCY^{+}13, MBM^{+}18, MH10, UFK^{+}10, WQD^{+}13, WØ15, XCR^{+}13, YZG16, YFC^{+}11, CMLY19, MLW^{+}15, SRXP14]. three-component [BSH^{+}10]. Three-dimensional [DBK^{+}13, LZYH19, MBSF14, BLH10, HFB^{+}11, MH10, WQD^{+}13, XCR^{+}13, YZG16]. three-scale [HCM^{+}15]. three-spined [MBM^{+}18]. threshold [CDR^{+}13, CD16, LKAY15, NA11, WLR18]. threshold-based [LKAY15]. thresholds [GK14, LLH^{+}13, LMGM11, NZL13, ROC^{+}13, VCE17]. throated [HBT^{+}18, MFB17]. throughflow [FSB13]. throughflows [TWP^{+}14]. Thunnus [PAGCV^{+}10]. Thymallus [PGS16]. thynnus [PAGCV^{+}10]. Tianjin [HLW15, HZZ^{+}18, ZLZ17b, NGWZ16, YLHZ14]. Tibet [SLZ^{+}17]. Tibetan [CBO16, SCG^{+}16, CSS^{+}19, GYL^{+}15]. Tick [WGTH16, BDG^{+}14, HM18a, HM19, HMA10a, LVLS14, WTG^{+}16]. Tick-borne [WGTH16]. ticks [WGT12, WGTH16]. tidal [CN10a, CTNH13, COK^{+}11, CDG15, GL10, LZW^{+}17, SQWS19, ZJL^{+}14, ZZZW15]. tide [DSW^{+}17]. tides [ZW10]. Tiger [TWS^{+}11, CLBG15, KSZW16]. tigers [LKG^{+}17]. tilapia [XCL^{+}11]. tillage [LSM19]. tilt [Md18]. timber [DLL18, KCKdOW12, MWML13b, PDD^{+}18]. Time [BLH10, LL17, MRES17, SBA^{+}13, WPB14, Ano17a, ACN14, BSH17, BLD10b, Bus17, CMSR19, CC10, CRA^{+}11, CGAB17, dSCdA19, DSK^{+}11, Dil10, GLT10, GDW12, HB13, HD16, HBL18, JA15a, KK15a, KMMS17, LJR14, LF11, LLTQ14, LLTQ15, MPR13, MGB13, NS15, RBM^{+}17, ROC^{+}13, RWP11, SCHGM16, SBC^{+}11, SP11, TL10, TMC^{+}16, VPS10, WQD^{+}13, WXB10, ZW12, ZZT13, ZSG^{+}18]. Time-delayed [LL17]. Time-dependent [SBA^{+}13, HD16]. time-dynamic [BSH17, LJR14]. time-lags [ROC^{+}13]. time-scaling [TL10]. time-series [RWP11]. timeline [BWML15]. times [Loe18, PRH17]. timescale [GO15b]. timescales [WBC^{+}12]. Timing [Jag14, BBRJ15, dSCdA19, DWWS15, SGG14, WBS13]. timothy [JBB^{+}12]. Tinsley [FTK^{+}14]. tips [XCG^{+}16]. tissue [PBP^{+}15]. Title [Ano10-28, Ano10-29, Ano10-30, Ano11-39, Ano11-40, Ano13-41, Ano14v, Ano14w, Ano14x, Ano14y, Ano14z, Ano14-27, Ano14-28, Ano14-29, Ano14-30, Ano14-31, Ano14-32, Ano14-33, Ano14-34, Ano14-35, Ano14-36, Ano14-37, Ano14-38, Ano14-39, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano15y, Ano15z, Ano15-27, Ano15-28, Ano15-29, Ano15-30, Ano15-31, Ano15-32, Ano15-33, Ano15-34, Ano16v, Ano16w, Ano16x, Ano16y, Ano16z, Ano16-27, Ano16-28, Ano16-29, Ano16-30, Ano16-31, Ano16-32, Ano16-33, Ano16-34, Ano16-35, Ano16-36, Ano16-37, Ano16-38, Ano16-39, Ano16-40, Ano16-41, Ano16-42, Ano16-43, Ano17-41, Ano17-42, Ano17-43, Ano17-44, Ano17-45, Ano17-46, Ano17-47, Ano17-48, Ano17-49, Ano17-50, Ano17-51, Ano21d]. Tixier [Och10]. TK [SBDS13]. TK/TD [SBDS13]. TK/TD-model [SBDS13]. TM [TQZ12, VB10]. toads [BKP10, Bar11, MBG17]. Tokyo [SSN^{+}18]. Tolerance [DP19]. Tolerance-fecundity [DP19]. tolerant [WTL17]. tomato [FKH^{+}15, PDG^{+}14]. Tonle [CSN^{+}19]. Tool [WB16, BLH10, BGK^{+}16, CVPR^{+}16, CLBSG11, CMSPJ11, DWD19, EMH10, HKG^{+}17, KTM^{+}16, LCM11, LSP^{+}10, LAEAG12, MAB19, Mul11, NW18, OFH^{+}19, PVB16, RPC^{+}17, RHS^{+}16, RDD^{+}14, SGB10, SGWV15, TVV13, ZWRK13, ZDG^{+}13, TTCH^{+}14, PSWM^{+}10, FRV^{+}16]. toolbox [LBSB^{+}14, SBC^{+}16a]. Toolkit [DFH^{+}14]. tools [GB15, LRB12, MG16, SZVV10, SCG^{+}13, TLW^{+}14, UFK^{+}10, vGKB10]. Toona [DSS11]. toothfish [MCMT^{+}16, PLS^{+}16]. top [BS15, BTB17, FE14b, GMTC^{+}17, MBD^{+}18, NGL13, PMBB^{+}16]. top-cited [BS15]. top-down [MBD^{+}18, PMBB^{+}16]. topic [BME14]. topo [TAR13]. topo-dynamical [TAR13]. topographic [Kos17, MAO19, MGGLBIBD13, MOM^{+}11, SMC11]. topographical [HS11]. Topological [NCMF10, NG11]. topology [BRPK14, DPPI14, WR12]. toposequences [GE17]. TOPSIS [GH13]. topsoil [LBE^{+}18]. torquans [LIM^{+}16]. tortoise [PTA^{+}12]. Tortrix [KSS11]. Tortugas [MZLCSZdP11, NCMF10]. ToSIA [PSWM^{+}10, LSP^{+}10]. total [BDL^{+}14, OGOWW16, WGD^{+}17, YCF15, YFL^{+}18]. tourism [PNHL14, SHCR10, WAM13]. tourism-based [WAM13]. tower [HCL^{+}14]. town [LLW^{+}15, ZFT^{+}10]. Towson [WTC^{+}18]. toxic [HHC15, JA15a, ZCSL18]. toxicant [HTL^{+}14, MBM^{+}18]. toxicant-induced [MBM^{+}18]. toxicity [GV11, GRB12]. toxicodynamic [Ash10]. Toxicokinetic [Ash10]. toxicokinetics [LBS^{+}14]. toxin [GV11, GBB^{+}11, HQF^{+}17]. toxins [FAMD^{+}12]. trace [GSPSP15, KKK^{+}19, GAF^{+}14]. tracer [LSR16]. Trachemys [BGA17]. Tracing [LCT^{+}17, Bai18]. Tracking [FDR^{+}17, RMvW13, WFA^{+}13, BBGJ13, KNPH^{+}17, MLK^{+}11, Pat15, ZSS17]. tracks [BCJ^{+}12]. Trade [BLD10b, GJB^{+}11, BBL11a, DZW^{+}11, DP19, GK14, HV15, LARZN13, LCPB^{+}10, MWML13b, RHD^{+}13, SDT10, XHC^{+}15, ZFZ^{+}14, dMLM^{+}17]. Trade-off [GJB^{+}11, BBL11a, DZW^{+}11, DP19, RHD^{+}13, ZFZ^{+}14]. Trade-offs [BLD10b, GK14, HV15, LARZN13, LCPB^{+}10, MWML13b, SDT10, dMLM^{+}17]. Tradeoffs [HCS^{+}10, Str15a, Str15b, dOAC18]. Traditional [Cho13, FMBD15, Ned15]. trail [SG16]. trail-laying [SG16]. training [RFV^{+}12]. Trait [GDTP13, ZMS19, CVS^{+}13, KCG18, LJ15a, PBD19, PLL^{+}19, RPM13, SJGB16]. Trait-based [ZMS19, CVS^{+}13, KCG18, LJ15a, PLL^{+}19, SJGB16]. trait-structured [RPM13]. Traits [GANB11, AWWD16, CDT10, DJH10, FL11, GDC11, GDC12, GEEC14, GPH17, HW12, HGAT17, KSI14, KAF^{+}14, MJL^{+}17, RKBG14, RDH^{+}13, SMG^{+}12, TKW17, TM12, TM18, XKI^{+}11]. traits-based [GEEC14]. Traitspace [LJ15a]. trajectories [Bus17, GHST19, POB18, WB10]. trajectory [BTBM17]. Trans [YCF15, HRSKW15, RXSK14]. Trans-boundary [YCF15, HRSKW15]. Trans-Himalayan [RXSK14]. Transcarpathian [HOŠ^{+}12]. transfer [BBDB13, DBP16, GSPSP15, GGR^{+}13]. Transferability [BRV11, DLvGGNC16, FHAG12, LPT^{+}18, MAMNL^{+}15, WZL^{+}16]. transferring [AGQP^{+}14]. transformation [Abe15, BP15, MNEG11, MB17, PRHK19, SAAD16, ZFT^{+}10, ZJB19]. transformations [BP15, BP16, BP19, Fat14]. Transforming [KHHN17]. transformities [Abe10, LLCR10, Pat14b]. Transformity [Til15, CXL^{+}19, GDC^{+}19, LB18b]. transgene [GDT^{+}14]. transgenic [TWL11]. transience [GvdWBS11]. Transient [CGAB17, MM19, Sin17, PP16, WAM13]. Transition [NSN17, Bae18, BCV13, BH13a, BH13b, DBG14b, DBG14a, FLY13, FDR^{+}17, For11, IKK^{+}12, KP13, KF16, NF11, Phi11, SHBG10, VSF10, YZ18]. transitional [QCLW18]. transitions [GPM^{+}16a, GVV17, JDZC14, PMTKC17, RWFD13, dAdSC15, vLTdBP16]. translocation [HBT^{+}18, TL15]. transmission [BSCG10, FBB^{+}18, HJ10a, HTVE18, POS^{+}19, PMP^{+}13, SSQ^{+}10]. transmitted [BGG^{+}16b]. transmitter [PR13]. transparency [SAI10]. transpiration [MEC14]. Transport [NSN17, ABR12b, ABR12c, BKN14, BWA17, CNI11, CGG^{+}18, GJM^{+}13, GPD^{+}13, JA15b, LZYH19, MK14, MWM^{+}16, MCMT^{+}16, OdlFDP15, QDH^{+}10, SH14, SEHD18, WFW^{+}17, ZSG^{+}18]. transported [BEZ19]. trap [CGB^{+}13]. trap-effects [CGB^{+}13]. trapping [CGS^{+}17]. Träsk [WFM16]. traveled [RDJ19]. trawlers [RPP^{+}15]. trawling [ADL^{+}10, DLH^{+}14, TSN17]. treatment [MXC13, Mur11, SWZ^{+}13, TA11, YqWB^{+}15, ZF11]. treatment-resistant [Mur11]. treatments [DAA^{+}12]. Trebonsko [HV15]. Tree [MBHW12, AK17, BlMG^{+}11, BBL^{+}17, BFNH14, BFH14, CHG^{+}15, CRD^{+}16, CLBSG11, CW11, DAF^{+}11, DFRT14, EFM^{+}10, EKB18, FAS^{+}19, FMF12, FCV^{+}12, GHST19, GVGZ^{+}21, GLR^{+}19, GGSP14, GBP^{+}10, GHS^{+}10, GOSH^{+}12, GCIB10, HMP19a, HHW11, KAF^{+}14, KHS^{+}19, KGD14, KPK^{+}17, KR14, LBZA19, LHP16, LS10, LL18a, MBG^{+}10, MK12, MWML13b, MdMS^{+}19, NZL13, NLMdL10, NMS10, NP15, NPH^{+}17, OBLJ13, OJ15, PR11, PMS11, RFZB12, SZSV13, SVH^{+}19, SFL^{+}19, SNZ^{+}19, TCSG15, TKH17, TYEH10, WK11, YLL^{+}17, dJGM^{+}18]. tree- [FAS^{+}19]. tree-grass [YLL^{+}17]. tree-ring [EFM^{+}10, MK12]. Treefrog [ACC^{+}10].

Treeline [KWJH16, MU17]. Trees [AWWD16, DHC14, DSK^{+}11, DTS^{+}12, EBL^{+}11, Gav15, KKR19, KMM^{+}13, LKDD15, MROGVR12, MCS^{+}18, PVL^{+}14, SZSV12a, SH14, VCSSC12]. trembling [ARM^{+}12]. tremuloides [ARM^{+}12]. trend [Hui11, LZ10, RDD^{+}18, WJvdL^{+}13, YYL13]. trends [CHR^{+}13, CL15b, JMSM12, LMSK12, LW17, PAS^{+}10, RKRB16]. Tri [BPT^{+}19, CH19, SFHK12]. Tri-Annotated [BPT^{+}19]. tri-trophic [CH19, SFHK12]. trial [JAKJ10]. tributary [MLW^{+}15]. trichotoma [JBB^{+}11]. tridentata [SLB14]. trip [VRM^{+}10]. tripartita [DZS12]. TRIPLEX [WPZ^{+}11, ZXD^{+}13]. TRIPLEX-Management [WPZ^{+}11]. TRITON [MJL^{+}14, MJL^{+}13]. tritrophic [BWH13]. Trophic [BSASSJ^{+}13, BPGSS^{+}18, CD17, GAV^{+}11, LBD10, MDL^{+}10, URUNAM12, ZRCNLI^{+}15, AS14, Ban11, BFP^{+}18, BS10b, CH19, CGEdM12, DATI10, DLH^{+}14, DUASLB^{+}12, DBVS19, DCZ^{+}15, ELPK12, ESLJ18, FAMD^{+}12, FE14b, FSML11, GLS10a, GDTP13, HKL^{+}11, JTB^{+}19, JHH^{+}12, mLcLJ15, mLcLJ16, LVV^{+}15, MTZ^{+}10, OBC^{+}21, Pea15, RHR^{+}15b, RDR^{+}17, SB11, SLW^{+}14, SFHK12, SS11, SPP18, TW18, TNHB12, TAR13, TTCC^{+}19, ZB15]. trophodynamic [RDD^{+}17, VPB^{+}17]. Trophodynamics [GPR^{+}13, VASL^{+}15]. Tropical [FAC^{+}19a, BBRJ15, Bre12, CPR10, CGGL16, DH12, FSML11, FBH^{+}14, FBdP^{+}16, GHST19, Gat18, GGTP^{+}17, GANB11, HBG^{+}16, KK12b, LS10, LCZ11, LYL^{+}17, LCDJ^{+}12, NCEGB16, PDD^{+}18, PTE^{+}12, PMS11, PMBB^{+}16, PGA^{+}11, SBRD10, SLZ^{+}13, TDH12, VRC17, VRL18, ZNS^{+}18, AH17]. tropics [SV13]. trout [ARV^{+}16, CSR15, FB13, KGM12, MMMGLN^{+}18]. truncatus [BWC17, MMCBP19]. trutta [CSR15, MMMGLN^{+}18]. TSD [PL10a]. tsetse [LDM15b, Pec12]. TSPM [YCF15]. Tuber [YKL^{+}12]. tuberculata [RPA12]. tuberosum [KYvdPS14]. Tucuruí [DBK^{+}13]. tuna [DFM12a, DFM12b, MROGVR12, PAGCV^{+}10]. tunas [FE14a]. tundra [WR18]. tuning [AG11, NYIH17, SMI^{+}19]. Tunisia [HBS^{+}16a]. turbid [MJB^{+}11a]. turbidity [LBE^{+}17]. turbines [KHKC^{+}18]. TurbSeed [HRWT12]. turbulence [YLC^{+}18]. turbulent [HRWT12]. turfgrass [MKTO19]. Turing [FR11]. Turkey [KAF^{+}14]. turn [Ber13]. turnover [JDGJ13, LW17, MCR^{+}15, SKK^{+}13]. Tursiops [BWC17, MMCBP19]. turtle [HGWL11, HVL^{+}19, KWGG19, PL10a, PRH17]. turtles [AB13, DBP16, FP13, MJL^{+}17, PP19, WHRR15]. Tuscany [MJP^{+}12, MCBM12, MCPB15]. Tuta [RSCS19]. Two [FDHL17, MCT10, MPF^{+}14, YWY10, AR11, AHM^{+}16a, BVC^{+}18, BND19, BMCG12, CLL14, CSN^{+}19, CRA^{+}11, CCL^{+}13, DZS12, FMBD15, FSM^{+}19a, FAPM12, GTC16, GBP^{+}10, qHLlZ^{+}16, HB14, HQF^{+}17, JAH^{+}10, JGB15, KSM^{+}18, KPS19, KR14, KRCVF16, LDM15a, LkSXhX17, LZJ^{+}19, MJLC14, OCB^{+}13, OLG^{+}13, PKG16, PDBP11, RBM^{+}17, RW17, Rin18, RDH^{+}13, SELB10, SFR^{+}17, SRV^{+}12, SFHK12, TKP14, Tul14, WBS13, WJL14, WCZ11, YHW^{+}12, ZMS19, ZW11, ZJM^{+}11, ZSJ15, dHR18, Kal19]. Two-dimensional [YWY10, AHM^{+}16a, LkSXhX17]. two-layered [JAH^{+}10]. two-patch [LDM15a]. two-sex [TKP14]. two-sidedness [WJL14]. two-species [YHW^{+}12]. two-toxin [HQF^{+}17]. two-year [qHLlZ^{+}16]. two-zone [WCZ11]. type [BGMJ^{+}12, CCL^{+}13, CN10b, HS13c, HS13b, HWE^{+}18, SJGB16, SA16, WZWL10]. types [BC13, GANB11, HMP19a, HBA16b, LM13, LAD11, LS10, ROG^{+}16]. Typha [ESA12]. typhoon [PCTW17]. typical [BCC^{+}11, JHH^{+}12, JC11b, MLW^{+}15, mNkGyT^{+}17, SLW^{+}14, SYY^{+}19]. typographus [OFH^{+}19]. Tyrannidae [LIM^{+}16]. Tyto [dÁGP14]. Tyuleniy [ZKSF17].

U.S. [BTPG10, LJ13, PAS^{+}10, ZLB14]. U.S.A. [Tat16a]. Ukrainian [BRPK14, HOŠ^{+}12]. Ulanowicz [Kat11]. Ulansuhai [MWY^{+}18]. Ulex [RPGA15]. Ulleung [LY16]. Ulrich [FM11]. ultraoligotrophic [HC11]. unbalanced [FMF12]. uncertain [TB12b]. uncertainties [COY^{+}10, CLNN14, FCG^{+}19, JKS^{+}10, KPK^{+}17, LLZ^{+}16, OLG^{+}13, PNT^{+}19, THC18, ZHC^{+}16]. Uncertainty [Ant14, GTC16, GLY^{+}15, Ing10, JJMW14, JAKJ10, PAS^{+}10, RHS^{+}15, RHD^{+}16, ALG16, CS17, CC14, CJR^{+}16, CAMCK^{+}12, CTMS10, CBA16, CWMC^{+}12, DAA^{+}12, DHS^{+}16, DPBS10, DBJ13, GB16a, GDT^{+}14, GLC^{+}15, HLK^{+}12, HL19, HT14, ISG13, JSL^{+}10, JLZ^{+}18, KB10, LKMCW14, LA11, LkSXhX17, LHZN11, LLCR11, Log16, Log17, Mar12, MTJF11a, MDKK15, MSF18, MSHC13, OOFVBC16, OKS^{+}11, PSB^{+}17, PMCKH13, RHW17, RMTM^{+}16, RAF17, RHZ^{+}18, RR12, RIMPBT19, ST13, SMT19a, SMT19b, Sma13, VPM^{+}18, WC13a, WBB^{+}21, Wil11a, Wil12, YTT^{+}11, YLHZ14, ZBS14]. Uncovering [DS12]. under-bark [VCSSC12]. under-ice [AHS^{+}16]. underestimate [Log19b]. underlying [CLT^{+}11, GGSP14, PMF^{+}13, Sil11]. underplanted [DSS11]. understand [ACE^{+}17, FBdP^{+}16, FCA18, LS10, RCBPMN10, WHH^{+}18]. Understanding [ALC12, AH17, BU11b, FKSK^{+}10, KPK^{+}17, LV17, LBH^{+}15, LWTZ18, PMF^{+}13, TB13, WFW^{+}17, XZCL16, CGGL16, JM15, LWC^{+}18, LQW^{+}19, RDH^{+}13, VPC11, WJS^{+}18, Wei19]. understory [KLS^{+}14, XKI^{+}11]. Underwater [CLD^{+}17, FSN^{+}15, PC17]. unexpected [FJH15]. unfished [EPM^{+}14]. ungulate [DDM^{+}17]. ungulates [DDB^{+}12]. unicellular [GRD^{+}14]. Unified [HBP18, TT17a, TA11, YLF^{+}15]. uniform [MH13]. uniformity [MdMS^{+}19]. uninhabited [NCK10]. Union [LCBNP19]. unique [Bin17, Fis21]. uniquely [MCB14]. uniqueness [mLcLJ15, mLcLJ16]. unit [LDC^{+}16, PPF11, SFM^{+}18]. United [Gli14, HSR^{+}15, GZK^{+}11, CRD^{+}16, LLT^{+}14, LLT^{+}17, RWFD13, RTE^{+}15, SHC^{+}15, SBF15, SWR^{+}16, SB15]. units [Mul11, TB11, TB12a]. universalities [BC12]. university [WHH^{+}18]. unknown [Agu11, Dil10]. unmarked [CGS^{+}17]. unmodeled [DAP18]. unobserved [BLFS18]. unravel [VPM^{+}18]. Unsuccessful [DCCS17]. unsupervised [ASL13]. Unveiling [CMM^{+}10]. unwittingly [RDV^{+}17]. update [Fat10, Fat12, GBD^{+}10, HGLZ13]. Updated [BU10]. Updating [GGV10]. upgrades [FN18]. upland [DBG14b, DBG14a]. upon [RLHD11]. upper [JJYM19, DWNN10, DVF^{+}19, PSBH16, dFMdSA18, KH11]. Uppsala [NPP^{+}18]. Upscaling [NPH^{+}17, GWC^{+}10, MGS16]. upstream [MWK19, YXD^{+}18, ZVS18]. uptake [AJDS16, HCL^{+}14, LNB17, LPD14, LB18a, MCH15, MBW14, NP15, NPH^{+}17, SAAD16, VRL14, ZLFI11]. upwelling [BCO17, OCCF17, RWP11, RBGB16]. Urban [CL15a, HLW15, NLM^{+}19, Wei19, ZF19, ACC^{+}10, Ano12t, BLOJ^{+}19, CWLL15, Che15a, DLZ^{+}17, DLD^{+}11, DWP13, EBA^{+}19, ESP^{+}17, FKWZ19, FDR^{+}17, FCZ13, GWHL15, GCV14, GWZ^{+}15, GLI^{+}21, Hal11, HCM^{+}15, HJ17, HZZ^{+}18, HM18b, IBD^{+}19, Ji15, JZG10, JMW11, JC11a, JDT^{+}17, KRE^{+}18, LRM^{+}14, LkSXhX17, LYSW15, LHL18, LWTZ18, LYCZ11, LYCZ13, LCP14, LYCU14, LL18b, MD16, MGR^{+}18, MJLC14, MGS15, mNkGyT^{+}17, TJYZ11, WZG15, WFM16, XHC^{+}15, XZSL17, XJY^{+}16, YKZZ14, YCF15, YZPL16, ZYFL10, ZLF^{+}11, ZYLY11, ZLC13, ZWG^{+}13, ZLF14, ZLF^{+}15, ZLFZ16, ZLZ17b, ZWF18, ZTL^{+}17, ZHCX15, ZMH19, MSS^{+}19]. urban-rural [MJLC14]. Urbanization [YWH15, CLHH19, LLQ^{+}15, TLL^{+}15, TOQW11, ZWG^{+}17, ZSKI17]. urbanized [SDY^{+}15]. urbanizing [LSZC10, WZG15]. urchin [RCP17]. urea [VRL14]. urease [TB19]. ursinus [ZKSF17]. Ursus [DFK^{+}11, MSH^{+}19]. urticae [dOMZ13]. Uruguay [dFMdSA18]. USA [BN10, Jør14a, CPS15, DVF^{+}19, DLH^{+}14, DWP11, FP17, FYWA12, FWRZ10, HSTS19, JMB^{+}11, KB10, KBS^{+}15, KPT11, KMP14, LdMS^{+}16, LXS^{+}16, MSB^{+}17, MR17, MJLC14, OC15, PANL11, RR10, SHX^{+}12, SGDP19, TOQW11, WMW^{+}12, WTG^{+}16, WTC^{+}18, WPB14, ZF11]. Use [BVC^{+}18, BFJL15, CCC12, JN14, LM12, LCD^{+}16, LSJG17, MGGLBIBD13, SZSV12b, TF13, ZSG12, BSC^{+}15a, BPR^{+}11, BPC^{+}17, BKMK10, BK13, CLH17, CZ11, CC11, CPS^{+}14, DK17, DMB^{+}18, DSM^{+}12, DPLC13, DGKSM13, DSA^{+}18, ER13, Fel12, FCFT16, GAP^{+}19, GK14, GMP13, GFP^{+}16, GLI^{+}21, GS19c, HLZ^{+}18, HHL17, Hoa11, HNWT10, IKH^{+}11, JBV^{+}15, JJYM19, KMRW16, LBA14, LBA15, LBE^{+}18, LY11, LCL^{+}13, LBK^{+}13, LWL^{+}17b, LLZ12, LOLA13, LZZW17, LBC^{+}15, LCDJ^{+}12, MVC^{+}17, MDOJ14, MLG^{+}18b, Oga11, PR13, PK16, RLHD11, SJGB16, SGA^{+}16, SDD^{+}13, SMH^{+}12, SLH^{+}11, SRTS18, SWR^{+}16, SDY^{+}15, TBL^{+}15, TB19, VR12, VGR^{+}11, WHB12, WO14, WBG^{+}14, WS15, WSE10, WHNR19, WFM16, XLS11, XLR11, YWH15, YZL12, YLW14, YZC14, YXD^{+}18, ZFT^{+}10, ZZFH15, ZZJ^{+}16, ZJP^{+}11, dFMdSA18, dLB19, dSD14, vDSC^{+}10, vVBHZ11, vVHZHvD13]. use/cover [LWL^{+}17b]. use/land [ADF^{+}17, XLR11]. useable [Ken17]. used [BTM^{+}10, CD16, CBA16, FML^{+}17, GWR11, HM17, Löp11, MBLC14, RPGSCF16]. useful [MTA19, PVB16, QL15, SBD^{+}19]. usefulness [HJN12]. users [CBC^{+}16]. uses [WXB10, ZQC15]. Using [ARM^{+}12, BS10c, CDR^{+}13, CPD^{+}11, CMGF14, DTS^{+}12, DLH^{+}14, DDR^{+}18, EBL^{+}11, FRV^{+}16, FCLFB18, FCFT16, Fre16, FP13, HAOS17, JBE^{+}15, JGX13, KGM12, KHHN17, KKI^{+}14, LSR16, LQW^{+}19, LL18b, MTZ^{+}10, MV11, MU10, MDH15, MCPB13, MCMT^{+}16, NSN17, OOFVBC16, RM13, SEHD18, SGRAdJ17, TPG14, TNHB12, Tul14, VSF10, WLY^{+}15, WV12, ZSG^{+}18, dMLM^{+}17, AGJ^{+}12, ABB^{+}10, AAO10, AS11, AYASR^{+}17, Ana17, AHNK13, AHS^{+}16, AAFH15, BDL^{+}13, BHD16, BCB^{+}11, BB10, BC12, BLBB13, BSH17, BBL^{+}17, BT16, BFNH14, BBGJ13, BMCG12, BPC^{+}17, BJR16, dÁGP14, BB18, BQF^{+}18, BBH^{+}17, BK13, BLAV18, BPD^{+}19, BRB14, BLCPB11, CF14, CLBG15, CC10, CBGK15, CZ11, COPP13, CN11, CMA^{+}13, COK^{+}11, CGAB17, CAMNN16, CAY^{+}10, CTMS10, CDL^{+}11, CBP^{+}17, CVG^{+}15, Con14, CSFR15, CRQW18, DJT^{+}19, DSK^{+}11]. using [DMDJ13, DSM^{+}12, DS21, DBP16, DG16, ETG^{+}18, EST^{+}19, FTK^{+}14, FSB13, FLH17, FIG^{+}10, FT16, FAPM12, FMLM11, FBKS17, GCDA16, GWHL15, GGV10, GKF^{+}14, GPCG15, GE17, GOSH^{+}12, GAF^{+}14, GNB^{+}14, GHS^{+}16b, GCW^{+}16, GCvDM12, HHC^{+}13, HBS^{+}16a, HCW^{+}17, HPJ15, HVL^{+}19, HSK17, HMS^{+}11, HS11, HTL^{+}14, HCL^{+}14, HJS^{+}16, HSHF^{+}13, HHC15, HVCP16, HMA10b, HWL17, HNWT10, HGLZ13, HBA16a, HBA16b, HAA^{+}21, HT14, IPSH14, IT17, JVY19, JR16, JG10, JHH^{+}12, JFS^{+}12, JMdW^{+}10, JDP^{+}12, KSS11, KW18, KCF^{+}13, KSZW16, KKL^{+}12, KPM^{+}14, KLS17, KC18, LRZ12, LCCL15, LJ15b, LJ15c, LLKY18, LWY^{+}11, LPV11, LLC11, LVZ^{+}11, LPX^{+}12, LLP17, LAK^{+}18, LCMG18, LRB12, MK13, MLK^{+}11, MAO19, MYHN17, MMS12, MKSL^{+}17, MK12, MJLC14, MBM^{+}18, MJF19, MRR^{+}14, MSHC13, MdMS^{+}19, MAMNL^{+}15, MDVG^{+}15, MSO10, MCP16, NP18, NOOM^{+}16, NMS^{+}14, NB11, NNP19]. using [NiSK21, NJCF11, OKS^{+}11, PSB^{+}17, PC17, PMS^{+}17, Pat12, PWGS^{+}17, PD15, Pér18, PWP^{+}19, PMRGF10, PKW^{+}10, PvVC13, PICR13, PSS16, PSKO16, Qia12, RVL^{+}18, RPSA12, RvS16, RWS^{+}16, RGHS11, RAR15, RBRK16, SCB13, SZV14, San17, SNS^{+}12, SHT^{+}21, SMI^{+}19, SØB14, SHM^{+}14, SCHGM16, SDEP18, SBC^{+}15, SGP18, SSM^{+}13, SFB^{+}15, SM14, SSW^{+}18, SGKR10, TSN17, TKW17, TB11, TB12a, TCSG15, TQZ12, URUNAM12, VRM^{+}10, VALC12, VB10, VASH^{+}16, WDL^{+}15, WTWS19, WLR^{+}11, WFA^{+}13, WBG^{+}14, WGD^{+}17, WCG^{+}11, Wil11a, WKH19, WX11, XZCL16, YKL^{+}12, YKZZ14, YLF^{+}15, YZPL16, YRGH16, YHW^{+}12, YLL^{+}12, ZLY16, ZH11, ZZJ^{+}16, ZZ17, ZCS13, ZZT13, ZHLZ17, ZSKI17, dJHMS19, dMSL^{+}16, vdKGV11, vdVBJS15]. utility [AMH16, KA17, LYCZ11, TWP^{+}14]. Utilization [BRL12, BFKS16]. UV [MRCW10]. UVAFME [WSL17]. UVAFME-VOC [WSL17].

v.2 [SGB10]. v1.0 [WSL17]. VA [JMB^{+}11]. vaccination [CESC19, SM11]. Vaccinium [BHD17]. valid [LM16]. validate [DLH^{+}14]. validated [ASV19, MCH15, TWL11]. Validating [CVPR^{+}16, DFRT14, PWK^{+}17]. Validation [BBL^{+}17, BH10, FSS^{+}17, HGN^{+}16, KCF17, KCF18, NB11, OCCF17, AVG14, BlMG^{+}11, BRM^{+}14, HD16, HKlL14, HSVL12, JSdSdL^{+}18, KHW^{+}16, KSAS12, KHW^{+}17, MJL^{+}13, MJL^{+}14, MPL^{+}16, NGWZ16, OGOWW16, RWP11, SW11, SM13, WB10]. validity [KRL11]. Vallès [GPM^{+}16a]. Valley [FCFT16, TW18, MMS12]. valuable [FMBD15, GWR11]. valuation [LLZ12]. value [BFPR16, BvHR^{+}15, BPD^{+}19, CSF11, FAC^{+}19b, GBL^{+}13, JJMW14, MRES17, OT15, PPF11, RRBI14, RAB^{+}10, TAGC^{+}16, VPB^{+}17, WEB11, XHC^{+}15, vGKB10]. valued [SDY^{+}15]. values [DLL18, GFC15, HCS^{+}10, Ing10, KGD14, RHR^{+}15a, SLH^{+}11]. Valuing [VLZ^{+}19, NPP^{+}18]. vapor [MKF^{+}11, QlSeZ^{+}10]. variability [AK17, Ash10, ABD14, BDL^{+}13, BA14a, BMD18, CMT^{+}11, CdMMF16, CAC^{+}16, CBA10, CRQW18, CQS15, EC10, FHAG12, FPG11, GNB^{+}14, HM12, HBG^{+}16, JVY19, KM12b, KK15b, LY16, LDO^{+}18, LS13, LC11, MAM^{+}12, MK12, MLH^{+}17, MB14, NZL13, PVBC14, PRP17, PICR13, RNE^{+}12, SC18, SPG11, Uch16, VASL^{+}15, WLS^{+}10, WLS^{+}11, ZWA19, ZZT13, ZR19]. variable [CGLA10, CMT^{+}11, GO15a, GRD^{+}14, HKM^{+}18, HJN12, JFS^{+}12, Log13b, MKL14, MMC^{+}19, MRWN19, MFI13, PBO19, SKHH19, WB14, WGE^{+}19, XLR11, ZBMC14, ZLD^{+}19]. variables [BBH^{+}17, CKMD14, Fis21, HL19, IGR^{+}18, KHJ^{+}19, LHR^{+}15, MDH15, PSKO16, SZSV13, WRB^{+}12, WSBP13, ZLY16]. Variance [RHvdTS14, Sma13, SBPZ12, WBB^{+}21]. Variance-based [RHvdTS14, SBPZ12]. Variation [LXS^{+}16, SBM^{+}12, TLL^{+}15, BSFR17, BPWB14, BRS^{+}18, DCCS17, EMH10, FCS13, FS18, GHH^{+}15, Gre12, HCP10, HSM^{+}12, yHAB12, LMSK12, MDB19, MCRB17, PL10a, SNS^{+}12, SW19, SFMA15, SEHD18, SML12, WGA^{+}11, WGA^{+}12, WTG^{+}16, WWC^{+}18, WHNR19, XCR^{+}13, ZBSF17]. Variations [Kal19, ChHC12, EFM^{+}10, FS14a, MOM^{+}11, Nak17, NiSK21, PVM12, PVM13, SBC^{+}16a, TMC^{+}16, UDRA13, WLY^{+}15, XWC^{+}11]. varies [MGB13]. variety [BS13]. variograms [PS13]. Various [AMRH19, LWY^{+}11, MKTO19, NGGFM19, OJ10, WGSR19, XSL18, YHK^{+}19]. Varroa [Vet12]. Varying [SSBJ12, CTNH13, DTC^{+}10, KSS11, KD13, LBE^{+}17, LMSK12, LB15, Sol19, TT17b, dLB19]. vector [BC12, EPA^{+}10a, HM18a, HM19, LBR16, MD16, MMVAHMC16, SZSV12b, TF13, ZYXT18]. vector-borne [LBR16]. vegetable [LHB^{+}18, SJS^{+}13]. Vegetation [BKB10, BvBB10, YCL^{+}14, AD13, BBDB13, BLD^{+}10a, BlMG^{+}11, BC13, BSS^{+}14, BFKS16, CRAP13, CHPS17, CLCC11, CN13, CSFR15, CV19, DSK^{+}11, DVCM^{+}10, DS21, FKSK^{+}10, FBDZ12, FPG^{+}10, GGM^{+}12, GMLRF10, GLRT16, HHC^{+}13, HOŠ^{+}12, HBL18, JWX18, JDGJ13, KMF^{+}21, KP13, KBS^{+}15, KF16, LJBS12, LKH^{+}18, LH15, LKG^{+}17, MBHW12, MKF^{+}11, MVC11, MPG18, NJM11, OC15, PRHH13, PBG^{+}17, PBFS14, PLL^{+}19, RPGSCF16, SS16, SBF15, SMH16, SV13, Tie16, Tul14, UR10, VB10, WJL^{+}13, WZGH18, Wol11, XGH^{+}17, YLZ15, YL18, YLL^{+}12, YFL^{+}18, vLTdBP16]. Vegetation-specific [YCL^{+}14]. vegetative [Hos11, IWW12]. vehicle [ACSR^{+}13, GJB^{+}11]. velocity [LBW18]. Venezuela [RDR^{+}13]. Venice [BSZ^{+}11, VZF^{+}11]. vent [HSvO^{+}18]. Ventotene [FBDR17]. Verhulst [MV17, Ned15]. verification [OGOWW16]. verify [CS17]. versatility [GLA11]. version [SYL^{+}13]. versus [BBH^{+}17, DJH10, LJ13, LCDJ^{+}12, MMM^{+}16, RUSL19, RAB^{+}10, Sto10, WB14, WEJ11]. vertebrate [BNGP19, PDV17, TV15]. Vertical [BM16, OAPP12, XHC^{+}15, BVJE18, BP16, BP17, GC14, LY16, MPM^{+}15, MUP^{+}13, MB17, MN12, SAI10]. vertically [BBH^{+}11, For14, LKLK15]. very [FRM^{+}19, LAC10, YHS^{+}11]. vessel [DMH17, WB10]. vessels [AWF10]. VGEs [ZCL^{+}16]. via [BA17, IHSM12, Jag14, VHM^{+}17, dSdCF10]. Viability [KCKdOW12, WAM13, CJM^{+}12, HCW^{+}17, KKR19, LN16, MCRB17, MSB^{+}17, MATS14, RJG^{+}13, ROG^{+}16, TWS^{+}11, WAGM17]. viable [VEK12]. Victoria [LWNG13, NOOM^{+}16, ZDG^{+}13]. victoriae [AR11]. view [DPR10, DUASLB^{+}12]. viewpoint [AG15]. villages [BMMC21, FMBD15]. vine [Gat18]. vineyard [Pie10a]. Viral [KGM12]. Virgin [CNI11, EKB18, MRG13a]. virginianus [WTG^{+}16]. virginica [LLC^{+}17]. viridana [KSS11]. virtual [BGGL16, HKVD16, OMH12, PC17, SDEP18, ZCL^{+}16]. virtually [Bin17]. virus [FBB^{+}18]. Viruses [KGM12]. vis [RSCS19]. vis-à-vis [RSCS19]. vision [GAB12]. visitation [OBSL15]. visitations [RW10]. visited [BRL12]. visual [FYR^{+}18, PC17, PC15]. Visualization [FKDG14, SSM^{+}13]. vital [DCCS17, PRL^{+}13]. vitellogenin [MPK18]. VMS [JBCÑ11, VRM^{+}10, WB10]. VOC [WSL17]. volatile [WSL17]. volatility [LW17, MDB19]. Volterra [For18a, GvdWBS11, TSM^{+}15, WSSM15]. volume [DBP16, PDD^{+}18]. volumetric [LKP13]. voluntary [BA17]. volunteer [MGSB11]. Voronoi [SvdR10]. vs [BVVS15, BOW^{+}12, BRF^{+}11, Con11, Con12, Gra14, KWH14, LTL15, LM19, Log19a, MWK19, OOFVBC16, Pér18, RHS^{+}15, RHD^{+}16, Ten15b, TS11]. vulgaris [CDT10, CDT10, FLLB11]. vulnerability [ASP^{+}15, CLH17, CW11, RCSS13, Sol19, SCBH18, SH18, ZLZ^{+}17a]. vulnerable [CHTC11, CSFR15, IPSH14, ZZA^{+}18].

W2 [ZSJ15]. WA [KR10a]. Wabash [MCP16]. wader [SDT10]. wahoo [MROGVR12]. walk [PvVC15, PPBD13, SA16]. Walker [PvVC13]. warbler [DHS^{+}16]. warm [GSF^{+}16, CBB^{+}10]. warm-water [GSF^{+}16]. warming [BH13a, BSH17, BSHG16, CSR15, CN13, DBS18, FAS^{+}19, GLR^{+}19, GHP12, Gra14, HHK19, HEJP17, KTK^{+}15, NGL13, OQH^{+}13, VPS10, ZJL^{+}14]. warning [Bae18, LSZC10, OFGH^{+}19]. wars [MR10]. Was [BPB14]. Washington [PAB13a]. wasp [GPTR13, PFP^{+}16]. wasps [SFHK12]. waste [GRRU15, KS18, MWM^{+}16, PAGCV^{+}10, RPC^{+}17, XZCL16, ZHCX15]. wasteland [Ber13, WWYO17]. wastewater [BKN14, dABLBAdOR17, GJÁ^{+}10, PLFO^{+}12, SWZ^{+}13]. wasting [HMPC11, PMP^{+}13]. watching [RIMPBT19]. Water [Jam16, Ju19, KS17a, KS17b, LZJ^{+}19, MOMC17, Mur12, NGGFM19, PRHK19, WRMB18, XYTS19, YZPL16, BA14a, BAB^{+}16, BC13, BT16, BDTV11, BKB10, BvBB10, BMU10, CZC^{+}15, CW17, CHR^{+}13, CZ11, CSR15, CN13, CDG15, DC10, DBG14b, DSS11, ER13, FKWZ19, Fel12, FCH^{+}17, FP17, FLB^{+}15, FAC^{+}19a, GSF^{+}16, mGZK^{+}10, GA15, GPCG15, GMP13, GWK^{+}12, GS10, GCL^{+}19, GEB10, HKL^{+}11, HCM^{+}15, HLZ^{+}18, HCL^{+}14, HV12, HGBA^{+}13, HC11, HhLY^{+}19, HBA16a, HBA16b, HYL^{+}17, HAA^{+}21, JZG10, JMW11, JLZ^{+}18, KPM^{+}14, KLS17, KJS13, LBZA19, LHL^{+}17, LY11, LWGH13, LLQ^{+}15, LZL^{+}18, LBW18, LHB^{+}18, LPV11, LKP13, LGY10, LHR^{+}15, LPD14, LB18a, LL18b, MTZ^{+}10, MBSW12, MBW14, MKF^{+}11, MGR^{+}18, MA10, MVC^{+}17, MGW^{+}13, MH10, MLH^{+}17, MB14, MMRS^{+}19, MCP16, NLM^{+}19, NGWZ16, NiSK21, OSS13, OdlFDP15, PAT17, PBFS14, PANL11, PPF11, QlSeZ^{+}10, RLB^{+}12]. water [RAF17, RZF^{+}13, RHZ^{+}18, RCCN11, RMLTCQ^{+}19, RPGSCF16, RPP^{+}15, SZSV12b, SFMA15, SMH16, SKTG11, TSB^{+}18, TBL^{+}15, TGB^{+}17, UR10, WMW^{+}12, WKZ^{+}19, WYY^{+}19, WO14, WXB10, WV12, WRSL11, WFSH15, Wol11, WZG15, XLS11, YqWB^{+}15, YWH15, YY11, YY13, YLL^{+}17, ZHC^{+}16, ZWRK13, ZLLC13, ZLZS14, ZSJ15, ZLJM17b, ZLJM17a, ZTL^{+}17, ZW10, ZJP^{+}11]. water-controlled [RPGSCF16]. water-energy [FKWZ19]. water-environment [ZHC^{+}16]. water-limited [CN13, YLL^{+}17]. waterbird [GDW12]. waterbirds [WRSL11]. watered [CCL^{+}13]. waterfowl [ATL18, BKN21, MRD^{+}12, MRES17, PSBH16]. Waterfowls [JJKJ11]. waterholes [FKSK^{+}10]. waters [GEB10, HSCS14, NvGZ^{+}16, RRHH10, SØB14, WLZC11, WNT^{+}13]. Watershed [DZW^{+}11, ER13, AAO10, AADA14, Ano12s, CL10, HKVR13, KJMA11, MCW13, NXZ18, PAT17, PPF11, YDZ^{+}21, ZF11]. watersheds [GTC16, MC11, ZJM^{+}13]. watsonii [GRD^{+}14]. wave [JGB15, KL12, LZW^{+}17, ZLFI11]. wave- [ZLFI11]. wavelet [SLB^{+}18]. wavelets [ROC^{+}13]. Waves [CLMA19]. way [Mur12, Vav10, WYPF15]. ways [MCT10]. Weak [MJC^{+}11, WLR18, BDG^{+}14, RWH^{+}14, DJH10]. weaken [CHZ18]. wear [SM10]. weather [CVV^{+}13, CCSF11, JCKG15, KCF^{+}13, SPG11, WGA^{+}11, WGA^{+}12, WGA^{+}13, WLS^{+}10, WLS^{+}11]. Weathers [Jør14b]. Web [HhLY^{+}19, ABAG14, BSH17, BSZ^{+}11, BPGSS^{+}18, BRPK14, CGGL16, CKS^{+}17, CPS^{+}19, CPS15, DCZ^{+}15, ESLJ18, FSML11, Fuj16, GLC^{+}15, GYL^{+}13, HBS^{+}16a, HCL^{+}16, KKH^{+}17, KSAS12, KEM^{+}16, KHL^{+}16, LJR14, LSR16, LBSB^{+}14, LLKY18, LLFL12, LVV^{+}15, MKL14, MDL^{+}10, MSFR15, NOOM^{+}16, NCMF10, NCEGB16, NMC^{+}14, NMS^{+}14, OGG^{+}16, OHS^{+}17, PJ17, PTL14, PICR13, PMBB^{+}16, RGU^{+}15, RMR^{+}16, SVD^{+}13, SBR^{+}19, SFHK12, SKM^{+}16, SCG^{+}13, SS11, TRD^{+}15, TNHB12, TCH^{+}13, VASH^{+}16, WTWS19, Pau14]. webs [AdMC^{+}13, BNMP11, BCV13, BPK^{+}12, CR12a, FE14b, KCC16, LCTH12, RBM^{+}17, SBVA^{+}13, SGAS14, TPL^{+}15, TAR13]. Webworld [AMG19]. Weddell [BTB17]. weed [BGMJ^{+}12, CDT10, GDC11, GDC12, Gat18, MSKS13, MJGC13, RPA^{+}18, SPWR17, SMKM19]. weeds [DZS12, MMC^{+}19]. weedy [CDT10]. Weibull [WXB10]. weight [JG10, Sim11]. weighted [WFA^{+}13]. Weights [YKL^{+}12, RPPW16]. well [SHC^{+}15]. wellbeing [ESP^{+}17]. Wenchuan [LLL^{+}10]. Wenyuhe [RZF^{+}13]. West [BCB^{+}11, KCF17, KCF18, ABGB15, BMMC21, BBC^{+}19, BTBM17, DA19, SS16]. Western [QDH^{+}10, ARM^{+}12, BKP10, Bar11, CCCP17, EGMW12, GTC16, HCS^{+}10, HHC15, JXL^{+}15, JHHS17, LMBL12, dSD14, ADPS10, LMBL12, MMM^{+}16, PKB^{+}18, TTL^{+}18, URR^{+}17, TCH^{+}13]. wet [LAC10, MXC13, ZCYL15]. Wetland [AK17, CCLX17, WMW^{+}12, YM21, ACK^{+}14, BCC^{+}11, CHTC11, CBGK15, CZW10, CCC11, CZ11, CL15b, CDG15, CGM^{+}14, DTC^{+}10, DLD^{+}11, HZT^{+}13, LLM^{+}13, MLMFN^{+}15, NXZ18, OSS13, PJ16, PLFO^{+}12, PK14, QCB^{+}11, RHB14, RZF^{+}13, RSC12, RSC13, SZSV12a, SZSV12b, SZSV13, SZV14, SWZ^{+}13, SXZ^{+}15, SGKR10, TXM^{+}10, WJL^{+}13, WCZ11, XTPT11, YDT^{+}13, ZC11, ZWJ^{+}12, ZHC^{+}16]. Wetland-tree [AK17]. Wetlands [HV15, Jør10a, CZCL18, CHVH13, DA17, JDGJ13, LHZS10, MLB^{+}16, MXC13, MB17, RSC12, RSC13, SKH^{+}13, Yan11, YC11, YY13, ZZA^{+}18, MGMR15]. wetlandscapes [ZZA^{+}18]. WH8501 [GRD^{+}14]. Whale [RIMPBT19, CLD^{+}17, PWGS^{+}17]. whales [CMA^{+}13, RMTM^{+}16, SGSR10, WCG^{+}11]. wheat [AWL^{+}13, ADPS10, GLS^{+}17, LYD^{+}17, PVBC14, PZM13, SMKM19, WPZ^{+}14, ZBS14]. wheat-belt [ADPS10]. Where [AG14, Nie16, RA13, BL15, DFH^{+}14, GHWG^{+}14, vdVJS16]. Whether [Gav15]. Which [BGD^{+}16, CHPS17, MFI13, WXB10]. while [HKM^{+}18]. Whistler [MSH^{+}19]. White [Bec10, LSJG17, ABK^{+}16, BTBM17, CESC19, LJ15b, RTE^{+}15, WTG^{+}16, WR13b]. White-fronted [LSJG17]. white-nose [CESC19, RTE^{+}15]. white-tailed [WTG^{+}16, WR13b]. whitebark [MSM^{+}17]. whitefish [FMG^{+}16, VVH12]. Who [SM10, TL15, dSJHG10]. whole [GS10, LBZA19, LBSB^{+}14, TTCH^{+}14, TM14, THBF^{+}17]. whole-ecosystem [LBSB^{+}14]. whole-profile [TTCH^{+}14]. whole-tree [LBZA19]. wholeness [Fat14]. wholeness-extending [Fat14]. WICKED [BFR17]. wide [AVL^{+}11, BPK^{+}12, DHS^{+}16, FGH^{+}10, GBB13, MvDL^{+}10, SHT^{+}21]. wide-ranging [MvDL^{+}10]. Widening [GPM^{+}16a]. width [CH12, SRXP14]. wild [CHvdW^{+}10, DS12, GDT^{+}14, HA18, PKB^{+}18, RCS^{+}19]. wildebeest [War10]. wilderness [SRP18]. Wildfire [SCB13, ZJG10, ABB^{+}10, ABD^{+}18, FAS^{+}19, KK10, KUK11, RCBPMN10, TTL^{+}18, WLS^{+}10, WLS^{+}11, ZMMHR10, MRC19]. wildland [MRC19]. Wildlife [ACSR^{+}13, SZSV12a, BG15, BFJZ^{+}11, BTPG10, CLT^{+}11, Flo18, LAK^{+}21, MSMM11, SHCR10, SK11, SvdR10]. wildlife-habitat [BTPG10]. Will [PW10b, Fat16b, GAG^{+}12, vdVJS16]. Willow [BSP12, HPS10, KØ13]. Wils [KKP^{+}17, LGDW12]. wilt [NPJ^{+}17]. Wind [JX17, DMB^{+}18, DYMYGO19a, DYMYGO19b, HRWT12, KHKC^{+}18, SBM^{+}11, SHB^{+}11a, SMKM19, TKN14, WGE^{+}19, ZWJ^{+}12]. wind-assisted [SBM^{+}11]. wind-borne [WGE^{+}19]. Window [Kat11]. WINDROT [HOKH18]. windrow [BT16]. Windthrow [AMR16]. Wing [LAR^{+}16]. winged [FDP^{+}10, KK12a]. Winners [KTV^{+}15]. winter [GLR^{+}19, GHS^{+}10, KKBZ17, LYD^{+}17, MCWG21, MAC^{+}12, PKT15, SMKM19]. wintering [EGMW12]. wintertime [CQS15]. within [ARM^{+}12, BA14a, BLD^{+}10a, BlMG^{+}11, BGC10, CF17, CCC12, CLNN14, DJA^{+}10, GNB^{+}14, Hal11, HM17, JBW^{+}17, KYvdPS14, KPTC14, LHO^{+}16, LWKS10, MK13, MYHN17, MJC^{+}11, MRR^{+}14, OC15, RCSS13, Tul14, UB12, dBGS^{+}14, dOAC18]. within-field [GNB^{+}14]. without [BS16a]. Wivenhoe [CRO13]. WOFOST [GCCB18]. WOFOST-based [GCCB18]. wolf [CWL^{+}16, JSW^{+}13, LDM15a, MAM^{+}10, SGDP19]. Wolong [LLL^{+}10]. Wolonghu [CZ11]. wolverines [MvDL^{+}10]. wolves [SGWV15]. Wonderforest [TB13]. Wood [TLRL11, LSP^{+}10, LSGT13, LBS^{+}14, PSWM^{+}10, dJGM^{+}18]. wooded [IBD^{+}19]. wooden [SBC^{+}11]. woodland [DHS^{+}16, HMA10a, LHP16, PGR^{+}16]. woodlands [PBG^{+}17]. woodpeckers [VSPHS15]. woodpigeon [OFK^{+}12]. woodpigeons [KKI^{+}14]. woody [HHUD14, RWFD13, WMF^{+}13]. word [Rau11]. work [BPU11, Jør15a, Jør15b]. worked [PTL14]. working [ZWRK13]. workshop [WLF^{+}16]. world [ABB^{+}10, CC11, Dol10, ISG13, KR10b]. worlds [AA12]. worldwide [JTB^{+}19]. worth [PDBP11]. Wuhan [CWLL15]. Wular [MSKS13]. Wuxi [LWTZ18]. Wuyishan [LYC^{+}15a, LYC^{+}15b]. WV [ZF11].

xenobiotics [GFP^{+}16]. Xiamen [LSZC10, YKZZ14]. Xiangxi [CMLY19]. Xilingol [SAW^{+}11, YSA^{+}14]. Xinanjiang [YDZ^{+}21]. Xinjiang [DYU^{+}12, FCG^{+}19, LZLC13].

Yang [XWW^{+}11]. Yangtze [GWZ^{+}15, GLS^{+}17, GYL^{+}13, MCL^{+}17, TJYZ11, WSW^{+}13]. Yasso07 [DFRT14]. year [BGGL16, GZK^{+}11, qHLlZ^{+}16, LJ13, NB11]. years [Ano17a, CVV^{+}13, CHK^{+}12, GAC^{+}14, JASW18, LBC^{+}15, NJCF11, VASH^{+}16, ZKSF17]. Yellow [DSW^{+}17, JHHS17, LLM^{+}13, LC11, XLS11, YYL13, Yan11]. yellow-billed [JHHS17]. yellowfin [KHW^{+}16]. yerba [GCR18, MRB14]. yerba-mate [GCR18, MRB14]. yield [BDL^{+}14, BS13, FYWA12, KG13, KYvdPS14, LG10, LKG10, LG12, LZZ^{+}14, MCBM12, SX11, Tat16a, Til15, VZF^{+}11, WC19, XBvdV^{+}14, ZCS13, GBP^{+}10]. Yield-SAFE [GBP^{+}10]. yields [AVL^{+}11, For18a, IGBSASSJ16, Pog11, PPDF14, SKM^{+}16]. Yinma [ZLZ^{+}17a]. yolk [JNHF18]. yolk-feeding [JNHF18]. York [HSR^{+}15, WGD^{+}17, NJCF11]. youngest [VVH12]. Yser [BVP^{+}13]. Yuan [WhFxGZ16]. Yushania [RBRK16, SGP18]. YY [EEB^{+}17]. YY-males [EEB^{+}17].

Zambezi [MMS12]. Zealand [JBB^{+}11, MP17, PNHL14, PKD11, RFV^{+}12]. Zebra [CPS15, Lin15b, KVP16]. Zebrafish [KNU^{+}15, HTL^{+}14, LLC11]. ZELIG [LAD11]. ZELIG-CFS [LAD11]. zero [CMT^{+}11, LBE^{+}13, LJ15c]. zero-inflated [CMT^{+}11, LBE^{+}13, LJ15c]. Zhangziying [LLW^{+}15]. Zika [ZYXT18]. Zoige [MLB^{+}16]. Zone [SPM^{+}18, SEHD18, TZS^{+}14, BH13a, BH13b, CL15b, DBG14b, DBG14a, DWP16, LSR13, OQH^{+}13, PVBC14, QCLW18, UDRA13, WCZ11, ZSKI17]. zones [Ber13, DBG14b, DBG14a, FLY13, LLQ^{+}15, OTA16, RV13]. zoning [LLP17]. zoonotic [MC11]. zooplankton [CRUF13, CN10b, JA15a, MF12, PAB13a, QDH^{+}10, SAI10, TM18, TBM^{+}14]. zooxanthellae [BMR^{+}18]. Zostera [Hos11, KR10a, KvTL^{+}18]. Zucc [Oga11].

References

Aitken:2012:FWE

[AA12] Stuart C. Aitken and Li An. Figured worlds: Environmental complexity and affective ecologies in Fanjingshan, China. Ecological Modelling, 229(??):5–15, March 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003048.

Ali:2013:SCB

[AAA13] Emad Ali, Mohammed Asif, and AbdelHamid Ajbar. Study of chaotic behavior in predator–prey interactions in a chemostat. Ecological Modelling, 259(??):10–15, June 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001324.

Ahmadi:2014:MMS

[AADA14] Mehdi Ahmadi, James C. Ascough, Kendall C. DeJonge, and Mazdak Arabi. Multisite-multivariable sensitivity analysis of distributed watershed models: Enhancing the perceptions from computationally frugal methods. Ecological Modelling, 279(??):54–67, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000982.

Arnold:2015:ASA

[AAFH15] S. Arnold, S. Attinger, K. Frank, and A. Hildebrandt. Assessing the structural adequacy of alternative ecohydrological models using a pattern-oriented approach. Ecological Modelling, 316(??):52–61, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003476.

Aitkenhead:2021:DTP

[AAJB21] M. J. Aitkenhead, F. Albanito, M. B. Jones, and H. I. J. Black. Development and testing of a process-based model (MOSES) for simulating soil processes, functions and ecosystem services. Ecological Modelling, 222(20–22):3795–3810, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004741.

A:2018:MAI

[AALP18] Shiganova Tamara A., Elena Alekseenko, Moskalenko Lidia, and Nival Paul. Modelling assessment of interactions in the Black Sea of the invasive ctenophores Mnemiopsis leidyi and Beroe ovata. Ecological Modelling, 376(??):1–14, May 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300619.

Adams:2011:ELI

[AAME11] Thomas Adams, Graeme Ackland, Glenn Marion, and Colin Edwards. Effects of local interaction and dispersal on the dynamics of size-structured populations. Ecological Modelling, 222(8):1414–1422, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000639.

Agostinho:2010:ALW

[AAO10] Feni Agostinho, Luís Alberto Ambrósio, and Enrique Ortega. Assessment of a large watershed in Brazil using Emergy Evaluation and Geographical Information System. Ecological Modelling, 221(8):1209–1220, April 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008655.

An:2015:AEE

[AAW^{+}15] Qier An, Haizhong An, Lang Wang, Xiangyun Gao, and Na Lv. Analysis of embodied exergy flow between Chinese industries based on network theory. Ecological Modelling, 318(??):26–35, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000393.

Ares:2010:MHF

[AB10] Jorge O. Ares and Mónica Bertiller. Modeling highfrequency position data of large herbivores with a phasestate model. Ecological Modelling, 221(19):2323–2329, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380010002814.

Armstrong:2013:AHB

[AB13] Doug P. Armstrong and Ronald J. Brooks. Application of hierarchical biphasic growth models to long-term data for snapping turtles. Ecological Modelling, 250 (??):119–125, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005248.

Arshad:2021:MBB

[ABAA21] Fatimah Mohamed Arshad, B. K. Bala, E. F. Alias, and Ibragimov Abdulla. Modelling boom and bust of cocoa production systems in Malaysia. Ecological Modelling, 309–310(??):22–32, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001500157X.

Alva-Basurto:2014:MEC

[ABAG14] Jorge Christian Alva-Basurto and Jesús Ernesto Arias-González. Modelling the effects of climate change on a Caribbean coral reef food web. Ecological Modelling, 289(??):1–14, October 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002968.

Agostinho:2015:ICE

[ABAG15] Feni Agostinho, Ana B. B. Bertaglia, Cecília M. V. B. Almeida, and Biagio F. Giannetti. Influence of cellulase enzyme production on the energetic–environmental performance of lignocellulosic ethanol. Ecological Modelling, 315(??):46–56, November 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004207.

Agostinho:2016:CIC

[ABAG16] Feni Agostinho, Ana B. B. Bertaglia, Cecília M. V. B. Almeida, and Biagio F. Giannetti. Corrigendum to “Influence of cellulose enzyme production on the energetic environmental performance of lignocellulosic ethanol” [Ecol. Modell. 315 (2015) 46–56]. Ecological Modelling, 340 (??):149–152, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001630360X.

Adou:2010:SWP

[ABB^{+}10] J. K. Adou, Y. Billaud, D. A. Brou, J.-P. Clerc, J.- L. Consalvi, A. Fuentes, A. Kaiss, F. Nmira, B. Porterie, L. Zekri, and N. Zekri. Simulating wildfire patterns using a small-world network model. Ecological Modelling, 221(11):1463–1471, June 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001109.

Abadi:2014:IAP

[ABB^{+}14] Fitsum Abadi, Christophe Barbraud, Dominique Besson, Joël Bried, Pierre-André Crochet, Karine Delord, Jaume Forcada, Vladimir Grosbois, Richard A. Phillips, Paul Sagar, Paul Thompson, Susan Waugh, Henri Weimerskirch, Andrew G. Wood, and Olivier Gimenez. Importance of accounting for phylogenetic dependence in multi-species mark–recapture studies. Ecological Modelling, 273(??):236–241, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005656.

Aerts:2014:QSC

[ABC^{+}14] Diederik Aerts, Jan Broekaert, Marek Czachor, Maciej Kuna, Barry Sinervo, and Sandro Sozzo. Quantum structure in competing lizard communities. Ecological Modelling, 281(??):38–51, June 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000945.

Arosa:2017:LTS

[ABC^{+}17] M. L. Arosa, R. Bastos, J. A. Cabral, H. Freitas, S. R. Costa, and M. Santos. Long-term sustainability of cork oak agro-forests in the Iberian Peninsula: a modelbased approach aimed at supporting the best management options for the montado conservation. Ecological Modelling, 343(??):68–79, January 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305531.

Azevedo:2014:IFI

[ABD14] Isabel C. Azevedo, Adriano A. Bordalo, and Pedro Duarte. Influence of freshwater inflow variability on the douro estuary primary productivity: a modelling study. Ecological Modelling, 272(??):1–15, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004390.

Ager:2018:AFS

[ABD^{+}18] Alan A. Ager, Ana M. G. Barros, Michelle A. Day, Haiganoush K. Preisler, Thomas A. Spies, and John Bolte. Analyzing fine-scale spatiotemporal drivers of wildfire in a forest landscape model. Ecological Modelling, 384(??):87–102, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302230.

Abel:2010:HTG

[Abe10] Thomas Abel. Human transformities in a global hierarchy: Energy and scale in the production of people and culture. Ecological Modelling, 221(17):2112–2117, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001000267X.

Abel:2013:EED

[Abe13] Thomas Abel. Emergy evaluation of DNA and culture in ‘information cycles’. Ecological Modelling, 251 (??):85–98, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005650.

Abel:2015:CDP

[Abe15] Thomas Abel. Convergence and divergence in the production of energy transformation hierarchies. Ecological Modelling, 315(??):4–11, November 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006012.

Auban:2015:MIN

[ABGB15] Joan Bernabeu Aubán, C. Michael Barton, Salvador Pardo Gordó, and Sean M. Bergin. Modeling initial Neolithic dispersal. The first agricultural groups in West Mediterranean. Ecological Modelling, 307(??):22–31, July 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500126X.

Amichev:2016:CSW

[ABK^{+}16] Beyhan Y. Amichev, Murray J. Bentham, Werner A. Kurz, Colin P. Laroque, Suren Kulshreshtha, Joseph M. Piwowar, and Ken C. J. Van Rees. Carbon sequestration by white spruce shelterbelts in Saskatchewan, Canada: 3PG and CBM-CFS3 model simulations. Ecological Modelling, 325(??):35–46, April 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016000107.

Arima:2013:HBM

[ABL^{+}13] Serena Arima, Alberto Basset, Giovanna Jona Lasinio, Alessio Pollice, and Ilaria Rosati. A hierarchical Bayesian model for the ecological status classification of lagoons. Ecological Modelling, 263(??):187–195, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002561.

Adeva:2012:SMA

[ABR12a] J. J. García Adeva, J. H. Botha, and M. Reynolds. A simulation modelling approach to forecast establishment and spread of Bactrocera fruit flies. Ecological Modelling, 227(??):93–108, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005680.

Adjou:2012:EMI

[ABR12b] Mohamed Adjou, Jørgen Bendtsen, and Katherine Richardson. Erratum to “Modeling the influence from ocean transport, mixing and grazing on phytoplankton diversity” [Ecol. Modell. 225 (2011) 19–27]. Ecological Modelling, 235–236(??):119, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002128.

Adjou:2012:MIO

[ABR12c] Mohamed Adjou, Jørgen Bendtsen, and Katherine Richardson. Modeling the influence from ocean transport, mixing and grazing on phytoplankton diversity. Ecological Modelling, 225(??):19–27, January 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005254.

Akopov:2017:ABM

[ABS17] Andranik S. Akopov, Levon A. Beklaryan, and Armen K. Saghatelyan. Agent-based modelling for ecological economics: a case study of the Republic of Armenia. Ecological Modelling, 346(??):99–118, February 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303295.

Ackleh:2010:MMS

[ACC^{+}10] Azmy S. Ackleh, Jacoby Carter, Lauren Cole, Tom Nguyen, Jay Monte, and Claire Pettit. Measuring and modeling the seasonal changes of an urban green treefrog (Hyla cinerea) population. Ecological Modelling, 221(2):281–289, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006656.

Augustine:2011:MMG

[ACC11] Ben C. Augustine, Philip H. Crowley, and John J. Cox. A mechanistic model of GPS collar location data: Implications for analysis and bias mitigation. Ecological Modelling, 222(19):3616–3625, October 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004583.

Ackleh:2016:MIF

[ACCM16] Azmy S. Ackleh, Jacoby Carter, Vinodh K. Chellamuthu, and Baoling Ma. A model for the interaction of frog population dynamics with Batrachochytrium dendrobatidis, Janthinobacterium lividum and temperature and its implication for chytridiomycosis management. Ecological Modelling, 320(??):158–169, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004378.

Ameztegui:2017:MSD

[ACDC17] Aitor Ameztegui, Antoine Cabon, Miquel De Cáceres, and Lluís Coll. Managing stand density to enhance the adaptability of Scots pine stands to climate change: a modelling approach. Ecological Modelling, 356(??):141–150, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016307931.

Ai:2013:MNP

[ACE^{+}13] Dexiecuo Ai, Chengjin Chu, M. D. Farnon Ellwood, Rui Hou, and Gang Wang. Migration and niche partitioning simultaneously increase species richness and rarity. Ecological Modelling, 258(??):33–39, June 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300149X.

Ainley:2017:HOL

[ACE^{+}17] David G. Ainley, Elizabeth L. Crockett, Joseph T. Eastman, William R. Fraser, Nadav Nur, Kristin O’Brien, Leo A. Salas, and Donald B. Siniff. How overfishing a large piscine mesopredator explains growth in Ross Sea penguin populations: a framework to better understand impacts of a controversial fishery. Ecological Modelling, 349(??):69–75, April 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016306263.

Arias:2014:IHC

[ACK^{+}14] Mauricio E. Arias, Thomas A. Cochrane, Matti Kummu, Hannu Lauri, Gordon W. Holtgrieve, Jorma Koponen, and Thanapon Piman. Impacts of hydropower and climate change on drivers of ecological productivity of Southeast Asia’s most important wetland. Ecological Modelling, 272(??):252–263, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004821.

Aerts:2013:SES

[ACKS13] Diederik Aerts, Marek Czachor, Maciej Kuna, and Sandro Sozzo. Systems, environments, and soliton rate equations: a non-Kolmogorovian framework for population dynamics. Ecological Modelling, 267(??):80–92, October 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003554.

Ameztegui:2015:MEC

[ACM15] Aitor Ameztegui, Lluís Coll, and Christian Messier. Modelling the effect of climate-induced changes in recruitment and juvenile growth on mixed-forest dynamics: the case of montane–subalpine Pyrenean ecotones. Ecological Modelling, 313(??):84–93, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002781.

Arce:2014:TSS

[ACN14] G. L. A. F. Arce, J. A. Carvalho, and L. F. C. Nascimento. A time series sequestration and storage model of atmospheric carbon dioxide. Ecological Modelling, 272(??):59–67, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013004353.

Ascensao:2013:WVC

[ACSR^{+}13] Fernando Ascensão, Anthony Clevenger, Margarida Santos-Reis, Paulo Urbano, and Nathan Jackson. Wildlife– vehicle collision mitigation: Is partial fencing the answer? An agent-based model approach. Ecological Modelling, 257(??):36–43, May 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013001294.

Alam:2012:PBD

[AD12] Md Jahangir Alam and Dushmanta Dutta. A processbased and distributed model for nutrient dynamics in river basin: Development, testing and applications. Ecological Modelling, 247(??):112–124, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003808.

Accatino:2013:HSF

[AD13] Francesco Accatino and Carlo De Michele. Humid savanna– forest dynamics: a matrix model with vegetation–fire interactions and seasonality. Ecological Modelling, 265 (??):170–179, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002883.

Anderson:2016:GPP

[AD16] Taylor M. Anderson and Suzana Dragićević. Geospatial pest-parasitoid agent based model for optimizing biological control of forest insect infestation. Ecological Modelling, 337(??):310–329, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302617.

Anderson:2018:NAB

[AD18] Taylor M. Anderson and Suzana Dragićević. Networkagent based model for simulating the dynamic spatial network structure of complex ecological systems. Ecological Modelling, 389(??):19–32, December 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303314.

Avgar:2013:EPI

[ADF13] Tal Avgar, Rob Deardon, and John M. Fryxell. An empirically parameterized individual based model of animal movement, perception, and memory. Ecological Modelling, 251(??):158–172, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005686.

Aquilue:2017:SAP

[ADF^{+}17] Núria Aquilué, Miquel De Cáceres, Marie-Josée Fortin, Andrew Fall, and Lluís Brotons. A spatial allocation procedure to model land-use/land-cover changes: Accounting for occurrence and spread processes. Ecological Modelling, 344(??):73–86, January 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306676.

Alexandridis:2017:QMF

[ADJ^{+}17] Nikolaos Alexandridis, Jeffrey M. Dambacher, Fred Jean, Nicolas Desroy, and Cédric Bacher. Qualitative modelling of functional relationships in marine benthic communities. Ecological Modelling, 360(??):300–312, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304513.

Antony:2010:ERS

[ADL^{+}10] P. J. Antony, S. Dhanya, P. S. Lyla, B. M. Kurup, and S. Ajmal Khan. Ecological role of stomatopods (mantis shrimps) and potential impacts of trawling in a marine ecosystem of the southeast coast of India. Ecological Modelling, 221(21):2604–2614, October 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003728.

Angelini:2013:AFW

[AdMC^{+}13] Ronaldo Angelini, Ronny José de Morais, Agostinho Carlos Catella, Emiko Kawakami Resende, and Simone Libralato. Aquatic food webs of the oxbow lakes in the pantanal: a new site for fisheries guaranteed by alternated control? Ecological Modelling, 253(??):82–96, March 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300015X.

Asseng:2010:RHS

[ADPS10] S. Asseng, A. Dray, P. Perez, and X. Su. Rainfall– human–spatial interactions in a salinity-prone agricultural region of the Western Australian wheat-belt. Ecological Modelling, 221(5):812–824, March 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008308.

Ashcroft:2011:EEF

[AFC11] Michael B. Ashcroft, Kristine O. French, and Laurie A. Chisholm. An evaluation of environmental factors affecting species distributions. Ecological Modelling, 222 (3):524–531, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005314.

Ashcroft:2012:SPH

[AFC12] Michael B. Ashcroft, Kristine O. French, and Laurie A. Chisholm. A simple post-hoc method to add spatial context to predictive species distribution models. Ecological Modelling, 228(??):17–26, March 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006089.

Anderson:2011:SST

[AG11] Robert P. Anderson and Israel Gonzalez. Species-specific tuning increases robustness to sampling bias in models of species distributions: an implementation with Maxent. Ecological Modelling, 222(15):2796–2811, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002195.

Allen:2014:HCG

[AG14] Timothy Allen and Mario Giampietro. Holons, creaons, genons, environs, in hierarchy theory: Where we have gone. Ecological Modelling, 293(??):31–41, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002993.

Alexandrov:2015:BAV

[AG15] G. A. Alexandrov and G. S. Golitsyn. Biological age from the viewpoint of the thermodynamic theory of ecological systems. Ecological Modelling, 313(??):103–108, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002719.

Anderson:2013:MCC

[AGB^{+}13] James J. Anderson, Eliezer Gurarie, Chloe Bracis, Brian J. Burke, and Kristin L. Laidre. Modeling climate change impacts on phenology and population dynamics of migratory marine species. Ecological Modelling, 264(??):83–97, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001300166X.

Abadi:2012:ESD

[AGJ^{+}12] Fitsum Abadi, Olivier Gimenez, Hans Jakober, Wolfgang Stauber, Raphaël Arlettaz, and Michael Schaub. Estimating the strength of density dependence in the presence of observation errors using integrated population models. Ecological Modelling, 242(??):1–9, September 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200227X.

Acevedo:2014:GTS

[AGQP^{+}14] Pelayo Acevedo, Pablo González-Quirós, José M. Prieto, Thomas R. Etherington, Christian Gortázar, and Ana Balseiro. Generalizing and transferring spatial models: a case study to predict Eurasian badger abundance in Atlantic Spain. Ecological Modelling, 275(??):1–8, March 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005905.

aGonzalez-Sanchis:2015:MAF

[aGSCMF15] Marí a González-Sanchis, Antonio D. Del Campo, Antonio J. Molina, and Tarcísio J. G. Fernandes. Modeling adaptive forest management of a semi-arid Mediterranean Aleppo pine plantation. Ecological Modelling, 308(??):34–44, July 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015001325.

Aguilera:2011:ISC

[Agu11] Anna G. Aguilera. The influence of soil community density on plant-soil feedbacks: an important unknown in plant invasion. Ecological Modelling, 222(18):3413–3420, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003668.

Ayanlade:2017:UCT

[AH17] Ayansina Ayanlade and Michael T. Howard. Understanding changes in a Tropical Delta: a multi-method narrative of landuse/landcover change in the Niger Delta. Ecological Modelling, 364(??):53–65, November 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304076.

Azimi:2013:MIR

[AHF^{+}13] M. Azimi, Gh. A. Heshmati, M. Farahpour, M. Faramarzi, and K. C. Abbaspour. Modeling the impact of rangeland management on forage production of sagebrush species in arid and semi-arid regions of Iran. Ecological Modelling, 250(??):1–14, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005194.

Alizad:2016:CTD

[AHM^{+}16a] Karim Alizad, Scott C. Hagen, James T. Morris, Peter Bacopoulos, Matthew V. Bilskie, John F. Weishampel, and Stephen C. Medeiros. A coupled, two-dimensional hydrodynamic-marsh model with biological feedback. Ecological Modelling, 327(??):29–43, May 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300011.

Angelini:2016:EMI

[AHM^{+}16b] Silvia Angelini, Richard Hillary, Elisabetta B. Morello, Éva E. Plagányi, Michela Martinelli, Chiara Manfredi, Igor Isajlović, and Alberto Santojanni. An ecosystem model of intermediate complexity to test management options for fisheries: a case study. Ecological Modelling, 319(??):218–232, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003440.

Anderson:2013:MIF

[AHNK13] Kurt E. Anderson, Lee R. Harrison, Roger M. Nisbet, and Allison Kolpas. Modeling the influence of flow on invertebrate drift across spatial scales using a 2D hydraulic model and a 1D population model. Ecological Modelling, 265(??):207–220, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002950.

Acheampong:2014:MDF

[AHS14] Emmanuel Acheampong, Inga Hense, and Michael A. St. John. A model for the description of feeding regulation by mesozooplankton under different conditions of temperature and prey nutritional status. Ecological Modelling, 272(??):84–97, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004389.

Armstrong:2016:MCC

[AHS^{+}16] Douglas E. Armstrong, Matthew J. Hennen, Christopher P. Saunders, Thomas Brandenburger, Michael L. Brown, and Cedric Neumann. Modeling common carp underice movement using hierarchical Markov simulation. Ecological Modelling, 334(??):44–50, August 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301405.

Aldrin:2017:SSB

[AHS^{+}17] M. Aldrin, R. B. Huseby, A. Stien, R. N. Grøntvedt, H. Viljugrein, and P. A. Jansen. A stage-structured Bayesian hierarchical model for salmon lice populations at individual salmon farms — estimated from multiple farm data sets. Ecological Modelling, 359(??):333–348, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017300352.

Azanu:2016:SMM

[AJDS16] David Azanu, Sven Erik JØrgensen, Godfred Darko, and Bjarne Styrishave. Simple metal model for predicting uptake and chemical processes in sewage-fed aquaculture ecosystem. Ecological Modelling, 319(??):130–136, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003294.

Allen:2017:WTG

[AK17] Scott T. Allen and Richard F. Keim. Wetland-tree growth responses to hydrologic variability derived from development and optimization of a non-linear radial growth model. Ecological Modelling, 354(??):49–61, June 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306706.

Asfaw:2019:ETR

[AKLB19] Manalebish Debalike Asfaw, Semu Mitiku Kassa, Edward M. Lungu, and Woldeamlak Bewket. Effects of temperature and rainfall in plant–herbivore interactions at different altitude. Ecological Modelling, 406(??):50–59, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301942.

Aertsen:2010:CRD

[AKvO^{+}10] Wim Aertsen, Vincent Kint, Jos van Orshoven, Kürşad Özkan, and Bart Muys. Comparison and ranking of different modelling techniques for prediction of site index in Mediterranean mountain forests. Ecological Modelling, 221(8):1119–1130, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010000463.

Augueres:2016:BRN

[AL16] Anne-Sophie Auguères and Michel Loreau. Biotic regulation of non-limiting nutrient pools and coupling of biogeochemical cycles. Ecological Modelling, 334(??):1–7, August 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016301429.

An:2012:UHD

[ALC12] Li An and David López-Carr. Understanding human decisions in coupled natural and human systems. Ecological Modelling, 229(??):1–4, March 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005151.

Akomeah:2019:CAE

[ALC19] Eric Akomeah, Karl-Erich Lindenschmidt, and Steven C. Chapra. Comparison of aquatic ecosystem functioning between eutrophic and hypereutrophic cold-region river-lake systems. Ecological Modelling, 393(??):25–36, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304150.

Azevedo:2017:ISM

[ALeSD17] Ana Azevedo, Ana Isabel Lillebø, João Lencart e Silva, and João Miguel Dias. Intertidal seagrass models: Insights towards the development and implementation of a desiccation module. Ecological Modelling, 354(??):20–25, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016308079.

Aagaard:2016:BAC

[ALG16] Kevin Aagaard, Julie L. Lockwood, and Edwin J. Green. A Bayesian approach for characterizing uncertainty in declaring a population collapse. Ecological Modelling, 328(??):78–84, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300436.

Aho:2010:NIM

[ALKB10] Ken A. Aho, Ryan A. Long, John G. Kie, and R. Terry Bowyer. A new index for measuring perpendicularity of animal movements in relation to patch boundaries. Ecological Modelling, 221(17):2003–2007, August 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002735.

Amponsah:2011:RFE

[ALL11] N. Y. Amponsah, O. Le Corre, and B. Lacarriere. Recycling flows in emergy evaluation: a mathematical paradox? Ecological Modelling, 222(17):3071–3081, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011003310.

Acevedo:2012:DLG

[AMF12] Miguel A. Acevedo, Mariano Marcano, and Robert J. Fletcher. A diffusive logistic growth model to describe forest recovery. Ecological Modelling, 244(??):13–19, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003390.

Abernethy:2019:RMS

[AMG19] Gavin M. Abernethy, Mark McCartney, and David H. Glass. The role of migration in a spatial extension of the webworld eco-evolutionary model. Ecological Modelling, 397(??):122–140, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018304149.

Alexander:2016:SEM

[AMH16] Karen A. Alexander, Sophie A. Meyjes, and Johanna J. Heymans. Spatial ecosystem modelling of marine renewable energy installations: Gauging the utility of ecospace. Ecological Modelling, 331(??):115–128, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300047.

Anastacio:2013:MET

[AML13] Pedro M. Anastácio, Bruna Marques, and Ana I. Lillebø. Modeling the effect of temperature, solar radiation and salinity on Bolboschoenus maritimus sequestration of mercury. Ecological Modelling, 256(??):31–42, May 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000975.

Almpanidou:2014:PIH

[AMM^{+}14] Vasiliki Almpanidou, Antonios D. Mazaris, Yorgos Mertzanis, Ioannis Avraam, Ioannis Antoniou, John D. Pantis, and Stefanos P. Sgardelis. Providing insights on habitat connectivity for male brown bears: a combination of habitat suitability and landscape graph-based models. Ecological Modelling, 286(??):37–44, August 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002142.

Anyomi:2016:WMO

[AMR16] K. A. Anyomi, S. J. Mitchell, and J. C. Ruel. Windthrow modelling in old-growth and multi-layered boreal forests. Ecological Modelling, 327(??):105–114, May 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300199.

Adhikari:2019:VMP

[AMRH19] Arjun Adhikari, Kumar P. Mainali, Imtiaz Rangwala, and Andrew J. Hansen. Various measures of potential evapotranspiration have species-specific impact on species distribution models. Ecological Modelling, 414(??): Article 108836, December 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019303448.

An:2012:MHD

[An12] Li An. Modeling human decisions in coupled human and natural systems: Review of agent-based models. Ecological Modelling, 229(??):25–36, March 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003802.

Anandhi:2017:CCM

[Ana17] Aavudai Anandhi. CISTA-A: Conceptual model using indicators selected by systems thinking for adaptation strategies in a changing climate: Case study in agro-ecosystems. Ecological Modelling, 345(??):41–55, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630730X.

Acheampong:2012:TAM

[ANMJ12] E. Acheampong, M. H. Nielsen, A. Mitra, and M. A. St. John. Towards an adaptive model for simulating growth of marine mesozooplankton: a macromolecular perspective. Ecological Modelling, 225(??):1–18, January 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005205.

Anonymous:2010:EBa

[Ano10a] Anonymous. Editorial Board. Ecological Modelling, 221(1):i, January 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000900773X.

Anonymous:2010:EBb

[Ano10b] Anonymous. Editorial Board. Ecological Modelling, 221 (13–14):i, July 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380010002206.

Anonymous:2010:EBc

[Ano10c] Anonymous. Editorial Board. Ecological Modelling, 221(16):CO2, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010002899.

Anonymous:2010:EBd

[Ano10d] Anonymous. Editorial Board. Ecological Modelling, 221 (21):i, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380010004291.

Anonymous:2010:EBe

[Ano10e] Anonymous. Editorial Board. Ecological Modelling, 221(23):i, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005041.

Anonymous:2010:EBf

[Ano10f] Anonymous. Editorial Board. Ecological Modelling, 221(24):i, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005351.

Anonymous:2010:PAa

[Ano10g] Anonymous. Pages 1089–1244 (24 April 2010). Ecological Modelling, 221(8):??, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PMc

[Ano10h] Anonymous. Pages 1245–1336 (10 May 2010). Ecological Modelling, 221(9):??, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PJa

[Ano10i] Anonymous. Pages 131–378 (24 January 2010). Ecological Modelling, 221(2):??, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PMd

[Ano10j] Anonymous. Pages 1337–1450 (24 May 2010). Ecological Modelling, 221(10):??, May 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PJb

[Ano10k] Anonymous. Pages 1451–1544 (10 June 2010). Ecological Modelling, 221(11):??, June 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PJc

[Ano10l] Anonymous. Pages 1545–1648 (24 June 2010). Ecological Modelling, 221(12):??, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PJd

[Ano10m] Anonymous. Pages 1649–1756 (10 July 2010). Ecological Modelling, 221(13–14):??, July 10, 2010. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PJe

[Ano10n] Anonymous. Pages 1757–1838 (24 July 2010). Ecological Modelling, 221(15):??, July 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PAb

[Ano10o] Anonymous. Pages 1839–2002 (10 August 2010). Ecological Modelling, 221(16):??, August 10, 2010. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PAc

[Ano10p] Anonymous. Pages 2003–2132 (24 August 2010). Ecological Modelling, 221(17):??, August 24, 2010. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PS

[Ano10q] Anonymous. Pages 2227–2414 (24 September 2010). Ecological Modelling, 221(19):??, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:POa

[Ano10r] Anonymous. Pages 2415–2508 (10 October 2010). Ecological Modelling, 221(20):??, October 10, 2010. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:POb

[Ano10s] Anonymous. Pages 2509–2624 (24 October 2010). Ecological Modelling, 221(21):??, October 24, 2010. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PNa

[Ano10t] Anonymous. Pages 2625–2738 (10 November 2010). Ecological Modelling, 221(22):??, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PNc

[Ano10u] Anonymous. Pages 2739–2858 (24 November 2010). Ecological Modelling, 221(23):??, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PD

[Ano10v] Anonymous. Pages 2859–2958 (15 December 2010). Ecological Modelling, 221(24):??, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PFa

[Ano10w] Anonymous. Pages 379–564 (10 February 2010). Ecological Modelling, 221(3):??, February 10, 2010. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PFb

[Ano10x] Anonymous. Pages 565–728 (24 February 2010). Ecological Modelling, 221(4):??, February 24, 2010. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PMa

[Ano10y] Anonymous. Pages 729–884 (10 March 2010). Ecological Modelling, 221(5):??, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PMb

[Ano10z] Anonymous. Pages 885–964 (24 March 2010). Ecological Modelling, 221(6):??, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2010:PNb

[Ano10-27] Anonymous. Publisher’s note. Ecological Modelling, 221(22):2625, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004163.

Anonymous:2010:Ta

[Ano10-28] Anonymous. Title. Ecological Modelling, 221(17):i, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010003157.

Anonymous:2010:Tb

[Ano10-29] Anonymous. Title. Ecological Modelling, 221(19):i, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001000356X.

Anonymous:2010:Tc

[Ano10-30] Anonymous. Title. Ecological Modelling, 221(20):i, October 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010003893.

Anonymous:2011:A

[Ano11a]  Anonymous.  Announcement.  Ecological Modelling,

222(1):220, January 10, 2011. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001000565X.

Anonymous:2011:EBa

[Ano11b] Anonymous. Editorial Board. Ecological Modelling, 222(1): iii, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380010005594.

Anonymous:2011:EBb

[Ano11c] Anonymous. Editorial Board. Ecological Modelling, 222 (4):i, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011000081.

Anonymous:2011:EBc

[Ano11d] Anonymous. Editorial Board. Ecological Modelling, 222 (6):i, March 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011000676.

Anonymous:2011:EBd

[Ano11e] Anonymous. Editorial Board. Ecological Modelling, 222(7):i, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011000834.

Anonymous:2011:EBe

[Ano11f] Anonymous. Editorial Board. Ecological Modelling, 222(9):i, May 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011001529.

Anonymous:2011:EBf

[Ano11g] Anonymous. Editorial Board. Ecological Modelling, 222 (10):i, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011001906.

Anonymous:2011:EBg

[Ano11h] Anonymous. Editorial Board. Ecological Modelling, 222 (11):i, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011002043.

Anonymous:2011:EBh

[Ano11i] Anonymous. Editorial Board. Ecological Modelling, 222 (12):iii, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011002341.

Anonymous:2011:EBi

[Ano11j] Anonymous. Editorial Board. Ecological Modelling, 222 (13):iii, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011002754.

Anonymous:2011:EBj

[Ano11k] Anonymous. Editorial Board. Ecological Modelling, 222 (14):iii, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011003176.

Anonymous:2011:EBk

[Ano11l] Anonymous. Editorial Board. Ecological Modelling, 222(17):iii, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004030.

Anonymous:2011:EBl

[Ano11m] Anonymous. Editorial Board. Ecological Modelling, 222(18):iii, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001100442X.

Anonymous:2011:EBm

[Ano11n] Anonymous. Editorial Board. Ecological Modelling, 222(19):iii, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004790.

Anonymous:2011:EBn

[Ano11o] Anonymous. Editorial Board. Ecological Modelling, 222 (23–24):iii, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005345.

Anonymous:2011:E

[Ano11p] Anonymous. ESEM 2011. Ecological Modelling, 222(2):II, January 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010006058.

Anonymous:2011:I

[Ano11q] Anonymous. ISEM 2011. Ecological Modelling, 222(2):I, January 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010006046.

Anonymous:2011:IA

[Ano11r] Anonymous. ISEM announcement. Ecological Modelling, 222(4):iv, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011000056.

Anonymous:2011:PJa

[Ano11s] Anonymous. Pages 1–220 (10 January 2011). Ecological Modelling, 222(1):??, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PMa

[Ano11t] Anonymous. Pages 1069–1208 (10 March 2011). Ecological Modelling, 222(5):??, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PMb

[Ano11u] Anonymous. Pages 1209–1262 (24 March 2011). Ecological Modelling, 222(6):??, March 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PAa

[Ano11v] Anonymous. Pages 1263–1342 (10 April 2011). Ecological Modelling, 222(7):??, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PAb

[Ano11w] Anonymous. Pages 1343–1556 (24 April 2011). Ecological Modelling, 222(8):??, April 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PMc

[Ano11x] Anonymous. Pages 1557–1656 (10 May 2011). Ecological Modelling, 222(9):??, May 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PMd

[Ano11y] Anonymous. Pages 1657–1792 (24 May 2011). Ecological Modelling, 222(10):??, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PJb

[Ano11z] Anonymous. Pages 1793–1912 (10 June 2011). Ecological Modelling, 222(11):??, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PJc

[Ano11-27] Anonymous. Pages 1913–2036 (24 June 2011). Ecological Modelling, 222(12):??, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PJd

[Ano11-28] Anonymous. Pages 2037–2168 (10 July 2011). Ecological Modelling, 222(13):??, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PJe

[Ano11-29] Anonymous. Pages 2169–2594 (24 July 2011). Ecological Modelling, 222(14):??, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PAc

[Ano11-30] Anonymous. Pages 2595–2866 (10 August 2011). Ecological Modelling, 222(15):??, August 10, 2011. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PSa

[Ano11-31] Anonymous. Pages 3011–3300 (10 September 2011). Ecological Modelling, 222(17):??, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PSb

[Ano11-32] Anonymous. Pages 3305–3456 (24 September 2011). Ecological Modelling, 222(18):??, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PO

[Ano11-33] Anonymous. Pages 3457–3650 (10 October 2011). Ecological Modelling, 222(19):??, October 10, 2011. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PFa

[Ano11-34] Anonymous. Pages 381–902 (10 February 2011). Ecological Modelling, 222(3):??, February 10, 2011. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PD

[Ano11-35] Anonymous. Pages 3811–3956 (10 December 2011). Ecological Modelling, 222(23–24):??, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:PFb

[Ano11-36] Anonymous. Pages 903–1068 (24 February 2011). Ecological Modelling, 222(4):??, February 24, 2011. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2011:SIA

[Ano11-37] Anonymous. Special issue announcement. Ecological Modelling, 222(4):iii, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011000044.

Anonymous:2011:STI

[Ano11-38] Anonymous. Special (thematic) issues. Ecological Modelling, 222(2):III, January 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001000606X.

Anonymous:2011:Ta

[Ano11-39] Anonymous. Title. Ecological Modelling, 222(8):i, April 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001207.

Anonymous:2011:Tb

[Ano11-40] Anonymous. Title. Ecological Modelling, 222(15):iii, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011003486.

Anonymous:2012:BTM

[Ano12a] Anonymous. Building and testing models of long-term agricultural intensification and population dynamics: a case study from the Leeward Kohala Field System, Hawai’i [Ecol. Model. 227 (2012) 18–20]. Ecological Modelling, 241(??):1, August 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/

S0304380012003055. Publisher note about mistaken printing of this article in an incorrect issue.

Anonymous:2012:EBa

[Ano12b] Anonymous. Editorial Board. Ecological Modelling, 224(1): iii, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011005746.

Anonymous:2012:EBb

[Ano12c] Anonymous. Editorial Board. Ecological Modelling, 225(??):iii, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005953.

Anonymous:2012:EBc

[Ano12d] Anonymous. Editorial Board. Ecological Modelling, 226(??):iii, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012000099.

Anonymous:2012:EBd

[Ano12e] Anonymous. Editorial Board. Ecological Modelling, 227(??):iii, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012000300.

Anonymous:2012:EBe

[Ano12f] Anonymous. Editorial Board. Ecological Modelling, 228 (??):iii, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380012000439.

Anonymous:2012:EBf

[Ano12g] Anonymous. Editorial Board. Ecological Modelling, 230 (??):iii, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380012000853.

Anonymous:2012:EBg

[Ano12h] Anonymous. Editorial Board. Ecological Modelling, 231 (??):iii, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380012001305.

Anonymous:2012:EBh

[Ano12i] Anonymous. Editorial Board. Ecological Modelling, 232 (??):iii, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380012001524.

Anonymous:2012:EBi

[Ano12j] Anonymous. Editorial Board. Ecological Modelling, 233 (??):iii, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001200186X.

Anonymous:2012:EBj

[Ano12k] Anonymous. Editorial Board. Ecological Modelling, 235–236(??):iii, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002165.

Anonymous:2012:EBk

[Ano12l] Anonymous. Editorial Board. Ecological Modelling, 237–238(??):iii, July 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002475.

Anonymous:2012:EBl

[Ano12m] Anonymous. Editorial Board. Ecological Modelling, 240 (??):iii, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380012002943.

Anonymous:2012:EBm

[Ano12n] Anonymous. Editorial Board. Ecological Modelling, 242(??):iii, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003584.

Anonymous:2012:EBn

[Ano12o] Anonymous. Editorial Board. Ecological Modelling, 243(??):iii, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003687.

Anonymous:2012:EBo

[Ano12p] Anonymous. Editorial Board. Ecological Modelling, 244(??):iii, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003900.

Anonymous:2012:EBp

[Ano12q] Anonymous. Editorial Board. Ecological Modelling, 246(??):iii, November 10, 2012. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004450.

Anonymous:2012:EBq

[Ano12r] Anonymous. Editorial Board. Ecological Modelling, 247 (??):iii, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380012004942.

Anonymous:2012:EHR

[Ano12s] Anonymous. Exploring the hydrologic relationships in a swamp-dominated watershed — a network-environ-analysis based approach. Ecological Modelling, 247(??):168–174, December 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012003791.

Anonymous:2012:MTB

[Ano12t] Anonymous. Modelling a thermodynamic-based comparative framework for urban sustainability: Incorporating economic and ecological losses into emergy analysis. Ecological Modelling, 247(??):233–240, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004231.

Anonymous:2012:PMc

[Ano12u] Anonymous. Pages 1–104 (24 May 2012). Ecological Modelling, 233(??):1–104, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PJc

[Ano12v] Anonymous. Pages 1–120 (24 June 2012). Ecological Modelling, 235–236(??):1–120, June 24, 2012. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PJa

[Ano12w] Anonymous. Pages 1–130 (10 January 2012). Ecological Modelling, 224(1):??, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PMa

[Ano12x] Anonymous. Pages 1–130 (10 March 2012). Ecological Modelling, 228(??):1–130, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PFa

[Ano12y] Anonymous. Pages 1–142 (10 February 2012). Ecological Modelling, 226(??):1–142, February 10, 2012. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PJd

[Ano12z] Anonymous. Pages 1–142 (10 July 2012). Ecological Modelling, 237–238(??):1–142, July 10, 2012. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PSb

[Ano12-27] Anonymous. Pages 1–142 (24 September 2012). Ecological Modelling, 243(??):1–142, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PAa

[Ano12-28] Anonymous. Pages 1–148 (10 April 2012). Ecological Modelling, 230(??):1–148, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PAb

[Ano12-29] Anonymous. Pages 1–152 (24 April 2012). Ecological Modelling, 231(??):1–152, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PAc

[Ano12-30] Anonymous. Pages 1–160 (10 August 2012). Ecological Modelling, 240(??):1–160, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PMb

[Ano12-31] Anonymous. Pages 1–160 (10 May 2012). Ecological Modelling, 232(??):1–160, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PO

[Ano12-32] Anonymous. Pages 1–162 (10 October 2012). Ecological Modelling, 244(??):1–162, October 10, 2012. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PFb

[Ano12-33] Anonymous. Pages 1–162 (24 February 2012). Ecological Modelling, 227(??):1–162, February 24, 2012. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PJb

[Ano12-34] Anonymous. Pages 1–176 (24 January 2012). Ecological Modelling, 225(??):1–176, January 24, 2012. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PSa

[Ano12-35] Anonymous. Pages 1–198 (10 September 2012). Ecological Modelling, 242(??):1–198, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PD

[Ano12-36] Anonymous. Pages 1–326 (December 2012). Ecological Modelling, 247(??):1–326, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PAAa

[Ano12-37] Anonymous. Pages A1–A2, 1–156 (10 November 2012). Ecological Modelling, 246(??):A1–, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2012:PN

[Ano12-38] Anonymous. Publisher’s note. Ecological Modelling, 246(??):A1, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004498.

Anonymous:2013:Aa

[Ano13a] Anonymous. Announcement. Ecological Modelling, 252 (??):I, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013000793.

Anonymous:2013:Ab

[Ano13b] Anonymous. Announcement. Ecological Modelling, 253 (??):I, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001300135X.

Anonymous:2013:EBa

[Ano13c] Anonymous. Editorial Board. Ecological Modelling, 248(??):iii, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005352.

Anonymous:2013:EBb

[Ano13d] Anonymous. Editorial Board. Ecological Modelling, 250(??):iii, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013000203.

Anonymous:2013:EBc

[Ano13e] Anonymous. Editorial Board. Ecological Modelling, 251(??):iii, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013000562.

Anonymous:2013:EBd

[Ano13f] Anonymous. Editorial Board. Ecological Modelling, 253 (??):iii, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013001166.

Anonymous:2013:EBe

[Ano13g] Anonymous. Editorial Board. Ecological Modelling, 254 (??):iii, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013001257.

Anonymous:2013:EBf

[Ano13h] Anonymous. Editorial Board. Ecological Modelling, 255 (??):iii, April 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013001415.

Anonymous:2013:EBg

[Ano13i] Anonymous. Editorial Board. Ecological Modelling, 256 (??):iii, May 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013001610.

Anonymous:2013:EBh

[Ano13j] Anonymous. Editorial Board. Ecological Modelling, 257 (??):iii, May 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001300183X.

Anonymous:2013:EBi

[Ano13k] Anonymous. Editorial Board. Ecological Modelling, 258 (??):iii, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013001956.

Anonymous:2013:EBj

[Ano13l] Anonymous. Editorial Board. Ecological Modelling, 259 (??):iii, June 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001300224X.

Anonymous:2013:EBk

[Ano13m] Anonymous. Editorial Board. Ecological Modelling, 260 (??):iii, July 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001300238X.

Anonymous:2013:EBl

[Ano13n] Anonymous. Editorial Board. Ecological Modelling, 261–262(??):iii, July 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002640.

Anonymous:2013:EBm

[Ano13o] Anonymous. Editorial Board. Ecological Modelling, 263 (??):iii, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013003062.

Anonymous:2013:EBn

[Ano13p] Anonymous. Editorial Board. Ecological Modelling, 265(??):iii, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013003657.

Anonymous:2013:EBo

[Ano13q] Anonymous. Editorial Board. Ecological Modelling, 266(??):iii, September 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013003906.

Anonymous:2013:EBp

[Ano13r] Anonymous. Editorial Board. Ecological Modelling, 267(??):iii, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004092.

Anonymous:2013:EBq

[Ano13s] Anonymous. Editorial Board. Ecological Modelling, 268(??):iii, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004262.

Anonymous:2013:EBr

[Ano13t] Anonymous. Editorial Board. Ecological Modelling, 270 (??):iii, December 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013004924.

Anonymous:2013:FAI

[Ano13u] Anonymous. First announcement: International Society for Ecological Modelling Conference (ISEM 2013). Ecological Modelling, 251(??):327, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000525.

Anonymous:2013:PMc

[Ano13v] Anonymous. Pages 1–112 (24 May 2013). Ecological Modelling, 257(??):1–112, May 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PMa

[Ano13w] Anonymous. Pages 1–116 (24 March 2013). Ecological Modelling, 253(??):1–116, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PN

[Ano13x] Anonymous. Pages 1–126 (10 November 2013). Ecological Modelling, 269(??):1–126, November 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PJb

[Ano13y] Anonymous. Pages 1–134 (10 June 2013). Ecological Modelling, 258(??):1–134, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:POb

[Ano13z] Anonymous. Pages 1–134 (24 October 2013). Ecological Modelling, 268(??):1–134, October 24, 2013. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:POa

[Ano13-27] Anonymous. Pages 1–180 (10 October 2013). Ecological Modelling, 267(??):1–180, October 10, 2013. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PSb

[Ano13-28] Anonymous. Pages 1–186 (24 September 2013). Ecological Modelling, 266(??):1–186, September 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PJa

[Ano13-29] Anonymous. Pages 1–214 (10 January 2013). Ecological Modelling, 248(??):1–214, January 10, 2013. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PSa

[Ano13-30] Anonymous. Pages 1–250 (10 September 2013). Ecological Modelling, 265(??):1–250, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PFb

[Ano13-31] Anonymous. Pages 1–328 (24 February 2013). Ecological Modelling, 251(??):1–328, February 24, 2013. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PAc

[Ano13-32] Anonymous. Pages 1–348 (10 August 2013). Ecological Modelling, 263(??):1–348, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PFa

[Ano13-33] Anonymous. Pages 1–402 (10 February 2013). Ecological Modelling, 250(??):1–402, February 10, 2013. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PJc

[Ano13-34] Anonymous. Pages 1–74 (24 June 2013). Ecological Modelling, 259(??):1–74, June 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PJd

[Ano13-35] Anonymous. Pages 1–76 (10 July 2013). Ecological Modelling, 260(??):1–76, July 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PD

[Ano13-36] Anonymous. Pages 1–76 (December 2013). Ecological Modelling, 270(??):1–76, December 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PMb

[Ano13-37] Anonymous. Pages 1–84 (10 May 2013). Ecological Modelling, 256(??):1–84, May 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PAb

[Ano13-38] Anonymous. Pages 1–86 (24 April 2013). Ecological Modelling, 255(??):1–86, April 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PAa

[Ano13-39] Anonymous. Pages 1–92 (10 April 2013). Ecological Modelling, 254(??):1–92, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:PJe

[Ano13-40] Anonymous. Pages 1–98 (24 July 2013). Ecological Modelling, 261–262(??):1–98, July 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2013:TP

[Ano13-41] Anonymous. Title page. Ecological Modelling, 269(??): iii, November 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380013004535.

Anonymous:2014:EB

[Ano14a] Anonymous. Editorial Board. Ecological Modelling, 272(??):iii, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013005516.

Anonymous:2014:PMa

[Ano14b] Anonymous. Pages 1–100 (10 March 2014). Ecological Modelling, 275(??):1–100, March 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PAb

[Ano14c] Anonymous. Pages 1–114 (24 April 2014). Ecological Modelling, 278(??):1–114, April 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PMc

[Ano14d] Anonymous. Pages 1–118 (10 May 2014). Ecological Modelling, 279(??):1–118, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PAa

[Ano14e] Anonymous. Pages 1–132 (10 April 2014). Ecological Modelling, 277(??):1–132, April 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PD

[Ano14f] Anonymous. Pages 1–136 (24 December 2014). Ecological Modelling, 294(??):1–136, December 24, 2014. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PO

[Ano14g] Anonymous. Pages 1–150 (10 October 2014). Ecological Modelling, 289(??):1–150, October 10, 2014. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PS

[Ano14h] Anonymous. Pages 1–212 (24 September 2014). Ecological Modelling, 288(??):1–212, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PNa

[Ano14i] Anonymous. Pages 1–260 (10 November 2014). Ecological Modelling, 291(??):1–260, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PFa

[Ano14j] Anonymous. Pages 1–298 (10 February 2014). Ecological Modelling, 273(??):1–298, February 10, 2014. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PJa

[Ano14k] Anonymous. Pages 1–410 (24 January 2014). Ecological Modelling, 272(??):1–410, January 24, 2014. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PNb

[Ano14l] Anonymous. Pages 1–62 (24 November 2014). Ecological Modelling, 292(??):1–62, November 24, 2014. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PJb

[Ano14m] Anonymous. Pages 1–64 (10 June 2014). Ecological Modelling, 281(??):1–64, June 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PJe

[Ano14n] Anonymous. Pages 1–74 (24 July 2014). Ecological Modelling, 284(??):1–74, July 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PAd

[Ano14o] Anonymous. Pages 1–78 (24 August 2014). Ecological Modelling, 286(??):1–78, August 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PJc

[Ano14p] Anonymous. Pages 1–82 (24 June 2014). Ecological Modelling, 282(??):1–82, June 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PJd

[Ano14q] Anonymous. Pages 1–84 (10 July 2014). Ecological Modelling, 283(??):1–84, July 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PAc

[Ano14r] Anonymous. Pages 1–90 (10 August 2014). Ecological Modelling, 285(??):1–90, August 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PFb

[Ano14s] Anonymous. Pages 1–92 (24 February 2014). Ecological Modelling, 274(??):1–92, February 24, 2014. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PMb

[Ano14t] Anonymous. Pages 1–94 (24 March 2014). Ecological Modelling, 276(??):1–94, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PAAa

[Ano14u] Anonymous. Pages A1-A4, 1-60 (10 September 2014). Ecological Modelling, 287(??):A1–, September 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2014:PIOa

[Ano14v] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 273(??):iii, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013006078.

Anonymous:2014:PIOb

[Ano14w] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 274(??):iii, February 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000295.

Anonymous:2014:PIOc

[Ano14x] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 275(??):iii, March 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000519.

Anonymous:2014:PIOd

[Ano14y] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 276(??):iii, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000775.

Anonymous:2014:PIOe

[Ano14z] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 277(??):iii, April 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001070.

Anonymous:2014:PIOf

[Ano14-27] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 278(??):iii, April 24, 2014. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001203.

Anonymous:2014:PIOg

[Ano14-28] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 279(??):iii, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001367.

Anonymous:2014:PIOh

[Ano14-29] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 281(??):iii, June 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001641.

Anonymous:2014:PIOi

[Ano14-30] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 282(??):iii, June 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001768.

Anonymous:2014:PIOj

[Ano14-31] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 283(??):iii, July 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001400204X.

Anonymous:2014:PIOk

[Ano14-32] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 284(??):iii, July 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002208.

Anonymous:2014:PIOl

[Ano14-33] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 285(??):iii, August 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002324.

Anonymous:2014:PIOm

[Ano14-34] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 286(??):iii, August 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001400249X.

Anonymous:2014:PIOn

[Ano14-35] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 287(??):iii, September 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002804.

Anonymous:2014:PIOo

[Ano14-36] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 288(??):iii, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400310X.

Anonymous:2014:PIOp

[Ano14-37] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 291(??):iii, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004037.

Anonymous:2014:PIOq

[Ano14-38] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 292(??):iii, November 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004372.

Anonymous:2014:PIOr

[Ano14-39] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 294(??):iii, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005043.

Anonymous:2014:SIP

[Ano14-40] Anonymous. Special issue of PMA’12. Ecological Modelling, 290(??):??, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PAb

[Ano15a] Anonymous. Pages 1–106 (10 April 2015). Ecological Modelling, 301(??):1–106, April 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PD

[Ano15b] Anonymous. Pages 1–106 (10 December 2015). Ecological Modelling, 317(??):1–106, December 10, 2015. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PMc

[Ano15c] Anonymous. Pages 1–110 (10 May 2015). Ecological Modelling, 303(??):1–110, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PJa

[Ano15d] Anonymous. Pages 1–136 (24 January 2015). Ecological Modelling, 296(??):1–136, January 24, 2015. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PMb

[Ano15e] Anonymous. Pages 1–136 (24 March 2015). Ecological Modelling, 300(??):1–136, March 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PMa

[Ano15f] Anonymous. Pages 1–156 (10 March 2015). Ecological Modelling, 299(??):1–156, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PN

[Ano15g] Anonymous. Pages 1–230 (24 November 2015). Ecological Modelling, 316(??):1–230, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PO

[Ano15h] Anonymous. Pages 1–376 (10 October 2015). Ecological Modelling, 313(??):1–376, October 10, 2015. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PSb

[Ano15i] Anonymous. Pages 1–416 (24 September 2015). Ecological Modelling, 312(??):1–416, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PAc

[Ano15j] Anonymous. Pages 1–54 (24 April 2015). Ecological Modelling, 302(??):1–54, April 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PJd

[Ano15k] Anonymous. Pages 1–62 (24 July 2015). Ecological Modelling, 308(??):1–62, July 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PJb

[Ano15l] Anonymous. Pages 1–64 (10 June 2015). Ecological Modelling, 305(??):1–64, June 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PJc

[Ano15m] Anonymous. Pages 1–66 (10 July 2015). Ecological Modelling, 307(??):1–66, July 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PMd

[Ano15n] Anonymous. Pages 1–84 (24 May 2015). Ecological Modelling, 304(??):1–84, May 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PSa

[Ano15o] Anonymous. Pages 1–88 (10 September 2015). Ecological Modelling, 311(??):1–88, September 10, 2015. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PAAb

[Ano15p] Anonymous. Pages A1-A2, 1-144 (24 October 2015). Ecological Modelling, 314(??):A1–, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PAAa

[Ano15q] Anonymous. Pages A1-A2, 1-246 (10 February 2015). Ecological Modelling, 297(??):A1–, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2015:PIOa

[Ano15r] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 296(??):iii, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005808.

Anonymous:2015:PIOb

[Ano15s] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 297(??):iii, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014006243.

Anonymous:2015:PIOc

[Ano15t] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 299(??):iii, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000289.

Anonymous:2015:PIOd

[Ano15u] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 300(??):iii, March 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001500054X.

Anonymous:2015:PIOe

[Ano15v] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 301(??):iii, April 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000903.

Anonymous:2015:PIOf

[Ano15w] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 302(??):iii, April 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000794.

Anonymous:2015:PIOg

[Ano15x] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 303(??):iii, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001040.

Anonymous:2015:PIOh

[Ano15y] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 304(??):iii, May 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001155.

Anonymous:2015:PIOi

[Ano15z] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 305(??):iii, June 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001416.

Anonymous:2015:PIOj

[Ano15-27] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 307(??):iii, July 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001805.

Anonymous:2015:PIOk

[Ano15-28] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 308(??):iii, July 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001500191X.

Anonymous:2015:PIOl

[Ano15-29] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 311(??):iii, September 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002446.

Anonymous:2015:PIOm

[Ano15-30] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 312(??):iii, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002926.

Anonymous:2015:PIOn

[Ano15-31] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 313(??):iii, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003397.

Anonymous:2015:PIOo

[Ano15-32] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 314(??):iii, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003610.

Anonymous:2015:PIOp

[Ano15-33] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 316(??):iii, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004251.

Anonymous:2015:PIOq

[Ano15-34] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 317(??):iii, December 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004457.

Anonymous:2016:PAb

[Ano16a] Anonymous. Pages 1–100 (10 August 2016). Ecological Modelling, 333(??):1–100, August 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PJd

[Ano16b] Anonymous. Pages 1–112 (24 July 2016). Ecological Modelling, 332(??):1–112, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PMc

[Ano16c] Anonymous. Pages 1–114 (10 May 2016). Ecological Modelling, 327(??):1–114, May 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PJb

[Ano16d] Anonymous. Pages 1–120 (10 June 2016). Ecological Modelling, 329(??):1–120, June 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PFa

[Ano16e] Anonymous. Pages 1–130 (10 February 2016). Ecological Modelling, 321(??):1–130, February 10, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PMa

[Ano16f] Anonymous. Pages 1–132 (10 March 2016). Ecological Modelling, 323(??):1–132, March 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PSa

[Ano16g] Anonymous. Pages 1–138 (10 September 2016). Ecological Modelling, 335(??):1–138, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PFb

[Ano16h] Anonymous. Pages 1–138 (24 February 2016). Ecological Modelling, 322(??):1–138, February 24, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PMd

[Ano16i] Anonymous. Pages 1–148 (24 May 2016). Ecological Modelling, 328(??):1–148, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:POb

[Ano16j] Anonymous. Pages 1–148 (24 October 2016). Ecological Modelling, 338(??):1–148, October 24, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PNb

[Ano16k] Anonymous. Pages 1–152 (24 November 2016). Ecological Modelling, 340(??):1–152, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PNa

[Ano16l] Anonymous. Pages 1–156 (10 November 2016). Ecological Modelling, 339(??):1–156, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PDb

[Ano16m] Anonymous. Pages 1–198 (24 December 2016). Ecological Modelling, 342(??):1–198, December 24, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:POa

[Ano16n] Anonymous. Pages 1–350 (10 October 2016). Ecological Modelling, 337(??):1–350, October 10, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PJa

[Ano16o] Anonymous. Pages 1–406 (24 January 2016). Ecological Modelling, 320(??):1–406, January 24, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PMb

[Ano16p] Anonymous. Pages 1–54 (24 March 2016). Ecological Modelling, 324(??):1–54, March 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PDa

[Ano16q] Anonymous. Pages 1–62 (10 December 2016). Ecological Modelling, 341(??):1–62, December 10, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PJc

[Ano16r] Anonymous. Pages 1–62 (24 June 2016). Ecological Modelling, 330(??):1–62, June 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PSb

[Ano16s] Anonymous. Pages 1–82 (24 September 2016). Ecological Modelling, 336(??):1–82, September 24, 2016. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PAa

[Ano16t] Anonymous. Pages 1–88 (10 April 2016). Ecological Modelling, 325(??):1–88, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PAc

[Ano16u] Anonymous. Pages 1–90 (24 August 2016). Ecological Modelling, 334(??):1–90, August 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2016:PIIa

[Ano16v] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 336(??):iii, September 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302277.

Anonymous:2016:PIIb

[Ano16w] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 337(??):iii, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302733.

Anonymous:2016:PIIc

[Ano16x] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 338(??):iii, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303362.

Anonymous:2016:PIId

[Ano16y] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 339(??):iii, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303921.

Anonymous:2016:PIIe

[Ano16z] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 340(??):iii, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304550.

Anonymous:2016:PIIf

[Ano16-27] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 341(??):iii, December 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305324.

Anonymous:2016:PIIg

[Ano16-28] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 342(??):iii, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305981.

Anonymous:2016:PIOa

[Ano16-29] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 320(??):iii, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005384.

Anonymous:2016:PIOb

[Ano16-30] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 321(??):iii, February 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005657.

Anonymous:2016:PIOc

[Ano16-31] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 322(??):iii, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005864.

Anonymous:2016:PIOd

[Ano16-32] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 323(??):iii, March 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016000211.

Anonymous:2016:PIOe

[Ano16-33] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 324(??):iii, March 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016000351.

Anonymous:2016:PIOf

[Ano16-34] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 325(??):iii, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300151.

Anonymous:2016:PIOg

[Ano16-35] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 327(??):iii, May 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300540.

Anonymous:2016:PIOh

[Ano16-36] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 328(??):iii, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300734.

Anonymous:2016:PIOi

[Ano16-37] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 329(??):iii, June 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300849.

Anonymous:2016:PIOj

[Ano16-38] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 330(??):iii, June 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301119.

Anonymous:2016:PIOk

[Ano16-39] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 331(??):iii, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301260.

Anonymous:2016:PIOl

[Ano16-40] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 332(??):iii, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301570.

Anonymous:2016:PIOm

[Ano16-41] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 333(??):iii, August 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301715.

Anonymous:2016:PIOn

[Ano16-42] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 334(??):iii, August 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301831.

Anonymous:2016:PIOo

[Ano16-43] Anonymous. Prelim (ii oriii) title page. Ecological Modelling, 335(??):iii, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302162.

Anonymous:2017:YSM

[Ano17a] Anonymous. 75 years since Monod: It is time to increase the complexity of our predictive ecosystem models (opinion). Ecological Modelling, 346(??):77–87, February 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303209.

Anonymous:2017:Aa

[Ano17b] Anonymous. Announcement. Ecological Modelling, 349 (??):I, April 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380017301709.

Anonymous:2017:Ab

[Ano17c] Anonymous. Announcement. Ecological Modelling, 350 (??):I, April 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380017301990.

Anonymous:2017:Ac

[Ano17d] Anonymous. Announcement. Ecological Modelling, 351 (??):I, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001730217X.

Anonymous:2017:IFCf

[Ano17e] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 359(??):ifc, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303460.

Anonymous:2017:IFCg

[Ano17f] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 360(??):ifc, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303630.

Anonymous:2017:IFCh

[Ano17g] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 361(??):ifc, October 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303824.

Anonymous:2017:IFCi

[Ano17h] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 362(??):ifc, October 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303976.

Anonymous:2017:IFCj

[Ano17i] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 363(??):ifc, November 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304118.

Anonymous:2017:IFCk

[Ano17j] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 364(??):ifc, November 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304258.

Anonymous:2017:IFCl

[Ano17k] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 365(??):ifc, December 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304593.

Anonymous:2017:IFCm

[Ano17l] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 366(??):ifc, December 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304520.

Anonymous:2017:IFCa

[Ano17m] Anonymous. Inside front cover — editorial board page/cover image legend if applicable. Ecological Modelling, 354(??):ifc, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017302557.

Anonymous:2017:IFCb

[Ano17n] Anonymous. Inside front cover — editorial board page/cover image legend if applicable. Ecological Modelling, 355(??):ifc, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017302685.

Anonymous:2017:IFCc

[Ano17o] Anonymous. Inside front cover — editorial board page/cover image legend if applicable. Ecological Modelling, 356(??):ifc, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017302879.

Anonymous:2017:IFCd

[Ano17p]  Anonymous.  Inside front cover — editorial board

page/cover image legend if applicable. Ecological Modelling, 357(??):ifc, August 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303058.

Anonymous:2017:IFCe

[Ano17q]  Anonymous.  Inside front cover — editorial board

page/cover image legend if applicable. Ecological Modelling, 358(??):ifc, August 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303162.

Anonymous:2017:PAb

[Ano17r] Anonymous. Pages 1–108 (24 April 2017). Ecological Modelling, 350(??):1–108, April 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PFb

[Ano17s] Anonymous. Pages 1–118 (24 February 2017). Ecological Modelling, 346(??):1–118, February 24, 2017. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:POb

[Ano17t] Anonymous. Pages 1–120 (24 October 2017). Ecological Modelling, 362(??):1–120, October 24, 2017. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PJd

[Ano17u] Anonymous. Pages 1–126 (10 July 2017). Ecological Modelling, 355(??):1–126, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PNb

[Ano17v] Anonymous. Pages 1–132 (24 November 2017). Ecological Modelling, 364(??):1–132, November 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PMb

[Ano17w] Anonymous. Pages 1–134 (24 March 2017). Ecological Modelling, 348(??):1–134, March 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PDa

[Ano17x] Anonymous. Pages 1–140 (10 December 2017). Ecological Modelling, 365(??):1–140, December 10, 2017. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PMc

[Ano17y] Anonymous. Pages 1–140 (10 May 2017). Ecological Modelling, 351(??):1–140, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PJa

[Ano17z] Anonymous. Pages 1–142 (10 January 2017). Ecological Modelling, 343(??):1–142, January 10, 2017. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PMd

[Ano17-27] Anonymous. Pages 1–152 (24 May 2017). Ecological Modelling, 352(??):1–152, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PJe

[Ano17-28] Anonymous. Pages 1–172 (24 July 2017). Ecological Modelling, 356(??):1–172, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PFa

[Ano17-29] Anonymous. Pages 1–176 (10 February 2017). Ecological Modelling, 345(??):1–176, February 10, 2017. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PJc

[Ano17-30] Anonymous. Pages 1–176 (24 June 2017). Ecological Modelling, 354(??):1–176, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:POa

[Ano17-31] Anonymous. Pages 1–196 (10 October 2017). Ecological Modelling, 361(??):1–196, October 10, 2017. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PNa

[Ano17-32] Anonymous. Pages 1–236 (10 November 2017). Ecological Modelling, 363(??):1–236, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PSb

[Ano17-33] Anonymous. Pages 1–436 (24 September 2017). Ecological Modelling, 360(??):1–436, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PSa

[Ano17-34] Anonymous. Pages 1–468 (10 September 2017). Ecological Modelling, 359(??):1–468, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PAc

[Ano17-35] Anonymous. Pages 1–58 (10 August 2017). Ecological Modelling, 357(??):1–58, August 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PAd

[Ano17-36] Anonymous. Pages 1–58 (24 August 2017). Ecological Modelling, 358(??):1–58, August 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PDb

[Ano17-37] Anonymous. Pages 1–68 (24 December 2017). Ecological Modelling, 366(??):1–68, December 24, 2017. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PAa

[Ano17-38] Anonymous. Pages 1–86 (10 April 2017). Ecological Modelling, 349(??):1–86, April 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PMa

[Ano17-39] Anonymous. Pages 1–94 (10 March 2017). Ecological Modelling, 347(??):1–94, March 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PJb

[Ano17-40] Anonymous. Pages 1–94 (24 January 2017). Ecological Modelling, 344(??):1–94, January 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2017:PIIa

[Ano17-41] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 343(??):iii, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306809.

Anonymous:2017:PIIb

[Ano17-42] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 344(??):iii, January 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307803.

Anonymous:2017:PIIc

[Ano17-43] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 345(??):iii, February 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308328.

Anonymous:2017:PIId

[Ano17-44] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 346(??):iii, February 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300248.

Anonymous:2017:PIIe

[Ano17-45] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 347(??):iii, March 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300522.

Anonymous:2017:PIIf

[Ano17-46] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 348(??):iii, March 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301126.

Anonymous:2017:PIIg

[Ano17-47] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 349(??):iii, April 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301667.

Anonymous:2017:PIIh

[Ano17-48] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 350(??):iii, April 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301953.

Anonymous:2017:PIIi

[Ano17-49] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 351(??):iii, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017302119.

Anonymous:2017:PIIj

[Ano17-50] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 352(??):iii, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017302235.

Anonymous:2017:PIIk

[Ano17-51] Anonymous. Prelim (ii or iii) title page. Ecological Modelling, 353(??):iii, June 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302375.

Anonymous:2018:EBa

[Ano18a] Anonymous. Editorial Board. Ecological Modelling, 368(??):ii, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018300048.

Anonymous:2018:EBb

[Ano18b] Anonymous. Editorial Board. Ecological Modelling, 369(??):ii, February 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018300206.

Anonymous:2018:EBc

[Ano18c] Anonymous. Editorial Board. Ecological Modelling, 370(??):ii, February 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018300292.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Ecological Modelling, 371 (??):ii, March 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018300474.

Anonymous:2018:EBe

[Ano18e] Anonymous. Editorial Board. Ecological Modelling, 372 (??):ii, March 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018300577.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Ecological Modelling, 373 (??):ii, April 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001830070X.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. Ecological Modelling, 374 (??):ii, April 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018300826.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. Ecological Modelling, 375 (??):ii, May 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018300930.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. Ecological Modelling, 376 (??):ii, May 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018301030.

Anonymous:2018:EBj

[Ano18j] Anonymous. Editorial Board. Ecological Modelling, 377 (??):ii, June 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018301194.

Anonymous:2018:EBk

[Ano18k] Anonymous. Editorial Board. Ecological Modelling, 378 (??):ii, June 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018301364.

Anonymous:2018:EBl

[Ano18l] Anonymous. Editorial Board. Ecological Modelling, 379 (??):ii, July 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018301509.

Anonymous:2018:EBm

[Ano18m] Anonymous. Editorial Board. Ecological Modelling, 380 (??):ii, July 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018301625.

Anonymous:2018:EBn

[Ano18n] Anonymous. Editorial Board. Ecological Modelling, 381 (??):ii, August 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018301790.

Anonymous:2018:EBo

[Ano18o] Anonymous. Editorial Board. Ecological Modelling, 382 (??):ii, August 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001830190X.

Anonymous:2018:EBp

[Ano18p] Anonymous. Editorial Board. Ecological Modelling, 383(??):ii, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302072.

Anonymous:2018:EBq

[Ano18q] Anonymous. Editorial Board. Ecological Modelling, 384(??):ii, September 24, 2018. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302539.

Anonymous:2018:EBr

[Ano18r] Anonymous. Editorial Board. Ecological Modelling, 385(??):ii, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302734.

Anonymous:2018:EBs

[Ano18s] Anonymous. Editorial Board. Ecological Modelling, 386(??):ii, October 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302965.

Anonymous:2018:EBt

[Ano18t] Anonymous. Editorial Board. Ecological Modelling, 387(??):ii, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303685.

Anonymous:2018:EBu

[Ano18u] Anonymous. Editorial Board. Ecological Modelling, 388(??):ii, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303545.

Anonymous:2018:EBv

[Ano18v] Anonymous. Editorial Board. Ecological Modelling, 389(??):ii, December 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303831.

Anonymous:2018:EBw

[Ano18w] Anonymous. Editorial Board. Ecological Modelling, 390(??):ii, December 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303776.

Anonymous:2018:IFC

[Ano18x] Anonymous. Inside front cover — Editorial Board page. Ecological Modelling, 367(??):ifc, January 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304702.

Anonymous:2018:PFa

[Ano18y] Anonymous. Pages 1–100 (10 February 2018). Ecological Modelling, 369(??):1–100, February 10, 2018. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:POb

[Ano18z] Anonymous. Pages 1–114 (24 October 2018). Ecological Modelling, 386(??):1–114, October 24, 2018. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PMa

[Ano18-27] Anonymous. Pages 1–122 (10 March 2018). Ecological Modelling, 371(??):1–122, March 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PNb

[Ano18-28] Anonymous. Pages 1–148 (24 November 2018). Ecological Modelling, 388(??):1–148, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PSa

[Ano18-29] Anonymous. Pages 1–170 (10 September 2018). Ecological Modelling, 383(??):1–170, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:POa

[Ano18-30] Anonymous. Pages 1–226 (10 October 2018). Ecological Modelling, 385(??):1–226, October 10, 2018. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PNa

[Ano18-31] Anonymous. Pages 1–238 (10 November 2018). Ecological Modelling, 387(??):1–238, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PSb

[Ano18-32] Anonymous. Pages 1–370 (24 September 2018). Ecological Modelling, 384(??):1–370, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PJb

[Ano18-33] Anonymous. Pages 1–424 (24 January 2018). Ecological Modelling, 368(??):1–424, January 24, 2018. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PDa

[Ano18-34] Anonymous. Pages 1–50 (10 December 2018). Ecological Modelling, 389(??):1–50, December 10, 2018. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PJf

[Ano18-35] Anonymous. Pages 1–52 (24 July 2018). Ecological Modelling, 380(??):1–52, July 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PMc

[Ano18-36] Anonymous. Pages 1–54 (10 May 2018). Ecological Modelling, 375(??):1–54, May 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PJd

[Ano18-37] Anonymous. Pages 1–58 (24 June 2018). Ecological Modelling, 378(??):1–58, June 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PAb

[Ano18-38] Anonymous. Pages 1–60 (24 April 2018). Ecological Modelling, 374(??):1–60, April 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PJc

[Ano18-39] Anonymous. Pages 1–66 (10 June 2018). Ecological Modelling, 377(??):1–66, June 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PFb

[Ano18-40] Anonymous. Pages 1–66 (24 February 2018). Ecological Modelling, 370(??):1–66, February 24, 2018. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PAa

[Ano18-41] Anonymous. Pages 1–68 (10 April 2018). Ecological Modelling, 373(??):1–68, April 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PJa

[Ano18-42] Anonymous. Pages 1–68 (10 January 2018). Ecological Modelling, 367(??):1–68, January 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PMb

[Ano18-43] Anonymous. Pages 1–70 (24 March 2018). Ecological Modelling, 372(??):1–70, March 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PJe

[Ano18-44] Anonymous. Pages 1–72 (10 July 2018). Ecological Modelling, 379(??):1–72, July 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PMd

[Ano18-45] Anonymous. Pages 1–76 (24 May 2018). Ecological Modelling, 376(??):1–76, May 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PAc

[Ano18-46] Anonymous. Pages 1–78 (10 August 2018). Ecological Modelling, 381(??):1–78, August 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PAd

[Ano18-47] Anonymous. Pages 1–82 (24 August 2018). Ecological Modelling, 382(??):1–82, August 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2018:PDb

[Ano18-48] Anonymous. Pages 1–88 (24 December 2018). Ecological Modelling, 390(??):1–88, December 24, 2018. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Da

[Ano19a] Anonymous. 1 December 2019. Ecological Modelling, 413 (??):??, December 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Na

[Ano19b] Anonymous. 1 November 2019. Ecological Modelling, 411 (??):??, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Oa

[Ano19c] Anonymous. 1 October 2019. Ecological Modelling, 409(??): ??, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Sa

[Ano19d] Anonymous. 1 September 2019. Ecological Modelling, 407 (??):??, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Db

[Ano19e] Anonymous. 15 December 2019. Ecological Modelling, 414 (??):??, December 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Nb

[Ano19f] Anonymous. 15 November 2019. Ecological Modelling, 412 (??):??, November 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Ob

[Ano19g] Anonymous. 15 October 2019. Ecological Modelling, 410 (??):??, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:Sb

[Ano19h] Anonymous. 15 September 2019. Ecological Modelling, 408 (??):??, September 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:EBa

[Ano19i] Anonymous. Editorial Board. Ecological Modelling, 391(??):ii, January 10, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303946.

Anonymous:2019:EBb

[Ano19j] Anonymous. Editorial Board. Ecological Modelling, 392(??):ii, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001930047X.

Anonymous:2019:EBc

[Ano19k] Anonymous. Editorial Board. Ecological Modelling, 393 (??):ii, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019300134.

Anonymous:2019:EBd

[Ano19l] Anonymous. Editorial Board. Ecological Modelling, 394(??):ii, February 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001930033X.

Anonymous:2019:EBe

[Ano19m] Anonymous. Editorial Board. Ecological Modelling, 395 (??):ii, March 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019300535.

Anonymous:2019:EBf

[Ano19n] Anonymous. Editorial Board. Ecological Modelling, 396 (??):ii, March 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019300626.

Anonymous:2019:EBg

[Ano19o] Anonymous. Editorial Board. Ecological Modelling, 397 (??):ii, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019300729.

Anonymous:2019:EBh

[Ano19p] Anonymous. Editorial Board. Ecological Modelling, 398 (??):ii, April 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019300948.

Anonymous:2019:EBi

[Ano19q] Anonymous. Editorial Board. Ecological Modelling, 399 (??):ii, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019301139.

Anonymous:2019:EBj

[Ano19r] Anonymous. Editorial Board. Ecological Modelling, 400 (??):ii, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001930136X.

Anonymous:2019:EBk

[Ano19s] Anonymous. Editorial Board. Ecological Modelling, 401 (??):ii, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019301528.

Anonymous:2019:EBl

[Ano19t] Anonymous. Editorial Board. Ecological Modelling, 402 (??):ii, June 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019301991.

Anonymous:2019:EBm

[Ano19u] Anonymous. Editorial Board. Ecological Modelling, 403 (??):ii, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019301772.

Anonymous:2019:EBn

[Ano19v] Anonymous. Editorial Board. Ecological Modelling, 404 (??):ii, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019301875.

Anonymous:2019:EBo

[Ano19w] Anonymous. Editorial Board. Ecological Modelling, 405 (??):ii, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001930208X.

Anonymous:2019:EBp

[Ano19x] Anonymous. Editorial Board. Ecological Modelling, 406 (??):ii, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019302376.

Anonymous:2019:EBq

[Ano19y] Anonymous. Editorial Board. Ecological Modelling, 407(??):Article 108750, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302583.

Anonymous:2019:EBr

[Ano19z] Anonymous. Editorial Board. Ecological Modelling, 408(??):Article 108769, September 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302777.

Anonymous:2019:EBs

[Ano19-27] Anonymous. Editorial Board. Ecological Modelling, 409 (??):Article 108789, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302972.

Anonymous:2019:EBt

[Ano19-28] Anonymous. Editorial Board. Ecological Modelling, 410 (??):Article 108811, October 15, 2019. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019303199.

Anonymous:2019:EBu

[Ano19-29] Anonymous. Editorial Board. Ecological Modelling, 411(??):Article 108850, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303588.

Anonymous:2019:EBv

[Ano19-30] Anonymous. Editorial Board. Ecological Modelling, 412(??):Article 108842, November 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303503.

Anonymous:2019:EBw

[Ano19-31] Anonymous. Editorial Board. Ecological Modelling, 413(??):Article 108866, December 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303746.

Anonymous:2019:EBx

[Ano19-32] Anonymous. Editorial Board. Ecological Modelling, 414(??):Article 108885, December 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930393X.

Anonymous:2019:PAc

[Ano19-33] Anonymous. Pages 1–106 (1 August 2019). Ecological Modelling, 405(??):1–106, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PJd

[Ano19-34] Anonymous. Pages 1–106 (24 June 2019). Ecological Modelling, 402(??):1–106, June 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PJf

[Ano19-35] Anonymous. Pages 1–112 (24 July 2019). Ecological Modelling, 404(??):1–112, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PAa

[Ano19-36] Anonymous. Pages 1–140 (1 April 2019). Ecological Modelling, 397(??):1–140, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PAd

[Ano19-37] Anonymous. Pages 1–144 (24 August 2019). Ecological Modelling, 406(??):1–144, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PJc

[Ano19-38] Anonymous. Pages 1–154 (1 June 2019). Ecological Modelling, 401(??):1–154, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PFa

[Ano19-39] Anonymous. Pages 1–156 (1 February 2019). Ecological Modelling, 393(??):1–156, February 1, 2019. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PJb

[Ano19-40] Anonymous. Pages 1–278 (24 January 2019). Ecological Modelling, 392(??):1–278, January 24, 2019. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PMa

[Ano19-41] Anonymous. Pages 1–50 (1 March 2019). Ecological Modelling, 395(??):1–50, March 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PJa

[Ano19-42] Anonymous. Pages 1–62 (10 January 2019). Ecological Modelling, 391(??):1–62, January 10, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PMd

[Ano19-43] Anonymous. Pages 1–72 (24 May 2019). Ecological Modelling, 400(??):1–72, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PMc

[Ano19-44] Anonymous. Pages 1–76 (1 May 2019). Ecological Modelling, 399(??):1–76, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PFb

[Ano19-45] Anonymous. Pages 1–78 (24 February 2019). Ecological Modelling, 394(??):1–78, February 24, 2019. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PJe

[Ano19-46] Anonymous. Pages 1–84 (1 July 2019). Ecological Modelling, 403(??):1–84, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PAb

[Ano19-47] Anonymous. Pages 1–84 (24 April 2019). Ecological Modelling, 398(??):1–84, April 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PMb

[Ano19-48] Anonymous. Pages 1–86 (24 March 2019). Ecological Modelling, 396(??):1–86, March 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2019:PN

[Ano19-49] Anonymous. Publisher’s note. Ecological Modelling, 407(??):Article 108751, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302595.

Anonymous:2021:EB

[Ano21a] Anonymous. Editorial Board. Ecological Modelling, 222 (20–22):iii, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011005096.

Anonymous:2021:PA

[Ano21b] Anonymous. Pages 1–170 (10–24 August 2015). Ecological Modelling, 309–310(??):1–170, ???? 2021. CODEN EC-MODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:PON

[Ano21c] Anonymous. Pages 3651–3810 (October–November 2011). Ecological Modelling, 222(20–22):??, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic).

Anonymous:2021:PTP

[Ano21d] Anonymous. Preliminary title page. Ecological Modelling, 309–310(??):iii, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002264.

Antonarakis:2014:UIF

[Ant14] A. S. Antonarakis. Uncertainty in initial forest structure and composition when predicting carbon dynamics in a temperate forest. Ecological Modelling, 291(??):134–141, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003780.

Aznar:2013:MCH

[AOD^{+}13] Victoria Romeo Aznar, Marcelo Otero, María Sol De Majo, Sylvia Fischer, and Hernán G. Solari. Modeling the complex hatching and development of Aedes aegypti in temperate climates. Ecological Modelling, 253(??):44–55, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005704.

Arundel:2012:MHB

[AOW12] Jonathan Arundel, Benjamin P. Oldroyd, and Stephan Winter. Modelling honey bee queen mating as a measure of feral colony density. Ecological Modelling, 247(??):48–57, December 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012003869.

Arundel:2013:MEH

[AOW13] Jonathan Arundel, Benjamin P. Oldroyd, and Stephan Winter. Modelling estimates of honey bee (Apis spp.) colony density from drones. Ecological Modelling, 267(??):1–10, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013003530.

Arenas:2015:AMB

[APWT15] Antonio Arenas, Marcela Politano, Larry Weber, and Mark Timko. Analysis of movements and behavior of smolts swimming in hydropower reservoirs. Ecological Modelling, 312(??):292–307, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002124.

Acevedo:2011:BDB

[AR11] Pelayo Acevedo and Raimundo Real. Biogeographical differences between the two Capra pyrenaica subspecies, C. p. victoriae and C. p. hispanica, inhabiting the Iberian Peninsula: Implications for conservation. Ecological Modelling, 222(3):814–823, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005478.

Anyomi:2012:UHG

[ARM^{+}12] Kenneth A. Anyomi, Frédéric Raulier, Daniel Mailly, Martin P. Girardin, and Yves Bergeron. Using height growth to model local and regional response of trembling aspen (Populus tremuloides Michx.) to climate within the boreal forest of western Québec. Ecological Modelling, 243(??):123–132, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002992.

Arbault:2014:FGS

[ARTBB14] Damien Arbault, Benedetto Rugani, Ligia Tiruta-Barna, and Enrico Benetto. A first global and spatially explicit emergy database of rivers and streams based on highresolution GIS-maps. Ecological Modelling, 281(??):52–64, June 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014001288.

Ayllon:2016:IGM

[ARV^{+}16] Daniel Ayllón, Steven F. Railsback, Simone Vincenzi, Jürgen Groeneveld, Ana Almodóvar, and Volker Grimm. InSTREAM-Gen: Modelling eco-evolutionary dynamics of trout populations under anthropogenic environmental change. Ecological Modelling, 326(??):36–53, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003324.

Alarcon:2011:MCG

[AS11] Vladimir J. Alarcon and Gretchen F. Sassenrath. Modeling cotton (Gossypium spp.) leaves and canopy using computer aided geometric design (CAGD). Ecological Modelling, 222(12):1951–1963, June 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001748.

Arseneault:2012:ICG

[AS12] Justin E. Arseneault and Mike R. Saunders. Incorporating canopy gap-induced growth responses into spatially implicit growth model projections. Ecological Modelling, 237–238(??):120–131, July 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200169X.

Arreguin-Sanchez:2014:MRA

[AS14] Francisco Arreguín-Sánchez. Measuring resilience in aquatic trophic networks from supply–demand-of-energy relationships. Ecological Modelling, 272(??):271–276, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004857.

Ashauer:2010:TTM

[Ash10] Roman Ashauer. Toxicokinetic–toxicodynamic modelling in an individual based context — consequences of parameter variability. Ecological Modelling, 221(9):1325–1328, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010000542.

Areerachakul:2013:IUS

[ASL13] Sirilak Areerachakul, Peraphon Sophatsathit, and Chidchanok Lursinsap. Integration of unsupervised and supervised neural networks to predict dissolved oxygen concentration in canals. Ecological Modelling, 261–262(??):1–7, July 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001919.

Aretano:2015:MEV

[ASP^{+}15] Roberta Aretano, Teodoro Semeraro, Irene Petrosillo, Antonella De Marco, Maria Rita Pasimeni, and Giovanni Zurlini. Mapping ecological vulnerability to fire for effective conservation management of natural protected areas. Ecological Modelling, 295(??):163–175, January 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004323.

Arancio:2014:PLC

[ASS14] Marc Arancio, Marc Sourisseau, and Sami Souissi. Processes leading to the coexistence of a host and its parasitoid in homogeneous environments: the role of an infected dormant stage. Ecological Modelling, 279(??):78–88, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014001008.

Attorre:2019:TMP

[ASV19] F. Attorre, E. Sciubba, and M. Vitale. A thermodynamic model for plant growth, validated with Pinus sylvestris data. Ecological Modelling, 391(??):53–62, January 10, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303594.

Alderman:2019:MIN

[ATCJ^{+}19] R. Alderman, G. N. Tuck, C. Castillo-Jordán, M. Haddon, and A. E. Punt. Macquarie Island’s northern giant petrels and the impacts of pest eradication on population abundance. Ecological Modelling, 393(??):66–75, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303727.

Aagaard:2018:TIE

[ATL18] Kevin J. Aagaard, Wayne E. Thogmartin, and Eric V. Lonsdorf. Temperature-influenced energetics model for migrating waterfowl. Ecological Modelling, 378(??):46–58, June 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018301121.

Augusiak:2014:MVE

[AVG14] Jacqueline Augusiak, Paul J. Van den Brink, and Volker Grimm. Merging validation and evaluation of ecological models to ‘evaludation’: a review of terminology and a practical approach. Ecological Modelling, 280(??):117–128, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005450.

Adam:2011:EMD

[AVL^{+}11] M. Adam, L. G. J. Van Bussel, P. A. Leffelaar, H. Van Keulen, and F. Ewert. Effects of modelling detail on simulated potential crop yields under a wide range of climatic conditions. Ecological Modelling, 222(1):131–143, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004564.

Ainsworth:2015:TCM

[AW15] C. H. Ainsworth and C. J. Walters. Ten common mistakes made in Ecopath with Ecosim modelling. Ecological Modelling, 308(??):14–17, July 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001301.

Adhikari:2016:CCI

[AW16] Arjun Adhikari and Joseph D. White. Climate change impacts on regenerating shrubland productivity. Ecological Modelling, 337(??):211–220, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302459.

Acosta:2010:FER

[AWF10] Hernando Acosta, Dongrui Wu, and Barrie M. Forrest. Fuzzy experts on recreational vessels, a risk modelling approach for marine invasions. Ecological Modelling, 221(5):850–863, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008321.

Adam:2013:SAR

[AWL^{+}13] M. Adam, J. Wery, P. A. Leffelaar, F. Ewert, M. Corbeels, and H. Van Keulen. A systematic approach for re-assembly of crop models: an example to simulate pea growth from wheat growth. Ecological Modelling, 250(??):258–268, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005145.

Anteau:2014:MPA

[AWSS14] Michael J. Anteau, Mark T. Wiltermuth, Mark H. Sherfy, and Terry L. Shaffer. Measuring and predicting abundance and dynamics of habitat for piping plovers on a large reservoir. Ecological Modelling, 272(??):16–27, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004225.

Accatino:2016:TGF

[AWWD16] Francesco Accatino, Kerstin Wiegand, David Ward, and Carlo De Michele. Trees, grass, and fire in humid savannas — the importance of life history traits and spatial processes. Ecological Modelling, 320(??):135–144, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004366.

Alvarez-Yela:2017:IAA

[AYASR^{+}17] Astrid Catalina Alvarez-Yela, María Camila Alvarez-Silva, Silvia Restrepo, Johana Husserl, María Mercedes Zambrano, Giovanna Danies, Jorge M. Gómez, and Andrés Fernando González Barrios. Influence of agricultural activities in the structure and metabolic functionality of paramo soil samples in Colombia studied using a metagenomics analysis in dynamic state. Ecological Modelling, 351(??):63–76, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016305002.

Bailey:2014:STV

[BA14a] Ryan T. Bailey and Mehdi Ahmadi. Spatial and temporal variability of in-stream water quality parameter influence on dissolved oxygen and nitrate within a regional stream network. Ecological Modelling, 277(??):87–96, April 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000581.

Bone:2014:MMS

[BA14b] Christopher Bone and Mark Altaweel. Modeling micro-scale ecological processes and emergent patterns of mountain pine beetle epidemics. Ecological Modelling, 289(??):45–58, October 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003007.

Bindewald:2017:AMG

[BA17] Eckart Bindewald and Shady S. Atallah. Achieving multiple goals via voluntary efforts and motivation asymmetry. Ecological Modelling, 354(??):37–48, June 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304161.

Bala:2014:SEH

[BAA^{+}14] B. K. Bala, Fatimah M. Arshad, E. F. Alias, S. F. Sidique, K. M. Noh, M. K. Rowshon, Q. M. M. Islam, and M. M.

Islam. Sustainable exploitation of Hilsa fish (Tenualosa ilisha) population in Bangladesh: Modeling and policy implications. Ecological Modelling, 283(??):19–30, July 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014001574.

Baggio:2018:DCG

[BAAB18] Rafael A. Baggio, Sabrina B. L. Araujo, Daniel Ayllón, and Walter A. Boeger. Dams cause genetic homogenization in populations of fish that present homing behavior: Evidence from a demogenetic individual-based model. Ecological Modelling, 384(??):209–220, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302242.

Baird:2016:BRS

[BAB^{+}16] Mark E. Baird, Matthew P. Adams, Russell C. Babcock, Kadija Oubelkheir, Mathieu Mongin, Karen A. Wild-Allen, Jennifer Skerratt, Barbara J. Robson, Katherina Petrou, Peter J. Ralph, Katherine R. O’Brien, Alex B. Carter, Jessie C. Jarvis, and Michael A. Rasheed. A biophysical representation of seagrass growth for application in a complex shallow-water biogeochemical model. Ecological Modelling, 325(??):13–27, April 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005797.

Baeza:2018:MCT

[Bae18] Andres Baeza. Modelling the critical transition from Chilean evergreen forest to savanna: Early warning signals and livestock management. Ecological Modelling, 388 (??):115–123, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303193.

Bailey:2018:RRT

[Bai18] Brian N. Bailey. A reverse ray-tracing method for modelling the net radiative flux in leaf-resolving plant canopy simulations. Ecological Modelling, 368(??):233–245, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304842.

Brun:2011:EHS

[BAJ^{+}11] Mélanie Brun, Christophe Abraham, Marc Jarry, Jacques Dumas, Frédéric Lange, and Etienne Prévost. Estimating an homogeneous series of a population abundance indicator despite changes in data collection procedure: a hierarchical Bayesian modelling approach. Ecological Modelling, 222(5):1069–1079, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006332.

Banas:2011:ACT

[Ban11] Neil S. Banas. Adding complex trophic interactions to a size-spectral plankton model: Emergent diversity patterns and limits on predictability. Ecological Modelling, 222(15):2663–2675, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002985.

Behera:2018:BMM

[BAP18] Bunushree Behera, Nazimdhine Aly, and Balasubramanian P. Biophysical modeling of microalgal cultivation in open ponds. Ecological Modelling, 388(??):61–71, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303211.

Bartelt:2011:EME

[Bar11] Paul E. Bartelt. Erratum to “Modeling energetics, habitat suitability, and movements of western toads, Anaxyrus (=Bufo) boreas, across present and future landscapes” [Ecol. Modell. 221 (2010) 2675–2686]. Ecological Modelling, 222(1):219, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001000520X.

Brochier:2015:IAH

[BAT^{+}15] Timothée Brochier, Pierre Auger, Ndiaga Thiam, Momar Sow, Sidiya Diouf, Hans Sloterdijk, and Patrice Brehmer. Implementation of artificial habitats: Inside or outside the marine protected areas? Insights from a mathematical approach. Ecological Modelling, 297(??):98–106, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005328.

Baral:2010:EAU

[BB10] Anil Baral and Bhavik R. Bakshi. Energy analysis using US economic input–output models with applications to life cycles of gasoline and corn ethanol. Ecological Modelling, 221(15):1807–1818, July 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002127.

Boria:2018:ELS

[BB18] Robert A. Boria and Jessica L. Blois. The effect of large sample sizes on ecological niche models: Analysis using a North American rodent, Peromyscus maniculatus. Ecological Modelling, 386(??):83–88, October 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302825.

Bodini:2012:CEG

[BBA12] Antonio Bodini, Cristina Bondavalli, and Stefano Allesina. Cities as ecosystems: Growth, development and implications for sustainability. Ecological Modelling, 245 (??):185–198, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012000981.

Banks:2017:MBB

[BBB^{+}17] H. T. Banks, J. E. Banks, Riccardo Bommarco, A. N. Laubmeier, N. J. Myers, Maj Rundlöf, and Kristen Tillman. Modeling bumble bee population dynamics with delay differential equations. Ecological Modelling, 351(??):14–23, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017301606.

Bisson:2019:WAM

[BBC^{+}19] Anne Bisson, Simon Boudsocq, Céline Casenave, Sébastien Barot, Raphaël J. Manlay, Jonathan Vayssières, Dominique Masse, and Tanguy Daufresne. West African mixed farming systems as meta-ecosystems: a sourcesink modelling approach. Ecological Modelling, 412(??): Article 108803, November 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019303114.

Beget:2013:PSR

[BBDB13] M. E. Beget, V. A. Bettachini, C. M. Di Bella, and F. Baret. SAILHFlood: a radiative transfer model for flooded vegetation. Ecological Modelling, 257(??):25–35, May 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001129.

Bidegain:2013:LPC

[BBGJ13] Gorka Bidegain, Javier Francisco Bárcena, Andrés García, and José Antonio Juanes. LARVAHS: Predicting clam larval dispersal and recruitment using habitat suitabilitybased particle tracking model. Ecological Modelling, 268(??):78–92, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001300375X.

Braakhekke:2011:SVE

[BBH^{+}11] Maarten C. Braakhekke, Christian Beer, Marcel R. Hoosbeek, Markus Reichstein, Bart Kruijt, Marion Schrumpf, and Pavel Kabat. SOMPROF: a vertically explicit soil organic matter model. Ecological Modelling, 222 (10):1712–1730, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011000962.

Brandt:2017:CCE

[BBH^{+}17] Laura A. Brandt, Allison M. Benscoter, Rebecca Harvey, Carolina Speroterra, David Bucklin, Stephanie S. Romañach, James I. Watling, and Frank J. Mazzotti. Comparison of climate envelope models developed using expert-selected variables versus statistical selection. Ecological Modelling, 345(??):10–20, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307293.

Benadi:2013:CSS

[BBHP13] Gita Benadi, Nico Blüthgen, Thomas Hovestadt, and Hans-Joachim Poethke. Contrasting specialization–stability relationships in plant–animal mutualistic systems. Ecological Modelling, 258(??):65–73, June 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001518.

Barve:2011:CRA

[BBJV^{+}11] Narayani Barve, Vijay Barve, Alberto Jiménez-Valverde, Andrés Lira-Noriega, Sean P. Maher, A. Townsend Peterson, Jorge Soberón, and Fabricio Villalobos. The crucial role of the accessible area in ecological niche modeling and species distribution modeling. Ecological Modelling, 222(11):1810–1819, June 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000780.

Biesinger:2011:PLP

[BBL11a] Zy Biesinger, Benjamin M. Bolker, and William J. Lindberg. Predicting local population distributions around a central shelter based on a predation risk-growth trade-off. Ecological Modelling, 222(8):1448–1455, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000767.

Bornhofen:2011:ECL

[BBL11b] S. Bornhofen, S. Barot, and C. Lattaud. The evolution of CSR life-history strategies in a plant model with explicit physiology and architecture. Ecological Modelling, 222(1):1–10, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004850.

Beyer:2017:VFS

[BBL^{+}17] Robert Beyer, Dominik Bayer, Véronique Letort, Hans Pretzsch, and Paul-Henry Cournède. Validation of a functionalstructural tree model using terrestrial lidar data. Ecological Modelling, 357(??):55–57, August 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301722.

Banks:2015:OSF

[BBRJ15] J. E. Banks, H. T. Banks, K. Rinnovatore, and C. M. Jackson. Optimal sampling frequency and timing of threatened tropical bird populations: a modeling approach. Ecological Modelling, 303(??):70–77, May 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000617.

Bevan:2011:TFM

[BC11] Andrew Bevan and James Conolly. Terraced fields and Mediterranean landscape structure: an analytical case study from Antikythera, Greece. Ecological Modelling, 222(7):1303–1314, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006824.

Barberis:2012:DIG

[BC12] L. Barberis and C. A. Condat. Describing interactive growth using vector universalities. Ecological Modelling, 227(??):56–63, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005990.

Bellot:2013:HEH

[BC13] Juan Bellot and Esteban Chirino. Hydrobal: an ecohydrological modelling approach for assessing water balances in different vegetation types in semi-arid areas. Ecological Modelling, 266(??):30–41, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300327X.

Baklouti:2011:SPD

[BCB^{+}11] M. Baklouti, C. Chevalier, M. Bouvy, D. Corbin, M. Pagano, M. Troussellier, and R. Arfi. A study of plankton dynamics under osmotic stress in the Senegal River Estuary, West Africa, using a 3D mechanistic model. Ecological Modelling, 222(15):2704–2721, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002596.

Bai:2011:SDE

[BCC^{+}11] Junhong Bai, Baoshan Cui, Bin Chen, Kejiang Zhang, Wei Deng, Haifeng Gao, and Rong Xiao. Spatial distribution and ecological risk assessment of heavy metals in surface sediments from a typical plateau lake wetland, China. Ecological Modelling, 222(2):301–306, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000900831X.

Blanco:2014:KBH

[BCC^{+}14] María Valeria Blanco, Diego Ignacio Cattoni, Pedro Carriquiriborde, José Raúl Grigera, and Osvaldo Chara. Kinetics of bioaccumulation of heavy metals in Odontesthes bonariensis is explained by a single and common mechanism. Ecological Modelling, 274(??):50–56, February 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005802.

Bregaglio:2012:ESG

[BCD12] Simone Bregaglio, Giovanni Cappelli, and Marcello Donatelli. Evaluating the suitability of a generic fungal infection model for pest risk assessment studies. Ecological Modelling, 247(??):58–63, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004012.

Brown:2016:GEBb

[BCDU16] Mark T. Brown, Daniel E. Campbell, Christopher De Vilbiss, and Sergio Ulgiati. The geobiosphere emergy baseline: a synthesis. Ecological Modelling, 339(??):92–95, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300916.

Bonhomme:2017:CCP

[BCI^{+}17] Vincent Bonhomme, Mathieu Castets, Thomas Ibanez, Hubert Géraux, Christelle Hély, and Cédric Gaucherel. Configurational changes of patchy landscapes dynamics. Ecological Modelling, 363(??):1–7, November 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300959.

Breed:2012:SSM

[BCJ^{+}12] Greg A. Breed, Daniel P. Costa, Ian D. Jonsen, Patrick W. Robinson, and Joanna Mills-Flemming. State-space methods for more completely capturing behavioral dynamics from animal tracks. Ecological Modelling, 235–236(??):49–58, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012001378.

Blanco:2014:MSD

[BCL^{+}14] Aníbal M. Blanco, Guillermo R. Chantre, Mariela V. Lodovichi, J. Alberto Bandoni, Ricardo L. López, Mario R. Vigna, Ramón Gigón, and Mario R. Sabbatini. Modeling seed dormancy release and germination for predicting Avena fatua L. field emergence: a genetic algorithm approach. Ecological Modelling, 272(??):293–300, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004808.

Berrios:2017:EEB

[BCO17] Fernando Berrios, Daniel E. Campbell, and Marco Ortiz. Emergy evaluation of benthic ecosystems influenced by upwelling in northern Chile: Contributions of the ecosystems to the regional economy. Ecological Modelling, 359(??):146–164, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302314.

Bae:2015:MBC

[BCP15] Mi-Jung Bae, Tae-Soo Chon, and Young-Seuk Park. Modeling behavior control of golden apple snails at different temperatures. Ecological Modelling, 306(??):86–94, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001400489X.

Bastianoni:2016:ISO

[BCP16] Simone Bastianoni, Luca Coscieme, and Federico M. Pulselli. The input-state-output model and related indicators to investigate the relationships among environment, society and economy. Ecological Modelling, 325(??):84–88, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004840.

Brown:2015:ESE

[BCT15] Mark T. Brown, Daniel E. Campbell, and David R. Tilley. Emergy synthesis 8 ∼ emergy and environmental accounting: Theories, applications, and methodologies. Ecological Modelling, 315(??):1–3, November 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003816.

Battle:2016:IDR

[BCTL16] Laurie Battle, Hao-Yen Chang, Chyng-Shyan Tzeng, and Hsing-Juh Lin. The impact of dam removal and climate change on the abundance of the Formosan landlocked salmon. Ecological Modelling, 339(??):23–32, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302915.

Brown:2016:GEBa

[BCUF16] Mark T. Brown, Daniel E. Campbell, Sergio Ulgiati, and Pier Paolo Franzese. The geobiosphere emergy baseline: a synthesis. Ecological Modelling, 339(??):89–91, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303969.

Bellingeri:2013:IER

[BCV13] Michele Bellingeri, Davide Cassi, and Simone Vincenzi. Increasing the extinction risk of highly connected species causes a sharp robust-to-fragile transition in empirical food webs. Ecological Modelling, 251(??):1–8, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005789.

Bosi:2018:EMM

[BD18] Stefano Bosi and David Desmarchelier. An economic model of metapopulation dynamics. Ecological Modelling, 387(??):196–204, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303090.

Bianchini:2011:MPA

[BdCS11] Irineu Bianchini and Marcela Bianchessi da Cunha Santino. Model parameterization for aerobic decomposition of plant resources drowned during man-made lakes formation. Ecological Modelling, 222(7):1263–1271, April 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000512.

Bord:2014:BEA

[BDG^{+}14] S. Bord, P. Druilhet, P. Gasqui, D. Abrial, and G. Vourc’h. Bayesian estimation of abundance based on removal sampling under weak assumption of closed population with catchability depending on environmental conditions. Application to tick abundance. Ecological Modelling, 274 (??):72–79, February 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013005826.

Baey:2013:MIV

[BDL^{+}13] Charlotte Baey, Anne Didier, Sébastien Lemaire, Fabienne Maupas, and Paul-Henry Cournède. Modelling the interindividual variability of organogenesis in sugar beet populations using a hierarchical segmented model. Ecological Modelling, 263(??):56–63, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002172.

Baey:2014:PFC

[BDL^{+}14] Charlotte Baey, Anne Didier, Sébastien Lemaire, Fabienne Maupas, and Paul-Henry Cournède. Parametrization of five classical plant growth models applied to sugar beet and comparison of their predictive capacity on root yield and total biomass. Ecological Modelling, 290(??):11–20, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005395.

Brizhik:2011:RWI

[BDTV11] Larissa S. Brizhik, Emilio Del Giudice, Alberto Tedeschi, and Vladimir L. Voeikov. The role of water in the information exchange between the components of an ecosystem. Ecological Modelling, 222(16):2869–2877, August 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002973.

Banaru:2019:IEE

[BDV^{+}19] Daniela Bănaru, Fréderic Diaz, Philippe Verley, Rose Campbell, Jonathan Navarro, Christophe Yohia, Ricardo Oliveros-Ramos, Capucine Mellon-Duval, and Yunne-Jai Shin. Implementation of an end-to-end model of the Gulf of Lions ecosystem (NW Mediterranean Sea). I. Parameterization, calibration and evaluation. Ecological Modelling, 401(??):1–19, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301024.

Beyer:2016:SCP

[BDWP16] Hawthorne L. Beyer, Yann Dujardin, Matthew E. Watts, and Hugh P. Possingham. Solving conservation planning problems with integer linear programming. Ecological Modelling, 328(??):14–22, May 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300217.

Bouderbala:2019:SSH

[BEBA19] Ilhem Bouderbala, Nadjia El Saadi, Alassane Bah, and Pierre Auger. A simulation study on how the resource competition and anti-predator cooperation impact the motilephytoplankton groups’ formation under predation stress. Ecological Modelling, 391(??):16–28, January 10, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303429.

Beck:2010:GCE

[Bec10] M. Bruce Beck. Grand challenges for environmental modeling: US NSF White Paper published. Ecological Modelling, 221(3):563, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007200.

Bentley:2012:CDP

[Ben12] Richard P. Bentley. Consumption driven population dynamics (CDPD): Further explorations. Ecological Modelling, 246(??):131–146, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003481.

Bernal-Escobar:2015:SDS

[BEPC15] Adriana Bernal-Escobar, Esteban Payan, and Juan M. Cordovez. Sex dependent spatially explicit stochastic dispersal modeling as a framework for the study of jaguar conservation and management in South America. Ecological Modelling, 299(??):40–50, March 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006085.

Berenguer:2013:SBM

[Ber13] J. Segarra Berenguer. A simple bistable model for reforestation in semi-arid zones, or how to turn a wasteland into a forest. Ecological Modelling, 266(??):58–67, September 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003360.

Beven:2015:WWS

[Bev15] Keith Beven. What we see now: Event-persistence and the predictability of hydro-eco-geomorphological systems. Ecological Modelling, 298(??):4–15, February 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003561.

Bampoh:2019:ERI

[BEZ19] Daniel Bampoh, Julia E. Earl, and Patrick A. Zollner. Examining the relative influence of animal movement patterns and mortality models on the distribution of animal transported subsidies. Ecological Modelling, 412(??):Article 108824, November 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303321.

Borrett:2011:RET

[BF11] S. R. Borrett and M. A. Freeze. Reconnecting environs to their environment. Ecological Modelling, 222(14):2393–2403, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010005752.

Babin-Fenske:2011:ABS

[BFA11] Jennifer Babin-Fenske and Madhur Anand. Agent-based simulation of effects of stress on forest tent caterpillar (Malacosoma disstria Hübner) population dynamics. Ecological Modelling, 222(14):2561–2569, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002225.

Bohn:2014:CRT

[BFH14] Friedrich J. Bohn, Karin Frank, and Andreas Huth. Of climate and its resulting tree growth: Simulating the productivity of temperate forests. Ecological Modelling, 278(??):9–17, April 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000647.

Banitz:2011:ABB

[BFJ^{+}11] Thomas Banitz, Ingo Fetzer, Karin Johst, Lukas Y. Wick, Hauke Harms, and Karin Frank. Assessing biodegradation benefits from dispersal networks. Ecological Modelling, 222(14):2552–2560, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010003455.

Burkhard:2015:UEI

[BFJL15] Benjamin Burkhard, Brian D. Fath, Sven E. Jørgensen, and B. Larry Li. Use of ecological indicators in models. Ecological Modelling, 295(??):1–4, January 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004852.

Bennett:2011:MRW

[BFJZ^{+}11] Victoria J. Bennett, Esteban Fernández-Juricic, Patrick A. Zollner, Matthew J. Beard, Lynne Westphal, and Cherie L. LeBlanc Fisher. Modelling the responses of wildlife to human disturbance: an evaluation of alternative management scenarios for black-crowned night-herons. Ecological Modelling, 222(15):2770–2779, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002559.

Bunting:2016:USM

[BFKS16] Erin L. Bunting, Timothy Fullman, Gregory Kiker, and Jane Southworth. Utilization of the SAVANNA model to analyze future patterns of vegetation cover in Kruger National Park under changing climate. Ecological Modelling, 342(??):147–160, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304148.

Burkhard:2011:AAC

[BFM11] Benjamin Burkhard, Brian D. Fath, and Felix Müller. Adapting the adaptive cycle: Hypotheses on the development of ecosystem properties and services. Ecological Modelling, 222(16):2878–2890, August 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002961.

Bialozyt:2014:PSS

[BFNH14] Ronald Bialozyt, Sebastian Flinkerbusch, Marc Niggemann, and Eckhard W. Heymann. Predicting the seed shadows of a neotropical tree species dispersed by primates using an agent-based model with internal decision making for movements. Ecological Modelling, 278(??):74–84, April 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001400088X.

Baudrot:2018:ECT

[BFP^{+}18] Virgile Baudrot, Clémentine Fritsch, Antoine Perasso, Malay Banerjee, and Francis Raoul. Effects of contaminants and trophic cascade regulation on food chain stability: Application to cadmium soil pollution on small mammals — raptor systems. Ecological Modelling, 382(??):33–42, August 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018301455.

Beaudry:2016:IAO

[BFPR16] Frederic Beaudry, Michael C. Ferris, Anna M. Pidgeon, and Volker C. Radeloff. Identifying areas of optimal multispecies conservation value by accounting for incompatibilities between species. Ecological Modelling, 332(??):74–82, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016301041.

Benjamin-Fink:2017:RMD

[BFR17] Nicole Benjamin-Fink and Brian K. Reilly. A road map for developing and applying object-oriented Bayesian networks to “WICKED” problems. Ecological Modelling, 360(??):27–44, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307955.

Borrett:2011:ERI

[BFS11] S. R. Borrett, M. A. Freeze, and A. K. Salas. Equivalence of the realized input and output oriented indirect effects metrics in ecological network analysis. Ecological Modelling, 222(13):2142–2148, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001980.

Borrett:2014:ISI

[BFW14] Stuart R. Borrett, Brian D. Fath, and Stuart J. Whipple. Introduction to the special issue “Systems Ecology: a Network Perspective and Retrospective”. Ecological Modelling, 293(??):1–3, December 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004712.

Buse:2011:ICD

[BG11] Jörn Buse and Eva Maria Griebeler. Incorporating classified dispersal assumptions in predictive distribution models — a case study with grasshoppers and bush-crickets. Ecological Modelling, 222(13):2130–2141, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002183.

Belsare:2015:MBA

[BG15] Aniruddha V. Belsare and Matthew E. Gompper. A modelbased approach for investigation and mitigation of disease spillover risks to wildlife: Dogs, foxes and canine distemper in central India. Ecological Modelling, 296(??):102–112, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014005298.

Barraquand:2019:IMD

[BG19] Frédéric Barraquand and Olivier Gimenez. Integrating multiple data sources to fit matrix population models for interacting species. Ecological Modelling, 411(??): Article 108713, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302133.

Banha:2017:ERO

[BGA17] Filipe Banha, Mafalda Gama, and Pedro Manuel Anastácio. The effect of reproductive occurrences and human descriptors on invasive pet distribution modelling: Trachemys scripta elegans in the Iberian Peninsula. Ecological Modelling, 360(??):45–52, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300868.

Beyer:2010:SNP

[BGC10] Kathleen Beyer, Rodolphe E. Gozlan, and Gordon H. Copp. Social network properties within a fish assemblage invaded by non-native sunbleak Leucaspius delineatus. Ecological Modelling, 221(17):2118–2122, August 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002681.

Buzzelli:2015:AOS

[BGD^{+}15] Christopher Buzzelli, Patricia Gorman, Peter H. Doering, Zhiqiang Chen, and Yongshan Wan. The application of oyster and seagrass models to evaluate alternative inflow scenarios related to Everglades restoration. Ecological Modelling, 297(??):154–170, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005274.

Beaumont:2016:WSD

[BGD^{+}16] Linda J. Beaumont, Erin Graham, Daisy Englert Duursma, Peter D. Wilson, Abigail Cabrelli, John B. Baumgartner, Willow Hallgren, Manuel Esperón-Rodríguez, David A. Nipperess, Dan L. Warren, Shawn W. Laffan, and Jeremy VanDerWal. Which species distribution models are more (or less) likely to project broad-scale, climate-induced shifts in species ranges? Ecological Modelling, 342 (??):135–146, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305221.

Bennetsen:2016:SDM

[BGG16a] Elina Bennetsen, Sacha Gobeyn, and Peter L. M. Goethals. Species distribution models grounded in ecological theory for decision support in river management. Ecological Modelling, 325(??):1–12, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005918.

Bonnell:2016:SPP

[BGG^{+}16b] Tyler R. Bonnell, Ria R. Ghai, Tony L. Goldberg, Raja Sengupta, and Colin A. Chapman. Spatial patterns of persistence for environmentally transmitted parasites: Effects of regional climate and local landscape. Ecological Modelling, 338(??):78–89, October 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302629.

Bevacqua:2016:CAI

[BGGL16] Daniele Bevacqua, Isabelle Grechi, Michel Génard, and Francoise Lescourret. The consequences of aphid infestation on fruit production become evident in a multi-year perspective: Insights from a virtual experiment. Ecological Modelling, 338(??):11–16, October 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302666.

Becher:2016:BMB

[BGK^{+}16] M. A. Becher, V. Grimm, J. Knapp, J. Horn, G. Twiston-Davies, and J. L. Osborne. BEESCOUT: a model of bee scouting behaviour and a software tool for characterizing nectar/pollen landscapes for BEEHAVE. Ecological Modelling, 340(??):126–133, November 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630415X.

Banos-Gonzalez:2015:DIS

[BGMFES15] Isabel Banos-González, Julia Martínez-Fernández, and Miguel Ángel Esteve-Selma. Dynamic integration of sustainability indicators in insular socio-ecological systems. Ecological Modelling, 306(??):130–144, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003937.

Bourgeois:2012:IWS

[BGMJ^{+}12] A. Bourgeois, S. Gaba, N. Munier-Jolain, B. Borgy, P. Monestiez, and S. Soubeyrand. Inferring weed spatial distribution from multi-type data. Ecological Modelling, 226(??):92–98, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004959.

Beckage:2011:GFF

[BGP11] Brian Beckage, Louis J. Gross, and William J. Platt. Grass feedbacks on fire stabilize savannas. Ecological Modelling, 222(14):2227–2233, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000470.

Bryhn:2010:PFE

[BGPJ10] Andreas C. Bryhn, Cyrille Girel, Gérard Paolini, and Stéphan Jacquet. Predicting future effects from nutrient abatement and climate change on phosphorus concentrations in Lake Bourget, France. Ecological Modelling, 221(10):1440–1450, May 24, 2010. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001080.

Brush:2010:ISI

[BH10] Mark J. Brush and Lora A. Harris. Introduction to the special issue of Ecological Modelling: “Advances in Modeling Estuarine and Coastal Ecosystems: Approaches, Validation, and Applications”. Ecological Modelling, 221(7):965–968, April 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010000220.

Bendtsen:2013:EGW

[BH13a] Jørgen Bendtsen and Jørgen L. S. Hansen. Effects of global warming on hypoxia in the Baltic Sea–North Sea transition zone. Ecological Modelling, 264(??):17–26, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002906.

Bendtsen:2013:MLC

[BH13b] Jørgen Bendtsen and Jørgen L. S. Hansen. A model of life cycle, connectivity and population stability of benthic macro-invertebrates in the North Sea/Baltic Sea transition zone. Ecological Modelling, 267(??):54–65, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003578.

Beckmann:2017:IPE

[BH17] Aike Beckmann and Inga Hense. The impact of primary and export production on the formation of the secondary nitrite maximum: a model study. Ecological Modelling, 359(??):25–33, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302946.

Bose:2015:MSD

[BHC^{+}15] Arun K. Bose, Brian D. Harvey, K. David Coates, Suzanne Brais, and Yves Bergeron. Modelling stand development after partial harvesting in boreal mixedwoods of eastern Canada. Ecological Modelling, 300(??):123–136, March 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015000149.

Bagdon:2016:MES

[BHD16] Benjamin A. Bagdon, Ching-Hsun Huang, and Stephen Dewhurst. Managing for ecosystem services in northern Arizona ponderosa pine forests using a novel simulation-to-optimization methodology. Ecological Modelling, 324(??): 11–27, March 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380015005803.

Bajcz:2017:GSM

[BHD17] Alex W. Bajcz, David Hiebeler, and Francis A. Drummond. Grid–Set–Match, an agent-based simulation model, predicts fruit set for the lowbush blueberry (Vaccinium angustifolium) agroecosystem. Ecological Modelling, 361 (??):80–94, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017301011.

Becher:2010:BTT

[BHHM10] Matthias A. Becher, Hanno Hildenbrandt, Charlotte K. Hemelrijk, and Robin F. A. Moritz. Brood temperature, task division and colony survival in honeybees: a model. Ecological Modelling, 221(5):769–776, March 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007984.

Buonocore:2014:AEC

[BHPF14] Elvira Buonocore, Tiina Häyhä, Alessandro Paletto, and Pier Paolo Franzese. Assessing environmental costs and impacts of forestry activities: a multi-method approach to environmental accounting. Ecological Modelling, 271(??):10–20, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013000884.

Billionnet:2011:SPR

[Bil11] Alain Billionnet. Solving the probabilistic reserve selection problem. Ecological Modelling, 222(3):546–554, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005508.

Bindewald:2017:SSI

[Bin17] Eckart Bindewald. A survey suggests individual priorities are virtually unique: Implications for group dynamics, goal achievement and ecology. Ecological Modelling, 362(??):69–79, October 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016304173.

Berglund:2012:OSM

[BJJ12] Moa Berglund, Martin Nilsson Jacobi, and Per R. Jonsson. Optimal selection of marine protected areas based on connectivity and habitat quality. Ecological Modelling, 240(??):105–112, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001779.

Bate:2016:MIC

[BJK^{+}16] Andrew M. Bate, Glyn Jones, Adam Kleczkowski, Alan MacLeod, Rebecca Naylor, Jon Timmis, Julia Touza, and Piran C. L. White. Modelling the impact and control of an infectious disease in a plant nursery with infected plant material inputs. Ecological Modelling, 334(??):27–43, August 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016301399.

Bila:2011:QMM

[BJPB11] J. Bila, J. Jura, J. Pokorny, and I. Bukovsky. Qualitative modeling and monitoring of selected ecosystem functions. Ecological Modelling, 222(19):3640–3650, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004194.

Bonneau:2016:SEC

[BJR16] Mathieu Bonneau, Fred A. Johnson, and Christina M. Romagosa. Spatially explicit control of invasive species using a reaction–diffusion model. Ecological Modelling, 337(??):15–24, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301934.

Bristow:2013:MUE

[BK13] David N. Bristow and Christopher A. Kennedy. Maximizing the use of energy in cities using an open systems network approach. Ecological Modelling, 250(??):155–164, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005303.

Brolsma:2010:VCMa

[BKB10] R. J. Brolsma, D. Karssenberg, and M. F. P. Bierkens. Vegetation competition model for water and light limitation. I: Model description, one-dimensional competition and the influence of groundwater. Ecological Modelling, 221(10):1348–1363, May 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001079.

Biswas:2018:DMM

[BKM^{+}18] Jatish C. Biswas, Naveen Kalra, M. Maniruzzaman, A. K. Choudhury, M. A. H. S. Jahan, M. B. Hossain, S. Ishtiaque, M. M. Haque, and Wais Kabir. Development of mungbean model (MungGro) and its application for climate change impact analysis in Bangladesh. Ecological Modelling, 384(??):1–9, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301972.

Bobrovsky:2010:MDS

[BKMK10] Maxim Bobrovsky, Alexander Komarov, Alexey Mikhailov, and Larisa Khanina. Modelling dynamics of soil organic matter under different historical land-use management techniques in European Russia. Ecological Modelling, 221(6):953–959, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000900859X.

Berlin:2014:NMT

[BKN14] M. Berlin, G. Suresh Kumar, and Indumathi M. Nambi. Numerical modelling on transport of nitrogen from wastewater and fertilizer applied on paddy fields. Ecological Modelling, 278(??):85–99, April 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000933.

Baveco:2021:LSM

[BKN21] Johannes M. Baveco, Harold Kuipers, and Bart A. Nolet. A large-scale multi-species spatial depletion model for overwintering waterfowl. Ecological Modelling, 222 (20–22):3773–3784, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004728.

Bartelt:2010:MAE

[BKP10] Paul E. Bartelt, Robert W. Klaver, and Warren P. Porter. Modeling amphibian energetics, habitat suitability, and movements of western toads, Anaxyrus (= Bufo) boreas, across present and future landscapes. Ecological Modelling, 221(22):2675–2686, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003492.

Bassett:2017:CDM

[BKV17] Alastair Bassett, Andrew L. Krause, and Robert A. Van Gorder. Continuous dispersal in a model of predator– prey-subsidy population dynamics. Ecological Modelling, 354(??):115–122, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307724.

Baveye:2015:MGL

[BL15] Philippe C. Baveye and Magdeline Laba. Moving away from the geostatistical lamppost: Why, where, and how does the spatial heterogeneity of soils matter? Ecological Modelling, 298(??):24–38, February 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001707.

Bruzzone:2018:IHI

[BLAV18] Octavio A. Bruzzone, Guillermo A. Logarzo, María B. Aguirre, and Eduardo G. Virla. Intra-host interspecific larval parasitoid competition solved using modelling and Bayesian statistics. Ecological Modelling, 385 (??):114–123, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302461.

Beyer:2017:LDB

[BLB^{+}17] Robert Beyer, Véronique Letort, Dominik Bayer, Hans Pretzsch, and Paul-Henry Cournède. Leaf density-based modelling of phototropic crown dynamics and long-term predictive application to European beech. Ecological Modelling, 347(??):63–71, March 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308092.

BenTouhami:2013:SPP

[BLBB13] Haythem Ben Touhami, Romain Lardy, Vincent Barra, and Gianni Bellocchi. Screening parameters in the pasture simulation model using the Morris method. Ecological Modelling, 266(??):42–57, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003372.

Byron:2011:CEC

[BLCPB11] Carrie Byron, Jason Link, Barry Costa-Pierce, and David Bengtson. Calculating ecological carrying capacity of shellfish aquaculture using mass-balance modeling: Narragansett Bay, Rhode Island. Ecological Modelling, 222 (10):1743–1755, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001177.

Bellassen:2010:MFM

[BLD^{+}10a] V. Bellassen, G. Le Maire, J. F. Dhôte, P. Ciais, and N. Viovy. Modelling forest management within a global vegetation model — Part 1: Model structure and general behaviour. Ecological Modelling, 221(20):2458–2474, October 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010003480.

Bettridge:2010:TOB

[BLD10b] Caroline Bettridge, J. Lehmann, and R. I. M. Dunbar. Trade-offs between time, predation risk and life history, and their implications for biogeography: a systems modelling approach with a primate case study. Ecological Modelling, 221(5):777–790, March 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008011.

Baldwin:2018:BSS

[BLFS18] Justin W. Baldwin, Katie Leap, John T. Finn, and Jennifer R. Smetzer. Bayesian state-space models reveal unobserved off-shore nocturnal migration from Motus data. Ecological Modelling, 386(??):38–46, October 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302692.

Bailleul:2010:SFP

[BLH10] Frédéric Bailleul, Véronique Lesage, and Mike O. Hammill. Spherical first passage time: a tool to investigate arearestricted search in three-dimensional movements. Ecological Modelling, 221(13–14):1665–1673, July 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001869.

Beck:2017:PSI

[BLLJ17] Marcus W. Beck, John C. Lehrter, Lisa L. Lowe, and Brandon M. Jarvis. Parameter sensitivity and identifiability for a biogeochemical model of hypoxia in the northern Gulf of Mexico. Ecological Modelling, 363(??):17–30, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017303782.

Bellassen:2011:MFM

[BlMG^{+}11] V. Bellassen, G. le Maire, O. Guin, J. F. Dhôte, P. Ciais, and N. Viovy. Modelling forest management within a global vegetation model — Part 2: Model validation from a tree to a continental scale. Ecological Modelling, 222(1):57–75, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004527.

Braham:2019:SCU

[BLOJ^{+}19] William W. Braham, Jae Min Lee, Evan Oskierko-Jeznacki, Barry Silverman, and Nasrin Khansari. Spatial concentration of urban assets in the Philadelphia region: an emergy synthesis. Ecological Modelling, 401(??):52–68, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301218.

Blum:2015:MIP

[BLRN15] Moshe Blum, Itamar M. Lensky, Polychronis Rempoulakis, and David Nestel. Modeling insect population fluctuations with satellite land surface temperature. Ecological Modelling, 311(??):39–47, September 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002021.

Bird:2017:AFM

[BLW^{+}17] Tomas Bird, Jarod Lyon, Simon Wotherspoon, Ruth King, and Michael McCarthy. Accounting for false mortality in telemetry tag applications. Ecological Modelling, 355(??):116–125, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017300698.

Bardi:2011:MEN

[BLY11] Ugo Bardi, Alessandro Lavacchi, and Leigh Yaxley. Modelling EROEI and net energy in the exploitation of non renewable resources. Ecological Modelling, 223(1):54–58, December 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003012.

Boyles:2010:ECH

[BM10] Justin G. Boyles and Andrew E. McKechnie. Energy conservation in hibernating endotherms: Why “suboptimal” temperatures are optimal. Ecological Modelling, 221(12):1644–1647, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001638.

Barry:2011:EAD

[BM11] Ronald P. Barry and Julie McIntyre. Estimating animal densities and home range in regions with irregular boundaries and holes: a lattice-based alternative to the kernel density estimator. Ecological Modelling, 222(10):1666–1672, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000974.

Bengfort:2016:VMH

[BM16] Michael Bengfort and Horst Malchow. Vertical mixing and hysteresis in the competition of buoyant and nonbuoyant plankton prey species in a shallow lake. Ecological Modelling, 323(??):51–60, March 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005773.

Bittebiere:2012:MCB

[BMCG12] A. K. Bittebiere, C. Mony, B. Clément, and M. Garbey. Modeling competition between plants using an individual based model: Methods and effects on the growth of two species with contrasted growth forms. Ecological Modelling, 234(??):38–50, June 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003085.

Brias:2018:IAR

[BMD18] Antoine Brias, Jean-Denis Mathias, and Guillaume Deffuant. Inter-annual rainfall variability may foster lake regime shifts: an example from Lake Bourget in France. Ecological Modelling, 389(??):11–18, December 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303272.

Borrett:2014:RNE

[BME14] Stuart R. Borrett, James Moody, and Achim Edelmann. The rise of network ecology: Maps of the topic diversity and scientific collaboration. Ecological Modelling, 293 (??):111–127, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001136.

Bastianoni:2011:CEE

[BMF^{+}11a] Simone Bastianoni, Fabiana Morandi, Tommaso Flaminio, Riccardo M. Pulselli, and Elisa B. P. Tiezzi. Corrigendum to “Emergy and emergy algebra explained by means of ingenuous set theory” [Ecol. Modell. 222 (August (16)) (2011) 2903–2907]. Ecological Modelling, 222(18):3455, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004522.

Bastianoni:2011:EEA

[BMF^{+}11b] Simone Bastianoni, Fabiana Morandi, Tommaso Flaminio, Riccardo M. Pulselli, and Elisa B. P. Tiezzi. Emergy and emergy algebra explained by means of ingenuous set theory. Ecological Modelling, 222(16):2903–2907, August 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002936.

Belem:2021:CMA

[BMMC21] Mahamadou Belem, Raphaël J. Manlay, Jean-Pierre Müller, and Jean-Luc Chotte. CaTMAS: a multi-agent model for simulating the dynamics of carbon resources of West African villages. Ecological Modelling, 222(20–22):3651–3661, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001100456X.

Bedoya:2011:CBR

[BMN11] David Bedoya, Elias S. Manolakos, and Vladimir Novotny. Characterization of biological responses under different environmental conditions: a hierarchical modeling approach. Ecological Modelling, 222(3):532–545, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001000548X.

Bonneau:2019:OSA

[BMP^{+}19] Mathieu Bonneau, Julien Martin, Nathalie Peyrard, Leroy Rodgers, Christina M. Romagosa, and Fred A. Johnson. Optimal spatial allocation of control effort to manage invasives in the face of imperfect detection and misclassification. Ecological Modelling, 392(??):108–116, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304009.

Baird:2018:MMC

[BMR^{+}18] Mark E. Baird, Mathieu Mongin, Farhan Rizwi, Line K. Bay, Neal E. Cantin, Monika Soja-Woźniak, and Jennifer Skerratt. A mechanistic model of coral bleaching due to temperature-mediated light-driven reactive oxygen build-up in zooxanthellae. Ecological Modelling, 386(??):20–37, October 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302485.

Bauduin:2016:OCS

[BMSLC16] Sarah Bauduin, Eliot McIntire, Martin-Hugues St-Laurent, and Steve Cumming. Overcoming challenges of sparse telemetry data to estimate caribou movement. Ecological Modelling, 335(??):24–34, September 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301673.

Brown:2010:EAA

[BMU10] Mark T. Brown, Amaya Martínez, and Javier Uche. Energy analysis applied to the estimation of the recovery of costs for water services under the European Water Framework Directive. Ecological Modelling, 221(17):2123–2132, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002711.

Brush:2010:MRM

[BN10] Mark J. Brush and Scott W. Nixon. Modeling the role of macroalgae in a shallow sub-estuary of Narragansett Bay, RI (USA). Ecological Modelling, 221(7):1065–1079, April 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007480.

Blum:2018:PHH

[BNC^{+}18] Moshe Blum, David Nestel, Yafit Cohen, Eitan Goldshtein, David Helman, and Itamar M. Lensky. Predicting Heliothis (Helicoverpa armigera) pest population dynamics with an age-structured insect population model driven by satellite data. Ecological Modelling, 369(??):1–12, February 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017305021.

Bertram:2019:CTM

[BND19] Jason Bertram, Erica A. Newman, and Roderick C. Dewar. Comparison of two maximum entropy models highlights the metabolic structure of metacommunities as a key determinant of local community assembly. Ecological Modelling, 407(??):Article 108720, September 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302200.

Bogoni:2019:MFP

[BNGP19] Juliano André Bogoni, Ana Beatriz Navarro, Maurício Eduardo Graipel, and Nivaldo Peroni. Modeling the frugivory of a plant with inconstant productivity and solid interaction with relictual vertebrate biota. Ecological Modelling, 408(??):Article 108728, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302285.

Bauer:2013:MEH

[BNL^{+}13] Carolyn M. Bauer, Gösta Nachman, Sara M. Lewis, Lynn F. Faust, and J. Michael Reed. Modeling effects of harvest on firefly population persistence. Ecological Modelling, 256(??):43–52, May 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001051.

Baeta:2011:MEE

[BNMP11] Alexandra Baeta, Nathalie Niquil, João C. Marques, and Joana Patrício. Modelling the effects of eutrophication, mitigation measures and an extreme flood event on estuarine benthic food webs. Ecological Modelling, 222 (6):1209–1221, March 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006599.

Buckley:2010:PSP

[BNP10] F. M. Buckley, S. Nicol, and P. K. Pollett. Preface to the selected papers on modelling and control of metapopulation networks. Ecological Modelling, 221(21):2512–2514, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002759.

Bagarello:2014:DCE

[BO14] Fabio Bagarello and Francesco Oliveri. Dynamics of closed ecosystems described by operators. Ecological Modelling, 275(??):89–99, March 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005875.

Bregaglio:2016:DEN

[BOF^{+}16] Simone Bregaglio, Francesca Orlando, Emanuela Forni, Tommaso De Gregorio, Simone Falzoi, Chiara Boni, Michele Pisetta, and Roberto Confalonieri. Development and evaluation of new modelling solutions to simulate hazelnut (Corylus avellana L.) growth and development. Ecological Modelling, 329(??):86–99, June 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300655.

Boria:2014:SFR

[BOGA14] Robert A. Boria, Link E. Olson, Steven M. Goodman, and Robert P. Anderson. Spatial filtering to reduce sampling bias can improve the performance of ecological niche models. Ecological Modelling, 275(??):73–77, March 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005917.

Bradley:2012:SDV

[BOW^{+}12] Bethany A. Bradley, Aaryn D. Olsson, Ophelia Wang, Brett G. Dickson, Lori Pelech, Steven E. Sesnie, and Luke J. Zachmann. Species detection vs. habitat suitability: Are we biasing habitat suitability models with remotely sensed data? Ecological Modelling, 244(??):57–64, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002980.

Buckley:2010:AMS

[BP10] F. M. Buckley and P. K. Pollett. Analytical methods for a stochastic mainland–island metapopulation model. Ecological Modelling, 221(21):2526–2530, October 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001134.

Bartsev:2015:CMS

[BP15] Sergey I. Bartsev and Aleksei A. Pochekutov. A continual model of soil organic matter transformations based on a scale of transformation rate. Ecological Modelling, 302(??):25–28, April 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000502.

Bartsev:2016:VDS

[BP16] Sergey I. Bartsev and Aleksei A. Pochekutov. The vertical distribution of soil organic matter predicted by a simple continuous model of soil organic matter transformations. Ecological Modelling, 328(??):95–98, May 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300497.

Bartsev:2017:QDV

[BP17] Sergey I. Bartsev and Aleksei A. Pochekutov. Quantitative description of vertical organic matter distribution in real soil profiles by means a simple continuous model. Ecological Modelling, 360(??):219–222, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303307.

Bartsev:2019:EMD

[BP19] Sergey I. Bartsev and Aleksei A. Pochekutov. An elementary multistage discrete model of soil organic matter transformations with a continuous scale of stability. Ecological Modelling, 393(??):61–65, February 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830423X.

Blamey:2014:WOP

[BPB14] Laura K. Blamey, Éva E. Plagányi, and George M. Branch. Was overfishing of predatory fish responsible for a lobster-induced regime shift in the Benguela? Ecological Modelling, 273(??):140–150, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005401.

Bellot:2018:IEP

[BPB^{+}18] Benoit Bellot, Sylvain Poggi, Jacques Baudry, Yoann Bourhis, and Nicolas Parisey. Inferring ecological processes from population signatures: a simulation-based heuristic for the selection of sampling strategies. Ecological Modelling, 385(??):12–25, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830214X.

Blancher:2017:ENM

[BPC^{+}17] Eldon C. Blancher, Richard A. Park, Jonathan S. Clough, Scott P. Milroy, W. Monty Graham, Chet F. Rakocinski, J. Read Hendon, Jerry D. Wiggert, and Robert Leaf. Establishing nearshore marine secondary productivity baseline estimates for multiple habitats in coastal Mississippi and Alabama using AQUATOX 3.1 NME for use in the Deepwater Horizon natural resource damage assessment. Ecological Modelling, 359(??):49–68, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305695.

Buonocore:2019:MME

[BPD^{+}19] Elvira Buonocore, Flavio Picone, Luigia Donnarumma, Giovanni F. Russo, and Pier Paolo Franzese. Modeling matter and energy flows in marine ecosystems using emergy and eco-exergy methods to account for natural capital value. Ecological Modelling, 392(??):137–146, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001830406X.

Bueno-Pardo:2018:TWS

[BPGSS^{+}18] Juan Bueno-Pardo, Eva García-Seoane, Ana I. Sousa, João P. Coelho, Mariana Morgado, Silja Frankenbach, João Ezequiel, Nuno Vaz, Victor Quintino, Ana M. Rodrigues, Sérgio Leandro, António Luis, João Serôdio, Marina R. Cunha, António J. Calado, Ana Lillebø, Jose E. Rebelo, and Henrique Queiroga. Trophic web structure and ecosystem attributes of a temperate coastal lagoon (Ria de Aveiro, Portugal). Ecological Modelling, 378(??):13–25, June 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300991.

Buzhdygan:2012:DSW

[BPK^{+}12] Oksana Y. Buzhdygan, Bernard C. Patten, Caner Kazanci, Qianqian Ma, and Svitlana S. Rudenko. Dynamical and system-wide properties of linear flow-quantified food webs. Ecological Modelling, 245(??):176–184, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001044.

Bidegain:2016:MDD

[BPK^{+}16] G. Bidegain, E. N. Powell, J. M. Klinck, T. Ben-Horin, and E. E. Hofmann. Microparasitic disease dynamics in benthic suspension feeders: Infective dose, non-focal hosts, and particle diffusion. Ecological Modelling, 328(??):44–61, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001630031X.

Bourhis:2015:PBF

[BPM^{+}15] Yoann Bourhis, Sylvain Poggi, Youcef Mammeri, Anne-Marie Cortesero, Anne Le Ralec, and Nicolas Parisey. Perception-based foraging for competing resources: Assessing pest population dynamics at the landscape scale from heterogeneous resource distribution. Ecological Modelling, 312(??):211–221, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002355.

Bourhis:2017:FLG

[BPM^{+}17] Yoann Bourhis, Sylvain Poggi, Youcef Mammeri, Ronan Le Cointe, Anne-Marie Cortesero, and Nicolas Parisey. Foraging as the landscape grip for population dynamics — a mechanistic model applied to crop protection. Ecological Modelling, 354(??):26–36, June 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302885.

Buenau:2012:SDF

[BPN12] Kate E. Buenau, Nichole N. Price, and Roger M. Nisbet. Size dependence, facilitation, and microhabitats mediate space competition between coral and crustose coralline algae in a spatially explicit model. Ecological Modelling, 237–238(??):23–33, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001792.

Bert:2011:ABM

[BPR^{+}11] Federico E. Bert, Guillermo P. Podestá, Santiago L. Rovere, Ángel N. Menéndez, Michael North, Eric Tatara, Carlos E. Laciana, Elke Weber, and Fernando Ruiz Toranzo. An agent based model to simulate structural and land use changes in agricultural systems of the Argentine Pampas. Ecological Modelling, 222(19):3486–3499, October 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004248.

Butterworth:2015:PMA

[BPR^{+}15] D. S. Butterworth, É. E. Plagányi, W. M. L. Robinson, N. Moosa, and C. L. de Moor. Penguin modelling approach queried. Ecological Modelling, 316(??):78–80, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003452.

Bean:2013:P

[BPRT13] N. G. Bean, P. K. Pollett, J. V. Ross, and P. G. Taylor. Preface. Ecological Modelling, 249(??):1–2, January 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003833.

Battaglia:2011:MID

[BPS^{+}11] M. Battaglia, E. A. Pinkard, P. J. Sands, J. L. Bruce, and A. Quentin. Modelling the impact of defoliation and leaf damage on forest plantation function and production. Ecological Modelling, 222(17):3193–3202, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003656.

Bonilla:2019:CPT

[BPT^{+}19] Silvia H. Bonilla, Fábio Papalardo, Celso A. Tassinari, Jose B. Sacomano, and Fabio Romeu de Carvalho. Contribution of the paraconsistent tri-annotated logic to emergy accounting and decision making. Ecological Modelling, 393(??):98–106, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018304265.

Brown:2011:AGW

[BPU11] Mark T. Brown, Gaetano Protano, and Sergio Ulgiati. Assessing geobiosphere work of generating global reserves of coal, crude oil, and natural gas. Ecological Modelling, 222(3):879–887, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006113.

Boyd:2014:MMP

[BPWB14] Charlotte Boyd, André E. Punt, Henri Weimerskirch, and Sophie Bertrand. Movement models provide insights into variation in the foraging effort of central place foragers. Ecological Modelling, 286(??):13–25, August 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001598.

Boult:2018:IBM

[BQF^{+}18] Victoria L. Boult, Tristan Quaife, Vicki Fishlock, Cynthia J. Moss, Phyllis C. Lee, and Richard M. Sibly. Individualbased modelling of elephant population dynamics using remote sensing to estimate food availability. Ecological Modelling, 387(??):187–195, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303065.

Belarde:2016:NPO

[BR16] Tyler A. Belarde and Steven F. Railsback. New predictions from old theory: Emergent effects of multiple stressors in a model of piscivorous fish. Ecological Modelling, 326(??):54–62, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500318X.

Burns:2014:QDI

[BRB14] Thomas P. Burns, Kenneth A. Rose, and Antoinette L. Brenkert. Quantifying direct and indirect effects of perturbations using model ecosystems. Ecological Modelling, 293(??):69–80, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000027.

Breuer:2012:SEF

[Bre12] Lutz Breuer. Simulating ecosystem functioning of tropical mountainous cloud forests in southern Ecuador. Ecological Modelling, 239(??):1–2, July 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002578.

Budroni:2011:DVS

[BRF^{+}11] M. A. Budroni, F. Rossi, E. Farris, R. Filigheddu, and M. Rustici. Dispersal vs. stochasticity: Competition for persistence in a reaction-diffusion model with strong Allee dynamics. Ecological Modelling, 222(16):2891–2896, August 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001100295X.

Braun:2015:MMD

[BRHD15] A. Braun, H. J. Rosner, R. Hagensieker, and S. Dieball. Multi-method dynamical reconstruction of the ecological impact of copper mining on Chinese historical landscapes. Ecological Modelling, 303(??):42–54, May 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000691.

Buzhdygan:2016:EIB

[BRKP16] Oksana Y. Buzhdygan, Svitlana S. Rudenko, Caner Kazanci, and Bernard C. Patten. Effect of invasive black locust (Robinia pseudoacacia L.) on nitrogen cycle in floodplain ecosystem. Ecological Modelling, 319 (??):170–177, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003312.

Benhamou:2012:BUD

[BRL12] Simon Benhamou and Louise Riotte-Lambert. Beyond the utilization distribution: Identifying home range areas that are intensively exploited or repeatedly visited. Ecological Modelling, 227(??):112–116, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100603X.

Bert:2014:LCV

[BRM^{+}14] Federico E. Bert, Santiago L. Rovere, Charles M. Macal, Michael J. North, and Guillermo P. Podestá. Lessons from a comprehensive validation of an agent based-model: the experience of the Pampas Model of Argentinean agricultural systems. Ecological Modelling, 273(??):284–298, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005723.

Bhattacharyya:2018:ECI

[BRMW18] Joydeb Bhattacharyya, Daniel L. Roelke, Rika M. W. Muhl, and Frances G. Withrow. Exploitative competition of invaders differentially influences the diversity of neutral, lumpy and intransitive phytoplankton assemblages in spatially heterogeneous environments. Ecological Modelling, 370(??):59–66, February 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017305033.

Buzhdygan:2014:FWT

[BRPK14] O. Y. Buzhdygan, S. S. Rudenko, B. C. Patten, and S. S. Kostyshyn. Food-web topology of Ukrainian mountain grasslands: Comparative properties and relations to ecosystem parameters. Ecological Modelling, 293 (??):128–138, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002403.

Bajocco:2010:KFI

[BRR10] S. Bajocco, L. Rosati, and C. Ricotta. Knowing fire incidence through fuel phenology: a remotely sensed approach. Ecological Modelling, 221(1):59–66, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009000234.

Boyd:2018:GAI

[BRS^{+}18] Robin Boyd, Shovonlal Roy, Richard Sibly, Robert Thorpe, and Kieran Hyder. A general approach to incorporating spatial and temporal variation in individual-based models of fish populations with application to Atlantic mackerel. Ecological Modelling, 382(??):9–17, August 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301327.

Barbosa:2011:ETE

[BRV11] A. Márcia Barbosa, Raimundo Real, and J. Mario Vargas. Erratum to “Transferability of environmental favourability models in geographic space: the case of the Iberian desman (Galemys pyrenaicus) in Portugal and Spain” [Ecol. Model. 220 (2009) 747–754]. Ecological Modelling, 222(4):1067, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006538.

Boschetti:2012:AAC

[BRW^{+}12] Fabio Boschetti, Claire Richert, Iain Walker, Jennifer Price, and Leo Dutra. Assessing attitudes and cognitive styles of stakeholders in environmental projects involving computer modelling. Ecological Modelling, 247(??):98–111, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003766.

Bodin:2010:RIH

[BS10a] Örjan Bodin and Santiago Saura. Ranking individual habitat patches as connectivity providers: Integrating network analysis and patch removal experiments. Ecological Modelling, 221(19):2393–2405, September 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002863.

Borrett:2010:ERH

[BS10b] S. R. Borrett and A. K. Salas. Evidence for resource homogenization in 50 trophic ecosystem networks. Ecological Modelling, 221(13–14):1710–1716, July 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001900.

Buchanan:2010:UMM

[BS10c] Rebecca A. Buchanan and John R. Skalski. Using multistate mark-recapture methods to model adult salmonid migration in an industrialized river. Ecological Modelling, 221(4):582–589, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007960.

Barnes:2013:IMC

[BS13] Belinda Barnes and Harvinder Sidhu. The impact of marine closed areas on fishing yield under a variety of management strategies and stock depletion levels. Ecological Modelling, 269(??):113–125, November 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004055.

Barbosa:2015:CTC

[BS15] Fabiana G. Barbosa and Fabiana Schneck. Characteristics of the top-cited papers in species distribution predictive models. Ecological Modelling, 313(??):77–83, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001500263X.

Bell:2016:DMC

[BS16a] David M. Bell and Daniel R. Schlaepfer. On the dangers of model complexity without ecological justification in species distribution modeling. Ecological Modelling, 330(??):50–59, June 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300783.

Brazhnik:2016:SNS

[BS16b] Ksenia Brazhnik and H. H. Shugart. SIBBORK: a new spatially-explicit gap model for boreal forest. Ecological Modelling, 320(??):182–196, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500438X.

Brazhnik:2017:MSS

[BS17] Ksenia Brazhnik and H. H. Shugart. Model sensitivity to spatial resolution and explicit light representation for simulation of boreal forests in complex terrain. Ecological Modelling, 352(??):90–107, May 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305233.

Bae:2018:AMA

[BS18] Soonyim Bae and Dongil Seo. Analysis and modeling of algal blooms in the Nakdong River, Korea. Ecological Modelling, 372(??):53–63, March 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300401.

Boit:2019:EDE

[BS19] Alice Boit and Matthew Spencer. Equivalence and dissimilarity of ecosystem states. Ecological Modelling, 396(??):12–22, March 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300237.

Bayle-Sempere:2013:TSE

[BSASSJ^{+}13] Just T. Bayle-Sempere, Francisco Arreguín-Sánchez, Pablo Sanchez-Jerez, Luis A. Salcido-Guevara, Damián Fernandez-Jover, and Manuel J. Zetina-Rejón. Trophic structure and energy fluxes around a Mediterranean fish farm. Ecological Modelling, 248(??):135–147, January 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004395.

Bagnara:2015:BOL

[BSC^{+}15a] Maurizio Bagnara, Matteo Sottocornola, Alessandro Cescatti, Stefano Minerbi, Leonardo Montagnani, Damiano Gianelle, and Federico Magnani. Bayesian optimization of a light use efficiency model for the estimation of daily gross primary productivity in a range of Italian forest ecosystems. Ecological Modelling, 306(??):57–66, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004451.

Bhowmick:2015:CSI

[BSC^{+}15b] Amiya Ranjan Bhowmick, Bapi Saha, Joydev Chattopadhyay, Santanu Ray, and Sabyasachi Bhattacharya. Cooperation in species: Interplay of population regulation and extinction through global population dynamics database. Ecological Modelling, 312(??):150–165, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002203.

Bonnell:2010:ABM

[BSCG10] Tyler R. Bonnell, Raja R. Sengupta, Colin A. Chapman, and Tony L. Goldberg. An agent-based model of red colobus resources and disease dynamics implicates key resource sites as hot spots of disease transmission. Ecological Modelling, 221(20):2491–2500, October 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003753.

Banerjee:2017:TVK

[BSFR17] Arnab Banerjee, Ursula M. Scharler, Brian D. Fath, and Santanu Ray. Temporal variation of keystone species and their impact on system performance in a South African estuarine ecosystem. Ecological Modelling, 363 (??):207–220, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001630792X.

Brewin:2010:TCM

[BSH^{+}10] Robert J. W. Brewin, Shubha Sathyendranath, Takafumi Hirata, Samantha J. Lavender, Rosa M. Barciela, and Nick J. Hardman-Mountford. A three-component model of phytoplankton size class for the Atlantic Ocean. Ecological Modelling, 221(11):1472–1483, June 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001092.

Bentley:2017:IPI

[BSH17] Jacob W. Bentley, Natalia Serpetti, and Johanna Jacomina Heymans. Investigating the potential impacts of ocean warming on the Norwegian and Barents Seas ecosystem using a time-dynamic food-web model. Ecological Modelling, 360(??):94–107, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307104.

Boyle:2016:DCW

[BSHG16] Maria Boyle, Lisa Schwanz, Jim Hone, and Arthur Georges. Dispersal and climate warming determine range shift in model reptile populations. Ecological Modelling, 328(??):34–43, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300345.

Borrett:2018:BRE

[BSMA18] Stuart R. Borrett, Laura Sheble, James Moody, and Evan C. Anway. Bibliometric review of ecological network analysis: 2010–2016. Ecological Modelling, 382(??):63–82, August 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018301431.

Bryhn:2013:MBF

[BSO13] Andreas C. Bryhn, Henrik Ragnarsson Stabo, and Jens Olsson. Modelling the biomass of functional groups of fish in an archipelago bay of the Baltic Sea. Ecological Modelling, 269(??):86–97, November 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004158.

Bakian:2012:ESE

[BSP12] Amanda V. Bakian, Kimberly A. Sullivan, and Eben H. Paxton. Elucidating spatially explicit behavioral landscapes in the willow flycatcher. Ecological Modelling, 232(??):119–132, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012000816.

Blanco:2014:FBV

[BSS^{+}14] Carolina Casagrande Blanco, Simon Scheiter, Enio Sosinski, Alessandra Fidelis, Madhur Anand, and Valério D. Pillar. Feedbacks between vegetation and disturbance processes promote long-term persistence of forest–grassland mosaics in south Brazil. Ecological Modelling, 291(??):224–232, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003615.

Bal:2018:CSD

[BSW18] Guillaume Bal, Mark D. Scheuerell, and Eric J. Ward. Characterizing the strength of density dependence in atrisk species through Bayesian model averaging. Ecological Modelling, 381(??):1–9, August 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301297.

Brigolin:2011:IMA

[BSZ^{+}11] D. Brigolin, C. Savenkoff, M. Zucchetta, F. Pranovi, P. Franzoi, P. Torricelli, and R. Pastres. An inverse model for the analysis of the Venice lagoon food web. Ecological Modelling, 222(14):2404–2413, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001979.

Bhattacharjee:2016:IMW

[BT16] Natalia V. Bhattacharjee and Ernest W. Tollner. Improving management of windrow composting systems by modeling runoff water quality dynamics using recurrent neural network. Ecological Modelling, 339(??):68–76, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303106.

Beltran:2017:ABB

[BTB17] Roxanne S. Beltran, J. Ward Testa, and Jennifer M. Burns. An agent-based bioenergetics model for predicting impacts of environmental change on a top marine predator, the Weddell seal. Ecological Modelling, 351(??):36–50, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303696.

Braccini:2017:MPT

[BTBM17] Matias Braccini, Stephen Taylor, Barry Bruce, and Rory McAuley. Modelling the population trajectory of West Australian white sharks. Ecological Modelling, 360(??): 363–377, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303447.

Baasch:2010:ETS

[BTM^{+}10] David M. Baasch, Andrew J. Tyre, Joshua J. Millspaugh, Scott E. Hygnstrom, and Kurt C. Vercauteren. An evaluation of three statistical methods used to model resource selection. Ecological Modelling, 221(4):565–574, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007194.

Boykin:2010:AGA

[BTPG10] Kenneth G. Boykin, Bruce C. Thompson, and Suzanne Propeck-Gray. Accuracy of gap analysis habitat models in predicting physical features for wildlife-habitat associations in the southwest U.S. Ecological Modelling, 221 (23):2769–2775, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004473.

Barton:2019:ITP

[BTS19] Madeleine G. Barton, John S. Terblanche, and Brent J. Sinclair. Incorporating temperature and precipitation extremes into process-based models of African lepidoptera changes the predicted distribution under climate change. Ecological Modelling, 394(??):53–65, February 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304289.

Brown:2010:UEE

[BU10] Mark T. Brown and Sergio Ulgiati. Updated evaluation of exergy and emergy driving the geobiosphere: a review and refinement of the emergy baseline. Ecological Modelling, 221(20):2501–2508, October 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003285.

Brown:2011:CES

[BU11a] Mark T. Brown and Sergio Ulgiati. Can emergy sustainability index be improved? A response to Harizaj. Ecological Modelling, 222(12):2034–2035, June 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100144X.

Brown:2011:UGE

[BU11b] Mark T. Brown and Sergio Ulgiati. Understanding the global economic crisis: a biophysical perspective. Ecological Modelling, 223(1):4–13, December 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002997.

Brown:2016:AGE

[BU16a] Mark T. Brown and Sergio Ulgiati. Assessing the global environmental sources driving the geobiosphere: a revised emergy baseline. Ecological Modelling, 339 (??):126–132, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300904.

Brown:2016:EAG

[BU16b] Mark T. Brown and Sergio Ulgiati. Emergy assessment of global renewable sources. Ecological Modelling, 339 (??):148–156, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300692.

Barton:2012:LFP

[BUB^{+}12] C. Michael Barton, Isaac I. T. Ullah, Sean M. Bergin, Helena Mitasova, and Hessam Sarjoughian. Looking for the future in the past: Long-term change in socioecological systems. Ecological Modelling, 241(??):42–53, August 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000786.

Buschke:2017:BTT

[Bus17] Falko T. Buschke. Biodiversity trajectories and the time needed to achieve no net loss through averted-loss biodiversity offsets. Ecological Modelling, 352(??):54–57, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304719.

Boschetti:2015:HMC

[BV15] Fabio Boschetti and Mathew A. Vanderklift. How the movement characteristics of large marine predators influence estimates of their abundance. Ecological Modelling, 313(??):223–236, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002847.

Brolsma:2010:VCMb

[BvBB10] R. J. Brolsma, L. P. H. van Beek, and M. F. P. Bierkens. Vegetation competition model for water and light limitation. II: Spatial dynamics of groundwater and vegetation. Ecological Modelling, 221(10):1364–1377, May 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000840.

Bagnara:2018:BCS

[BVC^{+}18] Maurizio Bagnara, Marcel Van Oijen, David Cameron, Damiano Gianelle, Federico Magnani, and Matteo Sottocornola. Bayesian calibration of simple forest models with multiplicative mathematical structure: a case study with two light use efficiency models in an alpine forest. Ecological Modelling, 371(??):90–100, March 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302636.

Babey:2017:SSD

[BVGR^{+}17] Tristan Babey, Laure Vieublé-Gonod, Alain Rapaport, Marc Pinheiro, Patricia Garnier, and Jean-Raynald de Dreuzy. Spatiotemporal simulations of 2,4-D pesticide degradation by microorganisms in 3D soil-core experiments. Ecological Modelling, 344(??):48–61, January 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306664.

Bredemeier:2015:ENC

[BvHR^{+}15] Birte Bredemeier, Christina von Haaren, Stefan Rüter, Michael Reich, and Thomas Meise. Evaluating the nature conservation value of field habitats: a model approach for targeting agri-environmental measures and projecting their effects. Ecological Modelling, 295(??):113–122, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003883.

Bandara:2018:HRM

[BVJE18] Kanchana Bandara, Øystein Varpe, Rubao Ji, and Ketil Eiane. A high-resolution modeling study on diel and seasonal vertical migrations of high-latitude copepods. Ecological Modelling, 368(??):357–376, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304969.

Baetens:2013:IBM

[BVP^{+}13] J. M. Baetens, S. Van Nieuland, I. S. Pauwels, B. De Baets, A. M. Mouton, and P. L. M. Goethals. An individual-based model for the migration of pike (Esox lucius) in the river yser, Belgium. Ecological Modelling, 258(??):40–52, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001336.

Bicout:2015:MDM

[BVVS15] Dominique J. Bicout, Mathias Vautrin, Cecile Vignolles, and Philippe Sabatier. Modeling the dynamics of mosquito breeding sites vs. rainfall in Barkedji area, Senegal. Ecological Modelling, 317(??):41–49, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004147.

Brolly:2012:MMM

[BW12] Matthew Brolly and Iain H. Woodhouse. A “Matchstick Model” of microwave backscatter from a forest. Ecological Modelling, 237–238(??):74–87, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001809.

Boulange:2017:MCM

[BWA17] Julien Boulange, Hirozumi Watanabe, and Shinpei Akai. A Markov chain Monte Carlo technique for parameter estimation and inference in pesticide fate and transport modeling. Ecological Modelling, 360(??):270–278, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017300315.

Bejarano:2017:DBM

[BWC17] Adriana C. Bejarano, Randall S. Wells, and Daniel P. Costa. Development of a bioenergetic model for estimating energy requirements and prey biomass consumption of the bottlenose dolphin Tursiops truncatus. Ecological Modelling, 356(??):162–172, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300595.

Bahlai:2013:MMT

[BWH13] Christine A. Bahlai, Ross M. Weiss, and Rebecca H. Hallett. A mechanistic model for a tritrophic interaction involving soybean aphid, its host plants, and multiple natural enemies. Ecological Modelling, 254(??):54–70, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000410.

Buchheister:2015:SBE

[BWML15] Andre Buchheister, Michael J. Wilberg, Thomas J. Miller, and Robert J. Latour. Simulating bottom-up effects on predator productivity and consequences for the rebuilding timeline of a depleted population. Ecological Modelling, 311(??):48–62, September 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001982.

Boschetti:2016:MAT

[BWP16] Fabio Boschetti, Iain Walker, and Jennifer Price. Modelling and attitudes towards the future. Ecological Modelling, 322(??):71–81, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005311.

Bacalso:2016:ERI

[BWRA16] Regina Therese M. Bacalso, Matthias Wolff, Rina Maria Rosales, and Nygiel B. Armada. Effort reallocation of illegal fishing operations: a profitable scenario for the municipal fisheries of Danajon Bank, Central Philippines.

Ecological Modelling, 331(??):5–16, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300035.

Bani-Yaghoub:2018:MQL

[BYR18] Majid Bani-Yaghoub and Aaron Reed. A methodology to quantify the long-term changes in social networks of competing species. Ecological Modelling, 368(??):147–157, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017304386.

Chen:2016:EIC

[CAC^{+}16] Bin Chen, M. Altaf Arain, Jing M. Chen, Holly Croft, Robert F. Grant, Werner A. Kurz, Pierre Bernier, Luc Guindon, David Price, and Ziyu Wang. Evaluating the impacts of climate variability and cutting and insect defoliation on the historical carbon dynamics of a boreal black spruce forest landscape in eastern Canada. Ecological Modelling, 321(??):98–109, February 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005335.

Canelles:2019:SLM

[CADB19] Q. Canelles, N. Aquilué, A. Duane, and L. Brotons. From stand to landscape: modelling post-fire regeneration and species growth. Ecological Modelling, 404(??):103–111, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301668.

Cabral:2013:CSR

[CAL13] Reniel B. Cabral, Porfirio M. Aliño, and May T. Lim. A coupled stock-recruitment-age-structured model of the North Sea cod under the influence of depensation. Ecological Modelling, 253(??):1–8, March 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000148.

Campbell:2015:EAE

[Cam15] Elliott T. Campbell. Emergy analysis of emerging methods of fossil fuel production. Ecological Modelling, 315 (??):57–68, November 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005547.

Campbell:2016:EBE

[Cam16] Daniel E. Campbell. Emergy baseline for the Earth: a historical review of the science and a new calculation. Ecological Modelling, 339(??):96–125, November 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005785.

Chu-Agor:2012:SFF

[CAMCK^{+}12] M. L. Chu-Agor, R. Muñoz-Carpena, G. A. Kiker, M. E. Aiello-Lammens, H. R. Akçakaya, M. Convertino, and I. Linkov. Simulating the fate of Florida snowy plovers with sea-level rise: Exploring research and management priorities with a global uncertainty and sensitivity analysis perspective. Ecological Modelling, 224(1):33–47, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005138.

Chudzinska:2016:DBP

[CAMNN16] Magda Chudzińska, Daniel Ayllón, Jesper Madsen, and Jacob Nabe-Nielsen. Discriminating between possible foraging decisions using pattern-oriented modelling: the case of pink-footed geese in mid-Norway during their spring migration. Ecological Modelling, 320(??):299–315, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004639.

Chuvieco:2010:DFF

[CAY^{+}10] Emilio Chuvieco, Inmaculada Aguado, Marta Yebra, Héctor Nieto, Javier Salas, M. Pilar Martín, Lara Vilar, Javier Martínez, Susana Martín, Paloma Ibarra, Juan de la Riva, Jaime Baeza, Francisco Rodríguez, Juan R. Molina, Miguel A. Herrera, and Ricardo Zamora. Development of a framework for fire risk assessment using remote sensing and geographic information system technologies. Ecological Modelling, 221(1):46–58, January 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008005735.

Chen:2015:LBS

[CB15] Shi Chen and Forrest Sheng Bao. Linking body size and energetics with predation strategies: a game theoretic modeling framework. Ecological Modelling, 316(??):81–86, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003555.

Cilek:2019:BCM

[CB19] Ahmet Cilek and Suha Berberoglu. Biotope conservation in a Mediterranean agricultural land by incorporating crop modelling. Ecological Modelling, 392(??):52–66, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303909.

Confalonieri:2010:PIQ

[CBA10] R. Confalonieri, S. Bregaglio, and M. Acutis. A proposal of an indicator for quantifying model robustness based on the relationship between variability of errors and of explored conditions. Ecological Modelling, 221(6):960–964, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009008333.

Confalonieri:2012:QPS

[CBA12] R. Confalonieri, S. Bregaglio, and M. Acutis. Quantifying plasticity in simulation models. Ecological Modelling, 225(??):159–166, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005643.

Confalonieri:2016:QUC

[CBA16] Roberto Confalonieri, Simone Bregaglio, and Marco Acutis. Quantifying uncertainty in crop model predictions due to the uncertainty in the observations used for calibration. Ecological Modelling, 328(??):72–77, May 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300424.

Confalonieri:2010:CSA

[CBB^{+}10] R. Confalonieri, G. Bellocchi, S. Bregaglio, M. Donatelli, and M. Acutis. Comparison of sensitivity analysis techniques: a case study with the rice model WARM. Ecological Modelling, 221(16):1897–1906, August 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002371.

Cheng:2019:MEI

[CBB19] Yiwei Cheng, Nicholas J. Bouskill, and Eoin L. Brodie. Model exploration of interactions between algal functional diversity and productivity in chemostats to represent open ponds systems across climate gradients. Ecological Modelling, 406(??):121–132, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301723.

Cartwright:2016:CCE

[CBC^{+}16] Samantha J. Cartwright, Katharine M. Bowgen, Catherine Collop, Kieran Hyder, Jacob Nabe-Nielsen, Richard Stafford, Richard A. Stillman, Robert B. Thorpe, and Richard M. Sibly. Communicating complex ecological models to non-scientist end users. Ecological Modelling, 338(??):51–59, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001630254X.

Chakraborty:2017:PRF

[CBCB17] Biman Chakraborty, Amiya Ranjan Bhowmick, Joydev Chattopadhyay, and Sabyasachi Bhattacharya. Physiological responses of fish under environmental stress and extension of growth (curve) models. Ecological Modelling, 363(??):172–186, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303691.

Chatterjee:2015:AEF

[CBGK15] Kajal Chatterjee, Abhirup Bandyopadhyay, Amitava Ghosh, and Samarjit Kar. Assessment of environmental factors causing wetland degradation, using Fuzzy Analytic Network Process: a case study on Keoladeo National Park, India. Ecological Modelling, 316(??):1–13, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001500335X.

Cheung:2016:LSD

[CBO16] Alan Kwok Lun Cheung, Gary Brierley, and David O’Sullivan. Landscape structure and dynamics on the Qinghai–Tibetan Plateau. Ecological Modelling, 339 (??):7–22, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016302599.

Clough:2017:ENM

[CBP^{+}17] Jonathan S. Clough, Eldon C. Blancher, Richard A. Park, Scott P. Milroy, W. Monty Graham, Chet F. Rakocinski, J. Read Hendon, Jerry D. Wiggert, and Robert Leaf. Establishing nearshore marine injuries for the Deepwater Horizon natural resource damage assessment using AQUATOX. Ecological Modelling, 359(??):258–268, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305580.

Chaalali:2016:SDE

[CBR^{+}16] Aurélie Chaalali, Grégory Beaugrand, Virginie Raybaud, Géraldine Lassalle, Blanche Saint-Béat, François Le Loc’h, Laurent Bopp, Samuele Tecchio, Georges Safi, Marina Chifflet, Jérémy Lobry, and Nathalie Niquil. From species distributions to ecosystem structure and function: a methodological perspective. Ecological Modelling, 334(??):78–90, August 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016301508.

Chaparro:2010:DME

[CC10] Mauro A. E. Chaparro and Graciela A. Canziani. A discrete model for estimating the development time from egg to infecting larva of Ostertagia ostertagi parametrized using a fuzzy rule-based system. Ecological Modelling, 221 (21):2582–2589, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010003479.

Chen:2011:GGE

[CC11] G. Q. Chen and Z. M. Chen. Greenhouse gas emissions and natural resources use by the world economy: Ecological input–output modeling. Ecological Modelling, 222(14):2362–2376, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006320.

Campillo:2012:SAI

[CC12] F. Campillo and N. Champagnat. Simulation and analysis of an individual-based model for graph-structured plant dynamics. Ecological Modelling, 234(??):93–105, June 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001200124X.

Chen:2014:DAR

[CC14] Yuting Chen and Paul-Henry Cournède. Data assimilation to reduce uncertainty of crop model prediction with convolution particle filtering. Ecological Modelling, 290 (??):165–177, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000738.

Chen:2011:CEB

[CCC11] Z. M. Chen, B. Chen, and G. Q. Chen. Cosmic exergy based ecological assessment for a wetland in Beijing. Ecological Modelling, 222(2):322–329, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004941.

Ciric:2012:USA

[CCC12] C. Ciric, P. Ciffroy, and S. Charles. Use of sensitivity analysis to identify influential and non-influential parameters within an aquatic ecosystem model. Ecological Modelling, 246(??):119–130, November 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003031.

Cafarelli:2017:HNL

[CCCP17] Barbara Cafarelli, Crescenza Calculli, Daniela Cocchi, and Elettra Pignotti. Hierarchical non-linear mixedeffects models for estimating growth parameters of western Mediterranean solitary coral populations. Ecological Modelling, 346(??):1–9, February 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308225.

Chaparro:2013:PSO

[CCF13a] M. A. E. Chaparro, G. A. Canziani, and C. A. Fiel. Parasitic stage of Ostertagia ostertagi: a mathematical model for the livestock production region of Argentina. Ecological Modelling, 265(??):56–63, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002871.

Chen:2013:ERA

[CCF13b] Shaoqing Chen, Bin Chen, and Brian D. Fath. Ecological risk assessment on the system scale: a review of state-of-the-art models and future perspectives. Ecological Modelling, 250(??):25–33, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005170.

Colon:2015:BAA

[CCG15] C. Colon, D. Claessen, and M. Ghil. Bifurcation analysis of an agent-based model for predator–prey interactions. Ecological Modelling, 317(??):93–106, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004111.

Chen:2013:AOT

[CCL^{+}13] Qiuwen Chen, Duan Chen, Ruonan Li, Jinfeng Ma, and Koen Blanckaert. Adapting the operation of two cascaded reservoirs for ecological flow requirement of a de-watered river channel due to diversion-type hydropower stations. Ecological Modelling, 252(??):266–272, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001159.

Chen-Charpentier:2014:MCF

[CCL14] B. Chen-Charpentier and M. C. A. Leite. A model for coupling fire and insect outbreak in forests. Ecological Modelling, 286(??):26–36, August 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001896.

Chen:2017:WMF

[CCLX17] Bin Chen, Lifan Chen, Ming Lu, and Bing Xu. Wetland mapping by fusing fine spatial and hyperspectral resolution images. Ecological Modelling, 353(??):95–106, June 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001730011X.

Caceres:2013:CEG

[CCS13] Manuel O. Cáceres and Iris Cáceres-Saez. Calculating effective growth rate from a random Leslie model: Application to incidental mortality analysis. Ecological Modelling, 251(??):312–322, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000045.

Chaparro:2011:EPI

[CCSF11] M. A. E. Chaparro, G. A. Canziani, C. A. Saumell, and C. A. Fiel. Estimation of pasture infectivity according to weather conditions through a fuzzy parametrized model for the free-living stage of Ostertagia ostertagi. Ecological Modelling, 222(11):1820–1832, June 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001426.

Cawley:2015:CMM

[CCW^{+}15a] Gavin C. Cawley, Kevin Cowtan, Robert G. Way, Peter Jacobs, and Ari Jokimäki. Corrigendum to “On a minimal model for estimating climate sensitivity” [Ecol. Model. 297 (2015) 20–25]. Ecological Modelling, 312 (??):102–103, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002008.

Cawley:2015:MME

[CCW^{+}15b] Gavin C. Cawley, Kevin Cowtan, Robert G. Way, Peter Jacobs, and Ari Jokimäki. On a minimal model for estimating climate sensitivity. Ecological Modelling, 297(??):20–25, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004876.

Councill:2015:RIS

[CD15] Elizabeth L. Councill and David J. Die. The relative importance of subpopulation connectivity and the age distribution of mortality in exploited marine fish populations. Ecological Modelling, 312(??):247–255, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500201X.

Collins:2016:CTU

[CD16] O. C. Collins and K. J. Duffy. Consumption threshold used to investigate stability and ecological dominance in consumer-resource dynamics. Ecological Modelling, 319(??):155–162, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001581.

Canning:2017:TCD

[CD17] A. D. Canning and R. G. Death. Trophic cascade direction and flow determine network flow stability. Ecological Modelling, 355(??):18–23, July 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305622.

Crawford:2015:PMR

[CDG15] Michael Crawford, Stephen Davies, and Alan Griffith. Predicting metapopulation responses of a tidal wetland annual to environmental stochasticity and water dispersal through an individual-based model. Ecological Modelling, 316(??):217–229, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003828.

Chen:2017:BSI

[CDK^{+}17] C. C.-M. Chen, C. C. Drovandi, J. M. Keith, K. Anthony, M. J. Caley, and K. L. Mengersen. Bayesian semiindividual based model with approximate Bayesian computation for parameters calibration: Modelling Crown-of-Thorns populations on the Great Barrier Reef. Ecological Modelling, 364(??):113–123, November 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301291.

Clark:2011:IBM

[CDL^{+}11] James R. Clark, Stuart J. Daines, Timothy M. Lenton, Andrew J. Watson, and Hywel T. P. Williams. Individual-based modelling of adaptation in marine microbial populations using genetically defined physiological parameters. Ecological Modelling, 222(23–24):3823–3837, December 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004868.

Crimmins:2013:EEF

[CDM13] Shawn M. Crimmins, Solomon Z. Dobrowski, and Alison R. Mynsberge. Evaluating ensemble forecasts of plant species distributions under climate change. Ecological Modelling, 266(??):126–130, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003384.

Cavalcanti:2016:PBM

[CdMMF16] J. Rafael Cavalcanti, David da Motta-Marques, and Carlos Ruberto Fragoso. Process-based modeling of shallow lake metabolism: Spatio-temporal variability and relative importance of individual processes. Ecological Modelling, 323(??):28–40, March 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005323.

Campos:2010:EEH

[CdOR10] Paulo R. A. Campos, Viviane M. de Oliveira, and Alexandre Rosas. Epistasis and environmental heterogeneity in the speciation process. Ecological Modelling, 221 (21):2546–2554, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010003789.

Cao:2013:UPM

[CDR^{+}13] Yong Cao, R. Edward DeWalt, Jason L. Robinson, Tari Tweddale, Leon Hinz, and Massimo Pessino. Using Maxent to model the historic distributions of stonefly species in Illinois streams: the effects of regularization and threshold selections. Ecological Modelling, 259(??):30–39, June 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001695.

Colbach:2010:IKL

[CDT10] Nathalie Colbach, Henri Darmency, and Yann Tricault. Identifying key life-traits for the dynamics and gene flow in a weedy crop relative: Sensitivity analysis of the Gene Sys simulation model for weed beet (Beta vulgaris vulgaris). Ecological Modelling, 221(2):225–237, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006644.

Carlotti:2010:MEC

[CEC10] F. Carlotti, L. Eisenhauer, and A. Calbet. Modelling the effect of constant and fluctuating food supply on egg production rates of Acartia grani. Ecological Modelling, 221(3):495–502, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007212.

Cornwell:2019:MVS

[CESC19] Eva Cornwell, David C. Elzinga, Shelby Stowe, and Alex Capaldi. Modeling vaccination strategies to control whitenose syndrome in little brown bat colonies. Ecological Modelling, 407(??):Article 108724, September 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302248.

Caskenette:2018:PES

[CEWW18] Amanda Caskenette, Eva C. Enders, Doug Watkinson, and Dale Wrubleski. Partial exclusion of spawning Cyprinus carpio to improve coastal marsh habitat may come at the cost of increased carp population growth. Ecological Modelling, 385(??):58–64, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302400.

Carson:2014:SES

[CF14] Stuart Carson and Joanna Mills Flemming. Seal encounters at sea: a contemporary spatial approach using R-INLA. Ecological Modelling, 291(??):175–181, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003597.

Cenek:2017:AAB

[CF17] Martin Cenek and Maxwell Franklin. An adaptable agentbased model for guiding multi-species Pacific salmon fisheries management within a SES framework. Ecological Modelling, 360(??):132–149, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307347.

Choquet:2017:TSE

[CGAB17] Rémi Choquet, Alexandre Garnier, Edem Awuve, and Aurélien Besnard. Transient state estimation using continuous-time processes applied to opportunistic capture– recapture data. Ecological Modelling, 361(??):157–163, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001730354X.

Choquet:2013:ESD

[CGB^{+}13] R. Choquet, Y. Guédon, A. Besnard, M. Guillemain, and R. Pradel. Estimating stop over duration in the presence of trap-effects. Ecological Modelling, 250(??):111–118, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012005273.

Correa-Gonzalez:2014:PRR

[CGdCCPCPM14] Juan Carlos Correa-González, Ma. del Carmen Chávez-Parga, José Apolinar Cortés, and Ricardo Miguel Pérez-Munguía. Photosynthesis, respiration and reaeration in a stream with complex dissolved oxygen pattern and temperature dependence. Ecological Modelling, 273 (??):220–227, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013005668.

Colleter:2012:MTF

[CGEdM12] Mathieu Colléter, Didier Gascuel, Jean-Marc Ecoutin, and Luis Tito de Morais. Modelling trophic flows in ecosystems to assess the efficiency of marine protected area (MPA), a case study on the coast of Sénégal. Ecological Modelling, 232(??):1–13, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000397.

Chianese:2018:SRA

[CGG^{+}18] E. Chianese, A. Galletti, G. Giunta, T. C. Landi, L. Marcellino, R. Montella, and A. Riccio. Spatiotemporally resolved ambient particulate matter concentration by fusing observational data and ensemble chemical transport model simulations. Ecological Modelling, 385(??):173–181, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302606.

Chea:2016:TEU

[CGGL16] Ratha Chea, Chuanbo Guo, Gaël Grenouillet, and Sovan Lek. Toward an ecological understanding of a flood-pulse system lake in a tropical ecosystem: Food web structure and ecosystem health. Ecological Modelling, 323(??):1–11, March 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005463.

Carlman:2015:MMS

[CGL15] Inga Carlman, Erik Grönlund, and Anna Longueville. Models and methods as support for sustainable decisionmaking with focus on legal operationalisation. Ecological Modelling, 306(??):95–100, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004797.

Cabral:2010:EVF

[CGLA10] Reniel B. Cabral, Rollan C. Geronimo, May T. Lim, and Porfirio M. Aliño. Effect of variable fishing strategy on fisheries under changing effort and pressure: an agent-based model application. Ecological Modelling, 221 (2):362–369, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006425.

Currie:2014:ENC

[CGM^{+}14] William S. Currie, Deborah E. Goldberg, Jason Martina, Radka Wildova, Emily Farrer, and Kenneth J. Elgersma. Emergence of nutrient-cycling feedbacks related to plant size and invasion success in a wetland community– ecosystem model. Ecological Modelling, 282(??):69–82, June 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000441.

Crisci:2012:RSM

[CGP12] C. Crisci, B. Ghattas, and G. Perera. A review of supervised machine learning algorithms and their applications to ecological data. Ecological Modelling, 240(??):113–122, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012001081.

Chauvenet:2017:QBD

[CGS^{+}17] Alienor L. M. Chauvenet, Robin M. A. Gill, Graham C. Smith, Alastair I. Ward, and Giovanna Massei. Quantifying the bias in density estimated from distance sampling and camera trapping of unmarked individuals. Ecological Modelling, 350(??):79–86, April 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001730145X.

Caruso:2015:MEN

[CGV^{+}15] N. Caruso, M. Guerisoli, E. M. Luengos Vidal, D. Castillo, E. B. Casanave, and M. Lucherini. Modelling the ecological niche of an endangered population of Puma concolor: First application of the GNESFA method to an elusive carnivore. Ecological Modelling, 297(??):11–19, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005626.

Caetano:2011:OPR

[CGY11] Marco Antonio Leonel Caetano, Douglas Francisco Marcolino Gherardi, and Takashi Yoneyama. An optimized policy for the reduction of CO_{2} emission in the Brazilian legal Amazon. Ecological Modelling, 222(15):2835–2840, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011002614.

Chaianunporn:2012:CER

[CH12] Thotsapol Chaianunporn and Thomas Hovestadt. Concurrent evolution of random dispersal and habitat niche width in host-parasitoid systems. Ecological Modelling, 247(??):241–250, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001200470X.

Coates:2014:IMR

[CH14] Julia H. Coates and Kevin A. Hovel. Incorporating movement and reproductive asynchrony into a simulation model of fertilization success for a marine broadcast spawner. Ecological Modelling, 283(??):8–18, July 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001562.

Crowley:2015:MAC

[CH15] Philip H. Crowley and Kevin R. Hopper. Mechanisms for adaptive cohort splitting. Ecological Modelling, 308(??):1–13, July 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001295.

Chaianunporn:2019:DEM

[CH19] Thotsapol Chaianunporn and Thomas Hovestadt. Dispersal evolution in metacommunities of tri-trophic systems. Ecological Modelling, 395(??):28–38, March 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300262.

Charney:2012:EEO

[Cha12] Noah D. Charney. Evaluating expert opinion and spatial scale in an amphibian model. Ecological Modelling, 242(??):37–45, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002566.

Chateau:2015:IAS

[CHCC15] Pierre-Alexandre Château, Yuan-Chao Angelo Huang, Chaolun Allen Chen, and Yang-Chi Chang. Integrated assessment of sustainable marine cage culture through system dynamics modeling. Ecological Modelling, 299 (??):140–146, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014006322.

Chen:2010:DPI

[Che10] Shi Chen. A dynamic programming implemented resource competition game theoretic model. Ecological Modelling, 221(16):1847–1851, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002152.

Chen:2015:IEM

[Che15a] Bin Chen. Integrated ecological modelling for sustainable urban metabolism and management. Ecological Modelling, 318(??):1–4, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004676.

Chen:2015:IDL

[Che15b] Youhua Chen. Impacts of dispersal limitation on temporal biodiversity patterns in a neutral model. Ecological Modelling, 305(??):22–28, June 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001118.

Chen:2016:LSC

[Che16] Youhua Chen. Local speciation can be incorporated into neutral theory of biodiversity. Ecological Modelling, 325(??):67–70, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016000272.

Chen:2017:SEM

[Che17] Assaf Chen. Spatially explicit modelling of agricultural dynamics in semi-arid environments. Ecological Modelling, 363(??):31–47, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303915.

Clasen:2015:WFA

[CHG^{+}15] Christian Clasen, Marco Heurich, Laurent Glaesener, Eckhard Kennel, and Thomas Knoke. What factors affect the survival of tree saplings under browsing, and how can a loss of admixed tree species be forecast? Ecological Modelling, 305(??):1–9, June 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000952.

Chiao:2012:ESE

[ChHC12] Ling-Yun Chiao, Chih hao Hsieh, and Tai-Sheng Chiu. Exploring spatiotemporal ecological variations by the multiscale interpolation. Ecological Modelling, 246(??):26–33, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S030438001200364X.

Costanza:2012:SES

[CHK^{+}12] Jennifer K. Costanza, Jiri Hulcr, Frank H. Koch, Todd Earnhardt, Alexa J. McKerrow, Rob R. Dunn, and Jaime A. Collazo. Simulating the effects of the southern pine beetle on regional dynamics 60 years into the future. Ecological Modelling, 244(??):93–103, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003407.

Ceia-Hasse:2018:PPL

[CHNdÁP18] Ana Ceia-Hasse, Laetitia M. Navarro, Luís Borda de Água, and Henrique M. Pereira. Population persistence in landscapes fragmented by roads: Disentangling isolation, mortality, and the effect of dispersal. Ecological Modelling, 375(??):45–53, May 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307505.

Cho:2013:ERM

[Cho13] Cheol-Joo Cho. An exploration of reliable methods of estimating emergy requirements at the regional scale: Traditional emergy analysis, regional thermodynamic input– output analysis, or the conservation rule-implicit method. Ecological Modelling, 251(??):288–296, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000082.

Carus:2017:WFP

[CHPS17] Jana Carus, Maike Heuner, Maike Paul, and Boris Schröder. Which factors and processes drive the spatio-temporal dynamics of brackish marshes? — insights from development and parameterisation of a mechanistic vegetation model. Ecological Modelling, 363(??):122–136, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017300923.

Carvalho:2013:HDH

[CHR^{+}13] Diogo Carvalho, Pedro Horta, Helena Raposeira, Mário Santos, António Luís, and João Alexandre Cabral. How do hydrological and climatic conditions influence the diversity and behavioural trends of water birds in small Mediterranean reservoirs? A community-level modelling approach. Ecological Modelling, 257(??):80–87, May 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000963.

Cai:2011:IOS

[CHTC11] Y. P. Cai, G. H. Huang, Q. Tan, and B. Chen. Identification of optimal strategies for improving eco-resilience to floods in ecologically vulnerable regions of a wetland. Ecological Modelling, 222(2):360–369, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008424.

Chalak:2010:REW

[CHvdW^{+}10] Morteza Chalak, Lia Hemerik, Wopke van der Werf, Arjan Ruijs, and Ekko C. van Ierland. On the risk of extinction of a wild plant species through spillover of a biological control agent: Analysis of an ecosystem compartment model. Ecological Modelling, 221(16):1934–1943, August 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002425.

Coletti:2013:HCC

[CHVH13] Janaine Z. Coletti, Christoph Hinz, Ryan Vogwill, and Matthew R. Hipsey. Hydrological controls on carbon metabolism in wetlands. Ecological Modelling, 249(??):3–18, January 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003353.

Chen:2018:AEH

[CHZ18] Lifan Chen, Zheng Y. X. Huang, and Shurong Zhou. The Allee effect in hosts can weaken the dilution effect of host diversity on parasitoid infections. Ecological Modelling, 382(??):43–50, August 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301443.

Chung:2014:IPP

[CIHL14] S. W. Chung, J. Imberger, M. R. Hipsey, and H. S. Lee. The influence of physical and physiological processes on the spatial heterogeneity of a Microcystis bloom in a stratified reservoir. Ecological Modelling, 289(??):133–149, October 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003378.

Cho:2011:RCG

[CJC11] Woon-Seok Cho, Sven Erik Jørgensen, and Tae-Soo Chon. The role of conjugation in the gene–individual–population relationships in increasing eco-exergy. Ecological Modelling, 222(3):407–418, February 10, 2011. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005284.

Campillo:2011:SMC

[CJLV11] F. Campillo, M. Joannides, and I. Larramendy-Valverde. Stochastic modeling of the chemostat. Ecological Modelling, 222(15):2676–2689, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002572.

Cormont:2012:CPS

[CJM^{+}12] Anouk Cormont, René Jochem, Agnieszka Malinowska, Jana Verboom, Michiel F. WallisDeVries, and Paul Opdam. Can phenological shifts compensate for adverse effects of climate change on butterfly metapopulation viability? Ecological Modelling, 227(??):72–81, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005850.

Cheung:2016:SUP

[CJR^{+}16] William W. L. Cheung, Miranda C. Jones, Gabriel Reygondeau, Charles A. Stock, Vicky W. Y. Lam, and Thomas L. Frölicher. Structural uncertainty in projecting global fisheries catches under climate change. Ecological Modelling, 325(??):57–66, April 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016000053.

Crabtree:2012:MAM

[CK12] Stefani A. Crabtree and Timothy A. Kohler. Modelling across millennia: Interdisciplinary paths to ancient socioecological systems. Ecological Modelling, 241(??):2–4, August 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012001032.

Cosme:2015:EFM

[CKH15] Nuno Cosme, Marja Koski, and Michael Z. Hauschild. Exposure factors for marine eutrophication impacts assessment based on a mechanistic biological model. Ecological Modelling, 317(??):50–63, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004123.

Cord:2014:CSC

[CKMD14] Anna F. Cord, Doris Klein, Franz Mora, and Stefan Dech. Comparing the suitability of classified land cover data and remote sensing variables for modeling distribution patterns of plants. Ecological Modelling, 272(??):129–140, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004407.

Chertov:2017:PRMa

[CKS^{+}17] Oleg Chertov, Alexander Komarov, Cindy Shaw, Sergey Bykhovets, Pavel Frolov, Vladimir Shanin, Pavel Grabarnik, Irina Priputina, Elena Zubkova, and Maxim Shashkov. Romul Hum — a model of soil organic matter formation coupling with soil biota activity. II. Parameterisation of the soil food web biota activity. Ecological Modelling, 345(??):125–139, February 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306354.

Chen:2010:SEB

[CL10] Yushun Chen and Lian-Shin Lin. Structural equationbased latent growth curve modeling of watershed attributeregulated stream sensitivity to reduced acidic deposition. Ecological Modelling, 221(17):2086–2094, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002620.

Campillo:2012:EPS

[CL12a] F. Campillo and C. Lobry. Effect of population size in a predator–prey model. Ecological Modelling, 246 (??):1–10, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003432.

Cushman:2012:MTP

[CL12b] Samuel A. Cushman and Erin L. Landguth. Multi-taxa population connectivity in the Northern Rocky Mountains. Ecological Modelling, 231(??):101–112, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000798.

Chen:2015:UNN

[CL15a] Bin Chen and Yi Lu. Urban nexus: a new paradigm for urban studies. Ecological Modelling, 318(??):5–7, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004688.

Chen:2015:HTW

[CL15b] Minmin Chen and Junguo Liu. Historical trends of wetland areas in the agriculture and pasture interlaced zone: a case study of the Huangqihai Lake Basin in northern China. Ecological Modelling, 318(??):168–176, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006188.

Chen:2012:ABM

[CLA^{+}12] Xiaodong Chen, Frank Lupi, Li An, Ryan Sheely, Andrés Viña, and Jianguo Liu. Agent-based modeling of the effects of social norms on enrollment in payments for ecosystem services. Ecological Modelling, 229(??):16–24, March 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011003401.

Caron-Lormier:2011:MEE

[CLBD^{+}11] Geoffrey Caron-Lormier, David A. Bohan, Richard Dye, Cathy Hawes, Roger W. Humphry, and Alan Raybould. Modelling an ecosystem: the example of agro-ecosystems. Ecological Modelling, 222(5):1163–1173, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001000637X.

Carter:2015:MTP

[CLBG15] Neil Carter, Simon Levin, Adam Barlow, and Volker Grimm. Modeling tiger population and territory dynamics using an agent-based approach. Ecological Modelling, 312(??):347–362, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002574.

Collin:2011:MTC

[CLBSG11] A. Collin, A. Lamorlette, D. Bernardin, and O. Séro-Guillaume. Modelling of tree crowns with realistic morphological features: New reconstruction methodology based on iterated function system tool. Ecological Modelling, 222(3):503–513, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005302.

Chuang:2011:AMC

[CLCC11] Chin-Wei Chuang, Chao-Yuan Lin, Chang-Hai Chien, and Wen-Chieh Chou. Application of Markov-chain model for vegetation restoration assessment at landslide areas caused by a catastrophic earthquake in Central Taiwan. Ecological Modelling, 222(3):835–845, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006125.

Chion:2017:UAI

[CLD^{+}17] Clément Chion, Dominic Lagrois, Jérôme Dupras, Samuel Turgeon, Ian H. McQuinn, Robert Michaud, Nadia Ménard, and Lael Parrott. Underwater acoustic impacts of shipping management measures: Results from a socialecological model of boat and whale movements in the St. Lawrence River Estuary (Canada). Ecological Modelling, 354(??):72–87, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305750.

Cordova-Lepe:2018:PCL

[CLDRJ18] Fernando Córdova-Lepe, Rodrigo Del Valle, and Rodrigo Ramos-Jiliberto. The process of connectivity loss during habitat fragmentation and their consequences on population dynamics. Ecological Modelling, 376(??):68–75, May 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300358.

Chang:2017:ISE

[CLH17] Yu-Tsun Chang, Ying-Chieh Lee, and Shu-Li Huang. Integrated spatial ecosystem model for simulating land use change and assessing vulnerability to flooding. Ecological Modelling, 362(??):87–100, October 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303757.

Chiu:2019:HDP

[CLHH19] Hao-Wei Chiu, Ying-Chieh Lee, Shu-Li Huang, and Ya-Cheng Hsieh. How does peri-urbanization teleconnect remote areas? An emergy approach. Ecological Modelling, 403(??):57–69, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019301012.

Campbell:2014:EEG

[CLL14] Daniel E. Campbell, Hongfang Lu, and Bin-Le Lin. Emergy evaluations of the global biogeochemical cycles of six biologically active elements and two compounds. Ecological Modelling, 271(??):32–51, January 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000409.

Courbaud:2015:AEM

[CLL^{+}15] B. Courbaud, V. Lafond, G. Lagarrigues, G. Vieilledent, T. Cordonnier, F. Jabot, and F. de Coligny. Applying ecological model evaludation: Lessons learned with the forest dynamics model samsara2. Ecological Modelling, 314(??):1–14, October 24, 2015. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002884.

Caubel:2012:GRF

[CLLB12] J. Caubel, M. Launay, C. Lannou, and N. Brisson. Generic response functions to simulate climate-based processes in models for the development of airborne fungal crop pathogens. Ecological Modelling, 242(??):92–104, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012002335.

C:2019:WDC

[CLMA19] Y. C. Daza C., M. F. Laguna, J. A. Monjeau, and G. Abramson. Waves of desertification in a competitive ecosystem. Ecological Modelling, 396(??):42–49, March 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300389.

Cremona:2014:HFD

[CLNN14] Fabien Cremona, Alo Laas, Peeter Nõges, and Tiina Nõges. High-frequency data within a modeling framework: On the benefit of assessing uncertainties of lake metabolism. Ecological Modelling, 294(??):27–35, December 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004281.

Chion:2011:ECP

[CLT^{+}11] Clément Chion, P. Lamontagne, S. Turgeon, L. Parrott, J.-A. Landry, D. J. Marceau, C. C. A. Martins, R. Michaud, N. Ménard, G. Cantin, and S. Dionne. Eliciting cognitive processes underlying patterns of human– wildlife interactions for agent-based modelling. Ecological Modelling, 222(14):2213–2226, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000950.

Chen:2019:NGC

[CLW^{+}19] Tao Chen, Min Lv, Yigui Wu, Jinzhao He, Guanghua Huang, Dapeng Wang, and Huawei Ma. A new growth curve model for giant freshwater prawn Macrobrachium rosenbergii in a prawn–plant symbiotic system. Ecological Modelling, 411(??):Article 108801, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303096.

Chen:2015:QAH

[CLWH15] Aifang Chen, Ruiyun Li, Honglin Wang, and Bin He. Quantitative assessment of human appropriation of aboveground net primary production in China. Ecological Modelling, 312(??):54–60, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002148.

Chiquet:2013:DAS

[CMA^{+}13] Ross A. Chiquet, Baoling Ma, Azmy S. Ackleh, Nabendu Pal, and Natalia Sidorovskaia. Demographic analysis of sperm whales using matrix population models. Ecological Modelling, 248(??):71–79, January 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005029.

Cisneros-Montemayor:2012:EMM

[CMCASS12] Andrés M. Cisneros-Montemayor, Villy Christensen, Francisco Arreguín-Sánchez, and U. Rashid Sumaila. Ecosystem models for management advice: an analysis of recreational and commercial fisheries policies in Baja California Sur, Mexico. Ecological Modelling, 228(??):8–16, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011006090.

Couce-Montero:2015:ESS

[CMCC15] Lorena Couce-Montero, Villy Christensen, and José J. Castro. Effects of small-scale and recreational fisheries on the gran canaria ecosystem. Ecological Modelling, 312 (??):61–76, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002185.

Coquillard:2012:OPP

[CMD12] P. Coquillard, A. Muzy, and F. Diener. Optimal phenotypic plasticity in a stochastic environment minimises the cost/ benefit ratio. Ecological Modelling, 242(??):28–36, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012002402.

Coro:2015:IDQ

[CMEP15] Gianpaolo Coro, Chiara Magliozzi, Anton Ellenbroek, and Pasquale Pagano. Improving data quality to build a robust distribution model for Architeuthis dux. Ecological Modelling, 305(??):29–39, June 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001222.

Colomer:2014:UBM

[CMGF14] M. Àngels Colomer, Albert Montori, Eder García, and Cristian Fondevilla. Using a bioinspired model to determine the extinction risk of Calotriton asper populations as a result of an increase in extreme rainfall in a scenario of climatic change. Ecological Modelling, 281(??):1–14, June 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014001033.

Castanho:2014:FMD

[CMH14] M. J. P. Castanho, R. P. Mateus, and K. D. Hein. Fuzzy model of Drosophila mediopunctata population dynamics. Ecological Modelling, 287(??):9–15, September 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002154.

Chuo:2019:EIP

[CMLY19] Mingying Chuo, Jun Ma, Defu Liu, and Zhengjian Yang. Effects of the impounding process during the flood season on algal blooms in Xiangxi Bay in the Three Gorges Reservoir, China. Ecological Modelling, 392(??):236–249, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304058.

Carrasco:2010:UHA

[CMM^{+}10] L. R. Carrasco, J. D. Mumford, A. MacLeod, T. Harwood, G. Grabenweger, A. W. Leach, J. D. Knight, and R. H. A. Baker. Unveiling human-assisted dispersal mechanisms in invasive alien insects: Integration of spatial stochastic simulation and phenology models. Ecological Modelling, 221(17):2068–2075, August 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002644.

Colomer:2011:BIC

[CMSPJ11] M. Àngels Colomer, Antoni Margalida, Delfí Sanuy, and Mario J. Pérez-Jiménez. A bio-inspired computing model as a new tool for modeling ecosystems: the avian scavengers as a case study. Ecological Modelling, 222(1):33–47, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004837.

Cardil:2019:ARS

[CMSR19] Adrián Cardil, Santiago Monedero, Carlos Alberto Silva, and Joaquín Ramirez. Adjusting the rate of spread of fire simulations in real-time. Ecological Modelling, 395(??):39–44, March 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300377.

Calama:2011:MST

[CMT^{+}11] Rafael Calama, Sven Mutke, José Tomé, Javier Gordo, Gregorio Montero, and Margarida Tomé. Modelling spatial and temporal variability in a zero-inflated variable: the case of stone pine (Pinus pinea L.) cone production. Ecological Modelling, 222(3):606–618, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004916.

Coro:2016:EAL

[CMV^{+}16] Gianpaolo Coro, Chiara Magliozzi, Edward Vanden Berghe, Nicolas Bailly, Anton Ellenbroek, and Pasquale Pagano. Estimating absence locations of marine species from data of scientific surveys in OBIS. Ecological Modelling, 323(??):61–76, March 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005761.

Cerco:2010:MMM

[CN10a] Carl F. Cerco and Mark R. Noel. Monitoring, modeling, and management impacts of bivalve filter feeders in the oligohaline and tidal fresh regions of the Chesapeake Bay system. Ecological Modelling, 221(7):1054–1064, April 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009005249.

Cropp:2010:PCZ

[CN10b] Roger Cropp and John Norbury. Parameterising competing zooplankton for survival in plankton functional type models. Ecological Modelling, 221(16):1852–1864, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002437.

Chikumbo:2011:ETR

[CN11] Oliver Chikumbo and Ian Nicholas. Efficient thinning regimes for Eucalyptus fastigata: Multi-objective stand-level optimisation using the island model genetic algorithm. Ecological Modelling, 222(10):1683–1695, May 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001116.

Claesson:2013:CEG

[CN13] Jonas Claesson and Jonas Nycander. Combined effect of global warming and increased CO_{2}-concentration on vegetation growth in water-limited conditions. Ecological Modelling, 256(??):23–30, May 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000872.

Cropp:2018:LOM

[CN18] Roger Cropp and John Norbury. Linking obligate mutualism models in an extended consumer-resource framework. Ecological Modelling, 374(??):1–13, April 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830053X.

Cropp:2019:CCC

[CN19] Roger Cropp and J. Norbury. Carrying capacity — a capricious construct. Ecological Modelling, 401(??):20–26, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301073.

Cantarello:2011:SPE

[CNH^{+}11] Elena Cantarello, Adrian C. Newton, Ross A. Hill, Natalia Tejedor-Garavito, Guadalupe Williams-Linera, Fabiola López-Barrera, Robert H. Manson, and Duncan J. Golicher. Simulating the potential for ecological restoration of dryland forests in Mexico under different disturbance regimes. Ecological Modelling, 222(5):1112–1128, March 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000685X.

Cherubin:2011:FTC

[CNI11] L. M. Cherubin, R. S. Nemeth, and N. Idrisi. Flow and transport characteristics at an Epinephelus guttatus (red hind grouper) spawning aggregation site in St. Thomas (US Virgin Islands). Ecological Modelling, 222 (17):3132–3148, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003115.

Cazelles:2013:SFD

[COBF13] K. Cazelles, W. Otten, P. C. Baveye, and R. E. Falconer. Soil fungal dynamics: Parameterisation and sensitivity analysis of modelled physiological processes, soil architecture and carbon distribution. Ecological Modelling, 248(??):165–173, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200419X.

Choi:2011:CHP

[COK^{+}11] Jong-Kuk Choi, Hyun-Joo Oh, Bon Joo Koo, Joo-Hyung Ryu, and Saro Lee. Crustacean habitat potential mapping in a tidal flat using remote sensing and GIS. Ecological Modelling, 222(8):1522–1533, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006575.

Convertino:2011:NMC

[Con11] M. Convertino. Neutral metacommunity clustering and SAR: River basin vs. 2-D landscape biodiversity patterns. Ecological Modelling, 222(11):1863–1879, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001384. See corrigen-

dum [Con12].

Convertino:2012:CNM

[Con12] M. Convertino. Corrigendum: Neutral metacommunity clustering and SAR: River basin vs. 2-D landscape biodiversity patterns [Ecol. Model. 222 (June (11)) (2011) 1863–1879]. Ecological Modelling, 226(??):139, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005667. See [Con11].

Confalonieri:2014:CSA

[Con14] R. Confalonieri. CoSMo: a simple approach for reproducing plant community dynamics using a single instance of generic crop simulators. Ecological Modelling, 286(??):1–10, August 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002002.

Cheng:2013:PMR

[COPP13] Kun Cheng, Stephen M. Ogle, William J. Parton, and Genxing Pan. Predicting methanogenesis from rice paddies using the DAYCENT ecosystem model. Ecological Modelling, 261–262(??):19–31, July 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300207X.

Cho:2010:EEP

[COY^{+}10] Jaeil Cho, Taikan Oki, Pat J.-F. Yeh, Shinjiro Kanae, and Wonsik Kim. The effect of estimated PAR uncertainties on the physiological processes of biosphere models. Ecological Modelling, 221(12):1575–1579, June 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001432.

Costa:2015:FST

[CPAG15] Hugo Costa, Nuno B. Ponte, Eduardo B. Azevedo, and Artur Gil. Fuzzy set theory for predicting the potential distribution and cost-effective monitoring of invasive species. Ecological Modelling, 316(??):122–132, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003567.

Cassagne:2011:UFP

[CPD^{+}11] Nathalie Cassagne, François Pimont, Jean-Luc Dupuy, Rodman R. Linn, Anders Mårell, Chloé Oliveri, and Eric Rigolot. Using a fire propagation model to assess the efficiency of prescribed burning in reducing the fire hazard. Ecological Modelling, 222(8):1502–1514, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000615.

Coro:2013:CSE

[CPE13] Gianpaolo Coro, Pasquale Pagano, and Anton Ellenbroek. Combining simulated expert knowledge with neural networks to produce ecological niche models for Latimeria chalumnae. Ecological Modelling, 268(??):55–63, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013003980.

Coscieme:2013:TBC

[CPJ^{+}13] Luca Coscieme, Federico M. Pulselli, Sven E. Jørgensen, Simone Bastianoni, and Nadia Marchettini. Thermodynamicsbased categorization of ecosystems in a socio-ecological context. Ecological Modelling, 258(??):1–8, June 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001348.

Calama:2013:MER

[CPMP13] Rafael Calama, Jaime Puértolas, Guillermo Madrigal, and Marta Pardos. Modeling the environmental response of leaf net photosynthesis in Pinus pinea L. natural regeneration. Ecological Modelling, 251(??):9–21, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005777.

Cerezo:2010:LLI

[CPR10] Alexis Cerezo, Susana Perelman, and Chandler S. Robbins. Landscape-level impact of tropical forest loss and fragmentation on bird occurrence in eastern Guatemala. Ecological Modelling, 221(3):512–526, February 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007467.

Collalti:2014:PBM

[CPS^{+}14] Alessio Collalti, Lucia Perugini, Monia Santini, Tommaso Chiti, Angelo Nolè, Giorgio Matteucci, and Riccardo Valentini. A process-based model to simulate growth in forests with complex structure: Evaluation and use of 3D-CMCC Forest Ecosystem Model in a deciduous forest in Central Italy. Ecological Modelling, 272(??):362–378, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004456.

Colvin:2015:FWM

[CPS15] Michael E. Colvin, Clay L. Pierce, and Timothy W. Stewart. A food web modeling analysis of a Midwestern, USA eutrophic lake dominated by non-native common carp and zebra mussels. Ecological Modelling, 312(??):26–40, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002136.

Coll:2019:PMS

[CPS^{+}19] M. Coll, M. Grazia Pennino, J. Steenbeek, J. Sole, and J. M. Bellido. Predicting marine species distributions: Complementarity of food-web and Bayesian hierarchical modelling approaches. Ecological Modelling, 405(??):86–101, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001930170X.

Chen:2016:DSS

[CpW16] Ning Chen and Xin ping Wang. Driver-system state interaction in regime shifts: a model study of desertification in drylands. Ecological Modelling, 339(??):1–6, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303027.

Cossarini:2015:CSC

[CQS15] G. Cossarini, S. Querin, and C. Solidoro. The continental shelf carbon pump in the northern Adriatic Sea (Mediterranean Sea): Influence of wintertime variability. Ecological Modelling, 314(??):118–134, October 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003300.

Creelman:2011:NDS

[CR11] Chance Creelman and Dave Risk. Network design for soil CO_{2} monitoring of the northern North American region. Ecological Modelling, 222(18):3421–3428, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003681. See cor-

rigendum [CR12b].

Carscallen:2012:SRS

[CR12a] W. Mather A. Carscallen and Tamara N. Romanuk. Structure and robustness to species loss in Arctic and Antarctic ice-shelf meta-ecosystem webs. Ecological Modelling, 245(??):208–218, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001433.

Creelman:2012:CND

[CR12b] Chance Creelman and Dave Risk. Corrigendum to “Network design for soil CO_{2} monitoring of the northern North American region” [Ecol. Model. 222 (18) (2011) 3421–3428]. Ecological Modelling, 226(??):141, February 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005618. See [CR11].

Craig:2010:IIS

[Cra10] Peter D. Craig. Imposed and inherent scales in cellular automata models of habitat. Ecological Modelling, 221 (20):2425–2434, October 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010003510.

Charouki:2011:MOC

[CRA^{+}11] Najib Charouki, Nadia Raïssi, Pierre Auger, Rachid Mchich, and Hmida Atmani. A management oriented competitive model with two time scales: the case of sardine fishery along the Atlantic coast between Cantin Cape and Blanc Cape. Ecological Modelling, 222(6):1253–1261, March 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000032.

Carmona:2013:GMP

[CRAP13] Carlos P. Carmona, Achim Röder, Francisco M. Azcárate, and Begoña Peco. Grazing management or physiography? Factors controlling vegetation recovery in Mediterranean grasslands. Ecological Modelling, 251(??):73–84, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012005716.

Cao:2016:SEF

[CRB16] Hongqing Cao, Friedrich Recknagel, and Michael Bartkow. Spatially-explicit forecasting of cyanobacteria assemblages in freshwater lakes by multi-objective hybrid evolutionary algorithms. Ecological Modelling, 342(??):97–112, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304938.

Campbell:2011:IBM

[CRBC11] Matthew D. Campbell, Kenneth Rose, Kevin Boswell, and James Cowan. Individual-based modeling of an artificial reef fish community: Effects of habitat quantity and degree of refuge. Ecological Modelling, 222(23–24):3895–3909, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004947.

Clough:2016:CTF

[CRD^{+}16] Brian J. Clough, Matthew B. Russell, Grant M. Domke, Christopher W. Woodall, and Philip J. Radtke. Comparing tree foliage biomass models fitted to a multispecies, felled-tree biomass dataset for the United States. Ecological Modelling, 333(??):79–91, August 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301065.

Cao:2013:EFR

[CRO13] Hongqing Cao, Friedrich Recknagel, and Philip T. Orr. Enhanced functionality of the redesigned hybrid evolutionary algorithm HEA demonstrated by predictive modelling of algal growth in the Wivenhoe Reservoir, Queensland (Australia). Ecological Modelling, 252(??):32–43, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004747.

Contreras-Reyes:2018:BMI

[CRQW18] Javier E. Contreras-Reyes, Freddy O. López Quintero, and Rodrigo Wiff. Bayesian modeling of individual growth variability using back-calculation: Application to pink cusk-eel (Genypterus blacodes) off Chile. Ecological Modelling, 385(??):145–153, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302370.

Colombo:2015:EBA

[CRTS15] Emanuela Colombo, Matteo V. Rocco, Claudia Toro, and Enrico Sciubba. An exergy-based approach to the joint economic and environmental impact assessment of possible photovoltaic scenarios: a case study at a regional level in Italy. Ecological Modelling, 318(??):64–74, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001400564X.

Castellani:2013:MBP

[CRUF13] Marco Castellani, Rune Rosland, Agurtzane Urtizberea, and Øyvind Fiksen. A mass-balanced pelagic ecosystem model with size-structured behaviourally adaptive zooplankton and fish. Ecological Modelling, 251(??):54–63, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001200573X.

Cruz-Rivera:2013:OST

[CRVS13] Erica Cruz-Rivera, Olga Vasilieva, and Mikhail Svinin. Optimal short-term policies for protection of single biological species from local extinction. Ecological Modelling, 263(??):273–280, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002755.

Chefaoui:2017:AUP

[CS17] Rosa M. Chefaoui and Ester A. Serrão. Accounting for uncertainty in predictions of a marine species: Integrating population genetics to verify past distributions. Ecological Modelling, 359(??):229–239, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304215.

Chen:2018:OBA

[CS18] Bingzhang Chen and S. Lan Smith. Optimality-based approach for computationally efficient modeling of phytoplankton growth, chlorophyll-to-carbon, and nitrogen-to-carbon ratios. Ecological Modelling, 385(??):197–212, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302655.

Christensen:2011:CEV

[CSF11] Villy Christensen, Jeroen Steenbeek, and Pierre Failler. A combined ecosystem and value chain modeling approach for evaluating societal cost and benefit of fishing. Ecological Modelling, 222(3):857–864, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005016.

Cunha-Santino:2017:MAC

[CSFB17] Marcela B. Cunha-Santino, Ângela T. Fushita, and Irineu Bianchini. A modeling approach for a cascade of reservoirs in the Juquiá-Guaçu River (Atlantic Forest, Brazil). Ecological Modelling, 356(??):48–58, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305506.

Conlisk:2015:PIF

[CSFR15] Erin Conlisk, Alexandra D. Syphard, Janet Franklin, and Helen M. Regan. Predicting the impact of fire on a vulnerable multi-species community using a dynamic vegetation model. Ecological Modelling, 301(??):27–39, April 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015000605.

Cherubini:2011:EBF

[CSH11] Francesco Cherubini, Anders H. Strømman, and Edgar Hertwich. Effects of boreal forest management practices on the climate impact of CO_{2} emissions from bioenergy. Ecological Modelling, 223(1):59–66, December 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003693.

Cherubini:2013:BCF

[CSH13] Francesco Cherubini, Anders H. Strømman, and Edgar Hertwich. Biogenic CO_{2} fluxes from bioenergy and climate — a response. Ecological Modelling, 253(??):79–81, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000343.

Cucco:2012:MSB

[CSL^{+}12] Andrea Cucco, Matteo Sinerchia, Christel Lefrançois, Paolo Magni, Michol Ghezzo, Georg Umgiesser, Angelo Perilli, and Paolo Domenici. A metabolic scope based model of fish response to environmental changes. Ecological Modelling, 237–238(??):132–141, July 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001925.

Chan:2019:MST

[CSN^{+}19] Bunyeth Chan, Ratha Sor, Peng Bun Ngor, Christophe Baehr, and Sovan Lek. Modelling spatial and temporal dynamics of two small mud carp species in the tonle sap floodpulse ecosystem. Ecological Modelling, 392(??):82–91, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018303892.

Chambers:2017:FBG

[CSOS17] Mark S. Chambers, Leesa A. Sidhu, Ben O’Neill, and Nokuthaba Sibanda. Flexible von Bertalanffy growth models incorporating Bayesian splines. Ecological Modelling, 355(??):1–11, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016306470.

Cianfrani:2015:SMF

[CSR15] C. Cianfrani, Héctor F. Satizábal, and C. Randin. A spatial modelling framework for assessing climate change impacts on freshwater ecosystems: Response of brown trout (Salmo trutta L.) biomass to warming water temperature. Ecological Modelling, 313(??):1–12, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002720.

Coll:2016:MCS

[CSS^{+}16] Marta Coll, Jeroen Steenbeek, Jordi Sole, Isabel Palomera, and Villy Christensen. Modelling the cumulative spatial–temporal effects of environmental drivers and fishing in a NW Mediterranean marine ecosystem. Ecological Modelling, 331(??):100–114, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630093X.

Chertov:2017:PRMb

[CSS^{+}17] Oleg Chertov, Cindy Shaw, Maxim Shashkov, Alexander Komarov, Sergey Bykhovets, Vladimir Shanin, Pavel Grabarnik, Pavel Frolov, Olga Kalinina, Irina Priputina, and Elena Zubkova. Romul Hum model of soil organic matter formation coupled with soil biota activity. III. Parameterisation of earthworm activity. Ecological Modelling, 345(??):140–149, February 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302368.

Chai:2019:CFP

[CSS^{+}19] Xi Chai, Peili Shi, Minghua Song, Ning Zong, Yongtao He, Guangshai Zhao, and Xianzhou Zhang. Carbon flux phenology and net ecosystem productivity simulated by a bioclimatic index in an alpine steppe-meadow on the Tibetan Plateau. Ecological Modelling, 394(??):66–75, February 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304356.

Chen:2011:MPC

[CT11] Fei Chen and William D. Taylor. A model of phosphorus cycling in the epilimnion of oligotrophic and mesotrophic lakes. Ecological Modelling, 222(5):1103–1111, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006563.

Corre:2012:ECE

[CT12] O. Le Corre and L. Truffet. Exact computation of emergy based on a mathematical reinterpretation of the rules of emergy algebra. Ecological Modelling, 230(??):101–113, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011006065.

Campbell:2016:RBR

[CT16] Elliott T. Campbell and David R. Tilley. Relationships between renewable emergy storage or flow and biodiversity: a modeling investigation. Ecological Modelling, 340(??):134–148, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302848.

Cailly:2012:CDA

[CTB^{+}12] Priscilla Cailly, Annelise Tran, Thomas Balenghien, Grégory L’Ambert, Céline Toty, and Pauline Ezanno. A climatedriven abundance model to assess mosquito control strategies. Ecological Modelling, 227(??):7–17, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005229.

Collard:2018:SOH

[CTC^{+}18] B. Collard, P. Tixier, D. Carval, C. Lavigne, and T. Delattre. Spatial organisation of habitats in agricultural plots affects per-capita predator effect on conservation biological control: an individual based modelling study. Ecological Modelling, 388(??):124–135, November 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303235.

Clancy:2010:QPU

[CTMS10] Damian Clancy, Jason E. Tanner, Stephen McWilliam, and Matthew Spencer. Quantifying parameter uncertainty in a coral reef model using Metropolis-coupled Markov chain Monte Carlo. Ecological Modelling, 221(10):1337–1347, May 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010000724.

Cerco:2013:MPT

[CTNH13] Carl F. Cerco, Tammy Threadgill, Mark R. Noel, and Scott Hinz. Modeling the pH in the tidal fresh Potomac River under conditions of varying hydrology and loads. Ecological Modelling, 257(??):101–112, May 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000951.

Crisci:2017:MMA

[CTP^{+}17] Carolina Crisci, Rafael Terra, Juan Pablo Pacheco, Badih Ghattas, Mario Bidegain, Guillermo Goyenola, Juan José Lagomarsino, Gustavo Méndez, and Néstor Mazzeo. Multimodel approach to predict phytoplankton biomass and composition dynamics in a eutrophic shallow lake governed by extreme meteorological events. Ecological Modelling, 360(??):80–93, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304422.

Cerepnalkoski:2012:IPF

[ČTT^{+}12] Darko Čerepnalkoski, Katerina Taškova, Ljupčo Todorovski, Nataša Atanasova, and Sašo Džeroski. The influence of parameter fitting methods on model structure selection in automated modeling of aquatic ecosystems. Ecological Modelling, 245(??):136–165, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002724.

Consolo:2019:SSD

[CV19] Giancarlo Consolo and Giovanna Valenti. Secondary seed dispersal in the Klausmeier model of vegetation for sloped semi-arid environments. Ecological Modelling, 402(??):66–75, June 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001930078X.

Colleter:2015:GOA

[CVG^{+}15] Mathieu Colléter, Audrey Valls, Jérôme Guitton, Didier Gascuel, Daniel Pauly, and Villy Christensen. Global overview of the applications of the Ecopath with Ecosim modeling approach using the EcoBase models repository. Ecological Modelling, 302(??):42–53, April 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000447.

Coro:2018:FOI

[CVM^{+}18] Gianpaolo Coro, Luis Gonzalez Vilas, Chiara Magliozzi, Anton Ellenbroek, Paolo Scarponi, and Pasquale Pagano. Forecasting the ongoing invasion of Lagocephalus sceleratus in the Mediterranean Sea. Ecological Modelling, 371(??):37–49, March 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018300164.

Canal-Verges:2014:ERE

[CVPH^{+}14] Paula Canal-Vergés, Michael Potthoff, Flemming Thorbjørn Hansen, Nikolaj Holmboe, Erik Kock Rasmussen, and Mogens R. Flindt. Eelgrass re-establishment in shallow estuaries is affected by drifting macroalgae — evaluated by agentbased modeling. Ecological Modelling, 272(??):116–128, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013004377.

Canal-Verges:2016:VGT

[CVPR^{+}16] Paula Canal-Vergés, Jens K. Petersen, Erik K. Rasmussen, Anders Erichsen, and Mogens R. Flindt. Validating GIS tool to assess eelgrass potential recovery in the Limfjorden (Denmark). Ecological Modelling, 338(??):135–148, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001630151X.

Castellani:2013:SSM

[CVS^{+}13] Marco Castellani, Selina Våge, Espen Strand, T. Frede Thingstad, and Jarl Giske. The scaled subspaces method: a new trait-based approach to model communities of populations with largely inhomogeneous density. Ecological Modelling, 251(??):173–186, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005728.

Campioli:2013:CDR

[CVV^{+}13] Matteo Campioli, Hans Verbeeck, Joris Van den Bossche, Jian Wu, Andreas Ibrom, Ettore D’Andrea, Giorgio Matteucci, Roeland Samson, Kathy Steppe, and André Granier. Can decision rules simulate carbon allocation for years with contrasting and extreme weather conditions? A case study for three temperate beech forests. Ecological Modelling, 263(??):42–55, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002160.

Coops:2011:EVF

[CW11] Nicholas C. Coops and Richard H. Waring. Estimating the vulnerability of fifteen tree species under changing climate in Northwest North America. Ecological Modelling, 222(13):2119–2129, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001700.

Carroll:2017:IDM

[CW17] Rosemary W. H. Carroll and John J. Warwick. The importance of dynamic mercury water column concentrations on body burdens in a planktivorous fish: a bioenergetic and mercury mass balance perspective. Ecological Modelling, 364(??):66–76, November 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307372.

Coro:2015:CDS

[CWA^{+}15] Gianpaolo Coro, Thomas J. Webb, Ward Appeltans, Nicolas Bailly, André Cattrijsse, and Pasquale Pagano. Classifying degrees of species commonness: North Sea fish as a case study. Ecological Modelling, 312(??):272–280, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002392.

Cormont:2013:HPM

[CWJ^{+}13] Anouk Cormont, G. W. Wieger Wamelink, René Jochem, Michiel F. WallisDeVries, and Ruut M. A. Wegman. Host plant-mediated effects of climate change on the occurrence of the Alcon blue butterfly (Phengaris alcon). Ecological Modelling, 250(??):329–337, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005601.

Chapron:2016:EWC

[CWL^{+}16] Guillaume Chapron, Camilla Wikenros, Olof Liberg, Petter Wabakken, Øystein Flagstad, Cyril Milleret, Johan Månsson, Linn Svensson, Barbara Zimmermann, Mikael Åkesson, and Håkan Sand. Estimating wolf (Canis lupus) population size from number of packs and an individual based model. Ecological Modelling, 339(??):33–44, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016303179.

Cao:2013:ESM

[CWLC13] Shan Cao, Jiadao Wang, Dangguo Li, and Darong Chen. Ecological and social modeling for migration and adhesion pattern of a benthic diatom. Ecological Modelling, 250(??):269–278, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005546.

Chen:2015:DFM

[CWLL15] Kunlun Chen, Xu Wang, Dan Li, and Zhaohua Li. Driving force of the morphological change of the urban lake ecosystem: a case study of Wuhan, 1990–2013. Ecological Modelling, 318(??):204–209, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002665.

Cresswell:2012:MPC

[CWM12] Katherine A. Cresswell, John R. Wiedenmann, and Marc Mangel. A model of parental conflict: Predicting provisioning behavior of penguin partners in response to local changes in krill. Ecological Modelling, 246(??):68–78, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003249.

Convertino:2012:EUP

[CWMC^{+}12] Matteo Convertino, Paul Welle, Rafael Muñoz-Carpena, Gregory A. Kiker, Ma.L. Chu-Agor, Richard A. Fischer, and Igor Linkov. Epistemic uncertainty in predicting shorebird biogeography affected by sea-level rise. Ecological Modelling, 240(??):1–15, August 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001780.

Chen:2019:DAD

[CXL^{+}19] Dan Chen, Haoqiang Xi, Zhaohui Luo, Xinchun Cao, Jing Chen, and Chunying Zhou. Discussion of the approaches and dimensions of human transformity through an educational case. Ecological Modelling, 403(??):20–22, July 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301231.

Chi:2019:EAL

[CXW19] Yuan Chi, Zuolun Xie, and Jing Wang. Establishing archipelagic landscape ecological network with full connectivity at dual spatial scales. Ecological Modelling, 399(??):54–65, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300997.

Chen:2013:MMS

[CY13] Bin Chen and Zhifeng Yang. Modelling for multi-scale ecosystems in the context of global climate change. Ecological Modelling, 252(??):1–2, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000616.

Chen:2011:EEF

[CZ11] H. Chen and Y. W. Zhao. Evaluating the environmental flows of China’s wolonghu wetland and land use changes using a hydrological model, a water balance model, and remote sensing. Ecological Modelling, 222(2):253–260, January 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009008667.

Cao:2015:ILA

[CZC^{+}15] XiaoLei Cao, ZuHao Zhou, XiangDong Chen, WeiWei Shao, and ZiRu Wang. Improving leaf area index simulation of IBIS model and its effect on water carbon and energy — a case study in Changbai Mountain broadleaved forest of China. Ecological Modelling, 303(??):97–104, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001500068X.

Chen:2018:EEC

[CZCL18] Quan Chen, Qian Zhao, Pimao Chen, and Hongfang Lu. Effect of exotic cordgrass Spartina alterniflora on the eco-exergy based thermodynamic health of the macrobenthic faunal community in mangrove wetlands. Ecological Modelling, 385(??):106–113, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302564.

Chen:2014:AMP

[CZR^{+}14] Qiuwen Chen, Chengcheng Zhang, Friedrich Recknagel, Jing Guo, and Koen Blanckaert. Adaptation and multiple parameter optimization of the simulation model SALMO as prerequisite for scenario analysis on a shallow eutrophic lake. Ecological Modelling, 273(??):109–116, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005425.

Chen:2010:ERA

[CZW10] G. Q. Chen, L. Zeng, and Z. Wu. An ecological risk assessment model for a pulsed contaminant emission into a wetland channel flow. Ecological Modelling, 221(24): 2927–2937, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004138.

Dick:2017:FMF

[DA17] D. D. C. Dick and D. Ayllón. FloMan-MF: Floodplain management for the Moor Frog — a simulation model for amphibian conservation in dynamic wetlands. Ecological Modelling, 348(??):110–124, March 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306147.

DiLeone:2019:EKB

[DA19] A. M. Gray DiLeone and C. H. Ainsworth. Effects of Karenia brevis harmful algal blooms on fish community structure on the West Florida shelf. Ecological Modelling, 392(??):250–267, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304101.

DeJonge:2012:GSU

[DAA^{+}12] Kendall C. DeJonge, James C. Ascough, Mehdi Ahmadi, Allan A. Andales, and Mazdak Arabi. Global sensitivity and uncertainty analysis of a dynamic agroecosystem model under different irrigation treatments. Ecological Modelling, 231(??):113–125, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000531.

Brito:2017:MER

[dABLBAdOR17] Izabella de Andrade Brito, Ellie Anne López-Barrera, Sabrina Borges Lino Araújo, and Ciro Alberto de Oliveira Ribeiro. Modeling the exposure risk of the silver catfish Rhamdia quelen (Teleostei, Heptapteridae) to wastewater. Ecological Modelling, 347(??):40–49, March 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308250.

dosAnjos:2015:ISG

[dAdSC15] Lucas dos Anjos and Michel Iskin da S. Costa. Increase in specific growth rate can generate alternate transitions between stability and instability in single population dynamics. Ecological Modelling, 316(??):191–193, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003841.

deAndrade:2016:FMP

[dAEPE16] Lúcio Pereira de Andrade, Rogério Pinto Espíndola, Gilberto Carvalho Pereira, and Nelson Francisco Favilla Ebecken. Fuzzy modeling of plankton networks. Ecological Modelling, 337(??):149–155, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302381.

DiTraglia:2011:CAA

[DAF^{+}11] Mario Di Traglia, Fabio Attorre, Fabio Francesconi, Roberto Valenti, and Marcello Vitale. Is cellular automata algorithm able to predict the future dynamical shifts of tree species in Italy under climate change scenarios? A methodological approach. Ecological Modelling, 222(4):925–934, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006587.

Borda-de-Agua:2014:MIR

[dÁGP14] Luís Borda de Água, Clara Grilo, and Henrique M. Pereira. Modeling the impact of road mortality on barn owl (Tyto alba) populations using age-structured models. Ecological Modelling, 276(??):29–37, March 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000076.

Duarte:2018:FMM

[DAP18] Adam Duarte, Michael J. Adams, and James T. Peterson. Fitting N-mixture models to count data with unmodeled heterogeneity: Bias, diagnostics, and alternative approaches. Ecological Modelling, 374(??):51–59, April 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300607.

deAvila-Simas:2019:MBP

[dÁSMRT^{+}19] Sunshine de Ávila-Simas, Marcelo M. Morato, David A. Reynalte-Tataje, Hector B. Silveira, Evoy Zaniboni-Filho, and Julio E. Normey-Rico. Model-based predictive control for the regulation of the golden mussel Limnoperna fortunei (Dunker, 1857). Ecological Modelling, 406(??):84–97, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019301838.

Darwall:2010:LFL

[DATI10] William R. T. Darwall, Edward H. Allison, George F. Turner, and Kenneth Irvine. Lake of flies, or lake of fish? A trophic model of Lake Malawi. Ecological Modelling, 221(4):713–727, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007479.

DiVittorio:2010:DOA

[DAW^{+}10] Alan V. Di Vittorio, Ryan S. Anderson, Joseph D. White, Norman L. Miller, and Steven W. Running. Development and optimization of an Agro-BGC ecosystem model for C_{4} perennial grasses. Ecological Modelling, 221 (17):2038–2053, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010002656.

DeVilbiss:2015:NMC

[DB15] Christopher D. De Vilbiss and Mark T. Brown. New method to compute the emergy of crustal minerals. Ecological Modelling, 315(??):108–115, November 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001374.

Dimitrov:2014:TZEb

[DBG14a] Dimitre D. Dimitrov, Jagtar S. Bhatti, and Robert F. Grant. The transition zones (ecotone) between boreal forests and peatlands: Ecological controls on ecosystem productivity along a transition zone between upland black spruce forest and a poor forested fen in central Saskatchewan. Ecological Modelling, 291(??):96–108, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003573.

Dimitrov:2014:TZEa

[DBG14b] Dimitre D. Dimitrov, Jagtar S. Bhatti, and Robert F. Grant. The transition zones (ecotone) between boreal forests and peatlands: Modelling water table along a transition zone between upland black spruce forest and poor forested fen in central Saskatchewan. Ecological Modelling, 274(??):57–70, February 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005796.

deBruijn:2014:TMR

[dBGS^{+}14] Arjan de Bruijn, Eric J. Gustafson, Brian R. Sturtevant, Jane R. Foster, Brian R. Miranda, Nathanael I. Lichti, and Douglass F. Jacobs. Toward more robust projections of forest landscape dynamics under novel environmental conditions: Embedding PnET within LANDIS-II. Ecological Modelling, 287(??):44–57, September 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002415.

Dkhili:2017:SOS

[DBH^{+}17] Jamila Dkhili, Uta Berger, Lalla Mina Idrissi Hassani, Saïd Ghaout, Ronny Peters, and Cyril Piou. Self-organized spatial structures of locust groups emerging from local interaction. Ecological Modelling, 361(??):26–40, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017303423.

Dzotsi:2013:DUS

[DBJ13] K. A. Dzotsi, B. Basso, and J. W. Jones. Development, uncertainty and sensitivity analysis of the simple SALUS crop model in DSSAT. Ecological Modelling, 260(??):62–76, July 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013001749.

Deus:2013:TDM

[DBK^{+}13] Ricardo Deus, David Brito, Isabella Ascione Kenov, Marcelo Lima, Vanessa Costa, Adaelson Medeiros, Ramiro Neves, and C. N. Alves. Three-dimensional model for analysis of spatial and temporal patterns of phytoplankton in Tucuruí reservoir, Pará, Brazil. Ecological Modelling, 253(??):28–43, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005157.

Dudley:2016:CCI

[DBP16] Peter N. Dudley, Riccardo Bonazza, and Warren P. Porter. Climate change impacts on nesting and internesting leatherback sea turtles using 3D animated computational fluid dynamics and finite volume heat transfer. Ecological Modelling, 320(??):231–240, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004901.

DeAngelis:2012:MPN

[DBR12] Antonella De Angelis, Sofia Bajocco, and Carlo Ricotta. Modelling the phenological niche of large fires with remotely sensed NDVI profiles. Ecological Modelling, 228 (??):106–111, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001200004X.

Diop:2018:MIR

[DBS18] Bassirou Diop, Fabian Blanchard, and Nicolas Sanz. Mangrove increases resiliency of the French Guiana shrimp fishery facing global warming. Ecological Modelling, 387 (??):27–37, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302837.

DeVilbiss:2016:CGE

[DBSA16] C. De Vilbiss, M. T. Brown, E. Siegel, and S. Arden. Computing the geobiosphere emergy baseline: a novel approach. Ecological Modelling, 339(??):133–139, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016301648.

Diaz:2019:IEE

[DBVS19] Frédéric Diaz, Daniela Bănaru, Philippe Verley, and Yunne-Jai Shin. Implementation of an end-to-end model of the Gulf of Lions ecosystem (NW Mediterranean Sea). II. Investigating the effects of high trophic levels on nutrients and plankton dynamics and associated feedbacks. Ecological Modelling, 405(??):51–68, August 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301693.

Dalyander:2010:IFB

[DC10] P. Soupy Dalyander and Carl F. Cerco. Integration of a fish bioenergetics model into a spatially explicit water quality model: Application to menhaden in Chesapeake Bay. Ecological Modelling, 221(16):1922–1933, August 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002413.

Deere:2017:UDA

[DCCS17] Jacques A. Deere, Tim Coulson, Sarah Cubaynes, and Isabel M. Smallegange. Unsuccessful dispersal affects life history characteristics of natal populations: the role of dispersal related variation in vital rates. Ecological Modelling, 366(??):37–47, December 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304374.

Deckmyn:2011:SCC

[DCMK11] G. Deckmyn, M. Campioli, B. Muys, and H. Kraigher. Simulating C cycles in forest soils: Including the active role of micro-organisms in the ANAFORE forest model. Ecological Modelling, 222(12):1972–1985, June 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001347.

Das:2018:MMP

[DCMS18] Tanaya Das, Saranya Chakraborti, Joydeep Mukherjee, and Goutam Kumar Sen. Mathematical modelling for phytoplankton distribution in Sundarbans Estuarine System, India. Ecological Modelling, 368(??):111–120, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304817.

Du:2015:CTS

[DCZ^{+}15] Jianguo Du, William W. L. Cheung, Xinqing Zheng, Bin Chen, Jianji Liao, and Wenjia Hu. Comparing trophic structure of a subtropical bay as estimated from massbalance food web model and stable isotope analysis. Ecological Modelling, 312(??):175–181, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002331.

Dhondt:2012:DDS

[DDB^{+}12] Bram D’hondt, Sophie D’hondt, Dries Bonte, Rein Brys, and Maurice Hoffmann. A data-driven simulation of endozoochory by ungulates illustrates directed dispersal. Ecological Modelling, 230(??):114–122, April 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000257.

Dennhardt:2015:MAM

[DDBK15] Andrew J. Dennhardt, Adam E. Duerr, David Brandes, and Todd E. Katzner. Modeling autumn migration of a rare soaring raptor identifies new movement corridors in central appalachia. Ecological Modelling, 303(??):19–29, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015000666.

DeJager:2017:SUH

[DDM^{+}17] Nathan R. De Jager, Patrick J. Drohan, Brian M. Miranda, Brian R. Sturtevant, Susan L. Stout, Alejandro A. Royo, Eric J. Gustafson, and Mark C. Romanski. Simulating ungulate herbivory across forest landscapes: a browsing extension for LANDIS-II. Ecological Modelling, 350(??):11–29, April 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016304380.

Dyer:2014:BRL

[DDR^{+}14] A. G. Dyer, A. Dorin, V. Reinhardt, J. E. Garcia, and M. G. P. Rosa. Bee reverse-learning behavior and intra-colony differences: Simulations based on behavioral experiments reveal benefits of diversity. Ecological Modelling, 277(??):119–131, April 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400043X.

DiFebbraro:2018:UMC

[DDR^{+}18] Mirko Di Febbraro, Manuela D’Amen, Pasquale Raia, Davide De Rosa, Anna Loy, and Antoine Guisan. Using macroecological constraints on spatial biodiversity predictions under climate change: the modelling method matters. Ecological Modelling, 390(??):79–87, December 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303600.

DelBiancoFaria:2011:MIP

[DdSCG11] Lucas Del Bianco Faria, Michel Iskin da S. Costa, and Wesley A. Conde Godoy. A model for intraguild predation dynamics between immature stages. Ecological Modelling, 222(17):3295–3299, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002523.

Duthie:2013:IHA

[DF13] A. Bradley Duthie and Matthew R. Falcy. The influence of habitat autocorrelation on plants and their seed-eating pollinators. Ecological Modelling, 251(??):260–270, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000021.

Dunstan:2011:MBG

[DFD11] Piers K. Dunstan, Scott D. Foster, and Ross Darnell. Model based grouping of species across environmental gradients. Ecological Modelling, 222(4):955–963, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006393.

Duffy:2017:ISM

[DFF17] Catriona Duffy, Rowan Fealy, and Reamonn M. Fealy. An improved simulation model to describe the temperaturedependent population dynamics of the grain aphid, Sitobion avenae. Ecological Modelling, 354(??):140–171, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017302132.

Drielsma:2014:BFT

[DFH^{+}14] Michael Drielsma, Simon Ferrier, Gary Howling, Glenn Manion, Subhashni Taylor, and Jamie Love. The biodiversity forecasting toolkit: Answering the ‘how much’, ‘what’, and ‘where’ of planning for biodiversity persistence. Ecological Modelling, 274(??):80–91, February 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005760.

Doko:2011:IHP

[DFK^{+}11] Tomoko Doko, Hiromichi Fukui, Andre Kooiman, A. G. Toxopeus, Tomohiro Ichinose, Wenbo Chen, and A. K. Skidmore. Identifying habitat patches and potential ecological corridors for remnant Asiatic black bear (Ursus thibetanus japonicus) populations in Japan. Ecological Modelling, 222(3):748–761, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006101.

Dai:2016:NSF

[DFL^{+}16] Dongchen Dai, Hongwei Fang, Songheng Li, Guojian He, Lei Huang, and Wenqi Peng. Numerical simulation of fish movement behavior for habitat assessment by Eulerian-Eulerian-habitat-selection (EEHS) method. Ecological Modelling, 337(??):156–167, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302393.

Dueri:2012:MSTa

[DFM12a] Sibylle Dueri, Blaise Faugeras, and Olivier Maury. Modelling the skipjack tuna dynamics in the Indian Ocean with APECOSM-E: Part 1. Model formulation. Ecological Modelling, 245(??):41–54, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000750.

Dueri:2012:MSTb

[DFM12b] Sibylle Dueri, Blaise Faugeras, and Olivier Maury. Modelling the skipjack tuna dynamics in the Indian Ocean with APECOSM-E: Part 2: Parameter estimation and sensitivity analysis. Ecological Modelling, 245(??):55–64, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000762.

deFreitas:2018:LUC

[dFMdSA18] Marcos Wellausen Dias de Freitas, Pablo Muñoz, João Roberto dos Santos, and Diógenes Salas Alves. Land use and cover change modelling and scenarios in the Upper Uruguay Basin (Brazil). Ecological Modelling, 384(??):128–144, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302151.

Didion:2014:VTL

[DFRT14] M. Didion, B. Frey, N. Rogiers, and E. Thürig. Validating tree litter decomposition in the yasso07 carbon model. Ecological Modelling, 291(??):58–68, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003652.

Ding:2018:NOC

[DFW^{+}18] Ziyu Ding, Guohua Fang, Xin Wen, Qiaofeng Tan, Xianfeng Huang, Xiaohui Lei, Yu Tian, and Jin Quan. A novel operation chart for cascade hydropower system to alleviate ecological degradation in hydrological extremes. Ecological Modelling, 384(??):10–22, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301984.

Das:2011:REM

[DG11] Kalyan Das and Nurul Huda Gazi. Random excitations in modelling of algal blooms in estuarine systems. Ecological Modelling, 222(14):2495–2501, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006277.

Dunham:2016:EMC

[DG16] Kylee Dunham and James B. Grand. Effects of model complexity and priors on estimation using sequential importance sampling/resampling for species conservation. Ecological Modelling, 340(??):28–36, November 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303155.

Dukowska:2013:FNP

[DGKSM13] Małgorzata Dukowska, Maria Grzybkowska, Andrzej Kruk, and Eliza Szczerkowska-Majchrzak. Food niche partitioning between perch and ruffe: Combined use of a self-organising map and the IndVal index for analysing fish diet. Ecological Modelling, 265(??):221–229, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003153.

Dislich:2012:MIS

[DH12] Claudia Dislich and Andreas Huth. Modelling the impact of shallow landslides on forest structure in tropical montane forests. Ecological Modelling, 239(??):40–53, July 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001822.

Dorsey:2018:SFD

[DH18] Joseph W. Dorsey and Leon C. Hardy. Sustainability factors in dynamical systems modeling: Simulating the non-linear aspects of multiple equilibria. Ecological Modelling, 368(??):69–77, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300182.

DaSilva:2014:LIE

[DHC14] David Da Silva, Liqi Han, and Evelyne Costes. Light interception efficiency of apple trees: a multiscale computational study based on MAppleT. Ecological Modelling, 290(??):45–53, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013005784.

Deslauriers:2016:TFB

[DHC16] David Deslauriers, Laura B. Heironimus, and Steven R. Chipps. Test of a foraging-bioenergetics model to evaluate growth dynamics of endangered pallid sturgeon (Scaphirhynchus albus). Ecological Modelling, 336 (??):1–12, September 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301971.

Demsar:2015:SPA

[DHHB15] Jure Demšar, Charlotte K. Hemelrijk, Hanno Hildenbrandt, and Iztok Lebar Bajec. Simulating predator attacks on schools: Evolving composite tactics. Ecological Modelling, 304(??):22–33, May 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000745.

Downs:2012:AHR

[DHL^{+}12] Joni A. Downs, Justin H. Heller, Rebecca Loraamm, Dana Oppenheim Stein, Cassandra McDaniel, and Dave Onorato. Accuracy of home range estimators for homogeneous and inhomogeneous point patterns. Ecological Modelling, 225(??):66–73, January 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005308.

dAcampora:2018:CDM

[dHR18] Bárbara H. A. d’Acampora, Ester Higueras, and Emilia Román. Combining different metrics to measure the ecological connectivity of two mangrove landscapes in the Municipality of Florianópolis, Southern Brazil. Ecological Modelling, 384(??):103–110, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302126.

Duarte:2016:SRW

[DHS^{+}16] Adam Duarte, Jeff S. Hatfield, Todd M. Swannack, Michael R. J. Forstner, M. Clay Green, and Floyd W. Weckerly. Simulating range-wide population and breeding habitat dynamics for an endangered woodland warbler in the face of uncertainty. Ecological Modelling, 320(??):52–61, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004408.

Dillingham:2010:GTM

[Dil10] Peter W. Dillingham. Generation time and the maximum growth rate for populations with age-specific fecundities and unknown juvenile survival. Ecological Modelling, 221(6):895–899, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008382.

Dzotsi:2010:MSP

[DJA^{+}10] K. A. Dzotsi, J. W. Jones, S. G. K. Adiku, J. B. Naab, U. Singh, C. H. Porter, and A. J. Gijsman. Modeling soil and plant phosphorus within DSSAT. Ecological Modelling, 221(23):2839–2849, November 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004187.

deJalon:2018:FSM

[dJGM^{+}18] S. García de Jalón, A. Graves, G. Moreno, J. H. N. Palma, J. Crous-Durán, S. Kay, and P. J. Burgess. Forage-SAFE: a model for assessing the impact of tree cover on wood pasture profitability. Ecological Modelling, 372(??):24–32, March 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300371.

Dislich:2010:WEC

[DJH10] Claudia Dislich, Karin Johst, and Andreas Huth. What enables coexistence in plant communities? Weak versus strong species traits and the role of local processes. Ecological Modelling, 221(19):2227–2236, September 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002802.

deJager:2019:MSD

[dJHMS19] Monique de Jager, Geerten M. Hengeveld, Wolf M. Mooij, and Elisabeth Slooten. Modelling the spatial dynamics of Maui dolphins using individual-based models. Ecological Modelling, 402(??):59–65, June 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301462.

Dur:2013:IBM

[DJMBD^{+}13] Gaël Dur, Raquel Jiménez-Melero, Delphine Beyrend-Dur, Jiang-Shiou Hwang, and Sami Souissi. Individual-based model of the phenology of egg-bearing copepods: Application to Eurytemora affinis from the Seine estuary, France. Ecological Modelling, 269(??):21–36, November 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003992.

Das:2010:MES

[DJS10] Anindita Das, Dubravko Justić, and Erick Swenson. Modeling estuarine-shelf exchanges in a deltaic estuary: Implications for coastal carbon budgets and hypoxia. Ecological Modelling, 221(7):978–985, April 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009000751.

David:2019:MPD

[DJT^{+}19] Viviane David, Sandrine Joachim, Cleo Tebby, Jean-Marc Porcher, and Rémy Beaudouin. Modelling population dynamics in mesocosms using an individual-based model coupled to a bioenergetics model. Ecological Modelling, 398(??):55–66, April 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300778.

Dang:2017:IBS

[DK17] Anh Nguyet Dang and Akiyuki Kawasaki. Integrating biophysical and socio-economic factors for land-use and landcover change projection in agricultural economic regions. Ecological Modelling, 344(??):29–37, January 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306639.

Domisch:2013:CSA

[DKJH13] Sami Domisch, Mathias Kuemmerlen, Sonja C. Jähnig, and Peter Haase. Choice of study area and predictors affect habitat suitability projections, but not the performance of species distribution models of stream biota. Ecological Modelling, 257(??):1–10, May 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001063.

Dorr:2014:TMS

[DKM^{+}14] Gary J. Dorr, Daryl M. Kempthorne, Lisa C. Mayo, W. Alison Forster, Jerzy A. Zabkiewicz, Scott W. McCue, John A. Belward, Ian W. Turner, and Jim Hanan. Towards a model of spray–canopy interactions: Interception, shatter, bounce and retention of droplets on horizontal leaves. Ecological Modelling, 290(??):94–101, October 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005383.

deLeon:2019:EUB

[dLB19] María A. Ponce de León and Brian N. Bailey. Evaluating the use of Beer’s law for estimating light interception in canopy architectures with varying heterogeneity and anisotropy. Ecological Modelling, 406(??):133–143, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019301474.

Deygout:2013:HML

[DLCR13] C. Deygout, A. Lesne, F. Campillo, and A. Rapaport. Homogenised model linking microscopic and macroscopic dynamics of a biofilm: Application to growth in a plug flow reactor. Ecological Modelling, 250(??):15–24, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005224.

Duan:2011:EEI

[DLD^{+}11] N. Duan, X. D. Liu, J. Dai, C. Lin, X. H. Xia, R. Y. Gao, Y. Wang, S. Q. Chen, J. Yang, and J. Qi. Evaluating the environmental impacts of an urban wetland park based on emergy accounting and life cycle assessment: a case study in Beijing. Ecological Modelling, 222(2):351–359, January 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004230.

Deehr:2014:USI

[DLH^{+}14] Rebecca A. Deehr, Joseph J. Luczkovich, Kevin J. Hart, Lisa M. Clough, Beverly J. Johnson, and Jeffrey C. Johnson. Using stable isotope analysis to validate effective trophic levels from Ecopath models of areas closed and open to shrimp trawling in Core Sound, NC, USA. Ecological Modelling, 282(??):1–17, June 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400129X.

Dong:2018:DAF

[DLL18] Lingbo Dong, Wei Lu, and Zhaogang Liu. Developing alternative forest spatial management plans when carbon and timber values are considered: a real case from northeastern China. Ecological Modelling, 385(??):45–57, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302448.

DeCubber:2019:APE

[DLL^{+}19] Lola De Cubber, Sébastien Lefebvre, Théo Lancelot, Lionel Denis, and Sylvie Marylène Gaudron. Annelid polychaetes experience metabolic acceleration as other lophotrochozoans: Inferences on the life cycle of Arenicola marina with a Dynamic Energy Budget model. Ecological Modelling, 411(??):Article 108773, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302819.

Duque-Lazo:2016:TSD

[DLvGGNC16] J. Duque-Lazo, H. van Gils, T. A. Groen, and R. M. Navarro-Cerrillo. Transferability of species distribution models: the case of Phytophthora cinnamomi in Southwest Spain and Southwest Australia. Ecological Modelling, 320(??):62–70, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001500441X.

Drielsma:2017:BGB

[DLW^{+}17] Michael J. Drielsma, Jamie Love, Kristen J. Williams, Glenn Manion, Hanieh Saremi, Tom Harwood, and Janeen Robb. Bridging the gap between climate science and regional-scale biodiversity conservation in south-eastern Australia. Ecological Modelling, 360(??):343–362, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308171.

Dong:2017:HRE

[DLZ^{+}17] Liang Dong, Hanwei Liang, Liguo Zhang, Zhaowen Liu, Zhiqiu Gao, and Mingming Hu. Highlighting regional ecoindustrial development: Life cycle benefits of an urban industrial symbiosis and implications in China. Ecological Modelling, 361(??):164–176, October 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302818.

DiTullio:2018:SUM

[DMB^{+}18] Giacomo R. Di Tullio, Patrizio Mariani, Guido Benassai, Diana Di Luccio, and Luisa Grieco. Sustainable use of marine resources through offshore wind and mussel farm colocation. Ecological Modelling, 367(??):34–41, January 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017302338.

Desforges:2019:QFL

[DMB^{+}19] Jean-Pierre Desforges, Gonçalo M. Marques, Larissa T. Beumer, Marianna Chimienti, John Blake, Janice E. Rowell, Jan Adamczewski, Niels Martin Schmidt, and Floris M. van Beest. Quantification of the full lifecycle bioenergetics of a large mammal in the high Arctic. Ecological Modelling, 401(??):27–39, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301164.

Dupont:2011:MIB

[DMBS11] H. Dupont, J. B. Mihoub, N. Becu, and F. Sarrazin. Modelling interactions between scavenger behaviour and farming practices: Impacts on scavenger population and ecosystem service efficiency. Ecological Modelling, 222 (4):982–992, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006381.

Devi:2013:FPC

[DMDJ13] B. S. Shanthala Devi, M. S. R. Murthy, Bijan Debnath, and C. S. Jha. Forest patch connectivity diagnostics and prioritization using graph theory. Ecological Modelling, 251(??):279–287, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000057.

Dowling:2017:QEV

[DMH17] Natalie A. Dowling, Marc Mangel, and Malcolm Haddon. Quantifying the effect of vessel interference on catch rates: a theoretical approach. Ecological Modelling, 359(??):293–300, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307098.

Merino-de-Miguel:2010:MRA

[dMHGA10] Silvia Merino de Miguel, Margarita Huesca, and Federico González-Alonso. Modis reflectance and active fire data for burn mapping and assessment at regional level. Ecological Modelling, 221(1):67–74, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000900636X.

Dubois:2015:MDO

[DML15] Laetitia Dubois, Jérôme Mathieu, and Nicolas Loeuille. The manager dilemma: Optimal management of an ecosystem service in heterogeneous exploited landscapes. Ecological Modelling, 301(??):78–89, April 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000228.

deMutsert:2017:UEM

[dMLM^{+}17] Kim de Mutsert, Kristy Lewis, Scott Milroy, Joe Buszowski, and Jeroen Steenbeek. Using ecosystem modeling to evaluate trade-offs in coastal management: Effects of large-scale river diversions on fish and fisheries. Ecological Modelling, 360(??):14–26, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308535.

Diamond:2013:SEG

[DMR13] Sandra L. Diamond, Cheryl A. Murphy, and Kenneth A. Rose. Simulating the effects of global climate change on Atlantic croaker population dynamics in the mid-Atlantic region. Ecological Modelling, 264(??):98–114, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002512.

deMutsert:2016:EEH

[dMSL^{+}16] Kim de Mutsert, Jeroen Steenbeek, Kristy Lewis, Joe Buszowski, James H. Cowan, and Villy Christensen. Exploring effects of hypoxia on fish and fisheries in the northern Gulf of Mexico using a dynamic spatially explicit ecosystem model. Ecological Modelling, 331(??):142–150, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004913.

Dou:2019:MIL

[DMZ^{+}19] Ming Dou, Xiaokuan Ma, Yan Zhang, Yongyong Zhang, and Yaxin Shi. Modeling the interaction of light and nutrients as factors driving lake eutrophication. Ecological Modelling, 400(??):41–52, May 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301206.

Ditmer:2018:DEG

[DNFG18] Mark A. Ditmer, Karen V. Noyce, John R. Fieberg, and David L. Garshelis. Delineating the ecological and geographic edge of an opportunist: the American black bear exploiting an agricultural landscape. Ecological Modelling, 387(??):205–219, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302874.

deOliveira:2018:ITW

[dOAC18] Viviane M. de Oliveira, André Amado, and Paulo R. A. Campos. The interplay of tradeoffs within the framework of a resource-based modelling. Ecological Modelling, 384(??):249–260, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830231X.

Dedman:2015:MAH

[DOB^{+}15] Simon Dedman, Rick Officer, Deirdre Brophy, Maurice Clarke, and David G. Reid. Modelling abundance hotspots for data-poor Irish Sea rays. Ecological Modelling, 312 (??):77–90, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002070.

Dolgonosov:2010:RHG

[Dol10] B. M. Dolgonosov. On the reasons of hyperbolic growth in the biological and human world systems. Ecological Modelling, 221(13–14):1702–1709, July 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001882.

deOliveira:2013:IBM

[dOMZ13] A. C. S. de Oliveira, S. G. F. Martins, and M. S. Zacarias. An individual-based model for the interaction of the mite Tetranychus urticae (Koch, 1836) with its predator Neoseiulus californicus (McGregor, 1954) (Acari: Tetranychidae, phytoseiidae). Ecological Modelling, 255(??):11–20, April 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000318.

Dissanayake:2012:ISR

[DÖWB12] Sahan T. M. Dissanayake, Hayri Önal, James D. Westervelt, and Harold E. Balbach. Incorporating species relocation in reserve design models: an example from Ft. Benning, GA. Ecological Modelling, 224(1):65–75, January 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003991.

Donohue:2015:MMS

[DP15] John G. Donohue and Petri T. Piiroinen. Mathematical modelling of seasonal migration with applications to climate change. Ecological Modelling, 299(??):79–94, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014006097.

Droz:2019:TFT

[DP19] Michel Droz and Andrzej Pȩkalski. Tolerance-fecundity trade-off on a homogeneous habitat. Ecological Modelling, 411(??):Article 108796, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303047.

Delattre:2010:GFM

[DPBK10] Thomas Delattre, Jean-Baptiste Pichancourt, Francoise Burel, and Pavel Kindlmann. Grassy field margins as potential corridors for butterflies in agricultural landscapes: a simulation study. Ecological Modelling, 221(2):370–377, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006632.

Durao:2010:ASU

[DPBS10] R. M. Durão, M. J. Pereira, C. Branquinho, and A. Soares. Assessing spatial uncertainty of the Portuguese fire risk through direct sequential simulation. Ecological Modelling, 221(1):27–33, January 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006127.

Dietrich:2013:RLI

[DPLC13] Jan Philipp Dietrich, Alexander Popp, and Hermann Lotze-Campen. Reducing the loss of information and gaining accuracy with clustering methods in a global land-use model. Ecological Modelling, 263(??):233–243, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002603.

DellaCroce:2014:SES

[DPPI14] Patrick Della Croce, Geoffrey C. Poole, Robert A. Payn, and Clemente Izurieta. Simulating the effects of stream network topology on the spread of introgressive hybridization across fish populations. Ecological Modelling, 279(??):68–77, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000994.

Dercole:2010:TCF

[DPR10] Fabio Dercole, Charlotte Prieu, and Sergio Rinaldi. Technological change and fisheries sustainability: the point of view of adaptive dynamics. Ecological Modelling, 221 (3):379–387, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007091.

Douma:2016:PMA

[DPV^{+}16] J. C. Douma, M. Pautasso, R. C. Venette, C. Robinet, L. Hemerik, M. C. M. Mourits, J. Schans, and W. van der Werf. Pathway models for analysing and managing the introduction of alien plant pests’an overview and categorization. Ecological Modelling, 339(??):58–67, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001630312X.

Denis:2011:HTH

[DPvBvA11] Damien Denis, Jean-Sébastien Pierre, Joan van Baaren, and Jacques J. M. van Alphen. How temperature and habitat quality affect parasitoid lifetime reproductive success — a simulation study. Ecological Modelling, 222(9):1604–1613, May 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011001049.

Danuso:2014:SSE

[DR14] Francesco Danuso and Alvaro Rocca. SEMoLa: a simple and easy modelling language. Ecological Modelling, 285(??):54–77, August 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001938.

Dercole:2017:DEG

[DR17] Fabio Dercole and Fabio Della Rossa. A deterministic ecogenetic model for the short-term evolution of exploited fish stocks. Ecological Modelling, 343(??):80–100, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016305853.

Duraes:2016:SRC

[DRS^{+}16] Lucas Durães, Fabio O. Roque, Tadeu Siqueira, Anderson M. Santos, Magno A. Borges, and Renan S. Rezende. Simulating the role of connectivity in shaping stream insect metacommunities under colonization cycle dynamics. Ecological Modelling, 334(??):19–26, August 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630148X.

Drewes:2012:USD

[DS12] Annette D. Drewes and Janet Silbernagel. Uncovering the spatial dynamics of wild rice lakes, harvesters and management across Great Lakes landscapes for shared regional conservation. Ecological Modelling, 229(??):97–107, March 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004753.

daSilva:2018:CEN

[dS18] D. J. Pamplona da Silva. Crossing-effect in non-isolated and non-symmetric systems of patches. Ecological Modelling, 384(??):168–172, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302291.

Djebou:2021:RVG

[DS21] Dagbegnon C. Sohoulande Djebou and Vijay P. Singh. Retrieving vegetation growth patterns from soil moisture, precipitation and temperature using maximum entropy. Ecological Modelling, 309–310(??):10–21, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001593.

Dunnett:2018:MOL

[DSA^{+}18] A. Dunnett, P. B. Shirsath, P. K. Aggarwal, P. Thornton, P. K. Joshi, B. D. Pal, A. Khatri-Chhetri, and J. Ghosh. Multi-objective land use allocation modelling for prioritizing climate-smart agricultural interventions. Ecological Modelling, 381(??):23–35, August 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830125X.

Dietsch:2018:EPR

[DSBM^{+}18] Alia M. Dietsch, Kristina M. Slagle, Sharon Baruch-Mordo, Stewart W. Breck, and Lana M. Ciarniello. Education is not a panacea for reducing human–black bear conflicts. Ecological Modelling, 367(??):10–12, January 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304477.

Duncan:2015:CFR

[DSBS15] Brean W. Duncan, Paul A. Schmalzer, David R. Breininger, and Eric D. Stolen. Comparing fuels reduction and patch mosaic fire regimes for reducing fire spread potential: a spatial modeling approach. Ecological Modelling, 314(??):90–99, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003191.

Costa:2018:MHE

[dSCdA18] Michel Iskin da S. Costa and Lucas dos Anjos. Multiple hydra effect in a predator–prey model with Allee effect and mutual interference in the predator. Ecological Modelling, 373(??):22–24, April 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300528.

Costa:2019:OEO

[dSCdA19] Michel Iskin da Silveira Costa and Lucas dos Anjos. Order of events: Optimal harvest fraction in a discrete time model of a spatially structured single population protected by a marine reserve can be overestimated due to an imprecise modelling of harvest timing. Ecological Modelling, 411(??):Article 108799, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303072.

Dragicevic:2013:SND

[DSD13] Arnaud Z. Dragicevic and Bernard Sinclair-Desgagné. Sustainable network dynamics. Ecological Modelling, 270(??):43–53, December 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004328.

deSouza:2014:USD

[dSD14] Rodrigo Antônio de Souza and Paulo De Marco. The use of species distribution models to predict the spatial distribution of deforestation in the western Brazilian Amazon. Ecological Modelling, 291(??):250–259, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003342.

deSouza:2018:ISM

[dSD18] Rodrigo Antônio de Souza and Paulo De Marco. Improved spatial model for Amazonian deforestation: an empirical assessment and spatial bias analysis. Ecological Modelling, 387(??):1–9, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302849.

deSouza:2010:FSI

[dSdCF10] Lucirene Aguiar de Souza and Carlos Edwar de Carvalho Freitas. Fishing sustainability via inclusion of man in predator–prey models: a case study in Lago Preto, Manacapuru, Amazonas. Ecological Modelling, 221 (4):703–712, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009002865.

Guarino:2012:OAM

[dSGGBW12] Ernestino de Souza Gomes Guarino, Ana Márcia Barbosa, and Jorge Luiz Waechter. Occurrence and abundance models of threatened plant species: Applications to mitigate the impact of hydroelectric power dams. Ecological Modelling, 230(??):22–33, April 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200018X.

Davison:2019:SEC

[DSJ19] Raziel Davison, Marc Stadman, and Eelke Jongejans. Stochastic effects contribute to population fitness differences. Ecological Modelling, 408(??):Article 108760, September 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019302686.

dosSantos:2010:IEH

[dSJHG10] Francisca Soares dos Santos, Karin Johst, Andreas Huth, and Volker Grimm. Interacting effects of habitat destruction and changing disturbance rates on biodiversity: Who is going to survive? Ecological Modelling, 221(23):2776–2783, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000400X.

Debeljak:2011:ATS

[DSK^{+}11] Marko Debeljak, Geoffrey R. Squire, Dragi Kocev, Cathy Hawes, Mark W. Young, and Sašo Džeroski. Analysis of time series data on agroecosystem vegetation using predictive clustering trees. Ecological Modelling, 222(14):2524–2529, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010005818.

Dietrich:2012:MAL

[DSM^{+}12] Jan Philipp Dietrich, Christoph Schmitz, Christoph Müller, Marianela Fader, Hermann Lotze-Campen, and Alexander Popp. Measuring agricultural land-use intensity — a global analysis using a model-assisted approach. Ecological Modelling, 232(??):109–118, May 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001093.

Dordel:2011:RBS

[DSS11] Julia Dordel, Brad Seely, and Suzanne W. Simard. Relationships between simulated water stress and mortality and growth rates in underplanted Toona ciliata Roem. in subtropical Argentinean plantations. Ecological Modelling, 222(17):3226–3235, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003073.

Du:2017:CNM

[DSW^{+}17] Xiangjun Du, Fengjing Shao, Shunyao Wu, Rencheng Sun, and Changying Wang. Complex network modeling for mechanisms of red tide occurrence: a case study in Bohai Sea and North Yellow Sea of China. Ecological Modelling, 361(??):41–48, October 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300820.

DeAngelis:2010:FPD

[DTC^{+}10] Donald L. DeAngelis, Joel C. Trexler, Chris Cosner, Adam Obaza, and Fred Jopp. Fish population dynamics in a seasonally varying wetland. Ecological Modelling, 221(8):1131–1137, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008679.

Debeljak:2012:URD

[DTS^{+}12] Marko Debeljak, Aneta Trajanov, Daniela Stojanova, Florence Leprince, and Sašo Džeroski. Using relational decision trees to model out-crossing rates in a multi-field setting. Ecological Modelling, 245(??):75–83, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001810.

Diaz-Uribe:2012:IET

[DUASLB^{+}12] J. Gabriel Díaz-Uribe, Francisco Arreguín-Sánchez, Diego Lercari-Bernier, Víctor H. Cruz-Escalona, Manuel J. Zetina-Rejón, Pablo del Monte-Luna, and Susana Martínez-Aguilar. An integrated ecosystem trophic model for the North and Central Gulf of California: an alternative view for endemic species conservation. Ecological Modelling, 230(??):73–91, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000208.

Diaz-Varela:2010:STA

[DVCM^{+}10] Ramón Alberto Díaz-Varela, Roberto Colombo, Michele Meroni, María Silvia Calvo-Iglesias, Armando Buffoni, and Antonio Tagliaferri. Spatio-temporal analysis of alpine ecotones: a spatial explicit model targeting altitudinal vegetation shifts. Ecological Modelling, 221(4):621–633, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000900756X.

Duarte:2014:MSL

[DVD^{+}14] B. Duarte, J. M. Valentim, J. M. Dias, H. Silva, J. C. Marques, and I. Caçador. Modelling sea level rise (SLR) impacts on salt marsh detrital outwelling C and N exports from an estuarine coastal lagoon to the ocean (Ria de Aveiro, Portugal). Ecological Modelling, 289(??):36–44, October 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003020.

DeJager:2019:SEM

[DVF^{+}19] Nathan R. De Jager, Molly Van Appledorn, Timothy J. Fox, Jason J. Rohweder, Lyle J. Guyon, Andrew R. Meier, Robert J. Cosgriff, and Benjamin J. Vandermyde. Spatially explicit modelling of floodplain forest succession: interactions among flood inundation, forest successional processes, and other disturbances in the Upper Mississippi River floodplain, USA. Ecological Modelling, 405(??):15–32, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001930167X.

Dion:2011:LEF

[DVL11] Elise Dion, Louis VanSchalkwyk, and Eric F. Lambin. The landscape epidemiology of foot-and-mouth disease in South Africa: a spatially explicit multi-agent simulation. Ecological Modelling, 222(13):2059–2072, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001499.

deVillemereuil:2011:CFR

[dVLS11] Pierre B. de Villemereuil and Andrés López-Sepulcre. Consumer functional responses under intra- and interspecific interference competition. Ecological Modelling, 222(3):419–426, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005521.

Duboz:2010:AEA

[DVT^{+}10] Raphaël Duboz, David Versmisse, Morgane Travers, Eric Ramat, and Yunne-Jai Shin. Application of an evolutionary algorithm to the inverse parameter estimation of an individual-based model. Ecological Modelling, 221(5):840–849, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008102.

Wang:2015:SOA

[dWcWbZ15] Shi dong Wang, Xin chuang Wang, and He bing Zhang. Simulation on optimized allocation of land resource based on DE-CA model. Ecological Modelling, 314(??):135–144, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003178.

Davis:2019:DIE

[DWD19] Melanie J. Davis, Isa Woo, and Susan E. W. De La Cruz. Development and implementation of an empirical habitat change model and decision support tool for estuarine ecosystems. Ecological Modelling, 410(??):Article 108722, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302224.

Daraio:2010:MFI

[DWNN10] Joseph A. Daraio, Larry J. Weber, Teresa J. Newton, and John M. Nestler. A methodological framework for integrating computational fluid dynamics and ecological models applied to juvenile freshwater mussel dispersal in the Upper Mississippi River. Ecological Modelling, 221(2):201–214, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006619.

Duncan:2011:SNF

[DWP11] Brean W. Duncan, John F. Weishampel, and Seth H. Peterson. Simulating a natural fire regime on an Atlantic coast barrier island complex in Florida, USA. Ecological Modelling, 222(9):1639–1650, May 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000986.

Dugdale:2013:BMP

[DWP13] Richard C. Dugdale, Frances P. Wilkerson, and Alexander E. Parker. A biogeochemical model of phytoplankton productivity in an urban estuary: the importance of ammonium and freshwater flow. Ecological Modelling, 263(??):291–307, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002779.

Dugdale:2016:ECG

[DWP16] Richard C. Dugdale, Frances P. Wilkerson, and Alexander E. Parker. The effect of clam grazing on phytoplankton spring blooms in the low-salinity zone of the San Francisco Estuary: a modelling approach. Ecological Modelling, 340(??):1–16, November 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303684.

DeGregorio:2015:IEC

[DWWS15] Brett A. DeGregorio, James D. Westervelt, Patrick J. Weatherhead, and Jinelle H. Sperry. Indirect effect of climate change: Shifts in ratsnake behavior alter intensity and timing of avian nest predation. Ecological Modelling, 312(??):239–246, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002379.

Diaz-Yanez:2019:CMD

[DYMYGO19a] Olalla Díaz-Yáñez, Blas Mola-Yudego, and José Ramón González-Olabarria. Corrigendum to “Modelling damage occurrence by snow and wind in forest ecosystems” [Ecol. Modell. 408 (2019) 108741]. Ecological Modelling, 414(??):Article 108846, December 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303540.

Diaz-Yanez:2019:MDO

[DYMYGO19b] Olalla Díaz-Yáñez, Blas Mola-Yudego, and José Ramón González-Olabarria. Modelling damage occurrence by snow and wind in forest ecosystems. Ecological Modelling, 408(??):Article 108741, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302492.

Dong:2012:IHA

[DYU^{+}12] Xiaobin Dong, Weikun Yang, Sergio Ulgiati, Maochao Yan, and Xinshi Zhang. The impact of human activities on natural capital and ecosystem services of natural pastures in North Xinjiang, China. Ecological Modelling, 225(??):28–39, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005266.

Danuso:2012:MAE

[DZS12] Francesco Danuso, Giuseppe Zanin, and Ivan Sartorato. A modelling approach for evaluating phenology and adaptation of two congeneric weeds (Bidens frondosa and Bidens tripartita). Ecological Modelling, 243(??):33–41, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002815.

Deng:2011:ATB

[DZW^{+}11] Xiangzheng Deng, Yonghong Zhao, Feng Wu, Yingzhi Lin, Qi Lu, and Jing Dai. Analysis of the trade-off between economic growth and the reduction of nitrogen and phosphorus emissions in the Poyang Lake Watershed, China. Ecological Modelling, 222(2):330–336, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000445X.

Engen:2011:DEH

[EAB11] Steinar Engen, Kaare Aagaard, and Terje Bongard. Disentangling the effects of heterogeneity, stochastic dynamics and sampling in a community of aquatic insects. Ecological Modelling, 222(8):1387–1393, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000305.

Eliassen:2016:SEA

[EAJG16] Sigrunn Eliassen, Bjørn Snorre Andersen, Christian Jørgensen, and Jarl Giske. From sensing to emergent adaptations: Modelling the proximate architecture for decision-making. Ecological Modelling, 326(??):90–100, April 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003907.

Elliot:2019:SOU

[EBA^{+}19] Thomas Elliot, Alexandre Bertrand, Javier Babí Almenar, Claudio Petucco, Vânia Proença, and Benedetto Rugani. Spatial optimisation of urban ecosystem services through integrated participatory and multi-objective integer linear programming. Ecological Modelling, 409(??): Article 108774, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302820.

Everaert:2011:UCT

[EBL^{+}11] Gert Everaert, Pieter Boets, Koen Lock, Sašo Džeroski, and Peter L. M. Goethals. Using classification trees to analyze the impact of exotic species on the ecological assessment of polder lakes in Flanders, Belgium. Ecological Modelling, 222(14):2202–2212, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004084.

Ezard:2010:HSE

[EC10] Thomas H. G. Ezard and Tim Coulson. How sensitive are elasticities of long-run stochastic growth to how environmental variability is modelled? Ecological Modelling, 221(2):191–200, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009006383.

Erickson:2014:DPM

[ECO^{+}14] Richard A. Erickson, Stephen B. Cox, Jessica L. Oates, Todd A. Anderson, Christopher J. Salice, and Kevin R. Long. A Daphnia population model that considers pesticide exposure and demographic stochasticity. Ecological Modelling, 275(??):37–47, March 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013006029.

Erickson:2017:IPM

[EEB^{+}17] Richard A. Erickson, Eric A. Eager, Marybeth K. Brey, Michael J. Hansen, and Patrick M. Kocovsky. An integral projection model with YY-males and application to evaluating grass carp control. Ecological Modelling, 361(??):14–25, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303526.

Etterson:2011:MFB

[EEFE^{+}11] Matthew A. Etterson, Susan N. Ellis-Felege, David Evers, Gilles Gauthier, Joseph A. Grzybowski, Brady J. Mattsson, Laura R. Nagy, Brian J. Olsen, Craig M. Pease, Max Post van der Burg, and Aaron Potvien. Modeling fecundity in birds: Conceptual overview, current models, and considerations for future developments. Ecological Modelling, 222(14):2178–2190, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005545.

Erickson:2018:SDI

[EEK^{+}18] Richard A. Erickson, Eric A. Eager, Patrick M. Kocovsky, David C. Glover, Jahn L. Kallis, and K. R. Long. A spatially discrete, integral projection model and its application to invasive carp. Ecological Modelling, 387(??):163–171, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018303028.

Eglin:2010:LIS

[EFM^{+}10] Thomas Eglin, Christophe Francois, Alice Michelot, Nicolas Delpierre, and Claire Damesin. Linking intra-seasonal variations in climate and tree-ring δ ^{13}C: a functional modelling approach. Ecological Modelling, 221(15):1779–1797, July 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010002097.

ElHamidi:2012:PMC

[EGA12] A. El Hamidi, M. Garbey, and N. Ali. A PDE model of clonal plant competition with nonlinear diffusion. Ecological Modelling, 234(??):83–92, June 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001196.

Engelman:2012:MEC

[EGMW12] Catherine A. Engelman, William E. Grant, Miguel A. Mora, and Marc Woodin. Modelling effects of chemical exposure on birds wintering in agricultural landscapes: the western burrowing owl (Athene cunicularia hypugaea) as a case study. Ecological Modelling, 224(1):90–102, January 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005023.

Ehrich:2015:REE

[EH15] Melinda K. Ehrich and Lora A. Harris. A review of existing eastern oyster filtration rate models. Ecological Modelling, 297(??):201–212, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005961.

Ehrich:2019:CRE

[EH19] Melinda K. Ehrich and Lora A. Harris. Corrigendum to “A review of existing eastern oyster filtration rate models” [Ecological Modelling 297 (2015) 201–212]. Ecological Modelling, 404(??):112, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301450.

Esposito:2010:IMF

[EIG^{+}10] S. Esposito, G. Incerti, F. Giannino, D. Russo, and S. Mazzoleni. Integrated modelling of foraging behaviour, energy budget and memory properties. Ecological Modelling, 221(9):1283–1291, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010000487.

Engel:2018:PTC

[EKB18] Markus Engel, Michael Körner, and Uta Berger. Plastic tree crowns contribute to small-scale heterogeneity in virgin beech forests — an individual-based modeling approach. Ecological Modelling, 376(??):28–39, May 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300772.

Elser:2012:LRL

[ELPK12] James J. Elser, Irakli Loladze, Angela L. Peace, and Yang Kuang. Lotka re-loaded: Modeling trophic interactions under stoichiometric constraints. Ecological Modelling, 245(??):3–11, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000749.

Elsheikh:2010:TNP

[Els10] E. M. Elsheikh. Towards a new physical theory of biotic evolution and development. Ecological Modelling, 221(8):1108–1118, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008680.

Ebrahimi:2010:HSG

[EMH10] A. Ebrahimi, T. Milotić, and M. Hoffmann. A herbivore specific grazing capacity model accounting for spatiotemporal environmental variation: a tool for a more sustainable nature conservation and rangeland management. Ecological Modelling, 221(6):900–910, March 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008394.

Erickson:2015:PCD

[ENC^{+}15] Adam Erickson, Craig Nitschke, Nicholas Coops, Steven Cumming, and Gordon Stenhouse. Past-century decline in forest regeneration potential across a latitudinal and elevational gradient in Canada. Ecological Modelling, 313(??):94–102, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002768.

Eynaud:2011:IEH

[ENM11] Yoan Eynaud, Roger M. Nisbet, and Erik B. Muller. Impact of excess and harmful radiation on energy budgets in scleractinian corals. Ecological Modelling, 222(7):1315–1322, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011000263.

Erickson:2010:DMD

[EPA^{+}10a] Richard A. Erickson, Steven M. Presley, Linda J. S. Allen, Kevin R. Long, and Stephen B. Cox. A dengue model with a dynamic Aedes albopictus vector population. Ecological Modelling, 221(24):2899–2908, December 15, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004497.

Erickson:2010:SSA

[EPA^{+}10b] Richard A. Erickson, Steven M. Presley, Linda J. S. Allen, Kevin R. Long, and Stephen B. Cox. A stagestructured, Aedes albopictus population model. Ecological Modelling, 221(9):1273–1282, May 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000578.

Eddy:2014:LKO

[EPM^{+}14] Tyler D. Eddy, Tony J. Pitcher, Alison B. MacDiarmid, Tamsen T. Byfield, Jamie C. Tam, Timothy T. Jones, James J. Bell, and Jonathan P. A. Gardner. Lobsters as keystone: Only in unfished ecosystems? Ecological Modelling, 275(??):48–72, March 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005851.

Eldridge:2010:OSH

[ER10] Peter M. Eldridge and Daniel L. Roelke. Origins and scales of hypoxia on the Louisiana shelf: Importance of seasonal plankton dynamics and river nutrients and discharge. Ecological Modelling, 221(7):1028–1042, April 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009003226.

Erol:2013:WEM

[ER13] Ayten Erol and Timothy O. Randhir. Watershed ecosystem modeling of land-use impacts on water quality. Ecological Modelling, 270(??):54–63, December 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004341.

Espinosa-Romero:2011:ERM

[ERGW^{+}11a] Maria J. Espinosa-Romero, Edward J. Gregr, Carl Walters, Villy Christensen, and Kai M. A. Chan. Erratum to “Representing mediating effects and species reintroductions in Ecopath with Ecosim” [Ecol. Model. 222 (2011) 1569–1579]. Ecological Modelling, 222(13):2168, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001967.

Espinosa-Romero:2011:RME

[ERGW^{+}11b] Maria J. Espinosa-Romero, Edward J. Gregr, Carl Walters, Villy Christensen, and Kai M. A. Chan. Representing mediating effects and species reintroductions in Ecopath with Ecosim. Ecological Modelling, 222(9):1569–1579, May 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000755.

Eid:2012:MGD

[ESA12] Ebrahem M. Eid, Kamal H. Shaltout, and Takashi Asaeda. Modeling growth dynamics of Typha domingensis (Pers.) Poir. ex Steud. in Lake Burullus, Egypt. Ecological Modelling, 243(??):63–72, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002736.

Endredi:2018:FWD

[ESLJ18] Anett Endrédi, Vera Senánszky, Simone Libralato, and Ferenc Jordán. Food web dynamics in trophic hierarchies. Ecological Modelling, 368(??):94–103, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001730282X.

Endreny:2017:IMU

[ESP^{+}17] T. Endreny, R. Santagata, A. Perna, C. De Stefano, R. F. Rallo, and S. Ulgiati. Implementing and managing urban forests: a much needed conservation strategy to increase ecosystem services and urban wellbeing. Ecological Modelling, 360(??):328–335, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300960.

Evans:2019:QEA

[EST^{+}19] Luke C. Evans, Richard M. Sibly, Pernille Thorbek, Ian Sims, Tom H. Oliver, and Richard J. Walters. Quantifying the effectiveness of agri-environment schemes for a grassland butterfly using individual-based models. Ecological Modelling, 411(??):Article 108798, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303060.

Edholm:2018:MII

[ETG^{+}18] Christina J. Edholm, Brigitte Tenhumberg, Chris Guiver, Yu Jin, Stuart Townley, and Richard Rebarber. Management of invasive insect species using optimal control theory.

Ecological Modelling, 381(??):36–45, August 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301285.

Earl:2014:EAM

[EZ14] Julia E. Earl and Patrick A. Zollner. Effects of animal movement strategies and costs on the distribution of active subsidies across simple landscapes. Ecological Modelling, 283(??):45–52, July 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001720.

F:2011:DAS

[F.11] María I. Cabrera F. Deterministic approach to the study of the interaction predator–prey in a chemostat with predator mutual interference. Implications for the paradox of enrichment. Ecological Modelling, 222(3):598–605, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004795.

Fourriere:2019:EFS

[FAC^{+}19a] Manon Fourriére, Juan José Alvarado, Jorge Cortés, Marc H. Taylor, Arturo Ayala-Bocos, Juan Carlos Azofeifa-Solano, Randall Arauz, Maike Heidemeyer, Andrés López-Garro, Ilena Zanella, and Matthias Wolff. Energy flow structure and role of keystone groups in shallow water environments in Isla del Coco, Costa Rica, Eastern Tropical Pacific. Ecological Modelling, 396(??):74–85, March 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300183.

Fuente:2019:ECC

[FAC^{+}19b] G. De La Fuente, V. Asnaghi, M. Chiantore, S. Thrush, P. Povero, P. Vassallo, M. Petrillo, and C. Paoli. The effect of Cystoseira canopy on the value of midlittoral habitats in NW Mediterranean, an emergy assessment. Ecological Modelling, 404(??):1–11, July 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301425.

Fabina:2010:SPP

[FAG10] Nicholas S. Fabina, Karen C. Abbott, and R. Tucker Gilman. Sensitivity of plant–pollinator–herbivore communities to changes in phenology. Ecological Modelling, 221(3):453–458, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007066.

Fairweather:2010:FMS

[Fai10] John Fairweather. Farmer models of socio-ecologic systems: Application of causal mapping across multiple locations. Ecological Modelling, 221(3):555–562, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007121.

Feng:2012:PTT

[FAMD^{+}12] Zhilan Feng, Jorge A. Alfaro-Murillo, Donald L. DeAngelis, Jennifer Schmidt, Matthew Barga, Yiqiang Zheng, Muhammad Hanis B. Ahmad Tamrin, Mark Olson, Tim Glaser, Knut Kielland, F. Stuart Chapin, and John Bryant. Plant toxins and trophic cascades alter fire regime and succession on a boreal forest landscape. Ecological Modelling, 244(??):79–92, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003018.

Freitas:2012:MRE

[FAPM12] Simone R. Freitas, Marcos M. Alexandrino, Renata Pardini, and Jean Paul Metzger. A model of road effect using line integrals and a test of the performance of two new road indices using the distribution of small mammals in an Atlantic forest landscape. Ecological Modelling, 247(??):64–70, December 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012003821.

Foster:2019:ITS

[FAS^{+}19] Adrianna C. Foster, Amanda H. Armstrong, Jacquelyn K. Shuman, Herman H. Shugart, Brendan M. Rogers, Michelle C. Mack, Scott J. Goetz, and K. Jon Ranson. Importance of tree- and species-level interactions with wildfire, climate, and soils in interior Alaska: Implications for forest change under a warming climate. Ecological Modelling, 409(??):Article 108765, October 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930273X.

Fath:2010:NEM

[Fat10] Brian D. Fath. Notes from the editor, mid-2010 update. Ecological Modelling, 221(21):2509–2511, October 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003790.

Fath:2012:EPU

[Fat12] Brian D. Fath. Editorial policy and update. Ecological Modelling, 232(??):v–viii, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000956.

Fath:2014:SSP

[Fat14] Brian D. Fath. Sustainable systems promote wholenessextending transformations: the contributions of systems thinking. Ecological Modelling, 293(??):42–48, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000362.

Fath:2015:NA

[Fat15a] Brian D. Fath. New announcements. Ecological Modelling, 297(??):A1, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005870.

Frisk:2015:EES

[FAT^{+}15b] C. Frisk, K. H. Andersen, A. Temming, J. P. Herrmann, K. S. Madsen, and G. Kraus. Environmental effects on sprat (Sprattus sprattus) physiology and growth at the distribution frontier: a bioenergetic modelling approach. Ecological Modelling, 299(??):130–139, March 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005997.

Fath:2016:E

[Fat16a] Brian D. Fath. Editorial. Ecological Modelling, 319(??): 1–2, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380015005001.

Fath:2016:SEJ

[Fat16b]  Brian D. Fath.  Sven Erik Jørgensen (1934–2016):

Indomitable will and intellect. Ecological Modelling, 330(??):60–61, June 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300886.

Frank:2013:SBT

[FB13] Béatrice M. Frank and Philippe V. Baret. Simulating brown trout demogenetics in a river/nursery brook system: the individual-based model DemGenTrout. Ecological Modelling, 248(??):184–202, January 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004838.

Ferraro:2015:LTS

[FB15] D. O. Ferraro and P. Benzi. A long-term sustainability assessment of an Argentinian agricultural system based on emergy synthesis. Ecological Modelling, 306(??):121–129, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014002981.

Fortin:2011:ECD

[FBA^{+}11] J. G. Fortin, M. A. Bolinder, F. Anctil, T. Kätterer, O. Andrén, and L. E. Parent. Effects of climatic data low-pass filtering on the ICBM temperature- and moisturebased soil biological activity factors in a cool and humid climate. Ecological Modelling, 222(17):3050–3060, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003449.

Ferguson:2018:EMV

[FBB^{+}18] Paige F. B. Ferguson, Rachel Breyta, Ilana Brito, Gael Kurath, and Shannon L. LaDeau. An epidemiological model of virus transmission in salmonid fishes of the Columbia River basin. Ecological Modelling, 377(??):1–15, June 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300784.

Fischer:2016:LLA

[FBdP^{+}16] Rico Fischer, Friedrich Bohn, Mateus Dantas de Paula, Claudia Dislich, Jürgen Groeneveld, Alvaro G. Gutiérrez, Martin Kazmierczak, Nikolai Knapp, Sebastian Lehmann, Sebastian Paulick, Sandro Pütz, Edna Rödig, Franziska Taubert, Peter Köhler, and Andreas Huth. Lessons learned from applying a forest gap model to understand ecosystem and carbon dynamics of complex tropical forests. Ecological Modelling, 326(??):124–133, April 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005505.

Franzese:2017:NCA

[FBDR17] Pier Paolo Franzese, Elvira Buonocore, Luigia Donnarumma, and Giovanni F. Russo. Natural capital accounting in marine protected areas: the case of the Islands of Ventotene and S. Stefano (Central Italy). Ecological Modelling, 360(??):290–299, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001730248X.

Fu:2012:MRG

[FBDZ12] Lintao Fu, Tianli Bo, Guozhen Du, and Xiaojing Zheng. Modeling the responses of grassland vegetation coverage to grazing disturbance in an alpine meadow. Ecological Modelling, 247(??):221–232, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004383.

Finn:2014:ANM

[FBH^{+}14] J. T. Finn, J. W. Brownscombe, C. R. Haak, S. J. Cooke, R. Cormier, T. Gagne, and A. J. Danylchuk. Applying network methods to acoustic telemetry data: Modeling the movements of tropical marine fishes. Ecological Modelling, 293(??):139–149, December 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013006017.

Fullman:2017:PSL

[FBKS17] Timothy J. Fullman, Erin L. Bunting, Gregory A. Kiker, and Jane Southworth. Predicting shifts in large herbivore distributions under climate change and management using a spatially-explicit ecosystem model. Ecological Modelling, 352(??):1–18, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630607X.

Fovet:2010:MPI

[FBL^{+}10] O. Fovet, G. Belaud, X. Litrico, S. Charpentier, C. Bertrand, A. Dauta, and C. Hugodot. Modelling periphyton in irrigation canals. Ecological Modelling, 221(8):1153–1161, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010000232.

Franzese:2014:EAE

[FBU14] Pier Paolo Franzese, Mark T. Brown, and Sergio Ulgiati. Environmental accounting: Emergy, systems ecology, and ecological modelling. Ecological Modelling, 271(??):1–3, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013004742.

Forzieri:2011:SDR

[FC11] Giovanni Forzieri and Filippo Catani. Scale-dependent relations in land cover biophysical dynamics. Ecological Modelling, 222(17):3285–3290, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003437.

Figueiredo:2012:DMC

[FC12] Joana Figueiredo and Sean R. Connolly. Dispersalmediated coexistence under recruitment limitation and displacement competition. Ecological Modelling, 243 (??):133–142, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001200302X.

Freitas:2018:CGA

[FCA18] Simone R. Freitas, Everton Constantino, and Marcos M. Alexandrino. Computational geometry applied to develop new metrics of road and edge effects and their performance to understand the distribution of small mammals in an Atlantic forest landscape. Ecological Modelling, 388(??):24–30, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302679.

Fondevilla:2016:UNP

[FCFT16] Cristian Fondevilla, M. Àngels Colomer, Federico Fillat, and Ulrike Tappeiner. Using a new PDP modelling approach for land-use and land-cover change predictions: a case study in the Stubai Valley (Central Alps). Ecological Modelling, 322(??):101–114, February 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005487.

Fang:2019:IUC

[FCG^{+}19] Xia Fang, Zhi Chen, Xulin Guo, Shihua Zhu, Tong Liu, Chaofan Li, and Biao He. Impacts and uncertainties of climate/CO_{2} change on net primary productivity in Xinjiang, China (2000–2014): a modelling approach. Ecological Modelling, 408(??):Article 108742, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302509.

Feng:2017:DFA

[FCH^{+}17] Le Feng, Bin Chen, Tasawar Hayat, Ahmed Alsaedi, and Bashir Ahmad. Dynamic forecasting of agricultural water footprint based on Markov Chain — a case study of the Heihe River basin. Ecological Modelling, 353(??):150–157, June 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016306536.

Fois:2018:USD

[FCLFB18] Mauro Fois, Alba Cuena-Lombraña, Giuseppe Fenu, and Gianluigi Bacchetta. Using species distribution models at local scale to guide the search of poorly known species: Review, methodological issues and future directions. Ecological Modelling, 385(??):124–132, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830259X.

Flynn:2013:MLV

[FCS13] Kyle F. Flynn, Steven C. Chapra, and Michael W. Suplee. Modeling the lateral variation of bottom-attached algae in rivers. Ecological Modelling, 267(??):11–25, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003566.

Fu:2012:BCS

[FCV^{+}12] Yongshuo H. Fu, Matteo Campioli, Marcel Van Oijen, Gaby Deckmyn, and Ivan A. Janssens. Bayesian comparison of six different temperature-based budburst models for four temperate tree species. Ecological Modelling, 230(??):92–100, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001200021X.

Feng:2013:SDM

[FCZ13] Y. Y. Feng, S. Q. Chen, and L. X. Zhang. System dynamics modeling for urban energy consumption and CO_{2} emissions: a case study of Beijing, China. Ecological Modelling, 252(??):44–52, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004735.

Fondi:2019:SEM

[FD19] Marco Fondi and Francesca Di Patti. A synthetic ecosystem for the multi-level modelling of heterotroph-phototroph metabolic interactions. Ecological Modelling, 399(??):13–22, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001930081X.

Fort:2017:TEA

[FDHL17] Hugo Fort, Francisco Dieguez, Virginia Halty, and Juan Manuel Soares Lima. Two examples of application of ecological modeling to agricultural production: Extensive livestock farming and overyielding in grassland mixtures. Ecological Modelling, 357(??):23–34, August 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305452.

Fukuda:2011:EMF

[FDM^{+}11] Shinji Fukuda, Bernard De Baets, Ans M. Mouton, Willem Waegeman, Jun Nakajima, Takahiko Mukai, Kazuaki Hiramatsu, and Norio Onikura. Effect of model formulation on the optimization of a genetic Takagi–Sugeno fuzzy system for fish habitat suitability evaluation. Ecological Modelling, 222(8):1401–1413, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100055X.

Fabre:2010:HBM

[FDP^{+}10] Frédéric Fabre, Charles-Antoine Dedryver, Manuel Plantegenest, Maurice Hullé, and Etienne Rivot. Hierarchical Bayesian modelling of plant colonisation by winged aphids:

Inferring dispersal processes by linking aerial and field count data. Ecological Modelling, 221(15):1770–1778, July 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010001924.

Fang:2017:CFU

[FDR^{+}17] Kai Fang, Liang Dong, Jingzheng Ren, Qifeng Zhang, Ling Han, and Huizhen Fu. Carbon footprints of urban transition: Tracking circular economy promotions in Guiyang, China. Ecological Modelling, 365(??):30–44, December 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017302958.

Ferriss:2014:CFC

[FE14a] Bridget E. Ferriss and Timothy E. Essington. Can fish consumption rate estimates be improved by linking bioenergetics and mercury mass balance models? Application to tunas. Ecological Modelling, 272(??):232–241, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004778.

Ferriss:2014:DTS

[FE14b] Bridget E. Ferriss and Timothy E. Essington. Does trophic structure dictate mercury concentrations in top predators? A comparative analysis of pelagic food webs in the Pacific Ocean. Ecological Modelling, 278(??):18–28, April 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000726.

Felzer:2012:CNW

[Fel12] Benjamin S. Felzer. Carbon, nitrogen, and water response to climate and land use changes in Pennsylvania during the 20th and 21st centuries. Ecological Modelling, 240(??):49–63, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002232.

Freire-Gonzalez:2011:MEE

[FG11] Jaume Freire-González. Methods to empirically estimate direct and indirect rebound effect of energy-saving technological changes in households. Ecological Modelling, 223(1):32–40, December 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004595.

Fath:2014:E

[FGG^{+}14] Brian D. Fath, William E. Grant, Volker Grimm, Alexander S. Komarov, and Santanu Ray. Editorial. Ecological Modelling, 287(??):A1–A4, September 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002889.

Feng:2010:MPD

[FGH^{+}10] Hongqiang Feng, Fred Gould, Yunxin Huang, Yuying Jiang, and Kongming Wu. Modeling the population dynamics of cotton bollworm Helicoverpa armigera (Hübner) (Lepidoptera: Noctuidae) over a wide area in northern China. Ecological Modelling, 221(15):1819–1830, July 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001894.

Filotas:2010:EPI

[FGPR10] Elise Filotas, Martin Grant, Lael Parrott, and Per Arne Rikvold. The effect of positive interactions on community structure in a multi-species metacommunity model along an environmental gradient. Ecological Modelling, 221(6):885–894, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008357.

Fahse:2011:SAO

[FH11] Lorenz Fahse and Marco Heurich. Simulation and analysis of outbreaks of bark beetle infestations and their management at the stand level. Ecological Modelling, 222(11):1833–1846, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011001372.

Fernandez:2012:DAM

[FHAG12] Miguel Fernández, Healy Hamilton, Otto Alvarez, and Qinghua Guo. Does adding multi-scale climatic variability improve our capacity to explain niche transferability in invasive species? Ecological Modelling, 246(??):60–67, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012003742.

Fritsch:2015:MAC

[FHC15] Coralie Fritsch, Jérôme Harmand, and Fabien Campillo. A modeling approach of the chemostat. Ecological Modelling, 299(??):1–13, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005948.

Fitzgerald:2012:MSA

[FHG12] Katherine Fitzgerald, Nicole Heller, and Deborah M. Gordon. Modeling the spread of the Argentine ant into natural areas: Habitat suitability and spread from neighboring sites. Ecological Modelling, 247(??):262–272, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003857.

Fiechter:2013:BPE

[FHL^{+}13] J. Fiechter, R. Herbei, W. Leeds, J. Brown, R. Milliff, C. Wikle, A. Moore, and T. Powell. A Bayesian parameter estimation method applied to a marine ecosystem model for the coastal Gulf of Alaska. Ecological Modelling, 258(??):122–133, June 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300152X.

Fiechter:2012:AME

[Fie12] Jerome Fiechter. Assessing marine ecosystem model properties from ensemble calculations. Ecological Modelling, 242(??):164–179, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002372.

Fernandes:2010:FRP

[FIG^{+}10] Jose A. Fernandes, Xabier Irigoien, Nerea Goikoetxea, Jose A. Lozano, Iñaki Inza, Aritz Pérez, and Antonio Bode. Fish recruitment prediction, using robust supervised classification methods. Ecological Modelling, 221(2):338–352, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009006437.

Fischer:2021:HDU

[Fis21] Andreas Fischer. How to determine the unique contributions of input-variables to the nonlinear regression function of a multilayer perceptron. Ecological Modelling, 309–310(??):60–63, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001660.

Fernandez:2015:LES

[FJH15] Pilar Fernández, Diego Jordano, and Juan Fernández Haeger. Living on the edge in species distribution models: the unexpected presence of three species of butterflies in a protected area in southern Spain. Ecological Modelling, 312(??):335–346, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002380.

Fourcaud:2014:ESI

[FKDG14] Thierry Fourcaud, Mengzhen Kang, Yves Dumont, and Yan Guo. Editorial of the special issue of the 4th International Symposium on Plant Growth Modeling, Simulation, Visualization and Applications (PMA’12). Ecological Modelling, 290(??):1–2, October 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003676.

Fan:2015:KDD

[FKH^{+}15] Xing-Rong Fan, Meng-Zhen Kang, Ep Heuvelink, Philippe de Reffye, and Bao-Gang Hu. A knowledge-and-data-driven modeling approach for simulating plant growth: a case study on tomato growth. Ecological Modelling, 312 (??):363–373, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002550.

Fruh:2018:MPD

[FKK^{+}18] Linus Früh, Helge Kampen, Antje Kerkow, Günter A. Schaub, Doreen Walther, and Ralf Wieland. Modelling the potential distribution of an invasive mosquito species: comparative evaluation of four machine learning methods and their combinations. Ecological Modelling, 388 (??):136–144, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302801.

Franz:2010:UER

[FKSK^{+}10] Mathias Franz, Stephanie Kramer-Schadt, Werner Kilian, Christian Wissel, and Jürgen Groeneveld. Understanding the effects of rainfall on elephant–vegetation interactions around waterholes. Ecological Modelling, 221 (24):2909–2917, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004588.

Fan:2019:WEN

[FKWZ19] Jing-Li Fan, Ling-Si Kong, Hang Wang, and Xian Zhang. A water-energy nexus review from the perspective of urban metabolism. Ecological Modelling, 392(??):128–136, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304071.

Fibich:2011:DBI

[FL11] Pavel Fibich and Jan Lepš. Do biodiversity indices behave as expected from traits of constituent species in simulated scenarios? Ecological Modelling, 222(13):2049–2058, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011002213.

Franzese:2019:EAM

[FLA19] Pier Paolo Franzese, Gengyuan Liu, and Salvatore Aricò. Environmental accounting models and nature conservation strategies. Ecological Modelling, 397(??):36–38, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300365.

Forio:2015:BBN

[FLB^{+}15] Marie Anne Eurie Forio, Dries Landuyt, Elina Bennetsen, Koen Lock, Thi Hanh Tien Nguyen, Minar Naomi Damanik Ambarita, Peace Liz Sasha Musonge, Pieter Boets, Gert Everaert, Luis Dominguez-Granda, and Peter L. M. Goethals. Bayesian belief network models to analyse and predict ecological water quality in rivers. Ecological Modelling, 312 (??):222–238, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002227.

Fay:2017:AEO

[FLH17] Gavin Fay, Jason S. Link, and Jonathan A. Hare. Assessing the effects of ocean acidification in the Northeast US using an end-to-end marine ecosystem model. Ecological Modelling, 347(??):1–10, March 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308237.

Fargue-Lelievre:2011:IFT

[FLLB11] A. Fargue-Lelièvre, D. Le Cœur, and J. Baudry. Integrating farming techniques in an ecological matrix model: Implementation on the primrose (Primula vulgaris). Ecological Modelling, 222(4):1002–1015, February 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006344.

Fay:2013:TSF

[FLLG13] Gavin Fay, Scott I. Large, Jason S. Link, and Robert J. Gamble. Testing systemic fishing responses with ecosystem indicators. Ecological Modelling, 265(??):45–55, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002780.

Flores:2014:MLP

[Flo14] J. C. Flores. Modelling late Pleistocene megafaunal extinction and critical cases: a simple prey–predator perspective. Ecological Modelling, 291(??):218–223, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003810.

Flores:2018:DFD

[Flo18] J. C. Flores. Decreasing fractal dimensions as a strategy for oceanic wildlife conservation: Application to species with large migration patterns. Ecological Modelling, 384(??):30–33, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830200X.

Fan:2014:AIC

[FLX^{+}14] Juntao Fan, Junsheng Li, Rui Xia, Lile Hu, Xiaopu Wu, and Guo Li. Assessing the impact of climate change on the habitat distribution of the giant panda in the qinling mountains of China. Ecological Modelling, 274(??):12–20, February 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013005711.

Fan:2013:LCC

[FLY13] Ze-Meng Fan, Jing Li, and Tian-Xiang Yue. Land-cover changes of biome transition zones in Loess Plateau of China. Ecological Modelling, 252(??):129–140, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200405X.

Fayle:2010:RRD

[FM10] Tom M. Fayle and Andrea Manica. Reducing over-reporting of deterministic co-occurrence patterns in biotic communities. Ecological Modelling, 221(19):2237–2242, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010002826.

Fayle:2011:BNM

[FM11] Tom M. Fayle and Andrea Manica. Bias in null model analyses of species co-occurrence: a response to gotelli and ulrich (2011). Ecological Modelling, 222(7):1340–1341, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000329.

Fennell:2012:PSS

[FMA^{+}12] Mark Fennell, James E. Murphy, Cristina Armstrong, Tommy Gallagher, and Bruce Osborne. Plant spread simulator: a model for simulating large-scale directed dispersal processes across heterogeneous environments. Ecological Modelling, 230(??):1–10, April 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000191.

Falkowski:2015:HVC

[FMBD15] Tomasz B. Falkowski, Isaias Martinez-Bautista, and Stewart A. W. Diemont. How valuable could traditional ecological knowledge education be for a resource-limited future?: an emergy evaluation in two Mexican villages. Ecological Modelling, 300(??):40–49, March 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006139.

Freeman:2012:EES

[FMF12] Elizabeth A. Freeman, Gretchen G. Moisen, and Tracey S. Frescino. Evaluating effectiveness of down-sampling for stratified designs and unbalanced prevalence in random forest models of tree species distributions in Nevada. Ecological Modelling, 233(??):1–10, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001147.

Ficker:2016:SSP

[FMG^{+}16] Harald Ficker, Rupert Mazzucco, Hubert Gassner, Josef Wanzenböck, and Ulf Dieckmann. Stocking strategies for a pre-alpine whitefish population under temperature stress. Ecological Modelling, 320(??):170–176, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004603.

Fuentes:2017:SRE

[FML^{+}17] Mariana M. P. B. Fuentes, Jonathan Monsinjon, Milagros Lopez, Paulo Lara, Alexsandro Santos, Maria A. G. dei Marcovaldi, and Marc Girondot. Sex ratio estimates for species with temperature-dependent sex determination differ according to the proxy used. Ecological Modelling, 365(??):55–67, December 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301850.

Frisk:2011:ABG

[FMLM11] M. G. Frisk, T. J. Miller, R. J. Latour, and S. J. D. Martell. Assessing biomass gains from marsh restoration in Delaware Bay using Ecopath with Ecosim. Ecological Modelling, 222(1):190–200, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004217.

Feddema:2013:MHS

[FMS13] Johannes J. Feddema, Joy Nystrom Mast, and Melissa Savage. Modeling high-severity fire, drought and climate change impacts on ponderosa pine regeneration. Ecological Modelling, 253(??):56–69, March 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000124.

Fern:2019:IBR

[FMW^{+}19] Rachel R. Fern, Michael L. Morrison, Hsiao-Hsuan Wang, William E. Grant, and Tyler A. Campbell. Incorporating biotic relationships improves species distribution models: Modeling the temporal influence of competition in conspecific nesting birds. Ecological Modelling, 408(??):Article 108743, September 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302510.

Fernandez:2015:ESS

[FN15] Daniel Fernández and Miguel Nakamura. Estimation of spatial sampling effort based on presence-only data and accessibility. Ecological Modelling, 299(??):147–155, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014006309.

Fitzpatrick:2018:ADR

[FN18] Kimberly B. Fitzpatrick and Thomas M. Neeson. Aligning dam removals and road culvert upgrades boosts conservation return-on-investment. Ecological Modelling, 368 (??):198–204, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304799.

Frisman:2011:CDP

[FNR11] E. Y. Frisman, G. P. Neverova, and O. L. Revutskaya. Complex dynamics of the population with a simple age structure. Ecological Modelling, 222(12):1943–1950, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001955.

Foran:2011:LCT

[For11] Barney Foran. Low carbon transition options for Australia. Ecological Modelling, 223(1):72–80, December 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002705.

Forrester:2014:SLL

[For14] David I. Forrester. A stand-level light interception model for horizontally and vertically heterogeneous canopies. Ecological Modelling, 276(??):14–22, March 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000064.

Fort:2018:PSY

[For18a] Hugo Fort. On predicting species yields in multispecies communities: Quantifying the accuracy of the linear Lotka– Volterra generalized model. Ecological Modelling, 387 (??):154–162, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303053.

Fort:2018:QPC

[For18b] Hugo Fort. Quantitative predictions from competition theory with an incomplete knowledge of model parameters tested against experiments across diverse taxa. Ecological Modelling, 368(??):104–110, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304441.

Fuentes:2013:UMM

[FP13] M. M. P. B. Fuentes and W. P. Porter. Using a microclimate model to evaluate impacts of climate change on sea turtles. Ecological Modelling, 251(??):150–157, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000033.

Feng:2017:PLE

[FP17] Xiao Feng and Monica Papeş. Physiological limits in an ecological niche modeling framework: a case study of water temperature and salinity constraints of freshwater bivalves invasive in USA. Ecological Modelling, 346(??):48–57, February 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016306883.

Fyllas:2010:SRV

[FPG^{+}10] Nikolaos M. Fyllas, Patrizia I. Politi, Alexandros Galanidis, Panayiotis G. Dimitrakopoulos, and Margarita Arianoutsou. Simulating regeneration and vegetation dynamics in Mediterranean coniferous forests. Ecological Modelling, 221(11):1494–1504, June 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001377.

Fulford:2011:EMH

[FPG11] R. S. Fulford, M. S. Peterson, and P. O. Grammer. An ecological model of the habitat mosaic in estuarine nursery areas: Part I — interaction of dispersal theory and habitat variability in describing juvenile fish distributions. Ecological Modelling, 222(17):3203–3215, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003711.

Fahse:2018:EAM

[FPO18] Lorenz Fahse, Phillip Papastefanou, and Mathias Otto. Estimating acute mortality of lepidoptera caused by the cultivation of insect-resistant Bt maize — the LepiX model. Ecological Modelling, 371(??):50–59, March 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300152.

Fulford:2014:EMH

[FPWG14] R. S. Fulford, M. S. Peterson, W. Wu, and P. O. Grammer. An ecological model of the habitat mosaic in estuarine nursery areas: Part II — projecting effects of sea level rise on fish production. Ecological Modelling, 273(??):96–108, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005310.

Fasani:2011:FPI

[FR11] Stefano Fasani and Sergio Rinaldi. Factors promoting or inhibiting Turing instability in spatially extended prey–predator systems. Ecological Modelling, 222(18):

3449–3452, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003723.

Fretzer:2016:UEA

[Fre16] Sarah Fretzer. Using the Ecopath approach for environmental impact assessment — a case study analysis. Ecological Modelling, 331(??):160–172, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004536.

Fenocchi:2019:AOD

[FRM^{+}19] Andrea Fenocchi, Michela Rogora, Giuseppe Morabito, Aldo Marchetto, Stefano Sibilla, and Claudia Dresti. Applicability of a one-dimensional coupled ecological-hydrodynamic numerical model to future projections in a very deep large lake (Lake Maggiore, Northern Italy/Southern Switzerland). Ecological Modelling, 392(??):38–51, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303739.

Frossard:2018:CNR

[FRP18] Victor Frossard, Frédéric Rimet, and Marie-Elodie Perga. Causal networks reveal the dominance of bottom-up interactions in large, deep lakes. Ecological Modelling, 368 (??):136–146, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304829.

Flindt:2016:UGT

[FRV^{+}16] Mogens R. Flindt, Erik K. Rasmussen, Thomas Valdemarsen, Anders Erichsen, Hanne Kaas, and Paula Canal-Vergés. Using a GIS-tool to evaluate potential eelgrass reestablishment in estuaries. Ecological Modelling, 338 (??):122–134, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016302472.

Fan:2014:NSS

[FS14a] Wei Fan and Jinbao Song. A numerical study of the seasonal variations of nutrients in the Changjiang River estuary and its adjacent sea area. Ecological Modelling, 291(??):69–81, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003639.

Feugier:2014:HFC

[FS14b] François G. Feugier and Akiko Satake. Hyperbolic features of the circadian clock oscillations can explain linearity in leaf starch dynamics and adaptation of plants to diverse light and dark cycles. Ecological Modelling, 290(??):110–120, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013005474.

Fust:2018:IST

[FS18] Pascal Fust and Eva Schlecht. Integrating spatio-temporal variation in resource availability and herbivore movements into rangeland management: RaMDry — an agent-based model on livestock feeding ecology in a dynamic, heterogeneous, semi-arid environment. Ecological Modelling, 369(??):13–41, February 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016304227.

Fath:2013:DNM

[FSB13] Brian D. Fath, Ursula M. Scharler, and Dan Baird. Dependence of network metrics on model aggregation and throughflow calculations: Demonstration using the Sylt–Rømø Bight Ecosystem. Ecological Modelling, 252 (??):214–219, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002827.

Fitzgerald:2019:EMO

[FSM^{+}19a] Colm J. Fitzgerald, Samuel Shephard, Paul McLoone, Fiona L. Kelly, and Keith D. Farnsworth. Evaluating management options for two fisheries that conflict through predator–prey interactions of target species. Ecological Modelling, 410(??):Article 108740, October 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302480.

Frighetto:2019:STP

[FSM19b] Daiane Frighetto Frighetto, Gustavo Maia Souza, and Alexandre Molter. Spatio-temporal population control applied to management of aquatic plants. Ecological Modelling, 398(??):77–84, April 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303120.

Fetahi:2011:FWS

[FSML11] Tadesse Fetahi, Michael Schagerl, Seyoum Mengistou, and Simone Libralato. Food web structure and trophic interactions of the tropical highland Lake Hayq, Ethiopia. Ecological Modelling, 222(3):804–813, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005259.

Febrina:2015:MPA

[FSN^{+}15] Rina Febrina, Masahiko Sekine, Hiroyuki Noguchi, Koichi Yamamoto, Ariyo Kanno, Takaya Higuchi, and Tsuyoshi Imai. Modeling the preference of ayu (Plecoglossus altivelis) for underwater sounds to determine the migration path in a river. Ecological Modelling, 299(??):102–113, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014006164.

Foster:2017:VAF

[FSS^{+}17] Adrianna C. Foster, Jacquelyn K. Shuman, Herman H. Shugart, Kathleen A. Dwire, Paula J. Fornwalt, Jason Sibold, and Jose Negron. Validation and application of a forest gap model to the southern Rocky Mountains. Ecological Modelling, 351(??):109–128, May 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306482.

Fei:2012:PMT

[FSV^{+}12] Teng Fei, Andrew K. Skidmore, Valentijn Venus, Tiejun Wang, Bert Toxopeus, Meng Bian, and Yaolin Liu. Predicting micro thermal habitat of lizards in a dynamic thermal environment. Ecological Modelling, 231(??):126–133, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000804.

Ferrarini:2010:NAA

[FT10] Alessandro Ferrarini and Marcello Tomaselli. A new approach to the analysis of adjacencies: Potentials for landscape insights. Ecological Modelling, 221(16):1889–1896, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001000236X.

Forrester:2016:AST

[FT16] David I. Forrester and Xiaolu Tang. Analysing the spatial and temporal dynamics of species interactions in mixedspecies forests and the effects of stand density using the 3-PG model. Ecological Modelling, 319(??):233–254, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003166.

Focks:2014:ICF

[FtHvdB^{+}14] Andreas Focks, Mechteld ter Horst, Erik van den Berg, Hans Baveco, and Paul J. van den Brink. Integrating chemical fate and population-level effect models for pesticides at landscape scale: New options for risk assessment. Ecological Modelling, 280(??):102–116, May 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004614.

Fand:2014:PIC

[FTK^{+}14] Babasaheb B. Fand, Henri E. Z. Tonnang, Mahesh Kumar, Santanu K. Bal, Naveen P. Singh, D. V. K. N. Rao, Ankush L. Kamble, Dhananjay D. Nangare, and Paramjit S. Minhas. Predicting the impact of climate change on regional and seasonal abundance of the mealybug Phenacoccus solenopsis Tinsley (Hemiptera: Pseudococcidae) using temperature-driven phenology model linked to GIS. Ecological Modelling, 288(??):62–78, September 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002671.

Fu:2018:RBC

[FTTV^{+}18] Caihong Fu, Morgane Travers-Trolet, Laure Velez, Arnaud Grüss, Alida Bundy, Lynne J. Shannon, Elizabeth A. Fulton, Ekin Akoglu, Jennifer E. Houle, Marta Coll, Philippe Verley, Johanna J. Heymans, Emma John, and Yunne-Jai Shin. Risky business: the combined effects of fishing and changes in primary productivity on fish communities. Ecological Modelling, 368(??):265–276, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302661.

Fukano:2013:SDC

[FTYI13] Yuya Fukano, Yuuya Tachiki, Tetsukazu Yahara, and Yoh Iwasa. Soil disturbances can suppress the invasion of alien plants under plant–soil feedback. Ecological Modelling, 260(??):42–49, July 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001798.

Fujiwara:2016:IDD

[Fuj16] Masami Fujiwara. Incorporating demographic diversity into food web models: Effects on community structure and dynamics. Ecological Modelling, 322(??):10–18, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005475.

Franck:2011:HSN

[FvBM^{+}11] Siegfried Franck, Werner von Bloh, Christoph Müller, Alberte Bondeau, and B. Sakschewski. Harvesting the sun: New estimations of the maximum population of planet Earth. Ecological Modelling, 222(12):2019–2026, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001670.

Fulton:2010:HMM

[FWRZ10] John W. Fulton, Chad R. Wagner, Megan E. Rogers, and Gregory F. Zimmerman. Hydraulic modeling of mussel habitat at a bridge-replacement site, Allegheny River, Pennsylvania, USA. Ecological Modelling, 221(3):540–554, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009007054.

Fleming:2018:CAV

[FYR^{+}18] Alyson H. Fleming, Tina Yack, Jessica V. Redfern, Elizabeth A. Becker, Thomas J. Moore, and Jay Barlow. Combining acoustic and visual detections in habitat models of Dall’s porpoise. Ecological Modelling, 384(??):198–208, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302199.

Folberth:2012:IID

[FYWA12] Christian Folberth, Hong Yang, Xiuying Wang, and Karim C. Abbaspour. Impact of input data resolution and extent of harvested areas on crop yield estimates in largescale agricultural modeling for maize in the USA. Ecological Modelling, 235–236(??):8–18, June 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001627.

Ghezzehei:2015:SDR

[GA15] Teamrat A. Ghezzehei and Ammar A. Albalasmeh. Spatial distribution of rhizodeposits provides built-in water potential gradient in the rhizosphere. Ecological Modelling, 298(??):53–63, February 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005262.

Giannetti:2012:CES

[GAB12] Biagio F. Giannetti, Cecilia M. V. B. Almeida, and Silvia H. Bonilla. Can emergy sustainability index be improved? Complementary insights for extending the vision. Ecological Modelling, 244(??):158–161, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200107X.

Gilhespy:2014:FYD

[GAC^{+}14] Sarah L. Gilhespy, Steven Anthony, Laura Cardenas, David Chadwick, Agustin del Prado, Changsheng Li, Thomas Misselbrook, Robert M. Rees, William Salas, Alberto Sanz-Cobena, Pete Smith, Emma L. Tilston, Cairistiona F. E. Topp, Sylvia Vetter, and Jagadeesh B. Yeluripati. First 20 years of DNDC (DeNitrification DeComposition): Model evolution. Ecological Modelling, 292(??):51–62, November 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004190.

Grimm:2014:TBM

[GAF^{+}14] Volker Grimm, Jacqueline Augusiak, Andreas Focks, Béatrice M. Frank, Faten Gabsi, Alice S. A. Johnston, Chun Liu, Benjamin T. Martin, Mattia Meli, Viktoriia Radchuk, Pernille Thorbek, and Steven F. Railsback. Towards better modelling and decision support: Documenting model development, testing, and analysis using TRACE. Ecological Modelling, 280(??):129–139, May 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000611.

Giannini:2012:PSR

[GAG^{+}12] Tereza C. Giannini, André L. Acosta, Carlos A. Garófalo, Antonio M. Saraiva, Isabel Alves dos Santos, and Vera L. Imperatriz-Fonseca. Pollination services at risk: Bee habitats will decrease owing to climate change in Brazil. Ecological Modelling, 244(??):127–131, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003250.

Gottschalk:2011:IGS

[GAHE11] Thomas K. Gottschalk, Birgit Aue, Stefan Hotes, and Klemens Ekschmitt. Influence of grain size on species–habitat models. Ecological Modelling, 222(18):3403–3412, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003784.

Gottlicher:2011:OPS

[GANB11] Dietrich Göttlicher, Janina Albert, Thomas Nauss, and Jörg Bendix. Optical properties of selected plants from a tropical mountain ecosystem — traits for plant functional types to parametrize a land surface model. Ecological Modelling, 222(3):493–502, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004928.

Gaglio:2019:MPP

[GAP^{+}19] M. Gaglio, V. Aschonitis, L. Pieretti, L. Santos, E. Gissi, G. Castaldelli, and E. A. Fano. Modelling past, present and future ecosystem services supply in a protected floodplain under land use and climate changes. Ecological Modelling, 403(??):23–34, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301644.

Garcia:2017:CAM

[Gar17] Oscar García. Cohort aggregation modelling for complex forest stands: Spruce–aspen mixtures in British Columbia. Ecological Modelling, 343(??):109–122, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016306160.

Gatti:2018:MWV

[Gat18] Roberto Cazzolla Gatti. Modelling weed and vine disturbance in tropical forests after selective logging and clearcutting. Ecological Modelling, 375(??):13–19, May 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300383.

Gubiani:2011:TMN

[GAV^{+}11] Éder A. Gubiani, Ronaldo Angelini, Ludgero C. G. Vieira, Luiz C. Gomes, and Angelo A. Agostinho. Trophic models in neotropical reservoirs: Testing hypotheses on the relationship between aging and maturity. Ecological Modelling, 222(23–24):3838–3848, December 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004923.

Gavrikov:2015:WRT

[Gav15] Vladimir L. Gavrikov. Whether respiration in trees can scale isometrically with bole surface area: a test of hypothesis. Ecological Modelling, 312(??):318–321, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002598.

Grunewald:2015:EAM

[GB15] Karsten Grunewald and Olaf Bastian. Ecosystem assessment and management as key tools for sustainable landscape development: a case study of the Ore Mountains region in Central Europe. Ecological Modelling, 295(??):151–162, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003949.

Gao:2016:IDU

[GB16a] Lei Gao and Brett A. Bryan. Incorporating deep uncertainty into the elementary effects method for robust global sensitivity analysis. Ecological Modelling, 321(??):1–9, February 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004949.

Grimm:2016:NGE

[GB16b] Volker Grimm and Uta Berger. Next-generation ecological modelling: a special issue dedicated to Donald DeAngelis on the occasion of his 70th birthday. Ecological Modelling, 326(??):1–3, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016000041.

Grimm:2016:RAD

[GB16c] Volker Grimm and Uta Berger. Robustness analysis: Deconstructing computational models for ecological theory and applications. Ecological Modelling, 326(??):162–167, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003245.

Grimm:2016:SRE

[GB16d] Volker Grimm and Uta Berger. Structural realism, emergence, and predictions in next-generation ecological modelling: Synthesis from a special issue. Ecological Modelling, 326(??):177–187, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016000077.

Gueguen:2011:MPS

[GBB^{+}11] Marielle Guéguen, Régis Baron, Michèle Bardouil, Philippe Truquet, Hansy Haberkorn, Patrick Lassus, Laurent Barillé, and Zouher Amzil. Modelling of paralytic shellfish toxin biotransformations in the course of Crassostrea gigas detoxification kinetics. Ecological Modelling, 222(18): 3394–3402, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003772.

Gilioli:2013:MMA

[GBB13] Gianni Gilioli, Antonella Bodini, and Johann Baumgärtner. Metapopulation modelling and area-wide pest management strategies evaluation. An application to the pine processionary moth. Ecological Modelling, 260(??):1–10, July 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001774.

Grimm:2010:OPR

[GBD^{+}10] Volker Grimm, Uta Berger, Donald L. DeAngelis, J. Gary Polhill, Jarl Giske, and Steven F. Railsback. The ODD protocol: a review and first update. Ecological Modelling, 221(23):2760–2768, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000414X.

Gaudio:2015:CEA

[GBG^{+}15] Noémie Gaudio, Salim Belyazid, Xavier Gendre, Arnaud Mansat, Manuel Nicolas, Simon Rizzetto, Harald Sverdrup, and Anne Probst. Combined effect of atmospheric nitrogen deposition and climate change on temperate forest soil biogeochemistry: a modeling approach. Ecological Modelling, 306(??):24–34, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004682.

Gordon:2013:SVC

[GBL^{+}13] Ascelin Gordon, Lucy Bastin, William T. Langford, Alex M. Lechner, and Sarah A. Bekessy. Simulating the value of collaboration in multi-actor conservation planning. Ecological Modelling, 249(??):19–25, January 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003341.

Graves:2010:ICP

[GBP^{+}10] A. R. Graves, P. J. Burgess, J. Palma, K. J. Keesman, W. van der Werf, C. Dupraz, H. van Keulen, F. Herzog, and M. Mayus. Implementation and calibration of the parameter-sparse Yield-SAFE model to predict production and land equivalent ratio in mixed tree and crop systems under two contrasting production situations in Europe. Ecological Modelling, 221(13–14):1744–1756, July 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001420.

Galli:2016:MRC

[GBP^{+}16] Giovanni Galli, Lorenzo Bramanti, Cristina Priori, Sergio Rossi, Giovanni Santangelo, Georgios Tsounis, and Cosimo Solidoro. Modelling red coral (Corallium rubrum) growth in response to temperature and nutrition. Ecological Modelling, 337(??):137–148, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302332.

Gregorini:2013:MDG

[GBR^{+}13a] Pablo Gregorini, Pierre C. Beukes, Alvaro J. Romera, Gil Levy, and Mark D. Hanigan. A model of diurnal grazing patterns and herbage intake of a dairy cow, MINDY: Model description. Ecological Modelling, 270(??):11–29, December 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004201.

Gustafsson:2013:IAH

[GBR13b] Malin S. M. Gustafsson, Mark E. Baird, and Peter J. Ralph. The interchangeability of autotrophic and heterotrophic nitrogen sources in scleractinian coral symbiotic relationships: a numerical study. Ecological Modelling, 250(??):183–194, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005285.

Gregorini:2015:DIR

[GBW^{+}15] Pablo Gregorini, Pierre Beukes, Garry Waghorn, David Pacheco, and Mark Hanigan. Development of an improved representation of rumen digesta outflow in a mechanistic and dynamic model of a dairy cow, Molly. Ecological Modelling, 313(??):293–306, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002987.

Gertseva:2011:PDS

[GC11] Vladlena V. Gertseva and Jason M. Cope. Population dynamics of splitnose rockfish (Sebastes diploproa) in the Northeast Pacific Ocean. Ecological Modelling, 222(4):973–981, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006526.

Greenwood:2014:SNM

[GC14] Jim Greenwood and Peter Craig. A simple numerical model for predicting vertical distribution of phytoplankton on the continental shelf. Ecological Modelling, 273(??):165–172, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005413.

Garcia-Callejas:2016:EMD

[GCA16] David García-Callejas and Miguel B. Araújo. The effects of model and data complexity on predictions from species distributions models. Ecological Modelling, 326(??):4–12, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002513.

Govind:2011:SDM

[GCB^{+}11] Ajit Govind, Jing Ming Chen, Pierre Bernier, Hank Margolis, Luc Guindon, and Andre Beaudoin. Spatially distributed modeling of the long-term carbon balance of a boreal landscape. Ecological Modelling, 222(15):2780–2795, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011002158.

Gamito:2010:MEE

[GCC^{+}10] Sofia Gamito, Paula Chainho, José Lino Costa, João Paulo Medeiros, Maria José Costa, and João Carlos Marques. Modelling the effects of extreme events on the dynamics of the amphipod Corophium orientale. Ecological Modelling, 221(3):459–466, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007170.

Gilardelli:2018:SWB

[GCCB18] Carlo Gilardelli, Roberto Confalonieri, Giovanni Alessandro Cappelli, and Gianni Bellocchi. Sensitivity of WOFOSTbased modelling solutions to crop parameters under climate change. Ecological Modelling, 368(??):1–14, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304453.

Gray:2012:MIS

[GCCJ12] Steven Gray, Alex Chan, Dan Clark, and Rebecca Jordan. Modeling the integration of stakeholder knowledge in social–ecological decision-making: Benefits and limitations to knowledge diversity. Ecological Modelling, 229(??):88–96, March 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004716.

Gama:2016:PGH

[GCDA16] M. Gama, D. Crespo, M. Dolbeth, and P. Anastácio. Predicting global habitat suitability for Corbicula fluminea using species distribution models: the importance of different environmental datasets. Ecological Modelling, 319(??):163–169, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002501.

Guidi:2010:RCN

[GCIB10] Lionel Guidi, Vincent Calcagno, Frédéric Ibanez, and Grégory Beaugrand. Reply to comment on “A new procedure to optimize the selection of groups in a classification tree: Applications for ecological data”. Ecological Modelling, 221(23):2739–2740, November 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004114.

Govind:2015:DME

[GCK^{+}15] Ajit Govind, Sharon Cowling, Jyothi Kumari, Nithya Rajan, and Amen Al-Yaari. Distributed modeling of ecohydrological processes at high spatial resolution over a landscape having patches of managed forest stands and crop fields in SW Europe. Ecological Modelling, 297(??):126–140, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004888.

Guo:2019:MFC

[GCL^{+}19] Chuanbo Guo, Yushun Chen, Han Liu, Yin Lu, Xiao Qu, Hui Yuan, Sovan Lek, and Songguang Xie. Modelling fish communities in relation to water quality in the impounded lakes of China’s South-to-North Water Diversion Project.

Ecological Modelling, 397(??):25–35, April 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300298.

Gurkan:2013:STD

[GCM^{+}13] Zeren Gurkan, Asbjørn Christensen, Marie Maar, Eva Friis Møller, Kristine Skovgaard Madsen, Peter Munk, and Henrik Mosegaard. Spatio-temporal dynamics of growth and survival of lesser sandeel early life-stages in the North Sea: Predictions from a coupled individual-based and hydrodynamic–biogeochemical model. Ecological Modelling, 250(??):294–306, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005467.

Guedon:2018:MYM

[GCR18] Yann Guédon, Evelyne Costes, and Miroslava Rakocevic. Modulation of the yerba-mate metamer production phenology by the cultivation system and the climatic factors. Ecological Modelling, 384(??):188–197, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302254.

Gonzalez:2015:CBM

[GCTS15] Darinka Gonzalez, João Alexandre Cabral, Laura Torres, and Mário Santos. A cohort-based modelling approach for managing olive moth Prays oleae (Bernard, 1788) populations in olive orchards. Ecological Modelling, 296(??):46–56, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004815.

Giordano:2014:FEH

[GCV14] P. Giordano, P. Caputo, and A. Vancheri. Fuzzy evaluation of heterogeneous quantities: Measuring urban ecological efficiency. Ecological Modelling, 288(??):112–126, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002695.

Gurkan:2012:GSL

[GCvDM12] Zeren Gurkan, Asbjørn Christensen, Mikael van Deurs, and Henrik Mosegaard. Growth and survival of larval and early juvenile lesser sandeel in patchy prey field in the North Sea: an examination using individual-based modeling. Ecological Modelling, 232(??):78–90, May 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200097X.

Guan:2016:ASH

[GCW^{+}16] Qingqing Guan, Jun Chen, Zhicheng Wei, Yuxia Wang, Masae Shiyomi, and Yungui Yang. Analyzing the spatial heterogeneity of number of plant individuals in grassland community by using power law model. Ecological Modelling, 320(??):316–321, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004974.

Gimenez:2019:EHN

[GDBB19] J. C. Giménez, L. Diaz-Balteiro, and M. Bertomeu. The effective hydrological neighborhood: a new concept to formulate harvest area constraints. Ecological Modelling, 404(??):83–90, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300869.

Gardarin:2011:PGR

[GDC11] Antoine Gardarin, Carolyne Dürr, and Nathalie Colbach. Prediction of germination rates of weed species: Relationships between germination speed parameters and species traits. Ecological Modelling, 222(3):626–636, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005466.

Gardarin:2012:MDE

[GDC12] Antoine Gardarin, Carolyne Dürr, and Nathalie Colbach. Modeling the dynamics and emergence of a multispecies weed seed bank with species traits. Ecological Modelling, 240(??):123–138, August 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002244.

Griffon:2014:AES

[GdC14] Sébastien Griffon and François de Coligny. AMAPstudio: an editing and simulation software suite for plants architecture modelling. Ecological Modelling, 290(??):3–10, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001300536X.

Giannetti:2019:HOS

[GDC^{+}19] Biagio F. Giannetti, Maria De Fatima D.F.B.Marcilio, Luca Coscieme, Feni Agostinho, Gengyuan Liu, and Cecilia M. V. B. Almeida. Howard Odum’s “Self-organization, transformity and information”: Three decades of empirical evidence. Ecological Modelling, 407(??):Article 108717, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302170.

Guallar:2016:ANN

[GDDFT16] Carles Guallar, Maximino Delgado, Jorge Diogène, and Margarita Fernández-Tejedor. Artificial neural network approach to population dynamics of harmful algal blooms in Alfacs Bay (NW Mediterranean): Case studies of Karlodinium and Pseudo-nitzschia. Ecological Modelling, 338(??):37–50, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016302526.

Giovanelli:2010:MSR

[GdSHA10] João G. R. Giovanelli, Marinez Ferreira de Siqueira, Célio F. B. Haddad, and João Alexandrino. Modeling a spatially restricted distribution in the neotropics: How the size of calibration area affects the performance of five presence-only methods. Ecological Modelling, 221(2):215–224, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006620.

Garnier:2014:SCM

[GDT^{+}14] Aurelie Garnier, Henri Darmency, Yann Tricault, Anne-Marie Chèvre, and Jane Lecomte. A stochastic cellular model with uncertainty analysis to assess the risk of transgene invasion after crop-wild hybridization: Oilseed rape and wild radish as a case study. Ecological Modelling, 276(??):85–94, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000386.

Ghermandi:2010:NLE

[GdTCFG10] L. Ghermandi, M. I. de Torres Curth, J. Franzese, and S. Gonzalez. Non-linear ecological processes, fires, environmental heterogeneity and shrub invasion in northwestern Patagonia. Ecological Modelling, 221(1):113–121, January 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009005262.

Giacomini:2013:TCF

[GDTP13] Henrique C. Giacomini, Donald L. DeAngelis, Joel C. Trexler, and Miguel Petrere. Trait contributions to fish community assembly emerge from trophic interactions in an individual-based model. Ecological Modelling, 251(??):32–43, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005698.

Guillaumet:2012:TOP

[GDW12] Alban Guillaumet, Brian Dorr, and Guiming Wang. Towards optimized population control efficiency in space and time: a modelling framework adapted to a colonial waterbird. Ecological Modelling, 235–236(??):95–101, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001445.

Gollany:2017:SSO

[GE17] Hero T. Gollany and Abdelhamid A. Elnaggar. Simulating soil organic carbon changes across toposequences under dryland agriculture using CQESTR. Ecological Modelling, 355(??):97–104, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305130.

Gutierrez-Estrada:2010:HAP

[GEB10] Juan C. Gutiérrez-Estrada and David T. Bilton. A heuristic approach to predicting water beetle diversity in temporary and fluctuating waters. Ecological Modelling, 221(11):1451–1462, June 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001419.

Gardner:2014:TBM

[GEEC14] Robert H. Gardner, Katharina A. M. Engelhardt, Andrew J. Elmore, and Dan Cadol. A traits-based model of species diversity. Ecological Modelling, 288(??):178–194, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002749.

Gutierrez-Estrada:2013:CFR

[GEPCB13] Juan C. Gutiérrez-Estrada, Inmaculada Pulido-Calvo, and David T. Bilton. Consistency of fuzzy rules in an ecological context. Ecological Modelling, 251(??):187–198, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005935.

Goebel:2013:MPD

[GEZ^{+}13] N. L. Goebel, C. A. Edwards, J. P. Zehr, M. J. Follows, and S. G. Morgan. Modeled phytoplankton diversity and productivity in the California Current system. Ecological Modelling, 264(??):37–47, August 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005455.

Gronlund:2015:DVE

[GFC15] Erik Grönlund, Morgan Fröling, and Inga Carlman. Donor values in emergy assessment of ecosystem services. Ecological Modelling, 306(??):101–105, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004803.

Grechi:2016:EML

[GFP^{+}16] Laura Grechi, Antonio Franco, Luca Palmeri, Alberto Pivato, and Alberto Barausse. An ecosystem model of the Lower Po River for use in ecological risk assessment of xenobiotics. Ecological Modelling, 332(??):42–58, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016300679.

Gong:2017:DEC

[GFW^{+}17] Minghao Gong, Zhiyong Fan, Junyan Wang, Gang Liu, and Chen Lin. Delineating the ecological conservation redline based on the persistence of key species: Giant pandas (Ailuropoda melanoleuca) inhabiting the qinling mountains. Ecological Modelling, 345(??):56–62, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304343.

Garreta:2012:PBC

[GGM^{+}12] V. Garreta, J. Guiot, F. Mortier, J. Chadœuf, and C. Hély. Pollen-based climate reconstruction: Calibration of the vegetation–pollen processes. Ecological Modelling, 235–236(??):81–94, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001470.

Govind:2013:ECA

[GGR^{+}13] Ajit Govind, Dominique Guyon, Jean-Louis Roujean, Nathalie Yauschew-Raguenes, Jyothi Kumari, Jan Pisek, and Jean-Pierre Wigneron. Effects of canopy architectural parameterizations on the modeling of radiative transfer mechanism. Ecological Modelling, 251(??):114–126, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012005510.

Genet:2014:IMU

[GGSP14] Astrid Genet, Pavel Grabarnik, Olga Sekretenko, and David Pothier. Incorporating the mechanisms underlying intertree competition into a random point process model to improve spatial tree pattern analysis in forestry. Ecological Modelling, 288(??):143–154, September 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002701.

Gonzalez:2017:MBT

[GGTP^{+}17] Tania Marisol González, Juan David González-Trujillo, John R. B. Palmer, Joan Pino, and Dolors Armenteras. Movement behavior of a tropical mammal: the case of Tapirus terrestris. Ecological Modelling, 360(??): 223–229, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016308195.

Gaston:2010:UCS

[GGV10] Aitor Gastón and Juan I. García-Viñas. Updating coarsescale species distribution models using small fine-scale samples. Ecological Modelling, 221(21):2576–2581, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010003716.

Gaston:2011:MSD

[GGV11] Aitor Gastón and Juan I. García-Viñas. Modelling species distributions with penalised logistic regressions: a comparison with maximum entropy models. Ecological Modelling, 222(13):2037–2041, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011002237.

Gaston:2013:EPP

[GGV13] Aitor Gastón and Juan I. García-Viñas. Evaluating the predictive performance of stacked species distribution models applied to plant species selection in ecological restoration. Ecological Modelling, 263(??):103–108, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002445.

Gras:2011:IBM

[GGVB11] Anna Gras, Marta Ginovart, Joaquim Valls, and Philippe C. Baveye. Individual-based modelling of carbon and nitrogen dynamics in soils: Parameterization and sensitivity analysis of microbial components. Ecological Modelling, 222(12):1998–2010, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001165.

Gregorius:2014:RMC

[GGZ14] Hans-Rolf Gregorius, Elizabeth M. Gillet, and Martin Ziehe. Relating measures of compositional differentiation among communities to conceptual models of migration and selection. Ecological Modelling, 279(??):24–35, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000866.

Ge:2016:SPG

[GGZ^{+}16] Zhen-Ming Ge, Hai-Qiang Guo, Bin Zhao, Chao Zhang, Heli Peltola, and Li-Quan Zhang. Spatiotemporal patterns of the gross primary production in the salt marshes with rapid community change: a coupled modeling approach. Ecological Modelling, 321(??):110–120, February 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005207.

Gao:2013:IPM

[GH13] Lei Gao and Atakelty Hailu. Identifying preferred management options: an integrated agent-based recreational fishing simulation model with an AHP-TOPSIS evaluation method. Ecological Modelling, 249(??):75–83, January 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003274.

Galic:2019:MGP

[GHD^{+}19] Nika Galic, Allyson G. Hindle, John P. DeLong, Karen Watanabe, Valery Forbes, and C. Loren Buck. Modeling genomes to phenomes to populations in a changing climate: the need for collaborative networks. Ecological Modelling, 406(??):80–83, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302042.

Goto:2015:SVF

[GHH^{+}15] Daisuke Goto, Martin J. Hamel, Jeremy J. Hammen, Matthew L. Rugg, Mark A. Pegg, and Valery E. Forbes. Spatiotemporal variation in flow-dependent recruitment of long-lived riverine fish: Model development and evaluation. Ecological Modelling, 296(??):79–92, January 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004992.

Gatti:2017:BA

[GHK17] Roberto Cazzolla Gatti, Wim Hordijk, and Stuart Kauffman. Biodiversity is autocatalytic. Ecological Modelling, 346(??):70–76, February 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305749.

Ghommem:2012:INA

[GHP12] Mehdi Ghommem, Muhammad R. Hajj, and Ishwar K. Puri. Influence of natural and anthropogenic carbon dioxide sequestration on global warming. Ecological Modelling, 235–236(??):1–7, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001718.

Grechi:2010:MCP

[GHS^{+}10] Isabelle Grechi, Nadine Hilgert, Benoît Sauphanor, Rachid Senoussi, and Françoise Lescourret. Modelling coupled peach tree–aphid population dynamics and their control by winter pruning and nitrogen fertilization. Ecological Modelling, 221(19):2363–2373, September 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000284X.

Gao:2016:SDM

[GHS^{+}16a] Wei Gao, Bongghi Hong, Dennis P. Swaney, Robert W. Howarth, and Huaicheng Guo. A system dynamics model for managing regional N inputs from human activities. Ecological Modelling, 322(??):82–91, February 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005529.

Gruss:2016:MSE

[GHS^{+}16b] Arnaud Grüss, William J. Harford, Michael J. Schirripa, Laure Velez, Skyler R. Sagarese, Yunne-Jai Shin, and Philippe Verley. Management strategy evaluation using the individual-based, multispecies modeling approach OS-MOSE. Ecological Modelling, 340(??):86–105, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304124.

Gaoue:2019:CRT

[GHST19] Orou G. Gaoue, Carol C. Horvitz, Ulrich K. Steiner, and Shripad Tuljapurkar. Climate, rather than human disturbance, is the main driver of age-specific mortality trajectories in a tropical tree. Ecological Modelling, 400(??):34–40, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019301048.

Gabsi:2014:CDM

[GHWG^{+}14] Faten Gabsi, Monika Hammers-Wirtz, Volker Grimm, Andreas Schäffer, and Thomas G. Preuss. Coupling different mechanistic effect models for capturing individualand population-level effects of chemicals: Lessons from a case where standard risk assessment failed. Ecological Modelling, 280(??):18–29, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003025.

Giannantoni:2014:RES

[Gia14] Corrado Giannantoni. The relevance of emerging solutions for thinking, decision making and acting. The case of Smart Grids. Ecological Modelling, 271(??):62–71, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001907.

Garcia:2010:ARS

[GJÁ^{+}10] Andrés García, José A. Juanes, César Álvarez, José A. Revilla, and Raúl Medina. Assesment of the response of a shallow macrotidal estuary to changes in hydrological and wastewater inputs through numerical modelling. Ecological Modelling, 221(8):1194–1208, April 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000207.

Grosman:2011:TBR

[GJB^{+}11] Paul D. Grosman, Jochen A. G. Jaeger, Pascale M. Biron, Christian Dussault, and Jean-Pierre Ouellet. Tradeoff between road avoidance and attraction by roadside salt pools in moose: an agent-based model to assess measures for reducing moose-vehicle collisions. Ecological Modelling, 222(8):1423–1435, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000548.

Groeneveld:2015:HBC

[GJK^{+}15] Jürgen Groeneveld, Karin Johst, So Kawaguchi, Bettina Meyer, Mathias Teschke, and Volker Grimm. How biological clocks and changing environmental conditions determine local population growth and species distribution in Antarctic krill (Euphausia superba): a conceptual model. Ecological Modelling, 303(??):78–86, May 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000654.

Garcia:2013:DFE

[GJM^{+}13] Tatiana Garcia, P. Ryan Jackson, Elizabeth A. Murphy, Albert J. Valocchi, and Marcelo H. Garcia. Development of a fluvial egg drift simulator to evaluate the transport and dispersion of Asian carp eggs in rivers. Ecological Modelling, 263(??):211–222, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300255X.

Ghosh:2014:SUP

[GK14] Bapan Ghosh and T. K. Kar. Sustainable use of prey species in a prey–predator system: Jointly determined ecological thresholds and economic trade-offs. Ecological Modelling, 272(??):49–58, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004420.

Gieder:2014:BNA

[GKF^{+}14] Katherina D. Gieder, Sarah M. Karpanty, James D. Fraser, Daniel H. Catlin, Benjamin T. Gutierrez, Nathaniel G. Plant, Aaron M. Turecek, and E. Robert Thieler. A Bayesian network approach to predicting nest presence of the federally-threatened piping plover (Charadrius melodus) using barrier island features. Ecological Modelling, 276(??):38–50, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000398.

Guichard:2010:MPA

[GKKW10] S. Guichard, D. J. Kriticos, J. M. Kean, and S. P. Worner. Modelling pheromone anemotaxis for biosecurity surveillance: Moth movement patterns reveal a downwind component of anemotaxis. Ecological Modelling, 221(23):2801–2807, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004254.

Glynatsi:2018:EGT

[GKL18] Nikoleta E. Glynatsi, Vincent Knight, and Tamsin E. Lee. An evolutionary game theoretic model of rhino horn devaluation. Ecological Modelling, 389(??):33–40, December 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303260.

Gong:2013:MCC

[GKW^{+}13] Jinnan Gong, Seppo Kellomäki, Kaiyun Wang, Chao Zhang, Narasinha Shurpali, and Pertti J. Martikainen. Modeling CO_{2} and CH_{4} flux changes in pristine peatlands of Finland under changing climate conditions. Ecological Modelling, 263(??):64–80, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002330.

Gecek:2010:TCC

[GL10] Sunčana Geček and Tarzan Legović. Towards carrying capacity assessment for aquaculture in the Bolinao Bay, Philippines: a numerical study of tidal circulation. Ecological Modelling, 221(10):1394–1412, May 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000761.

Garcia-Lopez:2011:MPV

[GLA11] Javier M. García-López and Carmen Allué. Modelling phytoclimatic versatility as a large scale indicator of adaptive capacity to climate change in forest ecosystems. Ecological Modelling, 222(8):1436–1447, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000585.

Guesnet:2015:IFW

[GLC^{+}15] Vanessa Guesnet, Géraldine Lassalle, Aurélie Chaalali, Kelly Kearney, Blanche Saint-Béat, Battle Karimi, Boutheina Grami, Samuele Tecchio, Nathalie Niquil, and Jérémy Lobry. Incorporating food-web parameter uncertainty into Ecopath-derived ecological network indicators. Ecological Modelling, 313(??):29–40, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002495.

Gao:2018:EPN

[GLC^{+}18] Yan Gao, Gengyuan Liu, Marco Casazza, Yan Hao, Yan Zhang, and Biagio F. Giannetti. Economy-pollution nexus model of cities at river basin scale based on multiagent simulation: a conceptual framework. Ecological Modelling, 379(??):22–38, July 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301157.

Glick:2014:MCH

[Gli14] Henry B. Glick. Modeling cougar habitat in the Northeastern United States. Ecological Modelling, 285(??):78–89, August 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014000891.

Guan:2021:MUL

[GLI^{+}21] DongJie Guan, HaiFeng Li, Takuro Inohae, Weici Su, Tadashi Nagaie, and Kazunori Hokao. Modeling urban land use change by the integration of cellular automaton and Markov model. Ecological Modelling, 222(20–22):3761–3772, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004698.

Gauzere:2019:SAT

[GLR^{+}19] Julie Gauzere, Camille Lucas, Ophélie Ronce, Hendrik Davi, and Isabelle Chuine. Sensitivity analysis of tree phenology models reveals increasing sensitivity of their predictions to winter chilling temperature and photoperiod with warming climate. Ecological Modelling, 411(??): Article 108805, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019303138.

Guo:2016:RSA

[GLRT16] Tong Guo, Dirk Lohmann, Gregor Ratzmann, and Britta Tietjen. Response of semi-arid savanna vegetation composition towards grazing along a precipitation gradient — the effect of including plant heterogeneity into an ecohydrological savanna model. Ecological Modelling, 325(??):47–56, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016000119.

Gamez:2010:OCL

[GLS10a] M. Gámez, I. López, and A. Shamandy. Open- and closed-loop equilibrium control of trophic chains. Ecological Modelling, 221(16):1839–1846, August 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002139.

Griffith:2010:HAB

[GLS10b] Cameron S. Griffith, Byron L. Long, and Jeanne M. Sept. HOMINIDS: an agent-based spatial simulation model to evaluate behavioral patterns of early Pleistocene hominids. Ecological Modelling, 221(5):738–760, March 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007558.

Gu:2017:EGD

[GLS^{+}17] Yunqian Gu, Gang Li, Yutong Sun, Weihong Luo, Xue Liu, Wei Zhang, Chunjie Qi, Yang Zhao, Kailei Tang, Yan Zhang, Liping Shao, Yan Xiong, Chuanfei Si, and Chunjiang Zhao. The effects of global dimming on the wheat crop grown in the Yangtze Basin of China simulated by SUCROS LL, a processbased model. Ecological Modelling, 350(??):42–54, April 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017301497.

Giannini:2011:ENS

[GLSA^{+}11] T. C. Giannini, R. Lira-Saade, R. Ayala, A. M. Saraiva, and I. Alves dos Santos. Ecological niche similarities of Peponapis bees and non-domesticated Cucurbita species. Ecological Modelling, 222(12):2011–2018, June 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001682.

Ghermandi:2010:IAT

[GLT10] Luciana Ghermandi, Rosa Lasaponara, and Luciano Telesca. Intra-annual time dynamical patterns of fire sequences observed in Patagonia (Argentina). Ecological Modelling, 221(1):94–97, January 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438000800570X.

Guo:2015:UEM

[GLY^{+}15] Chuanbo Guo, Sovan Lek, Shaowen Ye, Wei Li, Jiashou Liu, and Zhongjie Li. Uncertainty in ensemble modelling of large-scale species distribution: Effects from species characteristics and model techniques. Ecological Modelling, 306(??):67–75, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003792.

Gautestad:2010:SMH

[GM10] Arild O. Gautestad and Ivar Mysterud. Spatial memory, habitat auto-facilitation and the emergence of fractal home range patterns. Ecological Modelling, 221(23): 2741–2750, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004096.

Gobeyn:2019:EAS

[GMC^{+}19] Sacha Gobeyn, Ans M. Mouton, Anna F. Cord, Andrea Kaim, Martin Volk, and Peter L. M. Goethals. Evolutionary algorithms for species distribution modelling: a review in the context of machine learning. Ecological Modelling, 392(??):179–195, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304010.

Gimenez:2019:IAS

[GMK^{+}19] Olivier Gimenez, Lorena Mansilla, M. Javier Klaich, Mariano A. Coscarella, Susana N. Pedraza, and Enrique A. Crespo. Inferring animal social networks with imperfect detection. Ecological Modelling, 401(??):69–74, June 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301309.

Giusti:2010:ASD

[GMLRF10] Elisabetta Giusti, Stefano Marsili-Libelli, Monia Renzi, and Silvano Focardi. Assessment of spatial distribution of submerged vegetation in the Orbetello lagoon by means of a mathematical model. Ecological Modelling, 221(11):1484–1493, June 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001353.

Glavan:2013:FOI

[GMP13] Matjaž Glavan, Vesna Miličić, and Marina Pintar. Finding options to improve catchment water quality — lessons learned from historical land use situations in a Mediterranean catchment in Slovenia. Ecological Modelling, 261–262(??):58–73, July 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002081.

Grabarnik:2011:CTM

[GMS11] Pavel Grabarnik, Mari Myllymäki, and Dietrich Stoyan. Correct testing of mark independence for marked point patterns. Ecological Modelling, 222(23–24):3888–3894, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001100490X.

Goedegebuure:2017:BBF

[GMTC^{+}17] Merel Goedegebuure, Jessica Melbourne-Thomas, Stuart P. Corney, Mark A. Hindell, and Andrew J. Constable. Beyond big fish: the case for more detailed representations of top predators in marine ecosystem models. Ecological Modelling, 359(??):182–192, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304896.

Grossowicz:2017:DEB

[GMvV17] Michal Grossowicz, Gonçalo M. Marques, and George A. K. van Voorn. A Dynamic Energy Budget (DEB) model to describe population dynamics of the marine cyanobacterium Prochlorococcus marinus. Ecological Modelling, 359(??):320–332, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301473.

Grossel:2014:SSV

[GNB^{+}14] A. Grossel, B. Nicoullaud, H. Bourennane, P. Rochette, C. Guimbaud, M. Chartier, V. Catoire, and C. Hénault. Simulating the spatial variability of nitrous oxide emission from cropped soils at the within-field scale using the NOE model. Ecological Modelling, 288(??):155–165, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014002750.

Giam:2015:NBM

[GO15a] Xingli Giam and Julian D. Olden. A new R^{2}-based metric to shed greater insight on variable importance in artificial neural networks. Ecological Modelling, 313(??):307–313, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002835.

Gothlich:2015:DCD

[GO15b] Lena Göthlich and Andreas Oschlies. Disturbance characteristics determine the timescale of competitive exclusion in a phytoplankton model. Ecological Modelling, 296(??):126–135, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005316.

Gonzalez-Olivares:2012:CEP

[GORJ12] Eduardo González-Olivares and Rodrigo Ramos-Jiliberto. Comments to “The effect of prey refuge in a simple predator–prey model” [Ecol. Model. 222 (September (18)) (2011) 3453–3454]. Ecological Modelling, 232(??):158–160, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011006077. See [MLW11].

Grechi:2012:DIM

[GOSH^{+}12] Isabelle Grechi, Mohamed-Mahmoud Ould-Sidi, Nadine Hilgert, Rachid Senoussi, Benoît Sauphanor, and Françoise Lescourret. Designing integrated management scenarios using simulation-based and multi-objective optimization: Application to the peach tree–Myzus persicae aphid system. Ecological Modelling, 246(??):47–59, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003547.

Giannini:2013:ILS

[GPA^{+}13] T. C. Giannini, C. E. Pinto, A. L. Acosta, M. Taniguchi, A. M. Saraiva, and I. Alves dos Santos. Interactions at large spatial scale: the case of Centris bees and floral oil producing plants in South America. Ecological Modelling, 258(??):74–81, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001506.

Gurney:2014:EMP

[GPC14] Leigh J. Gurney, Evgeny A. Pakhomov, and Villy Christensen. An ecosystem model of the Prince Edward Island archipelago. Ecological Modelling, 294(??):117–136, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004232.

Girard:2015:ETS

[GPCG15] Philippe Girard, Lael Parrott, Charles-André Caron, and David M. Green. Effects of temperature and surface water availability on spatiotemporal dynamics of stream salamanders using pattern-oriented modelling. Ecological Modelling, 296(??):12–23, January 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004645.

Guber:2013:EMR

[GPD^{+}13] Andrey K. Guber, Yakov A. Pachepsky, Thanh H. Dao, Daniel R. Shelton, and Ali M. Sadeghi. Evaluating manure release parameters for nonpoint contaminant transport model KINEROS2/STWIR. Ecological Modelling, 263(??):126–138, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002585.

Gatti:2017:CBT

[GPH17] Paul Gatti, Pierre Petitgas, and Martin Huret. Comparing biological traits of anchovy and sardine in the Bay of Biscay: a modelling approach with the Dynamic Energy Budget. Ecological Modelling, 348(??):93–109, March 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304379.

Gagnon:2015:MSI

[GPJL15] Karine Gagnon, Stephanie J. Peacock, Yu Jin, and Mark A. Lewis. Modelling the spread of the invasive alga Codium fragile driven by long-distance dispersal of buoyant propagules. Ecological Modelling, 316(??):111–121, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003749.

Gaiser:2013:MBE

[GPK^{+}13] Thomas Gaiser, Ute Perkons, Paul Martin Küpper, Timo Kautz, Daniel Uteau-Puschmann, Frank Ewert, Andreas Enders, and Gunther Krauss. Modeling biopore effects on root growth and biomass production on soils with pronounced sub-soil clay accumulation. Ecological Modelling, 256(??):6–15, May 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001002.

Galan:2016:WAE

[GPM^{+}16a] E. Galán, R. Padró, I. Marco, E. Tello, G. Cunfer, G. I. Guzmán, M. González de Molina, F. Krausmann, S. Gingrich, V. Sacristán, and D. Moreno-Delgado. Widening the analysis of energy return on investment (EROI) in agroecosystems: Socio-ecological transitions to industrialized farm systems (the Vallès County, Catalonia, c. 1860 and 1999). Ecological Modelling, 336(??):13–25, September 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016301922.

Gilioli:2016:MFP

[GPM16b] Gianni Gilioli, Sara Pasquali, and Enrico Marchesini. A modelling framework for pest population dynamics and management: an application to the grape berry moth. Ecological Modelling, 320(??):348–357, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004962.

Guiet:2016:MCS

[GPM16c] Jérôme Guiet, Jean-Christophe Poggiale, and Olivier Maury. Modelling the community size-spectrum: recent developments and new directions. Ecological Modelling, 337(??):4–14, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301958.

Gruss:2019:BBB

[GPP^{+}19] Arnaud Grüss, Maria L. D. Palomares, Jorrit H. Poelen, Josephine R. Barile, Casey D. Aldemita, Shelumiel R. Ortiz, Nicolas Barrier, Yunne-Jai Shin, James Simons, and Daniel Pauly. Building bridges between global information systems on marine organisms and ecosystem models. Ecological Modelling, 398(??):1–19, April 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300432.

Goldsworthy:2013:TEG

[GPR^{+}13] Simon D. Goldsworthy, Brad Page, Paul J. Rogers, Cathy Bulman, Annelise Wiebkin, Lachlan J. McLeay, Luke Einoder, Alastair M. M. Baylis, Michelle Braley, Robin Caines, Keryn Daly, Charlie Huveneers, Kristian Peters, Andrew D. Lowther, and Tim M. Ward. Trophodynamics of the eastern Great Australian Bight ecosystem: Ecological change associated with the growth of Australia’s largest fishery. Ecological Modelling, 255(??):38–57, April 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000331.

Gilioli:2013:MLL

[GPTR13] G. Gilioli, S. Pasquali, S. Tramontini, and F. Riolo. Modelling local and long-distance dispersal of invasive chestnut gall wasp in Europe. Ecological Modelling, 263 (??):281–290, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002731.

Grant:2018:PMB

[GPZB18] Tyler J. Grant, Hazel R. Parry, Myron P. Zalucki, and Steven P. Bradbury. Predicting monarch butterfly (Danaus plexippus) movement and egg-laying with a spatially-explicit agent-based model: the role of monarch perceptual range and spatial memory. Ecological Modelling, 374(??):37–50, April 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018300644.

Garcia-Quismondo:2018:MPD

[GQNMR18] Manuel García-Quismondo, Ian C. T. Nisbet, Carolyn Mostello, and J. Michael Reed. Modeling population dynamics of roseate terns (Sterna dougallii) in the Northwest Atlantic Ocean. Ecological Modelling, 368(??):298–311, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017304945.

Garcia-Quismondo:2017:ERN

[GQRC^{+}17] Manuel García-Quismondo, J. Michael Reed, Frances S. Chew, Miguel A. Martínez del Amor, and Mario J. Pérez-Jiménez. Evolutionary response of a native butterfly to concurrent plant invasions: Simulation of population dynamics. Ecological Modelling, 360(??):410–424, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016308201.

Grant:2014:NMD

[Gra14] R. F. Grant. Nitrogen mineralization drives the response of forest productivity to soil warming: Modelling in ecosys vs. measurements from the Harvard soil heating experiment. Ecological Modelling, 288(??):38–46, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002646.

Grover:2012:MPC

[GRB12] James P. Grover, Daniel L. Roelke, and Bryan W. Brooks. Modeling of plankton community dynamics characterized by algal toxicity and allelopathy: a focus on historical Prymnesium parvum blooms in a Texas reservoir. Ecological Modelling, 227(??):147–161, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006004.

Grimaud:2014:MDC

[GRD^{+}14] Ghjuvan Micaelu Grimaud, Sophie Rabouille, Anthony Dron, Antoine Sciandra, and Olivier Bernard. Modelling the dynamics of carbon–nitrogen metabolism in the unicellular diazotrophic cyanobacterium Crocosphaera watsonii WH8501, under variable light regimes. Ecological Modelling, 291(??):121–133, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003524.

Gregorius:2012:ABV

[Gre12] Hans-Rolf Gregorius. Assessing biological variation from the perspective of diversity. Ecological Modelling, 224(1):

1–10, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011005175.

Greig:2018:STM

[GRL18] Clara Greig, Colin Robertson, and André E. B. Lacerda. Spectral-temporal modelling of bamboo-dominated forest succession in the Atlantic forest of Southern Brazil. Ecological Modelling, 384(??):316–332, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302333.

Gala:2015:DWP

[GRRU15] Alba Bala Gala, Marco Raugei, Maddalena Ripa, and Sergio Ulgiati. Dealing with waste products and flows in life cycle assessment and emergy accounting: Methodological overview and synergies. Ecological Modelling, 315(??):69–76, November 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000976.

Gonzalez-Rivero:2011:RSC

[GRYM11] Manuel González-Rivero, Laith Yakob, and Peter J. Mumby. The role of sponge competition on coral reef alternative steady states. Ecological Modelling, 222(11):1847–1853, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011001438.

Gregoire:2010:CNO

[GS10] M. Grégoire and K. Soetaert. Carbon, nitrogen, oxygen and sulfide budgets in the Black Sea: a biogeochemical model of the whole water column coupling the oxic and anoxic parts. Ecological Modelling, 221(19):2287–2301, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002760.

Golubtsov:2019:ANI

[GS19a] Peter Golubtsov and Stein Ivar Steinshamn. Analytical and numerical investigation of optimal harvest with a continuously age-structured model. Ecological Modelling, 392(??):67–81, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303983.

Gomez-Sanz:2019:SSE

[GS19b] Valentín Gómez-Sanz. Site- scale ecological marginality: Evaluation model and application to a case study. Ecological Modelling, 408(??):Article 108739, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302479.

Gupta:2019:PBF

[GS19c] Rajit Gupta and Laxmi Kant Sharma. The process-based forest growth model 3-PG for use in forest management: a review. Ecological Modelling, 397(??):55–73, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300213.

Giannetti:2019:FSS

[GSA^{+}19] Biagio F. Giannetti, Fábio Sevegnani, Cecília M. V. B. Almeida, Feni Agostinho, Roberto R. Moreno García, and Gengyuan Liu. Five sector sustainability model: a proposal for assessing sustainability of production systems. Ecological Modelling, 406(??):98–108, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302169.

Garcia:2016:NPC

[GSF^{+}16] Clement Garcia, Richard A. Stillman, Rodney M. Forster, Tiago Silva, and Julie Bremner. Nuclear power and coastal birds: Predicting the ecological consequences of warm-water outflows. Ecological Modelling, 342(??):60–81, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304628.

Garcia:2010:AEP

[GSJ^{+}10] Andrés García, María Luisa Sámano, José A. Juanes, Raúl Medina, José A. Revilla, and César Álvarez. Assessment of the effects of a port expansion on algae appearance in a costal bay through mathematical modelling. Application to San Lorenzo Bay (North Spain). Ecological Modelling, 221(10):1413–1426, May 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000736.

Garneau:2015:MTM

[GSPSP15] Cyril Garneau, S. Sauvage, A. Probst, and J. M. Sánchez-Pérez. Modelling of trace metal transfer in a large river under different hydrological conditions (the Garonne River in southwest France). Ecological Modelling, 306(??):195–204, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004268.

Gao:2017:MSM

[GSQ17] Hailong Gao, Qianyun Shi, and Xin Qian. A multispecies modelling approach to select appropriate submerged macrophyte species for ecological restoration in Gonghu Bay, Lake Taihu, China. Ecological Modelling, 360 (??):179–188, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307748.

Grueters:2014:MFD

[GSS^{+}14] U. Grueters, T. Seltmann, H. Schmidt, H. Horn, A. Pranchai, A. G. Vovides, R. Peters, J. Vogt, F. Dahdouh-Guebas, and U. Berger. The mangrove forest dynamics model meso-FON. Ecological Modelling, 291(??):28–41, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001400341X.

Gonzalez-Salazar:2013:CRC

[GSSM13] Constantino González-Salazar, Christopher R. Stephens, and Pablo A. Marquet. Comparing the relative contributions of biotic and abiotic factors as mediators of species’ distributions. Ecological Modelling, 248(??):57–70, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005078.

Grimm:2014:PME

[GT14] Volker Grimm and Pernille Thorbek. Population models for ecological risk assessment of chemicals: Short introduction and summary of a special issue. Ecological Modelling, 280(??):1–4, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400060X.

Garcia:2016:UCA

[GTC16] Elizabeth S. Garcia, Christina L. Tague, and Janet S. Choate. Uncertainty in carbon allocation strategy and ecophysiological parameterization influences on carbon and streamflow estimates for two western US forested watersheds. Ecological Modelling, 342(??):19–33, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304823.

Gowda:2019:FMR

[GTM^{+}19] Juan H. Gowda, Florencia Tiribelli, Mónica Mermoz, Thomas Kitzberger, and Juan Manuel Morales. Fragmentation modulates the response of dichotomous landscapes to fire and seed dispersal. Ecological Modelling, 392(??):22–30, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303375.

Gotelli:2011:RBN

[GU11] Nicholas J. Gotelli and Werner Ulrich. Over-reporting bias in null model analysis: a response to fayle and manica (2010). Ecological Modelling, 222(7):1337–1339, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006137.

Gledhill:2011:MET

[GV11] Michelle Gledhill and Robert W. Van Kirk. Modeling effects of toxin exposure in fish on long-term population size, with an application to selenium toxicity in bluegill (Lepomis macrochirus). Ecological Modelling, 222(19):3587–3597, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004558.

Glaum:2015:PCN

[GV15] Paul Glaum and John Vandermeer. Potential for and consequences of naturalized Bt products: Qualitative dynamics from indirect intransitivities. Ecological Modelling, 299(??):121–129, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014006127.

Guala:2010:DHM

[GVC10] Sebastián D. Guala, Flora A. Vega, and Emma F. Covelo. The dynamics of heavy metals in plant–soil interactions. Ecological Modelling, 221(8):1148–1152, April 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000244.

Geijzendorffer:2011:SDT

[GvdWBS11] I. R. Geijzendorffer, W. van der Werf, F. J. J. A. Bianchi, and R. P. O. Schulte. Sustained dynamic transience in a Lotka–Volterra competition model system for grassland species. Ecological Modelling, 222(15):2817–2824, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003097.

Garay:2016:SCC

[GVGC16] József Garay, Zoltán Varga, Manuel Gámez, and Tomas Cabello. Sib cannibalism can be adaptive for kin. Ecological Modelling, 334(??):51–59, August 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301636.

Garcia-Valdes:2021:ECS

[GVGZ^{+}21] Raúl García-Valdés, Nicholas J. Gotelli, Miguel A. Zavala, Drew W. Purves, and Miguel B. Araújo. Effects of climate, species interactions, and dispersal on decadal colonization and extinction rates of Iberian tree species. Ecological Modelling, 309–310(??):118–127, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001337.

Gielen:2016:CDA

[GVV^{+}16] Sienke Gielen, Jordi Vives i Batlle, Caroline Vincke, May Van Hees, and Hildergarde Vandenhove. Concentrations and distributions of Al, Ca, Cl, k, Mg and Mn in a Scots pine forest in Belgium. Ecological Modelling, 324(??):1–10, March 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005906.

Groen:2017:DSH

[GVV17] Thomas A. Groen, Claudius A. D. M. Van de Vijver, and Frank Van Langevelde. Do spatially homogenising and heterogenising processes affect transitions between alternative stable states? Ecological Modelling, 365 (??):119–128, December 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304234.

Gibbons:2010:SAR

[GWC^{+}10] J. M. Gibbons, A. T. A. Wood, J. Craigon, S. J. Ramsden, and N. M. J. Crout. Semi-automatic reduction and upscaling of large models: a farm management example. Ecological Modelling, 221(4):590–598, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007522.

Guo:2018:RLH

[GWF^{+}18] Tong Guo, Hanna Weise, Sebastian Fiedler, Dirk Lohmann, and Britta Tietjen. The role of landscape heterogeneity in regulating plant functional diversity under different precipitation and grazing regimes in semi-arid savannas. Ecological Modelling, 379(??):1–9, July 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301261.

Gao:2015:ABU

[GWHL15] Jie Gao, Rusong Wang, Jinlou Huang, and Min Liu. Application of BMP to urban runoff control using SUS-TAIN model: Case study in an industrial area. Ecological Modelling, 318(??):177–183, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002677.

Gong:2012:MWT

[GWK^{+}12] Jinnan Gong, Kaiyun Wang, Seppo Kellomäki, Chao Zhang, Pertti J. Martikainen, and Narasinha Shurpali. Modeling water table changes in boreal peatlands of Finland under changing climate conditions. Ecological Modelling, 244(??):65–78, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003213.

Gaikwad:2011:ENM

[GWR11] Jitendra Gaikwad, Peter D. Wilson, and Shoba Ranganathan. Ecological niche modeling of customary medicinal plant species used by Australian Aborigines to identify species-rich and culturally valuable areas for conservation. Ecological Modelling, 222(18):3437–3443, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003759.

Gu:2015:EFA

[GWZ^{+}15] Qiwei Gu, Hongqi Wang, Yinan Zheng, Jingwen Zhu, and Xiaoke Li. Ecological footprint analysis for urban agglomeration sustainability in the middle stream of the Yangtze River. Ecological Modelling, 318(??):86–99, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003282.

Guo:2013:NIF

[GYL^{+}13] Chuanbo Guo, Shaowen Ye, Sovan Lek, Jiashou Liu, Tanglin Zhang, Jin Yuan, and Zhongjie Li. The need for improved fishery management in a shallow macrophytic lake in the Yangtze River basin: Evidence from the food web structure and ecosystem analysis. Ecological Modelling, 267(??):138–147, October 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300358X.

Gao:2015:RSM

[GYL^{+}15] Yanni Gao, Guirui Yu, Shenggong Li, Huimin Yan, Xianjin Zhu, Qiufeng Wang, Peili Shi, Liang Zhao, Yingnian Li, Fawei Zhang, Yanfen Wang, and Junhui Zhang. A remote sensing model to estimate ecosystem respiration in Northern China and the Tibetan Plateau. Ecological Modelling, 304(??):34–43, May 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000940.

Ge:2011:CCC

[GZK^{+}11] Zhen-Ming Ge, Xiao Zhou, Seppo Kellomäki, Heli Peltola, and Kai-Yun Wang. Climate, canopy conductance and leaf area development controls on evapotranspiration in a boreal coniferous forest over a 10-year period: a united model assessment. Ecological Modelling, 222(9):1626–1638, May 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001037.

Gao:2018:CAN

[GZS^{+}18] Chengkang Gao, Shuaibing Zhang, Kaihui Song, Hongming Na, Fan Tian, Menghui Zhang, and Wengang Gao. Conjoint analysis of nitrogen, phosphorus and sulfur metabolism: a case study of Liaoning Province, China. Ecological Modelling, 390(??):70–78, December 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303430.

Ghisellini:2014:MES

[GZVU14] Patrizia Ghisellini, Amalia Zucaro, Silvio Viglia, and Sergio Ulgiati. Monitoring and evaluating the sustainability of Italian agricultural system. An emergy decomposition analysis. Ecological Modelling, 271(??):132–148, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000987.

Hajiahmadi:2018:FRA

[HA18] Delnia Hajiahmadi and Jamil Amanollahi. Fuzzy risk assessment modelling of wild animal life in bijar protected area. Ecological Modelling, 387(??):49–60, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302898.

Huang:2021:ANC

[HAA^{+}21] Suo Huang, M. Altaf Arain, Vivek K. Arora, Fengming Yuan, Jason Brodeur, and Matthias Peichl. Analysis of nitrogen controls on carbon and water exchanges in a conifer forest using the CLASS-CTEM ^{N+} model. Ecological Modelling, 222(20–22):3743–3760, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004686.

Hall:2011:PAS

[Hal11] Myrna H. P. Hall. A preliminary assessment of socioecological metabolism for three neighborhoods within a rust belt urban ecosystem. Ecological Modelling, 223(1):20–31, December 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004364.

Heinrichs:2017:UDP

[HAOS17] Julie A. Heinrichs, Cameron L. Aldridge, Michael S. O’Donnell, and Nathan H. Schumaker. Using dynamic population simulations to extend resource selection analyses and prioritize habitats for conservation. Ecological Modelling, 359(??):449–459, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301424.

Harizaj:2011:CES

[Har11] Pëllumb Harizaj. Can the emergy sustainability index be improved? Ecological Modelling, 222(12):2031–2033, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001451.

Hatten:2014:MMM

[Hat14] James R. Hatten. Mapping and monitoring Mount Graham red squirrel habitat with Lidar and Landsat imagery. Ecological Modelling, 289(??):106–123, October 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003317.

Hense:2010:RCL

[HB10a] Inga Hense and Aike Beckmann. The representation of cyanobacteria life cycle processes in aquatic ecosystem models. Ecological Modelling, 221(19):2330–2338, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002838.

Hense:2010:MCS

[HB10b] Inga Hense and Hans Burchard. Modelling cyanobacteria in shallow coastal seas. Ecological Modelling, 221 (2):238–244, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006140.

Harris:2012:BGB

[HB12] Lora A. Harris and Mark J. Brush. Bridging the gap between empirical and mechanistic models of aquatic primary production with the metabolic theory of ecology: an example from estuarine ecosystems. Ecological Modelling, 233(??):83–89, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001408.

Harris:2013:FCT

[HB13] Keith J. Harris and Paul G. Blackwell. Flexible continuoustime modelling for heterogeneous animal movement. Ecological Modelling, 255(??):29–37, April 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000471.

Hines:2014:CNN

[HB14] David E. Hines and Stuart R. Borrett. A comparison of network, neighborhood, and node levels of analyses in two models of nitrogen cycling in the Cape Fear River Estuary. Ecological Modelling, 293(??):210–220, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005607.

Hense:2015:TID

[HB15] Inga Hense and Aike Beckmann. A theoretical investigation of the diatom cell size reduction–restitution cycle. Ecological Modelling, 317(??):66–82, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500410X.

Huang:2016:AGT

[HBA16a] Suo Huang, Paul Bartlett, and M. Altaf Arain. An analysis of global terrestrial carbon, water and energy dynamics using the carbon–nitrogen coupled CLASS-CTEM N^{+} model. Ecological Modelling, 336(??):36–56, September 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016301995.

Huang:2016:ANC

[HBA16b] Suo Huang, Paul Bartlett, and M. Altaf Arain. Assessing nitrogen controls on carbon, water and energy exchanges in major plant functional types across North America using a carbon and nitrogen coupled ecosystem model. Ecological Modelling, 323(??):12–27, March 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005530.

Hynes:2015:IBM

[HBB15] Annette M. Hynes, Brad J. Blythe, and Brian J. Binder. An individual-based model for the analysis of Prochlorococcus diel cycle behavior. Ecological Modelling, 301(??):1–15, April 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001500023X.

Hall:2018:ISI

[HBF18] Charles A. S. Hall, Andrew S. Brainard, and Brian D. Fath. Introduction to special issue of ecological modelling on teaching systems ecology. Ecological Modelling, 388(??):10–12, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303144.

Hiltner:2016:IPV

[HBG^{+}16] Ulrike Hiltner, Achim Bräuning, Aster Gebrekirstos, Andreas Huth, and Rico Fischer. Impacts of precipitation variability on the dynamics of a dry tropical montane forest. Ecological Modelling, 320(??):92–101, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004524.

Hidy:2012:DBB

[HBH^{+}12] D. Hidy, Z. Barcza, L. Haszpra, G. Churkina, K. Pintér, and Z. Nagy. Development of the Biome-BGC model for simulation of managed herbaceous ecosystems. Ecological Modelling, 226(??):99–119, February 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100528X.

Huber:2018:GSA

[HBL18] Nica Huber, Harald Bugmann, and Valentine Lafond. Global sensitivity analysis of a dynamic vegetation model: Model sensitivity depends on successional time, climate and competitive interactions. Ecological Modelling, 368 (??):377–390, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304982.

Hartzell:2018:URC

[HBP18] Samantha Hartzell, Mark S. Bartlett, and Amilcare Porporato. Unified representation of the C3, C4, and CAM photosynthetic pathways with the Photo3 model. Ecological Modelling, 384(??):173–187, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302175.

Haller-Bull:2019:OFG

[HBR19] Vanessa Haller-Bull and Elena Rovenskaya. Optimizing functional groups in ecosystem models: Case study of the Great Barrier Reef. Ecological Modelling, 411(??): Article 108806, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001930314X.

Halouani:2016:MFW

[HBS^{+}16a] Ghassen Halouani, Frida Ben Rais Lasram, Yunne-Jai Shin, Laure Velez, Philippe Verley, Tarek Hattab, Ricardo Oliveros-Ramos, Frédéric Diaz, Frédéric Ménard, Melika Baklouti, Arnaud Guyennon, Mohamed Salah Romdhane, and François Le Loc’h. Modelling food web structure using an end-to-end approach in the coastal ecosystem of the Gulf of Gabes (Tunisia). Ecological Modelling, 339(??):45–57, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016303118.

Heinrichs:2016:HDL

[HBS16b] Julie A. Heinrichs, Darren J. Bender, and Nathan H. Schumaker. Habitat degradation and loss as key drivers of regional population extinction. Ecological Modelling, 335(??):64–73, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001630179X.

Herkt:2016:HRM

[HBSF16] K. Matthias B. Herkt, Günter Barnikel, Andrew K. Skidmore, and Jakob Fahr. A high-resolution model of bat diversity and endemism for continental Africa. Ecological Modelling, 320(??):9–28, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004184.

Herzog:2018:DPG

[HBT^{+}18] Sebastian K. Herzog, Jens Bürger, Alejandra J. Troncoso, Renzo R. Vargas, Tjalle Boorsma, and Rodrigo W. Soria-Auza. Deterministic population growth models and conservation translocation as a management strategy for the critically endangered blue-throated macaw (Ara glaucogularis): a critique of maestri et al. Ecological Modelling, 388(??):145–148, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302114.

Homa:2011:MIC

[HC11] Elizabeth S. Homa and Steven C. Chapra. Modeling the impacts of calcite precipitation on the epilimnion of an ultraoligotrophic, hard-water lake. Ecological Modelling, 222(1):76–90, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004825.

Hannah:2019:DFP

[HCE^{+}19] Lee Hannah, Christopher Costello, Vittoria Elliot, Brandon Owashi, So Nam, Rodrigo Oyanedel, Ratha Chea, Ouk Vibol, Chheng Phen, and Gavin McDonald. Designing freshwater protected areas (FPAs) for indiscriminate fisheries. Ecological Modelling, 393(??):127–134, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304174.

Chen:2018:EIC

[hCjHMW18] You hua Chen, Sun jun Huang, Ashok K. Mishra, and X. Henry Wang. Effects of input capacity constraints on food quality and regulation mechanism design for food safety management. Ecological Modelling, 385(??):89–95, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018301091.

Haider:2018:ROA

[HCK18] Zulqarnain Haider, Hadi Charkhgard, and Changhyun Kwon. A robust optimization approach for solving problems in conservation planning. Ecological Modelling, 368 (??):288–297, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304933.

He:2014:OWU

[HCL^{+}14] Liming He, Jing M. Chen, Jane Liu, Gang Mo, Stéphane Bélair, Ting Zheng, Rong Wang, Bin Chen, Holly Croft, M. Altaf Arain, and Alan G. Barr. Optimization of water uptake and photosynthetic parameters in an ecosystem model using tower flux data. Ecological Modelling, 294(??):94–104, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004438.

Heymans:2016:BPE

[HCL^{+}16] Johanna Jacomina Heymans, Marta Coll, Jason S. Link, Steven Mackinson, Jeroen Steenbeek, Carl Walters, and Villy Christensen. Best practice in ecopath with Ecosim food-web models for ecosystem-based management. Ecological Modelling, 331(??):173–184, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500575X.

Han:2015:EWU

[HCM^{+}15] M. Y. Han, G. Q. Chen, M. T. Mustafa, T. Hayat, Ling Shao, J. S. Li, X. H. Xia, and Xi Ji. Embodied water for urban economy: a three-scale input–output analysis for Beijing 2010. Ecological Modelling, 318(??):19–25, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002215.

Heinanen:2018:IMA

[HCM^{+}18] Stefan Heinänen, Magda Ewa Chudzinska, Jonas Brandi Mortensen, Theophilus Zhi En Teo, Kjell Rong Utne, Lise Doksæter Sivle, and Frank Thomsen. Integrated modelling of Atlantic mackerel distribution patterns and movements: a template for dynamic impact assessments. Ecological Modelling, 387(??):118–133, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302795.

Henry:2010:MMS

[HCP10] Mickaël Henry, Jean François Cosson, and Jean Marc Pons. Modelling multi-scale spatial variation in species richness from abundance data in a complex neotropical bat assemblage. Ecological Modelling, 221(17):2018–2027, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002632.

Hauer:2010:TBF

[HCS^{+}10] Grant Hauer, Steve Cumming, Fiona Schmiegelow, Wiktor Adamowicz, Marian Weber, and Robert Jagodzinski. Tradeoffs between forestry resource and conservation values under alternate policy regimes: a spatial analysis of the western Canadian boreal plains. Ecological Modelling, 221(21):2590–2603, October 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003534.

Halsey:2017:PPV

[HCW^{+}17] Samniqueka J. Halsey, Scott Cinel, Jared Wilson, Timothy J. Bell, and Marlin Bowles. Predicting population viability of a monocarpic perennial dune thistle using individual-based models. Ecological Modelling, 359 (??):363–371, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303277.

Han:2017:EII

[HCZ^{+}17] Dongyan Han, Yong Chen, Chongliang Zhang, Yiping Ren, Ying Xue, and Rong Wan. Evaluating impacts of intensive shellfish aquaculture on a semi-closed marine ecosystem. Ecological Modelling, 359(??):193–200, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017301849.

Hathurusingha:2014:PMT

[HD14] Priyantha I. Hathurusingha and Kenneth R. Davey. A predictive model for taste taint accumulation in recirculating aquaculture systems (RAS) farmed-fish — demonstrated with geosmin (GSM) and 2-methylisoborneol (MIB). Ecological Modelling, 291(??):242–249, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400386X.

Hathurusingha:2016:EVT

[HD16] Priyantha I. Hathurusingha and Kenneth R. Davey. Experimental validation of a time-dependent model for chemical taste taint accumulation as geosmin (GSM) and 2methylisoborneol (MIB) in commercial RAS farmed barramundi (Lates calcarifer). Ecological Modelling, 340 (??):17–27, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016303672.

Hu:2016:RSM

[HDL^{+}16] Dawei Hu, Xiaojie Du, Leyuan Li, Yi Sun, Jinhui Zhang, and Fanlu Bai. Robust stability mechanism of an artificial ecosystem based on biological mutations and synergies driven by ecological information. Ecological Modelling, 319(??):208–217, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002690.

Heikinheimo:2011:IBC

[Hei11] Outi Heikinheimo. Interactions between cod, herring and sprat in the changing environment of the Baltic Sea: a dynamic model analysis. Ecological Modelling, 222(10):1731–1742, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001128.

Humphreys:2017:BGA

[HEJP17] John M. Humphreys, James B. Elsner, Thomas H. Jagger, and Stephanie Pau. A Bayesian geostatistical approach to modeling global distributions of Lygodium microphyllum under projected climate warming. Ecological Modelling, 363(??):192–206, November 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304064.

Hellweger:2010:RCC

[Hel10] Ferdi L. Hellweger. Resonating circadian clocks enhance fitness in cyanobacteria in silico. Ecological Modelling, 221(12):1620–1629, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001602.

Hayha:2014:ESA

[HF14] Tiina Häyhä and Pier Paolo Franzese. Ecosystem services assessment: a review under an ecological-economic and systems perspective. Ecological Modelling, 289 (??):124–132, October 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003299.

He:2011:ATD

[HFB^{+}11] Guojian He, Hongwei Fang, Sen Bai, Xiaobo Liu, Minghong Chen, and Jing Bai. Application of a three-dimensional eutrophication model for the Beijing Guanting Reservoir, China. Ecological Modelling, 222(8):1491–1501, April 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006551.

Han:2019:SDC

[HFH^{+}19] Yong Han, Hongwei Fang, Lei Huang, Songheng Li, and Guojian He. Simulating the distribution of Corbicula fluminea in Lake Taihu by benthic invertebrate biomass dynamic model (BIBDM). Ecological Modelling, 409(??): Article 108730, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302303.

Henry:2012:SAH

[HFMM^{+}12] Mickaël Henry, Marie Fröchen, Julie Maillet-Mezeray, Elisabeth Breyne, Fabrice Allier, Jean-François Odoux, and Axel Decourtye. Spatial autocorrelation in honeybee foraging activity reveals optimal focus scale for predicting agro-environmental scheme efficiency. Ecological Modelling, 225(??):103–114, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005448.

Hamda:2014:SDD

[HFSL14] Natnael T. Hamda, Valery E. Forbes, John D. Stark, and Ryszard Laskowski. Stochastic density-dependent matrix model for extrapolating individual-level effects of chemicals to the population: Case study on effects of Cd on Folsomia candida. Ecological Modelling, 280(??):53–64, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004626.

Hayha:2011:EEP

[HFU11] Tiina Häyhä, Pier Paolo Franzese, and Sergio Ulgiati. Economic and environmental performance of electricity production in Finland: a multicriteria assessment framework. Ecological Modelling, 223(1):81–90, December 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004984.

Huang:2017:IEK

[HG17] Jiacong Huang and Junfeng Gao. An improved ensemble Kalman filter for optimizing parameters in a coupled phosphorus model for lowland polders in Lake Taihu Basin, China. Ecological Modelling, 357(??):14–22, August 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303830.

Herberich:2017:HNS

[HGAT17] Maximiliane Marion Herberich, Sebastian Gayler, Madhur Anand, and Katja Tielbörger. Hydrological niche segregation of plant functional traits in an individual-based model. Ecological Modelling, 356(??):14–24, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001730008X.

Holguin-Gonzalez:2013:IHP

[HGBA^{+}13] Javier E. Holguin-Gonzalez, Pieter Boets, Andres Alvarado, Felipe Cisneros, María C. Carrasco, Guido Wyseure, Ingmar Nopens, and Peter L. M. Goethals. Integrating hydraulic, physicochemical and ecological models to assess the effectiveness of water quality management strategies for the River Cuenca in Ecuador. Ecological Modelling, 254(??):1–14, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013000380.

Hermans:2010:EFE

[HGE^{+}10] C. M. L. Hermans, I. R. Geijzendorffer, F. Ewert, M. J. Metzger, P. H. Vereijken, G. B. Woltjer, and A. Verhagen. Exploring the future of European crop production in a liberalised market, with specific consideration of climate change and the regional competitiveness. Ecological Modelling, 221(18):2177–2187, September 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001663.

Huang:2013:SPU

[HGLZ13] Jiacong Huang, Junfeng Gao, Jutao Liu, and Yinjun Zhang. State and parameter update of a hydrodynamicphytoplankton model using ensemble Kalman filter. Ecological Modelling, 263(??):81–91, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002469.

Hunt:2016:VAP

[HGN^{+}16] Natalie D. Hunt, Stith T. Gower, Knute Nadelhoffer, Kate Lajtha, Kimberly Townsend, and Kristofor R. Brye. Validation of an agroecosystem process model (AGRO-BGC) on annual and perennial bioenergy feedstocks. Ecological Modelling, 321(??):23–34, February 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005141.

Haas:2015:DDN

[HGPF15] Marcelo B. Haas, Björn Guse, Matthias Pfannerstill, and Nicola Fohrer. Detection of dominant nitrate processes in ecohydrological modeling with temporal parameter sensitivity analysis. Ecological Modelling, 314(??):62–72, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003154.

Hossard:2015:LCS

[HGSJ15] Laure Hossard, Marie Gosme, Véronique Souchère, and Marie-Hélène Jeuffroy. Linking cropping system mosaics to disease resistance durability. Ecological Modelling, 307(??):1–9, July 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001271.

Hamann:2011:MFM

[HGWL11] Mark Hamann, Alana Grech, Eric Wolanski, and Jonathan Lambrechts. Modelling the fate of marine turtle hatchlings. Ecological Modelling, 222(8):1515–1521, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000603.

Han:2012:MSS

[HH12] Hongjuan Han and Weiping Hu. Modeling the species shift from Potamogeton malaianus Miq. to Potamogeton maackianus A. Bennett in the experiment pond. Ecological Modelling, 242(??):10–19, September 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002384.

Halofsky:2013:APC

[HHC^{+}13] Jessica E. Halofsky, Miles A. Hemstrom, David R. Conklin, Joshua S. Halofsky, Becky K. Kerns, and Dominique Bachelet. Assessing potential climate change effects on vegetation using a linked model approach. Ecological Modelling, 266(??):131–143, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003281.

Henrichs:2015:IBO

[HHC15] Darren W. Henrichs, Robert D. Hetland, and Lisa Campbell. Identifying bloom origins of the toxic dinoflagellate Karenia brevis in the western Gulf of Mexico using a spatially explicit individual-based model. Ecological Modelling, 313(??):251–258, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002872.

Hanberry:2012:SSM

[HHD12] B. B. Hanberry, H. S. He, and D. C. Dey. Sample sizes and model comparison metrics for species distribution models. Ecological Modelling, 227(??):29–33, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005837.

Huang:2010:SME

[HHF^{+}10] Ching-Yu Huang, Paul F. Hendrix, Timothy J. Fahey, Patrick J. Bohlen, and Peter M. Groffman. A simulation model to evaluate the impacts of invasive earthworms on soil carbon dynamics. Ecological Modelling, 221 (20):2447–2457, October 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001000311X.

Hinners:2019:MPA

[HHK19] Jana Hinners, Inga Hense, and Anke Kremp. Modelling phytoplankton adaptation to global warming based on resurrection experiments. Ecological Modelling, 400(??):27–33, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019301036.

He:2017:EIL

[HHL17] Jianhua He, Junlong Huang, and Chun Li. The evaluation for the impact of land use change on habitat quality: a joint contribution of cellular automata scenario simulation and habitat quality assessment model. Ecological Modelling, 366(??):58–67, December 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304222.

Hu:2019:RPE

[HhLY^{+}19] Wen Hu, Chun hua Li, Chun Ye, Ji Wang, Wei wei Wei, and Yong Deng. Research progress on ecological models in the field of water eutrophication: CiteSpace analysis based on data from the ISI Web of Science database. Ecological Modelling, 410(??):Article 108779, October 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930287X.

Hafs:2014:QRW

[HHUD14] Andrew W. Hafs, Lee R. Harrison, Ryan M. Utz, and Thomas Dunne. Quantifying the role of woody debris in providing bioenergetically favorable habitat for juvenile salmon. Ecological Modelling, 285(??):30–38, August 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001963.

Hjelkrem:2017:SAB

[HHvO^{+}17] Anne-Grete Roer Hjelkrem, Mats Höglind, Marcel van Oijen, Jürgen Schellberg, Thomas Gaiser, and Frank Ewert. Sensitivity analysis and Bayesian calibration for testing robustness of the BASGRA model in different environments. Ecological Modelling, 359(??):80–91, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016307281.

Herrmann:2011:IAM

[HHW11] Isabelle Herrmann, Torsten Herrmann, and Sven Wagner. Improvements in anisotropic models of single tree effects in Cartesian coordinates. Ecological Modelling, 222(7):1333–1336, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011000287.

He:2010:STT

[HJ10a] Hongming He and C. Y. Jim. Simulation of thermodynamic transmission in green roof ecosystem. Ecological Modelling, 221(24):2949–2958, December 15, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004576.

Holmes:2010:MCR

[HJ10b] G. Holmes and R. W. Johnstone. Modelling coral reef ecosystems with limited observational data. Ecological Modelling, 221(8):1173–1183, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000499.

Han:2017:SSD

[HJ17] Yu Han and Haifeng Jia. Simulating the spatial dynamics of urban growth with an integrated modeling approach: a case study of Foshan, China. Ecological Modelling, 353(??):107–116, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301028.

Hof:2012:UEP

[HJN12] Anouschka R. Hof, Roland Jansson, and Christer Nilsson. The usefulness of elevation as a predictor variable in species distribution modelling. Ecological Modelling, 246(??):86–90, November 10, 2012. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003778.

He:2016:FCN

[HJS^{+}16] Hongxing He, Per-Erik Jansson, Magnus Svensson, Astrid Meyer, Leif Klemedtsson, and Åsa Kasimir. Factors controlling nitrous oxide emission from a spruce forest ecosystem on drained organic soil, derived using the CoupModel. Ecological Modelling, 321(??):46–63, February 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005153.

Hallstan:2012:CCT

[HJWG12] Simon Hallstan, Richard K. Johnson, Eva Willén, and Ulf Grandin. Comparison of classification-then-modelling and species-by-species modelling for predicting lake phytoplankton assemblages. Ecological Modelling, 231(??):11–19, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000385.

Hopton:2017:STE

[HKG^{+}17] Matthew E. Hopton, Arunprakash T. Karunanithi, Ahjond S. Garmestani, Denis White, Jerry R. Choate, and Heriberto Cabezas. A supplementary tool to existing approaches for assessing ecosystem community structure. Ecological Modelling, 355(??):64–69, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017300844.

Han:2011:ITN

[HKL^{+}11] Jeong-Ho Han, Hema K. Kumar, Jae Hoon Lee, Chang-Ik Zhang, Se-Wha Kim, Jung-Ho Lee, Sang Don Kim, and Kwang-Guk An. Integrative trophic network assessments of a lentic ecosystem by key ecological approaches of water chemistry, trophic guilds, and ecosystem health assessments along with an ECOPATH model. Ecological Modelling, 222(19):3457–3472, October 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004182.

Helser:2014:MEF

[HKlL14] Thomas E. Helser, Craig R. Kastelle, and Han lin Lai. Modeling environmental factors affecting assimilation of bomb-produced Δ ^{14}C in the North Pacific Ocean: Implications for age validation studies. Ecological Modelling, 277(??):108–118, April 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000453.

Hodgson:2018:CSV

[HKM^{+}18] Emma E. Hodgson, Isaac C. Kaplan, Kristin N. Marshall, Jerry Leonard, Timothy E. Essington, D. Shallin Busch, Elizabeth A. Fulton, Chris J. Harvey, Albert J. Hermann, and Paul McElhany. Consequences of spatially variable ocean acidification in the California Current: Lower pH drives strongest declines in benthic species in southern regions while greatest economic impacts occur in northern regions. Ecological Modelling, 383(??):106–117, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018301856.

Holmstrom:2017:MBS

[HKN17] Emma Holmström, Matts Karlsson, and Urban Nilsson. Modeling birch seed supply and seedling establishment during forest regeneration. Ecological Modelling, 352(??):31–39, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016304811.

Hovestadt:2010:IPM

[HKP10] Thomas Hovestadt, Alexander Kubisch, and Hans-Joachim Poethke. Information processing in models for densitydependent emigration: a comparison. Ecological Modelling, 221(3):405–410, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007510.

Huang:2016:NLM

[HKVD16] Minyi Huang, Xiaoquan Kong, Sara Varela, and Renyan Duan. The Niche Limitation Method (NicheLim), a new algorithm for generating virtual species to study biogeography. Ecological Modelling, 320(??):197–202, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004615.

Hesse:2013:CSA

[HKVR13] Cornelia Hesse, Valentina Krysanova, Tobias Vetter, and Julia Reinhardt. Comparison of several approaches representing terrestrial and in-stream nutrient retention and decomposition in watershed modelling. Ecological Modelling, 269(??):70–85, November 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004171.

Hall:2018:CES

[HKW^{+}18] Charles A. S. Hall, Frances Knickmeyer, Adrian Wiegman, Andrew Brainard, Avriel Rose Diaz, Carolyn Huynh, and Jerry Mead. A class exercise for systems ecology: Synthesis of stream energetics and testing Allen’s paradox. Ecological Modelling, 369(??):42–65, February 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303332.

Holt:2019:LVF

[HL19] Johnson Holt and Adrian W. Leach. Linguistic variables as fuzzy sets to model uncertainty in the combined efficacy of multiple phytosanitary measures in pest risk analysis. Ecological Modelling, 406(??):73–79, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301620.

Hillary:2012:MLS

[HLK^{+}12] Richard M. Hillary, Polina Levontin, Sakari Kuikka, Samu Manteniemi, Iago Mosqueira, and Laurie Kell. Multi-level stock–recruit analysis: Beyond steepness and into model uncertainty. Ecological Modelling, 242(??):69–80, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012002542.

Han:2019:DSI

[HLL^{+}19] Zhi-Quan Han, Tong Liu, Hua-Feng Liu, Xiao-Ran Hao, Wei Chen, and Bai-Lian Li. Derivation of species interactions strength in a plant community with game theory. Ecological Modelling, 394(??):27–33, February 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304290.

Helser:2012:BHM

[HlLB12] Thomas E. Helser, Han lin Lai, and Bryan A. Black. Bayesian hierarchical modeling of Pacific geoduck growth increment data and climate indices. Ecological Modelling, 247(??):210–220, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004358.

Holst:2013:CIM

[HLLO13a] Niels Holst, Andreas Lang, Gabor Lövei, and Mathias Otto. Corrigendum to “Increased mortality is predicted of Inachis io larvae caused by Bt-maize pollen in European farmland” [Ecol. Model. 250 (2013) 126–133]. Ecological Modelling, 265(??):250, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003165.

Holst:2013:IMP

[HLLO13b] Niels Holst, Andreas Lang, Gabor Lövei, and Mathias Otto. Increased mortality is predicted of Inachis io larvae caused by Bt-maize pollen in European farmland. Ecological Modelling, 250(??):126–133, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005315.

Herbei:2010:PPI

[HLLP^{+}10] Radu Herbei, W. Berry Lyons, Johanna Laybourn-Parry, Christopher Gardner, John C. Priscu, and Diane M. McKnight. Physiochemical properties influencing biomass abundance and primary production in Lake Hoare, Antarctica.

Ecological Modelling, 221(8):1184–1193, April 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008618.

Han:2014:ACI

[HLS^{+}14] Zhi-Quan Han, Tong Liu, QinMing Sun, Ru Li, Jiang-Bo Xie, and Bai-Lian Li. Application of compound interest laws in biology: Reunification of existing models to develop seed bank dynamics model of annual plants. Ecological Modelling, 278(??):67–73, April 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000672.

Han:2015:UES

[HLW15] Baolong Han, Hongxiao Liu, and Rusong Wang. Urban ecological security assessment for cities in the Beijing–tianjin– hebei metropolitan region based on fuzzy and entropy methods. Ecological Modelling, 318(??):217–225, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014006280.

Han:2018:GDW

[HLZ^{+}18] Qifei Han, Chaofan Li, Chengyi Zhao, Yaoqi Zhang, and Shoubo Li. Grazing decreased water use efficiency in Central Asia from 1979 to 2011. Ecological Modelling, 388(??):72–79, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001830317X.

Heikkinen:2010:THS

[HM10] Juha Heikkinen and Raisa Mäkipää. Testing hypotheses on shape and distribution of ecological response curves. Ecological Modelling, 221(3):388–399, February 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007169.

Hiebeler:2012:QST

[HM12] David E. Hiebeler and Isaac J. Michaud. Quantifying spatial and temporal variability of spatially correlated disturbances. Ecological Modelling, 240(??):64–73, August 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001652.

Holloway:2017:QSM

[HM17] Paul Holloway and Jennifer A. Miller. A quantitative synthesis of the movement concepts used within species distribution modelling. Ecological Modelling, 356(??):91–103, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016307517.

Halsey:2018:SAB

[HM18a] Samniqueka J. Halsey and James R. Miller. A spatial agent-based model of the disease vector Ixodes scapularis to explore host-tick associations. Ecological Modelling, 387(??):96–106, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303016.

Hu:2018:DEU

[HM18b] Guangwen Hu and Xianzhong Mu. Dominants in evolution of urban energy metabolism: a case study of Beijing. Ecological Modelling, 385(??):26–34, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302436.

Halsey:2019:CSA

[HM19] Samniqueka J. Halsey and James R. Miller. Corrigendum to “A spatial agent-based model of the disease vector Ixodes scapularis to explore host-tick associations” [Ecol. Model. 387 (2018) 96–106]. Ecological Modelling, 393(??):152, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304307.

Hoch:2010:IHM

[HMA10a] T. Hoch, Y. Monnet, and A. Agoulon. Influence of host migration between woodland and pasture on the population dynamics of the tick Ixodes ricinus: a modelling approach. Ecological Modelling, 221(15):1798–1806, July 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002103.

Hossain:2010:EEA

[HMA10b] Md. Monir Hossain, Takashi Matsuishi, and George Arhonditsis. Elucidation of ecosystem attributes of an oligotrophic lake in Hokkaido, Japan, using Ecopath with Ecosim (EwE). Ecological Modelling, 221(13–14):1717–1730, July 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010001705.

Halvorsen:2016:HIC

[HMD^{+}16] Rune Halvorsen, Sabrina Mazzoni, John Wirkola Dirksen, Erik Næsset, Terje Gobakken, and Mikael Ohlson. How important are choice of model selection method and spatial autocorrelation of presence data for distribution modelling by MaxEnt? Ecological Modelling, 328(??):108–118, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016300503.

Habel:2019:NIB

[HMP19a] Henrike Häbel, Mari Myllymäki, and Arne Pommerening. New insights on the behaviour of alternative types of individual-based tree models for natural forests. Ecological Modelling, 406(??):23–32, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300821.

Hamda:2019:ASE

[HMP^{+}19b] Natnael T. Hamda, Benjamin Martin, Jamilynn B. Poletto, Dennis E. Cocherell, Nann A. Fangue, Joel Van Eenennaam, Ethan A. Mora, and Eric Danner. Applying a simplified energy-budget model to explore the effects of temperature and food availability on the life history of green sturgeon (Acipenser medirostris). Ecological Modelling, 395(??):1–10, March 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019300195.

Habib:2011:MLE

[HMPC11] Thomas J. Habib, Evelyn H. Merrill, M. J. Pybus, and David W. Coltman. Modelling landscape effects on density– contact rate relationships of deer in eastern Alberta: Implications for chronic wasting disease. Ecological Modelling, 222(15):2722–2732, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002687.

Hagemann:2010:CMM

[HMS^{+}10] Ulrike Hagemann, Martin T. Moroni, Cindy H. Shaw, Werner A. Kurz, and Franz Makeschin. Comparing measured and modelled forest carbon stocks in highboreal forests of harvest and natural-disturbance origin in Labrador, Canada. Ecological Modelling, 221 (5):825–839, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008114.

Hashimoto:2011:SMS

[HMS^{+}11] Shoji Hashimoto, Tomoaki Morishita, Tadashi Sakata, Shigehiro Ishizuka, Shinji Kaneko, and Masamichi Takahashi. Simple models for soil CO_{2}, CH_{4}, and N_{2}O fluxes calibrated using a Bayesian approach and multi-site data. Ecological Modelling, 222(7):1283–1292, April 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000457.

Hasumi:2014:MGC

[HN14] Hiroyasu Hasumi and Toshi Nagata. Modeling the global cycle of marine dissolved organic matter and its influence on marine productivity. Ecological Modelling, 288 (??):9–24, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002580.

Huang:2010:MNP

[HNWT10] Ni Huang, Zheng Niu, Chaoyang Wu, and Michelle Coreena Tappert. Modeling net primary production of a fastgrowing forest using a light use efficiency model. Ecological Modelling, 221(24):2938–2948, December 15, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004552.

Hoang:2011:MDC

[Hoa11] Viet-Ngu Hoang. Measuring and decomposing changes in agricultural productivity, nitrogen use efficiency and cumulative exergy efficiency: Application to OECD agriculture. Ecological Modelling, 222(1):164–175, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005193.

Honkaniemi:2018:IDA

[HOKH18] Juha Honkaniemi, Risto Ojansuu, Risto Kasanen, and Kari Heliövaara. Interaction of disturbance agents on Norway spruce: a mechanistic model of bark beetle dynamics integrated in simulation framework WINDROT. Ecological Modelling, 388(??):45–60, November 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303107.

Holtsmark:2013:BFM

[Hol13] Bjart Holtsmark. Boreal forest management and its effect on atmospheric CO_{2}. Ecological Modelling, 248 (??):130–134, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005066.

Hosokawa:2011:MML

[Hos11] Shinya Hosokawa. A model for measuring leaf growth rate of vegetative shoot in the seagrass, Zostera marina. Ecological Modelling, 222(1):105–111, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004862.

Hruska:2012:LTF

[HOŠ^{+}12] J. Hruška, F. Oulehle, P. Šamonil, J. Šebesta, K. Tahovská, R. Hleb, J. Houška, and J. Šikl. Long-term forest soil acidification, nutrient leaching and vegetation development: Linking modelling and surveys of a primeval spruce forest in the Ukrainian Transcarpathian Mountains. Ecological Modelling, 244(??):28–37, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003043.

Hoover:2013:EHFa

[HPC13a] Carie Hoover, Tony Pitcher, and Villy Christensen. Effects of hunting, fishing and climate change on the Hudson Bay marine ecosystem: I. Re-creating past changes 1970– 2009. Ecological Modelling, 264(??):130–142, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000859.

Hoover:2013:EHFb

[HPC13b] Carie Hoover, Tony Pitcher, and Villy Christensen. Effects of hunting, fishing and climate change on the Hudson Bay marine ecosystem: II. Ecosystem model future projections. Ecological Modelling, 264(??):143–156, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000379.

Hochard:2010:IMS

[HPG^{+}10] Sébastien Hochard, Christel Pinazo, Christian Grenz, Jessica L. Burton Evans, and Olivier Pringault. Impact of microphytobenthos on the sediment biogeochemical cycles: a modeling approach. Ecological Modelling, 221 (13–14):1687–1701, July 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001870.

Henschke:2018:MLC

[HPGM18] Natasha Henschke, Evgeny A. Pakhomov, Jürgen Groeneveld, and Bettina Meyer. Modelling the life cycle of Salpa thompsoni. Ecological Modelling, 387(??):17–26, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302862.

Hamilton:2015:HSM

[HPJ15] Serena H. Hamilton, Carmel A. Pollino, and Anthony J. Jakeman. Habitat suitability modelling of rare species using Bayesian networks: Model evaluation under limited data. Ecological Modelling, 299(??):64–78, March 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006103.

Hermoso:2015:PCR

[HPO^{+}15] Virgilio Hermoso, Francis Pantus, Jon Olley, Simon Linke, James Mugodo, and Patrick Lea. Prioritising catchment rehabilitation for multi objective management: an application from SE-Queensland, Australia. Ecological Modelling, 316(??):168–175, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003804.

Hochard:2012:BPC

[HPRNP12] Sébastien Hochard, Christel Pinazo, Emma Rochelle-Newall, and Olivier Pringault. Benthic pelagic coupling in a shallow oligotrophic ecosystem: Importance of microphytobenthos and physical forcing. Ecological Modelling, 247(??):307–318, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004048.

Hatten:2010:MDH

[HPS10] James R. Hatten, Eben H. Paxton, and Mark K. Sogge. Modeling the dynamic habitat and breeding population of southwestern willow flycatcher. Ecological Modelling, 221 (13–14):1674–1686, July 10, 2010. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001845.

Huntzinger:2012:NAC

[HPW^{+}12] D. N. Huntzinger, W. M. Post, Y. Wei, A. M. Michalak, T. O. West, A. R. Jacobson, I. T. Baker, J. M. Chen, K. J. Davis, D. J. Hayes, F. M. Hoffman, A. K. Jain, S. Liu, A. D. McGuire, R. P. Neilson, Chris Potter, B. Poulter, David Price, B. M. Raczka, H. Q. Tian, P. Thornton, E. Tomelleri, N. Viovy, J. Xiao, W. Yuan, N. Zeng, M. Zhao, and R. Cook. North American carbon program (NACP) regional interim synthesis: Terrestrial biospheric model intercomparison. Ecological Modelling, 232(??):144–157, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000725.

Huang:2017:MEI

[HQF^{+}17] Yunxin Huang, Yun Qin, Hongqiang Feng, Peng Wan, and Zhaohua Li. Modeling the evolution of insect resistance to one- and two-toxin Bt-crops in spatially heterogeneous environments. Ecological Modelling, 347(??):72–84, March 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017300066.

Hoeppner:2011:IBM

[HR11] Susanne S. Hoeppner and Kenneth A. Rose. Individualbased modeling of flooding and salinity effects on a coastal swamp forest. Ecological Modelling, 222(19):3541–3558, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004340.

Hainz-Renetzeder:2015:APS

[HRSKW15] C. Hainz-Renetzeder, A. Schneidergruber, M. Kuttner, and T. Wrbka. Assessing the potential supply of landscape services to support ecological restoration of degraded landscapes: a case study in the Austrian–Hungarian trans-boundary region of Lake Neusiedl. Ecological Modelling, 295(??):196–206, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003287.

Horn:2012:TMW

[HRWT12] S. Horn, A. Raabe, H. Will, and O. Tackenberg. Turb-Seed — a model for wind dispersal of seeds in turbulent currents based on publicly available climate data. Ecological Modelling, 237–238(??):1–10, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001755.

Haywood:2011:MEF

[HS11] Andrew Haywood and Christine Stone. Mapping eucalypt forest susceptible to dieback associated with bell miners (Manorina melanophys) using laser scanning, SPOT 5 and ancillary topographical data. Ecological Modelling, 222(5):1174–1184, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006782.

Haythorne:2013:IPG

[HS13a] Sean Haythorne and Andrew Skabar. An improved pattern-guided evolution approach for the development of adaptive individual-based ecological models. Ecological Modelling, 252(??):72–82, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004413.

Heinle:2013:IPC

[HS13b] A. Heinle and T. Slawig. Impact of parameter choice on the dynamics of NPZD type ecosystem models. Ecological Modelling, 267(??):93–101, October 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003748.

Heinle:2013:IDN

[HS13c] A. Heinle and T. Slawig. Internal dynamics of NPZD type ecosystem models. Ecological Modelling, 254(??):33–42, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013000392.

Hossain:2014:DSG

[HSCS14] M. Shahadat Hossain, Subrata Sarker, Sayedur Rahman Chowdhury, and S. M. Sharifuzzaman. Discovering spawning ground of Hilsa shad (Tenualosa ilisha) in the coastal waters of Bangladesh. Ecological Modelling, 282(??):59–68, June 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001400115X.

Huntingford:2015:CAN

[HSD^{+}15] Chris Huntingford, D. Mark Smith, William J. Davies, Richard Falk, Stephen Sitch, and Lina M. Mercado. Combining the [ABA] and net photosynthesis-based model equations of stomatal conductance. Ecological Modelling, 300(??):81–88, March 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000174.

Hedger:2013:MCL

[HSHF^{+}13] R. D. Hedger, L. E. Sundt-Hansen, T. Forseth, O. H. Diserud, O. Ugedal, and A. G. Finstad. Modelling the complete life-cycle of Atlantic salmon (Salmo salar L.) using a spatially explicit individual-based approach. Ecological Modelling, 248(??):119–129, January 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005030.

Hararuk:2017:COM

[HSK17] Oleksandra Hararuk, Cindy Shaw, and Werner A. Kurz. Constraining the organic matter decay parameters in the CBM-CFS3 using Canadian National Forest Inventory data and a Bayesian inversion technique. Ecological Modelling, 364(??):1–12, November 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300418.

Hallgren:2019:SDM

[HSLC^{+}19] W. Hallgren, F. Santana, S. Low-Choy, Y. Zhao, and B. Mackey. Species distribution models can be highly sensitive to algorithm configuration. Ecological Modelling, 408(??):Article 108719, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302194.

Hong:2012:ERV

[HSM^{+}12] Bongghi Hong, Dennis P. Swaney, Carl-Magnus Mörth, Erik Smedberg, Hanna Eriksson Hägg, Christoph Humborg, Robert W. Howarth, and Fayçal Bouraoui. Evaluating regional variation of net anthropogenic nitrogen and phosphorus inputs (NANI/NAPI), major drivers, nutrient retention pattern and management implications in the multinational areas of Baltic Sea basin. Ecological Modelling, 227(??):117–135, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005849.

Hetherington:2015:MCC

[HSR^{+}15] Amy Lee Hetherington, Rebecca L. Schneider, Lars G. Rudstam, Gideon Gal, Arthur T. DeGaetano, and M. Todd Walter. Modeling climate change impacts on the thermal dynamics of polymictic Oneida Lake, New York, United States. Ecological Modelling, 300(??):1–11, March 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006310.

Hatten:2019:SMS

[HSTS19] James R. Hatten, Gary L. Slater, Jerrmaine L. Treadwell, and Matthew R. Stevenson. A spatial model of streaked horned lark breeding habitat in the Columbia River, USA. Ecological Modelling, 409(??):Article 108734, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302406.

Holm:2012:GMD

[HSVL12] Jennifer A. Holm, H. H. Shugart, S. J. Van Bloem, and G. R. Larocque. Gap model development, validation, and application to succession of secondary subtropical dry forests of Puerto Rico. Ecological Modelling, 233(??):70–82, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001214.

Husson:2018:MIH

[HSvO^{+}18] Bérengère Husson, Jozée Sarrazin, Dick van Oevelen, Pierre-Marie Sarradin, Karline Soetaert, and Alain Menesguen. Modelling the interactions of the hydrothermal mussel Bathymodiolus azoricus with vent fluid. Ecological Modelling, 377(??):35–50, June 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830098X.

Hudson:2014:AUE

[HT14] Amy Hudson and David R. Tilley. Assessment of uncertainty in emergy evaluations using Monte Carlo simulations. Ecological Modelling, 271(??):52–61, January 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002810.

Heymans:2016:RSN

[HT16] Johanna J. Heymans and Maciej T. Tomczak. Regime shifts in the Northern Benguela ecosystem: Challenges for management. Ecological Modelling, 331(??):151–159, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005128.

Hefley:2013:FPG

[HTB13] Trevor J. Hefley, Andrew J. Tyre, and Erin E. Blankenship. Fitting population growth models in the presence of measurement and detection error. Ecological Modelling, 263(??):244–250, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002536.

Hefley:2017:RFP

[HTB17] Trevor J. Hefley, Andrew J. Tyre, and Erin E. Blankenship. Reprint of: Fitting population growth models in the presence of measurement and detection error. Ecological Modelling, 359(??):461–467, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004882.

Hazlerigg:2014:PRT

[HTL^{+}14] Charles R. E. Hazlerigg, Charles R. Tyler, Kai Lorenzen, James R. Wheeler, and Pernille Thorbek. Population relevance of toxicant mediated changes in sex ratio in fish: an assessment using an individual-based zebrafish (Danio rerio) model. Ecological Modelling, 280(??):76–88, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013006030.

Hoch:2018:BGP

[HTVE18] Thierry Hoch, Suzanne Touzeau, Anne-France Viet, and Pauline Ezanno. Between-group pathogen transmission: From processes to modeling. Ecological Modelling, 383 (??):138–149, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018301832.

Hu:2016:RML

[Hu16] Weiping Hu. A review of the models for Lake Taihu and their application in lake environmental management. Ecological Modelling, 319(??):9–20, January 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003348.

Hui:2011:FPT

[Hui11] Cang Hui. Forecasting population trend from the scaling pattern of occupancy. Ecological Modelling, 222 (3):442–446, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005806.

Hellmann:2012:ICC

[HV12] Fritz Hellmann and Jan E. Vermaat. Impact of climate change on water management in Dutch peat polders. Ecological Modelling, 240(??):74–83, August 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002256.

Harmackova:2015:MRE

[HV15] Zuzana V. Harmáčková and David Vačkář. Modelling regulating ecosystem services trade-offs across landscape scenarios in Třeboňsko Wetlands Biosphere Reserve, Czech Republic. Ecological Modelling, 295(??):207–215, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004694.

Hoglind:2016:PBS

[HVCP16] Mats Höglind, Marcel Van Oijen, David Cameron, and Tomas Persson. Process-based simulation of growth and overwintering of grassland using the BASGRA model. Ecological Modelling, 335(??):1–15, September 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301661.

Hancock:2019:OFM

[HVL^{+}19] Joana Hancock, Sara Vieira, Hipólito Lima, Vanessa Schmitt, Jaconias Pereira, Rui Rebelo, and Marc Girondot. Overcoming field monitoring restraints in estimating marine turtle internesting period by modelling individual nesting behaviour using capture-mark-recapture data. Ecological Modelling, 402(??):76–84, June 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301589.

Herben:2012:CLE

[HW12] Tomáš Herben and Radka Wildová. Community-level effects of plant traits in a grassland community examined by multispecies model of clonal plant growth. Ecological Modelling, 234(??):60–69, June 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003607.

Henzler:2018:SSF

[HWE^{+}18] Julia Henzler, Hanna Weise, Neal J. Enright, Susanne Zander, and Britta Tietjen. A squeeze in the suitable fire interval: Simulating the persistence of fire-killed plants in a Mediterranean-type ecosystem under drier conditions. Ecological Modelling, 389(??):41–49, December 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303338.

Hu:2018:RSC

[HWH^{+}18] Dawei Hu, Kai Wang, Jingfei Hu, Xinming Xu, and Yufei Long. Robust stability of closed artificial ecosystem cultivating cabbage realized by ecological thermodynamics and dissipative structure system. Ecological Modelling, 380(??):1–7, July 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018301315.

Hou:2017:RBP

[HWL17] Jian Hou, Huiqing Wang, and Zongshan Li. Relationship between plants and soil resource patterns on forest land at different scales using a new theoretical model. Ecological Modelling, 349(??):33–40, April 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630326X.

Hashimoto:2011:NSI

[HWS11] S. Hashimoto, M. Wattenbach, and P. Smith. A new scheme for initializing process-based ecosystem models by scaling soil carbon pools. Ecological Modelling, 222 (19):3598–3602, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004285.

Hao:2014:AID

[HYG14] Lu Hao, Li-Zhe Yang, and Jing-Min Gao. The application of information diffusion technique in probabilistic analysis to grassland biological disasters risk. Ecological Modelling, 272(??):264–270, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300481X.

Huang:2017:PBW

[HYL^{+}17] Yongmei Huang, Xiaona Yu, Engui Li, Huiying Chen, Liu Li, Xiuchen Wu, and Xiaoyan Li. A process-based water balance model for semi-arid ecosystems: a case study of psammophytic ecosystems in Mu Us Sandland, Inner Mongolia, China. Ecological Modelling, 353(??):77–85, June 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017300121.

Han:2017:SEE

[HZSH17] Tianfang Han, Chuntao Zhang, Yan Sun, and Xiaomin Hu. Study on environment-economy-society relationship model of Liaohe River Basin based on multi-agent simulation. Ecological Modelling, 359(??):135–145, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303490.

He:2013:MPE

[HZT^{+}13] Wei He, Yanyan Zhang, Rong Tian, Hongxing Hu, Bin Chen, Lam Kin Chen, and Fuliu Xu. Modeling the purification effects of the constructed Sphagnum wetland on phosphorus and heavy metals in Dajiuhu Wetland Reserve, China. Ecological Modelling, 252(??):23–31, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005108.

Hao:2018:MSA

[HZZ^{+}18] Yan Hao, Menghui Zhang, Yan Zhang, Chenling Fu, and Zhongming Lu. Multi-scale analysis of the energy metabolic processes in the Beijing–Tianjin–Hebei (Jing–Jin–Ji) urban agglomeration. Ecological Modelling, 369(??):66–76, February 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017304210.

Innangi:2019:SBC

[IBD^{+}19] Michele Innangi, Rosario Balestrieri, Tiziana Danise, Francesco d’Alessandro, and Antonietta Fioretto. From soil to bird community: a partial least squares approach to investigate a natural wooded area surrounded by urban patchwork (Astroni crater, southern Italy). Ecological Modelling, 394(??):1–10, February 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930002X.

Iturbide:2015:FSD

[IBH^{+}15] Maialen Iturbide, Joaquín Bedia, Sixto Herrera, Oscar del Hierro, Miriam Pinto, and Jose Manuel Gutiérrez. A framework for species distribution modelling with improved pseudo-absence generation. Ecological Modelling, 312(??):166–174, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001500215X.

Ibarra:2014:CDE

[IFC14] Diego A. Ibarra, Katja Fennel, and John J. Cullen. Coupling 3-d Eulerian bio-physics (ROMS) with individualbased shellfish ecophysiology (SHELL-E): a hybrid model for carrying capacity and environmental impacts of bivalve aquaculture. Ecological Modelling, 273(??):63–78, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005000.

Izquierdo-Gomez:2016:MPD

[IGBSASSJ16] David Izquierdo-Gomez, Just T. Bayle-Sempere, Francisco Arreguín-Sánchez, and Pablo Sánchez-Jerez. Modeling population dynamics and small-scale fisheries yields of fish farming escapes in Mediterranean coastal areas. Ecological Modelling, 331(??):56–67, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016000314.

Inomata:2018:SSS

[IGR^{+}18] Sandrelly Oliveira Inomata, Alba Maria Guadalupe Orellana Gonzalez, Rodrigo Máximo Sánchez Román, Lucirene Aguiar de Souza, and Carlos Edwar de Carvalho Freitas. Sustainability of small-scale fisheries in the middle Negro River (Amazonas — Brazil): a model with operational and biological variables. Ecological Modelling, 368(??):312–320, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302739.

Bogosian:2012:HRN

[IHSM12] Victor Bogosian III, Eric C. Hellgren, Michael W. Sears, and Raymond W. Moody. High-resolution niche models via a correlative approach: Comparing and combining correlative and process-based information. Ecological Modelling, 237–238(??):63–73, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001901.

Iwata:2011:SPT

[IKH^{+}11] Shigehide Iwata, Kazuyuki Kobayashi, Shinichiro Higa, Jin Yoshimura, and Kei ichi Tainaka. A simple population theory for mutualism by the use of lattice gas model. Ecological Modelling, 222(13):2042–2048, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002171.

Islam:2012:NMT

[IKK^{+}12] Md. Nazrul Islam, Daisuke Kitazawa, Naoki Kokuryo, Shigeru Tabeta, Takamitsu Honma, and Nobuyuki Komatsu. Numerical modeling on transition of dominant algae in Lake Kitaura, Japan. Ecological Modelling, 242 (??):146–163, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002347.

Isomaa:2013:BCG

[IKL13] Marleena Isomaa, Veijo Kaitala, and Jouni Laakso. Baltic cod (Gadus morhua callarias) recovery potential under different environment and fishery scenarios. Ecological Modelling, 266(??):118–125, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002998.

Isomaa:2014:DII

[IKL14] Marleena Isomaa, Veijo Kaitala, and Jouni Laakso. Determining the impact of initial age structure on the recovery of a healthy over-harvested population. Ecological Modelling, 286(??):45–52, August 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001975.

Izaurralde:2017:SMD

[IMW^{+}17] R. César Izaurralde, William B. McGill, Jimmy R. Williams, Curtis D. Jones, Robert P. Link, David H. Manowitz, D. Elisabeth Schwab, Xuesong Zhang, G. Philip Robertson, and Neville Millar. Simulating microbial denitrification with EPIC: Model description and evaluation. Ecological Modelling, 359(??):349–362, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303265.

Ireland:2016:SAB

[IN16] Timothy Ireland and Rachael Miller Neilan. A spatial agentbased model of feral cats and analysis of population and nuisance controls. Ecological Modelling, 337(??):123–136, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001630237X.

Ingwersen:2010:UCE

[Ing10] Wesley W. Ingwersen. Uncertainty characterization for emergy values. Ecological Modelling, 221(3):445–452, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009007182.

Ibrahim:2014:CIR

[IPSH14] Lara Ibrahim, Thomas G. Preuss, Andreas Schaeffer, and Udo Hommen. A contribution to the identification of representative vulnerable fish species for pesticide risk assessment in Europe — a comparison of population resilience using matrix models. Ecological Modelling, 280(??):65–75, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013003827.

Ives:2013:BBE

[IS13] Matthew C. Ives and James P. Scandol. BIOMAS: a bio-economic modelling and assessment system for fisheries management strategy evaluation. Ecological Modelling, 249(??):42–49, January 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003316.

Ives:2013:BEM

[ISG13] M. C. Ives, J. P. Scandol, and J. Greenville. A bioeconomic management strategy evaluation for a multispecies, multi-fleet fishery facing a world of uncertainty. Ecological Modelling, 256(??):69–84, May 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000513.

Ihde:2017:AMS

[IT17] Thomas F. Ihde and Howard M. Townsend. Accounting for multiple stressors influencing living marine resources in a complex estuarine ecosystem using an Atlantis model. Ecological Modelling, 365(??):1–9, December 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308614.

Ikegami:2012:ELD

[IWW12] Makihiko Ikegami, Dennis F. Whigham, and Marinus J. A. Werger. Effects of local density of clonal plants on their sexual and vegetative propagation strategies in a lattice structure model. Ecological Modelling, 234(??):51–59, June 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001421.

Javidi:2015:DAT

[JA15a] Mohammad Javidi and Bashir Ahmad. Dynamic analysis of time fractional order phytoplankton–toxic phytoplankton– zooplankton system. Ecological Modelling, 318(??):8–18, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380015002653.

Jonsson:2015:MCT

[JA15b] Anders Jonsson and Sebastian Agerberg. Modelling of E. coli transport in an oligotrophic river in northern Scandinavia. Ecological Modelling, 306(??):145–151, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004906.

Jager:2014:TOC

[Jag14] Henriette I. Jager. Thinking outside the channel: Timing pulse flows to benefit salmon via indirect pathways. Ecological Modelling, 273(??):117–127, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005437.

Jonard:2010:MFF

[JAH^{+}10] Mathieu Jonard, Laurent Augusto, Emmanuel Hanert, David L. Achat, Mark R. Bakker, Christian Morel, Alain Mollier, and Sylvain Pellerin. Modeling forest floor contribution to phosphorus supply to maritime pine seedlings in two-layered forest soils. Ecological Modelling, 221(6):927–935, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008631.

Juston:2010:UAC

[JAKJ10] John Juston, Olof Andrén, Thomas Kätterer, and Per-Erik Jansson. Uncertainty analyses for calibrating a soil carbon balance model to agricultural field trial data in Sweden and Kenya. Ecological Modelling, 221(16):1880–1888, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002358.

James:2016:RLO

[Jam16] R. Thomas James. Recalibration of the Lake Okeechobee Water Quality Model (LOWQM) to extreme hydrometeorological events. Ecological Modelling, 325(??):71–83, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016000260.

Junior:2018:MDG

[JASW18] Antonio F. Nogueira Júnior, Lilian Amorim, Serge Savary, and Laetitia Willocquet. Modelling the dynamics of grapevine growth over years. Ecological Modelling, 369 (??):77–87, February 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304994.

Jaremo:2011:ILH

[JB11] Johannes Järemo and Göran Bengtsson. On the importance of life history and age structure in biological invasions. Ecological Modelling, 222(3):485–492, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006174.

James:2011:ASM

[JBB^{+}11] Alex James, Richard Brown, Britta Basse, Graeme W. Bourdôt, Shona L. Lamoureaux, Mick Roberts, and David J. Saville. Application of a spatial meta-population model with stochastic parameters to the management of the invasive grass Nassella trichotoma in North Canterbury, New Zealand. Ecological Modelling, 222(4):1030–1037, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001000640X.

Jing:2012:RSP

[JBB^{+}12] Qi Jing, Gilles Bélanger, Vern Baron, Helge Bonesmo, Perttu Virkajärvi, and David Young. Regrowth simulation of the perennial grass timothy. Ecological Modelling, 232(??):64–77, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012000920.

Joo:2011:OAN

[JBCÑ11] Rocío Joo, Sophie Bertrand, Alexis Chaigneau, and Miguel Ñiquen. Optimization of an artificial neural network for identifying fishing set positions from VMS data: an example from the Peruvian anchovy purse seine fishery. Ecological Modelling, 222(4):1048–1059, February 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004539.

Jamshidi:2015:UHA

[JBE^{+}15] Shahrad Jamshidi, Jocelyn E. Behm, Damien Eveillard, E. Toby Kiers, and Philippe Vandenkoornhuyse. Using hybrid automata modelling to study phenotypic plasticity and allocation strategies in the plant mycorrhizal mutualism. Ecological Modelling, 311(??):11–19, September 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001726.

Jager:2014:DEB

[JBH^{+}14] Tjalling Jager, Alpar Barsi, Natnael T. Hamda, Benjamin T. Martin, Elke I. Zimmer, and Virginie Ducrot. Dynamic energy budgets in population ecotoxicology: Applications and outlook. Ecological Modelling, 280(??):140–147, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003190.

Jacobs:2015:MRR

[JBV^{+}15] Sander Jacobs, Benjamin Burkhard, Toon Van Daele, Jan Staes, and Anik Schneiders. ‘The matrix reloaded’: a review of expert knowledge use for mapping ecosystem services.

Ecological Modelling, 295(??):21–30, January 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004141.

Johnston:2017:IEM

[JBW^{+}17] John M. Johnston, M. Craig Barber, Kurt Wolfe, Mike Galvin, Mike Cyterski, and Rajbir Parmar. An integrated ecological modeling system for assessing impacts of multiple stressors on stream and riverine ecosystem services within river basins. Ecological Modelling, 354(??):104–114, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001630446X.

Jiang:2011:IUE

[JC11a] M. M. Jiang and B. Chen. Integrated urban ecosystem evaluation and modeling based on embodied cosmic exergy. Ecological Modelling, 222(13):2149–2165, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002201.

Ju:2011:EEE

[JC11b] L. P. Ju and B. Chen. Embodied energy and emergy evaluation of a typical biodiesel production chain in China. Ecological Modelling, 222(14):2385–2392, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003765.

Jiang:2015:POM

[JCKG15] Weimin Jiang, Chris Cornelisen, Ben Knight, and Mark Gibbs. A pattern-oriented model for assessing effects of weather and freshwater discharge on black coral (Antipathes fiordensis) distribution in a fjord. Ecological Modelling, 304(??):59–68, May 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000861.

Ju:2010:NFL

[JD10] Shu Ju and Donald L. DeAngelis. Nutrient fluxes at the landscape level and the R∗ rule. Ecological Modelling, 221(2):141–146, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006401.

Jager:2018:CIL

[JD18] Henriette I. Jager and Donald L. DeAngelis. The confluences of ideas leading to, and the flow of ideas emerging from, individual-based modeling of riverine fishes. Ecological Modelling, 384(??):341–352, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302187.

Juston:2013:MPC

[JDGJ13] John M. Juston, Thomas A. DeBusk, Kevin A. Grace, and Scott D. Jackson. A model of phosphorus cycling to explore the role of biomass turnover in submerged aquatic vegetation wetlands for Everglades restoration. Ecological Modelling, 251(??):135–149, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005674.

Jones:2012:MCF

[JDP^{+}12] Miranda C. Jones, Stephen R. Dye, John K. Pinnegar, Rachel Warren, and William W. L. Cheung. Modelling commercial fish distributions: Prediction and assessment using different approaches. Ecological Modelling, 225(??):133–145, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005217.

Jiang:2016:DNG

[JDT^{+}16] Jiang Jiang, Donald L. DeAngelis, Su-Yean Teh, Ken W. Krauss, Hongqing Wang, Haidong Li, Thomas J. Smith, and Hock-Lye Koh. Defining the next generation modeling of coastal ecotone dynamics in response to global change.

Ecological Modelling, 326(??):168–176, April 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001647.

Jiang:2017:MPI

[JDT^{+}17] Weiguo Jiang, Yue Deng, Zhenghong Tang, Xuan Lei, and Zheng Chen. Modelling the potential impacts of urban ecosystem changes on carbon storage under different scenarios by linking the CLUE-S and the InVEST models. Ecological Modelling, 345(??):30–40, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307554.

Jiang:2014:PAI

[JDZC14] Jiang Jiang, Donald L. DeAngelis, Bo Zhang, and Joel E. Cohen. Population age and initial density in a patchy environment affect the occurrence of abrupt transitions in a birth-and-death model of Taylor’s law. Ecological Modelling, 289(??):59–65, October 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003044.

Jensen:2012:RDP

[Jen12] Christopher X. Jon Jensen. Ratio dependent predator– prey theory: Aged, mellowed, and distilled. Ecological Modelling, 246(??):103–104, November 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003754.

Jiang:2012:OHP

[JFS^{+}12] Liping Jiang, Xing Fang, Heinz G. Stefan, Peter C. Jacobson, and Donald L. Pereira. Oxythermal habitat parameters and identifying cisco refuge lakes in Minnesota under future climate scenarios using variable benchmark periods. Ecological Modelling, 232(??):14–27, May 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000907.

Jankowski:2010:CMS

[JG10] Paweł Jankowski and Dariusz Gozdowski. Canonical modelling of spring barley growth using dry matter weight data from a field experiment. Ecological Modelling, 221 (2):161–172, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006322.

Jones:2015:MEW

[JGB15] Timothy Jones, Jonathan P. A. Gardner, and James J. Bell. Modelling the effect of wave forces on subtidal macroalgae: a spatial evaluation of predicted disturbance for two habitatforming species. Ecological Modelling, 313(??):149–161, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002756.

Jakoby:2014:POP

[JGF14] Oliver Jakoby, Volker Grimm, and Karin Frank. Patternoriented parameterization of general models for ecological application: Towards realistic evaluations of management approaches. Ecological Modelling, 275(??):78–88, March 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005887.

Junguo:2013:UIB

[JGX13] Hu Junguo, Zhou Guomo, and Xu Xiaojun. Using an improved back propagation neural network to study spatial distribution of sunshine illumination from sensor network data. Ecological Modelling, 266(??):86–96, September 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003220.

Jia:2012:MTS

[JHH^{+}12] Peiqiao Jia, Menghong Hu, Zhongjun Hu, Qigen Liu, and Zhen Wu. Modeling trophic structure and energy flows in a typical macrophyte dominated shallow lake using the mass balanced model. Ecological Modelling, 233(??):26–30, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001068.

Johnson:2017:IWY

[JHHS17] Matthew J. Johnson, James R. Hatten, Jennifer A. Holmes, and Patrick B. Shafroth. Identifying western yellow-billed cuckoo breeding habitat with a dual modelling approach. Ecological Modelling, 347(??):50–62, March 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307980.

Johnston:2014:EBA

[JHT^{+}14] A. S. A. Johnston, M. E. Hodson, P. Thorbek, T. Alvarez, and R. M. Sibly. An energy budget agent-based model of earthworm populations and its application to study the effects of pesticides. Ecological Modelling, 280(??):5–17, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004419.

Ji:2015:TPU

[Ji15] Xi Ji. Taking the pulse of urban economy: From the perspective of systems ecology. Ecological Modelling, 318 (??):36–48, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000988.

Jeong:2011:WHM

[JJKJ11] Kwang-Seuk Jeong, Ji-Deok Jang, Dong-Kyun Kim, and Gea-Jae Joo. Waterfowls habitat modeling: Simulation of nest site selection for the migratory little tern (Sterna albifrons) in the Nakdong estuary. Ecological Modelling, 222(17):3149–3156, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003139.

Johnson:2014:URV

[JJMW14] Fred A. Johnson, Gitte H. Jensen, Jesper Madsen, and Byron K. Williams. Uncertainty, robustness, and the value of information in managing an expanding Arctic goose population. Ecological Modelling, 273(??):186–199, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005309.

Jones:2010:MIT

[JJS10] Dominic Jones, Henrik Jeldtoft Jensen, and Paolo Sibani. Mutual information in the tangled nature model. Ecological Modelling, 221(3):400–404, February 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007236.

Jin:2019:ELU

[JJYM19] Xin Jin, Yanxiang Jin, Donghai Yuan, and Xufeng Mao. Effects of land-use data resolution on hydrologic modelling, a case study in the upper reach of the Heihe River, Northwest China. Ecological Modelling, 404(??):61–68, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300808.

Janssen:2010:AUE

[JKS^{+}10] J. A. E. B. Janssen, M. S. Krol, R. M. J. Schielen, A. Y. Hoekstra, and J.-L. de Kok. Assessment of uncertainties in expert knowledge, illustrated in fuzzy rule-based models. Ecological Modelling, 221(9):1245–1251, May 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000505.

Jonsson:2018:ECS

[JL18] Anna Maria Jönsson and Fredrik Lagergren. Effects of climate and soil conditions on the productivity and defence capacity of Picea abies in Sweden — an ecosystem model assessment. Ecological Modelling, 384(??):154–167, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001830228X.

Jorgensen:2010:FEI

[JLN10] Sven Erik Jørgensen, Alessandro Ludovisi, and Søren Nors Nielsen. The free energy and information embodied in the amino acid chains of organisms. Ecological Modelling, 221(19):2388–2392, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002693.

Jiang:2018:PUS

[JLZ^{+}18] Long Jiang, Yiping Li, Xu Zhao, Martin R. Tillotson, Wencai Wang, Shuangshuang Zhang, Linda Sarpong, Qhtan Asmaa, and Baozhu Pan. Parameter uncertainty and sensitivity analysis of water quality model in Lake Taihu, China. Ecological Modelling, 375(??):1–12, May 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300735.

Janzen:2015:SAM

[JM15] Daryl Janzen and Jeffrey J. McDonnell. A stochastic approach to modelling and understanding hillslope runoff connectivity dynamics. Ecological Modelling, 298 (??):64–74, February 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003068.

Johnston:2011:IMF

[JMB^{+}11] John M. Johnston, Daniel J. McGarvey, M. Craig Barber, Gerry Laniak, Justin Babendreier, Rajbir Parmar, Kurt Wolfe, Stephen R. Kraemer, Michael Cyterski, Chris Knightes, Brenda Rashleigh, Luis Suarez, and Robert Ambrose. An integrated modeling framework for performing environmental assessments: Application to ecosystem services in the Albemarle–Pamlico basins (NC and VA, USA). Ecological Modelling, 222(14):2471–2484, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001736.

Johnson:2010:MCR

[JMdW^{+}10] Sandra Johnson, Kerrie Mengersen, Alta de Waal, Kelly Marnewick, Deon Cilliers, Ann Marie Houser, and Lorraine Boast. Modelling cheetah relocation success in southern Africa using an iterative Bayesian network development cycle. Ecological Modelling, 221(4):641–651, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007947.

Jimenez-Munoz:2012:RTS

[JMSM12] Juan C. Jiménez-Muñoz, José A. Sobrino, and Cristian Mattar. Recent trends in solar exergy and net radiation at global scale. Ecological Modelling, 228(??):59–65, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011006156.

Jia:2011:CME

[JMW11] Haifeng Jia, Hongtao Ma, and Mingjie Wei. Calculation of the minimum ecological water requirement of an urban river system and its deployment: a case study in Beijing central region. Ecological Modelling, 222(17):3271–3276, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011003061.

Jorgensen:2014:UEE

[JN14] S. E. Jørgensen and S. N. Nielsen. Use of eco-exergy in ecological networks. Ecological Modelling, 293(??): 202–209, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002440.

Jorgensen:2015:HN

[JN15a] S. E. Jørgensen and S. N. Nielsen. Hierarchical networks. Ecological Modelling, 295(??):59–65, January 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002944.

Jorgensen:2015:CCM

[JN15b] Sven Erik Jørgensen and Søren Nors Nielsen. A carbon cycling model developed for the renewable Energy Danish Island, Samsø. Ecological Modelling, 306(??):106–120, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014002725.

Jutras:2011:DOC

[JNC^{+}11] Marie-France Jutras, Mina Nasr, Mark Castonguay, Christopher Pit, Joseph H. Pomeroy, Todd P. Smith, Cheng fu Zhang, Charles D. Ritchie, Fan-Rui Meng, Thomas A. Clair, and Paul A. Arp. Dissolved organic carbon concentrations and fluxes in forest catchments and streams: DOC-3 model. Ecological Modelling, 222(14):2291–2313, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001724.

Jorgensen:2016:RPS

[JNF16] Sven E. Jørgensen, Søren Nors Nielsen, and Brian D. Fath. Recent progress in systems ecology. Ecological Modelling, 319(??):112–118, January 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003701.

Jager:2018:SEB

[JNHF18] Tjalling Jager, Raymond Nepstad, Bjørn Henrik Hansen, and Julia Farkas. Simple energy-budget model for yolkfeeding stages of Atlantic cod (Gadus morhua). Ecological Modelling, 385(??):213–219, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302667.

Johnson:2010:IDL

[Joh10] Leah R. Johnson. Implications of dispersal and life history strategies for the persistence of linyphiid spider populations. Ecological Modelling, 221(8):1138–1147, April 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008722.

Johanson:2017:SSE

[JOHW17] Arne N. Johanson, Andreas Oschlies, Wilhelm Hasselbring, and Boris Worm. SPRAT: a spatially-explicit marine ecosystem model based on population balance equations. Ecological Modelling, 349(??):11–25, April 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300704.

Jorgensen:2010:PTS

[Jør10a] Sven Erik Jørgensen. Preface to the thematic section on Lake and Wetlands Models. Ecological Modelling, 221(4):652–653, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008345.

Jorgensen:2010:RRD

[Jør10b] Sven Erik Jørgensen. A review of recent developments in lake modelling. Ecological Modelling, 221(4):689–692, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438000900708X.

Jorgensen:2011:SMH

[Jör11a] Sven Erik Jörgensen. Surface Modeling: High Accuracy and High Speed Methods. CRC Press, published February (2011), 711 pp., hardbound, price $129.95. Ecological Modelling, 222(17):3300, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003309.

Jorgensen:2011:E

[Jor11b] Sven Erik Jorgensen. Editorial. Ecological Modelling, 222(16):2867–2868, August 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011002900.

Jorgensen:2014:BRD

[Jør14a] Sven Erik Jørgensen. Book review:Discovery of the Life-Organizing Principle: In Search of the Fundamentals Laws of Life, E. M. Elsheik. iUniverse LLC, Bloomington, USA (2014), 217 pp., paper back $19.99, hard cover $29.99, Ebook $3.99. Ecological Modelling, 291(??):151, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003834.

Jorgensen:2014:FES

[Jør14b] Sven Erik Jørgensen. Fundamentals of Ecosystem Science, K. C. Weathers, D. L. Strayer, G. E. Likens. Academic Press and Elsevier (2013), 312 pp., Paperback, Price: $61.42, color figures. Ecological Modelling, 291(??):182, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003846.

Jorgensen:2015:ENM

[Jør15a] Sven Erik Jørgensen. Erratum to “New method to calculate the work energy of information and organisms” [Ecol. Model. 295 (2015) 18–20]. Ecological Modelling, 302(??):54, April 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015000150.

Jorgensen:2015:NMC

[Jør15b] Sven Erik Jørgensen. New method to calculate the work energy of information and organisms. Ecological Modelling, 295(??):18–20, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004153.

Jorgensen:2016:E

[Jør16] Sven Erik Jørgensen. Editorial. Ecological Modelling, 319(??):3–8, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015004998.

Jager:2016:MGN

[JR16] Tjalling Jager and Elisa Ravagnan. Modelling growth of northern krill (Meganyctiphanes norvegica) using an energybudget approach. Ecological Modelling, 325(??):28–34, April 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016000089.

Jordan:2012:CRP

[JS12] Ferenc Jordán and Marco Scotti. From C:P ratios to polar meta-ecosystems. Ecological Modelling, 245(??):1–2, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012004073.

Jian:2014:EFD

[JSB^{+}14] Yun Jian, Sonia Silvestri, Enrica Belluco, Andrea Saltarin, Giovanni Chillemi, and Marco Marani. Environmental forcing and density-dependent controls of Culex pipiens abundance in a temperate climate (Northeastern Italy). Ecological Modelling, 272(??):301–310, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004869.

Jimenez:2019:PMF

[JSCS19] Laura Jiménez, Jorge Soberón, J. Andrés Christen, and Desireé Soto. On the problem of modeling a fundamental niche from occurrence data. Ecological Modelling, 397(??):74–83, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019300407.

Jose-Silva:2018:IVE

[JSdSdL^{+}18] Leandro José-Silva, Reginaldo Carvalho dos Santos, Bruna Martins de Lima, Mendelson Lima, José Francisco de Oliveira-Júnior, Paulo Eduardo Teodoro, Pedro V. Eisenlohr, and Carlos Antonio da Silva Junior. Improving the validation of ecological niche models with remote sensing analysis. Ecological Modelling, 380(??):22–30, July 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301303.

Jager:2015:CLH

[JSH15] Tjalling Jager, Iurgi Salaberria, and Bjørn Henrik Hansen. Capturing the life history of the marine copepod Calanus sinicus into a generic bioenergetics framework. Ecological Modelling, 299(??):114–120, March 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006176.

Jablonska-Sabuka:2015:DME

[JSKK15] Matylda Jabłońska-Sabuka, Rishee Kalaria, and Tuomo Kauranne. A dynamical model for epidemic outbursts by Begomovirus population clusters. Ecological Modelling, 297(??):60–68, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005663.

Janse:2010:ECP

[JSL^{+}10] J. H. Janse, M. Scheffer, L. Lijklema, L. Van Liere, J. S. Sloot, and W. M. Mooij. Estimating the critical phosphorus loading of shallow lakes with the ecosystem model PCLake: Sensitivity, calibration and uncertainty. Ecological Modelling, 221(4):654–665, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009005250.

Jonzen:2013:SBA

[JSW^{+}13] Niclas Jonzén, Håkan Sand, Petter Wabakken, Jon E. Swenson, Jonas Kindberg, Olof Liberg, and Guillaume Chapron. Sharing the bounty — adjusting harvest to predator return in the Scandinavian human–wolf–bear–moose system. Ecological Modelling, 265(??):140–148, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002809.

Janssen:2019:PPB

[JTB^{+}19] Annette B. G. Janssen, Sven Teurlincx, Arthur H. W. Beusen, Mark A. J. Huijbregts, Jasmijn Rost, Aafke M. Schipper, Laura M. S. Seelen, Wolf M. Mooij, and Jan H. Janse. PCLake+: a process-based ecological model to assess the trophic state of stratified and non-stratified freshwater lakes worldwide. Ecological Modelling, 396(??):23–32, March 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300201.

Jarnevich:2017:MEM

[JTM^{+}17] Catherine S. Jarnevich, Marian Talbert, Jeffery Morisette, Cameron Aldridge, Cynthia S. Brown, Sunil Kumar, Daniel Manier, Colin Talbert, and Tracy Holcombe. Minimizing effects of methodological decisions on interpretation and prediction in species distribution studies: an example with background selection. Ecological Modelling, 363(??):48–56, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016308213.

Ju:2019:RBD

[Ju19] Yiyi Ju. Revealing the bilateral dependencies and policy implication of food production of Japan and China: From the perspective of food-energy-water nexus. Ecological Modelling, 391(??):29–39, January 10, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303636.

Jacome:2019:PFI

[JVY19] Gabriel Jácome, Paulina Vilela, and ChangKyoo Yoo. Present and future incidence of dengue fever in Ecuador nationwide and coast region scale using species distribution modeling for climate variability’s effect. Ecological Modelling, 400(??):60–72, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301176.

Jones:2017:MMI

[JWLS17] Hannah Jones, Andrew White, Peter Lurz, and Craig Shuttleworth. Mathematical models for invasive species management: Grey squirrel control on Anglesey. Ecological Modelling, 359(??):276–284, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300364.

Jia:2018:RTC

[JWX18] Binghao Jia, Yuanyuan Wang, and Zhenghui Xie. Responses of the terrestrial carbon cycle to drought over China: Modeling sensitivities of the interactive nitrogen and dynamic vegetation. Ecological Modelling, 368(??):52–68, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304647.

Jeger:2015:MDP

[JX15] M. J. Jeger and X. M. Xu. Modelling the dynamics of a plant pathogen and a biological control agent in relation to flowering pattern and populations present on leaves. Ecological Modelling, 313(??):13–28, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002641.

Jiang:2017:WES

[JX17] Long Jiang and Meng Xia. Wind effects on the spring phytoplankton dynamics in the middle reach of the Chesapeake Bay. Ecological Modelling, 363(??):68–80, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016305713.

Jiang:2015:BMA

[JXL^{+}15] Long Jiang, Meng Xia, Stuart A. Ludsin, Edward S. Rutherford, Doran M. Mason, Jose Marin Jarrin, and Kevin L. Pangle. Biophysical modeling assessment of the drivers for plankton dynamics in dreissenid-colonized western Lake Erie. Ecological Modelling, 308(??):18–33, July 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001349.

Jager:2012:SDE

[JZ12] Tjalling Jager and Elke I. Zimmer. Simplified Dynamic Energy Budget model for analysing ecotoxicity data. Ecological Modelling, 225(??):74–81, January 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005412.

Jamali-Zghal:2014:MRA

[JZCL14] N. Jamali-Zghal, O. Le Corre, and B. Lacarrière. Mineral resource assessment: Compliance between emergy and exergy respecting Odum’s hierarchy concept. Ecological Modelling, 272(??):208–219, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004468.

Jia:2010:DMS

[JZG10] Haifeng Jia, Yansong Zhang, and Yu Guo. The development of a multi-species algal ecodynamic model for urban surface water systems and its application. Ecological Modelling, 221(15):1831–1838, July 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002115.

Kazanci:2017:EUT

[KA17] C. Kazanci and M. R. Adams. Ecological utility theory: Solving a series convergence issue. Ecological Modelling, 358(??):19–24, August 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307049.

Kirch:2012:BTM

[KAC^{+}12a] P. V. Kirch, G. Asner, O. A. Chadwick, J. Field, T. Ladefoged, C. Lee, C. Puleston, S. Tuljapurkar, and P. M. Vitousek. Building and testing models of long-term agricultural intensification and population dynamics: a case study from the Leeward Kohala Field System, Hawai’i. Ecological Modelling, 227(??):18–28, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005825.

Kirch:2012:RBT

[KAC^{+}12b] P. V. Kirch, G. Asner, O. A. Chadwick, J. Field, T. Ladefoged, C. Lee, C. Puleston, S. Tuljapurkar, and P. M. Vitousek. Reprint: Building and testing models of long-term agricultural intensification and population dynamics: a case study from the Leeward Kohala Field System, Hawai’i. Ecological Modelling, 241(??):54–64, August 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003067.

Kint:2014:ETM

[KAF^{+}14] V. Kint, W. Aertsen, N. M. Fyllas, A. Trabucco, E. Janssen, K. Özkan, and B. Muys. Ecological traits of Mediterranean tree species as a basis for modelling forest dynamics in the Taurus mountains, Turkey. Ecological Modelling, 286(??):53–65, August 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002130.

Kamp:2019:MME

[KAJØ19] Andreas Kamp, Morten Ambye-Jensen, and Hanne Østergård. Modelling matter and energy flows of local, refined grassclover protein feed as alternative to imported soy meal. Ecological Modelling, 410(??):Article 108738, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302467.

Kaliszewicz:2019:VHR

[Kal19] Anita Kaliszewicz. Variations of hydra reproductive strategies arising from its modular structure. Two aspects of the modular reproductive effect. Ecological Modelling, 393(??):52–60, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018304277.

Kolasa:2012:PRH

[KASS12] Jurek Kolasa, Craig R. Allen, Jan Sendzimir, and Craig A. Stow. Predictions and retrodictions of the hierarchical representation of habitat in heterogeneous environments. Ecological Modelling, 245(??):199–207, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001469.

Katzmair:2011:BRT

[Kat11] Harald Katzmair. Book review: A Third Window. Natural Life beyond Newton and Darwin, R. E. Ulanowicz. Templeton Press (2009), ISBN-13: 978-1-59947-154-9. Ecological Modelling, 222(13):2166–2167, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001712.

Kaufman:2010:ENA

[KB10] Anthony G. Kaufman and Stuart R. Borrett. Ecosystem network analysis indicators are generally robust to parameter uncertainty in a phosphorus model of Lake Sidney Lanier, USA. Ecological Modelling, 221(8):1230–1238, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009008643.

Kuparinen:2011:TPP

[KB11] Anna Kuparinen and Mats Björklund. Theory put into practice: an R implementation of the infinite-dimensional model. Ecological Modelling, 222(12):2027–2030, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001931.

Koenig:2018:IFD

[KB18] Shantel J. Koenig and Darren J. Bender. Increasing the function in distance-based functional connectivity assessments: a modified spatial interaction model (SIM) approach. Ecological Modelling, 386(??):47–58, October 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830276X.

Kininmonth:2011:DCR

[KBB^{+}11] Stuart Kininmonth, Maria Beger, Michael Bode, Eric Peterson, Vanessa M. Adams, Dan Dorfman, Daniel R. Brumbaugh, and Hugh P. Possingham. Dispersal connectivity and reserve selection for marine conservation. Ecological Modelling, 222(7):1272–1282, April 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000445.

Kohler:2012:MPP

[KBC^{+}12] Timothy A. Kohler, R. Kyle Bocinsky, Denton Cockburn, Stefani A. Crabtree, Mark D. Varien, Kenneth E. Kolm, Schaun Smith, Scott G. Ortman, and Ziad Kobti. Modelling prehispanic Pueblo societies in their ecosystems. Ecological Modelling, 241(??):30–41, August 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000038.

Knudby:2010:NAM

[KBL10] Anders Knudby, Alexander Brenning, and Ellsworth LeDrew. New approaches to modelling fish–habitat relationships. Ecological Modelling, 221(3):503–511, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009007546.

Kursun:2015:LCE

[KBMM15] Berrin Kursun, Bhavik R. Bakshi, Manoj Mahata, and Jay F. Martin. Life cycle and emergy based design of energy systems in developing countries: Centralized and localized options. Ecological Modelling, 305(??):40–53, June 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001106.

King:2015:EPE

[KBS^{+}15] David A. King, Dominique M. Bachelet, Amy J. Symstad, Ken Ferschweiler, and Michael Hobbins. Estimation of potential evapotranspiration from extraterrestrial radiation, air temperature and humidity to assess future climate change effects on the vegetation of the Northern Great Plains, USA. Ecological Modelling, 297(??):86–97, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014005559.

Kim:2018:PSF

[KC18] Seung Ki Kim and Sung-Uk Choi. Prediction of suitable feeding habitat for fishes in a stream using physical habitat simulations. Ecological Modelling, 385(??):65–77, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302497.

Keesen:2017:SPS

[KCAP17] F. Keesen, A. Castro e Silva, E. Arashiro, and C. F. S. Pinheiro. Simulations of populations of Sapajus robustus in a fragmented landscape. Ecological Modelling, 344(??):38–47, January 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016306627.

Komarov:2017:PRM

[KCB^{+}17] Alexander Komarov, Oleg Chertov, Sergey Bykhovets, Cindy Shaw, Marina Nadporozhskaya, Pavel Frolov, Maxim Shashkov, Vladimir Shanin, Pavel Grabarnik, Irina Priputina, and Elena Zubkova. Romul Hum model of soil organic matter formation coupled with soil biota activity. I. Problem formulation, model description, and testing. Ecological Modelling, 345(??):113–124, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303088.

Kostic:2016:ISI

[KCC16] Vladimir R. Kostić, Ljiljana Cvetković, and Dragana Lj. Cvetković. Improved stability indicators for empirical food webs. Ecological Modelling, 320(??):1–8, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004159.

Keane:2013:ERF

[KCF^{+}13] Robert E. Keane, Geoffrey J. Cary, Mike D. Flannigan, Russell A. Parsons, Ian D. Davies, Karen J. King, Chao Li, Ross A. Bradstock, and Malcolm Gill. Exploring the role of fire, succession, climate, and weather on landscape dynamics using comparative modeling. Ecological Modelling, 266(??):172–186, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300313X.

Kleemann:2017:VAE

[KCF17] Janina Kleemann, Enrico Celio, and Christine Fürst. Validation approaches of an expert-based Bayesian Belief Network in Northern Ghana, West Africa. Ecological Modelling, 365(??):10–29, December 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304167.

Kleemann:2018:RVA

[KCF18] Janina Kleemann, Enrico Celio, and Christine Fürst. Reprint of “Validation approaches of an expert-based Bayesian Belief Network in Northern Ghana, West Africa”. Ecological Modelling, 371(??):101–118, March 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017305045.

Klauschies:2018:BDB

[KCG18] Toni Klauschies, Renato Mendes Coutinho, and Ursula Gaedke. A beta distribution-based moment closure enhances the reliability of trait-based aggregate models for natural populations and communities. Ecological Modelling, 381(??):46–77, August 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018300425.

Klimas:2012:VCT

[KCKdOW12] Christie A. Klimas, Wendell P. Cropper, Karen A. Kainer, and Lúcia H. de Oliveira Wadt. Viability of combined timber and non-timber harvests for one species: a Carapa guianensis case study. Ecological Modelling, 246 (??):147–156, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003535.

Kang:2014:SFS

[KCLL14] Fenni Kang, Paul-Henry Cournède, Jérémie Lecoeur, and Véronique Letort. SUNLAB: a functional–structural model for genotypic and phenotypic characterization of the sunflower crop. Ecological Modelling, 290(??):21–33, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001400091X.

Kowalewski:2015:API

[KCP^{+}15] Lucas K. Kowalewski, Christopher J. Chizinski, Larkin A. Powell, Kevin L. Pope, and Mark A. Pegg. Accuracy or precision: Implications of sample design and methodology on abundance estimation. Ecological Modelling, 316(??):185–190, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003798.

Katsev:2013:MDS

[KD13] Sergei Katsev and Maria Dittrich. Modeling of decadal scale phosphorus retention in lake sediment under varying redox conditions. Ecological Modelling, 251(??):246–259, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005741.

Kuan:2018:SAS

[KDHC^{+}18] A. Carmen Kuan, Gloria DeGrandi-Hoffman, Robert J. Curry, Kristina V. Garber, Andrew R. Kanarek, Marcia N. Snyder, Kurt L. Wolfe, and S. Thomas Purucker. Sensitivity analyses for simulating pesticide impacts on honey bee colonies. Ecological Modelling, 376(??):15–27, May 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300632.

Keane:2010:MMF

[KDK^{+}10] Robert E. Keane, Stacy A. Drury, Eva C. Karau, Paul F. Hessburg, and Keith M. Reynolds. A method for mapping fire hazard and risk across multiple scales and its application in fire management. Ecological Modelling, 221(1):2–18, January 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380008005073.

Kielbassa:2010:ATD

[KDMPC10] J. Kielbassa, M. L. Delignette-Muller, D. Pont, and S. Charles. Application of a temperature-dependent von Bertalanffy growth model to bullhead (Cottus gobio). Ecological Modelling, 221(20):2475–2481, October 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003261.

Kibler:2017:GFS

[KDT^{+}17] Steven R. Kibler, Eric D. Davenport, Patricia A. Tester, D. Ransom Hardison, William C. Holland, and R. Wayne Litaker. Gambierdiscus and Fukuyoa species in the greater Caribbean: Regional growth projections for ciguateraassociated dinoflagellates. Ecological Modelling, 360(??): 204–218, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001730340X.

Koen:2017:EDC

[KdVRdW17] Hildegarde Koen, J. P. de Villiers, Henk Roodt, and Alta de Waal. An expert-driven causal model of the rhino poaching problem. Ecological Modelling, 347(??):29–39, March 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016307621.

Kleczkowski:2017:PBE

[KEHG17] Adam Kleczkowski, Ciaran Ellis, Nick Hanley, and David Goulson. Pesticides and bees: Ecological-economic modelling of bee populations on farmland. Ecological Modelling, 360(??):53–62, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306846.

Keitt:2012:PNI

[Kei12] Timothy H. Keitt. Productivity, nutrient imbalance and fragility in coupled producer–decomposer systems. Ecological Modelling, 245(??):12–18, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002414.

Koehn:2016:DHT

[KEM^{+}16] Laura E. Koehn, Timothy E. Essington, Kristin N. Marshall, Isaac C. Kaplan, William J. Sydeman, Amber I. Szoboszlai, and Julie A. Thayer. Developing a high taxonomic resolution food web model to assess the functional role of forage fish in the California Current ecosystem. Ecological Modelling, 335(??):87–100, September 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301909.

Kenny:2017:MNS

[Ken17] Daniel C. Kenny. Modeling of natural and social capital on farms: Toward useable integration. Ecological Modelling, 356(??):1–13, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304082.

Kennedy:2019:EDP

[Ken19] Maureen C. Kennedy. Experimental design principles to choose the number of Monte Carlo replicates for stochastic ecological models. Ecological Modelling, 394 (??):11–17, February 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018304332.

King:2016:CET

[KF16] Elizabeth G. King and Trenton E. Franz. Combining ecohydrologic and transition probability-based modeling to simulate vegetation dynamics in a semi-arid rangeland. Ecological Modelling, 329(??):41–63, June 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300485.

Kendall:2019:PPC

[KFDL^{+}19] Bruce E. Kendall, Masami Fujiwara, Jasmin Diaz-Lopez, Sandra Schneider, Jakob Voigt, and Sören Wiesner. Persistent problems in the construction of matrix population models. Ecological Modelling, 406(??):33–43, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301085.

Kienast:2015:SLM

[KFvSH15] Felix Kienast, Jacqueline Frick, Maarten J. van Strien, and Marcel Hunziker. The Swiss landscape monitoring program — a comprehensive indicator set to measure landscape change. Ecological Modelling, 295(??):136–150, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003858.

Karcher:2019:SEP

[KFWM19] Oskar Kärcher, Karin Frank, Ariane Walz, and Danijela Markovic. Scale effects on the performance of nichebased models of freshwater fish distributions. Ecological Modelling, 405(??):33–42, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301711.

Kar:2013:IMS

[KG13] T. K. Kar and Bapan Ghosh. Impacts of maximum sustainable yield policy to prey–predator systems. Ecological Modelling, 250(??):134–142, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005522.

Kolmanic:2014:FST

[KGD14] Simon Kolmanič, Nikola Guid, and Jurij Diaci. Forest-MAS — a single tree based secondary succession model employing Ellenberg indicator values. Ecological Modelling, 279(??):100–113, May 10, 2014. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001400101X.

Kilburn:2012:UMC

[KGM12] R. Kilburn, A. Gregory, and A. G. Murray. Using a Markovchain Monte-Carlo modelling approach to identify the relative risk to farmed Scottish rainbow trout (Oncorhynchus mykiss) in a multi-sector industry of viral haemorrhagic septicaemia viruses from introduction and emergent sources. Ecological Modelling, 237–238(??):34–42, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001688.

Keenan:2011:OCI

[KGS11] Trevor F. Keenan, Rüdiger Grote, and Santiago Sabaté. Overlooking the canopy: the importance of canopy structure in scaling isoprenoid emissions from the leaf to the landscape. Ecological Modelling, 222(3):737–747, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005946.

Kulkarni:2015:CCD

[KGYG15] Madhura V. Kulkarni, Peter M. Groffman, Joseph B. Yavitt, and Christine L. Goodale. Complex controls of denitrification at ecosystem, landscape and regional scales in northern hardwood forests. Ecological Modelling, 298(??):39–52, February 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001549.

Keller:2011:MSA

[KH11] David P. Keller and Raleigh R. Hood. Modeling the seasonal autochthonous sources of dissolved organic carbon and nitrogen in the upper Chesapeake Bay. Ecological Modelling, 222(5):1139–1162, March 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006800.

Koleva:2015:PDC

[KH15] Kameliya Z. Koleva and Ferdi L. Hellweger. From protein damage to cell aging to population fitness in E. coli: Insights from a multi-level agent-based model. Ecological Modelling, 301(??):62–71, April 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000435.

Kopainsky:2017:TFS

[KHHN17] Birgit Kopainsky, Gerid Hager, Hugo Herrera, and Progress H. Nyanga. Transforming food systems at local levels: Using participatory system dynamics in an interactive manner to refine small-scale farmers’ mental models. Ecological Modelling, 362(??):101–110, October 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303721.

Kim:2019:DSV

[KHJ^{+}19] Hyo Gyeom Kim, Sungwon Hong, Kwang-Seuk Jeong, Dong-Kyun Kim, and Gea-Jae Joo. Determination of sensitive variables regardless of hydrological alteration in artificial neural network model of chlorophyll a: Case study of Nakdong River. Ecological Modelling, 398(??):67–76, April 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300584.

Kleyheeg-Hartman:2018:PBC

[KHKC^{+}18] J. C. Kleyheeg-Hartman, K. L. Krijgsveld, M. P. Collier, M. J. M. Poot, A. R. Boon, T. A. Troost, and S. Dirksen. Predicting bird collisions with wind turbines: Comparison of the new empirical flux collision model with the SOSS band model. Ecological Modelling, 387(??):144–153, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302308.

Kong:2016:CFW

[KHL^{+}16] Xiangzhen Kong, Wei He, Wenxiu Liu, Bin Yang, Fuliu Xu, Sven Erik Jørgensen, and Wolf M. Mooij. Changes in food web structure and ecosystem functioning of a large, shallow Chinese lake during the 1950s, 1980s and 2000s. Ecological Modelling, 319(??):31–41, January 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003014.

Klauberg:2019:CFE

[KHS^{+}19] Carine Klauberg, Andrew T. Hudak, Carlos Alberto Silva, Sarah A. Lewis, Peter R. Robichaud, and Terrie B. Jain. Characterizing fire effects on conifers at tree level from airborne laser scanning and high-resolution, multispectral satellite data. Ecological Modelling, 412(??): Article 108820, November 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001930328X.

Kastelle:2016:IBP

[KHW^{+}16] Craig R. Kastelle, Thomas E. Helser, Stephen G. Wischniowski, Timothy Loher, Betty J. Goetz, and Lisa A. Kautzi. Incorporation of bomb-produced ^{14}C into fish otoliths: a novel approach for evaluating age validation and bias with an application to yellowfin sole and northern rockfish. Ecological Modelling, 320(??):79–91, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004354.

Klosterhalfen:2017:MSC

[KHW^{+}17] A. Klosterhalfen, M. Herbst, L. Weihermüller, A. Graf, M. Schmidt, A. Stadler, K. Schneider, J.-A. Subke, J. A. Huisman, and H. Vereecken. Multi-site calibration and validation of a net ecosystem carbon exchange model for croplands. Ecological Modelling, 363(??):137–156, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017300054.

Kong:2013:PRE

[KJH^{+}13] Xiang-Zhen Kong, Sven Erik Jørgensen, Wei He, Ning Qin, and Fu-Liu Xu. Predicting the restoration effects by a structural dynamic approach in Lake Chaohu, China. Ecological Modelling, 266(??):73–85, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003256.

Kemanian:2011:ISC

[KJMA11] Armen R. Kemanian, Stefan Julich, Valipuram S. Manoranjan, and Jeffrey R. Arnold. Integrating soil carbon cycling with that of nitrogen and phosphorus in the watershed model SWAT: Theory and model testing. Ecological Modelling, 222(12):1913–1921, June 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001402.

Kumar:2013:EMW

[KJS13] R. Kumar, M. K. Jat, and V. Shankar. Evaluation of modeling of water ecohydrologic dynamics in soil–root system. Ecological Modelling, 269(??):51–60, November 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004195.

Kerimoglu:2017:MPG

[KJVL^{+}17] Onur Kerimoglu, Stéphan Jacquet, Brigitte Vinçon-Leite, Bruno J. Lemaire, Frédéric Rimet, Frédéric Soulignac, Dominique Trévisan, and Orlane Anneville. Modelling the plankton groups of the deep, peri-alpine Lake Bourget. Ecological Modelling, 359(??):415–433, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303131.

Keane:2010:EEB

[KK10] Robert E. Keane and Eva Karau. Evaluating the ecological benefits of wildfire by integrating fire and ecosystem simulation models. Ecological Modelling, 221(8):1162–1172, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010000475.

Karsai:2011:CFH

[KK11] Istvan Karsai and George Kampis. Connected fragmented habitats facilitate stable coexistence dynamics. Ecological Modelling, 222(3):447–455, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005922.

Kreakie:2012:IDD

[KK12a] Betty J. Kreakie and Timothy H. Keitt. Integration of distance, direction and habitat into a predictive migratory movement model for blue-winged teal (Anas discors). Ecological Modelling, 224(1):25–32, January 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005047.

Kumagai:2012:IDR

[KK12b] Tomo’omi Kumagai and Tomonori Kume. Influences of diurnal rainfall cycle on CO_{2} exchange over bornean tropical rainforests. Ecological Modelling, 246(??):91–98, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003420.

Kanapaux:2013:DTO

[KK13] William Kanapaux and Gregory A. Kiker. Development and testing of an object-oriented model for adaptively managing human disturbance of least tern (Sternula antillarum) nesting habitat. Ecological Modelling, 268(??):64–77, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013003852.

Klug:2015:OEI

[KK15a] Hermann Klug and Alexander Kmoch. Operationalizing environmental indicators for real time multi-purpose decision making and action support. Ecological Modelling, 295(??):66–74, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001902.

Kumar:2015:MMS

[KK15b] Vijay Kumar and Beena Kumari. Mathematical modelling of the seasonal variability of plankton and forage fish in the Gulf of Kachchh. Ecological Modelling, 313(??):237–250, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002963.

Kalmykov:2016:EMP

[KK16] Vyacheslav L. Kalmykov and Lev V. Kalmykov. On ecological modelling problems in the context of resolving the biodiversity paradox. Ecological Modelling, 329(??):1–4, June 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016300643.

Klimek-Kopyra:2017:BCP

[KKBZ17] Agnieszka Klimek-Kopyra, Magdalena Bacior, and Tadeusz Zaja̧c. Biodiversity as a creator of productivity and interspecific competitiveness of winter cereal species in mixed cropping. Ecological Modelling, 343(??):123–130, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302861.

Kaplan:2017:MFW

[KKH^{+}17] Isaac C. Kaplan, Laura E. Koehn, Emma E. Hodgson, Kristin N. Marshall, and Timothy E. Essington. Modeling food web effects of low sardine and anchovy abundance in the California Current. Ecological Modelling, 359(??):1–24, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016308262.

Kulakowska:2014:UIB

[KKI^{+}14] K. A. Kułakowska, T. M. Kułakowski, I. R. Inglis, G. C. Smith, P. J. Haynes, P. Prosser, P. Thorbek, and R. M. Sibly. Using an individual-based model to select among alternative foraging strategies of woodpigeons: Data support a memory-based model with a flocking mechanism. Ecological Modelling, 280(??):89–101, May 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004481.

Kumar:2019:PSM

[KKK^{+}19] Manoj Kumar, Naveen Kalra, Peter Khaiter, N. H. Ravindranath, Varsha Singh, Hukum Singh, Subrat Sharma, and Shahryar Rahnamayan. PhenoPine: a simulation model to trace the phenological changes in Pinus roxhburghii in response to ambient temperature rise. Ecological Modelling, 404(??):12–20, July 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301681.

Kim:2012:RPC

[KKL^{+}12] Sohee Kim, Sinkyu Kang, Jong-Hwan Lim, Jung-Hwa Chun, and Joo-Han Sung. Regional parameterization of canopy onset models using MODIS and flowering onset data. Ecological Modelling, 247(??):190–198, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004371.

Klaich:2011:EDA

[KKP^{+}11] M. Javier Klaich, Paul G. Kinas, Susana N. Pedraza, Mariano A. Coscarella, and Enrique A. Crespo. Estimating dyad association probability under imperfect and heterogeneous detection. Ecological Modelling, 222(15):2642–2650, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001645.

Koo:2017:SKF

[KKP^{+}17] Kyung Ah Koo, Woo-Seok Kong, Seon Uk Park, Joon Ho Lee, Jaeuk Kim, and Huicheul Jung. Sensitivity of Korean fir (Abies koreana Wils.), a threatened climate relict species, to increasing temperature at an island subalpine area. Ecological Modelling, 353(??):5–16, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300662.

Kappler:2019:EPV

[KKR19] R. H. Kappler, K. S. Knight, and K. V. Root. Evaluating the population viability of green ash trees (Fraxinus pennsylvanica) before and after the emerald ash borer beetle (Agrilus planipennis) invasion. Ecological Modelling, 400(??):53–59, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001930122X.

Kim:2010:PMP

[KL10] Dong-Soon Kim and Joon-Ho Lee. A population model for the peach fruit moth, Carposina sasakii Matsumura (Lepidoptera: Carposinidae), in a Korean orchard system. Ecological Modelling, 221(2):268–280, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009006590.

Krivtsov:2012:DBH

[KL12] Vladimir Krivtsov and Brian Linfoot. Disruption to benthic habitats by moorings of wave energy installations: a modelling case study and implications for overall ecosystem functioning. Ecological Modelling, 245(??):121–124, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012001056.

Kunegel-Lion:2019:WMF

[KLGL19] Melodie Kunegel-Lion, Devin W. Goodsman, and Mark A. Lewis. When managers forage for pests: Implementing the functional response in pest management. Ecological Modelling, 396(??):59–73, March 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303363.

Kuusinen:2014:MMA

[KLS^{+}14] Nea Kuusinen, Petr Lukeš, Pauline Stenberg, Janne Levula, Eero Nikinmaa, and Frank Berninger. Measured and modelled albedos in Finnish boreal forest stands of different species, structure and understory. Ecological Modelling, 284(??):10–18, July 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001884.

Kim:2017:ABP

[KLS17] Jaeyoung Kim, Tongeun Lee, and Dongil Seo. Algal bloom prediction of the lower Han River, Korea using the EFDC hydrodynamic and water quality model. Ecological Modelling, 366(??):27–36, December 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304428.

Kazanci:2012:EEN

[KM12a] C. Kazanci and Q. Ma. Extending ecological network analysis measures to dynamic ecosystem models. Ecological Modelling, 242(??):180–188, September 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002426.

Kim:2012:ECD

[KM12b] Hae-Cheol Kim and Paul A. Montagna. Effects of climatedriven freshwater inflow variability on macrobenthic secondary production in Texas lagoonal estuaries: a modeling study. Ecological Modelling, 235–236(??):67–80, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001200138X.

Keane:2021:RCD

[KMF^{+}21] Robert E. Keane, Donald McKenzie, Donald A. Falk, Erica A. H. Smithwick, Carol Miller, and Lara-Karena B. Kellogg. Representing climate, disturbance, and vegetation interactions in landscape models. Ecological Modelling, 309–310(??):33–47, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001500160X.

Klimenko:2019:SHP

[KMKL19] Alexandra I. Klimenko, Yury G. Matushkin, Nikolay A. Kolchanov, and Sergey A. Lashin. Spatial heterogeneity promotes antagonistic evolutionary scenarios in microbial community explained by ecological stratification: a simulation study. Ecological Modelling, 399(??):66–76, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300687.

Kumagai:2013:SCC

[KMM^{+}13] Tomo’omi Kumagai, Ryan G. Mudd, Yoshiyuki Miyazawa, Wen Liu, Thomas W. Giambelluca, Nakako Kobayashi, Tiva Khan Lim, Mayuko Jomura, Kazuho Matsumoto, Maoyi Huang, Qi Chen, Alan Ziegler, and Song Yin. Simulation of canopy CO_{2} /H_{2}O fluxes for a rubber (Hevea brasiliensis) plantation in central Cambodia: the effect of the regular spacing of planted trees. Ecological Modelling, 265(??):124–135, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003001.

Kvamsdal:2017:BBC

[KMMS17] Sturla Kvamsdal, José M. Maroto, Manuel Morán, and Leif K. Sandal. A bridge between continuous and discretetime bioeconomic models: Seasonality in fisheries. Ecological Modelling, 364(??):124–131, November 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304192.

Koo:2014:PRS

[KMP14] Kyung Ah Koo, Marguerite Madden, and Bernard C. Patten. Projection of red spruce (Picea rubens Sargent) habitat suitability and distribution in the Southern Appalachian Mountains, USA. Ecological Modelling, 293 (??):91–101, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002737.

Kaszta:2016:BFS

[KMRW16] Żaneta Kaszta, Jorgelina Marino, Abel Ramoelo, and Eléonore Wolff. Bulk feeder or selective grazer: African buffalo space use patterns based on fine-scale remotely sensed data on forage quality and quantity. Ecological Modelling, 323(??):115–122, March 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005748.

Kupisch:2012:TDC

[KMS^{+}12] Moritz Kupisch, Sylvia Moenickes, Jeanette Schlief, Marieke Frassl, and Otto Richter. Temperature-dependent consumerresource dynamics: a coupled structured model for Gammarus pulex (L.) and leaf litter. Ecological Modelling, 247(??):157–167, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004036.

Kupisch:2013:ETD

[KMS^{+}13] Moritz Kupisch, Sylvia Moenickes, Jeanette Schlief, Marieke Frassl, and Otto Richter. Erratum to “Temperaturedependent consumer-resource dynamics: a coupled structured model for Gammarus pulex (L.) and leaf litter [Ecol. Model. 247 (2012) 157–167]. Ecological Modelling, 251(??):323–326, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200508X.

Kocev:2010:LHM

[KNM^{+}10] Dragi Kocev, Andreja Naumoski, Kosta Mitreski, Svetislav Krstić, and Sašo Džeroski. Learning habitat models for the diatom community in Lake Prespa. Ecological Modelling, 221(2):330–337, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009006103.

Korner-Nievergelt:2017:IMR

[KNPH^{+}17] Fränzi Korner-Nievergelt, Céline Prévot, Steffen Hahn, Lukas Jenni, and Felix Liechti. The integration of mark re-encounter and tracking data to quantify migratory connectivity. Ecological Modelling, 344(??):87–94, January 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016306925.

Kim:2015:ZDR

[KNU^{+}15] Hungsoo Kim, Tuyen Van Nguyen, Takashi Uehara, Muyoung Heo, and Tae-Soo Chon. Zebrafish (Danio rerio) movement in addressing stress to conflicting stimuli, foods and predators. Ecological Modelling, 306(??):257–267, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004785.

Kamp:2013:HMC

[KØ13] Andreas Kamp and Hanne Østergård. How to manage co-product inputs in emergy accounting exemplified by willow production for bioenergy. Ecological Modelling, 253(??):70–78, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013000100.

Kasimova:2013:OSA

[KOBK13] R. G. Kasimova, Yu. V. Obnosov, F. B. Baksht, and A. R. Kacimov. Optimal shape of an anthill dome: Bejan’s constructal law revisited. Ecological Modelling, 250 (??):384–390, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005595.

Kosicki:2017:STM

[Kos17] Jakub Z. Kosicki. Should topographic metrics be considered when predicting species density of birds on a large geographical scale? A case of Random Forest approach. Ecological Modelling, 349(??):76–85, April 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306755.

Kramer:2010:ASR

[KP10] K. Kramer and H. H. T. Prins. Allometric scaling of resource acquisition by ruminants in dynamic and heterogeneous environments. Ecological Modelling, 221(21):2555–2564, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010003522.

Kim:2013:PSM

[KP13] Daehyun Kim and Jonathan D. Phillips. Predicting the structure and mode of vegetation dynamics: an application of graph theory to state-and-transition models. Ecological Modelling, 265(??):64–73, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002858.

Koo:2017:PCC

[KPK^{+}17] Kyung Ah Koo, Seon Uk Park, Woo-Seok Kong, Seungbum Hong, Inyoung Jang, and Changwan Seo. Potential climate change effects on tree distributions in the Korean Peninsula: Understanding model & climate uncertainties. Ecological Modelling, 353(??):17–27, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305440.

Kjelland:2015:IMA

[KPLS15] Michael E. Kjelland, Candice D. Piercy, Tahirih Lackey, and Todd M. Swannack. An integrated modeling approach for elucidating the effects of different management strategies on Chesapeake Bay oyster metapopulation dynamics. Ecological Modelling, 308(??):45–62, July 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001234.

Kim:2014:ICR

[KPM^{+}14] MinHyeok Kim, Namyong Park, R. I. (Bob) McKay, Haisoo Shin, Yun-Geun Lee, Kwang-Seuk Jeong, and Dong-Kyun Kim. Improvement of complex and refractory ecological models: Riverine water quality modelling using evolutionary computation. Ecological Modelling, 291 (??):205–217, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003585.

Korner:2014:MEB

[KPMJ14] Katrin Körner, Hans Pfestorf, Felix May, and Florian Jeltsch. Modelling the effect of belowground herbivory on grassland diversity. Ecological Modelling, 273 (??):79–85, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013005243.

Kjelland:2017:BGT

[KPS17] Michael E. Kjelland, Candice D. Piercy, and Todd M. Swannack. Beyond graphs and tables: Enhancing explanatory power of complex environmental simulations through 3D printed model output. Ecological Modelling, 360(??):244–251, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001730337X.

Krishna:2019:CTC

[KPS19] Shubham Krishna, Markus Pahlow, and Markus Schartau. Comparison of two carbon-nitrogen regulatory models calibrated with mesocosm data. Ecological Modelling, 411(??):Article 108711, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302029.

Kim:2010:ESB

[KPSC10] Jung-Woo Kim, Yakov A. Pachepsky, Daniel R. Shelton, and Cary Coppock. Effect of streambed bacteria release on E. coli concentrations: Monitoring and modeling with the modified SWAT. Ecological Modelling, 221(12):1592–1604, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010001390.

Koo:2011:AEF

[KPT11] Kyung-Ah Koo, Bernard C. Patten, and Robert O. Teskey. Assessing environmental factors in red spruce (Picea rubens sarg.) growth in the Great Smoky Mountains National Park, USA: From conceptual model, envirogram, to simulation model. Ecological Modelling, 222(3):824–834, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006253.

Koo:2014:CCE

[KPTC14] Kyung Ah Koo, Bernard C. Patten, Robert O. Teskey, and Irena F. Creed. Climate change effects on red spruce decline mitigated by reduction in air pollution within its shrinking habitat range. Ecological Modelling, 293(??):81–90, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003536.

Kairis:2010:SLR

[KR10a] Peter A. Kairis and John M. Rybczyk. Sea level rise and eelgrass (Zostera marina) production: a spatially explicit relative elevation model for Padilla Bay, WA. Ecological Modelling, 221(7):1005–1016, April 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009000799.

Kissinger:2010:IEA

[KR10b] Meidad Kissinger and William E. Rees. An interregional ecological approach for modelling sustainability in a globalizing world — reviewing existing approaches and emerging directions. Ecological Modelling, 221(21):2615–2623, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010003431.

Krug:2014:MET

[KR14] Rainer M. Krug and David M. Richardson. Modelling the effect of two biocontrol agents on the invasive alien tree Acacia cyclops — flowering, seed production and agent survival. Ecological Modelling, 278(??):100–113, April 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000714.

Kubicek:2016:MMF

[KR16] Andreas Kubicek and Hauke Reuter. Mechanics of multiple feedbacks in benthic coral reef communities. Ecological Modelling, 329(??):29–40, June 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300473.

Kuusemae:2016:MSE

[KRCVF16] Kadri Kuusemäe, Erik Kock Rasmussen, Paula Canal-Vergés, and Mogens R. Flindt. Modelling stressors on the eelgrass recovery process in two Danish estuaries. Ecological Modelling, 333(??):11–42, August 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301053.

Keena:2018:CCD

[KRE^{+}18] Naomi Keena, Marco Raugei, Mohamed Aly Etman, Daniel Ruan, and Anna Dyson. Clark’s crow: a design plugin to support emergy analysis decision making towards sustainable urban ecologies. Ecological Modelling, 367(??):42–57, January 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017302764.

Ko:2011:MDE

[KRL11] Chia-Ying Ko, Terry L. Root, and Pei-Fen Lee. Movement distances enhance validity of predictive models. Ecological Modelling, 222(4):947–954, February 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006502.

Kim:2016:SCD

[KRL^{+}16] Youngil Kim, Nigel T. Roulet, Changsheng Li, Steve Frolking, Ian B. Strachan, Changhui Peng, Cristian R. Teodoru, Yves T. Prairie, and Alain Tremblay. Simulating carbon dioxide exchange in boreal ecosystems flooded by reservoirs. Ecological Modelling, 327(??):1–17, May 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016000259.

Kogler:2017:CWS

[KS17a] F. Kögler and D. Söffker. Corrigendum to “Water (stress) models and deficit irrigation: System-theoretical description and causality mapping” [Ecol. Model. 361 (2017) 135– 156]. Ecological Modelling, 363(??):234–235, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304052.

Kogler:2017:WSM

[KS17b] F. Kögler and D. Söffker. Water (stress) models and deficit irrigation: System-theoretical description and causality mapping. Ecological Modelling, 361(??):135–156, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017300686.

Kangas:2018:IET

[KS18] Patrick Kangas and Gary E. Seibel. An industrial ecology teaching exercise on cycling e-waste. Ecological Modelling, 371(??):119–122, March 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303617.

Kearney:2012:CPE

[KSAS12] Kelly A. Kearney, Charles Stock, Kerim Aydin, and Jorge L. Sarmiento. Coupling planktonic ecosystem and fisheries food web models for a pelagic ecosystem: Description and validation for the subarctic Pacific. Ecological Modelling, 237–238(??):43–62, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200172X.

Kuemmerlen:2014:ICP

[KSG^{+}14] Mathias Kuemmerlen, Britta Schmalz, Björn Guse, Qinghua Cai, Nicola Fohrer, and Sonja C. Jähnig. Integrating catchment properties in small scale species distribution models of stream macroinvertebrates. Ecological Modelling, 277(??):77–86, April 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000635.

Kautz:2014:ITD

[KSI14] Markus Kautz, Reinhard Schopf, and Muhammad Ali Imron. Individual traits as drivers of spatial dispersal and infestation patterns in a host–bark beetle system. Ecological Modelling, 273(??):264–276, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300570X.

Kim:2018:MII

[KSM^{+}18] Yongeun Kim, Jino Son, Hyoung-Ho Mo, Yun-Sik Lee, and Kijong Cho. Modeling the influence of initial density and copper exposure on the interspecific competition of two algal species. Ecological Modelling, 383(??):160–170, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018301418.

Kapeller:2011:MSP

[KSS11] Stefan Kapeller, Hilke Schroeder, and Silvio Schueler. Modelling the spatial population dynamics of the green oak leaf roller (Tortrix viridana) using density dependent competitive interactions: Effects of herbivore mortality and varying host-plant quality. Ecological Modelling, 222(7):1293–1302, April 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011000275.

Keshtkar:2013:ABN

[KSSA13] A. R. Keshtkar, A. Salajegheh, A. Sadoddin, and M. G. Allan. Application of Bayesian networks for sustainability assessment in catchment modeling and management (case study: the Hablehrood River catchment). Ecological Modelling, 268(??):48–54, October 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003967.

Kienzle:2016:MLE

[KSZW16] Marco Kienzle, David Sterling, Shijie Zhou, and You-Gan Wang. Maximum likelihood estimation of natural mortality and quantification of temperature effects on catchability of brown tiger prawn (Penaeus esculentus) in Moreton Bay (Australia) using logbook data. Ecological Modelling, 322(??):1–9, February 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005293.

Keohane:2019:SBM

[KTGH19] Isaac P. Keohane, Jacqueline Tay, Rasika K. Gawde, and Raleigh R. Hood. Sensitivity of a biogeochemical model to the formulation of oyster filtration. Ecological Modelling, 403(??):70–84, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301322.

Kibler:2015:EOW

[KTK^{+}15] Steven R. Kibler, Patricia A. Tester, Kenneth E. Kunkel, Stephanie K. Moore, and R. Wayne Litaker. Effects of ocean warming on growth and distribution of dinoflagellates associated with ciguatera fish poisoning in the Caribbean. Ecological Modelling, 316(??):194–210, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500383X.

Kluger:2016:CCS

[KTM^{+}16] Lotta C. Kluger, Marc H. Taylor, Jaime Mendo, Jorge Tam, and Matthias Wolff. Carrying capacity simulations as a tool for ecosystem-based management of a scallop aquaculture system. Ecological Modelling, 331(??):44–55, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003919.

Kaloveloni:2015:WLC

[KTV^{+}15] Aggeliki Kaloveloni, Thomas Tscheulin, Ante Vujić, Snežana Radenković, and Theodora Petanidou. Winners and losers of climate change for the genus Merodon (Diptera: Syrphidae) across the Balkan Peninsula. Ecological Modelling, 313(??):201–211, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002811.

Katsumata:2018:DDP

[KUI^{+}18] Yuki Katsumata, Takashi Uehara, Hiromu Ito, Jin Yoshimura, Kei ichi Tainaka, and Genki Ichinose. Density-dependent population model of effective release policy for ayu fish. Ecological Modelling, 388(??):80–87, November 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303181.

Kuter:2011:BDK

[KUK11] Semih Kuter, Nurunnisa Usul, and Nazan Kuter. Bandwidth determination for kernel density analysis of wildfire events at forest sub-district scale. Ecological Modelling, 222(17):3033–3040, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003395.

Kininmonth:2010:DCS

[KvOP10] Stuart Kininmonth, Madeleine J. H. van Oppen, and Hugh P. Possingham. Determining the community structure of the coral Seriatopora hystrix from hydrodynamic and genetic networks. Ecological Modelling, 221(24): 2870–2880, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001000459X.

Kumar:2016:SZM

[KVP16] Rajeev Kumar, Divya Varkey, and Tony Pitcher. Simulation of zebra mussels (Dreissena polymorpha) invasion and evaluation of impacts on Mille Lacs Lake, Minnesota: an ecosystem model. Ecological Modelling, 331(??):68–76, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016300072.

Krapivin:2017:SRC

[KVS17] Vladimir F. Krapivin, Costas A. Varotsos, and Vladimir Yu. Soldatov. Simulation results from a coupled model of carbon dioxide and methane global cycles. Ecological Modelling, 359(??):69–79, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303101.

Kuusemae:2018:ABM

[KvTL^{+}18] Kadri Kuusemäe, Miriam von Thenen, Troels Lange, Erik Kock Rasmussen, Michael Pothoff, Ana I. Sousa, and Mogens R. Flindt. Agent based modelling (ABM) of eelgrass (Zostera marina) seedbank dynamics in a shallow Danish estuary. Ecological Modelling, 371(??):60–75, March 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300115.

Kremer:2010:SPE

[KVU^{+}10] James N. Kremer, Jamie M. P. Vaudrey, Dave S. Ullman, Deanna L. Bergondo, Nicole LaSota, Chris Kincaid, Dan L. Codiga, and Mark J. Brush. Simulating property exchange in estuarine ecosystem models at ecologically appropriate scales. Ecological Modelling, 221(7):1080–1088, April 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009008606.

Kazama:2018:EPD

[KW18] So Kazama and Kengo Watanabe. Estimation of periphyton dynamics in a temperate catchment using a distributed nutrient-runoff model. Ecological Modelling, 367(??):1–9, January 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304489.

Kocmoud:2019:PDE

[KWGG19] Amanda R. Kocmoud, Hsiao-Hsuan Wang, William E. Grant, and Benny J. Gallaway. Population dynamics of the endangered Kemp’s Ridley sea turtle following the 2010 oil spill in the Gulf of Mexico: Simulation of potential cause-effect relationships. Ecological Modelling, 392(??):159–178, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304022.

Koralewski:2019:CGP

[KWGW19] Tomasz E. Koralewski, John K. Westbrook, William E. Grant, and Hsiao-Hsuan Wang. Coupling general physical environmental process models with specific questiondriven ecological simulation models. Ecological Modelling, 405(??):102–105, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300651.

Kazmierczak:2014:NVN

[KWH14] Martin Kazmierczak, Thorsten Wiegand, and Andreas Huth. A neutral vs. non-neutral parametrizations of a physiological forest gap model. Ecological Modelling, 288 (??):94–102, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002270.

Kruse:2016:TDS

[KWJH16] Stefan Kruse, Mareike Wieczorek, Florian Jeltsch, and Ulrike Herzschuh. Treeline dynamics in Siberia under changing climates as inferred from an individual-based model for Larix. Ecological Modelling, 338(??):101–121, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302800.

Karam:2013:DEN

[KWS^{+}13] Sarah L. Karam, Peter J. Weisberg, Robert M. Scheller, Dale W. Johnson, and W. Wally Miller. Development and evaluation of a nutrient cycling extension for the LANDIS-II landscape simulation model. Ecological Modelling, 250(??):45–57, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005182.

Khan:2014:EMA

[KYvdPS14] Muhammad S. Khan, Xinyou Yin, Peter E. L. van der Putten, and Paul C. Struik. An ecophysiological model analysis of yield differences within a set of contrasting cultivars and an F1 segregating population of potato (Solanum tuberosum L.) grown under diverse environments. Ecological Modelling, 290(??):146–154, October 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005620.

Kim:2013:IRI

[KZR^{+}13] Dong-Kyun Kim, Weitao Zhang, Yerubandi R. Rao, Sue Watson, Shan Mugalingam, Tanya Labencki, Maria Dittrich, Andrew Morley, and George B. Arhonditsis. Improving the representation of internal nutrient recycling with phosphorus mass balance models: a case study in the Bay of Quinte, Ontario, Canada. Ecological Modelling, 256(??):53–68, May 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013001014.

Lennox:2011:SSL

[LA11] Gareth D. Lennox and Paul R. Armsworth. Suitability of short or long conservation contracts under ecological and socio-economic uncertainty. Ecological Modelling, 222(15):2856–2866, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002699.

Li:2010:EGE

[LAB^{+}10] Hong Li, Mijail Arias, Anouk Blauw, Hans Los, Arthur E. Mynett, and Steef Peters. Enhancing generic ecological model for short-term prediction of Southern North Sea algal dynamics with remote sensing images. Ecological Modelling, 221(20):2435–2446, October 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000308X.

Labrum:2011:AEE

[Lab11] Matthew J. Labrum. Allee effects and extinction debt. Ecological Modelling, 222(5):1205–1207, March 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006794.

Lopes:2010:MEP

[LAC10] José Fortes Lopes, M. A. Almeida, and M. A. Cunha. Modelling the ecological patterns of a temperate lagoon in a very wet spring season. Ecological Modelling, 221 (19):2302–2322, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010002784.

Larocque:2011:DGM

[LAD11] Guy R. Larocque, Louis Archambault, and Claude Delisle. Development of the gap model ZELIG-CFS to predict the dynamics of North American mixed forest types with complex structures. Ecological Modelling, 222(14):2570–2583, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004485.

Lopez-Alfaro:2012:IBM

[LAEAG12] Claudia López-Alfaro, Cristián F. Estades, Dennis K. Aldridge, and Robin M. A. Gill. Individual-based modeling as a decision tool for the conservation of the endangered huemul deer (Hippocamelus bisulcus) in southern Chile. Ecological Modelling, 244(??):104–116, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003225.

Liukkonen:2018:MMS

[LAK^{+}18] Lauri Liukkonen, Daniel Ayllón, Mervi Kunnasranta, Marja Niemi, Jacob Nabe-Nielsen, Volker Grimm, and Anna-Maija Nyman. Modelling movements of Saimaa ringed seals using an individual-based approach. Ecological Modelling, 368(??):321–335, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001730488X.

Laguna:2021:MML

[LAK^{+}21] M. F. Laguna, G. Abramson, M. N. Kuperman, J. L. Lanata, and J. A. Monjeau. Mathematical model of livestock and wildlife: Predation and competition under environmental disturbances. Ecological Modelling, 309–310(??):110–117, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015001714.

Langille:2016:NAI

[LAR^{+}16] Aaron B. Langille, Ellen M. Arteca, Geraldine D. Ryan, Lisa M. Emiljanowicz, and Jonathan A. Newman. North American invasion of Spotted-Wing Drosophila (Drosophila suzukii): a mechanistic model of population dynamics. Ecological Modelling, 336(??):70–81, September 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301946.

Lopez-Alfaro:2013:EHR

[LARZN13] Claudia López-Alfaro, Charles T. Robbins, Andreas Zedrosser, and Scott E. Nielsen. Energetics of hibernation and reproductive trade-offs in brown bears. Ecological Modelling, 270(??):1–10, December 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004213.

Lasaponara:2010:E

[Las10] Rosa Lasaponara. Editorial. Ecological Modelling, 221(1):1, January 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009006449.

Lee:2011:EPR

[LB11] Seungjun Lee and Mark T. Brown. An energetic perspective on the relationship between disturbance and ecosystem productivity. Ecological Modelling, 222(19):3577–3586, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004327.

Lutnesky:2015:SMP

[LB15] Marvin M. F. Lutnesky and Thomas R. Brown. Simulation of movement that potentially maximizes assessment, presence, and defense in territorial animals with varying movement strategies. Ecological Modelling, 313(??):50–58, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002525.

Lee:2017:BEA

[LB17] Jae Min Lee and William W. Braham. Building emergy analysis of Manhattan: Density parameters for highdensity and high-rise developments. Ecological Modelling, 363(??):157–171, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307451.

Lowman:2018:PCB

[LB18a] Lauren E. L. Lowman and Ana P. Barros. Predicting canopy biophysical properties and sensitivity of plant carbon uptake to water limitations with a coupled eco-hydrological framework. Ecological Modelling, 372(??):33–52, March 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300346.

Lupinacci:2018:EAD

[LB18b] Daniel M. Lupinacci and Silvia H. Bonilla. Exploring approaches and dimensions of human transformity through an educational case. Ecological Modelling, 368 (??):336–343, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016306317.

Larocque:2014:IMS

[LBA14] Guy R. Larocque, Jagtar Bhatti, and André Arsenault. Integrated modelling software platform development for effective use of ecosystem models. Ecological Modelling, 288 (??):195–202, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002932.

Larocque:2015:IMS

[LBA15] Guy R. Larocque, Jagtar Bhatti, and André Arsenault. Integrated modelling software platform development for effective use of ecosystem models. Ecological Modelling, 306(??):318–325, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003809.

Luo:2014:CCP

[LBC14] Qunying Luo, Michael Bange, and Loretta Clancy. Cotton crop phenology in a new temperature regime. Ecological Modelling, 285(??):22–29, August 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001999.

Lo:2015:LUC

[LBC^{+}15] Yueh-Hsin Lo, Juan A. Blanco, Rosa M. Canals, Ester González de Andrés, Leticia San Emeterio, J. Bosco Imbert, and Federico J. Castillo. Land use change effects on carbon and nitrogen stocks in the pyrenees during the last 150 years: a modeling approach. Ecological Modelling, 312(??):322–334, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002549.

Lercari:2010:TMS

[LBD10] Diego Lercari, Leandro Bergamino, and Omar Defeo. Trophic models in sandy beaches with contrasting morphodynamics: Comparing ecosystem structure and biomass flow. Ecological Modelling, 221(23):2751–2759, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004229.

Lo:2019:CFP

[LBdA^{+}19] Yueh-Hsin Lo, Juan A. Blanco, Ester González de Andrés, J. Bosco Imbert, and Federico J. Castillo. CO_{2} fertilization plays a minor role in long-term carbon accumulation patterns in temperate pine forests in the southwestern Pyrenées. Ecological Modelling, 407(??):Article 108737, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302455.

Lecomte:2013:MHA

[LBE^{+}13] J. B. Lecomte, H. P. Benoît, M. P. Etienne, L. Bel, and E. Parent. Modeling the habitat associations and spatial distribution of benthic macroinvertebrates: a hierarchical Bayesian model for zero-inflated biomass data. Ecological Modelling, 265(??):74–84, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003013.

Larsen:2017:MMC

[LBE^{+}17] Trine C. Larsen, Nicola K. Browne, Anders C. Erichsen, Karenne Tun, and Peter A. Todd. Modelling for management: Coral photo-physiology and growth potential under varying turbidity regimes. Ecological Modelling, 362(??):1–12, October 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306378.

Leite:2018:MSP

[LBE^{+}18] Marcos V. M. Leite, Lenka Bobuľská, Suéllen P. Espíndola, Maria R. C. Campos, Lucas C. B. Azevedo, and Adão S. Ferreira. Modeling of soil phosphatase activity in land use ecosystems and topsoil layers in the Brazilian Cerrado. Ecological Modelling, 385(??):182–188, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302631.

Lorscheid:2019:CGP

[LBGM19] Iris Lorscheid, Uta Berger, Volker Grimm, and Matthias Meyer. From cases to general principles: a call for theory development through agent-based modeling. Ecological Modelling, 393(??):153–156, February 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303296.

Lausch:2015:UQL

[LBH^{+}15] Angela Lausch, Thomas Blaschke, Dagmar Haase, Felix Herzog, Ralf-Uwe Syrbe, Lutz Tischendorf, and Ulrich Walz. Understanding and quantifying landscape structure — a review on relevant process characteristics, data models and landscape metrics. Ecological Modelling, 295(??):31–41, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003974.

Li:2013:EEI

[LBK^{+}13] Fengqing Li, Mi-Jung Bae, Yong-Su Kwon, Namil Chung, Soon-Jin Hwang, Sang-Jung Park, Hae-Kyung Park, Dong Soo Kong, and Young-Seuk Park. Ecological exergy as an indicator of land-use impacts on functional guilds in river ecosystems. Ecological Modelling, 252(??):53–62, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004711.

Lanier:2016:IBF

[LBM16] Wendy E. Lanier, Larissa L. Bailey, and Erin Muths. Integrating biology, field logistics, and simulations to optimize parameter estimation for imperiled species. Ecological Modelling, 335(??):16–23, September 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301764.

Lau:2018:RSH

[LBN^{+}18] Matthew K. Lau, Benjamin Baiser, Amanda Northrop, Nicholas J. Gotelli, and Aaron M. Ellison. Regime shifts and hysteresis in the pitcher-plant microecosystem. Ecological Modelling, 382(??):1–8, August 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830139X.

Laperriere:2016:CSM

[LBR16] Vincent Laperrière, Katharina Brugger, and Franz Rubel. Cross-scale modeling of a vector-borne disease, from the individual to the metapopulation: the seasonal dynamics of sylvatic plague in Kazakhstan. Ecological Modelling, 342(??):34–48, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016304902.

Landry:2017:FDS

[LBR^{+}17] C. Landry, F. Bonnot, V. Ravigné, J. Carlier, D. Rengifo, J. Vaillant, and C. Abadie. A foliar disease simulation model to assist the design of new control methods against black leaf streak disease of banana. Ecological Modelling, 359(??):383–397, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301412.

Liu:2014:ITI

[LBS^{+}14] Chun Liu, Agnieszka J. Bednarska, Richard M. Sibly, Roger C. Murfitt, Peter Edwards, and Pernille Thorbek. Incorporating toxicokinetics into an individual-based model for more realistic pesticide exposure estimates: a case study of the wood mouse. Ecological Modelling, 280(??):30–39, May 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013004365.

Lassalle:2014:TED

[LBSB^{+}14] Géraldine Lassalle, Pierre Bourdaud, Blanche Saint-Béat, Sébastien Rochette, and Nathalie Niquil. A toolbox to evaluate data reliability for whole-ecosystem models: Application on the Bay of Biscay continental shelf food-web model. Ecological Modelling, 285(??):13–21, August 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001835.

Li:2018:QPL

[LBW18] Yan Li, Vicki S. Blazer, and Tyler Wagner. Quantifying population-level effects of water temperature, flow velocity and chemical-induced reproduction depression: a simulation study with smallmouth bass. Ecological Modelling, 384(??):63–74, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302205.

Lewis:2014:MBB

[LBWW14] D. L. Lewis, S. W. Breck, K. R. Wilson, and C. T. Webb. Modeling black bear population dynamics in a human-dominated stochastic environment. Ecological Modelling, 294(??):51–58, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004098.

Lampa:2019:EWT

[LBZA19] Martin Erlandsson Lampa, Salim Belyazid, Giuliana Zanchi, and Cecilia Akselsson. Effects of whole-tree harvesting on soil, soil water and tree growth — a dynamic modelling exercise in four long-term experiments. Ecological Modelling, 414(??):Article 108832, December 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303400.

Liu:2011:ICC

[LC11] Qiang Liu and Baoshan Cui. Impacts of climate change/ variability on the streamflow in the Yellow River Basin, China. Ecological Modelling, 222(2):268–274, January 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009008096.

Logsdon:2013:QAE

[LC13] Rebecca A. Logsdon and Indrajeet Chaubey. A quantitative approach to evaluating ecosystem services. Ecological Modelling, 257(??):57–65, May 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000902.

Lloret-Climent:2014:CIB

[LCAJGFNS14] Miguel Lloret-Climent, Rocco Amorós-Jiménez, Lucía González-Franco, and Josué-Antonio Nescolarde-Selva. Coverage and invariance for the biological control of pests in Mediterranean greenhouses. Ecological Modelling, 292 (??):37–44, November 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004116.

Lewandowska-Czarnecka:2019:EEA

[LCBNP19] Anna Lewandowska-Czarnecka, Luz Selene Buller, Andrzej Nienartowicz, and Agnieszka Piernik. Energy and emergy analysis for assessing changes in Polish agriculture since the accession to the European union. Ecological Modelling, 412(??):Article 108819, November 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303278.

Li:2014:DSC

[LCC^{+}14] Yuesong Li, Xinjun Chen, Changsheng Chen, Jianzhong Ge, Rubao Ji, Rucheng Tian, Pengfei Xue, and Liuxiong Xu. Dispersal and survival of chub mackerel (Scomber japonicus) larvae in the East China Sea. Ecological Modelling, 283(??):70–84, July 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001604.

Li:2015:DAE

[LCCL15] Weiming Li, Qiuwen Chen, Desuo Cai, and Ruonan Li. Determination of an appropriate ecological hydrograph for a rare fish species using an improved fish habitat suitability model introducing landscape ecology index. Ecological Modelling, 311(??):31–38, September 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002069.

Lennox:2016:USA

[LCD^{+}16] Robert J. Lennox, Steven J. Cooke, Ola H. Diserud, Torgeir B. Havn, Martin R. Johansen, Eva B. Thorstad, Frederick G. Whoriskey, and Ingebrigt Uglem. Use of simulation approaches to evaluate the consequences of catch-and-release angling on the migration behaviour of adult Atlantic salmon (Salmo salar). Ecological Modelling, 333(??):43–50, August 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301077.

Lopez-Carr:2012:SVP

[LCDJ^{+}12] David López-Carr, Jason Davis, Marta M. Jankowska, Laura Grant, Anna Carla López-Carr, and Matthew Clark. Space versus place in complex human–natural systems: Spatial and multi-level models of tropical land use and cover change (LUCC) in Guatemala. Ecological Modelling, 229(??):64–75, March 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004388.

Liu:2019:MLC

[LCFD19] Guohua Liu, Xiaoqiu Chen, Yongshuo Fu, and Nicolas Delpierre. Modelling leaf coloration dates over temperate China by considering effects of leafy season climate. Ecological Modelling, 394(??):34–43, February 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304319.

Lesnoff:2012:SAR

[LCH12] Matthieu Lesnoff, Christian Corniaux, and Pierre Hiernaux. Sensitivity analysis of the recovery dynamics of a cattle population following drought in the Sahel region. Ecological Modelling, 232(??):28–39, May 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000944.

Li:2013:EAN

[LCL^{+}13] Jia Li, Yaoping Cui, Jiyuan Liu, Wenjiao Shi, and Yaochen Qin. Estimation and analysis of net primary productivity by integrating MODIS remote sensing data with a light use efficiency model. Ecological Modelling, 252(??):3–10, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005649.

Ledo:2011:IMC

[LCM11] Alicia Ledo, Sonia Condés, and Fernando Montes. Intertype mark correlation function: a new tool for the analysis of species interactions. Ecological Modelling, 222 (3):580–587, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005958.

Louvrier:2018:AMH

[LCMG18] Julie Louvrier, Thierry Chambert, Eric Marboutin, and Olivier Gimenez. Accounting for misidentification and heterogeneity in occupancy studies using hidden Markov models. Ecological Modelling, 387(??):61–69, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302904.

Loudermilk:2011:LPP

[LCML11] E. L. Loudermilk, W. P. Cropper, R. J. Mitchell, and H. Lee. Longleaf pine (Pinus palustris) and hardwood dynamics in a fire-maintained ecosystem: a simulation approach. Ecological Modelling, 222(15):2733–2750, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002626.

Liu:2014:AEG

[LCP14] Wen Liu, Weiping Chen, and Chi Peng. Assessing the effectiveness of green infrastructures on urban flooding reduction: a community scale study. Ecological Modelling, 291(??):6–14, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003391.

Liu:2015:ISS

[LCP15] Wen Liu, Weiping Chen, and Chi Peng. Influences of setting sizes and combination of green infrastructures on community’s stormwater runoff reduction. Ecological Modelling, 318(??):236–244, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005651.

Lotze-Campen:2010:SGB

[LCPB^{+}10] Hermann Lotze-Campen, Alexander Popp, Tim Beringer, Christoph Müller, Alberte Bondeau, Stefanie Rost, and Wolfgang Lucht. Scenarios of global bioenergy production: the trade-offs between agricultural expansion, intensification and trade. Ecological Modelling, 221 (18):2188–2196, September 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006395.

Lorrilliere:2012:EDI

[LCR12] Romain Lorrillière, Denis Couvet, and Alexandre Robert. The effects of direct and indirect constraints on biological communities. Ecological Modelling, 224(1):103–110, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001100500X.

Liu:2011:PSR

[LCS11] Xuehua Liu, Xiangnan Cheng, and Andrew K. Skidmore. Potential solar radiation pattern in relation to the monthly distribution of giant pandas in Foping Nature Reserve, China. Ecological Modelling, 222(3):645–652, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005533.

Liu:2017:TIR

[LCT^{+}17] Gengyuan Liu, Marco Casazza, Xin Tian, Yan Zhang, Yan Hao, Lixiao Zhang, Sai Liang, and Biagio F. Giannetti. Tracing the inter-regional coal flows and environmental impacts in jing-Jin-ji region. Ecological Modelling, 364(??):98–112, November 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303216.

Liu:2012:FTM

[LCTH12] Wei-Chung Liu, Hsuan-Wien Chen, Tsung-Hsi Tsai, and Hsien-Kuei Hwang. A fish tank model for assembling food webs. Ecological Modelling, 245(??):166–175, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001639.

Li:2013:EPS

[LCY^{+}13] Xia Li, Baoshan Cui, Qichun Yang, Yan Lan, Tingting Wang, and Zhen Han. Effects of plant species on macrophyte decomposition under three nutrient conditions in a eutrophic shallow lake, North China. Ecological Modelling, 252(??):121–128, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004061.

Lin:2011:SOT

[LCZ11] Hua Lin, Min Cao, and Yiping Zhang. Self-organization of tropical seasonal rain forest in southwest China. Ecological Modelling, 222(15):2812–2816, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003467.

Louca:2016:RCM

[LD16] Stilianos Louca and Michael Doebeli. Reaction-centric modeling of microbial ecosystems. Ecological Modelling, 335(??):74–86, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301910.

Liu:2014:RDF

[LDB14] Rongsong Liu, Donald L. DeAngelis, and John P. Bryant. Ratio-dependent functional response emerges from optimal foraging on a complex landscape. Ecological Modelling, 292(??):45–50, November 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004104.

Lin:2016:IGU

[LDC^{+}16] Laurence Lin, Lisa Davis, Sagy Cohen, Elise Chapman, and Jennifer W. Edmonds. The influence of geomorphic unit spatial distribution on nitrogen retention and removal in a large river. Ecological Modelling, 336(??):26–35, September 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016301983.

Lobianco:2021:IIS

[LDCB21] Antonello Lobianco, Philippe Delacote, Sylvain Caurla, and Ahmed Barkaoui. The importance of introducing spatial heterogeneity in bio-economic forest models: Insights gleaned from FFSM++. Ecological Modelling, 309–310(??):82–92, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001635.

Lane:2014:EPR

[LDG14] Murray Lane, Les Dawes, and Peter Grace. The essential parameters of a resource-based carrying capacity assessment model: an Australian case study. Ecological Modelling, 272(??):220–231, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004717.

Lacarriere:2015:EAB

[LDJZC15] Bruno Lacarrière, Kévin Ruben Deutz, Nadia Jamali-Zghal, and Olivier Le Corre. Emergy assessment of the benefits of closed-loop recycling accounting for material losses. Ecological Modelling, 315(??):77–87, November 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000344.

Li:2012:SDM

[LDL12] Fu Jia Li, Suo Cheng Dong, and Fei Li. A system dynamics model for analyzing the eco-agriculture system with policy recommendations. Ecological Modelling, 227(??):34–45, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005874.

Lacitignola:2015:DST

[LDM15a] Deborah Lacitignola, Fasma Diele, and Carmela Marangi. Dynamical scenarios from a two-patch predator–prey system with human control — implications for the conservation of the wolf in the Alta Murgia National Park. Ecological Modelling, 316(??):28–40, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003336.

Lin:2015:ABM

[LDM15b] Shengpan Lin, Mark H. DeVisser, and Joseph P. Messina. An agent-based model to simulate tsetse fly distribution and control techniques: a case study in Nguruman, Kenya. Ecological Modelling, 314(??):80–89, October 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500321X.

Lewis:2016:EEM

[LdMS^{+}16] K. A. Lewis, K. de Mutsert, J. Steenbeek, H. Peele, J. H. Cowan, and J. Buszowski. Employing ecosystem models and geographic information systems (GIS) to investigate the response of changing marsh edge on historical biomass of estuarine nekton in Barataria Bay, Louisiana, USA. Ecological Modelling, 331(??):129–141, July 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300059.

Lercari:2018:LTS

[LDO^{+}18] Diego Lercari, Omar Defeo, Leonardo Ortega, Luis Orlando, Ignacio Gianelli, and Eleonora Celentano. Long-term structural and functional changes driven by climate variability and fishery regimes in a sandy beach ecosystem. Ecological Modelling, 368(??):41–51, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304507.

Likasiri:2014:MME

[LDP14] Chulin Likasiri, Eakkapong Duangdai, and Radom Pongvuthithum. Mathematical model on the effects of global climate change and decreasing forest cover on seasonal rainfall in Northern Thailand. Ecological Modelling, 272(??):388–393, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013004985.

Langton:2014:SMC

[LDS14] R. Langton, I. M. Davies, and B. E. Scott. A simulation model coupling the behaviour and energetics of a breeding central place forager to assess the impact of environmental changes. Ecological Modelling, 273(??):31–43, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005292.

Long:2018:EER

[LDYL18] Yin Long, Liang Dong, Yoshikuni Yoshida, and Zhaoling Li. Evaluation of energy-related household carbon footprints in metropolitan areas of Japan. Ecological Modelling, 377(??):16–25, June 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303927.

Lecerf:2017:MEE

[Lec17] Antoine Lecerf. Methods for estimating the effect of litterbag mesh size on decomposition. Ecological Modelling, 362(??):65–68, October 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303745.

Lee:2014:RPC

[Lee14] Seungjun Lee. Resource pulses can increase power acquisition of an ecosystem. Ecological Modelling, 271 (??):21–31, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005662.

Lemos:2016:ASR

[Lem16] Ricardo T. Lemos. An alternative stock-recruitment function for age-structured models. Ecological Modelling, 341 (??):14–26, December 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016304318.

Lester:2011:ESC

[LF11] Rebecca E. Lester and Peter G. Fairweather. Ecosystem states: Creating a data-derived, ecosystem-scale ecological response model that is explicit in space and time. Ecological Modelling, 222(15):2690–2703, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002717.

Liu:2015:FSB

[LFH^{+}15] Zhu Liu, Kuishuang Feng, Klaus Hubacek, Sai Liang, Laura Diaz Anadon, Chao Zhang, and Dabo Guan. Four system boundaries for carbon accounts. Ecological Modelling, 318(??):118–125, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000496.

Legovic:2010:IMS

[LG10] Tarzan Legović and Sunčana Geček. Impact of maximum sustainable yield on independent populations. Ecological Modelling, 221(17):2108–2111, August 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000270X.

Legovic:2012:IMS

[LG12] Tarzan Legović and Sunčana Geček. Impact of maximum sustainable yield on mutualistic communities. Ecological Modelling, 230(??):63–72, April 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000178.

Lomas:2017:EAE

[LG17] Pedro L. Lomas and Mario Giampietro. Environmental accounting for ecosystem conservation: Linking societal and ecosystem metabolisms. Ecological Modelling, 346(??):10–19, February 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001630566X.

Lu:2012:AHS

[LGDW12] Chun Yan Lu, Wei Gu, Ai Hua Dai, and Hai Yan Wei. Assessing habitat suitability based on geographic information system (GIS) and fuzzy: a case study of Schisandra sphenanthera Rehd. et Wils. in Qinling Mountains, China. Ecological Modelling, 242(??):105–115, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012002748.

LaScala-Gruenewald:2019:SPP

[LGMLD19] Diana E. LaScala-Gruenewald, Rohan S. Mehta, Yu Liu, and Mark W. Denny. Sensory perception plays a larger role in foraging efficiency than heavy-tailed movement strategies. Ecological Modelling, 404(??):69–82, July 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300845.

Liu:2010:EIL

[LGY10] Yong Liu, Huaicheng Guo, and Pingjian Yang. Exploring the influence of lake water chemistry on chlorophyll a: a multivariate statistical model analysis. Ecological Modelling, 221(4):681–688, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009001884.

Lehmann:2015:FCD

[LH15] Sebastian Lehmann and Andreas Huth. Fast calibration of a dynamic vegetation model with minimum observation data. Ecological Modelling, 301(??):98–105, April 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000320.

Liang:2018:DWN

[LHB^{+}18] Hao Liang, Kelin Hu, William D. Batchelor, Wei Qin, and Baoguo Li. Developing a water and nitrogen management model for greenhouse vegetable production in China: Sensitivity analysis and evaluation. Ecological Modelling, 367(??):24–33, January 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304416.

Liao:2012:TAH

[LHH12] Wenjie Liao, Reinout Heijungs, and Gjalt Huppes. Thermodynamic analysis of human–environment systems: a review focused on industrial ecology. Ecological Modelling, 228(??):76–88, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012000051.

Lee:2017:MAE

[LHL^{+}17] Ingyu Lee, Hyundong Hwang, Jungwoo Lee, Nayoung Yu, Jinhuck Yun, and Hyunook Kim. Modeling approach to evaluation of environmental impacts on river water quality: a case study with Galing River, Kuantan, Pahang, Malaysia. Ecological Modelling, 353(??):167–173, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300674.

Li:2018:ENA

[LHL18] Jizhe Li, Guohe Huang, and Lirong Liu. Ecological network analysis for urban metabolism and carbon emissions based on input–output tables: a case study of Guangdong province. Ecological Modelling, 383(??):118–126, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018301698.

Luo:2015:ESE

[LHLW15] Xu Luo, Hong S. He, Yu Liang, and Zhiwei Wu. Evaluating simulated effects of succession, fire, and harvest for LANDIS PRO forest landscape model. Ecological Modelling, 297(??):1–10, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005584.

Lamonica:2016:MMD

[LHO^{+}16] Dominique Lamonica, Ulysse Herbach, Frédéric Orias, Bernard Clément, Sandrine Charles, and Christelle Lopes. Mechanistic modelling of daphnid-algae dynamics within a laboratory microcosm. Ecological Modelling, 320 (??):213–230, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015004494.

Li:2016:MAC

[LHP16] Guangqi Li, Sandy P. Harrison, and I. Colin Prentice. A model analysis of climate and CO_{2} controls on tree growth and carbon allocation in a semi-arid woodland. Ecological Modelling, 342(??):175–185, December 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305257.

Liu:2015:ECV

[LHR^{+}15] Min Liu, Honglin He, Xiaoli Ren, Xiaomin Sun, Guirui Yu, Shijie Han, Huimin Wang, and Guoyi Zhou. The effects of constraining variables on parameter optimization in carbon and water flux modeling over different forest ecosystems. Ecological Modelling, 303(??):30–41, May 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000460.

Liang:2015:SSP

[LHW^{+}15] Yu Liang, Hong S. He, Wen J. Wang, Jacob S. Fraser, Zhi-Wei Wu, and Jiawei Xu. The site-scale processes affect species distribution predictions of forest landscape models. Ecological Modelling, 300(??):89–101, March 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000198.

Li:2010:MRE

[LHZ10] Junmin Li, Ping Huang, and Renduo Zhang. Modeling the refuge effect of submerged macrophytes in ecological dynamics of shallow lakes: a new model of fish functional response. Ecological Modelling, 221(17):2076–2085, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002449.

Li:2011:ISR

[LHZN11] Y. P. Li, G. H. Huang, N. Zhang, and S. L. Nie. An inexact-stochastic with recourse model for developing regional economic-ecological sustainability under uncertainty. Ecological Modelling, 222(2):370–379, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008400.

Li:2010:CBM

[LHZS10] Tingting Li, Yao Huang, Wen Zhang, and Changchun Song. CH4MOD_{wetland}: a biogeophysical model for simulating methane emissions from natural wetlands. Ecological Modelling, 221(4):666–680, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009003858.

Lopez:2016:MMP

[LIM^{+}16] Leonardo López, Agustin Izquierdo, Darío Manzoli, Pablo Beldoménico, and Leonardo Giovanini. A myiasis model for Philornis torquans (Diptera: Muscidae) and Pitangus sulphuratus (Passeriformes: Tyrannidae). Ecological Modelling, 328(??):62–71, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300102.

Link:2010:ARE

[Lin10] Jason S. Link. Adding rigor to ecological network models by evaluating a set of pre-balance diagnostics: a plea for PREBAL. Ecological Modelling, 221(12):1580–1591, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010001468.

Lin:2015:TEF

[Lin15a] Henry Lin. Thermodynamic entropy fluxes reflect ecosystem characteristics and succession. Ecological Modelling, 298(??):75–86, February 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004931.

Lindim:2015:MIZ

[Lin15b] C. Lindim. Modeling the impact of zebra mussels (Dreissena polymorpha) on phytoplankton and nutrients in a lowland river. Ecological Modelling, 301(??):17–26, April 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015000241.

Link:2016:CAR

[Lin16] Jason S. Link. Corrigendum to “Adding rigor to ecological network models by evaluating a set of pre-balance diagnostics: a plea for PREBAL” [Ecol. Model. 221 (2010) 1582–1593]. Ecological Modelling, 337(??):348–349, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016302356.

Levin:2010:ASA

[LJ10] Gregor Levin and Martin Rudbeck Jepsen. Abolition of set-aside schemes, associated impacts on habitat structure and modelling of potential effects of cross-farm regulation. Ecological Modelling, 221(22):2728–2737, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004102.

Li:2013:MSB

[LJ13] Yan Li and Yan Jiao. Modeling seabird bycatch in the U.S. Atlantic pelagic longline fishery: Fixed year effect versus random year effect. Ecological Modelling, 260(??):36–41, July 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013001786.

Li:2014:FBS

[LJ14] Zhaofeng Li and Yichuan Jiang. Friction based social force model for social foraging of sheep flock. Ecological Modelling, 273(??):55–62, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005280.

Laughlin:2015:TCT

[LJ15a] Daniel C. Laughlin and Chaitanya Joshi. Theoretical consequences of trait-based environmental filtering for the breadth and shape of the niche: New testable hypotheses generated by the traitspace model. Ecological Modelling, 307(??):10–21, July 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001246.

Li:2015:ESS

[LJ15b] Yan Li and Yan Jiao. Evaluation of stocking strategies for endangered white abalone using a hierarchical demographic model. Ecological Modelling, 299(??):14–22, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014006048.

Li:2015:MSP

[LJ15c] Yan Li and Yan Jiao. Modeling spatial patterns of rare species using eigenfunction-based spatial filters: an example of modified delta model for zero-inflated data. Ecological Modelling, 299(??):51–63, March 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006115.

Lalechere:2017:NEP

[LJAD17] Etienne Lalechère, Franck Jabot, Frédéric Archaux, and Guillaume Deffuant. Non-equilibrium plant metapopulation dynamics challenge the concept of ancient/recent forest species. Ecological Modelling, 366(??):48–57, December 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016307645.

Logan:2012:LCF

[LJB12] J. David Logan, John Janovy, and Brittany E. Bunker. The life cycle and fitness domain of gregarine (Apicomplexa) parasites. Ecological Modelling, 233(??):31–40, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001251.

Lagergren:2012:ISD

[LJBS12] Fredrik Lagergren, Anna Maria Jönsson, Kristina Blennow, and Benjamin Smith. Implementing storm damage in a dynamic vegetation model for regional applications in Sweden. Ecological Modelling, 247(??):71–82, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200422X.

Losapio:2015:SDR

[LJCG15] Gianalberto Losapio, Ferenc Jordán, Marco Caccianiga, and Mauro Gobbi. Structure-dynamic relationship of plant– insect networks along a primary succession gradient on a glacier foreland. Ecological Modelling, 314(??):73–79, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003208.

Lagarrigues:2015:ABC

[LJLC15] Guillaume Lagarrigues, Franck Jabot, Valentine Lafond, and Benoit Courbaud. Approximate Bayesian computation to recalibrate individual-based models with population data: Illustration with a forest simulation model. Ecological Modelling, 306(??):278–286, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004475.

Livi:2011:IKS

[LJLO11] Carmen Maria Livi, Ferenc Jordán, Paola Lecca, and Thomas A. Okey. Identifying key species in ecosystems with stochastic sensitivity analysis. Ecological Modelling, 222(14):2542–2551, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004965.

Langseth:2014:EAM

[LJR14] Brian J. Langseth, Michael L. Jones, and Stephen C. Riley. The effect of adjusting model inputs to achieve mass balance on time-dynamic simulations in a food-web model of Lake Huron. Ecological Modelling, 273(??):44–54, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005267.

Lloyd-Jones:2014:GGM

[LJWN14] Luke R. Lloyd-Jones, You-Gan Wang, and Warwick J. Nash. Generalised growth models for aquatic species with an application to blacklip abalone (Haliotis rubra). Ecological Modelling, 272(??):311–322, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004791.

Linhoss:2013:DAS

[LKAL^{+}13] Anna C. Linhoss, Gregory A. Kiker, Matthew E. Aiello-Lammens, Ma. Librada Chu-Agor, Matteo Convertino, Rafael Muñoz-Carpena, Richard Fischer, and Igor Linkov. Decision analysis for species preservation under sea-level rise. Ecological Modelling, 263(??):264–272, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002767.

Liu:2015:TTB

[LKAY15] Junguo Liu, Giri Kattel, Hans Peter H. Arp, and Hong Yang. Towards threshold-based management of freshwater ecosystems in the context of climate change. Ecological Modelling, 318(??):265–274, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004256.

Levatic:2015:CSM

[LKDD15] Jurica Levatić, Dragi Kocev, Marko Debeljak, and Sašo Džeroski. Community structure models are improved by exploiting taxonomic rank with predictive clustering trees. Ecological Modelling, 306(??):294–304, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400492X.

Legovic:2010:MSY

[LKG10] Tarzan Legović, Jasminka Klanjšček, and Sunčana Geček. Maximum sustainable yield and species extinction in ecosystems. Ecological Modelling, 221(12):1569–1574, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010001699.

Li:2017:EHD

[LKG^{+}17] Zhilin Li, Aili Kang, Jiayin Gu, Yangang Xue, Yi Ren, Zhiwen Zhu, Peiqi Liu, Jianzhang Ma, and Guangshun Jiang. Effects of human disturbance on vegetation, prey and Amur tigers in Hunchun Nature Reserve, China. Ecological Modelling, 353(??):28–36, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303489.

Leauthaud:2018:MGF

[LKH^{+}18] C. Leauthaud, L. Kergoat, P. Hiernaux, M. Grippa, W. Musila, S. Duvail, and J. Albergel. Modelling the growth of floodplain grasslands to explore the impact of changing hydrological conditions on vegetation productivity. Ecological Modelling, 387(??):220–237, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303119.

Laine-Kaulio:2014:ERM

[LKKK^{+}14] Hanne Laine-Kaulio, Harri Koivusalo, Alexander S. Komarov, Mari Lappalainen, Samuli Launiainen, and Ari Laurén. Extending the ROMUL model to simulate the dynamics of dissolved and sorbed C and N compounds in decomposing boreal mor. Ecological Modelling, 272 (??):277–292, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004687.

Launiainen:2015:CBF

[LKLK15] Samuli Launiainen, Gabriel G. Katul, Ari Lauren, and Pasi Kolari. Coupling boreal forest CO_{2}, H_{2}O and energy flows by a vertically structured forest canopy — soil model with separate bryophyte layer. Ecological Modelling, 312(??):385–405, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002562.

Lagerwall:2014:GUS

[LKMCW14] Gareth Lagerwall, Gregory Kiker, Rafael Muñoz-Carpena, and Naiming Wang. Global uncertainty and sensitivity analysis of a spatially distributed ecological model. Ecological Modelling, 275(??):22–30, March 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005899.

Linkosalo:2013:NDR

[LKP13] Tapio Linkosalo, Pasi Kolari, and Jukka Pumpanen. New decomposition rate functions based on volumetric soil water content for the ROMUL soil organic matter dynamics model. Ecological Modelling, 263(??):109–118, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002500.

Li-kun:2017:DTD

[LkSXhX17] Yang Li-kun, Peng Sen, Zhao Xin-hua, and Li Xia. Development of a two-dimensional eutrophication model in an urban lake (China) and the application of uncertainty analysis. Ecological Modelling, 345(??):63–74, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307189.

Lafuite:2017:TDB

[LL17] A. S. Lafuite and M. Loreau. Time-delayed biodiversity feedbacks and the sustainability of social-ecological systems. Ecological Modelling, 351(??):96–108, May 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307128.

Lister:2018:MST

[LL18a] Andrew J. Lister and Laura P. Leites. Modeling and simulation of tree spatial patterns in an oak-hickory forest with a modular, hierarchical spatial point process framework. Ecological Modelling, 378(??):37–45, June 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301108.

Luo:2018:UEM

[LL18b] Xi Luo and Xuyong Li. Using the EFDC model to evaluate the risks of eutrophication in an urban constructed pond from different water supply strategies. Ecological Modelling, 372(??):1–11, March 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300413.

Laplanche:2019:AIH

[LLB^{+}19] Christophe Laplanche, Pedro M. Leunda, Laurie Boithias, José Ardaíz, and Francis Juanes. Advantages and insights from a hierarchical Bayesian growth and dynamics model based on salmonid electrofishing removal data. Ecological Modelling, 392(??):8–21, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303417.

Liu:2011:ABC

[LLC11] Yuedan Liu, Sang-Hee Lee, and Tae-Soo Chon. Analysis of behavioral changes of zebrafish (Danio rerio) in response to formaldehyde using self-organizing map and a hidden Markov model. Ecological Modelling, 222(14):2191–2201, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011000779.

Lavaud:2017:IES

[LLC^{+}17] Romain Lavaud, Megan K. La Peyre, Sandra M. Casas, Cédric Bacher, and Jérôme F. La Peyre. Integrating the effects of salinity on the physiology of the eastern oyster, Crassostrea virginica, in the northern Gulf of Mexico through a Dynamic Energy Budget model. Ecological Modelling, 363(??):221–233, November 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304027.

Li:2021:MEC

[LLC^{+}21] L. Li, G. Luo, X. Chen, Y. Li, G. Xu, H. Xu, and J. Bai. Modelling evapotranspiration in a Central Asian desert ecosystem. Ecological Modelling, 222(20–22):3680–3691, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004601.

Li:2010:EAI

[LLCR10] Linjun Li, Hongfang Lu, Daniel E. Campbell, and Hai Ren. Emergy algebra: Improving matrix methods for calculating transformities. Ecological Modelling, 221(3):411–422, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009006681.

Li:2011:MEU

[LLCR11] Linjun Li, Hongfang Lu, Daniel E. Campbell, and Hai Ren. Methods for estimating the uncertainty in emergy tableform models. Ecological Modelling, 222(15):2615–2622, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011002535.

Long:2012:IEN

[LLFL12] Zachary T. Long, Shawn J. Leroux, Thibault Faninger, and Michel Loreau. Interactive effects of nutrient enrichment and the manipulation of intermediate hosts by parasites on infection prevalence and food web structure. Ecological Modelling, 228(??):1–7, March 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006016.

Liao:2013:MPP

[LLH^{+}13] Jinbao Liao, Zhenqing Li, David E. Hiebeler, Magdy El-Bana, Gaby Deckmyn, and Ivan Nijs. Modelling plant population size and extinction thresholds from habitat loss and habitat fragmentation: Effects of neighbouring competition and dispersal strategy. Ecological Modelling, 268(??):9–17, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003761.

Liang:2017:IRS

[LLH^{+}17] Lu Liang, Xuecao Li, Yanbo Huang, Yuchu Qin, and Huabing Huang. Integrating remote sensing, GIS and dynamic models for landscape-level simulation of forest insect disturbance. Ecological Modelling, 354(??):1–10, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017302053.

Lim:2018:PFW

[LLKY18] R. B. H. Lim, J. H. Liew, J. T. B. Kwik, and D. C. J. Yeo. Predicting food web responses to biomanipulation using Bayesian Belief Network: Assessment of accuracy and applicability using in-situ exclosure experiments. Ecological Modelling, 384(??):308–315, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302229.

Li:2010:ALC

[LLL^{+}10] Hailong Li, Dihua Li, Ting Li, Qing Qiao, Jian Yang, and Hemin Zhang. Application of least-cost path model to identify a giant panda dispersal corridor network after the Wenchuan earthquake — case study of Wolong Nature Reserve in China. Ecological Modelling, 221(6):944–952, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009008369.

Liu:2017:DNR

[LLL17] Haoqi Liu, Weide Li, and Guanghui Lv. The design of nature reserves in the face of habitat loss. Ecological Modelling, 358(??):50–58, August 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301382.

Liu:2019:HNH

[LLL19] Haoqi Liu, Weide Li, and Guanghui Lv. How nonrandom habitat loss affects nature reserve planning strategies. Ecological Modelling, 397(??):39–46, April 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304253.

Li:2013:RSE

[LLM^{+}13] Zhouyuan Li, Xuehua Liu, Tianxiao Ma, De Kejia, Qingping Zhou, Bingquan Yao, and Tianlin Niu. Retrieval of the surface evapotranspiration patterns in the alpine grassland–wetland ecosystem applying SEBAL model in the source region of the Yellow River, China. Ecological Modelling, 270(??):64–75, December 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300433X.

Liu:2017:EZR

[LLP17] Xuehua Liu, Lin Liu, and Yu Peng. Ecological zoning for regional sustainable development using an integrated modeling approach in the Bohai Rim, China. Ecological Modelling, 353(??):158–166, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304240.

Li:2015:RBL

[LLQ^{+}15] Yi Li, Yangfan Li, Salman Qureshi, Martin Kappas, and Klaus Hubacek. On the relationship between landscape ecological patterns and water quality across gradient zones of rapid urbanization in coastal China. Ecological Modelling, 318(??):100–108, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000472.

Liao:2013:ESP

[LLQN13] Jinbao Liao, Zhenqing Li, Jan J. Quets, and Ivan Nijs. Effects of space partitioning in a plant species diversity model. Ecological Modelling, 251(??):271–278, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000136.

Li:2014:CCN

[LLT^{+}14] Zhengpeng Li, Shuguang Liu, Zhengxi Tan, Norman B. Bliss, Claudia J. Young, Tristram O. West, and Stephen M. Ogle. Comparing cropland net primary production estimates from inventory, a satellite-based model, and a process-based model in the Midwest of the United States. Ecological Modelling, 277(??):1–12, April 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000465.

Li:2017:SEM

[LLT^{+}17] Zhengpeng Li, Shuguang Liu, Zhengxi Tan, Terry L. Sohl, and Yiping Wu. Simulating the effects of management practices on cropland soil organic carbon changes in the Temperate Prairies Ecoregion of the United States from 1980 to 2012. Ecological Modelling, 365(??):68–79, December 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016305154.

Li:2014:ETS

[LLTQ14] Linjun Li, Hongfang Lu, David R. Tilley, and Guoyu Qiu. Effect of time scale on accounting for renewable emergy in ecosystems located in humid and arid climates. Ecological Modelling, 287(??):1–8, September 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002166.

Li:2015:RET

[LLTQ15] Linjun Li, Hongfang Lu, David R. Tilley, and Guoyu Qiu. Reprint of “Effect of time scale on accounting for renewable emergy in ecosystems located in humid and arid climates”. Ecological Modelling, 315(??):88–95, November 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003439.

Li:2015:ARH

[LLW^{+}15] Xue Li, Cong Lin, Yang Wang, Lingying Zhao, Na Duan, and Xudong Wu. Analysis of rural household energy consumption and renewable energy systems in zhangziying town of Beijing. Ecological Modelling, 318(??):184–193, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002082.

Li:2013:EGP

[LLY^{+}13] Xianglan Li, Shunlin Liang, Guirui Yu, Wenping Yuan, Xiao Cheng, Jiangzhou Xia, Tianbao Zhao, Jinming Feng, Zhuguo Ma, Mingguo Ma, Shaomin Liu, Jiquan Chen, Changliang Shao, Shenggong Li, Xudong Zhang, Zhiqiang Zhang, Shiping Chen, Takeshi Ohta, Andrej Varlagin, Akira Miyata, Kentaro Takagi, Nobuko Saiqusa, and Tomomichi Kato. Estimation of gross primary production over the terrestrial ecosystems in China. Ecological Modelling, 261–262(??):80–92, July 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001890.

Liu:2012:ESV

[LLZ12] Yong Liu, Jinchang Li, and Hong Zhang. An ecosystem service valuation of land use change in Taiyuan City, China. Ecological Modelling, 225(??):127–132, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005461.

Li:2015:ECI

[LLZ^{+}15] Yuzhao Li, Yong Liu, Lei Zhao, Alan Hastings, and Huaicheng Guo. Exploring change of internal nutrients cycling in a shallow lake: a dynamic nutrient driven phytoplankton model. Ecological Modelling, 313(??):137–148, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002744.

Li:2016:ELC

[LLZ^{+}16] Zhengpeng Li, Shuguang Liu, Xuesong Zhang, Tristram O. West, Stephen M. Ogle, and Naijun Zhou. Evaluating land cover influences on model uncertainties — a case study of cropland carbon dynamics in the mid-continent intensive campaign region. Ecological Modelling, 337(??):176–187, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016302447.

Laitila:2012:UML

[LM12] Jussi Laitila and Atte Moilanen. Use of many low-level conservation targets reduces high-level conservation performance. Ecological Modelling, 247(??):40–47, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004218.

Laitila:2013:ADM

[LM13] Jussi Laitila and Atte Moilanen. Approximating the dispersal of multi-species ecological entities such as communities, ecosystems or habitat types. Ecological Modelling, 259(??):24–29, June 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013001531.

Lorscheid:2016:DCC

[LM16] Iris Lorscheid and Matthias Meyer. Divide and conquer: Configuring submodels for valid and efficient analyses of complex simulation models. Ecological Modelling, 326(??):152–161, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005451.

Logofet:2019:BVCa

[LM19] Dmitrii O. Logofet and Alexander A. Maslov. Bilberry vs. cowberry in a Scots pine boreal forest: Exclusion or coexistence in a post-fire succession? Ecological Modelling, 401(??):134–143, June 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303259.

Lozano-Montes:2012:EES

[LMBL12] Hector M. Lozano-Montes, Russ Babcock, and Neil R. Loneragan. Exploring the effects of spatial closures in a temperate marine ecosystem in Western Australia: a case study of the western rock lobster (Panulirus cygnus) fishery. Ecological Modelling, 245(??):31–40, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002360.

Linhoss:2012:FPD

[LMCA^{+}12] Anna Cathey Linhoss, Rafael Muñoz-Carpena, Micheal S. Allen, Gregory Kiker, and Ketlhatlogile Mosepele. A flood pulse driven fish population model for the Okavango Delta, Botswana. Ecological Modelling, 228(??):27–38, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011006107.

Lebon:2014:DAC

[LMDG14] A. Lebon, L. Mailleret, Y. Dumont, and F. Grognard. Direct and apparent compensation in plant–herbivore interactions. Ecological Modelling, 290(??):192–203, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014001148.

Lintz:2011:QET

[LMGM11] Heather E. Lintz, Bruce McCune, Andrew N. Gray, and Katherine A. McCulloh. Quantifying ecological thresholds from response surfaces. Ecological Modelling, 222 (3):427–436, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005776.

Leo:2016:SEE

[LMGW16] Jennifer P. Leo, Thomas J. Minello, William E. Grant, and Hsiao-Hsuan Wang. Simulating environmental effects on brown shrimp production in the northern Gulf of Mexico. Ecological Modelling, 330(??):24–40, June 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300461.

Liu:2017:CHS

[LML17] Fang Liu, William J. McShea, and Diqiang Li. Correlating habitat suitability with landscape connectivity: a case study of Sichuan golden monkey in China. Ecological Modelling, 353(??):37–46, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303805.

Leach:2016:MIB

[LMR16] Katie Leach, W. Ian Montgomery, and Neil Reid. Modelling the influence of biotic factors on species distribution patterns. Ecological Modelling, 337(??):96–106, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302216.

Lashin:2012:CMG

[LMSK12] Sergey A. Lashin, Yury G. Matushkin, Valentin V. Suslov, and Nikolay A. Kolchanov. Computer modeling of genome complexity variation trends in prokaryotic communities under varying habitat conditions. Ecological Modelling, 224(1):124–129, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005230.

Larsen:2013:NMA

[LMT13] Janus Larsen, Christian Mohn, and Karen Timmermann. A novel model approach to bridge the gap between box models and classic 3D models in estuarine systems. Ecological Modelling, 266(??):19–29, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003268.

Larocque:2011:CCE

[LMY^{+}11] G. R. Larocque, D. Mailly, T.-X. Yue, M. Anand, C. Peng, C. Kazanci, M. Etterson, P. Goethals, S. E. Jørgensen, J. R. Schramski, E. J. B. McIntire, D. J. Marceau, B. Chen, G. Q. Chen, Z. F. Yang, B. Novotna, N. Luckai, J. S. Bhatti, J. Liu, A. Munson, A. M. Gordon, and J. C. Ascough II. Common challenges for ecological modelling: Synthesis of facilitated discussions held at the symposia organized for the 2009 conference of the International Society for Ecological Modelling in Quebec City, Canada, (October 6–9, 2009). Ecological Modelling, 222(14):2456–2468, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006836.

LaRue:2011:MPH

[LN11] Michelle A. LaRue and Clayton K. Nielsen. Modelling potential habitat for cougars in midwestern North America. Ecological Modelling, 222(3):897–900, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006228.

LaRue:2016:PVR

[LN16] Michelle A. LaRue and Clayton K. Nielsen. Population viability of recolonizing cougars in midwestern North America. Ecological Modelling, 321(??):121–129, February 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004573.

Lin:2017:AUK

[LNB17] L. Lin, J. S. Norman, and J. E. Barrett. Ammonia-uptake kinetics and domain-level contributions of bacteria and archaea to nitrification in temperate forest soils. Ecological Modelling, 362(??):111–119, October 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303769.

Loehle:2012:CCM

[Loe12] C. Loehle. A conditional choice model of habitat selection explains the source-sink paradox. Ecological Modelling, 235–236(??):59–66, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001640.

Loehle:2014:MME

[Loe14] Craig Loehle. A minimal model for estimating climate sensitivity. Ecological Modelling, 276(??):80–84, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000404.

Loehle:2018:DRT

[Loe18] Craig Loehle. Disequilibrium and relaxation times for species responses to climate change. Ecological Modelling, 384(??):23–29, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302035.

Logofet:2013:CMR

[Log13a] Dmitrii O. Logofet. Calamagrostis model revisited: Matrix calibration as a constraint maximization problem. Ecological Modelling, 254(??):71–79, April 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000458.

Logofet:2013:PMV

[Log13b] Dmitrii O. Logofet. Projection matrices in variable environments: λ_{1} in theory and practice. Ecological Modelling, 251(??):307–311, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000112.

Logofet:2016:EFL

[Log16] Dmitrii O. Logofet. Estimating the fitness of a local discrete-structured population: From uncertainty to an exact number. Ecological Modelling, 329(??):112–120, June 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016300448.

Logofet:2017:AMM

[Log17] Dmitrii O. Logofet. Aggregation may or may not eliminate reproductive uncertainty. Ecological Modelling, 363 (??):187–191, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303575.

Logofet:2019:BVCb

[Log19a] Dmitrii O. Logofet. Bilberry vs. cowberry in a Scots pine boreal forest. II. Alternate modes of prediction. Ecological Modelling, 405(??):43–50, August 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301310.

Logofet:2019:DAO

[Log19b] Dmitrii O. Logofet. Does averaging overestimate or underestimate population growth? It depends. Ecological Modelling, 411(??):Article 108744, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302522.

Liu:2013:CSD

[LOLA13] Xiaoping Liu, Jinpei Ou, Xia Li, and Bin Ai. Combining system dynamics and hybrid particle swarm optimization for land use allocation. Ecological Modelling, 257(??):11–24, May 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013001300.

Loptien:2011:SSS

[Löp11] Ulrike Löptien. Steady states and sensitivities of commonly used pelagic ecosystem model components. Ecological Modelling, 222(8):1376–1386, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000627.

Liu:2017:SIC

[LP17] Xuehua Liu and Young-Seuk Park. Special issue of China–Korea joint seminars on multi-disciplinary and multimethod approaches toward sustainable human and nature interactions. Ecological Modelling, 353(??):1–4, June 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017302430.

Larocque:2011:SPPb

[LPA^{+}11] Guy R. Larocque, Changhui Peng, Madhur Anand, Lubos Matejicek, and Alexander Komarov. Selected papers presented in oral and poster sessions at the 2009 conference of the International Society for Ecological Modelling (ISEM) in Quebec City, Canada, October 6–9, 2009. Ecological Modelling, 222(14):2469–2470, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002018.

Lobet:2014:MAD

[LPD14] Guillaume Lobet, Loïc Pagès, and Xavier Draye. A modeling approach to determine the importance of dynamic regulation of plant hydraulic conductivities on the water uptake dynamics in the soil-plant-atmosphere system. Ecological Modelling, 290(??):65–75, October 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005735.

Latombe:2011:LEI

[LPF11] Guillaume Latombe, Lael Parrott, and Daniel Fortin. Levels of emergence in individual based models: Coping with scarcity of data and pattern redundancy. Ecological Modelling, 222(9):1557–1568, May 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001013.

Lessio:2015:MMF

[LPPA15] Federico Lessio, Alessandro Portaluri, Francesco Paparella, and Alberto Alma. A mathematical model of flavescence dorée epidemiology. Ecological Modelling, 312 (??):41–53, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002112.

Liang:2010:EAP

[LPS^{+}10] Liyin Liang, Shaolin Peng, Junmei Sun, Leiyi Chen, and Yuexiu Cao. Estimation of annual potential evapotranspiration at regional scale based on the effect of moisture on soil respiration. Ecological Modelling, 221(22): 2668–2674, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004059.

Liston:2016:MSH

[LPSC16] Glen E. Liston, Craig J. Perham, Richard T. Shideler, and April N. Cheuvront. Modeling snowdrift habitat for polar bear dens. Ecological Modelling, 320(??):114–134, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004329.

Liang:2018:EPA

[LPT^{+}18] Wanwan Liang, Monica Papeş, Liem Tran, Jerome Grant, Robert Washington-Allen, Scott Stewart, and Gregory Wiggins. The effect of pseudo-absence selection method on transferability of species distribution models in the context of non-adaptive niche shift. Ecological Modelling, 388(??):1–9, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303156.

Lindim:2011:AST

[LPV11] C. Lindim, J. L. Pinho, and J. M. P. Vieira. Analysis of spatial and temporal patterns in a large reservoir using water quality and hydrodynamic modeling. Ecological Modelling, 222(14):2485–2494, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003741.

Liu:2012:SRE

[LPX^{+}12] Zelin Liu, Changhui Peng, Wenhua Xiang, Xiangwen Deng, DaLun Tian, Meifang Zhao, and Guirui Yu. Simulations of runoff and evapotranspiration in Chinese fir plantation ecosystems using artificial neural networks. Ecological Modelling, 226(??):71–76, February 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005655.

Liang:2019:UBC

[LQW^{+}19] Zhongyao Liang, Song S. Qian, Sifeng Wu, Huili Chen, Yong Liu, Yanhong Yu, and Xuan Yi. Using Bayesian change point model to enhance understanding of the shifting nutrients-phytoplankton relationship. Ecological Modelling, 393(??):120–126, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304198.

Lee:2016:DRH

[LR16] Tamsin E. Lee and David L. Roberts. Devaluing rhino horns as a theoretical game. Ecological Modelling, 337 (??):73–78, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016302228.

Ludovisi:2012:TEE

[LRB12] Alessandro Ludovisi, Leonilde Roselli, and Alberto Basset. Testing the effectiveness of exergy-based tools on a seasonal succession in a coastal lagoon by using a size distribution approach. Ecological Modelling, 245(??):125–135, October 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000774.

LaBeau:2014:EFU

[LRM^{+}14] Meredith B. LaBeau, Dale M. Robertson, Alex S. Mayer, Bryan C. Pijanowski, and David A. Saad. Effects of future urban and biofuel crop expansions on the riverine export of phosphorus to the Laurentian Great Lakes. Ecological Modelling, 277(??):27–37, April 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000593.

Langseth:2012:MSI

[LRZ12] Brian J. Langseth, Mark Rogers, and Hongyan Zhang. Modeling species invasions in Ecopath with Ecosim: an evaluation using Laurentian Great Lakes models. Ecological Modelling, 247(??):251–261, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004267.

Liedloff:2010:PTC

[LS10] Adam C. Liedloff and Carl S. Smith. Predicting a ‘tree change’ in Australia’s tropical savannas: Combining different types of models to understand complex ecosystem behaviour. Ecological Modelling, 221(21):2565–2575, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010003777.

Lin:2013:CDI

[LS13] Yangchen Lin and William J. Sutherland. Color and degree of interspecific synchrony of environmental noise affect the variability of complex ecological networks. Ecological Modelling, 263(??):162–173, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002573.

Li:2018:INC

[LS18] Jie Li and Xuzhu Shen. An improved neutral community model for temporal observations in microbial communities. Ecological Modelling, 388(??):108–114, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303247.

Lavallee:2019:SIB

[LSA^{+}19] François Lavallée, Charline Smadi, Isabelle Alvarez, Björn Reineking, François-Marie Martin, Fanny Dommanget, and Sophie Martin. A stochastic individual-based model for the growth of a stand of Japanese knotweed including mowing as a management technique. Ecological Modelling, 413(??):Article 108828, December 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303369.

Liu:2013:LPE

[LSGT13] Chun Liu, Richard M. Sibly, Volker Grimm, and Pernille Thorbek. Linking pesticide exposure and spatial dynamics: an individual-based model of wood mouse (Apodemus sylvaticus) populations in agricultural landscapes. Ecological Modelling, 248(??):92–102, January 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004826.

Le:2015:IIA

[LSH15] Hai Dinh Le, Carl Smith, and John Herbohn. Identifying interactions among reforestation success drivers: a case study from the Philippines. Ecological Modelling, 316(??):62–77, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003531.

Lai:2018:BGB

[LSJ^{+}18] Tin-Yu Lai, Jani Salminen, Jukka-Pekka Jäppinen, Saija Koljonen, Laura Mononen, Emmi Nieminen, Petteri Vihervaara, and Soile Oinonen. Bridging the gap between ecosystem service indicators and ecosystem accounting in Finland.

Ecological Modelling, 377(??):51–65, June 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302491.

Lai:2019:EBG

[LSJ^{+}19] Tin-Yu Lai, Jani Salminen, Jukka-Pekka Jäppinen, Saija Koljonen, Laura Mononen, Emmi Nieminen, Petteri Vihervaara, and Soile Oinonen. Erratum to “Bridging the gap between ecosystem service indicators and ecosystem accounting in Finland” [Ecol. Model. 377 (2018) 51–65]. Ecological Modelling, 408(??):Article 108707, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301954.

Li:2017:DRE

[LSJG17] Xueyan Li, Yali Si, Luyan Ji, and Peng Gong. Dynamic response of East Asian greater white-fronted geese to changes of environment during migration: Use of multitemporal species distribution model. Ecological Modelling, 360(??):70–79, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307712.

Luce:2010:PBS

[LSL10] James J. Luce, Russell Steele, and Michel F. Lapointe. A physically based statistical model of sand abrasion effects on periphyton biomass. Ecological Modelling, 221 (2):353–361, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006413.

Lutz:2019:OME

[LSM19] Femke Lutz, Jetse J. Stoorvogel, and Christoph Müller. Options to model the effects of tillage on N_{2}O emissions at the global scale. Ecological Modelling, 392(??):212–225, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304034.

Lindner:2010:TTS

[LSP^{+}10] Marcus Lindner, Tommi Suominen, Taru Palosuo, Jordi Garcia-Gonzalo, Peter Verweij, Sergey Zudin, and Risto Päivinen. ToSIA — a tool for sustainability impact assessment of forest-wood-chains. Ecological Modelling, 221 (18):2197–2205, September 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009005651.

Lukes:2013:RBF

[LSR13] Petr Lukeš, Pauline Stenberg, and Miina Rautiainen. Relationship between forest density and albedo in the boreal zone. Ecological Modelling, 261–262(??):74–79, July 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002135.

Larsen:2016:MPU

[LSR16] Lars-Henrik Larsen, Kjetil Sagerup, and Silje Ramsvatn. The mussel path — using the contaminant tracer, Ecotracer, in Ecopath to model the spread of pollutants in an Arctic marine food web. Ecological Modelling, 331(??):77–85, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015004895.

Lausch:2015:DPB

[LSS^{+}15] Angela Lausch, Christoph Salbach, Andreas Schmidt, Daniel Doktor, Ines Merbach, and Marion Pause. Deriving phenology of barley with imaging hyperspectral remote sensing. Ecological Modelling, 295(??):123–135, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004670.

Lausch:2015:DML

[LST15] Angela Lausch, Andreas Schmidt, and Lutz Tischendorf. Data mining and linked open data — new perspectives for data analysis in environmental research. Ecological Modelling, 295(??):5–17, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004335.

Li:2010:EWM

[LSZC10] Yangfan Li, Xiang Sun, Xiaodong Zhu, and Huhua Cao. An early warning method of landscape ecological security in rapid urbanizing coastal areas and its application in Xiamen, China. Ecological Modelling, 221(19):2251–2260, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002322.

LeCorre:2015:EPC

[LT15] O. Le Corre and L. Truffet. Emergy paths computation from interconnected energy system diagram. Ecological Modelling, 313(??):181–200, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002793.

Louis:2016:MCF

[LTGD16] Marceau Louis, Etienne Toffin, Jean-Claude Gregoire, and Jean-Louis Deneubourg. Modelling collective foraging in endemic bark beetle populations. Ecological Modelling, 337(??):188–199, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302502.

Li:2013:CES

[LTL^{+}13] Linjun Li, David R. Tilley, Hongfang Lu, Hai Ren, and Guoyu Qiu. Comparison of an energy systems mini-model to a process-based eco-physiological model for simulating forest growth. Ecological Modelling, 263(??):32–41, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002342.

LeCorre:2015:OTB

[LTL15] O. Le Corre, L. Truffet, and C. Lahlou. Odum– Tennenbaum–Brown calculus vs emergy and co-emergy analysis. Ecological Modelling, 302(??):9–12, April 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001500040X. See reply

[Ten15b].

Langhammer:2019:ALG

[LTL^{+}19] Maria Langhammer, Jule Thober, Martin Lange, Karin Frank, and Volker Grimm. Agricultural landscape generators for simulation models: a review of existing solutions and an outline of future directions. Ecological Modelling, 393(??):135–151, February 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304216.

LaMorgia:2017:UHC

[LV17] Valentina La Morgia and Ezio Venturino. Understanding hybridization and competition processes between hare species: Implications for conservation and management on the basis of a mathematical model. Ecological Modelling, 364(??):13–24, November 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017301539.

Laforge:2015:PFM

[LVB^{+}15] Michel P. Laforge, Eric Vander Wal, Ryan K. Brook, Erin M. Bayne, and Philip D. McLoughlin. Processfocussed, multi-grain resource selection functions. Ecological Modelling, 305(??):10–21, June 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000964.

Li:2014:IDM

[LVLS14] Sen Li, Sophie O. Vanwambeke, Alain M. Licoppe, and Niko Speybroeck. Impacts of deer management practices on the spatial dynamics of the tick Ixodes ricinus: a scenario analysis. Ecological Modelling, 276(??):1–13, March 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000088.

Lin:2015:CSH

[LVPF15] Henry Lin, Hans-Jörg Vogel, Jonathan Phillips, and Brian D. Fath. Complexity of soils and hydrology in ecosystems. Ecological Modelling, 298(??):1–3, February 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005894.

Lopes:2015:ASL

[LVV^{+}15] José Fortes Lopes, Nuno Vaz, Leandro Vaz, Juan A. Ferreira, and João Miguel Dias. Assessing the state of the lower level of the trophic web of a temperate lagoon, in situations of light or nutrient stress: a modeling study. Ecological Modelling, 313(??):59–76, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002586.

Liu:2011:ECE

[LVZ^{+}11] Jinxun Liu, James E. Vogelmann, Zhiliang Zhu, Carl H. Key, Benjamin M. Sleeter, David T. Price, Jing M. Chen, Mark A. Cochrane, Jeffery C. Eidenshink, Stephen M. Howard, Norman B. Bliss, and Hong Jiang. Estimating California ecosystem carbon change using process model and land cover disturbance data: 1951–2000. Ecological Modelling, 222(14):2333–2341, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001943.

Lin:2014:DDN

[LW14] Laurence Lin and J. R. Webster. Detritus decomposition and nutrient dynamics in a forested headwater stream. Ecological Modelling, 293(??):58–68, December 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013006005.

Loehle:2017:DPT

[LW17] Craig Loehle and Philip Weatherford. Detecting population trends with historical data: Contributions of volatility, low detectability, and metapopulation turnover to potential sampling bias. Ecological Modelling, 362(??):13–18, October 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017303794.

Liang:2018:PME

[LWC^{+}18] Zhongyao Liang, Sifeng Wu, Huili Chen, Yanhong Yu, and Yong Liu. A probabilistic method to enhance understanding of nutrient limitation dynamics of phytoplankton. Ecological Modelling, 368(??):404–410, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304465.

Li:2013:ARB

[LWGH13] Yu Li, Anya M. Waite, Gideon Gal, and Matthew R. Hipsey. An analysis of the relationship between phytoplankton internal stoichiometry and water column N:P ratios in a dynamic lake environment. Ecological Modelling, 252(??):196–213, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003006.

Lenz-Wiedemann:2010:DTC

[LWKS10] V. I. S. Lenz-Wiedemann, C. W. Klar, and K. Schneider. Development and test of a crop growth model for application within a global change decision support system. Ecological Modelling, 221(2):314–329, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000900667X.

Li:2017:DCD

[LWL^{+}17a] Hu Li, Ligang Wang, Jianzheng Li, Maofang Gao, Jing Zhang, Jianfeng Zhang, Jianjun Qiu, Jia Deng, Changsheng Li, and Steve Frolking. The development of China– DNDC and review of its applications for sustaining Chinese agriculture. Ecological Modelling, 348(??):1–13, March 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303258.

Li:2017:LUC

[LWL^{+}17b] Zhouyuan Li, Wenzhao Wu, Xuehua Liu, Brian D. Fath, Hailian Sun, Xinchao Liu, Xinru Xiao, and Jun Cao. Land use/cover change and regional climate change in an arid grassland ecosystem of Inner Mongolia, China. Ecological Modelling, 353(??):86–94, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302630.

Liu:2013:SDM

[LWNG13] Canran Liu, Matt White, Graeme Newell, and Peter Griffioen. Species distribution modelling for conservation planning in Victoria, Australia. Ecological Modelling, 249(??):68–74, January 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003286.

Li:2018:UMU

[LWTZ18] Yanxian Li, Xinjing Wang, Xin Tian, and Yan Zhang. Understanding the mechanism of urban material metabolism with ecological network analysis: an experimental study of Wuxi, China. Ecological Modelling, 367(??):58–67, January 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001730323X.

Lin:2011:MEF

[LWY^{+}11] Yu-Pin Lin, Cheng-Long Wang, Hsiao-Hsuan Yu, Chung-Wei Huang, Yung-Chieh Wang, Yu-Wen Chen, and Wei-Yao Wu. Monitoring and estimating the flow conditions and fish presence probability under various flow conditions at reach scale using genetic algorithms and kriging methods. Ecological Modelling, 222(3):762–775, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006241.

Li:2016:VPF

[LXS^{+}16] Qianyu Li, Jianyang Xia, Zheng Shi, Kun Huang, Zhenggang Du, Guanghui Lin, and Yiqi Luo. Variation of parameters in a flux-based ecosystem model across 12 sites of terrestrial ecosystems in the conterminous USA. Ecological Modelling, 336(??):57–69, September 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630196X.

Li:2011:QSW

[LY11] Y. Li and Z. F. Yang. Quantifying the sustainability of water use systems: Calculating the balance between network efficiency and resilience. Ecological Modelling, 222(10):1771–1780, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001086.

Lee:2016:IVP

[LY16] Soonmi Lee and Sinjae Yoo. Interannual variability of the phytoplankton community by the changes in vertical mixing and atmospheric deposition in the Ulleung Basin, East Sea: a modelling study. Ecological Modelling, 322(??):31–47, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500544X.

Liu:2015:DMS

[LYC^{+}15a] Gengyuan Liu, Zhifeng Yang, Bin Chen, Shan Gao, Meirong Su, and Yan Zhang. Designing a multi-species spatially explicit nature reserve network construction framework based on extinction probability: a case study of wuyishan city. Ecological Modelling, 318(??):109–117, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005936.

Liu:2015:ENP

[LYC^{+}15b] Gengyuan Liu, Zhifeng Yang, Bin Chen, Lixiao Zhang, Yan Zhang, and Meirong Su. An ecological network perspective in improving reserve design and connectivity: a case study of Wuyishan Nature Reserve in China. Ecological Modelling, 306(??):185–194, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004700.

Liu:2014:EBD

[LYCU14] Gengyuan Liu, Zhifeng Yang, Bin Chen, and Sergio Ulgiati. Emergy-based dynamic mechanisms of urban development, resource consumption and environmental impacts. Ecological Modelling, 271(??):90–102, January 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004146.

Liu:2011:END

[LYCZ11] G. Y. Liu, Z. F. Yang, B. Chen, and Y. Zhang. Ecological network determination of sectoral linkages, utility relations and structural characteristics on urban ecological economic system. Ecological Modelling, 222(15):2825–2834, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011003127.

Liu:2013:MTB

[LYCZ13] Gengyuan Liu, Zhifeng Yang, Bin Chen, and Lixiao Zhang. Modelling a thermodynamic-based comparative framework for urban sustainability: Incorporating economic and ecological losses into emergy analysis. Ecological Modelling, 252(??):280–287, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000501.

Li:2017:EDM

[LYD^{+}17] Zhuoting Li, J. Y. Yang, C. F. Drury, X. M. Yang, W. D. Reynolds, Xiaogang Li, and Chunsheng Hu. Evaluation of the DNDC model for simulating soil temperature, moisture and respiration from monoculture and rotational corn, soybean and winter wheat in Canada. Ecological Modelling, 360(??):230–243, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308456.

Li:2015:HBC

[LYJ^{+}15] Xi Li, Jagadeesh Yeluripati, Edward O. Jones, Yoshitaka Uchida, and Ryusuke Hatano. Hierarchical Bayesian calibration of nitrous oxide (N_{2}O) and nitrogen monoxide (NO) flux module of an agro-ecosystem model: ECOSSE. Ecological Modelling, 316(??):14–27, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003269.

Liu:2017:ESL

[LYL^{+}17] Shiliang Liu, Yijie Yin, Xuehua Liu, Fangyan Cheng, Juejie Yang, Junran Li, Shikui Dong, and Annah Zhu. Ecosystem services and landscape change associated with plantation expansion in a tropical rainforest region of Southwest China. Ecological Modelling, 353(??):129–138, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300680.

Larocque:2011:SPPa

[LYP^{+}11] Guy R. Larocque, Tian-Xiang Yue, Changhui Peng, Jagtar Bhatti, Matthew A. Etterson, Caner Kazanci, and Peter Goethals. Selected papers presented in symposia organized at the 2009 conference of the International Society for Ecological Modelling (ISEM) in Quebec City, Canada, October 6–9, 2009. Ecological Modelling, 222(14):2169–2171, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011002006.

Li:2015:EUS

[LYSW15] Feng Li, Yaping Ye, Bowen Song, and Rusong Wang. Evaluation of urban suitable ecological land based on the minimum cumulative resistance model: a case study from Changzhou, China. Ecological Modelling, 318 (??):194–203, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004177.

Liang:2010:GMF

[LZ10] Jingjing Liang and Mo Zhou. A geospatial model of forest dynamics with controlled trend surface. Ecological Modelling, 221(19):2339–2352, September 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002851.

Lundstrom:2017:PEB

[LZB17] Niklas L. P. Lundström, Hong Zhang, and Åke Brännström. Pareto-efficient biological pest control enable high efficacy at small costs. Ecological Modelling, 364(??):89–97, November 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017300856.

Lindh:2014:PDD

[LZF^{+}14] Magnus Lindh, Lai Zhang, Daniel Falster, Oskar Franklin, and Åke Brännström. Plant diversity and drought: the role of deep roots. Ecological Modelling, 290(??):85–93, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002452.

Liao:2019:WFE

[LZJ^{+}19] Xiawei Liao, Xu Zhao, Yu Jiang, Yu Liu, Yujun Yi, and Martin R. Tillotson. Water footprint of the energy sector in China’s two megalopolises. Ecological Modelling, 391(??):9–15, January 10, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303326.

Li:2018:MSE

[LZL^{+}18] Xing Li, Qian Zhang, Yu Liu, Jia Song, and Feng Wu. Modeling social–economic water cycling and the water– land nexus: a framework and an application. Ecological Modelling, 390(??):40–50, December 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303399.

Li:2013:MCD

[LZLC13] Chaofan Li, Chi Zhang, Geping Luo, and Xi Chen. Modeling the carbon dynamics of the dryland ecosystems in Xinjiang, China from 1981 to 2007 — the spatiotemporal patterns and climate controls. Ecological Modelling, 267(??):148–157, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013002913.

Li:2017:IEM

[LZS^{+}17] Xi Li, Yi Zheng, Zan Sun, Yong Tian, Chunmiao Zheng, Jie Liu, Shaomin Liu, and Ziwei Xu. An integrated ecohydrological modeling approach to exploring the dynamic interaction between groundwater and phreatophytes. Ecological Modelling, 356(??):127–140, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306731.

Liu:2017:NMT

[LZW^{+}17] Haifei Liu, Jie Zhang, Hongda Wang, Yu Ding, and Yujun Yi. Numerical modeling of the tidal wave run-up and the eelgrass habitat at the Laizhou Bay. Ecological Modelling, 360(??):378–386, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306184.

Liu:2019:TDN

[LZYH19] Haidong Liu, Zhongquan C. Zheng, Bryan Young, and Ted D. Harris. Three-dimensional numerical modeling of the cyanobacterium Microcystis transport and its population dynamics in a large freshwater reservoir. Ecological Modelling, 398(??):20–34, April 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300420.

Li:2014:ARS

[LZZ^{+}14] Yan Li, Qingguo Zhou, Jian Zhou, Gaofeng Zhang, Chong Chen, and Jing Wang. Assimilating remote sensing information into a coupled hydrology-crop growth model to estimate regional maize yield in arid regions. Ecological Modelling, 291(??):15–27, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003408.

Li:2016:EGP

[LZZP16] Xiran Li, Zaichun Zhu, Hui Zeng, and Shilong Piao. Estimation of gross primary production in China (1982– 2010) with multiple ecosystem models. Ecological Modelling, 324(??):33–44, March 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016000065.

Liu:2017:NTS

[LZZW17] Dongya Liu, Xinqi Zheng, Chunxiao Zhang, and Hongbin Wang. A new temporal–spatial dynamics method of simulating land-use change. Ecological Modelling, 350(??):1–10, April 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016307141.

Martin:2010:IMC

[MA10] C. Martin and E. Ayesa. An integrated Monte Carlo methodology for the calibration of water quality models. Ecological Modelling, 221(22):2656–2667, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004035.

Mestre:2019:GGB

[MAB19] F. Mestre, F. Ascensão, and A. M. Barbosa. gDefrag: a graph-based tool to help defragmenting landscapes divided by linear infrastructures. Ecological Modelling, 392(??):1–5, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303351.

Martins:2012:WEA

[MAC^{+}12] Irene Martins, Maria L. Arosa, Ricardo S. Ceia, Sandra Parejo, Jaime A. Ramos, and Christian Damgaard. The winter energetics of the Azores bullfinch and the implications for the restoration of its native laurel forest habitat. Ecological Modelling, 231(??):80–86, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000269.

Masi:2014:PRF

[MAC14] M. D. Masi, C. H. Ainsworth, and D. Chagaris. A probabilistic representation of fish diet compositions from multiple data sources: a Gulf of Mexico case study. Ecological Modelling, 284(??):60–74, July 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001860.

Mateo:2019:LOH

[MAFG^{+}19] Rubén G. Mateo, María José Aroca-Fernández, Aitor Gastón, Virgilio Gómez-Rubio, Santiago Saura, and Juan Ignacio García-Viñas. Looking for an optimal hierarchical approach for ecologically meaningful niche modelling. Ecological Modelling, 409(??):Article 108735, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019302431.

Musiani:2010:HHS

[MAM^{+}10] Marco Musiani, Sk. Morshed Anwar, Gregory J. McDermid, Mark Hebblewhite, and Danielle J. Marceau. How humans shape wolf behavior in Banff and Kootenay National Parks, Canada. Ecological Modelling, 221(19):2374–2387, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010003078.

Martin:2012:ISH

[MAM^{+}12] Beatriz Martín, Juan Carlos Alonso, Carlos A. Martín, Carlos Palacín, Marina Magaña, and Javier Alonso. Influence of spatial heterogeneity and temporal variability in habitat selection: a case study on a great bustard metapopulation. Ecological Modelling, 228(??):39–48, March 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006120.

Moreno-Amat:2015:IMC

[MAMNL^{+}15] Elena Moreno-Amat, Rubén G. Mateo, Diego Nieto-Lugilde, Naia Morueta-Holme, Jens-Christian Svenning, and Ignacio García-Amorena. Impact of model complexity on crosstemporal transferability in Maxent species distribution models: an assessment using paleobotanical data. Ecological Modelling, 312(??):308–317, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002483.

Malahlela:2019:MSD

[MAO19] Oupa E. Malahlela, Clement Adjorlolo, and Jane M. Olwoch. Mapping the spatial distribution of Lippia javanica (Burm. f.) Spreng using Sentinel-2 and SRTM-derived topographic data in malaria endemic environment. Ecological Modelling, 392(??):147–158, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304083.

Meli:2013:PLC

[MAP^{+}13] Mattia Meli, Apolline Auclerc, Annemette Palmqvist, Valery E. Forbes, and Volker Grimm. Population-level consequences of spatially heterogeneous exposure to heavy metals in soil: an individual-based model of springtails. Ecological Modelling, 250(??):338–351, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005479.

Marcot:2012:MEP

[Mar12] Bruce G. Marcot. Metrics for evaluating performance and uncertainty of Bayesian network models. Ecological Modelling, 230(??):50–62, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000245.

Marcot:2017:CQT

[Mar17] Bruce G. Marcot. Common quandaries and their practical solutions in Bayesian network modeling. Ecological Modelling, 358(??):1–9, August 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308134.

Matter:2013:IDN

[Mat13] Stephen F. Matter. Increases and decreases in network synchrony associated with local population extinction and recolonization. Ecological Modelling, 266(??):68–72, September 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013003141.

Mateus:2017:MSC

[Mat17] M. Mateus. Milking spherical cows — yet another facet of model complexity. Ecological Modelling, 354(??):172–175, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017300935.

Muko:2014:IBM

[MATS14] Soyoka Muko, Seiji Arakaki, Reiko Tamai, and Kazuhiko Sakai. An individual-based model for population viability analysis of the brooding coral Seriatopora hystrix. Ecological Modelling, 277(??):68–76, April 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000684.

Mongin:2014:IEP

[MB14] Mathieu Mongin and Mark Baird. The interacting effects of photosynthesis, calcification and water circulation on carbon chemistry variability on a coral reef flat: a modelling study. Ecological Modelling, 284(??):19–34, July 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001859.

Mohammed:2017:MHM

[MB17] A. Mohammed and A. O. Babatunde. Modelling heavy metals transformation in vertical flow constructed wetlands. Ecological Modelling, 354(??):62–71, June 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300042.

Movedi:2019:DGC

[MBA^{+}19] Ermes Movedi, Gianni Bellocchi, Giovanni Argenti, Livia Paleari, Fosco Vesely, Nicolina Staglianò, Camilla Dibari, and Roberto Confalonieri. Development of generic crop models for simulation of multi-species plant communities in mown grasslands. Ecological Modelling, 401(??):111–128, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019300894.

Malavert:2017:TDR

[MBBA17] Cristian Malavert, Diego Batlla, and Roberto L. Benech-Arnold. Temperature-dependent regulation of induction into secondary dormancy of Polygonum aviculare L. seeds: a quantitative analysis. Ecological Modelling, 352(??):128–138, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017302065.

Maiorano:2012:CMA

[MBD^{+}12] Andrea Maiorano, Simone Bregaglio, Marcello Donatelli, Davide Fumagalli, and Antonio Zucchini. Comparison of modelling approaches to simulate the phenology of the European corn borer under future climate scenarios. Ecological Modelling, 245(??):65–74, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001615.

Maar:2018:RSP

[MBD^{+}18] Marie Maar, Momme Butenschön, Ute Daewel, Anja Eggert, Wei Fan, Solfrid S. Hjøllo, Marc Hufnagl, Martin Huret, Rubao Ji, Geneviève Lacroix, Myron A. Peck, Hagen Radtke, Sévrine Sailley, Matteo Sinerchia, Morten D. Skogen, Morgane Travers-Trolet, Tineke A. Troost, and Karen van de Wolfshaar. Responses of summer phytoplankton biomass to changes in top-down forcing: Insights from comparative modelling. Ecological Modelling, 376(??):54–67, May 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018300851.

Mazzoleni:2010:MEL

[MBG^{+}10] Stefano Mazzoleni, Giuliano Bonanomi, Francesco Giannino, Guido Incerti, Stefan C. Dekker, and Max Rietkerk. Modelling the effects of litter decomposition on tree diversity patterns. Ecological Modelling, 221(23):2784–2792, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004023.

Marchand:2017:SMM

[MBG17] Philippe Marchand, Morgan Boenke, and David M. Green. A stochastic movement model reproduces patterns of site fidelity and long-distance dispersal in a population of Fowler’s toads (Anaxyrus fowleri). Ecological Modelling, 360(??):63–69, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630850X.

Manusch:2012:TMD

[MBHW12] Corina Manusch, Harald Bugmann, Caroline Heiri, and Annett Wolf. Tree mortality in dynamic vegetation models — a key feature for accurately simulating forest properties. Ecological Modelling, 243(??):101–111, September 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002803.

Mammeri:2014:HCD

[MBLC14] Y. Mammeri, J. B. Burie, M. Langlais, and A. Calonnec. How changes in the dynamic of crop susceptibility and cultural practices can be used to better control the spread of a fungal pathogen at the plot scale? Ecological Modelling, 290(??):178–191, October 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001021.

Mintram:2018:API

[MBM^{+}18] Kate S. Mintram, A. Ross Brown, Samuel K. Maynard, Chun Liu, Sarah-Jane Parker, Charles R. Tyler, and Pernille Thorbek. Assessing population impacts of toxicant-induced disruption of breeding behaviours using an individualbased model for the three-spined stickleback. Ecological Modelling, 387(??):107–117, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302916.

Marvuglia:2013:SSC

[MBRR13] Antonino Marvuglia, Enrico Benetto, Gordon Rios, and Benedetto Rugani. SCALE: Software for CALculating emergy based on life cycle inventories. Ecological Modelling, 248(??):80–91, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004796.

Mao:2014:TDM

[MBSF14] Zhun Mao, Franck Bourrier, Alexia Stokes, and Thierry Fourcaud. Three-dimensional modelling of slope stability in heterogeneous montane forest ecosystems. Ecological Modelling, 273(??):11–22, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004845.

Maclean:2012:HRM

[MBSW12] Ilya M. D. Maclean, Jonathan J. Bennie, Amanda J. Scott, and Robert J. Wilson. A high-resolution model of soil and surface water conditions. Ecological Modelling, 237–238(??):109–119, July 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001457.

McDonald:2012:DTB

[MBUM12] C. P. McDonald, V. Bennington, N. R. Urban, and G. A. McKinley. 1-D test-bed calibration of a 3-D Lake Superior biogeochemical model. Ecological Modelling, 225(??):115–126, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005631.

Manusch:2014:SSP

[MBW14] Corina Manusch, Harald Bugmann, and Annett Wolf. Sensitivity of simulated productivity to soil characteristics and plant water uptake along drought gradients in the Swiss Alps. Ecological Modelling, 282(??):25–34, June 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014001306.

McBride:2011:NHM

[MC11] Graham B. McBride and Steven C. Chapra. New hydroepidemiological models of indicator organisms and zoonotic pathogens in agricultural watersheds. Ecological Modelling, 222(13):2093–2102, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001100216X.

Moore:2018:PDM

[MCA18] Christopher M. Moore, Samantha A. Catella, and Karen C. Abbott. Population dynamics of mutualism and intraspecific density dependence: How θ-logistic density dependence affects mutualistic positive feedback. Ecological Modelling, 368(??):191–197, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304775.

Morandi:2014:STA

[MCB14] Fabiana Morandi, Daniel E. Campbell, and Simone Bastianoni. Set theory applied to uniquely define the inputs to territorial systems in emergy analyses. Ecological Modelling, 271(??):149–157, January 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300032X.

Maselli:2012:SOF

[MCBM12] Fabio Maselli, Marta Chiesi, Lorenzo Brilli, and Marco Moriondo. Simulation of olive fruit yield in Tuscany through the integration of remote sensing and ground data. Ecological Modelling, 244(??):1–12, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003183.

McCann:2016:EAS

[McC16] Michael J. McCann. Evidence of alternative states in freshwater lakes: a spatially-explicit model of submerged and floating plants. Ecological Modelling, 337(??):298–309, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016302484.

Miquelajauregui:2019:STR

[MCG19] Yosune Miquelajauregui, Steven G. Cumming, and Sylvie Gauthier. Short-term responses of boreal carbon stocks to climate change: a simulation study of black spruce forests. Ecological Modelling, 409(??):Article 108754, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302625.

Malerba:2015:EVN

[MCH15] Martino E. Malerba, Sean R. Connolly, and Kirsten Heimann. An experimentally validated nitrate–ammonium– phytoplankton model including effects of starvation length and ammonium inhibition on nitrate uptake. Ecological Modelling, 317(??):30–40, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003877.

Mei:2017:HPY

[MCL^{+}17] Zhigang Mei, Mao Chen, Yongtao Li, Shiang-Lin Huang, Jie Haung, Yi Han, Binting Zhu, Chunsheng Li, Kexiong Wang, and Ding Wang. Habitat preference of the Yangtze finless porpoise in a minimally disturbed environment. Ecological Modelling, 353(??):47–53, June 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016308602.

Mori:2016:USA

[MCMT^{+}16] Mao Mori, Stuart P. Corney, Jessica Melbourne-Thomas, Dirk C. Welsford, Andreas Klocker, and Philippe E. Ziegler. Using satellite altimetry to inform hypotheses of transport of early life stage of Patagonian toothfish on the Kerguelen Plateau. Ecological Modelling, 340(??):45–56, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303325.

Moroz:2016:CPM

[MCN16] Irene M. Moroz, Roger Cropp, and John Norbury. Chaos in plankton models: Foraging strategy and seasonal forcing. Ecological Modelling, 332(??):103–111, July 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301156.

Muenich:2016:EPW

[MCP16] Rebecca Logsdon Muenich, Indrajeet Chaubey, and Mark Pyron. Evaluating potential water quality drivers of a fish regime shift in the Wabash River using the SWAT model. Ecological Modelling, 340(??):116–125, November 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304033.

Morandi:2013:ULS

[MCPB13] Fabiana Morandi, Daniel E. Campbell, Riccardo M. Pulselli, and Simone Bastianoni. Using the language of sets to describe nested systems in emergy evaluations. Ecological Modelling, 265(??):85–98, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002901.

Morandi:2015:EEH

[MCPB15] Fabiana Morandi, Daniel E. Campbell, Federico M. Pulselli, and Simone Bastianoni. Emergy evaluation of hierarchically nested systems: application to EU27, Italy and Tuscany and consequences for the meaning of emergy indicators. Ecological Modelling, 315(??):12–27, November 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001313.

McCormack:2015:SFE

[MCR^{+}15] M. Luke McCormack, Elizabeth Crisfield, Brett Raczka, Frank Schnekenburger, David M. Eissenstat, and Erica A. H. Smithwick. Sensitivity of four ecological models to adjustments in fine root turnover rate. Ecological Modelling, 297(??):107–117, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005869.

Maldonado-Chaparro:2017:CIV

[MCRB17] Adriana A. Maldonado-Chaparro, Dwight W. Read, and Daniel T. Blumstein. Can individual variation in phenotypic plasticity enhance population viability? Ecological Modelling, 352(??):19–30, May 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308043.

Mathys:2018:DDS

[MCS^{+}18] A. S. Mathys, N. C. Coops, S. W. Simard, R. H. Waring, and S. N. Aitken. Diverging distribution of seedlings and mature trees reflects recent climate change in British Columbia. Ecological Modelling, 384(??):145–153, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302138.

Marchi:2010:BIT

[MCT10] Michela Marchi, Ilaria Corsi, and Enzo Tiezzi. Biological invasions and their threat to ecosystems: Two ways to thermodynamic euthanasia. Ecological Modelling, 221(5):882–883, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009008072.

Mao:2013:EHR

[MCW13] Xufeng Mao, Lijuan Cui, and Changhai Wang. Exploring the hydrologic relationships in a swamp-dominated watershed — a network-environ-analysis based approach. Ecological Modelling, 252(??):273–279, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300029X.

Ma:2021:OEA

[MCWG21] Shaoxiu Ma, Galina Churkina, Ralf Wieland, and Arthur Gessler. Optimization and evaluation of the ANTHRO-BGC model for winter crops in Europe. Ecological Modelling, 222(20–22):3662–3679, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004571.

Maneerat:2016:SAB

[MD16] Somsakun Maneerat and Eric Daudé. A spatial agentbased simulation model of the dengue vector Aedes aegypti to explore its population dynamics in urban areas. Ecological Modelling, 333(??):66–78, August 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301387.

Membiela:2018:HAM

[Md18] F. Agustín Membiela and Matías G. dell’Erba. A hydrodynamic analytical model of fish tilt angle: Implications regarding acoustic target strength modelling. Ecological Modelling, 387(??):70–82, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301959.

Malanson:2019:IVE

[MDB19] George P. Malanson, R. Justin DeRose, and Matthew F. Bekker. Individual variation and ecotypic niches in simulations of the impact of climatic volatility. Ecological Modelling, 411(??):Article 108782, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930290X.

McCullough:2018:DMO

[MDF^{+}18] Ian M. McCullough, Hilary A. Dugan, Kaitlin J. Farrell, Ana M. Morales-Williams, Zutao Ouyang, Derek Roberts, Facundo Scordo, Sarah L. Bartlett, Samantha M. Burke, Jonathan P. Doubek, Flora E. Krivak-Tetley, Nicholas K. Skaff, Jamie C. Summers, Kathleen C. Weathers, and Paul C. Hanson. Dynamic modeling of organic carbon fates in lake ecosystems. Ecological Modelling, 386(??):71–82, October 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302783.

Mouton:2010:ERP

[MDG10] Ans M. Mouton, Bernard De Baets, and Peter L. M. Goethals. Ecological relevance of performance criteria for species distribution models. Ecological Modelling, 221(16):1995–2002, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002334.

Meineri:2015:UGB

[MDH15] Eric Meineri, C. Johan Dahlberg, and Kristoffer Hylander. Using Gaussian Bayesian networks to disentangle direct and indirect associations between landscape physiography, environmental variables and species distribution. Ecological Modelling, 313(??):127–136, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500277X.

Marzloff:2011:EAS

[MDJ^{+}11] Martin P. Marzloff, Jeffrey M. Dambacher, Craig R. Johnson, L. Richard Little, and Stewart D. Frusher. Exploring alternative states in ecological systems with a qualitative analysis of community feedback. Ecological Modelling, 222(15):2651–2662, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001773.

Mellert:2015:MSC

[MDKK15] Karl H. Mellert, Veronika Deffner, Helmut Küchenhoff, and Christian Kölling. Modeling sensitivity to climate change and estimating the uncertainty of its impact: a probabilistic concept for risk assessment in forestry. Ecological Modelling, 316(??):211–216, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003774.

Milessi:2010:TMB

[MDL^{+}10] Andrés C. Milessi, Calliari Danilo, Rodríguez-Graña Laura, Conde Daniel, Sellanes Javier, and Lorena Rodríguez-Gallego. Trophic mass-balance model of a subtropical coastal lagoon, including a comparison with a stable isotope analysis of the food-web. Ecological Modelling, 221(24):2859–2869, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004503.

Marva:2015:CSC

[MdlP15] Marcos Marvá and Rafael Bravo de la Parra. Coexistence and superior competitor exclusion in the Leslie– Gower competition model with fast dispersal. Ecological Modelling, 306(??):247–256, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005572.

Mohan:2019:OIT

[MdMS^{+}19] Midhun Mohan, Bruno Araujo Furtado de Mendonça, Carlos Alberto Silva, Carine Klauberg, Acauã Santos de Saboya Ribeiro, Emanuel José Gomes de Araújo, Marco Antonio Monte, and Adrián Cardil. Optimizing individual tree detection accuracy and measuring forest uniformity in coconut (Cocos nucifera L.) plantations using airborne laser scanning. Ecological Modelling, 409(??):Article 108736, October 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302443.

Meiyappan:2014:SMA

[MDOJ14] Prasanth Meiyappan, Michael Dalton, Brian C. O’Neill, and Atul K. Jain. Spatial modeling of agricultural land use change at global scale. Ecological Modelling, 291 (??):152–174, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003640.

Mouton:2012:ISE

[MDV^{+}12] A. M. Mouton, A. Dillen, T. Van den Neucker, D. Buysse, M. Stevens, and J. Coeck. Impact of sampling efficiency on the performance of data-driven fish habitat models. Ecological Modelling, 245(??):94–102, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200110X.

Mosnier:2015:IPP

[MDVG^{+}15] A. Mosnier, T. Doniol-Valcroze, J.-F. Gosselin, V. Lesage, L. N. Measures, and M. O. Hammill. Insights into processes of population decline using an integrated population model: the case of the St. Lawrence Estuary beluga (Delphinapterus leucas). Ecological Modelling, 314(??):15–31, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003129.

Maclean:2012:MDA

[MDW^{+}12] Heather Maclean, Denis Dochain, Geoffrey Waters, Michael Stasiak, Mike Dixon, and Dominique Van Der Straeten. A model development approach to ensure identifiability of a simple mass balance model for photosynthesis and respiration in a plant growth chamber. Ecological Modelling, 246(??):105–118, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200350X.

Muller:2014:LPS

[MEC14] Johannes Müller, André Eschenröder, and Olaf Christen. LEAFC3-N photosynthesis, stomatal conductance, transpiration and energy balance model: Finite mesophyll conductance, drought stress, stomata ratio, optimized solution algorithms, and code. Ecological Modelling, 290(??):134–145, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005358.

Meron:2012:PFA

[Mer12]  Ehud Meron.  Pattern-formation approach to mod-

elling spatially extended ecosystems. Ecological Modelling, 234(??):70–82, June 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003279.

Montagnes:2012:PAA

[MF12] David J. S. Montagnes and Andy Fenton. Prey-abundance affects zooplankton assimilation efficiency and the outcome of biogeochemical models. Ecological Modelling, 243(??):1–7, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002268.

McDermott:2016:IRH

[MF16] Shana M. McDermott and David C. Finnoff. Impact of repeated human introductions and the Allee effect on invasive species spread. Ecological Modelling, 329(??):100–111, June 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300606.

Maestri:2017:EMS

[MFB17] María Laura Maestri, Rosana Ferrati, and Igor Berkunsky. Evaluating management strategies in the conservation of the critically endangered blue-throated macaw (Ara glaucogularis). Ecological Modelling, 361(??):74–79, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017303435.

Mattei:2015:MMP

[MFD^{+}15] M. R. Mattei, L. Frunzo, B. D’Acunto, G. Esposito, and F. Pirozzi. Modelling microbial population dynamics in multispecies biofilms including anammox bacteria. Ecological Modelling, 304(??):44–58, May 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000630.

Martinez-Fernandez:2013:SMI

[MFESBG^{+}13] Julia Martínez-Fernández, Miguel Angel Esteve-Selma, Isabel Baños-González, Francisca Carreño, and Angeles Moreno. Sustainability of Mediterranean irrigated agrolandscapes. Ecological Modelling, 248(??):11–19, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001200484X.

Mutshinda:2013:WEF

[MFI13] Crispin M. Mutshinda, Zoe V. Finkel, and Andrew J. Irwin. Which environmental factors control phytoplankton populations? A Bayesian variable selection approach. Ecological Modelling, 269(??):1–8, November 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003803.

Mattei:2018:DRA

[MFS18] F. Mattei, S. Franceschini, and M. Scardi. A depth-resolved artificial neural network model of marine phytoplankton primary production. Ecological Modelling, 382(??):51–62, August 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018301467.

MacPherson:2016:IBE

[MG16] Brian MacPherson and Robin Gras. Individual-based ecological models: Adjunctive tools or experimental systems? Ecological Modelling, 323(??):106–114, March 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005815.

Menesguen:2018:MBIa

[MG18] Alain Ménesguen and Thomas Grégoris. Modelling benthic invasion by the colonial gastropod Crepidula fornicata and its competition with the bivalve Pecten maximus. 1. A new 0D model for population dynamics of colony-forming species. Ecological Modelling, 368(??):277–287, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304921.

Miguet:2013:BHS

[MGB13] Paul Miguet, Cédric Gaucherel, and Vincent Bretagnolle. Breeding habitat selection of skylarks varies with crop heterogeneity, time and spatial scale, and reveals spatial and temporal crop complementation. Ecological Modelling, 266(??):10–18, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003244.

McGill:2017:EME

[MGCL17] Lillian M. McGill, Brandon S. Gerig, Dominic T. Chaloner, and Gary A. Lamberti. An ecosystem model for evaluating the effects of introduced Pacific salmon on contaminant burdens of stream-resident fish. Ecological Modelling, 355(??):39–48, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307943.

Marescot:2012:RMP

[MGD^{+}12] Lucile Marescot, Olivier Gimenez, Christophe Duchamp, Eric Marboutin, and Guillaume Chapron. Reducing matrix population models with application to social animal species. Ecological Modelling, 232(??):91–96, May 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000932.

Martin-Garcia:2013:UTP

[MGGLBIBD13] Laura Martín-García, Gustavo González-Lorenzo, Isabel T. Brito-Izquierdo, and Jacinto Barquín-Diez. Use of topographic predictors for macrobenthic community mapping in the Marine Reserve of La Palma (Canary Islands, Spain). Ecological Modelling, 263(??):19–31, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002093.

Mandal:2015:SMP

[MGMR15] Sudipto Mandal, Abhishek Roy Goswami, Subhra Kumar Mukhopadhyay, and Santanu Ray. Simulation model of phosphorus dynamics of an eutrophic impoundment — East Calcutta Wetlands, a Ramsar site in India. Ecological Modelling, 306(??):226–239, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003354.

Maragno:2018:FSA

[MGR^{+}18] Denis Maragno, Mattias Gaglio, Martina Robbi, Federica Appiotti, Elisa Anna Fano, and Elena Gissi. Fine-scale analysis of urban flooding reduction from green infrastructure: an ecosystem services approach for the management of water flows. Ecological Modelling, 386(??):1–10, October 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302643.

Mohajeri:2015:STM

[MGS15] Nahid Mohajeri, Agust Gudmundsson, and Jean-Louis Scartezzini. Statistical-thermodynamics modelling of the built environment in relation to urban ecology. Ecological Modelling, 307(??):32–47, July 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001258.

Mekonnen:2016:SMN

[MGS16] Zelalem A. Mekonnen, Robert F. Grant, and Christopher Schwalm. Sensitivity of modeled NEP to climate forcing and soil at site and regional scales: Implications for upscaling ecosystem models. Ecological Modelling, 320(??):241–257, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004627.

Mony:2011:LSP

[MGSB11] C. Mony, M. Garbey, M. Smaoui, and M.-L. Benot. Large scale parameter study of an individual-based model of clonal plant with volunteer computing. Ecological Modelling, 222(4):935–946, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005557.

McNair:2013:NME

[MGW^{+}13] James N. McNair, Leon C. Gereaux, Anthony D. Weinke, Meagan R. Sesselmann, Scott T. Kendall, and Bopaiah A. Biddanda. New methods for estimating components of lake metabolism based on free-water dissolved-oxygen dynamics. Ecological Modelling, 263(??):251–263, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001300272X.

Mulindwa:2021:SSM

[MGW^{+}21] Henry Mulindwa, Esau Galukande, Maria Wurzinger, Julie Ojango, Ally Mwai Okeyo, and Johann Sölkner. Stochastic simulation model of ankole pastoral production system: Model development and evaluation. Ecological Modelling, 222(20–22):3692–3700, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004613.

Mora:2013:SER

[MGWW13] Miguel A. Mora, William E. Grant, Leann Wilkins, and Hsiao-Hsuan Wang. Simulated effects of reduced spring flow from the Edwards aquifer on population size of the fountain darter (Etheostoma fonticola). Ecological Modelling, 250(??):235–243, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005509.

Ge:2010:ECC

[mGZK^{+}10] Zhen ming Ge, Xiao Zhou, Seppo Kellomäki, Kai yun Wang, Heli Peltola, Hannu Väisänen, and Harri Strandman. Effects of changing climate on water and nitrogen availability with implications on the productivity of Norway spruce stands in Southern Finland. Ecological Modelling, 221(13–14):1731–1743, July 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001626.

Missaghi:2010:EAT

[MH10] Shahram Missaghi and Miki Hondzo. Evaluation and application of a three-dimensional water quality model in a shallow lake with complex morphometry. Ecological Modelling, 221(11):1512–1525, June 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000773.

Murphy:2011:AEN

[MH11] David J. Murphy and Charles A. S. Hall. Adjusting the economy to the new energy realities of the second half of the age of oil. Ecological Modelling, 223(1):67–71, December 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001100370X.

Maleewong:2013:ANR

[MH13] Montri Maleewong and Sion Hasadsri. Analytical and numerical results of dissolved oxygen and biochemical oxygen demand in non-uniform open channel. Ecological Modelling, 252(??):11–22, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005534.

Moe:2016:CCC

[MHC16] S. Jannicke Moe, Sigrid Haande, and Raoul-Marie Couture. Climate change, cyanobacteria blooms and ecological status of lakes: a Bayesian network approach. Ecological Modelling, 337(??):330–347, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302460.

Menesguen:2018:MBIb

[MHG18] Alain Ménesguen, Aloïs Hachet, and Thomas Grégoris. Modelling benthic invasion by the colonial gastropod Crepidula fornicata and its competition with the bivalve Pecten maximus. 2. Coupling the 0D model of colonyforming species to a connectivity matrix for a realistic distributed simulation of benthic invasion. Ecological Modelling, 375(??):30–44, May 10, 2018. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018300760.

Ma:2019:ERI

[MHL^{+}19] Ping Ma, Xiao-Hui Han, Yue Lin, John Moore, Yao-Xin Guo, and Ming Yue. Exploring the relative importance of biotic and abiotic factors that alter the self-thinning rule: Insights from individual-based modelling and machinelearning. Ecological Modelling, 397(??):16–24, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300390.

Matus-Hernandez:2019:LDE

[MHMRAHHS19] Miguel Ángel Matus-Hernández, Raúl Octavio Martínez-Rincón, Rosa Judith Aviña-Hernández, and Norma Yolanda Hernández-Saavedra. Landsat-derived environmental factors to describe habitat preferences and spatiotemporal distribution of phytoplankton. Ecological Modelling, 408(??):Article 108759, September 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302674.

Moe:2019:EAP

[MHRB19] S. Jannicke Moe, Dag Ø. Hjermann, Elisa Ravagnan, and Renée K. Bechmann. Effects of an aquaculture pesticide (diflubenzuron) on non-target shrimp populations: Extrapolation from laboratory experiments to the risk of population decline. Ecological Modelling, 413(??): Article 108833, December 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019303412.

Marley:2017:DHE

[MHS^{+}17] Jessa Marley, Andrea Hyde, Joseph H. Salkeld, Marie-Caroline Prima, Lael Parrott, Susan E. Senger, and Rebecca C. Tyson. Does human education reduce conflicts between humans and bears? An agent-based modelling approach. Ecological Modelling, 343(??):15–24, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016305798.

Marchand:2015:TMB

[MHTC15] Philippe Marchand, Alexandra N. Harmon-Threatt, and Ignacio Chapela. Testing models of bee foraging behavior through the analysis of pollen loads and floral density data. Ecological Modelling, 313(??):41–49, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002689.

Moritz:2015:SMR

[MHY^{+}15] Mark Moritz, Ian M. Hamilton, Andrew J. Yoak, Paul Scholte, Jeff Cronley, Paul Maddock, and Hongyang Pi. Simple movement rules result in ideal free distribution of mobile pastoralists. Ecological Modelling, 305(??):54–63, June 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001210.

Melica:2014:LDD

[MIC14] Valentina Melica, Sergio Invernizzi, and Gabriella Caristi. Logistic density-dependent growth of an Aurelia aurita polyps population. Ecological Modelling, 291(??):1–5, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380014003366.

Muci:2012:CCA

[MJÁ^{+}12] Adrialy L. Muci, Milko A. Jorquera, Ándres I. Ávila, Zed Rengel, David E. Crowley, and María de la Luz Mora. A combination of cellular automata and agent-based models for simulating the root surface colonization by bacteria. Ecological Modelling, 247(??):1–10, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003845.

Marchi:2011:DML

[MJB^{+}11a] Michela Marchi, Sven Erik Jørgensen, Eloy Bécares, Ilaria Corsi, Nadia Marchettini, and Simone Bastianoni. Dynamic model of Lake Chozas (León, NW Spain) — decrease in eco-exergy from clear to turbid phase due to introduction of exotic crayfish. Ecological Modelling, 222(16):3002–3010, August 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011002249.

Marchi:2011:RRO

[MJB^{+}11b] Michela Marchi, Sven Erik Jørgensen, Eloy Bécares, Ilaria Corsi, Nadia Marchettini, and Simone Bastianoni. Resistance and re-organization of an ecosystem in response to biological invasion: Some hypotheses. Ecological Modelling, 222(16):2992–3001, August 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011002250.

Min:2011:WIE

[MJC^{+}11] Yong Min, Xiaogang Jin, Jie Chang, Changhui Peng, Baojing Gu, Ying Ge, and Yang Zhong. Weak indirect effects inherent to nitrogen biogeochemical cycling within anthropogenic ecosystems: a network environ analysis. Ecological Modelling, 222(17):3277–3284, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003619.

Mishra:2019:MSO

[MJF19] Gaurav Mishra, Abhishek Jangir, and Rosa Francaviglia. Modeling soil organic carbon dynamics under shifting cultivation and forests using rothc model. Ecological Modelling, 396(??):33–41, March 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300286.

Munier-Jolain:2013:MLI

[MJGC13] N. M. Munier-Jolain, S. H. M. Guyot, and N. Colbach. A 3D model for light interception in heterogeneous crop:weed canopies: Model structure and evaluation. Ecological Modelling, 250(??):101–110, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005261.

Milligan:2017:PEM

[MJK^{+}17] William R. Milligan, Marjorie T. Jones, Lee B. Kats, Timothy A. Lucas, and Courtney L. Davis. Predicting the effects of manual crayfish removal on California newt persistence in Santa Monica Mountain streams. Ecological Modelling, 352(??):139–151, May 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302988.

Marzloff:2013:SAP

[MJL^{+}13] Martin P. Marzloff, Craig R. Johnson, L. Rich Little, Jean-Christophe Soulié, Scott D. Ling, and Stewart D. Frusher. Sensitivity analysis and pattern-oriented validation of TRITON, a model with alternative community states: Insights on temperate rocky reefs dynamics. Ecological Modelling, 258(??):16–32, June 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001099.

Marzloff:2014:CSA

[MJL^{+}14] Martin P. Marzloff, Craig R. Johnson, L. Rich Little, Jean-Christophe Soulié, Scott D. Ling, and Stewart D. Frusher. Corrigendum to “Sensitivity analysis and patternoriented validation of TRITON, a model with alternative community states: Insights on temperate rocky reefs dynamics” [Ecol. Model. 258 (2013) 16–32]. Ecological Modelling, 274(??):71, February 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300584X.

Marn:2017:EEG

[MJL^{+}17] Nina Marn, Marko Jusup, Tarzan Legović, S. A. L. M. Kooijman, and Tin Klanjšček. Environmental effects on growth, reproduction, and life-history traits of loggerhead turtles. Ecological Modelling, 360(??):163–178, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016308596.

Meyer:2014:DSD

[MJLC14] Spencer R. Meyer, Michelle L. Johnson, Robert J. Lilieholm, and Christopher S. Cronan. Development of a stakeholderdriven spatial modeling framework for strategic landscape planning using Bayesian networks across two urbanrural gradients in Maine, USA. Ecological Modelling, 291(??):42–57, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003056.

Marchi:2012:MCC

[MJP^{+}12] Michela Marchi, Sven Erik Jørgensen, Federico Maria Pulselli, Nadia Marchettini, and Simone Bastianoni. Modelling the carbon cycle of Siena Province (Tuscany, central Italy). Ecological Modelling, 225(??):40–60, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005278.

Mao:2017:AEB

[MJY17] Xuhui Mao, Haifeng Jia, and Shaw L. Yu. Assessing the ecological benefits of aggregate LID-BMPs through modelling. Ecological Modelling, 353(??):139–149, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305932.

Metsaranta:2012:IAV

[MK12] J. M. Metsaranta and W. A. Kurz. Inter-annual variability of ecosystem production in boreal jack pine forests (1975– 2004) estimated from tree-ring data using CBM-CFS3. Ecological Modelling, 224(1):111–123, January 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005187.

Ma:2013:AIE

[MK13] Q. Ma and C. Kazanci. Analysis of indirect effects within ecosystem models using pathway-based methodology. Ecological Modelling, 252(??):238–245, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002220.

Miedziejko:2014:IPC

[MK14] Ewa M. Miedziejko and Andrzej Kȩdziora. Impact of plant canopy structure on the transport of ecosystem entropy. Ecological Modelling, 289(??):15–25, October 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002956.

Menezes:2019:GIF

[MK19] Jorge F. S. Menezes and Burt P. Kotler. The generalized ideal free distribution model: Merging current ideal free distribution models into a central framework. Ecological Modelling, 397(??):47–54, April 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300225.

Manzoni:2011:MVA

[MKF^{+}11] Stefano Manzoni, Gabriel Katul, Philip A. Fay, H. Wayne Polley, and Amilcare Porporato. Modeling the vegetation– atmosphere carbon dioxide and water vapor interactions along a controlled CO_{2} gradient. Ecological Modelling, 222(3):653–665, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005764.

Macpherson:2017:EIP

[MKH^{+}17] Morag F. Macpherson, Adam Kleczkowski, John R. Healey, Christopher P. Quine, and Nick Hanley. The effects of invasive pests and pathogens on strategies for forest diversification. Ecological Modelling, 350(??):87–99, April 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303726.

Marshall:2014:NTF

[MKL14] K. N. Marshall, I. C. Kaplan, and P. S. Levin. New target fisheries lead to spatially variable food web effects in an ecosystem model of the California Current. Ecological Modelling, 289(??):96–105, October 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003305.

Matinzadeh:2017:DII

[MKSL^{+}17] Mohammad Mehdi Matinzadeh, Jahangir Abedi Koupai, Adnan Sadeghi-Lari, Hamed Nozari, and Mohammad Shayannejad. Development of an innovative integrated model for the simulation of nitrogen dynamics in farmlands with drainage systems using the system dynamics approach. Ecological Modelling, 347(??):11–28, March 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305051.

Matsubara:2019:GEG

[MKTO19] Takashi Matsubara, Yoshiko Kosugi, Satoru Takanashi, and Kiyotoshi Otsuka. Gas exchange and growth/decline model of C_{3} turfgrass fields under various light conditions. Ecological Modelling, 397(??):107–121, April 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303089.

Minarro:2018:SSE

[MLAT^{+}18] Sara Miñarro, Johannes Leins, Esteban Acevedo-Trejos, Elizabeth A. Fulton, and Hauke Reuter. SEAMAN-CORE: a spatially explicit simulation model for assisting the local MANagement of COral REefs. Ecological Modelling, 384(??):296–307, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301996.

Ma:2016:CSO

[MLB^{+}16] Kun Ma, Junguo Liu, Juraj Balkovič, Rastislav Skalský, Ligia B. Azevedo, and Florian Kraxner. Changes in soil organic carbon stocks of wetlands on China’s Zoige Plateau from 1980 to 2010. Ecological Modelling, 327(??):18–28, May 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016000284.

Ma:2016:FSC

[MLC^{+}16] Wu Ma, Jingjing Liang, Jonathan R. Cumming, Eungul Lee, Amy B. Welsh, James V. Watson, and Mo Zhou. Fundamental shifts of central hardwood forests under climate change. Ecological Modelling, 332(??):28–41, July 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300941.

Lai:2015:TOB

[mLcLJ15] Shu mei Lai, Wei chung Liu, and Ferenc Jordán. A trophic overlap-based measure for species uniqueness in ecological networks. Ecological Modelling, 299(??):95–101, March 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014006206.

Lai:2016:CTO

[mLcLJ16] Shu mei Lai, Wei chung Liu, and Ferenc Jordán. Corrigendum to “A trophic overlap-based measure for species uniqueness in ecological networks” [Ecol. Model. 299 (March) (2015) 95–101]. Ecological Modelling, 336(??):82, September 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016301004.

Marilleau:2018:CAB

[MLG18a] Nicolas Marilleau, Christophe Lang, and Patrick Giraudoux. Coupling agent-based with equation-based models to study spatially explicit megapopulation dynamics. Ecological Modelling, 384(??):34–42, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302163.

Moudry:2018:UGD

[MLG^{+}18b] Vítězslav Moudrý, Vincent Lecours, Kateřina Gdulová, Lukáš Gábor, Lucie Moudrá, Jan Kropáček, and Jan Wild. On the use of global DEMs in ecological modelling and the accuracy of new bare-earth DEMs. Ecological Modelling, 383(??):3–9, September 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301571.

Mo:2017:STW

[MLH^{+}17] X. Mo, S. Liu, S. Hu, J. Xia, and Y. Shen. Sensitivity of terrestrial water and carbon fluxes to climate variability in semi-humid basins of Haihe River, China. Ecological Modelling, 353(??):117–128, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303799.

Maggini:2011:SBT

[MLK^{+}11] Ramona Maggini, Anthony Lehmann, Marc Kéry, Hans Schmid, Martin Beniston, Lukas Jenni, and Niklaus Zbinden. Are Swiss birds tracking climate change?: Detecting elevational shifts using response curve shapes. Ecological Modelling, 222(1):21–32, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004813.

Merikanto:2017:OHP

[MLK17] Ilona Merikanto, Jouni T. Laakso, and Veijo Kaitala. Outside-host predation as a biological control against an environmental opportunist disease. Ecological Modelling, 347(??):85–93, March 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017300078.

Martinez-Lopez:2015:OSS

[MLMFN^{+}15] Javier Martínez-López, Julia Martínez-Fernández, Babak Naimi, María F. Carreño, and Miguel A. Esteve. An open-source spatio-dynamic wetland model of plant community responses to hydrological pressures. Ecological Modelling, 306(??):326–333, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005973.

Mohammed:2018:FSD

[MLMH18] M. M. A. Mohammed, P. Landi, H. O. Minoarivelo, and C. Hui. Frugivory and seed dispersal: Extended bistable persistence and reduced clustering of plants. Ecological Modelling, 380(??):31–39, July 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301273.

Massoudieh:2011:STM

[MLS^{+}11] Arash Massoudieh, Erik Loboschefsky, Ted Sommer, Timothy Ginn, Kenneth Rose, and Frank Loge. Spatiotemporal modeling of striped-bass egg, larval movement, and fate in the San Francisco Bay–Delta. Ecological Modelling, 222(19):3513–3523, October 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004261.

Ma:2011:EPR

[MLW11] Zhihui Ma, Weide Li, and Shufan Wang. The effect of prey refuge in a simple predator–prey model. Ecological Modelling, 222(18):3453–3454, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003000. See com-

ments [GORJ12].

Ma:2015:MDC

[MLW^{+}15] Jun Ma, Defu Liu, Scott A. Wells, Hongwu Tang, Daobin Ji, and Zhengjian Yang. Modeling density currents in a typical tributary of the Three Gorges Reservoir, China. Ecological Modelling, 296(??):113–125, January 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005286.

Ma:2019:SRH

[MLY^{+}19] Diaoyuan Ma, Wenguang Luo, Guolu Yang, Jing Lu, and Yangzhen Fan. A study on a river health assessment method based on ecological flow. Ecological Modelling, 401(??):144–154, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304113.

Marois:2016:MPR

[MM16] Darryl E. Marois and William J. Mitsch. Modeling phosphorus retention at low concentrations in Florida Everglades mesocosms. Ecological Modelling, 319(??):42–62, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001500455X.

Maroto:2019:TDE

[MM19] José M. Maroto and Manuel Morán. Transient dynamics: Equilibrium, collapse, and extinction in age-structured models. The case of the Northern cod stock. Ecological Modelling, 398(??):35–43, April 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300675.

Moller:2011:PES

[MMC11] Jan Kloppenborg Møller, Henrik Madsen, and Jacob Carstensen. Parameter estimation in a simple stochastic differential equation for phytoplankton modelling. Ecological Modelling, 222(11):1793–1799, June 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001487.

Metcalfe:2019:MES

[MMC^{+}19] H. Metcalfe, A. E. Milne, K. Coleman, A. J. Murdoch, and J. Storkey. Modelling the effect of spatially variable soil properties on the distribution of weeds. Ecological Modelling, 396(??):1–11, March 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303648.

Martinez-Minaya:2019:DPB

[MMCBP19] Joaquín Martínez-Minaya, David Conesa, Haakon Bakka, and Maria Grazia Pennino. Dealing with physical barriers in bottlenose dolphin (Tursiops truncatus) distribution.

Ecological Modelling, 406(??):44–49, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302030.

Maeda:2017:BMF

[MMD^{+}17] Eduardo Eiji Maeda, Samu Mäntyniemi, Smaragda Despoti, Claudia Musumeci, Vassiliki Vassilopoulou, Konstantinos I. Stergiou, Marianna Giannoulaki, Alessandro Ligas, and Sakari Kuikka. A Bayesian model of fisheries discards with flexible structure and priors defined by experts. Ecological Modelling, 366(??):1–14, December 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304350.

Morales:2015:SFS

[MMGK15] Juan Manuel Morales, Mónica Mermoz, Juan Haridas Gowda, and Thomas Kitzberger. A stochastic fire spread model for north Patagonia based on fire occurrence maps. Ecological Modelling, 300(??):73–80, March 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000162.

Mo:2015:QMR

[MMK15] Yu Mo, Bahram Momen, and Michael S. Kearney. Quantifying moderate resolution remote sensing phenology of Louisiana coastal marshes. Ecological Modelling, 312 (??):191–199, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002197.

Maar:2011:EMA

[MML^{+}11] Marie Maar, Eva Friis Møller, Jesper Larsen, Kristine Skovgaard Madsen, Zhenwen Wan, Jun She, Lars Jonasson, and Thomas Neumann. Ecosystem modelling across a salinity gradient from the North Sea to the Baltic Sea. Ecological Modelling, 222(10):1696–1711, May 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100113X.

Maar:2016:ILV

[MMM^{+}16] Marie Maar, Stiig Markager, Kristine Skovgaard Madsen, Jørgen Windolf, Maren Moltke Lyngsgaard, Hans Estrup Andersen, and Eva Friis Møller. The importance of local versus external nutrient loads for chl a and primary production in the Western Baltic Sea. Ecological Modelling, 320(??):258–272, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004548.

Munoz-Mas:2021:CMP

[MMMCAHM21] R. Muñoz-Mas, F. Martínez-Capel, J. D. Alcaraz-Hernández, and A. M. Mouton. Can multilayer perceptron ensembles model the ecological niche of freshwater fish species? Ecological Modelling, 309–310(??):72–81, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001763.

Munoz-Mas:2018:CLB

[MMMGLN^{+}18] Rafael Muñoz-Mas, Patricia Marcos-Garcia, Antonio Lopez-Nicolas, Francisco J. Martínez-García, Manuel Pulido-Velazquez, and Francisco Martínez-Capel. Combining literature-based and data-driven fuzzy models to predict brown trout (Salmo trutta L.) spawning habitat degradation induced by climate change. Ecological Modelling, 386(??):98–114, October 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302813.

Mohd:2016:EDS

[MMPG16] Mohd Hafiz Mohd, Rua Murray, Michael J. Plank, and William Godsoe. Effects of dispersal and stochasticity on the presence–absence of multiple species. Ecological Modelling, 342(??):49–59, December 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305129.

Morales-Marin:2019:CSW

[MMRS^{+}19] L. A. Morales-Marín, P. Rokaya, P. R. Sanyal, J. Sereda, and K. E. Lindenschmidt. Changes in streamflow and water temperature affect fish habitat in the Athabasca River basin in the context of climate change. Ecological Modelling, 407(??):Article 108718, September 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302182.

Matawa:2012:EEL

[MMS12] Farai Matawa, Amon Murwira, and Karin S. Schmidt. Explaining elephant (Loxodonta africana) and buffalo (Syncerus caffer) spatial distribution in the Zambezi Valley using maximum entropy modelling. Ecological Modelling, 242(??):189–197, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001200230X.

Munoz-Mas:2016:RIP

[MMVAHMC16] Rafael Muñoz-Mas, Paolo Vezza, Juan Diego Alcaraz-Hernández, and Francisco Martínez-Capel. Risk of invasion predicted with support vector machines: a case study on northern pike (Esox lucius, L.) and bleak (Alburnus alburnus, L.). Ecological Modelling, 342(??):123–134, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016305397.

Mullon:2012:GTA

[MN12] Christian Mullon and Anna Nagurney. A game theoretical approach to the vertical coexistence of small and big fish. Ecological Modelling, 240(??):41–48, August 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002098.

Marin:2011:PLC

[MNEG11] Sandra L. Marín, Laura Nahuelhual, Cristian Echeverría, and William E. Grant. Projecting landscape changes in southern Chile: Simulation of human and natural processes driving land transformation. Ecological Modelling, 222(15):2841–2855, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002560.

Mose:2013:MDM

[MNHW^{+}13] Victor N. Mose, Tri Nguyen-Huu, David Western, Pierre Auger, and Charles Nyandwi. Modelling the dynamics of migrations for large herbivore populations in the Amboseli National Park, Kenya. Ecological Modelling, 254(??):43–49, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013000434.

Mazzucco:2021:AAI

[MNK^{+}21] Rupert Mazzucco, Tuyen Van Nguyen, Dong-Hwan Kim, Tae-Soo Chon, and Ulf Dieckmann. Adaptation of aquatic insects to the current flow in streams. Ecological Modelling, 309–310(??):143–152, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001702.

Na:2017:MBA

[mNkGyT^{+}17] Hong ming Na, Cheng kang Gao, Ming yan Tian, Zhao qian Qi, and Zhou Ye. MFA-based analysis of CO_{2} emissions from typical industry in urban — as a case of steel industry. Ecological Modelling, 365(??):45–54, December 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303605.

Mobus:2018:TST

[Mob18] George E. Mobus. Teaching systems thinking to general education ^{1} students. Ecological Modelling, 373(??):13–21, April 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017303939.

Murphy:2011:MMT

[MOM^{+}11] Paul N. C. Murphy, Jae Ogilvie, Fan-Rui Meng, Barry White, Jagtar S. Bhatti, and Paul A. Arp. Modelling and mapping topographic variations in forest soils at high resolution: a case study. Ecological Modelling, 222(14):2314–2332, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011000251.

Mesgaran:2017:WAS

[MOMC17] Mohsen B. Mesgaran, Andrea Onofri, Hamid R. Mashhadi, and Roger D. Cousens. Water availability shifts the optimal temperatures for seed germination: a modelling approach. Ecological Modelling, 351(??):87–95, May 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630521X.

Monte:2013:CNL

[Mon13] Luigi Monte. Characterisation of a nonlinear Leslie matrix model for predicting the dynamics of biological populations in polluted environments: Applications to radioecology. Ecological Modelling, 248(??):174–183, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005054.

McVinish:2013:IBH

[MP13] R. McVinish and P. K. Pollett. Interaction between habitat quality and an Allee-like effect in metapopulations. Ecological Modelling, 249(??):84–89, January 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003262.

Morales:2017:SSM

[MP17] Narkis S. Morales and George L. W. Perry. A spatial simulation model to explore the long-term dynamics of podocarp-tawa forest fragments, northern New Zealand. Ecological Modelling, 357(??):35–46, August 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306068.

Meli:2014:TPE

[MPF^{+}14] Mattia Meli, Annemette Palmqvist, Valery E. Forbes, Jürgen Groeneveld, and Volker Grimm. Two pairs of eyes are better than one: Combining individual-based and matrix models for ecological risk assessment of chemicals. Ecological Modelling, 280(??):40–52, May 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003839.

Moulin:2018:MVD

[MPG18] Thibault Moulin, Antoine Perasso, and François Gillet. Modelling vegetation dynamics in managed grasslands: Responses to drivers depend on species richness. Ecological Modelling, 374(??):22–36, April 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300668.

Montero-Pau:2014:MPG

[MPGCS14] Javier Montero-Pau, Carmen Gabaldón, María José Carmona, and Manuel Serra. Measuring the potential for growth in populations investing in diapause. Ecological Modelling, 272(??):76–83, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004493.

Messan:2018:EVA

[MPK18] Marisabel Rodriguez Messan, Robert E. Page, and Yun Kang. Effects of vitellogenin in age polyethism and population dynamics of honeybees. Ecological Modelling, 388 (??):88–107, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303077.

Manso:2012:MST

[MPKC12] Rubén Manso, Marta Pardos, Christopher R. Keyes, and Rafael Calama. Modelling the spatio-temporal pattern of primary dispersal in stone pine (Pinus pinea L.) stands in the Northern Plateau (Spain). Ecological Modelling, 226(??):11–21, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005709.

Minunno:2016:CVS

[MPL^{+}16] F. Minunno, M. Peltoniemi, S. Launiainen, M. Aurela, A. Lindroth, A. Lohila, I. Mammarella, K. Minkkinen, and A. Mäkelä. Calibration and validation of a semi-empirical flux ecosystem model for coniferous forests in the boreal region. Ecological Modelling, 341(??):37–52, December 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304744.

Marchi:2015:CDS

[MPM^{+}15] Michela Marchi, Riccardo Maria Pulselli, Nadia Marchettini, Federico Maria Pulselli, and Simone Bastianoni. Carbon dioxide sequestration model of a vertical greenery system. Ecological Modelling, 306(??):46–56, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014003925.

Mestre:2015:CDM

[MPP^{+}15] Frederico Mestre, Ricardo Pita, Joana Paupério, Filipa M. S. Martins, Paulo Célio Alves, António Mira, and Pedro Beja. Combining distribution modelling and noninvasive genetics to improve range shift forecasting. Ecological Modelling, 297(??):171–179, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005912.

Mancy:2013:DCT

[MPR13] Rebecca Mancy, Patrick Prosser, and Simon Rogers. Discrete and continuous time simulations of spatial ecological processes predict different final population sizes and interspecific competition outcomes. Ecological Modelling, 259(??):50–61, June 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013001701.

Maes:2011:DED

[MPT^{+}11] W. H. Maes, T. Pashuysen, A. Trabucco, F. Veroustraete, and B. Muys. Does energy dissipation increase with ecosystem succession? Testing the ecosystem exergy theory combining theoretical simulations and thermal remote sensing observations. Ecological Modelling, 222(23– 24):3917–3941, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004674.

Mazalov:2010:FWC

[MR10] V. V. Mazalov and A. N. Rettieva. Fish wars and cooperation maintenance. Ecological Modelling, 221(12):1545–1553, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010001456.

McLaughlin:2017:BNA

[MR17] Douglas B. McLaughlin and Kenneth H. Reckhow. A Bayesian network assessment of macroinvertebrate responses to nutrients and other factors in streams of the Eastern Corn Belt Plains, Ohio, USA. Ecological Modelling, 345(??):21–29, February 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307578.

Matsunaga:2014:MSR

[MRB14] Fabio Takeshi Matsunaga, Miroslava Rakocevic, and Jacques Duílio Brancher. Modeling the 3D structure and rhythmic growth responses to environment in dioecious yerba-mate. Ecological Modelling, 290(??):34–44, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013005346.

Monedero:2019:PFS

[MRC19] Santiago Monedero, Joaquin Ramirez, and Adrián Cardil. Predicting fire spread and behaviour on the fireline. Wildfire analyst pocket: a mobile app for wildland fire prediction. Ecological Modelling, 392(??):103–107, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304046.

Mangel:2010:MEU

[MRCW10] Marc Mangel, Kate Richerson, Katherine A. Cresswell, and John R. Wiedenmann. Modelling the effects of UV radiation on the survival of Antarctic krill (Euphausia superba Dana) in the face of limited data. Ecological Modelling, 221(17):2095–2101, August 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002723.

Mattsson:2012:MFI

[MRD^{+}12] B. J. Mattsson, M. C. Runge, J. H. Devries, G. S. Boomer, J. M. Eadie, D. A. Haukos, J. P. Fleskes, D. N. Koons, W. E. Thogmartin, and R. G. Clark. A modeling framework for integrated harvest and habitat management of North American waterfowl: Case-study of northern pintail metapopulation dynamics. Ecological Modelling, 225(??):146–158, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011005242.

Miller:2017:TFC

[MRES17] Matt L. Miller, Kevin M. Ringelman, John M. Eadie, and Jeffrey C. Schank. Time to fly: a comparison of marginal value theorem approximations in an agent-based model of foraging waterfowl. Ecological Modelling, 351(??):77–86, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016303180.

Mandal:2013:IML

[MRG13a] Sudipto Mandal, Santanu Ray, and Phani Bhusan Ghosh. Impact of mangrove litterfall on nitrogen dynamics of virgin and reclaimed islands of Sundarban mangrove ecosystem, India. Ecological Modelling, 252(??):153–166, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001200347X.

Mukherjee:2013:SDM

[MRG13b] Joyita Mukherjee, Santanu Ray, and Phani Bhusan Ghosh. A system dynamic modeling of carbon cycle from mangrove litter to the adjacent hooghly estuary, India. Ecological Modelling, 252(??):185–195, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003365.

Manzoor:2016:GTI

[MRM16] Talha Manzoor, Elena Rovenskaya, and Abubakr Muhammad. Game-theoretic insights into the role of environmentalism and social-ecological relevance: a cognitive model of resource consumption. Ecological Modelling, 340(??):74–85, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016303866.

Mestre:2017:MAP

[MRM^{+}17] Frederico Mestre, Benjamin B. Risk, António Mira, Pedro Beja, and Ricardo Pita. A metapopulation approach to predict species range shifts under different climate change and landscape connectivity scenarios. Ecological Modelling, 359(??):406–414, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307669.

Marin:2012:SEM

[MRN12] Víctor H. Marín, Luis C. Rodríguez, and Hermann M. Niemeyer. A socio-ecological model of the opuntia scrublands in the Peruvian Andes. Ecological Modelling, 227 (??):136–146, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005989.

Martinez-Rincon:2012:CPG

[MROGVR12] Raul O. Martínez-Rincón, Sofía Ortega-García, and Juan G. Vaca-Rodríguez. Comparative performance of generalized additive models and boosted regression trees for statistical modeling of incidental catch of wahoo (Acanthocybium solandri) in the Mexican tuna purse-seine fishery. Ecological Modelling, 233(??):20–25, May 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001135.

Mohanty:2011:MMG

[MRP^{+}11] M. Mohanty, K. Sammi Reddy, M. E. Probert, R. C. Dalal, A. Subba Rao, and N. W. Menzies. Modelling N mineralization from green manure and farmyard manure from a laboratory incubation study. Ecological Modelling, 222(3):719–726, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005909.

Marshalek:2014:IHS

[MRP14] Elaina C. Marshalek, Benjamin S. Ramage, and Matthew D. Potts. Integrating harvest scheduling and reserve design to improve biodiversity conservation. Ecological Modelling, 287(??):27–35, September 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002129.

Mittanck:2014:ESW

[MRR^{+}14] Cody M. Mittanck, Paul C. Rogers, R. Douglas Ramsey, Dale L. Bartos, and Ronald J. Ryel. Exploring succession within aspen communities using a habitat-based modeling approach. Ecological Modelling, 288(??):203–212, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002920.

Myrgiotis:2018:SAI

[MRTW18] Vasileios Myrgiotis, Robert M. Rees, Cairistiona F. E. Topp, and Mathew Williams. A systematic approach to identifying key parameters and processes in agroecosystem models. Ecological Modelling, 368(??):344–356, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304957.

Meyer:2019:ISP

[MRWN19] Hanna Meyer, Christoph Reudenbach, Stephan Wöllauer, and Thomas Nauss. Importance of spatial predictor variable selection in machine learning applications — moving from data reproduction to spatial prediction. Ecological Modelling, 411(??):Article 108815, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303230.

Mellino:2014:EGA

[MRZU14] Salvatore Mellino, Maddalena Ripa, Amalia Zucaro, and Sergio Ulgiati. An emergy–GIS approach to the evaluation of renewable resource flows: a case study of Campania Region, Italy. Ecological Modelling, 271(??):103–112, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013000069.

McLeod:2014:FCM

[MS14] Steven R. McLeod and Glen Saunders. Fertility control is much less effective than lethal baiting for controlling foxes. Ecological Modelling, 273(??):1–10, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004833.

Murray:2016:SMR

[MS16] Alexander G. Murray and Nabeil K. G. Salama. A simple model of the role of area management in the control of sea lice. Ecological Modelling, 337(??):39–47, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302125.

McHugh:2017:LMA

[MSB^{+}17] Peter A. McHugh, W. Carl Saunders, Nicolaas Bouwes, C. Eric Wall, Sara Bangen, Joseph M. Wheaton, Matthew Nahorniak, James R. Ruzycki, Ian A. Tattam, and Chris E. Jordan. Linking models across scales to assess the viability and restoration potential of a threatened population of steelhead (Oncorhynchus mykiss) in the Middle Fork John Day River, Oregon, USA. Ecological Modelling, 355(??):24–38, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016308468.

Morris:2014:EGS

[MSCH14a] David J. Morris, Douglas C. Speirs, Angus I. Cameron, and Michael R. Heath. Erratum to “Global sensitivity analysis of an end-to-end marine ecosystem model of the North Sea: Factors affecting the biomass of fish and benthos” [Ecol. Model. 273 (2014) 251–263]. Ecological Modelling, 279(??):114–117, May 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000957.

Morris:2014:GSA

[MSCH14b] David J. Morris, Douglas C. Speirs, Angus I. Cameron, and Michael R. Heath. Global sensitivity analysis of an end-to-end marine ecosystem model of the North Sea: Factors affecting the biomass of fish and benthos. Ecological Modelling, 273(??):251–263, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300567X.

McHuron:2018:SDM

[MSCM18] Elizabeth A. McHuron, Lisa K. Schwarz, Daniel P. Costa, and Marc Mangel. A state-dependent model for assessing the population consequences of disturbance on income-breeding mammals. Ecological Modelling, 385(??):133–144, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302576.

Melo:2018:EMS

[MSF18] L. C. Melo, R. Schneider, and M. Fortin. Estimating modeland sampling-related uncertainty in large-area growth predictions. Ecological Modelling, 390(??):62–69, December 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001830334X.

Mukherjee:2015:MSR

[MSFR15] Joyita Mukherjee, Ursula M. Scharler, Brian D. Fath, and Santanu Ray. Measuring sensitivity of robustness and network indices for an estuarine food web model under perturbations. Ecological Modelling, 306(??):160–173, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005006.

Maire:2013:PPF

[MSG^{+}13] Vincent Maire, Jean-François Soussana, Nicolas Gross, Bruno Bachelet, Loïc Pagès, Raphaël Martin, Tanja Reinhold, Christian Wirth, and David Hill. Plasticity of plant form and function sustains productivity and dominance along environment and competition gradients. A modeling experiment with G emini. Ecological Modelling, 254(??):80–91, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001664.

Macias:2014:RDC

[MSGG14] Diego Macías, Adolf Stips, and Elisa Garcia-Gorriz. The relevance of deep chlorophyll maximum in the open Mediterranean Sea evaluated through 3D hydrodynamicbiogeochemical coupled simulations. Ecological Modelling, 281(??):26–37, June 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001161.

Marley:2019:IBM

[MSH^{+}19] Jessa Marley, Joseph H. Salkeld, Tony Hamilton, Susan E. Senger, Rebecca C. Tyson, and Lael Parrott. Individualbased modelling of black bear (Ursus americanus) foraging in Whistler, BC: Reducing human–bear interactions. Ecological Modelling, 407(??):Article 108725, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001930225X.

Mocq:2013:AAS

[MSHC13] J. Mocq, A. St-Hilaire, and R. A. Cunjak. Assessment of Atlantic salmon (Salmo salar) habitat quality and its uncertainty using a multiple-expert fuzzy model applied to the Romaine River (Canada). Ecological Modelling, 265(??):14–25, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002834.

Masoodi:2013:PSA

[MSKS13] Ather Masoodi, Anand Sengupta, Fareed A. Khan, and Gyan P. Sharma. Predicting the spread of alligator weed (Alternanthera philoxeroides) in Wular Lake, India: a mathematical approach. Ecological Modelling, 263(??):119–125, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002457.

Mullon:2016:QPN

[MSM^{+}16] C. Mullon, F. Steinmetz, G. Merino, J. A. Fernandes, W. W. L. Cheung, M. Butenschön, and M. Barange. Quantitative pathways for Northeast Atlantic fisheries based on climate, ecological–economic and governance modelling scenarios. Ecological Modelling, 320(??):273–291, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004585.

McLane:2017:EBS

[MSM^{+}17] Adam J. McLane, Christina Semeniuk, Gregory J. Mc-Dermid, Diana F. Tomback, Teresa Lorenz, and Danielle Marceau. Energetic behavioural-strategy prioritization of Clark’s nutcrackers in whitebark pine communities: an agent-based modeling approach. Ecological Modelling, 354(??):123–139, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017300741.

McLane:2011:RAB

[MSMM11] Adam J. McLane, Christina Semeniuk, Gregory J. Mc-Dermid, and Danielle J. Marceau. The role of agentbased models in wildlife ecology and management. Ecological Modelling, 222(8):1544–1556, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000524.

Mouri:2010:EPA

[MSO10] Goro Mouri, Seirou Shinoda, and Taikan Oki. Estimating Plecoglossus altivelis altivelis migration using a mass balance model expressed by hydrological distribution parameters in a major limpid river basin in Japan. Ecological Modelling, 221(23):2808–2815, November 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004242.

Magnea:2013:MHA

[MSP^{+}13] Ulrika Magnea, Roberta Sciascia, Francesco Paparella, Rocco Tiberti, and Antonello Provenzale. A model for highaltitude alpine lake ecosystems and the effect of introduced fish. Ecological Modelling, 251(??):211–220, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005947.

Montalto:2014:TET

[MSR^{+}14] Valeria Montalto, Gianluca Sarà, Paolo Michele Ruti, Alessandro Dell’Aquila, and Brian Helmuth. Testing the effects of temporal data resolution on predictions of the effects of climate change on bivalves. Ecological Modelling, 278(??):1–8, April 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000623.

Marando:2019:RES

[MSS^{+}19] Federica Marando, Elisabetta Salvatori, Alessandro Sebastiani, Lina Fusaro, and Fausto Manes. Regulating ecosystem services and green infrastructure: assessment of Urban Heat Island effect mitigation in the municipality of Rome, Italy. Ecological Modelling, 392(??):92–102, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303995.

Merckx:2011:NMR

[MSV^{+}11] Bea Merckx, Maaike Steyaert, Ann Vanreusel, Magda Vincx, and Jan Vanaverbeke. Null models reveal preferential sampling, spatial autocorrelation and overfitting in habitat suitability modelling. Ecological Modelling, 222(3):588–597, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006216.

Meineri:2012:MAP

[MSV12] Eric Meineri, Olav Skarpaas, and Vigdis Vandvik. Modeling alpine plant distributions at the landscape scale: Do biotic interactions matter? Ecological Modelling, 231(??):1–10, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012000506.

Martin:2019:CGB

[MTA19] Thomas Martin, Pernille Thorbek, and Roman Ashauer. Common ground between growth models of rival theories: a useful illustration for beginners. Ecological Modelling, 407(??):Article 108712, September 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302121.

Matson:2017:NCP

[MTGD17] Sean E. Matson, Ian G. Taylor, Vladlena V. Gertseva, and Martin W. Dorn. Novel catch projection model for a commercial groundfish catch shares fishery. Ecological Modelling, 349(??):51–61, April 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304677.

Melbourne-Thomas:2011:CSU

[MTJF11a] J. Melbourne-Thomas, C. R. Johnson, and E. A. Fulton. Characterizing sensitivity and uncertainty in a multiscale model of a complex coral reef system. Ecological Modelling, 222(18):3320–3334, September 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100384X.

Melbourne-Thomas:2011:RSS

[MTJF11b] J. Melbourne-Thomas, C. R. Johnson, and E. A. Fulton. Regional-scale scenario analysis for the Meso-American Reef system: Modelling coral reef futures under multiple stressors. Ecological Modelling, 222(10):1756–1770, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001153.

Massad:2010:MSA

[MTL^{+}10] Raia-Silvia Massad, Andrée Tuzet, Benjamin Loubet, Alain Perrier, and Pierre Cellier. Model of stomatal ammonia compensation point (STAMP) in relation to the plant nitrogen and carbon metabolisms and environmental conditions. Ecological Modelling, 221(3):479–494, February 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007157.

Marin:2013:MSS

[MTPC13] Víctor H. Marín, Antonio Tironi, María Alejandra Paredes, and Manuel Contreras. Modeling suspended solids in a Northern Chilean Patagonia glacier-fed fjord: GLOF scenarios under climate change conditions. Ecological Modelling, 264(??):7–16, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200289X.

Ma:2010:UFE

[MTZ^{+}10] Hongguang Ma, Howard Townsend, Xinsheng Zhang, Maddy Sigrist, and Villy Christensen. Using a fisheries ecosystem model with a water quality model to explore trophic and habitat impacts on a fisheries stock: a case study of the blue crab population in the Chesapeake Bay. Ecological Modelling, 221(7):997–1004, April 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009000787.

McDonald:2010:UMS

[MU10] Cory P. McDonald and Noel R. Urban. Using a model selection criterion to identify appropriate complexity in aquatic biogeochemical models. Ecological Modelling, 221 (3):428–432, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007078.

Maria:2017:WIM

[MU17] Bobrowski Maria and Schickhoff Udo. Why input matters: Selection of climate data sets for modelling the potential distribution of a treeline species in the Himalayan region. Ecological Modelling, 359(??):92–102, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301801.

Muller:2011:SUM

[Mul11] Erik B. Muller. Synthesizing units as modeling tool for photosynthesizing organisms with photoinhibition and nutrient limitation. Ecological Modelling, 222(3):637–644, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005491.

Medrano:2013:CHB

[MUP^{+}13] E. Aparicio Medrano, R. E. Uittenbogaard, L. M. Dionisio Pires, B. J. H. van de Wiel, and H. J. H. Clercx. Coupling hydrodynamics and buoyancy regulation in Microcystis aeruginosa for its vertical distribution in lakes. Ecological Modelling, 248(??):41–56, January 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004784.

Murray:2011:SMA

[Mur11] Alexander G. Murray. A simple model to assess selection for treatment-resistant sea lice. Ecological Modelling, 222(11):1854–1862, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001396.

Murphy:2012:ECH

[Mur12] John T. Murphy. Exploring complexity with the hohokam water management simulation: a middle way for archaeological modeling. Ecological Modelling, 241(??):15–29, August 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011006144.

Murray:2013:ILB

[Mur13] Alexander G. Murray. Implications of leaky boundaries for compartmentalised control of pathogens: a modelling case study for bacterial kidney disease in Scottish salmon aquaculture. Ecological Modelling, 250(??):177–182, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005297.

MacFarland:2011:USE

[MV11] David M. MacFarland and Timothy R. Van Deelen. Using simulation to explore the functional relationships of terrestrial carnivore population indices. Ecological Modelling, 222(15):2761–2769, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002584.

Miskinis:2017:ASH

[MV17] Paulius Miškinis and Vaida Vasiliauskienė. The analytical solutions of the harvesting Verhulst’s evolution equation. Ecological Modelling, 360(??):189–193, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017303319.

Matejicek:2011:STM

[MVC11] Lubos Matejicek, Eva Vavrova, and Pavel Cudlin. Spatiotemporal modelling of ground vegetation development in mountain spruce forests. Ecological Modelling, 222 (14):2584–2592, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006150.

Maselli:2017:MAW

[MVC^{+}17] F. Maselli, F. P. Vaccari, M. Chiesi, S. Romanelli, and L. P. D’Acqui. Modelling and analyzing the water and carbon dynamics of Mediterranean macchia by the use of ground and remote sensing data. Ecological Modelling, 351(??):1–13, May 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016306469.

May:2010:STR

[MvDL^{+}10] Roel May, Jiska van Dijk, Arild Landa, Roy Andersen, and Reidar Andersen. Spatio-temporal ranging behaviour and its relevance to foraging strategies in wide-ranging wolverines. Ecological Modelling, 221(6):936–943, March 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009008709.

Medrano:2016:SDM

[MvdWU^{+}16] E. Aparicio Medrano, B. J. H. van de Wiel, R. E. Uittenbogaard, L. M. Dionisio Pires, and H. J. H. Clercx. Simulations of the diurnal migration of Microcystis aeruginosa based on a scaling model for physical-biological interactions. Ecological Modelling, 337(??):200–210, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302423.

Ma:2012:SAE

[MW12] Zhihui Ma and Shufan Wang. Stability analysis: the effect of prey refuge in predator–prey model. Ecological Modelling, 247(??):95–97, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004243.

Ma:2015:REP

[MW15] Zhihui Ma and Shufan Wang. Reconsider the effect of prey refuge in a predation model. Ecological Modelling, 317(??):64–65, December 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015004160.

McBride:2010:MPD

[MWB^{+}10] Marissa F. McBride, Kerrie A. Wilson, Jutta Burger, Yi-Chin Fang, Megan Lulow, David Olson, Mike O’Connell, and Hugh P. Possingham. Mathematical problem definition for ecological restoration planning. Ecological Modelling, 221(19):2243–2250, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002140.

Muthukrishnan:2015:ERL

[MWD^{+}15] Ranjan Muthukrishnan, Natalie M. West, Adam S. Davis, Nicholas R. Jordan, and James D. Forester. Evaluating the role of landscape in the spread of invasive species: the case of the biomass crop Miscanthus × giganteus. Ecological Modelling, 317(??):6–15, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003853.

Markovic:2019:SEP

[MWK19] Danijela Markovic, Ariane Walz, and Oskar Kärcher. Scale effects on the performance of niche-based models of freshwater fish distributions: Local vs. upstream area influences. Ecological Modelling, 411(??):Article 108818, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303266.

Moller:2016:MAT

[MWM^{+}16] Cathrine Christmas Møller, Johan J. Weisser, Sijaona Msigala, Robinson Mdegela, Sven Erik Jørgensen, and Bjarne Styrishave. Modelling antibiotics transport in a waste stabilization pond system in Tanzania. Ecological Modelling, 319(??):137–146, January 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004391.

Millington:2013:FGC

[MWML13a] James D. A. Millington, Michael B. Walters, Megan S. Matonis, and Jianguo Liu. Filling the gap: a compositional gap regeneration model for managed northern hardwood forests. Ecological Modelling, 253(??):17–27, March 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000288.

Millington:2013:MFM

[MWML13b] James D. A. Millington, Michael B. Walters, Megan S. Matonis, and Jianguo Liu. Modelling for forest management synergies and trade-offs: Northern hardwood tree regeneration, timber and deer. Ecological Modelling, 248(??):103–112, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004875.

Mao:2018:ENA

[MWY^{+}18] Xufeng Mao, Xiaoyan Wei, Donghai Yuan, Yanxiang Jin, and Xin Jin. An ecological-network-analysis based perspective on the biological control of algal blooms in Ulansuhai Lake, China. Ecological Modelling, 386(??):11–19, October 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302618.

Marimon:2013:SDM

[MXC13] Zachary A. Marimon, Zhemin Xuan, and Ni-Bin Chang. System dynamics modeling with sensitivity analysis for floating treatment wetlands in a stormwater wet pond. Ecological Modelling, 267(??):66–79, October 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003724.

Mao:2011:FAI

[MY11] Xufeng Mao and Zhifeng Yang. Functional assessment of interconnected aquatic ecosystems in the Baiyangdian Basin — an ecological-network-analysis based approach. Ecological Modelling, 222(23–24):3811–3820, December 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100487X.

Masuda:2017:CCA

[MYHN17] Yoshio Masuda, Yasuhiro Yamanaka, Takafumi Hirata, and Hideyuki Nakano. Competition and community assemblage dynamics within a phytoplankton functional group: Simulation using an eddy-resolving model to disentangle deterministic and random effects. Ecological Modelling, 343(??):1–14, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305841.

Myers:2014:FPR

[MYJ^{+}14] Jared T. Myers, Daniel L. Yule, Michael L. Jones, Henry R. Quinlan, and Eric K. Berglund. Foraging and predation risk for larval cisco (Coregonus artedi) in Lake Superior: a modelling synthesis of empirical survey data. Ecological Modelling, 294(??):71–83, December 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004244.

Morales-Zarate:2011:MER

[MZLCSZdP11] M. V. Morales-Zárate, S. E. Lluch-Cota, E. Serviere-Zaragoza, and S. Guzmán del Próo. Modeling an exploited rocky coastal ecosystem: Bahia Tortugas, Mexico. Ecological Modelling, 222(5):1185–1191, March 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006605.

Nenzen:2011:CTA

[NA11] H. K. Nenzén and M. B. Araújo. Choice of threshold alters projections of species range shifts under climate change. Ecological Modelling, 222(18):3346–3354, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003814.

Nakayama:2017:SDS

[Nak17] Tadanobu Nakayama. Scaled-dependence and seasonal variations of carbon cycle through development of an advanced eco-hydrologic and biogeochemical coupling model. Ecological Modelling, 356(??):151–161, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305634.

Ngugi:2011:VMF

[NB11] Michael R. Ngugi and Daniel B. Botkin. Validation of a multispecies forest dynamics model using 50-year growth from Eucalyptus forests in eastern Australia. Ecological Modelling, 222(17):3261–3270, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003644.

Ngugi:2013:RMC

[NBD^{+}13] Michael R. Ngugi, Daniel B. Botkin, David Doley, Mark Cant, and Jack Kelley. Restoration and management of callitris forest ecosystems in Eastern Australia: Simulation of attributes of growth dynamics, growth increment and biomass accumulation. Ecological Modelling, 263(??):152–161, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002548.

Nendel:2011:MMT

[NBK^{+}11] C. Nendel, M. Berg, K. C. Kersebaum, W. Mirschel, X. Specka, M. Wegehenkel, K. O. Wenkel, and R. Wieland. The MONICA model: Testing predictability for crop growth, soil moisture and nitrogen dynamics. Ecological Modelling, 222(9):1614–1625, May 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000998.

Nguyen:2019:CNC

[NBSGW19] Vuong Nguyen, Yvonne M. Buckley, Roberto Salguero-Gómez, and Glenda M. Wardle. Consequences of neglecting cryptic life stages from demographic models. Ecological Modelling, 408(??):Article 108723, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302236.

Navia:2016:SMT

[NCEGB16] Andrés Felipe Navia, Víctor Hugo Cruz-Escalona, Alan Giraldo, and Alberto Barausse. The structure of a marine tropical food web, and its implications for ecosystem-based fisheries management. Ecological Modelling, 328(??):23–33, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300321.

Nam:2010:CCU

[NCK10] Jungho Nam, Wonkeun Chang, and Daeseok Kang. Carrying capacity of an uninhabited island off the southwestern coast of Korea. Ecological Modelling, 221 (17):2102–2107, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010002747.

Nicol:2010:CDM

[NCLP10] Samuel C. Nicol, Iadine Chadès, Simon Linke, and Hugh P. Possingham. Conservation decision-making in large state spaces. Ecological Modelling, 221(21):2531–2536, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010000803.

Navia:2010:TAE

[NCMF10] Andrés F. Navia, Enric Cortés, and Paola A. Mejía-Falla. Topological analysis of the ecological importance of elasmobranch fishes: a food web study on the Gulf of Tortugas, Colombia. Ecological Modelling, 221(24):2918–2926, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001000462X.

Nagai:2016:CSR

[NdA16] Micael E. Nagai and Marcus A. M. de Aguiar. Coevolution in sexually reproducing populations of predators and prey. Ecological Modelling, 337(??):168–175, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001630240X.

Nedorezov:2015:PAD

[Ned15] L. V. Nedorezov. Paramecium aurelia dynamics: Nontraditional approach to estimation of model parameters (on an example of Verhulst and Gompertz models). Ecological Modelling, 317(??):1–5, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003889.

Neill:2011:ICR

[Nei11] Cathy Neill. Impacts of crop residue management on soil organic matter stocks: a modelling study. Ecological Modelling, 222(15):2751–2760, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002602.

Nicholson:2011:CST

[NF11] Ann E. Nicholson and M. Julia Flores. Combining state and transition models with dynamic Bayesian networks. Ecological Modelling, 222(3):555–566, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000551X.

Nagisetty:2019:DOM

[NFU19] Raja M. Nagisetty, Kyle F. Flynn, and Dylan Uecker. Dissolved oxygen modeling of effluent-dominated macrophyterich Silver Bow Creek. Ecological Modelling, 393(??):85– 97, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380018304204.

Niizato:2011:MTI

[NG11] Takayuki Niizato and Yukio-Pegio Gunji. Metric–topological interaction model of collective behavior. Ecological Modelling, 222(17):3041–3049, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003413.

Nieto:2019:WEA

[NGGFM19] P. J. García Nieto, E. García-Gonzalo, J. R. Alonso Fernández, and C. Díaz Muñiz. Water eutrophication assessment relied on various machine learning techniques: a case study in the Englishmen Lake (Northern Spain). Ecological Modelling, 404(??):91–102, July 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301061.

Nye:2013:RIW

[NGL13] Janet A. Nye, Robert J. Gamble, and Jason S. Link. The relative impact of warming and removing top predators on the Northeast US large marine biotic community. Ecological Modelling, 264(??):157–168, August 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004309.

Niccolucci:2011:TNE

[NGR^{+}11] V. Niccolucci, A. Galli, A. Reed, E. Neri, M. Wackernagel, and S. Bastianoni. Towards a 3D National Ecological Footprint Geography. Ecological Modelling, 222(16):2939–2944, August 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011002286.

Nieves-Gonzalez:2019:MMI

[NGRDTHRL19] Aniel Nieves-González, Claudia P. Ruiz-Diaz, Carlos Toledo-Hernández, and Juan S. Ramírez-Lugo. A mathematical model of the interactions between Acropora cervicornis and its environment. Ecological Modelling, 406(??):7–22, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019301413.

Niu:2016:SWE

[NGWZ16] Zhiguang Niu, Qianqian Gou, Xiujun Wang, and Ying Zhang. Simulation of a water ecosystem in a landscape lake in Tianjin with AQUATOX: Sensitivity, calibration, validation and ecosystem prognosis. Ecological Modelling, 335(??):54–63, September 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630165X.

Nielsen:2016:SOC

[Nie16] Søren Nors Nielsen. Second order cybernetics and semiotics in ecological systems — where complexity really begins. Ecological Modelling, 319(??):119–129, January 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003579.

Nukazawa:2021:ESH

[NiSK21] Kei Nukazawa, Jun ichi Shiraiwa, and So Kazama. Evaluations of seasonal habitat variations of freshwater fishes, fireflies, and frogs using a habitat suitability index model that includes river water temperature. Ecological Modelling, 222(20–22):3718–3726, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004649.

Nuttall:2011:IYD

[NJCF11] M. A. Nuttall, A. Jordaan, R. M. Cerrato, and M. G. Frisk. Identifying 120 years of decline in ecosystem structure and maturity of Great South Bay, New York using the Ecopath modelling approach. Ecological Modelling, 222(18):3335–3345, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003747.

Ngugi:2011:REB

[NJM11] Michael R. Ngugi, Robert W. Johnson, and William J. F. McDonald. Restoration of ecosystems for biodiversity and carbon sequestration: Simulating growth dynamics of brigalow vegetation communities in Australia. Ecological Modelling, 222(3):785–794, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000493X.

Niraula:2013:ICS

[NKSA13] Rewati Niraula, Latif Kalin, Puneet Srivastava, and Christopher J. Anderson. Identifying critical source areas of nonpoint source pollution with SWAT and GWLF. Ecological Modelling, 268(??):123–133, October 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004006.

Naghibi:2012:AIH

[NL12] Ali Naghibi and Barbara Lence. Assessing impacts of high flow events on fish population: Evaluation of risk-based performance measures. Ecological Modelling, 240(??):16–28, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002086.

Nawab:2019:UEW

[NLM^{+}19] Asim Nawab, Gengyuan Liu, Fanxin Meng, Yan Hao, and Yan Zhang. Urban energy-water nexus: Spatial and inter-sectoral analysis in a multi-scale economy. Ecological Modelling, 403(??):44–56, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301656.

Nanos:2010:IME

[NLMdL10] Nikos Nanos, Kajsa Larson, Matias Millerón, and Sara Sjöstedt de Luna. Inverse modeling for effective dispersal: Do we need tree size to estimate fecundity? Ecological Modelling, 221(20):2415–2424, October 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010003443.

Nielsen:2019:ESA

[NMA^{+}19] J. K. Nielsen, F. J. Mueter, M. D. Adkison, T. Loher, S. F. McDermott, and A. C. Seitz. Effect of study area bathymetric heterogeneity on parameterization and performance of a depth-based geolocation model for demersal fishes. Ecological Modelling, 402(??):18–34, June 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301280.

Neira:2014:ACSa

[NMC^{+}14] Sergio Neira, Coleen Moloney, Villy Christensen, Philippe Cury, Lynne Shannon, and Hugo Arancibia. Analysing changes in the southern Humboldt ecosystem for the period 1970–2004 by means of dynamic food web modelling. Ecological Modelling, 274(??):41–49, February 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004602.

Nisbet:2016:IEI

[NMdR16] Roger M. Nisbet, Benjamin T. Martin, and Andre M. de Roos. Integrating ecological insight derived from individual-based simulations and physiologically structured population models. Ecological Modelling, 326(??):101–112, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003762.

Nicholls:2010:COS

[NMS10] A. O. Nicholls, Sue McIntyre, and Jacqui Stol. Comments on optimizing the selection of the number of groups in a classification tree. Ecological Modelling, 221(9):1333–1335, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010000785.

Neira:2014:ACSb

[NMS^{+}14] Sergio Neira, Coleen Moloney, Lynne J. Shannon, Villy Christensen, Hugo Arancibia, and Astrid Jarre. Assessing changes in the southern Humboldt in the 20th century using food web models. Ecological Modelling, 278(??):52–66, April 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014000374.

Nathan:2019:SEI

[NMT^{+}19] Lucas R. Nathan, Nadya Mamoozadeh, Hayley R. Tumas, Samuel Gunselman, Keren Klass, Anya Metcalfe, Chris Edge, Lisette P. Waits, Paul Spruell, Erin Lowery, Ed Connor, Andrew R. Bearlin, Marie-Josée Fortin, and Erin Landguth. A spatially-explicit, individual-based demogenetic simulation framework for evaluating hybridization dynamics. Ecological Modelling, 401(??):40–51, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300900.

Nguyen:2019:SBM

[NNP19] Trung H. Nguyen, Duy Nong, and Keith Paustian. Surrogate-based multi-objective optimization of management options for agricultural landscapes using artificial neural networks. Ecological Modelling, 400(??):1–13, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300870.

Nabe-Nielsen:2014:ENC

[NNST^{+}14] Jacob Nabe-Nielsen, Richard M. Sibly, Jakob Tougaard, Jonas Teilmann, and Signe Sveegaard. Effects of noise and by-catch on a Danish harbour porpoise population. Ecological Modelling, 272(??):242–251, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004675.

Natugonza:2016:ESF

[NOOM^{+}16] Vianny Natugonza, Richard Ogutu-Ohwayo, Laban Musinguzi, Benedicto Kashindye, Steingrímur Jónsson, and Hreidar Thor Valtysson. Exploring the structural and functional properties of the Lake Victoria food web, and the role of fisheries, using a mass balance model. Ecological Modelling, 342(??):161–174, December 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630518X.

Nietfeld:2015:MRI

[NP15] H. Nietfeld and J. Prenzel. Modeling the reactive ion dynamics in the rhizosphere of tree roots growing in acid soils. I. Rhizospheric distribution patterns and root uptake of M_{b} cations as affected by root-induced pH and Al dynamics. Ecological Modelling, 307(??):48–65, July 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000678.

Nakayama:2018:IGM

[NP18] Tadanobu Nakayama and Gregory J. Pelletier. Impact of global major reservoirs on carbon cycle changes by using an advanced eco-hydrologic and biogeochemical coupling model. Ecological Modelling, 387(??):172–186, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001830303X.

Nietfeld:2017:MMN

[NPH^{+}17] H. Nietfeld, J. Prenzel, H. S. Helmisaari, A. Polle, and F. Beese. Modeling of mineral nutrient uptake of spruce tree roots as affected by the ion dynamics in the rhizosphere: Upscaling of model results to field plot scale. Ecological Modelling, 345(??):150–164, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303854.

Nguyen:2017:SEM

[NPJ^{+}17] Tuyen Van Nguyen, Young-Seuk Park, Chang-Sik Jeoung, Won-Il Choi, Yong-Kuk Kim, Il-Hyo Jung, Nanako Shigesada, Kohkichi Kawasaki, Fugo Takasu, and Tae-Soo Chon. Spatially explicit model applied to pine wilt disease dispersal based on host plant infestation. Ecological Modelling, 353(??):54–62, June 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306226.

Nikodinoska:2018:AVM

[NPP^{+}18] Natasha Nikodinoska, Alessandro Paletto, Fabio Pastorella, Madeleine Granvik, and Pier Paolo Franzese. Assessing, valuing and mapping ecosystem services at city level: the case of Uppsala (Sweden). Ecological Modelling, 368(??):411–424, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302478.

Nedorezov:2015:DLB

[NS15] L. V. Nedorezov and D. L. Sadykova. Dynamics of larch bud moth populations: Application of Moran–Ricker models with time lag. Ecological Modelling, 297(??):26–32, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005614.

Nagatani:2017:SPC

[NSIiT17] Takashi Nagatani, Kazunori Sato, Genki Ichinose, and Kei ichi Tainaka. Space promotes the coexistence of species: Effective medium approximation for rock-paper-scissors system. Ecological Modelling, 359(??):240–245, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017303241.

Nichols:2017:UOT

[NSN17] J. M. Nichols, J. A. Spendelow, and J. D. Nichols. Using optimal transport theory to estimate transition probabilities in metapopulation dynamics. Ecological Modelling, 359(??):311–319, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303253.

Nunn:2014:SRD

[NTK14] Charles L. Nunn, Peter H. Thrall, and Peter M. Kappeler. Shared resources and disease dynamics in spatially structured populations. Ecological Modelling, 272 (??):198–207, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004663.

Nielsen:2011:OOA

[NU11] Søren Nors Nielsen and Robert E. Ulanowicz. Ontic openness: an absolute necessity for all developmental processes. Ecological Modelling, 222(16):2908–2912, August 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002924.

Niu:2016:PCP

[NvGZ^{+}16] Lixia Niu, P. H. A. J. M. van Gelder, Changkuan Zhang, Yiqing Guan, and J. K. Vrijling. Physical control of phytoplankton bloom development in the coastal waters of Jiangsu (China). Ecological Modelling, 321 (??):75–83, February 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015004664.

Nilsson:2018:MTP

[NW18] Lino Nilsson and Anders Widerlund. Modelling tool for predicting and simulating nitrogen concentrations in coldclimate mining ponds. Ecological Modelling, 380(??):40–52, July 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018301236.

Neeson:2011:RNS

[NWAR11] Thomas M. Neeson, Michael J. Wiley, Sara A. Adlerstein, and Rick L. Riolo. River network structure shapes interannual feedbacks between adult sea lamprey migration and larval habitation. Ecological Modelling, 222(17): 3181–3192, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003620.

Neeson:2012:HRN

[NWAR12] Thomas M. Neeson, Michael J. Wiley, Sara A. Adlerstein, and Rick L. Riolo. How river network structure and habitat availability shape the spatial dynamics of larval sea lampreys. Ecological Modelling, 226(??):62–70, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005813.

Ni:2018:CWM

[NXZ18] Jingneng Ni, Jiuping Xu, and Mengxiang Zhang. Constructed wetland modelling for watershed ecosystem protection under a certain economic load: a case study at the Chaohu Lake watershed, China. Ecological Modelling, 368(??):180–190, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304805.

Neverova:2016:DPD

[NYF16] G. P. Neverova, I. P. Yarovenko, and E. Ya. Frisman. Dynamics of populations with delayed density dependent birth rate regulation. Ecological Modelling, 340 (??):64–73, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016303817.

Nakagawa:2015:CAP

[NYH15] Yoshiaki Nakagawa, Masayuki Yokozawa, and Toshihiko Hara. Competition among plants can lead to an increase in aggregation of smaller plants around larger ones. Ecological Modelling, 301(??):41–53, April 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000332.

Nakagawa:2017:ETP

[NYIH17] Yoshiaki Nakagawa, Masayuki Yokozawa, Akihiko Ito, and Toshihiko Hara. Effectively tuning plant growth models with different spatial complexity: a statistical perspective. Ecological Modelling, 361(??):95–112, October 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305166.

Nabel:2013:ICV

[NZL13] Julia E. M. S. Nabel, Natalie Zurbriggen, and Heike Lischke. Interannual climate variability and population density thresholds can have a substantial impact on simulated tree species’ migration. Ecological Modelling, 257(??):88–100, May 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013000999.

Ospina-Alvarez:2012:VME

[OAPP12] Andrés Ospina-Alvarez, Carolina Parada, and Isabel Palomera. Vertical migration effects on the dispersion and recruitment of European anchovy larvae: From spawning to nursery areas. Ecological Modelling, 231(??):65–79, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000567.

Ortiz:2021:MBT

[OBC^{+}21] Marco Ortiz, Fernando Berrios, Leonardo Campos, Roberto Uribe, Alejandro Ramirez, Brenda Hermosillo-Núñez, Jorge González, and Fabián Rodriguez-Zaragoza. Mass balanced trophic models and short-term dynamical simulations for benthic ecological systems of Mejillonesa and Antofagasta bays (SE Pacific): Comparative network structure and assessment of human impacts. Ecological Modelling, 309–310(??):153–162, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001362.

Olsson:2013:PTP

[OBLJ13] Cecilia Olsson, Kjell Bolmgren, Johan Lindström, and Anna Maria Jönsson. Performance of tree phenology models along a bioclimatic gradient in Sweden. Ecological Modelling, 266(??):103–117, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003219.

Olsson:2015:MPB

[OBSL15] Ola Olsson, Arvid Bolin, Henrik G. Smith, and Eric V. Lonsdorf. Modeling pollinating bee visitation rates in heterogeneous landscapes from foraging theory. Ecological Modelling, 316(??):133–143, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003725.

Opalinska:2015:MMB

[OC15] B. Opalinska and S. A. Cowling. Modelling the movement of biogenic silica from terrestrial vegetation to riverine systems within the continental USA. Ecological Modelling, 312(??):104–113, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002094.

Ortiz:2013:NPK

[OCB^{+}13] Marco Ortiz, Leonardo Campos, Fernando Berrios, Fabián Rodriguez, Brenda Hermosillo, and Jorge González. Network properties and keystoneness assessment in different intertidal communities dominated by two ecosystem engineer species (SE Pacific coast): a comparative analysis. Ecological Modelling, 250(??):307–318, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005169.

Ortega-Cisneros:2017:AMS

[OCCF17] Kelly Ortega-Cisneros, Kevern Cochrane, and Elizabeth A. Fulton. An Atlantis model of the southern Benguela upwelling system: Validation, sensitivity analysis and insights into ecosystem functioning. Ecological Modelling, 355(??):49–63, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305518.

Owens:2013:CIE

[OCD^{+}13] Hannah L. Owens, Lindsay P. Campbell, L. Lynnette Dornak, Erin E. Saupe, Narayani Barve, Jorge Soberón, Kate Ingenloff, Andrés Lira-Noriega, Christopher M. Hensz, Corinne E. Myers, and A. Townsend Peterson. Constraints on interpretation of ecological niche models by limited environmental ranges on calibration areas. Ecological Modelling, 263(??):10–18, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002159.

Ochoa:2010:CTP

[Och10] Andrés Ochoa. Comments on Tixier, P. et al., 2004. “SIMBA-POP: a cohort population model for long-term simulation of banana crop harvest” [Ecol. Model. 180, 407– 417]. Ecological Modelling, 221(12):1648, June 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001110.

Otjacques:2016:DCD

[ODK16] W. Otjacques, F. De Laender, and P. Kestemont. Discerning the causes of a decline in a common European fish, the roach (Rutilus rutilus L.): a modelling approach. Ecological Modelling, 322(??):92–100, February 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005542.

Ordonez:2015:MIB

[OdlFDP15] César Ordoñez, Alberto de la Fuente, and Paula Díaz-Palma. Modeling the influence of benthic primary production on oxygen transport through the water–sediment interface. Ecological Modelling, 311(??):1–10, September 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002045.

Oborny:2012:PGF

[OE12] Beáta Oborny and Péter Englert. Plant growth and foraging for a patchy resource: a credit model. Ecological Modelling, 234(??):20–30, June 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006053.

Oomen:2016:MRP

[OES16] Roelof J. Oomen, Frank Ewert, and Hennie A. Snyman. Modelling rangeland productivity in response to degradation in a semi-arid climate. Ecological Modelling, 322(??):54–70, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005189.

Orozco-Fuentes:2019:EWS

[OFGH^{+}19] S. Orozco-Fuentes, G. Griffiths, M. J. Holmes, R. Ettelaie, J. Smith, A. W. Baggaley, and N. G. Parker. Early warning signals in plant disease outbreaks. Ecological Modelling, 393(??):12–19, February 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303715.

Ogris:2019:RPM

[OFH^{+}19] Nikica Ogris, Mitja Ferlan, Tine Hauptman, Roman Pavlin, Andreja Kavčič, Maja Jurc, and Maarten de Groot. RITY — a phenology model of Ips typographus as a tool for optimization of its monitoring. Ecological Modelling, 410(??):Article 108775, October 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302832.

ORegan:2012:RWC

[OFK^{+}12] Suzanne M. O’Regan, Denis Flynn, Thomas C. Kelly, Michael J. A. O’Callaghan, Alexei V. Pokrovskii, and Dmitrii Rachinskii. The response of the woodpigeon (Columba palumbus) to relaxation of intraspecific competition: a hybrid modelling approach. Ecological Modelling, 224(1):54–64, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005035.

Ogawa:2011:TAC

[Oga11] Kazuharu Ogawa. Theoretical analysis of change in forest carbon use efficiency with stand development: a case study on Hinoki Cypress (Chamaecyparis obtusa (Sieb. et Zucc.) Endl.) plantation from the seedling stage. Ecological Modelling, 222(3):437–441, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000582X.

Ogawa:2017:MAR

[Oga17] Kazuharu Ogawa. Modeling age-related stand respiration changes in forest stands under the self-thinning law.

Ecological Modelling, 349(??):62–68, April 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305816.

Ogawa:2018:MCA

[Oga18] Kazuharu Ogawa. Mathematical consideration of the agerelated decline in leaf biomass in forest stands under the selfthinning law. Ecological Modelling, 372(??):64–69, March 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001830036X.

Ofir:2016:DCF

[OGG^{+}16] E. Ofir, G. Gal, M. Goren, J. Shapiro, and E. Spanier. Detecting changes to the functioning of a lake ecosystem following a regime shift based on static food-web models. Ecological Modelling, 320(??):145–157, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003890.

Ostojski:2016:IRS

[OGOWW16] Mieczysław S. Ostojski, Joanna Gȩbala, Paulina Orlińska-Woźniak, and Paweł Wilk. Implementation of robust statistics in the calibration, verification and validation step of model evaluation to better reflect processes concerning total phosphorus load occurring in the catchment. Ecological Modelling, 332(??):83–93, July 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301016.

Ofir:2017:PIL

[OHS^{+}17] E. Ofir, J. J. Heymans, J. Shapiro, M. Goren, E. Spanier, and G. Gal. Predicting the impact of Lake Biomanipulation based on food-web modeling — Lake Kinneret as a case study. Ecological Modelling, 348(??):14–24, March 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016308584.

Ogris:2010:SFN

[OJ10] Nikica Ogris and Maja Jurc. Sanitary felling of Norway spruce due to spruce bark beetles in Slovenia: a model and projections for various climate change scenarios. Ecological Modelling, 221(2):290–302, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009003251.

Olsson:2015:MFT

[OJ15] Cecilia Olsson and Anna Maria Jönsson. A model framework for tree leaf colouring in Europe. Ecological Modelling, 316(??):41–51, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003464.

Ohta:2012:AMB

[OJM^{+}12] Umika Ohta, Marko Jusup, Tsutomu Mano, Hifumi Tsuruga, and Hiroyuki Matsuda. Adaptive management of the brown bear population in Hokkaido, Japan. Ecological Modelling, 242(??):20–27, September 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002323.

Ortiz:2011:MSC

[OKS^{+}11] Carina Ortiz, Erik Karltun, Johan Stendahl, Annemieke I. Gärdenäs, and Göran I. Ågren. Modelling soil carbon development in Swedish coniferous forest soils — an uncertainty analysis of parameters and model estimates using the GLUE method. Ecological Modelling, 222(17):3020–3032, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003267.

Oli:2017:DLN

[OLC^{+}17] Madan K. Oli, W. J. Loughry, Hal Caswell, Carolina Perez-Heydrich, Colleen M. McDonough, and Richard W. Truman. Dynamics of leprosy in nine-banded armadillos: Net reproductive number and effects on host population dynamics.

Ecological Modelling, 350(??):100–108, April 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307050.

Ortiz:2013:SOC

[OLG^{+}13] Carina A. Ortiz, Jari Liski, Annemieke I. Gärdenäs, Aleksi Lehtonen, Mattias Lundblad, Johan Stendahl, Göran I. Ågren, and Erik Karltun. Soil organic carbon stock changes in Swedish forest soils — a comparison of uncertainties and their sources through a national inventory and two simulation models. Ecological Modelling, 251(??):221–231, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012005972.

Oliver:2011:IHR

[OMDP11] Lauren J. Oliver, Byron J. T. Morgan, Sarah M. Durant, and Nathalie Pettorelli. Individual heterogeneity in recapture probability and survival estimates in cheetah. Ecological Modelling, 222(3):776–784, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006265.

Oborny:2012:VPR

[OMH12] B. Oborny, C. Mony, and T. Herben. From virtual plants to real communities: a review of modelling clonal growth. Ecological Modelling, 234(??):3–19, June 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001172.

Ochoa-Ochoa:2016:UOV

[OOFVBC16] Leticia M. Ochoa-Ochoa, Oscar A. Flores-Villela, and Juan E. Bezaury-Creel. Using one vs. many, sensitivity and uncertainty analyses of species distribution models with focus on conservation area networks. Ecological Modelling, 320(??):372–382, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005165.

Omori:2012:SFF

[OOH^{+}12] Koji Omori, Hidejiro Ohnishi, Hideki Hamaoka, Tadao Kunihiro, Sayaka Ito, Michinobu Kuwae, Hiroki Hata, Todd W. Miller, and Keiichiro Iguchi. Speciation of fluvial forms from amphidromous forms of migratory populations. Ecological Modelling, 243(??):89–94, September 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002785.

Ouyang:2013:ICP

[OQH^{+}13] Wei Ouyang, Shasha Qi, Fanghua Hao, Xuelei Wang, Yushu Shan, and Siyang Chen. Impact of crop patterns and cultivation on carbon sequestration and global warming potential in an agricultural freeze zone. Ecological Modelling, 252(??):228–237, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002293.

Ophithakorn:2013:SMD

[OSS13] Thiwari Ophithakorn, Chaisri Suksaroj, and Thunwadee Tachapattaworakul Suksaroj. Simulation modelling of dissolved organic matter removal in a free water surface constructed wetland. Ecological Modelling, 258(??):82–90, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001567.

Oke:2015:DMM

[OT15] Oluwatobi A. Oke and Ken A. Thompson. Distribution models for mountain plant species: the value of elevation. Ecological Modelling, 301(??):72–77, April 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000381.

Ogbazghi:2016:MMS

[OTA16] Z. M. Ogbazghi, E. H. Tesfamariam, and J. G. Annandale. Modelling N mineralisation from sludge-amended soils across agro-ecological zones: a case study from South Africa.

Ecological Modelling, 322(??):19–30, February 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005517.

Ooba:2010:BMB

[OWMK10] Makoto Ooba, Qinxue Wang, Shogo Murakami, and Kunio Kohata. Biogeochemical model (BGC-ES) and its basin-level application for evaluating ecosystem services under forest management practices. Ecological Modelling, 221(16):1979–1994, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002590.

Perhar:2013:MZG

[PAB13a] Gurbir Perhar, George B. Arhonditsis, and Michael T. Brett. Modeling zooplankton growth in Lake Washington: a mechanistic approach to physiology in a eutrophication model. Ecological Modelling, 258(??):101–121, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001117.

Perry:2013:NER

[PAB^{+}13b] Joe N. Perry, Salvatore Arpaia, Detlef Bartsch, A. Nick E. Birch, Yann Devos, Achim Gathmann, Andrea Gennaro, Jozsef Kiss, Antoine Messéan, Sylvie Mestdagh, Marco Nuti, Jeremy B. Sweet, and Christoph C. Tebbe. No evidence requiring change in the risk assessment of Inachis io larvae. Ecological Modelling, 268(??):103–122, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003979.

Pereira:2019:ENS

[PAEE19] G. C. Pereira, L. P. Andrade, R. P. Espíndola, and N. F. F. Ebecken. Ecological networks simulation by fuzzy ecotoxicological rules. Ecological Modelling, 409(??): Article 108733, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302339.

Piedecausa:2010:STP

[PAGCV^{+}10] M. A. Piedecausa, F. Aguado-Giménez, J. Cerezo-Valverde, M. D. Hernández-Llorente, and B. García-García. Simulating the temporal pattern of waste production in farmed gilthead seabream (Sparus aurata), European seabass (Dicentrarchus labrax) and Atlantic bluefin tuna (Thunnus thynnus). Ecological Modelling, 221(4):634–640, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007571.

Powers:2011:MWS

[PANL11] S. E. Powers, J. C. Ascough, R. G. Nelson, and G. R. Larocque. Modeling water and soil quality environmental impacts associated with bioenergy crop production and biomass removal in the midwest USA. Ecological Modelling, 222(14):2430–2447, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001050.

Parsons:2010:CDN

[Par10] David Parsons. Comment on: the dual nature of ecosystem dynamics. Ecological Modelling, 221(22):2626, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010003856.

Pan:2010:UFS

[PAS^{+}10] Z. Pan, D. Andrade, M. Segal, J. Wimberley, N. McKinney, and E. Takle. Uncertainty in future soil carbon trends at a central U.S. site under an ensemble of GCM scenario climates. Ecological Modelling, 221(5):876–881, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007959.

Patten:2010:EES

[Pat10] Bernard C. Patten. Environs, eco-systems, and ecosystems: reply to a critique of ecosystems emerging, comment on D. Schizas, G. Stamou [Ecol. Model. 221 (2010) 1630– 1635]. Ecological Modelling, 221(12):1636–1643, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010001584.

Patterson:2012:APE

[Pat12] Murray G. Patterson. Are all processes equally efficient from an emergy perspective?: Analysis of ecological and economic networks using matrix algebra methods. Ecological Modelling, 226(??):77–91, February 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100545X.

Patten:2014:SEE

[Pat14a] Bernard C. Patten. Systems ecology and environmentalism: Getting the science right. Part I: Facets for a more holistic Nature Book of ecology. Ecological Modelling, 293(??):4–21, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014001914.

Patterson:2014:EMA

[Pat14b] Murray Patterson. Evaluation of matrix algebra methods for calculating transformities from ecological and economic network data. Ecological Modelling, 271(??):72–82, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001683.

Patten:2015:LTQ

[Pat15] Bernard C. Patten. Link tracking: Quantifying network flows from qualitative node–link digraphs. Ecological Modelling, 295(??):47–58, January 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004311.

Patten:2016:CHH

[Pat16a] Bernard C. Patten. The cardinal hypotheses of Holoecology: Facets for a general systems theory of the organism–environment relationship. Ecological Modelling, 319(??):63–111, January 10, 2016. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003488.

Patten:2016:SEE

[Pat16b] Bernard C. Patten. Systems ecology and environmentalism: Getting the science right. Part II: the Janus Enigma Hypothesis. Ecological Modelling, 335(??):101–138, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015004986.

Park:2017:SWQ

[PAT17] Jong-Yoon Park, Srinivasulu Ale, and W. Richard Teague. Simulated water quality effects of alternate grazing management practices at the ranch and watershed scales. Ecological Modelling, 360(??):1–13, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305208.

Pautasso:2014:JNE

[Pau14] Marco Pautasso. The jump in network ecology research between 1990 and 1991 is a Web of Science artefact. Ecological Modelling, 286(??):11–12, August 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002105.

Powell:2014:PDD

[PB14] James A. Powell and Barbara J. Bentz. Phenology and density-dependent dispersal predict patterns of mountain pine beetle (Dendroctonus ponderosae) impact. Ecological Modelling, 273(??):173–185, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005334.

Pages:2014:CEA

[PBB^{+}14] Loïc Pagès, Carole Bécel, Hassan Boukcim, Delphine Moreau, Christophe Nguyen, and Anne-Sophie Voisin. Calibration and evaluation of ArchiSimple, a simple model of root system architecture. Ecological Modelling, 290 (??):76–84, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013005619.

Picone:2017:INC

[PBD^{+}17] F. Picone, E. Buonocore, R. D’Agostaro, S. Donati, R. Chemello, and P. P. Franzese. Integrating natural capital assessment and marine spatial planning: a case study in the Mediterranean sea. Ecological Modelling, 361(??):1–13, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017302648.

Pedersen:2019:NBM

[PBD19] Rune Østergaard Pedersen, Anne Bonis, and Christian Damgaard. A nonlinear Bayesian model of trait selection forces. Ecological Modelling, 393(??):107–119, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304137.

Perissi:2017:DPO

[PBEL17] Ilaria Perissi, Ugo Bardi, Toufic El Asmar, and Alessandro Lavacchi. Dynamic patterns of overexploitation in fisheries. Ecological Modelling, 359(??):285–292, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016306767.

Peterman:2014:SDA

[PBFS14] Wendy Peterman, Dominique Bachelet, Ken Ferschweiler, and Timothy Sheehan. Soil depth affects simulated carbon and water in the MC2 dynamic global vegetation model. Ecological Modelling, 294(??):84–93, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004499.

Peringer:2017:DGV

[PBG^{+}17] Alexander Peringer, Alexandre Buttler, François Gillet, Ileana Pătru-Stupariu, Kiowa A. Schulze, Mihai-Sorin Stupariu, and Gert Rosenthal. Disturbance-grazer-vegetation interactions maintain habitat diversity in mountain pasturewoodlands. Ecological Modelling, 359(??):301–310, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001730100X.

Prokopkin:2014:ODM

[PBK14] I. G. Prokopkin, Y. V. Barkhatov, and E. B. Khromechek. A one-dimensional model for phytoflagellate distribution in the meromictic lake. Ecological Modelling, 288(??):1–8, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014002609.

Park:2015:EME

[PBL^{+}15] Young-Seuk Park, Christophe Baehr, Guy R. Larocque, José M. Sánchez-Pérez, and Sabine Sauvage. Ecological modelling for ecosystem sustainability. Ecological Modelling, 306(??):1–5, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001453.

Pascual:2019:ARV

[PBO19] Adrián Pascual, Felipe Bravo, and Cristóbal Ordoñez. Assessing the robustness of variable selection methods when accounting for co-registration errors in the estimation of forest biophysical and ecological attributes. Ecological Modelling, 403(??):11–19, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301632.

Port:2015:AES

[PBP^{+}15] Alexander Port, Karin R. Bryan, Conrad A. Pilditch, David P. Hamilton, and Kai Bischof. Algebraic equilibrium solution of tissue nitrogen quota in algae and the discrepancy between calibrated parameters and physiological properties. Ecological Modelling, 312(??):281–291, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002409.

Pedersen:2011:EMN

[PBT^{+}11] M. W. Pedersen, C. W. Berg, U. H. Thygesen, A. Nielsen, and H. Madsen. Estimation methods for nonlinear state-space models in ecology. Ecological Modelling, 222 (8):1394–1400, April 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011000299.

Pierucci:2015:EEA

[PC15] Andrea Pierucci and Andrés Cózar. An equation to estimate absolute population density from visual census of mobile animals. Ecological Modelling, 303(??):105–110, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015000733.

Paleari:2016:SAS

[PC16] Livia Paleari and Roberto Confalonieri. Sensitivity analysis of a sensitivity analysis: We are likely overlooking the impact of distributional assumptions. Ecological Modelling, 340(??):57–63, November 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303878.

Pais:2017:FBE

[PC17] Miguel Pessanha Pais and Henrique N. Cabral. Fish behaviour effects on the accuracy and precision of underwater visual census surveys. A virtual ecologist approach using an individual-based model. Ecological Modelling, 346(??):58–69, February 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016308055.

Pulselli:2011:ESC

[PCB11] Federico M. Pulselli, Luca Coscieme, and Simone Bastianoni. Ecosystem services as a counterpart of emergy flows to ecosystems. Ecological Modelling, 222(16):2924–2928, August 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011002304.

Parasiewicz:2012:MMP

[PCRP12] Piotr Parasiewicz, Elena Castelli, Joseph N. Rogers, and Ethan Plunkett. Multiplex modeling of physical habitat for endangered freshwater mussels. Ecological Modelling, 228(??):66–75, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011006119.

Pan:2017:MPB

[PCTW17] Gang Pan, Fei Chai, DanLing Tang, and Dongxiao Wang. Marine phytoplankton biomass responses to typhoon events in the South China Sea based on physical-biogeochemical model. Ecological Modelling, 356(??):38–47, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016306457.

Portalier:2016:SRE

[PCZ^{+}16] Sébastien M. J. Portalier, Mehdi Cherif, Lai Zhang, Gregor F. Fussmann, and Michel Loreau. Size-related effects of physical factors on phytoplankton communities. Ecological Modelling, 323(??):41–50, March 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005554.

Perez:2012:LLS

[PD12] Liliana Perez and Suzana Dragicevic. Landscape-level simulation of forest insect disturbance: Coupling swarm intelligent agents with GIS-based cellular automata model. Ecological Modelling, 231(??):53–64, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200049X.

Peng:2015:ICC

[PD15] Jing Peng and Li Dan. Impacts of CO_{2} concentration and climate change on the terrestrial carbon flux using six global climate–carbon coupled models. Ecological Modelling, 304(??):69–83, May 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000721.

Pauer:2014:AOD

[PDAL14] James J. Pauer, Phillip A. DePetro, Amy M. Anstead, and John C. Lehrter. Application of a one-dimensional model to explore the drivers and lability of carbon in the northern Gulf of Mexico. Ecological Modelling, 294(??):59–70, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004220.

Probert:2011:RAT

[PDBP11] William J. M. Probert, Martin Drechsler, Peter W. J. Baxter, and Hugh P. Possingham. Resource allocation in two species systems: Is it worth acknowledging species interactions? Ecological Modelling, 222(10):1781–1789, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001074.

Piponiot:2018:ATV

[PDD^{+}18] Camille Piponiot, Géraldine Derroire, Laurent Descroix, Lucas Mazzei, Ervan Rutishauser, Plinio Sist, and Bruno Hérault. Assessing timber volume recovery after disturbance in tropical forests — a new modelling framework. Ecological Modelling, 384(??):353–369, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301960.

Prudent:2014:RCM

[PDG^{+}14] Marion Prudent, Zhan Wu Dai, Michel Génard, Nadia Bertin, Mathilde Causse, and Philippe Vivin. Resource competition modulates the seed number–fruit size relationship in a genotype-dependent manner: a modeling approach in grape and tomato. Ecological Modelling, 290(??):54–64, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013004997.

Piqueira:2011:NLC

[PdM11] José Roberto C. Piqueira and Sérgio Henrique V. L. de Mattos. Note on LMC complexity measure. Ecological Modelling, 222(19):3603–3604, October 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004297.

Pepin:2017:EDS

[PDV17] Kim M. Pepin, Amy J. Davis, and Kurt C. VerCauteren. Efficiency of different spatial and temporal strategies for reducing vertebrate pest populations. Ecological Modelling, 365(??):106–118, December 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304349.

Peace:2015:ELN

[Pea15] Angela Peace. Effects of light, nutrients, and food chain length on trophic efficiencies in simple stoichiometric aquatic food chain models. Ecological Modelling, 312 (??):125–135, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002161.

Peck:2012:NHP

[Pec12] Steven L. Peck. Networks of habitat patches in tsetse fly control: Implications of metapopulation structure on assessing local extinction probabilities. Ecological Modelling, 246(??):99–102, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003444.

Penczak:2011:FAC

[Pen11] Tadeusz Penczak. Fish assemblages composition in a natural, then regulated, stream: a quantitative long-term study. Ecological Modelling, 222(13):2103–2118, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001694.

Peron:2018:PBV

[Pér18] Guillaume Péron. Process-based vs. ad-hoc methods to estimate mortality using carcass surveys data: a review and a note about evidence complacency. Ecological Modelling, 384(??):111–118, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302266.

Pavao:2019:CDC

[PES19] D. C. Pavão, R. B. Elias, and L. Silva. Comparison of discrete and continuum community models: Insights from numerical ecology and Bayesian methods applied to Azorean plant communities. Ecological Modelling, 402(??):93–106, June 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019301267.

Pchelkina:2012:COB

[PF12] Irina Pchelkina and Alexander L. Fradkov. Control of oscillatory behavior of multispecies populations. Ecological Modelling, 227(??):1–6, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100514X.

Patten:2018:NIC

[PF18] Bernard C. Patten and Brian D. Fath. Notes from an introductory course on field systems ecology. Ecological Modelling, 368(??):33–40, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303344.

Pauer:2016:MSE

[PFA^{+}16] James J. Pauer, Timothy J. Feist, Amy M. Anstead, Phillip A. DePetro, Wilson Melendez, John C. Lehrter, Michael C. Murrell, Xiaomi Zhang, and Dong S. Ko. A modeling study examining the impact of nutrient boundaries on primary production on the Louisiana continental shelf. Ecological Modelling, 328(??):136–147, May 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300230.

Paparella:2016:BCC

[PFP^{+}16] Francesco Paparella, Chiara Ferracini, Alessandro Portaluri, Alberto Manzo, and Alberto Alma. Biological control of the chestnut gall wasp with T. sinensis: a mathematical model. Ecological Modelling, 338(??):17–36, October 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302678.

Pretzsch:2015:RSM

[PFR15] Hans Pretzsch, David I. Forrester, and Thomas Rötzer. Representation of species mixing in forest growth models. A review and perspective. Ecological Modelling, 313 (??):276–292, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003002.

Prince:2014:ASG

[PFSF14] Dallan R. Prince, Marianne E. Fletcher, Chen Shen, and Thomas H. Fletcher. Application of L-systems to geometrical construction of chamise and juniper shrubs. Ecological Modelling, 273(??):86–95, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005371.

Putz:2011:FDT

[PGA^{+}11] S. Pütz, J. Groeneveld, L. F. Alves, J. P. Metzger, and A. Huth. Fragmentation drives tropical forest fragments to early successional states: a modelling study for Brazilian Atlantic forests. Ecological Modelling, 222(12):1986–1997, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001100175X.

Perkin:2013:SFD

[PGATS13] Joshuah S. Perkin, Keith B. Gido, Ola Al-Ta’ani, and Caterina Scoglio. Simulating fish dispersal in stream networks fragmented by multiple road crossings. Ecological Modelling, 257(??):44–56, May 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001087.

Portell:2014:ISI

[PGG14] Xavier Portell, Anna Gras, and Marta Ginovart. INDISIM-Saccha, an individual-based model to tackle Saccharomyces cerevisiae fermentations. Ecological Modelling, 279(??):12–23, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000921.

Prada:2019:DDG

[PGGMMA19] Marta Prada, Marta González-García, Juan Majada, and Celia Martínez-Alonso. Development of a dynamic growth model for sweet chestnut coppice: a case study in Northwest Spain. Ecological Modelling, 409(??):Article 108761, October 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019302698.

Pilli:2013:ACC

[PGK^{+}13] Roberto Pilli, Giacomo Grassi, Werner A. Kurz, Carolyn E. Smyth, and Viorel Blujdea. Application of the CBM-CFS3 model to estimate Italy’s forest carbon budget, 1995– 2020. Ecological Modelling, 266(??):144–171, September 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003529.

Penczak:2012:ISO

[PGKG12] T. Penczak, Ł. Głowacki, A. Kruk, and W. Galicka. Implementation of a self-organizing map for investigation of impoundment impact on fish assemblages in a large, lowland river: Long-term study. Ecological Modelling, 227(??):64–71, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005886.

Peringer:2016:LSS

[PGR^{+}16] Alexander Peringer, François Gillet, Gert Rosenthal, Ioana Stoicescu, Ileana Pătru-Stupariu, Mihai-Sorin Stupariu, and Alexandre Buttler. Landscape-scale simulation experiments test Romanian and Swiss management guidelines for mountain pasture-woodland habitat diversity. Ecological Modelling, 330(??):41–49, June 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300795.

Pletterbauer:2016:EBD

[PGS16] Florian Pletterbauer, Wolfram Graf, and Stefan Schmutz. Effect of biotic dependencies in species distribution models: the future distribution of Thymallus thymallus under consideration of Allogamus auricollis. Ecological Modelling, 327(??):95–104, May 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016000296.

Piroddi:2010:ELF

[PGV10] Chiara Piroddi, Bearzi Giovanni, and Christensen Villy. Effects of local fisheries and ocean productivity on the northeastern Ionian Sea ecosystem. Ecological Modelling, 221(11):1526–1544, June 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001365.

Pauwels:2019:EMR

[PGW19] Valentijn R. N. Pauwels, Adrien Guyot, and Jeffrey P. Walker. Evaluating model results in scatter plots: a critique. Ecological Modelling, 411(??):Article 108802, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019303102.

Petrosyan:2019:MCI

[PGZ^{+}19] Varos G. Petrosyan, Victor V. Golubkov, Nikolay A. Zavyalov, Lyudmila A. Khlyap, Natalia N. Dergunova, and Fedor A. Osipov. Modelling of competitive interactions between native Eurasian (Castor fiber) and alien North American (Castor canadensis) beavers based on long-term monitoring data (1934–2015). Ecological Modelling, 409(??):Article 108763, October 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302716.

Panayotova:2018:MIC

[PH18] Iordanka N. Panayotova and Lisa Horth. Modeling the impact of climate change on a rare color morph in fish. Ecological Modelling, 387(??):10–16, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302771.

Peng:2010:EMS

[PHC10] Shao-Lin Peng, Yu-Ping Hou, and Bao-Ming Chen. Establishment of Markov successional model and its application for forest restoration reference in Southern China. Ecological Modelling, 221(9):1317–1324, May 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000554.

Phillips:2011:PMS

[Phi11] Jonathan D. Phillips. Predicting modes of spatial change from state-and-transition models. Ecological Modelling, 222(3):475–484, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001000623X.

Phillips:2015:RC

[Phi15] Jonathan D. Phillips. The robustness of chronosequences. Ecological Modelling, 298(??):16–23, February 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000039.

Payn:2014:GOM

[PHP^{+}14] R. A. Payn, A. M. Helton, G. C. Poole, C. Izurieta, A. J. Burgin, and E. S. Bernhardt. A generalized optimization model of microbially driven aquatic biogeochemistry based on thermodynamic, kinetic, and stoichiometric ecological theory. Ecological Modelling, 294(??):1–18, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004189.

Politikos:2015:CMB

[PHP15] Dimitrios V. Politikos, Martin Huret, and Pierre Petitgas. A coupled movement and bioenergetics model to explore the spawning migration of anchovy in the Bay of Biscay. Ecological Modelling, 313(??):212–222, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002859.

Pacilly:2018:SCD

[PHvB^{+}18] Francine C. A. Pacilly, Gert Jan Hofstede, Edith T. Lammerts van Bueren, Geert J. T. Kessel, and Jeroen C. J. Groot. Simulating crop-disease interactions in agricultural landscapes to analyse the effectiveness of host resistance in disease control: the case of potato late blight. Ecological Modelling, 378(??):1–12, June 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830108X.

Prado:2013:ESV

[PICR13] Patricia Prado, Carles Ibáñez, Nuno Caiola, and Enrique Reyes. Evaluation of seasonal variability in the foodweb properties of coastal lagoons subjected to contrasting salinity gradients using network analyses. Ecological Modelling, 265(??):180–193, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002846.

Pieri:2010:MRBb

[Pie10a] Philippe Pieri. Modelling radiative balance in a row-crop canopy: Cross-row distribution of net radiation at the soil surface and energy available to clusters in a vineyard. Ecological Modelling, 221(5):802–811, March 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009005298.

Pieri:2010:MRBa

[Pie10b] Philippe Pieri. Modelling radiative balance in a row-crop canopy: Row–soil surface net radiation partition. Ecological Modelling, 221(5):791–801, March 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008060.

Perez:2015:ESH

[PJ15] Irene Pérez and Marco A. Janssen. The effect of spatial heterogeneity and mobility on the performance of social– ecological systems. Ecological Modelling, 296(??):1–11, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004839.

Pal:2016:MID

[PJ16] David S. Pal and Peter R. Jaffé. Modeling the inhibition of dissolved H_{2} on propionate fermentation and methanogenesis in wetland sediments. Ecological Modelling, 322(??):115–123, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005220.

Patonai:2017:AIF

[PJ17] Katalin Patonai and Ferenc Jordán. Aggregation of incomplete food web data may help to suggest sampling strategies. Ecological Modelling, 352(??):77–89, May 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305488.

Phillis:2012:SSH

[PK12] Yannis A. Phillis and Vassilis S. Kouikoglou. Systemof-systems hierarchy of biodiversity conservation problems. Ecological Modelling, 235–236(??):36–48, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001482.

Pickens:2014:LMT

[PK14] Bradley A. Pickens and Sammy L. King. Linking multitemporal satellite imagery to coastal wetland dynamics and bird distribution. Ecological Modelling, 285(??):1–12, August 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001400194X.

Paul:2016:IIS

[PK16] Prosenjit Paul and T. K. Kar. Impacts of invasive species on the sustainable use of native exploited species. Ecological Modelling, 340(??):106–115, November 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303763.

Pacioni:2018:SEM

[PKB^{+}18] Carlo Pacioni, Malcolm S. Kennedy, Oliver Berry, Danielle Stephens, and Nathan H. Schumaker. Spatially-explicit model for assessing wild dog control strategies in Western Australia. Ecological Modelling, 368(??):246–256, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017304878.

Pitt:2011:TLS

[PKD11] Joel P. W. Pitt, Darren J. Kriticos, and Michael B. Dodd. Temporal limits to simulating the future spread pattern of invasive species: Buddleja davidii in Europe and New Zealand. Ecological Modelling, 222(11):1880–1887, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001463.

Paul:2016:EFC

[PKG16] Prosenjit Paul, T. K. Kar, and Abhijit Ghorai. Ecotourism and fishing in a common ground of two interacting species. Ecological Modelling, 328(??):1–13, May 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300333.

Poethke:2016:EDD

[PKMH16] Hans Joachim Poethke, Alexander Kubisch, Oliver Mitesser, and Thomas Hovestadt. The evolution of density-dependent dispersal under limited information. Ecological Modelling, 338(??):1–10, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016302654.

Pekkarinen:2015:RMW

[PKT15] Antti-Juhani Pekkarinen, Jouko Kumpula, and Olli Tahvonen. Reindeer management and winter pastures in the presence of supplementary feeding and government subsidies. Ecological Modelling, 312(??):256–271, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002367.

Pinkard:2010:EST

[PKW^{+}10] E. A. Pinkard, D. J. Kriticos, T. J. Wardlaw, A. J. Carnegie, and Agathe Leriche. Estimating the spatio-temporal risk of disease epidemics using a bioclimatic niche model. Ecological Modelling, 221(23):2828–2838, November 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004126.

Parrott:2010:ETV

[PL10a] Amy Parrott and J. David Logan. Effects of temperature variation on TSD in turtle (C. picta) populations. Ecological Modelling, 221(10):1378–1393, May 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000075X.

Pongvuthithum:2010:ADS

[PL10b] Radom Pongvuthithum and Chulin Likasiri. Analytical discussions on species extinction in competitive communities due to habitat destruction. Ecological Modelling, 221(22):2634–2641, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004072.

Peters:2016:MLM

[PLB16] Ronny Peters, Yue Lin, and Uta Berger. Machine learning meets individual-based modelling: Self-organising feature maps for the analysis of below-ground competition among plants. Ecological Modelling, 326(??):142–151, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004925.

Penha-Lopes:2012:OCD

[PLFO^{+}12] Gil Penha-Lopes, Mogens R. Flindt, Bo Ommen, Erik Kristensen, Pedro Garret, and José Paula. Organic carbon dynamics in a constructed mangrove wastewater wetland populated with benthic fauna: a modelling approach. Ecological Modelling, 232(??):97–108, May 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000737.

Pliscoff:2014:EAS

[PLHG14] Patricio Pliscoff, Federico Luebert, Hartmut H. Hilger, and Antoine Guisan. Effects of alternative sets of climatic predictors on species distribution models and associated estimates of extinction risk: a test with plants in an arid environment. Ecological Modelling, 288(??):166–177, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014002713.

Pfeiffer:2019:GAR

[PLL^{+}19] Mirjam Pfeiffer, Liam Langan, Anja Linstädter, Carola Martens, Camille Gaillard, Jan C. Ruppert, Steven I. Higgins, Edwin I. Mudongo, and Simon Scheiter. Grazing and aridity reduce perennial grass abundance in semi-arid rangelands — insights from a trait-based dynamic vegetation model. Ecological Modelling, 395(??):11–22, March 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304241.

Pacella:2013:IDI

[PLR^{+}13] Stephen R. Pacella, Benoit Lebreton, Pierre Richard, Donald Phillips, Theodore H. DeWitt, and Nathalie Niquil. Incorporation of diet information derived from Bayesian stable isotope mixing models into mass-balanced marine ecosystem models: a case study from the Marennes–Oléron Estuary, France. Ecological Modelling, 267(??):127–137, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003736.

Peeters:2010:ILD

[PLS^{+}10] F. Peeters, J. Li, D. Straile, K.-O. Rothhaupt, and J. Vijverberg. Influence of low and decreasing food levels on Daphnia-algal interactions: Numerical experiments with a new Dynamic Energy Budget model. Ecological Modelling, 221(22):2642–2655, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004011.

Pommerening:2011:MBA

[PLS11] Arne Pommerening, Valerie LeMay, and Dietrich Stoyan. Model-based analysis of the influence of ecological processes on forest point pattern formation — a case study. Ecological Modelling, 222(3):666–678, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000579X.

Pinkerton:2016:IAP

[PLS^{+}16] Matt H. Pinkerton, Philip O’B. Lyver, Darren W. Stevens, Jeff Forman, Regina Eisert, and Sophie Mormede. Increases in Adélie penguins in the Ross Sea: Could the fishery for Antarctic toothfish be responsible? Ecological Modelling, 337(??):262–271, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302496.

Pommerening:2016:RPG

[PM16] Arne Pommerening and Anders Muszta. Relative plant growth revisited: Towards a mathematical standardisation of separate approaches. Ecological Modelling, 320 (??):383–392, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015004937.

Parolari:2017:BBC

[PMB^{+}17] Anthony J. Parolari, Megan L. Mobley, Allan R. Bacon, Gabriel G. Katul, Daniel deB. Richter, and Amilcare Porporato. Boom and bust carbon-nitrogen dynamics during reforestation. Ecological Modelling, 360(??):108–119, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017303320.

Pecchi:2019:SDM

[PMB^{+}19] Matteo Pecchi, Maurizio Marchi, Vanessa Burton, Francesca Giannetti, Marco Moriondo, Iacopo Bernetti, Marco Bindi, and Gherardo Chirici. Species distribution modelling to support forest management. A literature review. Ecological Modelling, 411(??):Article 108817, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303254.

Pujoni:2016:EFW

[PMBB^{+}16] Diego Guimarães Florencio Pujoni, Paulina Maria Maia-Barbosa, Francisco Antônio Rodrigues Barbosa, Carlos Ruberto Fragoso, Jr., and Egbert H. van Nes. Effects of food web complexity on top-down control in tropical lakes. Ecological Modelling, 320(??):358–365, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004640.

Panagopoulos:2011:SPI

[PMBM11] Y. Panagopoulos, C. Makropoulos, E. Baltas, and M. Mimikou. SWAT parameterization for the identification of critical diffuse pollution source areas under data limitations. Ecological Modelling, 222(19):3500–3512, October 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100425X.

Perz:2013:EER

[PMCKH13] Stephen G. Perz, Rafael Muñoz-Carpena, Gregory Kiker, and Robert D. Holt. Evaluating ecological resilience with global sensitivity and uncertainty analysis. Ecological Modelling, 263(??):174–186, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002482.

Punt:2016:EIH

[PME^{+}16] André E. Punt, Alec D. MacCall, Timothy E. Essington, Tessa B. Francis, Felipe Hurtado-Ferro, Kelli F. Johnson, Isaac C. Kaplan, Laura E. Koehn, Phillip S. Levin, and William J. Sydeman. Exploring the implications of the harvest control rule for Pacific sardine, accounting for predator dynamics: a MICE model. Ecological Modelling, 337(??):79–95, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302034.

Pie:2013:UMU

[PMF^{+}13] Marcio R. Pie, Andreas L. S. Meyer, Carina R. Firkowski, Luiz F. Ribeiro, and Marcos R. Bornschein. Understanding the mechanisms underlying the distribution of microendemic montane frogs (Brachycephalus spp., terrarana: Brachycephalidae) in the Brazilian Atlantic Rainforest. Ecological Modelling, 250(??):165–176, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005212.

Patterson:2017:TMC

[PMH17] Murray Patterson, Garry McDonald, and Derrylea Hardy. Is there more in common than we think? Convergence of ecological footprinting, emergy analysis, life cycle assessment and other methods of environmental accounting. Ecological Modelling, 362(??):19–36, October 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300613.

Potapov:2011:SGM

[PMLL11] Alex Potapov, Jim R. Muirhead, Subhash R. Lele, and Mark A. Lewis. Stochastic gravity models for modeling lake invasions. Ecological Modelling, 222(4):964–972, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010003820.

Parsons:2011:LSM

[PMM11] Russell A. Parsons, William E. Mell, and Peter Mc-Cauley. Linking 3D spatial models of fuels and fire: Effects of spatial heterogeneity on fire behavior. Ecological Modelling, 222(3):679–691, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005831.

Potapov:2013:CWD

[PMP^{+}13] Alex Potapov, Evelyn Merrill, Margo Pybus, David Coltman, and Mark A. Lewis. Chronic wasting disease: Possible transmission mechanisms in deer. Ecological Modelling, 250(??):244–257, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005492.

Picard:2010:CSU

[PMRGF10] Nicolas Picard, Frédéric Mortier, Vivien Rossi, and Sylvie Gourlet-Fleury. Clustering species using a model of population dynamics and aggregation theory. Ecological Modelling, 221(2):152–160, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009006668.

Pouteau:2011:SBM

[PMS11] Robin Pouteau, Jean-Yves Meyer, and Benoît Stoll. A SVM-based model for predicting distribution of the invasive tree Miconia calvescens in tropical rainforests. Ecological Modelling, 222(15):2631–2641, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002638.

Pang:2017:HNA

[PMS^{+}17] Xi Pang, Ulla Mörtberg, Ola Sallnäs, Renats Trubins, Eva-Maria Nordström, and Hannes Böttcher. Habitat network assessment of forest bioenergy options using the landscape simulator LandSim — a case study of kronoberg, southern Sweden. Ecological Modelling, 345(??):99–112, February 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001630761X.

Priester:2017:ATD

[PMTKC17] C. Robert Priester, Jessica Melbourne-Thomas, Andreas Klocker, and Stuart Corney. Abrupt transitions in dynamics of a NPZD model across Southern Ocean fronts. Ecological Modelling, 359(??):372–382, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305737.

Pirotta:2014:AMD

[PNHL14] Enrico Pirotta, Leslie New, John Harwood, and David Lusseau. Activities, motivations and disturbance: an agentbased model of bottlenose dolphin behavioral dynamics and interactions with tourism in Doubtful Sound, New Zealand. Ecological Modelling, 282(??):44–58, June 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001537.

Parreira:2019:DUN

[PNT^{+}19] Micael Rosa Parreira, João Carlos Nabout, Geiziane Tessarolo, Matheus de Souza Lima-Ribeiro, and Fabrício Barreto Teresa. Disentangling uncertainties from niche modeling in freshwater ecosystems. Ecological Modelling, 391(??):1–8, January 10, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303612.

Peters:2018:NMT

[POB18] Ronny Peters, Adewole Olagoke, and Uta Berger. A new mechanistic theory of self-thinning: Adaptive behaviour of plants explains the shape and slope of self-thinning trajectories. Ecological Modelling, 390(??):1–9, December 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303284.

Picard:2010:CCS

[POBH10] Nicolas Picard, Dakis Ouédraogo, and Avner Bar-Hen. Choosing classes for size projection matrix models. Ecological Modelling, 221(19):2270–2279, September 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002796.

Podani:2018:EMP

[PÓF^{+}18] János Podani, Péter Ódor, Simone Fattorini, Giovanni Strona, Jani Heino, and Dénes Schmera. Exploring multiple presence-absence data structures in ecology. Ecological Modelling, 383(??):41–51, September 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301674.

Pogson:2011:MMY

[Pog11] Mark Pogson. Modelling Miscanthus yields with low resolution input data. Ecological Modelling, 222(23–24): 3849–3853, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004935.

Peace:2019:HIC

[POS^{+}19] Angela Peace, Suzanne M. O’Regan, Jennifer A. Spatz, Patrick N. Reilly, Rachel D. Hill, E. Davis Carter, Rebecca P. Wilkes, Thomas B. Waltzek, Debra L. Miller, and Matthew J. Gray. A highly invasive chimeric ranavirus can decimate tadpole populations rapidly through multiple transmission pathways. Ecological Modelling, 410(??): Article 108777, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302856.

Pagendam:2010:ROO

[PP10] D. E. Pagendam and P. K. Pollett. Robust optimal observation of a metapopulation. Ecological Modelling, 221 (21):2521–2525, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001146.

Piou:2012:DGI

[PP12] Cyril Piou and Etienne Prévost. A demo-genetic individualbased model for Atlantic salmon populations: Model structure, parameterization and sensitivity. Ecological Modelling, 231(??):37–52, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012000543.

Parolari:2016:FSC

[PP16] Anthony J. Parolari and Amilcare Porporato. Forest soil carbon and nitrogen cycles under biomass harvest: Stability, transient response, and feedback. Ecological Modelling, 329(??):64–76, June 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630062X.

Painter:2019:EIH

[PP19] K. J. Painter and A. Z. Plochocka. Efficiency of island homing by sea turtles under multimodal navigating strategies. Ecological Modelling, 391(??):40–52, January 10, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303624.

Paoli:2018:NCE

[PPB^{+}18] C. Paoli, P. Povero, E. Burgos, G. Dapueto, G. Fanciulli, F. Massa, P. Scarpellini, and P. Vassallo. Natural capital and environmental flows assessment in marine protected areas: the case study of Liguria region (NW Mediterranean Sea). Ecological Modelling, 368(??):121–135, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017302284.

Proulx:2013:IRR

[PPBD13] Catherine L. Proulx, Lorraine Proulx, and Gabriel Blouin-Demers. Improving the realism of random walk movement analyses through the incorporation of habitat bias. Ecological Modelling, 269(??):18–20, November 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300402X.

Punt:2014:EIO

[PPDF14] André E. Punt, Dusanka Poljak, Michael G. Dalton, and Robert J. Foy. Evaluating the impact of ocean acidification on fishery yields and profits: the example of red king crab in Bristol Bay. Ecological Modelling, 285(??):39–53, August 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014001987.

Pulselli:2011:CUE

[PPF11] Federico M. Pulselli, Nicoletta Patrizi, and Silvia Focardi. Calculation of the unit emergy value of water in an Italian watershed. Ecological Modelling, 222(16):2929–2938, August 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011002298.

Prowe:2012:CDP

[PPO12] A. E. Friederike Prowe, Markus Pahlow, and Andreas Oschlies. Controls on the diversity–productivity relationship in a marine ecosystem model. Ecological Modelling, 225(??):167–176, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005606.

Parry:2017:ELD

[PPZ^{+}17] H. R. Parry, C. A. Paull, M. P. Zalucki, A. R. Ives, A. Hulthen, and N. A. Schellhorn. Estimating the landscape distribution of eggs by Helicoverpa spp., with implications for Bt resistance management. Ecological Modelling, 365(??):129–140, December 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304337.

Peringer:2011:EPS

[PR11] Alexander Peringer and Gert Rosenthal. Establishment patterns in a secondary tree line ecotone. Ecological Modelling, 222(17):3120–3131, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100305X.

Pagendam:2013:OUG

[PR13] D. E. Pagendam and J. V. Ross. Optimal use of GPS transmitter for estimating species migration rate. Ecological Modelling, 249(??):37–41, January 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003328.

Prato:2012:IRN

[Pra12] Tony Prato. Increasing resilience of natural protected areas to future climate change: a fuzzy adaptive management approach. Ecological Modelling, 242(??):46–53, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002359.

Piacenza:2017:ABM

[PRH17] Susan E. Piacenza, Paul M. Richards, and Selina S. Heppell. An agent-based model to evaluate recovery times and monitoring strategies to increase accuracy of sea turtle population assessments. Ecological Modelling, 358(??):25–39, August 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016307396.

Pachzelt:2013:CPG

[PRHH13] Adrian Pachzelt, Anja Rammig, Steven Higgins, and Thomas Hickler. Coupling a physiological grazer population model with a generalized model for vegetation dynamics. Ecological Modelling, 263(??):92–102, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002494.

Park:2019:WRE

[PRHK19] Joseph Park, Jed Redwine, Troy D. Hill, and Kevin Kotun. Water resource and ecotone transformation in coastal ecosystems. Ecological Modelling, 405(??):69–85, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301607.

Pethybridge:2013:REA

[PRL^{+}13] H. Pethybridge, D. Roos, V. Loizeau, L. Pecquerie, and C. Bacher. Responses of European anchovy vital rates and population growth to environmental fluctuations: an individual-based modeling approach. Ecological Modelling, 250(??):370–383, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005558.

Pelak:2017:DSF

[PRP17] Norman Pelak, Roberto Revelli, and Amilcare Porporato. A dynamical systems framework for crop models: Toward optimal fertilization and irrigation strategies under climatic variability. Ecological Modelling, 365(??):80–92, December 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304325.

Pommerening:2013:WMV

[PS13] Arne Pommerening and Aila Särkkä. What mark variograms tell about spatial plant interactions. Ecological Modelling, 251(??):64–72, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005753.

Page:2017:CUP

[PSB^{+}17] Trevor Page, Paul J. Smith, Keith J. Beven, Ian D. Jones, J. Alex Elliott, Stephen C. Maberly, Eleanor B. Mackay, Mitzi De Ville, and Heidrun Feuchtmayr. Constraining uncertainty and process-representation in an algal community lake model using high frequency in-lake observations. Ecological Modelling, 357(??):1–13, August 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301345.

Peirce:2016:MAT

[PSBH16] J. P. Peirce, G. J. Sandland, B. Bennie, and R. J. Haro. Modeling and analysis of a temperature-driven outbreak of waterfowl disease in the Upper Mississippi River. Ecological Modelling, 320(??):71–78, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004330.

Pasquali:2019:EFM

[PSG19] S. Pasquali, C. Soresina, and G. Gilioli. The effects of fecundity, mortality and distribution of the initial condition in phenological models. Ecological Modelling, 402(??):45–58, June 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019301243.

Patten:2011:EEC

[PSJ11] Bernard C. Patten, Milan Straškraba, and Sven E. Jørgensen. Ecosystems emerging. 5: Constraints. Ecological Modelling, 222(16):2945–2972, August 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002274.

Puzachenko:2016:ATV

[PSKO16] Y. G. Puzachenko, R. B. Sandlersky, A. N. Krenke, and A. Olchev. Assessing the thermodynamic variables of landscapes in the southwest part of East European plain in Russia using the MODIS multispectral band measurements. Ecological Modelling, 319(??):255–274, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003026.

Price:2012:EEK

[PSM^{+}12] Jessica Price, Janet Silbernagel, Nicholas Miller, Randy Swaty, Mark White, and Kristina Nixon. Eliciting expert knowledge to inform landscape modeling of conservation scenarios. Ecological Modelling, 229(??):76–87, March 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004704.

Puzachenko:2016:AST

[PSS16] Y. G. Puzachenko, R. B. Sandlersky, and A. G. Sankovski. Analysis of spatial and temporal organization of biosphere using solar reflectance data from MODIS satellite. Ecological Modelling, 341(??):27–36, December 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304835.

Puzachenko:2011:ETP

[PSSH11] Y. G. Puzachenko, R. B. Sandlersky, and A. Svirejeva-Hopkins. Estimation of thermodynamic parameters of the biosphere, based on remote sensing. Ecological Modelling, 222(16):2913–2923, August 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011002912.

Palosuo:2010:ART

[PSWM^{+}10] Taru Palosuo, Tommi Suominen, Wendelin Werhahn-Mees, Jordi Garcia-Gonzalo, and Marcus Lindner. Assigning results of the Tool for Sustainability Impact Assessment (ToSIA) to products of a forest-wood-chain. Ecological Modelling, 221(18):2215–2225, September 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001651.

Pontin:2011:DFI

[PSWW11] D. R. Pontin, S. Schliebs, S. P. Worner, and M. J. Watts. Determining factors that influence the dispersal of a pelagic species: a comparison between artificial neural networks and evolutionary algorithms. Ecological Modelling, 222(10):1657–1665, May 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001098.

Perez:2012:ESI

[PTA^{+}12] Irene Pérez, Alicia Tenza, José Daniel Anadón, Julia Martínez-Fernández, Andrés Pedreño, and Andrés Giménez. Exurban sprawl increases the extinction probability of a threatened tortoise due to pet collections. Ecological Modelling, 245(??):19–30, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001238.

Plesca:2012:MIE

[PTE^{+}12] I. Plesca, E. Timbe, J.-F. Exbrayat, D. Windhorst, P. Kraft, P. Crespo, K. B. Vaché, H.-G. Frede, and L. Breuer. Model intercomparison to explore catchment functioning: Results from a remote montane tropical rainforest. Ecological Modelling, 239(??):3–13, July 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002663.

Pinnegar:2014:HDL

[PTL14] John K. Pinnegar, Maciej T. Tomczak, and Jason S. Link. How to determine the likely indirect food-web consequences of a newly introduced non-native species: a worked example. Ecological Modelling, 272(??):379–387, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004705.

Papaix:2014:CEC

[PTML14] Julien Papaïx, Suzanne Touzeau, Hervé Monod, and Christian Lannou. Can epidemic control be achieved by altering landscape connectivity in agricultural systems? Ecological Modelling, 284(??):35–47, July 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001951.

Pratheepa:2016:SIT

[PVB16] M. Pratheepa, Abraham Verghese, and H. Bheemanna. Shannon information theory a useful tool for detecting significant abiotic factors influencing the population dynamics of Helicoverpa armigera (Hübner) on cotton crop. Ecological Modelling, 337(??):25–28, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302022.

Pavlova:2014:MEC

[PVBC14] Vera N. Pavlova, Svetlana E. Varcheva, Raushan Bokusheva, and Pierluigi Calanca. Modelling the effects of climate variability on spring wheat productivity in the steppe zone of Russia and Kazakhstan. Ecological Modelling, 277(??):57–67, April 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400057X.

Peters:2014:CAR

[PVL^{+}14] Ronny Peters, Alejandra G. Vovides, Soledad Luna, Uwe Grüters, and Uta Berger. Changes in allometric relations of mangrove trees due to resource availability — a new mechanistic modelling approach. Ecological Modelling, 283(??):53–61, July 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001823.

Patara:2012:INA

[PVM12] Lavinia Patara, Marcello Vichi, and Simona Masina. Impacts of natural and anthropogenic climate variations on North Pacific plankton in an Earth system model. Ecological Modelling, 244(??):132–147, October 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002840.

Patara:2013:RIN

[PVM13] Lavinia Patara, Marcello Vichi, and Simona Masina. Reprint of: “Impacts of natural and anthropogenic climate variations on North Pacific plankton in an Earth System Model”. Ecological Modelling, 264(??):48–63, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002196.

Potgieter:2013:RDM

[PvVC13] L. Potgieter, J. H. van Vuuren, and D. E. Conlong. A reaction–diffusion model for the control of Eldana saccharina Walker in sugarcane using the sterile insect technique. Ecological Modelling, 250(??):319–328, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005571.

Potgieter:2015:RHA

[PvVC15] L. Potgieter, J. H. van Vuuren, and D. E. Conlong. The role of heterogeneous agricultural landscapes in the suppression of pest species following random walk dispersal patterns. Ecological Modelling, 306(??):240–246, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006024.

Padgham:2010:MSM

[PW10a] Mark Padgham and J. Angus Webb. Multiple structural modifications to dendritic ecological networks produce simple responses. Ecological Modelling, 221(21):2537–2545, October 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010002607.

Podur:2010:WCC

[PW10b] Justin Podur and Michael Wotton. Will climate change overwhelm fire management capacity? Ecological Modelling, 221(9):1301–1309, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010000529.

Parker:2016:ELT

[PWBG16] Denham Parker, Henning Winker, Anthony Bernard, and Albrecht Götz. Evaluating long-term monitoring of temperate reef fishes: a simulation testing framework to compare methods. Ecological Modelling, 333(??):1–10, August 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001630103X.

Pavanato:2017:EHW

[PWGS^{+}17] Heloise J. Pavanato, Leonardo L. Wedekin, Fernando R. Guilherme-Silveira, Márcia H. Engel, and Paul G. Kinas. Estimating humpback whale abundance using hierarchical distance sampling. Ecological Modelling, 358(??):10–18, August 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017302806.

Pierce:2017:DVM

[PWK^{+}17] Maria E. Pierce, Tom Warnke, Uwe Krumme, Tobias Helms, Cornelius Hammer, and Adelinde M. Uhrmacher. Developing and validating a multi-level ecological model of eastern Baltic cod (Gadus morhua) in the Bornholm Basin — a case for domain-specific languages. Ecological Modelling, 361(??):49–65, October 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306901.

Pethybridge:2019:CPB

[PWP^{+}19] Heidi R. Pethybridge, Mariska Weijerman, Holly Perrymann, Asta Audzijonyte, Javier Porobic, Vidette McGregor, Raphael Girardin, Cathy Bulman, Kelly Ortega-Cisneros, Matteo Sinerchia, Trevor Hutton, Hector Lozano-Montes, Mao Mori, Camilla Novaglio, Gavin Fay, Rebecca Gorton, and Elizabeth Fulton. Calibrating process-based marine ecosystem models: an example case using Atlantis. Ecological Modelling, 412(??):Article 108822, November 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303308.

Piyaratne:2013:APD

[PZM13] M. K. D. K. Piyaratne, Huiyan Zhao, and Qingxiang Meng. APHIDSim: a population dynamics model for wheat aphids based on swallowtail catastrophe theory. Ecological Modelling, 253(??):9–16, March 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000161.

Qu:2011:ERA

[QCB^{+}11] C. S. Qu, W. Chen, J. Bi, L. Huang, and F. Y. Li. Ecological risk assessment of pesticide residues in Taihu Lake wetland, China. Ecological Modelling, 222(2):287–292, January 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010003698.

Qiao:2018:SDP

[QCLW18] Jianmin Qiao, Qian Cao, Yupeng Liu, and Quanyuan Wu. Scale dependence and parameter sensitivity of the EPIC model in the agro-pastoral transitional zone of north China. Ecological Modelling, 390(??):51–61, December 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303442.

Qiu:2010:IHP

[QDH^{+}10] Z. F. Qiu, A. M. Doglioli, Z. Y. Hu, P. Marsaleix, and F. Carlotti. The influence of hydrodynamic processes on zooplankton transport and distributions in the North Western Mediterranean: Estimates from a Lagrangian model. Ecological Modelling, 221(23):2816–2827, November 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010003832.

Quillet:2013:ARP

[QFG^{+}13] Anne Quillet, Steve Frolking, Michelle Garneau, Julie Talbot, and Changhui Peng. Assessing the role of parameter interactions in the sensitivity analysis of a model of peatland dynamics. Ecological Modelling, 248(??):30–40, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004346.

Han:2016:TYL

[qHLlZ^{+}16] Zhi quan Han, Tong Liu, Xiao ling Zeng, Hua feng Liu, Xiao ran Hao, Yi neng Ouyang, Xin jun Zhao, and Bai lian Li. A two-year life history cycle model for autumn and spring seedling coexistence in an annual plant — an example of intraspecific niche differentiation. Ecological Modelling, 330(??):16–23, June 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300771.

Qian:2012:MCE

[Qia12] Song S. Qian. On model coefficient estimation using Markov chain Monte Carlo simulations: a potential problem and the solution. Ecological Modelling, 247(??):302–306, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004310.

Quijano:2015:SIE

[QL15] Juan C. Quijano and Henry Lin. Is spatially integrated entropy production useful to predict the dynamics of ecosystems? Ecological Modelling, 313(??):341–354, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002616.

Qin:2010:IIF

[QlSeZ^{+}10] Zhong Qin, Gao li Su, Jia en Zhang, Ying Ouyang, Qiang Yu, and Jun Li. Identification of important factors for water vapor flux and CO_{2} exchange in a cropland. Ecological Modelling, 221(4):575–581, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007534.

Qiu:2019:EEF

[QWJL19] Jun Qiu, Jia-Hua Wei, Hao Jiang, and Fang-Fang Li. Ecohydrological evaluation for fish spawning based on fluctuation identification algorithm (FIA). Ecological Modelling, 402(??):35–44, June 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019301486.

Rose:2013:MME

[RA13] Kenneth Alan Rose and J. Icarus Allen. Modeling marine ecosystem responses to global climate change: Where are we now and where should we be going? Ecological Modelling, 264(??):1–6, August 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002184.

Rogers:2010:SME

[RAB^{+}10] Mark W. Rogers, Micheal S. Allen, Paul Brown, Taylor Hunt, Wayne Fulton, and Brett A. Ingram. A simulation model to explore the relative value of stock enhancement versus harvest regulations for fishery sustainability. Ecological Modelling, 221(6):919–926, March 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000900862X.

Reigada:2012:PES

[RAdA12] Carolina Reigada, Sabrina B. L. Araujo, and M. A. M. de Aguiar. Patch exploitation strategies of parasitoids: the role of sex ratio and forager’s interference in structuring metapopulations. Ecological Modelling, 230(??):11–21, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000233.

Reder:2017:SUA

[RAF17] Klara Reder, Joseph Alcamo, and Martina Flörke. A sensitivity and uncertainty analysis of a continental-scale water quality model of pathogen pollution in African rivers. Ecological Modelling, 351(??):129–139, May 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301485.

Raimondo:2013:DDF

[Rai13] Sandy Raimondo. Density dependent functional forms drive compensation in populations exposed to stressors. Ecological Modelling, 265(??):149–157, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002986.

Richards:2010:IEA

[RAM10] M. L. A. Richards, M. J. Aitkenhead, and A. J. S. Mc-Donald. Improving the effectiveness of angular dispersion in plant neighbourhood models. Ecological Modelling, 221(13–14):1649–1654, July 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001407.

Ropero:2015:ASS

[RAR15] R. F. Ropero, P. A. Aguilera, and R. Rumí. Analysis of the socioecological structure and dynamics of the territory using a hybrid Bayesian network classifier. Ecological Modelling, 311(??):73–87, September 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002057.

Raugei:2011:EIA

[Rau11] Marco Raugei. Emergy indicators applied to human economic systems — a word of caution. Ecological Modelling, 222(23–24):3821–3822, December 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004972.

Ruzicka:2016:PCE

[RBGB16] James J. Ruzicka, Kenneth H. Brink, Dian J. Gifford, and Frank Bahr. A physically coupled end-to-end model platform for coastal ecosystems: Simulating the effects of climate change and changing upwelling characteristics on the Northern California Current ecosystem. Ecological Modelling, 331(??):86–99, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300060.

Russell:2013:DMH

[RBH13] Stephen Russell, Andrew B. Barron, and David Harris. Dynamic modelling of honey bee (Apis mellifera) colony growth and failure. Ecological Modelling, 265 (??):158–169, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002895.

Rakshit:2017:CSF

[RBM^{+}17] Nabyendu Rakshit, Arnab Banerjee, Joyita Mukherjee, Moitreyee Chakrabarty, Stuart R. Borrett, and Santanu Ray. Comparative study of food webs from two different time periods of hooghly matla estuarine system, India through network analysis. Ecological Modelling, 356(??):25–37, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016308560.

Roy:2016:MCP

[RBRK16] Arijit Roy, Sudeepto Bhattacharya, M. Ramprakash, and A. Senthil Kumar. Modelling critical patches of connectivity for invasive maling bamboo (Yushania maling) in darjeeling Himalayas using graph theoretic approach. Ecological Modelling, 329(??):77–85, June 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630045X.

Romero-Calcerrada:2010:SMS

[RCBPMN10] R. Romero-Calcerrada, F. Barrio-Parra, J. D. A. Millington, and C. J. Novillo. Spatial modelling of socioeconomic data to understand patterns of human-caused wildfire ignition risk in the SW of Madrid (central Spain). Ecological Modelling, 221(1):34–45, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009005638.

Rodriguez:2011:CAM

[RCC^{+}11] Daniel Rodríguez, José Ricardo Cure, José Miguel Cotes, Andrew Paul Gutierrez, and Fernando Cantor. A coffee agroecosystem model: I. Growth and development of the coffee plant. Ecological Modelling, 222(19):3626–3639, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011004200.

Romeiro:2011:LCC

[RCCN11] Neyva M. L. Romeiro, Rigoberto G. S. Castro, Eliandro R. Cirilo, and Paulo L. Natti. Local calibration of coliforms parameters of water quality problem at Igapó I Lake, Londrina, Paraná, Brazil. Ecological Modelling, 222(11):1888–1896, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011001414.

Rodriguez:2013:CAM

[RCG^{+}13] Daniel Rodríguez, José Ricardo Cure, Andrew Paul Gutierrez, José Miguel Cotes, and Fernando Cantor. A coffee agroecosystem model: II. Dynamics of coffee berry borer. Ecological Modelling, 248(??):203–214, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004814.

Rodriguez:2017:CAM

[RCGC17] Daniel Rodríguez, José Ricardo Cure, Andrew Paul Gutierrez, and José Miguel Cotes. A coffee agroecosystem model: III. Parasitoids of the coffee berry borer (Hypothenemus hampei). Ecological Modelling, 363(??):96–110, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017300200.

Rustici:2017:PES

[RCP17] M. Rustici, G. Ceccherelli, and L. Piazzi. Predator exploitation and sea urchin bistability: Consequence on benthic alternative states. Ecological Modelling, 344(??):1–5, January 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016306196.

Rutten:2019:FDW

[RCS^{+}19] Anneleen Rutten, Jim Casaer, Kristijn R. R. Swinnen, Marc Herremans, and Herwig Leirs. Future distribution of wild boar in a highly anthropogenic landscape: Models combining hunting bag and citizen science data. Ecological Modelling, 411(??):Article 108804, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303126.

Renton:2013:PMP

[RCSS13] Michael Renton, Samantha Childs, Rachel Standish, and Nancy Shackelford. Plant migration and persistence under climate change in fragmented landscapes: Does it depend on the key point of vulnerability within the lifecycle? Ecological Modelling, 249(??):50–58, January 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003304.

Richards:2011:MEC

[RCST11] Russell Richards, Milani Chaloupka, Marcello Sanò, and Rodger Tomlinson. Modelling the effects of ‘coastal’ acidification on copper speciation. Ecological Modelling, 222 (19):3559–3567, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004352.

Rougier:2014:GMT

[RDD^{+}14] Thibaud Rougier, Hilaire Drouineau, Nicolas Dumoulin, Thierry Faure, Guillaume Deffuant, Eric Rochard, and Patrick Lambert. The GR3D model, a tool to explore the global repositioning dynamics of diadromous fish distribution. Ecological Modelling, 283(??):31–44, July 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014001719.

Rodriguez:2017:AET

[RDD^{+}17] Ricardo A. Rodríguez, Janelle M. Duncan, Juan D. Delgado, Michael J. Vanni, Rodrigo Riera, Ada M. Herrera, and María J. González. Assessment of ecosystem trophodynamic power: a model based on the power equation for an oscillating string. Ecological Modelling, 362(??):80–86, October 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017303770.

Rodriguez:2018:AEE

[RDD^{+}18] Ricardo A. Rodríguez, Janelle M. Duncan, Juan D. Delgado, Michael J. Vanni, Rodrigo Riera, and María J. González. Additional empirical evidence on the intrinsic trend to stationarity in the long run and the nested relationship between abiotic, biotic and anthropogenic factors starting from the organic biophysics of ecosystems (OBEC). Ecological Modelling, 383(??):23–30, September 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301741.

Rodriguez:2013:ETT

[RDH^{+}13] Ricardo A. Rodríguez, Juan D. Delgado, Ada Ma. Herrera, Rodrigo Riera, Rafael Ma. Navarro, Carlos Melián, Lorenzo Dieguez, and Ángel Quirós. Effects of two traits of the ecological state equation on our understanding of species coexistence and ecosystem services. Ecological Modelling, 265(??):1–13, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002792.

Robinet:2019:MDT

[RDJ19] Christelle Robinet, Guillaume David, and Hervé Jactel. Modeling the distances traveled by flying insects based on the combination of flight mill and mark-release-recapture experiments. Ecological Modelling, 402(??):85–92, June 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301437.

Rodriguez:2013:MDM

[RDR^{+}13] Diego J. Rodríguez, Laura Delgado, Santiago Ramos, Vanessa Weinberger, and Yadira Rangel. A model for the dynamics of malaria in Paria Peninsula, Sucre State, Venezuela. Ecological Modelling, 259(??):1–9, June 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001543.

Rodriguez:2017:TDT

[RDR^{+}17] Ricardo A. Rodríguez, Janelle Duncan, Rodrigo Riera, Juan D. Delgado, Angel Quirós, Michael J. Vanni, Tomás Camarena, Jezahel V. Miranda, María E. Perdomo, Ada M. Herrera, and María J. González. Thermostatistical distribution of a trophic energy proxy: Extension for modelling energy pyramids at the inter-taxocene scale and under non-stationary conditions. Ecological Modelling, 361 (??):113–121, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017301035.

Rodriguez:2017:EAC

[RDV^{+}17] Ricardo A. Rodríguez, Janelle M. Duncan, Michael J. Vanni, Alexey V. Melkikh, Juan D. Delgado, Rodrigo Riera, Ada M. Herrera, Tomás Camarena, Ángel Quirós, José M. Fernández-Palacios, Jezahel V. Miranda, María E. Perdomo, María J. Fernández-Rodríguez, Antonia Jiménez-Rodríguez, Rüdiger Otto, Carlos G. Escudero, Rafael M. Navarro-Cerrillo, and María J. González. Exploring the analytical consequences of ecological subjects unwittingly neglected by the mainstream of evolutionary thought. Ecological Modelling, 355(??):70–83, July 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305269.

Roelke:2010:LRP

[RE10] Daniel L. Roelke and Peter M. Eldridge. Losers in the ‘rockpaper-scissors’ game: the role of non-hierarchical competition and chaos as biodiversity sustaining agents in aquatic systems. Ecological Modelling, 221(7):1017–1027, April 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009000908.

Reartes:2016:EEE

[REB^{+}16] S. B. Rodriguez Reartes, V. Estrada, R. Bazán, N. Larrosa, A. Cossavella, A. López, F. Busso, and M. S. Diaz. Evaluation of ecological effects of anthropogenic nutrient loading scenarios in Los molinos reservoir through a mathematical model. Ecological Modelling, 320(??):393–406, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001500513X.

Rempel:2011:ECC

[Rem11] Robert S. Rempel. Effects of climate change on moose populations: Exploring the response horizon through biometric and systems models. Ecological Modelling, 222 (18):3355–3365, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003826.

Rettieva:2012:SCS

[Ret12] A. N. Rettieva. Stable coalition structure in bioresource management problem. Ecological Modelling, 235– 236(??):102–118, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001226.

Ramsey:2012:ABA

[RFV^{+}12] David S. L. Ramsey, David M. Forsyth, Clare J. Veltman, Simon J. Nicol, Charles R. Todd, Robert B. Allen, Will J. Allen, Peter J. Bellingham, Sarah J. Richardson, Chris L. Jacobson, and Richard J. Barker. An approximate Bayesian algorithm for training fuzzy cognitive map models of forest responses to deer control in a New Zealand adaptive management experiment. Ecological Modelling, 240(??):93–104, August 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001950.

Rasche:2012:EGM

[RFZB12] Livia Rasche, Lorenz Fahse, Andreas Zingg, and Harald Bugmann. Enhancing gap model accuracy by modeling dynamic height growth and dynamic maximum tree height. Ecological Modelling, 232(??):133–143, May 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001111.

Roessiger:2013:HEP

[RGH^{+}13] Joerg Roessiger, Verena C. Griess, Fabian Härtl, Christian Clasen, and Thomas Knoke. How economic performance of a stand increases due to decreased failure risk associated with the admixing of species. Ecological Modelling, 255(??):58–69, April 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300046X.

Rees:2011:TEC

[RGHS11] Samuel G. Rees, Anne E. Goodenough, Adam G. Hart, and Richard Stafford. Testing the effectiveness of capture mark recapture population estimation techniques using a computer simulation with known population size. Ecological Modelling, 222(17):3291–3294, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003103.

Rodriguez-Gonzalez:2018:ISE

[RGRMRR18] Pablo T. Rodriguez-Gonzalez, Ramiro Rico-Martinez, and Vicente Rico-Ramirez. An integrated stochastic economicecological-social model with stratified-population. Ecological Modelling, 368(??):15–26, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304672.

Rakotomalala:2015:MEC

[RGU^{+}15] Christiane Rakotomalala, Karine Grangeré, Martin Ubertini, Martin Forêt, and Francis Orvain. Modelling the effect of Cerastoderma edule bioturbation on microphytobenthos resuspension towards the planktonic food web of estuarine ecosystem. Ecological Modelling, 316(??):155–167, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003737.

Rushdi:2015:RMB

[RH15] Ali Muhammad Ali Rushdi and Ahmad Kamal Hassan. Reliability of migration between habitat patches with heterogeneous ecological corridors. Ecological Modelling, 304(??):1–10, May 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000708.

Ruiz-Herrera:2017:CEP

[RH17] Alfonso Ruiz-Herrera. Carry-over effects: population abundance, ecological shifts, and the (dis-)appearance of oscillations. Ecological Modelling, 349(??):26–32, April 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303404.

Reiss:2014:ALD

[RHB14] Kelly Chinners Reiss, Erica Hernandez, and Mark T. Brown. Application of the landscape development intensity (LDI) index in wetland mitigation banking. Ecological Modelling, 271(??):83–89, January 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002329.

Rodriguez:2013:BDT

[RHD^{+}13] Ricardo A. Rodríguez, Ada Ma. Herrera, Juan D. Delgado, Rüdiger Otto, Ángel Quirós, Jacobo Santander, Jezahel V. Miranda, María J. Fernández, Antonia Jiménez-Rodríguez, Rodrigo Riera, Rafael Ma. Navarro, Ma. Elena Perdomo, José Ma. Fernández-Palacios, Carlos G. Escudero, José R. Arévalo, and Lorenzo Diéguez. Biomass-dispersal tradeoff and the functional meaning of species diversity. Ecological Modelling, 261–262(??):8–18, July 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001889.

Rodriguez:2016:RCU

[RHD^{+}16] Ricardo A. Rodríguez, Ada M. Herrera, Janelle Duncan, Rodrigo Riera, Ángel Quirós, María E. Perdomo, Antonia Jiménez-Rodríguez, José María Fernández-Palacios, Michael J. Vanni, Rüdiger Otto, Carlos G. Escudero, Tomás Camarena, Rafael M. Navarro-Cerrillo, Juan D. Delgado, and María J. González. Response to comments on “Uncertainty principle in niche assessment: a solution to the dilemma redundancy vs. competitive exclusion, and some analytical consequences”. Ecological Modelling, 341(??):1–4, December 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304252.

Ramanantoanina:2011:EDD

[RHO11] A. Ramanantoanina, C. Hui, and A. Ouhinou. Effects of density-dependent dispersal behaviours on the speed and spatial patterns of range expansion in predator– prey metapopulations. Ecological Modelling, 222(19): 3524–3530, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004273.

Rodriguez:2012:ESE

[RHO^{+}12] Ricardo A. Rodríguez, Ada M. Herrera, Rüdiger Otto, Juan D. Delgado, José M. Fernández-Palacios, and José R. Arévalo. Ecological state equation. Ecological Modelling, 224(1):18–24, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005059.

Rodriguez:2016:ESC

[RHQ^{+}16] Ricardo A. Rodríguez, Ada Ma. Herrera, Ángel Quirós, María J. Fernández-Rodríguez, Juan D. Delgado, Antonia Jiménez-Rodríguez, José Ma. Fernández-Palacios, Rüdiger Otto, Carlos G. Escudero, Tomás Camarena Luhrs, Jezahel V. Miranda, Rafael Ma. Navarro-Cerrillo, María E. Perdomo, and Rodrigo Riera. Exploring the spontaneous contribution of Claude E. Shannon to eco-evolutionary theory. Ecological Modelling, 327(??):57–64, May 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300096.

Rodriguez:2015:DSD

[RHR^{+}15a] Ricardo A. Rodríguez, Ada M. Herrera, Rodrigo Riera, Jacobo Santander, Jezahel V. Miranda, Ángel Quirós, María J. Fernández-Rodríguez, José M. Fernández-Palacios, Rüdiger Otto, Carlos G. Escudero, Antonia Jiménez-Rodríguez, Rafael M. Navarro-Cerrillo, María E. Perdomo, and Juan D. Delgado. Distribution of species diversity values: a link between classical and quantum mechanics in ecology. Ecological Modelling, 313(??):162–180, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002707.

Rodriguez:2015:TDT

[RHR^{+}15b] Ricardo A. Rodríguez, Ada Ma. Herrera, Rodrigo Riera, Juan D. Delgado, Ángel Quirós, María E. Perdomo, Jacobo Santander, Jezahel V. Miranda, María J. Fernández-Rodríguez, Antonia Jiménez-Rodríguez, José Ma. Fernández-Palacios, Rüdiger Otto, Carlos G. Escudero, and Rafael Ma.. Navarro-Cerrillo. Thermostatistical distribution of a trophic energy proxy with analytical consequences for evolutionary ecology, species coexistence and the maximum entropy formalism. Ecological Modelling, 296(??):24–35, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004864.

Rodriguez:2015:ECA

[RHR^{+}15c] Ricardo A. Rodríguez, Ada Ma. Herrera, Rodrigo Riera, Carlos G. Escudero, and Juan D. Delgado. Empirical clues about the fulfillment of quantum principles in ecology: Potential meaning and theoretical challenges. Ecological Modelling, 301(??):90–97, April 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000423.

Rodriguez:2015:UPN

[RHS^{+}15] Ricardo A. Rodríguez, Ada Ma. Herrera, Jacobo Santander, Jezahel V. Miranda, María J. Fernández-Rodríguez, Ángel Quirós, Rodrigo Riera, José Ma.. Fernández-Palacios, Rüdiger Otto, Carlos G. Escudero, Antonia Jiménez-Rodríguez, Rafael Ma.. Navarro-Cerrillo, María E. Perdomo, and Juan D. Delgado. Uncertainty principle in niche assessment: a solution to the dilemma redundancy vs. competitive exclusion, and some analytical consequences. Ecological Modelling, 316(??):87–110, November 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003543.

Rodriguez:2016:SNS

[RHS^{+}16] Ricardo A. Rodríguez, Ada Ma. Herrera, Jacobo Santander, Jezahel V. Miranda, María E. Perdomo, Ángel Quirós, Rodrigo Riera, and Brian D. Fath. From a stationary to a non-stationary ecological state equation: Adding a tool for ecological monitoring. Ecological Modelling, 320(??):44–51, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015004317.

Raj:2014:VBS

[RHvdTS14] R. Raj, N. A. S. Hamm, C. van der Tol, and A. Stein. Variance-based sensitivity analysis of BIOME-BGC for gross and net primary production. Ecological Modelling, 292(??):26–36, November 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003913.

Radinger:2017:AHU

[RHW17] Johannes Radinger, Franz Hölker, and Christian Wolter. Assessing how uncertainty and stochasticity affect the dispersal of fish in river networks. Ecological Modelling, 359(??):220–228, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305270.

Ren:2018:MUA

[RHZ^{+}18] Xiaoli Ren, Honglin He, Li Zhang, Fan Li, Min Liu, Guirui Yu, and Junhui Zhang. Modeling and uncertainty analysis of carbon and water fluxes in a broad-leaved Korean pine mixed forest based on model-data fusion. Ecological Modelling, 379(??):39–53, July 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830111X.

Rodriguez-Izquierdo:2019:MAC

[RIMPBT19] Emilio Rodríguez-Izquierdo, Yosune Miquelajauregui, Pablo Padilla, and Luis A. Bojórquez-Tapia. Modelling approach for crafting environmental regulations under deep uncertainty: Whale watching in ojo de liebre, Mexico. Ecological Modelling, 408(??):Article 108731, September 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302315.

Ringelman:2014:PFB

[Rin14] Kevin M. Ringelman. Predator foraging behavior and patterns of avian nest success: What can we learn from an agent-based model? Ecological Modelling, 272 (??):141–149, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004729.

Rinnan:2018:PPF

[Rin18] D. Scott Rinnan. Population persistence in the face of climate change and competition: a battle on two fronts. Ecological Modelling, 385(??):78–88, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302394.

Railsback:2011:POM

[RJ11] Steven F. Railsback and Matthew D. Johnson. Patternoriented modeling of bird foraging and pest control in coffee farms. Ecological Modelling, 222(18):3305–3319, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011003796.

Rinella:2017:MFI

[RJ17] Matthew J. Rinella and Jeremy J. James. A modelling framework for improving plant establishment during ecological restoration. Ecological Modelling, 361(??):177–183, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017303587.

Radchuk:2013:BSP

[RJG^{+}13] Viktoriia Radchuk, Karin Johst, Jürgen Groeneveld, Volker Grimm, and Nicolas Schtickzelle. Behind the scenes of population viability modeling: Predicting butterfly metapopulation dynamics under climate change. Ecological Modelling, 259(??):62–73, June 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001713.

Rammig:2010:ICF

[RJH^{+}10] A. Rammig, A. M. Jönsson, T. Hickler, B. Smith, L. Bärring, and M. T. Sykes. Impacts of changing frost regimes on Swedish forests: Incorporating cold hardiness in a regional ecosystem model. Ecological Modelling, 221(2):303–313, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009003354.

Robinson:2014:IST

[RKBG14] Natalie Robinson, Tomas Kadlec, M. Deane Bowers, and Robert P. Guralnick. Integrating species traits and habitat characteristics into models of butterfly diversity in a fragmented ecosystem. Ecological Modelling, 281(??):15–25, June 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014000659.

Riggs:2015:BFD

[RKC^{+}15] Robert A. Riggs, Robert E. Keane, Norm Cimon, Rachel Cook, Lisa Holsinger, John Cook, Timothy DelCurto, L. Scott Baggett, Donald Justice, David Powell, Martin Vavra, and Bridgett Naylor. Biomass and fire dynamics in a temperate forest-grassland mosaic: Integrating multi-species herbivory, climate, and fire with the FireBGCv2/GrazeBGC system. Ecological Modelling, 296(??):57–78, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004827.

Rafique:2015:ACA

[RKL^{+}15] Rashad Rafique, Sandeep Kumar, Yiqi Luo, Gerard Kiely, and Ghassem Asrar. An algorithmic calibration approach to identify globally optimal parameters for constraining the DayCent model. Ecological Modelling, 297 (??):196–200, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001400595X.

Rufener:2017:BSP

[RKNO17] Marie-Christine Rufener, Paul Gerhard Kinas, Marcelo Francisco Nóbrega, and Jorge Eduardo Lins Oliveira. Bayesian spatial predictive models for data-poor fisheries. Ecological Modelling, 348(??):125–134, March 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001730073X.

Reuter:2016:ETF

[RKRB16] Hauke Reuter, Maren Kruse, Alberto Rovellini, and Broder Breckling. Evolutionary trends in fish schools in heterogeneous environments. Ecological Modelling, 326(??):23–35, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015004172.

Ready:2010:PDM

[RKS^{+}10] Jonathan Ready, Kristin Kaschner, Andy B. South, Paul D. Eastwood, Tony Rees, Josephine Rius, Eli Agbayani, Sven Kullander, and Rainer Froese. Predicting the distributions of marine organisms at the global scale. Ecological Modelling, 221(3):467–478, February 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000900711X.

Rock:2012:MFH

[RKW12] K. S. Rock, M. J. Keeling, and D. A. Wood. Modelling the future of the Hawaiian honeycreeper: an ecological and epidemiological problem. Ecological Modelling, 235–236(??):26–35, June 24, 2012. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001706.

Rankin:2015:ASP

[RL15] Peter Sheldon Rankin and Ricardo T. Lemos. An alternative surplus production model. Ecological Modelling, 313(??):109–126, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002732.

Ramin:2012:BSP

[RLB^{+}12] Maryam Ramin, Tanya Labencki, Duncan Boyd, Dennis Trolle, and George B. Arhonditsis. A Bayesian synthesis of predictions from different models for setting water quality criteria. Ecological Modelling, 242(??):127–145, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012002530.

Reaney:2011:RBM

[RLHD11] Sim M. Reaney, Stuart N. Lane, A. Louise Heathwaite, and Lucy J. Dugdale. Risk-based modelling of diffuse land use impacts from rural landscapes upon salmonid fry abundance. Ecological Modelling, 222(4):1016–1029, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004175.

Rathbun:2011:DGM

[RLS11] Leah C. Rathbun, Valerie LeMay, and Nick Smith. Diameter growth models for mixed-species stands of coastal British Columbia including thinning and fertilization effects. Ecological Modelling, 222(14):2234–2248, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001992.

Richards:2013:UMC

[RM13] R. G. Richards and B. J. Mullins. Using microalgae for combined lipid production and heavy metal removal from leachate. Ecological Modelling, 249(??):59–67, January 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003298.

Ramos-Martin:2011:CWB

[RMGU11] Jesús Ramos-Martín, Mario Giampietro, and Sergio Ulgiati. Can we break the addiction to fossil energy? Ecological Modelling, 223(1):1–3, December 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005163.

Romagnoni:2015:EMA

[RMHE15] Giovanni Romagnoni, Steven Mackinson, Jiang Hong, and Anne Maria Eikeset. The ecospace model applied to the North Sea: Evaluating spatial predictions with fish biomass and fishing effort data. Ecological Modelling, 300(??):50–60, March 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014006292.

Ruiz:2013:SBM

[RML^{+}13] J. Ruiz, D. Macías, M. A. Losada, M. Díez-Minguito, and L. Prieto. A simple biogeochemical model for estuaries with high sediment loads: Application to the Guadalquivir River (SW Iberia). Ecological Modelling, 265 (??):194–206, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002962.

Raivonen:2015:SCE

[RML^{+}15] Maarit Raivonen, Päivi Mäkiranta, Annalea Lohila, Sari Juutinen, Timo Vesala, and Eeva-Stiina Tuittila. A simple CO_{2} exchange model simulates the seasonal leaf area development of peatland sedges. Ecological Modelling, 314(??):32–43, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003142.

Ruiz-Moreno:2019:MLM

[RMLTCQ^{+}19] Héctor Alejandro Ruiz-Moreno, Ana María López-Tamayo, Alejandro Caro-Quintero, Johana Husserl, and Andrés Fernando González Barrios. Metagenome level metabolic network reconstruction analysis reveals the microbiome in the Bogotá River is functionally close to the microbiome in produced water. Ecological Modelling, 399(??):1–12, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300560.

Rajabi:2018:SEA

[RMP^{+}18] Mohammadreza Rajabi, Ali Mansourian, Petter Pilesjö, Mohammad Reza Shirzadi, Reza Fadaei, and Javad Ramazanpour. A spatially explicit agent-based simulation model of a reservoir host of cutaneous leishmaniasis, Rhombomys opimus. Ecological Modelling, 370(??):33–49, February 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001730491X.

Reinaldo:2016:AED

[RMR^{+}16] Matías Ocampo Reinaldo, Andrés C. Milessi, María Alejandra Romero, Enrique Crespo, Matthias Wolff, and Raúl A. González. Assessing the effects of demersal fishing and conservation strategies of marine mammals over a Patagonian food web. Ecological Modelling, 331(??):31–43, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005104.

Ratnarajah:2016:PMI

[RMTM^{+}16] Lavenia Ratnarajah, Jessica Melbourne-Thomas, Martin P. Marzloff, Delphine Lannuzel, Klaus M. Meiners, Fanny Chever, Stephen Nicol, and Andrew R. Bowie. A preliminary model of iron fertilisation by baleen whales and Antarctic krill in the Southern Ocean: Sensitivity of primary productivity estimates to parameter uncertainty. Ecological Modelling, 320(??):203–212, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004652.

Roth:2013:TAF

[RMvW13] L. Roth, E. M. Muller, and R. van Woesik. Tracking Acropora fragmentation and population structure through thermal-stress events. Ecological Modelling, 263 (??):223–232, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002524.

Rogers:2012:MSV

[RNE^{+}12] J. Daniel Rogers, Teresa Nichols, Theresa Emmerich, Maciej Latek, and Claudio Cioffi-Revilla. Modeling scale and variability in human–environmental interactions in inner Asia. Ecological Modelling, 241(??):5–14, August 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005679.

Roach:2019:ESS

[RNS^{+}19] Ty N. F. Roach, James Nulton, Paolo Sibani, Forest Rohwer, and Peter Salamon. Emergent structure in a stochastic model of ecological evolution. Ecological Modelling, 401(??):129–133, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301000.

Recknagel:2013:ERT

[ROC^{+}13] Friedrich Recknagel, Ilia Ostrovsky, Hongqing Cao, Tamar Zohary, and Xiaoqing Zhang. Ecological relationships, thresholds and time-lags determining phytoplankton community dynamics of Lake Kinneret, Israel elucidated by evolutionary computation and wavelets. Ecological Modelling, 255(??):70–86, April 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013000860.

Radchuk:2016:SCR

[ROG^{+}16] Viktoriia Radchuk, Steffen Oppel, Jürgen Groeneveld, Volker Grimm, and Nicolas Schtickzelle. Simple or complex: Relative impact of data availability and model purpose on the choice of model types for population viability analyses.

Ecological Modelling, 323(??):87–95, March 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005736.

Ramirez:2015:EER

[ROSC15] Alejandro Ramírez, Marco Ortiz, Jeroen Steenbeek, and Villy Christensen. Evaluation of the effects on rockfish and kelp artisanal fisheries of the proposed Mejillones Peninsula marine protected area (northern Chile, SE Pacific coast). Ecological Modelling, 297(??):141–153, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005857.

Ross:2010:SRR

[RP10] J. V. Ross and P. K. Pollett. Simple rules for ranking and optimally managing metapopulations. Ecological Modelling, 221(21):2515–2520, October 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001122.

Ryrie:2011:HEP

[RP11] Susan C. Ryrie and I. Colin Prentice. Herbivores enable plant survival under nutrient limited conditions in a model grazing system. Ecological Modelling, 222 (3):381–397, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004953.

Ruiz:2012:MTC

[RPA12] Javier Ruiz, Laura Prieto, and Diana Astorga. A model for temperature control of jellyfish (Cotylorhiza tuberculata) outbreaks: a causal analysis in a Mediterranean coastal lagoon. Ecological Modelling, 233(??):59–69, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012001263.

Ricci:2018:HEL

[RPA^{+}18] Benoît Ricci, Sandrine Petit, Charlotte Allanic, Marie Langot, Nicolas Parisey, and Sylvain Poggi. How effective is large landscape-scale planning for reducing local weed infestations? A landscape-scale modelling approach. Ecological Modelling, 384(??):221–232, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302345.

Riera:2017:MTM

[RPC^{+}17] Rodrigo Riera, Óscar Pérez, Chris Cromey, Myriam Rodríguez, Eva Ramos, Omar Álvarez, Julián Domínguez, Óscar Monterroso, and Fernando Tuya. MACAROMOD: a tool to model particulate waste dispersion and benthic impact from offshore sea-cage aquaculture in the Macaronesian region. Ecological Modelling, 361(??):122–134, October 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017303599.

Rio:2015:TFE

[RPGA15] Jeremy Rio, Jean-Sébastien Pierre, Pascaline Le Gouar, and Anne Atlan. Temporal fluctuations in the environment and intra-specific polymorphism: a model simulating the flowering phenology of gorse (Ulex europaeus). Ecological Modelling, 306(??):76–85, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004669.

Ruiz-Perez:2016:CPM

[RPGSCF16] G. Ruiz-Pérez, M. González-Sanchis, A. D. Del Campo, and F. Francés. Can a parsimonious model implemented with satellite data be used for modelling the vegetation dynamics and water cycle in water-controlled environments? Ecological Modelling, 324(??):45–53, March 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016000090.

Record:2013:ECC

[RPM13] N. R. Record, A. J. Pershing, and F. Maps. Emergent copepod communities in an adaptive trait-structured model. Ecological Modelling, 260(??):11–24, July 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001750.

Russo:2015:MSM

[RPP^{+}15] Tommaso Russo, Jacopo Pulcinella, Antonio Parisi, Michela Martinelli, Andrea Belardinelli, Alberto Santojanni, Stefano Cataudella, Sabrina Colella, and Luca Anderlini. Modelling the strategy of mid-water trawlers targeting small pelagic fish in the Adriatic Sea and its drivers. Ecological Modelling, 300(??):102–113, March 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006061.

Roige:2016:SOM

[RPPW16] Mariona Roigé, Matthew Parry, Craig Phillips, and Susan Worner. Self-organizing maps for analysing pest profiles: Sensitivity analysis of weights and ranks. Ecological Modelling, 342(??):113–122, December 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305191.

Ramin:2012:EEN

[RPSA12] Maryam Ramin, Gurbir Perhar, Yuko Shimoda, and George B. Arhonditsis. Examination of the effects of nutrient regeneration mechanisms on plankton dynamics using aquatic biogeochemical modeling. Ecological Modelling, 240(??):139–155, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001913.

Rupp:2010:DIB

[RR10] Susan P. Rupp and Paul Rupp. Development of an individual-based model to evaluate elk (Cervus elaphus nelsoni) movement and distribution patterns following the Cerro Grande Fire in north central New Mexico, USA. Ecological Modelling, 221(12):1605–1619, June 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001596.

Rigosi:2012:PUE

[RR12] Anna Rigosi and Francisco J. Rueda. Propagation of uncertainty in ecological models of reservoirs: From physical to population dynamic predictions. Ecological Modelling, 247(??):199–209, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004334.

Raugei:2014:IEL

[RRBI14] Marco Raugei, Benedetto Rugani, Enrico Benetto, and Wesley W. Ingwersen. Integrating emergy into LCA: Potential added value and lingering obstacles. Ecological Modelling, 271(??):4–9, January 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005637.

Robinson:2015:LRS

[RRG^{+}15] Hugh S. Robinson, Toni Ruth, Justin A. Gude, David Choate, Rich DeSimone, Mark Hebblewhite, Kyran Kunkel, Marc R. Matchett, Michael S. Mitchell, Kerry Murphy, and Jim Williams. Linking resource selection and mortality modeling for population estimation of mountain lions in Montana. Ecological Modelling, 312(??):11–25, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002100.

Rodriguez:2019:DFD

[RRH^{+}19] Ricardo A. Rodríguez, Rodrigo Riera, Ada M. Herrera, Janelle M. Duncan, Michael J. Vanni, Juan D. Delgado, and María J. González. Degrees of freedom: Definitions and their minimum and most meaningful combination for the modelling of ecosystem dynamics with the help of physical principles. Ecological Modelling, 392(??):226–235, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018304095.

Ren:2010:EME

[RRHH10] Jeffrey S. Ren, Alex H. Ross, Mark G. Hadfield, and Barbara J. Hayden. An ecosystem model for estimating potential shellfish culture production in sheltered coastal waters. Ecological Modelling, 221(3):527–539, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007492.

Rodriguez-Rey:2013:SDM

[RRJVA13] Marta Rodríguez-Rey, Alberto Jiménez-Valverde, and Pelayo Acevedo. Species distribution models predict range expansion better than chance but not better than a simple dispersal model. Ecological Modelling, 256(??):1–5, May 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013000914.

Rankovic:2010:NNM

[RRR^{+}10] Vesna Ranković, Jasna Radulović, Ivana Radojević, Aleksandar Ostojić, and Ljiljana Čomić. Neural network modeling of dissolved oxygen in the Gruža reservoir, Serbia. Ecological Modelling, 221(8):1239–1244, April 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008692.

Ropero:2018:DED

[RRvdG18] R. F. Ropero, S. Renooij, and L. C. van der Gaag. Discretizing environmental data for learning Bayesian-network classifiers. Ecological Modelling, 368(??):391–403, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016308377.

Rogers:2012:MWS

[RSC12] Kerrylee Rogers, Neil Saintilan, and Craig Copeland. Modelling wetland surface elevation dynamics and its application to forecasting the effects of sea-level rise on estuarine wetlands. Ecological Modelling, 244(??):148–157, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002864.

Rogers:2013:RMW

[RSC13] Kerrylee Rogers, Neil Saintilan, and Craig Copeland. Reprint of modelling wetland surface elevation dynamics and its application to forecasting the effects of sealevel rise on estuarine wetlands. Ecological Modelling, 264(??):27–36, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002202.

Rossini:2019:NMA

[RSCS19] Luca Rossini, Maurizio Severini, Mario Contarini, and Stefano Speranza. A novel modelling approach to describe an insect life cycle vis-à-vis plant protection: description and application in the case study of Tuta absoluta. Ecological Modelling, 409(??):Article 108778, October 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302868.

Rose:2015:PBM

[RSD^{+}15] Kenneth A. Rose, Shaye Sable, Donald L. DeAngelis, Simeon Yurek, Joel C. Trexler, William Graf, and Denise J. Reed. Proposed best modeling practices for assessing the effects of ecosystem restoration on fish. Ecological Modelling, 300(??):12–29, March 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006334.

Ren:2012:EMO

[RSDP^{+}12] Jeffrey S. Ren, Jeanie Stenton-Dozey, David R. Plew, Jianguang Fang, and Mark Gall. An ecosystem model for optimising production in integrated multitrophic aquaculture systems. Ecological Modelling, 246(??):34–46, November 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012003493.

Rappoldt:2010:GME

[RSE10] Cornelis Rappoldt, Richard A. Stillman, and Bruno J. Ens. A geometrical model for the effect of interference on food intake. Ecological Modelling, 221(2):147–151, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006607.

Reeg:2017:MDI

[RSP^{+}17] Jette Reeg, Thorsten Schad, Thomas G. Preuss, Andreas Solga, Katrin Körner, Christine Mihan, and Florian Jeltsch. Modelling direct and indirect effects of herbicides on non-target grassland communities. Ecological Modelling, 348(??):44–55, March 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630504X.

Russell:2015:EST

[RTE^{+}15] Robin E. Russell, Wayne E. Thogmartin, Richard A. Erickson, Jennifer Szymanski, and Karl Tinsley. Estimating the short-term recovery potential of little brown bats in the eastern United States in the face of white-nose syndrome. Ecological Modelling, 314(??):111–117, October 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003221.

Roa-Ureta:2019:MLT

[RUSL19] Ruben H. Roa-Ureta, Miguel N. Santos, and Francisco Leitão. Modelling long-term fisheries data to resolve the attraction versus production dilemma of artificial reefs. Ecological Modelling, 407(??):Article 108727, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302273.

Romey:2013:SHB

[RV13] William L. Romey and Jose M. Vidal. Sum of heterogeneous blind zones predict movements of simulated groups. Ecological Modelling, 258(??):9–15, June 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001075.

Roitman:2015:ASC

[RV15] Iris Roitman and Jerome K. Vanclay. Assessing size–class dynamics of a neotropical gallery forest with stationary models. Ecological Modelling, 297(??):118–125, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005675.

Roubicek:2010:DCC

[RVB^{+}10] A. J. Roubicek, J. VanDerWal, L. J. Beaumont, A. J. Pitman, P. Wilson, and L. Hughes. Does the choice of climate baseline matter in ecological niche modelling? Ecological Modelling, 221(19):2280–2286, September 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003091.

Robinson:2010:CAS

[RvKM10] Todd P. Robinson, Rieks D. van Klinken, and Graciela Metternicht. Comparison of alternative strategies for invasive species distribution modeling. Ecological Modelling, 221(19):2261–2269, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002346.

Rahn:2018:EAS

[RVL^{+}18] Eric Rahn, Philippe Vaast, Peter Läderach, Piet van Asten, Laurence Jassogne, and Jaboury Ghazoul. Exploring adaptation strategies of coffee production to climate change using a process-based model. Ecological Modelling, 371(??):76–89, March 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300322.

Rinella:2011:EIS

[RVNB11] Matthew J. Rinella, Martin Vavra, Bridgett J. Naylor, and Jennifer M. Boyd. Estimating influence of stocking regimes on livestock grazing distributions. Ecological Modelling, 222(3):619–625, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005326.

Rammelt:2016:EEG

[RvS16] Crelis Ferdinand Rammelt and Maarten van Schie. Ecology and equity in global fisheries: Modelling policy options using theoretical distributions. Ecological Modelling, 337(??):107–122, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302344.

Rands:2010:EPD

[RW10] Sean A. Rands and Heather M. Whitney. Effects of pollinator density-dependent preferences on field margin visitations in the midst of agricultural monocultures: a modelling approach. Ecological Modelling, 221(9):1310–1316, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010000530.

Rasmussen:2017:MES

[RW17] Camilla Ruø Rasmussen and Jacob Weiner. Modelling the effect of size-asymmetric competition on size inequality: Simple models with two plants. Ecological Modelling, 343(??):101–108, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305658.

Russell:2013:EDW

[RWFD13] Matthew B. Russell, Christopher W. Woodall, Shawn Fraver, and Anthony W. D’Amato. Estimates of downed woody debris decay class transitions for forests across the eastern United States. Ecological Modelling, 251 (??):22–31, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005790.

Read:2014:STL

[RWH^{+}14] Jordan S. Read, Luke A. Winslow, Gretchen J. A. Hansen, Jamon Van Den Hoek, Paul C. Hanson, Louise C. Bruce, and Corey D. Markfort. Simulating 2368 temperate lakes reveals weak coherence in stratification phenology. Ecological Modelling, 291(??):142–150, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003664.

Ruzicka:2011:SPM

[RWP11] James J. Ruzicka, Thomas C. Wainwright, and William T. Peterson. A simple plankton model for the Oregon upwelling ecosystem: Sensitivity and validation against time-series ocean data. Ecological Modelling, 222(6):1222–1235, March 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000202.

Rashidi:2015:SSC

[RWS^{+}15] Parinaz Rashidi, Tiejun Wang, Andrew Skidmore, Anton Vrieling, Roshanak Darvishzadeh, Bert Toxopeus, Shadrack Ngene, and Patrick Omondi. Spatial and spatiotemporal clustering methods for detecting elephant poaching hotspots. Ecological Modelling, 297(??):180–186, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014005900.

Rashidi:2016:EPR

[RWS^{+}16] Parinaz Rashidi, Tiejun Wang, Andrew Skidmore, Hamed Mehdipoor, Roshanak Darvishzadeh, Shadrack Ngene, Anton Vrieling, and Albertus G. Toxopeus. Elephant poaching risk assessed using spatial and non-spatial Bayesian models. Ecological Modelling, 338(??):60–68, October 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302691.

Rose:2015:GFS

[RWvDM15] Hannah Rose, Tong Wang, Jan van Dijk, and Eric R. Morgan. GLOWORM-FL: a simulation model of the effects of climate and climate change on the free-living stages of gastro-intestinal nematode parasites of ruminants. Ecological Modelling, 297(??):232–245, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006073.

Reside:2011:ILR

[RWVK11] April E. Reside, Ian Watson, Jeremy VanDerWal, and Alex S. Kutt. Incorporating low-resolution historic species location data decreases performance of distribution models. Ecological Modelling, 222(18):3444–3448, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011003632.

Ranjitkar:2014:EFC

[RXSK14] Sailesh Ranjitkar, Jianchu Xu, Krishna Kumar Shrestha, and Roeland Kindt. Ensemble forecast of climate suitability for the Trans-Himalayan Nyctaginaceae species. Ecological Modelling, 282(??):18–24, June 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001276.

Ren:2013:MMW

[RZF^{+}13] Zongming Ren, Yang Zeng, Xiu’e Fu, Gaosheng Zhang, Linlin Chen, Jing Chen, Tae-Soo Chon, Yawei Wang, and Yuansong Wei. Modeling macrozooplankton and water quality relationships after wetland construction in the Wenyuhe River Basin, China. Ecological Modelling, 252(??):97–105, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001200436X.

Shcheglovitova:2013:EOC

[SA13] Mariya Shcheglovitova and Robert P. Anderson. Estimating optimal complexity for ecological niche models: a jackknife approach for species with small sample sizes. Ecological Modelling, 269(??):9–17, November 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004043.

Shimoda:2016:PFT

[SA16] Yuko Shimoda and George B. Arhonditsis. Phytoplankton functional type modelling: Running before we can walk? A critical evaluation of the current state of knowledge. Ecological Modelling, 320(??):29–43, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004196.

Sawyer:2016:GMP

[SAAD16] Jennifer M. Sawyer, Michael T. Arts, George Arhonditsis, and Miriam L. Diamond. A general model of polyunsaturated fatty acid (PUFA) uptake, loss and transformation in freshwater fish. Ecological Modelling, 323(??):96–105, March 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005566.

Serizawa:2010:EBT

[SAI10] Hiroshi Serizawa, Takashi Amemiya, and Kiminori Itoh. Effects of buoyancy, transparency and zooplankton feeding on surface maxima and deep maxima: Comprehensive mathematical model for vertical distribution in cyanobacterial biomass. Ecological Modelling, 221(17):2028–2037, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002619.

Sakanoue:2013:ILK

[Sak13] Seiichi Sakanoue. Integration of logistic and kinetics equations of population growth. Ecological Modelling, 261–262(??):93–97, July 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002111.

Soria-Auza:2010:IQC

[SAKB^{+}10] Rodrigo W. Soria-Auza, Michael Kessler, Kerstin Bach, Paola M. Barajas-Barbosa, Marcus Lehnert, Sebastian K. Herzog, and Jürgen Böhner. Impact of the quality of climate models for modelling species occurrences in countries with poor climatic documentation: a case study from Bolivia.

Ecological Modelling, 221(8):1221–1229, April 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000256.

Sanjerehei:2011:DAQ

[San11] Mohammad Mousaei Sanjerehei. Determination of an appropriate quadrat size and shape for detecting association between plant species. Ecological Modelling, 222 (10):1790–1792, May 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001141.

Sannigrahi:2017:MTE

[San17] Srikanta Sannigrahi. Modeling terrestrial ecosystem productivity of an estuarine ecosystem in the Sundarban Biosphere Region, India using seven ecosystem models. Ecological Modelling, 356(??):73–90, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306391.

Sanz:2019:CGE

[San19] Luis Sanz. Conditions for growth and extinction in matrix models with environmental stochasticity. Ecological Modelling, 411(??):Article 108797, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303059.

Schuwirth:2011:DMM

[SAR11] Nele Schuwirth, Vicenç Acuña, and Peter Reichert. Development of a mechanistic model (ERIMO-I) for analyzing the temporal dynamics of the benthic community of an intermittent Mediterranean stream. Ecological Modelling, 222(1):91–104, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004849.

Sainmont:2015:EAA

[SAT^{+}15] Julie Sainmont, Ken H. Andersen, Uffe H. Thygesen, Øyvind Fiksen, and André W. Visser. An effective algorithm for approximating adaptive behavior in seasonal environments. Ecological Modelling, 311(??):20–30, September 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001672.

Shiyomi:2011:GEM

[SAW^{+}11] Masae Shiyomi, Tsuyoshi Akiyama, Shiping Wang, Yiruhan, Ailikun, Yoshimichi Hori, Zuozhong Chen, Taisuke Yasuda, Kensuke Kawamura, and Yasuo Yamamura. A grassland ecosystem model of the Xilingol steppe, Inner Mongolia, China. Ecological Modelling, 222(13):2073–2083, July 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001657.

Salas:2011:EDI

[SB11] Andria K. Salas and Stuart R. Borrett. Evidence for the dominance of indirect effects in 50 trophic ecosystem networks. Ecological Modelling, 222(5):1192–1204, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010006514.

Sverdrup:2015:DAS

[SB15] Harald U. Sverdrup and Salim Belyazid. Developing an approach for Sweden, Switzerland, United States and France for setting critical loads based on biodiversity including management, pollution and climate change. Ecological Modelling, 306(??):35–45, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400444X.

Song:2013:TDS

[SBA^{+}13] Xiaodong Song, Brett A. Bryan, Auro C. Almeida, Keryn I. Paul, Gang Zhao, and Yin Ren. Time-dependent sensitivity of a process-based ecological model. Ecological Modelling, 265(??):114–123, September 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002974.

Sample:2019:QSS

[SBA^{+}19] Christine Sample, Joanna A. Bieri, Benjamin Allen, Yulia Dementieva, Alyssa Carson, Connor Higgins, Sadie Piatt, Shirley Qiu, Summer Stafford, Brady J. Mattsson, Darius J. Semmens, Wayne E. Thogmartin, and Jay E. Diffendorfer. Quantifying source and sink habitats and pathways in spatially structured populations: a generalized modelling approach. Ecological Modelling, 407(??):Article 108715, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302157.

Smith:2011:MAB

[SBC^{+}11] Amanda C. Smith, Jagtar S. Bhatti, Hua Chen, Mark E. Harmon, and Paul A. Arp. Modelling above- and below-ground mass loss and N dynamics in wooden dowels (LIDET) placed across North and Central America biomes at the decadal time scale. Ecological Modelling, 222(14):2276–2290, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004898.

Santos:2013:CCE

[SBC13] Mário Santos, Rita Bastos, and João Alexandre Cabral. Converting conventional ecological datasets in dynamic and dynamic spatially explicit simulations: Current advances and future applications of the Stochastic Dynamic Methodology (StDM). Ecological Modelling, 258(??):91–100, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001312.

Sokol:2015:LMB

[SBC^{+}15] Eric R. Sokol, Bryan L. Brown, Cayelan C. Carey, Brett M. Tornwall, Christopher M. Swan, and J. E. Barrett. Linking management to biodiversity in built ponds using metacommunity simulations. Ecological Modelling, 296(??):36–45, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004918.

Steenbeek:2016:PEP

[SBC^{+}16a] Jeroen Steenbeek, Joe Buszowski, Villy Christensen, Ekin Akoglu, Kerim Aydin, Nick Ellis, Dalai Felinto, Jerome Guitton, Sean Lucey, Kelly Kearney, Steven Mackinson, Mike Pan, Mark Platts, and Carl Walters. Ecopath with Ecosim as a model-building toolbox: Source code capabilities, extensions, and variations. Ecological Modelling, 319(??):178–189, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500280X.

Stella:2016:MSD

[SBC16b] Tommaso Stella, Simone Bregaglio, and Roberto Confalonieri. A model to simulate the dynamics of carbohydrate remobilization during rice grain filling. Ecological Modelling, 320(??):366–371, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005116.

Saha:2013:EAE

[SBCB13] Bapi Saha, Amiya Ranjan Bhowmick, Joydev Chattopadhyay, and Sabyasachi Bhattacharya. On the evidence of an Allee effect in herring populations and consequences for population survival: a model-based study. Ecological Modelling, 250(??):72–80, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005236.

Schuwirth:2019:HME

[SBD^{+}19] Nele Schuwirth, Florian Borgwardt, Sami Domisch, Martin Friedrichs, Mira Kattwinkel, David Kneis, Mathias Kuemmerlen, Simone D. Langhans, Javier Martínez-López, and Peter Vermeiren. How to make ecological models useful for environmental management. Ecological Modelling, 411(??):Article 108784, November 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302923.

Siegel:2016:CSE

[SBDA16] Eric Siegel, Mark T. Brown, Chris De Vilbiss, and Sam Arden. Calculating solar equivalence ratios of the four major heat-producing radiogenic isotopes in the Earth’s crust and mantle. Ecological Modelling, 339(??):140–147, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300989.

Schmitt:2013:MTT

[SBDS13] Walter Schmitt, Eric Bruns, Margit Dollinger, and Peter Sowig. Mechanistic TK/TD-model simulating the effect of growth inhibitors on Lemna populations. Ecological Modelling, 255(??):1–10, April 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000446.

Sheehan:2015:PMF

[SBF15] T. Sheehan, D. Bachelet, and K. Ferschweiler. Projected major fire and vegetation changes in the Pacific Northwest of the conterminous United States under selected CMIP5 climate futures. Ecological Modelling, 317(??):16–29, December 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003865.

Savage:2011:CMP

[SBM^{+}11] David Savage, Martin J. Barbetti, William J. MacLeod, Moin U. Salam, and Michael Renton. Can mechanistically parameterised, anisotropic dispersal kernels provide a reliable estimate of wind-assisted dispersal? Ecological Modelling, 222(10):1673–1682, May 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001104.

Saupe:2012:VND

[SBM^{+}12] E. E. Saupe, V. Barve, C. E. Myers, J. Soberón, N. Barve, C. M. Hensz, A. T. Peterson, H. L. Owens, and A. Lira-Noriega. Variation in niche and distribution model performance: the need for a priori assessment of key causal factors. Ecological Modelling, 237–238(??):11–22, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001676.

Song:2012:VBS

[SBPZ12] Xiaodong Song, Brett A. Bryan, Keryn I. Paul, and Gang Zhao. Variance-based sensitivity analysis of a forest growth model. Ecological Modelling, 247(??):135–143, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004024.

Schmolke:2019:SSP

[SBR^{+}19] Amelie Schmolke, Steven M. Bartell, Colleen Roy, Nicholas Green, Nika Galic, and Richard Brain. Species-specific population dynamics and their link to an aquatic food web: a hybrid modeling approach. Ecological Modelling, 405(??):1–14, August 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019301292.

Sierra:2010:MIT

[SBRD10] J. Sierra, N. Brisson, D. Ripoche, and M. Déqué. Modelling the impact of thermal adaptation of soil microorganisms and crop system on the dynamics of organic matter in a tropical soil under a climate change scenario. Ecological Modelling, 221(23):2850–2858, November 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004266.

Saint-Beat:2013:MFP

[SBVA^{+}13] B. Saint-Béat, A. F. Vézina, R. Asmus, H. Asmus, and N. Niquil. The mean function provides robustness to linear inverse modelling flow estimation in food webs: a comparison of functions derived from statistics and ecological theories. Ecological Modelling, 258(??):53–64, June 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000896.

Saitoh:2018:EVD

[SC18] Takashi Saitoh and Joel E. Cohen. Environmental variability and density dependence in the temporal Taylor’s law. Ecological Modelling, 387(??):134–143, November 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302588.

Sahragard:2021:EPH

[SC21] H. P. Sahragard and M. A. Zare Chahouki. An evaluation of predictive habitat models performance of plant species in Hoze soltan rangelands of Qom province. Ecological Modelling, 309–310(??):64–71, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001350.

Sadasivuni:2013:WRP

[SCB13] R. Sadasivuni, W. H. Cooke, and S. Bhushan. Wildfire risk prediction in Southeastern Mississippi using population interaction. Ecological Modelling, 251(??):297–306, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000070.

Strona:2018:IVN

[SCBH18] G. Strona, C. J. Carstens, P. S. A. Beck, and B. A. Han. The intrinsic vulnerability of networks to epidemics. Ecological Modelling, 383(??):91–97, September 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830173X.

Steenbeek:2013:BGB

[SCG^{+}13] Jeroen Steenbeek, Marta Coll, Leigh Gurney, Frédéric Mélin, Nicolas Hoepffner, Joe Buszowski, and Villy Christensen. Bridging the gap between ecosystem modeling tools and geographic information systems: Driving a food web model with external spatial–temporal data. Ecological Modelling, 263(??):139–151, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002597.

Sun:2016:ISO

[SCG^{+}16] Shaobo Sun, Baozhang Chen, Mengyu Ge, Junfeng Qu, Tao Che, Huifang Zhang, Xiaofeng Lin, Mingliang Che, Ziyuan Zhou, Lifeng Guo, and Bingyang Wang. Improving soil organic carbon parameterization of land surface model for cold regions in the Northeastern Tibetan Plateau, China. Ecological Modelling, 330(??):1–15, June 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300801.

Singh:2017:NPI

[SCH^{+}17] Shweta Singh, Jana E. Compton, Troy R. Hawkins, Daniel J. Sobota, and Ellen J. Cooter. A nitrogen physical inputoutput table (PIOT) model for Illinois. Ecological Modelling, 360(??):194–203, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302168.

Shi:2016:CTW

[SCHGM16] Pei-Jian Shi, Lei Chen, Cang Hui, and Henri D. Grissino-Mayer. Capture the time when plants reach their maximum body size by using the beta sigmoid growth equation. Ecological Modelling, 320(??):177–181, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004342.

Sanchez-Clavijo:2016:MEH

[SCHQA16] Lina M. Sánchez-Clavijo, Jessica Hearns, and Pedro F. Quintana-Ascencio. Modeling the effect of habitat selection mechanisms on population responses to landscape structure. Ecological Modelling, 328(??):99–107, May 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300631.

Sciubba:2011:RCE

[Sci11a] Enrico Sciubba. A revised calculation of the econometric factors α — and β for the Extended Exergy Accounting method. Ecological Modelling, 222(4):1060–1066, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005934.

Sciubba:2011:WDL

[Sci11b] Enrico Sciubba. What did Lotka really say? A critical reassessment of the “maximum power principle”. Ecological Modelling, 222(8):1347–1353, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000597.

Segoli:2013:ASO

[SDD^{+}13] M. Segoli, S. De Gryze, F. Dou, J. Lee, W. M. Post, K. Denef, and J. Six. AggModel: a soil organic matter model with measurable pools for use in incubation studies. Ecological Modelling, 263(??):1–9, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002147.

Simpkins:2018:APC

[SDEP18] Craig E. Simpkins, Todd E. Dennis, Thomas R. Etherington, and George L. W. Perry. Assessing the performance of common landscape connectivity metrics using a virtual ecologist approach. Ecological Modelling, 367(??):13–23, January 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001730443X.

Solbu:2018:HAS

[SDKE18] Erik B. Solbu, Ola H. Diserud, John A. Kålås, and Steinar Engen. Heterogeneity among species and community dynamics — Norwegian bird communities as a case study. Ecological Modelling, 388(??):13–23, November 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303041.

Strickland:2021:NSI

[SDS^{+}21] Christopher Strickland, Gerhard Dangelmayr, Patrick D. Shipman, Sunil Kumar, and Thomas J. Stohlgren. Network spread of invasive species and infectious diseases. Ecological Modelling, 309–310(??):1–9, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001611.

Spangenberg:2015:ELC

[SDSH15] J. H. Spangenberg, J. M. Douguet, J. Settele, and K. L. Heong. Escaping the lock-in of continuous insecticide spraying in rice: Developing an integrated ecological and socio-political DPSIR analysis. Ecological Modelling, 295 (??):188–195, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002592.

Sabatier:2010:MTO

[SDT10] R. Sabatier, L. Doyen, and M. Tichit. Modelling tradeoffs between livestock grazing and wader conservation in a grassland agroecosystem. Ecological Modelling, 221(9):1292–1300, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010000748.

Song:2015:ILU

[SDY^{+}15] Wei Song, Xiangzheng Deng, Yongwei Yuan, Zhan Wang, and Zhaohua Li. Impacts of land-use change on valued ecosystem service in rapidly urbanized North China plain. Ecological Modelling, 318(??):245–253, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015000484.

Solbu:2015:GWE

[SED15] Erik Blystad Solbu, Steinar Engen, and Ola Håvard Diserud. Guidelines when estimating temporal changes in density dependent populations. Ecological Modelling, 313(??):355–376, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002860.

Shi:2018:USD

[SEHD18] Yuning Shi, David M. Eissenstat, Yuting He, and Kenneth J. Davis. Using a spatially-distributed hydrologic biogeochemistry model with a nitrogen transport module to study the spatial variation of carbon processes in a critical zone observatory. Ecological Modelling, 380(??):8–21, July 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018301248.

Salacinska:2010:SAT

[SELB10] K. Salacinska, G. Y. El Serafy, F. J. Los, and A. Blauw. Sensitivity analysis of the two dimensional application of the generic ecological model (GEM) to algal bloom prediction in the North Sea. Ecological Modelling, 221(2):178–190, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009006358.

Samaranayaka:2010:MES

[SF10] Ari Samaranayaka and David Fletcher. Modelling environmental stochasticity in adult survival for a long-lived species. Ecological Modelling, 221(3):423–427, February 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006693.

Stoll:2015:AEI

[SFB^{+}15] Stefan Stoll, Mark Frenzel, Benjamin Burkhard, Mihai Adamescu, Algirdas Augustaitis, Cornelia Baeßler, Francisco J. Bonet, Maria Laura Carranza, Constantin Cazacu, Georgia L. Cosor, Ricardo Díaz-Delgado, Ulf Grandin, Peter Haase, Heikki Hämäläinen, Rob Loke, Jörg Müller, Angela Stanisci, Tomasz Staszewski, and Felix Müller. Assessment of ecosystem integrity and service gradients across Europe using the LTER Europe network. Ecological Modelling, 295(??):75–87, January 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003019.

Serpa:2012:MBP

[SFC^{+}12] Dalila Serpa, Pedro Pousão Ferreira, Miguel Caetano, Luís Cancela da Fonseca, Maria Teresa Dinis, and Pedro Duarte. Modelling of biogeochemical processes in fish earth ponds: Model development and calibration. Ecological Modelling, 247(??):286–301, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004322.

Seidl:2011:MND

[SFF^{+}11] Rupert Seidl, Paulo M. Fernandes, Teresa F. Fonseca, François Gillet, Anna Maria Jönsson, Katarína Merganičová, Sigrid Netherer, Alexander Arpaci, Jean-Daniel Bontemps, Harald Bugmann, Jose Ramon González-Olabarria, Petra Lasch, Céline Meredieu, Francisco Moreira, Mart-Jan Schelhaas, and Frits Mohren. Modelling natural disturbances in forest ecosystems: a review. Ecological Modelling, 222(4):903–924, February 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005272.

Soufbaf:2012:EPA

[SFHK12] Mahmoud Soufbaf, Yaghoub Fathipour, Cang Hui, and Javad Karimzadeh. Effects of plant availability and habitat size on the coexistence of two competing parasitoids in a tri-trophic food web of canola, diamondback moth and parasitic wasps. Ecological Modelling, 244(??):49–56, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012003237.

Sirakov:2019:SSH

[SFL^{+}19] Nikolay Sirakov, Bénédicte Fontez, Thérèse Libourel, Alessio dos Santos, Danielle Mitja, and Patrice Loisel. A stage-structured hierarchical Bayes model for the babassu palm tree population dynamics — estimated from anthropogenic open area data sets. Ecological Modelling, 400(??):14–26, May 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300857.

Sarker:2018:SSP

[SFM^{+}18] Subrata Sarker, Ulrike Feudel, Cédric L. Meunier, Peter Lemke, Partha S. Dutta, and Karen H. Wiltshire. To share or not to share? Phytoplankton species coexistence puzzle in a competition model incorporating multiple resource-limitation and synthesizing unit concepts. Ecological Modelling, 383(??):150–159, September 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301947.

Seiler:2015:EHI

[SFMA15] Lilian M. N. Seiler, Elisa Helena L. Fernandes, Flavio Martins, and Paulo Cesar Abreu. Evaluation of hydrologic influence on water quality variation in a coastal lagoon through numerical modeling. Ecological Modelling, 314(??):44–61, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003270.

Shi:2017:CTO

[SFR^{+}17] Pei-Jian Shi, Mei-Ling Fan, David A. Ratkowsky, Jian-Guo Huang, Hsin-I Wu, Lei Chen, Shui-Yuan Fang, and Chun-Xia Zhang. Comparison of two ontogenetic growth equations for animals and plants. Ecological Modelling, 349(??):1–10, April 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016306238.

Semenov:2010:PCS

[SFvB10] Alexander V. Semenov, Eelco Franz, and Ariena H. C. van Bruggen. COLIWAVE a simulation model for survival of E. coli O157:H7 in dairy manure and manure-amended soil. Ecological Modelling, 221(4):599–609, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007145.

Saraiva:2017:RBB

[SFvdM^{+}17] S. Saraiva, L. Fernandes, J. van der Meer, R. Neves, and S. A. L. M. Kooijman. The role of bivalves in the balgzand: First steps on an integrated modelling approach. Ecological Modelling, 359(??):34–48, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305701.

Soetaert:2011:EMB

[SG11a] Karline Soetaert and Marilaure Gregoire. Estimating marine biogeochemical rates of the carbonate pH system — a Kalman filter tested. Ecological Modelling, 222(12):1929–1942, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001359.

Sorman:2011:GBE

[SG11b] Alevgul H. Sorman and Mario Giampietro. Generating better energy indicators: Addressing the existence of multiple scales and multiple dimensions. Ecological Modelling, 223(1):41–53, December 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004996.

Sakiyama:2016:DAT

[SG16] Tomoko Sakiyama and Yukio-Pegio Gunji. Directional ambiguity in trail-laying algorithms. Ecological Modelling, 340(??):37–44, November 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001630374X.

Schwemmer:2016:MSS

[SGA^{+}16] Philipp Schwemmer, Franziska Güpner, Sven Adler, Knut Klingbeil, and Stefan Garthe. Modelling small-scale foraging habitat use in breeding Eurasian oystercatchers (Haematopus ostralegus) in relation to prey distribution and environmental predictors. Ecological Modelling, 320(??):322–333, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005086.

Salcido-Guevara:2014:EBC

[SGAS14] Luis A. Salcido-Guevara and Francisco Arreguín-Sánchez. Effects of biomass changes in the supply–demand balance of energy in aquatic food webs. Ecological Modelling, 276(??):64–79, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000428.

Speelman:2010:AVE

[SGB10] E. N. Speelman and L. E. García-Barrios. Agrodiversity v.2: an educational simulation tool to address some challenges for sustaining functional agrodiversity in agro-ecosystems. Ecological Modelling, 221(6):911–918, March 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008370.

Saqalli:2010:TIS

[SGBD10] M. Saqalli, B. Gérard, C. Bielders, and P. Defourny. Testing the impact of social forces on the evolution of Sahelian farming systems: a combined agent-based modeling and anthropological approach. Ecological Modelling, 221(22):2714–2727, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003996.

Stucchi:2019:RPN

[SGBG19] Luciano Stucchi, Luis Giménez-Benavides, and Javier Galeano. The role of parasitoids in a nursery-pollinator system: a population dynamics model. Ecological Modelling, 396(??):50–58, March 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019300250.

Shanin:2015:AK

[SGCB15] Vladimir Shanin, Pavel Grabarnik, Oleg Chertov, and Sergey Bykhovets. Alexander komarov (1945–2015). Ecological Modelling, 314(??):A1–A2, October 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003105.

Stricker:2019:MSH

[SGDP19] Heather K. Stricker, Thomas M. Gehring, Deahn Donner, and Tyler Petroelje. Multi-scale habitat selection model assessing potential gray wolf den habitat and dispersal corridors in Michigan, USA. Ecological Modelling, 397(??):84–94, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018304320.

Splawinski:2014:MPF

[SGG14] T. B. Splawinski, D. F. Greene, and S. Gauthier. A model of the post-fire recruitment of Picea mariana and Pinus banksiana as a function of salvage timing and intensity. Ecological Modelling, 282(??):35–43, June 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001318.

Speirs:2016:MEF

[SGH16] Douglas C. Speirs, Simon P. R. Greenstreet, and Michael R. Heath. Modelling the effects of fishing on the North Sea fish community size composition. Ecological Modelling, 321(??):35–45, February 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005177.

Strauss:2017:PHM

[SGHW^{+}17] Tido Strauss, Faten Gabsi, Monika Hammers-Wirtz, Pernille Thorbek, and Thomas G. Preuss. The power of hybrid modelling: an example from aquatic ecosystems. Ecological Modelling, 364(??):77–88, November 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300443.

Sutton-Grier:2010:ERB

[SGKR10] Ariana E. Sutton-Grier, Melissa A. Kenney, and Curtis J. Richardson. Examining the relationship between ecosystem structure and function using structural equation modelling: a case study examining denitrification potential in restored wetland soils. Ecological Modelling, 221(5):761–768, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009007972.

Silva:2014:SFB

[SGM^{+}14] Daniel P. Silva, Victor H. Gonzalez, Gabriel A. R. Melo, Mariano Lucia, Leopoldo J. Alvarez, and Paulo De Marco. Seeking the flowers for the bees: Integrating biotic interactions into niche models to assess the distribution of the exotic bee species Lithurgus huberi in South America. Ecological Modelling, 273(??):200–209, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005644.

Srivastava:2018:MIP

[SGP18] Vivek Srivastava, Verena C. Griess, and Hitendra Padalia. Mapping invasion potential using ensemble modelling. A case study on Yushania maling in the Darjeeling Himalayas. Ecological Modelling, 385(??):35–44, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302369.

Solana-Gutierrez:2017:UFC

[SGRAdJ17] Joaquín Solana-Gutiérrez, Gonzalo Rincón, Carlos Alonso, and Diego García de Jalón. Using fuzzy cognitive maps for predicting river management responses: a case study of the Esla River basin, Spain. Ecological Modelling, 360 (??):260–269, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017301060.

Srinivasan:2010:BGI

[SGSR10] Mridula Srinivasan, William E. Grant, Todd M. Swannack, and Jolly Rajan. Behavioral games involving a clever prey avoiding a clever predator: an individual-based model of dusky dolphins and killer whales. Ecological Modelling, 221(22):2687–2698, November 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003509.

Stenglein:2015:IBM

[SGWV15] Jennifer L. Stenglein, Jonathan H. Gilbert, Adrian P. Wydeven, and Timothy R. Van Deelen. An individual-based model for southern Lake Superior wolves: a tool to explore the effect of human-caused mortality on a landscape of risk. Ecological Modelling, 302(??):13–24, April 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000411.

Sellier:2014:PTT

[SH14] Damien Sellier and Jonathan J. Harrington. Phloem transport in trees: a generic surface model. Ecological Modelling, 290(??):102–109, October 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005693.

Su:2018:SAE

[SH18] Jia Su and Guangqiu Huang. Simulation and analysis of ecosystem vulnerability with cascading spread caused by dust migration based on object function GeoPetri net. Ecological Modelling, 379(??):54–72, July 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301224.

Scheller:2011:EFH

[SHB^{+}11a] Robert M. Scheller, Dong Hua, Paul V. Bolstad, Richard A. Birdsey, and David J. Mladenoff. The effects of forest harvest intensity in combination with wind disturbance on carbon dynamics in Lake States Mesic Forests. Ecological Modelling, 222(1):144–153, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004801.

Smajgl:2011:IED

[SHB11b] Alex Smajgl, Alan P. N. House, and James R. A. Butler. Implications of ecological data constraints for integrated policy and livelihoods modelling: an example from East Kalimantan, Indonesia. Ecological Modelling, 222(3):888–896, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010006204.

Sadler:2010:DST

[SHBG10] Rohan J. Sadler, Martin Hazelton, Matthias M. Boer, and Pauline F. Grierson. Deriving state-and-transition models from an image series of grassland pattern dynamics. Ecological Modelling, 221(3):433–444, February 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007133.

Schwalm:2015:HWD

[SHC^{+}15] C. R. Schwalm, D. N. Huntzinger, R. B. Cook, Y. Wei, I. T. Baker, R. P. Neilson, B. Poulter, Peter Caldwell, G. Sun, H. Q. Tian, and N. Zeng. How well do terrestrial biosphere models simulate coarse-scale runoff in the contiguous United States? Ecological Modelling, 303(??):87–96, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015000629.

Semeniuk:2010:LMA

[SHCR10] Christina A. D. Semeniuk, Wolfgang Haider, Andrew Cooper, and Kristina D. Rothley. A linked model of animal ecology and human behavior for the management of wildlife tourism. Ecological Modelling, 221(22):2699–2713, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001000373X.

Shaw:2014:ESE

[SHM^{+}14] C. H. Shaw, A. B. Hilger, J. Metsaranta, W. A. Kurz, G. Russo, F. Eichel, G. Stinson, C. Smyth, and M. Filiatrault. Evaluation of simulated estimates of forest ecosystem carbon stocks using ground plot data from Canada’s National Forest Inventory. Ecological Modelling, 272(??):323–347, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004699.

Snortheim:2017:MDH

[SHM^{+}17] Craig A. Snortheim, Paul C. Hanson, Katherine D. McMahon, Jordan S. Read, Cayelan C. Carey, and Hilary A. Dugan. Meteorological drivers of hypolimnetic anoxia in a eutrophic, north temperate lake. Ecological Modelling, 343(??):39–53, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001630583X.

Stokland:2011:SDM

[SHS11] Jogeir N. Stokland, Rune Halvorsen, and Bente Støa. Species distribution modelling — effect of design and sample size of pseudo-absence observations. Ecological Modelling, 222(11):1800–1809, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001062.

Shen:2012:MPB

[SHS^{+}12] Jian Shen, Bo Hong, Leonard Schugam, Yuan Zhao, and Jeff White. Modeling of polychlorinated biphenyls (PCBs) in the Baltimore Harbor. Ecological Modelling, 242(??):54–68, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002554.

Sawano:2021:DSF

[SHT^{+}21] Shinji Sawano, Norifumi Hotta, Nobuaki Tanaka, Yoshio Tsuboyama, and Masakazu Suzuki. Development of a simple forest evapotranspiration model using a processoriented model as a reference to parameterize data from a wide range of environmental conditions. Ecological Modelling, 309–310(??):93–109, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001623.

Shang:2012:ILM

[SHX^{+}12] ZongBo Shang, Hong S. He, Weimin Xi, Stephen R. Shifley, and Brian J. Palik. Integrating LANDIS model and a multicriteria decision-making approach to evaluate cumulative effects of forest management in the Missouri ozarks, USA. Ecological Modelling, 229(??):50–63, March 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004315.

Sillero:2011:WDE

[Sil11] Neftalí Sillero. What does ecological modelling model? A proposed classification of ecological niche models based on their underlying methods. Ecological Modelling, 222 (8):1343–1346, April 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011000500.

Simoy:2011:SIM

[Sim11] María Verónica Simoy. A simple individual model for simulating weight gain dynamics in response to intake rate and daily behavior. Ecological Modelling, 222(3):727–736, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010005910.

Sinha:2017:TES

[Sin17] Sumit Sinha. Transient evolution of suspended and benthic algae in a riverine ecosystem: a numerical study. Ecological Modelling, 348(??):78–92, March 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307219.

Shi:2014:EBE

[SIS^{+}14] Pei-Jian Shi, Tetsuroh Ishikawa, Hardev S. Sandhu, Cang Hui, Amit Chakraborty, Xian-Shi Jin, Katsunori Tachihara, and Bai-Lian Li. On the 3/4-exponent von Bertalanffy equation for ontogenetic growth. Ecological Modelling, 276(??):23–28, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000052.

Singer:2016:CDU

[SJB^{+}16] Alexander Singer, Karin Johst, Thomas Banitz, Mike S. Fowler, Jürgen Groeneveld, Alvaro G. Gutiérrez, Florian Hartig, Rainer M. Krug, Matthias Liess, Glenn Matlack, Katrin M. Meyer, Guy Pe’er, Viktoriia Radchuk, Ana-Johanna Voinopol-Sassu, and Justin M. J. Travis. Community dynamics under environmental change: How can next generation mechanistic models improve projections of species distributions? Ecological Modelling, 326(??):63–74, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005281.

Scherer:2016:MTB

[SJGB16] Cédric Scherer, Florian Jeltsch, Volker Grimm, and Niels Blaum. Merging trait-based and individual-based modelling: an animal functional type approach to explore the responses of birds to climatic and land use changes in semi-arid African savannas. Ecological Modelling, 326(??):75–89, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003117.

Straskraba:2014:EED

[SJP14] Milan Straškraba, Sven E. Jørgensen, and Bernard C. Patten. Ecosystems emerging: 6. Differentiation. Ecological Modelling, 278(??):29–51, April 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000350.

Schramski:2013:EPS

[SJS^{+}13] J. R. Schramski, K. L. Jacobsen, T. W. Smith, M. A. Williams, and T. M. Thompson. Energy as a potential systems-level indicator of sustainability in organic agriculture: Case study model of a diversified, organic vegetable production system. Ecological Modelling, 267 (??):102–114, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013003773.

Safuan:2013:ESN

[SJTS13] Hamizah M. Safuan, Zlatko Jovanoski, Isaac N. Towers, and Harvinder S. Sidhu. Exact solution of a nonautonomous logistic population model. Ecological Modelling, 251(??):99–102, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005960.

Schmickl:2010:ISR

[SK10] Thomas Schmickl and Istvan Karsai. The interplay of sex ratio, male success and density-independent mortality affects population dynamics. Ecological Modelling, 221(8):1089–1097, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010000219.

Singh:2011:RLR

[SK11] Aditya Singh and S. P. S. Kushwaha. Refining logistic regression models for wildlife habitat suitability modeling — a case study with muntjak and goral in the Central Himalayas, India. Ecological Modelling, 222(8):1354–1366, April 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000925.

Sajna:2014:MSF

[ŠK14] Nina Šajna and Primož Kušar. Modeling species fitness in competitive environments. Ecological Modelling, 275(??):31–36, March 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013005863.

Srinivasan:2015:EDR

[SK15] Venkatraman Srinivasan and Praveen Kumar. Emergent and divergent resilience behavior in catastrophic shift systems. Ecological Modelling, 298(??):87–105, February 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005814.

Skene:2013:EES

[Ske13] Keith R. Skene. The energetics of ecological succession: a logistic model of entropic output. Ecological Modelling, 250(??):287–293, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005583.

Sharifi:2013:CDE

[SKH^{+}13] Amirreza Sharifi, Latif Kalin, Mohamed M. Hantush, Sabahattin Isik, and Thomas E. Jordan. Carbon dynamics and export from flooded wetlands: a modeling approach. Ecological Modelling, 263(??):196–210, August 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002470.

Scheller:2019:LMV

[SKHH19] Robert Scheller, Alec Kretchun, Todd J. Hawbaker, and Paul D. Henne. A landscape model of variable socialecological fire regimes. Ecological Modelling, 401(??):85–93, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019301279.

Shanin:2013:CTM

[SKK^{+}13] Vladimir Shanin, Alexander Komarov, Yulia Khoraskina, Sergey Bykhovets, Tapio Linkosalo, and Raisa Mäkipää. Carbon turnover in mixed stands: Modelling possible shifts under climate change. Ecological Modelling, 251 (??):232–245, February 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005959.

Song:2013:SSD

[SKL^{+}13] Wonkyong Song, Eunyoung Kim, Dongkun Lee, Moungjin Lee, and Seong-Woo Jeon. The sensitivity of species distribution modeling to scale differences. Ecological Modelling, 248(??):113–118, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004772.

Stabler:2016:CEM

[SKM^{+}16] Moritz Stäbler, Alexander Kempf, Steven Mackinson, Jan Jaap Poos, Clement Garcia, and Axel Temming. Combining efforts to make maximum sustainable yields and good environmental status match in a food-web model of the southern North Sea. Ecological Modelling, 331(??):17–30, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300084.

Shanin:2011:MCN

[SKMB11] Vladimir N. Shanin, Alexander S. Komarov, Alexey V. Mikhailov, and Sergei S. Bykhovets. Modelling carbon and nitrogen dynamics in forest ecosystems of Central Russia under different climate change scenarios and forest management regimes. Ecological Modelling, 222(14):2262–2275, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010006149.

Strauss:2016:SLC

[SKPHW16] Tido Strauss, Devdutt Kulkarni, Thomas G. Preuss, and Monika Hammers-Wirtz. The secret lives of cannibals: Modelling density-dependent processes that regulate population dynamics in Chaoborus crystallinus. Ecological Modelling, 321(??):84–97, February 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005219.

Smyth:2011:IEW

[SKTG11] C. E. Smyth, W. A. Kurz, J. A. Trofymow, and C. I. D. E. T. Working Group. Including the effects of water stress on decomposition in the carbon budget model of the Canadian forest sector CBM-CFS3. Ecological Modelling, 222(5):1080–1091, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000654X.

Schlaepfer:2014:MRR

[SLB14] Daniel R. Schlaepfer, William K. Lauenroth, and John B. Bradford. Modeling regeneration responses of big sagebrush (Artemisia tridentata) to abiotic conditions. Ecological Modelling, 286(??):66–77, August 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002117.

Shuai:2018:SCL

[SLB^{+}18] Fangmin Shuai, Sovan Lek, Christophe Baehr, Young-Seuk Park, Yuefei Li, and Xinhui Li. Silver carp larva abundance in response to river flow rate revealed by cross-wavelet modelling. Ecological Modelling, 383(??):98–105, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018301935.

Szewczyk:2019:LMR

[SLD^{+}19] Tim M. Szewczyk, Tom Lee, Mark J. Ducey, Matthew E. Aiello-Lammens, Hayley Bibaud, and Jenica M. Allen. Local management in a regional context: Simulations with process-based species distribution models. Ecological Modelling, 413(??):Article 108827, December 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303357.

Siehoff:2011:PBM

[SLH^{+}11] Silvana Siehoff, Gottfried Lennartz, Ira C. Heilburg, Martina Roß-Nickoll, Hans Toni Ratte, and Thomas G. Preuss. Process-based modeling of grassland dynamics built on ecological indicator values for land use. Ecological Modelling, 222(23–24):3854–3868, December 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004881.

Simons-Legaard:2015:PAB

[SLLW15] Erin Simons-Legaard, Kasey Legaard, and Aaron Weiskittel. Predicting aboveground biomass with LANDIS-II: a global and temporal analysis of parameter sensitivity. Ecological Modelling, 313(??):325–332, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002823.

Smithwick:2014:IRR

[SLMS14] Erica A. H. Smithwick, Melissa S. Lucash, M. Luke Mc-Cormack, and Gajan Sivandran. Improving the representation of roots in terrestrial models. Ecological Modelling, 291(??):193–204, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003603.

Slone:2011:IAD

[Slo11] D. H. Slone. Increasing accuracy of dispersal kernels in grid-based population models. Ecological Modelling, 222(3):573–579, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010006289.

Sun:2019:DEM

[SLRC19] Ming Sun, Yunzhou Li, Yiping Ren, and Yong Chen. Developing and evaluating a management strategy evaluation framework for the Gulf of Maine cod (Gadus morhua). Ecological Modelling, 404(??):27–35, July 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301449.

Sallaba:2015:RNM

[SLSS15] Florian Sallaba, Dörte Lehsten, Jonathan Seaquist, and Martin T. Sykes. A rapid NPP meta-model for current and future climate and CO_{2} scenarios in Europe. Ecological Modelling, 302(??):29–41, April 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000459.

Shan:2014:MES

[SLW^{+}14] Kun Shan, Lin Li, Xiaoxiao Wang, Yanlong Wu, Lili Hu, Gongliang Yu, and Lirong Song. Modelling ecosystem structure and trophic interactions in a typical cyanobacterial bloom-dominated shallow Lake Dianchi, China. Ecological Modelling, 291(??):82–95, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003512.

Sagarese:2017:PTN

[SLW17] Skyler R. Sagarese, Matthew V. Lauretta, and John F. Walter. Progress towards a next-generation fisheries ecosystem model for the northern Gulf of Mexico. Ecological Modelling, 345(??):75–98, February 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306512.

Shi:2017:ERH

[SLY^{+}17] Xuan Shi, Jingling Liu, Xiaoguang You, Kun Bao, Bo Meng, and Bin Chen. Evaluation of river habitat integrity based on benthic macroinvertebrate-based multi-metric model. Ecological Modelling, 353(??):63–76, June 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302435.

Song:2013:EDS

[SLZ^{+}13] Qinghai Song, Hua Lin, Yiping Zhang, Zhenghong Tan, Junfu Zhao, Junbin Zhao, Xiang Zhang, Wenjun Zhou, Lei Yu, Lianyan Yang, Guirui Yu, and Xiaomin Sun. The effect of drought stress on self-organisation in a seasonal tropical rainforest. Ecological Modelling, 265(??):136–139, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013002949.

Sun:2017:IBB

[SLZ^{+}17] Qingling Sun, Baolin Li, Tao Zhang, Yecheng Yuan, Xizhang Gao, Jinsong Ge, Fei Li, and Zhijun Zhang. An improved Biome-BGC model for estimating net primary productivity of alpine meadow on the Qinghai-Tibet Plateau. Ecological Modelling, 350(??):55–68, April 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300972.

Sheng:2018:ECC

[SLZ^{+}18] Meiling Sheng, Junzhi Liu, A-Xing Zhu, David G. Rossiter, Liming Zhu, and Guoqiang Peng. Evaluation of CLM-Crop for maize growth simulation over Northeast China. Ecological Modelling, 377(??):26–34, June 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300899.

Schimit:2010:WSW

[SM10] P. H. T. Schimit and L. H. A. Monteiro. Who should wear mask against airborne infections? Altering the contact network for controlling the spread of contagious diseases. Ecological Modelling, 221(9):1329–1332, May 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000797.

Schimit:2011:VGB

[SM11] P. H. T. Schimit and L. H. A. Monteiro. A vaccination game based on public health actions and personal decisions. Ecological Modelling, 222(9):1651–1655, May 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001001.

Schimit:2012:EBR

[SM12] P. H. T. Schimit and L. H. A. Monteiro. On estimating the basic reproduction number in distinct stages of a contagious disease spreading. Ecological Modelling, 240(??):156–160, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012002104.

Sebastian:2013:DVP

[SM13] Carlos Ruiz Sebastián and Timothy R. McClanahan. Description and validation of production processes in the coral reef ecosystem model CAFFEE (coral–algae–fish-fisheries ecosystem energetics) with a fisheries closure and climatic disturbance. Ecological Modelling, 263(??):326–348, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002743.

Sun:2014:ANP

[SM14] Guodong Sun and Mu Mu. The analyses of the net primary production due to regional and seasonal temperature differences in eastern China using the LPJ model. Ecological Modelling, 289(??):66–76, October 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003032.

Smaldino:2013:MIU

[Sma13] Paul E. Smaldino. Measures of individual uncertainty for ecological models: Variance and entropy. Ecological Modelling, 254(??):50–53, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000422.

Suarez-Munoz:2019:PIA

[SMBGH^{+}19] María Suárez-Muñoz, Francisco Bonet-García, José A. Hódar, Javier Herrero, Mihai Tanase, and Lucía Torres-Muros. INSTAR: an Agent-Based Model that integrates existing knowledge to simulate the population dynamics of a forest pest. Ecological Modelling, 411(??): Article 108764, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302728.

Sabetraftar:2011:SMG

[SMC11] K. Sabetraftar, B. Mackey, and B. Croke. Sensitivity of modelled gross primary productivity to topographic effects on surface radiation: a case study in the Cotter River Catchment, Australia. Ecological Modelling, 222 (3):795–803, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005211.

Soussana:2012:GGM

[SMG^{+}12] Jean-François Soussana, Vincent Maire, Nicolas Gross, Bruno Bachelet, Loic Pagès, Raphaël Martin, David Hill, and Christian Wirth. Gemini: a grassland model simulating the role of plant traits for community dynamics and ecosystem functioning. Parameterization and evaluation. Ecological Modelling, 231(??):134–145, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000579.

Semeniuk:2012:IBE

[SMH^{+}12] C. A. D. Semeniuk, M. Musiani, M. Hebblewhite, S. Grindal, and D. J. Marceau. Incorporating behavioral–ecological strategies in pattern-oriented modeling of caribou habitat use in a highly industrialized landscape. Ecological Modelling, 243(??):18–32, September 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002761.

Simioni:2016:IVS

[SMH16] Guillaume Simioni, Guillaume Marie, and Roland Huc. Influence of vegetation spatial structure on growth and water fluxes of a mixed forest: Results from the NOTG 3D model. Ecological Modelling, 328(??):119–135, May 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300205.

Schratz:2019:HTP

[SMI^{+}19] Patrick Schratz, Jannes Muenchow, Eugenia Iturritxa, Jakob Richter, and Alexander Brenning. Hyperparameter tuning and performance assessment of statistical and machine-learning algorithms using spatial data. Ecological Modelling, 406(??):109–120, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302145.

Somerville:2019:MAG

[SMKM19] Gayle. J. Somerville, Bo Melander, Per Kudsk, and Solvejg K. Mathiassen. Modelling annual grass weed seed dispersal in winter wheat, when influenced by hedges and directional wind. Ecological Modelling, 410(??): Article 108729, October 15, 2019. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302297.

Sigourney:2012:CBN

[SML12] Douglas B. Sigourney, Stephan B. Munch, and Benjamin H. Letcher. Combining a Bayesian nonparametric method with a hierarchical framework to estimate individual and temporal variation in growth. Ecological Modelling, 247(??):125–134, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004206.

Stolpovsky:2011:IDD

[SMLH^{+}11] Konstantin Stolpovsky, Paula Martinez-Lavanchy, Hermann J. Heipieper, Philippe Van Cappellen, and Martin Thullner. Incorporating dormancy in dynamic microbial community models. Ecological Modelling, 222(17): 3092–3102, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003760.

Shi:2013:GOG

[SMS^{+}13] Pei-Jian Shi, Xing-Yuan Men, Hardev S. Sandhu, Amit Chakraborty, Bai-Lian Li, Fang Ou-Yang, Yu-Cheng Sun, and Feng Ge. The “general” ontogenetic growth model is inapplicable to crop growth. Ecological Modelling, 266(??):1–9, September 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013003207.

Sainte-Marie:2014:NPC

[SMSAN^{+}14] J. Sainte-Marie, L. Saint-André, Y. Nouvellon, J.-P. Laclau, O. Roupsard, G. le Maire, N. Delpierre, A. Henrot, and M. Barrandon. A new probabilistic canopy dynamics model (SLCD) that is suitable for evergreen and deciduous forest ecosystems. Ecological Modelling, 290(??):121–133, October 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000696.

Schwamborn:2019:ASU

[SMT19a] R. Schwamborn, T. K. Mildenberger, and M. H. Taylor. Assessing sources of uncertainty in length-based estimates of body growth in populations of fishes and macroinvertebrates with bootstrapped ELEFAN. Ecological Modelling, 393(??):37–51, February 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304125.

Schwamborn:2019:CCA

[SMT19b] R. Schwamborn, T. K. Mildenberger, and M. H. Taylor. Corrigendum: Corrigendum to “Assessing sources of uncertainty in length-based estimates of body growth in populations of fishes and macroinvertebrates with bootstrapped ELEFAN” [Ecol. Model. 393 (2019) 37–51]. Ecological Modelling, 394(??):76–77, February 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300031.

Sasai:2012:ASV

[SNS^{+}12] Takahiro Sasai, Saori Nakai, Yuko Setoyama, Keisuke Ono, Soushi Kato, Masayoshi Mano, Kazutaka Murakami, Akira Miyata, Nobuko Saigusa, Ramakrishna R. Nemani, and Kenlo N. Nasahara. Analysis of the spatial variation in the net ecosystem production of rice paddy fields using the diagnostic biosphere model, BEAMS. Ecological Modelling, 247(??):175–189, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004279.

Sujii:2019:GAS

[SNZ^{+}19] Patricia S. Sujii, Micael E. Nagai, Maria I. Zucchi, Pedro H. S. Brancalion, and Patrick M. A. James. A genetic approach for simulating persistence of reintroduced tree species populations in restored forests. Ecological Modelling, 403(??):35–43, July 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301590.

Seghetta:2014:EAU

[SØB14] Michele Seghetta, Hanne Østergård, and Simone Bastianoni. Energy analysis of using macroalgae from eutrophic waters as a bioethanol feedstock. Ecological Modelling, 288(??):25–37, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002439.

Solovjova:2019:ERM

[Sol19] N. V. Solovjova. Ecological risk modelling in developing resources of ecosystems characterized by varying vulnerability levels. Ecological Modelling, 406(??):60–72, August 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001930211X.

Stankowski:2010:SDM

[SP10] Philippe A. Stankowski and William H. Parker. Species distribution modelling: Does one size fit all? A phytogeographic analysis of Salix in Ontario. Ecological Modelling, 221(13–14):1655–1664, July 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001614.

Stankowski:2011:FDM

[SP11] Philippe A. Stankowski and William H. Parker. Future distribution modelling: a stitch in time is not enough. Ecological Modelling, 222(3):567–572, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005788.

Su:2012:PBS

[SP12] Zhenming Su and Randall M. Peterman. Performance of a Bayesian state-space model of semelparous species for stock-recruitment data subject to measurement error. Ecological Modelling, 224(1):76–89, January 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005199.

Sor:2017:ESP

[SPB^{+}17] Ratha Sor, Young-Seuk Park, Pieter Boets, Peter L. M. Goethals, and Sovan Lek. Effects of species prevalence on the performance of predictive models. Ecological Modelling, 354(??):11–19, June 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304367.

Stratton:2011:IDW

[SPG11] Tana Stratton, David T. Price, and Konrad Gajewski. Impacts of daily weather variability on simulations of the Canadian boreal forest. Ecological Modelling, 222(17):3250–3260, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003425.

Sankar:2018:SSO

[SPM^{+}18] S. Sankar, L. Polimene, L. Marin, N. N. Menon, A. Samuelsen, R. Pastres, and S. Ciavatta. Sensitivity of the simulated oxygen minimum zone to biogeochemical processes at an oligotrophic site in the Arabian Sea. Ecological Modelling, 372(??):12–23, March 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018300395.

Su:2018:RMT

[SPP18] Bei Su, Markus Pahlow, and A. E. Friederike Prowe. The role of microzooplankton trophic interactions in modelling a suite of mesocosm ecosystems. Ecological Modelling, 368(??):169–179, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304684.

Schwaiger:2018:GRA

[SPR^{+}18] Fabian Schwaiger, Werner Poschenrieder, Thomas Rötzer, Peter Biber, and Hans Pretzsch. Groundwater recharge algorithm for forest management models. Ecological Modelling, 385(??):154–164, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302412.

Somerville:2017:HDS

[SPWR17] Gayle J. Somerville, Stephen B. Powles, Michael J. Walsh, and Michael Renton. How do spatial heterogeneity and dispersal in weed population models affect predictions of herbicide resistance evolution? Ecological Modelling, 362(??):37–53, October 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303551.

Shen:2019:DDM

[SQWS19] Jian Shen, Qubin Qin, Ya Wang, and Mac Sisson. A datadriven modeling approach for simulating algal blooms in the tidal freshwater of James River in response to riverine nutrient loading. Ecological Modelling, 398(??):44–54, April 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300663.

Savage:2014:RDI

[SR14] David Savage and Michael Renton. Requirements, design and implementation of a general model of biological invasion. Ecological Modelling, 272(??):394–409, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004638.

Sah:2011:MSF

[SRHL11] Leena Sah, Diederik P. L. Rousseau, Christine M. Hooijmans, and Piet N. L. Lens. 3D model for a secondary facultative pond. Ecological Modelling, 222(9):1592–1603, May 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011001025.

Stanaway:2011:HBM

[SRM11] M. A. Stanaway, R. Reeves, and K. L. Mengersen. Hierarchical Bayesian modelling of plant pest invasions with human-mediated dispersal. Ecological Modelling, 222 (19):3531–3540, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004303.

Schulze:2018:IFL

[SRP18] Kiowa Alraune Schulze, Gert Rosenthal, and Alexander Peringer. Intermediate foraging large herbivores maintain semi-open habitats in wilderness landscape simulations. Ecological Modelling, 379(??):10–21, July 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301133.

Seidl:2012:IBP

[SRSS12] Rupert Seidl, Werner Rammer, Robert M. Scheller, and Thomas A. Spies. An individual-based process model to simulate landscape-scale forest ecosystem dynamics. Ecological Modelling, 231(??):87–100, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000919.

Strohm:2016:IMM

[SRT16] S. Strohm, M. L. Reid, and R. C. Tyson. Impacts of management on mountain pine beetle spread and damage: a process-rich model. Ecological Modelling, 337 (??):241–252, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016302514.

Smeti:2018:SES

[SRTS18] Evangelia Smeti, Daniel L. Roelke, George Tsirtsis, and Sofie Spatharis. Species extinctions strengthen the relationship between biodiversity and resource use efficiency. Ecological Modelling, 384(??):75–86, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302102.

Silva:2012:SCP

[SRV^{+}12] Brenner Silva, Kristin Roos, Ingo Voss, Nicolas König, Rütger Rollenbeck, Renate Scheibe, Erwin Beck, and Jörg Bendix. Simulating canopy photosynthesis for two competing species of an anthropogenic grassland community in the Andes of southern Ecuador. Ecological Modelling, 239(??):14–26, July 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000361.

Shan:2014:EOW

[SRXP14] Nan Shan, Xiao-Hong Ruan, Jing Xu, and Zha-Rong Pan. Estimating the optimal width of buffer strip for nonpoint source pollution control in the Three Gorges Reservoir Area, China. Ecological Modelling, 276(??):51–63, March 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000040.

Schizas:2010:BIC

[SS10] Dimitrios Schizas and George Stamou. Beyond identity crisis: the challenge of recontextualizing ecosystem delimitation. Ecological Modelling, 221(12):1630–1635, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010001389.

Stewart:2011:CBB

[SS11] Thomas. J. Stewart and W. Gary Sprules. Carbonbased balanced trophic structure and flows in the offshore Lake Ontario food web before (1987–1991) and after (2001–2005) invasion-induced ecosystem change. Ecological Modelling, 222(3):692–708, February 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005843.

Scheiter:2016:EMC

[SS16] Simon Scheiter and Patrice Savadogo. Ecosystem management can mitigate vegetation shifts induced by climate change in West Africa. Ecological Modelling, 332(??):19–27, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300953.

Salau:2012:VEC

[SSBJ12] Kehinde Salau, Michael L. Schoon, Jacopo A. Baggio, and Marco A. Janssen. Varying effects of connectivity and dispersal on interacting species dynamics. Ecological Modelling, 242(??):81–91, September 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002311.

Simoy:2015:ETP

[SSC15] M. I. Simoy, M. V. Simoy, and G. A. Canziani. The effect of temperature on the population dynamics of Aedes aegypti. Ecological Modelling, 314(??):100–110, October 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003130.

Shariati:2017:ESM

[SSD^{+}17] Mitra Shariati, Andrew K. Skidmore, Roshanak Darvishzadeh, Klaus-Michael Exo, Andrea Kölzsch, Larry Griffin, Julia Stahl, David Cabot, and Albertus G. Toxopeus. Expert system for modelling stopover site selection by barnacle geese. Ecological Modelling, 359(??):398–405, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016305439.

Small:2014:ENA

[SSF14] Gaston E. Small, Robert W. Sterner, and Jacques C. Finlay. An Ecological Network Analysis of nitrogen cycling in the Laurentian Great Lakes. Ecological Modelling, 293(??):150–160, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000854.

Shuman:2014:TIB

[SSK14] Jacquelyn K. Shuman, Herman H. Shugart, and Olga N. Krankina. Testing individual-based models of forest dynamics: Issues and an example from the boreal forests of Russia. Ecological Modelling, 293(??):102–110, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005279.

Singer:2018:CHS

[SSKJ18] Alexander Singer, Oliver Schweiger, Ingolf Kühn, and Karin Johst. Constructing a hybrid species distribution model from standard large-scale distribution data. Ecological Modelling, 373(??):39–52, April 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300437.

Stojkovic:2013:VFC

[SSM^{+}13] Milica Stojkovic, Vladica Simic, Djuradj Milosevic, Dejan Mancev, and Tadeusz Penczak. Visualization of fish community distribution patterns using the self-organizing map: a case study of the Great Morava River system (Serbia). Ecological Modelling, 248(??):20–29, January 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004802.

Sohma:2018:MCE

[SSN^{+}18] Akio Sohma, Hisashi Shibuki, Fumiyuki Nakajima, Atsushi Kubo, and Tomohiro Kuwae. Modeling a coastal ecosystem to estimate climate change mitigation and a model demonstration in Tokyo Bay. Ecological Modelling, 384(??):261–289, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830142X.

Silva:2010:EIO

[SSQ^{+}10] João Paulo Silva, Mário Santos, Luís Queirós, Domingos Leitão, Francisco Moreira, Márcia Pinto, Miguel Leqoc, and João Alexandre Cabral. Estimating the influence of overhead transmission power lines and landscape context on the density of little bustard Tetrax tetrax breeding populations. Ecological Modelling, 221(16):1954–1963, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010001857.

Shi:2016:DDD

[SSR16] Pei-Jian Shi, Hardev S. Sandhu, and Gadi V. P. Reddy. Dispersal distance determines the exponent of the spatial Taylor’s power law. Ecological Modelling, 335 (??):48–53, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301788.

Schippers:2011:CBC

[SSS^{+}11] Peter Schippers, Eric W. M. Stienen, Alex G. M. Schotman, Robbert P. H. Snep, and Pieter A. Slim. The consequences of being colonial: Allee effects in metapopulations of seabirds. Ecological Modelling, 222(17):3061–3070, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003024.

Schaubroeck:2012:IEN

[SSV^{+}12] Thomas Schaubroeck, Jeroen Staelens, Kris Verheyen, Bart Muys, and Jo Dewulf. Improved ecological network analysis for environmental sustainability assessment; a case study on a forest ecosystem. Ecological Modelling, 247(??):144–156, December 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004292.

Sun:2018:SAM

[SSW^{+}18] Zaijin Sun, Yuxian Shangguan, Yuan Wei, Benying Su, Changzhi Zhou, and Hong Hou. A study on antimony migration in soils using an artificial neural network model and a convection-dispersion diffusion model. Ecological Modelling, 389(??):1–10, December 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303223.

Shi:2014:STE

[SSY14] Yusheng Shi, Takahiro Sasai, and Yasushi Yamaguchi. Spatio-temporal evaluation of carbon emissions from biomass burning in Southeast Asia during the period 2001– 2010. Ecological Modelling, 272(??):98–115, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004596.

Schapaugh:2013:APU

[ST13] Adam W. Schapaugh and Andrew J. Tyre. Accounting for parametric uncertainty in Markov decision processes. Ecological Modelling, 254(??):15–21, April 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000306.

Simidjievski:2015:LEP

[STD15] Nikola Simidjievski, Ljupčo Todorovski, and Sašo Džeroski. Learning ensembles of population dynamics models and their application to modelling aquatic ecosystems. Ecological Modelling, 306(??):305–317, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003986.

Synodinos:2015:FDM

[STJ15] Alexis D. Synodinos, Britta Tietjen, and Florian Jeltsch. Facilitation in drylands: Modeling a neglected driver of savanna dynamics. Ecological Modelling, 304(??):11–21, May 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001500071X.

Stoddard:2010:CVB

[Sto10] Steven T. Stoddard. Continuous versus binary representations of landscape heterogeneity in spatially-explicit models of mobile populations. Ecological Modelling, 221 (19):2409–2414, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010003121.

Strona:2015:CSE

[Str15a] Giovanni Strona. Corrigendum to “A spatially explicit model to investigate how dispersal/colonization tradeoffs between short and long distance movement strategies affect species ranges” [Ecol. Model. 297 (2015) 80–85].

Ecological Modelling, 301(??):16, April 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000216.

Strona:2015:SEM

[Str15b] Giovanni Strona. A spatially explicit model to investigate how dispersal/colonization tradeoffs between short and long distance movement strategies affect species ranges. Ecological Modelling, 297(??):80–85, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005845.

Sapoukhina:2010:SMC

[STSA10] Natalia Sapoukhina, Yuri Tyutyunov, Ivan Sache, and Roger Arditi. Spatially mixed crops to control the stratified dispersal of airborne fungal diseases. Ecological Modelling, 221(23):2793–2800, November 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004151.

Stancampiano:2015:RDD

[SU15] L. Stancampiano and F. Usai. The role of densitydependent arrested larval stages on parasite dynamics and stability: Lessons from nematodes and donkeys. Ecological Modelling, 297(??):69–79, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005596.

Strauch:2013:SPG

[SV13] Michael Strauch and Martin Volk. SWAT plant growth modification for improved modeling of perennial vegetation in the tropics. Ecological Modelling, 269(??):98–112, November 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013004134.

Sakschewski:2014:FBP

[SvBH^{+}14] Boris Sakschewski, Werner von Bloh, Veronika Huber, Christoph Müller, and Alberte Bondeau. Feeding 10 billion people under climate change: How large is the production gap of current agricultural systems? Ecological Modelling, 288(??):103–111, September 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002683.

Sailley:2013:CFW

[SVD^{+}13] S. F. Sailley, M. Vogt, S. C. Doney, M. N. Aita, L. Bopp, E. T. Buitenhuis, T. Hashioka, I. Lima, C. Le Quéré, and Y. Yamanaka. Comparing food web structures and dynamics across a suite of global marine ecosystem models. Ecological Modelling, 261–262(??):43–57, July 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300210X.

Saraiva:2011:MFP

[SvdMKS11] S. Saraiva, J. van der Meer, S. A. L. M. Kooijman, and T. Sousa. Modelling feeding processes in bivalves: a mechanistic approach. Ecological Modelling, 222(3):514–523, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005181.

Stewart:2010:VDB

[SvdR10] Christopher William Stewart and Rodney van der Ree. A Voronoi diagram based population model for social species of wildlife. Ecological Modelling, 221(12):1554–1568, June 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001000164X.

Shearman:2019:MPF

[SVH^{+}19] Timothy M. Shearman, J. Morgan Varner, Sharon M. Hood, C. Alina Cansler, and J. Kevin Hiers. Modelling post-fire tree mortality: Can random forest improve discrimination of imbalanced data? Ecological Modelling, 414(??):Article 108855, December 15, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303631.

Shibu:2012:LTD

[SVL12] M. E. Shibu, H. Van Keulen, and P. A. Leffelaar. Longterm dynamics of soil C and N in intensive rice-based cropping systems of the Indo-gangetic plains (IGP): a modelling approach. Ecological Modelling, 232(??):40–63, May 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012000968.

Schindler:2015:PMS

[SvWP^{+}15] Stefan Schindler, Henrik von Wehrden, Kostas Poirazidis, Wesley M. Hochachka, Thomas Wrbka, and Vassiliki Kati. Performance of methods to select landscape metrics for modelling species richness. Ecological Modelling, 295(??):107–112, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002610.

Schellinck:2011:RAR

[SW11] Jen Schellinck and Tony White. A review of attraction and repulsion models of aggregation: Methods, findings and a discussion of model validation. Ecological Modelling, 222(11):1897–1911, June 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001360.

Sun:2015:ALH

[SW15] Yajun Sun and Mathew G. Wells. The application of life-history and predation allometry to population dynamics to predict the critical density of extinction. Ecological Modelling, 312(??):136–149, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001878.

Scharfe:2019:MIA

[SW19] Mirco Scharfe and Karen Helen Wiltshire. Modeling of intra-annual abundance distributions: Constancy and variation in the phenology of marine phytoplankton species over five decades at Helgoland Roads (North Sea). Ecological Modelling, 404(??):46–60, July 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300018.

Shao:2013:EBE

[SWC13] Ling Shao, Zi Wu, and G. Q. Chen. Exergy based ecological footprint accounting for China. Ecological Modelling, 252(??):83–96, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004401.

Stillman:2016:DSP

[SWGC16] Richard A. Stillman, Kevin A. Wood, and John D. Goss-Custard. Deriving simple predictions from complex models to support environmental decision-making. Ecological Modelling, 326(??):134–141, April 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001659.

Sohl:2016:DPF

[SWR^{+}16] Terry L. Sohl, Michael C. Wimberly, Volker C. Radeloff, David M. Theobald, and Benjamin M. Sleeter. Divergent projections of future land use in the United States arising from different models and scenarios. Ecological Modelling, 337(??):281–297, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302605.

Schmitt:2016:ISM

[SWS16] Melissa H. Schmitt, David Ward, and Adrian M. Shrader. Incorporating secondary metabolites, tannin-binding proteins, and diet breadth into carrying-capacity models for African elephants. Ecological Modelling, 332(??):8–18, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016300898.

Shang:2012:ISB

[SWZ^{+}12] Xiao Shang, Xinze Wang, Dalei Zhang, Weidong Chen, Xuechu Chen, and Hainan Kong. An improved SWAT-based computational framework for identifying critical source areas for agricultural pollution at the lake basin scale. Ecological Modelling, 226(??):1–10, February 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005801.

Shao:2013:EEA

[SWZ^{+}13] Ling Shao, Zi Wu, L. Zeng, Z. M. Chen, Y. Zhou, and G. Q. Chen. Embodied energy assessment for ecological wastewater treatment by a constructed wetland. Ecological Modelling, 252(??):63–71, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004693.

Shi:2011:ISH

[SX11] Hu Shi and Mo Xingguo. Interpreting spatial heterogeneity of crop yield with a process model and remote sensing. Ecological Modelling, 222(14):2530–2541, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006162.

Shen:2018:SAS

[SXNL18] Yang Shen, Zhichao Xu, Ivan Nijs, and Jinbao Liao. Spatial arrangement of size-different patches determines population dynamics in linear riverine systems. Ecological Modelling, 385(??):220–225, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001830262X.

Sun:2015:NIS

[SXZ^{+}15] Xiang Sun, Shu Xiong, Xiaojing Zhu, Xiaodong Zhu, Yangfan Li, and B. Larry Li. A new indices system for evaluating ecological-economic-social performances of wetland restorations and its application to Taihu Lake Basin, China. Ecological Modelling, 295(??):216–226, January 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004773.

Shi:2013:SPC

[SYL^{+}13] Mingjie Shi, Zong-Liang Yang, David M. Lawrence, Robert E. Dickinson, and Zachary M. Subin. Spin-up processes in the community land model version 4 with explicit carbon and nitrogen components. Ecological Modelling, 263(??):308–325, August 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013002123.

Su:2019:SFG

[SYY^{+}19] Kai Su, Qiang Yu, Depeng Yue, Qibin Zhang, Lan Yang, Zhili Liu, Teng Niu, and Xiaoting Sun. Simulation of a forest-grass ecological network in a typical desert oasis based on multiple scenes. Ecological Modelling, 413(??):Article 108834, December 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019303424.

Sciubba:2014:EBA

[SZ14] Enrico Sciubba and Federico Zullo. An exergy-based analysis of the co-evolution of different species sharing common resources. Ecological Modelling, 273(??):277–283, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001300478X.

Shen:2019:SPS

[SZNL19] Yang Shen, Chenghui Zeng, Ivan Nijs, and Jinbao Liao. Species persistence in spatially regular networks. Ecological Modelling, 406(??):1–6, August 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301826.

Sadeghi:2012:ACT

[SZSV12a] Roghayeh Sadeghi, Rahmat Zarkami, Karim Sabetraftar, and Patrick Van Damme. Application of classification trees to model the distribution pattern of a new exotic species Azolla filiculoides (Lam.) at Selkeh Wildlife Refuge, Anzali wetland, Iran. Ecological Modelling, 243(??):8–17, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012002839.

Sadeghi:2012:USV

[SZSV12b] Roghayeh Sadeghi, Rahmat Zarkami, Karim Sabetraftar, and Patrick Van Damme. Use of support vector machines (SVMs) to predict distribution of an invasive water fern Azolla filiculoides (Lam.) in Anzali wetland, southern Caspian Sea, Iran. Ecological Modelling, 244 (??):117–126, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003195.

Sadeghi:2013:AGA

[SZSV13] Roghayeh Sadeghi, Rahmat Zarkami, Karim Sabetraftar, and Patrick Van Damme. Application of genetic algorithm and greedy stepwise to select input variables in classification tree models for the prediction of habitat requirements of Azolla filiculoides (Lam.) in Anzali wetland, Iran. Ecological Modelling, 251(??):44–53, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005765.

Sadeghi:2014:MHP

[SZV14] Roghayeh Sadeghi, Rahmat Zarkami, and Patrick Van Damme. Modelling habitat preference of an alien aquatic fern, Azolla filiculoides (Lam.), in Anzali wetland (Iran) using data-driven methods. Ecological Modelling, 284(??):1–9, July 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014001847.

Sieber:2010:MBS

[SZVV10] Stefan Sieber, Peter Zander, Peter H. Verburg, and Martin Van Ittersum. Model-based systems to support impact assessment — methods, tools and applications. Ecological Modelling, 221(18):2133–2135, September 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003297.

Tzanakakis:2011:CEU

[TA11] V. A. Tzanakakis and A. N. Angelakis. Chemical exergy as a unified and objective indicator in the assessment and optimization of land treatment systems. Ecological Modelling, 222(17):3082–3091, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100367X.

Teresa:2014:MTA

[TAC^{+}14] Carone Maria Teresa, Guisan Antoine, Cianfrani Carmen, Simoniello Tiziana, Loy Anna, and Carranza Maria Laura. A multi-temporal approach to model endangered species distribution in Europe. The case of the Eurasian otter in Italy. Ecological Modelling, 274(??):21–28, February 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005759.

Turner:2016:RMD

[TAGC^{+}16] Katrine Grace Turner, Sharolyn Anderson, Mauricio Gonzales-Chang, Robert Costanza, Sasha Courville, Tommy Dalgaard, Estelle Dominati, Ida Kubiszewski, Sue Ogilvy, Luciana Porfirio, Nazmun Ratna, Harpinder Sandhu, Paul C. Sutton, Jens-Christian Svenning, Graham Mark Turner, Yann-David Varennes, Alexey Voinov, and Stephen Wratten. A review of methods, data, and models to assess changes in the value of ecosystem services from land degradation and restoration. Ecological Modelling, 319(??):190–207, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003233.

Torres-Alruiz:2013:TDP

[TAR13] Maria Daniela Torres-Alruiz and Diego J. Rodríguez. A topo-dynamical perspective to evaluate indirect interactions in trophic webs: New indexes. Ecological Modelling, 250(??):363–369, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005625.

Tatsumi:2016:EAM

[Tat16a] Kenichi Tatsumi. Effects of automatic multi-objective optimization of crop models on corn yield reproducibility in the U.S.A. Ecological Modelling, 322(??):124–137, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001500527X.

Tayyebi:2016:GBC

[TAT^{+}16b] Amin Tayyebi, Jamal J. Arsanjani, Amir H. Tayyebi, Hichem Omrani, and Hossein S. Moghadam. Group-based crop change planning: Application of SmartScape^{T M} spatial decision support system for resolving conflicts. Ecological Modelling, 333(??):92–100, August 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301466.

Tang:2011:PAB

[TB11] Wenwu Tang and David A. Bennett. Parallel agentbased modeling of spatial opinion diffusion accelerated using graphics processing units. Ecological Modelling, 222 (19):3605–3615, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004339.

Tang:2012:RPA

[TB12a] Wenwu Tang and David A. Bennett. Reprint of: Parallel agent-based modeling of spatial opinion diffusion accelerated using graphics processing units. Ecological Modelling, 229(??):108–118, March 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000580.

Taylor:2012:BRD

[TB12b] Rebecca L. Taylor and Gina K. Himes Boor. Beyond the robust design: Accounting for changing, uncertain states and sparse, biased detection in a multistate mark-recapture model. Ecological Modelling, 243(??):73–80, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002852.

Tamburino:2013:MWU

[TB13] Lucia Tamburino and Giangiacomo Bravo. Mice in wonderforest: Understanding mast seeding through individualbased modelling. Ecological Modelling, 250(??):34–44, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380012005121.

Tzemi:2019:RGG

[TB19] Domna Tzemi and James Breen. Reducing greenhouse gas emissions through the use of urease inhibitors: a farm level analysis. Ecological Modelling, 394(??):18–26, February 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018304344.

Thampi:2019:SEM

[TBA19] Vivek A. Thampi, Chris T. Bauch, and Madhur Anand. Socio-ecological mechanisms for persistence of native Australian grasses under pressure from nitrogen runoff and invasive species. Ecological Modelling, 413(??): Article 108830, December 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019303382.

Thebaud:2015:SOC

[TBJ^{+}15] Olivier Thébaud, Fabio Boschetti, Sarah Jennings, Anthony D. M. Smith, and Sean Pascoe. Of sets of offsets: Cumulative impacts and strategies for compensatory restoration. Ecological Modelling, 312(??):114–124, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001738.

Tron:2015:CDR

[TBL^{+}15] Stefania Tron, Gernot Bodner, Francesco Laio, Luca Ridolfi, and Daniel Leitner. Can diversity in root architecture explain plant water use efficiency? A modeling study. Ecological Modelling, 312(??):200–210, September 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002343.

Turner:2014:CPF

[TBM^{+}14] Evan L. Turner, Denise A. Bruesewitz, Rae F. Mooney, Paul A. Montagna, James W. McClelland, Alexey Sadovski, and Edward J. Buskey. Comparing performance of five nutrient phytoplankton zooplankton (NPZ) models in coastal lagoons. Ecological Modelling, 277(??):13–26, April 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000416.

Thompson:2016:MIP

[TCGC16] Robin N. Thompson, Richard C. Cobb, Christopher A. Gilligan, and Nik J. Cunniffe. Management of invading pathogens should be informed by epidemiology rather than administrative boundaries. Ecological Modelling, 324(??):28–32, March 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005827.

Torres:2013:FWS

[TCH^{+}13] María Ángeles Torres, Marta Coll, Johanna Jacomina Heymans, Villy Christensen, and Ignacio Sobrino. Foodweb structure of and fishing impacts on the Gulf of Cadiz ecosystem (South-western Spain). Ecological Modelling, 265(??):26–44, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013002822.

Taugourdeau:2015:CRI

[TCSG15] Olivier Taugourdeau, Yves Caraglio, Sylvie Sabatier, and Yann Guédon. Characterizing the respective importance of ontogeny and environmental constraints in forest tree development using growth phase duration distributions. Ecological Modelling, 300(??):61–72, March 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006358.

Taner:2011:IMP

[TCW11] Mehmet Ümit Taner, James N. Carleton, and Marjorie Wellman. Integrated model projections of climate change impacts on a North American lake. Ecological Modelling, 222(18):3380–3393, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001100398X.

Teixeira:2011:SOM

[TDC^{+}11] R. F. M. Teixeira, T. Domingos, A. P. S. V. Costa, R. Oliveira, L. Farropas, F. Calouro, A. M. Barradas, and J. P. B. G. Carneiro. Soil organic matter dynamics in Portuguese natural and sown rainfed grasslands. Ecological Modelling, 222(4):993–1001, February 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006186.

Trapp:2019:MIL

[TDF^{+}19] Stephanie E. Trapp, Casey C. Day, Elizabeth A. Flaherty, Patrick A. Zollner, and Winston P. Smith. Modeling impacts of landscape connectivity on dispersal movements of northern flying squirrels (Glaucomys sabrinus griseifrons). Ecological Modelling, 394(??):44–52, February 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304368.

Thinh:2012:EDL

[TDH12] Vu T. Thinh, Paul F. Doherty, and Kathryn P. Huyvaert. Effects of different logging schemes on bird communities in tropical forests: a simulation study. Ecological Modelling, 243(??):95–100, September 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002773.

Terra:2012:MCE

[TdS12] Talita Nogueira Terra and Rozely Ferreira dos Santos. Measuring cumulative effects in a fragmented landscape. Ecological Modelling, 228(??):89–95, March 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000026.

Tennenbaum:2015:ECE

[Ten15a] Stephen E. Tennenbaum. Emergy and co-emergy. Ecological Modelling, 315(??):116–134, November 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400427X.

Tennenbaum:2015:OTB

[Ten15b] Stephen E. Tennenbaum. Odum–Tennenbaum–Brown calculus vs emergy and co-emergy analysis: a reply. Ecological Modelling, 313(??):333–340, October 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003099. See [LTL15].

Thornley:2013:UGF

[TF13] John H. M. Thornley and James France. Use of growth functions to describe disease vector population dynamics — additional assumptions are required and are important. Ecological Modelling, 266(??):97–102, September 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003232.

Taubert:2012:RGM

[TFH12] Franziska Taubert, Karin Frank, and Andreas Huth. A review of grassland models in the biofuel context. Ecological Modelling, 245(??):84–93, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001731.

Tsang:2011:RIM

[TFMP11] Yin-Phan Tsang, Gary K. Felton, Glenn E. Moglen, and Michael Paul. Region of influence method improves macroinvertebrate predictive models in Maryland. Ecological Modelling, 222(19):3473–3485, October 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004236.

Tuzet:2017:MHF

[TGB^{+}17] Andrée Tuzet, André Granier, Pauline Betsch, Marianne Peiffer, and Alain Perrier. Modelling hydraulic functioning of an adult beech stand under non-limiting soil water and severe drought condition. Ecological Modelling, 348(??):56–77, March 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305117.

Teckentrup:2018:CCF

[TGKSJ18] Lisa Teckentrup, Volker Grimm, Stephanie Kramer-Schadt, and Florian Jeltsch. Community consequences of foraging under fear. Ecological Modelling, 383(??):80–90, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018301753.

Tanaka:2021:PNS

[TGM^{+}21] Yasuaki Tanaka, Andréa G. Grottoli, Yohei Matsui, Atsushi Suzuki, and Kazuhiko Sakai. Partitioning of nitrogen sources to algal endosymbionts of corals with long-term ^{15}N-labelling and a mixing model. Ecological Modelling, 309–310(??):163–169, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001684.

Tonnang:2017:ACI

[THBF^{+}17] Henri E. Z. Tonnang, Bisseleua D. B. Hervé, Lisa Biber-Freudenberger, Daisy Salifu, Sevgan Subramanian, Valentine B. Ngowi, Ritter Y. A. Guimapi, Bruce Anani, Francois M. M. Kakmeni, Hippolyte Affognon, Saliou Niassy, Tobias Landmann, Frank T. Ndjomatchoua, Sansao A. Pedro, Tino Johansson, Chrysantus M. Tanga, Paulin Nana, Komi M. Fiaboe, Samira F. Mohamed, Nguya K. Maniania, Lev V. Nedorezov, Sunday Ekesi, and Christian Borgemeister. Advances in crop insect modelling methods — towards a whole system approach. Ecological Modelling, 354(??):88–103, June 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001630549X.

Tai:2018:CMP

[THC18] Travis C. Tai, Christopher D. G. Harley, and William W. L. Cheung. Comparing model parameterizations of the biophysical impacts of ocean acidification to identify limitations and uncertainties. Ecological Modelling, 385(??):1–11, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302424.

Tautenhahn:2012:BDI

[THJ^{+}12] Susanne Tautenhahn, Hermann Heilmeier, Martin Jung, Anja Kahl, Jens Kattge, Antje Moffat, and Christian Wirth. Beyond distance-invariant survival in inverse recruitment modeling: a case study in Siberian Pinus sylvestris forests. Ecological Modelling, 233(??):90–103, May 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001160.

Thogmartin:2010:SAN

[Tho10a] Wayne E. Thogmartin. Sensitivity analysis of North American bird population estimates. Ecological Modelling, 221(2):173–177, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006334.

Topping:2010:OBB

[THO10b] Chris J. Topping, Toke T. Høye, and Carsten Riis Olesen. Opening the black box — development, testing and documentation of a mechanistically rich agent-based model. Ecological Modelling, 221(2):245–255, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009006346.

Topping:2010:POM

[THOA10] Christopher John Topping, Toke Thomas Høye, Peter Odderskær, and Nicholas J. Aebischer. A pattern-oriented modelling approach to simulating populations of grey partridge. Ecological Modelling, 221(5):729–737, March 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007509.

Trobajo:2014:MRM

[TIC^{+}14] Rosa Trobajo, Carles Ibañez, Ester Clavero, Joan Salvadó, and Sven Erik Jørgensen. Modelling the response of microalgae to CO_{2} addition. Ecological Modelling, 294(??):42–50, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014003998.

Tiezzi:2011:ETE

[Tie11] E. Tiezzi. Ecodynamics: Towards an evolutionary thermodynamics of ecosystems. Ecological Modelling, 222(16):2897–2902, August 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011002948.

Tietjen:2016:SRA

[Tie16] Britta Tietjen. Same rainfall amount different vegetation — how environmental conditions and their interactions influence savanna dynamics. Ecological Modelling, 326(??):13–22, April 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002628.

Tilley:2014:EOD

[Til14] David R. Tilley. Exploration of Odum’s dynamic emergy accounting rules for suggested refinements. Ecological Modelling, 279(??):36–44, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000908.

Tilley:2015:TDR

[Til15] David Tilley. Transformity dynamics related to maximum power for improved emergy yield estimations. Ecological Modelling, 315(??):96–107, November 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005535.

Tilley:2021:DAE

[Til21] David R. Tilley. Dynamic accounting of emergy cycling. Ecological Modelling, 222(20–22):3734–3742, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004662.

Tian:2011:UGS

[TJYZ11] Guangjin Tian, Jing Jiang, Zhifeng Yang, and Yaoqi Zhang. The urban growth, size distribution and spatio-temporal dynamic pattern of the Yangtze River Delta megalopolitan region, China. Ecological Modelling, 222(3):865–878, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010005235.

Thurnher:2017:MTG

[TKH17] Christopher Thurnher, Mario Klopf, and Hubert Hasenauer. MOSES — a tree growth simulator for modelling stand response in Central Europe. Ecological Modelling, 352(??):58–76, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016302903.

Trakhtenbrot:2014:MMS

[TKN14] A. Trakhtenbrot, G. G. Katul, and R. Nathan. Mechanistic modeling of seed dispersal by wind over hilly terrain. Ecological Modelling, 274(??):29–40, February 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005772.

Tahvonen:2014:OHA

[TKP14] Olli Tahvonen, Jouko Kumpula, and Antti-Juhani Pekkarinen. Optimal harvesting of an age-structured, twosex herbivore–plant system. Ecological Modelling, 272 (??):348–361, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004730.

Taherzadeh:2017:CWP

[TKW17] Niousha Taherzadeh, Onur Kerimoglu, and Kai W. Wirtz. Can we predict phytoplankton community size structure using size scalings of eco-physiological traits? Ecological Modelling, 360(??):279–289, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303393.

Telesca:2010:ATS

[TL10] Luciano Telesca and Rosa Lasaponara. Analysis of timescaling properties in forest-fire sequence observed in Italy. Ecological Modelling, 221(1):90–93, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009000696.

Todd:2015:WDY

[TL15] Charles R. Todd and Mark Lintermans. Who do you move? A stochastic population model to guide translocation strategies for an endangered freshwater fish in south-eastern Australia. Ecological Modelling, 311(??):63–72, September 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001970.

Tao:2015:VES

[TLL^{+}15] Yu Tao, Feng Li, Xusheng Liu, Dan Zhao, Xiao Sun, and Lianfang Xu. Variation in ecosystem services across an urbanization gradient: a study of terrestrial carbon stocks from Changzhou, China. Ecological Modelling, 318(??):210–216, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001994.

Tang:2017:IBN

[TLL^{+}17] P. Z. Tang, J. Z. Liu, H. W. Lu, Z. Wang, and L. He. Information-based network environ analysis for ecological risk assessment of heavy metals in soils. Ecological Modelling, 344(??):17–28, January 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305567.

Tan:2019:SPM

[TLLX19] Jianbo Tan, Ainong Li, Guangbin Lei, and Xinyao Xie. A SD-MaxEnt-CA model for simulating the landscape dynamic of natural ecosystem by considering socio-economic and natural impacts. Ecological Modelling, 410(??): Article 108783, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019302911.

Takashina:2017:EMR

[TLP17] Nao Takashina, Joung-Hun Lee, and Hugh P. Possingham. Effect of marine reserve establishment on noncooperative fisheries management. Ecological Modelling, 360 (??):336–342, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303368.

Tuomi:2011:WDM

[TLRL11] M. Tuomi, R. Laiho, A. Repo, and J. Liski. Wood decomposition model for boreal forests. Ecological Modelling, 222(3):709–718, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005855.

Todd:2016:ARE

[TLS^{+}16] Charles R. Todd, David B. Lindenmayer, Kasey Stamation, Silvana Acevedo-Cattaneo, Stephen Smith, and Linda F. Lumsden. Assessing reserve effectiveness: Application to a threatened species in a dynamic fire prone forest landscape. Ecological Modelling, 338(??):90–100, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302642.

Thomas:2014:NMG

[TLW^{+}14] Christopher J. Thomas, Jonathan Lambrechts, Eric Wolanski, Vincent A. Traag, Vincent D. Blondel, Eric Deleersnijder, and Emmanuel Hanert. Numerical modelling and graph theory tools to study ecological connectivity in the Great Barrier Reef. Ecological Modelling, 272 (??):160–174, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001300464X.

Tian:2017:OAN

[TLZ17] Wenchong Tian, Zhenliang Liao, and Jin Zhang. An optimization of artificial neural network model for predicting chlorophyll dynamics. Ecological Modelling, 364 (??):42–52, November 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017300777.

Tanaka:2012:FTH

[TM12] Yoshinari Tanaka and Hiroyuki Mano. Functional traits of herbivores and food chain efficiency in a simple aquatic community model. Ecological Modelling, 237–238(??):88–100, July 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012001949.

Taub:2014:ODA

[TM14] Frieda B. Taub and Anna K. McLaskey. Oxygen dynamics of aquatic closed ecological systems: Comparing the whole to a subsystem. Ecological Modelling, 293(??):49–57, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001400132X.

Tanioka:2018:EIA

[TM18] Tatsuro Tanioka and Katsumi Matsumoto. Effects of incorporating age-specific traits of zooplankton into a marine ecosystem model. Ecological Modelling, 368 (??):257–264, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017304866.

Tenan:2016:HMD

[TMC^{+}16] S. Tenan, C. Maffioletti, M. Caccianiga, C. Compostella, R. Seppi, and M. Gobbi. Hierarchical models for describing space-for-time variations in insect population size and sex-ratio along a primary succession. Ecological Modelling, 329(??):18–28, June 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300229.

Tardy:2018:IBC

[TMPF18] Olivia Tardy, Ariane Massé, Fanie Pelletier, and Daniel Fortin. Interplay between contact risk, conspecific density, and landscape connectivity: an individual-based modeling framework. Ecological Modelling, 373(??):25–38, April 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300504.

Tranmer:2018:HMF

[TMT^{+}18] Andrew W. Tranmer, Clelia L. Marti, Daniele Tonina, Rohan Benjankar, Dana Weigel, Leticia Vilhena, Claire Mc-Grath, Peter Goodwin, Matthew Tiedemann, Jim Mckean, and Jörg Imberger. A hierarchical modelling framework for assessing physical and biochemical characteristics of a regulated river. Ecological Modelling, 368(??):78–93, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304659.

Thiele:2017:CCI

[TNA^{+}17] Jan C. Thiele, Robert S. Nuske, Bernd Ahrends, Oleg Panferov, Matthias Albert, Kai Staupendahl, Udo Junghans, Martin Jansen, and Joachim Saborowski. Climate change impact assessment — a simulation experiment with Norway spruce for a forest district in Central Europe. Ecological Modelling, 346(??):30–47, February 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307153.

Tomczak:2012:EFD

[TNHB12] M. T. Tomczak, S. Niiranen, O. Hjerne, and T. Blenckner. Ecosystem flow dynamics in the Baltic proper — using a multi-trophic dataset as a basis for food–web modelling. Ecological Modelling, 230(??):123–147, April 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006028.

Townsend:2011:DDP

[TNTH11] Sunny E. Townsend, Scott Newey, Simon J. Thirgood, and Daniel T. Haydon. Dissecting the drivers of population cycles: Interactions between parasites and mountain hare demography. Ecological Modelling, 222(1):48–56, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004461.

Tenan:2014:BMS

[TOHT14] Simone Tenan, Robert B. O’Hara, Iris Hendriks, and Giacomo Tavecchia. Bayesian model selection: the steepest mountain to climb. Ecological Modelling, 283(??):62–69, July 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014001690.

Tian:2011:SSD

[TOQW11] Guangjin Tian, Yun Ouyang, Quan Quan, and Jianguo Wu. Simulating spatiotemporal dynamics of urbanization with multi-agent systems — a case study of the Phoenix metropolitan region, USA. Ecological Modelling, 222(5):1129–1138, March 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006848.

Tempel:2014:UIP

[TPG14] Douglas J. Tempel, M. Z. Peery, and R. J. Gutiérrez. Using integrated population models to improve conservation monitoring: California spotted owls as a case study. Ecological Modelling, 289(??):86–95, October 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003329.

Taffi:2015:BMS

[TPL^{+}15] Marianna Taffi, Nicola Paoletti, Pietro Liò, Sandra Pucciarelli, and Mauro Marini. Bioaccumulation modelling and sensitivity analysis for discovering key players in contaminated food webs: the case study of PCBs in the Adriatic Sea. Ecological Modelling, 306(??):205–215, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006036.

Tsakalakis:2018:DLC

[TPO^{+}18] Ioannis Tsakalakis, Markus Pahlow, Andreas Oschlies, Bernd Blasius, and Alexey B. Ryabov. Diel light cycle as a key factor for modelling phytoplankton biogeography and diversity. Ecological Modelling, 384(??):241–248, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302278.

Tsagaraki:2011:BCE

[TPT^{+}11] Tatiana M. Tsagaraki, George Petihakis, Kostas Tsiaras, George Triantafyllou, Manolis Tsapakis, Gerasimos Korres, George Kakagiannis, Constantin Frangoulis, and Ioannis Karakassis. Beyond the cage: Ecosystem modelling for impact evaluation in aquaculture. Ecological Modelling, 222(14):2512–2523, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006368.

Tian:2012:IRB

[TQZ12] Guangjin Tian, Zhi Qiao, and Yaoqi Zhang. The investigation of relationship between rural settlement density, size, spatial distribution and its geophysical parameters of China using Landsat TM images. Ecological Modelling, 231(??):25–36, April 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200052X.

Tecchio:2015:MHS

[TRD^{+}15] Samuele Tecchio, Armonie Tous Rius, Jean-Claude Dauvin, Jérémy Lobry, Géraldine Lassalle, Jocelyne Morin, Nicolas Bacq, Marie Cachera, Aurélie Chaalali, Maria Ching Villanueva, and Nathalie Niquil. The mosaic of habitats of the Seine estuary: Insights from food-web modelling and network analysis. Ecological Modelling, 312 (??):91–101, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001500232X.

Tipping:2012:NPS

[TRE^{+}12] E. Tipping, E. C. Rowe, C. D. Evans, R. T. E. Mills, B. A. Emmett, J. S. Chaplow, and J. R. Hall. N14C: a plant–soil nitrogen and carbon cycling model to simulate terrestrial ecosystem responses to atmospheric nitrogen deposition. Ecological Modelling, 247(??):11–26, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003997.

Tattoni:2012:CLD

[TRP12] Clara Tattoni, Franco Rizzolli, and Paolo Pedrini. Can Li-DAR data improve bird habitat suitability models? Ecological Modelling, 245(??):103–110, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001366.

Tsirtsis:2011:SSN

[TS11] George Tsirtsis and Sofie Spatharis. Simulating the structure of natural phytoplankton assemblages: Descriptive vs. mechanistic models. Ecological Modelling, 222 (12):1922–1928, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011001669.

Tashkova:2012:PEN

[TŠAD12] Katerina Tashkova, Jurij Šilc, Nataša Atanasova, and Sašo Džeroski. Parameter estimation in a nonlinear dynamic model of an aquatic ecosystem with meta-heuristic optimization. Ecological Modelling, 226(??):36–61, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005795.

Terry:2018:CMW

[TSB^{+}18] Julie A. Terry, Amir Sadeghian, Helen M. Baulch, Steven C. Chapra, and Karl-Erich Lindenschmidt. Challenges of modelling water quality in a shallow prairie lake with seasonal ice cover. Ecological Modelling, 384(??):43–52, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018302011.

Thierry:2015:LVM

[TSM^{+}15] Hugo Thierry, David Sheeren, Nicolas Marilleau, Nathalie Corson, Marion Amalric, and Claude Monteil. From the Lotka–Volterra model to a spatialised populationdriven individual-based model. Ecological Modelling, 306(??):287–293, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014004463.

Tabeta:2015:CMF

[TSN15] Shigeru Tabeta, Shota Suzuki, and Yoshiharu Nakamura. Combined modeling of fish behavior and fishing operations for conger eel fishery in Ise Bay. Ecological Modelling, 313(??):266–275, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002999.

Tabeta:2017:AFM

[TSN17] Shigeru Tabeta, Shota Suzuki, and Kenta Nakamura. Assessment of fishery management by using a fishery simulator for bottom otter trawling in Ise Bay. Ecological Modelling, 358(??):40–49, August 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308559.

Takahashi:2019:CPD

[TSP19] Daisuke Takahashi, Seung-Woo Sim, and Eun-Jin Park. Colonial population dynamics of Spartina alterniflora. Ecological Modelling, 395(??):45–50, March 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300274.

Treydte:2011:MLP

[TSSL11] A. C. Treydte, A. Salvatierra, J. Sauerborn, and M. Lamers. Modelling llama population development under environmental and market constraints in the Bolivian highlands. Ecological Modelling, 222(17):3157–3165, September 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003255.

Tepley:2012:AAS

[TT12] Alan J. Tepley and Enrique A. Thomann. Analytical approximation of a stochastic, spatial simulation model of fire and forest landscape dynamics. Ecological Modelling, 233(??):41–51, May 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001123.

Tjorve:2017:PFU

[TT17a] Kathleen M. C. Tjørve and Even Tjørve. A proposed family of unified models for sigmoidal growth. Ecological Modelling, 359(??):117–127, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300832.

Tuffa:2017:MBC

[TT17b] Samuel Tuffa and Anna C. Treydte. Modeling boran cattle populations under climate change and varying carrying capacity. Ecological Modelling, 352(??):113–127, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017302077.

Travers-Trolet:2019:ENT

[TTCC^{+}19] Morgane Travers-Trolet, Franck Coppin, Pierre Cresson, Philippe Cugier, Ricardo Oliveros-Ramos, and Philippe Verley. Emergence of negative trophic level-size relationships from a size-based, individual-based multispecies fish model. Ecological Modelling, 410(??):Article 108800, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303084.

Taghizadeh-Toosi:2014:CTS

[TTCH^{+}14] Arezoo Taghizadeh-Toosi, Bent T. Christensen, Nicholas J. Hutchings, Jonas Vejlin, Thomas Kätterer, Margaret Glendining, and Jørgen E. Olesen. C-TOOL: a simple model for simulating whole-profile carbon storage in temperate agricultural soils. Ecological Modelling, 292(??):11–25, November 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003950.

Tracy:2018:RSF

[TTL^{+}18] James L. Tracy, Antonio Trabucco, A. Michelle Lawing, J. Tomasz Giermakowski, Maria Tchakerian, Gail M. Drus, and Robert N. Coulson. Random subset feature selection for ecological niche models of wildfire activity in Western North America. Ecological Modelling, 383(??):52–68, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018301868.

Tulowiecki:2014:UVD

[Tul14] Stephen J. Tulowiecki. Using vegetation data within presettlement land survey records for species distribution modeling: a tale of two datasets. Ecological Modelling, 291(??):109–120, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003627.

Tompkins:2015:BMP

[TV15] Daniel M. Tompkins and Clare J. Veltman. Behaviourmanipulating parasites as adjuncts to vertebrate pest control. Ecological Modelling, 302(??):1–8, April 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000356.

Thorpe:2010:MSD

[TVF^{+}10] H. C. Thorpe, M. C. Vanderwel, M. M. Fuller, S. C. Thomas, and J. P. Caspersen. Modelling stand development after partial harvests: an empirically based, spatially explicit analysis for lowland black spruce. Ecological Modelling, 221(2):256–267, January 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009006589.

Tenan:2013:PAS

[TVI^{+}13] Simone Tenan, Andreu Rotger Vallespir, José Manuel Igual, Óscar Moya, J. Andrew Royle, and Giacomo Tavecchia. Population abundance, size structure and sex-ratio in an insular lizard. Ecological Modelling, 267(??):39–47, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013003608.

Tedesco:2012:PSE

[TVT12] Letizia Tedesco, Marcello Vichi, and David N. Thomas. Process studies on the ecological coupling between sea ice algae and phytoplankton. Ecological Modelling, 226 (??):120–138, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005400.

Torres:2013:ETE

[TVV13] César Torres, Antonio Valero, and Alicia Valero. Exergoecology as a tool for ecological modelling. The case of the US food production chain. Ecological Modelling, 255(??):21–28, April 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380013000483.

Tesfaye:2018:MTI

[TW18] Gashaw Tesfaye and Matthias Wolff. Modeling trophic interactions and the impact of an introduced exotic carp species in the Rift Valley Lake Koka, Ethiopia. Ecological Modelling, 378(??):26–36, June 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301145.

Tremblay:2014:ERM

[TWH^{+}14] Nelly Tremblay, Thorsten Werner, Kim Huenerlage, Friedrich Buchholz, Doris Abele, Bettina Meyer, and Thomas Brey. Euphausiid respiration model revamped: Latitudinal and seasonal shaping effects on krill respiration rates. Ecological Modelling, 291(??):233–241, November 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014003871.

Tyson:2011:MMP

[TWL11] Rebecca C. Tyson, J. Ben Wilson, and W. David Lane. A mechanistic model to predict transgenic seed contamination in bee-pollinated crops validated in an apple orchard. Ecological Modelling, 222(13):2084–2092, July 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011001761.

Tuominen:2014:CTE

[TWP^{+}14] Lindsey K. Tuominen, Stuart J. Whipple, Bernard C. Patten, Zekeriya Y. Karatas, and Caner Kazanci. Contribution of throughflows to the ecological interpretation of integral network utility. Ecological Modelling, 293 (??):187–201, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014000702.

Tian:2011:PVS

[TWS^{+}11] Yu Tian, Jianguo Wu, Andrew T. Smith, Tianming Wang, Xiaojun Kou, and Jianping Ge. Population viability of the Siberian tiger in a changing landscape: Going, going and gone? Ecological Modelling, 222(17):3166–3180, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001100336X.

Tian:2010:MER

[TXM^{+}10] Hanqin Tian, Xiaofeng Xu, Shili Miao, Erik Sindhoj, Bray J. Beltran, and Zhijian Pan. Modeling ecosystem responses to prescribed fires in a phosphorus-enriched Everglades wetland: I. Phosphorus dynamics and cattail recovery. Ecological Modelling, 221(9):1252–1266, May 10, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008710.

Toda:2010:MAT

[TYEH10] Motomu Toda, Masayuki Yokozawa, Seita Emori, and Toshihiko Hara. More asymmetric tree competition brings about more evapotranspiration and less runoff from the forest ecosystems: a simulation study. Ecological Modelling, 221(24):2887–2898, December 15, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004205.

Tian:2015:CCS

[TZD^{+}15] Huaiyu Tian, Sen Zhou, Lu Dong, Thomas P. Van Boeckel, Yao Pei, Qizhong Wu, Wenping Yuan, Yan Guo, Shanqian Huang, Wenhuan Chen, Xueliang Lu, Zhen Liu, Yuqi Bai, Tianxiang Yue, Bryan Grenfell, and Bing Xu. Climate change suggests a shift of H5N1 risk in migratory birds. Ecological Modelling, 306(??):6–15, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003822.

Tian:2014:IPA

[TZS^{+}14] Zhan Tian, Honglin Zhong, Laixiang Sun, Günther Fischer, Harrij van Velthuizen, and Zhuoran Liang. Improving performance of agro-ecological zone (AEZ) modeling by cross-scale model coupling: an application to japonica rice production in Northeast China. Ecological Modelling, 290(??):155–164, October 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005681.

Ulgiati:2011:MEE

[UAB^{+}11] S. Ulgiati, M. Ascione, S. Bargigli, F. Cherubini, P. P. Franzese, M. Raugei, S. Viglia, and A. Zucaro. Material, energy and environmental performance of technological and social systems under a life cycle assessment perspective. Ecological Modelling, 222(1):176–189, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004618.

Ulgiati:2012:RQT

[UB12] Sergio Ulgiati and Mark T. Brown. Resource quality, technological efficiency and factors of scale within the emergy framework: a response to Marco Raugei. Ecological Modelling, 227(??):109–111, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005862.

Uchmanski:2016:IVM

[Uch16] Janusz Uchmański. Individual variability and metapopulation dynamics: an individual-based model. Ecological Modelling, 334(??):8–18, August 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301478.

Uso-Domenech:2018:SMP

[UDNSLC^{+}18] Josep-Lluis Usó-Doménech, Josué-Antonio Nescolarde-Selva, Miguel Lloret-Climent, Lucía González-Franco, and Kristian Alonso-Stenberg. Spatial model of a pyrophite shrub in Mediterranean terrestrial ecosystems. Ecological Modelling, 384(??):333–340, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302357.

Uso-Domenech:2016:SLE

[UDNSLCG16] J. L. Usó-Doménech, J. A. Nescolarde-Selva, M. Lloret-Climent, and H. Gash. Semantics of language for ecosystems modelling: a model case. Ecological Modelling, 328(??):85–94, May 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300412.

Uria-Diez:2017:CPS

[UDP17] Jaime Uria-Diez and Arne Pommerening. Crown plasticity in Scots pine (Pinus sylvestris L.) as a strategy of adaptation to competition and environmental factors. Ecological Modelling, 356(??):117–126, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301515.

Urtizberea:2013:SEZ

[UDRA13] Agurtzane Urtizberea, Nicolas Dupont, Rune Rosland, and Dag L. Aksnes. Sensitivity of euphotic zone properties to CDOM variations in marine ecosystem models. Ecological Modelling, 256(??):16–22, May 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000926.

Uthes:2010:PRT

[UFK^{+}10] Sandra Uthes, Katharina Fricke, Hannes König, Peter Zander, Martin van Ittersum, Stefan Sieber, Katharina Helming, Annette Piorr, and Klaus Müller. Policy relevance of three integrated assessment tools — a comparison with specific reference to agricultural policies. Ecological Modelling, 221(18):2136–2152, September 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009005705.

Ulrich:2012:NMA

[UG12] Werner Ulrich and Nicholas J. Gotelli. A null model algorithm for presence–absence matrices based on proportional resampling. Ecological Modelling, 244(??):20–27, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012003201.

Ueyama:2011:SCS

[UKI^{+}11] Masahito Ueyama, Atsushi Kai, Kazuhito Ichii, Ken Hamotani, Yoshiko Kosugi, and Nobutaka Monji. The sensitivity of carbon sequestration to harvesting and climate conditions in a temperate cypress forest: Observations and modeling. Ecological Modelling, 222(17):3216–3225, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011002675.

Ulanowicz:2010:RCD

[Ula10] Robert E. Ulanowicz. Re: Comment on “The dual nature of ecosystem”. Ecological Modelling, 221(22):2627, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010003868.

Ulanowicz:2014:RNP

[Ula14] Robert E. Ulanowicz. Reckoning the nonexistent: Putting the science right. Ecological Modelling, 293(??):22–30, December 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380014001586.

Urban:2018:MMP

[UM18] Natasha A. Urban and Stephen F. Matter. Metapopulation mirages: Problems parsing process from pattern. Ecological Modelling, 375(??):20–29, May 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300656.

Ursino:2010:CEF

[UR10] Nadia Ursino and M. Cristina Rulli. Combined effect of fire and water scarcity on vegetation patterns in arid lands. Ecological Modelling, 221(19):2353–2362, September 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002875.

Urbani:2010:APB

[URJ10] Pasquinell Urbani and Rodrigo Ramos-Jiliberto. Adaptive prey behavior and the dynamics of intraguild predation systems. Ecological Modelling, 221(22):2628–2633, November 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010004047.

Upadhyay:2017:NTS

[URR^{+}17] Shashankaditya Upadhyay, Arijit Roy, M. Ramprakash, Jobin Idiculla, A. Senthil Kumar, and Sudeepto Bhattacharya. A network theoretic study of ecological connectivity in Western Himalayas. Ecological Modelling, 359 (??):246–257, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016308006.

Ullah:2012:TMC

[URUNAM12] Md. Hadayet Ullah, Md. Rashed-Un-Nabi, and Md. Abdulla Al-Mamun. Trophic model of the coastal ecosystem of the Bay of Bengal using mass balance Ecopath model. Ecological Modelling, 225(??):82–94, January 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005424.

Vacchiano:2018:RRH

[VAB^{+}18] Giorgio Vacchiano, Davide Ascoli, Fabio Berzaghi, Manuel Esteban Lucas-Borja, Thomas Caignard, Alessio Collalti, Paola Mairota, Ciprian Palaghianu, Christopher P. O. Reyer, Tanja G. M. Sanders, Eliane Schermer, Thomas Wohlgemuth, and Andrew Hacket-Pain. Reproducing reproduction: How to simulate mast seeding in forest models. Ecological Modelling, 376(??):40–53, May 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300887.

Volf:2011:DPM

[VAK^{+}11] Goran Volf, Nataša Atanasova, Boris Kompare, Robert Precali, and Nevenka Ožanić. Descriptive and prediction models of phytoplankton in the northern Adriatic. Ecological Modelling, 222(14):2502–2511, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000949.

Viana:2012:ECB

[VALC12] H. Viana, J. Aranha, D. Lopes, and Warren B. Cohen. Estimation of crown biomass of Pinus pinaster stands and shrubland above-ground biomass using forest inventory data, remotely sensed imagery and spatial prediction models. Ecological Modelling, 226(??):22–35, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011005692.

Vilas:2017:FDP

[VAO^{+}17] Maria P. Vilas, Matthew P. Adams, Carolyn E. Oldham, Clelia L. Marti, and Matthew R. Hipsey. Fragment dispersal and plant-induced dieback explain irregular ring-shaped pattern formation in a clonal submerged macrophyte. Ecological Modelling, 363(??):111–121, November 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303940.

Villasante:2016:MME

[VASH^{+}16] S. Villasante, F. Arreguín-Sánchez, J. J. Heymans, S. Libralato, C. Piroddi, V. Christensen, and M. Coll. Modelling marine ecosystems using the Ecopath with Ecosim food web approach: New insights to address complex dynamics after 30 years of developments. Ecological Modelling, 331(??):1–4, July 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016301430.

Vogler:2015:ELT

[VASL^{+}15] Rodolfo Vögler, Francisco Arreguín-Sánchez, Diego Lercari, Pablo del Monte-Luna, and Danilo Calliari. The effects of long-term climate variability on the trophodynamics of an estuarine ecosystem in southern South America. Ecological Modelling, 317(??):83–92, December 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004135.

Vavilin:2010:EIA

[Vav10] V. A. Vavilin. Equation for isotope accumulation in products and biomass as a way to reveal the pathways in mesophilic methanol methanization by microbial community. Ecological Modelling, 221(24):2881–2886, December 15, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004606.

Vavilin:2012:EED

[Vav12] V. A. Vavilin. Estimating evolution of δ ^{13}CH_{4} during methanogenesis in the boreal peatland ecosystems based on stoichiometric chemical reactions, microbial dynamics and stable carbon isotope fractionation. Ecological Modelling, 240(??):84–92, August 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002074.

Vila:2010:PFV

[VB10] José Pablo Solans Vila and Paulo Barbosa. Post-fire vegetation regrowth detection in the Deiva Marina region (Liguria-Italy) using Landsat TM and ETM+ data. Ecological Modelling, 221(1):75–84, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009001872.

Vuilleumier:2011:IEC

[VBPY11] S. Vuilleumier, A. Buttler, N. Perrin, and J. M. Yearsley. Invasion and eradication of a competitively superior species in heterogeneous landscapes. Ecological Modelling, 222(3):398–406, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005247.

Vanwindekens:2014:NAC

[VBS14] Frédéric M. Vanwindekens, Philippe V. Baret, and Didier Stilmant. A new approach for comparing and categorizing farmers’ systems of practice based on cognitive mapping and graph theory indicators. Ecological Modelling, 274(??):1–11, February 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005747.

Varga:2017:GEE

[VC17a] M. Varga and B. Csukas. Generation of extensible ecosystem models from a network structure and from locally executable programs. Ecological Modelling, 364(??):25–41, November 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304088.

Vinatier:2017:NMS

[VC17b] F. Vinatier and M. Chauvet. A neutral model for the simulation of linear networks in territories. Ecological Modelling, 363(??):8–16, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303800.

Vallina:2017:PFD

[VCD^{+}17] S. M. Vallina, P. Cermeno, S. Dutkiewicz, M. Loreau, and J. M. Montoya. Phytoplankton functional diversity increases ecosystem productivity and stability. Ecological Modelling, 361(??):184–196, October 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303246.

Velazquez-Castro:2017:RRM

[VCE17] Jorge Velazquez-Castro and Markus P. Eichhorn. Relative ranges of mating and dispersal modulate Allee thresholds in sessile species. Ecological Modelling, 359 (??):269–275, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016304926.

Vermunt:2012:CTE

[VCSSC12] Bradley Vermunt, Kim Cuddington, Stephanie Sobek-Swant, and Jill Crosthwaite. Cold temperature and emerald ash borer: Modelling the minimum under-bark temperature of ash trees in Canada. Ecological Modelling, 235–236(??):19–25, June 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001603.

vanDijk:2013:FQM

[vDHH^{+}13] Diana van Dijk, Rene Haijema, Eligius M. T. Hendrix, Rolf A. Groeneveld, and Ekko C. van Ierland. Fluctuating quota and management costs under multiannual adjustment of fish quota. Ecological Modelling, 265 (??):230–238, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013003128.

vanDijk:2014:SBL

[vDHH^{+}14] Diana van Dijk, Eligius M. T. Hendrix, Rene Haijema, Rolf A. Groeneveld, and Ekko C. van Ierland. On solving a bi-level stochastic dynamic programming model for analyzing fisheries policies: Fishermen behavior and optimal fish quota. Ecological Modelling, 272(??):68–75, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004444.

vandeKoppel:2011:SSE

[vdKGV11] Johan van de Koppel, Rohit Gupta, and Cornelis Vuik. Scaling-up spatially-explicit ecological models using graphics processors. Ecological Modelling, 222(17):3011–3019, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380011003371.

vanDelden:2010:IAA

[vDSC^{+}10] Hedwig van Delden, Tomasz Stuczynski, Pavel Ciaian, Maria Luisa Paracchini, Jelle Hurkens, Artur Lopatka, Yu e Shi, Oscar Gomez Prieto, Silvia Calvo, Jasper van Vliet, and Roel Vanhout. Integrated assessment of agricultural policies with dynamic land use change modelling. Ecological Modelling, 221(18):2153–2166, September 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010001687.

vanderVaart:2015:CEI

[vdVBJS15] Elske van der Vaart, Mark A. Beaumont, Alice S. A. Johnston, and Richard M. Sibly. Calibration and evaluation of individual-based models using approximate Bayesian computation. Ecological Modelling, 312(??):182–190, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002173.

vanderVaart:2016:PHM

[vdVJS16] Elske van der Vaart, Alice S. A. Johnston, and Richard M. Sibly. Predicting how many animals will be where: How to build, calibrate and evaluate individual-based models. Ecological Modelling, 326(??):113–123, April 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015003750.

Velez-Espino:2012:CIC

[VEK12] Luis A. Vélez-Espino and Marten A. Koops. Capacity for increase, compensatory reserves, and catastrophes as determinants of minimum viable population in freshwater fishes. Ecological Modelling, 247(??):319–326, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004905.

Vorpahl:2012:HCS

[VEMS12] Peter Vorpahl, Helmut Elsenbeer, Michael Märker, and Boris Schröder. How can statistical models help to determine driving factors of landslides? Ecological Modelling, 239(??):27–39, July 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005898.

Verdy:2010:MPP

[Ver10] Ariane Verdy. Modulation of predator–prey interactions by the Allee effect. Ecological Modelling, 221(8):1098–1107, April 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010000268.

Vetharaniam:2012:PRR

[Vet12] I. Vetharaniam. Predicting reproduction rate of Varroa. Ecological Modelling, 224(1):11–17, January 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005011.

Vasudev:2016:MCI

[VF16] Divya Vasudev and Robert J. Fletcher, Jr. Mate choice interacts with movement limitations to influence effective dispersal. Ecological Modelling, 327(??):65–73, May 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016300023.

Vihervaara:2019:IEE

[VFB19] Petteri Vihervaara, Pier Paolo Franzese, and Elvira Buonocore. Information, energy, and eco-exergy as indicators of ecosystem complexity. Ecological Modelling, 395(??):23–27, March 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019300249.

Verma:2016:MCE

[VFFP16] Manish Verma, Mark A. Friedl, Adrien Finzi, and Nathan Phillips. Multi-criteria evaluation of the suitability of growth functions for modeling remotely sensed phenology. Ecological Modelling, 323(??):123–132, March 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005712.

Verniest:2019:MAE

[VG19] Fabien Verniest and Sabine Greulich. Methods for assessing the effects of environmental parameters on biological communities in long-term ecological studies — a literature review. Ecological Modelling, 414(??):Article 108732, December 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019302327.

vonGeibler:2010:SAE

[vGKB10] Justus von Geibler, Kora Kristof, and Katrin Bienge. Sustainability assessment of entire forest value chains: Integrating stakeholder perspectives and indicators in decision support tools. Ecological Modelling, 221(18):2206– 2214, September 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380010001675.

Vincenot:2011:TCC

[VGR^{+}11] Christian Ernest Vincenot, Francesco Giannino, Max Rietkerk, Kazuyuki Moriya, and Stefano Mazzoleni. Theoretical considerations on the combined use of system dynamics and individual-based modeling in ecology. Ecological Modelling, 222(1):210–218, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005004.

Veresoglou:2012:MED

[VH12] Stavros D. Veresoglou and John M. Halley. A model that explains diversity patterns of arbuscular mycorrhizas. Ecological Modelling, 231(??):146–152, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000555.

Vimercati:2017:IAS

[VHDM17] Giovanni Vimercati, Cang Hui, Sarah J. Davies, and G. John Measey. Integrating age structured and landscape resistance models to disentangle invasion dynamics of a pond-breeding anuran. Ecological Modelling, 356(??):104–116, July 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307657.

Valbuena:2017:EAA

[VHM^{+}17] R. Valbuena, A. Hernando, J. A. Manzanera, E. B. Görgens, D. R. A. Almeida, F. Mauro, A. García-Abril, and D. A. Coomes. Enhancing of accuracy assessment for forest aboveground biomass estimates obtained from remote sensing via hypothesis testing and overfitting evaluation. Ecological Modelling, 366(??):15–26, December 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304362.

Villaescusa:2016:CDM

[VJR^{+}16] Juan Antonio Villaescusa, Sven Erik Jørgensen, Carlos Rochera, David Velázquez, Antonio Quesada, and Antonio Camacho. Carbon dynamics modelization and biological community sensitivity to temperature in an oligotrophic freshwater Antarctic lake. Ecological Modelling, 319(??):21–30, January 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015001192.

Verweij:2010:IPI

[VKdW^{+}10] P. J. F. M. Verweij, M. J. R. Knapen, W. P. de Winter, J. J. F. Wien, J. A. te Roller, S. Sieber, and J. M. L. Jansen. An IT perspective on integrated environmental modelling: the SIAT case. Ecological Modelling, 221(18):2167–2176, September 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010000451.

Vitale:2017:IIC

[VLF^{+}17] Marcello Vitale, Silvia Lorenzetti, Fabio Francesconi, Fabio Attorre, and Mario Di Traglia. The importance of interspecific competition in the actual and future distributions of plant species assessed by a 2-d grid agent modelling. Ecological Modelling, 360(??):399–409, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017301783.

vanLangevelde:2016:SSB

[vLTdBP16] Frank van Langevelde, Zewdu K. Tessema, Willem F. de Boer, and Herbert H. T. Prins. Soil seed bank dynamics under the influence of grazing as alternative explanation for herbaceous vegetation transitions in semi-arid rangelands. Ecological Modelling, 337(??):253–261, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302551.

Vallecillo:2019:ESA

[VLZ^{+}19] Sara Vallecillo, Alessandra La Notte, Grazia Zulian, Silvia Ferrini, and Joachim Maes. Ecosystem services accounts: Valuing the actual flow of nature-based recreation from ecosystems to people. Ecological Modelling, 392(??):196–211, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001830320X.

VanGroenendael:2012:ISI

[VMG^{+}12] J. Van Groenendael, C. Mony, M. Garbey, F. Campillo, and A. El Hamidi. Introduction to the special issue of ecological modelling: “Modelling clonal plant growth: From ecological concepts to mathematics”. Ecological Modelling, 234(??):1–2, June 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001202.

Valente:2019:FAN

[VMP^{+}19] Donatella Valente, Pier Paolo Miglietta, Donatella Porrini, Maria Rita Pasimeni, Giovanni Zurlini, and Irene Petrosillo. A first analysis on the need to integrate ecological aspects into financial insurance. Ecological Modelling, 392(??):117–127, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303971.

VanFrausum:2010:DPR

[VMSM10] Johan Van Frausum, Jack J. Middelburg, Karline Soetaert, and Filip J. R. Meysman. Different proxies for the reactivity of aquatic sediments towards oxygen: a model assessment. Ecological Modelling, 221(17):2054–2067, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002668.

Vuilleumier:2012:IPH

[VP12] Séverine Vuilleumier and Hugh P. Possingham. Interacting populations in heterogeneous environments. Ecological Modelling, 228(??):96–105, March 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006168.

Vilizzi:2012:BIE

[VPB^{+}12] L. Vilizzi, A. Price, L. Beesley, B. Gawne, A. J. King, J. D. Koehn, S. N. Meredith, D. L. Nielsen, and C. P. Sharpe. The belief index: an empirical measure for evaluating outcomes in Bayesian belief network modelling. Ecological Modelling, 228(??):123–129, March 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000166.

Vassallo:2017:AVN

[VPB^{+}17] P. Vassallo, C. Paoli, E. Buonocore, P. P. Franzese, G. F. Russo, and P. Povero. Assessing the value of natural capital in marine protected areas: a biophysical and trophodynamic environmental accounting model. Ecological Modelling, 355(??):12–17, July 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300984.

Vergara:2011:IBA

[VPC11] Pablo M. Vergara, Jaime Pizarro, and Sergio A. Castro. An island biogeography approach for understanding changes in compositional similarity at present scenario of biotic homogenization. Ecological Modelling, 222(12):1964–1971, June 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011001475.

VanLooy:2014:IMF

[VPC^{+}14] K. Van Looy, J. Piffady, C. Cavillon, T. Tormos, P. Landry, and Y. Souchon. Integrated modelling of functional and structural connectivity of river corridors for European otter recovery. Ecological Modelling, 273(??):228–235, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005462.

Vogel:2018:WED

[VPM^{+}18] L. E. Vogel, V. Pot, D. Makowski, P. Garnier, and P. C. Baveye. To what extent do uncertainty and sensitivity analyses help unravel the influence of microscale physical and biological drivers in soil carbon dynamics models? Ecological Modelling, 383(??):10–22, September 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301583.

Viola:2010:AGW

[VPS10] Flavio M. Viola, Susana L. D. Paiva, and Marcelo A. Savi. Analysis of the global warming dynamics from temperature time series. Ecological Modelling, 221(16):1964–1978, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010002401.

vanPoorten:2018:LBM

[vPTOW18] Brett T. van Poorten, Nathan Taylor, David O’Brien, and Carl J. Walters. A length-based mark-recapture model for estimating abundance and recruitment: Removing bias due to size-selective capture gear. Ecological Modelling, 381(??):10–22, August 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301406.

Voss:2018:QBS

[VQS^{+}18] Rudi Voss, Martin F. Quaas, Jörn O. Schmidt, Max T. Stoeven, Tessa B. Francis, Phillip S. Levin, Derek R. Armitage, Jaclyn S. Cleary, R. Russ Jones, Lynn C. Lee, Daniel K. Okamoto, Jennifer J. Silver, Thomas F. Thornton, Sherri C. Dressel, Alec D. MacCall, and André E. Punt. Quantifying the benefits of spatial fisheries management — an ecological-economic optimization approach. Ecological Modelling, 385(??):165–172, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302473.

VanderWal:2012:IBQ

[VR12] E. Vander Wal and A. R. Rodgers. An individual-based quantitative approach for delineating core areas of animal space use. Ecological Modelling, 224(1):48–53, January 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004911.

Virgili:2017:CHM

[VRA^{+}17] Auriane Virgili, Mélanie Racine, Matthieu Authier, Pascal Monestiez, and Vincent Ridoux. Comparison of habitat models for scarcely detected species. Ecological Modelling, 346(??):88–98, February 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016308146.

Vavilin:2017:MMF

[VRC17] Vasily Vavilin, Sergey Rytov, and Ralf Conrad. Modelling methane formation in sediments of tropical lakes focusing on syntrophic acetate oxidation: Dynamic and static carbon isotope equations. Ecological Modelling, 363(??):81–95, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017303903.

Vavilin:2014:NLD

[VRL14] V. A. Vavilin, S. V. Rytov, and L. Y. Lokshina. Nonlinear dynamics of nitrogen isotopic signature based on biological kinetic model of uptake and assimilation of ammonium, nitrate and urea by a marine diatom. Ecological Modelling, 279(??):45–53, May 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000970.

Vavilin:2018:DIE

[VRL18] Vasily Vavilin, Sergey Rytov, and Lyudmila Lokshina. Dynamic isotope equations for ^{13}CH_{4} and ^{13}CO_{2} describing methane formation with a focus on the effect of anaerobic respiration in sediments of some tropical lakes. Ecological Modelling, 386(??):59–70, October 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302680.

Vermard:2010:IFT

[VRM^{+}10] Youen Vermard, Etienne Rivot, Stéphanie Mahévas, Paul Marchal, and Didier Gascuel. Identifying fishing trip behaviour and estimating fishing effort from VMS data using Bayesian hidden Markov models. Ecological Modelling, 221(15):1757–1769, July 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001912.

Vanwindekens:2013:DBC

[VSB13] Frédéric M. Vanwindekens, Didier Stilmant, and Philippe V. Baret. Development of a broadened cognitive mapping approach for analysing systems of practices in social–ecological systems. Ecological Modelling, 250(??):352–362, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012005613.

Valdez:2018:IRA

[VSC18] L. D. Valdez, G. J. Sibona, and C. A. Condat. Impact of rainfall on Aedes aegypti populations. Ecological Modelling, 385(??):96–105, October 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302382.

Vance:2010:UIB

[VSF10] Richard R. Vance, Mark A. Steele, and Graham E. Forrester. Using an individual-based model to quantify scale transition in demographic rate functions: Deaths in a coral reef fish. Ecological Modelling, 221(16):1907–1921, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002164.

vanStrien:2016:HHC

[vSGR16] Maarten J. van Strien and Adrienne Grêt-Regamey. How is habitat connectivity affected by settlement and road network configurations? Results from simulating coupled habitat and human networks. Ecological Modelling, 342 (??):186–198, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305063.

Vergara:2015:HSD

[VSPHS15] Pablo M. Vergara, Santiago Saura, Christian G. Pérez-Hernández, and Gerardo E. Soto. Hierarchical spatial decisions in fragmented landscapes: Modeling the foraging movements of woodpeckers. Ecological Modelling, 300(??):114–122, March 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000186.

Vejre:2015:DCD

[VVA^{+}15] H. Vejre, J. P. Vesterager, P. S. Andersen, A. S. Olafsson, J. Brandt, and T. Dalgaard. Does cadastral division of areabased ecosystem services obstruct comprehensive management? Ecological Modelling, 295(??):176–187, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014004657.

vanVliet:2011:RKA

[vVBHZ11] Jasper van Vliet, Arnold K. Bregt, and Alex Hagen-Zanker. Revisiting kappa to account for change in the accuracy assessment of land-use change models. Ecological Modelling, 222(8):1367–1375, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000494.

Vanhatalo:2012:SDM

[VVH12] Jarno Vanhatalo, Lari Veneranta, and Richard Hudd. Species distribution modeling with Gaussian processes: a case study with the youngest stages of sea spawning whitefish (Coregonus lavaretus L. s.l.) larvae. Ecological Modelling, 228(??):49–58, March 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011006132.

vanVliet:2013:FSA

[vVHZHvD13] Jasper van Vliet, Alex Hagen-Zanker, Jelle Hurkens, and Hedwig van Delden. A fuzzy set approach to assess the predictive accuracy of land use simulations. Ecological Modelling, 261–262(??):32–42, July 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001762.

Vitense:2016:TIH

[VWTA16] Kelsey Vitense, Aaron J. Wirsing, Rebecca C. Tyson, and James J. Anderson. Theoretical impacts of habitat loss and generalist predation on predator–prey cycles. Ecological Modelling, 327(??):85–94, May 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300114.

Vincenzi:2011:ARF

[VZF^{+}11] Simone Vincenzi, Matteo Zucchetta, Piero Franzoi, Michele Pellizzato, Fabio Pranovi, Giulio A. De Leo, and Patrizia Torricelli. Application of a random forest algorithm to predict spatial distribution of the potential yield of Ruditapes philippinarum in the Venice lagoon, Italy. Ecological Modelling, 222(8):1471–1478, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000640.

Wildhaber:2017:FSA

[WAGM17] Mark L. Wildhaber, Janice L. Albers, Nicholas S. Green, and Edward H. Moran. A fully-stochasticized, agestructured population model for population viability analysis of fish: Lower Missouri River endangered pallid sturgeon example. Ecological Modelling, 359(??):434–448, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015003257.

Walz:2015:IMS

[Wal15] Ulrich Walz. Indicators to monitor the structural diversity of landscapes. Ecological Modelling, 295(??):88–106, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001400338X.

Wallace:2016:MEE

[Wal16] Rodrick Wallace. The metabolic economics of environmental adaptation. Ecological Modelling, 322(??):48–53, February 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015005499.

Wallentin:2017:SSS

[Wal17] Gudrun Wallentin. Spatial simulation: a spatial perspective on individual-based ecology — a review. Ecological Modelling, 350(??):30–41, April 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306640.

Wandersee:2012:PDM

[WALCY12] Sarah M. Wandersee, Li An, David López-Carr, and Yeqin Yang. Perception and decisions in modeling coupled human and natural systems: a case study from Fanjingshan National Nature Reserve, China. Ecological Modelling, 229(??):37–49, March 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004212.

Wei:2013:SAV

[WAM13] W. Wei, I. Alvarez, and S. Martin. Sustainability analysis: Viability concepts to consider transient and asymptotical dynamics in socio-ecological tourism-based systems. Ecological Modelling, 251(??):103–113, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200511X.

Wang:2019:ETC

[Wan19] Zhaoqi Wang. Estimating of terrestrial carbon storage and its internal carbon exchange under equilibrium state. Ecological Modelling, 401(??):94–110, June 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001930105X.

Ward:2010:POF

[War10] David Ward. A parsimonious optimal foraging model explaining mortality patterns in Serengeti wildebeest. Ecological Modelling, 221(19):2406–2408, September 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003108.

Walters:2016:EAP

[WAS^{+}16] Jeffrey P. Walters, David W. Archer, Gretchen F. Sassenrath, John R. Hendrickson, Jon D. Hanson, John M. Halloran, Peter Vadas, and Vladimir J. Alarcon. Exploring agricultural production systems and their fundamental components with system dynamics modelling. Ecological Modelling, 333(??):51–65, August 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301417.

Walker:2010:PVS

[WB10] E. Walker and N. Bez. A pioneer validation of a statespace model of vessel trajectories (VMS) with observers’ data. Ecological Modelling, 221(17):2008–2017, August 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010002589.

Wan:2014:CMS

[WB14] Zhenwen Wan and Hongsheng Bi. Comparing model scenarios of variable plankton N/P ratio versus the constant one for the application in the Baltic Sea. Ecological Modelling, 272(??):28–39, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300447X.

Whitworth:2016:IOC

[WB16] Kerry L. Whitworth and Darren S. Baldwin. Improving our capacity to manage hypoxic blackwater events in lowland rivers: the blackwater risk assessment tool. Ecological Modelling, 320(??):292–298, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004597.

Wang:2017:MBR

[WB17] Peng Wang and Gaboury Benoit. Modeling the biogeochemical role of photosynthetic sulfur bacteria in phosphorus cycling in a managed eutrophic lake. Ecological Modelling, 361(??):66–73, October 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307608.

Watling:2021:PMV

[WBB^{+}21] James I. Watling, Laura A. Brandt, David N. Bucklin, Ikuko Fujisaki, Frank J. Mazzotti, Stephanie S. Romañach, and Carolina Speroterra. Performance metrics and variance partitioning reveal sources of uncertainty in species distribution models. Ecological Modelling, 309–310(??):48–59, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001283.

Wang:2012:SDE

[WBC^{+}12] Tao Wang, Pierre Brender, Philippe Ciais, Shilong Piao, Miguel D. Mahecha, Frédéric Chevallier, Markus Reichstein, Catherine Ottlé, Fabienne Maignan, Altaf Arain, Gil Bohrer, Alessandro Cescatti, Gerard Kiely, Beverly Elizabeth Law, Merbold Lutz, Leonardo Montagnani, Eddy Moors, Bruce Osborne, Oleg Panferov, Dario Papale, and Francesco Primo Vaccari. State-dependent errors in a land surface model across biomes inferred from eddy covariance observations on multiple timescales. Ecological Modelling, 246(??):11–25, November 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003456.

Wells:2014:BBS

[WBG^{+}14] Adam G. Wells, Colby C. Blair, Edward O. Garton, Clifford G. Rice, Jon S. Horne, Janet L. Rachlow, and David O. Wallin. The Brownian bridge synoptic model of habitat selection and space use for animals using GPS telemetry data. Ecological Modelling, 273(??):242–250, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005449.

Wethey:2011:PIO

[WBHM11] David S. Wethey, Lindsay D. Brin, Brian Helmuth, and K. A. S. Mislan. Predicting intertidal organism temperatures with modified land surface models. Ecological Modelling, 222(19):3568–3576, October 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004376.

Wan:2013:CTL

[WBS13] Zhenwen Wan, Hongsheng Bi, and Jun She. Comparison of two light attenuation parameterization focusing on timing of spring bloom and primary production in the Baltic Sea. Ecological Modelling, 259(??):40–49, June 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001671.

Woodall:2014:MHI

[WBW14] Hannah Woodall, James M. Bullock, and Steven M. White. Modelling the harvest of an insect pathogen. Ecological Modelling, 287(??):16–26, September 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002282.

Wang:2013:ESGa

[WC13a] Gangsheng Wang and Shulin Chen. Evaluation of a soil greenhouse gas emission model based on Bayesian inference and MCMC: Model uncertainty. Ecological Modelling, 253(??):97–106, March 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004759.

Wang:2013:ESGb

[WC13b] Gangsheng Wang and Shulin Chen. Evaluation of a soil greenhouse gas emission model based on Bayesian inference and MCMC: Parameter identifiability and equifinality. Ecological Modelling, 253(??):107–116, March 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004760.

Williams:2013:MMD

[WC13c] B. J. Williams and B. Cole. Mining monitored data for decision-making with a Bayesian network model. Ecological Modelling, 249(??):26–36, January 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200333X.

Walters:2019:ENA

[WC19] Carl Walters and Villy Christensen. Effect of non-additivity in mortality rates on predictions of potential yield of forage fishes. Ecological Modelling, 410(??):Article 108776, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019302844.

Walters:2016:PSP

[WCF^{+}16] Carl Walters, Villy Christensen, Beth Fulton, Anthony D. M. Smith, and Ray Hilborn. Predictions from simple predator–prey theory about impacts of harvesting forage fishes. Ecological Modelling, 337(??):272–280, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302587.

Wiedenmann:2011:EER

[WCG^{+}11] John Wiedenmann, Katherine A. Cresswell, Jeremy Goldbogen, Jean Potvin, and Marc Mangel. Exploring the effects of reductions in krill biomass in the Southern Ocean on blue whales using a state-dependent foraging model. Ecological Modelling, 222(18):3366–3379, September 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003838.

Wang:2017:PIM

[WCL^{+}17] Hongqing Wang, Qin Chen, Megan K. La Peyre, Kelin Hu, and Jerome F. La Peyre. Predicting the impacts of Mississippi River diversions and sea-level rise on spatial patterns of eastern oyster growth rate and production. Ecological Modelling, 352(??):40–53, May 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304483.

Weller:2014:SDA

[WCS^{+}14] Florian Weller, Lee-Anne Cecchini, Lynne Shannon, Richard B. Sherley, Robert J. M. Crawford, Res Altwegg, Leanne Scott, Theodor Stewart, and Astrid Jarre. A system dynamics approach to modelling multiple drivers of the African penguin population on Robben Island, South Africa. Ecological Modelling, 277(??):38–56, April 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000477.

Wu:2011:EDT

[WCZ11] Z. Wu, G. Q. Chen, and L. Zeng. Environmental dispersion in a two-zone wetland. Ecological Modelling, 222 (3):456–474, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005867.

Wang:2015:MES

[WDL^{+}15] Ying Wang, Lijie Duan, Shiyu Li, Zeyu Zeng, and Pierre Failler. Modeling the effect of the seasonal fishing moratorium on the Pearl River Estuary using ecosystem simulation. Ecological Modelling, 312(??):406–416, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015002604.

Williams:2011:ARM

[WEB11] Byron K. Williams, Mitchell J. Eaton, and David R. Breininger. Adaptive resource management and the value of information. Ecological Modelling, 222(18): 3429–3436, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003735.

Weinberger:2019:UUE

[Wei19] Vanessa P. Weinberger. Understanding Urban Ecology: an Interdisciplinary Systems Approach, Myrna H. P. Hall, Stephen B. Balogh (Eds.). Springer International Publishing (2019). Ecological Modelling, 410(??):Article 108781, October 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019302893.

Wulff:2011:NVP

[WEJ11] Fred Wulff, Bradley D. Eyre, and Ron Johnstone. Nitrogen versus phosphorus limitation in a subtropical coastal embayment (Moreton Bay; Australia): Implications for management. Ecological Modelling, 222(1):120–130, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004540.

Wang:2010:CBF

[WFA10] Yonghe Wang, Mike Flannigan, and Kerry Anderson. Correlations between forest fires in British Columbia, Canada, and sea surface temperature of the Pacific Ocean. Ecological Modelling, 221(1):122–129, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380008005905.

Watts:2013:TSR

[WFA^{+}13] Michael J. Watts, Damien A. Fordham, H. Resit Akçakaya, Matthew E. Aiello-Lammens, and Barry W. Brook. Tracking shifting range margins using geographical centroids of metapopulations weighted by population density. Ecological Modelling, 269(??):61–69, November 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004031.

Wiederholt:2010:MIH

[WFDDR10] Ruscena Wiederholt, Eduardo Fernandez-Duque, Duane R. Diefenbach, and Rasanayagam Rudran. Modeling the impacts of hunting on the population dynamics of red howler monkeys (Alouatta seniculus). Ecological Modelling, 221(20):2482–2490, October 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010003273.

Wu:2016:NFF

[WFM16] Jiechen Wu, Daniel Franzén, and Maria E. Malmström. Nutrient flows following changes in source strengths, land use and climate in an urban catchment, Råcksta Träsk in Stockholm, Sweden. Ecological Modelling, 338(??):69–77, October 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001630268X.

Woillez:2016:HBM

[WFN^{+}16] Mathieu Woillez, Ronan Fablet, Tran-Thanh Ngo, Maxime Lalire, Pascal Lazure, and Hélène de Pontual. A HMMbased model to geolocate pelagic fish from high-resolution individual temperature and depth histories: European sea bass as a case study. Ecological Modelling, 321 (??):10–22, February 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005098.

Wildemeersch:2019:MMS

[WFS^{+}19] Matthias Wildemeersch, Oskar Franklin, Rupert Seidl, Joeri Rogelj, Inian Moorthy, and Stefan Thurner. Modelling the multi-scaled nature of pest outbreaks. Ecological Modelling, 409(??):Article 108745, October 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302534.

White:2015:HEM

[WFSH15] David J. White, Kuishuang Feng, Laixiang Sun, and Klaus Hubacek. A hydro-economic MRIO analysis of the Haihe River Basin’s water footprint and water stress. Ecological Modelling, 318(??):157–167, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000368.

Wang:2017:UTF

[WFW^{+}17] Chao Wang, Tao Feng, Peifang Wang, Jun Hou, and Jin Qian. Understanding the transport feature of bloom-forming Microcystis in a large shallow lake: a new combined hydrodynamic and spatially explicit agent-based modelling approach. Ecological Modelling, 343(??):25–38, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016305920.

Wang:2011:EWE

[WGA^{+}11] Z. Wang, R. F. Grant, M. A. Arain, B. N. Chen, N. Coops, R. Hember, W. A. Kurz, D. T. Price, G. Stinson, J. A. Trofymow, J. Yeluripati, and Z. Chen. Evaluating weather effects on interannual variation in net ecosystem productivity of a coastal temperate forest landscape: a model intercomparison. Ecological Modelling, 222(17):3236–3249, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380011003383. See cor-

rigendum [WGA^{+}12].

Wang:2012:CEW

[WGA^{+}12] Z. Wang, R. F. Grant, M. A. Arain, B. N. Chen, N. Coops, R. Hember, W. A. Kurz, D. T. Price, G. Stinson, J. A. Trofymow, J. Yeluripati, and Z. Chen. Corrigendum to “Evaluating weather effects on interannual variation in net ecosystem productivity of a coastal temperate forest landscape: a model intercomparison” [Ecol. Model. 222 (17) (2011) 3236–3249]. Ecological Modelling, 226(??):140, February 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001100562X. See [WGA^{+}11].

Wang:2013:IWS

[WGA^{+}13] Z. Wang, R. F. Grant, M. A. Arain, P. Y. Bernier, B. Chen, J. M. Chen, A. Govind, L. Guindon, W. A. Kurz, C. Peng, D. T. Price, G. Stinson, J. Sun, J. A. Trofymowe, and J. Yeluripati. Incorporating weather sensitivity in inventory-based estimates of boreal forest productivity: a meta-analysis of process model results. Ecological Modelling, 260(??):25–35, July 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001737.

Wen:2017:MPO

[WGD^{+}17] Bin Wen, Nickitas Georgas, Charles Dujardins, Anand Kumaraswamy, and Alan Cohn. Modeling pathogens for oceanic contact recreation advisories in the New York City area using total event simulations. Ecological Modelling, 365(??):93–105, December 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304209.

Wang:2019:IML

[WGE^{+}19] Hsiao-Hsuan Wang, William E. Grant, Norman C. Elliott, Michael J. Brewer, Tomasz E. Koralewski, John K. Westbrook, Tavvs M. Alves, and Gregory A. Sword. Integrated modelling of the life cycle and aeroecology of windborne pests in temporally-variable spatially-heterogeneous environment. Ecological Modelling, 399(??):23–38, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300833.

Wang:2012:FSD

[WGL^{+}12] Guiming Wang, Lowell L. Getz, Alicia V. Linzey, Norman A. Slade, and Michael H. Kesner. Fast and slow dynamics of northern small mammal populations. Ecological Modelling, 225(??):95–102, January 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005436.

Wedekind-Grunert:2019:BRN

[WGSR19] Sina Wedekind-Grunert, Boris Schröder, and Dania Richter. Basic reproduction number of Lyme disease spirochaetes

— modelling various genospecies-host associations in Central Europe. Ecological Modelling, 411(??):Article 108821, November 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0304380019303291.

Wang:2012:SCH

[WGT12] Hsiao-Hsuan Wang, W. E. Grant, and P. D. Teel. Simulation of climate–host–parasite–landscape interactions: a spatially explicit model for ticks (Acari: Ixodidae). Ecological Modelling, 243(??):42–62, September 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002797.

Wang:2016:TBI

[WGTH16] Hsiao-Hsuan Wang, W. E. Grant, P. D. Teel, and S. A. Hamer. Tick-borne infectious agents in nature: Simulated effects of changes in host density on spatialtemporal prevalence of infected ticks. Ecological Modelling, 323(??):77–86, March 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015005724.

Wang:2012:IEM

[WHB12] Szu-Hua Wang, Shu-Li Huang, and William W. Budd. Integrated ecosystem model for simulating land use allocation. Ecological Modelling, 227(??):46–55, February 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005916.

Wade:2016:PMM

[WHB^{+}16] M. J. Wade, J. Harmand, B. Benyahia, T. Bouchez, S. Chaillou, B. Cloez, J.-J. Godon, B. Moussa Boudjemaa, A. Rapaport, T. Sari, R. Arditi, and C. Lobry. Perspectives in mathematical modelling for microbial ecology. Ecological Modelling, 321(??):64–74, February 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015005190.

Wong:2013:ECE

[WHC^{+}13] H. Wong, J. Hillier, D. B. Clark, J. Smith, and P. Smith. An explicit and computationally efficient method to initialise first-order-based soil organic matter models — the geometric series solution (GSS). Ecological Modelling, 267(??):48–53, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013003591.

Wen:2016:AOC

[WhFxGZ16] Xin Wen, Guo hua Fang, Yu xue Guo, and Lei Zhou. Adapting the operation of cascaded reservoirs on Yuan River for fish habitat conservation. Ecological Modelling, 337(??):221–230, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302411.

Willamo:2018:LHU

[WHH^{+}18] R. Willamo, L. Helenius, C. Holmström, L. Haapanen, V. Sandström, E. Huotari, K. Kaarre, U. Värre, A. Nuotiomäki, J. Happonen, and L. Kolehmainen. Learning how to understand complexity and deal with sustainability challenges — a framework for a comprehensive approach and its application in university education. Ecological Modelling, 370(??):1–13, February 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302041.

Wood:2019:SVE

[WHNR19] Kevin A. Wood, Geoff M. Hilton, Julia L. Newth, and Eileen C. Rees. Seasonal variation in energy gain explains patterns of resource use by avian herbivores in an agricultural landscape: Insights from a mechanistic model. Ecological Modelling, 409(??):Article 108762, October 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302704.

Warden:2015:SEP

[WHRR15] Melissa L. Warden, Heather L. Haas, Kenneth A. Rose, and Paul M. Richards. A spatially explicit population model of simulated fisheries impact on loggerhead sea turtles (Caretta caretta) in the Northwest Atlantic Ocean. Ecological Modelling, 299(??):23–39, March 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005985.

Williams:2011:RSU

[Wil11a] Byron K. Williams. Resolving structural uncertainty in natural resources management using POMDP approaches. Ecological Modelling, 222(5):1092–1102, March 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006812.

Willis:2011:MSA

[Wil11b] Jay Willis. Modelling swimming aquatic animals in hydrodynamic models. Ecological Modelling, 222(23–24): 3869–3887, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004893.

Williams:2012:RUA

[Wil12] Byron K. Williams. Reducing uncertainty about objective functions in adaptive management. Ecological Modelling, 225(??):61–65, January 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011005291.

Weiss:2015:RSG

[WJ15] Lina Weiss and Florian Jeltsch. The response of simulated grassland communities to the cessation of grazing. Ecological Modelling, 303(??):1–11, May 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000587.

Weiping:2011:MCC

[WJF^{+}11] Hu Weiping, Sven Erik Jørgensen, Zhang Fabing, Chen Yonggen, Hu Zhixin, and Yang Longyuan. A model on the carbon cycling in Lake Taihu, China. Ecological Modelling, 222(16):2973–2991, August 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002262.

Wu:2011:MSC

[WJK11] Si Hong Wu, Per-Erik Jansson, and Pasi Kolari. Modeling seasonal course of carbon fluxes and evapotranspiration in response to low temperature and moisture in a boreal Scots pine ecosystem. Ecological Modelling, 222 (17):3103–3119, September 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011003036.

Wang:2013:MVD

[WJL^{+}13] Qiang Wang, Sven E. Jørgensen, Jianjian Lu, Søren N. Nielsen, and Jiarui Zhang. A model of vegetation dynamics of Spartina alterniflora and Phragmites australis in an expanding estuarine wetland: Biological interactions and sedimentary effects. Ecological Modelling, 250 (??):195–204, February 10, 2013. CODEN ECMODT.

ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012005091.

Wan:2014:MET

[WJL14] Nian-Feng Wan, Jie-Xian Jiang, and Bo Li. Modeling ecological two-sidedness for complex ecosystems. Ecological Modelling, 287(??):36–43, September 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001926.

Watermeyer:2018:FBM

[WJS^{+}18] K. E. Watermeyer, A. Jarre, L. J. Shannon, P. Mulumba, and J. Botha. A frame-based modelling approach to understanding changes in the distribution and abundance of sardine and anchovy in the southern Benguela. Ecological Modelling, 371(??):1–17, March 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017305008.

Wu:2013:MDT

[WJvdL^{+}13] J. Wu, P. E. Jansson, L. van der Linden, K. Pilegaard, C. Beier, and A. Ibrom. Modelling the decadal trend of ecosystem carbon fluxes demonstrates the important role of functional changes in a temperate deciduous forest. Ecological Modelling, 260(??):50–61, July 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001725.

Watt:2011:MBS

[WK11] Michael S. Watt and Miko U. F. Kirschbaum. Moving beyond simple linear allometric relationships between tree height and diameter. Ecological Modelling, 222(23– 24):3910–3916, December 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011004960.

Williams:2017:GMO

[WK17] Perry J. Williams and William L. Kendall. A guide to multi-objective optimization for ecological problems with an application to cackling goose management. Ecological Modelling, 343(??):54–67, January 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016305610.

Wieland:2017:AFS

[WKF^{+}17] Ralf Wieland, Antje Kerkow, Linus Früh, Helge Kampen, and Doreen Walther. Automated feature selection for a machine learning approach toward modeling a mosquito distribution. Ecological Modelling, 352(??):108–112, May 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016307906.

Williams:2019:SEM

[WKH19] Perry J. Williams, William L. Kendall, and Mevin B. Hooten. Selecting ecological models using multi-objective optimization. Ecological Modelling, 404(??):21–26, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301577.

Wang:2019:CWL

[WKZ^{+}19] Yuyu Wang, Yu-Chun Kao, Yangming Zhou, Huan Zhang, Xiubo Yu, and Guangchun Lei. Can water level management, stock enhancement, and fishery restriction offset negative effects of hydrological changes on the four major Chinese carps in China’s largest freshwater lake? Ecological Modelling, 403(??):1–10, July 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301255.

Wang:2013:MHS

[WL13] Fei Wang and Binliang Lin. Modelling habitat suitability for fish in the fluvial and lacustrine regions of a new eco-city. Ecological Modelling, 267(??):115–126, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003797.

Wang:2019:SEB

[WLC^{+}19] Xueqi Wang, Gengyuan Liu, Luca Coscieme, Biagio F. Giannetti, Yan Hao, Yan Zhang, and Mark T. Brown. Study on the emergy-based thermodynamic geography of the jing-Jin-ji region: Combined multivariate statistical data with DMSP-OLS nighttime lights data. Ecological Modelling, 397(??):1–15, April 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300419.

Wang:2012:FPE

[WLDL12] Y. Wang, S. Y. Li, L. J. Duan, and Y. Liu. Fishery policy exploration in the Pearl River Estuary based on an Ecosim model. Ecological Modelling, 230(??):34–43, April 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012000373.

Weijerman:2016:AEM

[WLF^{+}16] M. Weijerman, J. S. Link, E. A. Fulton, E. Olsen, H. Townsend, S. Gaichas, C. Hansen, M. Skern-Mauritzen, I. C. Kaplan, R. Gamble, G. Fay, M. Savina, C. Ainsworth, I. van Putten, R. Gorton, R. Brainard, K. Larsen, and T. Hutton. Atlantis ecosystem model summit: Report from a workshop. Ecological Modelling, 335(??):35–38, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016301776.

Weissenberger:2010:MCD

[WLH^{+}10] Sebastian Weissenberger, Marc Lucotte, Stéphane Houel, Nicolas Soumis, Éric Duchemin, and René Canuel. Modeling the carbon dynamics of the La Grande hydroelectric complex in northern Quebec. Ecological Modelling, 221(4):610–620, February 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009007455.

Watts:2011:NMM

[WLR^{+}11] Michael J. Watts, Yuxiao Li, Bayden D. Russell, Camille Mellin, Sean D. Connell, and Damien A. Fordham. A novel method for mapping reefs and subtidal rocky habitats using artificial neural networks. Ecological Modelling, 222(15):2606–2614, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002547.

Wang:2018:WDT

[WLR18] Hao Wang, Zexian Lu, and Aditya Raghavan. Weak dynamical threshold for the “strict homeostasis” assumption in ecological stoichiometry. Ecological Modelling, 384 (??):233–240, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018302321.

Wilson:2010:HBM

[WLS^{+}10] Adam M. Wilson, Andrew M. Latimer, John A. Silander, Alan E. Gelfand, and Helen de Klerk. A hierarchical Bayesian model of wildfire in a Mediterranean biodiversity hotspot: Implications of weather variability and global circulation. Ecological Modelling, 221(1):106–112, January 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009006371.

Wilson:2011:CHB

[WLS^{+}11] Adam M. Wilson, Andrew M. Latimer, John A. Silander, Alan E. Gelfand, and Helen de Klerk. Corrigendum to “A Hierarchical Bayesian model of wildfire in a Mediterranean biodiversity hotspot: Implications of weather variability and global circulation” [Ecol. Modell. 221 (January (1)) (2010) 106–112]. Ecological Modelling, 222(18):3456, September 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004546.

Wang:2015:UMS

[WLY^{+}15] Qing Wang, Xuehua Liu, Tianxiang Yue, Chenliang Wang, and John P. Wilson. Using models and spatial analysis to analyze spatio-temporal variations of food provision and food potential across China’s agro-ecosystems. Ecological Modelling, 306(??):152–159, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006152.

Wang:2011:AMP

[WLZC11] Wenjun Wang, Suling Liu, Shushen Zhang, and Jingwen Chen. Assessment of a model of pollution disaster in near-shore coastal waters based on catastrophe theory. Ecological Modelling, 222(2):307–312, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004783.

Wang:2014:CEP

[WM14] Shufan Wang and Zhihui Ma. Comments to the effect of prey refuge in a simple predator–prey model. Ecological Modelling, 287(??):58–59, September 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002579.

Wise:2012:QMF

[WMCM12] Laura Wise, Alberto G. Murta, Joao P. Carvalho, and Marta Mesquita. Qualitative modelling of fishermen’s behaviour in a pelagic fishery. Ecological Modelling, 228(??):112–122, March 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011005904.

Wang:2013:GSA

[WMF^{+}13] Fugui Wang, David J. Mladenoff, Jodi A. Forrester, Cindy Keough, and William J. Parton. Global sensitivity analysis of a modified CENTURY model for simulating impacts of harvesting fine woody biomass for bioenergy. Ecological Modelling, 259(??):16–23, June 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001579.

Webster:2013:MCD

[WMK^{+}13] K. L. Webster, J. W. McLaughlin, Y. Kim, M. S. Packalen, and C. S. Li. Modelling carbon dynamics and response to environmental change along a boreal fen nutrient gradient. Ecological Modelling, 248(??):148–164, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012005042.

Wang:2012:CBH

[WMW^{+}12] Hongqing Wang, Ehab A. Meselhe, Michael G. Waldon, Matthew C. Harwell, and Chunfang Chen. Compartmentbased hydrodynamics and water quality modeling of a Northern Everglades Wetland, Florida, USA. Ecological Modelling, 247(??):273–285, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004723.

Wallentin:2017:DHM

[WN17] Gudrun Wallentin and Christian Neuwirth. Dynamic hybrid modelling: Switching between AB and SD designs of a predator–prey model. Ecological Modelling, 345 (??):165–175, February 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016303714.

Watkins:2015:SEA

[WNF^{+}15] A. Watkins, J. Noble, R. J. Foster, B. J. Harmsen, and C. P. Doncaster. A spatially explicit agent-based model of the interactions between jaguar populations and their habitats. Ecological Modelling, 306(??):268–277, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005560.

Watson:2013:EMT

[WNT^{+}13] Reg A. Watson, Gabrielle B. Nowara, Sean R. Tracey, Elizabeth A. Fulton, Cathy M. Bulman, Graham J. Edgar, Neville S. Barrett, Jeremy M. Lyle, Stewart D. Frusher, and Colin D. Buxton. Ecosystem model of Tasmanian waters explores impacts of climate-change induced changes in primary productivity. Ecological Modelling, 264(??):115–129, August 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002281.

Watanabe:2014:DEA

[WO14] Marcos D. B. Watanabe and Enrique Ortega. Dynamic emergy accounting of water and carbon ecosystem services: a model to simulate the impacts of land-use change. Ecological Modelling, 271(??):113–131, January 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013001555.

Wright:2015:SRE

[WØ15] Christina Wright and Hanne Østergård. Scales of renewability exemplified by a case study of three Danish pig production systems. Ecological Modelling, 315(??):28–36, November 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015001696.

Wolf:2011:EPI

[Wol11] Annett Wolf. Estimating the potential impact of vegetation on the water cycle requires accurate soil water parameter estimation. Ecological Modelling, 222(15):2595–2605, August 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001100264X.

Winkler:2014:ICF

[WP14] Eckart Winkler and Markus Peintinger. Impact of changing flood regime on a lakeshore plant community: Longterm observations and individual-based simulation. Ecological Modelling, 273(??):151–164, February 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005322.

Whipple:2014:IED

[WPB14] Stuart J. Whipple, Bernard C. Patten, and Stuart R. Borrett. Indirect effects and distributed control in ecosystems: Comparative network environ analysis of a sevencompartment model of nitrogen storage in the Neuse River Estuary, USA: Time series analysis. Ecological Modelling, 293(??):161–186, December 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004165.

Walker:2015:ISC

[WPG15] Adam N. Walker, Jan-Jaap Poos, and Rolf A. Groeneveld. Invasive species control in a one-dimensional metapopulation network. Ecological Modelling, 316(??): 176–184, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015003786.

Wang:2011:DTM

[WPZ^{+}11] Weifeng Wang, Changhui Peng, S. Y. Zhang, Xiaolu Zhou, Guy R. Larocque, Daniel D. Kneeshaw, and Xiangdong Lei. Development of TRIPLEX-Management model for simulating the response of forest growth to pre-commercial thinning. Ecological Modelling, 222(14):2249–2261, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004904.

Wu:2014:BCM

[WPZ^{+}14] Wenqi Wu, M. K. D. K. Piyaratne, Huiyan Zhao, Chunlong Li, Zuqing Hu, and Xiangshun Hu. Butterfly catastrophe model for wheat aphid population dynamics: Construction, analysis and application. Ecological Modelling, 288(??):55–61, September 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400266X.

Wan:2013:MRT

[WQD^{+}13] Yongshan Wan, Chelsea Qiu, Peter Doering, Mayra Ashton, Detong Sun, and Teresa Coley. Modeling residence time with a three-dimensional hydrodynamic model: Linkage with chlorophyll a in a subtropical estuary. Ecological Modelling, 268(??):93–102, October 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004018.

White:2012:EST

[WR12] Denis White and Brenda Rashleigh. Effects of stream topology on ecological community results from neutral models. Ecological Modelling, 231(??):20–24, April 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000518.

Watkins:2013:EPI

[WR13a] Katherine Shepard Watkins and Kenneth A. Rose. Evaluating the performance of individual-based animal movement models in novel environments. Ecological Modelling, 250(??):214–234, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005480.

Weckel:2013:CCB

[WR13b] Mark Weckel and Robert F. Rockwell. Can controlled bow hunts reduce overabundant white-tailed deer populations in suburban ecosystems? Ecological Modelling, 250(??):143–154, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005200.

Watkins:2017:SIB

[WR17] Katherine Shepard Watkins and Kenneth A. Rose. Simulating individual-based movement in dynamic environments. Ecological Modelling, 356(??):59–72, July 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307888.

Wright:2018:TFC

[WR18] Kelseyann S. Wright and Adrian V. Rocha. A test of functional convergence in carbon fluxes from coupled C and N cycles in Arctic tundra. Ecological Modelling, 383(??):31–40, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301844.

Watling:2012:DBV

[WRB^{+}12] James I. Watling, Stephanie S. Romañach, David N. Bucklin, Carolina Speroterra, Laura A. Brandt, Leonard G. Pearlstine, and Frank J. Mazzotti. Do bioclimate variables improve performance of climate envelope models? Ecological Modelling, 246(??):79–85, November 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003468.

Wu:2011:DMO

[WRM^{+}11] Jianghua Wu, Nigel T. Roulet, Tim R. Moore, Peter Lafleur, and Elyn Humphreys. Dealing with microtopography of an ombrotrophic bog for simulating ecosystem-level CO_{2} exchanges. Ecological Modelling, 222(4):1038–1047, February 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010003704.

Withrow:2018:WCP

[WRMB18] Frances G. Withrow, Daniel L. Roelke, Rika M. W. Muhl, and Joydeb Bhattacharyya. Water column processes differentially influence richness and diversity of neutral, lumpy and intransitive phytoplankton assemblages. Ecological Modelling, 370(??):22–32, February 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300127.

Wen:2011:ICH

[WRSL11] Li Wen, Kerrylee Rogers, Neil Saintilan, and Joanne Ling. The influences of climate and hydrology on population dynamics of waterbirds in the lower Murrumbidgee River floodplains in Southeast Australia: Implications for environmental water management. Ecological Modelling, 222(1):154–163, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004874.

Wiggering:2015:CMS

[WS15] Hubert Wiggering and Uta Steinhardt. A conceptual model for site-specific agricultural land-use. Ecological Modelling, 295(??):42–46, January 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003901.

Westervelt:2013:MRF

[WSBP13] James D. Westervelt, Jinelle H. Sperry, Jennifer L. Burton, and John G. Palis. Modeling response of frosted flatwoods salamander populations to historic and predicted climate variables. Ecological Modelling, 268(??):18–24, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013003815.

Wszola:2019:SDC

[WSC^{+}19] Lyndsie S. Wszola, Victoria L. Simonsen, Lucía Corral, Christopher J. Chizinski, and Joseph J. Fontaine. Simulating detection-censored movement records for home range analysis planning. Ecological Modelling, 392 (??):268–278, January 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018303405.

Wilder:2010:USM

[WSE10] Shawn M. Wilder, Andrew V. Suarez, and Micky D. Eubanks. The use of simulation modeling to evaluate the mechanisms responsible for the nutritional benefits of food-for-protection mutualisms. Ecological Modelling, 221(11):1505–1511, June 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010001444.

Wang:2012:SHA

[WSE12] Hao Wang, Robert W. Sterner, and James J. Elser. On the “strict homeostasis” assumption in ecological stoichiometry. Ecological Modelling, 243(??):81–88, September 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001200275X.

Wang:2017:IBM

[WSL17] Bin Wang, Herman H. Shugart, and Manuel T. Lerdau. An individual-based model of forest volatile organic compound emissions — UVAFME-VOC v1.0. Ecological Modelling, 350(??):69–78, April 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303659.

Weller:2016:PPC

[WSS^{+}16] Florian Weller, Richard B. Sherley, Lynne J. Shannon, Astrid Jarre, Theodor Stewart, Leanne Scott, Res Altwegg, Lee-Anne Cecchini, Robert J. M. Crawford, Deon Geldenhuys, Katrin Ludynia, and Lauren J. Waller. Penguins’ perilous conservation status calls for complementary approach based on sound ecological principles: reply to Butterworth et al. (2015). Ecological Modelling, 337(??):1–3, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016302010.

Waters:2015:ELV

[WSSM15] Edward K. Waters, Harvinder S. Sidhu, Leesa A. Sidhu, and Geoffry N. Mercer. Extended Lotka–Volterra equations incorporating population heterogeneity: Derivation and analysis of the predator–prey case. Ecological Modelling, 297(??):187–195, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005924.

Wang:2013:OAR

[WSW^{+}13] Chao Wang, Qiyuan Sun, Peifang Wang, Jun Hou, and Aiyu Qu. An optimization approach to runoff regulation for potential estuarine eutrophication control: Model development and a case study of Yangtze Estuary, China. Ecological Modelling, 251(??):199–210, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000094.

Weller:2016:SDM

[WSW^{+}16] Florian Weller, Richard B. Sherley, Lauren J. Waller, Katrin Ludynia, Deon Geldenhuys, Lynne J. Shannon, and Astrid Jarre. System dynamics modelling of the Endangered African penguin populations on Dyer and Robben islands, South Africa. Ecological Modelling, 327(??):44–56, May 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016000302.

Wang:2010:TSB

[WSZT10] Jian Wang, Weiguo Song, Hongyang Zheng, and Luciano Telesca. Temporal scaling behavior of human-caused fires and their connection to relative humidity of the atmosphere. Ecological Modelling, 221(1):85–89, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009001914.

Tam:2012:OOS

[wTA12] Tze wai Tam and Put O. Ang. Object-oriented simulation of coral competition in a coral reef community. Ecological Modelling, 245(??):111–120, October 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001391.

Wang:2018:SIE

[WTC^{+}18] Hsiao-Hsuan Wang, Marianna Taffi, Clément Chion, Brenda Rashleigh, Tin Klanjšček, Lora Harris, Peter Goethals, and Brian D. Fath. Special issue: Ecological modelling global conference 2016: 20th biennial ISEM conference, 8–12 May 2016, Towson, Maryland, USA. Ecological Modelling, 383(??):1–2, September 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301716.

Wang:2016:SIW

[WTG^{+}16] Hsiao-Hsuan Wang, Pete D. Teel, William E. Grant, Greta Schuster, and A. A. Pérez de León. Simulated interactions of white-tailed deer (Odocoileus virginianus), climate variation and habitat heterogeneity on southern cattle tick (Rhipicephalus (Boophilus) microplus) eradication methods in south Texas, USA. Ecological Modelling, 342(??):82–96, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016305178.

Werle:2017:MSD

[WTL17] Rodrigo Werle, Brigitte Tenhumberg, and John L. Lindquist. Modeling shattercane dynamics in herbicide-tolerant grain sorghum cropping systems. Ecological Modelling, 343 (??):131–141, January 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016306342.

Wang:2019:REP

[WTWS19] Lin Wang, Ying Tang, Rui-Wu Wang, and Xiao-Ya Shang. Re-evaluating the ‘plankton paradox’ using an interlinked empirical data and a food web model. Ecological Modelling, 407(??):Article 108721, September 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302212.

Weterings:2015:DDA

[WUB15] Robbie Weterings, Chanin Umponstira, and Hannah L. Buckley. Density-dependent allometric functional response models. Ecological Modelling, 303(??):12–18, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015000599.

Weinberger:2012:UHM

[WV12] Stefan Weinberger and Mark Vetter. Using the hydrodynamic model DYRESM based on results of a regional climate model to estimate water temperature changes at Lake Ammersee. Ecological Modelling, 244(??):38–48, October 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002888.

Wang:2018:LVE

[WWC^{+}18] Sai Wang, Lin Wang, Hao-Yen Chang, Feng Li, Jin-Peng Tang, Xing-An Zhou, Xing Li, Shi-Mi Tian, Hsing-Juh Lin, and Yang Yang. Longitudinal variation in energy flow networks along a large subtropical river, China. Ecological Modelling, 387(??):83–95, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302886.

Wu:2017:STM

[WWYO17] Zijian Wu, Xiaofu Wu, Zhihui Yang, and Linnan Ouyang. A simple thermodynamic model for evaluating the ecological restoration effect on a manganese tailing wasteland. Ecological Modelling, 346(??):20–29, February 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304410.

Wu:2018:IER

[WWYO18] Zijian Wu, Xiaofu Wu, Zhihui Yang, and Linnan Ouyang. Internal energy ratios as ecological indicators for description of the phytoremediation process on a manganese tailing site. Ecological Modelling, 374(??):14–21, April 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300620.

Wu:2011:PAB

[WX11] Guozheng Wu and Zongxue Xu. Prediction of algal blooming using EFDC model: Case study in the Daoxiang Lake. Ecological Modelling, 222(6):1245–1252, March 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000020.

Watt:2010:DHT

[WXB10] Michael S. Watt, Vega Xu, and Mark Bloomberg. Development of a hydrothermal time seed germination model which uses the Weibull distribution to describe base water potential. Ecological Modelling, 221(9):1267–1272, May 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010000566.

Wang:2015:IPP

[WYHC15] Jinxia Wang, Yu Yang, Jikun Huang, and Kevin Chen. Information provision, policy support, and farmers’ adaptive responses against drought: an empirical study in the North China plain. Ecological Modelling, 318(??):275–282, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001400619X.

Wu:2015:NMW

[WYPF15] Wei Wu, Kevin M. Yeager, Mark S. Peterson, and Richard S. Fulford. Neutral models as a way to evaluate the Sea Level Affecting Marshes Model (SLAMM). Ecological Modelling, 303(??):55–69, May 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000642.

Wen:2012:LCD

[WYS12] Li Wen, Xihua Yang, and Neil Saintilan. Local climate determines the NDVI-based primary productivity and flooding creates heterogeneity in semi-arid floodplain ecosystem. Ecological Modelling, 242(??):116–126, September 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002396.

Wang:2019:TSC

[WYY^{+}19] Sisi Wang, Ke Yang, Donghai Yuan, Kongjian Yu, and Yijing Su. Temporal-spatial changes about the landscape pattern of water system and their relationship with food and energy in a mega city in China. Ecological Modelling, 401(??):75–84, June 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300791.

Wen:2014:AAN

[WZD^{+}14] Xuding Wen, Zhonghui Zhao, Xiangwen Deng, Wenhua Xiang, Dalun Tian, Wende Yan, Xiaolu Zhou, and Changhui Peng. Applying an artificial neural network to simulate and predict Chinese fir (Cunninghamia lanceolata) plantation carbon flux in subtropical China. Ecological Modelling, 294(??):19–26, December 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004219.

Wu:2015:PFU

[WZG15] Feng Wu, Jinyan Zhan, and İnci Güneralp. Present and future of urban water balance in the rapidly urbanizing Heihe River Basin, Northwest China. Ecological Modelling, 318(??):254–264, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400605X.

Wang:2018:AEH

[WZGH18] Junna Wang, Zhonglong Zhang, Blair Greimann, and Victor Huang. Application and evaluation of the HEC-RAS — riparian vegetation simulation module to the Sacramento River. Ecological Modelling, 368(??):158–168, January 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017304660.

Worthington:2016:LFM

[WZL^{+}16] Thomas A. Worthington, Tianjiao Zhang, Daniel R. Logue, Aaron R. Mittelstet, and Shannon K. Brewer. Landscape and flow metrics affecting the distribution of a federally-threatened fish: Improving management, model fit, and model transferability. Ecological Modelling, 342 (??):1–18, December 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016304434.

Wang:2010:PFP

[WZWL10] Weiming Wang, Lei Zhang, Hailing Wang, and Zhenqing Li. Pattern formation of a predator–prey system with ivlev-type functional response. Ecological Modelling, 221(2):131–140, January 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009006310.

Xiong:2014:CPI

[XBvdV^{+}14] Wei Xiong, Juraj Balkovič, Marijn van der Velde, Xuesong Zhang, R. César Izaurralde, Rastislav Skalský, Erda Lin, Nathan Mueller, and Michael Obersteiner. A calibration procedure to improve global rice yield simulations with EPIC. Ecological Modelling, 273(??):128–139, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005255.

Xuan:2014:MCI

[XC14] Zhemin Xuan and Ni-Bin Chang. Modeling the climateinduced changes of lake ecosystem structure under the cascade impacts of hurricanes and droughts. Ecological Modelling, 288(??):79–93, September 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002634.

Xiao:2016:MRI

[XCG^{+}16] Sa Xiao, Ragan M. Callaway, Ryan Graebner, Jose L. Hierro, and Daniel Montesinos. Modeling the relative importance of ecological factors in exotic invasion: the origin of competitors matters, but disturbance in the nonnative range tips the balance. Ecological Modelling, 335 (??):39–47, September 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016301752.

Xu:2011:ACC

[XCL^{+}11] Shannan Xu, Zuozhi Chen, Chunhou Li, Xiaoping Huang, and Shiyu Li. Assessing the carrying capacity of tilapia in an intertidal mangrove-based polyculture system of Pearl River Delta, China. Ecological Modelling, 222(3):846–856, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010006198.

Xu:2013:EII

[XCR^{+}13] Yi Xu, Fei Chai, Kenneth A. Rose, Miguel Ñiquen C., and Francisco P. Chavez. Environmental influences on the interannual variation and spatial distribution of Peruvian anchovy (Engraulis ringens) population dynamics from 1991 to 2007: a three-dimensional modeling study. Ecological Modelling, 264(??):64–82, August 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013000367.

Xue:2017:EMG

[XGH^{+}17] Bao-Lin Xue, Qinghua Guo, Tianyu Hu, Guoqiang Wang, Yongcai Wang, Shengli Tao, Yanjun Su, Jin Liu, and Xiaoqian Zhao. Evaluation of modeled global vegetation carbon dynamics: Analysis based on global carbon flux and aboveground biomass data. Ecological Modelling, 355(??):84–96, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016304914.

Xia:2015:VSG

[XHC^{+}15] X. H. Xia, Y. Hu, G. Q. Chen, A. Alsaedi, T. Hayat, and X. D. Wu. Vertical specialization, global trade and energy consumption for an urban economy: a value added export perspective for Beijing. Ecological Modelling, 318 (??):49–58, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005638.

Xiaoyi:2016:RHM

[XJY^{+}16] Wang Xiaoyi, Yao Junyang, Shi Yan, Su Tingli, Wang Li, and Xu Jiping. Research on hybrid mechanism modeling of algal bloom formation in urban lakes and reservoirs. Ecological Modelling, 332(??):67–73, July 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300667.

Xue:2011:ICS

[XKI^{+}11] Bao-Lin Xue, Tomo’omi Kumagai, Shin’ichi Iida, Taro Nakai, Kazuho Matsumoto, Hikaru Komatsu, Kyoichi Otsuki, and Takeshi Ohta. Influences of canopy structure and physiological traits on flux partitioning between understory and overstory in an eastern Siberian boreal larch forest. Ecological Modelling, 222(8):1479–1490, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000536.

Xu:2015:MER

[XLL15] Xiaoming Xu, Dejun Li, and Yiqi Luo. Modeled ecosystem responses to intra-annual redistribution and levels of precipitation in a prairie grassland. Ecological Modelling, 297(??):33–41, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005687.

Xie:2011:OML

[XLR11] Feng Xie, Yi Lin, and Wenwei Ren. Optimizing model for land use/land cover retrieval from remote sensing imagery based on variable precision rough sets. Ecological Modelling, 222(2):232–240, January 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004060.

Xie:2011:EGT

[XLS11] Tao Xie, Xinhui Liu, and Tao Sun. The effects of groundwater table and flood irrigation strategies on soil water and salt dynamics and reed water use in the Yellow River Delta, China. Ecological Modelling, 222(2):241–252, January 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010000517.

Xiao:2013:DII

[XM13] Sa Xiao and Richard Michalet. Do indirect interactions always contribute to net indirect facilitation? Ecological Modelling, 268(??):1–8, October 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003864.

Xu:2014:CER

[XPY^{+}14] Yanhong Xu, Hong Peng, Yinqun Yang, Wanshun Zhang, and Shuangling Wang. A cumulative eutrophication risk evaluation method based on a bioaccumulation model. Ecological Modelling, 289(??):77–85, October 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003330.

Xu:2013:MFM

[XQZ^{+}13] Fu-Liu Xu, Ning Qin, Ying Zhu, Wei He, Xiang-Zhen Kong, Michael T. Barbour, Qi-Shuang He, Yan Wang, Hui-Ling Ou-Yang, and Shu Tao. Multimedia fate modeling of polycyclic aromatic hydrocarbons (PAHs) in Lake Small Baiyangdian, Northern China. Ecological Modelling, 252(??):246–257, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012001767.

Xu:2018:PDV

[XSL18] Zhichao Xu, Yang Shen, and Jinbao Liao. Patch dynamics of various plant-animal interactions in fragmented landscapes. Ecological Modelling, 368(??):27–32, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304787.

Xu:2018:AIN

[XSN^{+}18] Xu Xu, Chen Sun, Fengtian Neng, Jing Fu, and Guanhua Huang. AHC: an integrated numerical model for simulating agroecosystem processes — model description and application. Ecological Modelling, 390(??):23–39, December 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018303387.

Xiaohui:2015:CMP

[XTLM15] Chen Xiaohui, Wang Tiansong, Samuele Lo Piano, and Kozo Mayumi. China’s metabolic patterns and their potential problems. Ecological Modelling, 318(??):75–85, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015001209.

Xu:2011:MER

[XTPT11] Xiaofeng Xu, Hanqin Tian, Zhijian Pan, and Cassondra R. Thomas. Modeling ecosystem responses to prescribed fires in a phosphorus-enriched Everglades wetland: II. Phosphorus dynamics and community shift in response to hydrological and seasonal scenarios. Ecological Modelling, 222(23–24):3942–3956, December 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004625.

Xu:2011:VES

[XWC^{+}11] Fu-Liu Xu, Jun-Jun Wang, Bin Chen, Ning Qin, Wen-Jing Wu, Wei He, Qi-Shuang He, and Yin Wang. The variations of exergies and structural exergies along eutrophication gradients in Chinese and Italian lakes. Ecological Modelling, 222(2):337–350, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008084.

Xu:2011:RLH

[XWW^{+}11] Fu-Liu Xu, Wen-Jing Wu, Jun-Jun Wang, Ning Qin, Yan Wang, Qi-Shuang He, Wei He, and Shu Tao. Residual levels and health risk of polycyclic aromatic hydrocarbons in freshwater fishes from Lake Small Bai-Yang-Dian, Northern China. Ecological Modelling, 222(2):275–286, January 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010005296.

Xu:2011:EHA

[XYCZ11] F. Xu, Z. F. Yang, B. Chen, and Y. W. Zhao. Ecosystem health assessment of the plant-dominated Baiyangdian Lake based on eco-exergy. Ecological Modelling, 222(1):201–209, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004989.

Xu:2013:ISP

[XYCZ13] F. Xu, Z. F. Yang, B. Chen, and Y. W. Zhao. Impact of submerged plants on ecosystem health of the plant-dominated Baiyangdian Lake, China. Ecological Modelling, 252(??):167–175, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012003419.

Xiao:2019:ICS

[XYTS19] Zhengyan Xiao, Meiqin Yao, Xiaotong Tang, and Luxi Sun. Identifying critical supply chains: an input-output analysis for Food–Energy–Water Nexus in China. Ecological Modelling, 392(??):31–37, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303880.

Xu:2016:NMA

[XYY^{+}16] Zhihao Xu, Xinan Yin, Zhifeng Yang, Yanpeng Cai, and Tao Sun. New model to assessing nutrient assimilative capacity in plant-dominated lakes: Considering ecological effects of hydrological changes. Ecological Modelling, 332(??):94–102, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016300928.

Xing:2017:IBM

[XZC^{+}17] Lei Xing, Chongliang Zhang, Yong Chen, Yunne-Jai Shin, Philippe Verley, Haiqing Yu, and Yiping Ren. An individual-based model for simulating the ecosystem dynamics of Jiaozhou Bay, China. Ecological Modelling, 360 (??):120–131, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380017301461.

Xu:2016:UMF

[XZCL16] Wanying Xu, Chuanbin Zhou, Aixin Cao, and Min Luo. Understanding the mechanism of food waste management by using stakeholder analysis and social network model: an industrial ecology perspective. Ecological Modelling, 337(??):63–72, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630206X.

Xia:2017:ASP

[XZSL17] Linlin Xia, Yan Zhang, Xiaoxi Sun, and Jinjian Li. Analyzing the spatial pattern of carbon metabolism and its response to change of urban form. Ecological Modelling, 355(??):105–115, July 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016307591.

Xiao:2013:IFN

[XZZ^{+}13] Sa Xiao, Liang Zhao, Jia-Lin Zhang, Xiang-Tai Wang, and Shu-Yan Chen. The integration of facilitation into the neutral theory of community assembly. Ecological Modelling, 251(??):127–134, February 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005984.

Yanez-Arenas:2014:PSA

[YAGMM^{+}14] Carlos Yañez-Arenas, Roger Guevara, Enrique Martínez-Meyer, Salvador Mandujano, and Jorge M. Lobo. Predicting species’ abundances from occurrence data: Effects of sample size and bias. Ecological Modelling, 294(??):36–41, December 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004293.

Yang:2011:MOO

[Yan11] Wei Yang. A multi-objective optimization approach to allocate environmental flows to the artificially restored wetlands of China’s Yellow River Delta. Ecological Modelling, 222(2):261–267, January 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010004199.

Yu:2016:SDS

[YBA^{+}16] Lin Yu, Salim Belyazid, Cecilia Akselsson, Gregory van der Heijden, and Giuliana Zanchi. Storm disturbances in a Swedish forest — a case study comparing monitoring and modelling. Ecological Modelling, 320(??):102–113, January 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015004561.

Yang:2011:SSW

[YC11] Zhifeng Yang and Bin Chen. Systematic studies on wetlands in China. Ecological Modelling, 222(2):221–223, January 2011. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010005260.

Yang:2013:GGE

[YC13] Q. Yang and G. Q. Chen. Greenhouse gas emissions of corn–ethanol production in China. Ecological Modelling, 252(??):176–184, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012003377.

Yang:2015:TBT

[YCF15] Siyuan Yang, Bin Chen, and Brian Fath. Trans-boundary total suspended particulate matter (TSPM) in urban ecosystems. Ecological Modelling, 318(??):59–63, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001400475X.

Yuan:2014:VSM

[YCL^{+}14] Wenping Yuan, Wenwen Cai, Shuguang Liu, Wenjie Dong, Jiquan Chen, M. Altaf Arain, Peter D. Blanken, Alessandro Cescatti, Georg Wohlfahrt, Teodoro Georgiadis, Lorenzo Genesio, Damiano Gianelle, Achim Grelle, Gerard Kiely, Alexander Knohl, Dan Liu, Michal V. Marek, Lutz Merbold, Leonardo Montagnani, Oleg Panferov, Mikko Peltoniemi, Serge Rambal, Antonio Raschi, Andrej Varlagin, and Jiangzhou Xia. Vegetation-specific model parameters are not required for estimating gross primary production. Ecological Modelling, 292(??):1–10, November 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003962.

Yang:2013:SAN

[YDD^{+}13] J. Y. Yang, C. F. Drury, R. De Jong, E. C. Huffman, X. M. Yang, and Keith Reid. Sensitivity analysis for nitrogen inputs, nitrogen outputs, and changes in biofuel crop acreages for predicting residual soil nitrogen and nitrate leaching in Canadian agricultural soils. Ecological Modelling, 267(??):26–38, October 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001300361X.

Yue:2015:SME

[YDL^{+}15] Tian-Xiang Yue, Zheng-Ping Du, Ming Lu, Ze-Meng Fan, Chen-Liang Wang, Yong-Zhong Tian, and Bing Xu. Surface modeling of ecosystem responses to climatic change in Poyang Lake Basin of China. Ecological Modelling, 306(??):16–23, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001400430X.

Yurek:2013:SMD

[YDT^{+}13] Simeon Yurek, Donald L. DeAngelis, Joel C. Trexler, Fred Jopp, and Douglas D. Donalson. Simulating mechanisms for dispersal, production and stranding of small forage fish in temporary wetland habitats. Ecological Modelling, 250(??):391–401, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200525X.

Yang:2021:CXM

[YDZ^{+}21] Shengtian Yang, Guotao Dong, Donghai Zheng, Honglin Xiao, Yunfei Gao, and Yang Lang. Coupling Xinanjiang model and SWAT to simulate agricultural non-point source pollution in Songtao watershed of Hainan, China. Ecological Modelling, 222(20–22):3701–3717, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004637.

Yue:2011:SMG

[YFC^{+}11] Tian-Xiang Yue, Ze-Meng Fan, Chuan-Fa Chen, Xiao-Fang Sun, and Bai-Lian Li. Surface modelling of global terrestrial ecosystems under three climate change scenarios. Ecological Modelling, 222(14):2342–2361, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006356.

Yuan:2018:MPE

[YFL^{+}18] Youquan Yuan, Jingjie Feng, Ran Li, Yinghan Huang, Juping Huang, and Zhenhua Wang. Modelling the promotion effect of vegetation on the dissipation of supersaturated total dissolved gas. Ecological Modelling, 386(??):89–97, October 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018302850.

Ying:2018:ECI

[YGL18] Zhixia Ying, Gang Ge, and Yongjie Liu. The effects of clonal integration on the responses of plant species to habitat loss and habitat fragmentation. Ecological Modelling, 384(??):290–295, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302217.

Huang:2012:MRV

[yHAB12] Cho ying Huang, Gregory P. Asner, and Nichole N. Barger. Modeling regional variation in net primary production of pinyon–juniper ecosystems. Ecological Modelling, 227(??):82–92, February 24, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380011006041.

Yun:2017:CMM

[YHJ^{+}17] Kyungdahm Yun, Jennifer Hsiao, Myung-Pyo Jung, In-Tae Choi, D. Michael Glenn, Kyo-Moon Shim, and Soo-Hyung Kim. Can a multi-model ensemble improve phenology predictions for climate change studies? Ecological Modelling, 362(??):54–64, October 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303563.

Yoshida:2019:ECF

[YHK^{+}19] Katsuhiko Yoshida, Kenji Hata, Kazuto Kawakami, Syuntaro Hiradate, Takeshi Osawa, and Naoki Kachi. Ecosystem changes following the eradication of invasive species: Evaluation of various eradication scenarios by computer simulation. Ecological Modelling, 413(??):Article 108831, December 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380019303394.

Yang:2011:ICD

[YHS^{+}11] Jian Yang, Hong S. He, Stephen R. Shifley, Frank R. Thompson, and Yangjian Zhang. An innovative computer design for modeling forest landscape change in very large spatial extents with fine resolutions. Ecological Modelling, 222(15):2623–2630, August 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011002651.

Yu:2012:CWP

[YHW^{+}12] Fei-Hai Yu, Tomáš Herben, Radka Wildová, Chad Hershock, and Deborah E. Goldberg. Can we predict performance and spatial structure of two-species mixtures using only single species information from monocultures? Ecological Modelling, 234(??):31–37, June 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011003280.

Yang:2019:SLD

[YHW19] Yinghui Yang, Bang-Sheng Han, and Qin Wang. Spatiotemporal landscape disturbance contributes to the suppression of competing invaders. Ecological Modelling, 393(??):76–84, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018304228.

Yu:2019:MEE

[YHZ19] Yantuan Yu, Jianhuan Huang, and Ning Zhang. Modeling the eco-efficiency of Chinese prefecture-level cities with regional heterogeneities: a comparative perspective. Ecological Modelling, 402(??):1–17, June 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019301097.

Yue:2011:PGE

[YJL11] Tian-Xiang Yue, Sven E. Jorgensen, and Guy R. Larocque. Progress in global ecological modelling. Ecological Modelling, 222(14):2172–2177, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010002772.

Yang:2015:AEN

[YJW15] Tianxiang Yang, Dong Jing, and Shoubing Wang. Applying and exploring a new modeling approach of functional connectivity regarding ecological network: a case study on the dynamic lines of space syntax. Ecological Modelling, 318(??):126–137, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005882.

Yang:2010:SCE

[YKCN10] Zhaoqing Yang, Tarang Khangaonkar, Maria Calvi, and Kurt Nelson. Simulation of cumulative effects of nearshore restoration projects on estuarine hydrodynamics. Ecological Modelling, 221(7):969–977, April 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438000800584X.

Yang:2012:LBG

[YKL^{+}12] Xue-Qing Yang, Gayantha R. L. Kodikara, Eike Luedeling, Xue-Fei Yang, Jun He, Pei gui Liu, and Jian-Chu Xu. Looking below the ground: Prediction of Tuber indicum habitat using the weights of evidence method. Ecological Modelling, 247(??):27–39, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200381X.

Yue:2016:DES

[YKvWK16] Chaofang Yue, Hans-Peter Kahle, Klaus von Wilpert, and Ulrich Kohnle. A dynamic environment-sensitive site index model for the prediction of site productivity potential under climate change. Ecological Modelling, 337(??):48–62, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016302046.

Yang:2014:ESD

[YKZZ14] Dewei Yang, William Tze Ming Kao, Guoqin Zhang, and Nanyang Zhang. Evaluating spatiotemporal differences and sustainability of Xiamen urban metabolism using emergy synthesis. Ecological Modelling, 272(??):40–48, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004432.

Yan:2017:GPS

[YL17] Zhigang Yan and Junqing Li. Giant panda survival crisis remains serious based on the ecosystem catastrophe model. Ecological Modelling, 359(??):128–134, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017300145.

You:2018:DST

[YL18] Xiaoguang You and Jingling Liu. Describing the spatial– temporal dynamics of groundwater-dependent vegetation (GDV): a theoretical methodology. Ecological Modelling, 383(??):127–137, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018301704.

Yang:2013:APR

[YLC13] Tao Yang, Jingling Liu, and Qiuying Chen. Assessment of plain river ecosystem function based on improved gray system model and analytic hierarchy process for the Fuyang River, Haihe River Basin, China. Ecological Modelling, 268(??):37–47, October 24, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013003785.

Yu:2018:MIH

[YLC^{+}18] Qian Yu, Zhaowei Liu, Yongcan Chen, Dejun Zhu, and Na Li. Modelling the impact of hydrodynamic turbulence on the competition between Microcystis and Chlorella for light. Ecological Modelling, 370(??):50–58, February 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300140.

Yang:2015:SEH

[YLF^{+}15] Xiaofan Yang, Chongxuan Liu, Yilin Fang, Ross Hinkle, Hong-Yi Li, Vanessa Bailey, and Ben Bond-Lamberty. Simulations of ecosystem hydrological processes using a unified multi-scale model. Ecological Modelling, 296 (??):93–101, January 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014005304.

You:2014:MRE

[YLHZ14] L. You, Y. P. Li, G. H. Huang, and J. L. Zhang. Modeling regional ecosystem development under uncertainty — a case study for New Binhai District of Tianjin. Ecological Modelling, 288(??):127–142, September 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002762.

Yuan:2012:IMP

[YLL^{+}12] Wenping Yuan, Shunlin Liang, Shuguang Liu, Ensheng Weng, Yiqi Luo, David Hollinger, and Haicheng Zhang. Improving model parameter estimation using coupling relationships between vegetation production and ecosystem respiration. Ecological Modelling, 240(??):29–40, August 10, 2012. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002116.

Ying:2017:MTG

[YLL^{+}17] Zhixia Ying, Jinbao Liao, Yongjie Liu, Shichang Wang, Hui Lu, Liang Ma, Dongdong Chen, and Zhenqing Li. Modelling tree-grass coexistence in water-limited ecosystems. Ecological Modelling, 360(??):387–398, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306093.

Yuan:2011:APF

[YLW^{+}11] Zengwei Yuan, Xin Liu, Huijun Wu, Ling Zhang, and Jun Bi. Anthropogenic phosphorus flow analysis of Lujiang County, Anhui Province, Central China. Ecological Modelling, 222(8):1534–1543, April 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000482.

Yang:2014:CWU

[YLW14] Yinghui Yang, Weide Li, and Gang Wang. Can we use disease to control biological invasion? — a theoretical research. Ecological Modelling, 277(??):97–107, April 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014000660.

Long:2011:NSI

[yLWhMhG11] Tian yu Long, Lei Wu, Guo hu Meng, and Wei hua Guo. Numerical simulation for impacts of hydrodynamic conditions on algae growth in Chongqing Section of Jialing River, China. Ecological Modelling, 222(1):112–119, January 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380010004990.

Yan:2018:SGP

[YLY^{+}18] Jinxia Yan, Jingling Liu, Xiaoguang You, Xuan Shi, and Lulu Zhang. Simulating the gross primary production and ecosystem respiration of estuarine ecosystem in North China with AQUATOX. Ecological Modelling, 373(??):1–12, April 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300516.

You:2015:EMR

[YLZ15] Xiaoguang You, Jingling Liu, and Lulu Zhang. Ecological modeling of riparian vegetation under disturbances: a review. Ecological Modelling, 318(??):293–300, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003087.

Yang:2021:WSN

[YM21] Zhifeng Yang and Xufeng Mao. Wetland system network analysis for environmental flow allocations in the Baiyangdian Basin, China. Ecological Modelling, 222 (20–22):3785–3794, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001100473X.

Yniguez:2021:CMP

[YMCV21] Aletta T. Yñiguez, John W. McManus, and Ligia Collado-Vides. Consequences of morphological plasticity and fragmentation on space occupation of coral reef macroalgae. Ecological Modelling, 309–310(??):128–142, ???? 2021. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015001751.

Yan:2015:IGS

[YqWB^{+}15] Hao Yan, Shao qiang Wang, Dave Billesbach, Walter Oechel, Gil Bohrer, Tilden Meyers, Timothy A. Martin, Roser Matamala, Richard P. Phillips, Faiz Rahman, Qin Yu, and Herman H. Shugart. Improved global simulations of gross primary product based on a new definition of water stress factor and a separate treatment of C3 and C4 plants. Ecological Modelling, 297(??):42–59, February 10, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014005602.

Yoak:2016:OFR

[YRGH16] Andrew J. Yoak, John F. Reece, Stanley D. Gehrt, and Ian M. Hamilton. Optimizing free-roaming dog control programs using agent-based models. Ecological Modelling, 341(??):53–61, December 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001630463X.

Yiruhan:2014:LTP

[YSA^{+}14] Yiruhan, Masae Shiyomi, Tsuyoshi Akiyama, Shiping Wang, Yasuo Yamamura, Yoshimichi Hori, and Ailikun. Long-term prediction of grassland production for five temporal patterns of precipitation during the growing season of plants based on a system model in Xilingol, Inner Mongolia, China. Ecological Modelling, 291(??):183–192, November 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014003548.

Yee:2011:AED

[YSB11] Susan Harrell Yee, Deborah L. Santavy, and Mace G. Barron. Assessing the effects of disease and bleaching on Florida keys corals by fitting population models to data. Ecological Modelling, 222(7):1323–1332, April 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000317.

Yang:2010:HEA

[YSH^{+}10] Zhaoqing Yang, Kathryn L. Sobocinski, Danelle Heatwole, Tarang Khangaonkar, Ronald Thom, and Roger Fuller. Hydrodynamic and ecological assessment of nearshore restoration: a modeling study. Ecological Modelling, 221(7):1043–1053, April 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380009004906.

Yoshimura:2011:EME

[YTT^{+}11] Jin Yoshimura, Yumi Tanaka, Tatsuya Togashi, Shigehide Iwata, and Kei ichi Tainaka. Erratum to “Mathematical equivalence of geometric mean fitness with probabilistic optimization under environmental uncertainty” [Ecol. Model. 220 (2009) 2611–2617]. Ecological Modelling, 222(3):901, February 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010005892.

Yang:2017:RAA

[YWA^{+}17] Yubin Yang, Lloyd T. Wilson, Frank H. Arthur, Jing Wang, and Canchun Jia. Regional analysis of bin aeration as an alternative to insecticidal control for post-harvest management of Sitophilus oryzae (L.) and Rhyzopertha dominica (F.). Ecological Modelling, 359(??):165–181, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017301564.

Yan:2015:UAW

[YWH15] Tingting Yan, Jinxia Wang, and Jikun Huang. Urbanization, agricultural water use, and regional and national crop production in China. Ecological Modelling, 318 (??):226–235, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014006346.

Yi:2010:TDH

[YWY10] Yujun Yi, Zhaoyin Wang, and Zhifeng Yang. Twodimensional habitat modeling of Chinese sturgeon spawning sites. Ecological Modelling, 221(5):864–875, March 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380009008023.

Yu:2018:DLU

[YXD^{+}18] Dan Yu, Ping Xie, Xiaohua Dong, Bob Su, Xiaonong Hu, Kai Wang, and Shijin Xu. The development of land use planning scenarios based on land suitability and its influences on eco-hydrological responses in the upstream of the Huaihe River basin. Ecological Modelling, 373(??):53–67, April 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018300334.

Yao:2021:MMA

[YXZ^{+}21] Jianyu Yao, Peng Xiao, Yunhuai Zhang, Min Zhan, and Jiangwei Cheng. A mathematical model of algal blooms based on the characteristics of complex networks theory. Ecological Modelling, 222(20–22):3727–3733, ???? 2021. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011004650.

Yin:2011:DCR

[YY11] X. A. Yin and Z. F. Yang. Development of a coupled reservoir operation and water diversion model: Balancing human and environmental flow requirements. Ecological Modelling, 222(2):224–231, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001000325X.

Yin:2013:ROM

[YY13] X. A. Yin and Z. F. Yang. A reservoir operating model for directing water supply to humans, wetlands, and cones of depression. Ecological Modelling, 252(??):114–120, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004255.

Yousefpour:2019:SES

[YYH19] Rasoul Yousefpour, Bin You, and Marc Hanewinkel. Simulation of extreme storm effects on regional forest soil carbon stock. Ecological Modelling, 399(??):39–53, May 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019300882.

Yan:2013:NTS

[YYL13] Yan Yan, Zhifeng Yang, and Qiang Liu. Nonlinear trend in streamflow and its response to climate change under complex ecohydrological patterns in the Yellow River Basin, China. Ecological Modelling, 252(??):220–227, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012002438.

Yang:2011:OAS

[YYQ11] Wei Yang, Zhifeng Yang, and Yan Qin. An optimization approach for sustainable release of e-flows for lake restoration and preservation: Model development and a case study of Baiyangdian Lake, China. Ecological Modelling, 222(14):2448–2455, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010006861.

Yan:2016:IIS

[YZ16] Chuan Yan and Zhibin Zhang. Interspecific interaction strength influences population density more than carrying capacity in more complex ecological networks. Ecological Modelling, 332(??):1–7, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300965.

Yan:2018:DST

[YZ18] Chuan Yan and Zhibin Zhang. Dome-shaped transition between positive and negative interactions maintains higher persistence and biomass in more complex ecological networks. Ecological Modelling, 370(??):14–21, February 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380018300139.

Yu:2018:MFP

[YZA^{+}18] Lin Yu, Giuliana Zanchi, Cecilia Akselsson, Håkan Wallander, and Salim Belyazid. Modeling the forest phosphorus nutrition in a southwestern Swedish forest site. Ecological Modelling, 369(??):88–100, February 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001730501X.

Yang:2014:LUC

[YZC14] Xin Yang, Xin-Qi Zheng, and Rui Chen. A land use change model: Integrating landscape pattern indexes and Markov-CA. Ecological Modelling, 283(??):1–7, July 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001550.

Yi:2016:GSA

[YZG16] Xuan Yi, Rui Zou, and Huaicheng Guo. Global sensitivity analysis of a three-dimensional nutrients-algae dynamic model for a large shallow lake. Ecological Modelling, 327(??):74–84, May 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001600017X.

Yang:2012:SML

[YZL12] Xin Yang, Xin-Qi Zheng, and Li-Na Lv. A spatiotemporal model of land use change based on ant colony optimization, Markov chain and cellular automata. Ecological Modelling, 233(??):11–19, May 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001184.

Yuan:2014:DDM

[YZLX14] Zhiyuan Yuan, Xinqi Zheng, Lina Lv, and Chunlu Xue. From design to digital model: a quantitative analysis approach to garden cities theory. Ecological Modelling, 289(??):26–35, October 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001400297X.

Yang:2016:WQA

[YZPL16] Likun Yang, Xinhua Zhao, Sen Peng, and Xia Li. Water quality assessment analysis by using combination of Bayesian and genetic algorithm approach in an urban lake, China. Ecological Modelling, 339(??):77–88, November 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016303647.

Zilverberg:2018:SDC

[ZAW^{+}18] Cody J. Zilverberg, Jay Angerer, Jimmy Williams, Loretta J. Metz, and Keith Harmoney. Sensitivity of diet choices and environmental outcomes to a selective grazing algorithm. Ecological Modelling, 390(??):10–22, December 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018303302.

Zarba:2015:CEC

[ZB15] Lucía Zarbá and Mark T. Brown. Cycling emergy: computing emergy in trophic networks. Ecological Modelling, 315(??):37–45, November 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015000757.

Zanchi:2014:MEM

[ZBAY14] Giuliana Zanchi, Salim Belyazid, Cecilia Akselsson, and Lin Yu. Modelling the effects of management intensification on multiple forest services: a Swedish case study. Ecological Modelling, 284(??):48–59, July 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014001872.

Zepeda-Benitez:2017:ICA

[ZBMBDU^{+}17] Viridiana Y. Zepeda-Benitez, Enrique Morales-Bojórquez, Juan Gabriel Díaz-Uribe, Manuel O. Nevárez-Martínez, Agustín Hernández-Herrera, and Juana López-Martínez. Implementation of catch-at-age model for the jumbo squid Dosidicus gigas. Ecological Modelling, 344(??):6–16, January 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016306123.

Zabel:2014:MTP

[ZBMC14] Richard W. Zabel, Brian J. Burke, Mary L. Moser, and Christopher C. Caudill. Modeling temporal phenomena in variable environments with parametric models: an application to migrating salmon. Ecological Modelling, 273(??):23–30, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004870.

Zhao:2014:SUA

[ZBS14] Gang Zhao, Brett A. Bryan, and Xiaodong Song. Sensitivity and uncertainty analysis of the APSIM-wheat model:

Interactions between cultivar, environmental, and management parameters. Ecological Modelling, 279(??):1–11, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380014000878.

Zhao:2017:ATV

[ZBSF17] Qing Zhao, G. Scott Boomer, Emily Silverman, and Kathy Fleming. Accounting for the temporal variation of spatial effect improves inference and projection of population dynamics models. Ecological Modelling, 360(??):252–259, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017300753.

Zeng:2011:EDH

[ZC11] L. Zeng and G. Q. Chen. Ecological degradation and hydraulic dispersion of contaminant in wetland. Ecological Modelling, 222(2):293–300, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007108.

Zhang:2016:WGA

[ZCL^{+}16] Chunxiao Zhang, Min Chen, Rongrong Li, Chaoyang Fang, and Hui Lin. What’s going on about geo-process modeling in virtual geographic environments (VGEs). Ecological Modelling, 319(??):147–154, January 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001500174X.

Zhang:2016:EFC

[ZCR16] Chongliang Zhang, Yong Chen, and Yiping Ren. The efficacy of fisheries closure in rebuilding depleted stocks: Lessons from size-spectrum modeling. Ecological Modelling, 332(??):59–66, July 24, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016300977.

Zhao:2013:ARS

[ZCS13] Yanxia Zhao, Sining Chen, and Shuanghe Shen. Assimilating remote sensing information with crop model using ensemble Kalman filter for improving LAI monitoring and yield estimation. Ecological Modelling, 270(??):30–42, December 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001300416X.

Zhang:2018:APA

[ZCSL18] Lulu Zhang, Jiansheng Cui, Tiance Song, and Yong Liu. Application of an AQUATOX model for direct toxic effects and indirect ecological effects assessment of polycyclic aromatic hydrocarbons (PAHs) in a plateau eutrophication lake, China. Ecological Modelling, 388(??):31–44, November 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380018303168.

Zhang:2019:RPH

[ZCY^{+}19] Helin Zhang, Dongdong Chen, Zhixia Ying, Feng Zhang, and Jinbao Liao. Robustness of the pollination-herbivory system with high-order interactions to habitat loss. Ecological Modelling, 414(??):Article 108826, December 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019303345.

Zhang:2015:ICS

[ZCYL15] Dejian Zhang, Xingwei Chen, Huaxia Yao, and Bingqing Lin. Improved calibration scheme of SWAT by separating wet and dry seasons. Ecological Modelling, 301(??):54–61, April 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001500037X.

Zouiten:2013:ATE

[ZDG^{+}13] Hala Zouiten, César Álvarez Díaz, Andrés García Gómez, José Antonio Revilla Cortezón, and Javier García Alba. An advanced tool for eutrophication modeling in coastal lagoons: Application to the Victoria lagoon in the north of Spain. Ecological Modelling, 265(??):99–113, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002937.

Zhang:2017:LIB

[ZDL^{+}17] Jingjing Zhang, Todd E. Dennis, Todd J. Landers, Elizabeth Bell, and George L. W. Perry. Linking individual-based and statistical inferential models in movement ecology: a case study with black petrels (Procellaria parkinsoni). Ecological Modelling, 360(??):425–436, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304860.

Zhao:2011:STO

[ZF11] Xiaobing Zhao and Jerald J. Fletcher. A spatial–temporal optimization approach to watershed management: Acid mine drainage treatment in the Cheat River watershed, WV, USA. Ecological Modelling, 222(9):1580–1591, May 10, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380011000433.

Zhang:2019:UMM

[ZF19] Yan Zhang and Brian D. Fath. Urban metabolism: Measuring sustainable cities through ecological modelling. Ecological Modelling, 392(??):6–7, January 24, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018303132.

Zhou:2013:DPP

[ZFG13] Can Zhou, Masami Fujiwara, and William E. Grant. Dynamics of a predator–prey interaction with seasonal reproduction and continuous predation. Ecological Modelling, 268(??):25–36, October 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013003840.

Zhang:2010:RES

[ZFT^{+}10] Jianjun Zhang, Meichen Fu, Jin Tao, Ying Huang, Ferri P. Hassani, and Zhongke Bai. Response of ecological storage and conservation to land use transformation: a case study of a mining town in China. Ecological Modelling, 221(10):1427–1439, May 24, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010000852.

Zhang:2014:TAO

[ZFZ^{+}14] Jianjun Zhang, Meichen Fu, Zhongya Zhang, Jin Tao, and Wei Fu. A trade-off approach of optimal land allocation between socio-economic development and ecological stability. Ecological Modelling, 272(??):175–187, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004754.

Zubizarreta-Gerendiain:2014:RBS

[ZGF14] A. Zubizarreta-Gerendiain and M. P. Fernández. Relative branch size in branch clusters modelled through a Markovian process. Ecological Modelling, 273(??):210–219, February 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013005590.

Zhang:2010:AEE

[ZGJ10] Jingjie Zhang, Zeren Gurkan, and Sven Erik Jørgensen. Application of eco-exergy for assessment of ecosystem health and development of structurally dynamic models. Ecological Modelling, 221(4):693–702, February 24, 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009007030.

Zhang:2011:ANP

[ZH11] H. Zhang and G. H. Huang. Assessment of non-point source pollution using a spatial multicriteria analysis approach. Ecological Modelling, 222(2):313–321, January 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009008412.

Zell:2013:ILB

[ZH13] Chris Zell and Jason A. Hubbart. Interdisciplinary linkages of biophysical processes and resilience theory: Pursuing predictability. Ecological Modelling, 248(??):1–10, January 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380012004899.

Zeng:2016:SBW

[ZHC^{+}16] X. T. Zeng, G. H. Huang, H. L. Chen, Y. P. Li, X. M. Kong, and Y. R. Fan. A simulation-based water-environment management model for regional sustainability in compound wetland ecosystem under multiple uncertainties. Ecological Modelling, 334(??):60–77, August 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016301491.

Zhou:2015:MCC

[ZHCX15] Chuanbin Zhou, Heping Huang, Aixin Cao, and Wanying Xu. Modeling the carbon cycle of the municipal solid waste management system for urban metabolism. Ecological Modelling, 318(??):150–156, December 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006000.

Zheng:2017:FFS

[ZHLZ17] Zhong Zheng, Wei Huang, Songnian Li, and Yongnian Zeng. Forest fire spread simulating model using cellular automaton with extreme learning machine. Ecological Modelling, 348(??):33–43, March 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016303064.

Zielinski:2014:MEB

[ZHV14] D. P. Zielinski, M. Hondzo, and V. R. Voller. Mathematical evaluation of behavioral deterrent systems to disrupt fish movement. Ecological Modelling, 272(??):150–159, January 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013004766.

Zhou:2019:ATD

[ZJB19] Can Zhou, Yan Jiao, and Joan Browder. K -aggregated transformation of discrete distributions improves modeling count data with excess ones. Ecological Modelling, 407(??):Article 108726, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019302261.

Zinck:2010:WLD

[ZJG10] Richard D. Zinck, Karin Johst, and Volker Grimm. Wildfire, landscape diversity and the Drossel–Schwabl model. Ecological Modelling, 221(1):98–105, January 2010. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009000210.

Zhang:2014:MCM

[ZJL^{+}14] Jiarui Zhang, Sven E. Jørgensen, Jianjian Lu, Søren N. Nielsen, and Qiang Wang. A model for the contribution of macrophyte-derived organic carbon in harvested tidal freshwater marshes to surrounding estuarine and oceanic ecosystems and its response to global warming. Ecological Modelling, 294(??):105–116, December 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004487.

Zhang:2016:CAM

[ZJL^{+}16] Na Zhang, Yong-Cai Jing, Cheng-Yu Liu, Yao Li, and Jing Shen. A cellular automaton model for grasshopper population dynamics in inner Mongolia steppe habitats. Ecological Modelling, 329(??):5–17, June 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016300618.

Zhang:2011:LTF

[ZJM^{+}11] Chengfu Zhang, Rob C. Jamieson, Fan-Rui Meng, Robert J. Gordon, Jagtar Bhatti, and Charles P.-A. Bourque. Longterm forest-floor litter dynamics in Canada’s boreal forest: Comparison of two model formulations. Ecological Modelling, 222(6):1236–1244, March 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000214.

Zhang:2013:SMD

[ZJM^{+}13] Chengfu Zhang, Rob C. Jamieson, Fan-Rui Meng, Robert J. Gordon, and Charles P. A. Bourque. Simulation of monthly dissolved organic carbon concentrations in small forested watersheds. Ecological Modelling, 250(??):205–213, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380012005443.

Zhu:2011:EEF

[ZJP^{+}11] Qiuan Zhu, Hong Jiang, Changhui Peng, Jinxun Liu, Xiaohua Wei, Xiuqin Fang, Shirong Liu, Guomo Zhou, and Shuquan Yu. Evaluating the effects of future climate change and elevated CO_{2} on the water use efficiency in terrestrial ecosystems of China. Ecological Modelling, 222(14):2414–2429, July 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010005223.

Zhang:2013:CSP

[ZKR^{+}13] Weitao Zhang, Dong-Kyun Kim, Yerubandi R. Rao, Sue Watson, Shan Mugalingam, Tanya Labencki, Maria Dittrich, Andrew Morley, and George B. Arhonditsis. Can simple phosphorus mass balance models guide management decisions? A case study in the Bay of Quinte, Ontario, Canada. Ecological Modelling, 257(??):66–79, May 24, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013001105.

Zhdanova:2017:WPN

[ZKSF17] Oksana L. Zhdanova, Alexey E. Kuzin, Elena I. Skaletskaya, and Efim Ya. Frisman. Why the population of the northern fur seals (Callorhinus ursinus) of tyuleniy Island does not recover following the harvest ban: Analysis of 56 years of observation data. Ecological Modelling, 363(??):57–67, November 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380017303873.

Zhang:2014:STH

[ZLB14] Tao Zhang, Jeremy W. Lichstein, and Richard A. Birdsey. Spatial and temporal heterogeneity in the dynamics of eastern U.S. forests: Implications for developing broad-scale forest dynamics models. Ecological Modelling, 279(??):89–99, May 10, 2014. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014000969.

Zhang:2013:CES

[ZLC13] Yan Zhang, Hong Liu, and Bin Chen. Comprehensive evaluation of the structural characteristics of an urban metabolic system: Model development and a case study of Beijing. Ecological Modelling, 252(??):106–113, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004280.

Zhang:2019:VEC

[ZLD^{+}19] Xiao Zhang, Hong-Yi Li, Zhiqun D. Deng, L. Ruby Leung, John R. Skalski, and Steven J. Cooke. On the variable effects of climate change on Pacific salmon. Ecological Modelling, 397(??):95–106, April 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380019300572.

Zhang:2011:AUE

[ZLF^{+}11] Yan Zhang, Shengsheng Li, Brian D. Fath, Zhifeng Yang, and Naijin Yang. Analysis of an urban energy metabolic system: Comparison of simple and complex model results. Ecological Modelling, 223(1):14–19, December 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011004224.

Zhang:2014:SAU

[ZLF14] Yan Zhang, Hong Liu, and Brian D. Fath. Synergism analysis of an urban metabolic system: Model development and a case study for Beijing, China. Ecological Modelling, 272(??):188–197, January 24, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013004651.

Zhang:2015:AUC

[ZLF^{+}15] Yan Zhang, Juan Li, Brian D. Fath, Hongmei Zheng, and Linlin Xia. Analysis of urban carbon metabolic processes and a description of sectoral characteristics: a case study of Beijing. Ecological Modelling, 316(??):144–154, November 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003713.

Zhang:2011:NMW

[ZLFI11] Zhenlin Zhang, Ryan Lowe, James Falter, and Greg Ivey. A numerical model of wave- and current-driven nutrient uptake by coral reef communities. Ecological Modelling, 222(8):1456–1470, April 24, 2011. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380011000469.

Zhang:2016:MUN

[ZLFZ16] Yan Zhang, Hanjing Lu, Brian D. Fath, and Hongmei Zheng. Modelling urban nitrogen metabolic processes based on ecological network analysis: a case of study in Beijing, China. Ecological Modelling, 337(??):29–38, October 10, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016302009.

Zhang:2016:CLC

[ZLG16] Xiaoling Zhang, Yong Liu, and Huaicheng Guo. Cross-lake comparisons of physical and biological settling of phosphorus: a phosphorus budget model with Bayesian hierarchical approach. Ecological Modelling, 337(??):231–240, October 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380016302538.

Zhang:2017:SBA

[ZLJM17a] Chengfu Zhang, Sheng Li, Rob C. Jamieson, and Fan-Rui Meng. Segment-based assessment of riparian buffers on stream water quality improvement by applying an integrated model. Ecological Modelling, 345(??):1–9, February 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S030438001630758X.

Zhang:2017:CSB

[ZLJM17b] Chengfu Zhang, Sheng Li, Rob C. Jamieson, and Fanrui Meng. Corrigendum to “Segment-based assessment of riparian buffers on stream water quality improvement by applying an integrated model” [Ecol. Model. 345 (2017) 1–9]. Ecological Modelling, 359(??):460, September 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302855.

Zhang:2013:MPS

[ZLLC13] Chi Zhang, Chaofan Li, Geping Luo, and Xi Chen. Modeling plant structure and its impacts on carbon and water cycles of the Central Asian arid ecosystem in the context of climate change. Ecological Modelling, 267(??):158–179, October 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013002925.

Zhang:2013:APA

[ZLLZ13] Lulu Zhang, Jingling Liu, Yi Li, and Yanwei Zhao. Applying AQUATOX in determining the ecological risk assessment of polychlorinated biphenyl contamination in Baiyangdian Lake, North China. Ecological Modelling, 265 (??):239–249, September 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001300286X.

Zhang:2015:MIM

[ZLW^{+}15] Fan Zhang, Changsheng Li, Zheng Wang, Stanley Glidden, Danielle S. Grogan, Xuxiang Li, Yan Cheng, and Steve Frolking. Modeling impacts of management on farmland soil carbon dynamics along a climate gradient in Northwest China during 1981–2000. Ecological Modelling, 312(??):1–10, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015002033.

Zhang:2013:IEH

[ZLY^{+}13] L. L. Zhang, J. L. Liu, Z. F. Yang, Y. Li, and Y. Yang. Integrated ecosystem health assessment of a macrophyte-dominated lake. Ecological Modelling, 252 (??):141–152, March 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001200378X.

Zeng:2016:NMS

[ZLY16] Yiwen Zeng, Bi Wei Low, and Darren C. J. Yeo. Novel methods to select environmental variables in MaxEnt: a case study using invasive crayfish. Ecological Modelling, 341(??):5–13, December 10, 2016. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380016304665.

Zhang:2017:EVA

[ZLZ^{+}17a] Feng Zhang, Xingpeng Liu, Jiquan Zhang, Rina Wu, Qiyun Ma, and Yanan Chen. Ecological vulnerability assessment based on multi-sources data and SD model in Yinma River Basin, China. Ecological Modelling, 349(??):41–50, April 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016304975.

Zhang:2017:ENA

[ZLZ17b] Yan Zhang, Yanxian Li, and Hongmei Zheng. Ecological network analysis of energy metabolism in the Beijing– Tianjin–Hebei (Jing–Jin–Ji) urban agglomeration. Ecological Modelling, 351(??):51–62, May 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016307876.

Zhang:2017:ANM

[ZLZ17c] Yan Zhang, Hanjing Lu, and Xiaolin Zhang. Analysis of nitrogen metabolism processes and a description of structure characteristics. Ecological Modelling, 357(??):47–54, August 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380017301795.

Zhang:2014:DTA

[ZLZS14] Z. Zhang, W. X. Lu, Y. Zhao, and W. B. Song. Development tendency analysis and evaluation of the water ecological carrying capacity in the siping area of Jilin Province in China based on system dynamics and analytic hierarchy process. Ecological Modelling, 275(??):9–21, March 10, 2014. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380013005838.

Zhu:2019:ENA

[ZMH19] Xueting Zhu, Xianzhong Mu, and Guangwen Hu. Ecological network analysis of urban energy metabolic system — a case study of Beijing. Ecological Modelling, 404(??):36–45, July 24, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301619.

Zhao:2012:MCG

[ZML^{+}12] Wenqing Zhao, Yali Meng, Wenfeng Li, Binglin Chen, Naiyin Xu, Youhua Wang, Zhiguo Zhou, and Derrick M. Oosterhuis. A model for cotton (Gossypium hirsutum L.) fiber length and strength formation considering temperature-radiation and N nutrient effects. Ecological Modelling, 243(??):112–122, September 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012002876.

Zamora:2010:MWP

[ZMMHR10] R. Zamora, J. R. Molina-Martínez, M. A. Herrera, and F. Rodríguez y Silva. A model for wildfire prevention planning in game resources. Ecological Modelling, 221(1):19–26, January 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380009004918.

Zakharova:2019:TBM

[ZMS19] L. Zakharova, K. M. Meyer, and M. Seifan. Trait-based modelling in ecology: a review of two decades of research. Ecological Modelling, 407(??):Article 108703, September 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380019301735.

Zweig:2018:DTP

[ZNS^{+}18] Christa L. Zweig, Susan Newman, Colin J. Saunders, Fred H. Sklar, and Wiley M. Kitchens. Deviations on a theme: Peat patterning in sub-tropical landscapes. Ecological Modelling, 371(??):25–36, March 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018300231.

Zohmann:2013:MHS

[ZPNM13] Margit Zohmann, Josef Pennerstorfer, and Ursula Nopp-Mayr. Modelling habitat suitability for alpine rock ptarmigan (Lagopus muta helvetica) combining object-based classification of IKONOS imagery and habitat suitability index modelling. Ecological Modelling, 254(??):22–32, April 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380013000355.

Zamora:2015:CBF

[ZPZ^{+}15] Maria A. Zamora, Andres Pinzón, Maria M. Zambrano, Silvia Restrepo, Linda J. Broadbelt, Matthew Moura, Johana Husserl Orjuela, and Andrés F. González Barrios. A comparison between functional frequency and metabolic flows framed by biogeochemical cycles in metagenomes: the case of “El Coquito” hot spring located at Colombia’s National Nevados Park. Ecological Modelling, 313 (??):259–265, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380015002975.

Zhang:2015:EEU

[ZQC15] Bo Zhang, H. Qiao, and B. Chen. Embodied energy uses by China’s four municipalities: a study based on multi-regional input–output model. Ecological Modelling, 318(??):138–149, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014004761.

Zhao:2019:DMM

[ZR19] Qing Zhao and J. Andrew Royle. Dynamic N-mixture models with temporal variability in detection probability. Ecological Modelling, 393(??):20–24, February 1, 2019. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304186.

Zhang:2015:SEM

[ZRC^{+}15] Xiaoqing Zhang, Friedrich Recknagel, Qiuwen Chen, Hongqing Cao, and Ruonan Li. Spatially-explicit modelling and forecasting of cyanobacteria growth in Lake Taihu by evolutionary computation. Ecological Modelling, 306(??):216–225, June 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002622.

Zetina-Rejon:2015:TMC

[ZRCNLI^{+}15] Manuel J. Zetina-Rejón, Erika Cabrera-Neri, Gladis A. López-Ibarra, N. Enrique Arcos-Huitrón, and Villy Christensen. Trophic modeling of the continental shelf ecosystem outside of Tabasco, Mexico: a network and modularity analysis. Ecological Modelling, 313(??):314–324, October 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380015003038.

Zheng:2012:GSA

[ZSF^{+}12] Wei Zheng, Honghua Shi, Guohong Fang, Long Hu, Shitao Peng, and Mingyuan Zhu. Global sensitivity analysis of a marine ecosystem dynamic model of the Sanggou Bay. Ecological Modelling, 247(??):83–94, December 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012004000.

Zarkami:2012:UFD

[ZSG12] Rahmat Zarkami, Roghayeh Sadeghi, and Peter Goethals. Use of fish distribution modelling for river management. Ecological Modelling, 230(??):44–49, April 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012000221.

Zhu:2018:URT

[ZSG^{+}18] Zhenduo Zhu, David T. Soong, Tatiana Garcia, Mina Shahed Behrouz, Steven E. Butler, Elizabeth A. Murphy, Matthew J. Diana, James J. Duncker, and David H. Wahl. Using reverse-time egg transport analysis for predicting Asian carp spawning grounds in the Illinois River. Ecological Modelling, 384(??):53–62, September 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302023.

Zhang:2015:IBS

[ZSJ15] Zhonglong Zhang, Bowen Sun, and Billy E. Johnson. Integration of a benthic sediment diagenesis module into the two dimensional hydrodynamic and water quality model — CE-QUAL-W2. Ecological Modelling, 297(??):213–231, February 10, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380014004943.

Zipper:2017:QIG

[ZSKI17] Samuel C. Zipper, Mehmet Evren Soylu, Christopher J. Kucharik, and Steven P. Loheide II. Quantifying indirect groundwater-mediated effects of urbanization on agroecosystem productivity using MODFLOW-AgroIBIS (MAGI), a complete critical zone model. Ecological Modelling, 359(??):201–219, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017300789.

Zaitouny:2017:MTF

[ZSS17] Ayham Zaitouny, Thomas Stemler, and Michael Small. Modelling and tracking the flight dynamics of flocking pigeons based on real GPS data (small flock). Ecological Modelling, 344(??):62–72, January 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306937.

Zubillaga:2018:HDD

[ZSSR18] María Zubillaga, Oscar Skewes, Nicolás Soto, and Jorge E. Rabinovich. How density-dependence and climate affect guanaco population dynamics. Ecological Modelling, 385 (??):189–196, October 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S030438001830245X.

Zhang:2010:LDN

[ZTJ^{+}10] Chengfu Zhang, John A. Trofymow, Rob C. Jamieson, Fan-Rui Meng, Robert Gordon, and Charles P.-A. Bourque. Litter decomposition and nitrogen mineralization from an annual to a monthly model. Ecological Modelling, 221 (16):1944–1953, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380010002176.

Zhao:2017:IDA

[ZTL^{+}17] X. Zhao, M. R. Tillotson, Y. W. Liu, W. Guo, A. H. Yang, and Y. F. Li. Index decomposition analysis of urban crop water footprint. Ecological Modelling, 348(??):25–32, March 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016305683.

Zielinski:2018:PIA

[ZVS18] D. P. Zielinski, VR. Voller, and P. W. Sorensen. A physiologically inspired agent-based approach to model upstream passage of invasive fish at a lock-and-dam. Ecological Modelling, 382(??):18–32, August 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018301558.

Zheng:2010:RBN

[ZW10] Lianyuan Zheng and Robert H. Weisberg. Rookery Bay and Naples Bay circulation simulations: Applications to tides and fresh water inflow regulation. Ecological Modelling, 221(7):986–996, April 2010. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380009000763.

Zhang:2011:IDA

[ZW11] Xu Zhang and Wendi Wang. Importance of dispersal adaptations of two competitive populations between patches. Ecological Modelling, 222(1):11–20, January 10, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004977.

Zeng:2012:NLT

[ZW12] Chunhua Zeng and Hua Wang. Noise and large time delay: Accelerated catastrophic regime shifts in ecosystems. Ecological Modelling, 233(??):52–58, May 24, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S030438001200141X.

Zhang:2019:IST

[ZWA19] Ping Zhang, Jianwen Wang, and Peter M. Atkinson. Identifying the spatio-temporal risk variability of avian influenza A H7N9 in China. Ecological Modelling, 414(??): Article 108807, December 15, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380019303151.

Zhang:2018:RSA

[ZWF18] Yan Zhang, Qiong Wu, and Brian D. Fath. Review of spatial analysis of urban carbon metabolism. Ecological Modelling, 371(??):18–24, March 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017302028.

Zhang:2013:HPM

[ZWG^{+}13] Chi Zhang, Jianguo Wu, Nancy B. Grimm, Melissa McHale, and Alexander Buyantuyev. A hierarchical patch mosaic ecosystem model for urban landscapes: Model development and evaluation. Ecological Modelling, 250 (??):81–100, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380012004887.

Zhao:2017:SDC

[ZWG^{+}17] Yabo Zhao, Shaojian Wang, Yuejing Ge, Qianqian Liu, and Xiaofeng Liu. The spatial differentiation of the coupling relationship between urbanization and the eco-environment in countries globally: a comprehensive assessment. Ecological Modelling, 360(??):313–327, September 24, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017303411.

Zeng:2012:EWC

[ZWJ^{+}12] L. Zeng, Y. H. Wu, P. Ji, B. Chen, Y. J. Zhao, G. Q. Chen, and Z. Wu. Effect of wind on contaminant dispersion in a wetland flow dominated by free-surface effect. Ecological Modelling, 237–238(??):101–108, July 10, 2012. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001937.

Zilverberg:2017:PBS

[ZWJ^{+}17] Cody J. Zilverberg, Jimmy Williams, Curtis Jones, Keith Harmoney, Jay Angerer, Loretta J. Metz, and William Fox. Process-based simulation of prairie growth. Ecological Modelling, 351(??):24–35, May 10, 2017. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016304951.

Zhang:2013:LHW

[ZWRK13] Weitao Zhang, Sue B. Watson, Yerubandi R. Rao, and Hedy J. Kling. A linked hydrodynamic, water quality and algal biomass model for a large, multi-basin lake: a working management tool. Ecological Modelling, 269 (??):37–50, November 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380013004183.

Zhang:2018:CLR

[ZWW^{+}18] Libing Zhang, Lulu Wu, Chengguo Wu, Juliang Jin, Shouliang Huo, and Yuliang Zhou. Construction of lake reference conditions for nutrient criteria based on system dynamics modelling. Ecological Modelling, 383(??):69–79, September 10, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001830156X.

Zhao:2013:ATM

[ZXD^{+}13] Meifang Zhao, Wenhua Xiang, Xiangwen Deng, Dalun Tian, Zhihong Huang, Xiaolu Zhou, Guirui Yu, Honglin He, and Changhui Peng. Application of TRIPLEX model for predicting Cunninghamia lanceolata and Pinus massoniana forest stand production in Hunan Province, southern China. Ecological Modelling, 250(??):58–71, February 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012005133.

Zhang:2010:ENA

[ZYFL10] Yan Zhang, Zhifeng Yang, Brian D. Fath, and Shengsheng Li. Ecological network analysis of an urban energy metabolic system: Model development, and a case study of four Chinese cities. Ecological Modelling, 221(16):1865–1879, August 10, 2010. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380010002577.

Zhang:2015:ENM

[ZYKS15] Zhibin Zhang, Chuan Yan, Charles J. Krebs, and Nils Chr. Stenseth. Ecological non-monotonicity and its effects on complexity and stability of populations, communities and ecosystems. Ecological Modelling, 312(??):374–384, September 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print),

1872-7026 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0304380015002537.

Zhang:2011:EAU

[ZYLY11] Yan Zhang, Zhifeng Yang, Gengyuan Liu, and Xiangyi Yu. Emergy analysis of the urban metabolism of Beijing. Ecological Modelling, 222(14):2377–2384, July 24, 2011. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380010004886.

Zheng:2018:AAD

[ZYXT18] Bo Zheng, Jianshe Yu, Zhiyong Xi, and Moxun Tang. The annual abundance of dengue and Zika vector Aedes albopictus and its stubbornness to suppression. Ecological Modelling, 387(??):38–48, November 10, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018302928.

Zhou:2017:SSM

[ZYZ17] Zheng-Xi Zhou, Ren-Cheng Yu, and Ming-Jiang Zhou. Seasonal succession of microalgal blooms from diatoms to dinoflagellates in the East China Sea: a numerical simulation study. Ecological Modelling, 360(??):150–162, September 24, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0304380016306871.

Zhang:2017:AEE

[ZZ17] Hua Zhang and Paul V. Zimba. Analyzing the effects of estuarine freshwater fluxes on fish abundance using artificial neural network ensembles. Ecological Modelling, 359(??):103–116, September 10, 2017. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380016306986.

Zamberletti:2018:CAW

[ZZA^{+}18] Patrizia Zamberletti, Marta Zaffaroni, Francesco Accatino, Irena F. Creed, and Carlo De Michele. Connectivity among wetlands matters for vulnerable amphibian populations in wetlandscapes. Ecological Modelling, 384 (??):119–127, September 24, 2018. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380018301686.

Zhang:2018:EID

[ZZB^{+}18] Sha Zhang, Jiahua Zhang, Yun Bai, Upama Ashish Koju, Tertsea Igbawua, Qing Chang, Da Zhang, and Fengmei Yao. Evaluation and improvement of the daily boreal ecosystem productivity simulator in simulating gross primary productivity at 41 flux sites across Europe. Ecological Modelling, 368(??):205–232, January 24, 2018. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380017304854.

Zhang:2015:ENA

[ZZF15] Yan Zhang, Hongmei Zheng, and Brian D. Fath. Ecological network analysis of an industrial symbiosis system: a case study of the shandong lubei eco-industrial park. Ecological Modelling, 306(??):174–184, June 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014002427.

Zhang:2015:CFL

[ZZFH15] Liang-Xia Zhang, De-Cheng Zhou, Jiang-Wen Fan, and Zhong-Min Hu. Comparison of four light use efficiency models for estimating terrestrial gross primary production. Ecological Modelling, 300(??):30–39, March 24, 2015. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015000137.

Zhang:2016:SMO

[ZZJ^{+}16] Honghui Zhang, Yongnian Zeng, Xiaobin Jin, Bangrong Shu, Yinkang Zhou, and Xuhong Yang. Simulating multiobjective land use optimization allocation using multi-agent system — a case study in Changsha, China. Ecological Modelling, 320(??):334–347, January 24, 2016. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380015004950.

Zhao:2019:EME

[ZZL^{+}19] Yunxia Zhao, Jihong Zhang, Fan Lin, Jeffrey S. Ren, Ke Sun, Yi Liu, Wenguang Wu, and Wei Wang. An ecosystem model for estimating shellfish production carrying capacity in bottom culture systems. Ecological Modelling, 393(??):1–11, February 1, 2019. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380018304162.

Zhou:2014:MES

[ZZLZ14] Decheng Zhou, Shuqing Zhao, Shuguang Liu, and Liangxia Zhang. Modeling the effects of the sloping land conversion program on terrestrial ecosystem carbon dynamics in the Loess Plateau: a case study with Ansai County, Shaanxi province, China. Ecological Modelling, 288(??):47–54, September 24, 2014. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (electronic). URL

http://www.sciencedirect.com/science/article/pii/
S0304380014002658.

Zhao:2013:PST

[ZZT13] Xiaodong Zhao, Hongjian Zhang, and Xiaolei Tao. Predicting the short-time-scale variability of chlorophyll a in the Elbe River using a Lagrangian-based multi-criterion analog model. Ecological Modelling, 250(??):279–286, February 10, 2013. CODEN ECMODT. ISSN 0304-3800 (print), 1872-

7026 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S030438001200556X.

Zhang:2013:EER

[ZZY^{+}13] Jinyun Zhang, Yan Zhang, Zhifeng Yang, Brian D. Fath, and Shengsheng Li. Estimation of energy-related carbon emissions in Beijing and factor decomposition analysis. Ecological Modelling, 252(??):258–265, March 10, 2013. CO-DEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380012001743.

Zhao:2015:RCT

[ZZZW15] Y. J. Zhao, L. Zeng, A. L. Zhang, and Y. H. Wu. Response of current, temperature, and algae growth to thermal discharge in tidal environment. Ecological Modelling, 318(??):283–292, December 24, 2015. CODEN ECMODT. ISSN 0304-3800 (print), 1872-7026 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0304380014006140.